A Bibliography of Scientific American: 1880–1889

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

02 May 2020 Version 1.04

Title word cross-reference

$200 [868].

10 [6134]. 100-Ton [226]. 110-Ton [8960]. 1812 [4574]. 1815 [9585]. 181st [10741]. 1880 [696]. 1881 [2418, 1561]. 1882 [3241, 3297, 3304]. 1883 [180, 604, 1152, 550, 1101, 1230, 3479, 4345]. 1887 [9263]. 1888 [9911]. 1889 [7487, 10183, 11097, 10985, 11215, 11473, 11752]. 1892 [12383, 12621]. 18lst [8291]. 19 [9628].

6 [3846]. 6th [4256].

A. [6807]. A.A.A.S. [5633]. Aard [514]. Abacus [3811]. Abbe [10648]. Abdominal [11776]. Abendroth [8390]. Aberdeen [12598]. Ability [6920]. Abroad [1178, 933]. Absorption [642, 3644, 10422]. Abutment [6458]. Academy [256, 346, 1076, 1108, 134, 257, 405, 1289, 1631, 2334, 3459, 4148, 6818, 7108, 7640]. Accelerating [7057]. Accession [9743]. Accessories [6887, 8622]. Accident [1179, 1257, 3295, 4005, 7884, 8355, 12329]. Accidents [2381]. Account [8892]. Accretion [8647]. Accumulators [2928, 6696, 9086]. Accuracy [3645]. Acid [1353, 865, 2873, 4386, 5435, 5339, 7981]. Acids [5412]. Acme [1339, 2330, 8139]. Acoustic [2776, 4999]. Across [5194, 8611, 8905, 6223]. Acting [14, 2111, 3374, 4901, 6036, 6008, 5687, 6386, 7495, 9179]. Action [590, 523, 2419, 7810, 7733]. Actions [12421]. Actual [1483]. Actuated [7371, 11815]. Adapted [6892]. Adders [4528]. Adding [4851, 7398, 8334]. Addition [3782]. Additional [10865]. Additions [12321]. Address [8740]. Addressing [11674]. Adelantado [10083]. Ader [3067]. Adhesive [8740]. Adirondack [4143]. Adjustable [538, 2475, 2859, 4098, 4529, 4605, 7340, 6741, 6382, 7645, 7619, 8010, 8870, 9042, 9657, 9720, 9455, 9513, 9589, 9702, 9958, 9987, 11384, 11284, 10551, 11598]. Adjusted [9485]. Adjuster [2011, 4730, 7468, 10259]. Adjusting [1014, 741, 3125]. Administration [11490]. Admiralty [8835]. Adulteration [427, 9219]. Adulterations [764, 1727, 4698, 4567]. Advancement [3358, 5634, 8637, 10095, 10120, 10070, 11412, 12599, 12132, 3195, 4279, 4895]. Advent [6499]. Advertisements [1254, 1280]. Advertising [11130]. Æchmeas [6212]. Æeolian [4220]. Aerated [7893]. Aeration [11088]. Aerial [3397, 4172, 4822, 4755, 6165, 6436, 8754, 11639, 5903]. Aerolite [6929]. Aeroplane [6459]. Aerostats [8526]. Ætna [2591]. Affected [2925]. Affecting [2234]. African [10673, 10467, 417, 5846, 8424]. After [3807, 5833, 6088]. Again [9928]. Against [1581, 4993, 6333, 8937, 9141, 11946, 12361, 7932]. Agassiz [9218]. Agates [2911]. Age [683]. Agent [2886, 7908]. Agglomerated [6660]. Ago [8463]. Agricultural [888, 70, 124, 1305, 1337, 190, 245, 275, 301, 324, 379, 536, 563, 580, 622, 640, 664, 676, 748, 769, 835, 941, 1043, 1162, 1220, 1248, 638, 122, 1390, 1417, 1514, 1541, 1641, 1661, 1785, 1799, 1835, 1858, 1975, 2045, 2132, 2203, 2268, 2323, 2372, 2449, 2504, 2618, 2669, 3539, 3560, 3579, 3586, 3610, 3631, 3659, 3681, 3703, 3726, 2737, 2762, 2867, 2889, 2913, 2931, 3012, 3053, 3070, 3115, 3153, 3179, 3201, 3229, 3286, 3312, 3332, 3382, 3420, 3442, 3467, 3488, 3515, 3196, 3751, 3770, 4400, 4426]. Agricultural [4451, 4479, 4507, 4533, 4559, 4588, 4613, 4638, 4661, 4686, 4711, 4735, 4760, 4785, 4810, 4833, 4857, 4883, 4905, 3795, 3815, 3836, 3859, 3882, 3906, 3932, 3952, 3994, 4016, 4044, 4065, 4106, 4132, 4154, 4175, 4200, 4246, 4265, 4292, 4320, 4348, 4373, 4929, 4955, 5626, 5649, 5670, 5715, 5739, 5764, 5788, 5813, 5837, 5858, 5884, 5908, 5932, 5955, 5978, 6020, 6047, 6076, 6102, 6126, 6153, 6174, 4983, 5008, 5033, 5055, 5080, 5128, 5152, 5175, 5199, 5224, 5249, 5274, 5346, 5373, 5398, 5423, 5449, 5477, 5505, 5528]. Agricultural [5555, 5582, 5604, 5832, 6214, 6242, 6821, 6843, 6868, 6897, 6923, 6943, 6964, 6995, 7021, 7045, 7070, 7099, 7128, 7154, 7209, 7233, 7260, 7279, 7306, 7329, 7352, 7379, 7404, 6265, 6291, 6317, 6347, 6371, 6394, 6416, 6445, 6464, 6488, 6514, 6537, 6564, 6586, 6606, 6627, 6653, 6675, 6699, 6722, 6746, 6772, 6796, 7427, 7453, 8271, 8298, 8323, 8378, 8407, 8432, 8457, 8485, 8514, 8543, 8575, 8601, 8657, 8748, 7501, 7527, 7548, 7573, 7599, 7628, 7685, 7712, 7766, 7793, 7898, 7923, 7945, 7971, 7995, 8019, 8042, 8092]. Agricultural [8143, 8169, 8198, 8220, 8245, 8771, 8794, 9596, 9619, 9642, 9666, 9684, 9709, 9733, 9785, 9809, 9833, 9889, 9915, 9944, 9973, 10002, 10029, 10062, 10085, 10114, 10142, 8825, 8882, 8914, 8941, 8967, 8997, 9052, 9076, 9104, 9130, 9187, 9212, 9239, 9268, 9298, 9326, 9350, 9377, 9406, 9468, 9495, 9523, 9549, 9578, 10168, 10186, 11268, 11292, 11317, 11342, 11369, 11400, 11430, 10210, 10231, 10265, 10294, 10369, 10393, 10427, 10454, 10478, 10511, 10582, 10607, 10627, 10675, 10747, 10812, 10855, 10910, 10929, 10958, 11018, 11040, 11067, 11138]. Agriculture [2333, 9308]. Aid [1088]. Aids [176, 1843, 11223]. Ailantus [1336]. Aillot [6990]. Ailsa [5121]. Aiming [7826]. Air [578, 843, 1153, 98, 749, 1126, 1364, 876, 1462, 1971, 2242, 1992, 2287, 1687, 2753, 3172, 3145, 3352, 3649, 4171, 4386, 4114, 3848, 4706, 4471, 4007, 3763, 5313, 6117, 6496, 7247, 7345, 6837, 6813, 6715, 6232, 8053, 8081, 7789, 8077, 9152, 8965, 9316, 9800, 8991, 9536, 8912, 9727, 10593, 11150, 10324, 10158, 10469, 10785, 11778, 12363, 12475, 11482, 11818, 12514, 148, 8165, 8141, 11802]. Airing [8316]. Ajax [2110]. Akron [1606]. Aladdin [7724, 7734]. Alarm [135, 1853, 2500, 2902, 3275, 2881, 2640, 2703, 4726, 4522, 5967, 5519, 6666, 6582, 8569, 8841, 10643, 10314, 10249, 10621, 12071, 12099, 12282, 12185]. Alaska [6202, 6289, 6314, 6345, 6252, 11790, 11604]. Albany [10180]. Albatross [784]. Albemarle [8476]. Albion [1662]. Alcohol [3503, 4821, 5184]. Ales [11423]. Alexander [286, 3044, 9218]. Alexandria [3596, 3597]. Alfred [9983]. Algeria [11427]. All-Around [9705]. All-Iron [10821]. Allard [8187]. Alleged [3598, 9518]. Allen [2288, 3745, 5313]. Allied [945]. Alligator [1692]. Alligators [11679]. Alloys [2457, 3430]. Almost [3916]. Alphabet [9172]. Alphonso [11197]. Alternating [4892]. Aluminum [8005]. Alva [12216, 9426]. Alvan [9593]. Amalgamator [3903, 7082, 9018]. Amateur [465, 1103, 1135, 1212, 1419, 2387, 9207]. Amateurs [5215, 9830, 9361]. Amazon [964]. Ambition [9360]. Amblystome [2299]. Ambulance [10026]. Ambulans} [11428]. Amending [229]. Amendment [3321, 5410, 9446, 10755]. Amendments [2656]. America [131, 1245, 1632, 2082, 1805, 2726, 3892, 4964, 5052, 5343, 6877, 8666, 9634]. American [856, 205, 12208, 7583, 428, 1234, 154, 178, 974, 42, 58, 1037, 1072, 1109, 1147, 1180, 1216, 1247, 1268, 1293, 1329, 99, 120, 138, 161, 186, 207, 239, 312, 382, 413, 441, 450, 477, 511, 535, 614, 641, 659, 720, 826, 857, 875, 959, 992, 1172, 887, 1031, 1296, 842, 923, 739, 803, 478, 252, 372, 1302, 1017, 1075, 907, 678, 2166, 1558, 1488, 1513, 1533, 1565, 1612, 1663, 1684, 1967, 1981, 2002, 2130, 2180, 2194, 2379, 2540, 2058, 1487, 1848, 1892, 3569, 3163]. American [3216, 2842, 2920, 2986, 3060, 3171, 3166, 3320, 4028, 4141, 4129, 4600, 3803, 4099, 4100, 5256, 5193, 5216, 5263, 5315, 5333, 5093, 5844, 5385, 5515, 7303, 6186, 7003, 7336, 7230, 8637, 7639, 7954, 7487, 10095, 9677, 9773, 9006, 10070, 10756, 11212, 10967, 11412, 12599, 12383, 12296, 11850, 12177, 11982, 12249, 12515, 12206, 11465, 12132, 11535, 2486, 6188, 1993, 3655, 9728, 3195, 4279, 4895, 6108, 126, 673, 1151, 3046, 5066, 5590, 5935, 5936, 9382, 8865, 12430, 12056]. Americans [3683]. Americanus} [8539]. Americas [12372]. Amiral [11779]. Amito [12451]. Ammonia [3088, 2886, 4190, 7093]. Ammunition [10737]. Among [5122, 5243, 1700, 5089, 5168, 5220]. Amphora [1811]. Amsterdam [4443]. Amusements [10263]. Analysis [3824]. Analyzing [2582]. Ancel [12005]. Anchor [6687, 6359, 8819, 9871, 11878]. Anchors [10276]. Ancient [621, 2469, 3721, 3552, 5759, 5846, 5464, 8453, 11656, 1252]. Andre [187, 2755]. Andromeda [6864]. Anemogene [8878]. Anemometer [6431]. Angeles [11656]. Angle [7168]. Animal [583, 410, 1637, 1605, 3480, 5102, 9148, 9279, 11383, 10604, 11879]. Animals [1032, 2444, 1826, 10074, 11147, 11542, 3536, 3013]. Annealing [5409, 12496]. Anniversary [9195]. Annual [7640, 9059, 10732]. Annular [3371]. Annunciator [11078]. Anomalies [6086]. Answers [3827]. Ant [15, 9008, 10138, 9293, 2863, 5712]. Ant-Eater [9008, 2863]. Antelope [11114, 10809]. Antelopes [417]. Anthony [4963]. Anthracite [1074, 1273, 1242, 1906]. Anthropology [10592]. Anti [602, 6915, 8592, 9851]. Anti-Clogging [9851]. Anti-Friction [6915, 8592]. Anti-Patent [602]. Antique [5241]. Antiquity [585, 4442, 5025, 6625]. Antiseptic [6474]. Antiseptics [3322]. Antlers [4343]. Ants [357, 1407, 4118]. Antwerp [7967]. Anvil [660, 2214, 1441]. Any [6711]. Anzin [7257]. Apartments [12363]. Apertures [8393]. Appalling [11829]. Apparatus [1014, 97, 943, 456, 689, 385, 1132, 383, 457, 897, 718, 1021, 506, 192, 507, 501, 915, 349, 643, 1657, 2018, 1443, 1881, 2424, 1572, 2198, 2170, 2567, 3326, 3376, 3088, 3481, 2816, 3455, 2961, 3186, 3093, 2874, 3556, 3169, 3505, 3745, 4208, 4314, 4171, 3985, 3854, 4357, 4468, 4731, 3947, 3944, 4096, 4059, 3986, 3828, 3929, 4769, 4038, 5145, 5146, 5618, 5577, 6015, 5874, 5386, 5517, 5471, 6737, 6440, 7093, 6301, 7144, 6893, 6406, 6916, 6311, 6889, 7267, 6232, 6387, 6482, 6235, 7251, 6481, 6765, 8393]. Apparatus [8286, 7938, 7893, 8188, 8362, 8106, 8626, 8742, 8336, 7673, 7842, 9336, 9846, 8819, 9941, 8988, 9173, 9517, 9885, 10077, 9659, 9761, 10593, 10670, 10416, 10645, 11304, 11113, 11112, 10872, 10161, 10359, 10785, 11266, 12080, 12628, 12475, 12093, 11931, 12280, 12115, 11970, 12263, 3032, 9774, 912, 1077, 9942, 11914, 11680, 7120, 9024, 9045, 9546, 10451, 10788, 5316, 12058]. Apparition [6649]. Appear [9399]. Appendage [11706]. Apple [3511]. Appliance [711, 459, 436, 2218, 10100, 1338]. Appliances [4359, 3737, 4051, 7056, 10345, 3547]. Application [549, 5380, 5809, 6862, 8100]. Applications [3049, 6919, 11641]. Applied [6073]. Apprenticeship [1033]. Approaches [12121, 12044]. Approximate [32, 632]. April [2691, 3826, 5026, 6274, 10274, 11537, 7641, 7537]. Apron [12159]. Apteryx [4854]. Aquarium [4879, 2948, 4528, 6562, 11679]. Aquariums [9830]. Aquatic [5498, 6719, 8295]. Aqueduct [2469, 11324, 10807, 10321, 12121, 2438, 7041]. Araometric [1686]. Arbor [4058, 5311]. Arc [9866, 12119, 9206]. Arch [2438, 5295, 8253, 12082]. Archeological [3892]. Archer [3833]. Archipelago [8405, 8318]. Architect [8865]. Architectural [7545]. Architecture [6188]. Arctic [2770, 5294]. Arena [8295]. Argand [3028]. Argentiferous [1795]. Argentine [2088]. Argol [4623]. Argols [4672]. Argus [7737, 4782]. Arizona [157, 4012]. Arm [4417, 9112, 12283]. Armadillo [10579]. Armament [10056]. Armaments [11995]. Armatures [3893]. Armed [8861, 9465]. Armor [7525, 8767, 12481]. Armor-Clad [8767]. Armorclad [9759]. Armored [9122, 10552, 12189, 11721]. Arms [2221, 3089, 5574]. Armstrong [11548]. Army [1232, 482, 5974, 5725, 7126, 8846, 724]. Arrangement [5578, 6222, 8152, 10444, 12217]. Arrangements [1465, 3377]. Arrester [4674, 7677, 7679, 10362]. Arrival [1259, 6594]. Arsenal [9925]. Arsenio [1319]. Art [1222, 345, 2166, 1689, 6534, 6710, 7624, 8729, 8730, 7708, 8624, 175]. Artesian [51, 1261, 4502, 9898, 10545, 11301, 11562]. Arthur [10616]. Article [7150]. Articles [2528, 4915]. Articulate [11073]. Artificial [1353, 1371, 36, 635, 340, 285, 1464, 2064, 1537, 3920, 6280, 7084, 11006, 11841, 12143, 12258]. Artillery [4163, 12556]. Artist [4187, 7623, 8847]. Artistic [2370, 6330, 7545, 9617]. Artists [11086]. Arts [6867, 7274]. Asa [10719]. Asbestos [1080, 6685]. Ascertain [2023]. Ascertaining [6671]. Ash [1543, 8737, 9120, 9658, 12097, 12182, 12560, 12365]. Aspect [11519]. Aspects [2691, 3619, 3453, 2876, 3123, 2771, 3087, 3355, 3277, 3142, 3826, 4774, 4624, 4117, 4006, 4079, 4724, 3919, 4234, 4492, 4360, 4900, 5026, 5895, 5442, 5257, 5356, 5988, 5140, 5548, 5776, 5679, 6162, 6274, 7324, 7192, 6640, 6523, 6783, 6413, 6758, 7005, 6909, 7390]. Aspirator [698]. Aspirators [11860]. Asquith [4543]. Assay [2510, 2457]. Assaying [1731]. Assignments [594]. Assist [10708]. Assistant [11587]. Association [1408, 3358, 8637, 10095, 10070, 11412, 12599, 12132, 3195, 4279, 4895, 178, 739, 803, 854, 823, 603, 2058, 3219, 3243, 7388, 12177, 11463, 12207]. Assured [131]. Asteroids [6251]. Astigmatic [12626]. Astronomer [2152]. Astronomical [766, 1294, 4028, 4141, 5589, 6832, 7078, 7569, 8839, 12249, 12096, 11061]. Asylums [9772]. Atalanta [6907]. Atavism [2938]. Atchison [8]. Athenian [3984]. Atlanta [2321, 4565, 10056, 9114, 9194, 9675]. Atlantic [7019, 7162, 6223, 7327, 7637, 7746, 8905, 11678, 374]. Atmosphere [5523, 5786, 8525, 10438]. Atmospheric [4009, 3899, 5099, 8066]. Atomizer [2604, 7319]. Attaching [9934, 5360]. Attachment [1129, 1039, 1262, 806, 100, 930, 1052, 2538, 1412, 2090, 3623, 3811, 4778, 4283, 5221, 5342, 5552, 6576, 7300, 7111, 6984, 7370, 7204, 6596, 6914, 8401, 7440, 8402, 7542, 8588, 8620, 8445, 8161, 8500, 8842, 10105, 8896, 9095, 9335, 8805, 9569, 9311, 9515, 9588, 10124, 9956, 9284, 9827, 9697, 9754, 10643, 10307, 11064, 10165, 10500, 10313, 11194, 10276, 10924, 12212, 11768, 11983, 11444, 12456, 12415, 11698, 12393, 12077, 12436, 5077]. Attachments [6480, 9507]. Attack [9059, 9675, 11352]. Attacks [8937]. Attempt [2755].

Attitudes [6148, 5833]. Attock [6682]. Attraction [9152, 10532]. Attractive [1820, 12124]. Audience [2498]. Audiphone [7937]. Auer [9295]. Auger [4602, 5876, 6689, 10982, 11418]. August [11410, 3619, 4774, 6134, 7324, 10012, 9628, 9263, 11276, 12517, 8102, 8641]. Augusta [11791]. Augustus [5090]. Aunrecognized [1934]. Aurania [4548]. Aurora [3692, 3920]. Ausable [3258]. Austral [3653, 4080]. Australasia [2696]. Australia [3452, 4502, 8568, 8908, 11721, 5710]. Australian [4499, 9400, 10289]. Austrian [6767]. Autographic [7387]. Autographs [9155]. Automatic [1129, 456, 209, 1182, 1331, 1107, 2318, 1385, 3395, 2961, 3007, 2751, 2988, 4414, 4440, 3927, 4359, 4280, 4488, 4168, 3872, 5967, 6122, 5965, 5439, 5773, 6044, 5832, 6261, 6501, 6694, 6433, 6500, 6691, 6911, 6309, 6405, 7246, 6645, 6934, 6234, 6528, 8401, 7645, 8139, 7593, 7677, 7419, 8038, 8474, 8080, 8338, 7701, 7912, 8060, 7492, 7648, 8480, 7521, 7889, 7679, 7730, 8807, 10044, 9560, 9939, 9748, 9726, 9760, 10018, 9312, 9011, 10944, 10292, 11355, 11419, 10804, 11359, 11029, 12097, 11833, 11479, 11764, 12300, 12476, 12499, 11681]. Automatic [11749, 12257, 12010, 5470]. Automatically [9912]. Automaton [4197]. Auxiliary [4491, 7893]. Avenue [7203]. Avenues [1963]. Average [7980]. Avoid [10918]. Awl [2708]. Awning [1666, 9485]. Awnings [9089]. Ax [9068, 9934]. Axle [3554, 4390, 5872, 5336, 4975, 5680, 6810, 9514, 9991, 10945, 10646, 11054, 11896, 11597, 12029]. Axles [1249]. Axolotl [5579]. Ayres [12420]. Aztec [6692].

B [11574, 12236]. B.R. [4393]. Babcock [341]. Babies [1559]. Baby [1876, 6260]. Babyboussa [467]. Bacchante [2567]. Back [4658, 6138, 6622, 6338, 6764, 7726, 9364, 9989, 9823, 11880]. Back-Washing [6764]. Backband [12162]. Backus [9229]. Backward [2395]. Bacteria [2630]. Bacterial [4625]. Bad [5043, 8730]. Bag [3192, 4467, 4219, 6167, 5751, 7036, 7368, 7200, 6512, 7226, 10381, 10412, 11673]. Bagging [10789]. Bagholder [4120]. Bahamas [782]. Baillehache [7301]. Baird [9519]. Bait [10598]. Bake [11003, 11647]. Bakeries [6962]. Bakers [9993]. Baking [2886, 5636]. Baku [4976, 5267, 8595]. Balance [1014, 1129, 2507, 4363, 7617, 8704, 11354, 10311, 9547, 2476, 9461, 9516]. Balanced [4497, 4316, 5823, 7061, 7205, 8420, 9365]. Balancing [4804, 5993, 6161]. Baldness [5774]. Bale [7680, 10307]. Baling [3249, 3127, 6163, 6525, 12536]. Ball [8217, 8599, 11980]. Ballast [9102]. Ballet [4974]. Balloon [4607, 5667, 8242, 7860, 11202]. Balloons [3759, 7126, 7164, 7695]. Ballot [9803]. Balls [8566]. Balsam [586]. Baltimore [10953, 12156, 12502]. Bamian [8482]. Band [1131, 2497, 2478, 2615, 3149, 4169, 4459, 6138, 5676, 6622, 9364, 10970, 11746, 12188, 11812]. Band-Saws [11812]. Banding [10569]. Bands [1592]. Banjo [8593]. Bank [5407, 9877, 10866, 10488]. Banksian [919]. Bar [2345, 2591, 3622, 2734, 5921, 5142, 5320, 6454, 7727, 9009, 10076, 10315, 12505, 11468, 11814]. Barb [4026]. Barbastelle [468]. Barbed [6277, 11726, 10436]. Barbs [4274].

Barcelona [9731]. Barff [6057]. Barge [1506]. Barker [2816, 9638]. Barn [6113]. Barnard [7669, 11780]. Barney [4488]. Barometer [1366, 892, 1586, 8531, 12056]. Barrack [7857]. Barrel [1458, 2611, 4628, 8281, 7566, 11309, 10954, 10157, 11772, 11499]. Barreled [2643]. Barrels [4866, 5170, 5187, 5143]. Barrett [9531]. Bars [6912, 7168, 12586]. Bartholdi [1207, 3965, 5240]. Bartlett [11908]. Barton [12497]. Base [7616]. Baseball [8729, 8612, 8624, 8116]. Basic [5947, 8584]. Basin [3948, 3938]. Basins [2036, 1984, 6036, 7831, 7442]. Basis [5537]. Basket [1548, 4476, 4553, 6121, 8760, 11591, 302]. Bass [3603, 12264]. Bass-Reliefs [3603]. Bath [691, 1467, 1687, 2508, 2839, 7266, 8512, 10785, 10947, 11599, 11802]. Bathing [718]. Batoidei [920]. Baton [9362]. Batrachian [8939]. Batteries [9745, 9559, 8767, 11178, 11667, 464, 11013]. Battery [1325, 121, 1421, 2249, 2091, 2482, 3394, 3742, 5001, 6341, 8936, 10107, 10827, 12217, 12343, 12439, 12632, 12602, 12325, 492, 9199, 9910]. Battery-Stored-Up [2249]. Batting [8116]. Battle [10552]. Battlefield [932]. Battles [9899]. Baudin [11779]. Baudre [7276]. Baxter [11110]. Bay [9140, 11761, 796, 5515]. Bayonets [7494]. be [2358, 1876, 3598, 4917, 6147, 7019, 6430, 6226, 7338, 6782, 8524, 8497, 11912]. Beach [10320, 12317, 12155]. Bead [5076, 6382]. Beam [9340, 12568]. Beams [12138]. Beans [269]. Bear [8512, 9022, 10984]. Bearing [5363, 6279, 8592, 11597, 11477, 12558]. Bearings [4568, 5967]. Beaten [6572]. Beater [637, 7294, 7372, 8284]. Beating [12567]. Beatrice [10494]. Beautiful [5219]. Beaver [4262, 12437]. Bed [2191, 2568, 3718, 4466, 7565, 8032, 9232, 10468, 11082, 11058]. Bed-Pans [3718]. Beds [156, 6786, 11308]. Bedstead [7619, 8010]. Bedsteads [7347, 11196]. Bee [5544, 5443, 6554]. Beecher [8838]. Beef [3242, 3646, 4274]. Beer [802, 1484, 1522, 6878, 11588]. Bees [11363]. Beet [6042, 9116, 9072]. Beetle [1137, 1617, 2420, 3535, 3114, 8196, 10623]. Beetling [9968]. Before [86, 4232, 5568, 4940, 5017, 10464, 7535, 9419]. Belgian [7377, 9616]. Bell [3122, 4656, 5245, 6741, 7029, 6624, 9567, 10017, 10439, 10603, 10757, 12054, 1379, 5962, 7413, 8353, 8386, 7932, 9419, 9358, 10241]. Bellows [7985, 9781]. Bells [2709, 6601, 8906]. Belt [1318, 1750, 3247, 2616, 2752, 4682, 3964, 5943, 6508, 6664, 7594, 9633, 9730, 10255, 10942, 10310, 10414, 12609]. Belted [9492]. Belting [12078]. Belts [3051, 5232, 5726, 9592]. Benbow [7066, 8279, 10765]. Bench [1948, 1878, 1665, 5214, 5870, 9451, 9046]. Bench-Plane [1878]. Benches [4921]. Bend [7966]. Bender [5297]. Bending [2505, 1774, 6039, 6974, 7706, 9607, 12624]. Benefits [7858]. Bengal [260]. Benjamin [5044, 908]. Benzole [809]. Berdan [6876, 9125]. Bergen [3786, 11168]. Bergh [10219]. Berlin [154, 864, 1289, 1723, 3720, 2948, 4528, 6562, 7737, 7790, 8371, 9886, 9071, 9099, 10025, 10673, 11679, 11816, 11839, 12083, 1334, 752]. Bernissart [4707]. Berrie [11643]. Berth [1988, 10943]. Besemer [12260]. Besides [8078]. Bessemer [2794]. Best [10918, 10864, 12320]. Bethlehem [4188]. Better [3187]. Between [684, 968, 6356, 9185, 11378, 1698, 1721, 4841, 8479, 8216, 11125, 11256]. Bevel [6956, 7618, 8809]. Beveling [2270, 7168]. Bevels [8230]. Bianchi [4975]. Bicycle [5868, 7088, 8016, 10026, 9371, 12098, 11985, 12399]. Bicycles [7993, 11064, 12050]. Bicyclette [9262]. Big [888, 1104, 3452, 7966]. Bilboa [10890]. Bill [1668, 5030, 7800, 10055, 630, 86, 2907, 9333, 10637, 10464, 10465]. Billed [919]. Billiard [805, 4337, 5124, 6643, 7091, 7652]. Bills [5408, 4940, 553, 4996, 5018, 5017, 5541, 5043, 4968, 5138, 7535]. Binary [642, 3391]. Binder [742, 11002, 11626]. Binding [5118, 6686, 8784]. Binns [10569]. Binocular [430]. Bird [2571, 1668, 4879, 5005, 6367, 7847, 8443, 8113, 9265, 10387, 12297, 2646]. Birds [2020, 2696, 8217, 9614, 9886, 11147, 9460]. Birmingham [8076, 11037]. Birthplace [10764]. Biscuit [12367]. Bishop [11763]. Bit [5876, 6833, 10106, 9570, 10225]. Black [4705, 2759, 9466]. Blackford [1597]. Blacking [824, 12140]. Blacksmith [8110]. Blackwell [2768]. Blade [9068]. Blake [136, 4302, 7231]. Blankets [9397]. Blast [451, 350, 3745, 5928, 7093, 7228, 12210]. Blasting [2211, 3670, 10206]. Bleaching [1784, 5242, 9941]. Blight [934]. Blind [1055, 2880, 7299, 7519, 8703, 9452, 9092, 1113]. Blinding [9569]. Blindness [1351, 375]. Block [9309, 10202, 10803, 10521, 11643, 11539, 11650]. Blocking [5386]. Blood [709, 1350, 3654]. Blooming [6865]. Bloomsdale [3222]. Blot [10122]. Blotter [5387, 8987]. Blow [5383, 12114]. Blow-Off [12114]. Blower [8110, 8876]. Blowers [7257, 7546]. Blowing [4102, 6937, 7536, 12518, 12135, 6286]. Blown [6958]. Blowpipe [609, 11952]. Blue [3269]. Boa [6994]. Board [320, 2172, 3578, 5388, 5685, 7984, 8700, 9451, 9847, 9509, 9254, 10046, 8962, 10378, 10952, 11568, 12493]. Boards [5641]. Boasting [2897]. Boat [871, 1224, 1288, 460, 2248, 1401, 2179, 3532, 3186, 3415, 2775, 2878, 4499, 4488, 4576, 4392, 4527, 5030, 5289, 5097, 7349, 7043, 7543, 8654, 9125, 10027, 9234, 8823, 8862, 8818, 9465, 9675, 11304, 10687, 10712, 11028, 10901, 11168, 10474, 12108, 12492, 11507, 11677]. Boat-Lowering [3186]. Boats [484, 4443, 4842, 3898, 5685, 6283, 6761, 8053, 7649, 8506, 8583, 9509, 9984, 11337, 11377, 10152, 11874]. Bob [8033]. Bob-Sleigh [8033]. Boddaert} [12297]. Bodies [6284, 10229, 10509]. Body [4880, 5270, 9876, 10976, 4700]. Bogie [6238, 8927]. Boiler [215, 341, 350, 488, 266, 961, 415, 685, 1078, 436, 2104, 2032, 2312, 2401, 1448, 2471, 2245, 2040, 2473, 1882, 2059, 2081, 2109, 2149, 2168, 2195, 2213, 2292, 2313, 2343, 2361, 2384, 2404, 2443, 2519, 2537, 2382, 1824, 2060, 3395, 3002, 2837, 2801, 2733, 2908, 2608, 2666, 2731, 3065, 3109, 4651, 3987, 4461, 3848, 4850, 4333, 5244, 5664, 6010, 5948, 6146, 5777, 4952, 5642, 6435, 7392, 6262, 6510, 7171, 6602, 6236, 7270, 6858, 6860, 6533, 6834, 8390, 8212, 8746, 8294, 7869, 8335, 7543, 7591, 7647, 8315]. Boiler [8131, 8354, 7654, 9851, 8900, 8910, 9957, 9281, 10072, 9587, 9960, 10023, 8901, 9283, 10248, 11312, 10722, 10740, 11392, 11998, 11741, 12136, 12322, 2451, 2243, 2423, 2501, 4261, 1966]. Boilers [456, 89, 707, 2499, 1879, 1460, 3035, 3224, 3068, 4726, 3805, 5093, 6582, 8313, 7673, 8844, 11174, 11111, 10529, 10763, 11193, 12164, 11767]. Bollee [7499]. Bolster [11950]. Bolt [1020, 6110, 6383, 7320, 6959, 6978, 10071, 9935, 10533, 12235]. Bolting [11855]. Bomb [1159]. Bombard [793]. Bombardment [3596]. Bombe [7349]. Bonds [11257]. Bone [4702, 4705, 5440]. Bones [8111, 1252]. Bonta [10080]. Book [1362, 93, 2296, 1735, 1808, 5191, 5098, 6062, 7085, 9847, 9513, 10684, 11033, 11570, 11623, 11744, 11500, 12276, 12382]. Book-Rack [1362]. Bookbinding [10443]. Books [2, 54, 1198, 1227, 1344, 81, 151, 171, 221, 279, 308, 336, 365, 394, 422, 447, 474, 519, 546, 597, 626, 650, 703, 839, 900, 1004, 1069, 1093, 1144, 1279, 1375, 1398, 1428, 1501, 1530, 1553, 1577, 1624, 1647, 1695, 1791, 1867, 1901, 1930, 1958, 2000, 2098, 2118, 2231, 2306, 2328, 2378, 2434, 2464, 2552, 2600, 2652, 2721, 2743, 2807, 2832, 2849, 2951, 2996, 3080, 3208, 3316, 3339, 3403, 3447, 3472, 3494, 3563, 3591, 3636, 3663, 3708, 3839, 3842, 4739, 4765, 4790, 4815, 4836]. Books [4862, 4888, 3864, 3888, 3910, 3956, 3976, 4000, 4048, 4070, 4136, 4251, 4270, 4299, 4325, 4353, 4377, 4403, 4429, 4454, 4482, 4510, 4538, 4619, 4645, 4689, 4911, 4932, 5766, 5791, 5816, 5861, 5888, 5911, 5937, 5999, 6024, 6050, 6079, 4959, 4988, 5012, 5037, 5058, 5085, 5108, 5179, 5203, 5251, 5278, 5326, 5351, 5402, 5426, 5453, 5480, 5510, 5532, 5560, 5672, 5697, 5718, 5742, 6192, 6217, 7237, 7282, 7308, 7355, 7382, 6322, 6350, 6397, 6420, 6467, 6491, 6516, 6567, 6609, 6630, 6677, 6727, 6749]. Books [6775, 6799, 6846, 6870, 6901, 6967, 7024, 7102, 7130, 7159, 7214, 6811, 7407, 7429, 8347, 8381, 8411, 8436, 8517, 8548, 8605, 8632, 8662, 8692, 8721, 7457, 7480, 7504, 7552, 7578, 7632, 7660, 7690, 7715, 7741, 7769, 7821, 7851, 7879, 7901, 7926, 7948, 7974, 8022, 8045, 8174, 8202, 8223, 8302, 9772, 8797, 8829, 9861, 9894, 9921, 9949, 9978, 10007, 10035, 10066, 10090, 8857, 8918, 8944, 8971, 9002, 9054, 9080, 9108, 9136, 9161, 9245, 9272, 9411, 9440, 9472, 9528, 9555, 9623, 9647, 9672, 9713, 9738]. Books [9766, 9838, 10122, 10171, 10192, 11024, 11045, 11070, 11120, 11186, 11208, 11230, 11249, 11272, 11296, 11346, 10215, 10236, 10300, 10330, 10352, 10400, 10432, 10459, 10515, 10540, 10561, 10586, 10612, 10631, 10656, 10679, 10729, 10751, 10774, 10794, 10914, 10934, 10963, 10991, 10254, 11436, 11583, 12528, 12573, 11614, 11635, 11663, 11686, 11711, 11759, 11785, 11891, 11943, 11962, 12016, 12040, 12063, 12200, 12222, 12244, 12270, 12291, 12311, 12404, 12445, 12467, 12487, 12511]. Boom [8783]. Boot [2054, 3571, 9936, 11856, 12348]. Booth [10554]. Boots [486, 8874, 9453]. Boracic [5435]. Borealis [3920]. Borers [8373]. Boring [1754, 4276, 4896, 3791, 4829, 4240, 4948, 5187, 5799, 8064, 8477, 8904, 11522]. Born [1113]. Bos [11136]. Bosom [6670, 10952]. Boss [1178]. Boston [2078, 2179, 9842, 12418]. Bot [9758]. Botanical [1693, 2021, 2200, 3036]. Botanizing [5733]. Bottle [1019, 582, 1013, 1884, 1681, 2511, 2364, 1728, 3623, 2879, 4501, 5068, 8017, 8896, 10555, 12366, 12079, 7149]. Bottle-Nosed [4501]. Bottler [4722]. Bottles [1111, 1412, 5854, 9697]. Bottom [4466, 11221]. Bottom-Trip [11221]. Bottoms [9233]. Boulogne [10831]. Boulogne-Sur-Mer [10831]. Bouquet [8162]. Bourgogne [7637]. Bourry [3602]. Bow [1356, 3830, 10078, 3511]. Box [236, 530, 2508, 2902, 2910, 2693, 3353, 2670, 4521, 4627, 5336, 5441, 5826, 5575, 5854, 6166, 5414, 5780, 6364, 6766, 6735, 7142, 7680, 8676, 7911, 8791, 9017, 9421, 9542, 8869, 9803, 10945, 10646, 11361, 10875, 10824, 10689, 12001, 12496, 12140]. Boxes [6861, 9663]. Boxing [7122]. Boy [11910]. Boynton [12098]. Boys [7243, 8640]. Brace [5687, 7736, 10106, 11196, 11129]. Bracket [4082, 8642, 7836, 10051, 11102, 11198, 12072]. Bradford [5418]. Brake [988, 749, 1971, 2017, 3112, 4543, 4440, 6934, 7397, 6578, 6621, 8392, 9335, 9337, 10018, 9372, 9700, 10134, 9037, 9906, 10804, 11359, 10202, 11105, 11778, 11539, 12024, 11443, 11564, 12004]. Brakes [5747, 10805]. Bran [655, 3760]. Branch [6507]. Branchiopod [1546]. Branding [3968]. Brandy [9913]. Brass [1383, 5412]. Braun [8965]. Brave [4795]. Brazil [502, 8854, 12380]. Brazilian [23]. Bread [4583, 4435, 5184, 10050]. Breadstuffs [3854]. Break [3144]. Breaker [2399]. Breakfast [2199]. Breaking [3510]. Breakwater [8766]. Breast [5162]. Breech [3721, 8682, 7474, 11244, 12033]. Breech-Loader [3721]. Breech-Loading [8682]. Breeds [8909]. Brewers [12248]. Brewery [2532]. Brewing [12384]. Brick [1246, 491, 2712, 5441, 6990, 8427, 10645, 11385, 11768, 12413]. Bricks [12001]. Brickwork [7988]. Bridal [12067]. Bridge [458, 732, 504, 554, 1698, 1721, 2057, 3656, 4005, 3980, 3987, 3916, 4330, 4622, 4441, 4341, 3960, 4752, 4599, 5112, 5819, 5063, 6207, 6635, 6623, 6988, 6521, 6458, 8253, 7799, 9865, 9140, 9998, 9318, 10636, 10741, 11394, 10463, 10890, 12253, 11911, 11809, 12398, 11736, 9547, 8291, 11895, 1841, 2416, 2768, 3650, 4420, 3967, 4255, 5853, 6682, 6298, 6412, 7146, 7434, 8959, 10616, 11338, 11253, 11048, 10697, 12227, 10137]. Bridgeport [11895]. Bridges [2397, 4772, 12612, 11742]. Bridging [9357]. Bridle [10312, 10225, 10980]. Brighton [10320]. Brims [9147]. Britannia [10651]. Britannic [9185]. British [854, 3219, 3451, 4307, 5974, 7241, 6544, 9533, 9085, 9489, 8835, 10756, 10467, 11721, 11492, 12580]. Broadway [1150, 1650, 5355]. Broiler [4340]. Broken [4548, 6406, 8353]. Bromide [7596, 12133]. Bronchial [542]. Bronze [251, 1811, 2794, 7315, 7092, 7140, 9617, 9502]. Brooklyn [3967, 11253, 10920, 1074, 1841, 1698, 1721, 4420, 7203, 7557, 8497, 12144, 12397, 11520]. Brooks [5514, 5186, 5486, 5550, 9572]. Brooms [10980]. Broths [2704]. Brougham [187]. Brown [5305, 12555]. Brown-Sequard [12555]. Bruijn [10138]. Bruijni [10138]. Bruning [2064]. Brunton [1010]. Brush [824, 879, 1656, 1842, 5848, 5689, 8705, 9535, 9459, 11701, 4828, 5771, 8726, 7720]. Brushes [5221, 10312, 10980, 10261]. Brushing [7470]. Brussels [9911]. Bubbles [6583, 8536]. Bucephala} [693]. Buchanan [8836, 8993]. Buck [6145, 7873, 10809]. Buckboard [11540]. Bucket [2290, 6171, 5189, 5572, 6041, 6434]. Buckle [771, 4609, 6138, 5196, 5365, 6986, 6740, 8532, 8160, 10413, 11496, 12583]. Budapesth [7062]. Buenos [12420]. Buffalo [4648, 11379, 6941]. Bug [4697]. Buggies [11777]. Buggy [3830, 6205, 8034]. Bugs [2803]. Builder [4139, 8865]. Builders [7859, 10618]. Building [820, 65, 3926, 4434, 4798, 5407, 6737, 9062, 9652, 10636, 10260, 10180, 11376, 12189, 11650, 12000, 12504, 243]. Buildings [4077, 5664, 5167, 5620, 9852, 10249, 11387, 10921, 11742]. Built [5567, 6756, 8164, 9401, 1993]. Built-Up [6756]. Bulkhead [7841]. Bulkheads [3922]. Bull [1951]. Bullfrog [7990]. Bullion [2457]. Bundle [7296, 11770]. Bungling [10995]. Bunk [9226]. Bunkers [4597]. Buoy [6044, 10642]. Buoying [182, 11407]. Bureau [11, 4303, 656]. Burgess [7319]. Burglar [135, 2881, 4010, 6666, 12071, 11974]. Burial [3735, 5474, 7313]. Buried [525, 8497]. Burmah [7896]. Burner [1024, 2905, 6255, 6306, 7889, 9295, 10045, 9630, 10906, 11675, 11652, 11497]. Burners [3028]. Burning [1176, 820, 3046, 7166, 7808, 9538, 10597]. Burst [634, 6830]. Bursting [3893, 5170, 1593]. Burtons [1164]. Burying [7399]. Bush [11397]. Business [1134, 1, 625, 649, 674, 702, 728, 759, 788, 816, 838, 872, 899, 927, 950, 979, 1003, 1028, 1044, 1068, 1091, 1118, 1143, 1168, 1197, 1226, 27, 1278, 1307, 1310, 1343, 53, 80, 105, 128, 150, 170, 200, 219, 247, 278, 307, 335, 364, 393, 421, 446, 473, 496, 518, 545, 570, 596, 1374, 1998, 2028, 2049, 2073, 2096, 2116, 2140, 2162, 2184, 2205, 2229, 2256, 2278, 2304, 2326, 2353, 2375, 2391, 2412, 2433, 2462, 2493, 2523, 2551, 1397, 1427, 1454, 1478, 1499, 1528, 1552]. Business [1575, 1600, 1622, 1645, 1673, 1694, 1718, 1742, 1764, 1789, 1815, 1837, 1865, 1899, 1928, 1956, 1978, 2578, 2598, 3136, 3160, 3183, 3207, 3237, 3264, 3289, 3315, 3338, 3363, 3385, 3402, 3426, 3448, 3474, 3497, 3520, 3545, 3562, 3590, 3613, 3635, 3662, 3686, 3707, 2624, 2650, 2678, 2698, 2719, 2741, 2764, 2787, 2806, 2830, 2848, 2869, 2893, 2916, 2934, 2950, 2976, 2995, 3016, 3039, 3056, 3078, 3102, 3118, 3731, 3755, 4324, 4352, 4378, 4404, 4430, 4455, 4483, 4512, 4537, 4562, 4594, 4618, 4644, 4664, 4691, 4713]. Business [4738, 4764, 4789, 4814, 4835, 4861, 4887, 3774, 3800, 3819, 3841, 3863, 3887, 3909, 3936, 3955, 3975, 3999, 4020, 4047, 4069, 4091, 4110, 4135, 4159, 4180, 4204, 4226, 4250, 4269, 4297, 4025, 6086, 4909, 4933, 5509, 5531, 5559, 5585, 5607, 5629, 5652, 5673, 5698, 5719, 5743, 5767, 5792, 5840, 5862, 5889, 5912, 5938, 5958, 5982, 5998, 6023, 6051, 6080, 6104, 6129, 4958, 6156, 4987, 5011, 5038, 5059, 5084, 5107, 5132, 5155, 5178, 5204, 5227, 5252, 5277, 5302, 5327, 5350, 5376, 5403, 5427, 5452, 5481, 5964, 5986]. Business [6176, 6193, 6776, 6800, 6823, 6847, 6871, 6900, 6925, 6947, 6968, 6998, 7025, 7048, 7073, 7103, 7131, 7158, 7185, 7213, 7236, 7262, 7283, 7309, 7332, 7356, 7383, 6218, 6245, 6267, 6294, 6323, 6351, 6374, 6396, 6421, 6448, 6466, 6492, 6517, 6540, 6568, 6588, 6610, 6631, 6656, 6678, 6702, 6726, 6750, 6931, 7408, 7430, 8046, 8070, 8095, 8124, 8147, 8173, 8201, 8222, 8249, 8275, 8303, 8325, 8348, 8382, 8412, 8437, 8459, 8490, 8518, 8549, 8578, 8606, 8633, 8661, 8693, 8722, 7458, 7481, 7505, 7530, 7553, 7579]. Business [7603, 7633, 7661, 7689, 7716, 7742, 7770, 7795, 7822, 7852, 7880, 7902, 7927, 7949, 7975, 7998, 8023, 8752, 8775, 9473, 9498, 9527, 9554, 9581, 9601, 9624, 9646, 9670, 9689, 9712, 9737, 9765, 9789, 9814, 9837, 9860, 9893, 9920, 9948, 9977, 10006, 10034, 10065, 10089, 8798, 10117, 8830, 8858, 8886, 8919, 8945, 8972, 9001, 9029, 9055, 9081, 9107, 9135, 9160, 9191, 9214, 9244, 9271, 9302, 9329, 9354, 9384, 9412, 9441, 10147, 10172, 10793, 10815, 10838, 10859, 10886, 10913, 10933, 10962, 10990, 11022, 11044, 11069, 11094, 11119, 11141]. Business [11164, 11185, 11207, 11229, 11248, 11271, 11295, 11320, 11345, 11373, 10191, 11403, 10214, 10235, 10269, 10299, 10329, 10353, 10373, 10399, 10430, 10458, 10482, 10514, 10539, 10560, 10585, 10611, 10632, 10657, 10680, 10704, 10728, 10750, 10773, 11435, 11458, 12039, 12062, 12086, 12104, 12128, 12151, 12173, 12201, 12221, 12243, 12269, 12290, 12310, 12333, 12355, 12376, 12405, 12425, 12444, 12465, 12488, 12509, 12529, 12551, 12574, 11485, 12594, 12616, 11512, 11529, 11558, 11582, 11613, 11634, 11661, 11685, 11712, 11732, 11757, 11786, 11805, 11824, 11845, 11866, 11890, 11918, 11941, 11960, 11989]. Business [12014]. Bustard [9400]. Butchers [7942]. Butler [1036]. Butte [7966]. Butter [406, 1942, 2529, 4521, 5123, 5310, 6663, 9016, 10340, 10425, 11724]. Butter-Flies [10425]. Butterworth [5044]. Butting [11009]. Button [2499, 7227, 7300, 6557, 6816, 6259, 6601, 9224, 9872, 10200]. Button-Set [2499]. Buttonhole [7124]. Buttons [2455, 5357, 7960]. Buy [933]. Buyers [3641].

C [3699]. C. [6401]. C.E. [8836]. Cab [792]. Cabbage [4625, 6733]. Cabinet [509, 3852, 10397, 10524, 10425, 11719]. Cable [4420, 4552, 4994, 5230, 5282, 6478, 7249, 6731, 6828, 8002, 8480, 7931, 9025, 9418, 11037, 11425, 10493, 11253, 11833, 11859, 11837]. Cables [2756, 4599, 6838, 8814, 11967]. Cactus [12419, 7207]. Cæsium [2856]. Cage [4879, 11901]. Cailletet [5577]. Caisson [7295]. Caissons [12429]. Cake [7063, 8926]. Cal [6118]. Calais [10598, 11133]. Calcite [3410]. Calculating [11016]. Calculator [7564]. Calendar [461, 4658, 5619]. Calf [216, 9994]. California [1261, 740, 206, 156, 49, 4695, 6016, 7966, 10196, 12307, 11572, 12230, 7920, 11761, 7207]. Caligraph [7446, 10397]. Caliper [4444, 6956]. Calipering [6708]. Calipers [2479, 3168, 4503, 4076, 6335, 12137, 11550]. Calk [11357]. Call [7018]. Calla [8065]. Called [9184]. Calming [10624]. Caloric [1040, 4023, 9445]. Cam [76]. Cambridge [7376]. Camel [6239, 5974]. Camera [6337, 6792, 6441, 9781, 9202, 9119, 9680, 10955]. Cameras [6790, 9512]. Campaign [3213]. Campbell [12514]. Cams [8681]. Can [1264, 2358, 1728, 3464, 4776, 7019, 8709, 8078, 9608, 11079, 11334, 11909, 11469, 11669, 12117, 966]. Canada [1463, 586]. Canadian [4633, 5493]. Canal [484, 4871, 5289, 7219, 7777, 11507, 957, 62, 2853, 4230, 4949, 5368, 6133, 7336, 6226, 6991, 6180, 7360, 7218, 7375, 7667, 7775, 9900, 9034, 8911, 9868, 9953, 10205, 10708, 11502, 12295, 11993]. Canals [7338]. Canceling [234, 2477]. Cancellation [8386]. Candelabra [7207]. Candlestick [1711, 4725, 4340]. Cane [2297, 1590, 4285, 5852, 7341, 7138, 7807, 9252, 10695, 11976]. Canes [4840].

Canister [5849]. Canning [2502]. Cannon [2261, 3006, 3299, 6028, 5169, 9060]. Cannonading [681]. Cannons [8454]. Canoe [163]. Canopy [9511, 10548]. Canopy-Supporting [9511]. Cans [529, 6301, 7491, 8620, 10132, 11080, 11594, 159]. Can’t [3187]. Can’t-Do-Its [3187]. Cantilever [4622, 7146]. Cap [938, 2014, 6410, 7039, 10338]. Capacity [10808]. Cape [957, 527, 606]. Capensis} [147]. Capillarity [8320]. Capillary [1204, 11913, 8430, 8449, 8296, 10532]. Capital [6198, 6452, 7486]. Capitol [10180]. Capra [10366]. Caps [9749]. Capsule [5875]. Capt. [1224, 1051, 11516]. Captain [11501]. Captive [7126, 7695]. Capture [166, 194]. Capturing [3416]. Car [1105, 1292, 118, 1249, 208, 92, 1191, 322, 1408, 2062, 2246, 2066, 1707, 2176, 1831, 1688, 1751, 1567, 3326, 3166, 3693, 2585, 2815, 2587, 3695, 3456, 3546, 2634, 3022, 4440, 4390, 4728, 3788, 4242, 4779, 4487, 5336, 5190, 5704, 5752, 6170, 5707, 5260, 5420, 5312, 5212, 5683, 5266, 5000, 5516, 5518, 5747, 5780, 5658, 7397, 6941, 6302, 6552, 6577, 7364, 6760, 7389, 6810, 6234, 7117, 6456, 7253, 6276, 6360, 6688, 7115, 6766, 6784, 7859, 8392, 7518, 7539, 7754, 8183, 8259, 8260, 8391, 8560, 8562, 7648, 8425, 8529]. Car [8059, 8185, 7804, 7730, 9335, 8758, 8759, 8924, 9337, 9420, 9025, 9722, 9171, 9726, 10018, 9372, 9700, 8899, 8932, 8957, 9145, 9174, 9483, 9563, 9780, 9848, 9995, 10020, 10125, 9539, 9561, 10134, 9937, 9932, 9011, 9676, 9169, 9770, 11051, 10863, 10221, 10804, 11359, 10646, 10181, 10250, 10319, 10495, 10523, 10784, 10969, 11034, 11081, 11217, 11281, 11310, 11417, 11425, 10391, 10762, 11151, 11278, 11054, 11192, 10163, 10389, 10618, 11202, 10896, 11896, 11443, 11526, 11593, 11670, 11837, 11882, 11885, 12003, 12584, 12434, 11835, 12051, 11477, 12112, 12452]. Car [11838, 11951, 12074, 12232, 11447, 12004, 12163, 11470]. Car-Axle [10646, 11896]. Car-Heater [10391]. Car-Starter [12051]. Caracas [8028]. Carbon [5928, 10638, 11013, 11745, 10422]. Carbonic [4386, 4032, 5339, 7981]. Carbonizer [1431]. Carbons [1510]. Card [6164]. Cardboard [5439]. Carding [7960]. Care [4164, 5437, 12360, 3536]. Careless [6147]. Carelessness [2748]. Carillon [8211]. Carnegie [10488]. Carnrick [10286]. Carolina [8400, 633]. Carp [402, 493, 1042, 6120, 1251]. Carpenter [39, 10893, 11884]. Carpenters [5639, 10048, 12137]. Carpentier [2610]. Carpet [2420, 3098, 3126, 4635, 6477, 7033, 9484, 9753, 9878]. Carpeting [66]. Carriage [235, 1828, 5665, 6260, 7499, 8572, 7760, 8732, 7707, 7447, 10094, 9321, 10710, 10923, 10423, 11454, 11570]. Carriages [7178, 7469, 9319, 8788, 11008, 10620, 11240, 10227]. Carrier [237, 4149, 6063, 6526, 8235, 9120, 8989, 9962, 11152, 12476, 11591, 11770, 12050]. Carry [8755, 7735]. Carrying [8286, 9227, 12280, 12163]. Cars [988, 1112, 930, 1250, 3431, 4420, 5950, 6454, 6696, 8080, 8282, 7727, 8683, 8743, 7915, 9531, 8805, 8894, 8925, 9032, 9699, 10100, 8848, 8898, 8866, 10601, 10599, 10252, 10161, 10335, 12058, 12024, 11775, 12161, 11898, 12436]. Carson [4868]. Cart [4338, 5549, 8108, 8933, 9929, 9988, 11727, 11504, 12111]. Carthage [5241].

Cartridge [1187, 7057, 10547, 10868]. Cartridges [10342, 10338]. Carts [4040]. Carving [6638]. Cascade [3549]. Case [867, 1175, 2088, 2272, 3061, 4797, 5003, 5409, 5500, 5286, 6763, 8469, 8261, 9172, 9701, 10737, 10822, 11282, 11744, 12347, 12432, 2539, 7413]. Cases [1941, 2080, 9286, 10241]. Cash [8989]. Casing [5968]. Cask [1585, 8812, 11798]. Caskets [9168]. Casks [5143, 12248]. Cassowary [4013]. Cast [2473, 4093, 3940, 4211, 5801, 5656, 5285, 6974, 6707, 6732, 6497, 8686, 7464, 11244, 12033]. Caster [11932]. Casting [21, 2062, 3740, 6028, 6135, 6378, 9502, 9124]. Castings [403, 4720, 4436]. Castor [12437]. Casualties [2211]. Cat [1950]. Catalogue [5307, 5935]. Catamaran [6055, 5784, 7017]. Catapult [3482]. Catarrh [542]. Catch [5597, 6011]. Catcher [4467, 5192, 8398, 11879]. Catching [10074, 10598]. Caterpillar [4800]. Catesbiana [7990]. Cathedral [7475]. Catskill [5112]. Catskills [1306]. Cattle [357, 1053, 1408, 1914, 2238, 3242, 5953, 7348, 6794, 7597, 9492, 9434, 10024, 9391, 10603]. Caucasian [1939]. Caudal [11706]. Caught [7277]. Cause [2312, 1771, 2748, 2672, 5208, 5948]. Caused [4115]. Causes [813, 1122, 1855]. Caustic [10422]. Cautery [3829]. Cavalry [7494]. Cave [1773, 1851, 1607, 1584, 2124, 12238]. Cavern [4412, 11801]. Caverns [2637]. Caves [8525, 5710]. Cayenne [3535]. Ceiling [437, 4237, 4121, 10180]. Celebrated [10505]. Celebration [9195, 8804]. Celestial [8980, 9113, 9220, 9505, 9869, 10153, 10305, 10408, 10709, 11074, 11171, 11494, 11994]. Cell [776]. Cellar [1859, 6066, 5898, 8558]. Cellars [5468, 11260]. Cells [12217]. Celluloid [2442]. Celtic [9185]. Cement [6403, 1059]. Cements [1908, 11619]. Cemetery [6793]. Census [3784, 4231]. Centenarian [11063, 11986]. Centenarians [11062]. Centenary [2236]. Centennial [3524, 8804, 9604, 11691, 4566]. Center [1039, 3642, 9509, 9342, 12598, 11569]. Centerboard [8395]. Centering [11113]. Centers [12450]. Central [320, 283, 2144, 4819, 5644, 7840, 10343, 11533, 1947, 3892, 5074, 5052, 8978, 10334, 11147, 10273]. Centreboard [8287]. Centrifugal [2444, 2297, 7732, 9941, 10954, 11605, 8511, 8655]. Centripetal [2444]. Century [7760, 7625, 7623]. Cephalopoda [2069]. Certain [8464]. Certification [33]. Chagres [62]. Chain [12369]. Chains [5256]. Chair [939, 2105, 2196, 4848, 4062, 5119, 6940, 6432, 7269, 7394, 7587, 9751, 10761, 10667, 10417, 10551, 11053, 10526, 10574, 12317, 11624]. Chairs [1737, 8738, 9940, 10548, 11598]. Chaise [10781]. ChamæPhenix [9782]. Chamber [1687, 3531, 8991]. Champion [1617, 4422, 5901, 7568]. Chance [795, 3166, 3452]. Chandelier [1086]. Chandler [9969]. Change [5780, 9900, 10390]. Changeable [5772]. Changed [7019]. Changes [1079, 855, 6379]. Changing [3326]. Channel [2753, 4209, 12253, 12398]. Channeling [8012, 12505]. Channels [60]. Channelways [8698, 7412]. Chaplin [11569]. Character [8639]. Charcoal [884, 1934, 3928, 10906]. Charger [5804]. Charges [943, 1560]. Charging [11222]. Chariot [7761]. Charity [11506]. Charles [2797, 7655, 2769, 3411, 9969]. Charlotte [8294, 8354]. Chart [2037, 5761, 10499, 10641]. Charts [7314]. Charybdis [2193]. Chase [6738, 9144]. Chasm [3258]. Chastant [658]. Chatel [3374]. Cheap [133, 68, 905, 677, 3675, 3782, 6341, 8238, 12124, 11524]. Cheapen [133]. Cheaper [3350]. Cheapness [5095]. Check [183, 3252, 6090, 5163, 6687, 6643, 7299, 6530, 7202, 6861, 8614, 8617, 8712, 8131, 9313, 8958, 10500, 10946, 10254, 10335]. Check-Hook [8958]. Check-Row [183, 10500]. Checks [9066, 10866, 11257]. Cheese [1370, 1457, 6886, 8791, 9315]. Cheiromeles [1472]. Chelydra [11263]. Chemical [428, 1234, 113, 2419, 2428, 2508, 2747, 4259, 6149, 7149, 6734, 7816, 10390, 7376]. Chemistry [4278, 12384]. Chemists [6892, 10995]. Chenot [3899]. Cherbourg [8242]. Cherries [6639]. Cherry [4820]. Chesley [11986]. Chess [4197]. Chest [5258, 8364, 8396, 11855]. Chevre [5900]. Chevreul [8135, 9682, 11630, 9466]. Chewing [5332]. Cheyne [2332]. Chicago [1853, 1580, 1627, 2394, 4055, 4051, 6837, 8714, 9571, 12429]. Chief [10242]. Child [3904, 8572, 8785, 12076]. Children [7469]. Chile [12484]. Chili [10985]. Chilian [5866]. Chime [6741]. Chimera [3855]. Chimney [938, 442, 7065, 7039, 7725, 8739, 10981, 10566, 11145, 12455]. Chimney-Piece [442]. Chimneys [7038, 8806, 11590]. Chimpanzee [5394]. Chimpanzees [5734]. China [4824, 7585, 12386]. Chinch [4697]. Chinch-Bug [4697]. Chinese [593, 541, 1140, 174, 2657, 9707, 10230, 11679, 4872, 752]. Chips [5113]. Chisel [11884]. Chisels [6089, 6004]. Chladni [8086]. Chlamydosaurus [10289]. Chloroform [1703]. Chocolate [2749]. Cholera [4671, 4823, 4460, 5880, 5599, 5601, 6475]. Chopper [2861, 4170, 7111, 8533]. Choppers [11568]. Christmas [9938]. Christopher [9731]. Chromatrope [11262]. Chronogbaph [3279]. Chronograph [2814]. Chronology [358]. Chrysanthemum [4370]. Chuck [2843, 3918, 5096, 11796]. Chucking [7838]. Churn [1056, 4523, 5851, 7118, 7439, 8155, 8210, 8646, 10128, 9482, 11238, 11154]. Chute [10870]. Cider [2569]. Cigar [4770, 6364]. Cigar-Cutter [6364]. Cigarette [4750, 9017]. Cigars [1011]. Ciliated [7119]. Cincinnati [1508, 5044, 973, 4952, 6262]. Cinder [4579]. Circle [7314]. Circuit [34, 3690, 4414, 10759]. Circular [1509, 2987, 3325, 3111, 4284, 7805, 10472, 10783]. Circulating [6837]. Cistern [1160, 10098]. Cisterns [1058, 6261, 6453, 11419, 5502]. Cities [897, 1963, 1937, 1402, 1494, 5089, 5122, 5168, 5220, 5243, 5382, 8152, 10830, 10487]. Citric [1353]. City [224, 986, 1582, 5208, 5147, 8399, 8817, 9388, 10021, 10466, 11326, 10852, 11048, 11678, 8291, 61, 452, 2282, 1630, 2089, 3597, 2982, 3043, 2746, 3498, 3369, 4210, 4867, 4994, 5167, 8650, 8111, 7799, 7747, 8073, 10308, 10343, 12598, 11618, 11653, 11736, 12068]. Civil [478, 4129]. Civilian [9224]. Clad [8767]. Clads [7178]. Claim [7004]. Claims [2999]. Clam [2857]. Clamond [3254]. Clamp [1948, 1468, 6006, 7172, 7832, 8467, 7586, 9202, 9041, 10312, 10823, 11080, 10522, 11305, 11421, 11628]. Clamping [1331]. Clamps [4921]. Clapp [7509, 7612]. Clark [9593, 10981]. Clasp [1750, 10383]. Class [1438, 11536]. Claw [5921, 5320, 2927]. Cleaner [266, 961, 1270, 2879, 5244, 5312, 5262, 5047, 5246, 6435, 7392, 7171, 6533, 8212, 8900, 9459, 11332, 11414, 11607, 11541, 12365, 11672, 11878]. Cleaning [2702, 6440, 9123]. Cleanser [7270]. Clearer [7835]. Clearing [9095]. Cleat [11103, 10693]. Cleopatra [526]. Cleveland [7953, 7720]. Clevis [1987, 4581, 9256, 12283, 12031]. Climatic [6180]. Climber [12451]. Climbing [457, 763]. Clinching [1128]. Clinical [12070]. Clip [6642, 8618, 11398]. Clock [1086, 513, 2223, 3894, 5663, 6890, 8473, 8367, 7894, 8509, 9317, 9805, 11766]. Clock-Dial [2223]. Clocks [990, 5759, 10643]. Clod [7643]. Clogging [9851]. Close [11472]. Closer [1190, 4414, 10285, 10759]. Closet [1847, 3276, 5271, 7915, 10689]. Closets [1240]. Closing [8393, 10047, 11765]. Cloth [2151, 4551, 4777, 9776, 10470, 11978]. Cloth-Cutting [2151]. Clothes [95, 1287, 4139, 5341, 5807, 7593, 7986, 8868, 9823, 9230, 11082, 10979, 10714]. Clothing [1368]. Cloud [6933]. Clouds [744, 1910, 5074]. Club [7683, 10778, 11602]. Clubs [11290]. Clutch [1356, 3190, 6183, 6254, 9040, 11425, 10805, 11681, 12480]. Clutching [11176]. Clyde [8037]. Co [5536, 5115, 7287, 11533, 4668, 4140, 10529]. Co-Operation [5115]. Co. [5732, 6401, 10150]. Coal [578, 1302, 1355, 2506, 1436, 3476, 2730, 2855, 4597, 4149, 5782, 8736, 8686, 7808, 7936, 7906, 7907, 8808, 9771, 10870, 12364, 1662]. Coaling [4208, 9719]. Coast [2398, 3006, 3045, 5645, 7362, 6486, 7584, 8085, 8667, 8780, 9389, 888]. Coasting [4498]. Coasts [3451]. Cobs [8111]. Cocaine [9448]. Cock [1182, 1884, 3374, 4578, 8359, 12114]. Cockatoo [919]. Cocking [9824, 2221]. Cocks [6576]. Coco [12591]. Cocoanut [12160]. Cocoon [12623]. Cod [957, 10544]. Codfish [2606]. Coendou [7097]. Coffee [319, 1451, 6059, 6060, 9219, 9038, 1642]. Cog [6222]. Cohesion [10198]. Cohoes [400]. Coiling [12624]. Coin [3034, 3367, 4150, 5733, 9591, 11596]. Col [7829]. Cold [1371, 110, 4114, 6089, 6004, 5384, 6388]. Colladon [5929]. Collapse [11849, 2451]. Collapsible [11798]. Collar [184, 5050, 6259, 7201, 8132, 8930, 10130, 9015, 10603, 10200]. Collars [2039, 9992, 8992]. Collect [1759]. Collecting [5210]. College [6961, 6185, 7933, 4089, 8817]. Colleges [1134]. Collet [8895]. Colliery [1036]. Collision [684, 3898, 8479, 9185, 9416, 11125, 11378, 10494, 10918, 8566]. Collisions [3085]. Cologne [10809, 10256]. Colonial [4443]. Colony [4262, 5515]. Color [1351, 955, 375, 1944, 3212, 7815]. Colorado [51, 9771, 11449, 1252]. Coloration [7908]. Colored [10919]. Coloring [2529, 3674, 4745, 9540, 4125]. Colors [503, 1618, 2046, 2250, 6579, 6975]. Colossal [7052, 8482, 8726]. Colossus [9085, 8357]. Colter [3245, 5781]. Columbia [354, 4031, 6288, 6404, 11923, 12462]. Columbus [9731]. Column [6340, 10686, 12116]. Columns [2057, 4870]. Comb [9459, 11336]. Comber [7151]. Combette [2086]. Combination [411, 1590, 2129, 3190, 2840, 4504, 4877, 6066, 5639, 5899, 5684, 5029, 6115, 8189, 9338, 8961, 9339, 9424, 9146, 9850, 9730, 10314, 10694, 10342, 12392]. Combinations [602]. Combined [965, 612, 892, 460, 2214, 1441, 2339, 3246, 2839, 3877, 4827, 4898, 4446, 4582, 4879, 5619, 6093, 5441, 5364, 5736, 7368, 6554, 6364, 6956, 6410, 6307, 6669, 6886, 7400, 6407, 7014, 7682, 7725, 8532, 7939, 7782, 8009, 7565, 7621, 7643, 9459, 9451, 9017, 8930, 9484, 10019, 9149, 9012, 9905, 10129, 9285, 9222, 8987, 9224, 11172, 10667, 10380, 11335, 10417, 11364, 10695, 10686, 12137]. Combining [912]. Combs [11052]. Combustible [2958]. Combustion [884, 4744, 6851, 10288]. Come [2911]. Comet [1722, 2418, 3064, 2725, 3298, 3297, 3304, 9585, 263, 2237, 2439, 2561, 3723, 4258, 4574, 5514, 5186, 5486, 5550, 6683, 9572]. Comets [2530, 6251, 6547, 7723, 3241, 7669]. Comfort [9289]. Commander [2332]. Commented [2174]. Commerce [907, 7610]. Commercial [2260, 1382, 8837]. Commission [1909]. Commissioner [5044, 10732, 1025, 2362, 2220, 3408, 5723, 6271, 8949, 11587, 11534, 12019]. Commissionership [254, 6755]. Committee [5463]. Common [13, 1920, 3270, 7560, 10227, 12521]. Communication [977, 323, 2430, 2847, 3014, 3037, 7183]. Communications [25, 52, 1141, 1196, 1253, 1340, 1372, 78, 125, 149, 167, 198, 304, 333, 362, 391, 444, 516, 543, 623, 672, 727, 785, 836, 896, 925, 1026, 1115, 1282, 1423, 1452, 1475, 1498, 1525, 1549, 1620, 1715, 1761, 1786, 1895, 1953, 1995, 2025, 2159, 2251, 2276, 2300, 2349, 2574, 2595, 2621, 2648, 2784, 2827, 2953, 3335, 3490, 3541, 3705, 3728, 3797, 3884, 3934, 3996, 4108, 4156, 4177, 4267, 4294, 4787, 4812, 4907, 6945]. Community [893]. Compact [7543]. Company [227, 4850, 6200, 8441, 11007, 5771, 6805, 10661]. Comparative [1510]. Compared [3673]. Comparison [6930]. Compartments [5578]. Compass [2854, 3460, 9965, 11871, 11965]. Competition [8291, 1408, 3022]. Complaints [11946]. Completed [3916]. Completion [4796, 4344, 6201, 8677]. Composing [9982]. Composite [5952, 5797, 6818]. Composition [9521]. Compound [4302, 4901, 6036, 6014, 6365, 7391, 6736, 7727, 9249, 8994, 11426, 11836, 3699]. Compounding [11914]. Compounds [2440]. Compressed [578, 1153, 1435, 2287, 2753, 3760, 6496, 6461, 8053, 9316, 9800, 9727, 10863, 10229]. Compressed-Air [2753]. Compressing [843]. Compression [2288, 4578]. Compressor [698, 4706, 11150]. Comptoir [3556]. Concentrating [1825]. Concentrator [8231]. Concerning [5018, 5723, 6299]. Concert [7244]. Conch [782]. Concord [737]. Concrete [716, 6737, 7762, 7653]. Condensation [8066]. Condensed [2509, 7608, 3504]. Condenser [4471, 4924, 8082, 9010, 12569, 11609]. Condensing [7967, 7654, 12287]. Condition [75, 1463, 7667, 9142, 12315, 9266]. Conditions [5354]. Condor [2226, 12169]. Conducting [101, 10736]. Conductivity [9547]. Conductor [7517, 11649, 12557]. Conductors [4794, 5118, 5753, 6559, 8953, 9023, 11544]. Conduit [6093, 8206, 8953, 9023]. Conduits [3985, 8757]. Coney [7052]. Confectioners [9993]. Conference [12252, 12275, 11738]. Confiscation [2900]. Congress [10592, 12208, 500, 10, 2015, 4232, 6029, 5568, 4940, 5017, 5541, 5064, 7535, 9928, 10464, 12372]. Congressional [2877, 5138]. Conical [6042]. Conn [11895]. Connected [6782]. Connecticut [1351, 5305, 7139, 11063, 11922]. Connecting [1779]. Connection [1938, 5287, 8988, 12568, 11594]. Connections [2838]. Conqueror [4469]. Conscientious [821]. Conservatories [6841]. Consideration [3982]. Considered [4464]. Consolidated [8441, 11007]. Constant [98, 4060]. Constantinople [9234]. Constitution [9604]. Constructing [5468, 9396]. Construction [1634, 6342, 7038, 8282, 7762, 7649, 9852, 9795, 10899, 12581, 11640]. Constructive [6920]. Consumer [12164]. Consuming [7197, 7367, 7758]. Consumption [2559, 9477]. Contagious [2559, 5472]. Containing [4191]. Continental [11956]. Continuous [97, 1969, 3506]. Continuous-Slide [1969]. Contraction [5820, 6088, 12069]. Contractor [3676, 7198]. Contracts [12091]. Contrast [7815]. Contribution [5440]. Contributions [9]. Control [605]. Controller [1316]. Controlling [71, 8683]. Controversy [1986, 8029]. Conveniences [8396]. Convenient [718, 4338, 7150, 8464, 9506, 9043, 9590, 9847]. Convention [3500, 4489, 5044, 10618, 1558, 10038, 10820, 11411, 10843]. Converse [630]. Convert [2927]. Convertible [2014, 7588, 8819, 8760, 10735]. Converting [4915, 12587]. Conveyer [8446]. Conveyers [12256]. Conveying [6406]. Cook [12265]. Cooke [9887]. Cooked [3084]. Cooker [4944, 7724, 10019, 7734]. Cooking [2013, 2963, 4923, 6310, 6368, 8842, 9365, 10248]. Cool [8558]. Cooled [9490]. Cooler [1942, 1707, 3246, 5415, 5321, 6307, 8077]. Cooling [1585, 6136, 5468, 10593, 11423, 12475]. Coop [6361, 7540, 12607]. Cooper [3868, 7175]. Cooperage [9903]. Copper [2510, 4937, 12179]. Copperhead [10744]. Copter [330]. Copy [7085, 10600]. Copying [2472, 7674, 10721]. Copyright [3321, 6084, 5410, 10684, 10662, 10464, 10465]. Copyrights [390, 471]. Coral [863, 3013]. Cord [2178, 2985, 5021, 6476, 9567, 10439]. Cords [10339]. Corea [11086]. Cores [3968]. Corinth [6133]. Cork [2480, 3052, 7204]. Corliss [10531, 10530]. Corn [353, 183, 995, 1701, 3221, 5069, 6090, 5569, 6977, 8614, 7542, 7729, 8111, 9252, 8872, 9704, 11416]. Cornelius [3377]. Cornell [6961, 7933, 6852, 8496]. Corners [9875]. Cornets [6091]. Cornice [6550]. Cornstalk [5617]. Corporation [8076]. Corps [5974]. Corpulence [3372]. Corrected [4287]. Correcting [6365]. Correspondence [897, 871, 568, 26, 104, 1167, 1309, 127, 169, 218, 276, 306, 392, 419, 472, 495, 517, 624, 701, 786, 978, 1067, 1090, 1117, 1395, 1426, 2324, 2350, 2373, 2389, 2411, 2461, 2489, 2549, 1526, 1550, 1599, 1621, 1643, 1670, 1739, 1762, 1814, 1864, 1897, 1926, 1977, 2027, 2048, 2138, 2227, 2254, 2302, 2577, 2649, 3469, 3491, 3588, 3612, 3633, 3684, 2718, 2739, 2763, 2892, 2915, 2933, 2973, 2993, 3134, 3156, 3205, 3234, 3262, 3336, 3400, 3423, 3730, 3753, 4375, 4402, 4428, 4453, 4481, 4509, 4535, 4561, 4593, 4643, 4663]. Correspondence [4688, 4737, 4762, 4813, 4859, 4885, 3772, 3798, 3817, 3838, 3861, 3885, 3974, 3997, 4018, 4067, 4090, 4157, 4178, 4202, 4224, 4249, 4268, 4295, 4322, 4350, 4908, 4931, 5508, 5530, 5558, 5584, 5606, 5628, 5651, 5696, 5717, 5741, 5790, 5815, 5839, 5860, 5887, 5910, 5957, 6022, 6049, 6078, 6128, 6155, 4957, 4986, 5010, 5036, 5057, 5083, 5131, 5154, 5177, 5202, 5226, 5276, 5301, 5325, 5349, 5375, 5401, 5425, 5451, 5479, 6191, 6216, 7023, 7047, 7072, 7101, 7157, 7184, 7212, 7281, 7331, 7354, 7381, 6293, 6321, 6349, 6419, 6447, 6490, 6539]. Correspondence [6566, 6608, 6629, 6655, 6701, 6725, 6748, 6798, 6845, 6899, 6946, 6966, 6997, 7406, 7456, 8044, 8123, 8146, 8172, 8200, 8248, 8274, 8346, 8410, 8435, 8489, 8547, 8577, 8604, 8631, 8660, 8691, 8720, 7479, 7503, 7529, 7551, 7577, 7602, 7631, 7688, 7714, 7740, 7768, 7820, 7850, 7878, 7900, 7925, 7947, 7973, 7997, 8021, 8751, 8774, 9471, 9526, 9553, 9580, 9600, 9622, 9669, 9688, 9788, 9813, 9836, 9859, 9892, 9919, 9947, 10005, 10033, 10064, 8796, 8828, 8856, 8885, 8917, 8970, 9000, 9079, 9134, 9159, 9190, 9213, 9243, 9270, 9301, 9353, 9383]. Correspondence [9410, 9439, 10146, 10190, 10989, 11021, 11043, 11093, 11118, 11140, 11183, 11205, 11247, 11294, 11372, 10213, 10268, 10298, 10328, 10372, 10398, 10481, 10538, 10610, 10630, 10655, 10727, 10772, 10792, 10814, 10837, 10858, 10885, 10932, 10961, 11528, 12549, 11433, 11456, 12331, 12353, 12402, 12423, 12507, 12571, 12614, 11510, 11556, 11580, 11611, 11659, 11683, 11709, 11755, 11822, 11843, 11888, 11916, 11939, 12012, 12060, 12102, 12127, 12150, 12198, 12219, 12241, 12267]. Correspondents [568, 3827]. Corrugated [2928, 4651]. Corsets [9477]. Corsica [8085]. Corundum [1099, 633]. Corvette [10388]. Corwin [12633]. Corymbosum [4370]. Cost [1483, 1677, 2792, 5095]. Costly [4696, 6545]. Costs [3917]. Cot [10622]. Cotta [4193]. Cottage [3675]. Cotter [6426]. Cotton [576, 577, 665, 879, 734, 1443, 2107, 2558, 2190, 1615, 2319, 2311, 2262, 2861, 3194, 3524, 3023, 4566, 4569, 4170, 4086, 4147, 4339, 6163, 7111, 7053, 6425, 8533, 8283, 7757, 9750, 9637, 9257, 9725, 9882, 9849, 9010, 12284, 11607, 11672, 6734, 881, 2609, 12032]. Could [793]. Council [453]. Count [811]. Counter [1588, 5965, 7520, 10713, 10471, 11352]. Counterbalance [4415]. Counterparts [3552]. Counting [7438]. Counting-House [7438]. Country [470, 4518, 8556]. Countrymen [11897]. County [1099, 3575, 7966]. Coupled [6238]. Coupler [1359, 4779, 6941, 6234, 7115, 6253, 7730, 9726, 10125, 9011, 11053, 11443]. Couplers [3693, 7859, 9676]. Coupling [1105, 208, 2400, 2176, 3696, 2585, 2815, 2634, 4728, 3788, 4216, 3943, 4447, 5190, 5704, 5752, 6170, 5000, 5516, 5518, 4946, 5117, 5824, 6017, 5829, 6302, 6552, 6577, 7364, 6276, 6360, 6688, 8618, 7518, 7539, 7754, 8183, 8259, 8260, 8391, 8560, 8562, 7648, 8425, 8529, 7672, 8758, 8759, 9567, 8899, 8932, 8957, 9145, 9174, 9483, 9563, 9780, 9848, 9995, 10020, 9564, 9260, 9225, 11051, 10439, 10181, 10250, 10319, 10495, 10523, 10784, 10969, 11034, 11081, 11217, 11281, 11310, 11417, 10599, 10666, 11362, 10763, 11896, 11526, 11593]. Coupling [11670, 11882, 11885, 12003, 12584, 11775, 12006, 11495, 11446, 11739, 12073, 12301, 12604]. Couplings [3166, 10365]. Coupon [11695]. Courier [10179]. Course [6480, 8817, 11923]. Court [1461, 1905, 1407, 34, 1403, 9419, 9818, 11947]. Courts [2548, 2572, 11220, 4799]. Couvreux [6171]. Cover [381, 2670, 5414, 6310, 6184, 6811, 8281, 8676, 8081, 7471, 7962, 9698, 9873, 10127, 11772, 12455, 12075]. Covered [7275]. Covering [1080, 101, 5338, 8814]. Coverings [11174, 10736]. Covers [3648, 5826, 11080]. Cow [5192, 6343, 9154]. Cowen [7992]. Cowl [7065, 8739, 10566]. Cows [7441]. Cox [6556]. Crab [48, 619, 5690, 387, 7448, 3511]. Crab-Apple [3511]. Crabs [644]. Cracker [8676]. Cradle [1734, 7246]. Craft [1507]. Craig [5121]. Cramp [10048]. Crane [1989, 1972, 3431, 3368, 4317, 3813, 6160, 5782, 7198, 6754, 6817, 7895, 8312, 9694, 9276, 9705, 11364, 10447, 10382, 12479, 11929, 12057, 12350]. Cranes [6918]. Crank [1361, 5469, 5431, 6405, 10040, 12049]. Cranks [6980]. Crate [237, 7642]. Crawfordsville [440]. Craze [2657]. Cream [4239, 12001]. Creek [5112]. Creeper [10735]. Creepers [6558]. Creeping [4528]. Creil [8216]. Cremating [2031]. Cremation [924, 268, 1384]. Crematory [10781]. Crests [1973]. Cretaceous [1252]. Crevasses [2926]. Crib [5871, 7619]. Cribbing [778]. Crickets [11427, 10565]. Crime [260]. Criminal [3020, 5945]. Criminals [12228, 11924]. Crimping [9453]. Crinoids [440]. Criticisms [4274]. Crocodile [3097]. Crocodiles [3569]. Croix [8678]. Crop [2260, 4462]. Crops [8011]. Crosby [10468]. Cross [1295, 4601, 8571, 11083]. Cross-Cut [11083]. Crosshead [8931]. Crossing [2631, 6838, 9560, 10621, 8291]. Crossings [831, 109, 6645]. Croton [7041]. Crow [5991]. Crowded [12450]. Crown [9237, 9266]. Crowned [4421, 9694]. Crude [5757, 11707, 12587]. Cruiser [4565, 6837, 6574, 8417, 8314, 8861, 9122, 9759, 9571, 9114, 9194, 11035, 10953, 12189, 12565, 12156, 12502]. Cruisers [4694, 4997, 8639]. Cruises [10388]. Crusher [1083, 2010, 3221, 5593, 7682, 7643, 9018, 136]. Crushing [11976]. Crustaceans [1546]. Crutch [4751, 5075, 11444]. Crystal [8680]. Crystalline [2923]. Crystallization [910, 586]. Crystallizing [11881]. Crystals [4196, 3307]. Cuban [10698]. Cube [1339]. Cubs [8512]. Cucumber [3463, 2672]. Cud [5332]. Cue [5124, 7652]. Cuff [5296, 6258, 7468, 10200, 10950]. Culinary [2886]. Cultivation [954, 1042]. Cultivator [2167, 4626, 4397, 5681, 7113, 8009, 8669, 7729, 9340, 9909, 8765, 9508, 9826, 9314, 11415, 12284, 11700, 11975, 12537, 12539]. Cultivators [100, 7111]. Cultural [178]. Culture [402, 493, 1434, 1805, 2055, 2726, 3107, 3329, 4699, 5544, 8556, 2685]. Culturist [1558]. Cunard [6484, 7510]. Cup [2946, 6035, 7080, 7645, 7614, 9316, 9288, 10827, 6877, 9634]. Cupola [3764, 5930, 5686]. Curbing [564]. Curbstone [6093]. Cure [4142, 5774]. Cures [7885]. Curio [4442]. Curiosities [3784, 4231, 4571, 5331]. Curiosity [11781]. Curious [166, 1204, 784, 735, 1922, 1594, 1856, 2803, 4555, 5003, 5500, 6931, 7316, 8647, 7919, 8768, 11159, 1889]. Curiously [2668, 12326].

Current [4892, 4027, 4060, 5637, 11328, 11831, 9640]. Currents [12450, 12258, 7120]. Curry [11336]. Currycombs [1439]. Curtain [2178, 6065, 8109, 9042, 9176, 11102, 10847, 12456, 11777]. Curtain-Cord [2178]. Curtis [10554]. Curve [6009]. Curves [8679]. Cushion [4337]. Cushioned [2320, 6883, 6408]. Cushioner [9115]. Cuspidor [6123]. Cut [2127, 2222, 3007, 2859, 4601, 4119, 4196, 4168, 5292, 5832, 6261, 6453, 6428, 7677, 7679, 7959, 9760, 9749, 10380, 11065, 10874, 11083, 11740, 12474]. Cut-Off [2127, 3007, 2859, 4168, 5292, 5832, 6261, 6453, 7959, 10380, 11065, 10874, 11740, 2222, 12474]. Cut-Out [7677, 7679]. Cutlery [5801]. Cuts [4418]. Cutter [770, 2319, 2405, 3622, 3454, 4238, 6164, 5617, 6005, 7124, 6364, 6573, 6812, 7339, 7749, 9315, 9252, 9009, 10076, 9828, 9660, 9849, 9993, 10970, 10470, 10868, 11108, 10688, 11548, 11746, 12188, 12367, 12008, 11811]. Cutting [1331, 2151, 1880, 3621, 3595, 3652, 4778, 5439, 5821, 5331, 5750, 6071, 5894, 7144, 7112, 7125, 7341, 6561, 8188, 8362, 7472, 7751, 7522, 7495, 9565, 12026, 12499, 11906]. Cuv. [11600]. Cuvier [10650]. Cycloid [8403]. Cycloidotrope [7229]. Cyclometer [7592]. Cyclones [953, 373]. Cyclorama [8319]. Cyclostyle [6483]. Cylinder [1182, 3327, 6815, 8359, 8653, 9347, 11522, 2243]. Cylinders [6161]. Cylindrical [6853]. Cynipidæ [3438]. Cyprien [481]. Cyprus [1303].

D [3241]. Dabbing [11052]. Dactyloptera [3306]. Daet [3148]. Daft [12045]. Daguerreotypy [9306]. Dairy [3244, 9955]. Dakota [8834, 11562]. Dam [5465, 11135, 12416, 12590, 3914, 7534, 11849]. Damage [12180]. Damp [5542, 7699, 7906]. Damper [508, 2340, 8038, 8806]. Dampers [5264]. Dana [9126]. Dandolo [3268]. Danenhower [3294]. Danger [8474, 11717]. Dangerous [1872, 3737, 11925]. Dangers [1772]. Daniel [9, 8979, 10386]. Daniell [776, 9199]. Darning [10951]. D’arrest [4258]. Darwin [2769]. Dash [11154]. Date [1632]. Davey [944]. Davis [8979, 7534, 7443]. Day [4902, 12409]. Daylight [1464]. Days [8863]. Dead [924, 986, 4572]. Deadly [2088]. Deaf [176, 4036, 1113]. Deafness [1771, 1654]. Dear [3646]. Death [3735, 3981, 6088, 7829, 11190, 11226, 11379, 11440, 5833, 6857]. Debris [740, 206]. Debt [2358]. December [3453, 4624, 5895, 7192, 9845, 11148, 12385, 7165, 7414]. Deception [1679, 1777, 3141, 5711]. Decided [4797, 10241]. Decision [34, 1403, 2417, 3408, 5962, 7242, 8194, 8267, 8309, 8195, 9843, 8815, 10486, 11947, 7828]. Decisions [725, 756, 775, 858, 917, 976, 1025, 1060, 1114, 1139, 1328, 195, 360, 390, 471, 494, 594, 993, 1277, 1304, 2548, 2572, 1389, 1416, 1664, 1894, 1924, 1946, 1968, 2403, 1446, 1516, 1614, 1635, 1685, 1753, 2244, 1534, 2220, 1905, 2619, 2674, 2858, 2965, 3175, 3198, 3223, 3251, 3303, 3574, 2711, 2754, 3719, 3856, 3988, 4057, 4085, 4520, 4753, 4011, 4035, 4942, 4967, 5234, 5646, 5668, 5737, 5830, 5906, 6012, 6070, 5723, 5975, 6189, 6560, 6720, 7323, 7465, 7561, 8063, 8087, 8450, 8499, 8597]. Decisions [8652, 7451, 7845, 8902, 8984, 9182, 9636, 9870, 9901, 10733, 11220, 11525, 11725, 11853, 12009, 8369, 8716]. Deck [9420]. Declared [10436]. Decline [3640, 7003, 6831].

Decomposition [10422]. Decorated [1619]. Decorating [2018, 387]. Decoration [8964]. Decorations [2426, 7595]. Deep [426, 1021, 501, 1504, 4276, 4418, 3792, 5340, 7777, 8697, 11899]. Deepest [7665]. Deer [3173, 4343, 6188]. Defeat [5043, 9333]. Defective [2959, 10180]. Defense [1156, 2398, 3006, 3045, 4997, 6545, 8667, 9196, 8863, 11099, 11968]. Defenses [842, 8780, 9389]. Deficient [4525]. Deflector [4579]. Deformities [5756]. Degree [3376]. Delaware [11489]. Delay [5964, 11641]. Delineator [12211]. Delivery [4749, 11050]. Delta [6169]. Delusions [1746]. Demagnetization [8178]. Demagnetizing [2570]. Demand [1217, 5185]. Demanded [3439]. Demoiselle [4317]. Demonstrating [1657]. Denain [7257]. Dense [5313]. Dental [1262, 9173]. Dentistry [1772, 12585]. Dentistry-Filling [12585]. Denver [3714, 6009]. Deodorizer [4259]. Department [10021, 6097, 7932, 9846]. Departure [12249]. Dephosphorization [295]. Deposit [4186]. Depositing [4382, 6905]. Depot [10343]. Deprez [8451]. Depths [5354, 7291]. Derailing [8805]. Dermal [8730]. Dernell [12283]. Derrick [636, 8706]. Derricks [573]. D’Escompte [3556]. Described [3900]. Description [1939]. Desert [2812, 8744, 205]. Design [5239, 6642, 7968, 7498, 8253, 9296, 9446, 11367]. Designed [6340]. Designing [12492, 11535]. Designs [910, 7545, 8981, 11881, 11978]. Desk [5259, 6598, 8588, 7466, 10397, 11056, 10897, 11595]. Desnos [3411]. Destroy [2755]. Destroyed [6120]. Destroyer [15, 7702, 5811]. Destroyers [3438, 4061]. Destroying [11949]. Destruction [935, 10437, 11849]. Destructive [2885]. Destructor [1431]. Detachable [3648, 6410, 6785, 6670, 7652, 8709, 9068]. Detacher [5638, 7809, 7839, 10251]. Detaching [2318, 9117, 11304, 11542]. Detecter [6527, 6979]. Detection [4825]. Detective [2902]. Detector [11764]. Determining [943, 3376, 9024]. Detonating [9637]. Devastation [10565]. Develop [4213]. Developer [12133]. Developing [11973]. Development [8727, 2539]. Developments [904, 3599]. Device [299, 347, 2217, 1826, 1414, 2845, 4756, 4310, 5993, 4921, 5290, 5492, 5826, 5825, 5360, 5616, 5658, 5805, 7247, 7318, 8683, 8701, 7681, 7562, 7441, 9506, 8931, 10074, 9117, 9253, 9750, 9536, 9631, 9168, 9089, 9959, 8844, 9612, 9014, 9363, 9633, 10098, 9776, 10047, 9565, 9453, 9259, 10206, 11082, 10568, 10291, 10364, 10309, 11077, 10951, 10199, 11130, 10548, 11363, 10318, 11176, 11360, 10316, 11243, 11834, 12023, 12026, 12434, 11472, 11542, 12457, 11571, 11603, 12435, 11948, 11765, 11696, 11973]. Devon [9434]. Dexterity [3668]. Diadems [4974]. Diagnosis [3878]. Diagonal [1963, 10260]. Dial [2223, 3647]. Diamond [1231, 4830, 4030, 5419, 4916, 7095, 8424, 11391, 7497]. Diamonds [36, 5728]. Diaphragm [3461]. Diaphragms [11508]. Diavolo [1951]. Dice [11361]. Dicentras [7303]. Dictionary [5564]. Dieulafoy [11729]. Difficulties [9034]. Digest [2006]. Digestion [4278, 4821]. Digger [72, 1682, 3946, 9909, 9606, 9398, 11601]. Diggers [217]. Digging [564].

Dikes [11627, 12000]. Dilatancy [7791]. Diligence [812]. Diminishing [7980]. Dining [12237]. Dinner [2290, 11799]. Dinoceras [9999]. Dinosaurs [5904]. Dioxide [10422]. Diphtheria [1155, 12046]. Diplograph [8675]. Dipper [687]. Dipterous [4700]. Direct [3374, 3654, 8526, 8756, 10941, 12586]. Directions [7512]. Directly [10135]. Dirt [9988, 6429]. Disaster [11124, 11829, 504, 554, 7778, 11887]. Disastrous [1179]. Discharge [744, 8613]. Discharger [7419]. Discharges [1032]. Discount [9877]. Discovered [10366]. Discoveries [2082, 1633, 3131, 5090, 5089, 5122, 5168, 5220, 5243, 12381, 12524]. Discovery [766, 397, 1938, 3892, 7723, 10438]. Disease [914, 1852, 4625, 8668, 1796, 6429]. Diseases [971, 1286, 3197, 5472]. Dish [772, 12389]. Disinfectant [2008]. Disinfectants [10059]. Disinfecter [6433, 7567]. Disinfecting [8742, 9885]. Disinfection [10059, 6429]. Disintegrating [2862, 4314]. Disintegration [8686]. Disintegrator [5593]. Disk [1566, 4631]. Dismantling [11792]. Dispatch [7349]. Displacement [97]. Display [9701]. Displaying [10203]. Disposal [12110, 12034]. Disposition [9745]. Dissertation [11]. Dissolving [10798]. Distance [1493, 3783, 8179, 8140, 9343]. Distances [5920, 6043]. Distillation [9913, 9718]. Distribute [11471]. Distributer [2272, 3570, 6411, 9067, 11565, 12433, 12538]. Distributing [9982]. Distribution [2282, 8036, 11717]. District [6404]. Ditch [9231]. Ditcher [11907, 12611]. Ditches [5465]. Ditching [7193, 11720]. Diverging [11225]. Dividers [2479, 4503, 6335, 11550]. Dividing [7618]. Diving [2198]. Division [1604]. Dix [12259]. Dnieper [6207]. Do [736, 3187, 6381]. Dobbins [1224]. Dock [2104, 4382, 7147, 6905, 12397, 11666]. Docks [3938]. Doctor [6430, 11934, 6227]. Doctors [5868]. Doctrine [5241]. Doe [2820]. Does [3915, 4917, 10819, 9399]. Dog [566, 3128, 5195, 12279, 11474]. Dogali [9681]. Dogs [8909, 11235, 10506]. Doings [3346]. Dollar [2546, 7030, 10639]. Dollars [3760]. Dolman [5264]. Dolphin [6574]. Dom [5853]. Domain [8004]. Domes [12238]. Domestic [7206, 7038]. Donald [7147]. Done [3187, 6809, 9592]. Door [92, 1683, 2565, 2663, 3694, 4653, 5595, 5288, 5266, 5659, 5163, 6530, 7202, 6618, 7320, 6304, 8561, 7804, 8062, 9311, 9935, 9313, 9775, 10054, 10017, 9201, 10761, 10285, 10341, 11306, 10925, 10875, 11218, 12235, 11800, 12282, 12232]. Doors [1020, 2224, 1831, 7253, 8080, 7841, 9485, 9233, 11538]. Dorcas} [10366]. Dose [1681]. Dose-Measuring [1681]. Double [806, 14, 1020, 1800, 2111, 2514, 2972, 4901, 4441, 6008, 5636, 5969, 5388, 6694, 6234, 6386, 6383, 6409, 6575, 8010, 8317, 7648, 8991, 9179, 9420, 9752, 11902, 11809, 8291]. Double-Acting [14, 2111, 6386, 9179]. Double-Deck [9420]. Double-Link [7648]. Double-Pointed [6694]. Double-Reciprocating [11902]. Doubler [8567]. Doubling [7781]. Doubt [8524]. Dough [12342]. Dove [3417]. Dovetailing [7912]. Dowd [946]. Dowling [6227]. Down [526]. Dowson [3876]. Dr [9295]. Dr. [9, 1314, 708, 2468, 2082, 1493, 3600, 5090, 5930, 5268, 5089, 5122, 5168, 5220, 5243, 6391, 12555]. Draft [10313, 10894]. Drag [6092, 4922, 8182, 9035, 12298]. Dragon [1833]. Drain [5898, 8570, 11390]. Drainage [9842, 11575, 1796]. Drainage-Trap [11575]. Drains [3621]. Drama [11223]. Dramatic [8622]. Draper [9651, 9706, 9, 3478, 2836]. Draught [98, 164, 611, 2218, 5524, 5436, 8423, 9574, 10280, 11566, 11832, 12319, 12164]. Draughtsman [376, 12564]. Draw [6454, 9140, 11809]. Drawbaugh [5962]. Drawbridge [6072, 6501, 8310, 7423, 12113, 11875]. Drawer [8617, 8712]. Drawers [5258, 5951, 6184]. Drawing [2173, 3578, 3407, 6649, 8106, 11259]. Dredge [4550, 4230]. Dredger [4949, 6648, 8317, 9064, 9462]. Dredges [7336]. Dredging [4357, 6118, 11761]. Drenching [6833]. Dressing [7915]. Drier [95, 63, 3246, 2774, 6307, 6891, 7593, 7986, 7699, 8008, 8083, 10979, 12389]. Drill [773, 660, 1270, 1886, 3457, 3899, 3918, 4342, 4240, 5459, 4916, 5162, 5023, 8808, 9777, 9046, 10878, 10337, 11128, 12505, 12433]. Drill-Hole [1270]. Drilling [3029, 4948, 5187, 5799, 6412, 7094, 8746, 9366, 9432, 10363, 11085, 11287]. Drills [4126, 7245, 9095]. Drinking [895, 8308]. Drive [61, 12307, 12369, 3962]. Driven [7557, 9388, 9843, 9166, 9819]. Driver [4526, 5780, 8187, 8114, 8079, 10043, 11424, 11904]. Driving [1128, 6030, 5615, 6684, 8626, 10164]. Drop [1412, 3247, 4004, 5048, 10397]. Dropper [7296, 9149]. Dropping [6662]. Drops [8100]. Druggists [4567]. Drugs [2582]. Drum [6280, 11358, 12387, 12261]. Dry [98, 2104, 2695, 3938, 4382, 7147, 7596, 8952, 12343, 12133, 12397, 11666]. Drying [98, 1544, 3531, 4551, 5145, 6002, 6363, 11385, 12476, 11768]. Dubious [3321]. Duc [2604]. Duck [3747]. Ducker [7857]. Ducking [1421]. Ducks [2158]. Due [6547]. Duguesclin [7827]. Duilio [226]. Dujour [4975]. Dump [322, 9929]. Dumping [4488, 5683, 5549, 7117, 8108, 7703, 8933, 9561, 11384, 10738, 12434, 12112, 12452]. Dunning [12236]. Duplex [1040, 5522, 9366]. Duplication [9504]. Dupré [1512]. During [11314]. Dusseldorf [5551]. Dust [2532, 4009, 4579, 4368, 6184, 8647, 7787, 9286, 356]. Dust-Pans [4368]. Duster [9909]. Dustless [9753]. Dutch [9492]. Duty [3164]. Dwarf [10579, 10809]. Dwelling [3100, 12124]. Dwight [9126]. Dyed [3762]. Dyeing [2262, 4623, 7842, 9941, 8821, 12052]. Dyestuffs [2064]. Dynamic [4524, 8965, 2407, 2458]. Dynamite [5169, 5051, 6973, 7002, 8417, 8330, 8861, 9613, 9465, 11035, 11234]. Dynamo [549, 1916, 1870, 2357, 2157, 3690, 2966, 4408, 3734, 4852, 9063, 9044, 11896, 11871, 12192, 12261, 11640, 10135, 8357, 8976, 9020, 10323, 10344, 8726]. Dynamo-Electric [549, 1870, 2157, 2966]. Dynamo-Telegraphy [1916]. Dynamometer [2533, 3392, 5970]. Dynamometric [3602]. Dynamos [10164].

Eads [8836, 1051, 1017, 5869, 9926, 8993]. Eagle [1041, 10184]. Ear [4285, 6280, 8258]. Early [3669, 3805, 6768, 7774, 8557, 8867]. Earrings [12023]. Ears [12023]. Earth [930, 2545, 2838, 2562]. Earthquake [385, 4555, 6134, 5432, 8400, 8128, 8787, 8880]. Easel [376, 11284]. Easily [2105, 7174, 9485]. East [3298, 3650, 11106, 2942]. Eastern [1632, 8769, 10758, 10943]. Easy [2513]. Eater [9008, 10138, 9293, 2863]. Eating [1221, 24]. Eaton [2426]. Eaves [8181]. Eaves-Trough [8181]. Eberhardt [12499]. Eccentrics [6980]. Echidna [10138]. Echteler [9296]. Eclipse [3371, 3576, 2898, 4078, 3846, 4256, 9628, 9263, 11215, 11748]. Ecliptic [3462]. Economic [3512]. Economical [9745, 11313, 11426]. Economizer [165, 4056]. Economy [32, 632, 3024, 2979, 4306]. Eddystone [2957]. Edge [41, 5309, 5020]. Edging [7054]. Edgings [9590]. Edible [5690]. Edición [12430]. Edinburgh [7523]. Edison [4408, 292, 255, 408, 462, 506, 1870, 2357, 3712, 4668, 5722, 8015, 10224, 11315, 10486, 12123, 12324, 12216, 177, 230]. Edition [126, 673, 8865]. Educated [3761]. Education [3527, 3823, 4358, 5591]. Educational [3349]. Edward [9717, 5796, 10081]. Eel [8510]. Effect [683, 1009, 709, 1350, 4821, 6137, 6711, 6481, 8400, 8727, 8388, 7733, 9464, 9899]. Effective [6545]. Effects [2617, 6972, 9637]. Effervescing [1884]. Efficiency [340, 5918, 8585]. Efficient [10075, 11313, 10688]. Egg [637, 237, 581, 5286, 6644, 7035, 7294, 7372, 8284, 9962, 10347]. Eggs [2033, 4680, 4528, 9227, 11571]. Egypt [3691, 4671]. Egyptian [8293, 621, 1259, 3163, 3213, 4473, 6239, 7814]. Ehret [2532]. Eider [2158]. Eiffel [11751, 11840, 12147, 12123, 11481]. Eight [8976, 9020]. Eighteen [1149]. Eighty [1149, 8463]. Eikonogen [12133]. Ejecting [10307]. Elastic [3149, 11999]. Elasticity [5232]. Elecampane [6474]. Electric [187, 549, 943, 32, 632, 988, 1032, 1150, 292, 1048, 1349, 77, 1096, 1327, 551, 372, 646, 425, 662, 688, 882, 431, 879, 904, 162, 745, 2127, 2248, 2223, 2085, 1874, 2041, 1606, 1689, 1650, 1877, 1870, 1935, 2202, 2065, 2470, 2157, 1723, 2560, 1842, 1405, 3295, 3712, 2655, 3720, 2821, 3532, 3643, 3189, 3462, 3690, 2772, 3049, 3413, 2709, 2628, 2756, 2966, 4828, 4726, 4408, 4668, 4522, 4012, 3878, 4191, 4288, 3734, 4632, 4074, 4852, 3939, 4607, 4515, 4241, 4184, 5118, 5771, 5288, 4974, 5848, 5444]. Electric [5754, 5708, 5855, 5027, 5753, 5667, 4992, 5879, 5316, 5393, 5732, 6931, 6893, 6597, 7256, 6781, 6527, 6892, 6711, 6894, 7007, 7107, 6210, 6311, 7267, 7388, 6601, 7317, 8367, 7695, 7861, 8206, 8510, 8192, 8134, 8648, 8569, 9063, 9086, 10057, 9817, 10073, 8906, 9044, 11110, 10710, 10778, 10923, 10638, 11015, 10489, 10758, 10891, 10651, 10996, 10486, 10757, 10759, 11112, 10176, 11157, 10474, 11980, 11717, 12045, 12228, 11858, 12523, 12471, 12521, 12612, 11618, 11563, 11766, 12337, 11997, 12439, 11831, 12229, 12477, 11996, 11463, 11640, 11518, 11639, 12600, 2665, 9640]. Electric [9155, 10204, 2637, 7720, 7120, 492, 11952, 10038, 11411]. Electric-Light [12471]. Electrical [408, 986, 661, 669, 1214, 34, 722, 1512, 2298, 1748, 2053, 3455, 3284, 3310, 3331, 3421, 3584, 3608, 3702, 2824, 2866, 3152, 3178, 3232, 3003, 3162, 4041, 4794, 4151, 5771, 5722, 5796, 5046, 5612, 5878, 5746, 5732, 6389, 6211, 8610, 8841, 9362, 9343, 10134, 9048, 11223, 10904, 11078, 11012, 10361, 10867, 10760, 11923, 11544, 12282, 10700, 1570, 177, 230, 5635, 10820]. Electrician [11224]. Electricity [744, 552, 12, 409, 1825, 2249, 1561, 2425, 1442, 1832, 3397, 3357, 3415, 3715, 3759, 3917, 3867, 5380, 6015, 6919, 6743, 7425, 7960, 8167, 7748, 7608, 8216, 8638, 9531, 8924, 11190, 11226, 10625, 11169, 11181, 11440, 10135, 10473, 2407, 2458, 2341]. Electricus [8510]. Electro [204, 1188, 2510, 1383, 1846, 1940, 1991, 2016, 2150, 2453, 3112, 4524, 4606, 3903, 6119, 6768, 6311, 6690, 7989, 8965, 9559, 9418, 10939, 12192, 10287]. Electro-Assay [2510]. Electro-Brass [1383]. Electro-Dynamic [8965]. Electro-Horticulture [204]. Electro-Magnet [10287]. Electro-Magnetic [3112, 6768]. Electro-Magnetophone [4606]. Electro-Massage [2453]. Electro-Metallurgy [1846, 1940, 1991, 2016, 2150]. Electro-Motor [9559, 9418, 10939]. Electro-Physiological [6311]. Electro-Plating [12192]. Electro-Pneumatic [6119]. Electro-Thermic [1188]. Electrolysis [10451]. Electrotyping [8988]. Elegant [1737]. Elephant [216, 967, 3748, 7052, 7874, 10853]. Elephants [11895]. Elephas [9728]. Elevated [224, 947, 1935, 2287, 3172, 5208, 5927, 5466, 7107, 8340, 9025, 9562, 11181]. Elevating [6765]. Elevation [6917]. Elevations [1294]. Elevator [484, 999, 3477, 4604, 3808, 5189, 6038, 5148, 5370, 6940, 7833, 9939, 11008, 11350, 10160, 12099]. Elevators [1638, 4391, 5989, 5597, 6011, 6336, 8118, 9795, 10944, 10718, 11983, 11750]. Elias [12092, 12473]. Ellipsograph [7524]. Elms [1387]. Embossing [4145]. Embroidering [1563]. Emergency [10222]. Emery [975, 4654, 5613, 8647, 8209, 9449]. Emotion [11146]. Emperor [10591]. Empire [7918]. Employ [1982]. Employed [5800]. Employee [844]. Employer [5800]. Employers [1098]. Employes [3982]. Employment [3390]. Emptier [8056]. Emulsion [10178]. Emulsions [3739]. Enactment [8804]. Enameled [12248]. Encke [6683]. Encouragement [654, 2003]. End [3019, 5521, 6007, 7342, 6915, 7589, 11285, 11854, 3720]. Ended [3213]. Endless [10504]. Ends [9363]. Enemy [4800, 7886]. Energy [13, 8036, 10638]. Enforcement [4525, 10745]. Engine [1040, 638, 1126, 327, 561, 433, 1448, 2503, 2126, 2222, 1775, 1391, 2156, 2695, 3191, 3007, 3196, 2798, 3352, 3071, 2706, 3305, 3327, 2813, 2859, 4901, 3987, 4167, 4023, 3812, 4101, 4168, 4542, 4472, 4195, 3746, 3766, 4570, 4084, 5391, 6146, 5543, 5049, 5337, 5832, 5972, 5028, 4938, 4972, 6014, 5611, 5016, 6602, 6238, 7326, 6837, 6982, 6485, 7015, 6739, 6581, 6817, 6532, 6275, 6305, 6528, 8139, 7967, 7591, 7521, 7789, 8420, 8343, 7731, 8118, 8672, 7654, 8133, 8568, 9249, 9760, 9445, 8994, 9310, 9091, 9369, 9070]. Engine [11364, 10802, 11426, 10530, 10409, 10246, 11011, 10424, 12098, 11999, 11727, 11902, 11476, 12328, 12049, 11606, 12215, 11704, 12287, 8449]. Engineer [883, 3164, 9297, 12178]. Engineering [16, 50, 1211, 1275, 1290, 1323, 1369, 94, 139, 193, 267, 331, 418, 438, 469, 489, 648, 663, 697, 721, 755, 781, 804, 827, 862, 886, 991, 1066, 1165, 1183, 1386, 1589, 1658, 1710, 1752, 1781, 1806, 1854, 1893, 2038, 2135, 2225, 2265, 2294, 2322, 2336, 2371, 2406, 2459, 2516, 2534, 2380, 3279, 2877, 2647, 2662, 3259, 3282, 3308, 3330, 3359, 3380, 3418, 3440, 3465, 3486, 3513, 3537, 3558, 3581, 3583, 3605, 3629, 3657, 3679, 2732, 3700, 3724, 2761, 2777, 2800, 2822, 2864, 2887, 2912, 2929]. Engineering [2991, 3010, 3027, 3075, 3091, 3116, 3155, 3176, 3199, 3227, 3749, 3793, 4424, 4449, 4477, 4505, 4531, 4557, 4586, 4611, 4636, 4659, 4684, 4709, 4733, 4758, 4783, 4808, 4831, 4855, 4881, 4903, 3834, 3857, 3880, 3930, 3950, 3971, 3992, 4014, 4042, 4063, 4104, 4130, 4152, 4173, 4198, 4221, 4244, 4263, 4290, 4318, 4346, 4371, 4398, 4927, 4953, 5580, 5602, 5624, 5647, 5669, 5693, 5714, 5762, 5787, 5812, 5835, 5857, 5882, 5907, 5931, 5954, 5977, 5995, 6018, 6045, 6074, 6100, 6124, 6151, 6172, 4981, 5006, 5031, 5053]. Engineering [5078, 5104, 5126, 5150, 5173, 5198, 5222, 5247, 5272, 5298, 5322, 5344, 5371, 5396, 5421, 5447, 5475, 5503, 5526, 5553, 6213, 6240, 6820, 6842, 6895, 6922, 6942, 6963, 7020, 7044, 7069, 7098, 7127, 7152, 7180, 7208, 7232, 7259, 7278, 7305, 7328, 7351, 7378, 7403, 6264, 6290, 6316, 6346, 6370, 6393, 6414, 6443, 6462, 6487, 6513, 6535, 6563, 6584, 6604, 6626, 6652, 6674, 6698, 6721, 6744, 6770, 6795, 7426, 7452, 8067, 8090, 8120, 8142, 8168, 8197, 8218, 8244, 8269, 8297, 8321, 8344, 8376, 8406, 8431, 8455, 8483]. Engineering [8513, 8541, 8574, 8600, 8628, 8656, 8688, 8717, 8747, 7477, 7500, 7526, 7547, 7572, 7598, 7627, 7657, 7684, 7710, 7738, 7764, 7792, 7817, 7848, 7875, 7897, 7922, 7944, 7970, 7994, 8018, 8041, 9898, 8770, 8792, 9494, 9522, 9548, 9577, 9595, 9618, 9641, 9665, 9683, 9708, 9732, 9762, 9784, 9808, 9832, 9855, 9888, 9914, 9943, 9971, 10000, 10028, 10061, 10084, 10113, 10141, 8824, 8852, 8881, 8913, 8940, 8966, 8996, 9026, 9050, 9074, 9103, 9129, 9156, 9186, 9211, 9238, 9267, 9324, 9348, 9375, 9404, 9435, 9467, 10167, 10185]. Engineering [10834, 10854, 10882, 10909, 10928, 10957, 10986, 11017, 11039, 11066, 11090, 11115, 11137, 11160, 11267, 11291, 11316, 11341, 11368, 11399, 11429, 10209, 10264, 10293, 10325, 10348, 10368, 10392, 10426, 10453, 10477, 10510, 10535, 10557, 10581, 10606, 10626, 10652, 10674, 10701, 10723, 10746, 10768, 10790, 10811, 11145, 11465, 2814]. Engineers [1031, 4129, 478, 1892]. Engines [71, 579, 777, 144, 1405, 1539, 3455, 2682, 2791, 2909, 2979, 4548, 3805, 4805, 3787, 4874, 5469, 5918, 5497, 5563, 5878, 5806, 6249, 6401, 6718, 6939, 7042, 6736, 8060, 8236, 8653, 7759, 7424, 9681, 10057, 9347, 10099, 8786, 9960, 9594, 9656, 10179, 10476, 10529, 11389, 11922, 12158, 12135, 12519]. England [3509, 3458, 906, 3166, 6356, 11315, 11256]. English [1782, 2583, 3847, 10365, 287, 3354, 3683, 3673, 5600, 6573, 7350, 7473, 12253, 12398]. Englishman [1296]. Engraved [4680]. Engraving [179, 2581, 9544]. Engravings [1591]. Enlarging [5146, 5471]. Ensilage [4875]. Enterprise [3526]. Entire [6340]. Entrance [3851, 11472]. Envelope [4827, 6168, 10938, 11548, 11674]. Envelopes [9014]. Ephemera [4902]. Epidemic [397, 680]. Epiornis [8822]. Equalizer [5524, 8423, 10280, 10894, 11566, 11832, 12319]. Equatorial [2927, 10577, 4667]. Equilibrium [8372]. Equitable [5568]. Equity [342]. Era [1632, 3430]. Eraser [12563]. Erection [3965, 10741]. Ericsson [1040, 5218, 11325, 11501, 12431, 11516]. Erie [3938, 6226]. Ernst [8820]. Error [5463]. Errors [6365]. Eruption [1125, 7813]. Eruptions [8598]. Escape [2593, 2644, 2694, 2947, 3483, 4466, 4552, 4439, 4630, 4676, 3874, 3902, 4367, 5103, 5334, 5496, 5660, 6095, 5116, 5141, 5926, 6691, 7008, 7034, 7369, 6181, 7393, 6976, 7194, 6507, 7565, 8733, 7415, 7983, 7646, 10041, 10573, 11153, 11420, 10553, 12007, 12028, 12349]. Escapement [3279, 6506, 11354]. Escritoire [9043]. Escutcheon [9755]. Esmeralda [5866]. Española [12430]. Esplanade [11473]. Establishment [3105, 9593]. Estimating [5920, 6043]. Estimation [1686]. Estrade [8554, 8743]. Etc [523, 756, 195, 390, 471, 594, 1389, 2263, 2457, 2674, 3175, 3223, 3251, 4149, 3988, 4726, 4340, 4804, 4116, 5619, 4921, 5646, 5830, 5573, 5143, 5828, 6091, 6686, 7172, 6363, 6892, 6406, 7168, 7341, 6980, 6336, 7525, 7642, 7807, 7699, 7831, 8111, 8708, 8447, 7442, 9590, 8961, 9701, 9014, 8806, 10048, 9010, 10040, 9698, 10050, 9749, 9993, 9697, 10206, 10803, 11174, 10866, 11423, 10254, 10207, 11257, 10825, 12363, 12137, 11881, 11903, 12248, 12498, 12002, 11794, 11880, 11594, 11590, 11698, 11450, 11013, 3674]. Ethics [4165]. Etna [6249, 7813]. Etruria [6484]. Eugene [8135, 11630]. Eugenie [6312]. Eureka [11391, 8582]. Europe [1974, 8787, 12515, 11465]. Eurypharyna [3792]. Evans [12183]. Evaporating [7419]. Evaporation [3412, 6136]. Evaporator [2819, 4448, 4748, 6363]. Evenings [7189]. Evils [110]. Evolution [1285, 5241, 12108]. Exactness [4649, 4966]. Examination [4772, 3626, 7119, 6975]. Examinations [6878, 6272]. Examiner [9877]. Examining [3871, 10416]. Excavating [4357, 4209]. Excavator [3507, 6171, 5900, 9723]. Excelsior [11031]. Exchange [1653]. Excitement [10819]. Excludes [8077]. Exclusiveness [287]. Execution [10625, 12228]. Exerciser [12181]. Exercises [9021]. Exercising [2845, 3986, 11602]. Exhaust [9392, 12304]. Exhauster [3573, 4631]. Exhaustion [8004]. Exhibit [5771, 5722, 5732, 5796, 5746]. Exhibiting [6555]. Exhibition [180, 696, 2064, 2321, 1561, 2034, 4549, 4443, 6401, 7096, 7487, 9911, 9653, 9662, 10183, 12383, 864, 1260, 2560, 2053, 3345, 2874, 4722, 4566, 4394, 4184, 5722, 5046, 5635, 5878, 5732, 7967, 7523, 9677, 9773, 10057, 9209, 10907, 12418, 11971, 12339, 11608, 11953, 12100, 12125, 12146, 12262, 12286, 12525, 12540, 12561, 12634, 12589, 1334, 752]. Exhibitions [2078, 3479]. Exhibitor [639, 8955, 12002]. Exhibits [154, 3509]. Existed [298]. Expanded [3061]. Expander [10575]. Expanding [7957]. Expansion [4878, 6959, 8653, 7759, 9594, 11704, 2247, 9344, 9291, 1034, 1655]. Expected [6499, 11810]. Expedition [859, 454, 2035, 3294, 3062, 5521, 6032, 10467, 12249]. Expeditions [5515]. Experiment [1314, 744, 2023, 2245, 1759, 1444, 3172, 2908, 8167, 9101, 9395, 9048, 11014, 11012, 11924, 12578, 1511, 10532, 12055]. Experimental [2401, 2733, 11015, 9729, 12276]. Experimentation [11699]. Experiments [1189, 318, 2315, 1985, 4163, 5948, 6237, 6211, 6329, 8557, 9637, 10112, 11190, 11226, 12555, 11913, 12544, 464, 10473, 9774, 8227, 8165, 8050, 9942, 10323, 10344, 10287, 10881, 11089, 11132, 11179, 11203, 11216, 11237, 11264, 11288, 11339, 11395, 11620, 11654, 11452, 11728, 11935, 11957, 12036, 12059, 11578, 1510, 6583, 8066, 8372, 7846, 8191, 8213, 8239, 8086, 8685, 8713, 10390, 12069]. Expert [1536, 3778]. Experts [4799]. Explode [1460]. Exploder [2665].

Exploration [4233, 5523]. Explorers [6109]. Explosion [2104, 2032, 2532, 2401, 2471, 2531, 2382, 2060, 3035, 2837, 2733, 6010, 5948, 4952, 6262, 6510, 6972, 6236, 8294, 8582, 8354, 10023, 10347, 11133, 11955, 11655, 2243, 2423, 2501, 1827, 1966, 1855]. Explosions [685, 1317, 1824, 3002, 5542, 10996, 11563, 12600]. Explosive [2527, 3269, 5779, 8845]. Explosives [1203, 3476, 9881]. Export [126, 673, 9033]. Exportation [252, 975]. Exports [4305]. Exposition [3547, 4051, 5771, 5796, 5746, 11097, 10985, 11473, 12621, 11752, 973, 3714, 4055, 5612, 5971, 7062, 6401, 6595, 6180, 6327, 6355, 6636, 8339, 10876, 12612, 12285, 12441, 11750]. Exposures [9431]. Express [10802]. Expulsion [8454]. Extensible [6836, 7807, 11838]. Extension [825, 2506, 1949, 7203, 7722, 8894, 10245, 11239]. Extensions [10096]. Extermination [3870]. Exterminator [9798, 11383]. Extinct [298, 9999, 12297]. Extinguisher [3028, 3897, 6122, 5827, 5881, 6911, 6209, 7981, 8807, 9722, 10904, 10163, 10711, 11029, 12300, 12114, 12341]. Extinguishing [1703, 3093, 4468, 8848, 11834]. Extra [833]. Extra-Sensitive [833]. Extracting [2128, 3088, 9336, 10503]. Extraction [1298]. Extractor [2297, 2480, 4849, 3809, 4575, 3763, 5574, 7956]. Extracts [5044]. Extraordinary [6010]. Eye [18, 4719, 5208, 8810]. Eye-Glasses [18]. Eyelet [6686]. Eyepiece [12626]. Eyes [1877, 6952, 8388]. Eyestones [2068].

F.R.S [7055]. Fabrication [8680]. Fabrics [1767, 3510, 3343, 2958, 3901]. Fabry [7669]. Face [1852]. Facilitate [9253, 10823, 10206]. Facilities [10865, 5063]. Facing [8064, 8477]. Fact [1889]. Factor [5566]. Factory [2518]. Facts [784, 1370, 6475, 11049]. Faculty [11934]. Faience [1736]. Fail [3502]. Failing [5411]. Failure [2312, 2639, 4213]. Fair [604, 1152, 550, 1101, 1230, 1909, 10967, 974, 856, 3216, 3509, 3458, 4600, 12296]. Faith [4843]. Falcon [1022]. Fall [916, 8537]. Falling [943, 843, 8599]. Falls [6546, 7218, 12067, 12462, 4963]. Family [1547, 12395]. Famous [4139]. Fan [1065, 940, 4631, 5342, 6500, 7256, 6819, 10013, 10601, 11797]. Fan-Tailed [940]. Fanaticism [1699]. Fanning [3828, 10975]. Fans [7943, 11389]. Farandole [4974]. Fare [6766]. Farm [606, 3222, 4040, 7417, 7780, 8013, 11330, 10446, 1773]. Farmers [3501]. Farming [4518, 5443, 11572, 3434, 1760]. farms [888]. Farwell [10259]. Fashioned [3274, 8907]. Fast [708, 399, 2468, 3320, 9984, 10594, 11126, 12515]. Fastener [92, 890, 1219, 1683, 2385, 1831, 2447, 2667, 2752, 3353, 3126, 4627, 4635, 4610, 3853, 4215, 5021, 6096, 5076, 6382, 6508, 6664, 7227, 6531, 6258, 6957, 6476, 8109, 8643, 7615, 8103, 8136, 9611, 9177, 10341, 11813]. Fastening [2535, 4921, 7342, 7253, 6619, 7913, 8447, 7756, 9986, 9873, 10339, 10336, 11648]. Fastest [6572, 8329]. Fasting [1314]. Fat [1686, 1876]. Fatal [2122]. Faucet [802, 4194, 5682, 10669, 10869]. Fault [11859]. Faure [2249]. Favor [8524]. Favors [7932]. Fe [8]. Feast [1798]. Feat [12438]. Feathered [7942]. Feathering [12234]. Feature [7607]. Features [6931]. February [4117, 4144, 5442, 5489, 6640, 10663, 11927, 7437, 7955]. Feed [456, 1756, 2239, 4171, 4312, 3809, 5269, 5922, 5519, 7753, 8471, 8910, 10019, 9750, 8843, 8954, 10248, 10497, 10849, 11673, 12569]. Feed-Water [2239]. Feeder [1914, 3395, 2984, 4339, 5573, 6858, 7343, 7869, 9487, 10970, 11312, 12188]. Feeding [1112, 3745, 4096, 7247, 6362, 7936, 7867, 7656]. Feeds [6087]. Feet [1140, 10981, 12206]. Felling [12457]. Felloe [4585]. Felly [233, 11398]. Felly-Clip [11398]. Felt [1080]. Fence [716, 936, 1778, 3167, 3026, 2603, 4495, 3810, 4026, 4530, 5070, 5172, 7084, 6277, 7346, 6551, 7006, 6230, 6479, 7417, 8619, 8467, 8361, 8735, 8929, 9880, 10570, 11107, 11356, 12497, 12547]. Fences [1634, 2542, 7346, 11649]. Fender [6067, 10253, 10620, 12074]. Ferdinand [62, 10845, 10166]. Fere [8228]. Ferments [3197]. Fern [4553]. Ferries [5355]. Ferrule [10442, 10162]. Ferry [4499, 11028, 11168]. Ferryboats [909]. Fertilizer [183, 3570, 9149, 9067, 11565, 12433, 12538]. Fever [658, 906, 132, 4115, 11214, 11409]. Fevers [297, 4995, 5065]. Few [6, 6887, 8165]. Fiber [2190, 2262, 9884, 12437]. Fibers [2702]. Field [398, 2419, 2101, 2145, 2724, 3617, 3526, 4922, 7857, 10506]. Fields [4830, 9771, 12484]. Fifteen [6918]. Fifth [9661]. Fifty [3804, 6146, 7326]. Fight [968, 6344]. Fighting [8890]. Figures [734, 1844, 356]. File [5705, 7782, 8105, 8784, 9401, 10055, 10822]. Files [4235, 5828]. Filigree [2910]. Filing [3743, 9258, 9568]. Fillers [2487]. Filling [8154, 8620, 10104, 12585]. Films [8567, 6583]. Filter [1160, 294, 2564, 2128, 2586, 4369, 4336, 3923, 6524, 8684, 9491, 9964, 10380, 10900, 1713]. Filtering [1058, 5536, 5502]. Filters [3822, 4192, 5536]. Filtration [11088]. Final [7273]. Financial [4591, 4616, 4641]. Finback [194]. Finder [9512]. Findlay [9195]. Fine [5953, 7808, 8591, 9616, 8788]. Finger [1591, 2845, 11199, 10873]. Finger-Supporting [2845]. Fingers [5290]. Finical [4094]. Finish [4094, 6707, 7436]. Finished [4167]. Finishes [8280]. Finishing [1443, 1767, 5412, 5019, 6056]. Finned [1523]. Fire [1036, 897, 1355, 767, 915, 1512, 1401, 1521, 3269, 2748, 2902, 3370, 3643, 2593, 2644, 3093, 3089, 3000, 3150, 2694, 2947, 3483, 4466, 4522, 4417, 4010, 4468, 4552, 4439, 4630, 4676, 3874, 3902, 4367, 4845, 6122, 5264, 5542, 5103, 5334, 5496, 5497, 5543, 5827, 5881, 5660, 6095, 5116, 5141, 5926, 6097, 4972, 5574, 6691, 7008, 7034, 7369, 6181, 7393, 6976, 7194, 7056, 6507, 7981, 7565, 8733, 7415, 7887, 7906, 7983, 8118, 7646, 8807, 9846, 9722, 10041, 10053, 10109, 8876, 9070, 9960, 10249, 10163, 10573, 10711, 11153]. Fire [11420, 11029, 11727, 12114, 12007, 12028, 12349, 3086, 9517, 10021]. Fire-Extinguishing [3093]. Fire-Hose [4845]. Fire-Place [10109]. Firearm [6981, 9013]. Firearms [188, 6597, 8475]. Fired [1233, 1448]. Fireless [9573, 3293]. Firemen [963, 12345]. Fireplace [6067, 6984, 6575, 10927, 12344]. Fireplaces [616, 4098]. Firepro [909]. Fireproof [820, 3296, 3439, 4237, 4121, 9937]. Fires [1122, 711, 657, 2101, 3439, 4077, 4802, 6940, 10489, 11834, 12361]. Fires-Causes [1122]. Firework [2429]. Fireworks [8622]. Firing [3598, 7254, 11667]. First [801, 88, 1245, 3532, 4288, 3900, 4095, 5543, 8494, 9006, 10833, 11536, 12147]. Firth [11911, 10137]. Fish [154, 178, 591, 1558, 2056, 1653, 1469, 1810, 1548, 2134, 2726, 3306, 3273, 3833, 4550, 4395, 3792, 4732, 3879, 4964, 6434, 7223, 7886, 7965, 9867, 10417, 11053, 11747, 2781, 1496, 1667, 7068, 302, 12483, 1194, 243, 1798, 10082, 4699, 5710]. Fish-Hooks [1469]. Fisheries [1504, 7139, 1334, 752, 782, 4549, 4394]. Fishers [7801]. Fishery [864, 6486, 4780]. Fishes [1523, 5340, 8405, 6920, 9370]. Fishing [964, 2056, 3505, 6511, 8540, 11200, 11087, 10162]. Fishing-Rod [10162]. Fishway [994]. Fishways [6769]. Fit [6003, 6637, 8645]. Fitch [1759]. Fits [6357]. Fitted [8081]. Fitter [512, 607, 828, 911]. Fitting [4756, 6235, 10742]. Fittings [3439]. Five [10447]. Fixation [8679]. Fixed [4076]. Fixing [3343]. Fixture [6065, 9176, 10101]. Flag [8755]. Flagon [1522]. Flame [8686]. Flames [10140, 11523]. Flanging [7016]. Flash [5291, 10257, 12541]. Flat [11925]. Flattening [3506]. Flavipes [7941]. Flax [1562]. Fleeming [7055]. Fleet [793, 1173, 12410]. Flesh [5848]. Flexible [4194, 6181, 8764]. Fliers [7402]. Flies [797, 3438, 4902, 10425]. Float [9756, 10668, 10999]. Floating [573, 2513, 3477, 5497, 6905, 9276, 11350, 11135, 11692, 12285, 9516]. Flood [3299, 6972, 7273, 6958, 8929, 2089, 6937]. Floods [3272, 3807, 5489, 8262]. Floor [965, 4237, 4121, 6635, 10048, 9041, 10890, 12259, 12568]. Flooring [4869, 7832, 11873]. Floors [3758]. Florida [4233, 11409, 8790, 11036]. Flotilla [9675]. Flour [655, 6978, 8364, 7424, 10104]. Flow [7435]. Flower [2106, 1492, 3417, 6121, 6839, 6339, 8426]. Flower-Like [1492]. Flowers [332, 7258, 10466, 10273, 10825]. Flows [12283]. Flue [5244, 7171, 9851, 11332, 11878, 2451]. Fluid [40, 5918, 9910]. Fluids [8372, 9291]. Flute [5470]. Fluttering [7811]. Fluviatilis} [2672]. Fly [2088, 1833, 8398, 7811, 9255, 10978, 1593, 9758]. Fly-Wheels [1593]. Flying [1367, 2085, 8939, 9493, 2408]. Flywheel [11176]. Focus [1358, 1440]. Focusing [882]. Fodder [4893]. Fog [1415, 1887, 4731, 3780]. Fogs [1035, 682]. Folding [407, 1288, 717, 1467, 1734, 4476, 4576, 4580, 5286, 6067, 5236, 5164, 5871, 6432, 8032, 9989, 9509, 9094, 9287, 9961, 9751, 10622, 10714, 10258, 11599, 11744, 11951, 11697, 11817]. Fontaine [2133, 2210, 2310]. Fonvielle [1214]. Food [427, 1653, 3306, 2582, 10286, 12360, 12589, 3504]. Foods [2728, 4698]. Foot [4708, 5362, 5547, 5685, 6460, 8645, 10126, 9998, 11308, 11903, 11568, 12320, 11651]. Foot-Shield [11903]. Football [11098]. Footed [3173, 6188]. Footprints [1860, 4868]. Force [694, 14, 1610, 3097, 4312, 4652, 8156, 7908, 11913, 12164, 8655]. Forces [8553]. Foreign [5981, 6320, 8855, 9381, 9409]. Foreigners [5159]. Foreman [4331]. Forest [2101]. Forests [1009]. Forfeit [4213]. Forge [6918]. Forges [6813, 8741, 10015, 7257]. Forging [4004, 4383, 7054, 7622, 10648, 11327]. Fork [2061, 8154, 11849]. Forks [6311]. Form [3656, 3394, 5245, 8504, 8754, 2665]. Formation [1302, 7919, 8679]. Former [9149]. Forming [4939, 6791, 9936, 10132]. Forms [1492, 4842, 6089].

Formulæ [2429, 6930]. Forster [11199]. Fort [10452, 10983]. Fortable [11112]. Forte [7325]. Forth [10137, 6298, 6412, 11338, 11911]. Forts [6545, 7525, 9417, 11792]. Forty [1177, 3106]. Forty-Nine [1177]. Forward [2395]. Fossil [11245]. Fossils [1252]. Foster [11828]. Foundation [7665, 8959, 10240]. Foundations [11999]. Founder [5590]. Founding [5114]. Foundlings [1702]. Foundry [3764]. Fountain [5221, 8901, 11701, 8268, 1847, 5929]. Fountains [3623, 12237, 12441]. Four [2643, 3328, 3927, 4186, 4673, 6238]. Four-Barreled [2643]. Four-Coupled [6238]. Four-Mile [4186]. Four-Roller [3927]. Four-Sided [3328]. Fourteen [3354, 6160]. Fourteen-Story [3354]. Fowl [6361]. Fox [2948, 5294]. Foxes [9493]. Fracture [226]. Frager [3083]. Frame [1563, 3225, 2664, 4601, 4388, 6171, 5663, 5366, 6786, 7365, 8593, 8473, 7443, 8014, 8621, 9511, 9284, 9175, 10714, 11549]. Frames [2011, 7315, 7704, 9989, 11881]. Framing [6651]. France [2922, 4519, 4463, 5074, 6010, 11133, 11256, 10556, 374]. Francois [811]. Frankel [5970]. Frankfort [5708]. Franklin [5515, 164, 859, 6510]. Frauds [3217]. Fraudulent [2999]. Frederic [9638]. Frederick [9969, 12516, 11780]. Free [9746]. Freezer [10998]. Freezing [5003, 5500, 6333, 6308, 7701, 8263]. Fregilupus [12297]. Freight [2066, 3693, 3695, 4440, 5266, 5747, 5785, 6454, 6456, 7253, 7804, 8819, 11192, 12232]. French [798, 9885, 9033, 10876, 75, 3673, 3358, 4029, 6099, 7349, 6780, 8085, 7637, 8314, 9759, 10183, 11178, 11377, 11473, 11779, 12441, 12589, 12169]. Fresh [111, 1890, 2671]. Fresh-Water [1890]. Fret [2590]. Frey [3696]. Friction [4568, 6030, 6098, 6279, 6183, 6254, 6915, 6913, 8592, 7940, 9040, 10164, 12480]. Frictional [12183]. Friend [11027]. Frilled [9450]. Frog [8540, 12209, 12504, 668]. Frogless [8358]. Front [2639, 6597, 11653]. Front-Sight [6597]. Fronts [10260]. Frost [3299, 7733, 9171]. Fruit [689, 63, 1269, 1730, 2515, 2819, 4748, 5145, 6363, 8081, 8157, 8332, 9200, 11080, 10418, 12134]. Fruits [7642, 8708, 9913]. Frying [401, 4311, 6204]. Fuel [1701, 1435, 1436, 2331, 3357, 4056, 6660, 8508, 12164, 12460, 10135]. Fuels [3370]. Full [2057]. Full-Sized [2057]. Fullerton [9519]. Fulminating [2440]. Fumigating [5601]. Fumigator [11383]. Funds [1961]. Funeral [3737]. Fungi [344]. Fungoid [6716]. Funnel [7248]. Furling [9223]. Furnace [451, 350, 2031, 3506, 2801, 2984, 2897, 3324, 3531, 4705, 6117, 5166, 6037, 7345, 7197, 7367, 7808, 8313, 8503, 7752, 7758, 8912, 9818, 10440, 11265, 11949, 12166]. Furnaces [2040, 3745, 4651, 3848, 4096, 7093, 7247, 6813, 7936, 9229]. Furniture [4657, 7373, 7595, 9754]. Furring [8991]. Further [724]. Fuse [11015]. Fuses [9749]. Fusible [4461, 8844]. Fusion [1566]. Future [552, 425, 284, 2107, 2583, 11491, 11969, 10983].

G [1512, 11631]. Ga [2321]. Gab [7571]. Gabert [3507]. Gaffney [2471]. Gains [2794]. Galago [618]. Galatea [8138]. Gale [8583, 10152]. Gall [3438]. Gallery [1689, 6442, 7708, 11971, 12147]. Galloway [1879]. Galls [3875]. Galvani [3255]. Galvanic [1325, 3742]. Galvanization [2503]. Galvanized [5727]. Galvano [3829]. Galvanometer [2338, 6852, 9100, 12145, 931, 8451]. Galvanometers [4409]. Galveston [7776]. Game [1588, 5965, 4925, 5125, 8235, 8962, 12093]. Games [6140]. Gamgee [1680, 1708, 1725, 2421, 1845]. Gang [321, 1754]. Gap [7495]. Garabit [11242]. Garbage [2031, 1431]. Garden [2179, 3438, 4061, 4678, 5551, 7624, 7498, 8158, 7763, 8011, 9886, 10025, 10821, 10810, 10809, 10256, 11816, 7737, 7790, 9071, 9099, 11839, 12083]. Gardens [10673, 10908]. Garfield [2147, 2122, 3345, 6343]. Garment [8711, 8161, 9429]. Garments [2636, 5803, 10364, 10318]. Gas [32, 632, 523, 228, 556, 1024, 165, 69, 487, 1317, 677, 1483, 1448, 1936, 2425, 1768, 3088, 3254, 3669, 3573, 2798, 2855, 2979, 4190, 3877, 4825, 3876, 4853, 4781, 4032, 4472, 4195, 3917, 5713, 5757, 5972, 5566, 5878, 5121, 5655, 5798, 6576, 6979, 6485, 6939, 7337, 6255, 6306, 6782, 6332, 7938, 8076, 8686, 8727, 8441, 7889, 8051, 8112, 7485, 8027, 9195, 8841, 10045, 9310, 9367, 10597, 10904, 10452, 10229, 10445, 11177, 11007, 10565, 11997, 11675, 11454, 11931, 12164, 11606, 12215, 8951, 8466]. Gas-Fired [1448]. Gas-Propelled [11454]. Gases [1432, 5577, 9344]. Gaskill [5563]. Gasometer [7563]. Gastebosteide [243]. Gastroscope [2202]. Gate [64, 1190, 1489, 2751, 2659, 3572, 3697, 3851, 4494, 5707, 5595, 5467, 4992, 6007, 6501, 7342, 7196, 6954, 6880, 7060, 8310, 7589, 7780, 8153, 9560, 9912, 9562, 8816, 9167, 9933, 9228, 10944, 11285, 11330, 10201, 10415, 10420, 10498, 11101, 12113, 11854, 11648, 12414, 6937, 11324]. Gates [10047]. Gatherer [1269, 9200]. Gatling [6840, 8538]. Gauge [801, 39, 225, 2658, 4444, 4038, 4493, 5445, 7122, 8402, 7472, 7728, 8468, 9514, 9756, 9283, 10245, 11010, 10567, 10691, 10686, 11863]. Gauges [834, 3481, 4076, 5306, 6695, 6427, 866]. Gauging [4194]. Gazers [9586]. Gear [777, 2405, 2703, 4778, 4315, 5292, 5359, 7891, 8088, 7837, 10099, 9656, 10568, 10164, 10874, 10690, 10507, 11241, 12326, 12499, 11740, 12281]. Geared [3145, 6409]. Gearing [5358, 10103, 10902, 12183]. Gears [5946]. Gecko [1994]. Geese [9853]. Gelatino [7596]. Gem [1487, 11693]. Gen [7313]. Gen. [11351]. General [1175, 3189, 3499, 2684, 1909, 7288]. Generator [1321, 2065, 8394, 12164]. Generators [8468]. Generosity [3001]. Genesta [6573]. Geneva [2023, 9793]. Genius [3300]. Gentilli [2961]. Geographical [2015, 6140]. Geography [10379]. Geological [1306, 2024, 11212, 7583]. Geologically [4464]. George [11691, 1823, 9638, 10531, 12265]. Gerard [4892]. Germ [12046]. German [8164, 8846, 10388, 12556]. German-Built [8164]. Germany [1098, 3323, 4463, 5708, 5551, 10591, 12285, 9266]. Germinal [2103]. Germs [4009, 8668]. Gerstle [12498]. Get [1702]. Giant [4681, 4039, 9881, 3379]. Gibbon [9071]. Giffard [3651, 3170, 4114]. Gig [11059]. Giganteus} [10650, 8790]. Gigantic [1537, 4937, 7258, 8370, 11245]. Ginning [7053]. Gins [4339]. Girard [12373]. Girder [4082]. Girdle [146]. Girdled [3925]. Girls [1990]. Girth [2632]. Glacial [1632]. Glance [2395, 2805]. Glasgow [9662, 10907]. Glass [398, 1124, 1243, 1936, 2235, 3506, 4102, 4196, 5758, 6960, 7472, 7485, 9877, 12248, 12518, 12542, 9544, 2923, 9520, 2270]. Glasses [18, 8810, 11698]. Glassware [621]. Globe [300, 4196, 12501]. Globes [5567]. Glorietta [11761]. Glossing [3407]. Glossograph [2961]. Gloucester [7801]. Glove [2447, 2367, 6531, 9611]. Glucose [2004, 4698, 3781, 7699, 12248]. Glutinous [1085]. Gluttons [589]. Glycerine [1366, 8531, 8035]. Glyptodon [10956]. Gmelin [5930]. Go [7067]. Goat [10366]. Goblets [1592]. Gold [397, 1964, 2750, 3367, 4683, 5944, 8936, 12567]. Goldbutt [4806]. Golden [1041]. Golding [3272]. Gone [4795]. Good [1354, 527, 606, 1877, 2569, 3915, 3758, 7591, 8730, 8524, 12164]. Goodridge [3965]. Goods [133, 577, 1443, 3871, 10203]. Gopher [15, 96, 3969]. Gordon [3734]. Gorgon [11114]. Gorilla [5369, 10787]. Gorton [10529]. Gothard [85, 2284]. Gould [9854]. Goupil [6459]. Gout [11966]. Government [1171, 2315, 1704, 1982, 7029, 9504, 12091, 605, 656, 3063, 2878, 5159, 8386, 8823]. Governor [1385, 2197, 3602, 3765, 5337, 6982, 6275, 7834, 8564, 10102, 10324, 11605, 12158, 12094]. Gower [323]. Grader [6115, 11907, 12351]. Grading [7193]. Graduating [2170]. Graft [9205]. Grafting [986]. Grain [913, 1560, 1831, 2260, 2774, 3248, 4604, 3787, 5593, 6307, 6497, 6891, 8338, 7699, 8446, 8111, 9095, 9795, 10292, 11350, 10789, 12080]. Grain-Car [1831]. Grammared [4917]. Granada [7475]. Grand [6600, 11690]. Grande [6009]. Granite [4870]. Granny [832]. Grant [1909, 7288, 7313, 9296]. Granting [2794, 11641]. Grape [4334, 4769, 10503]. Grapes [2671]. Graphical [6930]. Graphy [2250]. Grapple [4445]. Grappling [1542]. Grass [1755]. Grate [2345, 2591, 6066, 5166, 6912, 8503, 7784, 8876, 10440, 10738, 11468]. Grate-Bar [2345]. Grate-Fire [8876]. Grater [353, 9088]. Grates [4335]. Grave [1922]. Gravel [11774]. Gravimotor [5235]. Gravity [2730, 6183, 10830, 9024, 9546]. Gray [10025, 10719]. Grease [4171, 9316]. Great [708, 228, 1087, 855, 451, 601, 1698, 2380, 3535, 3523, 3170, 4005, 4796, 3987, 3916, 4330, 4622, 3938, 4830, 3845, 3989, 3960, 5677, 7314, 7336, 7257, 6994, 6412, 6210, 6780, 6852, 7291, 6208, 6458, 8537, 8128, 8441, 8399, 8558, 9502, 9927, 9417, 9087, 9615, 9639, 10437, 10981, 11158, 10223, 10808, 10435, 10406, 12598, 11911, 11984, 12479, 12620, 11533, 3485, 10386, 2863, 8678, 2348, 263, 205, 2418, 2237, 2321, 2078, 3122, 3298, 3297, 3304, 4344, 8769, 10802, 10758, 11410, 10943, 10876, 10343]. Greatly [10938]. Grebes [3072]. Greely [6032, 6109, 5149]. Green [4384]. Greenhouses [3215, 6841]. Greigsville [12543]. Greindl [6143]. Grenade [5827, 5881, 10053]. Grenet [11013]. Grevin [6894]. Greyhounds [7162]. Grid [6905]. Griffiths [7509, 7612]. Grille [7625]. Grinder [5441, 6615, 8209, 7864, 9828, 10878]. Grinding [1592, 5188, 5678, 5867, 5831, 7148, 7422, 8111, 9342, 10783]. Grindstone [12316]. Grindstones [2642, 8074, 10895]. Grip [5879, 6478, 7249, 8480, 10493, 11794]. Gripping [6838]. Griscom [2468]. Grit [3575]. Grizzly [8512]. Groove [1651]. Grooving [5641]. Ground [2687]. Grounding [3674]. Group [10452]. Grouse [11340]. Growth [694, 1387, 4678, 6711, 6716, 8768]. Grubber [10262]. Guacharo [2571]. Guard [4098, 4470, 5546, 6669, 7347, 7032, 7169, 8080, 9286, 10665, 10803, 12027, 12165, 11974]. Guayra [8028]. Guericke [9804]. Guide [2985, 3351, 5269, 5142, 9755, 11812]. Guiding [5290]. Guinea [5645, 7450, 10138]. Gulch [4718]. Gulling [3991]. Gum [10938]. Gummer [10546]. Gumming [9449]. Gun [226, 329, 1364, 937, 750, 384, 634, 2320, 2475, 2643, 4796, 3832, 4598, 5170, 5187, 5748, 5783, 5666, 5942, 5051, 4997, 6781, 6780, 6830, 6840, 7178, 6973, 7002, 8682, 8330, 8538, 7474, 8089, 9637, 8846, 9615, 8779, 9112, 8960, 10094, 9613, 11004, 11234, 11448, 11536, 12010, 12032, 9804]. Gunboat [10876, 12021, 11954]. Gunboats [6099, 5622]. Gunpowder [8454, 7872]. Guns [1850, 8591, 8163, 8861, 9783, 9488, 8891, 9447, 9465, 11225, 10765, 11479, 11968, 12417]. Gurami [1084]. Gutta [780]. Gutter [6550]. Gymnotus [8510]. Gyroscope [8416].

H [6401, 7289, 10531, 11574]. H.M.S. [2567, 7066, 8279, 8923, 10110, 9679, 10590]. H.P [10530]. Haas [540]. Habit [9448]. Habits [8419]. Hailstorms [1676]. Hair [5848, 7125, 8895]. Hairy [48, 3904, 7874]. Half [12409]. Hall [7583, 2426, 8342]. Halls [2498, 11729]. Halpin [12205]. Halske [7916]. Halter [778]. Hamath [9332]. Hamburg [8192]. Hame [1332, 4610, 4608, 6957, 6740]. Hamilton [2189]. Hamlin [10619]. Hammell [12265]. Hammer [560, 2668, 4289, 5969, 5547, 5293, 5048, 6504, 6883, 6408, 7704, 11335]. Hammer-Headed [4289]. Hammerless [937, 2643, 8682]. Hammers [3302, 3698]. Hammock [7275, 11309, 10574, 11771]. Hammocks [11603]. Hand [1161, 181, 531, 1131, 848, 1665, 2012, 1917, 1751, 1567, 2961, 2588, 2660, 4338, 4040, 4655, 4471, 4342, 5569, 5827, 5881, 5897, 5002, 5297, 6406, 6283, 6953, 7396, 6312, 6313, 6814, 7863, 7779, 7865, 8110, 9959, 9178, 10739, 10337, 10384, 11104, 11727, 12181, 11862]. Hand-Grenade [5827]. Hand-Protecting [6953]. Handed [3001]. Handle [261, 326, 1515, 2668, 4285, 5165, 5661, 6785, 7491, 10054, 9825, 10668, 11083, 12498, 11538]. Handles [6004]. Handling [1560, 4743, 7318, 8399, 10645, 10696]. Hands [3533]. Handy [2544]. Hang [9251]. Hanger [1210, 1613, 2663, 2636, 5781, 5808, 5778, 7231, 8181, 8161, 8988, 11858, 12523, 11800]. Hanging [5803, 10823, 10949]. Hangings [3674]. Hank [9931]. Hansom [9822]. Harbor [711, 1156, 3006, 9196, 573, 1231, 1560, 7776, 8341, 8863, 10598, 10707, 10800]. Hard [4869, 3942, 12264]. Hard-Scaled [12264]. Harden [5827, 5881]. Hardening [4773, 5409, 6137, 5285, 7464]. Hardens [631]. Harding [4145, 10713]. Hargreaves [10786]. Harlem [7434, 8959, 10741, 12121, 8291]. Harlequin [3535]. Harmonic [1097, 8566]. Harmonica [11283]. Harmony [6579]. Harness [771, 119, 4283, 7370, 9364, 9066, 12522, 12605]. Harp [4220]. Harrow [4675, 4703, 5364, 5545, 7340, 6714, 8009, 8528, 8530, 8870, 9090, 11109, 11546, 12008]. Hartford [4850]. Harvard [9651, 9706].

Harvest [1844]. Harvester [1266, 291]. Harvesters [1800, 7296]. Has-Done-It [3187]. Haskell [2261, 4796, 4598]. Hastening [9693]. Hastings [11619]. Hatcheling [1414]. Hatchway [2224, 11765]. Hatchways [416]. Hats [1205, 2636, 9457]. Haulage [9800]. Hauling [10504]. Haupt [10766]. Haven [2942]. Hawk [9454]. Hawley [10275]. Hay [2061, 3553, 3248, 4149, 6063, 5417, 7757, 7806, 7750, 9824, 9170, 9175, 10199, 10343]. Hay-Fork [2061]. Hayden [10166]. Head [110, 3096, 5143, 8738, 7566, 9202, 9660, 8786, 10157, 11447, 12135]. Head-Rest [11447]. Headed [4289, 5284, 9154]. Headless [4849]. Headquarters [10021]. Heads [2473, 4238, 6110, 10071]. Headway [9087]. Heal [11986]. Health [1368, 1852, 1404, 10378, 12177, 12493]. Healthiest [11897]. Healthy [1822]. Hear [4036]. Heard [3598]. Hearing [1843]. Heart [288]. Hearth [12166, 12033]. Heat [2282, 1677, 2123, 2077, 2617, 2792, 4669, 4413, 4943, 6411, 6917, 7934, 10638, 10921, 12396, 11926, 1511, 2562, 9291]. Heated [101, 4853, 9747]. Heater [1756, 2128, 2839, 4008, 3809, 5598, 8674, 9222, 8843, 9881, 9937, 9932, 10391, 11127, 10529, 11193, 12210, 12344, 12569, 12025, 11671, 11470]. Heaters [9957, 10163]. Heath [2082]. Heating [556, 1381, 2816, 3215, 4773, 4335, 3947, 5664, 5620, 6235, 8292, 7590, 8315, 8084, 9253, 9536, 9171, 8925, 9587, 8938, 10229, 11358, 11392, 10161, 12080, 11675, 12058]. Heavy [1217, 5748, 4916, 6830, 8667, 11667]. Hébert [10545]. Heel [12348]. Heels [6791, 8874, 9936]. Height [10981, 12206]. Hektograph [3618]. Held [7089]. Helicoidal [7444]. Hell [6937, 4992]. Helmet [1973, 4013]. Heloderma [3280]. Help [11948]. Helped [3501]. Helve [6883]. Helves [9934]. Hemens [4656]. Hemp [11967]. Henri [3651]. Henry [11351, 1104, 3478, 6457, 9651, 9706, 8838, 10219, 12461]. Her [2283, 10943]. Herat [6380]. Hercules [1137, 3114, 4702, 8083, 11929]. Hermella [1193]. Hermit [644]. Hero [8449]. Heroic [6378]. Heron [4681, 5030]. Herring [3722, 3379]. Herrings [3379]. Herrington [4749]. Herz [1493, 3600]. Hiddenite [1491]. Hide [5657]. Hides [4743]. Higgins [11191]. High [1294, 3191, 2746, 4077, 6527, 8554, 10229, 11275, 10463, 11389, 12450, 12185, 7799]. High-Service [2746]. Higher [6522]. Hilgard [8253]. Hill [763, 6064, 7542, 9149]. Hinge [1392, 2445, 8153, 8559, 9908, 10527, 12414, 11668]. Hinged [12498]. Hints [5139, 5963, 12518, 512, 607, 828, 911]. Hipp [3279]. Hiram [11255]. Hiring [8755]. Hirsch [3506]. History [1306, 947, 197, 74, 303, 355, 389, 620, 695, 754, 783, 970, 998, 1223, 1335, 1962, 1834, 1891, 1952, 2547, 3669, 3170, 2714, 3278, 3414, 3601, 2805, 3789, 3970, 4103, 4474, 4807, 4410, 5317, 7449, 8243, 8375, 1889, 12032, 5369, 7639, 7583]. Hitch [7095]. Hiver [6554]. Hiving [11363]. Hoax [3214]. Hobby [6439, 6910]. Hoboken [2730, 8002]. Hodson [3191]. Hoe [2012, 7829]. Hoffmann [6793]. Hog [3133]. Hoist [8186, 11307]. Hoisting [2424, 5755, 6406, 8819, 11008, 11311, 10529, 10343, 11467]. Hoists [6336]. Hold [531, 59, 12001]. Holdback [119]. Holder [1161, 890, 1942, 1656, 1469, 1735, 1801, 2009, 3252, 3578, 2641, 2642, 3192, 2964, 4682, 4446, 3831, 4219, 4366, 5191, 6167, 5296, 5899, 5705, 5638, 6694, 7085, 6557, 7036, 7368, 6530, 6644, 6953, 6647, 6334, 7083, 6503, 7226, 7272, 6759, 6339, 8162, 8055, 7469, 9590, 8934, 9513, 8849, 9608, 9534, 8986, 8903, 9907, 10049, 8875, 8928, 9697, 9537, 8952, 11199, 10381, 10412, 11033, 10600, 10950, 10472, 11283, 10159, 10254, 11032, 10572, 10895, 11007, 10335, 11701, 12501, 12316, 11500, 11723, 12139, 12318]. Holders [8076]. Holding [5290, 11080, 11082, 11767]. Hole [1270, 1682, 3791, 7300, 6689, 9755, 12165]. Holes [5459, 11055]. Holland [1218]. Holley [3044]. Hollow [4602]. Holly [10150]. Holstein [10024]. Holtz [1570]. Holy [2148]. Home [4583, 6109, 5621, 8745, 7651, 8964, 10968, 2486]. Home-Made [5621, 8745, 7651]. Homes [11219]. Homicide [948]. Hon. [5044]. Honey [7956]. Honors [371]. Hood [11543, 12566]. Hoof [10575, 12400]. Hook [380, 2318, 6557, 6986, 7223, 7064, 7680, 8532, 7965, 8644, 8958, 10101, 10094, 9510, 12162]. Hooks [1469, 9730]. Hooping [4866]. Hop [4260, 7467, 10968]. Hope [527, 606]. Hopkins [12276]. Hopper [9064, 11180]. Horace [6225]. Horizontal [14, 3007, 4901, 7326, 9091]. Horn [1415, 1887, 4116, 9184]. Hornbill [1920]. Hornbills [8627]. Horned [1473, 2820, 10208]. Horridum [3280]. Horse [965, 184, 2039, 2061, 4195, 5341, 6146, 5262, 6034, 5050, 5183, 4972, 7377, 7326, 6439, 6910, 7440, 8132, 7809, 7839, 8645, 7783, 10130, 9992, 10251, 11011, 12215, 9758]. Horseradish [5847]. Horses [1217, 778, 459, 4732, 6916, 9256, 9569, 9570, 9255, 12360, 11673]. Horseshoe [615, 3031, 6881, 6936, 7199, 6646, 8645, 10692, 10360, 177, 230]. Horseshoes [989, 5662, 11357]. Horticulture [204, 4118]. Hose [538, 235, 4845, 3943, 4097, 4447, 5389, 4946, 5117, 7393, 7992, 9429, 10821, 11387, 10810, 11778, 11775, 12006]. Hospital [7857]. Hospitals [7150, 7567]. Hostile [793, 4940, 4996]. Hot [1100, 350, 1687, 3352, 4284, 3848, 6117, 7345, 8912, 10469, 10785]. Hot-Air [1687]. Hot-Blast [350]. Hotchkiss [5666, 9783]. Hotel [8502, 10406, 10320]. Housatonic [12329]. House [3354, 3926, 5219, 7243, 7222, 7498, 7438, 10083, 9639, 8847, 11307, 12124, 11524, 11876, 654, 8352, 11324, 11224]. Household [344, 9658, 10998]. Houses [1465, 3100, 3274, 7037, 8316, 7847, 7450, 9117, 11925, 7304, 8118]. Housewife [1088]. Houston [5746]. Howe [1851]. Howell [10977]. Howgate [454]. Hub [209, 5360, 7122, 8421]. Hub-Boxing [7122]. Hudson [768, 845, 1154, 1257, 259, 1745, 3144, 3019, 3577, 3924, 4914, 5407, 9357, 9121, 12090, 1338]. Huge [1047, 4964, 6028]. Huller [3194]. Hulling [1615]. Hulls [8393, 9746]. Human [710, 2202, 1860, 4880, 1921, 4700]. Humane [1408]. Humanity [1699]. Humming [1668, 8745]. Hundred [1149, 2546, 4196, 3804, 9276]. Hunting [1022]. Hurts [3502]. Husker [11416]. Husking [2367]. Huston [1988]. Hut [5991]. Huttinger [3853]. Hyatt [5536]. Hydrants [6875]. Hydrating [5618]. Hydraulic [484, 38, 666, 1318, 1908, 1520, 2853, 3717, 4526, 5759, 5819, 5417, 5238, 6905, 6696, 7016, 6285, 7423, 10136, 10718, 10160, 10382, 1059]. Hydrocarbon [1218, 11652]. Hydrodynamic [6311]. Hydrography [7362]. Hydrophobia [7163]. Hydrostatics [8320]. Hyena [566]. Hygiene [1065]. Hypnoscope [6697]. Hyrax [667].

Ice [635, 1132, 192, 348, 452, 1859, 1873, 2010, 1845, 1492, 3009, 3529, 3595, 3716, 4114, 3741, 4095, 4462, 5313, 5473, 5809, 5386, 6406, 6558, 6620, 7224, 7919, 7751, 7421, 7953, 9986, 10060, 9320, 10735, 12001, 12283, 12395, 3626, 12069, 8006]. Ice-Machine [3716]. Ice-Making [192]. Ideas [1285]. Idle [3390, 7486]. Idlers [3165, 5538]. Igniting [9063]. Ignorance [5945]. Iguana [7277]. Iguanodons [4707]. II [5853, 8117]. Ill [5147]. Illuminated [9317]. Illuminating [1483, 2291, 4781, 7938, 8051, 9866]. Illumination [285, 1874, 1933, 11518, 4218]. Illusion [7316, 11159, 12388]. Illusions [5856]. Illustrated [1172]. Illustrating [744, 11266, 10323, 10344, 9045, 9546, 10788, 10390]. Illustration [9729]. Illustrator [6256]. Image [12305]. Imitate [577]. Imitation [6311]. Imitations [4696, 4116]. Immense [5916]. Immigration [9532]. Impact [8566]. Imperial [5853]. Implement [143, 6206, 8158, 7493, 10694, 10290, 10602]. Impolicy [602]. Importance [2260, 2886]. Important [34, 867, 1403, 2417, 3408, 3617, 5089, 5122, 5168, 5220, 5243, 10486, 12338, 11947]. Importation [853]. Imported [4625, 6733, 10833]. Imports [455]. Imposed [4385]. Impoverishment [4701]. Impromptu [7540]. Improve [7412]. Improved [1164, 583, 824, 1020, 1159, 491, 1362, 1019, 609, 1105, 39, 1053, 76, 1292, 1295, 71, 118, 1056, 1187, 1240, 611, 999, 662, 688, 407, 233, 994, 1024, 1325, 321, 329, 350, 771, 181, 1266, 1332, 261, 770, 582, 966, 439, 65, 1249, 939, 579, 1210, 143, 142, 191, 273, 100, 636, 64, 1209, 488, 1083, 1321, 184, 560, 1181, 1127, 43, 47, 115, 117, 485, 533, 1130, 860, 1061, 320, 714, 849, 777, 687, 1330, 272, 1971, 2242, 1992, 1543, 1881, 1666, 1948, 2497, 1878, 2296]. Improved [2091, 2178, 1914, 2320, 1451, 1779, 2319, 2477, 2129, 2246, 1987, 2111, 2290, 2479, 1683, 1544, 2565, 2173, 2340, 2065, 1949, 2564, 2066, 1415, 1730, 1610, 2106, 2239, 2197, 1392, 2424, 1665, 2224, 2012, 2061, 1917, 1520, 2191, 1545, 2215, 2485, 1941, 2177, 1904, 1636, 1918, 1682, 2295, 2337, 1468, 1568, 2568, 1989, 2400, 2126, 1887, 2385, 1469, 2445, 2365, 2484, 2563, 2346, 1849, 2273, 2017, 2196, 1609, 1807, 1912, 2241, 3412, 3196, 2751, 2801, 3186, 2615, 2616, 3622, 2585, 2815, 3221, 3464, 2987, 3098, 3325, 2587]. Improved [3029, 3695, 2593, 2644, 2819, 3093, 3477, 2984, 2586, 3167, 2798, 2659, 3572, 3697, 3031, 3352, 3009, 2687, 3146, 2693, 2611, 2635, 3301, 3245, 2909, 3071, 2588, 2661, 2735, 2592, 3508, 3276, 2860, 3047, 3624, 2990, 3456, 2884, 2983, 2985, 3454, 2706, 3305, 2667, 3225, 3534, 3025, 2686, 2818, 3253, 3546, 2707, 2846, 3242, 3327, 3090, 2594, 2843, 2614, 3554, 3026, 2796, 3220, 4706, 4851, 4337, 4169, 3788, 4170, 4023, 4086, 4147, 4582, 4242, 4397, 4382, 3764, 4339, 3874, 3902, 4367, 4748, 4845, 3810, 4311, 3822]. Improved [4369, 3742, 4604, 4471, 4260, 3831, 3741, 4239, 4058, 4315, 4361, 4388, 3785, 3813, 3808, 4705, 4281, 4062, 4805, 4363, 4192, 4470, 4438, 3896, 4217, 4365, 4846, 4216, 4336, 3873, 4475, 3812, 4101, 3945, 4603, 4008, 4083, 3744, 4123, 4236, 5680, 5876, 4999, 5162, 5870, 6145, 5474, 5320, 5000, 5516, 5518, 5782, 5663, 5781, 5431, 5075, 5681, 5851, 5259, 5524, 5465, 5659, 5388, 6038, 5027, 5779, 5682, 5342, 5245, 5660, 6095, 5166, 5070, 5172, 5547, 6098, 5467, 5783, 4946, 5117, 5666, 5002, 5363, 4971, 5237, 4973, 5022]. Improved [6142, 5849, 4980, 5640, 5641, 4947, 6069, 6035, 5171, 5576, 4979, 5729, 5924, 5045, 5686, 5950, 5023, 5549, 5545, 4978, 5873, 5311, 4944, 5047, 5750, 5049, 5050, 5293, 5598, 5976, 5905, 5265, 5973, 5546, 5902, 5525, 5097, 5824, 6017, 5165, 5024, 6002, 5850, 5071, 5735, 5073, 5874, 5211, 5987, 7319, 7231, 6556, 6226, 6916, 6431, 7082, 7270, 6858, 6865, 6525, 6578, 6816, 6277, 6764, 7204, 6276, 6360, 6688, 7039, 7115, 6258, 6766, 6550, 7249, 6811, 6425, 6977, 6784, 7033, 6342, 7295, 6648, 7320, 6335, 6575, 7094]. Improved [6304, 6959, 7294, 7372, 6982, 6836, 6978, 6254, 6976, 7194, 7223, 7346, 6551, 7006, 7248, 6954, 6891, 6485, 6840, 7148, 6957, 6312, 6313, 6285, 6814, 6714, 6881, 6936, 7199, 6558, 7373, 7032, 7059, 6328, 6754, 7293, 7268, 6808, 6231, 7271, 7322, 6735, 7166, 7200, 7344, 6385, 7195, 6661, 6762, 7374, 7151, 7080, 6529, 7091, 6738, 7015, 7086, 6617, 6668, 7269, 6718, 7087, 6889, 6860, 7142, 7228, 6259, 6763, 6739, 6504, 7064, 6303, 6641, 6713, 7123, 6789, 6935, 7011, 6228, 6229, 6693, 6987, 6203, 6455, 6665, 6369]. Improved [6384, 6879, 6581, 6817, 6615, 6334, 7170, 6835, 6278, 7145, 6712, 7114, 7395, 6938, 7167, 6524, 8480, 7521, 7789, 7889, 8064, 8017, 8033, 8420, 7985, 8160, 7468, 8425, 8529, 7438, 7653, 8059, 8185, 7439, 8155, 8210, 8211, 8646, 8669, 8588, 8108, 7699, 7703, 7466, 8706, 8209, 8032, 7804, 7832, 8467, 7983, 7780, 8503, 7965, 8013, 8446, 8153, 7422, 8112, 7783, 7590, 7863, 8528, 8530, 8559, 8404, 8475, 7788, 7420, 7917, 7892, 7614, 7810, 8336, 7706, 7911, 8114, 8258, 7803, 8649, 7676, 7698, 8184, 8365, 8621, 7779, 8208]. Improved [7678, 7647, 8012, 8643, 8315, 8644, 8674, 7541, 8057, 8535, 7490, 7516, 8472, 8734, 7843, 7416, 8039, 8137, 8363, 7993, 7749, 7864, 7785, 8234, 7519, 8241, 8422, 8707, 8477, 7833, 8421, 8061, 7755, 7890, 8104, 8586, 8587, 9513, 9569, 9726, 9760, 10018, 9514, 9871, 8892, 9312, 9589, 9702, 9958, 8809, 10106, 9567, 9229, 9570, 9016, 10071, 8904, 9452, 9633, 9957, 9009, 9372, 9700, 9823, 9852, 10076, 8899, 8932, 8957, 8989, 9145, 9174, 9202, 9483, 9563, 9780, 9848, 9995, 10020, 10125, 10098, 9445, 9176, 9795, 9828]. Improved [9660, 9146, 8926, 9776, 8872, 9038, 9257, 9704, 9725, 9882, 8808, 8994, 9119, 9391, 9511, 9539, 9849, 9878, 8812, 9850, 9010, 9230, 9994, 8765, 9508, 9826, 10128, 9311, 9935, 8933, 9313, 9561, 9929, 10047, 9775, 10054, 9231, 9877, 10017, 9777, 9035, 9201, 9962, 10134, 9562, 9937, 10041, 8929, 10053, 9961, 9986, 9255, 9491, 9512, 9880, 10109, 10126, 9487, 9661, 9013, 9964, 10045, 9310, 8876, 9611, 8816, 9367, 10103, 9167, 9933, 10102, 9658, 9824, 10130, 9064, 9783, 9090, 9170, 9091, 9908, 9254, 10046, 10073, 10060, 8995]. Improved [9931, 9486, 9755, 9421, 10055, 10040, 8849, 9630, 9542, 9093, 9368, 9721, 9991, 8871, 9507, 9515, 9540, 9698, 9462, 9223, 9309, 9904, 10050, 10097, 9610, 8893, 9541, 10078, 9018, 9288, 9608, 9276, 9796, 9143, 9534, 9039, 10131, 9588, 9232, 9144, 9227, 9606, 10015, 8986, 9992, 8985, 9480, 9724, 9019, 9566, 10016, 8851, 10133, 10099, 10052, 9940, 9723, 9373, 9393, 9801, 10043, 9779, 10077, 9659, 8869, 9281, 9538, 9872, 10051, 10072, 8958, 9115, 9564, 9932, 9286, 9398, 9177, 9258, 9568, 9228, 8903, 9907, 10049, 9609, 9015, 9458]. Improved [9799, 9153, 9428, 10042, 9369, 8897, 8983, 9260, 9982, 9825, 9423, 9341, 9778, 9402, 9427, 9874, 9481, 9261, 9037, 9876, 9906, 9798, 10013, 9990, 8813, 9797, 9180, 9314, 9756, 10101, 9323, 10014, 8982, 9422, 8952, 10713, 10275, 10945, 11064, 10551, 10737, 10804, 11359, 11130, 10416, 10645, 11006, 10165, 10444, 10324, 10158, 10604, 10982, 11418, 10202, 10225, 10952, 10980, 11196, 10439, 10603, 11304, 11354, 11312, 10803, 10381, 10412, 11033, 10311, 11003, 10942, 10521, 11058, 10310, 10340, 11221, 10533, 11002, 10646, 11113, 10906, 11102, 10181, 10200, 10250, 10319]. Improved [10470, 10495, 10523, 10566, 10784, 10870, 10969, 11034, 11081, 11217, 11281, 11310, 11336, 11417, 11425, 10548, 10979, 11053, 10493, 10391, 10600, 10950, 11416, 10547, 11016, 11328, 11079, 10338, 10500, 10762, 10893, 10946, 10471, 10472, 10868, 11151, 11278, 10874, 11103, 10805, 11054, 10383, 10522, 10693, 11238, 11305, 10313, 11361, 10285, 10280, 10894, 10341, 10878, 11363, 11259, 11306, 10363, 10318, 11390, 11085, 10925, 11078, 10757, 10759, 10245, 11285, 11239, 10361, 10867, 10601, 10622, 10714, 10822, 11176, 11192, 11332, 11398, 10163, 10573, 10711, 11112, 11153, 11420, 10570, 10440]. Improved [11330, 11350, 10418, 10339, 10497, 10975, 10978, 10446, 11200, 11135, 10258, 10849, 10900, 11308, 10442, 10669, 10927, 11107, 10445, 10201, 11059, 11177, 10902, 11010, 10262, 10415, 10420, 10498, 11101, 10469, 10251, 10337, 11358, 10160, 10575, 11283, 11311, 10692, 10384, 11104, 10360, 10527, 11109, 10850, 11060, 10243, 11414, 10687, 10712, 10875, 11421, 10782, 10278, 10526, 11382, 10502, 10443, 11407, 10499, 10641, 10668, 10999, 11005, 11030, 11331, 10247, 10283, 10410, 10534, 10550, 11258, 11108, 10336, 11334, 10549, 10414, 10866, 10827, 11385, 10949, 10290, 10228, 10877, 11337, 10389, 10421]. Improved [10524, 11194, 10282, 11131, 10716, 10226, 10157, 10621, 11423, 10738, 10897, 11128, 11287, 10409, 10690, 10449, 10671, 11381, 10879, 10525, 10244, 10496, 10571, 11279, 11386, 10528, 11001, 10362, 10413, 10722, 10740, 11105, 11198, 10599, 11152, 11282, 11415, 10203, 11360, 10507, 10546, 10567, 10691, 10159, 10254, 11032, 11307, 11392, 10419, 11175, 10317, 10279, 10898, 10973, 11076, 11280, 10647, 11173, 10974, 11155, 10284, 10315, 10872, 11241, 10411, 11422, 11218, 10869, 10277, 10903, 10971, 11129, 10666, 10316, 10572, 10696, 10895, 11333, 10871, 11084, 10508, 10948, 11362, 10717, 11195]. Improved [10281, 10640, 10922, 10503, 10847, 11055, 10252, 10848, 10824, 10976, 10689, 10620, 11356, 11127, 10501, 10182, 10644, 10715, 10826, 11000, 11329, 11393, 10576, 10806, 10828, 10926, 10972, 11261, 11388, 12497, 11597, 12496, 11681, 11879, 12475, 12058, 11749, 12143, 12096, 11482, 11818, 12182, 12560, 12029, 12071, 11539, 11650, 12140, 11570, 11591, 11770, 11772, 11855, 12114, 12188, 12367, 12505, 12162, 12210, 12413, 11647, 12366, 12536, 11623, 12024, 12348, 11724, 11496, 11499, 11540, 11626, 11985, 11444, 11443, 11526, 11593, 11670, 11837, 11882, 11885, 12003, 12284, 12455, 12584, 11550]. Improved [12434, 11605, 12392, 11909, 11596, 12160, 11836, 12456, 11607, 11835, 12051, 12076, 12070, 11740, 11477, 12364, 12607, 11700, 11975, 12537, 12539, 12235, 11472, 12112, 12369, 12452, 11542, 12389, 11566, 11832, 12319, 12457, 12113, 11800, 11720, 12052, 12342, 11571, 11799, 12564, 11902, 11603, 11875, 12298, 12435, 11575, 11948, 12099, 11831, 11838, 12282, 11854, 12304, 12477, 12002, 11568, 11599, 11744, 12320, 12568, 11951, 12074, 12232, 12480, 11565, 11949, 12007, 12028, 12349, 12344, 12395, 11873, 11673, 12134, 12234, 11697, 11817, 11747, 12569, 12547, 12093, 11468, 11812, 11648, 12414]. Improved [11931, 11774, 12027, 12158, 11794, 12316, 12094, 11652, 11765, 11775, 12006, 12185, 11467, 11447, 11538, 11543, 11771, 11668, 12522, 11696, 12563, 11576, 11541, 12165, 11743, 11719, 12365, 12415, 11497, 11998, 12078, 12583, 11649, 11796, 11901, 12279, 12341, 12138, 12545, 11722, 11977, 12118, 11857, 12184, 12559, 12603, 11500, 12624, 11976, 12280, 11795, 12142, 12391, 12500, 12630, 11972, 11830, 12520, 11469, 11669, 12117, 12323, 12394, 11646, 11814, 11950, 11495, 11609, 12031, 11601, 11627, 12416, 11723, 12139, 12186, 11625, 12072, 11547, 11498, 11549, 11478, 12606, 12558, 11905, 11505]. Improved [11702, 12231, 11504, 11624, 11628, 12111, 12209, 11907, 12351, 11546, 11521, 12478, 11645, 11651, 12370, 12629, 11741, 12115, 12474, 11769, 12030, 12346, 11815, 12159, 11880, 12054, 12136, 12609, 11594, 11672, 11777, 11932, 12008, 11595, 11904, 12433, 11476, 12328, 11813, 12345, 12025, 12318, 12605, 11474, 12299, 12362, 12368, 11577, 11592, 12161, 12538, 11567, 12303, 12211, 12371, 11621, 11877, 11981, 12191, 12213, 12233, 12278, 12454, 12625, 11445, 11776, 11742, 11446, 11739, 12073, 12301, 12604, 12611, 11590, 11970, 11773, 11694, 11979, 12141, 12166, 12048, 12582, 11564, 11522, 11811]. Improved [12135, 11451, 12281, 11503, 11930, 12519, 12390, 12053, 11974, 12347, 11933, 12495, 12193, 12302, 12546, 11475, 11622, 12453, 11552, 12022, 12095, 12190, 11884, 12322, 7149, 12562]. Improvement [805, 89, 18, 1112, 616, 416, 989, 141, 584, 217, 831, 463, 1263, 1271, 1291, 834, 1637, 2455, 2062, 1634, 1439, 2039, 2542, 1850, 2192, 2040, 1518, 2092, 2171, 1638, 3718, 2963, 2844, 3005, 3193, 3089, 3224, 3004, 3092, 3533, 4238, 4368, 4335, 4087, 4781, 4754, 4124, 3926, 4847, 4678, 5001, 5308, 6912, 6260, 6939, 7963, 9934, 8810, 10108, 11111, 10899, 10707, 12044, 11653]. Improvements [400, 847, 1871, 2154, 5092, 5925, 7053, 6600, 6951, 6854, 7989, 7546, 11253, 11886]. Improves [12501]. Improving [3940]. Impurities [895, 4171, 7418]. Inadequate [12579]. Inauguration [2957, 8307, 12588, 11691]. Incandescence [1048]. Incandescent [3254, 3189, 2971, 9295, 10176, 12229, 12277, 10473]. Incendiarism [12047]. Incendiary [958]. Inch [7474, 10246, 10577]. Incidents [8443]. Inclined [8002]. Inclinometer [10782]. Incombustible [9630]. Inconveniences [2470]. Increase [1174, 5136, 6452, 11492]. Increased [4841]. Increasing [2285, 4210, 5063]. Incubator [1545]. Independence [6856]. Independent [6837]. Index [1282, 4, 29, 652, 675, 705, 730, 761, 790, 841, 874, 902, 929, 952, 1006, 1030, 1046, 1071, 1094, 1120, 1146, 1229, 1256, 1283, 1312, 1346, 56, 83, 107, 129, 153, 173, 202, 222, 249, 281, 310, 338, 367, 396, 424, 449, 476, 498, 521, 548, 572, 599, 628, 818, 981, 1170, 1200, 1106, 1377, 1400, 2001, 2030, 2051, 2075, 2099, 2119, 2142, 2164, 2186, 2207, 2232, 2258, 2280, 2307, 2329, 2355, 2377, 2393, 2414, 2436, 2465, 2495, 2525, 2554, 1430, 1456, 1480, 1502, 1531, 1555]. Index [1578, 1602, 1625, 1648, 1675, 1697, 1720, 1744, 1766, 1792, 1817, 1839, 1868, 1902, 1931, 1959, 1980, 1505, 2956, 2580, 2601, 3120, 3138, 3159, 3184, 3210, 3238, 3266, 3291, 3318, 3341, 3365, 3387, 3405, 3428, 3445, 3471, 3493, 3518, 3542, 3565, 3593, 3615, 3638, 3665, 3688, 2626, 3710, 2653, 2680, 2700, 2722, 2744, 2766, 2789, 2809, 2833, 2851, 2871, 2895, 2918, 2936, 2954, 2978, 2998, 3018, 3041, 3058, 3081, 3103, 3757, 3776, 4432, 4457, 4485, 4513, 4540, 4564, 4596, 4621, 4647, 4666, 4692, 4715, 4741, 4767]. Index [4792, 4817, 4838, 4864, 4890, 3802, 3821, 3844, 3866, 3890, 3912, 3958, 3978, 4002, 4022, 4050, 4072, 4161, 4182, 4206, 4228, 4253, 4272, 4300, 4327, 4355, 4380, 4406, 4073, 4716, 4912, 4935, 5562, 5587, 5609, 5631, 5654, 5675, 5700, 5721, 5745, 5769, 5794, 5818, 5842, 5864, 5891, 5914, 5940, 5960, 5984, 6001, 6026, 6053, 6082, 6106, 6131, 6158, 4961, 4990, 5014, 5040, 5061, 5087, 5110, 5134, 5157, 5181, 5206, 5229, 5254, 5280, 5304, 5329, 5353, 5378, 5405, 5429, 5455, 5483, 5512, 5534, 5828, 6591, 6178, 6195]. Index [6802, 6825, 6849, 6873, 6903, 6927, 6949, 6970, 7000, 7027, 7050, 7075, 7105, 7133, 7161, 7187, 7216, 7239, 7264, 7285, 7311, 7334, 7358, 7385, 6220, 6247, 6269, 6296, 6325, 6353, 6376, 6399, 6423, 6450, 6469, 6494, 6519, 6542, 6570, 6590, 6612, 6633, 6658, 6680, 6704, 6729, 6752, 6778, 7410, 7432, 8071, 8097, 8126, 8149, 8176, 8204, 8225, 8251, 8277, 8305, 8327, 8350, 8384, 8414, 8439, 8461, 8492, 8520, 8551, 8580, 8608, 8635, 8664, 8695, 8724, 7460, 7483, 7507, 7532, 7555, 7581, 7605, 7635, 7663, 7692]. Index [7718, 7744, 7772, 7797, 7824, 7854, 7882, 7904, 7929, 7951, 7977, 8000, 8025, 8048, 7855, 8521, 9576, 8753, 8777, 9443, 9475, 9500, 9529, 9557, 9583, 9602, 9626, 9649, 9673, 9691, 9715, 9740, 9768, 9790, 9815, 9840, 9863, 9896, 9923, 9951, 9980, 10009, 10036, 8800, 10068, 10092, 10118, 8832, 8860, 8888, 8921, 8947, 8974, 9004, 9030, 9057, 9083, 9110, 9138, 9163, 9192, 9216, 9246, 9274, 9304, 9352, 9380, 9386, 9414, 9247, 10148, 10174, 10796, 10817, 10840, 10861, 10888, 10916, 10936, 10965, 10993, 11025, 11046, 11071, 11095, 11122]. Index [11143, 11166, 11188, 11210, 11232, 11251, 11273, 11298, 11322, 11348, 11374, 11405, 10194, 10217, 10238, 10271, 10302, 10332, 10355, 10375, 10402, 10433, 10461, 10483, 10517, 10542, 10563, 10588, 10614, 10634, 10659, 10682, 10705, 10730, 10753, 10776, 11438, 11460, 12065, 12088, 12106, 12130, 12153, 12175, 12203, 12224, 12246, 12272, 12293, 12313, 12335, 12357, 12378, 12407, 12427, 12447, 12468, 12490, 12512, 12513, 12531, 12553, 12576, 12596, 11487, 12618, 11514, 11531, 11560, 11585, 11616, 11637, 11664, 11688, 11714, 11734, 11788, 11807, 11826, 11847, 11867, 11893, 11920, 11944, 11963, 11991]. Index [12017, 12042, 12469, 11868]. India [1736, 1921, 780, 8262, 7638, 10448]. India-Rubber [780]. Indian [10505, 12483]. Indiana [5383, 10637, 8951]. Indians [9397]. Indicating [7906]. Indicator [661, 487, 433, 1512, 2904, 4438, 5976, 7732, 7916, 8565, 8875, 9807, 9965, 8898, 10361, 10867, 10801, 11175, 11479, 11983, 12186]. Indigo [2064, 5831].

Individual [7018]. Induced [4027]. Induction [2041, 2972, 2341, 2476]. Inductors [4191]. Industrial [131, 973, 591, 345, 1289, 1463, 2078, 1965, 5147, 5731, 5339, 5566, 8339, 7748, 9743, 9792, 9007, 10739]. Industries [42, 58, 1037, 1072, 1109, 1147, 1180, 1216, 1247, 1268, 1293, 1329, 99, 120, 138, 161, 186, 207, 239, 312, 382, 413, 441, 450, 477, 511, 535, 614, 641, 659, 720, 826, 857, 875, 959, 992, 1172, 887, 780, 1488, 1513, 1533, 1565, 1612, 1663, 1684, 1967, 1981, 2002, 2130, 2180, 2194, 2379, 2540, 2842, 2920, 2986, 3060, 3171, 3803, 4099, 5193, 5216, 5263, 5315, 5333, 5383, 6186, 8802]. Industry [1315, 1008, 905, 921, 2311, 1437, 2980, 5267, 8424, 8727, 8051, 12230, 3504]. Inertia [8879]. Inexpansibility [12069]. Inexpensive [9576, 9510, 11013, 8141, 9206]. Infants [2728]. Infected [1011]. Infection [9955]. Infernal [1348]. Influence [12, 2146, 2941, 4033]. Influences [11326]. Influx [1352]. Information [10891]. Infringement [1534]. Ingenious [2174]. Ingenuity [270]. Ingersoll [9366]. Ingot [10850]. Inhabitant [1594]. Inhalers [7562]. Injector [3391, 3170, 6813, 6580, 9456, 11749, 9729]. Injectors [316]. Injurious [11326]. Injury [4669]. Ink [1821, 2005, 7513, 12563]. Inking [10419, 11696]. Inkstand [461, 11576]. Inman [10852]. Innocent [3567, 2835, 3641]. Innocents [9928]. Inoculating [10853]. Inquiry [5457]. Insane [270, 4164, 9772, 10919]. Insanity [2285]. Insatiate [2899]. Insect [7702]. Insecticides [881, 3739]. Insectivorous [1726]. Insects [2363, 3895, 9141, 10178, 2885]. Insertible [6068]. Inserting [12023]. Inspection [707, 35, 4850, 4333, 4487]. Instance [2562]. Instantaneous [2083, 3748, 5318, 6792, 10801, 10842, 10905]. Instinct [12327]. Instinctive [7004]. Institute [974, 1749, 3216, 3458, 4600, 9677, 9773, 10967, 10920, 12296, 11520, 856, 2034, 7175]. Institutions [11506]. Instruction [2959]. Instrument [19, 211, 1432, 6182, 7497, 8230, 10049, 11259, 12049]. Instruments [12626]. Insulated [11174, 10760]. Insulating [8288]. Insulation [10864]. Insulator [1133, 2687, 8134, 10073, 11544]. Insurance [4850, 4590, 4615, 4640, 5400, 6200, 7455]. Intakes [12429]. Intelligence [13]. Intention [9187]. Intentions [3075, 4506, 4685, 10957, 10959, 10536, 11184, 11804, 11889]. Inter [7610]. Inter-State [7610]. Interest [1175, 10801]. Interesting [2245, 1486, 2473, 2733, 4075, 4151, 6211, 11012, 11718, 9774]. Interests [3244]. Interference [4797, 7721]. Interferences [1438]. Interior [5846, 7932]. Intermittent [7119]. International [696, 1653, 3479, 4303, 4184, 7523, 9911, 9792, 9209, 10684, 752, 1561, 2015, 4443, 5771, 5796, 5046, 5612, 5137, 5746, 6827, 9662, 10011, 10592, 12208, 12252, 12275, 12621]. Interoceanic [5994]. Interrupting [10636]. Intolerance [1700]. Invalid [2191, 9819, 10436, 10468]. Invalidation [2605]. Invalides [11473]. Invalids [2728]. Invented [7736]. Inventing [369]. Invention [267, 372, 88, 1352, 1102, 296, 277, 2107, 2145, 1675, 1562, 2889, 2913, 2794, 3284, 3310, 3331, 3421, 3581, 3584, 3605, 3608, 3702, 2822, 2866, 3155, 3232, 3170, 3527, 3076, 2890, 3129, 3130, 3203, 3231, 3260, 3607, 2825, 2724, 2631, 3085, 3452, 3202, 3285, 3311, 3333, 3609, 3661, 3760, 4398, 4376, 3754, 4917, 5457, 6710, 8378, 8092, 8198, 8245, 8513, 8717, 8091, 8219, 8270, 8322, 8377, 8456, 8484, 8542, 8718, 7907, 8794, 9596, 9619, 9666, 9889, 9944, 10085, 8882, 9076, 9130, 9298, 9523, 9578, 9762, 9888]. Invention [9971, 8824, 9211, 8793, 10001, 9051, 9075, 9349, 8803, 9277, 11400, 10265, 10427, 10454, 10511, 10582, 10607, 10627, 11018, 11040, 11067, 11267, 11291, 10209, 10325, 10392, 10477, 10557, 10581, 10626, 10701, 11910, 12338, 11519, 12550]. Inventions [888, 70, 124, 1305, 1337, 190, 245, 275, 301, 324, 379, 536, 563, 580, 622, 640, 664, 676, 748, 769, 835, 941, 1043, 1162, 1220, 1248, 16, 50, 1211, 1275, 1290, 1323, 1369, 94, 139, 193, 331, 418, 438, 469, 489, 648, 663, 697, 721, 755, 781, 804, 827, 862, 886, 991, 1066, 1165, 1183, 601, 4, 29, 652, 675, 705, 730, 761, 790, 841, 874, 902, 929, 952, 1006, 1030, 1046, 1071, 1094, 1120, 1146, 1229, 1256, 1283, 1312, 1346, 56, 83, 107, 129, 153]. Inventions [173, 202, 222, 249, 281, 310, 338, 367, 396, 424, 449, 476, 498, 521, 548, 572, 599, 628, 246, 305, 334, 363, 420, 445, 544, 569, 595, 700, 757, 815, 870, 926, 949, 1027, 1225, 1341, 79, 199, 758, 787, 837, 869, 898, 1002, 1089, 1116, 1142, 1308, 1342, 20, 22, 443, 466, 483, 490, 515, 534, 539, 559, 562, 588, 592, 610, 613, 647, 670, 671, 686, 692, 715, 747, 751, 774, 779, 807, 810, 829, 91, 830, 850, 891, 918, 922, 942, 962]. Inventions [972, 987, 1015, 1018, 1057, 1063, 1081, 1110, 1136, 1163, 1192, 1208, 1213, 1238, 1241, 1272, 1276, 1297, 1322, 103, 1333, 1357, 1363, 116, 137, 145, 160, 185, 196, 210, 213, 238, 242, 271, 274, 293, 325, 328, 359, 378, 386, 412, 414, 435, 699, 1001, 1215, 1265, 1300, 1326, 46, 617, 1390, 1417, 1514, 1541, 1641, 1661, 1785, 1799, 1835, 1858, 1975, 2045, 2132, 2203, 2268, 2323, 2372, 2449, 2504, 2003, 1386, 1589, 1658, 1710, 1752, 1781, 1806, 1854, 1893, 2038, 2135, 2225, 2265, 2294, 2322]. Inventions [2336, 2371, 2406, 2459, 2516, 2534, 1377, 1400, 2001, 2030, 2051, 2075, 2099, 2119, 2142, 2164, 2186, 2207, 2232, 2258, 2280, 2307, 2329, 2355, 2377, 2393, 2414, 2436, 2465, 2495, 2525, 2554, 1430, 1456, 1480, 1502, 1531, 1555, 1578, 1602, 1625, 1648, 1697, 1720, 1744, 1766, 1792, 1817, 1839, 1868, 1902, 1931, 1959, 1980, 1394, 1424, 2303, 2352, 2374, 2432, 2490, 2521, 2550, 1453, 1477, 1527, 1574, 1644, 1671, 1717, 1741, 1788, 1813, 1898, 1927, 1954, 1997, 2026, 2071, 2095, 2137, 2160, 2183, 2204, 2228]. Inventions [2255, 2277, 2187, 1373, 1396, 2252, 2301, 2325, 2351, 2390, 2431, 2460, 2491, 2522, 1425, 1476, 1551, 1573, 1598, 1672, 1716, 1740, 1763, 1787, 1812, 1836, 1862, 1896, 1925, 1955, 1976, 1996, 2072, 2094, 2114, 2139, 2161, 2182, 2253, 2492, 1388, 1413, 1830, 1857, 1861, 1883, 1888, 1923, 1943, 1970, 2019, 2022, 2047, 2067, 2070, 2084, 2113, 2125, 2136, 2153, 2175, 2181, 2201, 2216, 2219, 2240, 2271, 2274, 1422, 2289, 2293, 2317, 2342, 2347, 2366, 2369, 2383, 2388, 2402, 2410, 2448, 2456, 2483, 2488]. Inventions [2512, 2520, 2543, 2566, 2573, 1449, 1466, 1474, 1490, 1517, 1524, 1538, 1540, 1569, 1571, 1611, 1616, 1639, 1669, 1690, 1706, 1733, 1738, 1758, 1776, 1803, 1809, 1829, 1863, 2115, 2042, 2286, 3552, 3163, 2618, 2669, 3539, 3560, 3579, 3586, 3610, 3631, 3659, 3681, 3703, 3726, 2737, 2762, 2867, 2931, 3012, 3053, 3070, 3115, 3153, 3179, 3201, 3229, 3286, 3312, 3332, 3382, 3420, 3442, 3467, 3488, 3515, 2647, 2662, 3259, 3282, 3308, 3330, 3359, 3380, 3418, 3440, 3465, 3486, 3513, 3537, 3558, 3583, 3629, 3657, 3679]. Inventions [2732, 3700, 3724, 2761, 2777, 2800, 2824, 2864, 2887, 2912, 2929, 2991, 3010, 3027, 3091, 3116, 3152, 3176, 3178, 3199, 3227, 2580, 2601, 3120, 3138, 3159, 3184, 3210, 3238, 3266, 3291, 3318, 3341, 3365, 3387, 3405, 3428, 3445, 3471, 3493, 3518, 3542, 3565, 3593, 3615, 3638, 3665, 3688, 2626, 3710, 2653, 2680, 2700, 2722, 2744, 2766, 2789, 2809, 2833, 2851, 2871, 2895, 2918, 2936, 2954, 2978, 2998, 3018, 3041, 3058, 3081, 3103, 2576, 2597, 2623, 2677, 2716, 2738, 2786, 2974, 2994, 3038, 3077, 2575, 2596, 3425]. Inventions [3444, 3470, 3492, 3517, 3589, 3634, 3685, 3706, 2622, 2676, 2697, 2740, 2785, 2828, 2992, 3015, 3054, 3076, 3101, 3117, 3135, 3157, 3181, 3206, 3235, 3263, 3288, 3314, 3337, 3362, 3384, 3401, 2589, 2612, 2868, 2888, 2891, 2914, 2930, 2932, 2944, 2967, 2968, 2989, 3008, 3011, 3030, 3033, 3048, 3069, 3074, 3094, 3099, 3110, 3113, 3151, 3154, 2620, 3177, 3180, 3200, 3204, 3228, 3233, 3261, 3283, 3287, 3309, 3313, 3334, 3360, 3361, 3381, 3383, 3398, 3399, 3419, 3422, 3441, 3443, 3466, 2633, 3468, 3487]. Inventions [3489, 3514, 3516, 3538, 3540, 3559, 3561, 3580, 3582, 3585, 3587, 3606, 3611, 3627, 3628, 3630, 3632, 3658, 3660, 3680, 3682, 3701, 3704, 2638, 3725, 3727, 2645, 2673, 2688, 2692, 2705, 2715, 2729, 2736, 2758, 2760, 2778, 2780, 2783, 2795, 2804, 2817, 2823, 2826, 2841, 2865, 3348, 2829, 3055, 2793, 2981, 2717, 2949, 2975, 3678, 3424, 3230, 3751, 3770, 4400, 4426, 4451, 4479, 4507, 4533, 4559, 4588, 4613, 4638, 4661, 4686, 4711, 4735, 4760, 4785, 4810, 4833, 4857, 4883, 4905, 3795, 3815, 3836, 3859, 3882, 3906]. Inventions [3932, 3952, 3994, 4016, 4044, 4065, 4106, 4132, 4154, 4175, 4200, 4246, 4265, 4292, 4320, 4348, 4373, 3749, 3793, 4424, 4449, 4477, 4505, 4531, 4557, 4586, 4611, 4636, 4659, 4684, 4709, 4733, 4758, 4783, 4808, 4831, 4855, 4881, 4903, 3834, 3857, 3880, 3930, 3950, 3971, 3992, 4014, 4042, 4063, 4104, 4130, 4152, 4173, 4198, 4221, 4244, 4263, 4290, 4318, 4346, 4371, 3757, 3776, 4432, 4457, 4485, 4513, 4540, 4564, 4596, 4621, 4647, 4666, 4692, 4715, 4741, 4767, 4792, 4817, 4838, 4864, 4890, 3802, 3821, 3844, 3866]. Inventions [3890, 3912, 3958, 3978, 4002, 4022, 4072, 4161, 4182, 4206, 4228, 4253, 4272, 4300, 4327, 4355, 4380, 4406, 3729, 3773, 4788, 4860, 4886, 3799, 3818, 3840, 3862, 3886, 3908, 3954, 3998, 4019, 4046, 4068, 4109, 4134, 4158, 4179, 4203, 4225, 4248, 4296, 4323, 4351, 4376, 4536, 4763, 3750, 3752, 4043, 4045, 4064, 4066, 4105, 4107, 4131, 4133, 4153, 4155, 4174, 4176, 4199, 4201, 4222, 4223, 4245, 4247, 4264, 4266, 4291, 4293, 4319, 4321, 4347, 4349, 3769, 4372, 4374, 4399, 4401, 4425, 4427, 4450, 4452, 4478]. Inventions [4480, 4508, 4532, 4534, 4558, 4560, 4587, 4589, 4612, 4614, 4637, 4639, 4660, 4662, 4687, 3771, 4710, 4712, 4734, 4736, 4759, 4761, 4784, 4786, 4809, 4811, 4832, 4834, 4856, 4858, 4882, 4884, 4904, 4906, 3794, 3796, 3814, 3816, 3835, 3837, 3858, 3860, 3881, 3883, 3905, 3907, 3931, 3933, 3951, 3953, 3972, 3973, 3993, 3995, 4015, 4017, 4929, 4955, 5626, 5649, 5670, 5715, 5739, 5764, 5788, 5813, 5837, 5858, 5884, 5908, 5932, 5955, 5978, 6020, 6047, 6076, 6102, 6126, 6153, 6174, 4983, 5008, 5033, 5055, 5080]. Inventions [5128, 5152, 5175, 5199, 5224, 5249, 5274, 5346, 5373, 5398, 5423, 5449, 5477, 5505, 5528, 5555, 5582, 5604, 4927, 4953, 5580, 5602, 5624, 5647, 5669, 5693, 5714, 5762, 5787, 5812, 5835, 5857, 5882, 5907, 5931, 5954, 5977, 5995, 6018, 6045, 6074, 6100, 6124, 6151, 6172, 4981, 5006, 5031, 5053, 5078, 5104, 5126, 5150, 5173, 5198, 5222, 5247, 5272, 5298, 5322, 5344, 5371, 5396, 5421, 5447, 5475, 5503, 5526, 5553, 4912, 4935, 5562, 5587, 5609, 5631, 5654, 5675, 5700, 5721, 5745, 5769, 5794, 5818, 5842, 5864, 5891]. Inventions [5914, 5940, 5960, 5984, 6001, 6026, 6053, 6082, 6106, 6131, 6158, 4961, 4990, 5014, 5040, 5061, 5087, 5110, 5134, 5157, 5181, 5206, 5229, 5254, 5280, 5304, 5329, 5353, 5378, 5405, 5429, 5455, 5483, 5512, 5534, 4928, 4930, 5248, 5250, 5273, 5275, 5299, 5300, 5323, 5324, 5345, 5347, 5372, 5374, 5397, 5399, 5422, 5424, 5448, 5450, 5476, 5478, 5504, 5506, 5527, 5529, 5554, 5556, 4954, 5581, 5583, 5603, 5605, 5625, 5627, 5648, 5650, 5671, 5694, 5695, 5716, 5738, 5740, 5763, 5765, 5789, 5814, 5836, 5838, 5859]. Inventions [5883, 5885, 5909, 5933, 5956, 4956, 5979, 5996, 5997, 6019, 6021, 6046, 6048, 6075, 6077, 6101, 6103, 6125, 6127, 6152, 6154, 6173, 6175, 4982, 4984, 5007, 5009, 5032, 5034, 5054, 5056, 5079, 5081, 5105, 5106, 5127, 5129, 5151, 5153, 5174, 5176, 5200, 5201, 5223, 5225, 4965, 6214, 6242, 6821, 6843, 6868, 6897, 6923, 6943, 6964, 6995, 7021, 7045, 7070, 7099, 7128, 7154, 7209, 7233, 7260, 7279, 7306, 7329, 7352, 7379, 7404, 6265, 6291, 6317, 6347, 6371, 6394, 6416, 6445, 6464, 6488, 6514, 6537, 6564, 6586, 6606]. Inventions [6627, 6653, 6675, 6699, 6722, 6746, 6772, 6796, 6213, 6240, 6820, 6842, 6895, 6922, 6942, 6963, 7020, 7044, 7069, 7098, 7127, 7152, 7180, 7208, 7232, 7259, 7278, 7305, 7328, 7351, 7378, 7403, 6264, 6290, 6316, 6346, 6370, 6393, 6414, 6443, 6462, 6487, 6513, 6535, 6563, 6584, 6604, 6626, 6652, 6674, 6698, 6721, 6744, 6770, 6795, 6178, 6195, 6802, 6825, 6849, 6873, 6903, 6927, 6949, 6970, 7000, 7027, 7050, 7075, 7105, 7133, 7161, 7187, 7216, 7239, 7264, 7285, 7311, 7334, 7358, 7385, 6220, 6247, 6269, 6296, 6325]. Inventions [6353, 6376, 6399, 6423, 6450, 6469, 6494, 6519, 6542, 6570, 6590, 6612, 6633, 6658, 6680, 6704, 6729, 6752, 6778, 6190, 6215, 6654, 6676, 6700, 6723, 6745, 6747, 6771, 6773, 6797, 6822, 6844, 6869, 6896, 6898, 6924, 6944, 6965, 6996, 7022, 7046, 7071, 7100, 7129, 7153, 7155, 7181, 6241, 7210, 7234, 7261, 7280, 7307, 7330, 7353, 7380, 7405, 6243, 6266, 6292, 6318, 6348, 6372, 6395, 6415, 6417, 6444, 6446, 6463, 6465, 6489, 6515, 6536, 6538, 6565, 6585, 6587, 6605, 6607, 6628, 6272, 7427, 7453, 8271, 8298, 8323]. Inventions [8407, 8432, 8457, 8485, 8514, 8543, 8575, 8601, 8657, 8748, 7501, 7527, 7548, 7573, 7599, 7628, 7685, 7712, 7766, 7793, 7898, 7923, 7945, 7971, 7995, 8019, 8042, 8143, 8169, 8220, 7426, 7452, 8067, 8090, 8120, 8142, 8168, 8197, 8218, 8244, 8269, 8297, 8344, 8376, 8406, 8431, 8455, 8483, 8541, 8574, 8600, 8628, 8656, 8688, 8747, 7477, 7500, 7526, 7547, 7572, 7598, 7627, 7657, 7684, 7710, 7738, 7764, 7792, 7817, 7848, 7875, 7897, 7922, 7944, 7970, 7994, 8018, 8041, 7410, 7432, 8071, 8097, 8126, 8149, 8176, 8204]. Inventions [8225, 8251, 8277, 8305, 8327, 8350, 8384, 8414, 8439, 8461, 8492, 8520, 8551, 8580, 8608, 8635, 8664, 8695, 8724, 7460, 7483, 7507, 7532, 7555, 7581, 7605, 7635, 7663, 7692, 7718, 7744, 7772, 7797, 7824, 7854, 7882, 7904, 7929, 7951, 7977, 8000, 8025, 8048, 7638, 7428, 7454, 8043, 8068, 8093, 8121, 8144, 8170, 8199, 8221, 8246, 8272, 8299, 8324, 8345, 8379, 8408, 8433, 8458, 8486, 8515, 7478, 8544, 8576, 8602, 8629, 8658, 8689, 8719, 8749, 7502, 7528, 7549, 7574, 7600, 7629, 7658, 7686, 7711, 7713, 7739, 7765]. Inventions [7767, 7794, 7818, 7849, 7876, 7899, 7924, 7946, 7972, 7996, 8020, 7666, 8453, 8771, 9642, 9684, 9709, 9733, 9785, 9809, 9833, 9915, 9973, 10002, 10029, 10062, 10114, 10142, 8825, 8914, 8941, 8967, 8997, 9052, 9104, 9212, 9239, 9268, 9326, 9350, 9377, 9406, 9468, 9495, 9549, 8770, 8792, 9494, 9522, 9548, 9577, 9595, 9618, 9641, 9665, 9683, 9708, 9732, 9784, 9808, 9832, 9855, 9914, 9943, 10000, 10028, 10061, 10084, 10113, 10141, 8852, 8881, 8913, 8940, 8966, 8996, 9026, 9050, 9074, 9103, 9129, 9156, 9186, 9238, 9267, 9297]. Inventions [9324, 9348, 9375, 9404, 9435, 9467, 8753, 8777, 9443, 9475, 9500, 9529, 9557, 9583, 9602, 9626, 9649, 9673, 9691, 9715, 9740, 9768, 9790, 9815, 9840, 9863, 9896, 9923, 9951, 9980, 10009, 10036, 8800, 10068, 10092, 10118, 8832, 8860, 8888, 8921, 8947, 8974, 9004, 9030, 9057, 9083, 9110, 9138, 9163, 9192, 9216, 9246, 9274, 9304, 9352, 9380, 9386, 9414, 8772, 9351, 9376, 9378, 9405, 9407, 9436, 9437, 9469, 9496, 9524, 9550, 9579, 9597, 9620, 9643, 9667, 9685, 9710, 9734, 9763, 9786, 9810, 9834, 9856, 9890, 9916]. Inventions [8795, 9945, 9972, 9974, 10003, 10030, 10063, 10086, 10115, 10143, 8826, 8853, 8883, 8915, 8942, 8968, 8998, 9027, 9053, 9077, 9105, 9131, 9157, 9188, 9210, 9240, 9269, 9299, 9325, 9327, 10168, 10186, 11268, 11292, 11317, 11342, 11369, 11430, 10210, 10231, 10294, 10369, 10393, 10478, 10675, 10747, 10812, 10855, 10910, 10929, 10958, 11138, 10167, 10185, 10834, 10854, 10882, 10909, 10928, 10986, 11017, 11039, 11066, 11090, 11115, 11137, 11160, 11316, 11341, 11368, 11399, 11429, 10264, 10293, 10348, 10368, 10426, 10453, 10510, 10535, 10606, 10652, 10674, 10723, 10746, 10768]. Inventions [10790, 10811, 10148, 10174, 10796, 10817, 10840, 10861, 10888, 10916, 10936, 10965, 10993, 11025, 11046, 11071, 11095, 11122, 11143, 11166, 11188, 11210, 11232, 11251, 11273, 11298, 11322, 11348, 11374, 11405, 10194, 10217, 10238, 10271, 10302, 10332, 10355, 10375, 10402, 10433, 10461, 10483, 10517, 10542, 10563, 10588, 10614, 10634, 10659, 10682, 10705, 10730, 10753, 10776, 10169, 10187, 10813, 10835, 10856, 10883, 10911, 10930, 10987, 11019, 11041, 11068, 11091, 11116, 11139, 11161, 11269, 11293, 11318, 11343, 11370, 11401, 11431, 10211, 10232, 10266, 10295, 10326, 10349, 10370, 10394, 10428]. Inventions [10455, 10479, 10512, 10558, 10583, 10608, 10628, 10653, 10676, 10702, 10724, 10748, 10769, 10791, 11163, 11184, 11206, 11228, 11438, 11460, 12065, 12088, 12106, 12130, 12153, 12175, 12203, 12224, 12246, 12272, 12293, 12313, 12335, 12357, 12378, 12407, 12427, 12447, 12468, 12490, 12512, 12513, 12531, 12553, 12576, 12596, 11487, 12618, 11514, 11531, 11560, 11585, 11616, 11637, 11664, 11688, 11714, 11734, 11788, 11807, 11826, 11847, 11867, 11893, 11920, 11944, 11963, 11991, 12017, 12042, 11434, 11457, 12061, 12085, 12103, 12148, 12170, 12172, 12199, 12220, 12242, 12268, 12289, 12309, 12332]. Inventions [12354, 12375, 12403, 12424, 12443, 12464, 12486, 12508, 12527, 12550, 12572, 12593, 12615, 11484, 11511, 11554, 11557, 11581, 11612, 11633, 11660, 11684, 11710, 11731, 11756, 11784, 11804, 11823, 11844, 11865, 11889, 11917, 11940, 11959, 11988, 12013, 12038, 4050, 8321, 7096]. Inventor [883, 801, 371, 289, 3900, 8524, 8659, 9308, 11027]. Inventors [795, 819, 3166, 3165, 2607, 3550, 4489, 3941, 5568, 5139, 5963, 5485, 5747, 5094, 7550, 8603, 8630, 8690, 8750, 7576, 7877, 8069, 8094, 8122, 8145, 8171, 8247, 8273, 8301, 8409, 8488, 8516, 8546, 7463, 8773, 9408, 9438, 9470, 9525, 9552, 9599, 9621, 9645, 9668, 9687, 9711, 9736, 9764, 9787, 9812, 9835, 9858, 9891, 9918, 9946, 9976, 10004, 10032, 10088, 10116, 10145, 8827, 8884, 8916, 8943, 8969, 8999, 9028, 9078, 9106, 9133, 9158, 9242, 9300, 9328, 9379, 10170, 10189, 10654, 10678, 10703, 10726, 10749, 10771, 10836, 10857, 10884, 10912, 10931]. Inventors [10960, 10988, 11020, 11042, 11092, 11117, 10212, 11162, 11182, 11204, 11227, 11246, 11270, 11319, 11344, 10234, 11371, 11402, 11432, 10267, 10297, 10327, 10351, 10371, 10396, 10429, 10457, 10480, 10513, 10537, 10559, 10584, 10609, 10629, 10151, 11455, 11483, 11987, 12011, 12037, 12084, 12101, 12126, 12149, 12171, 12197, 12218, 12240, 12266, 12288, 12308, 12330, 12352, 12374, 12401, 11509, 12422, 12442, 12463, 12485, 12506, 12526, 12548, 12570, 11527, 12592, 12613, 12635, 11555, 11579, 11610, 11632, 11658, 11682, 11708, 11730, 11754, 11783, 11803, 11821, 11842, 11864, 11915, 11938, 11958, 5044]. Inverted [9347]. Investigating [9250]. Investigation [1432, 7029]. Investment [7861]. Investments [2396]. Iodide [12056]. Irene [10388]. Iridium [1724]. Iron [683, 574, 894, 455, 2505, 2542, 2365, 2473, 1778, 2399, 2481, 3146, 2924, 2734, 3066, 4284, 4191, 4186, 4853, 4146, 4720, 4168, 4366, 4516, 3940, 6057, 4915, 5801, 6149, 5797, 5727, 5114, 5656, 6135, 5285, 5113, 5775, 5565, 5384, 5282, 6974, 6707, 6732, 6388, 6497, 6953, 7178, 8005, 8686, 7541, 8057, 7969, 8666, 8995, 8850, 8977, 8938, 9831, 10821, 10661, 11967, 11933, 12587, 295]. Ironclad [3268, 9464]. Ironing [95, 7984, 7538, 9451, 9178, 9961, 9254, 10046, 9747]. Irons [1515, 9253]. Irradiation [9829]. Irrigating [5465]. Irrigating-Ditches [5465]. Irrigation [4695, 12034]. Irving [11763]. Isaac [9321]. Isatis [5294]. Island [2768, 7534, 10798, 3921, 1235, 1274, 7052]. Islands [12196]. Isochronal [9805]. Isolating [5472]. Isometric [3765]. Issued [3683]. Isthmian [9503]. Isthmus [1017, 109, 4053, 8611]. Italia [11300, 10485]. Italian [3268, 2878, 6249, 7623, 9615, 11300, 11850]. Italy [863, 3244]. Items [5018]. Ivory [4116].

J [12516, 11631]. Jabiru [3147]. Jablochkoff [3462]. Jack [1330, 4704, 5237, 7059, 8115, 9486, 9799, 9147, 11060, 12259, 11873, 12545, 12604, 6331]. Jackets [10129]. Jackscrews [141]. Jacob [12214]. Jacquard [8404]. Jacquelin [5900]. Jacques [11984]. Jacquet [3356]. Jade [5052]. Jails [4083]. Jam [8678]. James [10554, 9126, 7844, 8836, 8993, 10764, 11562]. Janeiro [4876]. January [4724, 5988, 11215, 11441, 12472, 7414, 7955]. Japan [3548, 4212, 6595]. Japanese [1222, 7847, 752]. Jar [1730, 3948, 8081, 7471, 9016, 10418]. Jardin [4127]. Jardiniere [4879]. Jarman [7421]. Jars [9873]. Javanese [4443]. Jaw [3097, 5540]. Jaws [9631]. Jay [9854]. Jeannette [2035, 3294, 2584, 3062, 5521]. Jelly [4964, 1496]. Jenkin [7055]. Jenkins [1660]. Jensen [8906]. Jerboa [12035]. Jerboas [2782, 11816]. Jersey [253, 2024, 1951, 4089, 6343, 9803]. Jessup [11602]. Jet [8527, 9536, 12108, 11657]. Jets [556, 9152]. Jetties [851, 6616]. Jewell [11801]. Jewelry [2154, 2910, 5444]. Jewels [9237]. Jews [1962]. Jib [9931]. Jigger [9796]. Jockey [7683]. John [7146, 11501, 1759, 3393, 3625, 2836, 9307, 9360, 11366, 12164, 11574]. Johnson [12226]. Johnston [291, 12387]. Johnstown [11849, 11887]. Joiners [10048]. Joins [9007]. Joint [5753, 7116, 7705, 9373, 9778, 10151, 12138, 9520]. Jointed [2172]. Jointer [2269]. Joseph [9359, 9296]. Josiah [9887]. Journal [2693, 5363, 9542]. Journals [1100, 5828]. Judge [2174, 8950]. Judges [6404]. Jugs [584, 1691]. Juice [2297]. Julien [9032]. July [3123, 4079, 5356, 6783, 10799, 11872, 8054, 8641]. Jumping [719, 3557, 3875]. June [2876, 4006, 5257, 6523, 10519, 11793, 8054, 7802]. Junk [4872]. Jupiter [1079, 264, 2710, 3779, 5877, 11718]. Justice [10242]. Jute [11493]. Juvet [300].

Kabath [2928]. Kalan [996]. Kaleidoscope [2215]. Kamschatka [9806]. Kane [12004]. Kangaroo [5623]. Kansas [525, 10452, 1252]. Kearney [10981]. Keel [6761]. Keely [1679, 1777, 3141, 5711]. Keep [2569, 6877, 9631]. Keepers [5288]. Keeping [6993]. Keg [7588]. Kensington [8428]. Kentucky [1860]. Kerosene [2527, 10178]. Kettle [714]. Kettles [2844, 4847]. Kettlewelli} [11365]. Key [741, 1219, 3281, 4215, 6788, 8500, 9755, 12165]. Key-Hole [12165]. Keyhole [7032]. Keying [7586]. Keyless [7568]. Keys [9612]. Keystone [9011]. Kid [2335]. Kill [3502, 10616]. Killer [8790, 10082]. Kiln [1544]. Kilns [11768]. Kindler [1521]. King [894, 3722, 863, 7179, 10435]. Kingfisher [9970]. Kingio [123]. Kinzua [2811]. Kitchen [10763, 11719]. Kite [8370]. Kites [10230]. Kiwi [4854]. Klinkerfues [2854]. Kluchefskaya [9806]. Kluchi [9806]. Knapsack [10243]. Kneader [12342]. Knee [10317]. Knickerbocker [7203]. Knife [261, 1244, 2295, 1880, 5441, 4971, 8058, 11414, 11541, 11743, 11625, 12562, 11451]. Knit [577]. Knives [4754, 6556, 10783]. Knob [5659, 8907]. Knockdown [2611, 7373]. Knowledge [1557, 7243, 6250, 9676, 12534]. Knurls [5539]. Kohler [3028]. Königstrasse [8371]. Koolokamba [5394]. Krakatoa [4500]. Krebs [5667]. Krugg [1691]. Krupp [9983, 9615]. Kuilenburg [458]. Kusu [2948].

L. [9831]. Labastide [3966]. Label [7083, 6402, 6404, 8849]. Labels [6084, 6033, 5723]. Labor [567, 1604, 3642, 3738, 6509, 8476, 4358]. Labor-Saving [567, 6509]. Laboratories [7149, 4029]. Laboratory [507, 3947, 7376, 8781, 9717, 1077, 11860, 11802, 10286]. Lace [770, 1640, 8447]. Lacing [10339, 12609]. Lack [4545, 12535]. Lactic [2873]. Ladder [95, 5164, 6141, 6181, 6836, 7322, 7415, 9287, 9312, 10052, 11239, 10279, 11817, 12214]. Ladders [12027]. Ladie [6791, 9457]. Ladle [1450]. Ladle-Shell [1450]. Ladronia} [44]. Lady [5515]. Laff [11600]. Lager [1484, 11588]. Lake [11302, 2023, 11678, 12429, 11676]. Lakes [1507]. Lamater [6401]. Lambertville [4168]. Lambrequin [9042]. Lambs [7656]. Lamp [187, 1164, 77, 255, 662, 688, 1181, 882, 431, 2454, 3254, 2971, 4437, 3897, 5754, 6073, 5655, 6911, 8642, 8620, 9630, 9093, 9938, 11382, 10534, 10176, 12300, 11858, 12523, 12229, 11497, 12341, 12184, 12072, 11590, 11671, 12119, 10473, 9206]. Lamps [1327, 2127, 3643, 4074, 6209, 9866, 10257, 177, 230]. Lance [1159]. Lancelet [4395]. Land [4701, 4037, 7643, 11761, 644]. Landing [846]. Langley [10111]. Lantern [510, 1969, 1913, 7321, 9997, 11386, 12608, 4218, 11480, 10605]. Lanterns [3005, 4087]. Laporte [5417]. Lard [1457, 6236]. Large [684, 194, 348, 1387, 1507, 6811, 8479, 9124, 9282, 9488, 10249, 10487, 11606, 9399]. Lartigue [9425]. Larvæ [9758, 4700]. Laryngoscope [5754, 7007]. Larynx [6257]. Last [1148, 3782, 4142, 7760, 10943]. Lasting [2467, 10502]. Latch [2339, 2565, 5901, 5595, 4973, 5438, 9012, 9775, 10054, 8816, 11306, 11722]. Latchinoff [2533]. Late [11516]. Latest [292, 5090, 5089, 5122, 5168, 5220, 5243, 7241, 8779, 12381]. Lath [6231]. Lathami} [9535]. Lathe [4543, 4746, 4240, 5431, 5750, 6087, 5096, 5308, 7170, 6357, 8481, 7522, 7495, 9748, 9361, 9342, 10246, 12415, 11796, 12279, 11862]. Lathes [1039, 2217, 4778]. Latrobe [11574]. Launch [8329, 11313, 11028, 10404, 12120]. Launches [9574]. Laundry [4853]. Lavater [1572]. Law [7, 654, 1049, 1482, 1960, 3321, 3271, 5410, 9446, 11440, 9045, 229, 342, 31, 4307, 5434, 10405, 11266]. Lawn [713, 41, 1485, 7293, 10278]. Lawrence [6521]. Laws [953, 2234, 2582, 6579, 982, 603, 2875, 2656, 3548, 5381, 5433, 7858, 9059, 8804, 10755, 10788]. Lawson [2908, 2501]. Laying [2756, 8282, 7763, 11970]. Layton [7708]. Lead [523, 1795, 5392, 7266, 9368, 10198]. Leading [5821]. Leaf [73, 1880, 4061, 5944, 4980, 12142]. Leak [11764]. Leakages [4825]. Leaks [11077]. Leander [11936]. Leaping [2782, 6534]. Learn [11235]. Learner [7497]. Leather [2335, 2980, 4696, 5232, 7594, 12078, 1251]. Leather-Link [7594]. Leaves [2702, 1138]. Lebel [11004]. Leclanche [121, 5001]. Lecture [3294]. Lectures [2332, 7933]. Lee [8700, 12187]. Leech [1760]. Leffel [3806]. Left [3001]. Left-Handed [3001]. Leg [5687, 8016, 11006, 11598, 12143]. Legal [7292]. Legislating [375, 1581]. Legislation [2629, 8555]. Legislative [2938, 2900, 2941]. Legislators [500]. Legs [3747]. Leidenfrost [8100]. Leikoes [8766]. Lemmings [11816]. Lemonade [4115]. Lemur [1714]. Lemurs [7790]. Lengths [8230]. Lens [1358, 6790, 12603]. Lepanto [3989, 3891]. Lesseps [62, 10845]. Lesson [2147, 2474, 10222]. Let [8956]. Let-off [8956]. Letter [5044, 5828, 6166, 9421, 10055, 8893, 10875, 10938, 11050]. Lettering [1757].

Letters [1769]. Levees [11627]. Level [2658, 6713, 6667, 9905, 10782, 11280, 8565]. Leveling [930, 1988, 4389, 5701, 6182, 7673, 8961]. Lever [4338, 5901, 6808, 7727, 10671]. Leverrier [12588]. Levesque [8675]. Levitt [3690]. Lewis [8990, 12440, 10766]. Liabilities [844]. Liberty [8387, 1207, 3523, 3961, 5240, 6594, 8307, 8280, 8677]. Library [7595]. Libyan [2812]. Licenses [7559]. Lick [814, 10196, 10578, 11748, 11061, 10577]. Lidgerwood [10529]. Lieut [5149]. Lieutenant [3294]. Life [1224, 639, 265, 2474, 3482, 2874, 2775, 4795, 4392, 4670, 6044, 5354, 5730, 5623, 4945, 6344, 7056, 6200, 8053, 8443, 7788, 9721, 10819, 11169, 11049, 11326, 10687, 10712, 10526, 11977, 11781]. Life-Preserving [10526]. Life-Saving [2474, 2874, 7056, 11977]. Lifeboat [9264]. Lift [2853, 5819, 8681]. Lifter [3786, 8985, 9423]. Lifting [1324, 4441, 4937, 5237, 7059, 8115, 9865, 11901, 12545]. Light [590, 292, 1349, 1096, 565, 879, 162, 745, 1123, 2282, 1606, 1689, 1877, 2202, 1677, 2123, 1497, 1405, 1444, 1459, 3295, 2792, 3413, 2628, 4828, 4408, 4668, 3878, 4720, 4304, 4515, 4992, 5436, 5144, 5732, 6892, 6711, 6389, 6223, 7388, 7695, 10057, 9361, 9574, 9194, 10919, 10941, 10489, 10486, 11112, 11980, 12471, 11795, 11996, 12578, 11852, 7119, 8567, 8264, 8337, 8623, 8050, 8976, 9020, 9128, 2637, 9045, 3790, 8596]. Light-Ships [6223]. Lighter [10904]. Lighthouse [574, 3005, 4873, 4914, 2957, 4992, 5121]. Lighting [32, 632, 1048, 904, 2425, 1842, 3712, 3669, 2655, 3189, 3690, 4012, 3917, 6931, 6893, 6894, 6340, 8316, 7861, 9531, 8924, 10038, 9032, 10820, 11411, 10758, 10891, 10651, 12277, 11463]. Lightning [1176, 767, 1258, 1016, 1233, 1095, 2386, 2838, 4310, 4880, 5291, 6782, 7677, 7679, 7512, 10553, 11649, 12541, 12627]. Lightning-Proof [1016]. Lights [1150, 1650, 4632, 9463]. Lignite [1435]. Like [2358, 1492]. Likely [9953]. Lily [8065]. Limbs [11841]. Lime [2128, 1497, 3670, 3809, 5618]. Lime-Extracting [2128]. Limestone [7166]. Liming [7268]. Line [1287, 750, 3320, 5459, 5949, 4997, 7121, 9752, 12515]. Line-Throwing [750]. Linen [1443]. Liner [10852, 12610]. Lines [1518, 3482, 2687]. Lining [5701, 11260]. Linings [3410, 6284]. Link [7123, 7042, 7648, 7594]. Lion [5712]. Lions [8978]. Lipped [7245]. Lippmann [7267]. Liquefaction [5577, 7250]. Liquid [5339, 8056]. Liquids [1884, 3412, 9727, 10509, 9024, 9546, 9344, 9150]. Liquors [3088, 10079]. Listed [7729]. Literary [197]. Litigation [7136]. Litigations [6906]. Little [679, 2792, 5209]. Livadia [1050, 1299]. Live [1112, 2960, 9867, 9208, 10787]. Liverpool [4841, 11676]. Lives [9334]. Living [5786, 9450]. Lizard [1994, 10208, 9520]. LL.D [7055]. Loa [1125]. Loader [3721, 6094, 7806, 10547]. Loading [5804, 6696, 8682, 7474, 9988, 10342, 10449, 11244, 12033]. Loan [2034]. Lobsters [1783]. Local [3499]. Location [814]. Locht [743]. Lock [2318, 2339, 1941, 2635, 2945, 2882, 3671, 4770, 4877, 4826, 4973, 5022, 4947, 5171, 5213, 5395, 5822, 5540, 6383, 7200, 6762, 7374, 6303, 6641, 6437, 7804, 8475, 8130, 8256, 9012, 10071, 9146, 9775, 9143, 9609, 10314, 11102, 11421, 10283, 10410, 10550, 11258, 10549, 10898, 10973, 11218, 12392, 12138, 11722, 12118, 11857, 12559, 11876, 11545]. Lock-Clamp [11421]. Lock-House [11876]. Lock-Out [4770]. Locking [380, 5438, 5197, 10568, 12583]. Lockport [10150]. Locks [5266, 6991, 7568, 8683, 9612, 10205]. Locomotion [986, 6263]. Locomotive [1245, 2133, 2210, 2310, 1972, 2287, 2144, 2331, 3164, 3172, 3435, 4632, 6160, 6142, 5192, 5785, 6768, 6754, 6471, 8037, 7840, 8554, 7704, 8013, 8131, 9800, 9573, 9542, 10124, 8927, 9703, 11111, 10808, 10844, 10833, 10649, 12045, 12365, 11998, 11955, 11655, 11716, 12057]. Locomotives [565, 2727, 2640, 6993, 6862, 6809, 6624, 7343, 8401, 9392, 8867, 10494, 11333, 12304, 12557, 3293]. Locust [4039, 6499]. Locusts [8908]. Lodge [10258]. Lodoicea [12591]. Log [2497, 2772, 2802, 3944, 5517, 11076, 8678]. Logic [605]. Logs [5676, 10206]. Londe [9431]. London [1035, 3150, 4549, 4752, 6988, 7096, 8242, 10787, 11809, 12227, 881]. Long [919, 1493, 3783, 5020, 8179, 8140, 11049, 3921]. Long-Billed [919]. Lontin [1214]. Looks [9115]. Loom [3143, 6312, 6313, 8154, 7786, 7987]. Loomis [12092]. Looms [8237, 8956, 10316]. Loop [5803, 9066]. Loose [5538, 5594, 6309, 6328, 7252]. Lord [6225]. Lorises [8318]. Loss [2770, 2584, 4830, 4545, 7510, 12514]. Losses [3000, 3713]. Lost [398, 9954]. Loud [6391]. Loud-Speaking [6391]. Louis [811, 7683, 11912, 5810]. Louvre [11729]. Lovell [9036]. Low [290, 5777, 6527, 9587, 12185]. Lowell [2174]. Lower [11908]. Lowering [3186, 9168]. Lowth [11453]. Lubricants [2427]. Lubricating [6044, 5492, 9991]. Lubricator [144, 2485, 4312, 5872, 5391, 5335, 7271, 7252, 10040, 9542, 9990]. Lucigen [10829]. Lucius [2064]. Luggage [12050]. Lumber [4193, 6115, 8008, 7644, 9927, 9635, 10437, 12598]. Luminous [2359, 4283, 12237, 12441, 3604, 10642]. Lump [1810]. Lunar [1406, 5761]. Luraz [2637]. Luscher [7224]. Lyall [3143]. Lycopodium [7846]. Lye [1298]. Lyman [3044, 2261, 4796, 6807]. Lynx [11839].

M [481, 7829, 12461]. M. [1512, 2610, 7230, 7163]. Macadamsfins [8099]. Maccoy [12631]. Macfarlane [4918]. Machine [209, 215, 1010, 798, 288, 491, 867, 115, 348, 885, 1331, 236, 1111, 214, 114, 1158, 1348, 1082, 21, 2505, 2054, 2085, 2315, 1870, 1992, 1850, 2177, 2484, 1592, 1879, 1774, 2478, 2151, 2157, 1754, 2467, 2570, 1564, 1709, 2270, 1845, 1418, 2264, 1447, 2753, 3717, 2615, 3029, 3009, 3090, 2796, 3220, 3510, 3674, 3407, 3343, 3436, 2590, 3595, 2702, 3132, 3484, 3716, 2966, 2799, 3437, 2970, 4551, 4408, 4892, 4102, 3734, 4114, 4851, 3741, 3896, 3968, 4209, 3871, 3901, 4332, 4126, 4673, 4896, 4750, 4852, 4151]. Machine [4416, 3791, 4113, 3990, 3743, 4829, 5313, 5439, 6086, 4948, 6039, 5875, 5852, 5755, 5641, 6002, 5073, 5831, 5942, 5310, 6058, 6071, 5802, 4926, 5923, 5338, 6990, 7398, 7174, 7193, 6210, 7125, 7016, 7396, 6764, 7094, 7148, 6285, 6620, 7086, 7228, 6789, 7224, 7168, 7341, 6791, 6651, 6461, 6284, 6553, 6436, 7401, 6955, 6599, 6281, 7143, 6672, 8338, 8334, 8746, 7470, 7618, 7939, 8014, 8494, 8064, 7653, 7751, 8404, 7917, 7706, 7803, 8012, 7843, 8477, 8061, 7421, 7960, 7921, 8670, 7622, 8159, 8448, 7445, 7838, 7515, 8285]. Machine [8589, 8031, 9968, 9063, 9178, 8904, 9231, 9366, 9824, 9540, 9258, 9568, 9982, 9323, 9342, 9036, 9282, 8779, 9432, 9936, 10132, 9956, 9096, 10022, 9044, 10075, 9747, 10127, 9925, 9831, 10255, 11172, 11065, 10504, 11016, 10363, 11311, 10949, 10421, 11287, 10449, 11381, 11173, 10783, 10207, 10789, 11031, 11674, 12499, 11681, 11720, 12052, 12395, 11467, 12624, 11830, 11905, 11521, 12299, 12362, 12048, 11522, 12495, 12193, 11475, 11471, 11968, 11462, 12306, 10135, 1570, 12010, 11971]. Machinery [819, 722, 1458, 1640, 1380, 1405, 3621, 4866, 3738, 4844, 7893, 7423, 8535, 7594, 7536, 10060, 10852, 10343, 2609]. Machines [549, 551, 1767, 2090, 3323, 4095, 3777, 5210, 5897, 5092, 7300, 6412, 6509, 6962, 6596, 7912, 7446, 7753, 7963, 7781, 9959, 9094, 9663, 10648, 10970, 10165, 10497, 11152, 11746, 12077, 11883]. Mackey [11029]. Maclaine [7302]. Macon [1099]. Macoushies [8715]. Macronus [11365]. Madagascar [9853]. Made [601, 713, 2105, 5539, 5621, 8745, 7651, 1510, 5357, 8697]. Madge [2283]. Madness [11213]. Magazine [8846, 9112, 9654, 11479, 12187]. Magazines [2506, 6686]. Magdalena [1993]. Magic [2298, 2316, 1913, 6819, 7321, 10505, 11396, 11397, 1885, 10605]. Magical [6649]. Magnesium [5144, 10257, 11852]. Magneso [3410]. Magneso-Calcite [3410]. Magnet [1073, 1320, 1104, 6690, 10435, 10287, 11699]. Magnetic [960, 2419, 3112, 5589, 6768, 9793, 10322, 12500, 9547, 2341, 8679, 4729]. Magnetism [813, 5382, 9101, 11620, 11654]. Magnetization [5635]. Magneto [2665]. Magneto-Electric [2665]. Magnetophone [4606]. Magnets [7989]. Magnifying [3067]. Mahdi [6197]. Maiche [2482]. Mail [4467, 5751, 7200]. Main [9842]. Maine [155, 738, 12189]. Mains [11997]. Maintain [1485]. Maintained [7338]. Major [10435]. Make [111, 1421, 2569, 1485, 6819, 7189, 9781, 9049, 12439, 9397, 3511, 5470]. Maker [5297, 7760]. Makers [3847, 4770, 12248]. Making [192, 577, 710, 677, 1483, 1936, 1919, 3529, 2696, 3436, 2712, 4583, 3918, 4036, 4411, 4491, 5916, 5757, 6639, 6461, 6224, 7938, 8427, 7653, 8591, 8163, 7940, 9231, 10060, 9036, 9663, 9960, 10760, 11031, 11966, 11745, 12623]. Malacca [467]. Malady [11098]. Malaisia [8939]. Malaisian [10058]. Malaria [1433, 2727, 3869]. Malarial [4995, 5065, 906]. Malay [8405, 8318]. Malignant [971]. Malleable [6732, 10661, 12587]. Mallet [212]. Malo [9318]. Mammal [9999]. Mammals [11245]. Mammoth [1401, 6505, 8163, 12419, 9728, 1773, 2124, 12238]. Man [1245, 1038, 3218, 6263, 6273, 9279, 11055, 11706]. Management [3020, 4332, 2885]. Manatee [8539, 11036]. Manchester [6991, 10057, 11502, 12295]. Maneuvering [12492]. Maneuvers [8085, 9533]. Mango [10505]. Manhole [9698]. Manila [9884]. Manipulator [10850]. Mannlicher [11448]. Mantel [8374, 10899, 10647]. Mantelpiece [2370].

Manual [3668, 8476]. Manufactories [6522]. Manufacture [1174, 557, 879, 921, 2558, 2192, 2442, 1640, 2309, 3620, 3051, 3396, 2805, 4781, 3928, 4794, 4074, 3787, 4769, 5460, 5809, 6660, 7596, 7485, 9913, 9488, 9072, 11087, 10359, 12586, 1360, 12032]. Manufactured [5536]. Manufacturers [287, 5091]. Manufactures [252, 2333, 4305]. Manufacturing [290, 1859, 4592, 4617, 4642, 4723, 4140, 5283, 7361, 8478, 9116, 10108, 10529, 10150, 11707, 11931]. Manure [6094]. Many [1944, 8396, 9987, 11219]. Many-Color [1944]. Maple [10154]. Maps [12002]. Marble [7963]. March [3087, 4234, 5548, 6758, 10779, 11464, 7437, 7537]. Marengo [4412]. Marent [4718]. Maria [12503]. Marine [1349, 1382, 3024, 5337, 9462, 9965, 9744, 11124, 12275, 3013, 11286, 893, 644, 11738]. Marinoni [4052]. Maritime [12418, 12252, 12339, 11490]. Mark [1482, 2220, 1770, 3271, 8195]. Marker [6643]. Market [4824]. Marking [5221, 7542]. Markings [3568]. Marks [756, 1025, 317, 10, 1389, 1794, 2674, 3175, 3251, 3988, 5646, 5830, 5723, 6402, 12009]. Marl [2487]. Marmosets [4634, 8573]. Marmots [11816]. Mars [3568, 3131]. Marsh [7655]. Marvin [7979]. Mary [7218]. Masdevallia [3855]. Mason [10668, 10999, 3981, 6673, 10619, 10343, 11466]. Mass [7969]. Massachusetts [6332, 10565]. Massage [2453]. Massive [4679]. Master [10618]. Mastodon [10650]. Mat [8062]. Match [6364, 7650, 9017]. Matches [1360]. Material [4798, 5943, 11535]. Materials [5678]. Maternal [12327]. Mather [6532]. Mathew [10698]. Mats [2639]. Matter [1202, 876, 1000, 1195]. Matters [4125]. Mauna [1125]. Maur [4729]. Mausoleum [11851]. Max [4951]. Maxim [882, 5942, 8779, 12010]. Maximum [2007, 5161]. May [2771, 3919, 3846, 4256, 5140, 6413, 10358, 11642, 7641, 7802]. Mayflower [8138]. McCartney [10386]. McCormick [11936]. Means [1404, 3527, 3867, 6116, 7608, 7756, 9223, 11407, 12524, 11895]. Measure [140, 5335, 9635, 8761]. Measurement [12239]. Measurements [7267, 6473, 177, 230]. Measurers [1856]. Measures [4303]. Measuring [1681, 3901, 5849, 5805, 6389, 8368, 8616, 9455, 10129, 9776, 9877, 9547, 11819, 7120]. Meat [3350, 3084, 4034, 12280]. Meat-Carrying [12280]. Mecca [2148]. Mechanic [2034, 6198, 6548, 7586, 6867, 7274]. Mechanical [1031, 20, 443, 483, 539, 562, 592, 610, 670, 692, 751, 774, 810, 829, 91, 891, 922, 942, 962, 1015, 1063, 1110, 1208, 1238, 1272, 1297, 1333, 1357, 137, 196, 213, 238, 274, 296, 328, 378, 414, 1657, 1591, 1388, 1413, 1830, 1857, 1883, 2022, 2067, 2084, 2113, 2125, 2175, 2216, 2240, 2271, 2289, 2317, 2347, 2369, 2388, 2402, 2448, 2488, 2520, 2566, 1474, 1517, 1538, 1571, 1611, 1738, 1758, 1803, 1911, 2446, 2604, 2612, 2888, 2930, 2968, 2989, 3008, 3033, 3069, 3099, 3113, 3151]. Mechanical [3177, 3200, 3228, 3309, 3360, 3381, 3398, 3419, 3441, 3466, 3487, 3514, 3538, 3559, 3582, 3585, 3606, 3627, 3630, 3658, 3680, 3701, 3725, 2645, 2688, 2705, 2736, 2760, 2778, 2795, 2823, 2865, 4274, 4649, 3750, 4043, 4064, 4105, 4131, 4153, 4174, 4199, 4222, 4245, 4264, 4291, 4319, 4347, 3769, 4372, 4399, 4425, 4450, 4478, 4506, 4532, 4558, 4587, 4612, 4637, 4660, 4685, 4710, 4734, 4759, 4784, 4809, 4832, 4856, 4882, 4904, 3794, 3814, 3835, 3858, 3881, 3905, 3931, 3951, 3972, 3993, 4015, 4727, 5306, 5640]. Mechanical [5564, 4928, 5248, 5273, 5299, 5323, 5345, 5372, 5397, 5422, 5448, 5476, 5504, 5527, 5554, 4954, 5581, 5603, 5625, 5648, 5694, 5738, 5763, 5836, 5883, 5996, 6019, 6046, 6075, 6101, 6125, 6152, 6173, 4982, 5007, 5032, 5054, 5079, 5105, 5127, 5151, 5174, 5200, 5223, 5494, 7344, 6533, 6745, 6771, 6896, 7153, 6241, 6415, 6444, 6463, 6536, 6585, 6605, 7520, 8091, 8219, 8270, 8322, 8377, 8456, 8484, 8542, 8718, 7711, 7765, 9903, 8821, 8793, 9376, 9405, 9436, 9972, 10001, 9051, 9075, 9325, 9349, 9102, 11030, 11331, 12391]. Mechanical [12494]. Mechanics [465, 1103, 1135, 1212, 1419, 2387, 3823, 4801, 5591, 6330, 6685, 6932, 6452, 9207]. Mechanism [2223, 2709, 5615, 5922, 7753, 8333, 9392, 10097, 8956, 10497, 10671, 11352, 11240]. Medals [2263]. Medical [1794, 4165, 7030]. Medicinal [1642]. Medicine [582, 1073, 1320, 8396, 11236, 11934]. Mediterranean [1298]. Medusæ [1890]. Medusle [102]. Meeting [854, 3358, 7388, 7640, 10070, 11412, 11996, 11463, 12207, 12132, 12109]. Meetings [3140]. Meigs [8340]. Meister [2064]. Melbourne [6185]. Melde [12055]. Melograph [2610]. Melting [3531, 4059, 7934, 8757, 11680]. Members [6818]. Memorandum [5098, 7089]. Memorial [9651, 9706]. Memory [10386]. Men [1032, 3389, 5092, 5691, 9652, 11996]. Menace [5488]. Mendenhall [12633]. Menlo [1874]. Mennig [8343]. Menominee [894]. Mental [1404]. Mer [12591, 10831]. Merchant [12410]. Mercurial [11657]. Mercury [12056, 8536, 8268]. Merino [878]. Merry [7067]. Merry-Go-Round [7067]. Mersey [7462, 8923]. Message [8495]. Messrs. [5667]. Metal [885, 1811, 2110, 5947, 6169, 7195, 6561, 8478, 8735, 9616, 9123, 10647, 12586]. Metallic [3644, 2856, 5413, 4998, 6358, 6502, 6549, 7914, 8257, 9309, 9904, 9610, 10247, 11972]. Metallized [1510]. Metallurgic [6037]. Metallurgical [2890, 3129, 3203, 3231, 3260, 3283, 2825]. Metallurgy [1846, 1940, 1991, 2016, 2150, 2102]. Metals [5702, 9544]. Metamorphosis [4343]. Meteoric [916, 7190, 7969]. Meteorite [37, 87]. Meteorites [343]. Meteorological [2536, 3556, 5589]. Meteorology [9]. Meteors [3409, 4517, 5749]. Meter [913, 1609, 3083, 3356, 8441, 10441, 9640]. Meters [487]. Method [1249, 182, 66, 377, 503, 1062, 1250, 1686, 1757, 2926, 4239, 4875, 5990, 6043, 7166, 6838, 8084, 8288, 10050, 9431, 9699, 9396, 8937, 10443, 11005, 10161, 10901, 12630, 3416, 11978]. Methods [5920, 6661]. Metropolitan [175, 5688, 6805, 11533]. Mexican [5579]. Mexico [574, 8854]. Miasm [297]. Mica [2169, 3188, 5952]. Mice [2782, 3925]. Mice-Girdled [3925]. Michel [8135, 11630]. Michigan [2899]. Micro [7937, 11863]. Micro-Audiphone [7937]. Micro-Surface [11863]. Micrographic [9761]. Micrometer [3168, 4113]. Microphone [1945]. Microscope [6887, 6256, 6867, 7274]. Microscopic [5146, 7119, 7496, 9047]. Microscopical [11038, 11057, 3626, 6975]. Middlings [7110, 7892]. Midland [6238]. Midwinter [1433]. Migration [205, 6286]. Miking [9291]. Milch [357]. Mildew [4334]. Mile [4186]. Miles [3598, 3804]. Military [7350, 7251, 9693, 9112, 9203, 8802, 10345, 10491, 10734, 11075, 11303, 12399, 12187]. Militia [2557]. Milk [612, 966, 406, 1686, 2509, 3148, 4239, 5415, 6886, 8668, 8709, 8077, 11214]. Milking [1915, 7441, 8873]. Mill [439, 3325, 3534, 3575, 3050, 3128, 2779, 3927, 4702, 4805, 4140, 5806, 6865, 6718, 6787, 7682, 7805, 7422, 7424, 9249, 9347, 9746, 10975, 10228, 10877, 11984, 11976, 11474, 12368, 2423]. Miller [3758, 8476]. Millers [453, 982]. Milling [1296, 7939, 7917, 12048, 1260]. Million [10639]. Million-Dollar [10639]. Millot [10136]. Mills [76, 2497, 3787, 5676, 5573, 5922, 5188, 5517, 6756, 7138, 6498, 8111, 12256]. Millstone [3250]. Milwaukee [7708]. Mind [7885, 11763, 4025]. Mine [1355, 6385, 11828]. Miner [4725, 4437, 9749]. Mineral [1318, 1491, 7668]. Minerals [253, 11254]. Miners [10337]. Mines [1099, 633, 426, 894, 253, 4937, 6979, 7906, 7907, 11360, 11254, 11901, 12179, 1662, 12543]. Mining [578, 740, 206, 155, 377, 1964, 3714, 3607, 7607, 9573]. Mink [8166]. Minneapolis [8253, 8339]. Minor [10176]. Minutes [4876]. Mirror [1161, 407, 2011, 6913, 8563, 12096, 11698]. Mirrors [1871, 11014, 1885]. Misapplied [2683]. Miscellaneous [22, 466, 490, 515, 534, 559, 588, 613, 647, 671, 686, 715, 747, 779, 807, 830, 850, 918, 972, 987, 1018, 1057, 1081, 1136, 1163, 1192, 1213, 1241, 1276, 1322, 103, 1363, 116, 145, 160, 185, 210, 242, 271, 293, 325, 359, 386, 412, 435, 1861, 1888, 1923, 1943, 1970, 2019, 2047, 2070, 2136, 2153, 2181, 2201, 2219, 2274, 1422, 2293, 2342, 2366, 2383, 2410, 2456, 2483, 2512, 2543, 2573, 1449, 1466, 1490, 1524, 1540, 1569, 1616, 1639, 1669, 1690, 1706, 1733, 1776, 1809, 1829]. Miscellaneous [2589, 2868, 2891, 2914, 2932, 2944, 2967, 3011, 3030, 3048, 3074, 3094, 3110, 3130, 3154, 2620, 3180, 3204, 3233, 3261, 3287, 3313, 3334, 3361, 3383, 3399, 3422, 3443, 2633, 3468, 3489, 3516, 3540, 3561, 3580, 3587, 3611, 3628, 3632, 3660, 3682, 3704, 2638, 3727, 2673, 2692, 2715, 2729, 2758, 2780, 2783, 2804, 2817, 2826, 2841, 3752, 4045, 4066, 4107, 4133, 4155, 4176, 4201, 4223, 4247, 4266, 4293, 4321, 4349, 4374, 4401, 4427, 4452, 4480, 4508, 4534, 4560, 4589, 4614, 4639, 4662, 4687, 3771, 4712, 4736]. Miscellaneous [4761, 4786, 4811, 4834, 4858, 4884, 4906, 3796, 3816, 3837, 3860, 3883, 3907, 3933, 3953, 3973, 3995, 4017, 4930, 5250, 5275, 5300, 5324, 5347, 5374, 5399, 5424, 5450, 5478, 5506, 5529, 5556, 5583, 5605, 5627, 5650, 5671, 5695, 5716, 5740, 5765, 5789, 5814, 5838, 5859, 5885, 5909, 5933, 5956, 4956, 5979, 5997, 6021, 6048, 6077, 6103, 6127, 6154, 6175, 4984, 5009, 5034, 5056, 5081, 5106, 5129, 5153, 5176, 5201, 5225, 6190, 6215, 6654, 6676, 6700, 6723, 6747, 6773, 6797, 6822, 6844, 6869, 6898, 6924, 6944]. Miscellaneous [6965, 6996, 7022, 7046, 7071, 7100, 7129, 7155, 7181, 7210, 7234, 7261, 7280, 7307, 7330, 7353, 7380, 7405, 6243, 6266, 6292, 6318, 6348, 6372, 6395, 6417, 6446, 6465, 6489, 6515, 6538, 6565, 6587, 6607, 6628, 7428, 7454, 8043, 8068, 8093, 8121, 8144, 8170, 8199, 8221, 8246, 8272, 8299, 8324, 8345, 8379, 8408, 8433, 8458, 8486, 8515, 7478, 8544, 8576, 8602, 8629, 8658, 8689, 8719, 8749, 7502, 7528, 7549, 7574, 7600, 7629, 7658, 7686, 7713, 7739, 7767, 7794, 7818, 7849, 7876, 7899, 7924, 7946, 7972, 7996, 8020]. Miscellaneous [8772, 9351, 9378, 9407, 9437, 9469, 9496, 9524, 9550, 9579, 9597, 9620, 9643, 9667, 9685, 9710, 9734, 9763, 9786, 9810, 9834, 9856, 9890, 9916, 8795, 9945, 9974, 10003, 10030, 10063, 10086, 10115, 10143, 8826, 8853, 8883, 8915, 8942, 8968, 8998, 9027, 9053, 9077, 9105, 9131, 9157, 9188, 9210, 9240, 9269, 9299, 9327, 10169, 10187, 10813, 10835, 10856, 10883, 10911, 10930, 10959, 10987, 11019, 11041, 11068, 11091, 11116, 11139, 11161, 11269, 11293, 11318, 11343, 11370, 11401, 11431, 10211, 10232, 10266, 10295, 10326, 10349, 10370, 10394, 10428]. Miscellaneous [10455, 10479, 10512, 10536, 10558, 10583, 10608, 10628, 10653, 10676, 10702, 10724, 10748, 10769, 10791, 10595]. Misinstruction [2330]. Misrepresentation [2941]. Mississippi [1173, 851, 1506, 3272, 2906, 6951, 6616, 11886]. Missouri [258]. Mitchell [12503]. Miter [5575, 6735]. Mitrailleuse [3966]. Mixer [4032, 8926, 11355, 12342, 12391]. Mixing [4059, 7562]. Mlle [11934]. Mode [1202, 6671, 8464, 7656, 9867]. Model [895, 8159, 10424]. Moderator [5713]. Modern [763, 3552, 4542, 5464, 10345, 10177]. Modification [3271]. Modified [776, 12306]. Moholi [618]. Moist [6813]. Moistener [6168, 7009]. Moistuer [3376]. Mole [96, 1729, 4396]. Molecular [4054, 12421]. Moles [527]. Mollie [6343]. Mollusks [864]. Molten [12586]. Momentum [9150]. Monarch [7118]. Moncrieff [7178]. Money [3767, 9632]. Monitor [8444, 11156]. Monkey [6562, 10698]. Monkeys [8129]. Monograms [11881]. Monster [4230, 8228, 1252, 1496]. Montejo [10083]. Montgomery [8950]. Monthly [673]. Montreal [3140]. Monument [3345, 2755, 6201, 9296]. Monumental [6390]. Moon [1537, 3124]. Moore [3575]. Mop [9515]. Moral [2146, 2683]. Morbus [5880]. Morgue [8840]. Mormolyce [73]. Morningside [8650]. Morris [12440, 2853, 10672]. Morse [10081]. Mortar [8228, 9507]. Mortars [1908]. Mortem [6148]. Mortisection [2921]. Mortiser [1780]. Mortising [209, 2177, 10165]. Morton [6457]. Mosaic [5241]. Mosquito [11099]. Mosquitoes [3869]. Moss [3553]. Most [5122, 5220, 5243, 2885, 5090, 5089, 5168]. Motay [481]. Moth [4061, 9664]. Mother [216, 1298, 6287]. Mother-Of-Pearl [6287]. Moths [4757, 11379]. Motion [1202, 1680, 1708, 1725, 1704, 1746, 3143, 3174, 3480, 5825, 6116, 6743, 7781, 8237, 9518, 9096, 10022, 10902, 11052, 11782]. Motions [2444, 1605, 8154, 10509, 11314]. Motive [1153, 3649, 6731, 8526]. Motor [944, 1214, 410, 40, 434, 608, 67, 1236, 1678, 1807, 1679, 1777, 1495, 2972, 3397, 3462, 3141, 3049, 2690, 3877, 4524, 4288, 4007, 5218, 5711, 4970, 5238, 7205, 7107, 7206, 7420, 9559, 9418, 9151, 11110, 10786, 11328, 10501, 10939, 12260, 11795, 10204]. Motors [4191, 4275, 4650, 6496, 8585, 9770, 11333]. Mougin [10983]. Mould [65, 3740, 5388, 6886, 12413]. Moulder [3328]. Moulding [432, 8907, 11005]. Mound [2696]. Mound-Making [2696]. Mount [7813, 12359, 12194]. Mountain [1904, 4384, 9783, 10574]. Mountains [3549, 10520, 11449, 8]. Mouse [1221, 3173, 7683, 6188]. Mouse-Eating [1221]. Mouth [5754, 5520, 12263]. Movable [5042, 6635, 7915, 10444, 11178, 10890]. Movement [1911, 2446, 5434, 5640, 5494, 7344, 7042, 8473, 8648, 11030, 11331, 12494]. Movements [385]. Mover [1191]. Moving [7727, 9292, 10943, 10320, 10406, 12420]. Mower [41, 8871, 10278]. Mowers [3622, 7144, 8188, 8362, 9009, 10076]. Mowing [713, 9956]. Mr. [630, 1017, 2147, 3272, 3377, 9717, 8818, 12123, 12276, 2501, 177, 230]. Mt [12566]. Mt. [9996]. Much [6381, 7486, 6430]. Mucilage [10555, 12079]. Mud [11600]. Muffler [7866]. Muhlenberg [10766]. Mulberry [4573]. Mule [7095, 8790]. Muller [3690]. Multichange [3006]. Multicharge [4598, 9060]. Multicycle [9203]. Multiple [3690, 3050, 3457, 9432, 11327]. Multiple-Spindle [3457]. Multipolar [12261]. Multnomah [12462]. Multum [8850]. Multum-In-Parvo [8850]. Mummy [6952, 8117]. Murder [11213]. Muscular [2363, 6088]. Museum [175, 1941, 1965, 5369, 7639, 7583, 6894, 6994, 6442, 8428]. Mushroom [3270, 1773]. Mushrooms [8289]. Music [4212, 4980, 8055, 9362, 10499, 10641, 11500, 12142]. Musical [19, 5120, 6861, 10140]. Musicians [10873]. Mussel [166]. Must [8524]. Mustard [361]. Muybridge [404]. Muzzle [7348]. Myiasis [2088]. Mysterious [1966, 1855].

N.C [8294]. N.C. [1099, 8354]. N.J. [5536, 10981, 11678]. N.Y [4393]. N.Y. [2556, 4828, 6941, 7203, 10180, 10920, 10150, 12543]. N.Z. [5194]. Nîmes [10321]. Nail [3847, 9541, 9036]. Nails [1128, 5233, 5284, 6428, 10696]. Naked [1510]. Nampa [12305]. Nasal [542]. Nation [8558]. National [174, 256, 346, 1076, 1108, 1631, 2334, 3459, 3547, 3479, 3243, 4051, 6818, 7388, 7108, 7640, 11088, 11463, 12207, 12339]. Native [11086]. Natural [74, 303, 355, 389, 620, 695, 754, 783, 970, 998, 1223, 1335, 1042, 1962, 1834, 1891, 1952, 2547, 2046, 2250, 2714, 3278, 3414, 3601, 3226, 3789, 3970, 4103, 4474, 4807, 4410, 5369, 5566, 5317, 7337, 7639, 8645, 7485, 7571, 7449, 8243, 8375, 8027, 9195, 10597, 7583, 1889, 8951]. Naturalists [6892]. Nature [8207, 589, 6734]. Nautilus [8862]. Navajo [9397]. Naval [842, 2398, 1382, 2874, 3413, 8085, 7474, 9533, 9707, 8802, 10617, 11490, 12449, 11491, 11969]. Navies [1974, 6625]. Navigable [6715]. Navigation [682, 3397, 4172, 4822, 11810, 12601, 5903]. Navy [1074, 12144, 12397, 12251, 6099, 6544, 6767, 8356, 8442, 9085, 9615, 9122, 8891, 9571, 9802, 9966, 9478, 12315, 12321, 11492, 12169]. Near [3720, 4418, 6187, 5710]. Nearing [8677]. Nebraska [10494, 11048]. Nebulæ [12524]. Necessary [6545]. Neck [10336]. Necklace [6952]. Necktie [8930, 12520]. Need [4723, 5747]. Needed [5564, 7053, 6854, 10865, 10938]. Needham [12473]. Needing [3641]. Needle [4848, 4732, 6603, 11108, 526]. Needles [11243]. Needs [3108, 8442, 12315]. Neglect [3020]. Neglecting [110]. Neith [5592]. Nemertes [997]. Nepaul [3133]. Nerve [7908]. Nest [1594, 8113, 11547, 3699, 7068, 12504, 243]. Nest-Building [12504, 243]. Nests [800, 2275, 3072, 7847, 2646, 6390]. Net [9255, 10978, 11977]. Netherland [8479]. Nets [11087]. Netting [9323]. Nevada [426, 3218].

Never [298, 6147]. Newark [5536, 9140, 5536]. Newly [9867, 10366]. Newport [9637]. Newspaper [3105, 7782]. Newton [9321, 11266, 10788]. Niagara [2416, 4622, 6546]. Nicaragua [7219, 9868, 11993]. Nickel [726, 251, 2102, 12179]. Nicking [4721]. Night [5901, 9118, 10534, 11202, 7165, 7031, 7641, 8102, 7414, 7437, 7955, 8054, 8641, 7537, 7802, 8180]. Nile [3576, 4941, 7473]. Nine [1177]. Nippers [7112]. Nitrate [12484]. Nitro [8035]. Nitro-Glycerine [8035]. Nitrogen [5731]. No [42, 58, 1037, 1072, 1109, 1216, 1247, 1268, 1293, 1329, 99, 120, 138, 186, 207, 239, 312, 382, 413, 441, 450, 477, 511, 535, 614, 641, 659, 720, 826, 857, 875, 959, 992, 1488, 1513, 1533, 1565, 1612, 1663, 1684, 1967, 1981, 2002, 2130, 2180, 2194, 2379, 2540, 2842, 2920, 2986, 3060, 3171, 3803, 5193, 5216, 5263, 5315, 5333, 6186, 8116, 4191, 12439]. No.38 [161]. No.49 [1180]. No.63 [1147]. No.86 [4099]. Noble [11052, 7151]. Nomenclature [2747]. Non [101, 2527, 2958, 7701, 9793, 9225, 10736, 10322]. Non-Combustible [2958]. Non-Conducting [101, 10736]. Non-Explosive [2527]. Non-Freezing [7701]. Non-Magnetic [9793, 10322]. Non-Rattling [9225]. Nordenfelt [3832, 7043, 10027, 9234]. Norfolk [11666]. Normal [7244]. Norremberg [8567]. North [1632, 633, 7327, 8389]. Northeast [5645]. Northern [4344, 10802, 10184, 724, 7179, 9463, 11324]. Norway [6486, 11655]. Nose [5756, 11673]. Nosed [4501, 4396, 6562]. Noses [4030]. Notable [3509, 7666]. Note [945, 9877]. Notebook [8875]. Notes [74, 303, 355, 389, 620, 695, 754, 783, 970, 998, 1223, 1335, 3, 28, 729, 760, 789, 817, 840, 873, 901, 928, 951, 980, 1005, 1029, 1045, 1070, 1092, 1119, 1145, 1169, 1199, 1228, 1255, 1281, 1311, 1345, 55, 82, 106, 152, 172, 201, 220, 248, 280, 309, 337, 366, 395, 423, 448, 475, 497, 520, 547, 571, 598, 627, 651, 704, 5, 30, 57, 84, 108, 130, 203, 223, 282, 339, 499, 522, 600, 629, 653, 706, 731, 762, 791, 903, 1201, 1313, 1347]. Notes [2032, 1693, 2021, 2200, 1834, 1891, 1952, 2547, 1376, 1399, 2029, 2050, 2074, 2097, 2117, 2141, 2163, 2185, 2206, 2230, 2257, 2279, 2305, 2327, 2354, 2376, 2392, 2413, 2435, 2463, 2494, 2524, 2553, 1429, 1455, 1479, 1500, 1529, 1554, 1576, 1601, 1623, 1646, 1674, 1696, 1719, 1743, 1765, 1790, 1816, 1838, 1866, 1900, 1929, 1957, 1979, 1999, 2212, 1378, 1481, 2415, 2437, 2466, 2496, 2526, 2555, 1503, 1532, 1556, 1579, 1603, 1626, 1649, 1818, 1840, 1869, 1903, 1932, 2052, 2076, 2100, 2120, 2143, 2165, 2208]. Notes [2233, 2259, 2281, 2308, 2356, 2059, 2081, 2109, 2149, 2168, 2195, 2213, 2292, 2313, 2343, 2361, 2384, 2404, 2443, 2519, 2537, 2156, 1770, 3036, 2714, 3278, 3414, 3601, 2579, 2599, 3158, 3182, 3209, 3236, 3265, 3290, 3317, 3340, 3364, 3386, 3404, 3427, 3446, 3473, 3495, 3519, 3543, 3564, 3592, 3614, 3637, 3664, 3687, 3709, 2625, 2651, 2679, 2699, 2720, 2742, 2765, 2788, 2808, 2831, 2850, 2870, 2894, 2917, 2935, 2952, 2977, 2997, 3017, 3040, 3057, 3079, 3119, 3137, 2602, 2627, 3239, 3267, 3292, 3319, 3342]. Notes [3366, 3388, 3406, 3429, 3449, 3475, 3496, 3521, 3544, 3566, 3594, 3616, 3639, 3666, 3689, 3711, 2654, 2681, 2701, 2723, 2745, 2767, 2790, 2810, 2834, 2852, 2872, 2896, 2919, 2937, 2955, 3042, 3059, 3082, 3104, 3121, 3139, 3161, 3185, 3211, 2608, 2666, 2731, 3065, 3109, 3782, 4592, 4617, 4642, 3789, 3970, 4103, 4474, 4807, 3732, 3756, 4326, 4354, 4379, 4405, 4431, 4456, 4484, 4511, 4539, 4563, 4595, 4620, 4646, 4665, 4690, 4714, 4740, 4766, 4791, 4816, 4837, 4863, 4889, 3775, 3801, 3820, 3843, 3865]. Notes [3889, 3911, 3935, 3957, 3977, 4001, 4021, 4049, 4071, 4092, 4111, 4137, 4160, 4181, 4205, 4227, 4252, 4271, 4298, 4128, 3733, 3913, 4742, 4768, 4793, 4818, 4839, 4865, 4891, 3937, 3959, 3979, 4003, 4112, 4138, 4162, 4183, 4207, 4229, 4254, 4273, 4301, 4328, 4356, 4381, 4407, 4433, 4458, 4486, 4514, 4541, 4693, 4717, 4910, 4934, 5561, 5586, 5608, 5630, 5653, 5674, 5699, 5720, 5744, 5768, 5793, 5817, 5841, 5863, 5890, 5913, 5939, 5959, 5983, 6000, 6025, 6052, 6081, 6105, 6130, 6157, 4960, 4989, 5013, 5039]. Notes [5060, 5086, 5109, 5133, 5156, 5180, 5205, 5228, 5253, 5279, 5303, 5328, 5352, 5377, 5404, 5428, 5454, 5482, 5511, 5533, 4913, 4936, 5770, 5795, 5843, 5865, 5892, 5915, 5941, 5961, 5985, 6027, 6054, 6083, 6107, 6132, 6159, 4962, 4991, 5015, 5041, 5062, 5088, 5111, 5135, 5158, 5182, 5207, 5255, 5281, 5330, 5379, 5406, 5430, 5456, 5484, 5513, 5535, 5588, 5610, 5632, 6832, 7078, 6177, 6194, 6801, 6824, 6848, 6872, 6902, 6926, 6948, 6969, 6999, 7026, 7049, 7074, 7104, 7132, 7160, 7186, 7215, 7238, 7263]. Notes [7284, 7310, 7333, 7357, 7384, 6219, 6246, 6268, 6295, 6324, 6352, 6375, 6398, 6422, 6449, 6468, 6493, 6518, 6541, 6569, 6589, 6611, 6632, 6657, 6679, 6703, 6728, 6751, 6777, 6179, 6196, 6803, 6826, 6850, 6874, 6904, 6928, 6950, 6971, 7001, 7028, 7051, 7076, 7106, 7134, 7188, 7217, 7240, 7265, 7286, 7312, 7335, 7359, 7386, 6221, 6248, 6270, 6297, 6326, 6354, 6377, 6400, 6424, 6451, 6470, 6495, 6520, 6543, 6571, 6592, 6613, 6634, 6659, 6681, 6705, 6730, 6753, 6779, 6863, 6888, 6921, 6989, 7013, 7040]. Notes [7079, 7109, 7141, 7176, 7191, 7221, 7449, 8243, 8375, 7409, 7431, 8096, 8125, 8148, 8175, 8203, 8224, 8250, 8276, 8304, 8326, 8349, 8383, 8413, 8438, 8460, 8491, 8519, 8550, 8579, 8607, 8634, 8663, 8694, 8723, 7459, 7482, 7506, 7531, 7554, 7580, 7604, 7634, 7662, 7691, 7717, 7743, 7771, 7796, 7823, 7853, 7881, 7903, 7928, 7950, 7976, 7999, 8024, 8047, 7411, 7433, 8072, 8127, 8150, 8177, 8205, 8226, 8252, 8278, 8306, 8328, 8351, 8385, 8440, 8462, 8522, 8552, 8581, 8609, 8636, 8665, 8696, 8725, 7461]. Notes [7484, 7508, 7533, 7556, 7582, 7606, 7636, 7664, 7693, 7719, 7745, 7773, 7798, 7825, 7856, 7883, 7905, 7930, 7952, 7978, 8001, 8026, 8049, 7476, 7488, 8498, 8594, 8625, 8651, 8687, 8699, 8731, 7544, 7613, 7670, 7696, 7812, 7830, 7862, 7888, 7910, 7935, 7964, 7982, 8007, 8030, 8052, 8075, 8119, 8193, 8214, 8229, 8255, 8331, 8839, 8776, 8799, 9625, 9648, 9671, 9690, 9714, 9739, 9767, 9839, 9862, 9895, 9922, 9950, 9979, 10008, 10067, 10091, 8831, 8859, 8887, 8920, 8946, 8973, 9003, 9056, 9082, 9109, 9137, 9162, 9215]. Notes [9273, 9303, 9330, 9355, 9385, 9413, 9442, 9474, 9499, 9556, 9582, 8778, 8801, 9530, 9558, 9584, 9603, 9627, 9650, 9674, 9692, 9716, 9741, 9769, 9791, 9816, 9841, 9864, 9897, 9924, 9952, 9981, 10010, 10037, 10069, 10093, 10119, 8833, 8889, 8922, 8948, 8975, 9005, 9031, 9058, 9084, 9111, 9139, 9164, 9193, 9217, 9248, 9275, 9305, 9331, 9356, 9387, 9415, 9444, 9476, 9501, 8789, 9181, 9197, 9221, 9280, 9390, 9479, 9543, 9629, 9655, 9678, 9696, 9794, 9883, 9902, 9985, 10123, 10820, 10491, 10595, 10734, 11038, 11057]. Notes [11075, 11303, 10617, 10173, 10193, 10887, 10915, 10935, 10964, 10992, 11023, 11121, 11142, 11165, 11187, 11209, 11231, 11250, 11297, 11321, 11347, 11404, 10216, 10237, 10270, 10301, 10331, 10354, 10374, 10401, 10431, 10460, 10516, 10541, 10562, 10587, 10613, 10633, 10658, 10681, 10752, 10775, 10795, 10816, 10839, 10860, 10149, 10175, 10797, 10818, 10841, 10862, 10889, 10917, 10937, 10966, 10994, 11026, 11047, 11072, 11096, 11123, 11144, 11167, 11189, 11211, 11233, 11252, 11274, 11299, 11323, 11349, 11375, 11406, 10195, 10218, 10239, 10272, 10303, 10333, 10356, 10376, 10403, 10434]. Notes [10462, 10484, 10518, 10543, 10564, 10589, 10615, 10635, 10660, 10683, 10706, 10731, 10754, 10777, 10155, 10197, 10306, 10492, 10596, 10664, 10685, 10780, 10880, 11100, 11134, 11149, 11277, 11380, 10843, 12449, 11437, 11459, 12087, 12105, 12129, 12152, 12174, 12202, 12223, 12245, 12271, 12292, 12312, 12334, 12356, 12377, 12406, 12426, 12446, 12466, 12489, 12510, 12530, 12552, 12575, 12595, 12617, 11486, 11513, 11530, 11559, 11584, 11615, 11636, 11662, 11687, 11713, 11733, 11758, 11787, 11806, 11825, 11846, 11892, 11919, 11942, 11961, 11990, 12015, 12041, 12064, 12493, 11439, 11461, 12066, 12089]. Notes [12107, 12131, 12154, 12176, 12204, 12225, 12247, 12273, 12294, 12314, 12336, 12358, 12379, 12408, 12428, 12448, 12470, 12491, 12532, 12554, 12577, 12597, 12619, 11488, 11515, 11532, 11561, 11586, 11617, 11638, 11665, 11689, 11715, 11735, 11760, 11789, 11808, 11827, 11848, 11869, 11894, 11921, 11945, 11964, 11992, 12018, 12043, 11442, 11553, 11573, 11589, 11629, 11753, 11861, 11900, 12081, 12122, 12195, 12412, 9047, 724, 1232, 5317, 7514]. Notions [6555]. Novel [749, 410, 742, 1055, 1287, 1264, 995, 411, 513, 1269, 717, 1190, 848, 853, 66, 1062, 1250, 608, 733, 299, 347, 188, 1243, 67, 1236, 240, 189, 690, 17, 528, 509, 326, 351, 529, 352, 861, 532, 1184, 1054, 2511, 1735, 2364, 1711, 1707, 2480, 2037, 2238, 1826, 1548, 1734, 1521, 2515, 2345, 1495, 2570, 1705, 1391, 1801, 1470, 1882, 2344, 2368, 1709, 3623, 2881, 2752, 3571, 3126, 2663, 3694, 2947, 2904, 2690, 3096, 2641, 3530, 2773, 2802, 2709, 2880, 4007, 4340, 4652, 4097, 4527, 4416, 4122]. Novel [5578, 5163, 5926, 5990, 6144, 5367, 5966, 6150, 4950, 6359, 6222, 6890, 6646, 6619, 6503, 6438, 7566, 7894, 8509, 7441, 7524, 7607, 8426, 8504, 7656, 7731, 9048, 8754, 8962, 9997, 8937, 8955, 9204, 8935, 10450, 10463, 11924, 11545, 12492]. Novelties [6, 3458, 10908]. Novelty [112, 6272, 10261]. November [4517, 3355, 3409, 4492, 5776, 7005, 7190, 9695, 10997, 12255, 7165, 7031]. Nozzle [538, 4845, 4097, 8078, 11469]. Nuisance [2607]. Nullification [2875, 5408, 5381, 5433, 7800, 9061]. Number [1148, 1121, 8893]. Nurse [6147]. Nursery [939]. Nursing [1013]. Nut [1127, 2635, 2882, 3671, 3132, 4058, 4826, 4947, 5197, 5395, 6762, 7374, 10283, 10410, 10550, 11258, 11545]. Nutrition [1023]. Nuts [6853]. Nutshell [9993].

O [11820, 9195]. Oak [3052, 3438, 4061]. Oakland [6118]. Oar [6069, 9143, 10549]. Oatmeal [214]. Obach [4409]. Obelisk [846, 1947, 1259, 7177]. Oberstadt [3531]. Object [1065, 6790]. Objects [5052]. Obscura [6337]. Observation [7496]. Observations [1294, 658, 11718, 1232]. Observatory [745, 7968, 10196, 11936, 814, 3525, 4667, 4729, 4303, 4803, 5589, 9651, 9706, 10578, 11748, 11061, 10577]. Observed [2088]. Obstructions [6149, 8286, 10124, 11489]. Obtaining [8230]. Ocarina [5692]. Occurrence [8680]. Ocean [3320, 6806, 8527, 11125, 11378, 11859, 11967, 11678, 12601, 11737, 12409, 12483]. Ochorowicz [6391]. October [3277, 4360, 5679, 6909, 9605, 10892, 12157, 7031, 8180]. Odometer [10131]. Odors [1629]. Off [2127, 3007, 2859, 4168, 5292, 5832, 6261, 6453, 7959, 10380, 11065, 10874, 12114, 11740, 2222, 6486, 8085, 9760, 8956, 12474]. Office [320, 283, 4434, 4345, 4819, 555, 2079, 3106, 3108, 3344, 4075, 4189, 6029, 5964, 5491, 6379, 6272, 8495, 11027, 11946]. Offices [3046, 5066]. Official [1906, 7775, 12021]. Officials [3500]. Offset [11240]. Offspring [3536]. Ohio [1606, 2034, 2209, 2382, 4952, 5490, 6262, 7337, 7534, 7720]. Ohm [7292]. Oil [1176, 681, 767, 1016, 1233, 349, 879, 1008, 227, 2571, 1918, 1542, 1705, 2946, 3346, 3599, 4569, 4310, 4575, 4490, 6035, 5121, 5004, 5113, 6411, 7080, 6614, 7298, 7645, 7614, 8620, 8078, 8595, 7445, 9336, 9288, 9608, 10624, 10900, 11334, 11469, 11669, 12117, 361]. Oil-Stove [1918]. Oil-Testing [349]. Oiler [6309, 6405, 6980, 9039, 12075, 5142]. Oiling [5594]. Oils [3396, 6440]. Olber [9585]. Olbers [9572]. Old [156, 2080, 3046, 3274, 3982, 4798, 3758, 4894, 5319, 7037, 6472, 7991, 9205, 8907, 8788, 10377, 159, 10832]. Oldest [9450]. One [1121, 1438, 3187, 3760, 4195, 4095, 4893, 6460, 9276, 9899, 12215, 589, 177, 230]. One-Horse [4195]. One-Third [12215]. Ontario [12179]. open [12033]. Opened [3980]. Opener [1264, 1728, 3464, 4776, 6304, 11079, 11909]. Opening [819, 1723, 2078, 1749, 2560, 3046, 3960, 7121, 8701, 11666, 12227]. Opera [11698]. Operate [7220]. Operated [7174, 9703]. Operating [377, 2217, 7423, 8333, 9089, 9959, 10291]. Operation [5115]. Operations [7273]. Operative [12585]. Operator [2880]. Operculums [2068]. Opinions [2174]. Opium [3620]. Opportunity [2107]. Opposition [7666]. Optical [2533, 5856, 7316, 6804, 11159, 12626, 12388]. Optometer [537, 1636, 6365]. Orange [4573, 11106, 2885]. Ordinary [9940]. Ordnance [8667, 9118]. Ore [462, 1186, 3301, 2897, 2779, 4186, 5148, 6955, 8231, 8360, 9018, 9796, 12500]. Oregon [7510, 7778, 7511, 7609, 11690]. Ores [2510, 1795, 10416]. Organ [2267, 5239, 8648, 12394]. Organic [876]. Organisms [5786, 7119]. Organization [6522]. Organs [7810, 7991, 10619]. Origin [878, 5457, 8003]. Original [7135]. Orlando [10110]. Orleans [5971, 6401, 6595, 6180, 6327, 6355, 6636, 7828, 2639, 2628, 8190, 6010]. Ornamental [3851, 3948, 6392]. Ornamentation [5614, 6866]. Ornamenting [885]. Ornithorhynchus [4977]. Ornoxious [9333]. Osage [4573]. Oscillating [9656]. Oscillations [4387]. Osprey [1420, 2486]. Ostrich [606, 4518, 10347, 11572]. Other [2358, 3745, 5905, 8288, 9226, 11073, 11389, 3013, 9881, 11126]. Othniel [7655]. Otis [11750]. Ottawa [3914]. Otter [996]. Otto [5930, 4195, 5878, 12215, 9804]. Our [1172, 455, 1315, 505, 1582, 1504, 2260, 1382, 3501, 2980, 6120, 5307, 7338, 8640, 9059, 10122, 9062, 9571, 9802, 9966, 9478, 10833, 12382, 9460, 5411, 5917, 7980]. Out-Door [10761]. Outlet [2036]. Outlets [1659]. Outlining [8762]. Oven [98, 799, 4777, 11646]. Overhead [5160, 9534]. Overlay [7675]. Overthrow [4026]. Owner [606]. Ox [10078]. Oxides [3644]. Oxygen [1784, 5731, 7250]. Oyster [166, 2212, 2055, 12323, 3434, 7448]. Oysters [935, 429, 753, 1235, 1274]. Ozone [955, 1629, 6908, 10079, 8236]. Ozoniferous [1596].

P [11780]. PA [3035, 4341, 732, 11899]. Pacific [4344, 888, 12264]. Pack [7095]. Package [5123, 9227, 9591, 9632, 11596]. Packages [7913]. Packers [10104]. Packing [43, 1660, 5854, 7911, 7914, 7705, 7726, 9663, 10311, 12606]. Pad [5098, 7089, 9992, 10692, 10525]. Paddle [8236, 10414, 11131, 12234]. Paddle-Belt [10414]. Padlock [8184, 11702]. Pads [6915]. Page [34, 370]. Paid [6430]. Pail [4979, 8258, 11799, 12323]. Pain [8003]. Paint [2359, 3250, 3758, 7436, 10312, 10228]. Painted [388]. Painters [8058]. Painting [8589]. Paintings [4024]. Pair [9208]. Palace [8006, 9320, 12589]. Paleontological [6442]. Palestine [4871]. Pallas [10184]. Pallet [11385, 12394]. Palm [1008, 9235, 9782]. Palmer [11884]. Palmers [7447]. Palms [11656]. Palustris} [3699]. Pan [1047, 3412, 4311, 5636, 5916, 6886, 6204, 7787, 8985, 11003, 11647, 12365]. Panama [1205, 4230, 4053, 5368, 7336, 6180, 7360, 7137, 7290, 7667, 7775, 9900, 9034, 8911, 9953, 10205, 10708]. Panda [244]. Panel [746, 2426]. Panolastite [5072]. Pans [401, 3718, 4368, 7419, 9120, 12498]. Pantelephone [743]. Panthers [5551]. Pantograph [9481]. Pantographs [11480]. Pants [984, 2538]. Paper [1087, 1331, 1748, 3503, 3578, 3674, 3192, 4100, 4673, 4658, 3963, 5439, 5619, 5736, 6164, 5705, 6166, 6649, 7226, 8388, 7513, 8987, 9395, 9049, 9540, 10823, 10949, 12476, 12133, 12546]. Paper-Bag [3192]. Papers [1319, 2309, 6686]. Parabolic [8291]. Paradise [3879, 9614, 9265, 10387, 7068]. Paradox [2969, 8599]. Paradoxes [2428]. Paradoxus} [11340]. Paraffine [5689]. Parallel [9631]. Parallels [4804]. Parasite [2901, 2857]. Parasites [1496, 1194]. Parasitic [1667]. Parasol [3648, 5366, 7469]. Parasols [4124]. Parenty [7921]. Parer [12160]. Paris [1561, 6442, 11971, 323, 2560, 2053, 3556, 4667, 4876, 4819, 5369, 5688, 5949, 7399, 6994, 6340, 8295, 8216, 8107, 9846, 9517, 8840, 9209, 10545, 11097, 10985, 10520, 11506, 12612, 11934, 11750, 11608, 11953, 12100, 12125, 12146, 12262, 12286, 12525, 12540, 12561, 12634]. Parisian [2682, 2791]. Park [4729, 8443, 8650, 2942, 1874, 1947, 3086, 6546, 7313, 9345, 8978, 10334, 11147, 10273]. Parkman [12473]. Parliament [7304]. Parlor [5239]. Parrot [919]. Parsons [9887]. Part [6260]. Parthenogenesis [2675]. Parts [10742]. Party [5268]. Parvo [8850]. Pascal [10699]. Passage [654, 4307, 10465]. Passenger [684, 3096, 5747, 6471, 10389, 11983, 12458]. Passengers [5601]. Past [555, 3344, 4189]. Pasteur [7230, 5810, 7163]. Pasting [9094]. Patent [602, 867, 1175, 500, 5, 30, 57, 84, 108, 130, 203, 223, 282, 339, 499, 522, 600, 629, 653, 706, 731, 762, 791, 903, 1201, 1313, 1347, 1049, 984, 2358, 1378, 1481, 2415, 2437, 2466, 2496, 2526, 2555, 1503, 1532, 1556, 1579, 1603, 1626, 1649, 1818, 1840, 1869, 1903, 1932, 2052, 2076, 2100, 2120, 2143, 2165, 2208, 2233, 2259, 2281, 2308, 2356, 1960, 1380, 2188, 1961, 2604, 3501, 2602, 2627, 3239, 3267, 3292, 3319, 3342, 3366, 3388, 3406, 3429, 3449, 3475, 3496, 3521, 3544]. Patent [3566, 3594, 3616, 3639, 3666, 3689, 3711, 2654, 2681, 2701, 2723, 2745, 2767, 2790, 2810, 2834, 2852, 2872, 2896, 2919, 2937, 2955, 3042, 3059, 3082, 3104, 3121, 3139, 3161, 3185, 3211, 2629, 3212, 3433, 4530, 4307, 3733, 3913, 4742, 4768, 4793, 4818, 4839, 4865, 4891, 3937, 3959, 3979, 4003, 4112, 4138, 4162, 4183, 4207, 4229, 4254, 4273, 4301, 4328, 4356, 4381, 4407, 4433, 4458, 4486, 4514, 4541, 4693, 4717, 4213, 4140, 5434, 4940, 5137, 5951, 4913, 4936, 5770, 5795, 5843, 5865, 5892, 5915]. Patent [5941, 5961, 5985, 6027, 6054, 6083, 6107, 6132, 6159, 4962, 4991, 5015, 5041, 5062, 5088, 5111, 5135, 5158, 5182, 5207, 5255, 5281, 5330, 5379, 5406, 5430, 5456, 5484, 5513, 5535, 5588, 5610, 5632, 5138, 6179, 6196, 6803, 6826, 6850, 6874, 6904, 6928, 6950, 6971, 7001, 7028, 7051, 7076, 7106, 7134, 7188, 7217, 7240, 7265, 7286, 7312, 7335, 7359, 7386, 6221, 6248, 6270, 6297, 6326, 6354, 6377, 6400, 6424, 6451, 6470, 6495, 6520, 6543, 6571, 6592, 6613, 6634, 6659, 6681, 6705, 6730, 6753, 6779]. Patent [7800, 7411, 7433, 8072, 8127, 8150, 8177, 8205, 8226, 8252, 8278, 8306, 8328, 8351, 8385, 8440, 8462, 8522, 8552, 8581, 8609, 8636, 8665, 8696, 8725, 7461, 7484, 7508, 7533, 7556, 7582, 7606, 7636, 7664, 7693, 7719, 7745, 7773, 7798, 7825, 7856, 7883, 7905, 7930, 7952, 7978, 8001, 8026, 8049, 8555, 9166, 9819, 9446, 9928, 8778, 8801, 9530, 9558, 9584, 9603, 9627, 9650, 9674, 9692, 9716, 9741, 9769, 9791, 9816, 9841, 9864, 9897, 9924, 9952, 9981, 10010, 10037, 10069, 10093, 10119, 8833, 8889]. Patent [8922, 8948, 8975, 9005, 9031, 9058, 9084, 9111, 9139, 9164, 9193, 9217, 9248, 9275, 9305, 9331, 9356, 9387, 9415, 9444, 9476, 9501, 10096, 9061, 10569, 10241, 10637, 10154, 10149, 10175, 10797, 10818, 10841, 10862, 10889, 10917, 10937, 10966, 10994, 11026, 11047, 11072, 11096, 11123, 11144, 11167, 11189, 11211, 11233, 11252, 11274, 11299, 11323, 11349, 11375, 11406, 10195, 10218, 10239, 10272, 10303, 10333, 10356, 10376, 10403, 10434, 10462, 10484, 10518, 10543, 10564, 10589, 10615, 10635, 10660, 10683, 10706, 10731, 10754, 10777, 10377, 11982, 11726]. Patent [12229, 12387, 11439, 11461, 12066, 12089, 12107, 12131, 12154, 12176, 12204, 12225, 12247, 12273, 12294, 12314, 12336, 12358, 12379, 12408, 12428, 12448, 12470, 12491, 12532, 12554, 12577, 12597, 12619, 11488, 11515, 11532, 11561, 11586, 11617, 11638, 11665, 11689, 11715, 11735, 11760, 11789, 11808, 11827, 11848, 11869, 11894, 11921, 11945, 11964, 11992, 12018, 12043, 229, 34, 630, 982, 86, 342, 31, 555, 553, 603, 1628, 2079, 3106, 2907, 2875, 3108, 2656, 3548, 3344, 3962, 4075, 4345, 4189, 4799, 6029, 5964, 4996, 5018, 5381, 5433, 5017]. Patent [5541, 5463, 5043, 5491, 4968, 6379, 6272, 7858, 8353, 7932, 7535, 8495, 9059, 8804, 9333, 9818, 10755, 10436, 10405, 11027, 11191, 11946]. Patented [2154, 2528, 3424, 3777, 9867, 11163, 11184, 11206, 11228, 11434, 11457, 12061, 12085, 12103, 12148, 12170, 12172, 12199, 12220, 12242, 12268, 12289, 12309, 12332, 12354, 12375, 12403, 12424, 12443, 12464, 12486, 12508, 12527, 12550, 12572, 12593, 12615, 11484, 11511, 11554, 11557, 11581, 11612, 11633, 11660, 11684, 11710, 11731, 11756, 11784, 11804, 11823, 11844, 11865, 11889, 11917, 11940, 11959, 11988, 12013, 12038]. Patentee [2835]. Patentees [1982, 3063]. Patents [133, 736, 725, 756, 775, 858, 917, 976, 1025, 1060, 1114, 1139, 1328, 1171, 369, 5, 30, 57, 84, 108, 130, 203, 223, 282, 339, 499, 522, 600, 629, 653, 706, 731, 762, 791, 903, 1201, 1313, 1347, 480, 370, 168, 250, 311, 368, 905, 195, 360, 390, 471, 494, 594, 993, 1277, 1304, 2054, 2548, 2572, 1389, 1416, 1664, 1894, 1924, 1946, 1968, 2403, 2187, 1505, 1794, 2006, 2396, 1378, 1481, 2415, 2437, 2466, 2496, 2526, 2555, 1503, 1532, 1556, 1579, 1603, 1626, 1649]. Patents [1818, 1840, 1869, 1903, 1932, 2052, 2076, 2100, 2120, 2143, 2165, 2208, 2233, 2259, 2281, 2308, 2356, 1446, 1516, 1614, 1635, 1685, 1753, 2244, 2234, 1905, 2794, 2619, 2674, 2858, 2965, 3175, 3198, 3223, 3251, 3303, 3574, 3683, 2999, 2605, 2602, 2627, 3239, 3267, 3292, 3319, 3342, 3366, 3388, 3406, 3429, 3449, 3475, 3496, 3521, 3544, 3566, 3594, 3616, 3639, 3666, 3689, 3711, 2654, 2681, 2701, 2723, 2745, 2767, 2790, 2810, 2834, 2852, 2872, 2896, 2919, 2937, 2955, 3042, 3059, 3082, 3104, 3121]. Patents [3139, 3161, 3185, 3211, 2711, 2754, 3719, 3433, 3084, 3856, 3988, 4057, 4085, 4520, 4753, 4026, 3733, 3913, 4742, 4768, 4793, 4818, 4839, 4865, 4891, 3937, 3959, 3979, 4003, 4112, 4138, 4162, 4183, 4207, 4229, 4254, 4273, 4301, 4328, 4356, 4381, 4407, 4433, 4458, 4486, 4514, 4541, 4693, 4717, 4011, 4035, 5408, 4942, 4967, 5234, 5646, 5668, 5737, 5830, 5906, 6012, 6070, 5044, 5981, 4993, 5064, 5159, 4913, 4936, 5770, 5795, 5843, 5865, 5892, 5915, 5941, 5961, 5985, 6027, 6054, 6083, 6107, 6132]. Patents [6159, 4962, 4991, 5015, 5041, 5062, 5088, 5111, 5135, 5158, 5182, 5207, 5255, 5281, 5330, 5379, 5406, 5430, 5456, 5484, 5513, 5535, 5588, 5610, 5632, 4969, 5893, 5975, 6189, 6560, 6720, 7323, 6320, 7029, 6179, 6196, 6803, 6826, 6850, 6874, 6904, 6928, 6950, 6971, 7001, 7028, 7051, 7076, 7106, 7134, 7188, 7217, 7240, 7265, 7286, 7312, 7335, 7359, 7386, 6221, 6248, 6270, 6297, 6326, 6354, 6377, 6400, 6424, 6451, 6470, 6495, 6520, 6543, 6571, 6592, 6613, 6634, 6659, 6681, 6705, 6730, 6753, 6779]. Patents [7465, 7561, 8063, 8087, 8194, 8267, 8309, 8450, 8499, 8597, 8652, 8727, 7722, 7411, 7433, 8072, 8127, 8150, 8177, 8205, 8226, 8252, 8278, 8306, 8328, 8351, 8385, 8440, 8462, 8522, 8552, 8581, 8609, 8636, 8665, 8696, 8725, 7461, 7484, 7508, 7533, 7556, 7582, 7606, 7636, 7664, 7693, 7719, 7745, 7773, 7798, 7825, 7856, 7883, 7905, 7930, 7952, 7978, 8001, 8026, 8049, 8098, 8415, 8493, 7451, 7845, 7721, 8815, 8902, 8984, 9182, 9636, 9870, 9901, 8855, 9381, 9409, 8854, 8778, 8801, 9530, 9558, 9584]. Patents [9603, 9627, 9650, 9674, 9692, 9716, 9741, 9769, 9791, 9816, 9841, 9864, 9897, 9924, 9952, 9981, 10010, 10037, 10069, 10093, 10119, 8833, 8889, 8922, 8948, 8975, 9005, 9031, 9058, 9084, 9111, 9139, 9164, 9193, 9217, 9248, 9275, 9305, 9331, 9356, 9387, 9415, 9444, 9476, 9501, 10733, 10151, 10149, 10175, 10797, 10818, 10841, 10862, 10889, 10917, 10937, 10966, 10994, 11026, 11047, 11072, 11096, 11123, 11144, 11167, 11189, 11211, 11233, 11252, 11274, 11299, 11323, 11349, 11375, 11406, 10195, 10218, 10239, 10272, 10303, 10333, 10356].

Patents [10376, 10403, 10434, 10462, 10484, 10518, 10543, 10564, 10589, 10615, 10635, 10660, 10683, 10706, 10731, 10754, 10777, 11641, 11525, 11725, 11853, 12009, 11640, 11439, 11461, 12066, 12089, 12107, 12131, 12154, 12176, 12204, 12225, 12247, 12273, 12294, 12314, 12336, 12358, 12379, 12408, 12428, 12448, 12470, 12491, 12532, 12554, 12577, 12597, 12619, 11488, 11515, 11532, 11561, 11586, 11617, 11638, 11665, 11689, 11715, 11735, 11760, 11789, 11808, 11827, 11848, 11869, 11894, 11921, 11945, 11964, 11992, 12018, 12043, 8369, 8716, 254, 1025, 2362, 2220, 3408, 5723, 6755, 6271]. Patents [8386, 8949, 10732, 11587, 11534, 12019]. Paths [7763]. Patients [7230]. Patrol [1580]. Pattern [136, 2591, 9565]. Patterns [5437, 5702, 10129]. Paul [9320, 8006]. Paul’s [3122]. Pavement [1651, 11505]. Pavements [986]. Pavilion [10985]. Pavilions [5472]. Paving [11643]. Pay [736, 3915]. Paying [3452, 11948]. Peacemaker [8444]. Peach [9019]. Peacock [4317]. Peacocks [7737]. Pear [934]. Pearl [6287, 10555, 12079]. Peas [6553]. Peculiar [5487]. Peculiarities [2069]. Pedo [608]. Pedo-Motor [608]. Pedro [5853]. Peerless [561, 3766]. Peg [9797, 9097]. Pegging [2708, 9799]. Peirce [908]. Pelayo [11551]. Pelecanoides} [3792]. Pelican [3991]. Pen [1053, 485, 240, 2173, 4658, 6694, 8986, 9752, 11199, 11723, 12139]. Penberthy [9456]. Pencil [2009, 1447, 2988, 5736, 11199, 11693]. Pencils [11450]. Pendant [8500, 8763, 12347]. Pendulum [3145, 6506, 8186, 9292, 7096]. Penelope [6313]. Penguin [894, 9008, 11180]. Pennsylvania [1355, 228, 6929, 10636, 12493, 2423]. Pens [4411, 6139]. People [12535]. Percentograph [5495]. Percha [780]. Percolation [1881]. Père [10781]. Perfidy [1628]. Perforate [12542]. Perforating [1564, 7675]. Perforation [6930]. Perforator [10044]. Performance [6809]. Perils [3939]. Periophthalmus [5499]. Peripatus [147, 3416]. Permanent [1583]. Permeability [11619]. Permutation [5171]. Perpetual [1680, 1708, 1725, 1704, 1746, 9518, 9096, 10022]. Perplexed [7859]. Perreaux [3375, 4113]. Personal [1, 625, 649, 674, 702, 728, 759, 788, 816, 838, 872, 899, 927, 950, 979, 1003, 1028, 1044, 1068, 1091, 1118, 1143, 1168, 1197, 1226, 27, 1278, 1307, 1310, 1343, 53, 80, 105, 128, 150, 170, 200, 219, 247, 278, 307, 335, 364, 393, 421, 446, 473, 496, 518, 545, 570, 596, 1374, 1998, 2028, 2049, 2073, 2096, 2116, 2140, 2162, 2184, 2205, 2229, 2256, 2278, 2304, 2326, 2353, 2375, 2391, 2412, 2433, 2462, 2493, 2523, 2551, 1397, 1427, 1454, 1478, 1499, 1528, 1552, 1575]. Personal [1600, 1622, 1645, 1673, 1694, 1718, 1742, 1764, 1789, 1815, 1837, 1865, 1899, 1928, 1956, 1978, 2578, 2598, 3136, 3160, 3183, 3207, 3237, 3264, 3289, 3315, 3338, 3363, 3385, 3402, 3426, 3448, 3474, 3497, 3520, 3545, 3562, 3590, 3613, 3635, 3662, 3686, 3707, 2624, 2650, 2678, 2698, 2719, 2741, 2764, 2787, 2806, 2830, 2848, 2869, 2893, 2916, 2934, 2950, 2976, 2995, 3016, 3039, 3056, 3078, 3102, 3118, 3731, 3755, 4324, 4352, 4378, 4404, 4430, 4455, 4483, 4512, 4537, 4562, 4594, 4618, 4644, 4664, 4691, 4713, 4738]. Personal [4764, 4789, 4814, 4835, 4861, 4887, 3774, 3800, 3819, 3841, 3863, 3887, 3909, 3936, 3955, 3975, 3999, 4020, 4047, 4069, 4091, 4110, 4135, 4159, 4180, 4204, 4226, 4250, 4269, 4297, 4909, 4933, 5509, 5531, 5559, 5585, 5607, 5629, 5652, 5673, 5698, 5719, 5743, 5767, 5792, 5840, 5862, 5889, 5912, 5938, 5958, 5982, 5998, 6023, 6051, 6080, 6104, 6129, 4958, 6156, 4987, 5011, 5038, 5059, 5084, 5107, 5132, 5155, 5178, 5204, 5227, 5252, 5277, 5302, 5327, 5350, 5376, 5403, 5427, 5452, 5481, 6176, 6193, 6776, 6800, 6823]. Personal [6847, 6871, 6900, 6925, 6947, 6968, 6998, 7025, 7048, 7073, 7103, 7131, 7158, 7185, 7213, 7236, 7262, 7283, 7309, 7332, 7356, 7383, 6218, 6245, 6267, 6294, 6323, 6351, 6374, 6396, 6421, 6448, 6466, 6492, 6517, 6540, 6568, 6588, 6610, 6631, 6656, 6678, 6702, 6726, 6750, 6250, 7408, 7430, 8046, 8070, 8095, 8124, 8147, 8173, 8201, 8222, 8249, 8275, 8303, 8325, 8348, 8382, 8412, 8437, 8459, 8490, 8518, 8549, 8578, 8606, 8633, 8661, 8693, 8722, 7458, 7481, 7505, 7530, 7553, 7579, 7603, 7633, 7661, 7689, 7716, 7742]. Personal [7770, 7795, 7822, 7852, 7880, 7902, 7927, 7949, 7975, 7998, 8023, 8752, 8775, 9473, 9498, 9527, 9554, 9581, 9601, 9624, 9646, 9670, 9689, 9712, 9737, 9765, 9789, 9814, 9837, 9860, 9893, 9920, 9948, 9977, 10006, 10034, 10065, 10089, 8798, 10117, 8830, 8858, 8886, 8919, 8945, 8972, 9001, 9029, 9055, 9081, 9107, 9135, 9160, 9191, 9214, 9244, 9271, 9302, 9329, 9354, 9384, 9412, 9441, 10147, 10172, 10793, 10815, 10838, 10859, 10886, 10913, 10933, 10962, 10990, 11022, 11044, 11069, 11094, 11119, 11141, 11164, 11185, 11207, 11229, 11248, 11271]. Personal [11295, 11320, 11345, 11373, 10191, 11403, 10214, 10235, 10269, 10299, 10329, 10353, 10373, 10399, 10430, 10458, 10482, 10514, 10539, 10560, 10585, 10611, 10632, 10657, 10680, 10704, 10728, 10750, 10773, 11435, 11458, 12039, 12086, 12104, 12128, 12151, 12173, 12201, 12221, 12243, 12269, 12290, 12310, 12333, 12355, 12376, 12405, 12425, 12444, 12465, 12488, 12509, 12529, 12551, 12574, 11485, 12594, 12616, 11512, 11529, 11558, 11582, 11613, 11634, 11661, 11685, 11712, 11732, 11757, 11786, 11805, 11824, 11845, 11866, 11890, 11918, 11941, 11960, 11989, 12014, 12062]. Peru [12180]. Peruke [2820]. Peruke-Horned [2820]. Peruvian [1348]. Pest [10832]. Pet [9360]. Peter [3868, 7289]. Petrel [1064]. Petroleum [935, 1156, 1939, 1703, 1984, 2121, 3021, 4464, 4413, 4304, 4894, 5757, 4976, 5267, 6862, 8293, 8508, 8595, 8582, 11133, 11707, 3739]. Pettenkofer [4951]. Phantasmoscope [2316]. Phantom [6009]. Pharaoh [1872]. Phase [6906]. Phases [5623]. Pheasant [4782, 5419]. Phelps [6743]. Phenomena [1204, 10176]. Phenomenon [735, 5490, 6975]. Philadelphia [5722, 877, 5771, 5796, 5878, 5746, 5732, 6731, 10940, 12120, 12565, 2243]. Philip [11351]. Philosophy [737]. Phoeniceus [3699]. Phonograph [11315, 10490, 12250]. Phos [3013]. Phos-Animals [3013]. Phosphorescence [1938]. Phosphorescent [2452, 3013]. Phosphorus [1360]. Phote [6593]. Photo [179, 2250, 5471, 6859, 9761]. Photo-Engraving [179]. Photo-Enlarging [5471]. Photo-Graphy [2250]. Photo-Micrographic [9761]. Photo-Tricycle [6859]. Photograph [5144, 11549]. Photographic [503, 2508, 4128, 5215, 6792, 6863, 6888, 6921, 6989, 7013, 7040, 7079, 7109, 7141, 7176, 7191, 7221, 6441, 7476, 7488, 8498, 8594, 8625, 8651, 8687, 8699, 8731, 7514, 7544, 7613, 7670, 7696, 7812, 7830, 7862, 7888, 7910, 7935, 7964, 7982, 8007, 8030, 8052, 8075, 8119, 8193, 8214, 8229, 8255, 8331, 9512, 9431, 8789, 9181, 9197, 9221, 9280, 9390, 9479, 9543, 9629, 9655, 9678, 9696, 9794, 9883, 9902, 9985, 10123, 9761, 10851, 10155, 10197, 10306, 10492, 10596, 10664, 10685, 10780, 10880, 11100, 11134, 11149, 11277, 11380, 12133, 11442]. Photographic [11553, 11573, 11589, 11629, 11753, 11861, 11900, 12081, 12122, 12195, 12412]. Photographing [1591, 3174, 5291, 6257, 1804]. Photography [2083, 2046, 3603, 3748, 4127, 6073, 5318, 6960, 7569, 9292, 10842, 12438, 12524, 10623]. Photometer [6593]. Photophone [765, 822, 1379, 1432]. Photophonic [1633]. Phyllirhoe [693]. Phylloxera} [49]. Physical [735, 1404, 1432, 4029, 8553, 12431, 10390]. Physician [1652]. Physicians [1700]. Physics [2474, 3929, 10881, 11089, 11132, 11179, 11203, 11216, 11237, 11264, 11288, 11339, 11395, 11452, 11728, 11935, 11957, 12036, 12059]. Physiological [6311]. Pianets [4006]. Pianists [12181]. Piano [6338, 7276, 7325, 12072]. Piano-Forte [7325]. Pianofortes [7991]. Pianos [5925, 6600, 10619, 11903]. Pick [439, 2337]. Picker [2515, 4802, 6425, 8157, 9257, 12134]. Pickerel [3949, 668]. Pickering [8557]. Picket [8735]. Pickett [1607, 1584]. Picking [665, 2609]. Pictet [3529, 3716]. Picture [1210, 93, 2011, 8955, 9284]. Pictures [503, 510, 1583, 9879]. Pie [7063]. Piece [442, 5520, 12390]. Piefke [3923]. Piercing [12023]. Pierpont [10111]. Piers [2730]. Pig [2399]. Pigeon [4421]. Pigs [1559]. Pike [3949]. Pikes [3379]. Pilatus [12359, 9996]. Pile [4526, 4060, 3829, 5357, 8079]. Pilgrim [4081]. Pill [798]. Pillar [2927]. Pillow [11950]. Pin [5469, 8783]. Pinch [7727, 11814]. Pine [852, 5614, 9184]. Pink [4719]. Pinking [1418]. Pins [4139, 6405, 6426, 8447, 10040]. Pioneer [9693, 2502]. Pipe [1339, 191, 4529, 4445, 4447, 5265, 5446, 5287, 5798, 6005, 7231, 6617, 7225, 8613, 8649, 7522, 7705, 9368, 9566, 10016, 9015, 9198, 10445, 11155, 10871, 11084, 10736, 10763, 10688, 12005, 12304, 12624, 11495, 11979]. Pipes [1385, 1768, 3435, 4756, 3740, 6149, 5968, 5905, 5392, 6333, 6782, 7887, 8806, 11174, 11421, 10487, 11764]. Pirates [10756]. Piston [2503, 2126, 2661, 3508, 7914, 7726, 10071, 8786, 10311, 12606]. Pitch [2172, 7244, 9657]. Pitch-Board [2172]. Pitcher [9183]. Pitching [8729, 8624]. Pitman [6596, 9480]. Pits [1298, 12238]. Pittsburg [732, 1365, 3035, 4341, 8027]. Pivot [6913]. Pivoted [4442]. Pivoting [6171]. Place [7313, 10109, 10545]. Placing [3961, 8844]. Plague [3452, 8908, 11427]. Plain [2243]. Plan [479, 111, 3965, 5136, 6793, 11644]. Planchette [6366]. Plane [1878, 3351, 5214, 11335]. Planer [746, 3146, 2735, 8995]. Planet [3568, 3131, 5592]. Planets [2691, 3619, 3453, 2876, 3123, 2771, 3087, 3355, 3277, 3142, 3073, 3826, 4774, 4624, 4117, 4079, 4724, 3919, 4234, 4492, 4360, 4900, 5026, 5895, 5442, 5257, 5356, 5988, 5140, 5548, 5776, 5679, 6162, 6274, 7324, 7192, 6640, 6523, 6783, 6413, 6758, 7005, 6909, 7390, 10012, 9845, 9695, 9605, 10121, 10274, 11276, 10663, 10519, 10799, 10358, 10779, 10997, 10892, 11413, 11537, 12517, 12385, 11927, 11441, 11793, 11872, 12472, 11464, 11642, 12255, 12157, 12622]. Planing [6620, 8014, 7803, 9831]. Plans [7287, 8835]. Plant [142, 174, 2190, 3676, 2757, 5219, 5121, 6603, 6742, 8702, 9183, 11169, 11157, 11547, 11781]. Plant-Trimming [2757]. Planter [183, 4147, 4282, 5069, 6090, 5569, 6407, 6977, 7011, 8283, 8614, 8872, 9704, 9725, 9882, 9341, 10946, 11498, 12231, 12538]. Planters [7542, 10500]. Plantes [4127]. Planting [5852, 9750, 9820]. Plants [590, 971, 1726, 1596, 3644, 3895, 6711, 9149, 10602, 11148, 11471]. Plasterer [9454]. Plastering [114, 5413]. Plate [215, 233, 921, 2270, 2481, 2735, 7086, 8014, 7706, 10017, 9576, 8893, 10014, 9607, 8952, 10417, 11053, 10875, 10421, 12481, 12161]. Plate-glass [2270]. Plated [10108]. Plates [2312, 4777, 7596, 12133, 11013, 6975, 11745, 8086]. Platform [4315, 8080, 8671, 7837, 8866, 11746]. Platforms [8767]. Plating [726, 1383, 2063, 3023, 12192]. Platinum [10438]. Player [4197]. Players [7325]. Plea [1704, 12410]. Pleasure [713, 11337]. Pliers [9367]. Pliocene [3218]. Plongeon [5090, 5268, 5089, 5122, 5168, 5220, 5243]. Plot [4993]. Plow [321, 47, 3245, 2860, 3025, 2756, 4361, 4281, 4677, 4493, 5364, 5493, 5388, 5924, 5390, 6064, 4920, 7400, 6814, 6529, 6935, 8009, 7698, 7490, 7516, 9285, 9338, 9986, 8851, 9723, 11415, 12031, 11645, 11592]. Plows [10313]. Plug [2775, 6036, 7831, 9851]. Plugs [4461, 8844, 10207]. Plumber [3696, 7752, 8811]. Plume [5005]. Plunge [10107]. Plunger [6409, 11478]. Plungers [10097]. Plunges [2514]. Pneumatic [1331, 990, 1270, 2264, 6119, 5822, 4919, 5949, 7002, 6300, 6233, 6717, 7810, 8330, 8564, 8861, 10094, 9807, 9613, 9465, 12226, 12631, 12256]. Pneumatics [8191, 8213, 8239]. Pneumonia [6227, 6908]. Poaching [11571]. Pocket [2295, 2636, 4754, 3966, 5635, 6410, 7091, 10282, 11625, 12562, 12432, 11450, 10955]. Point [47, 9340, 732]. Pointed [6139, 6694]. Points [11415]. Poison [3270, 11471]. Poisoned [3100, 4743]. Poisoning [5727]. Poisonous [401, 2441, 6059, 6060, 12483]. Poke [9148]. Polar [454, 9022]. Polarization [1747]. Polarized [1444, 8567, 8264, 8337, 8596, 8623, 8050, 9128]. Pole [4423, 6141, 7014, 8813, 8990]. Poles [7440]. Polia [997]. Police [1012, 1853, 1580]. Policy [7]. Polish [7436]. Polished [4870]. Polishing [115, 1158, 8448]. Politics [4969]. Polyparium [11428]. Pond [1759, 9831]. Ponds [1042]. Pons [5514, 5186, 5486, 5550]. Pontoon [11048]. Poodle [940]. Pool [7091]. Poor [12460]. Pop [11982]. Population [480, 2333]. Porcelain [593, 442, 2199, 1782, 2044, 2518, 5284]. Porcupine [23, 4633, 9293, 9099]. Porous [10827]. Port [111]. Portable [718, 579, 799, 691, 561, 2288, 1847, 2536, 2264, 1539, 2695, 2909, 3071, 3276, 2839, 3553, 3248, 3368, 4034, 4084, 6113, 5316, 6058, 5589, 6532, 7275, 6441, 7173, 7081, 7857, 8741, 9232, 10015, 9724, 9169, 11426, 10762, 10785, 12416, 10605]. Porter [5590]. Portraits [6818, 10851]. Portugal [8766]. Position [5485, 9845, 11082, 10274, 11276, 11148, 10663, 10519, 10799, 10358, 10779, 10997, 10892, 11413, 11537, 12517, 12385, 11927, 11441, 11793, 11872, 12472, 11464, 11642, 12255, 12157, 12622]. Positions [10012, 9695, 9605, 10121].

Positive [3143, 7592]. Possibilities [2311]. Possible [3526, 11236]. Post [716, 1682, 3810, 3791, 5118, 5070, 5172, 6148, 7084, 7346, 7293, 6689, 6230, 7417, 8114, 8079, 9880, 9093, 10570]. Postage [3217, 3713, 3782]. Postal [3782, 4150, 6300, 10865, 10649, 11898]. Posts [1634, 2603, 7616, 9871]. Pot [1451, 6339, 8426, 9038, 11773]. Potassium [1298]. Potato [217, 72, 3946, 4277, 6407, 9606, 11601, 11498]. Potatoes [2131]. Potter [819]. Pottery [1315, 2018]. Pottsville [2423]. Pouched [3969]. Poughkeepsie [9357, 10697]. Poultry [7867]. Pounder [5666]. Pounds [5231]. Pourtales [811]. Powder [1784, 2794, 5169, 9881, 10524]. Powders [2886]. Powell [11366]. Power [1153, 965, 181, 1131, 1677, 2363, 1780, 2144, 2556, 2588, 2792, 3649, 2970, 4708, 4275, 4491, 4342, 4195, 4747, 4775, 4287, 4329, 4547, 4650, 3867, 6146, 5362, 6034, 5897, 5547, 5411, 5183, 4972, 5644, 6406, 7396, 6731, 7220, 6561, 6408, 6273, 6672, 7840, 8526, 7783, 7868, 8216, 8638, 9343, 10103, 9866, 9115, 10739, 11194, 11011, 11973, 12215, 12155]. Powerful [2734, 7730, 10346, 12632]. Practicable [7137]. Practical [6043, 6548, 6263, 8100, 7512, 9613, 12178]. Practice [4570, 4938, 11353]. Prairie [8834]. Prairies [3485]. Prall [1381]. Pratt [10920, 11520]. Praxinoscope [3555]. Precautions [3807]. Precious [7954]. Precipitating [1062]. Precipitation [11088]. Prehensile [7097]. Prehistoric [10984]. Premium [9751]. Preparation [2481, 2873, 10425]. Prepared [2335]. Preparing [1443, 4462]. Present [2046, 2250, 4570, 4938, 9142, 11519]. Preservation [2033, 1768, 7177]. Preservatives [406]. Preserver [639, 265, 4945, 7788, 9721]. Preserving [689, 2603, 4875, 6057, 9867, 10526]. President [11691, 1909, 2122, 8495]. Press [1267, 1917, 1944, 3249, 3247, 3127, 3553, 3051, 3396, 4086, 4260, 4052, 4240, 6163, 6042, 6098, 5729, 5973, 5417, 6525, 6615, 7674, 7757, 7651, 7467, 9491, 10419, 11327, 10721, 12536, 11577]. Presses [181, 666, 2588, 5269, 8402, 7868, 10044, 10307, 10739, 10528]. Pressure [661, 3481, 3050, 2925, 3985, 4276, 4383, 4038, 5019, 6056, 5713, 5777, 5231, 7291, 6232, 6695, 7728, 8564, 9587, 12535, 8066]. Pressures [11150]. Prestidigitation [9021]. Prevailing [2940]. Prevent [778]. Preventing [2108]. Prevention [1122, 685, 657, 1286, 3272, 5774]. Preventives [3085]. Prey [3416]. Price [487, 12186]. Primary [9745]. Primigenius [11136, 9728]. Prince [9266]. Princeton [3240]. Principle [7987, 8879, 10323, 10344, 8268]. Principles [1657]. Printer [6738, 9144, 11743]. Printing [798, 1944, 1821, 4052, 5269, 7301, 8402, 7651, 10044, 9989, 9309, 10528, 10419, 11549, 12193]. Prints [4880]. Prismatic [7872]. Prison [3715]. Private [1042, 10378]. Prize [5747, 10506, 291]. Proa [44]. Probable [5043]. Probably [8353]. Problem [206, 452, 1822, 4185, 6507, 8558, 12110]. Problematic [5592]. Problems [2684, 2607]. Process [677, 921, 2472, 3529, 2924, 4145, 4769, 6057, 5386, 5242, 8584, 7509, 7612, 8977, 12587, 9544, 12034]. Processes [2063, 9116, 9072]. Procession [5855]. Prodigy [10386]. Produced [910, 9152]. Producer [3876]. Produces [3357]. Producing [503, 12258, 10135, 11978]. Product [319, 3504]. Production [2805, 3066, 5731, 5339, 8666, 10941]. Products [945, 9955, 12589]. Proechidna [10138]. Prof [8557]. Prof. [373, 10719]. Professional [12178]. Professor [1104, 1845, 7055]. Profilograph [1237]. Profit [4699]. Profitable [7189]. Profits [710, 3107]. Progress [285, 1075, 879, 1154, 1049, 1284, 1581, 1974, 1841, 1842, 1698, 1745, 1960, 3650, 3577, 3498, 3367, 3066, 3892, 4165, 4212, 4966, 5114, 6300, 7136, 6743, 8496, 7748, 8556, 10490, 11518, 11995, 12601, 3434]. Project [2380, 5725, 11908, 7219]. Projected [3320, 11256, 8599]. Projectile [733]. Projecting [3555]. Projects [4210]. Prominence [5167]. Promising [969]. Promoting [4678]. Promotion [9278]. Promotions [2079]. Proof [1016, 4010, 9171]. Propagation [2606]. Propelled [3415, 11454, 11703]. Propeller [796, 40, 6165, 5576, 6283, 8365, 8040, 9657, 9396, 11337, 10974, 11833, 12558, 5440]. Propeller-Shaft [12558]. Propellers [7302]. Propelling [2703, 5990, 12630]. Property [2358, 5893, 5537, 7004, 9792]. Proportion [480]. Proportions [5436]. Proposal [6805]. Proposed [656, 31, 1008, 604, 2006, 2394, 2386, 2656, 3524, 2628, 2629, 3548, 4752, 5410, 6521, 7487, 7722, 8118, 7434, 8555, 10890, 10939, 10407, 12398, 11492, 11465]. Propulsion [3759, 4919, 8527, 8582, 8756, 10901, 12521, 11507]. Prospects [502, 1007, 2046, 2250, 3329, 4968, 7361, 11490, 11517]. Prospectus [9382]. Prosperity [905, 5488]. Protecting [4310, 6953, 8937, 9754, 10364]. Protection [1095, 1206, 132, 2639, 3641, 2793, 9792, 8977, 9141, 12627, 881]. Protective [963, 8711]. Protector [884, 1808, 3247, 2687, 5706, 6742, 6559, 6333, 8570, 8702, 10109, 9457, 9875, 8782, 10338, 11856, 12366, 11627, 12302, 12546]. Protests [5091]. Protractor [6956, 12564]. Providence [753]. Pruning [143, 7222, 6392, 7493, 10290]. Public [1630, 2794, 2605, 4410, 8004, 10407, 12177, 2179]. Publication [12445]. Publications [2, 54, 1198, 1227, 1344, 81, 151, 171, 221, 279, 308, 336, 365, 394, 422, 447, 474, 519, 546, 597, 626, 650, 703, 839, 900, 1004, 1069, 1093, 1144, 1279, 1375, 1398, 1428, 1501, 1530, 1553, 1577, 1624, 1647, 1695, 1791, 1867, 1901, 1930, 1958, 2000, 2098, 2118, 2231, 2306, 2328, 2378, 2434, 2464, 2552, 2600, 2652, 2721, 2743, 2807, 2832, 2849, 2951, 2996, 3080, 3208, 3316, 3339, 3403, 3447, 3472, 3494, 3563, 3591, 3636, 3663, 3708, 3839, 3842, 4739, 4765, 4790, 4815, 4836]. Publications [4862, 4888, 3864, 3888, 3910, 3956, 3976, 4000, 4048, 4070, 4136, 4251, 4270, 4299, 4325, 4353, 4377, 4403, 4429, 4454, 4482, 4510, 4538, 4619, 4645, 4689, 4911, 4932, 5766, 5791, 5816, 5861, 5888, 5911, 5937, 5999, 6024, 6050, 6079, 4959, 4988, 5012, 5037, 5058, 5085, 5108, 5179, 5203, 5251, 5278, 5326, 5351, 5402, 5426, 5453, 5480, 5510, 5532, 5560, 5672, 5697, 5718, 5742, 6192, 6217, 7237, 7282, 7308, 7355, 7382, 6322, 6350, 6397, 6420, 6467, 6491, 6516, 6567, 6609, 6630, 6677, 6727, 6749]. Publications [6775, 6799, 6846, 6870, 6901, 6967, 7024, 7102, 7130, 7159, 7214, 7407, 7429, 8347, 8381, 8411, 8436, 8517, 8548, 8605, 8632, 8662, 8692, 8721, 7457, 7480, 7504, 7552, 7578, 7632, 7660, 7690, 7715, 7741, 7769, 7821, 7851, 7879, 7901, 7926, 7948, 7974, 8022, 8045, 8174, 8202, 8223, 8302, 8797, 8829, 9861, 9894, 9921, 9949, 9978, 10007, 10035, 10066, 10090, 8857, 8918, 8944, 8971, 9002, 9054, 9080, 9108, 9136, 9161, 9245, 9272, 9411, 9440, 9472, 9528, 9555, 9623, 9647, 9672, 9713, 9738, 9766, 9838]. Publications [10171, 10192, 11024, 11045, 11070, 11120, 11186, 11208, 11230, 11249, 11272, 11296, 11346, 10215, 10236, 10300, 10330, 10352, 10400, 10432, 10459, 10515, 10540, 10561, 10586, 10612, 10631, 10656, 10679, 10729, 10751, 10774, 10794, 10914, 10934, 10963, 10991, 11436, 11583, 12528, 12573, 11614, 11635, 11663, 11686, 11711, 11759, 11785, 11891, 11943, 11962, 12016, 12040, 12063, 12200, 12222, 12244, 12270, 12291, 12311, 12404, 12467, 12487, 12511]. Published [673]. Publishing [10684]. Puebla [2982]. Pug [10877]. Pull [10017]. Puller [117, 2319, 5261, 9458, 10346]. Pulley [1287, 4100, 5594, 6309, 7395, 7252, 7957, 9040]. Pulleys [4804, 5492, 5538, 5946, 6328, 6199, 8992]. Pullman [5147]. Pulp [3503, 6042]. Pulper [8332]. Pulping [1082]. Pulsatory [11905]. Pulvebizer [9285]. Pulverizer [3047, 3624, 3903, 6233, 6812, 7676, 9090]. Pulverizing [9588]. Pump [14, 1209, 865, 1756, 2242, 2111, 1610, 2514, 3461, 3145, 2661, 4628, 4652, 4472, 6008, 6143, 5809, 6409, 6668, 7298, 7918, 8081, 8156, 8368, 9179, 9491, 9480, 10133, 9065, 10324, 10716, 11609, 1713, 8165, 8141]. Pumping [1040, 666, 1775, 2746, 5563, 6401, 6837, 7015, 8336, 7536]. Pumps [377, 4302, 4924, 7546, 11478]. Punch [234, 2214, 2155, 2592, 4582, 6388, 10866, 10879, 11695]. Punching [1267, 3717, 3990, 7868, 9282, 11172, 11327]. Purchase [5159]. Purchaser [2835]. Purchasers [3567]. Purifier [6533, 7892, 12363]. Purifiers [5573]. Purifying [7110, 11088]. Purity [12340]. Purple [881]. Purposes [313, 2291, 2508, 3279, 2655, 7251, 8591, 11707, 11675, 2814]. Push [6601]. Pushing [8931]. Puzzle [1178, 9290, 9430, 10450]. Puzzles [6140]. Pygmy [3133, 7597]. Pyrethrum [4370]. Pyromagnetic [10135].

Quadricycle [4037]. Quadrumana [8265]. Quadruple [6668, 8653]. Quaker [8111]. Qualities [1934, 1642]. Quality [683]. Quantity [5183]. Quarantine [1173]. Quarries [5305, 9102]. Quartz [2487, 2990]. Queer [3346]. Queries [3, 28, 729, 760, 789, 817, 840, 873, 901, 928, 951, 980, 1005, 1029, 1045, 1070, 1092, 1119, 1145, 1169, 1199, 1228, 1255, 1281, 1311, 1345, 55, 82, 106, 152, 172, 201, 220, 248, 280, 309, 337, 366, 395, 423, 448, 475, 497, 520, 547, 571, 598, 627, 651, 704, 1376, 1399, 2029, 2050, 2074, 2097, 2117, 2141, 2163, 2185, 2206, 2230, 2257, 2279, 2305, 2327, 2354, 2376, 2392, 2413, 2435, 2463, 2494, 2524, 2553, 1429, 1455, 1479, 1500, 1529, 1554, 1576, 1601, 1623, 1646, 1674]. Queries [1696, 1719, 1743, 1765, 1790, 1816, 1838, 1866, 1900, 1929, 1957, 1979, 1999, 2579, 2599, 3158, 3182, 3209, 3236, 3265, 3290, 3317, 3340, 3364, 3386, 3404, 3427, 3446, 3473, 3495, 3519, 3543, 3564, 3592, 3614, 3637, 3664, 3687, 3709, 2625, 2651, 2679, 2699, 2720, 2742, 2765, 2788, 2808, 2831, 2850, 2870, 2894, 2917, 2935, 2952, 2977, 2997, 3017, 3040, 3057, 3079, 3119, 3137, 3732, 3756, 4326, 4354, 4379, 4405, 4431, 4456, 4484, 4511, 4539, 4563, 4595, 4620, 4646, 4665, 4690, 4714, 4740, 4766, 4791, 4816, 4837]. Queries [4863, 4889, 3775, 3801, 3820, 3843, 3865, 3889, 3911, 3935, 3957, 3977, 4001, 4021, 4049, 4071, 4092, 4111, 4137, 4160, 4181, 4205, 4227, 4252, 4271, 4298, 4910, 4934, 5561, 5586, 5608, 5630, 5653, 5674, 5699, 5720, 5744, 5768, 5793, 5817, 5841, 5863, 5890, 5913, 5939, 5959, 5983, 6000, 6025, 6052, 6081, 6105, 6130, 6157, 4960, 4989, 5013, 5039, 5060, 5086, 5109, 5133, 5156, 5180, 5205, 5228, 5253, 5279, 5303, 5328, 5352, 5377, 5404, 5428, 5454, 5482, 5511, 5533, 6177, 6194, 6801, 6824, 6848, 6872, 6902, 6926]. Queries [6948, 6969, 6999, 7026, 7049, 7074, 7104, 7132, 7160, 7186, 7215, 7238, 7263, 7284, 7310, 7333, 7357, 7384, 6219, 6246, 6268, 6295, 6324, 6352, 6375, 6398, 6422, 6449, 6468, 6493, 6518, 6541, 6569, 6589, 6611, 6632, 6657, 6679, 6703, 6728, 6751, 6777, 7409, 7431, 8096, 8125, 8148, 8175, 8203, 8224, 8250, 8276, 8304, 8326, 8349, 8383, 8413, 8438, 8460, 8491, 8519, 8550, 8579, 8607, 8634, 8663, 8694, 8723, 7459, 7482, 7506, 7531, 7554, 7580, 7604, 7634, 7662, 7691, 7717, 7743, 7771, 7796, 7823, 7853, 7881, 7903]. Queries [7928, 7950, 7976, 7999, 8024, 8047, 8776, 8799, 9625, 9648, 9671, 9690, 9714, 9739, 9767, 9839, 9862, 9895, 9922, 9950, 9979, 10008, 10067, 10091, 8831, 8859, 8887, 8920, 8946, 8973, 9003, 9056, 9082, 9109, 9137, 9162, 9215, 9273, 9303, 9330, 9355, 9385, 9413, 9442, 9474, 9499, 9556, 9582, 10173, 10193, 10887, 10915, 10935, 10964, 10992, 11023, 11121, 11142, 11165, 11187, 11209, 11231, 11250, 11297, 11321, 11347, 11404, 10216, 10237, 10270, 10301, 10331, 10354, 10374, 10401, 10431, 10460, 10516, 10541, 10562, 10587, 10613, 10633, 10658, 10681, 10752]. Queries [10775, 10795, 10816, 10839, 10860, 11437, 11459, 12087, 12105, 12129, 12152, 12174, 12202, 12223, 12245, 12271, 12292, 12312, 12334, 12356, 12377, 12406, 12426, 12446, 12466, 12489, 12510, 12530, 12552, 12575, 12595, 12617, 11486, 11513, 11530, 11559, 11584, 11615, 11636, 11662, 11687, 11713, 11733, 11758, 11787, 11806, 11825, 11846, 11892, 11919, 11942, 11961, 11990, 12015, 12041, 12064]. Question [4075, 5748, 10662]. Questions [3045]. Quick [3125, 9960, 11050, 11667]. Quick-Firing [11667]. Quieting [10309]. Quilter [10226]. Quilting [4388, 7443, 8621]. Quinnipiac [429].

R.I. [9637]. Rabbit [3452, 12307]. Race [298, 525, 9634, 8905, 6827, 10011]. Racine [163]. Racing [7746, 9489]. Rack [1362, 2037, 6512, 7063, 6555, 7255, 9455, 8868, 8982]. Racquet [9970]. Radial [1886, 11741, 4261]. Radiant [1511]. Radiated [11926]. Radiation [383]. Radiators [5620, 10846, 12320]. Raffard [3392]. Raft [5730, 9927, 10437, 11408]. Rafters [8230]. Rafting [8783]. Rag [4732]. Ragona [5589]. Rail [273, 2535, 3786, 4655, 4062, 4805, 5806, 5466, 7269, 6718, 6619, 7116, 10105, 9373, 10667, 10665, 11053, 10449, 10879, 10525, 11836, 12478]. Railroad [1250, 2381, 2286, 3172, 2884, 2983, 3348, 3169, 3085, 3813, 4554, 4998, 5407, 6645, 7087, 6502, 6880, 7060, 7116, 7058, 6482, 6759, 8361, 7884, 7445, 9912, 9531, 9926, 9904, 9393, 9801, 9032, 10124, 9652, 9996, 10100, 10244, 10496, 11001, 10227, 11181, 10335, 12024, 12435, 11972, 12044, 11645, 11809, 11449, 12257, 8, 4344, 7107, 12329]. Railroads [1132, 224, 4656, 5879]. Rails [5927, 9069, 10803, 11314, 12586]. Railway [1051, 988, 408, 669, 372, 425, 831, 1017, 268, 847, 1284, 85, 603, 1771, 1904, 2287, 1723, 1875, 2397, 3020, 3500, 3720, 3456, 3547, 2658, 3526, 2631, 2981, 3389, 3784, 4231, 4055, 4041, 3808, 4062, 4441, 4051, 4648, 4082, 3736, 5708, 5869, 5027, 5688, 5230, 4926, 5466, 6478, 7350, 6385, 6358, 6549, 7220, 6854, 6743, 6599, 6983, 6187, 8080, 8683, 8743, 8015, 8208, 8340, 8333, 8571, 8355, 7570, 8232, 8590, 7558, 7747, 9560, 8894, 8925, 9562, 9724, 9723, 9373, 9779, 8848, 9930, 11037, 10667, 10365, 10275, 11253]. Railway [10247, 10389, 10621, 10525, 10571, 11279, 11386, 10252, 10720, 10708, 10304, 10343, 11911, 12373, 12477, 11775, 12118, 12478, 12030, 12161, 12163, 11470, 4384, 5994, 6009, 6238, 6731, 8400, 8028, 8002, 7931, 9425, 10636, 10802, 12359]. Railway-Crossing [2631]. Railways [947, 109, 2470, 1873, 4470, 4994, 5208, 5282, 7585, 1935]. Rain [943, 1316, 712, 2314, 3872, 5571, 8684, 10380]. Rainfall [1009, 4053, 5464, 7980, 10039]. Rainfalls [1062]. Raisin [12230]. Raising [4239, 4080, 4435, 10050, 9727, 9742]. Rake [2686, 9170]. Ram [1520, 6249]. Rameses [8117]. Rams [8418]. Ran [1245]. Rana [7990]. Range [9222]. Ranges [2963, 6368, 11543]. Rapid [5666, 5686, 5549, 7287, 10830, 12482]. Rapidly [8292]. Rare [8129, 8573]. Rarefied [3649]. Ratchet [5023, 5661, 7736, 10671]. Raton [8]. Rattlesnake [10590]. Rattling [9225]. Raub [2144]. Raw [5657, 12474]. Rays [920]. Razor [12370]. Razors [923, 3193]. Re [4525, 10745]. Re-Enforcement [4525, 10745]. Reader [11763]. Readers [5917]. Reading [2959, 6480, 10897, 11899]. Ready [9591, 12250]. Ready-Reckoning [9591]. Real [1640]. Reality [3480]. Really [9447]. Reamer [11391]. Reamers [5867]. Reapers [3622, 7144, 8188, 8362, 9009, 10076]. Reappearance [4574]. Reasonable [812]. Rebacking [4024]. Received [25, 52, 1141, 1196, 1253, 1340, 1372, 78, 125, 149, 167, 198, 304, 333, 362, 391, 444, 516, 543, 623, 672, 727, 785, 836, 896, 925, 977, 1026, 1115, 1282, 1423, 1452, 1475, 1498, 1525, 1549, 1620, 1715, 1761, 1786, 1895, 1953, 1995, 2025, 2159, 2251, 2276, 2300, 2349, 2430, 2574, 2595, 2621, 2648, 2784, 2827, 2847, 2953, 3014, 3037, 3335, 3490, 3541, 3705, 3728, 3797, 3884, 3934, 3996, 4108, 4156, 4177, 4267, 4294, 4787, 4812, 4907, 6945, 7183]. Receiver [1445, 2689, 4605, 4243]. Receiving [9720, 10528]. Receiving-Table [10528]. Recently [3424, 8555, 11163, 11184, 11206, 11228, 11434, 11457, 12061, 12085, 12103, 12148, 12170, 12172, 12199, 12220, 12242, 12268, 12289, 12309, 12332, 12354, 12375, 12403, 12424, 12443, 12464, 12486, 12508, 12527, 12550, 12572, 12593, 12615, 11484, 11511, 11554, 11557, 11581, 11612, 11633, 11660, 11684, 11710, 11731, 11756, 11784, 11804, 11823, 11844, 11865, 11889, 11917, 11940, 11959, 11988, 12013, 12038]. Receptacle [2344, 8736]. Recipes [11966]. Reciprocating [7779, 10097, 11902]. Reckoning [9591].

Reclaiming [11761]. Reclamation [8028]. Reclining [5119, 7587]. Recognized [3300]. Record [542]. Recorder [5432, 2883]. Recording [385, 1445, 6809, 9965, 12628, 931]. Recovery [7093]. Rectangular [912, 11914]. Red [1360, 3779]. Reddening [3226]. Rediscovery [2759]. Reducing [11466]. Reed [1303, 2267, 10286]. Reef [1231]. Reel [2346, 2344, 5807, 5389, 7081, 6511, 9368, 10821, 11200]. Reeled [2922]. Reels [10162]. Reese [1509, 1566]. Reference [1106]. Refineries [5024]. Refiners [12248]. Reflection [9045]. Reflector [2454]. Reform [2188, 4723]. Refraction [6365]. Refreshments [2441]. Refrigerating [381, 1992, 5950, 11381]. Refrigeration [642, 3529]. Refrigerator [1568, 3877, 3785, 5045, 6232, 8709, 8708, 9490, 12346]. Refuge [6757]. Refuse [1562, 11949]. Regal [9664]. Regard [1406, 1905]. Regenerative [5655, 9818]. Regenerator [6736]. Regente [8452]. Regia [6315]. Region [1486]. Register [3275, 7035, 8502, 8507, 8892, 10469]. Registering [383, 1586, 2087, 3460, 5970, 6695, 6593, 8338, 9635, 10789]. Registers [4448]. Registration [317, 6402, 6404]. Regnard [2971, 3256]. Regulating [3985]. Regulation [60]. Regulator [164, 508, 189, 2239, 3256, 5519, 5989, 4924, 5992, 8038, 8471, 8910, 8954, 10846, 10905, 11831]. Reims [10556]. Rein [1801, 3252, 2641, 4446, 5703, 7272, 8782]. Rein-Holder [2641]. Reina [8452]. Reinholder [11769]. Reins [531]. Reis [4243, 7135, 7363]. Reissue [1403, 1409, 3061]. Reissued [1905, 2999]. Rejected [2080]. Related [8909]. Relating [725, 756, 775, 858, 917, 976, 1025, 1060, 1114, 1139, 1328, 195, 360, 390, 471, 494, 594, 993, 1277, 1304, 2548, 2572, 1389, 1416, 1664, 1894, 1924, 1946, 1968, 2403, 1446, 1516, 1614, 1635, 1685, 1753, 2244, 2619, 2674, 2858, 2965, 3175, 3198, 3223, 3251, 3303, 3574, 2711, 2754, 3719, 3856, 3988, 4057, 4085, 4520, 4753, 4011, 4035, 4942, 4967, 5234, 5646, 5668, 5737, 5830, 5906, 6012, 6070, 5975, 6189, 6560, 6720, 7323, 7465, 7561, 8063, 8087, 8194, 8267, 8309, 8450, 8499, 8597, 8652]. Relating [7451, 7845, 8815, 8902, 8984, 9182, 9636, 9870, 9901, 10733, 11525, 11725, 11853, 12009, 8369, 8716]. Relation [1368, 905, 2333, 2262, 3063, 7038]. Relations [4799]. Relative [3917, 6473]. Releasing [459]. Relief [2035, 5515, 8564]. Reliefs [3603]. Relieve [9256]. Remarkable [1317, 1125, 679, 227, 3105, 3599, 4556, 3983, 5090, 5490, 5309, 5749, 6716, 7884, 8355, 7435, 10023, 9416, 10039, 10553, 11124, 11408, 10720, 11955, 11655, 11737, 2539, 3416, 2562]. Remedies [4843, 6706, 8207]. Remedy [6733]. Remington [5048, 4972, 7671, 11201, 12187]. Reminiscences [9306]. Remittances [479]. Removable [5068, 10157, 11357, 11083]. Removal [1231, 2089, 7273, 7909, 8523, 11639]. Remove [7418]. Removing [1132, 4171, 5210, 10124]. Renard [5667]. Renewal [2416]. Renewing [9777]. Renovating [10318]. Reopened [2469]. Reorganization [5725, 10407]. Repairing [4599, 7175, 8429, 9069, 10240, 1185]. Repeating [6981, 11766]. Replacer [6760, 6784, 9169, 10762]. Replacing [8805, 12209]. Reply [12020]. Report [1906, 1536, 4345, 7775, 10732, 12019, 12600]. Reports [4078]. Representatives [654]. Reproduced [4127]. Reproduction [2263, 11073].

Republic [10985, 2088]. Repulsion [113]. Required [7220]. Requirements [1284]. Requiring [12439]. Rescue [5149]. Research [3617, 9717]. Researches [7163]. Reservation [2003]. Reservoir [11423, 12590]. Reservoirs [8429, 12581]. Resharpening [4235]. Residence [3377, 7623, 10367]. Residences [1820]. Resignation [2362]. Resin [6614]. Resistance [12239, 12116]. Resolved [8524]. Resources [11790]. Respiration [340]. Respirations [12628]. Responsibility [1098]. Rest [2296, 3096, 4584, 8738, 9847, 10126, 10942, 11447]. Restaurants [6509]. Restoration [1382, 5385, 10556]. Restricted [9532]. Rests [2217]. Result [851, 4555]. Results [3062, 6180]. Retain [9612]. Retarded [2539]. Retenoning [5802]. Retention [2562]. Retina [1583]. Retort [1218]. Retsof [12543]. Return [5845, 9585]. Rev. [11631]. Revaccination [1983]. Reversible [2339, 1587, 5760, 7042, 10043, 8938, 12629]. Reversing [4805, 5806, 6336, 10099, 12393]. Revised [4287]. Revival [907, 4820, 7219, 10096, 9061]. Revivifying [4705]. Reviving [2080]. Revolution [197, 11236, 11810]. Revolutions [6671]. Revolver [5215, 7254]. Revolvers [5804]. Revolving [636, 3089, 3095, 5545, 6062, 7113, 8469, 8366, 9482, 11929]. Reward [13, 3760]. Rheometric [3032]. Rheostat [12116]. Rhinoceros [9345]. Rhododendron [6650]. Ribbons [9590]. Rice [1434, 9746]. Rich [9771]. Richard [7829, 1586]. Richly [1619]. Ricks [3248]. Rid [1702]. Ridging [6620]. Riding [100, 5868]. Riffles [2862]. Rifle [11244, 12187, 12033]. Rifles [868, 9654]. Rigged [7067]. Right [5537, 7327]. Righting [1224, 9264]. Rights [1982, 3433, 4213, 5568]. Rim [5901]. Ring [3281, 5297, 10311, 12456]. Ringed [4528]. Ringer [2238]. Rings [9395]. Rink [635]. Rio [6009, 4876]. Rip [6362]. Ripening [10079]. Ripping [5077]. Risks [1771, 3643, 4385]. Rival [11493]. River [1096, 399, 768, 845, 156, 1154, 753, 60, 985, 1257, 1745, 3144, 2639, 3019, 3650, 3577, 3924, 3914, 6572, 6769, 7434, 9121, 8959, 10741, 12067, 12121, 12090, 11653, 11644, 8291, 11562, 5710, 1338, 259, 2209, 2906, 4622, 4914, 6951, 6521, 7534, 10890, 12462, 11886, 1993, 8678]. Rivers [1848]. Riverside [7313]. Rivet [11767]. Riveter [2288]. Riveting [38, 2499, 1879, 2264, 4926, 8233]. Riviera [8880]. Roach [3393, 9307, 9360]. Road [889, 1688, 1391, 2043, 3530, 5873, 6828, 9025, 9153, 8867, 10446, 10649, 11504, 12111, 11907, 12351]. Roadrunner [7920]. Roads [470, 1873, 12521]. Roadside [9820]. Roadway [6623]. Robbery [5537]. Robert [3872, 4918, 12587]. Robinson [9822]. Roburite [8845]. Rochelle [3358]. Rochester [3667, 4828]. Rock [773, 2089, 3899, 4418, 4342, 6071, 6937, 6972, 7273, 6958, 8680, 8373, 8808, 9366, 10077, 11128, 11287, 11180, 2942]. Rocker [4309, 9940, 12055]. Rocking [7246, 7394, 7587, 10738, 11624]. Rocky [11449]. Rod [1779, 4389, 6511, 9480, 10162, 11478, 12606]. Rodgers [2770]. Rods [2838, 6782, 7512, 7496, 11745]. Rolled [12586]. Roller [64, 3378, 3927, 4061, 5922, 5778, 5596, 6279, 6914, 7643, 9153, 9963, 11651]. Rolling [5758, 7086, 6787, 8767, 9318, 12612, 11984, 2423]. Rolls [9607]. Roman [5241, 10321, 10556]. Romance [3480]. Rondout [4255]. Roof [3248, 7172]. Roofed [470]. Roofing [12115]. Roofs [7038]. Room [983, 3085]. Rooms [3987, 7150, 12476, 12363]. Root [8390]. Roots [1023, 7536, 7546]. Rope [719, 1468, 4899, 5923, 6187, 10052, 10504, 11628]. Rope-Clamp [1468]. Rosaline [3212]. Rose [11396, 11397]. Rosin [945]. Rot [4277]. Rotary [1209, 327, 3191, 4281, 4052, 3740, 7297, 6305, 7678, 8705, 8672, 10409, 10220, 10385, 11546]. Rotating [4978]. Rotation [264]. Rouen [11394]. Rougerie [8878]. Round [4030, 7067]. Rounding [9147]. Route [4871, 7137, 9868]. Routing [11521]. Row [183, 3086, 6090, 5685, 6687, 8614, 10500, 10946]. Rowing [6387]. Royal [9235, 2152, 8352, 11224]. Royalties [1380]. Rubber [780, 1174, 7726, 9592, 11355, 12079]. Rubbing [8670]. Rudder [6202, 6289, 6314, 6345, 6208]. Rudderless [6889]. Rudge [9262]. Ruelle [4705]. Rufus [5590]. Ruin [10556]. Ruined [5089, 5122, 5168, 5220, 5243]. Rule [8524, 9905, 10893, 12137]. Ruled [6332]. Rules [3652]. Ruling [9752]. Rumor [11965]. Running [1250, 5359, 6260, 6862, 7891, 9699, 10690, 11181, 11314, 12281]. Running-Part [6260]. Ruptures [11776]. Rural [3433, 1796]. Russell [3625]. Russia [6356]. Russian [1522, 3845, 4527, 6121, 7178, 10520]. Rust [6403]. Rutherfurd [12440].

S [6807]. Saber [12459]. Saber-Toothed [12459]. Saccharin [12411]. Sack [10104]. Sack-Filling [10104]. Sacred [4473]. Sad [4366, 6953, 7541, 8057, 9253, 8938]. Sad-Iron [4366]. Saddle [2632, 5691, 7088, 11059, 8240]. Saddletree [8958]. Safe [985, 3410, 2962, 4034, 3767, 7979, 11218]. Safeguards [12361]. Safes [4010]. Safest [10594]. Safety [1182, 871, 1127, 459, 1519, 2221, 2634, 3111, 4359, 4461, 4365, 4391, 3780, 5546, 5989, 5597, 6011, 6739, 6455, 6253, 6861, 6255, 6306, 6582, 8139, 8390, 7907, 7570, 7841, 8447, 8501, 8500, 9939, 9932, 10100, 9827, 9066, 8866, 9632, 8992, 10276, 10896, 11257, 11982, 12629, 11932, 12605, 1338, 1130]. Sail [871, 6006, 5314, 7067, 7090]. Sailing [1411, 7314]. Sailor [6831]. Sailors [9142]. Sails [2055]. Saint [4729, 9318]. Salamander [1085]. Salamanders [9208]. Sale [3777, 8835]. Salida [11254]. Salpetriere [7425]. Salt [1298, 11302, 6866, 12543]. Salts [1298]. Sample [4978, 7142]. Samuel [10111]. Sand [1166, 4470, 3942, 7228, 7343, 11340]. Sanding [12435]. Sandpapering [8448, 12362]. Sandstone [1860]. Sandy [10094]. Sanitary [1465, 1965, 3377, 3807, 7442]. Sanitology [1629]. Santa [8]. Sargasso [1594]. Sargent [9782]. Sargentii [9782]. Sartell [11470]. Sash [890, 2985, 2667, 4363, 4730, 5021, 6383, 7173, 6476, 7365, 8643, 7755, 8704, 7615, 8103, 9177, 9907, 11102, 10973, 11813, 12318, 12213]. Sash-Cord [2985, 5021]. Sashes [8701]. Satisfactory [11517]. Saucepan [7962]. Savage [12205]. Saver [69]. Saving [567, 2474, 2874, 5115, 7056, 6509, 11977]. Savings [5724, 10488]. Saw [1131, 2497, 2478, 1509, 3325, 3534, 2664, 4601, 4284, 4169, 3787, 3743, 4140, 4721, 5676, 4921, 6092, 6145, 5311, 6068, 5517, 5100, 6228, 6229, 6987, 6362, 7805, 8182, 7444, 8472, 8039, 7489, 9035, 9258, 9568, 10259, 10472, 10546, 10691, 10903, 12298, 11474]. Saw-Clamps [4921]. Sawing [2615, 2590, 3896, 3942, 7396, 7963, 8031, 11065, 10497, 12299, 11475]. Sawmill [2987, 4459, 11240]. Sawmills [3944]. Saws [3111, 9506, 9449, 11083, 11812]. Sawyer [162]. Say [4567, 5868]. Scaffold [10051, 11198, 11032, 12303]. Scaffold-Supporting [12303]. Scaffolding [4679]. Scaffolds [7172]. Scalder [10019]. Scale [1107, 140, 3034, 4150, 10471, 10282, 10182, 10178, 9821]. Scaled [12264]. Scales [2904, 12186]. Scallops [1393]. Scaly [1523]. Scaly-Finned [1523]. Scarabæus [4473]. Scarf [2014]. Scenery [11690]. Scenes [11170]. Schaeberle [2561]. School [737, 345, 3811, 4544, 5418, 6598, 9589, 11595, 8476]. Schooled [3761]. Schooling [3640]. Schools [794, 4410, 6282, 11056]. Schultze [11934]. Science [739, 803, 1088, 823, 794, 2058, 3358, 4442, 4745, 3768, 3850, 4546, 5634, 5025, 8308, 10120, 8877, 8963, 9073, 9127, 9236, 9294, 9322, 9346, 9374, 9403, 9433, 9967, 5440, 256, 346, 1076, 2334, 8637, 10095, 10070, 11412, 12599, 12276, 12132, 3195, 4279, 4895]. Sciences [1108, 134, 257, 405, 1631, 3459, 4148, 6818, 7108, 7640]. Scientific [45, 3140, 3062, 4308, 9504, 10112, 10359, 10263, 10156, 12608, 12109, 12544, 8116, 8227, 8215, 9942, 11820, 126, 673, 1151, 3046, 5066, 5590, 5935, 5936, 9382, 8865, 12430, 12056]. Scissors [140, 1470, 2009, 5077, 12212]. Scoop [10844]. Scoring [8074]. Scott [3625, 10452]. Scourer [4971]. Scouring [11135]. Scraper [2043, 5760, 8764, 11384, 11567]. Scraping [6003, 8058, 8311]. Scratch [5445]. Screamer [1473]. Screen [1054, 6884, 8732, 11774, 12453]. Screw [315, 1912, 4565, 5331, 6098, 5750, 5845, 6110, 5458, 6040, 7302, 8187, 8040, 7495, 7622, 9657, 10043, 8992, 10974, 11422, 11168, 11424, 11904, 12371, 11906]. Screw-Threading [12371]. Screwdriver [12278, 12454]. Screws [4257, 5821, 9396]. Scroll [6229]. Scrub [7341]. Scylla [2193]. Scythian [7761]. Sea [1021, 501, 996, 2557, 1504, 1950, 1535, 3463, 3792, 4732, 4387, 3780, 5846, 5354, 5340, 7291, 7981, 7611, 9719, 8978, 8864, 10624, 10918, 10184, 12317, 12155, 269, 1594]. Sea-Sounding [1021]. Seagrave [3525]. Seal [967, 6303, 8256, 9609, 10898, 11257, 11577]. Seal-Press [11577]. Sealed [6301]. Seals [9022, 10025, 3921, 808]. Seamanship [7003]. Seaport [11912, 374]. Search [859, 2770, 3413, 9194, 12578]. Seas [1252]. Seasickness [4142]. Seat [1990, 4309, 5420, 5685, 6114, 5709, 6641, 6598, 6882, 8059, 8422, 8003, 10105, 9589, 10947, 11835, 11595]. Seating [4848]. Seats [7389, 9777, 11447]. Sebastes [3306]. Second [1721, 11266]. Secondary [3394, 12602, 12325]. Seconds [4876]. Secor [8582]. Secret [2513]. Sectional [5101, 6230, 8818, 10424]. Sections [752]. Secure [7907, 9225]. Securer [3694, 4653, 7173, 8561, 9201]. Securing [1249, 5826, 9014, 9537, 9879]. Seed [879, 1615, 3222, 3194, 4569, 4217, 7011, 9750, 9259, 11607, 11672, 12433, 12538]. Seeder [7400]. Seeders [9588]. Seeds [3557, 9336, 10670, 10503, 3875]. Seeing [409]. Seeking [3064]. Seen [1296, 3576, 3525, 3723, 5490]. Self [1224, 380, 741, 1988, 2087, 2221, 1410, 3460, 4194, 4385, 6036, 5687, 5594, 6362, 6593, 7495, 9482, 9123, 8848, 8938, 9988, 9635, 9264, 10419, 11222, 10642]. Self-Acting [6036, 5687, 7495]. Self-Adjusting [741]. Self-Charging [11222]. Self-Cleaning [9123]. Self-cocking [2221]. Self-Extinguishing [8848]. Self-Feeding [6362]. Self-Gauging [4194]. Self-Heating [8938]. Self-Imposed [4385]. Self-Inking [10419]. Self-Leveling [1988]. Self-Loading [9988]. Self-Locking [380]. Self-Luminous [10642]. Self-Oiling [5594]. Self-Registering [2087, 3460, 6593, 9635]. Self-Righting [1224, 9264]. Self-Ventilating [9482]. Self-Winding [1410]. Sell [933]. Semi [11426]. Semi-Portable [11426]. Semicircular [6139]. Senate [86, 3567, 10465]. Senegal [3147]. Seniority [9278]. Sense [7560]. Sensitive [833, 11523]. Separable [9872]. Separating [2702, 6955, 7110, 7681, 9884]. Separation [2439, 5113]. Separator [462, 960, 1186, 1705, 3301, 2750, 4683, 3872, 8360, 7418, 12364, 12500]. Separators [3787]. September [3142, 4900, 6162, 7390, 10121, 11413, 12622, 8102, 8180]. Sequard [12555]. Series [1172, 1177]. Serious [1257, 3502]. Serpents [1872, 3870, 11705]. Service [1768, 3499, 3150, 2746, 7591, 10872, 10649, 12274, 12458, 6239]. Service-Pipes [1768]. Serving [5923]. Set [767, 315, 2499, 2199, 6987, 9541, 9749, 8992]. Setting [3944, 5728, 5517, 5461, 8084, 11415]. Setting-Up [5461]. Settled [1986]. Seventeenth [7625]. Sevres [2199, 2093, 2517, 2518, 1595]. Sewage [2799, 11106, 12110, 12034]. Sewer [523, 65, 5265, 7203, 7843, 9842, 11981]. Sewerage [1582]. Sewers [1263, 1659, 7399, 9698, 11055]. Sewing [848, 2054, 2090, 3323, 4416, 6086, 5897, 7300, 6284, 6596, 9959, 10127]. Sewing-Machine [2054, 6086]. Seychellarum [12591]. Seychelles [12196]. Seymour [12516]. Shackle [12118]. Shade [665, 6065, 5808, 5778, 9042, 10101, 9141, 11382, 11393]. Shades [9701, 12002]. Shading [240]. Shadow [8113]. Shaft [2400, 4423, 5431, 5829, 5361, 6693, 6203, 6671, 9564, 8903, 12558, 11503]. Shafting [1613]. Shafts [5701, 5726, 6111, 7014, 6915, 6308, 12049]. Shake [9520]. Shall [59, 4213, 6226, 7338, 6782, 6430]. Sham [4333, 9899, 8890]. Shams [8890]. Shanks [6014]. Shape [4030]. Shaping [6789, 7618, 8159]. Shark [4289, 4556, 9450, 3604]. Sharks [3512]. Sharpener [5736, 11693]. Sharpening [4126, 9506, 12026]. Shaving [530, 8563]. Shavings [4096]. Shaw [7990]. Shawl [533]. Shear [2592, 3253, 2734]. Shearing [3373, 3990, 6285, 8470, 9428, 9282, 11172]. Shears [2214, 1637, 1755, 3003, 2818, 2757, 4582, 4629, 7222, 6561]. Sheep [884, 878, 877, 3003, 2818, 4494, 8470, 9428, 10673, 5811]. Sheet [5952, 6561, 8735, 10647]. Sheets [6809]. Shelf [11284]. Shell [4550, 4849, 5779, 5574, 7139, 8035, 10868, 1450]. Shellac [157, 4116]. Shelling [6553]. Shelves [8289, 10444]. Sheraton [4657]. Sheridan [11351]. Sheriff [2358, 3777]. Shield [3111, 10873, 11903, 12345]. Shields [4997]. Shifting [60, 9633]. Shingle [7753, 9610]. Shingling [7836, 11335, 10567].

Ship [1051, 1017, 109, 1875, 2802, 3460, 3989, 4469, 4871, 5866, 5869, 5047, 5994, 5246, 6991, 7220, 8159, 8611, 7558, 8965, 9926, 9503, 10158, 11076, 11300, 10708, 10552, 10304, 11779, 11482, 11818, 12514, 11502, 12295, 11721, 11993, 11551, 12254]. Shipbuilding [5385]. Shipman [5611]. Shipping [6763, 7642, 9142, 11798]. Ships [1974, 2108, 3673, 4208, 6545, 6889, 7241, 6223, 6715, 6544, 9365, 9417, 9087, 8755, 9844, 10177, 11126, 11850, 12481, 11834, 12620, 12458]. Shipwrecks [3451]. Shirt [6670]. Shoal [10867]. Shoe [867, 2054, 1380, 6531, 6578, 9936, 10896, 11856, 12024, 12348]. Shoebill [4088]. Shoemaker [9799]. Shoes [486, 6557, 6791, 8874, 9453]. Shop [1990, 4332, 5209, 5724, 4943, 6684, 6856, 6706, 7560, 9361, 7030]. Shore [4648, 11968]. Short [1440, 2961]. Short-Hand [2961]. Shorten [10819]. Shortening [12409]. Shot [2272, 8469, 10524]. Shot-Case [2272]. Should [1876]. Shovel [47, 3025, 4920, 6505, 8137, 10385, 11567]. Shovels [10220]. Show [1597, 2670, 3768, 3850, 5414, 11282, 154, 877]. Show-Box [2670]. Shower [691]. Showers [7190]. Shown [7096]. Shows [4308]. Shrinking [5067]. Shutter [2385, 2563, 4146, 5416, 6665, 9119, 12191]. Shutter-Worker [2563]. Shuttle [4416, 7786, 9115]. Sibley [6961, 7933, 11255]. Sick [7150, 12363]. Sickness [1535]. Side [1157, 2632, 3768, 3850, 6064, 5501, 10315, 11777, 12387]. Side-Saddle [2632]. Sided [3328, 6819]. Sides [4673]. Siege [2655]. Siemens [3437, 5655, 7916, 8820, 9818]. Sieve [528, 12077, 7735]. Sifter [1543, 8737, 8364, 9658, 12097, 12182, 12560]. Sight [188, 2237, 5783, 6597, 9118]. Sign [351]. Signal [4731, 4554, 6072, 5027, 6645, 8474, 8208, 9393, 9801, 11056, 10521, 10621, 10571, 11279, 11386, 11875, 12477, 12054, 12274]. Signaling [3169, 4041, 6482, 9659, 9744, 11360]. Signals [1935, 3299, 4656, 8333, 7611, 11852]. Signet [863]. Signing [9604]. Signs [1757, 3073]. Silent [4195, 11170]. Silex [7276]. Silhouettes [1572]. Silk [1727, 1805, 1437, 2922, 3407, 3023, 3107, 3329, 3762, 8556, 2685]. Silks [958]. Silkworm [4573]. Silkworms [4757]. Sill [716, 10715, 12161]. Silo [5973, 11621]. Silos [4875]. Silver [314, 5460, 9617, 7179, 10357, 10475]. Silvering [1871, 2924]. Simple [135, 677, 1219, 2507, 1444, 1913, 3394, 4776, 4678, 5471, 6671, 6601, 7646, 7756, 8288, 7650, 7961, 9873, 9453, 10075, 9041, 9100, 9510, 10107, 9761, 11266, 10924, 11154, 10721, 11243, 10688, 10998, 10346, 12439, 12544, 2927, 1713, 9774, 262, 2883, 8165, 8050, 9942, 9544, 10204, 10881, 11089, 11132, 11179, 11203, 11216, 11237, 11264, 11288, 11339, 11395, 11452, 11728, 11935, 11957, 12036, 12059, 11802]. Simplex [944]. Simplified [1570]. Sims [12324]. Sing [11114]. Singhalese [9575]. Single [4975, 5466, 6741, 7245, 6580, 8734, 9752, 9910]. Single-Bell [6741]. Singular [1938, 1492, 7733]. Sinking [2108, 6308, 6855, 7511, 7609, 8959]. Siphon [6237, 11222, 10441, 12625, 8430, 8449, 11657]. Siphons [1819]. Sir [2797, 9321, 9359]. Sire [6769]. Siren [7317, 9744, 11819]. Sirups [4698]. Site [1152, 12621]. Sites [3825]. Six [158, 3804, 5901, 5666, 8653, 10577].

Six-Lever [5901]. Sixteen [10246, 11011]. Sixteenth [7623]. Sixty [7895]. Size [576, 6378]. Sized [2057]. Sizing [576]. Skate [1411, 3378, 6879, 10735]. Skates [6914]. Skating [635, 7090]. Skeg [8445]. Skeins [3407]. Skeleton [11136]. Sketches [4394]. Sketching [4801, 9847, 12211]. Skiagraph [2613, 7366]. Skid [6456, 7620]. Skill [5185]. Skin [986]. Skins [591]. Skipjacks [6013]. Sky [4144, 7165, 7031, 7641, 8102, 7414, 7437, 7955, 8054, 8641, 7537, 7802, 8180]. Slate [690, 1447]. Slates [3811]. Slats [7299]. Slaughter [3389, 11307]. Sled [995, 460, 2773, 6144, 6384, 11105, 10971, 11703]. Sleds [12630]. Sleeper [7169]. Sleeping [2134, 7915, 12363]. Sleeve [6259, 10159, 10284]. Sleigh [4846, 4897, 8033, 8363, 9226, 10317]. Slender [1833]. Slide [1969, 3007, 4497, 4316, 4496, 5823, 6091, 7061, 9997, 9963, 11286]. Slides [5146]. Sliding [5416, 7495, 9228, 12373, 11538]. Sling [11221]. Sloop [7746]. Slot [3457]. Slotting [2484, 4629, 7939, 11173]. Slow [9478, 9844, 10594]. Small [35, 479, 67, 2793, 3716, 2966, 4190, 4275, 4650, 6014, 6496, 6939, 8585, 8700, 7559, 7654, 9509, 9044, 464]. Smallest [7254]. Smallpox [2360, 1797]. Smelting [3324, 6135]. Smith [1078]. Smoke [2607, 7197, 7367, 7758, 7608, 7517, 12164]. Smoke-Consuming [7367]. Smokeless [1436]. Smokers [10016]. Smoking [7921, 9198, 11155, 11979]. Snake [24, 3148, 4166, 12067, 12400, 9184, 9520]. Snakes [505, 10448]. Snap [7064, 8644]. Snapping [1692, 8465]. Snow [4498, 5493, 6814, 6935, 6505, 7490, 7516, 7435, 7909, 8523, 7934, 8757, 8834, 10385, 11645, 11592, 11567]. Snows [956]. So-Called [9184]. Soap [530, 6503, 6583]. Social [823]. Societies [11465]. Society [478, 1892, 428, 1234, 4028, 4141, 11212]. Socket [4703]. Sod [6812, 12008]. Soda [557, 12484, 10422]. Sodium [1298]. Sodom [12590]. Soemmering [5393]. Soft [1782]. Softening [5874, 5285]. Solar [383, 2086, 1910, 2898, 4078, 4844, 3846, 4256, 6917, 9989, 9263]. Sold [2358]. Soldering [90, 2365, 1515, 1442, 5656, 8261, 11680]. Sole [4806, 11856]. Solid [3021]. Solids [9344]. Solstice [2939]. Some [722, 2380, 2803, 3045, 3827, 4571, 5953, 7037, 8453, 8788, 11377, 11718, 12386, 464, 7402, 9370, 10287, 11578, 177, 230, 895]. Something [2639]. Soot [7916]. Sorghum [4546]. Sorter [914]. Soudan [4941]. Sound [2023, 1747, 8557, 2883, 10745, 11578, 8086, 8685, 8713, 3921]. Sounder [6788]. Sounding [1021, 501]. Sounds [1432, 11073]. Soups [2704]. Source [1352]. Sources [5435]. South [1965, 5443, 417, 2082, 1848, 3349, 8400, 8428, 11849]. Southern [263, 6016]. Southwest [1434]. Sower [4217]. Sowing [9259]. Spacing [1757]. Spain [9731, 10890]. Span [10463, 11809]. Spanish [4364, 8452, 9122, 8823, 11197, 12417, 11551]. Spar [10253]. Spark [4674, 10362, 12557, 9155]. Speaking [93, 5520, 6391]. Special [4985, 5035, 5082, 5130, 5348, 5507, 5557, 5886, 5934, 5980, 6244, 6319, 6373, 6418, 6724, 6774, 7156, 7182, 7211, 7235, 7575, 7601, 7630, 7659, 7687, 7819, 8300, 8380, 8434, 8487, 8545, 9132, 9189, 9241, 9497, 9551, 9598, 9644, 9686, 9735, 9811, 9857, 9917, 9975, 10031, 10087, 10144, 10188, 10233, 10296, 10350, 10395, 10456, 10677, 10725, 10770]. Specialists [2684]. Specially [5917]. Species [3677]. Specific [9024, 9546]. Specimens [5953, 8573]. Spectacle [4330]. Spectacled [11289]. Specters [148]. Spectrophonic [1633]. Spectroscope [68, 10578]. Speech [2583, 11073]. Speed [189, 3191, 3347, 4841, 5976, 6031, 6806, 7732, 8554, 9807, 11275, 10361, 11175, 11389, 12580]. Speeds [5772]. Speing [5687]. Spencer [9519]. Spezia [4163]. Spheroidal [7870]. Sphinx [7814]. Spicules [7953]. Spider [1221, 619, 2314, 6367, 10058, 12623, 387]. Spiders [12327]. Spike [2983, 7087, 7058]. Spiked [7620]. Spikes [3436]. Spindle [2707, 3457, 10848, 12233, 11883]. Spindles [10569]. Spinning [1124, 2090, 3225, 5210, 11883]. Spiral [4939, 8187, 11424, 1185, 10810]. Spirally [10580]. Spirals [8481]. Spirit [6713, 6667, 11280]. Spiritualism [9250]. Spittoon [2964]. Splendid [9592]. Split [7123]. Spoke [4585]. Sponge [4192, 11701]. Spongilla [2672]. Spontaneous [884, 4744, 6851, 10288, 10509]. Spool [8928]. Spoon [585, 9697]. Spoons [5460]. Sporting [384, 8591]. Spot [2007, 3779]. Spots [1910, 5161]. Spotted [723, 2299]. Spout [10104, 11594]. Spouts [1166]. Spread [1155, 2420]. Spring [540, 2568, 2445, 2241, 1755, 1919, 3698, 3830, 4874, 5358, 5501, 6114, 5016, 6642, 6618, 6786, 6913, 6427, 7565, 8734, 8257, 8671, 7617, 7644, 8130, 7958, 7697, 8615, 8895, 9286, 9369, 9537, 10201, 10315, 11880, 11930, 12075]. Spring-Making [1919]. Springs [611, 4939, 5490, 11254, 1185]. Sprinkler [142]. Sprout [117]. Spy [1243]. Spy-Glass [1243]. Squadron [9707, 12580]. Square [2846, 6426, 7416, 7726, 10132, 10948, 11569]. Squares [2172]. Squaring [3253, 8961]. Squirrel [1367, 2408]. St [3122, 7683, 11912, 4963, 8006]. St. [85, 2284, 6521, 7146, 7218, 9320, 11984, 8678]. Stable [965]. Stables [10378]. Stacker [7750, 10277]. Stacking [10199]. Staff [3250]. Stag [1617, 8196]. Stages [5042]. Stair [12451]. Staircase [5236, 9639]. Stairs [66]. Stairways [4655]. Stake [6687, 9422]. Stalk [2319, 2686, 9849]. Stalls [11390]. Stamp [76, 2477, 3217, 3713, 6168, 7009, 7682, 7438, 11695, 12368]. Stamper [2990, 8936]. Stamps [7318]. Stanchion [9391]. Stand [717, 2106, 6123, 4978, 6062, 6839, 7782, 7867, 8105, 8673, 7826, 9987, 10571, 11279, 10203, 10508, 10825, 12582, 2927]. Standard [1608, 5093, 5458, 6389, 6471, 7445, 12303, 866, 9199]. Stands [11876]. Stanley [5101, 6457, 8818, 12461, 12381]. Star [4396, 6864, 9586, 6933, 4188, 8110, 8741, 12610]. Starch [1443, 7699]. Starlit [4144]. Stars [1618, 2545]. Starter [1292, 5260, 8185, 9337, 9700, 9539, 12051]. Starvation [709, 1350]. State [2046, 2250, 2234, 2528, 5283, 7610, 7721, 8837, 7870]. Staten [1235, 1274]. States [5732, 7019, 12380, 948, 1505, 1437, 1608, 2960, 3329, 4694, 4565, 6837, 7668, 9743, 9389, 9604, 9007, 11220, 11691, 12208, 11705, 2685]. Static [8167, 10473]. Station [1289, 2536, 3712, 4438, 5644, 9637, 8898, 11533, 4729, 12274]. Stationary [1539, 3305, 2859, 6528, 7831, 9091]. Stations [12602]. Statistics [1354]. Statuary [9049]. Statue [1207, 6378, 9731, 3523, 3965, 3961, 5240, 6594, 8307, 8280, 8387, 8677, 12588]. Statues [8482]. Status [5138]. Statute [10122]. Stave [11309]. Stay [11058, 10284]. Steak [212]. Steak-Tendering [212]. Steam [456, 1080, 801, 488, 707, 1321, 560, 43, 144, 415, 1236, 561, 1078, 433, 316, 645, 436, 2013, 2357, 2111, 2040, 1989, 2126, 1887, 1972, 2514, 1882, 2059, 2081, 2109, 2149, 2168, 2195, 2213, 2292, 2313, 2343, 2361, 2384, 2404, 2443, 2519, 2537, 1405, 2156, 2055, 2500, 3376, 3374, 2695, 3455, 3007, 3224, 2706, 3546, 2813, 3435, 3498, 3375, 2608, 2666, 2731, 3065, 3109, 3215, 3302, 3068, 3698, 3551, 3257, 2925, 3368, 4499, 4302, 4422, 4726, 4306, 3805, 4850, 4167, 4542, 4570, 4777, 4577, 4565, 5844]. Steam [5469, 5969, 5497, 5543, 5563, 4944, 5049, 5293, 5598, 5902, 5900, 5665, 6055, 4938, 4924, 5642, 5784, 5517, 5611, 5231, 5570, 6572, 6918, 6754, 6860, 6504, 6505, 7297, 7371, 6624, 7017, 6275, 6235, 7010, 6582, 6305, 6834, 6815, 7499, 7992, 8526, 7918, 8313, 8329, 7887, 8420, 7647, 8315, 8674, 7673, 8343, 7940, 7866, 7707, 7731, 8508, 8527, 7447, 7705, 7909, 7567, 8133, 8394, 7934, 8312, 8468, 9152, 8756, 9171, 8925, 9957, 8994, 9281, 10072, 9398, 9153, 10042, 9574, 9587, 9070, 8786, 9960, 9114, 9319, 9747, 9321, 9283]. Steam [11313, 10722, 10740, 10599, 11392, 11241, 10846, 10220, 10385, 10424, 10743, 10423, 11810, 12515, 11727, 12158, 11815, 12136, 11476, 12328, 12025, 12519, 12164, 12601, 11703, 512, 607, 828, 911, 1827, 8066, 8191, 8213, 8239, 1034, 1655]. Steam-Actuated [11815]. Steam-Boiler [2059, 2081, 2537]. Steam-Propelled [11703]. Steamboat [1179, 1759, 6992]. Steamboats [1848, 3004, 7559]. Steamer [399, 679, 1242, 1365, 1906, 2770, 4081, 4923, 5343, 5101, 6572, 8479, 8236, 7473, 8452, 7609, 11133, 11197, 11922, 1993, 7510]. Steamers [684, 1096, 35, 3296, 4364, 8164, 8527, 11125, 11378, 10594]. Steaming [10248, 10808, 12080]. Steamship [2682, 2791, 3653, 6484, 6208, 7637, 9742, 10179, 11791]. Steamships [1271, 3320, 3922, 10651]. Steel [631, 184, 455, 122, 313, 2505, 2312, 1936, 2617, 2924, 3528, 3066, 4004, 4773, 4411, 4093, 4571, 4436, 3942, 4516, 4211, 4915, 5797, 5820, 6137, 5050, 6112, 5919, 5565, 7146, 6709, 8005, 8584, 8464, 7509, 7612, 7464, 8253, 8666, 8062, 9447, 8818, 10863, 11244, 12248, 12609, 12587, 12033, 1993, 2451]. Steering [643, 2567, 2703, 6889, 6252, 8088, 10568, 10507, 11241]. Steering-Gear [11241]. Stem [8763, 9827]. Stems [1303]. Stencil [2472]. Steno [9098]. Steno-Telegraphy [9098]. Step [88, 95, 1038, 5164, 9287, 9312, 10279, 11838, 11951, 11697, 11817, 12027]. Step-Ladder [11817]. Stephen [12194]. Stephenson [1823, 2236]. Steps [7799]. Stereoscopic [510]. Stern [6099, 7473]. Stern-Wheel [6099]. Stethoscope [8241]. Stevens [1749]. Stick [8089]. Stiletto [6907, 6602, 11677]. Still [4190, 4280]. Stirring [4059]. Stirrup [8501, 12629]. Stock [1112, 860, 2320, 2475, 2960, 2640, 4584, 5658, 7201, 7701, 9420, 9175, 10849, 10411, 3022]. Stock-Alarm [2640]. Stockholm [5370]. Stockings [10951]. Stokers [12345]. Stomach [2202, 9758]. Stone [916, 1083, 115, 2581, 4490, 5305, 7084, 7444, 8012, 7963, 7799, 9062, 11158, 12026, 12505, 12420]. Stone-Cutting [12026]. Stoner [9019, 1107]. Stones [3575, 4670, 7954, 12163]. Stool [612, 1915, 8873, 9751, 10127]. Stools [6338, 11880]. Stop [3435, 4391, 5870, 6914, 8060, 8154, 7781, 8237, 9452, 9953, 11151, 10925, 10869]. Stop-Valve [3435]. Stoppage [7953]. Stopper [1019, 8017, 8136]. Stoppers [3623, 3968]. Stopping [3455, 2926, 11077]. Storage [2386, 2077, 2121, 3394, 4329, 4547]. Store [10872]. Stored [2249]. Storehouse [8708]. Storing [1825, 1832, 4747, 4775, 4650, 4287]. Storm [1461, 3642, 3299, 3983, 10223, 12159, 11410]. Storm-Center [3642]. Storms [2123, 4033, 10309]. Stormy [1064]. Story [3354, 2646]. Stove [1052, 241, 403, 1918, 4756, 5905, 5287, 6369, 6206, 8112, 7590, 7784, 9365, 8806, 9538, 9015, 8897, 8983, 8848, 11278, 11177, 12005, 11651]. Stovepipe [463, 825, 3853, 9778, 11878]. Stoves [1291, 2506, 2218, 2963, 6411, 6368, 8842, 9722, 9487, 10221, 10742, 10924, 11543]. Straight [5020, 5309]. Straightening [6111, 10421]. Strain [9256]. Strainer [612, 2036, 6835, 7959]. Strange [680, 2531, 12047, 1667, 1798]. Strap [533, 11362, 10824, 11694]. Straps [5616]. Strasbourg [4286]. Strassburg [3894]. Straub [11644]. Straw [3454, 9538, 11152]. Streams [290]. Street [1164, 1181, 2470, 1933, 3500, 3546, 3498, 3551, 3917, 5780, 5370, 8805, 9698, 9699, 9770, 11333, 11898, 8291, 10741]. Streets [7909, 8523, 7934, 10260]. Strength [852, 5392, 12431]. Strengthening [6458]. Stress [2041]. Stretcher [2616, 3098, 3126, 6477, 7033, 6526, 9878, 9394, 11622]. Stretching [10255]. Strike [4771]. Strikes [2940]. Striking [734, 2709, 6662, 11766]. Strings [1356]. Strip [6061, 9484, 9233, 10644, 10826, 11000, 11329]. Striped [997]. Strong [542, 11716]. Strop [12370]. Struck [1258, 10544]. Structural [313]. Structure [1000, 1195, 2262, 3549, 11621]. Struggles [289]. Stuart [286]. Stub [10254]. Stud [6816, 9872]. Student [6932]. Studies [6263]. Study [1605, 6625, 8567]. Stuffing [8869]. Stump [5261, 9458, 10346]. Stumps [10206]. Sturgeon [1547, 4780, 1301]. Style [432, 1780, 9289, 9757]. Stylographic [485]. Sub [12602, 11644]. Sub-River [11644]. Sub-Stations [12602]. Subaqueous [1964]. Subjects [6084]. Submarine [745, 2874, 6893, 6761, 7043, 6829, 8654, 8444, 10027, 9234, 8862, 10474]. Submitted [8291]. Substances [2452, 4314]. Substitute [1361, 6960]. Substitution [4516]. Substitutions [764]. Suburban [1820, 7287, 8847, 10367]. Subway [8206, 10996, 11563]. Subways [8610, 10940, 11618, 12600]. Success [6616, 10639]. Successful [289, 4460, 10406, 12156, 2885]. Suction [6235]. Sudbury [12179]. Sudden [855]. Suez [7375]. Suffer [12535]. Suffocation [2855]. Sugar [3503, 3050, 2805, 4698, 4769, 4840, 4893, 5024, 5916, 7138, 8535, 9116, 9252, 9072, 10154, 12248, 11976]. Suggested [5308]. Suggestion [5160, 11507]. Suggestions [5094]. Suit [963, 5677, 8386, 7932, 9843, 11191, 11640, 5962].

Suits [9358, 10486]. Sulky [2368, 5617, 8240]. Sulphur [5490]. Sumac [954]. Summer [2939, 7498, 11524]. Sumptuous [8374]. Sun [6073, 646, 2123, 1910, 2007, 3371, 3647, 5218, 5161, 6340, 10438, 11748]. Sunfish [2348]. Sunken [9742]. Sunshade [11877]. Superb [10387]. Superficial [11193]. Superga [6187]. Superheating [5948]. Superior [9447]. Superiority [678]. Supersede [11462]. Supplement [5307, 5936]. Supplies [3825]. Supply [452, 1508, 3498, 3551, 4210, 4867, 5136, 8073, 9388, 9062, 11908, 12579, 12340, 11870, 12068, 4525, 8151, 9165, 12581, 12535]. Supplying [3068]. Support [4313, 4419, 5703, 5361, 5834, 6693, 6205, 8034, 8617, 8712, 8397, 9042, 9720, 9876, 12005, 11623, 11771, 11503, 11883]. Supporter [3571, 7621, 9429]. Supporting [2845, 6915, 9511, 12303]. Supports [3656]. Supposed [820]. Suppression [1438]. Supremacy [131]. Supreme [1403, 9419, 9818, 11947]. Sure [11379]. Surface [5413, 8082, 10509, 11863]. Surfaces [1080, 885, 101, 6003]. Surfacing [4948]. Surgery [7325]. Surgical [3322, 3878, 10049]. Surplus [59, 1961]. Survey [2024, 7584, 9868]. Surveyor [4389]. Suspended [2454, 2962, 3939, 9025]. Suspender [10413, 11450]. Suspenders [8257]. Suspending [11603]. Suspense [158]. Suspension [1698, 2416, 4599, 9998, 11999]. Sustained [34, 9166, 10241]. Swage [6228, 8472]. Swallowers [4308]. Swarmer [6554]. Swearing [3849]. Sweat [17]. Sweeper [8335, 9753]. Sweetest [12411]. Sweets [12411]. Swimmers [6472]. Swimming [299, 347]. Swindling [654]. Swing [9371, 10551, 10246, 12076]. Swinging [1489, 8287, 11290, 12213]. Swiss [5434]. Switch [320, 2884, 5546, 7912, 8358, 7570, 8673, 9724, 9169, 9703, 10762, 10244, 10496, 10571, 11279, 11876, 12257]. Switch-Stands [11876]. Switchboard [11533]. Switches [831, 12118]. Switzerland [10405]. Swivel [1469, 8927]. Swivel-Holder [1469]. Sword [1668, 4308]. Swordfish [968]. Swords [7494]. Sydney [5710, 696, 3723]. Sylvester [10081]. Symbolical [3073]. Symington [8867]. Synanceia [11600]. Syracuse [10661, 11797]. Syrrhaptes [11340]. System [642, 946, 323, 38, 1033, 564, 587, 666, 283, 1506, 1904, 1493, 1793, 1842, 2394, 1381, 3501, 3690, 2712, 2943, 4749, 4545, 4819, 5230, 5879, 5896, 6635, 6717, 6743, 7557, 8015, 8140, 8478, 8232, 9531, 9842, 9032, 9770, 10224, 11106, 11178, 11088, 10379, 10921, 10830, 11980, 12183, 11837, 12482, 12627, 12533, 12000, 1802, 8290, 1853, 1580]. Systems [3600, 4362].

T [2846]. T-Square [2846]. Table [95, 1244, 1949, 2671, 3095, 4580, 5758, 4925, 5125, 5966, 6556, 6362, 7538, 8470, 8894, 9961, 9428, 9169, 10259, 10528, 11719]. Tables [805, 2171, 7091, 12237]. Tablet [8710, 8987]. Tachyscope [12337]. Tackle [1324]. Tacoma [4597]. Tactics [12556]. Tag [4827, 8740]. Taguan [1367]. Tail [12390]. Tailed [940, 7097, 9970]. Tailor [10948]. Tails [2530]. Take [8956]. Take-Up [8956]. Taking [526, 1572, 5144]. Talcott [9730]. Talegallus [9535]. Talk [11792]. Tall [11145]. Tampico [4873]. Tamping [6599]. Tancarville [4949]. Tangent [6852, 12145]. Tank [1016, 6983, 11367, 12115]. Tanks [1176, 681, 767, 1233, 4310, 8429]. Tanner [1314, 708]. Tap [529, 6040, 11422]. Tape [2409, 8761]. Taper [6637]. Tapir [2713, 10256]. Tapping [3132]. Tar [3476, 7093]. Tarante [9865]. Tarantula [6016]. Tardily [3300]. Tare [12322]. Target [2087, 8366]. Tarpon [6511]. Tarsiers [8318]. Taste [2672]. Tattooed [10139]. Tax [8837]. Taxation [2528]. Taxi [8730]. Tay [504, 554, 7694]. Taylor [4140, 6528]. Tchebichef [6281]. Tchikoleff [1327]. Tea [1246, 714, 2167, 4847, 9219]. Teaching [794, 10379]. Technical [5418, 8817]. Technological [1965]. Technology [1749]. Teera [4193]. Teeth [4703, 6498, 7681, 9986, 12585]. Tehuantepec [1875, 7290, 7558]. Telegraph [358, 1097, 1133, 315, 1518, 2146, 1751, 1402, 2269, 2838, 2821, 3095, 3299, 6093, 5949, 5393, 7387, 7301, 6788, 7497]. Telegrapher [4771]. Telegraphing [3645, 6743, 8557]. Telegraphy [549, 551, 1075, 1722, 1916, 3367, 6804, 8015, 9653, 9098, 10224, 12482]. Telemeter [12533]. Telephone [1188, 284, 1061, 320, 1189, 318, 932, 1580, 2417, 1518, 1493, 1793, 1712, 2174, 1409, 3067, 2776, 3243, 2689, 4605, 3900, 4727, 4185, 4243, 3804, 3849, 4797, 5962, 5677, 4999, 5834, 5643, 7242, 7399, 6391, 7018, 6906, 7135, 7136, 7251, 7413, 8353, 8386, 7932, 8140, 7828, 8352, 8029, 8534, 9720, 10080, 9358, 10241, 12439, 11453, 12207, 11533, 262, 1471, 4819, 9419, 10843]. Telephones [1262, 1012, 2092, 7363, 7774]. Telephonic [323, 283, 1445, 1986, 3214, 3600]. Telephoning [3783]. Telephony [8179, 8864, 1802]. Telescope [1440, 3845, 4803, 10639, 11619, 10577, 3240]. Telescopes [6299, 6790, 430]. Telescopic [5877]. Telescoping [9778]. Tellurian [4122]. Telpher [7121]. Temminck [11263]. Temperature [855, 3256, 2925, 7019, 10229, 2124]. Tempering [3528, 8464]. Temple [1384]. Temporary [1654, 2567, 8088]. Ten [1974, 3066, 7474, 12057, 12350]. Ten-Ton [12057]. Tenant [3463]. Tendency [1905]. Tendering [212]. Tennis [7293]. Tenons [11830]. Tension [10316, 12450, 11681]. Tent [861, 241, 9427, 8990, 10258, 12005, 11472, 11445]. Tentering [4551]. Terebella [1193]. Teremakau [5194]. Termes [7941]. Terminus [4393, 4648, 11324]. Termite [10832]. Terraced [12034]. Terrestrial [813]. Terrible [2382, 9759]. Territory [12598]. Terror [11156]. Terry [5498]. Tessie [481]. Test [2473, 11244, 12033, 11895, 12049]. Tester [7035]. Testimony [3778]. Testing [215, 349, 2315, 1795, 2057, 3481, 3090, 2749, 3854, 4038, 3940, 6301, 7494, 7445, 9925, 10670, 10255, 5316, 1289]. Tests [77, 255, 4211, 6112, 10177]. Tetanus [5540]. Tetanus-Lock [5540]. Tether [9402, 10411]. Teutonic [12610]. Texas [1318, 8266, 10552]. Textile [3202, 3230, 3285, 3311, 3333, 3609, 3661, 4336, 3901]. Theater [3439, 5042]. Theaters [4442, 10593]. Theatrical [11146]. Their [2131, 2275, 3072, 3085, 357, 3272, 3847]. Thelyphonus [8790]. Theories [712]. Theory [373, 1406, 3272, 11353]. There [3124, 983]. Thermic [1188, 4970]. Thermo [6593, 10786, 492]. Thermo-Electric [492]. Thermo-Motor [10786]. Thermometer [833, 892, 8569, 9757, 12070, 11646, 9821]. Thermometers [1907, 6695].

Thermophones [1732]. Thermoscope [352]. Thermoscopic [9461]. Thibet [5195]. Thick [11275]. Thill [2594, 4216, 5824, 6017, 6334, 9260, 9225, 10666, 11446, 11739, 12073, 12301, 12604]. Thimble [11387]. Thin [6975, 8567]. Third [12215]. Thirty [8863, 10577]. Thirty-Six [10577]. Thistle [9489, 9545]. Thomas [10357, 10475, 12216, 12633]. Thompson [4243, 5242, 10830]. Thomson [2797, 3763, 5746]. Thorenberg [11157]. Thoroughbred [6794]. Thoughts [12386]. Thousand [3760, 5231, 6460]. Thrasher [968]. Thrashing [10970, 11152, 11746, 12077]. Thread [5005, 6603, 10981]. Threaded [4257]. Threading [11243, 12371]. Threads [3343, 5458]. Three [2546, 3327, 3187, 3747, 4196, 4487, 6819, 8909, 12372]. Three-Cylinder [3327]. Three-Sided [6819]. Thrifty [6200]. Throttle [10291]. Throwing [750, 3482]. Thrown [5169]. Thrust [6915]. Thunderer [634]. Thurston [7445]. Tice [373]. Ticket [234, 2155, 2344, 6759, 10335, 11695, 12432]. Tidal [1659]. Tide [5522]. Tie [1295, 4998, 6358, 6502, 6549, 9904, 9779, 10667, 10417, 10275, 10247, 11001, 11972, 12478]. Tied [7913]. Ties [8571]. Tiger [5318, 12459]. Tight [5578]. Tightener [2178, 2450, 4585, 6278, 6985, 8619, 10310, 12022]. Tightening [5616]. Tilden [8728]. Tile [2291, 3273, 6230, 8570, 8991, 12611, 2781]. Tilly [11828]. Timber [6651, 9226]. Time [300, 264, 1856, 1608, 6085, 7564, 8507, 11852]. Timepieces [33]. Times [11376]. Timing [9431]. Tin [529, 738, 921, 2481, 2063, 10132, 159]. Tin-Plating [2063]. Tinned [2481]. Tinsmith [5596]. Tip [5124, 7652, 8986, 12079]. Tire [2450, 3744, 6278, 6985, 11265]. Tires [556, 8084, 7756]. Tissandier [4607]. Titan [7096]. Toads [7626]. Toaster [9874]. Tobacco [1880, 4519, 7225, 9149, 10016, 9341, 10871, 11084, 11979]. Tobacco-Leaf [1880]. Toboggan [8234, 8791, 9963, 11286]. Together [10742]. Toilet [509, 1590, 2840, 10694, 10508]. Tomb [9854]. Tombstone [5217]. Tombstones [9879]. Ton [226, 3368, 6160, 6918, 7895, 9276, 8960, 10447, 10382, 12057, 12350]. Tong [123]. Tong-Tsing-Yo [123]. Tongs [1339, 191, 581, 1849, 5446, 6642, 7195, 9702, 9566, 10696, 11590]. Tongue [1800, 4419]. Tonguing [5641]. Tonkin [4465]. Tonnant [8314]. Tons [314]. Too [6381, 7486, 9289]. Tool [567, 848, 90, 326, 1128, 2544, 2129, 2217, 1542, 2541, 1919, 2642, 4504, 4004, 3852, 5639, 5899, 5684, 5165, 6115, 5319, 5661, 6615, 6479, 8189, 7779, 7864, 8233, 8311, 8961, 9339, 9424, 9850, 9749, 8762, 10342, 11085, 10502, 10895, 12316, 12631, 11767]. Tools [122, 2166, 2266, 1587, 6058, 5728, 10384, 12026]. Tooth [6068, 5100, 7489, 10261]. Toothed [936, 11195, 12459]. Toothed-Wheel [936]. Toothing [4721]. Top [983, 5871, 5850, 6368, 8034, 8745, 7961, 10105, 10112, 11481, 8511, 8227, 8215]. Topeka [8]. Tops [6205]. Topsails [9223]. Torchlight [5855]. Tornado [258, 11928]. Tornadoes [1676]. Toronto [12109]. Torpedo [2500, 2878, 3898, 5097, 6876, 7349, 6885, 6472, 7649, 7886, 8654, 8506, 8583, 9125, 9637, 9464, 8937, 8823, 9465, 9675, 11225, 10977, 10590, 11377, 10152, 11126, 11353, 10474, 12205, 12492, 12324, 11874, 11677]. Torpedoes [5773, 6767, 8463, 8418]. Torricelli [8268]. Torsion [4211]. Total [2898, 3846, 4256, 9628, 9263, 11215, 11748]. Toverenæ [6650]. Towage [4899]. Towel [7255]. Tower [2628, 7393, 6460, 7968, 8537, 8266, 11367, 6988, 11751, 11840, 12147, 12206, 12123, 11481]. Towle [2189]. Town [10404]. Towns [10830]. Townsmen [11897]. Toxicophidia [11705]. Toy [3767, 5367, 6150, 6438, 10156]. Toys [45, 1872, 8877, 8963, 9073, 9127, 9236, 9294, 9322, 9346, 9374, 9403, 9433, 9967]. Trace [4609, 5365, 6096, 5638, 6740]. Tracheotome [2903]. Tracing [9565]. Track [1651, 2658, 4441, 5312, 6288, 6993, 8282, 8232, 8590, 7835, 11970, 11809]. Tracks [10124, 12435]. Traction [3746, 4874, 5016, 6581, 6817, 8814, 9418, 9369, 11364]. Trade [756, 1025, 502, 1007, 317, 10, 1389, 1794, 1482, 2220, 1770, 2674, 3175, 3251, 3271, 3988, 5646, 5830, 5723, 6282, 6402, 6757, 8195, 12009]. Trademark [6404]. Trademarks [6033]. Trades [6331, 10684, 11235]. Trafalgar [9679]. Traffic [10636]. Train [5466, 6743, 9653]. Training [1404, 6916]. Trains [6715, 11181, 11314, 7375]. Tram [5194]. Trammel [9958]. Tramway [5927, 8192]. Tramways [11794, 3293]. Transactions [1892]. Transatlantic [12458]. Transcontinental [7860]. Transfer [3367, 11837]. Transferring [3431]. Transformation [2630, 8553]. Transformations [2174]. Transformers [8036]. Transfusion [3654]. Transit [7287, 8611, 9503, 10830, 3525, 3432, 3522]. Transits [6480, 3450]. Transmission [3392, 3649, 4656, 3867, 8216, 8638, 9343]. Transmitter [1712, 8534, 262, 1471]. Transmitting [5825, 6116, 5644, 10103, 10902]. Transom [9423]. Transparent [351]. Transplanting [10602, 11773]. Transportation [3242, 4749, 3736, 7777, 9930]. Trap [340, 583, 96, 1729, 4422, 4878, 4166, 4577, 5102, 5265, 5068, 5571, 5570, 6434, 7010, 7958, 9171, 10042, 8811, 11419, 10604, 10281, 11575, 11747]. Traps [4033]. Travelers [8740, 8934, 8837]. Traveling [797, 2014, 4515, 8312, 7961, 8079, 12479, 12350]. Tray [8785]. Treadle [806, 7993, 11194]. Treasure [4083]. Treating [6042]. Treatment [4572, 4460, 5756, 7163, 6709]. Tree [1336, 694, 4282, 5706, 6407, 7450, 8744, 8768, 10505, 7626]. Trees [1258, 1937, 3925, 6392, 9205, 9938, 9141, 10273, 12457, 1921, 934]. Trellis [8011]. Trestle [4718]. Trevelyan [12055]. Trial [1074, 4598, 9613, 10939, 12021, 11937]. Trials [6941, 10919, 12580, 12156]. Triangulation [7362]. Trichechus [8539]. Trichinosis [5462]. Trick [7650, 10505]. Tricks [10377]. Tricycle [3375, 3378, 4031, 4475, 5498, 6288, 6712, 6859, 6910, 10640, 10922, 10743, 11678]. Tricycles [3873, 8508]. Trimmer [1918]. Trimming [2757, 4673, 9363]. Trip [8242, 11221, 11202, 12409, 11937]. Triple [4970, 6481, 8065, 7759, 9594, 11426, 11704]. Tripod [9202, 12211]. Trireme [3984]. Triton [723]. Trocadero [4803]. Trombone [6091]. Tromp [8781].

Tropical [8196]. Trotting [2368]. Troubled [5004]. Troubles [5208, 12471]. Trough [7701, 8181, 10849, 11552]. Trousers [9394, 11694]. Trout [1597, 3149, 330]. Trubia [12479]. Truck [231, 2246, 2587, 4242, 5212, 5002, 5024, 7368, 7863, 7865, 11104, 11333]. Trucks [3326, 7036, 10896]. Truing [10207]. Trumpet [4285]. Trunk [4898, 5616, 4945, 7114, 6437, 9825, 9204, 9875, 11129, 11362]. Truss [5525, 7621, 11776, 11742]. Trust [8728]. Truths [2103]. Try [11569]. Tsing [123]. Tub [1467, 2839, 6663, 8812, 10340, 11599]. Tube [341, 5520, 5282, 6224, 6580, 7749, 8078, 11741, 12226]. Tubercle [2901]. Tubes [1879, 6300, 8478, 9363]. Tubing [10580]. Tubs [2036, 10947]. Tubular [4261]. Tug [611, 2594, 4608, 6740]. Tugboat [1827]. Tugs [6203]. Tumble [3485]. Tuning [6311, 9797, 9097]. Tuning-Forks [6311]. Tunnel [259, 2753, 3926, 4209, 4829, 6835, 8371, 11256, 11644, 768, 845, 1154, 1257, 85, 1745, 2284, 3144, 3019, 3577, 3924, 7966, 7462, 7953, 9121, 12121, 12090, 12429, 1338]. Tunneling [1010, 946, 5380, 7843, 8263]. Tunny [3655]. Turbine [4755, 3806, 5099]. Turbot [4806]. Turin [6187]. Turkey [9535]. Turner [4980, 12142]. Turning [1039, 5702, 7170, 8481, 9748, 10384, 11830]. Turntable [5621]. Turpentine [945, 6614]. Turret [4469]. Turtle [1692, 2546, 388, 8465]. Tweddell [38]. Tweezers [2192]. Twelve [5005, 5785, 9265, 10382]. Twelve-Wheel [5785]. Twelve-Wired [9265]. Twenty [3368, 10447]. Twenty-Five [10447]. Twenty-Ton [3368]. Twick [7643]. Twin [6876, 7302, 11606, 11656]. Twine [4219, 6167, 8254, 9534, 11108, 10572]. Twist [4126]. Twisting [2346]. Two [553, 343, 314, 1824, 3035, 4084, 5856, 5749, 6288, 7145, 7440, 7723, 10025, 9154, 9334, 10717, 10908, 12141, 2646]. Two-Horse [7440]. Two-Story [2646]. Two-Wheeled [12141]. Tympan [7785]. Tyndall [1511]. Type [21, 8238, 8494, 8342, 7675, 7671, 9006, 9450, 9982, 8779, 11028, 10672, 10463, 10851, 11696, 12254, 12393, 10700]. Types [8538, 12169]. Typesetting [11462]. Typewriter [11201, 12167]. Typhoid [11214]. Typical [524].

U.S. [34, 6574, 8981, 9637, 9194, 10552, 10953, 12565, 11954, 12502]. U.S.N. [9114]. Ultramarine [5831]. Umbrella [3648, 4313, 5366, 8397, 12582]. Umbrellas [4124]. Umbromania [10767]. Unburned [8454]. Unconstitutional [8837]. Uncut [4119]. Under-Ground [2687]. Underground [1985, 4241, 4362, 8288, 7747, 9800, 8953, 9023, 8757, 10487]. Undesirable [12091]. Undulating [12166]. Unebi’kan [9954]. Uneven [5067]. Unexplored [6507]. Unicorn [880]. Uniform [9224]. Uniformity [2747, 5919, 4943]. Union [5565, 9743, 9792, 9007, 2438, 4362, 5137, 6717]. Unipolar [3437]. Unique [8290, 6580]. United [9604, 11691, 948, 1505, 1437, 1608, 2960, 3329, 4694, 4565, 5732, 6837, 7668, 9743, 9389, 9007, 11220, 12380, 11705, 2685]. Universal [3507, 2843, 5319, 7826, 9481, 12133, 12499, 12048, 11862, 11752]. University [10196, 11936, 4286, 6852, 7376, 6961, 8496, 7933]. Unloading [6696, 6765, 10199, 12436]. Unsinkable [9264]. Unsuccessful [7463].

Untimely [956]. Unwholesome [2004]. Unwin [11052]. Unwise [1652]. Upon [1009, 2174, 5733, 9059, 2617, 4821, 4033, 8388, 10122]. Upper [3296]. Uppers [3652]. Upright [2837, 4240, 6002, 5925, 6600, 6883, 6834]. Upsetting [5775]. Upward [1881]. Urn [2044, 2093]. Urus [11136]. Use [1349, 1171, 1042, 1303, 313, 1591, 3322, 2886, 2979, 4623, 7126, 6940, 8388, 7567, 8027, 9744, 10761, 10951, 10947, 12317, 11948, 11743, 11621, 12256, 11926, 1077, 2476, 11860, 12032, 11802, 2605, 6892, 8526, 8118]. Used [2753, 4240]. Useful [436, 4657, 9334]. Uses [591, 3963, 5339, 5798, 9987, 10739, 10760, 11389, 159, 1642]. Utah [11302, 11301]. Utensil [10503]. Utensils [6310]. Utilization [2131, 1557, 1494, 6917, 12460, 4118]. Utilized [9065]. Utilizing [1562, 12155].

Va [8476, 11666]. Vaccination [1983, 1797]. Vacua [8296]. Vacuum [506, 1047, 3412, 5916, 5809, 12263, 8191, 8213, 8239, 9804]. Valises [6785]. Valkyrie [11762]. Valley [11562, 6857]. Value [2103, 4965, 8744, 3512]. Valve [71, 777, 3007, 3508, 3435, 4497, 4365, 4316, 4496, 5823, 6036, 5292, 6119, 5071, 6091, 7061, 7391, 6455, 7297, 6336, 7371, 6582, 6815, 7834, 8131, 7442, 9777, 9123, 9656, 10874, 10207, 11982, 11740, 11815, 12135, 12519, 11466, 1130]. Valve-Gear [777]. Valves [1129, 1660, 2859, 4167, 6875, 7866, 9777, 10291, 11258]. Vampires [2899]. Vanderbilt [3377, 9426]. Vanderveer [10166]. Vane [8935]. Vapor [2905, 5918, 8582, 10785]. Vaporizer [4259]. Vaporizing [7562]. Vapors [1432]. Vari [1714, 12083]. Variable [1358, 11150]. Variations [6031]. Variegation [1138]. Varius [12297]. Vark [514]. Vase [593, 541, 1619, 1595, 2517, 10699]. Vat [7268]. Vault [11218]. Vaults [4083, 8613, 9698]. Vegetable [4314, 9088, 11811, 11451, 4125]. Vegetables [7642]. Vegetation [12, 10334, 7816]. Vehicle [3554, 3530, 5873, 5850, 7145, 6621, 6882, 8734, 8422, 8505, 7837, 7697, 7700, 10105, 9037, 9876, 9906, 9990, 8813, 9226, 10665, 10717, 10847, 10848, 12141, 11564, 12281, 11503, 11930]. Vehicles [4423, 5358, 5501, 7014, 8618, 8671, 7891, 9117, 10014, 9251, 10202, 10690, 10315, 11999, 11854, 11697, 11877, 11933]. Veil [12067]. Velocipede [849, 1239, 645, 2179, 1828, 1567, 2614, 3257, 4897, 4950, 6719, 12053]. Velocipedes [5615]. Velocities [11819]. Velocity [2023]. Venetian [7519, 10304]. Venomous [505, 3870, 11705]. Ventilated [2508]. Ventilating [2816, 2943, 5468, 5896, 10013, 9482, 11055]. Ventilation [587, 486, 2498]. Ventilator [938, 232, 558, 1666, 3672, 5898, 6884, 7725, 7678, 9261, 8990, 10566, 10252, 10576, 11981, 12095]. Venus [146, 3525, 3432, 3450, 3522, 5877]. Vermin [9798]. Vernier [3168]. verrucosa} [11600]. Versus [1699, 2055, 2835]. Vertebrates [9450]. Vertical [3812, 4101, 7838, 7424, 9607, 11522]. Vessel [8504, 8754, 10590]. Vessels [182, 3347, 5993, 5990, 5896, 4919, 5436, 6014, 6342, 7375, 8393, 8286, 7525, 8395, 8700, 8445, 8819, 8981, 8937, 9930, 11077, 10253, 11103, 11407, 11833, 11671, 11517, 8287]. Vesuvius [11035, 10476, 10404, 11937]. Viaduct [4963, 2811, 7694, 11242]. Vibrating [6815, 8241, 7496]. Vibration [2397]. Vibrations [4054, 1804, 912, 9521, 11914, 11508]. Vibratory [11314]. Vicinity [10223]. Victor [12167]. Victoria [6315, 8266, 9822, 10651, 11791]. Victorian [5622]. Victories [2283]. Vienna [4184]. Views [3433, 5877, 6792, 6404, 11604, 12566]. Villa [11956]. Vine [969]. Vines [9537]. Violin [9797, 9097, 12390]. Viper [11009, 11289]. Vipers [6286]. Virchow [2704]. Virginia [11936]. Virot [10423]. Virulent [1286]. Vise [660, 532, 1441, 1665, 2273, 2660, 3125, 3945, 4603, 5735, 5297, 9631]. Vises [5552]. Vision [6579, 430]. Visit [1289, 11465]. Vivisection [2921, 6329]. Voice [1804]. Volatile [361]. Volcanic [8598]. Volcano [9806]. Volcanoes [1406]. Volney [10343]. Voltameter [2247]. Volumetric [2457]. Vortex [11782]. Voyage [7860, 11737]. Vs [655, 8418, 9418, 3869, 4573, 4030, 7290, 11967]. Vulcan [6408]. Vulcanite [975]. Vulcanizer [8707]. Vyrnwy [11676].

W [9831, 10343]. W. [1327, 11763]. W.S [4393]. W.T. [4597]. Wad [329]. Waddell [7147]. Waddington [10474]. Waesland [8479]. Wages [575, 4463]. Wagon [1330, 889, 2017, 2241, 4704, 4419, 5120, 5314, 5270, 6007, 5359, 5709, 6934, 7703, 8084, 7756, 7672, 9991, 8982, 9422, 11384, 10805, 10446, 10976]. Wagons [4315, 6114, 7440, 7589, 7644, 11285, 10227]. Wah [244]. Wahl [3412]. Waits [10242]. Wales [1965, 5443]. Walking [6281, 8089]. Wall [1319, 437, 241, 1994, 2309, 4083, 5760, 9754, 10823, 12546]. Walls [4145, 6737, 11387, 10921]. Walnut [9664]. Waltham [10322]. Want [3349, 4917]. Wanted [1102, 1562, 2286, 2167, 2639, 3348, 2631, 2981, 4166, 7638, 12338, 11536]. War [1974, 3673, 3347, 3989, 4364, 5866, 7241, 6544, 6356, 7164, 8452, 8981, 9417, 9930, 8803, 9277, 9844, 11300, 10177, 11850, 12481, 11779, 12620, 11721, 11551, 12254, 11491, 11517, 11535, 3163, 3691]. Ward [8838]. Wardrobe [4898, 5029, 8161, 10101]. Ware [10108]. Warehouse [231]. Warfare [6829, 11353]. Warmer [11308]. Warner [8952]. Warnings [1461]. Warping [5073]. Warrior [6692]. Wars [11969]. Warwick [7037]. was [5514, 9954]. Wash [6036, 7831, 7442, 8901, 12322, 7149]. Wash-Bottle [7149]. Washboard [12302]. Washer [772, 690, 9230, 12364]. Washing [1111, 1709, 2796, 3220, 7174, 6764, 7401, 7143, 7515, 8285, 10077, 10075, 11905, 12495]. Washingtion [11691]. Washington [2438, 6201, 7640, 12598, 12082, 11736]. Wasp [6673]. Wasps [2275]. Waste [1385, 1494, 5210, 7077, 9010, 10900, 10869]. Watch [741, 2086, 1410, 3533, 4196, 5297, 5992, 8500, 9793, 9286, 9092, 9680, 10322, 12347, 6108]. Watchcase [8615]. Watches [1014, 678, 2570, 5635, 8500, 8763, 9827, 11354, 8178]. Water [456, 340, 341, 843, 834, 290, 111, 1166, 1385, 1462, 1942, 2023, 1756, 1890, 2239, 1638, 1609, 1847, 1822, 2331, 2235, 1508, 2556, 3083, 3276, 3356, 3226, 3667, 2672, 3068, 2925, 3124, 3369, 4171, 3985, 4280, 3824, 4276, 4033, 4008, 3947, 4032, 3809, 3948, 4210, 4037, 3872, 3825, 4867, 6149, 5519, 5874, 5578, 5411, 5136, 5183, 5571, 5271, 5321, 4997, 7205, 6386, 6527, 6533, 6805, 6559, 6333, 6983, 6782, 6917, 6332, 7012, 7893, 7492, 7777, 7557, 8585, 8537, 8471, 8399, 8292, 7833, 8151, 8684, 7953, 7873, 8468]. Water [8565, 8073, 8266, 9388, 8843, 8954, 9756, 9165, 9490, 9283, 9151, 9065, 9332, 11419, 10380, 11367, 11106, 10867, 10689, 11127, 10844, 11088, 10441, 10809, 10565, 10686, 11908, 12260, 12581, 12180, 12569, 12185, 12474, 12579, 12190, 12429, 11973, 12535, 12340, 11676, 11870, 12068, 4525, 8290, 7735, 7870, 10451, 10422, 12069, 8107]. Water-Tight [5578]. Waterhouse [10905]. Watering [1112, 687, 1826, 5658, 7701, 11552]. Waterproof [7513, 11260]. Waters [985, 895, 5004, 7179, 7668]. Waterspouts [1676]. Watertown [2060, 2556, 9925]. Waterwheels [7196]. Waterworks [2394, 1627]. Watson [7289]. Watt [7844, 10764]. Wave [2970, 6044, 12155]. Waves [9065, 10309]. Way [9251, 2927]. Weak [9844]. Weakness [710]. Weaner [9994]. Wear [10014, 11933]. Weather [332, 3499, 2854, 6061, 9484, 8935, 9233, 11275, 10644, 10826, 11000, 11329]. Weaver [2020]. Weaving [1124, 2422]. Web [12193, 12623]. Webbed [6199]. Weber [7317]. Webs [2314]. Webster [11191]. Wedge [7339, 10162, 8599]. Weed [3485]. Weeder [6669]. Week [314]. Weeke [12000]. Weeks [158]. Weigher [10292]. Weighing [2904, 8338, 10282, 10182, 10789, 8466]. Weight [4658, 5619, 5736, 6672, 8987]. Weighted [615]. Weights [4303]. Weiss [9770]. Welded [10580]. Welding [7704, 9817]. Well [61, 228, 2148, 3369, 3962, 5572, 6041, 6005, 7557, 7571, 9898, 9843, 9166, 9819, 10545, 11656]. Wells [51, 1261, 564, 1542, 4502, 4276, 7337, 7298, 8293, 9388, 11301, 10452, 11899, 4525, 11562, 8697, 8595, 3723]. Were [4196]. Werner [8820]. Wesley [11366]. West [4648, 3720]. Western [1157, 1096, 603, 4362, 6717]. Weston [5796, 9717]. Whale [194, 968, 4501, 7327, 6486, 8479, 10139, 2759, 832]. Whalebone [5295]. Whales [7709]. Whaling [7709]. Wharf [10382]. Wharve [12233]. Wheat [7418]. Wheel [936, 2405, 2316, 2969, 4755, 4654, 3744, 4874, 5637, 5522, 5576, 5902, 6099, 5785, 6810, 6386, 7395, 6305, 7012, 7618, 8647, 7473, 8477, 8421, 8616, 7700, 9661, 9369, 9314, 10136, 11054, 11131, 11195, 10620, 12326, 11477, 12234, 12190]. Wheelbarrow [1184]. Wheeled [4487, 4084, 7145, 10717, 12141]. Wheelers [1157]. Wheels [1014, 975, 1249, 2062, 6161, 5492, 5613, 4937, 5657, 5016, 6222, 9449, 9332, 10863, 10665, 1593]. Where [1769, 2911]. Whey [3504]. Which [4196]. Whiffletree [1519, 8587]. While [6743, 1982]. Whip [4446, 10695]. Whistle [6741]. Whistling [7077]. Whitcomb [3505]. White [251, 1811, 1407, 3173, 5614, 6649, 7092, 7140, 6562, 8388, 7726, 6188, 10574, 12610]. White-Footed [3173, 6188]. White-Nosed [6562]. Whittled [7943]. Whitworth [9359]. Who [59, 7736, 1245]. Wholesale [4567]. Wholesome [3781]. Whose [1460]. Wick [1918, 8078]. Wide [12163]. Widths [3964]. Wilcox [341]. Wild [9279, 10366]. Wilderness [4143]. Will [6147, 6830, 8497, 9096, 11820]. Will-O’-the-Wisp [11820]. Willard [12287]. William [2836, 3981, 12236].

Williamsport [746]. Wilson [3084]. Wimshurst [12306]. Wind [1807, 4491, 3983, 4747, 4775, 4287, 4329, 4547, 4650, 5028]. Winding [1410, 8568, 9827, 11011]. Windlass [4975, 5474, 5572]. Windmill [272, 4236, 8586, 11367, 10501, 10926, 11388]. Window [118, 1054, 2196, 3506, 2880, 4730, 5021, 5416, 5076, 6382, 7167, 7365, 6884, 8701, 7755, 7485, 8703, 7625, 9701, 10925, 10715, 11393, 10576, 10825, 12213, 11974, 12453, 12095]. Windows [9485, 12074]. Wine [9033]. Wing [11389]. Winkle [1450]. Winter [3550, 4678, 7189, 9460]. Wire [800, 936, 225, 2269, 3026, 4899, 6093, 5894, 5805, 5194, 5338, 6428, 6277, 7346, 7081, 6187, 6526, 6339, 6479, 8619, 7444, 8062, 8589, 8760, 9323, 10436, 10824, 11356, 11726, 11622, 12022, 866, 1077, 11978]. Wired [9265]. Wires [1985, 1402, 3295, 2838, 3939, 4241, 5160, 6687, 7399, 6855, 8152, 8134, 8288, 8497, 10073, 10760, 11997, 11639]. Wireworms [6013]. Wisconsin [7708]. Wisp [11820]. Within [8863, 6933]. Without [3370, 4498, 3929, 6015, 5927, 1677, 8084, 10636, 10638, 11202]. Woehler [3735]. Wolf [5600, 4214]. Wolfe [12257]. Women [1140]. Wonderful [6864, 7402]. Wood [1912, 2487, 2422, 3503, 3484, 4869, 5799, 5487, 7084, 6638, 8031, 7871, 9455, 9748, 9718, 11568, 11475, 11631]. Wood-Sawing [11475]. Wood-Working [3484]. Woodcock [12168]. Woodpecker [7207]. Woodpeckers [2112, 2821]. Woodwork [1206]. Woodworkers [11010]. Woodworking [8061]. Wool [577, 877, 914, 3373, 7151, 7871, 9010, 11905]. Worcester [3043]. Work [555, 575, 3344, 3550, 3915, 3941, 4921, 5306, 5491, 6133, 6381, 6809, 7760, 8101, 8419, 9617, 9504, 9616, 9308, 9361, 9096, 9592, 9953, 11795, 11516, 8357, 11061, 4916]. Work-Benches [4921]. Worker [2563, 4309, 6665, 11724, 12191]. Workers [821]. Working [3484, 5310, 5487, 5384, 6273, 9912, 9102, 12534]. Workingman [4544]. Workingman’s [6185]. Workman [895]. Workmanship [5987]. Workmen [75]. Works [1630, 3296, 2746, 3667, 4190, 4828, 7492, 8107, 9842, 8911, 10981, 10661, 10407, 10150, 12484, 8290, 4168, 8037, 7557, 11106, 11984, 12429, 11676, 6108, 7720]. Workshop [1749, 7844]. World [1909, 4965, 7665, 8980, 9113, 9220, 9505, 9869, 10153, 10305, 10408, 10709, 11074, 11171, 11170, 11494, 11994, 12020, 604, 550, 1101, 1230, 3105, 10832]. Worm [482, 2405, 2409, 4625, 6733, 724, 1232]. Worn [2668, 9777, 12326]. Worrell [8083]. Worth [868]. Wound [2122]. Wrapper [2364]. Wrapping [5338]. Wreck [10720]. Wrecks [11692]. Wrench [4529, 4123, 5211, 5461, 6808, 6617, 6938, 8649, 7890, 8104, 8681, 9339, 9566, 9180, 10445, 10806, 10828, 10972, 11261, 12278]. Wretchedness [3847]. Wringer [4232]. Wringers [3092]. Wrist [7089]. Wrist-Held [7089]. Writer [8238, 8342, 7671, 9006, 10672, 10700]. Writers [11696, 12393]. Writing [211, 2005, 5290, 7446, 8494, 8710, 7513]. Written [1769]. Wrought [5384, 5282]. Wyoming [1984]. Wyville [2797].

XII [11197].

Yacht [4139, 6572, 6827, 7746, 8582, 10011, 9489, 9575, 8834, 8905, 9545, 9426]. Yachting [5473]. Yachts [5844]. Yankee [524]. Yard [1074, 12144, 12397, 12251]. Yarn [3162]. Year [1354, 555, 1974, 3344, 3066, 4189]. Years [3106, 8463, 9682]. Yeast [6461]. Yellow [658, 132, 852, 2123, 11409, 1360]. Yield [451]. Yo [123]. Yoke [10336]. York [61, 573, 134, 257, 405, 180, 452, 1231, 1152, 1560, 2282, 1935, 1630, 2089, 2746, 3105, 3498, 3369, 4148, 3967, 4393, 4648, 4210, 4867, 4994, 5167, 6097, 7107, 8650, 7931, 8341, 8151, 8118, 7799, 7747, 8073, 8863, 10070, 9165, 11253, 10707, 10800, 10343, 12581, 11653, 12251, 12274, 12535, 12020, 11736, 12068, 1259, 635, 793, 1150, 345, 111, 1650, 2287, 1698, 1721, 3046, 3350, 3189, 3019, 3393, 3551, 3367, 4697, 4418, 4841, 4434, 3917, 5497, 5136, 7056, 6828, 6282, 8698, 8206, 8441, 8610, 8387, 8817, 9196, 10021, 8847]. York [10308, 10223, 10404, 10852, 11007, 11908, 11440, 12471, 11618, 12044, 12579, 12340, 11639, 11928, 12082, 11870, 8291]. York-Town [10404]. Yorktown [10476, 12021, 11954]. Young [2713, 4634, 5734, 5551, 6452, 10256, 512, 607, 828, 911, 2863]. Youth [2147]. Yucatan [5090, 5089, 5122, 5168, 5220, 5243].

Zarabatana [8715]. Zealand [8598]. Zebra [3677, 4214]. Zeromotor [2421]. Zirconia [11254]. Zodiacal [1459]. Zoo [1221]. Zoogyroscope [404]. Zoological [112, 7737, 7790, 9071, 9099, 10673, 11839, 12083, 5551, 9886, 10025, 10908, 10809, 10256, 11816]. Zschiesche [5238].

References

Anonymous:1880:BPa

[1] Anonymous. Business and personal. Scientific American, 42(1):11, January 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

       (electronic). URL http://www.nature.com/scientificamerican/
       journal/v42/n1/pdf/scientificamerican01031880-11.pdf.

Anonymous:1880:BNBa

[2] Anonymous. Books: New books and publications. Scientific American, 42(1):11, January 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

       scientificamerican/journal/v42/n1/pdf/scientificamerican01031880-
       11a.pdf.

Anonymous:1880:NQa

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n1/pdf/scientificamerican01031880-11b.pdf.

Anonymous:1880:IIa

[4] Anonymous. Index of inventions. Scientific American, 42(1):12, January 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n1/pdf/scientificamerican01031880-12a.pdf.

Anonymous:1880:PNPa

[5] Anonymous. Patent notes: Patents. Scientific American, 42(1):12, January 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n1/pdf/scientificamerican01031880-12b.pdf.

Anonymous:1880:FN

[6] Anonymous. A few novelties. Scientific American, 42(1):1, January 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n1/pdf/scientificamerican01031880-1.pdf.

Anonymous:1880:NPL

[7] Anonymous. A new policy in law. Scientific American, 42(1):2, January 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n1/pdf/scientificamerican01031880-2.pdf.

Anonymous:1880:ATS

[8] Anonymous. The Atchison, Topeka and Santa Fe Railroad over the Raton Mountains. Scientific American, 42(1):2, January 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n1/pdf/scientificamerican01031880-2a.pdf.

Anonymous:1880:DDD

[9] Anonymous. Dr. Daniel Draper’s contributions to meteorology. Scientific American, 42(1):2–3, January 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v42/n1/pdf/scientificamerican01031880-
2b.pdf.

Anonymous:1880:TMC

[10] Anonymous. Trade marks in Congress. Scientific American, 42(1):3, January 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n1/pdf/scientificamerican01031880-3.pdf.

Anonymous:1880:DB

[11] Anonymous. A dissertation bureau. Scientific American, 42(1):3, January 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n1/pdf/scientificamerican01031880-3a.pdf.

Anonymous:1880:IEV

[12] Anonymous. Influence of electricity on vegetation. Scientific American, 42(1):3, January 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n1/pdf/scientificamerican01031880-
3b.pdf.

Anonymous:1880:CRI

[13] Anonymous. The common reward of intelligence and energy. Scientific American, 42(1):3–4, January 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n1/pdf/scientificamerican01031880-
3c.pdf.

Anonymous:1880:HDA

[14] Anonymous. Horizontal double-acting force pump. Scientific American, 42(1):4, January 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n1/pdf/scientificamerican01031880-
4.pdf.

Anonymous:1880:GAD

[15] Anonymous. Gopher and ant destroyer. Scientific American, 42(1):4, January 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n1/pdf/scientificamerican01031880-4a.pdf.

Anonymous:1880:EIa

[16] Anonymous. Engineering inventions. Scientific American, 42(1):4–5, January 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n1/pdf/scientificamerican01031880-4b.pdf.

Anonymous:1880:NSa

[17] Anonymous. Novel hat sweat. Scientific American, 42(1):5, January 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n1/pdf/scientificamerican01031880-5.pdf.

Anonymous:1880:IEG

[18] Anonymous. Improvement in eye-glasses. Scientific American, 42(1):5–6, January 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n1/pdf/scientificamerican01031880-
5a.pdf.

Anonymous:1880:NMI

[19] Anonymous. New musical instrument. Scientific American, 42(1):6, January 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n1/pdf/scientificamerican01031880-6.pdf.

Anonymous:1880:MIa

[20] Anonymous. Mechanical inventions. Scientific American, 42(1):6–7, January 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n1/pdf/scientificamerican01031880-6a.pdf.

Anonymous:1880:TCM

[21] Anonymous. Type casting machine. Scientific American, 42(1):8, January 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n1/pdf/scientificamerican01031880-8.pdf.

Anonymous:1880:MIb

[22] Anonymous. Miscellaneous inventions. Scientific American, 42(1):8, January 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n1/pdf/scientificamerican01031880-8a.pdf.

Anonymous:1880:BPb

[23] Anonymous. The Brazilian porcupine. Scientific American, 42(1):9, January 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n1/pdf/scientificamerican01031880-9.pdf.

Anonymous:1880:SES

[24] Anonymous. Snake eating snake. Scientific American, 42(1):9–10, January 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n1/pdf/scientificamerican01031880-9a.pdf.

Anonymous:1880:CRa

[25] Anonymous. Communications received. Scientific American, 42(1):12, January 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n1/pdf/scientificamerican01031880-12.pdf.

Anonymous:1880:Ca

[26] Anonymous. Correspondence. Scientific American, 42(10):148, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n10/pdf/scientificamerican03061880-148.pdf.

Anonymous:1880:BPc

[27] Anonymous. Business and personal. Scientific American, 42(10):155, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n10/pdf/scientificamerican03061880-155.pdf.

Anonymous:1880:NQb

[28] Anonymous. Notes and queries. Scientific American, 42(10):155, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n10/pdf/scientificamerican03061880-155a.pdf.

Anonymous:1880:IIb

[29] Anonymous. Index of inventions. Scientific American, 42(10):155–156, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n10/pdf/scientificamerican03061880-155c.pdf.

Anonymous:1880:PNPb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n10/pdf/scientificamerican03061880-156.pdf.

Anonymous:1880:PNPc

[31] Anonymous. The proposed new Patent Law. Scientific American, 42(10):144, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n10/pdf/scientificamerican03061880-
144.pdf.

Anonymous:1880:AEGa

[32] Anonymous. Approximate economy of gas and electric lighting. Scientific American, 42(10):144, March 6, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v42/
n10/pdf/scientificamerican03061880-144a.pdf.

Anonymous:1880:CTa

[33] Anonymous. The certification of timepieces. Scientific American, 42(10):145, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n10/pdf/scientificamerican03061880-
145.pdf.

Anonymous:1880:IDU

[34] Anonymous. Important decision by the U.S. Circuit Court — the Page Electrical Patent sustained. Scientific American, 42(10):145, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n10/pdf/scientificamerican03061880-145a.pdf.

Anonymous:1880:ISSa

[35] Anonymous. The inspection of small steamers. Scientific American, 42(10):145, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n10/pdf/scientificamerican03061880-
145b.pdf.

Anonymous:1880:ADa

[36] Anonymous. Artificial diamonds. Scientific American, 42(10):145, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n10/pdf/scientificamerican03061880-145c.pdf.

Anonymous:1880:NMa

[37] Anonymous. A new meteorite. Scientific American, 42(10):145, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n10/pdf/scientificamerican03061880-145d.pdf.

Anonymous:1880:HRTa

[38] Anonymous. Hydraulic riveting — Tweddell system. Scientific American, 42(10):146, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n10/pdf/scientificamerican03061880-
146.pdf.

Anonymous:1880:ICG

[39] Anonymous. Improved carpenter’s gauge. Scientific American, 42(10):147, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n10/pdf/scientificamerican03061880-
147.pdf.

Anonymous:1880:NFP

[40] Anonymous. New fluid propeller or motor. Scientific American, 42(10):147, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n10/pdf/scientificamerican03061880-
147a.pdf.

Anonymous:1880:NLE

[41] Anonymous. New lawn edge mower. Scientific American, 42(10):147, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n10/pdf/scientificamerican03061880-147b.pdf.

Anonymous:1880:AINa

[42] Anonymous. American industries, no. 33. Scientific American, 42(10):149, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n10/pdf/scientificamerican03061880-
149.pdf.

Anonymous:1880:ISPa

[43] Anonymous. Improved steam packing. Scientific American, 42(10):150, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n10/pdf/scientificamerican03061880-150.pdf.

Anonymous:1880:PLa

[44] Anonymous. Proa Ladronia. Scientific American, 42(10):150, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n10/pdf/scientificamerican03061880-150a.pdf.

Anonymous:1880:ST

[45] Anonymous. Scientific toys. Scientific American, 42(10):151, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n10/pdf/scientificamerican03061880-151.pdf.

Anonymous:1880:RIa

[46] Anonymous. Recent inventions. Scientific American, 42(10):151–152, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n10/pdf/scientificamerican03061880-151a.pdf.

Anonymous:1880:ISPb

[47] Anonymous. Improved shovel plow point. Scientific American, 42(10):152, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n10/pdf/scientificamerican03061880-
152.pdf.

Anonymous:1880:HC

[48] Anonymous. The hairy crab. Scientific American, 42(10):153, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n10/pdf/scientificamerican03061880-153.pdf.

Anonymous:1880:PCa

[49] Anonymous. The Phylloxera in California. Scientific American, 42(10):153–154, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n10/pdf/scientificamerican03061880-
153a.pdf.

Anonymous:1880:EIb

[50] Anonymous. Engineering inventions. Scientific American, 42(10):154, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n10/pdf/scientificamerican03061880-154.pdf.

Anonymous:1880:AWCa

[51] Anonymous. Artesian wells for Colorado. Scientific American, 42(10):154, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n10/pdf/scientificamerican03061880-
154a.pdf.

Anonymous:1880:CRb

[52] Anonymous. Communications received. Scientific American, 42 (10):155, March 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n10/pdf/scientificamerican03061880-
155b.pdf.

Anonymous:1880:BPd

[53] Anonymous. Business and personal. Scientific American, 42(11):171, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n11/pdf/scientificamerican03131880-171.pdf.

Anonymous:1880:BNBb

[54] Anonymous. Books: New books and publications. Scientific American, 42(11):171, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n11/pdf/scientificamerican03131880-
171a.pdf.

Anonymous:1880:NQc

[55] Anonymous. Notes and queries. Scientific American, 42(11):171, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n11/pdf/scientificamerican03131880-171b.pdf.

Anonymous:1880:IIc

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n11/pdf/scientificamerican03131880-171d.pdf.

Anonymous:1880:PNPd

[57] Anonymous. Patent notes: Patents. Scientific American, 42(11):172, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n11/pdf/scientificamerican03131880-172.pdf.

Anonymous:1880:AINb

[58] Anonymous. American industries, no. 34. Scientific American, 42(11):159, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n11/pdf/scientificamerican03131880-
159.pdf.

Anonymous:1880:WSH

[59] Anonymous. Who shall hold the surplus? Scientific American, 42(11):160, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n11/pdf/scientificamerican03131880-
160.pdf.

Anonymous:1880:RSR

[60] Anonymous. Regulation of shifting river channels. Scientific American, 42(11):160, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n11/pdf/scientificamerican03131880-
160a.pdf.

Anonymous:1880:DWN

[61] Anonymous. The drive well in New York City. Scientific American, 42(11):161, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n11/pdf/scientificamerican03131880-
161.pdf.

Anonymous:1880:FLC

[62] Anonymous. Ferdinand De Lesseps and the Chagres Canal. Scientific American, 42(11):161, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n11/pdf/scientificamerican03131880-
161a.pdf.

Anonymous:1880:NFD

[63] Anonymous. A new fruit drier. Scientific American, 42(11):162, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n11/pdf/scientificamerican03131880-162.pdf.

Anonymous:1880:IRGa

[64] Anonymous. Improved roller gate. Scientific American, 42(11):162, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n11/pdf/scientificamerican03131880-162a.pdf.

Anonymous:1880:IMSa

[65] Anonymous. Improved mould for sewer building. Scientific American, 42(11):162, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n11/pdf/scientificamerican03131880-
162b.pdf.

Anonymous:1880:NMC

[66] Anonymous. Novel method of carpeting stairs. Scientific American, 42(11):162, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n11/pdf/scientificamerican03131880-
162c.pdf.

Anonymous:1880:NSMa

[67] Anonymous. A novel small motor. Scientific American, 42(11):163–164, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n11/pdf/scientificamerican03131880-163.pdf.

Anonymous:1880:CSa

[68] Anonymous. Cheap spectroscope. Scientific American, 42(11):164, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n11/pdf/scientificamerican03131880-164.pdf.

Anonymous:1880:NGS

[69] Anonymous. New gas saver. Scientific American, 42(11):164–165, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n11/pdf/scientificamerican03131880-164a.pdf.

Anonymous:1880:AIa

[70] Anonymous. Agricultural inventions. Scientific American, 42(11):165, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n11/pdf/scientificamerican03131880-165.pdf.

Anonymous:1880:ICV

[71] Anonymous. Improved controlling valve for engines. Scientific American, 42(11):166, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n11/pdf/scientificamerican03131880-
166.pdf.

Anonymous:1880:NPD

[72] Anonymous. New potato digger. Scientific American, 42(11):166–167, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n11/pdf/scientificamerican03131880-166a.pdf.

Anonymous:1880:LM

[73] Anonymous. The leaf Mormolyce. Scientific American, 42(11):167, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n11/pdf/scientificamerican03131880-167.pdf.

Anonymous:1880:NHNa

[74] Anonymous. Natural history notes. Scientific American, 42(11): 167–169, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n11/pdf/scientificamerican03131880-
167a.pdf.

Anonymous:1880:CFW

[75] Anonymous. The condition of French workmen. Scientific American, 42(11):169, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n11/pdf/scientificamerican03131880-
169.pdf.

Anonymous:1880:ICSa

[76] Anonymous. Improved cam for stamp mills. Scientific American, 42(11):169–170, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n11/pdf/scientificamerican03131880-
169a.pdf.

Anonymous:1880:ELTa

[77] Anonymous. Electric lamp tests. Scientific American, 42(11):170, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n11/pdf/scientificamerican03131880-170.pdf.

Anonymous:1880:CRc

[78] Anonymous. Communications received. Scientific American, 42 (11):171, March 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n11/pdf/scientificamerican03131880-
171c.pdf.

Anonymous:1880:IRIa

[79] Anonymous. Inventions: Recent inventions. Scientific American, 42(12):180, March 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n12/pdf/scientificamerican03201880-
180b.pdf.

Anonymous:1880:BPe

[80] Anonymous. Business and personal. Scientific American, 42(12):186, March 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n12/pdf/scientificamerican03201880-186.pdf.

Anonymous:1880:BNBc

[81] Anonymous. Books: New books and publications. Scientific American, 42(12):186, March 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n12/pdf/scientificamerican03201880-
186a.pdf.

Anonymous:1880:NQd

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n12/pdf/scientificamerican03201880-186b.pdf.

Anonymous:1880:IId

[83] Anonymous. Index of inventions. Scientific American, 42(12):187–188, March 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n12/pdf/scientificamerican03201880-187.pdf.

Anonymous:1880:PNPe

[84] Anonymous. Patent notes: Patents. Scientific American, 42(12):188, March 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n12/pdf/scientificamerican03201880-188.pdf.

Anonymous:1880:SGR

[85] Anonymous. The St. Gothard Railway Tunnel. Scientific American, 42(12):176, March 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n12/pdf/scientificamerican03201880-
176.pdf.

Anonymous:1880:PBN

[86] Anonymous. The Patent Bill now before the Senate. Scientific American, 42(12):176–177, March 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n12/pdf/scientificamerican03201880-
176a.pdf.

Anonymous:1880:NMb

[87] Anonymous. The new meteorite. Scientific American, 42(12):177, March 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n12/pdf/scientificamerican03201880-177.pdf.

Anonymous:1880:FSI

[88] Anonymous. The first step in invention. Scientific American, 42 (12):177–178, March 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n12/pdf/scientificamerican03201880-
177a.pdf.

Anonymous:1880:IBa

[89] Anonymous. Improvement in boilers. Scientific American, 42(12):178, March 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n12/pdf/scientificamerican03201880-178.pdf.

Anonymous:1880:NSTa

[90] Anonymous. New soldering tool. Scientific American, 42(12):178, March 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n12/pdf/scientificamerican03201880-178a.pdf.

Anonymous:1880:MIc

[91] Anonymous. Mechanical inventions. Scientific American, 42(12):178, March 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n12/pdf/scientificamerican03201880-178b.pdf.

Anonymous:1880:NCD

[92] Anonymous. New car door fastener. Scientific American, 42(12):179, March 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n12/pdf/scientificamerican03201880-179.pdf.

Anonymous:1880:SPB

[93] Anonymous. Speaking picture book. Scientific American, 42(12):179, March 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n12/pdf/scientificamerican03201880-179a.pdf.

Anonymous:1880:EIc

[94] Anonymous. Engineering inventions. Scientific American, 42(12):179, March 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n12/pdf/scientificamerican03201880-179b.pdf.

Anonymous:1880:ITCa

[95] Anonymous. Ironing table, clothes drier, and step ladder. Scientific American, 42(12):179, March 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n12/pdf/scientificamerican03201880-
179c.pdf.

Anonymous:1880:NMG

[96] Anonymous. New mole and gopher trap. Scientific American, 42 (12):179–180, March 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n12/pdf/scientificamerican03201880-
179d.pdf.

Anonymous:1880:ACD

[97] Anonymous. Apparatus for continuous displacement. Scientific American, 42(12):180, March 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n12/pdf/scientificamerican03201880-
180.pdf.

Anonymous:1880:DOC

[98] Anonymous. Drying oven with constant draught of dry air. Scientific American, 42(12):180, March 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n12/pdf/scientificamerican03201880-
180a.pdf.

Anonymous:1880:AINc

[99] Anonymous. American industries. — no. 35. Scientific American, 42(12):181, March 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n12/pdf/scientificamerican03201880-
181.pdf.

Anonymous:1880:IRA

[100] Anonymous. Improved riding attachment for cultivators. Scientific American, 42(12):182, March 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n12/pdf/scientificamerican03201880-
182.pdf.

Anonymous:1880:NCCa

[101] Anonymous. Non-conducting covering for heated surfaces. Scientific American, 42(12):182, March 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n12/pdf/scientificamerican03201880-
182a.pdf.

Anonymous:1880:M

[102] Anonymous. Medusle. Scientific American, 42(12):183–184, March 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n12/pdf/scientificamerican03201880-183.pdf.

Anonymous:1880:MId

[103] Anonymous. Miscellaneous inventions. Scientific American, 42 (12):184–185, March 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n12/pdf/scientificamerican03201880-
184.pdf.

Anonymous:1880:Cb

[104] Anonymous. Correspondence. Scientific American, 42(13):197, March 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n13/pdf/scientificamerican03271880-197.pdf.

Anonymous:1880:BPf

[105] Anonymous. Business and personal. Scientific American, 42(13):202, March 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n13/pdf/scientificamerican03271880-202.pdf.

Anonymous:1880:NQe

[106] Anonymous. Notes and queries. Scientific American, 42(13):202–203, March 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n13/pdf/scientificamerican03271880-202a.pdf.

Anonymous:1880:IIe

[107] Anonymous. Index of inventions. Scientific American, 42(13):203, March 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n13/pdf/scientificamerican03271880-203a.pdf.

Anonymous:1880:PNPf

[108] Anonymous. Patent notes: Patents. Scientific American, 42(13):204, March 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n13/pdf/scientificamerican03271880-204.pdf.

Anonymous:1880:SRI

[109] Anonymous. Ship railways for Isthmus crossings. Scientific American, 42(13):192, March 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n13/pdf/scientificamerican03271880-
192.pdf.

Anonymous:1880:ENC

[110] Anonymous. Evils of neglecting cold in the head. Scientific American, 42(13):192, March 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n13/pdf/scientificamerican03271880-
192a.pdf.

Anonymous:1880:PMN

[111] Anonymous. A plan to make New York a fresh water port. Scientific American, 42(13):192, March 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n13/pdf/scientificamerican03271880-
192b.pdf.

Anonymous:1880:ZN

[112] Anonymous. A zoological novelty. Scientific American, 42(13):192–193, March 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n13/pdf/scientificamerican03271880-192c.pdf.

Anonymous:1880:CRd

[113] Anonymous. Chemical repulsion. Scientific American, 42(13):193, March 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n13/pdf/scientificamerican03271880-193.pdf.

Anonymous:1880:NPMa

[114] Anonymous. New plastering machine. Scientific American, 42(13):194, March 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n13/pdf/scientificamerican03271880-194.pdf.

Anonymous:1880:ISPc

[115] Anonymous. Improved stone polishing machine. Scientific American, 42(13):194, March 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n13/pdf/scientificamerican03271880-
194a.pdf.

Anonymous:1880:MIe

[116] Anonymous. Miscellaneous inventions. Scientific American, 42(13):195, March 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n13/pdf/scientificamerican03271880-195.pdf.

Anonymous:1880:ISPd

[117] Anonymous. Improved sprout puller. Scientific American, 42(13):195, March 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n13/pdf/scientificamerican03271880-195a.pdf.

Anonymous:1880:ICW

[118] Anonymous. Improved car window. Scientific American, 42(13):195, March 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n13/pdf/scientificamerican03271880-195b.pdf.

Anonymous:1880:NHHa

[119] Anonymous. New holdback for harness. Scientific American, 42(13):195, March 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n13/pdf/scientificamerican03271880-
195c.pdf.

Anonymous:1880:AINd

[120] Anonymous. American industries — no. 36. Scientific American, 42(13):196–197, March 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n13/pdf/scientificamerican03271880-
196.pdf.

Anonymous:1880:NLB

[121] Anonymous. The new Leclanche battery. Scientific American, 42(13):198, March 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n13/pdf/scientificamerican03271880-
198.pdf.

Anonymous:1880:SAT

[122] Anonymous. Steel in agricultural tools. Scientific American, 42(13):198, March 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n13/pdf/scientificamerican03271880-198a.pdf.

Anonymous:1880:KTT

[123] Anonymous. The Kingio and the Tong-Tsing-Yo. Scientific American, 42(13):199–201, March 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n13/pdf/scientificamerican03271880-
199.pdf.

Anonymous:1880:AIb

[124] Anonymous. Agricultural inventions. Scientific American, 42(13):201, March 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n13/pdf/scientificamerican03271880-201.pdf.

Anonymous:1880:CRe

[125] Anonymous. Communications received. Scientific American, 42 (13):203, March 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n13/pdf/scientificamerican03271880-
203.pdf.

Anonymous:1880:SAEa

[126] Anonymous. The Scientific American export edition. Scientific American, 42(13):203–204, March 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n13/pdf/scientificamerican03271880-
203b.pdf.

Anonymous:1880:Cc

[127] Anonymous. Correspondence. Scientific American, 42(14):211–212, April 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n14/pdf/scientificamerican04031880-211a.pdf.

Anonymous:1880:BPg

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n14/pdf/scientificamerican04031880-219.pdf.

Anonymous:1880:IIf

[129] Anonymous. Index of inventions. Scientific American, 42(14):219, April 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n14/pdf/scientificamerican04031880-219b.pdf.

Anonymous:1880:PNPg

[130] Anonymous. Patent notes: Patents. Scientific American, 42(14):220, April 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n14/pdf/scientificamerican04031880-220.pdf.

Anonymous:1880:AISa

[131] Anonymous. America’s industrial supremacy assured. Scientific American, 42(14):208, April 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n14/pdf/scientificamerican04031880-
208.pdf.

Anonymous:1880:PYF

[132] Anonymous. Protection from yellow fever. Scientific American, 42(14):208–209, April 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n14/pdf/scientificamerican04031880-
208a.pdf.

Anonymous:1880:CPC

[133] Anonymous. Cheap patents cheapen goods. Scientific American, 42(14):209, April 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n14/pdf/scientificamerican04031880-
209.pdf.

Anonymous:1880:NYAa

[134] Anonymous. New York Academy of Sciences. Scientific American, 42(14):209, April 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n14/pdf/scientificamerican04031880-
209a.pdf.

Anonymous:1880:SBA

[135] Anonymous. Simple burglar alarm. Scientific American, 42(14):210, April 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n14/pdf/scientificamerican04031880-210.pdf.

Anonymous:1880:NPB

[136] Anonymous. The “New Pattern” Blake Crusher. Scientific American, 42(14):210, April 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n14/pdf/scientificamerican04031880-
210a.pdf.

Anonymous:1880:MIf

[137] Anonymous. Mechanical inventions. Scientific American, 42(14):210, April 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n14/pdf/scientificamerican04031880-210b.pdf.

Anonymous:1880:AINe

[138] Anonymous. American industries — no. 37. Scientific American, 42(14):211, April 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n14/pdf/scientificamerican04031880-
211.pdf.

Anonymous:1880:EId

[139] Anonymous. Engineering inventions. Scientific American, 42(14):213, April 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n14/pdf/scientificamerican04031880-213.pdf.

Anonymous:1880:SSM

[140] Anonymous. Scissors and scale measure. Scientific American, 42 (14):213, April 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n14/pdf/scientificamerican04031880-
213a.pdf.

Anonymous:1880:IJa

[141] Anonymous. Improvement in jackscrews. Scientific American, 42 (14):213, April 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n14/pdf/scientificamerican04031880-
213b.pdf.

Anonymous:1880:IPS

[142] Anonymous. Improved plant sprinkler. Scientific American, 42(14):213, April 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n14/pdf/scientificamerican04031880-213c.pdf.

Anonymous:1880:IPI

[143] Anonymous. Improved pruning implement. Scientific American, 42(14):213, April 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n14/pdf/scientificamerican04031880-
213d.pdf.

Anonymous:1880:NLS

[144] Anonymous. New lubricator for steam engines. Scientific American, 42(14):214, April 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n14/pdf/scientificamerican04031880-
214.pdf.

Anonymous:1880:MIg

[145] Anonymous. Miscellaneous inventions. Scientific American, 42 (14):214–215, April 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n14/pdf/scientificamerican04031880-
214a.pdf.

Anonymous:1880:VG

[146] Anonymous. Venus’ girdle. Scientific American, 42(14):215, April 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n14/pdf/scientificamerican04031880-215.pdf.

Anonymous:1880:PCb

[147] Anonymous. Peripatus Capensis. Scientific American, 42(14):215, April 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n14/pdf/scientificamerican04031880-215a.pdf.

Boyle:1880:SA

[148] Chas. B. Boyle. Specters in the air. Scientific American, 42(14):216–218, April 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n14/pdf/scientificamerican04031880-216.pdf.

Anonymous:1880:CRf

[149] Anonymous. Communications received. Scientific American, 42 (14):219, April 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n14/pdf/scientificamerican04031880-
219a.pdf.

Anonymous:1880:BPh

[150] Anonymous. Business and personal. Scientific American, 42(15):234, April 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n15/pdf/scientificamerican04101880-234.pdf.

Anonymous:1880:BNBd

[151] Anonymous. Books: New books and publications. Scientific American, 42(15):234, April 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n15/pdf/scientificamerican04101880-
234a.pdf.

Anonymous:1880:NQf

[152] Anonymous. Notes and queries. Scientific American, 42(15):234–235, April 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n15/pdf/scientificamerican04101880-234b.pdf.

Anonymous:1880:IIg

[153] Anonymous. Index of inventions. Scientific American, 42(15):235–236, April 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n15/pdf/scientificamerican04101880-235a.pdf.

Anonymous:1880:AEB

[154] Anonymous. American exhibits at the Berlin Fish Show. Scientific American, 42(15):224, April 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n15/pdf/scientificamerican04101880-
224.pdf.

Anonymous:1880:MMb

[155] Anonymous. Mining in Maine. Scientific American, 42(15):224, April 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n15/pdf/scientificamerican04101880-224a.pdf.

Anonymous:1880:ORB

[156] Anonymous. The old river beds of California. Scientific American, 42(15):225, April 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n15/pdf/scientificamerican04101880-
225.pdf.

Anonymous:1880:ASa

[157] Anonymous. Arizona shellac. Scientific American, 42(15):225, April 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n15/pdf/scientificamerican04101880-225a.pdf.

Anonymous:1880:ASW

[158] Anonymous. Another six weeks of suspense. Scientific American, 42(15):225–226, April 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n15/pdf/scientificamerican04101880-
225b.pdf.

Roberts:1880:NUO

[159] A. W. Roberts. New uses for old tin cans. Scientific American, 42(15):226–227, April 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n15/pdf/scientificamerican04101880-
226.pdf.

Anonymous:1880:MIh

[160] Anonymous. Miscellaneous inventions. Scientific American, 42(15):228, April 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n15/pdf/scientificamerican04101880-228.pdf.

Anonymous:1880:AINf

[161] Anonymous. American industries,no.38. Scientific American, 42 (15):229, April 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n15/pdf/scientificamerican04101880-
229.pdf.

Anonymous:1880:SEL

[162] Anonymous. The Sawyer electric light. Scientific American, 42 (15):230–231, April 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n15/pdf/scientificamerican04101880-
230.pdf.

Anonymous:1880:RCa

[163] Anonymous. The Racine canoe. Scientific American, 42(15):231, April 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n15/pdf/scientificamerican04101880-231.pdf.

Anonymous:1880:FDR

[164] Anonymous. The Franklin draught regulator. Scientific American, 42(15):231, April 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n15/pdf/scientificamerican04101880-
231a.pdf.

Anonymous:1880:NGE

[165] Anonymous. A new gas economizer. Scientific American, 42 (15):231–232, April 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n15/pdf/scientificamerican04101880-
231b.pdf.

Anonymous:1880:CCO

[166] Anonymous. Curious capture by an oyster and a mussel. Scientific American, 42(15):233, April 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n15/pdf/scientificamerican04101880-
233.pdf.

Anonymous:1880:CRg

[167] Anonymous. Communications received. Scientific American, 42 (15):235, April 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n15/pdf/scientificamerican04101880-
235.pdf.

Anonymous:1880:Pa

[168] Anonymous. Patents. Scientific American, 42(15):236, April 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n15/pdf/scientificamerican04101880-236.pdf.

Anonymous:1880:Cd

[169] Anonymous. Correspondence. Scientific American, 42(16):249, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n16/pdf/scientificamerican04171880-249.pdf.

Anonymous:1880:BPi

[170] Anonymous. Business and personal. Scientific American, 42(16):250, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n16/pdf/scientificamerican04171880-250a.pdf.

Anonymous:1880:BNBe

[171] Anonymous. Books: New books and publications. Scientific American, 42(16):250, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n16/pdf/scientificamerican04171880-
250b.pdf.

Anonymous:1880:NQg

[172] Anonymous. Notes and queries. Scientific American, 42(16):250, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n16/pdf/scientificamerican04171880-250c.pdf.

Anonymous:1880:IIh

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n16/pdf/scientificamerican04171880-251.pdf.

Anonymous:1880:NPCa

[174] Anonymous. The national plant of the Chinese. Scientific American, 42(16):240, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n16/pdf/scientificamerican04171880-
240.pdf.

Anonymous:1880:MMA

[175] Anonymous. The Metropolitan Museum of Art. Scientific American, 42(16):240, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n16/pdf/scientificamerican04171880-
240a.pdf.

Anonymous:1880:ADb

[176] Anonymous. Aids for the deaf. Scientific American, 42(16):240–241, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n16/pdf/scientificamerican04171880-240b.pdf.

Morton:1880:SEMa

[177] Henry Morton, Alfred M. Mayer, and B. F. Thomas. Some electrical measurements of one of Mr. Edison’s horseshoe lamps. Scientific American, 42(16):241, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n16/pdf/scientificamerican04171880-
241.pdf.

Anonymous:1880:AFC

[178] Anonymous. The American Fish Cultural Association. Scientific American, 42(16):241, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n16/pdf/scientificamerican04171880-
241a.pdf.

Anonymous:1880:PE

[179] Anonymous. Photo-engraving. Scientific American, 42(16):241, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n16/pdf/scientificamerican04171880-241b.pdf.

Anonymous:1880:NYE

[180] Anonymous. The New York Exhibition of 1883. Scientific American, 42(16):241–242, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n16/pdf/scientificamerican04171880-
241c.pdf.

Anonymous:1880:IHP

[181] Anonymous. Improved hand and power presses. Scientific American, 42(16):242, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n16/pdf/scientificamerican04171880-
242.pdf.

Anonymous:1880:NMBa

[182] Anonymous. New method of buoying vessels. Scientific American, 42(16):243, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n16/pdf/scientificamerican04171880-
243.pdf.

Anonymous:1880:NCR

[183] Anonymous. New check-row, corn planter, and fertilizer. Scientific American, 42(16):243, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n16/pdf/scientificamerican04171880-
243a.pdf.

Anonymous:1880:ISHa

[184] Anonymous. Improved steel horse collar. Scientific American, 42(16):243, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n16/pdf/scientificamerican04171880-
243b.pdf.

Anonymous:1880:MIi

[185] Anonymous. Miscellaneous inventions. Scientific American, 42(16):243, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n16/pdf/scientificamerican04171880-243c.pdf.

Anonymous:1880:AINg

[186] Anonymous. American industries, no. 39. Scientific American, 42(16):244, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n16/pdf/scientificamerican04171880-
244.pdf.

Anonymous:1880:ABE

[187] Anonymous. The Andre–Brougham electric lamp. Scientific American, 42(16):245, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n16/pdf/scientificamerican04171880-
245.pdf.

Anonymous:1880:NSF

[188] Anonymous. Novel sight for firearms. Scientific American, 42(16):245, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n16/pdf/scientificamerican04171880-245a.pdf.

Anonymous:1880:NSR

[189] Anonymous. Novel speed regulator. Scientific American, 42(16):245, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n16/pdf/scientificamerican04171880-245b.pdf.

Anonymous:1880:AIc

[190] Anonymous. Agricultural inventions. Scientific American, 42(16):246, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n16/pdf/scientificamerican04171880-246.pdf.

Anonymous:1880:IPT

[191] Anonymous. Improved pipe tongs. Scientific American, 42(16):246, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n16/pdf/scientificamerican04171880-246a.pdf.

Anonymous:1880:IMA

[192] Anonymous. Ice-making apparatus. Scientific American, 42(16):246, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n16/pdf/scientificamerican04171880-246b.pdf.

Anonymous:1880:EIe

[193] Anonymous. Engineering inventions. Scientific American, 42(16):246, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n16/pdf/scientificamerican04171880-246c.pdf.

Anonymous:1880:CLFa

[194] Anonymous. Capture of a large finback whale. Scientific American, 42(16):247–248, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n16/pdf/scientificamerican04171880-
247.pdf.

Anonymous:1880:RDRa

[195] Anonymous. Recent decisions relating to patents, etc. Scientific American, 42(16):248, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n16/pdf/scientificamerican04171880-
248.pdf.

Anonymous:1880:MIj

[196] Anonymous. Mechanical inventions. Scientific American, 42(16):248, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n16/pdf/scientificamerican04171880-248a.pdf.

Anonymous:1880:LRH

[197] Anonymous. The literary revolution. — history. Scientific American, 42(16):250, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n16/pdf/scientificamerican04171880-
250.pdf.

Anonymous:1880:CRh

[198] Anonymous. Communications received. Scientific American, 42 (16):250, April 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n16/pdf/scientificamerican04171880-
250d.pdf.

Anonymous:1880:IRIb

[199] Anonymous. Inventions: Recent inventions. Scientific American, 42(17):259, April 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n17/pdf/scientificamerican04241880-
259.pdf.

Anonymous:1880:BPj

[200] Anonymous. Business and personal. Scientific American, 42(17):267, April 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n17/pdf/scientificamerican04241880-267.pdf.

Anonymous:1880:NQh

[201] Anonymous. Notes and queries. Scientific American, 42(17):267, April 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n17/pdf/scientificamerican04241880-267a.pdf.

Anonymous:1880:IIi

[202] Anonymous. Index of inventions. Scientific American, 42(17):267–268, April 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n17/pdf/scientificamerican04241880-267b.pdf.

Anonymous:1880:PNPh

[203] Anonymous. Patent notes: Patents. Scientific American, 42(17):269, April 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n17/pdf/scientificamerican04241880-269.pdf.

Anonymous:1880:EH

[204] Anonymous. Electro-horticulture. Scientific American, 42(17):256, April 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n17/pdf/scientificamerican04241880-256.pdf.

Anonymous:1880:MGA

[205] Anonymous. The migration of the Great American Desert. Scientific American, 42(17):256–257, April 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v42/n17/pdf/scientificamerican04241880-
256a.pdf.

Anonymous:1880:MDPa

[206] Anonymous. The mining debris problem in California. Scientific American, 42(17):257, April 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n17/pdf/scientificamerican04241880-
257.pdf.

Anonymous:1880:AINh

[207] Anonymous. American industries, no. 40. Scientific American, 42 (17):258–259, April 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n17/pdf/scientificamerican04241880-
258.pdf.

Anonymous:1880:NCCb

[208] Anonymous. New car coupling. Scientific American, 42(17):260, April 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n17/pdf/scientificamerican04241880-260.pdf.

Anonymous:1880:AHM

[209] Anonymous. Automatic hub mortising machine. Scientific American, 42(17):260, April 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n17/pdf/scientificamerican04241880-
260a.pdf.

Anonymous:1880:MIk

[210] Anonymous. Miscellaneous inventions. Scientific American, 42(17):260, April 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n17/pdf/scientificamerican04241880-260b.pdf.

Anonymous:1880:NWI

[211] Anonymous. A new writing instrument. Scientific American, 42 (17):261, April 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n17/pdf/scientificamerican04241880-
261.pdf.

Anonymous:1880:NSTb

[212] Anonymous. New steak-tendering mallet. Scientific American, 42(17):261, April 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n17/pdf/scientificamerican04241880-
261a.pdf.

Anonymous:1880:MIl

[213] Anonymous. Mechanical inventions. Scientific American, 42(17):261, April 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n17/pdf/scientificamerican04241880-261b.pdf.

Anonymous:1880:NOM

[214] Anonymous. New oatmeal machine. Scientific American, 42(17):261, April 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n17/pdf/scientificamerican04241880-261c.pdf.

Anonymous:1880:BPT

[215] Anonymous. Boiler plate testing machine. Scientific American, 42(17):262, April 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n17/pdf/scientificamerican04241880-
262.pdf.

Anonymous:1880:ECM

[216] Anonymous. The elephant calf and its mother. Scientific American, 42(17):263–264, April 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n17/pdf/scientificamerican04241880-
263.pdf.

Anonymous:1880:IPD

[217] Anonymous. Improvement in potato diggers. Scientific American, 42(17):265–266, April 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n17/pdf/scientificamerican04241880-
265.pdf.

Anonymous:1880:Ce

[218] Anonymous. Correspondence. Scientific American, 42(18):281, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n18/pdf/scientificamerican05011880-281.pdf.

Anonymous:1880:BPk

[219] Anonymous. Business and personal. Scientific American, 42(18):282, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n18/pdf/scientificamerican05011880-282.pdf.

Anonymous:1880:NQi

[220] Anonymous. Notes and queries. Scientific American, 42(18):282–283, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n18/pdf/scientificamerican05011880-282a.pdf.

Anonymous:1880:BNBf

[221] Anonymous. Books: New books and publications. Scientific American, 42(18):283, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n18/pdf/scientificamerican05011880-
283.pdf.

Anonymous:1880:IIj

[222] Anonymous. Index of inventions. Scientific American, 42(18):283–284, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n18/pdf/scientificamerican05011880-283a.pdf.

Anonymous:1880:PNPi

[223] Anonymous. Patent notes: Patents. Scientific American, 42(18):285, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n18/pdf/scientificamerican05011880-285.pdf.

Anonymous:1880:ECR

[224] Anonymous. Elevated city railroads. Scientific American, 42(18):272, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n18/pdf/scientificamerican05011880-272.pdf.

Anonymous:1880:WG

[225] Anonymous. The wire gauge. Scientific American, 42(18):272, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n18/pdf/scientificamerican05011880-272a.pdf.

Anonymous:1880:FDT

[226] Anonymous. The fracture of the Duilio’s 100-ton gun. Scientific American, 42(18):272–273, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n18/pdf/scientificamerican05011880-
272b.pdf.

Anonymous:1880:ROC

[227] Anonymous. A remarkable oil company. Scientific American, 42 (18):273, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n18/pdf/scientificamerican05011880-
273.pdf.

Anonymous:1880:GGW

[228] Anonymous. The great gas well of Pennsylvania. Scientific American, 42(18):273, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n18/pdf/scientificamerican05011880-
273a.pdf.

Anonymous:1880:APL

[229] Anonymous. Amending the Patent Law. Scientific American, 42(18):273, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n18/pdf/scientificamerican05011880-
273b.pdf.

Morton:1880:SEMb

[230] Henry Morton, A. M. Mayer, and B. F. Thomas. Some electrical measurements of one of Mr. Edison’s horseshoe lamps. Scientific American, 42(18):273–274, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n18/pdf/scientificamerican05011880-
273c.pdf.

Anonymous:1880:NWTa

[231] Anonymous. A new warehouse truck. Scientific American, 42(18):274, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n18/pdf/scientificamerican05011880-274.pdf.

Anonymous:1880:NVa

[232] Anonymous. A new ventilator. Scientific American, 42(18):274, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n18/pdf/scientificamerican05011880-274a.pdf.

Anonymous:1880:IFP

[233] Anonymous. Improved felly plate. Scientific American, 42(18):274–275, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n18/pdf/scientificamerican05011880-274b.pdf.

Anonymous:1880:NTCa

[234] Anonymous. New ticket or canceling punch. Scientific American, 42(18):275, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n18/pdf/scientificamerican05011880-
275.pdf.

Anonymous:1880:NHC

[235] Anonymous. New hose carriage. Scientific American, 42(18):275–276, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n18/pdf/scientificamerican05011880-275a.pdf.

Anonymous:1880:NBM

[236] Anonymous. New box machine. Scientific American, 42(18):276, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n18/pdf/scientificamerican05011880-276.pdf.

Anonymous:1880:NEC

[237] Anonymous. New egg carrier and crate. Scientific American, 42(18):276, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n18/pdf/scientificamerican05011880-276a.pdf.

Anonymous:1880:MIm

[238] Anonymous. Mechanical inventions. Scientific American, 42(18):276, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n18/pdf/scientificamerican05011880-276b.pdf.

Anonymous:1880:AINi

[239] Anonymous. American industries — no. 41. Scientific American, 42(18):277–278, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n18/pdf/scientificamerican05011880-
277.pdf.

Anonymous:1880:NSPa

[240] Anonymous. A novel shading pen. Scientific American, 42(18):278, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n18/pdf/scientificamerican05011880-278.pdf.

Anonymous:1880:NWTb

[241] Anonymous. New wall tent and stove. Scientific American, 42(18):278, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n18/pdf/scientificamerican05011880-278a.pdf.

Anonymous:1880:MIn

[242] Anonymous. Miscellaneous inventions. Scientific American, 42(18):278, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n18/pdf/scientificamerican05011880-278b.pdf.

Roberts:1880:NBF

[243] A. W. Roberts. Nest-building fish, Gastebosteide. Scientific American, 42(18):279–280, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n18/pdf/scientificamerican05011880-
279.pdf.

Anonymous:1880:PWa

[244] Anonymous. The panda, or wah. Scientific American, 42(18):280, May 1, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n18/pdf/scientificamerican05011880-280.pdf.

Anonymous:1880:AId

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n18/pdf/scientificamerican05011880-281a.pdf.

Anonymous:1880:INIa

[246] Anonymous. Inventions: New inventions. Scientific American, 42(19):291–292, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n19/pdf/scientificamerican05081880-
291a.pdf.

Anonymous:1880:BPl

[247] Anonymous. Business and personal. Scientific American, 42(19):299, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n19/pdf/scientificamerican05081880-299.pdf.

Anonymous:1880:NQj

[248] Anonymous. Notes and queries. Scientific American, 42(19):299, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n19/pdf/scientificamerican05081880-299a.pdf.

Anonymous:1880:IIk

[249] Anonymous. Index of inventions. Scientific American, 42(19):299, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n19/pdf/scientificamerican05081880-299b.pdf.

Anonymous:1880:Pb

[250] Anonymous. Patents. Scientific American, 42(19):301, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n19/pdf/scientificamerican05081880-301.pdf.

Anonymous:1880:WNB

[251] Anonymous. White nickel bronze. Scientific American, 42(19):288, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n19/pdf/scientificamerican05081880-288.pdf.

Anonymous:1880:EAM

[252] Anonymous. The exportation of American manufactures. Scientific American, 42(19):288, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n19/pdf/scientificamerican05081880-
288a.pdf.

Anonymous:1880:NJM

[253] Anonymous. New Jersey mines and minerals. Scientific American, 42(19):288–289, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n19/pdf/scientificamerican05081880-
288b.pdf.

Anonymous:1880:CPa

[254] Anonymous. The Commissionership of Patents. Scientific American, 42(19):289, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n19/pdf/scientificamerican05081880-
289.pdf.

Anonymous:1880:ELTb

[255] Anonymous. The Edison lamp tests. Scientific American, 42(19):289, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n19/pdf/scientificamerican05081880-289a.pdf.

Anonymous:1880:NASa

[256] Anonymous. The National Academy of Science. Scientific American, 42(19):289, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n19/pdf/scientificamerican05081880-
289b.pdf.

Anonymous:1880:NYAb

[257] Anonymous. New York Academy of Sciences. Scientific American, 42(19):289, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n19/pdf/scientificamerican05081880-
289c.pdf.

Anonymous:1880:TMa

[258] Anonymous. Tornado in Missouri. Scientific American, 42(19):289, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n19/pdf/scientificamerican05081880-289d.pdf.

Anonymous:1880:TUH

[259] Anonymous. The tunnel under the Hudson River. Scientific American, 42(19):290–291, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n19/pdf/scientificamerican05081880-
290.pdf.

Anonymous:1880:CBa

[260] Anonymous. Crime in Bengal. Scientific American, 42(19):291, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n19/pdf/scientificamerican05081880-291.pdf.

Anonymous:1880:IKH

[261] Anonymous. Improved knife handle. Scientific American, 42(19):292, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n19/pdf/scientificamerican05081880-292.pdf.

Hopkins:1880:STT

[262] Geo. M. Hopkins. Simple telephone transmitter. Scientific American, 42(19):292–293, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n19/pdf/scientificamerican05081880-
292a.pdf.

Anonymous:1880:GSCa

[263] Anonymous. The Great Southern Comet. Scientific American, 42(19):293, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n19/pdf/scientificamerican05081880-
293.pdf.

Anonymous:1880:WTJ

[264] Anonymous. What is the time of Jupiter’s rotation. Scientific American, 42(19):293–294, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n19/pdf/scientificamerican05081880-
293a.pdf.

Anonymous:1880:NLP

[265] Anonymous. New life preserver. Scientific American, 42(19):294, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n19/pdf/scientificamerican05081880-294.pdf.

Anonymous:1880:NBCa

[266] Anonymous. New boiler cleaner. Scientific American, 42(19):294, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n19/pdf/scientificamerican05081880-294a.pdf.

Anonymous:1880:EIf

[267] Anonymous. Engineering invention. Scientific American, 42(19): 294–295, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n19/pdf/scientificamerican05081880-
294b.pdf.

Anonymous:1880:RCb

[268] Anonymous. Railway cremation. Scientific American, 42(19):295, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n19/pdf/scientificamerican05081880-295.pdf.

Roberts:1880:SB

[269] A. W. Roberts. Sea beans. Scientific American, 42(19):295, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n19/pdf/scientificamerican05081880-295a.pdf.

Anonymous:1880:IIl

[270] Anonymous. Insane ingenuity. Scientific American, 42(19):295, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n19/pdf/scientificamerican05081880-295b.pdf.

Anonymous:1880:MIo

[271] Anonymous. Miscellaneous inventions. Scientific American, 42(19):296, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n19/pdf/scientificamerican05081880-296.pdf.

Anonymous:1880:IWa

[272] Anonymous. Improved windmill. Scientific American, 42(19):296, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n19/pdf/scientificamerican05081880-296a.pdf.

Anonymous:1880:IR

[273] Anonymous. Improved rail. Scientific American, 42(19):296, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n19/pdf/scientificamerican05081880-296b.pdf.

Anonymous:1880:MIp

[274] Anonymous. Mechanical inventions. Scientific American, 42(19): 296–298, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n19/pdf/scientificamerican05081880-
296c.pdf.

Anonymous:1880:AIe

[275] Anonymous. Agricultural inventions. Scientific American, 42(19):298, May 8, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n19/pdf/scientificamerican05081880-298.pdf.

Anonymous:1880:Cf

[276] Anonymous. Correspondence. Scientific American, 42(2):20–21, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n2/pdf/scientificamerican01101880-20a.pdf.

Anonymous:1880:NIa

[277] Anonymous. New invention. Scientific American, 42(2):26, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n2/pdf/scientificamerican01101880-26.pdf.

Anonymous:1880:BPm

[278] Anonymous. Business and personal. Scientific American, 42(2):27, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n2/pdf/scientificamerican01101880-27.pdf.

Anonymous:1880:BNBg 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n2/pdf/scientificamerican01101880-
27a.pdf.

Anonymous:1880:NQk

[280] Anonymous. Notes and queries. Scientific American, 42(2):27, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n2/pdf/scientificamerican01101880-27b.pdf.

Anonymous:1880:IIm

[281] Anonymous. Index of inventions. Scientific American, 42(2):28, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n2/pdf/scientificamerican01101880-28.pdf.

Anonymous:1880:PNPj

[282] Anonymous. Patent notes: Patents. Scientific American, 42(2):28, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n2/pdf/scientificamerican01101880-28a.pdf.

Anonymous:1880:TCO

[283] Anonymous. The telephonic central office system. Scientific American, 42(2):15, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n2/pdf/scientificamerican01101880-
15.pdf.

Anonymous:1880:FT

[284] Anonymous. The future of the telephone. Scientific American, 42(2):16, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n2/pdf/scientificamerican01101880-
16.pdf.

Anonymous:1880:PAI

[285] Anonymous. Progress of artificial illumination. Scientific American, 42(2):16–17, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n2/pdf/scientificamerican01101880-
16a.pdf.

Anonymous:1880:ASb

[286] Anonymous. Alexander Stuart. Scientific American, 42(2):17, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n2/pdf/scientificamerican01101880-17.pdf.

Anonymous:1880:EEM

[287] Anonymous. Exclusiveness of English manufacturers. Scientific American, 42(2):17, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n2/pdf/scientificamerican01101880-
17a.pdf.

Anonymous:1880:HM

[288] Anonymous. The heart as a machine. Scientific American, 42 (2):17, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n2/pdf/scientificamerican01101880-
17b.pdf.

Anonymous:1880:SSI

[289] Anonymous. The struggles of a successful inventor. Scientific American, 42(2):17, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n2/pdf/scientificamerican01101880-
17c.pdf.

Anonymous:1880:LWM

[290] Anonymous. Low water on manufacturing streams. Scientific American, 42(2):17, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n2/pdf/scientificamerican01101880-
17d.pdf.

Anonymous:1880:JHP

[291] Anonymous. The Johnston Harvester Prize. Scientific American, 42(2):18–19, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n2/pdf/scientificamerican01101880-
18.pdf.

Anonymous:1880:ELE

[292] Anonymous. Edison’s latest electric light. Scientific American, 42(2):19, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n2/pdf/scientificamerican01101880-
19.pdf.

Anonymous:1880:MIq

[293] Anonymous. Miscellaneous inventions. Scientific American, 42 (2):19–20, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n2/pdf/scientificamerican01101880-
19a.pdf.

Anonymous:1880:NFa

[294] Anonymous. A new filter. Scientific American, 42(2):20, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n2/pdf/scientificamerican01101880-20.pdf.

Keil:1880:DI

[295] Maurice Keil. On the dephosphorization of iron. Scientific American, 42(2):21, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n2/pdf/scientificamerican01101880-
21.pdf.

Anonymous:1880:MIr

[296] Anonymous. Mechanical invention. Scientific American, 42(2):21, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n2/pdf/scientificamerican01101880-21a.pdf.

Anonymous:1880:MF

[297] Anonymous. Miasm and fevers. Scientific American, 42(2):21–22, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n2/pdf/scientificamerican01101880-21b.pdf.

Anonymous:1880:AER

[298] Anonymous. Another extinct race that never existed. Scientific American, 42(2):22, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n2/pdf/scientificamerican01101880-
22.pdf.

Anonymous:1880:NSDa

[299] Anonymous. Novel swimming device. Scientific American, 42 (2):22, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n2/pdf/scientificamerican01101880-
22a.pdf.

Anonymous:1880:JTG

[300] Anonymous. Juvet’s time globe. Scientific American, 42(2):22, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n2/pdf/scientificamerican01101880-22b.pdf.

Anonymous:1880:AIf

[301] Anonymous. Agricultural inventions. Scientific American, 42 (2):23, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n2/pdf/scientificamerican01101880-
23.pdf.

Hovey:1880:BF

[302] H. C. Hovey. The basket fish. Scientific American, 42(2):23–24, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n2/pdf/scientificamerican01101880-23a.pdf.

Anonymous:1880:NHNb

[303] Anonymous. Natural history notes. Scientific American, 42(2): 24–26, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n2/pdf/scientificamerican01101880-
24.pdf.

Anonymous:1880:CRj

[304] Anonymous. Communications received. Scientific American, 42 (2):27, January 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n2/pdf/scientificamerican01101880-
27c.pdf.

Anonymous:1880:INIb

[305] Anonymous. Inventions: New inventions. Scientific American, 42 (20):307, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n20/pdf/scientificamerican05151880-
307a.pdf.

Anonymous:1880:Cg

[306] Anonymous. Correspondence. Scientific American, 42(20):308, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n20/pdf/scientificamerican05151880-308a.pdf.

Anonymous:1880:BPn

[307] Anonymous. Business and personal. Scientific American, 42(20):314, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n20/pdf/scientificamerican05151880-314.pdf.

Anonymous:1880:BNBh

[308] Anonymous. Books: New books and publications. Scientific American, 42(20):314, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n20/pdf/scientificamerican05151880-
314a.pdf.

Anonymous:1880:NQl

[309] Anonymous. Notes and queries. Scientific American, 42(20):314–315, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n20/pdf/scientificamerican05151880-314b.pdf.

Anonymous:1880:IIn

[310] Anonymous. Index of inventions. Scientific American, 42(20):315, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n20/pdf/scientificamerican05151880-315a.pdf.

Anonymous:1880:Pd

[311] Anonymous. Patents. Scientific American, 42(20):317, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n20/pdf/scientificamerican05151880-317.pdf.

Anonymous:1880:AINj

[312] Anonymous. American industries — no. 42. Scientific American, 42(20):303, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n20/pdf/scientificamerican05151880-
303.pdf.

Anonymous:1880:USS

[313] Anonymous. The use of steel for structural purposes. Scientific American, 42(20):304, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n20/pdf/scientificamerican05151880-
304.pdf.

Anonymous:1880:TTS

[314] Anonymous. Two tons of silver per week. Scientific American, 42(20):304, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n20/pdf/scientificamerican05151880-
304a.pdf.

Anonymous:1880:TSS

[315] Anonymous. The telegraph set screw. Scientific American, 42(20):304, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n20/pdf/scientificamerican05151880-304b.pdf.

Anonymous:1880:SI

[316] Anonymous. Steam injectors. Scientific American, 42(20):304, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n20/pdf/scientificamerican05151880-304c.pdf.

Anonymous:1880:RTM

[317] Anonymous. The registration of trade marks. Scientific American, 42(20):304–305, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n20/pdf/scientificamerican05151880-
304d.pdf.

Anonymous:1880:RTE

[318] Anonymous. Recent telephone experiments. Scientific American, 42(20):305, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n20/pdf/scientificamerican05151880-
305.pdf.

Anonymous:1880:CPb

[319] Anonymous. The coffee product. Scientific American, 42(20):305, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n20/pdf/scientificamerican05151880-305a.pdf.

Anonymous:1880:ITCb

[320] Anonymous. Improved telephone central office switch board. Scientific American, 42(20):305, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n20/pdf/scientificamerican05151880-
305b.pdf.

Anonymous:1880:IGP

[321] Anonymous. Improved gang plow. Scientific American, 42(20):306, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n20/pdf/scientificamerican05151880-306.pdf.

Anonymous:1880:NDC

[322] Anonymous. New dump car. Scientific American, 42(20):306–307, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n20/pdf/scientificamerican05151880-306a.pdf.

Anonymous:1880:GST

[323] Anonymous. The Gower system of telephonic communication, Paris. Scientific American, 42(20):307, May 15, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v42/
n20/pdf/scientificamerican05151880-307.pdf.

Anonymous:1880:AIg

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n20/pdf/scientificamerican05151880-308.pdf.

Anonymous:1880:MIs

[325] Anonymous. Miscellaneous inventions. Scientific American, 42(20):309, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n20/pdf/scientificamerican05151880-309.pdf.

Anonymous:1880:NTH

[326] Anonymous. Novel tool handle. Scientific American, 42(20):310, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n20/pdf/scientificamerican05151880-310.pdf.

Anonymous:1880:NRE

[327] Anonymous. New rotary engine. Scientific American, 42(20):310, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n20/pdf/scientificamerican05151880-310a.pdf.

Anonymous:1880:MIt

[328] Anonymous. Mechanical inventions. Scientific American, 42(20):310, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n20/pdf/scientificamerican05151880-310b.pdf.

Anonymous:1880:IGW

[329] Anonymous. Improved gun wad. Scientific American, 42(20):310, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n20/pdf/scientificamerican05151880-310c.pdf.

Beard:1880:CT

[330] Daniel C. Beard. A copter on trout. Scientific American, 42 (20):311–312, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n20/pdf/scientificamerican05151880-
311.pdf.

Anonymous:1880:EIg

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n20/pdf/scientificamerican05151880-312.pdf.

Anonymous:1880:FW

[332] Anonymous. Flowers and the weather. Scientific American, 42(20):313, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n20/pdf/scientificamerican05151880-313.pdf.

Anonymous:1880:CRk

[333] Anonymous. Communications received. Scientific American, 42 (20):315, May 15, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n20/pdf/scientificamerican05151880-
315.pdf.

Anonymous:1880:INIc

[334] Anonymous. Inventions: New inventions. Scientific American, 42 (21):322, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n21/pdf/scientificamerican05221880-
322a.pdf.

Anonymous:1880:BPo

[335] Anonymous. Business and personal. Scientific American, 42(21):331, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n21/pdf/scientificamerican05221880-331.pdf.

Anonymous:1880:BNBi

[336] Anonymous. Books: New books and publications. Scientific American, 42(21):331, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n21/pdf/scientificamerican05221880-
331a.pdf.

Anonymous:1880:NQm

[337] Anonymous. Notes and queries. Scientific American, 42(21):331–332, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n21/pdf/scientificamerican05221880-331b.pdf.

Anonymous:1880:IIo

[338] Anonymous. Index of inventions. Scientific American, 42(21):332, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n21/pdf/scientificamerican05221880-332a.pdf.

Anonymous:1880:PNPk

[339] Anonymous. Patent notes: Patents. Scientific American, 42(21):333, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n21/pdf/scientificamerican05221880-333.pdf.

Anonymous:1880:AREa

[340] Anonymous. Artificial respiration, the efficiency of the water trap. Scientific American, 42(21):319, May 22, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v42/
n21/pdf/scientificamerican05221880-319.pdf.

Anonymous:1880:BWW

[341] Anonymous. The Babcock & Wilcox water tube boiler. Scientific American, 42(21):319, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n21/pdf/scientificamerican05221880-
319a.pdf.

Anonymous:1880:PLE

[342] Anonymous. Patent Law and equity. Scientific American, 42(21):320, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n21/pdf/scientificamerican05221880-320.pdf.

Anonymous:1880:TMb

[343] Anonymous. Two meteorites. Scientific American, 42(21):320, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n21/pdf/scientificamerican05221880-320a.pdf.

Anonymous:1880:HFa

[344] Anonymous. Household fungi. Scientific American, 42(21):320–321, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n21/pdf/scientificamerican05221880-320b.pdf.

Anonymous:1880:NIS

[345] Anonymous. A new industrial school of art in New York. Scientific American, 42(21):321, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n21/pdf/scientificamerican05221880-
321.pdf.

Anonymous:1880:NASb

[346] Anonymous. The National Academy of Science. Scientific American, 42(21):321, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n21/pdf/scientificamerican05221880-
321a.pdf.

Anonymous:1880:NSDb

[347] Anonymous. Novel swimming device. Scientific American, 42(21):322, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n21/pdf/scientificamerican05221880-322.pdf.

Anonymous:1880:LIM

[348] Anonymous. A large ice machine. Scientific American, 42(21):322, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n21/pdf/scientificamerican05221880-322b.pdf.

Anonymous:1880:NOT

[349] Anonymous. New oil-testing apparatus. Scientific American, 42 (21):323, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n21/pdf/scientificamerican05221880-
323.pdf.

Anonymous:1880:IHBa

[350] Anonymous. Improved hot-blast boiler furnace. Scientific American, 42(21):323–324, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n21/pdf/scientificamerican05221880-
323a.pdf.

Anonymous:1880:NTS

[351] Anonymous. Novel transparent sign. Scientific American, 42(21):324, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n21/pdf/scientificamerican05221880-324.pdf.

Anonymous:1880:NTa

[352] Anonymous. A novel thermoscope. Scientific American, 42(21):324, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n21/pdf/scientificamerican05221880-324a.pdf.

Anonymous:1880:NCG

[353] Anonymous. New corn grater. Scientific American, 42(21):324, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n21/pdf/scientificamerican05221880-324b.pdf.

Anonymous:1880:Ch

[354] Anonymous. The Columbia. Scientific American, 42(21):326, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n21/pdf/scientificamerican05221880-326.pdf.

Anonymous:1880:NHNc

[355] Anonymous. Natural history notes. Scientific American, 42(21):327, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n21/pdf/scientificamerican05221880-327.pdf.

Noble:1880:DF

[356] Burr Noble. Dust figures. Scientific American, 42(21):327, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n21/pdf/scientificamerican05221880-327a.pdf.

Anonymous:1880:ATM

[357] Anonymous. Ants and their milch cattle. Scientific American, 42(21):327–328, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n21/pdf/scientificamerican05221880-
327b.pdf.

Anonymous:1880:CTb

[358] Anonymous. The chronology of the telegraph. Scientific American, 42(21):329, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n21/pdf/scientificamerican05221880-
329.pdf.

Anonymous:1880:MIu

[359] Anonymous. Miscellaneous inventions. Scientific American, 42(21):330, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n21/pdf/scientificamerican05221880-330.pdf.

Anonymous:1880:RDRb

[360] Anonymous. Recent decisions relating to patents. Scientific American, 42(21):330, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n21/pdf/scientificamerican05221880-
330a.pdf.

Dudley:1880:VOM

[361] Wm. L. Dudley. The volatile oil of mustard. Scientific American, 42(21):330, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n21/pdf/scientificamerican05221880-
330b.pdf.

Anonymous:1880:CRl

[362] Anonymous. Communications received. Scientific American, 42 (21):332, May 22, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n21/pdf/scientificamerican05221880-
332.pdf.

Anonymous:1880:INId

[363] Anonymous. Inventions: New inventions. Scientific American, 42 (22):346, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n22/pdf/scientificamerican05291880-
346.pdf.

Anonymous:1880:BPp

[364] Anonymous. Business and personal. Scientific American, 42(22):347, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n22/pdf/scientificamerican05291880-347.pdf.

Anonymous:1880:BNBj

[365] Anonymous. Books: New books and publications. Scientific American, 42(22):347, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n22/pdf/scientificamerican05291880-
347a.pdf.

Anonymous:1880:NQn

[366] Anonymous. Notes and queries. Scientific American, 42(22):347, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n22/pdf/scientificamerican05291880-347b.pdf.

Anonymous:1880:IIp

[367] Anonymous. Index of inventions. Scientific American, 42(22):347–348, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n22/pdf/scientificamerican05291880-347d.pdf.

Anonymous:1880:Pf

[368] Anonymous. Patents. Scientific American, 42(22):349, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n22/pdf/scientificamerican05291880-349.pdf.

Anonymous:1880:PI

[369] Anonymous. Patents for not inventing. Scientific American, 42(22):336, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n22/pdf/scientificamerican05291880-336.pdf.

Anonymous:1880:PPa

[370] Anonymous. The Page patents. Scientific American, 42(22):336, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n22/pdf/scientificamerican05291880-336a.pdf.

Anonymous:1880:HI

[371] Anonymous. Honors to an inventor. Scientific American, 42(22):336, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n22/pdf/scientificamerican05291880-336b.pdf.

Anonymous:1880:ERA

[372] Anonymous. The electric railway an American invention. Scientific American, 42(22):336–337, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v42/n22/pdf/scientificamerican05291880-
336c.pdf.

Anonymous:1880:PTT

[373] Anonymous. Prof. Tice’s theory of cyclones. Scientific American, 42(22):337, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n22/pdf/scientificamerican05291880-
337.pdf.

Catlin:1880:NAS

[374] George L. Catlin. New Atlantic seaport in France. Scientific American, 42(22):337, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n22/pdf/scientificamerican05291880-
337a.pdf.

Anonymous:1880:LCB

[375] Anonymous. Legislating on color blindness. Scientific American, 42(22):337, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n22/pdf/scientificamerican05291880-
337b.pdf.

Anonymous:1880:NDE

[376] Anonymous. New draughtsman’s easel. Scientific American, 42(22):338, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n22/pdf/scientificamerican05291880-338.pdf.

Anonymous:1880:NMO

[377] Anonymous. New method of operating mining pumps. Scientific American, 42(22):338, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n22/pdf/scientificamerican05291880-
338a.pdf.

Anonymous:1880:MIv

[378] Anonymous. Mechanical inventions. Scientific American, 42(22): 338–339, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n22/pdf/scientificamerican05291880-
338b.pdf.

Anonymous:1880:AIh

[379] Anonymous. Agricultural inventions. Scientific American, 42(22):339, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n22/pdf/scientificamerican05291880-339.pdf.

Anonymous:1880:NSL

[380] Anonymous. New self-locking hook. Scientific American, 42(22):339, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n22/pdf/scientificamerican05291880-339a.pdf.

Anonymous:1880:NRC

[381] Anonymous. New refrigerating cover. Scientific American, 42(22):339, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n22/pdf/scientificamerican05291880-339b.pdf.

Anonymous:1880:AINk

[382] Anonymous. American industries no. 43. Scientific American, 42 (22):340–341, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n22/pdf/scientificamerican05291880-
340.pdf.

Anonymous:1880:ARS

[383] Anonymous. Apparatus for registering solar radiation. Scientific American, 42(22):341–342, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n22/pdf/scientificamerican05291880-
341.pdf.

Anonymous:1880:NSGa

[384] Anonymous. New sporting gun. Scientific American, 42(22):342, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n22/pdf/scientificamerican05291880-342.pdf.

Anonymous:1880:AREb

[385] Anonymous. Apparatus for recording earthquake movements. Scientific American, 42(22):342, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n22/pdf/scientificamerican05291880-
342a.pdf.

Anonymous:1880:MIw

[386] Anonymous. Miscellaneous inventions. Scientific American, 42(22):342, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n22/pdf/scientificamerican05291880-342b.pdf.

Roberts:1880:DSC

[387] A. W. Roberts. The decorating spider crab. Scientific American, 42(22):343, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n22/pdf/scientificamerican05291880-
343.pdf.

Seiss:1880:PT

[388] C. Few Seiss. The painted turtle. Scientific American, 42(22):343–344, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n22/pdf/scientificamerican05291880-343a.pdf.

Anonymous:1880:NHNd

[389] Anonymous. Natural history notes. Scientific American, 42(22):344, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n22/pdf/scientificamerican05291880-344.pdf.

Anonymous:1880:RDRc

[390] Anonymous. Recent decisions relating to patents, copyrights, etc. Scientific American, 42(22):344–345, May 29, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v42/
n22/pdf/scientificamerican05291880-344a.pdf.

Anonymous:1880:CRm

[391] Anonymous. Communications received. Scientific American, 42 (22):347, May 29, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n22/pdf/scientificamerican05291880-
347c.pdf.

Anonymous:1880:Ci

[392] Anonymous. Correspondence. Scientific American, 42(23):357–358, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n23/pdf/scientificamerican06051880-357a.pdf.

Anonymous:1880:BPq

[393] Anonymous. Business and personal. Scientific American, 42(23):363, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n23/pdf/scientificamerican06051880-363.pdf.

Anonymous:1880:BNBk

[394] Anonymous. Books: New books and publications. Scientific American, 42(23):363, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n23/pdf/scientificamerican06051880-
363a.pdf.

Anonymous:1880:NQo

[395] Anonymous. Notes and queries. Scientific American, 42(23):363, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n23/pdf/scientificamerican06051880-363b.pdf.

Anonymous:1880:IIq

[396] Anonymous. Index of inventions. Scientific American, 42(23):363–364, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n23/pdf/scientificamerican06051880-363c.pdf.

Anonymous:1880:EGD

[397] Anonymous. An epidemic of gold discovery. Scientific American, 42(23):352, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n23/pdf/scientificamerican06051880-
352.pdf.

Anonymous:1880:ALF

[398] Anonymous. Another lost field glass. Scientific American, 42(23):352, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n23/pdf/scientificamerican06051880-352a.pdf.

Anonymous:1880:FRS

[399] Anonymous. A fast river steamer. Scientific American, 42(23):352, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n23/pdf/scientificamerican06051880-352b.pdf.

Anonymous:1880:ICa

[400] Anonymous. Improvements at Cohoes. Scientific American, 42 (23):352–353, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n23/pdf/scientificamerican06051880-
352c.pdf.

Anonymous:1880:PFP

[401] Anonymous. Poisonous frying pans. Scientific American, 42(23):353, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n23/pdf/scientificamerican06051880-353.pdf.

Anonymous:1880:CCa

[402] Anonymous. Carp culture. Scientific American, 42(23):353, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n23/pdf/scientificamerican06051880-353a.pdf.

Anonymous:1880:SCa

[403] Anonymous. Stove castings. Scientific American, 42(23):353, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n23/pdf/scientificamerican06051880-353b.pdf.

Anonymous:1880:MZ

[404] Anonymous. Muybridge’s zoogyroscope. Scientific American, 42 (23):353, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n23/pdf/scientificamerican06051880-
353c.pdf.

Anonymous:1880:NYAc

[405] Anonymous. New York Academy of Sciences. Scientific American, 42(23):353, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n23/pdf/scientificamerican06051880-
353d.pdf.

Anonymous:1880:MBP

[406] Anonymous. Milk and butter preservatives. Scientific American, 42(23):353–354, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n23/pdf/scientificamerican06051880-
353e.pdf.

Anonymous:1880:IFM

[407] Anonymous. Improved folding mirror. Scientific American, 42(23):354, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n23/pdf/scientificamerican06051880-354.pdf.

Anonymous:1880:ENE

[408] Anonymous. Edison’s new electrical railway. Scientific American, 42(23):354–355, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n23/pdf/scientificamerican06051880-
354a.pdf.

Anonymous:1880:SEa

[409] Anonymous. Seeing by electricity. Scientific American, 42(23):355, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n23/pdf/scientificamerican06051880-355.pdf.

Anonymous:1880:NAM

[410] Anonymous. Novel animal motor. Scientific American, 42(23):355, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n23/pdf/scientificamerican06051880-355a.pdf.

Anonymous:1880:NCa

[411] Anonymous. A novel combination. Scientific American, 42(23):355–356, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n23/pdf/scientificamerican06051880-355b.pdf.

Anonymous:1880:MIx

[412] Anonymous. Miscellaneous inventions. Scientific American, 42 (23):356–357, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n23/pdf/scientificamerican06051880-
356.pdf.

Anonymous:1880:AINl

[413] Anonymous. American industries — no 44. Scientific American, 42(23):357, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n23/pdf/scientificamerican06051880-
357.pdf.

Anonymous:1880:MIy

[414] Anonymous. Mechanical inventions. Scientific American, 42(23):358, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n23/pdf/scientificamerican06051880-358.pdf.

Anonymous:1880:NSB

[415] Anonymous. New steam boiler. Scientific American, 42(23):358, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n23/pdf/scientificamerican06051880-358a.pdf.

Anonymous:1880:IHa

[416] Anonymous. Improvement in hatchways. Scientific American, 42 (23):358–359, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n23/pdf/scientificamerican06051880-
358b.pdf.

Anonymous:1880:SAA

[417] Anonymous. South African antelopes. Scientific American, 42 (23):359–360, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n23/pdf/scientificamerican06051880-
359.pdf.

Anonymous:1880:EIh

[418] Anonymous. Engineering inventions. Scientific American, 42(23): 360–363, June 5, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n23/pdf/scientificamerican06051880-
360.pdf.

Anonymous:1880:Cj

[419] Anonymous. Correspondence. Scientific American, 42(24):373, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n24/pdf/scientificamerican06121880-373.pdf.

Anonymous:1880:INIe

[420] Anonymous. Inventions: New inventions. Scientific American, 42(24):376, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n24/pdf/scientificamerican06121880-
376.pdf.

Anonymous:1880:BPr

[421] Anonymous. Business and personal. Scientific American, 42(24):379, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n24/pdf/scientificamerican06121880-379.pdf.

Anonymous:1880:BNBl

[422] Anonymous. Books: New books and publications. Scientific American, 42(24):379, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n24/pdf/scientificamerican06121880-
379a.pdf.

Anonymous:1880:NQp

[423] Anonymous. Notes and queries. Scientific American, 42(24):379, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n24/pdf/scientificamerican06121880-379b.pdf.

Anonymous:1880:IIr

[424] Anonymous. Index of inventions. Scientific American, 42(24):380, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n24/pdf/scientificamerican06121880-380.pdf.

Anonymous:1880:FER

[425] Anonymous. The future of the electric railway. Scientific American, 42(24):368, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n24/pdf/scientificamerican06121880-
368.pdf.

Anonymous:1880:DMN

[426] Anonymous. Deep mines in Nevada. Scientific American, 42(24):368, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n24/pdf/scientificamerican06121880-368a.pdf.

Anonymous:1880:FAa

[427] Anonymous. Food adulteration. Scientific American, 42(24):368, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n24/pdf/scientificamerican06121880-368b.pdf.

Anonymous:1880:ACSa

[428] Anonymous. The American Chemical Society. Scientific American, 42(24):368–369, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n24/pdf/scientificamerican06121880-
368c.pdf.

Anonymous:1880:OQ

[429] Anonymous. Oysters at Quinnipiac. Scientific American, 42(24):369, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n24/pdf/scientificamerican06121880-369.pdf.

Boyle:1880:BVT

[430] Charles B. Boyle. Binocular vision in telescopes. Scientific American, 42(24):370–371, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n24/pdf/scientificamerican06121880-
370.pdf.

Anonymous:1880:NEL

[431] Anonymous. New electric lamp. Scientific American, 42(24):371–372, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n24/pdf/scientificamerican06121880-371.pdf.

Anonymous:1880:NSMb

[432] Anonymous. New style of moulding. Scientific American, 42(24):372, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n24/pdf/scientificamerican06121880-372.pdf.

Anonymous:1880:SEI

[433] Anonymous. The steam engine indicator. Scientific American, 42(24):372, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n24/pdf/scientificamerican06121880-
372a.pdf.

Anonymous:1880:NMd

[434] Anonymous. A new motor. Scientific American, 42(24):373, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n24/pdf/scientificamerican06121880-373a.pdf.

Anonymous:1880:MIz

[435] Anonymous. Miscellaneous inventions. Scientific American, 42(24):373, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n24/pdf/scientificamerican06121880-373b.pdf.

Anonymous:1880:USB

[436] Anonymous. A useful steam boiler appliance. Scientific American, 42(24):374, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n24/pdf/scientificamerican06121880-
374.pdf.

Anonymous:1880:NCW

[437] Anonymous. New ceiling and wall. Scientific American, 42(24):374, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n24/pdf/scientificamerican06121880-374a.pdf.

Anonymous:1880:EIi

[438] Anonymous. Engineering inventions. Scientific American, 42(24):374, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n24/pdf/scientificamerican06121880-374b.pdf.

Anonymous:1880:IMP

[439] Anonymous. Improved mill pick. Scientific American, 42(24):374–375, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n24/pdf/scientificamerican06121880-374c.pdf.

Hovey:1880:CC

[440] H. C. Hovey. The crinoids of Crawfordsville. Scientific American, 42(24):375–376, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n24/pdf/scientificamerican06121880-
375.pdf.

Anonymous:1880:AINm

[441] Anonymous. American industries — no 45. Scientific American, 42(24):376, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n24/pdf/scientificamerican06121880-
376a.pdf.

Anonymous:1880:PCP

[442] Anonymous. Porcelain chimney-piece. Scientific American, 42(24):377, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n24/pdf/scientificamerican06121880-377.pdf.

Anonymous:1880:MIba

[443] Anonymous. Mechanical inventions. Scientific American, 42(24):378, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n24/pdf/scientificamerican06121880-378.pdf.

Anonymous:1880:CRn

[444] Anonymous. Communications received. Scientific American, 42 (24):379, June 12, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n24/pdf/scientificamerican06121880-
379c.pdf.

Anonymous:1880:INIf

[445] Anonymous. Inventions: New inventions. Scientific American, 42(25):386, June 19, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n25/pdf/scientificamerican06191880-
386b.pdf.

Anonymous:1880:BPs

[446] Anonymous. Business and personal. Scientific American, 42(25):395, June 19, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n25/pdf/scientificamerican06191880-395.pdf.

Anonymous:1880:BNBm

[447] Anonymous. Books: New books and publications. Scientific American, 42(25):395, June 19, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n25/pdf/scientificamerican06191880-
395a.pdf.

Anonymous:1880:NQq

[448] Anonymous. Notes and queries. Scientific American, 42(25):395–396, June 19, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n25/pdf/scientificamerican06191880-395b.pdf.

Anonymous:1880:IIs

[449] Anonymous. Index of inventions. Scientific American, 42(25):396, June 19, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n25/pdf/scientificamerican06191880-396.pdf.

Anonymous:1880:AINn

[450] Anonymous. American industries — no 46. Scientific American, 42(25):383–384, June 19, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n25/pdf/scientificamerican06191880-
383.pdf.

Anonymous:1880:GYB

[451] Anonymous. Great yield of a blast furnace. Scientific American, 42(25):384, June 19, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n25/pdf/scientificamerican06191880-
384.pdf.

Anonymous:1880:PIS

[452] Anonymous. The problem of the ice supply for New York City. Scientific American, 42(25):384–385, June 19, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v42/
n25/pdf/scientificamerican06191880-384a.pdf.

Anonymous:1880:MC

[453] Anonymous. The Millers in council. Scientific American, 42(25):385, June 19, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n25/pdf/scientificamerican06191880-385.pdf.

Anonymous:1880:HPE

[454] Anonymous. The Howgate Polar Expedition. Scientific American, 42(25):385, June 19, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n25/pdf/scientificamerican06191880-
385a.pdf.

Anonymous:1880:OIS

[455] Anonymous. Our iron and steel imports. Scientific American, 42 (25):385–386, June 19, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n25/pdf/scientificamerican06191880-
385b.pdf.

Anonymous:1880:AFW

[456] Anonymous. Automatic feed water apparatus for steam boilers. Scientific American, 42(25):386, June 19, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v42/
n25/pdf/scientificamerican06191880-386.pdf.

Anonymous:1880:CAa

[457] Anonymous. Climbing apparatus. Scientific American, 42(25):386, June 19, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n25/pdf/scientificamerican06191880-386a.pdf.

Anonymous:1880:KB

[458] Anonymous. The Kuilenburg bridge. Scientific American, 42(25):387, June 19, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n25/pdf/scientificamerican06191880-387.pdf.

Anonymous:1880:SAR

[459] Anonymous. Safety appliance for releasing horses. Scientific American, 42(25):387, June 19, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n25/pdf/scientificamerican06191880-
387a.pdf.

Anonymous:1880:SBC

[460] Anonymous. Sled and boat combined. Scientific American, 42(25):387, June 19, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n25/pdf/scientificamerican06191880-387b.pdf.

Anonymous:1880:CIa

[461] Anonymous. Calendar inkstand. Scientific American, 42(25):387–388, June 19, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n25/pdf/scientificamerican06191880-387c.pdf.

Anonymous:1880:EOS

[462] Anonymous. Edison’s ore separator. Scientific American, 42(25):388, June 19, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n25/pdf/scientificamerican06191880-388.pdf.

Anonymous:1880:ISa

[463] Anonymous. Improvement in stovepipe. Scientific American, 42 (25):388, June 19, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n25/pdf/scientificamerican06191880-
388a.pdf.

Baxter:1880:SES

[464] M. L. Baxter. Some experiments with small batteries. Scientific American, 42(25):389, June 19, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n25/pdf/scientificamerican06191880-
389.pdf.

Anonymous:1880:AMa

[465] Anonymous. Amateur mechanics. Scientific American, 42(25):390, June 19, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n25/pdf/scientificamerican06191880-390.pdf.

Anonymous:1880:MIbb

[466] Anonymous. Miscellaneous inventions. Scientific American, 42(25):390, June 19, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n25/pdf/scientificamerican06191880-390a.pdf.

Anonymous:1880:BM

[467] Anonymous. The Babyboussa of Malacca. Scientific American, 42(25):391, June 19, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n25/pdf/scientificamerican06191880-
391.pdf.

Anonymous:1880:Ba

[468] Anonymous. The Barbastelle. Scientific American, 42(25):391–392, June 19, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n25/pdf/scientificamerican06191880-391a.pdf.

Anonymous:1880:EIj

[469] Anonymous. Engineering inventions. Scientific American, 42(25):392, June 19, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n25/pdf/scientificamerican06191880-392.pdf.

Anonymous:1880:RCR

[470] Anonymous. Roofed country roads. Scientific American, 42(25): 392–393, June 19, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n25/pdf/scientificamerican06191880-
392a.pdf.

Anonymous:1880:RDRd

[471] Anonymous. Recent decisions relating to patents, copyrights, etc. Scientific American, 42(25):394, June 19, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v42/
n25/pdf/scientificamerican06191880-394.pdf.

Anonymous:1880:Ck

[472] Anonymous. Correspondence. Scientific American, 42(26):405, June 26, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n26/pdf/scientificamerican06261880-405.pdf.

Anonymous:1880:BPu

[473] Anonymous. Business and personal. Scientific American, 42(26):409, June 26, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n26/pdf/scientificamerican06261880-409.pdf.

Anonymous:1880:BNBn

[474] Anonymous. Books: New books and publications. Scientific American, 42(26):409, June 26, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n26/pdf/scientificamerican06261880-
409a.pdf.

Anonymous:1880:NQr

[475] Anonymous. Notes and queries. Scientific American, 42(26):409–410, June 26, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n26/pdf/scientificamerican06261880-409b.pdf.

Anonymous:1880:IIt

[476] Anonymous. Index of inventions. Scientific American, 42(26):410–412, June 26, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n26/pdf/scientificamerican06261880-410.pdf.

Anonymous:1880:AINo

[477] Anonymous. American industries — no 47. Scientific American, 42(26):399–400, June 26, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n26/pdf/scientificamerican06261880-
399.pdf.

Anonymous:1880:ASCa

[478] Anonymous. The American Society of Civil Engineers. Scientific American, 42(26):400, June 26, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n26/pdf/scientificamerican06261880-
400.pdf.

Anonymous:1880:NPSa

[479] Anonymous. A new plan for small remittances. Scientific American, 42(26):400, June 26, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n26/pdf/scientificamerican06261880-
400a.pdf.

Anonymous:1880:PPP

[480] Anonymous. The proportion of patents to population. Scientific American, 42(26):400, June 26, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n26/pdf/scientificamerican06261880-
400b.pdf.

Anonymous:1880:CMT

[481] Anonymous. Cyprien M. Tessie Du Motay. Scientific American, 42(26):400–401, June 26, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n26/pdf/scientificamerican06261880-
400c.pdf.

Anonymous:1880:AW

[482] Anonymous. The Army worm. Scientific American, 42(26):401, June 26, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n26/pdf/scientificamerican06261880-401.pdf.

Anonymous:1880:MIbc

[483] Anonymous. Mechanical inventions. Scientific American, 42(26): 402–403, June 26, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n26/pdf/scientificamerican06261880-
402.pdf.

Anonymous:1880:HEC

[484] Anonymous. Hydraulic elevator for canal boats. Scientific American, 42(26):402–403, June 26, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n26/pdf/scientificamerican06261880-
402a.pdf.

Anonymous:1880:ISPe

[485] Anonymous. Improved stylographic pen. Scientific American, 42 (26):403, June 26, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n26/pdf/scientificamerican06261880-
403.pdf.

Anonymous:1880:VBS

[486] Anonymous. Ventilation of boots and shoes. Scientific American, 42(26):403, June 26, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n26/pdf/scientificamerican06261880-
403a.pdf.

Anonymous:1880:PIG

[487] Anonymous. Price indicator for gas meters. Scientific American, 42(26):403–404, June 26, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n26/pdf/scientificamerican06261880-
403b.pdf.

Anonymous:1880:ISBa

[488] Anonymous. Improved steam boiler. Scientific American, 42(26):404, June 26, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n26/pdf/scientificamerican06261880-404.pdf.

Anonymous:1880:EIk

[489] Anonymous. Engineering inventions. Scientific American, 42(26): 404–405, June 26, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n26/pdf/scientificamerican06261880-
404a.pdf.

Anonymous:1880:MIbd

[490] Anonymous. Miscellaneous inventions. Scientific American, 42(26):405, June 26, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n26/pdf/scientificamerican06261880-405a.pdf.

Anonymous:1880:IBM

[491] Anonymous. Improved brick machine. Scientific American, 42 (26):406–407, June 26, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n26/pdf/scientificamerican06261880-
406.pdf.

Niaudet:1880:NTE

[492] M. A. Niaudet. New thermo-electric battery. Scientific American, 42(26):407, June 26, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n26/pdf/scientificamerican06261880-
407.pdf.

Anonymous:1880:CCb

[493] Anonymous. The carp and its culture. Scientific American, 42 (26):407–408, June 26, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n26/pdf/scientificamerican06261880-
407a.pdf.

Anonymous:1880:RDRe

[494] Anonymous. Recent decisions relating to patents. Scientific American, 42(26):408–409, June 26, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n26/pdf/scientificamerican06261880-
408.pdf.

Anonymous:1880:Cl

[495] Anonymous. Correspondence. Scientific American, 42(3):37, January 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n3/pdf/scientificamerican01171880-37a.pdf.

Anonymous:1880:BPv

[496] Anonymous. Business and personal. Scientific American, 42(3):44, January 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n3/pdf/scientificamerican01171880-44.pdf.

Anonymous:1880:NQs

[497] Anonymous. Notes and queries. Scientific American, 42(3):44, January 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n3/pdf/scientificamerican01171880-44a.pdf.

Anonymous:1880:IIu

[498] Anonymous. Index of inventions. Scientific American, 42(3):44, January 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n3/pdf/scientificamerican01171880-44c.pdf.

Anonymous:1880:PNPl

[499] Anonymous. Patent notes: Patents. Scientific American, 42(3):45, January 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n3/pdf/scientificamerican01171880-45.pdf.

Anonymous:1880:PLC

[500] Anonymous. Patent legislators in Congress. Scientific American, 42(3):32, January 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n3/pdf/scientificamerican01171880-
32.pdf.

Anonymous:1880:NDS

[501] Anonymous. A new deep sea sounding apparatus. Scientific American, 42(3):32, January 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n3/pdf/scientificamerican01171880-
32a.pdf.

Anonymous:1880:PTB

[502] Anonymous. Prospects of trade in Brazil. Scientific American, 42(3):32–33, January 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n3/pdf/scientificamerican01171880-
32b.pdf.

Anonymous:1880:NMPa

[503] Anonymous. New method of producing photographic pictures in colors. Scientific American, 42(3):33, January 17, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v42/
n3/pdf/scientificamerican01171880-33.pdf.

Anonymous:1880:TBDa

[504] Anonymous. The Tay Bridge Disaster. Scientific American, 42 (3):33, January 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n3/pdf/scientificamerican01171880-
33a.pdf.

Anonymous:1880:OVS

[505] Anonymous. Our venomous snakes. Scientific American, 42(3):33, January 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n3/pdf/scientificamerican01171880-33b.pdf.

Anonymous:1880:EVA

[506] Anonymous. Edison’s vacuum apparatus. Scientific American, 42 (3):34–35, January 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n3/pdf/scientificamerican01171880-
34.pdf.

Anonymous:1880:LA

[507] Anonymous. Laboratory apparatus. Scientific American, 42(3):35, January 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n3/pdf/scientificamerican01171880-35.pdf.

Anonymous:1880:NDR

[508] Anonymous. New damper regulator. Scientific American, 42(3):36, January 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n3/pdf/scientificamerican01171880-36.pdf.

Anonymous:1880:NTCb

[509] Anonymous. Novel toilet cabinet. Scientific American, 42(3):36, January 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n3/pdf/scientificamerican01171880-36a.pdf.

Anonymous:1880:SLP

[510] Anonymous. Stereoscopic lantern pictures. Scientific American, 42(3):36–37, January 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n3/pdf/scientificamerican01171880-
36b.pdf.

Anonymous:1880:AINp

[511] Anonymous. American industries — no. 28. Scientific American, 42(3):37, January 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n3/pdf/scientificamerican01171880-
37.pdf.

Baldwin:1880:HYSa

[512] Wm. J. Baldwin. Hints to the young steam fitter. Scientific American, 42(3):38–39, January 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n3/pdf/scientificamerican01171880-
38.pdf.

Anonymous:1880:NCb

[513] Anonymous. A novel clock. Scientific American, 42(3):39, January 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n3/pdf/scientificamerican01171880-39.pdf.

Anonymous:1880:AV

[514] Anonymous. The aard vark. Scientific American, 42(3):39–43, January 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n3/pdf/scientificamerican01171880-39a.pdf.

Anonymous:1880:MIbe

[515] Anonymous. Miscellaneous inventions. Scientific American, 42 (3):43, January 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n3/pdf/scientificamerican01171880-
43.pdf.

Anonymous:1880:CRo

[516] Anonymous. Communications received. Scientific American, 42 (3):44, January 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n3/pdf/scientificamerican01171880-
44b.pdf.

Anonymous:1880:Cm

[517] Anonymous. Correspondence. Scientific American, 42(4):52, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n4/pdf/scientificamerican01241880-52.pdf.

Anonymous:1880:BPw

[518] Anonymous. Business and personal. Scientific American, 42(4):59, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n4/pdf/scientificamerican01241880-59a.pdf.

Anonymous:1880:BNBo

[519] Anonymous. Books: New books and publications. Scientific American, 42(4):59, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n4/pdf/scientificamerican01241880-
59b.pdf.

Anonymous:1880:NQt

[520] Anonymous. Notes and queries. Scientific American, 42(4):59, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n4/pdf/scientificamerican01241880-59c.pdf.

Anonymous:1880:IIv

[521] Anonymous. Index of inventions. Scientific American, 42(4):60, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n4/pdf/scientificamerican01241880-60.pdf.

Anonymous:1880:PNPm

[522] Anonymous. Patent notes: Patents. Scientific American, 42(4):60, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n4/pdf/scientificamerican01241880-60a.pdf.

Anonymous:1880:ASG

[523] Anonymous. Action of sewer gas on lead, etc. Scientific American, 42(4):48, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n4/pdf/scientificamerican01241880-
48.pdf.

Anonymous:1880:TY

[524] Anonymous. The typical Yankee. Scientific American, 42(4):48, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n4/pdf/scientificamerican01241880-48a.pdf.

Anonymous:1880:BRK

[525] Anonymous. A buried race in Kansas. Scientific American, 42 (4):48, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n4/pdf/scientificamerican01241880-
48b.pdf.

Anonymous:1880:TCN

[526] Anonymous. Taking down Cleopatra’s Needle. Scientific American, 42(4):48–49, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n4/pdf/scientificamerican01241880-
48c.pdf.

Anonymous:1880:MCG

[527] Anonymous. Moles at the Cape of Good Hope. Scientific American, 42(4):49–50, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n4/pdf/scientificamerican01241880-
49.pdf.

Anonymous:1880:NSc

[528] Anonymous. Novel sieve. Scientific American, 42(4):50, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n4/pdf/scientificamerican01241880-50.pdf.

Anonymous:1880:NTT

[529] Anonymous. Novel tap for tin cans. Scientific American, 42(4):50, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n4/pdf/scientificamerican01241880-50a.pdf.

Anonymous:1880:SSB

[530] Anonymous. Soap and shaving box. Scientific American, 42 (4):50–51, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n4/pdf/scientificamerican01241880-
50b.pdf.

Anonymous:1880:NHHb

[531] Anonymous. New hand hold for reins. Scientific American, 42 (4):51, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n4/pdf/scientificamerican01241880-
51.pdf.

Anonymous:1880:NVb

[532] Anonymous. Novel vise. Scientific American, 42(4):51, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n4/pdf/scientificamerican01241880-51a.pdf.

Anonymous:1880:ISSb

[533] Anonymous. Improved shawl strap. Scientific American, 42(4):51, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n4/pdf/scientificamerican01241880-51b.pdf.

Anonymous:1880:MIbf

[534] Anonymous. Miscellaneous inventions. Scientific American, 42 (4):51–52, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n4/pdf/scientificamerican01241880-
51c.pdf.

Anonymous:1880:AINq

[535] Anonymous. American industries — no. 29. Scientific American, 42(4):53, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n4/pdf/scientificamerican01241880-
53.pdf.

Anonymous:1880:AIi

[536] Anonymous. Agricultural inventions. Scientific American, 42 (4):54, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n4/pdf/scientificamerican01241880-
54.pdf.

Anonymous:1880:NO

[537] Anonymous. New optometer. Scientific American, 42(4):54, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n4/pdf/scientificamerican01241880-54a.pdf.

Anonymous:1880:NAHa

[538] Anonymous. New adjustable hose nozzle. Scientific American, 42(4):54, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n4/pdf/scientificamerican01241880-
54b.pdf.

Anonymous:1880:MIbg

[539] Anonymous. Mechanical inventions. Scientific American, 42(4):54, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n4/pdf/scientificamerican01241880-54c.pdf.

Anonymous:1880:SH

[540] Anonymous. The spring Haas. Scientific American, 42(4):55, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n4/pdf/scientificamerican01241880-55.pdf.

Anonymous:1880:CV

[541] Anonymous. Chinese vase. Scientific American, 42(4):55–58, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n4/pdf/scientificamerican01241880-55a.pdf.

Anonymous:1880:NBCb

[542] Anonymous. Nasal and bronchial catarrh — a strong record. Scientific American, 42(4):59, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n4/pdf/scientificamerican01241880-
59.pdf.

Anonymous:1880:CRp

[543] Anonymous. Communications received. Scientific American, 42 (4):59, January 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n4/pdf/scientificamerican01241880-
59d.pdf.

Anonymous:1880:INIg

[544] Anonymous. Inventions: New inventions. Scientific American, 42(5):66, January 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n5/pdf/scientificamerican01311880-
66.pdf.

Anonymous:1880:BPx

[545] Anonymous. Business and personal. Scientific American, 42(5):75, January 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n5/pdf/scientificamerican01311880-75.pdf.

Anonymous:1880:BNBp

[546] Anonymous. Books: New books and publications. Scientific American, 42(5):75, January 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n5/pdf/scientificamerican01311880-
75a.pdf.

Anonymous:1880:NQu

[547] Anonymous. Notes and queries. Scientific American, 42(5):75, January 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n5/pdf/scientificamerican01311880-75b.pdf.

Anonymous:1880:IIw

[548] Anonymous. Index of inventions. Scientific American, 42(5):75–76, January 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n5/pdf/scientificamerican01311880-75c.pdf.

Anonymous:1880:ADEa

[549] Anonymous. Application of dynamo-electric machines to telegraphy. Scientific American, 42(5):63–64, January 31, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v42/
n5/pdf/scientificamerican01311880-63.pdf.

Anonymous:1880:WFa

[550] Anonymous. The World’s Fair of 1883. Scientific American, 42 (5):64, January 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n5/pdf/scientificamerican01311880-
64.pdf.

Anonymous:1880:EMT

[551] Anonymous. Electric machines in telegraphy. Scientific American, 42(5):64, January 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n5/pdf/scientificamerican01311880-
64a.pdf.

Anonymous:1880:FE

[552] Anonymous. The future of electricity. Scientific American, 42 (5):64–65, January 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n5/pdf/scientificamerican01311880-
64b.pdf.

Anonymous:1880:TMP

[553] Anonymous. Two more Patent Bills. Scientific American, 42(5):65, January 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n5/pdf/scientificamerican01311880-65.pdf.

Anonymous:1880:TBDb

[554] Anonymous. The Tay Bridge Disaster. Scientific American, 42 (5):65, January 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n5/pdf/scientificamerican01311880-
65a.pdf.

Anonymous:1880:PYW

[555] Anonymous. The past year’s work in the Patent Office. Scientific American, 42(5):65, January 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n5/pdf/scientificamerican01311880-
65b.pdf.

Anonymous:1880:HTG

[556] Anonymous. Heating tires by gas jets. Scientific American, 42 (5):66, January 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n5/pdf/scientificamerican01311880-
66a.pdf.

Anonymous:1880:MSa

[557] Anonymous. The manufacture of soda. Scientific American, 42 (5):67, January 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n5/pdf/scientificamerican01311880-
67.pdf.

Anonymous:1880:NVc

[558] Anonymous. A new ventilator. Scientific American, 42(5):67, January 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n5/pdf/scientificamerican01311880-67a.pdf.

Anonymous:1880:MIbh

[559] Anonymous. Miscellaneous inventions. Scientific American, 42 (5):67–68, January 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n5/pdf/scientificamerican01311880-
67b.pdf.

Anonymous:1880:ISHb

[560] Anonymous. Improved steam hammer. Scientific American, 42 (5):68, January 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n5/pdf/scientificamerican01311880-
68.pdf.

Anonymous:1880:PPS

[561] Anonymous. The Peerless portable steam engine. Scientific American, 42(5):68, January 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n5/pdf/scientificamerican01311880-
68a.pdf.

Anonymous:1880:MIbi

[562] Anonymous. Mechanical inventions. Scientific American, 42(5): 68–69, January 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n5/pdf/scientificamerican01311880-
68b.pdf.

Anonymous:1880:AIj

[563] Anonymous. Agricultural inventions. Scientific American, 42 (5):70, January 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n5/pdf/scientificamerican01311880-
70.pdf.

Anonymous:1880:NSDc

[564] Anonymous. New system of digging and curbing wells. Scientific American, 42(5):70, January 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n5/pdf/scientificamerican01311880-
70a.pdf.

Anonymous:1880:LL

[565] Anonymous. Light locomotives. Scientific American, 42(5):70–71, January 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n5/pdf/scientificamerican01311880-70b.pdf.

Anonymous:1880:HD

[566] Anonymous. The hyena dog. Scientific American, 42(5):71–74, January 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n5/pdf/scientificamerican01311880-71.pdf.

Anonymous:1880:LST

[567] Anonymous. A labor-saving tool. Scientific American, 42(5):74, January 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n5/pdf/scientificamerican01311880-74.pdf.

Anonymous:1880:CCc

[568] Anonymous. Correspondence: Correspondents. Scientific American, 42(6):83, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n6/pdf/scientificamerican02071880-
84.pdf.

Anonymous:1880:INIh

[569] Anonymous. Inventions: New inventions. Scientific American, 42(6):85, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n6/pdf/scientificamerican02071880-
85a.pdf.

Anonymous:1880:BPy

[570] Anonymous. Business and personal. Scientific American, 42 (6):91, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n6/pdf/scientificamerican02071880-
91.pdf.

Anonymous:1880:NQv (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n6/pdf/scientificamerican02071880-
91a.pdf.

Anonymous:1880:IIx

[572] Anonymous. Index of inventions. Scientific American, 42(6):92, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n6/pdf/scientificamerican02071880-92.pdf.

Anonymous:1880:FDN

[573] Anonymous. Floating derricks of New York Harbor. Scientific American, 42(6):79–80, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n6/pdf/scientificamerican02071880-
79.pdf.

Anonymous:1880:ILM

[574] Anonymous. Iron lighthouse for Mexico. Scientific American, 42(6):80, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n6/pdf/scientificamerican02071880-
80.pdf.

Anonymous:1880:WW

[575] Anonymous. Work and wages. Scientific American, 42(6):80, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n6/pdf/scientificamerican02071880-80a.pdf.

Anonymous:1880:CSC

[576] Anonymous. Cotton size and cotton sizing. Scientific American, 42(6):80–81, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n6/pdf/scientificamerican02071880-
80b.pdf.

Anonymous:1880:MKC

[577] Anonymous. Making knit cotton goods to imitate wool. Scientific American, 42(6):81, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n6/pdf/scientificamerican02071880-
81.pdf.

Anonymous:1880:CAC

[578] Anonymous. Compressed air in coal mining. Scientific American, 42(6):81, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n6/pdf/scientificamerican02071880-
81a.pdf.

Anonymous:1880:IPE

[579] Anonymous. Improved portable engines. Scientific American, 42 (6):82, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n6/pdf/scientificamerican02071880-
82.pdf.

Anonymous:1880:AIk

[580] Anonymous. Agricultural inventions. Scientific American, 42(6): 82–83, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n6/pdf/scientificamerican02071880-
82a.pdf.

Anonymous:1880:NET

[581] Anonymous. New egg tongs. Scientific American, 42(6):83, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n6/pdf/scientificamerican02071880-83.pdf.

Anonymous:1880:IMB

[582] Anonymous. Improved medicine bottle. Scientific American, 42 (6):83, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n6/pdf/scientificamerican02071880-
83a.pdf.

Anonymous:1880:IAT

[583] Anonymous. Improved animal trap. Scientific American, 42 (6):83, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n6/pdf/scientificamerican02071880-
83b.pdf.

Anonymous:1880:IJb

[584] Anonymous. Improvement in jugs. Scientific American, 42(6):83, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n6/pdf/scientificamerican02071880-83c.pdf.

Anonymous:1880:ASd

[585] Anonymous. The antiquity of the spoon. Scientific American, 42(6):84–85, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n6/pdf/scientificamerican02071880-
84a.pdf.

Hopkins:1880:CCB

[586] Geo. M. Hopkins. On the crystallization of Canada balsam. Scientific American, 42(6):85, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n6/pdf/scientificamerican02071880-
85.pdf.

Anonymous:1880:NSV

[587] Anonymous. New system of ventilation. Scientific American, 42 (6):86, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n6/pdf/scientificamerican02071880-
86.pdf.

Anonymous:1880:MIbj

[588] Anonymous. Miscellaneous inventions. Scientific American, 42 (6):86–87, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n6/pdf/scientificamerican02071880-
86a.pdf.

Beard:1880:ONG

[589] Daniel C. Beard. One of nature’s gluttons. Scientific American, 42(6):87–88, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n6/pdf/scientificamerican02071880-
87.pdf.

Anonymous:1880:ALP

[590] Anonymous. The action of light on plants. Scientific American, 42(6):88, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n6/pdf/scientificamerican02071880-
88.pdf.

Anonymous:1880:IUF

[591] Anonymous. The industrial uses of fish skins. Scientific American, 42(6):88–89, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n6/pdf/scientificamerican02071880-
88a.pdf.

Anonymous:1880:MIbk

[592] Anonymous. Mechanical inventions. Scientific American, 42(6): 89, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n6/pdf/scientificamerican02071880-
89.pdf.

Anonymous:1880:CPV

[593] Anonymous. Chinese porcelain vase. Scientific American, 42 (6):89, February 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n6/pdf/scientificamerican02071880-
89a.pdf.

Anonymous:1880:RDRf

[594] Anonymous. Recent decisions relating to patents, assignments, etc. Scientific American, 42(6):90–91, February 7, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v42/
n6/pdf/scientificamerican02071880-90.pdf.

Anonymous:1880:INIi

[595] Anonymous. Inventions: New inventions. Scientific American, 42 (7):98–99, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n7/pdf/scientificamerican02141880-
98a.pdf.

Anonymous:1880:BPz

[596] Anonymous. Business and personal. Scientific American, 42(7): 107, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n7/pdf/scientificamerican02141880-
107.pdf.

Anonymous:1880:BNBq

[597] Anonymous. Books: New books and publications. Scientific American, 42(7):107, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n7/pdf/scientificamerican02141880-
107a.pdf.

Anonymous:1880:NQw

[598] Anonymous. Notes and queries. Scientific American, 42(7):107, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n7/pdf/scientificamerican02141880-107b.pdf.

Anonymous:1880:IIy

[599] Anonymous. Index of inventions. Scientific American, 42(7): 107–108, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n7/pdf/scientificamerican02141880-
107d.pdf.

Anonymous:1880:PNPn

[600] Anonymous. Patent notes: Patents. Scientific American, 42 (7):108, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n7/pdf/scientificamerican02141880-
108.pdf.

Anonymous:1880:HGI

[601] Anonymous. How great inventions are made. Scientific American, 42(7):96, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n7/pdf/scientificamerican02141880-
96.pdf.

Anonymous:1880:IAP

[602] Anonymous. The impolicy of anti-patent combinations. Scientific American, 42(7):96–97, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v42/n7/pdf/scientificamerican02141880-
96a.pdf.

Anonymous:1880:WRA

[603] Anonymous. The Western Railway Association and the Patent Laws. Scientific American, 42(7):97, February 14, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v42/
n7/pdf/scientificamerican02141880-97.pdf.

Anonymous:1880:PWF

[604] Anonymous. The proposed World’s Fair in 1883. Scientific American, 42(7):97, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n7/pdf/scientificamerican02141880-
97a.pdf.

Anonymous:1880:LGC

[605] Anonymous. The logic of Government control. Scientific American, 42(7):97, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n7/pdf/scientificamerican02141880-
97b.pdf.

Anonymous:1880:OCG

[606] Anonymous. The owner of the Cape of Good Hope and his ostrich farm. Scientific American, 42(7):97, February 14, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v42/
n7/pdf/scientificamerican02141880-97c.pdf.

Baldwin:1880:HYSb

[607] W. M. J. Baldwin. Hints to the young steam fitter. Scientific American, 42(7):98, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n7/pdf/scientificamerican02141880-
98.pdf.

Anonymous:1880:NPMb

[608] Anonymous. Novel pedo-motor. Scientific American, 42(7):99, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n7/pdf/scientificamerican02141880-99.pdf.

Anonymous:1880:IBb

[609] Anonymous. Improved blowpipe. Scientific American, 42(7):99, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n7/pdf/scientificamerican02141880-99a.pdf.

Anonymous:1880:MIbl

[610] Anonymous. Mechanical inventions. Scientific American, 42(7): 99–100, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n7/pdf/scientificamerican02141880-
99b.pdf.

Anonymous:1880:IDT

[611] Anonymous. Improved draught tug springs. Scientific American, 42(7):100, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n7/pdf/scientificamerican02141880-
100.pdf.

Anonymous:1880:CMS

[612] Anonymous. Combined milk strainer and stool. Scientific American, 42(7):100, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n7/pdf/scientificamerican02141880-
100a.pdf.

Anonymous:1880:MIbm

[613] Anonymous. Miscellaneous inventions. Scientific American, 42 (7):100, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n7/pdf/scientificamerican02141880-
100b.pdf.

Anonymous:1880:AINr

[614] Anonymous. American industries — no. 30. Scientific American, 42(7):101–102, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n7/pdf/scientificamerican02141880-
101.pdf.

Anonymous:1880:NWH

[615] Anonymous. New weighted horseshoe. Scientific American, 42 (7):102, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n7/pdf/scientificamerican02141880-
102.pdf.

Anonymous:1880:IFa

[616] Anonymous. Improvement in fireplaces. Scientific American, 42 (7):102, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n7/pdf/scientificamerican02141880-
102a.pdf.

Anonymous:1880:RIb

[617] Anonymous. Recent inventions. Scientific American, 42(7):102, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n7/pdf/scientificamerican02141880-102b.pdf.

Anonymous:1880:MG

[618] Anonymous. The Moholi Galago. Scientific American, 42(7):103, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n7/pdf/scientificamerican02141880-103.pdf.

Anonymous:1880:SCb

[619] Anonymous. Spider crab. Scientific American, 42(7):103, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n7/pdf/scientificamerican02141880-103a.pdf.

Anonymous:1880:NHNe

[620] Anonymous. Natural history notes. Scientific American, 42(7): 103–104, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n7/pdf/scientificamerican02141880-
103b.pdf.

Anonymous:1880:AEGb

[621] Anonymous. Ancient Egyptian glassware. Scientific American, 42 (7):105–106, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n7/pdf/scientificamerican02141880-
105.pdf.

Anonymous:1880:AIl

[622] Anonymous. Agricultural inventions. Scientific American, 42(7): 106, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n7/pdf/scientificamerican02141880-
106.pdf.

Anonymous:1880:CRq

[623] Anonymous. Communications received. Scientific American, 42 (7):107, February 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n7/pdf/scientificamerican02141880-
107c.pdf.

Anonymous:1880:Cn

[624] Anonymous. Correspondence. Scientific American, 42(8):116, February 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n8/pdf/scientificamerican02211880-116.pdf.

Anonymous:1880:BPba

[625] Anonymous. Business and personal. Scientific American, 42(8): 123, February 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n8/pdf/scientificamerican02211880-
123.pdf.

Anonymous:1880:BNBr

[626] Anonymous. Books: New books and publications. Scientific American, 42(8):123, February 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n8/pdf/scientificamerican02211880-
123a.pdf.

Anonymous:1880:NQx

[627] Anonymous. Notes and queries. Scientific American, 42(8):123– 124, February 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n8/pdf/scientificamerican02211880-
123b.pdf.

Anonymous:1880:IIz

[628] Anonymous. Index of inventions. Scientific American, 42(8):124, February 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n8/pdf/scientificamerican02211880-124.pdf.

Anonymous:1880:PNPo

[629] Anonymous. Patent notes: Patents. Scientific American, 42 (8):125, February 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n8/pdf/scientificamerican02211880-
125.pdf.

Anonymous:1880:MCP

[630] Anonymous. Mr. Converse’s Patent Bill. Scientific American, 42(8):112, February 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n8/pdf/scientificamerican02211880-
112.pdf.

Anonymous:1880:HSH

[631] Anonymous. “How steel hardens”. Scientific American, 42(8): 112–113, February 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n8/pdf/scientificamerican02211880-
112a.pdf.

Anonymous:1880:AEGc

[632] Anonymous. Approximate economy of gas and electric lighting. Scientific American, 42(8):113, February 21, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v42/
n8/pdf/scientificamerican02211880-113.pdf.

Anonymous:1880:CMN

[633] Anonymous. The corundum mines of North Carolina. Scientific American, 42(8):113, February 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n8/pdf/scientificamerican02211880-
113a.pdf.

Anonymous:1880:WTG

[634] Anonymous. Why the thunderer’s gun burst. Scientific American, 42(8):113, February 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n8/pdf/scientificamerican02211880-
113b.pdf.

Anonymous:1880:AISb

[635] Anonymous. Artificial ice skating rink, New York.,and more. Scientific American, 42(8):113, February 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v42/n8/pdf/scientificamerican02211880-
113c.pdf.

Anonymous:1880:IRD

[636] Anonymous. Improved revolving derrick. Scientific American, 42 (8):114, February 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n8/pdf/scientificamerican02211880-
114.pdf.

Anonymous:1880:NEB

[637] Anonymous. New egg beater. Scientific American, 42(8):115, February 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n8/pdf/scientificamerican02211880-115.pdf.

Anonymous:1880:NAEa

[638] Anonymous. New agricultural engine. Scientific American, 42 (8):115, February 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n8/pdf/scientificamerican02211880-
115a.pdf.

Anonymous:1880:LPE

[639] Anonymous. Life preserver exhibitor. Scientific American, 42 (8):115, February 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n8/pdf/scientificamerican02211880-
115b.pdf.

Anonymous:1880:AIm

[640] Anonymous. Agricultural inventions. Scientific American, 42(8): 115–116, February 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n8/pdf/scientificamerican02211880-
115c.pdf.

Anonymous:1880:AINs

[641] Anonymous. American industries — no 31. Scientific American, 42(8):117, February 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n8/pdf/scientificamerican02211880-
117.pdf.

Anonymous:1880:BAS

[642] Anonymous. The binary absorption system of refrigeration. Scientific American, 42(8):118, February 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v42/n8/pdf/scientificamerican02211880-
118.pdf.

Anonymous:1880:NSAa

[643] Anonymous. New steering apparatus. Scientific American, 42 (8):118, February 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n8/pdf/scientificamerican02211880-
118a.pdf.

Roberts:1880:MLH

[644] A. W. Roberts. Marine and land hermit crabs. Scientific American, 42(8):119–121, February 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n8/pdf/scientificamerican02211880-
119.pdf.

Anonymous:1880:SV

[645] Anonymous. The steam velocipede. Scientific American, 42(8): 121, February 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n8/pdf/scientificamerican02211880-
121.pdf.

Anonymous:1880:ESa

[646] Anonymous. The electric Sun. Scientific American, 42(8):121, February 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n8/pdf/scientificamerican02211880-121a.pdf.

Anonymous:1880:MIbn

[647] Anonymous. Miscellaneous inventions. Scientific American, 42(8): 121–122, February 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n8/pdf/scientificamerican02211880-
121b.pdf.

Anonymous:1880:EIl

[648] Anonymous. Engineering inventions. Scientific American, 42(8): 122, February 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n8/pdf/scientificamerican02211880-
122.pdf.

Anonymous:1880:BPbb

[649] Anonymous. Business and personal. Scientific American, 42(9): 139, February 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n9/pdf/scientificamerican02281880-
139.pdf.

Anonymous:1880:BNBs

[650] Anonymous. Books: New books and publications. Scientific American, 42(9):139, February 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n9/pdf/scientificamerican02281880-
139a.pdf.

Anonymous:1880:NQy

[651] Anonymous. Notes and queries. Scientific American, 42(9):139, February 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n9/pdf/scientificamerican02281880-139b.pdf.

Anonymous:1880:IIba

[652] Anonymous. Index of inventions. Scientific American, 42(9): 139–140, February 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n9/pdf/scientificamerican02281880-
139d.pdf.

Anonymous:1880:PNPp

[653] Anonymous. Patent notes: Patents. Scientific American, 42 (9):140, February 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n9/pdf/scientificamerican02281880-
140a.pdf.

Anonymous:1880:PHRa

[654] Anonymous. Passage by the House of Representatives of a law for the encouragement of swindling. Scientific American, 42(9):128, February 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n9/pdf/scientificamerican02281880-128.pdf.

Anonymous:1880:FVB

[655] Anonymous. Flour vs bran. Scientific American, 42(9):128–129, February 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n9/pdf/scientificamerican02281880-128a.pdf.

Anonymous:1880:NGB

[656] Anonymous. A new Government Bureau proposed. Scientific American, 42(9):129, February 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n9/pdf/scientificamerican02281880-
129.pdf.

Anonymous:1880:PFa

[657] Anonymous. The prevention of fires. Scientific American, 42 (9):129, February 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n9/pdf/scientificamerican02281880-
129a.pdf.

Anonymous:1880:COY

[658] Anonymous. Chastant’s observations on yellow fever. Scientific American, 42(9):129, February 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n9/pdf/scientificamerican02281880-
129b.pdf.

Anonymous:1880:AINt

[659] Anonymous. American industries, no 32. Scientific American, 42 (9):130–131, February 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n9/pdf/scientificamerican02281880-
130.pdf.

Anonymous:1880:NVAa

[660] Anonymous. New vise and anvil drill. Scientific American, 42 (9):131, February 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n9/pdf/scientificamerican02281880-
131.pdf.

Anonymous:1880:EPI

[661] Anonymous. Electrical pressure indicator. Scientific American, 42(9):131, February 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n9/pdf/scientificamerican02281880-
131a.pdf.

Anonymous:1880:IELa

[662] Anonymous. Improved electric lamp. Scientific American, 42 (9):131, February 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n9/pdf/scientificamerican02281880-
131b.pdf.

Anonymous:1880:EIm

[663] Anonymous. Engineering inventions. Scientific American, 42(9): 131–133, February 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n9/pdf/scientificamerican02281880-
131c.pdf.

Anonymous:1880:AIo

[664] Anonymous. Agricultural inventions. Scientific American, 42(9): 133, February 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n9/pdf/scientificamerican02281880-
133.pdf.

Anonymous:1880:NCP

[665] Anonymous. New cotton picking shade. Scientific American, 42 (9):134, February 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n9/pdf/scientificamerican02281880-
134.pdf.

Anonymous:1880:PSH

[666] Anonymous. Pumping system for hydraulic presses. Scientific American, 42(9):134–135, February 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n9/pdf/scientificamerican02281880-
134a.pdf.

Anonymous:1880:H

[667] Anonymous. The hyrax. Scientific American, 42(9):135, February 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n9/pdf/scientificamerican02281880-135.pdf.

Seiss:1880:PF

[668] C. Few Seiss. The pickerel frog. Scientific American, 42(9): 135–137, February 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n9/pdf/scientificamerican02281880-
135a.pdf.

Anonymous:1880:ER

[669] Anonymous. Electrical railway. Scientific American, 42(9):137, February 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v42/n9/pdf/scientificamerican02281880-137.pdf.

Anonymous:1880:MIbo

[670] Anonymous. Mechanical inventions. Scientific American, 42(9): 137–138, February 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n9/pdf/scientificamerican02281880-
137a.pdf.

Anonymous:1880:MIbp

[671] Anonymous. Miscellaneous inventions. Scientific American, 42 (9):138, February 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n9/pdf/scientificamerican02281880-
138.pdf.

Anonymous:1880:CRr

[672] Anonymous. Communications received. Scientific American, 42 (9):139, February 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v42/n9/pdf/scientificamerican02281880-
139c.pdf.

Anonymous:1880:SAEb

[673] Anonymous. The Scientific American export edition published monthly. Scientific American, 42(9):140, February 28, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v42/
n9/pdf/scientificamerican02281880-140.pdf.

Anonymous:1880:BPbc

[674] Anonymous. Business and personal. Scientific American, 43(1):12, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n1/pdf/scientificamerican07031880-12.pdf.

Anonymous:1880:IIbb

[675] Anonymous. Index of inventions. Scientific American, 43(1):12–14, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n1/pdf/scientificamerican07031880-12a.pdf.

Anonymous:1880:AIp

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n1/pdf/scientificamerican07031880-1.pdf.

Anonymous:1880:SCP

[677] Anonymous. Simple and cheap process of gas making. Scientific American, 43(1):1–2, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n1/pdf/scientificamerican07031880-
1a.pdf.

Anonymous:1880:SAWa

[678] Anonymous. The superiority of American watches. Scientific American, 43(1):2, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n1/pdf/scientificamerican07031880-
2.pdf.

Anonymous:1880:RLS

[679] Anonymous. A remarkable little steamer. Scientific American, 43(1):2, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n1/pdf/scientificamerican07031880-2a.pdf.

Anonymous:1880:SEb

[680] Anonymous. A strange epidemic. Scientific American, 43(1):2–3, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n1/pdf/scientificamerican07031880-2b.pdf.

Anonymous:1880:COT

[681] Anonymous. Cannonading of oil tanks. Scientific American, 43(1):3, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n1/pdf/scientificamerican07031880-3.pdf.

Anonymous:1880:NFb

[682] Anonymous. Navigation in fogs. Scientific American, 43(1):3, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n1/pdf/scientificamerican07031880-3a.pdf.

Anonymous:1880:EAQ

[683] Anonymous. Effect of age on the quality of iron. Scientific American, 43(1):3, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n1/pdf/scientificamerican07031880-
3b.pdf.

Anonymous:1880:CBL

[684] Anonymous. A collision between large passenger steamers. Scientific American, 43(1):3, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n1/pdf/scientificamerican07031880-
3c.pdf.

Anonymous:1880:PBE

[685] Anonymous. Prevention of boiler explosions. Scientific American, 43(1): 4, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n1/pdf/scientificamerican07031880-4.pdf.

Anonymous:1880:MIbq

[686] Anonymous. Miscellaneous inventions. Scientific American, 43(1):4–5, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n1/pdf/scientificamerican07031880-4a.pdf.

Anonymous:1880:IWD

[687] Anonymous. Improved watering dipper. Scientific American, 43(1):5, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n1/pdf/scientificamerican07031880-5.pdf.

Anonymous:1880:IELb

[688] Anonymous. Improved electric lamp. Scientific American, 43(1):5, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n1/pdf/scientificamerican07031880-5a.pdf.

Anonymous:1880:APF

[689] Anonymous. Apparatus for preserving fruit. Scientific American, 43(1): 5, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n1/pdf/scientificamerican07031880-5b.pdf.

Anonymous:1880:NSW

[690] Anonymous. Novel slate washer. Scientific American, 43(1):5, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n1/pdf/scientificamerican07031880-5c.pdf.

Anonymous:1880:NPSb

[691] Anonymous. New portable shower bath. Scientific American, 43(1):6, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n1/pdf/scientificamerican07031880-6.pdf.

Anonymous:1880:MIbr

[692] Anonymous. Mechanical inventions. Scientific American, 43(1):7, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n1/pdf/scientificamerican07031880-7.pdf.

Anonymous:1880:PBa

[693] Anonymous. Phyllirhoe Bucephala. Scientific American, 43(1):7, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n1/pdf/scientificamerican07031880-7a.pdf.

Anonymous:1880:FTG

[694] Anonymous. The force of tree growth. Scientific American, 43(1):7, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n1/pdf/scientificamerican07031880-7b.pdf.

Anonymous:1880:NHNf

[695] Anonymous. Natural history notes. Scientific American, 43(1):7, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n1/pdf/scientificamerican07031880-7c.pdf.

Anonymous:1880:SIE

[696] Anonymous. Sydney International Exhibition — 1879–1880. Scientific American, 43(1):8–9, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n1/pdf/scientificamerican07031880-
8.pdf.

Anonymous:1880:EIn

[697] Anonymous. Engineering inventions. Scientific American, 43(1):10–11, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n1/pdf/scientificamerican07031880-10.pdf.

Anonymous:1880:ACb

[698] Anonymous. Aspirator and compressor. Scientific American, 43(1):11, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n1/pdf/scientificamerican07031880-11.pdf.

Anonymous:1880:NIb

[699] Anonymous. New inventions. Scientific American, 43(1):11, July 3, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n1/pdf/scientificamerican07031880-11a.pdf.

Anonymous:1880:INIj

[700] Anonymous. Inventions: New inventions. Scientific American, 43 (10):147–148, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
147b.pdf.

Anonymous:1880:Co

[701] Anonymous. Correspondence. Scientific American, 43(10):148, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n10/pdf/scientificamerican09041880-148b.pdf.

Anonymous:1880:BPbd

[702] Anonymous. Business and personal. Scientific American, 43 (10):155, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
155.pdf.

Anonymous:1880:BNBt

[703] Anonymous. Books: New books and publications. Scientific American, 43(10):155, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
155a.pdf.

Anonymous:1880:NQz

[704] Anonymous. Notes and queries. Scientific American, 43(10): 155, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
155b.pdf.

Anonymous:1880:IIbc

[705] Anonymous. Index of inventions. Scientific American, 43(10): 155–156, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
155d.pdf.

Anonymous:1880:PNPq

[706] Anonymous. Patent notes: Patents. Scientific American, 43 (10):157, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
157.pdf.

Anonymous:1880:ISBb

[707] Anonymous. The inspection of steam boilers. Scientific American, 43(10):144, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
144.pdf.

Anonymous:1880:DTG

[708] Anonymous. Dr. Tanner’s great fast. Scientific American, 43(10): 144–145, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
144a.pdf.

Anonymous:1880:ESBa

[709] Anonymous. Effect of starvation on the blood. Scientific American, 43(10):145, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
145.pdf.

Anonymous:1880:MPH

[710] Anonymous. Making profits out of human weakness. Scientific American, 43(10):145, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
145a.pdf.

Anonymous:1880:NAHb

[711] Anonymous. New appliance for harbor fires. Scientific American, 43(10):145, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
145b.pdf.

Anonymous:1880:RT

[712] Anonymous. Rain theories. Scientific American, 43(10):145, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n10/pdf/scientificamerican09041880-145c.pdf.

Anonymous:1880:LMM

[713] Anonymous. Lawn mowing made a pleasure. Scientific American, 43(10):146, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
146.pdf.

Anonymous:1880:ITK

[714] Anonymous. Improved tea kettle. Scientific American, 43(10): 146, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
146a.pdf.

Anonymous:1880:MIbs

[715] Anonymous. Miscellaneous inventions. Scientific American, 43(10): 146–147, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
146b.pdf.

Anonymous:1880:CFP

[716] Anonymous. Concrete fence post and sill. Scientific American, 43(10):147, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
147.pdf.

Anonymous:1880:NFS

[717] Anonymous. Novel folding stand. Scientific American, 43(10): 147, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
147a.pdf.

Anonymous:1880:CPBa

[718] Anonymous. Convenient portable bathing apparatus. Scientific American, 43(10):148, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
148.pdf.

Anonymous:1880:RJ

[719] Anonymous. Rope jumping. Scientific American, 43(10):148, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n10/pdf/scientificamerican09041880-148a.pdf.

Anonymous:1880:AINu

[720] Anonymous. American industries — no. 55. Scientific American, 43 (10):149–150, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
149.pdf.

Anonymous:1880:EIo

[721] Anonymous. Engineering inventions. Scientific American, 43(10): 150, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
150.pdf.

Anonymous:1880:SNE

[722] Anonymous. Some new electrical machinery. Scientific American, 43 (10):150–151, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
150a.pdf.

Seiss:1880:ST

[723] C. Few Seiss. The spotted triton. Scientific American, 43(10): 151, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
151.pdf.

Riley:1880:FNN

[724] C. V. Riley. Further notes on the northern Army worm. Scientific American, 43(10):152, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
152.pdf.

Anonymous:1880:DRPa

[725] Anonymous. Decisions relating to patents. Scientific American, 43 (10):152–153, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
152a.pdf.

Anonymous:1880:NPa

[726] Anonymous. Nickel plating. Scientific American, 43(10):153– 154, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
153.pdf.

Anonymous:1880:CRs

[727] Anonymous. Communications received. Scientific American, 43 (10):155, September 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n10/pdf/scientificamerican09041880-
155c.pdf.

Anonymous:1880:BPbe

[728] Anonymous. Business and personal. Scientific American, 43 (11):171, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n11/pdf/scientificamerican09111880-
171.pdf.

Anonymous:1880:NQba

[729] Anonymous. Notes and queries. Scientific American, 43(11): 171, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n11/pdf/scientificamerican09111880-
171a.pdf.

Anonymous:1880:IIbd

[730] Anonymous. Index of inventions. Scientific American, 43(11): 171–172, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n11/pdf/scientificamerican09111880-
171b.pdf.

Anonymous:1880:PNPr

[731] Anonymous. Patent notes: Patents. Scientific American, 43 (11):173, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n11/pdf/scientificamerican09111880-
173.pdf.

Anonymous:1880:PBP

[732] Anonymous. Point Bridge, Pittsburg, Pa. Scientific American, 43 (11):159–160, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n11/pdf/scientificamerican09111880-
159.pdf.

Anonymous:1880:NPe

[733] Anonymous. A novel projectile. Scientific American, 43(11): 160, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n11/pdf/scientificamerican09111880-
160.pdf.

Anonymous:1880:SFA

[734] Anonymous. Striking figures about cotton. Scientific American, 43(11):160, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n11/pdf/scientificamerican09111880-
160a.pdf.

Anonymous:1880:CPP

[735] Anonymous. A curious physical phenomenon. Scientific American, 43(11):160, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n11/pdf/scientificamerican09111880-
160b.pdf.

Anonymous:1880:DPP

[736] Anonymous. Do patents pay! Scientific American, 43(11): 160–161, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n11/pdf/scientificamerican09111880-
160c.pdf.

Anonymous:1880:CSP

[737] Anonymous. The Concord School of Philosophy. Scientific American, 43(11):161, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n11/pdf/scientificamerican09111880-
161.pdf.

Anonymous:1880:TMd

[738] Anonymous. Tin in Maine. Scientific American, 43(11):161, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n11/pdf/scientificamerican09111880-161a.pdf.

Anonymous:1880:ASAa

[739] Anonymous. The American Science Association. Scientific American, 43(11):161, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n11/pdf/scientificamerican09111880-
161b.pdf.

Anonymous:1880:MDCa

[740] Anonymous. Mining debris in California. Scientific American, 43 (11):161–162, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n11/pdf/scientificamerican09111880-
161c.pdf.

Anonymous:1880:SAWb

[741] Anonymous. Self-adjusting watch key. Scientific American, 43 (11):162, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n11/pdf/scientificamerican09111880-
162.pdf.

Anonymous:1880:NBa

[742] Anonymous. A novel binder. Scientific American, 43(11):162, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n11/pdf/scientificamerican09111880-162a.pdf.

Anonymous:1880:LP

[743] Anonymous. De Locht’s pantelephone. Scientific American, 43 (11):162, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n11/pdf/scientificamerican09111880-
162b.pdf.

Anonymous:1880:EIDa

[744] Anonymous. Experiment illustrating discharge of electricity from clouds. Scientific American, 43(11):163, September 11, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v43/
n11/pdf/scientificamerican09111880-163.pdf.

Anonymous:1880:SOE

[745] Anonymous. Submarine observatory and electric light. Scientific American, 43(11):163–164, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v43/n11/pdf/scientificamerican09111880-
163a.pdf.

Anonymous:1880:WPP

[746] Anonymous. The Williamsport panel planer. Scientific American, 43(11):164, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n11/pdf/scientificamerican09111880-
164.pdf.

Anonymous:1880:MIbt

[747] Anonymous. Miscellaneous inventions. Scientific American, 43(11): 164–165, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n11/pdf/scientificamerican09111880-
164a.pdf.

Anonymous:1880:AIq

[748] Anonymous. Agricultural inventions. Scientific American, 43 (11):165, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n11/pdf/scientificamerican09111880-
165.pdf.

Anonymous:1880:NAB

[749] Anonymous. Novel air brake. Scientific American, 43(11):166, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n11/pdf/scientificamerican09111880-166.pdf.

Anonymous:1880:NLT

[750] Anonymous. New line-throwing gun. Scientific American, 43 (11):166, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n11/pdf/scientificamerican09111880-
166a.pdf.

Anonymous:1880:MIbu

[751] Anonymous. Mechanical inventions. Scientific American, 43(11): 166–167, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n11/pdf/scientificamerican09111880-
166b.pdf.

Roberts:1880:JCS

[752] A. W. Roberts. The Japanese and Chinese sections of the Berlin International Fisheries Exhibition. Scientific American, 43(11):167, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n11/pdf/scientificamerican09111880-167.pdf.

Anonymous:1880:PRO

[753] Anonymous. Providence river oysters. Scientific American, 43(11): 168–169, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n11/pdf/scientificamerican09111880-
168.pdf.

Anonymous:1880:NHNg

[754] Anonymous. Natural history notes. Scientific American, 43(11): 169, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n11/pdf/scientificamerican09111880-
169.pdf.

Anonymous:1880:EIp

[755] Anonymous. Engineering inventions. Scientific American, 43(11): 169, September 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n11/pdf/scientificamerican09111880-
169a.pdf.

Anonymous:1880:DRPb

[756] Anonymous. Decisions relating to patents, trade marks, etc. Scientific American, 43(11):170–171, September 11, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v43/
n11/pdf/scientificamerican09111880-170.pdf.

Anonymous:1880:INIk

[757] Anonymous. Inventions: New inventions. Scientific American, 43 (12):179, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
179a.pdf.

Anonymous:1880:IRIc

[758] Anonymous. Inventions: Recent inventions. Scientific American, 43(12):181, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
181.pdf.

Anonymous:1880:BPbf

[759] Anonymous. Business and personal. Scientific American, 43 (12):187, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
187.pdf.

Anonymous:1880:NQbb

[760] Anonymous. Notes and queries. Scientific American, 43(12): 187–188, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
187a.pdf.

Anonymous:1880:IIbe

[761] Anonymous. Index of inventions. Scientific American, 43(12): 188, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
188a.pdf.

Anonymous:1880:PNPs

[762] Anonymous. Patent notes: Patents. Scientific American, 43 (12):189, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
189.pdf.

Anonymous:1880:MHC

[763] Anonymous. Modern hill climbing. Scientific American, 43(12): 175, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
175.pdf.

Anonymous:1880:ASe

[764] Anonymous. Adulterations and substitutions. Scientific American, 43(12):176, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
176.pdf.

Anonymous:1880:Pg

[765] Anonymous. The photophone. Scientific American, 43(12):176, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n12/pdf/scientificamerican09181880-176a.pdf.

Anonymous:1880:ADc

[766] Anonymous. An astronomical discovery. Scientific American, 43 (12):176, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
176b.pdf.

Anonymous:1880:HOT

[767] Anonymous. How are the oil tanks set on fire by the lightning? Scientific American, 43(12):176–177, September 18, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v43/
n12/pdf/scientificamerican09181880-176c.pdf.

Anonymous:1880:HRTb

[768] Anonymous. The Hudson River Tunnel. Scientific American, 43 (12):177, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
177.pdf.

Anonymous:1880:AIr

[769] Anonymous. Agricultural inventions. Scientific American, 43 (12):178, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
178.pdf.

Anonymous:1880:ILC

[770] Anonymous. Improved lace cutter. Scientific American, 43(12): 178, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
178a.pdf.

Anonymous:1880:IHBb

[771] Anonymous. Improved harness buckle. Scientific American, 43 (12):178, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
178b.pdf.

Anonymous:1880:NDW

[772] Anonymous. New dish washer. Scientific American, 43(12): 178, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
178c.pdf.

Anonymous:1880:NRD

[773] Anonymous. A new rock drill. Scientific American, 43(12):179, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n12/pdf/scientificamerican09181880-179.pdf.

Anonymous:1880:MIbv

[774] Anonymous. Mechanical inventions. Scientific American, 43(12): 179–180, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
179b.pdf.

Anonymous:1880:DRPc

[775] Anonymous. Decisions relating to patents. Scientific American, 43 (12):180, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
180.pdf.

Anonymous:1880:MDCb

[776] Anonymous. A modified Daniell cell. Scientific American, 43(12): 181–182, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
181a.pdf.

Anonymous:1880:IVG

[777] Anonymous. Improved valve-gear for engines. Scientific American, 43(12):182, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
182.pdf.

Anonymous:1880:HPH

[778] Anonymous. Halter to prevent horses from cribbing. Scientific American, 43(12):182, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
182a.pdf.

Anonymous:1880:MIbw

[779] Anonymous. Miscellaneous inventions. Scientific American, 43 (12):182, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
182b.pdf.

Anonymous:1880:IRGb

[780] Anonymous. The India-rubber and gutta percha industries. Scientific American, 43(12):183–185, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
183.pdf.

Anonymous:1880:EIq

[781] Anonymous. Engineering inventions. Scientific American, 43(12): 185–186, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
185.pdf.

Weed:1880:CFB

[782] W. H. Weed. The conch fisheries of the Bahamas. Scientific American, 43(12):186, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
186.pdf.

Anonymous:1880:NHNh

[783] Anonymous. Natural history notes. Scientific American, 43(12): 186, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
186a.pdf.

Anonymous:1880:CFA

[784] Anonymous. Curious facts about the albatross. Scientific American, 43(12):186–187, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
186b.pdf.

Anonymous:1880:CRt

[785] Anonymous. Communications received. Scientific American, 43 (12):188, September 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n12/pdf/scientificamerican09181880-
188.pdf.

Anonymous:1880:Cq

[786] Anonymous. Correspondence. Scientific American, 43(13):201, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n13/pdf/scientificamerican09251880-201.pdf.

Anonymous:1880:IRId

[787] Anonymous. Inventions: Recent inventions. Scientific American, 43 (13):201–202, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n13/pdf/scientificamerican09251880-
201b.pdf.

Anonymous:1880:BPbg

[788] Anonymous. Business and personal. Scientific American, 43 (13):203, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n13/pdf/scientificamerican09251880-
203.pdf.

Anonymous:1880:NQbc

[789] Anonymous. Notes and queries. Scientific American, 43(13): 203, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n13/pdf/scientificamerican09251880-
203a.pdf.

Anonymous:1880:IIbf

[790] Anonymous. Index of inventions. Scientific American, 43(13): 203–204, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n13/pdf/scientificamerican09251880-
203b.pdf.

Anonymous:1880:PNPt

[791] Anonymous. Patent notes: Patents. Scientific American, 43 (13):205, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n13/pdf/scientificamerican09251880-
205.pdf.

Anonymous:1880:NCe

[792] Anonymous. New cab. Scientific American, 43(13):191, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n13/pdf/scientificamerican09251880-191.pdf.

Anonymous:1880:CHF

[793] Anonymous. Could a hostile fleet bombard New York? Scientific American, 43(13):192, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n13/pdf/scientificamerican09251880-
192.pdf.

Anonymous:1880:STS

[794] Anonymous. Science teaching in schools. Scientific American, 43 (13):192–193, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n13/pdf/scientificamerican09251880-
192a.pdf.

Anonymous:1880:CIb

[795] Anonymous. A chance for inventors. — the. Scientific American, 43(13):193, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n13/pdf/scientificamerican09251880-
193.pdf.

Anonymous:1880:BPbh

[796] Anonymous. The De Bay propeller. Scientific American, 43 (13):193, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n13/pdf/scientificamerican09251880-
193a.pdf.

Anonymous:1880:TF

[797] Anonymous. Traveling flies. Scientific American, 43(13):193, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n13/pdf/scientificamerican09251880-193b.pdf.

Anonymous:1880:FPP

[798] Anonymous. French pill printing machine. Scientific American, 43 (13):194, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n13/pdf/scientificamerican09251880-
194.pdf.

Anonymous:1880:NPO

[799] Anonymous. New portable oven. Scientific American, 43(13): 194, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n13/pdf/scientificamerican09251880-
194a.pdf.

Anonymous:1880:HWN

[800] Anonymous. Hens’ wire nests. Scientific American, 43(13):194– 195, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n13/pdf/scientificamerican09251880-
194b.pdf.

Anonymous:1880:FIS

[801] Anonymous. The first inventor of the steam gauge. Scientific American, 43(13):195–196, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n13/pdf/scientificamerican09251880-
195.pdf.

Anonymous:1880:NBF

[802] Anonymous. New beer faucet. Scientific American, 43(13):196, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n13/pdf/scientificamerican09251880-196.pdf.

Anonymous:1880:ASAb

[803] Anonymous. The American Science Association. Scientific American, 43(13):196–197, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n13/pdf/scientificamerican09251880-
196a.pdf.

Anonymous:1880:EIr

[804] Anonymous. Engineering inventions. Scientific American, 43(13): 197, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n13/pdf/scientificamerican09251880-
197.pdf.

Anonymous:1880:IBT

[805] Anonymous. Improvement in billiard tables. Scientific American, 43(13):198, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n13/pdf/scientificamerican09251880-
198.pdf.

Anonymous:1880:DTA

[806] Anonymous. Double treadle attachment. Scientific American, 43 (13):198, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n13/pdf/scientificamerican09251880-
198a.pdf.

Anonymous:1880:MIbx

[807] Anonymous. Miscellaneous inventions. Scientific American, 43 (13):198, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n13/pdf/scientificamerican09251880-
198b.pdf.

Roberts:1880:Sa

[808] A. W. Roberts. Seals. Scientific American, 43(13):199–200, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n13/pdf/scientificamerican09251880-199.pdf.

Anonymous:1880:Bb

[809] Anonymous. Benzole. Scientific American, 43(13):200, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n13/pdf/scientificamerican09251880-200.pdf.

Anonymous:1880:MIby

[810] Anonymous. Mechanical inventions. Scientific American, 43(13): 200–201, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n13/pdf/scientificamerican09251880-
200a.pdf.

Anonymous:1880:CLFb

[811] Anonymous. Count Louis François De Pourtales. Scientific American, 43(13):201, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n13/pdf/scientificamerican09251880-
201a.pdf.

Anonymous:1880:RD

[812] Anonymous. Reasonable diligence. Scientific American, 43(13): 202, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n13/pdf/scientificamerican09251880-
202.pdf.

Anonymous:1880:CTM

[813] Anonymous. The causes of terrestrial magnetism. Scientific American, 43(13):202, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n13/pdf/scientificamerican09251880-
202a.pdf.

Anonymous:1880:LLO

[814] Anonymous. The location of the Lick Observatory. Scientific American, 43(13):202, September 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n13/pdf/scientificamerican09251880-
202b.pdf.

Anonymous:1880:INIl

[815] Anonymous. Inventions: New inventions. Scientific American, 43(14):210, October 2, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n14/pdf/scientificamerican10021880-
210.pdf.

Anonymous:1880:BPbi

[816] Anonymous. Business and personal. Scientific American, 43(14):219, October 2, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n14/pdf/scientificamerican10021880-219.pdf.

Anonymous:1880:NQbd

[817] Anonymous. Notes and queries. Scientific American, 43(14):219, October 2, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n14/pdf/scientificamerican10021880-219a.pdf.

Anonymous:1880:Ia

[818] Anonymous. Index. Scientific American, 43(14):219–220, October 2, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n14/pdf/scientificamerican10021880-219c.pdf.

Anonymous:1880:PMO

[819] Anonymous. Potters’ machinery — an opening for inventors. Scientific American, 43(14):208, October 2, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n14/pdf/scientificamerican10021880-
208.pdf.

Anonymous:1880:BSF

[820] Anonymous. Burning of a supposed fireproof building. Scientific American, 43(14):208, October 2, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n14/pdf/scientificamerican10021880-
208a.pdf.

Anonymous:1880:CW

[821] Anonymous. Conscientious workers. Scientific American, 43(14): 208–209, October 2, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n14/pdf/scientificamerican10021880-
208b.pdf.

Anonymous:1880:Ph

[822] Anonymous. The photophone. Scientific American, 43(14):209, October 2, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n14/pdf/scientificamerican10021880-209.pdf.

Anonymous:1880:SSA

[823] Anonymous. The Social Science Association. Scientific American, 43(14):209, October 2, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n14/pdf/scientificamerican10021880-
209a.pdf.

Anonymous:1880:IBBa

[824] Anonymous. Improved blacking brush. Scientific American, 43(14):210, October 2, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n14/pdf/scientificamerican10021880-210a.pdf.

Anonymous:1880:NES

[825] Anonymous. New extension stovepipe. Scientific American, 43(14):210, October 2, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n14/pdf/scientificamerican10021880-210b.pdf.

Anonymous:1880:AINv

[826] Anonymous. American industries. — no. 56. Scientific American, 43(14):211–212, October 2, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n14/pdf/scientificamerican10021880-
211.pdf.

Anonymous:1880:EIs

[827] Anonymous. Engineering inventions. Scientific American, 43(14): 212–213, October 2, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n14/pdf/scientificamerican10021880-
212.pdf.

Baldwin:1880:HYSc

[828] W. M. J. Baldwin. Hints to the young steam fitter. Scientific American, 43(14):213, October 2, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n14/pdf/scientificamerican10021880-
213.pdf.

Anonymous:1880:MIbz

[829] Anonymous. Mechanical inventions. Scientific American, 43(14):213, October 2, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n14/pdf/scientificamerican10021880-213a.pdf.

Anonymous:1880:MIca

[830] Anonymous. Miscellaneous inventions. Scientific American, 43(14):214, October 2, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n14/pdf/scientificamerican10021880-214.pdf.

Anonymous:1880:IRS

[831] Anonymous. Improvement in railway switches and crossings. Scientific American, 43(14):214–215, October 2, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v43/n14/pdf/scientificamerican10021880-
214a.pdf.

Roberts:1880:GW

[832] A. W. Roberts. The granny whale. Scientific American, 43(14):215–216, October 2, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n14/pdf/scientificamerican10021880-215.pdf.

Anonymous:1880:EST

[833] Anonymous. An extra-sensitive thermometer. Scientific American, 43(14):216–217, October 2, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n14/pdf/scientificamerican10021880-
216.pdf.

Anonymous:1880:IWG

[834] Anonymous. Improvement in water gauges. Scientific American, 43(14):217–218, October 2, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n14/pdf/scientificamerican10021880-
217.pdf.

Anonymous:1880:AIt

[835] Anonymous. Agricultural inventions. Scientific American, 43(14):218, October 2, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n14/pdf/scientificamerican10021880-218.pdf.

Anonymous:1880:CRu

[836] Anonymous. Communications received. Scientific American, 43 (14):219, October 2, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n14/pdf/scientificamerican10021880-
219b.pdf.

Anonymous:1880:IRIe 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n15/pdf/scientificamerican10091880-
226a.pdf.

Anonymous:1880:BPbj

[838] Anonymous. Business and personal. Scientific American, 43(15):235, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n15/pdf/scientificamerican10091880-235a.pdf.

Anonymous:1880:BNBu

[839] Anonymous. Books: New books and publications. Scientific American, 43(15):235, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n15/pdf/scientificamerican10091880-
235b.pdf.

Anonymous:1880:NQbe

[840] Anonymous. Notes and queries. Scientific American, 43(15):235, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n15/pdf/scientificamerican10091880-235c.pdf.

Anonymous:1880:IIbg

[841] Anonymous. Index of inventions. Scientific American, 43(15):236, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n15/pdf/scientificamerican10091880-236.pdf.

Anonymous:1880:AND

[842] Anonymous. American Naval defenses. Scientific American, 43(15):224, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n15/pdf/scientificamerican10091880-224.pdf.

Anonymous:1880:CAF

[843] Anonymous. Compressing air by falling water. Scientific American, 43(15):224, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n15/pdf/scientificamerican10091880-
224a.pdf.

Anonymous:1880:EL

[844] Anonymous. Employees’ liabilities. Scientific American, 43(15):224, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n15/pdf/scientificamerican10091880-224b.pdf.

Anonymous:1880:HRTc

[845] Anonymous. The Hudson River Tunnel. Scientific American, 43(15):225, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n15/pdf/scientificamerican10091880-
225.pdf.

Anonymous:1880:LO

[846] Anonymous. The landing of the obelisk. Scientific American, 43(15):225, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n15/pdf/scientificamerican10091880-
225a.pdf.

Anonymous:1880:RIc

[847] Anonymous. Railway improvements. Scientific American, 43(15):226, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n15/pdf/scientificamerican10091880-226.pdf.

Anonymous:1880:NHS

[848] Anonymous. Novel hand sewing tool. Scientific American, 43(15):227, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n15/pdf/scientificamerican10091880-227.pdf.

Anonymous:1880:IV

[849] Anonymous. Improved velocipede. Scientific American, 43(15):227, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n15/pdf/scientificamerican10091880-227a.pdf.

Anonymous:1880:MIcb

[850] Anonymous. Miscellaneous inventions. Scientific American, 43(15):227, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n15/pdf/scientificamerican10091880-227b.pdf.

Anonymous:1880:RMJ

[851] Anonymous. A result of the Mississippi jetties. Scientific American, 43(15):228, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n15/pdf/scientificamerican10091880-
228.pdf.

Anonymous:1880:SYP

[852] Anonymous. Strength of yellow pine. Scientific American, 43(15):228, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n15/pdf/scientificamerican10091880-228a.pdf.

Anonymous:1880:NIc

[853] Anonymous. A novel importation. Scientific American, 43(15):228, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n15/pdf/scientificamerican10091880-228b.pdf.

Anonymous:1880:BAM

[854] Anonymous. The British Association meeting. Scientific American, 43(15):228, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n15/pdf/scientificamerican10091880-
228c.pdf.

Anonymous:1880:GSCb

[855] Anonymous. Great and sudden changes of temperature. Scientific American, 43(15):228, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n15/pdf/scientificamerican10091880-
228d.pdf.

Anonymous:1880:FAI

[856] Anonymous. The Fair of the American Institute. Scientific American, 43(15):228–229, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n15/pdf/scientificamerican10091880-
228e.pdf.

Anonymous:1880:AINw

[857] Anonymous. American industries. — no. 57. Scientific American, 43(15):229, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n15/pdf/scientificamerican10091880-
229.pdf.

Anonymous:1880:DRPd

[858] Anonymous. Decisions relating to patents. Scientific American, 43(15):229, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n15/pdf/scientificamerican10091880-
229a.pdf.

Anonymous:1880:FSE

[859] Anonymous. The Franklin Search Expedition. Scientific American, 43(15):229, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n15/pdf/scientificamerican10091880-
229b.pdf.

Anonymous:1880:ISc

[860] Anonymous. Improved die stock. Scientific American, 43(15):230, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n15/pdf/scientificamerican10091880-230.pdf.

Anonymous:1880:NTb

[861] Anonymous. A novel tent. Scientific American, 43(15):230, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n15/pdf/scientificamerican10091880-230a.pdf.

Anonymous:1880:EIt

[862] Anonymous. Engineering inventions. Scientific American, 43(15): 230–231, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n15/pdf/scientificamerican10091880-
230b.pdf.

Anonymous:1880:CSK

[863] Anonymous. The coral signet of the King of Italy. Scientific American, 43(15):231, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n15/pdf/scientificamerican10091880-
231.pdf.

Anonymous:1880:MBF

[864] Anonymous. The mollusks at the Berlin Fishery Exhibition. Scientific American, 43(15):231–232, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v43/n15/pdf/scientificamerican10091880-
231a.pdf.

Anonymous:1880:NAP

[865] Anonymous. A new acid pump. Scientific American, 43(15):232–233, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n15/pdf/scientificamerican10091880-232.pdf.

Griswold:1880:SWG

[866] M. W. Griswold. Standard wire gauges. Scientific American, 43(15):233, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n15/pdf/scientificamerican10091880-233.pdf.

Anonymous:1880:ISM

[867] Anonymous. An important shoe machine patent case. Scientific American, 43(15):234, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n15/pdf/scientificamerican10091880-
234.pdf.

Anonymous:1880:WR

[868] Anonymous. $200,000 worth of rifles. Scientific American, 43(15):235, October 9, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n15/pdf/scientificamerican10091880-235.pdf.

Anonymous:1880:IRIf

[869] Anonymous. Inventions: Recent inventions. Scientific American, 43(16):242, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n16/pdf/scientificamerican10161880-
242b.pdf.

Anonymous:1880:INIm

[870] Anonymous. Inventions: New inventions. Scientific American, 43 (16):244–245, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n16/pdf/scientificamerican10161880-
244a.pdf.

Anonymous:1880:CCN

[871] Anonymous. Correspondence: Correspondence, a new safety sail boat, and more. Scientific American, 43(16):245, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n16/pdf/scientificamerican10161880-245.pdf.

Anonymous:1880:BPbk

[872] Anonymous. Business and personal. Scientific American, 43 (16):251, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n16/pdf/scientificamerican10161880-
251.pdf.

Anonymous:1880:NQbf

[873] Anonymous. Notes and queries. Scientific American, 43(16):251, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n16/pdf/scientificamerican10161880-251a.pdf.

Anonymous:1880:IIbh

[874] Anonymous. Index of inventions. Scientific American, 43(16): 251–252, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n16/pdf/scientificamerican10161880-
251c.pdf.

Anonymous:1880:AINx

[875] Anonymous. American industries — no. 58. Scientific American, 43(16):239, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n16/pdf/scientificamerican10161880-
239.pdf.

Anonymous:1880:OMA

[876] Anonymous. Organic matter in the air. Scientific American, 43 (16):240, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n16/pdf/scientificamerican10161880-
240.pdf.

Anonymous:1880:PSW

[877] Anonymous. The Philadelphia Sheep and Wool Show. Scientific American, 43(16):240, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n16/pdf/scientificamerican10161880-
240a.pdf.

Anonymous:1880:OMS

[878] Anonymous. Origin of the Merino sheep. Scientific American, 43 (16):240–241, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n16/pdf/scientificamerican10161880-
240b.pdf.

Anonymous:1880:PCS

[879] Anonymous. Progress of cotton seed oil manufacture, progress of the brush electric light, and more. Scientific American, 43(16):241, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n16/pdf/scientificamerican10161880-241.pdf.

Anonymous:1880:U

[880] Anonymous. The unicorn. Scientific American, 43(16):241, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n16/pdf/scientificamerican10161880-241a.pdf.

Riley:1880:IPC

[881] C. V. Riley. Insecticides for the protection of cotton, London purple, and more. Scientific American, 43(16):241–242, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n16/pdf/scientificamerican10161880-241b.pdf.

Anonymous:1880:MNF

[882] Anonymous. Maxim’s new focusing electric lamp. Scientific American, 43(16):242, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n16/pdf/scientificamerican10161880-
242.pdf.

Anonymous:1880:EIu

[883] Anonymous. Engineer and inventor. Scientific American, 43 (16):242, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n16/pdf/scientificamerican10161880-
242a.pdf.

Anonymous:1880:NSPb

[884] Anonymous. New sheep protector, spontaneous combustion of charcoal, and more. Scientific American, 43(16):243, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n16/pdf/scientificamerican10161880-243.pdf.

Anonymous:1880:MOM

[885] Anonymous. Machine for ornamenting metal surfaces. Scientific American, 43(16):243, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n16/pdf/scientificamerican10161880-
243a.pdf.

Anonymous:1880:EIv

[886] Anonymous. Engineering inventions. Scientific American, 43(16): 243, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n16/pdf/scientificamerican10161880-
243b.pdf.

Anonymous:1880:AIu

[887] Anonymous. American industries. Scientific American, 43(16):244, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n16/pdf/scientificamerican10161880-244.pdf.

Anonymous:1880:AIBa

[888] Anonymous. Agricultural inventions, big farms on the Pacific Coast, and more. Scientific American, 43(16):245, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n16/pdf/scientificamerican10161880-245a.pdf.

Anonymous:1880:NRW

[889] Anonymous. A new road wagon. Scientific American, 43(16):246, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n16/pdf/scientificamerican10161880-246.pdf.

Anonymous:1880:NSH

[890] Anonymous. New sash holder and fastener. Scientific American, 43(16):246, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n16/pdf/scientificamerican10161880-
246a.pdf.

Anonymous:1880:MIcc

[891] Anonymous. Mechanical inventions. Scientific American, 43(16): 246, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n16/pdf/scientificamerican10161880-
246b.pdf.

Anonymous:1880:NCBa

[892] Anonymous. A new combined barometer and thermometer. Scientific American, 43(16):246, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v43/n16/pdf/scientificamerican10161880-
246c.pdf.

Roberts:1880:MC

[893] A. W. Roberts. A marine community. Scientific American, 43 (16):247–248, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n16/pdf/scientificamerican10161880-
247.pdf.

Anonymous:1880:KPM

[894] Anonymous. The king penguin, the Menominee iron mines, and more. Scientific American, 43(16):248–249, October 16, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v43/
n16/pdf/scientificamerican10161880-248.pdf.

Anonymous:1880:SID

[895] Anonymous. On some impurities of drinking waters, the model workman, and more. Scientific American, 43(16):250, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n16/pdf/scientificamerican10161880-250.pdf.

Anonymous:1880:CRv

[896] Anonymous. Communications received. Scientific American, 43 (16):251, October 16, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n16/pdf/scientificamerican10161880-
251b.pdf.

Anonymous:1880:CCF

[897] Anonymous. Correspondence: Correspondence, fire apparatus in cities, and more. Scientific American, 43(17):265, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n17/pdf/scientificamerican10231880-265.pdf.

Anonymous:1880:IRIg

[898] Anonymous. Inventions: Recent inventions. Scientific American, 43(17):265, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n17/pdf/scientificamerican10231880-
265a.pdf.

Anonymous:1880:BPbl

[899] Anonymous. Business and personal. Scientific American, 43 (17):267, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n17/pdf/scientificamerican10231880-
267.pdf.

Anonymous:1880:BNBv

[900] Anonymous. Books: New books and publications. Scientific American, 43(17):267, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n17/pdf/scientificamerican10231880-
267a.pdf.

Anonymous:1880:NQbg

[901] Anonymous. Notes and queries. Scientific American, 43(17):267, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n17/pdf/scientificamerican10231880-267b.pdf.

Anonymous:1880:IIbi

[902] Anonymous. Index of inventions. Scientific American, 43(17): 267–268, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n17/pdf/scientificamerican10231880-
267d.pdf.

Anonymous:1880:PNPu

[903] Anonymous. Patent notes: Patents. Scientific American, 43 (17):269, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n17/pdf/scientificamerican10231880-
269.pdf.

Anonymous:1880:RDE

[904] Anonymous. Recent developments in electric lighting. Scientific American, 43(17):255–256, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n17/pdf/scientificamerican10231880-
255.pdf.

Anonymous:1880:RCP

[905] Anonymous. The relation of cheap patents to industry prosperity. Scientific American, 43(17):256, October 23, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v43/
n17/pdf/scientificamerican10231880-256.pdf.

Anonymous:1880:MFN

[906] Anonymous. “Malarial” fever in New England. Scientific American, 43(17):256–257, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n17/pdf/scientificamerican10231880-
256a.pdf.

Anonymous:1880:RAC

[907] Anonymous. The revival of American commerce. Scientific American, 43(17):257, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n17/pdf/scientificamerican10231880-
257.pdf.

Anonymous:1880:BPbm

[908] Anonymous. Benjamin Peirce. Scientific American, 43(17):257, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n17/pdf/scientificamerican10231880-257a.pdf.

Anonymous:1880:FF

[909] Anonymous. Firepro of ferryboats. Scientific American, 43(17):257, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n17/pdf/scientificamerican10231880-257b.pdf.

Anonymous:1880:DPC

[910] Anonymous. Designs produced by crystallization. Scientific American, 43(17):257, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n17/pdf/scientificamerican10231880-
257c.pdf.

Baldwin:1880:HYSd

[911] William J. Baldwin. Hints to the young steam fitter. Scientific American, 43(17):258–259, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n17/pdf/scientificamerican10231880-
258.pdf.

Hopkins:1880:ACR

[912] George M. Hopkins. Apparatus for combining rectangular vibrations. Scientific American, 43(17):259, October 23, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v43/
n17/pdf/scientificamerican10231880-259.pdf.

Anonymous:1880:NGM

[913] Anonymous. New grain meter. Scientific American, 43(17):259–260, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n17/pdf/scientificamerican10231880-259a.pdf.

Anonymous:1880:WSD

[914] Anonymous. The wool sorter’s disease. Scientific American, 43 (17):261, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n17/pdf/scientificamerican10231880-
261.pdf.

Anonymous:1880:NFAa

[915] Anonymous. A new fire apparatus. Scientific American, 43(17):261, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n17/pdf/scientificamerican10231880-261a.pdf.

Anonymous:1880:FMS

[916] Anonymous. Fall of a meteoric stone. Scientific American, 43 (17):261, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n17/pdf/scientificamerican10231880-
261b.pdf.

Anonymous:1880:DRPe

[917] Anonymous. Decisions relating to patents. Scientific American, 43(17):261, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n17/pdf/scientificamerican10231880-
261c.pdf.

Anonymous:1880:MIcd

[918] Anonymous. Miscellaneous inventions. Scientific American, 43 (17):262–263, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n17/pdf/scientificamerican10231880-
262.pdf.

Anonymous:1880:LBP

[919] Anonymous. Long-billed parrot and Banksian cockatoo. Scientific American, 43(17):263–264, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v43/n17/pdf/scientificamerican10231880-
263.pdf.

Roberts:1880:BR

[920] A. W. Roberts. Batoidei, or rays. Scientific American, 43(17): 264–265, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n17/pdf/scientificamerican10231880-
264.pdf.

Anonymous:1880:TPI

[921] Anonymous. Tin plate industry — the process of manufacture. Scientific American, 43(17):266, October 23, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v43/
n17/pdf/scientificamerican10231880-266.pdf.

Anonymous:1880:MIce

[922] Anonymous. Mechanical inventions. Scientific American, 43(17): 266, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n17/pdf/scientificamerican10231880-
266a.pdf.

Anonymous:1880:AR

[923] Anonymous. American razors. Scientific American, 43(17):266, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n17/pdf/scientificamerican10231880-266b.pdf.

Anonymous:1880:CDa

[924] Anonymous. Cremation of the dead. Scientific American, 43 (17):266–267, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n17/pdf/scientificamerican10231880-
266c.pdf.

Anonymous:1880:CRw

[925] Anonymous. Communications received. Scientific American, 43 (17):267, October 23, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n17/pdf/scientificamerican10231880-
267c.pdf.

Anonymous:1880:INIn

[926] Anonymous. Inventions: New inventions. Scientific American, 43 (18):274–275, October 30, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n18/pdf/scientificamerican10301880-
274b.pdf.

Anonymous:1880:BPbn

[927] Anonymous. Business and personal. Scientific American, 43 (18):283, October 30, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n18/pdf/scientificamerican10301880-
283.pdf.

Anonymous:1880:NQbh

[928] Anonymous. Notes and queries. Scientific American, 43(18):283, October 30, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n18/pdf/scientificamerican10301880-283a.pdf.

Anonymous:1880:IIbj

[929] Anonymous. Index of inventions. Scientific American, 43(18):283, October 30, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n18/pdf/scientificamerican10301880-283b.pdf.

Anonymous:1880:LAE

[930] Anonymous. Leveling attachment for earth cars. Scientific American, 43(18):271, October 30, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n18/pdf/scientificamerican10301880-
271.pdf.

Hopkins:1880:NRG

[931] M. Hopkins. New recording galvanometer. Scientific American, 43 (18):271–272, October 30, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n18/pdf/scientificamerican10301880-
271a.pdf.

Anonymous:1880:TB

[932] Anonymous. The telephone on the battlefield. Scientific American, 43(18):272, October 30, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n18/pdf/scientificamerican10301880-
272.pdf.

Anonymous:1880:WWB

[933] Anonymous. What we buy and sell abroad. Scientific American, 43(18):272–273, October 30, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n18/pdf/scientificamerican10301880-
272a.pdf.

Burrill:1880:BPT

[934] T. J. Burrill. Blight of pear trees. Scientific American, 43(18):273, October 30, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n18/pdf/scientificamerican10301880-273.pdf.

Anonymous:1880:DOP

[935] Anonymous. Destruction of oysters by petroleum. Scientific American, 43(18):273–274, October 30, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n18/pdf/scientificamerican10301880-
273a.pdf.

Anonymous:1880:TWW

[936] Anonymous. Toothed-wheel wire fence. Scientific American, 43 (18):274, October 30, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n18/pdf/scientificamerican10301880-
274.pdf.

Anonymous:1880:NHG

[937] Anonymous. New hammerless gun. Scientific American, 43(18):274, October 30, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n18/pdf/scientificamerican10301880-274a.pdf.

Anonymous:1880:NCCc

[938] Anonymous. New chimney cap and ventilator. Scientific American, 43(18):275, October 30, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n18/pdf/scientificamerican10301880-
275.pdf.

Anonymous:1880:INC

[939] Anonymous. Improved nursery chair. Scientific American, 43 (18):275–276, October 30, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n18/pdf/scientificamerican10301880-
275a.pdf.

Anonymous:1880:FTP

[940] Anonymous. The fan-tailed poodle. Scientific American, 43(18): 276, October 30, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n18/pdf/scientificamerican10301880-
276.pdf.

Anonymous:1880:AIv

[941] Anonymous. Agricultural inventions. Scientific American, 43(18): 276–277, October 30, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n18/pdf/scientificamerican10301880-
276a.pdf.

Anonymous:1880:MIcf

[942] Anonymous. Mechanical inventions. Scientific American, 43(18): 277–278, October 30, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n18/pdf/scientificamerican10301880-
277.pdf.

Anonymous:1880:ADEb

[943] Anonymous. Apparatus for determining the electric charges of falling rain. Scientific American, 43(18):278, October 30, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v43/
n18/pdf/scientificamerican10301880-278.pdf.

Anonymous:1880:DSM

[944] Anonymous. Davey’s simplex motor. Scientific American, 43(18): 278–279, October 30, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n18/pdf/scientificamerican10301880-
278a.pdf.

Anonymous:1880:NTR

[945] Anonymous. Note on turpentine, rosin, and allied products. Scientific American, 43(18):279, October 30, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v43/n18/pdf/scientificamerican10301880-
279.pdf.

Anonymous:1880:DTS

[946] Anonymous. The Dowd tunneling system. Scientific American, 43 (18):279–280, October 30, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n18/pdf/scientificamerican10301880-
279a.pdf.

Anonymous:1880:HER

[947] Anonymous. History of elevated railways. Scientific American, 43(18):281, October 30, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n18/pdf/scientificamerican10301880-
281.pdf.

Anonymous:1880:HUS

[948] Anonymous. Homicide in the United States. Scientific American, 43(18):282, October 30, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n18/pdf/scientificamerican10301880-
282.pdf.

Anonymous:1880:INIo

[949] Anonymous. Inventions: New inventions. Scientific American, 43 (19):294, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
294.pdf.

Anonymous:1880:BPbo

[950] Anonymous. Business and personal. Scientific American, 43 (19):299, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
299.pdf.

Anonymous:1880:NQbi

[951] Anonymous. Notes and queries. Scientific American, 43(19): 299, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
299a.pdf.

Anonymous:1880:IIbk

[952] Anonymous. Index of inventions. Scientific American, 43(19): 299–300, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
299c.pdf.

Anonymous:1880:LC

[953] Anonymous. The laws of cyclones. Scientific American, 43(19): 288, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
288.pdf.

Anonymous:1880:CSb

[954] Anonymous. The cultivation of the sumac. Scientific American, 43(19):288, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
288a.pdf.

Anonymous:1880:COa

[955] Anonymous. The color of ozone. Scientific American, 43(19): 288, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
288b.pdf.

Anonymous:1880:US

[956] Anonymous. Untimely snows. Scientific American, 43(19):288– 289, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
288c.pdf.

Anonymous:1880:CCCa

[957] Anonymous. The Cape Cod Canal. Scientific American, 43 (19):289, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
289.pdf.

Anonymous:1880:ISd

[958] Anonymous. Incendiary silks. Scientific American, 43(19):289, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n19/pdf/scientificamerican11061880-289a.pdf.

Anonymous:1880:AINy

[959] Anonymous. American industries — no. 69. Scientific American, 43 (19):290–291, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
290.pdf.

Anonymous:1880:MSb

[960] Anonymous. The magnetic separator. Scientific American, 43 (19):291, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
291.pdf.

Anonymous:1880:NBCc

[961] Anonymous. A new boiler cleaner. Scientific American, 43(19): 291, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
291a.pdf.

Anonymous:1880:MIcg

[962] Anonymous. Mechanical inventions. Scientific American, 43(19): 292, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
292.pdf.

Anonymous:1880:NPSc

[963] Anonymous. New protective suit for firemen. Scientific American, 43(19):292–293, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
292a.pdf.

Anonymous:1880:FAb

[964] Anonymous. Fishing on the Amazon. Scientific American, 43 (19):293, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
293.pdf.

Anonymous:1880:CHP

[965] Anonymous. Combined horse power and stable floor. Scientific American, 43(19):294, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
294a.pdf.

Anonymous:1880:IMC

[966] Anonymous. Improved milk can. Scientific American, 43(19): 294–295, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
294b.pdf.

Anonymous:1880:ESc

[967] Anonymous. The elephant seal. Scientific American, 43(19): 295, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
295.pdf.

Anonymous:1880:FBT

[968] Anonymous. Fight between a thrasher, swordfish, and a whale. Scientific American, 43(19):295, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
295a.pdf.

Anonymous:1880:PV

[969] Anonymous. A promising vine. Scientific American, 43(19):295, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n19/pdf/scientificamerican11061880-295b.pdf.

Anonymous:1880:NHNi

[970] Anonymous. Natural history notes. Scientific American, 43(19): 295–296, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
295c.pdf.

Anonymous:1880:MDPb

[971] Anonymous. Malignant diseases of plants. Scientific American, 43(19):296, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
296.pdf.

Anonymous:1880:MIch

[972] Anonymous. Miscellaneous inventions. Scientific American, 43 (19):297, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
297.pdf.

Anonymous:1880:CIE

[973] Anonymous. The Cincinnati Industrial Exposition. Scientific American, 43(19):298, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
298.pdf.

Anonymous:1880:AIFa

[974] Anonymous. The American Institute Fair. Scientific American, 43(19):298, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
298a.pdf.

Anonymous:1880:EVE

[975] Anonymous. Exportation of vulcanite emery wheels. Scientific American, 43(19):298, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
298b.pdf.

Anonymous:1880:DRPf

[976] Anonymous. Decisions relating to patents. Scientific American, 43(19):298, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
298c.pdf.

Anonymous:1880:CRx

[977] Anonymous. Communication received. Scientific American, 43 (19):299, November 6, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n19/pdf/scientificamerican11061880-
299b.pdf.

Anonymous:1880:Cu

[978] Anonymous. Correspondence. Scientific American, 43(2):21, July 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n2/pdf/scientificamerican07101880-21a.pdf.

Anonymous:1880:BPbp

[979] Anonymous. Business and personal. Scientific American, 43(2):27, July 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n2/pdf/scientificamerican07101880-27.pdf.

Anonymous:1880:NQbj

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n2/pdf/scientificamerican07101880-27a.pdf.

Anonymous:1880:Id

[981] Anonymous. Index. Scientific American, 43(2):27–28, July 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n2/pdf/scientificamerican07101880-27b.pdf.

Anonymous:1880:MPL

[982] Anonymous. The Millers and the Patent Laws. Scientific American, 43(2):16, July 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n2/pdf/scientificamerican07101880-
16.pdf.

Anonymous:1880:TRT

[983] Anonymous. “There’s room at the top”. Scientific American, 43(2):16–17, July 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n2/pdf/scientificamerican07101880-
16a.pdf.

Anonymous:1880:PPb

[984] Anonymous. Patent pants. Scientific American, 43(2):17, July 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n2/pdf/scientificamerican07101880-17.pdf.

Anonymous:1880:RWS

[985] Anonymous. Are river waters safe. Scientific American, 43(2):17, July 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n2/pdf/scientificamerican07101880-17a.pdf.

Anonymous:1880:EPC

[986] Anonymous. Electrical pavements for city locomotion, skin grafting from the dead and more. Scientific American, 43(2):17, July 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n2/pdf/scientificamerican07101880-17b.pdf.

Anonymous:1880:MIci

[987] Anonymous. Miscellaneous inventions. Scientific American, 43(2):18, July 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n2/pdf/scientificamerican07101880-18.pdf.

Anonymous:1880:EBR

[988] Anonymous. Electric brake for railway cars. Scientific American, 43(2):18, July 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n2/pdf/scientificamerican07101880-
18a.pdf.

Anonymous:1880:IHc

[989] Anonymous. Improvement in horseshoes. Scientific American, 43(2):18, July 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n2/pdf/scientificamerican07101880-18b.pdf.

Anonymous:1880:PCc

[990] Anonymous. Pneumatic clocks. Scientific American, 43(2):19, July 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n2/pdf/scientificamerican07101880-19.pdf.

Anonymous:1880:EIw

[991] Anonymous. Engineering inventions. Scientific American, 43(2):19–20, July 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n2/pdf/scientificamerican07101880-19a.pdf.

Anonymous:1880:AINz

[992] Anonymous. American industries — no. 48. Scientific American, 43(2):20–21, July 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n2/pdf/scientificamerican07101880-
20.pdf.

Anonymous:1880:RDRg

[993] Anonymous. Recent decisions relating to patents. Scientific American, 43(2):21, July 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n2/pdf/scientificamerican07101880-
21.pdf.

Anonymous:1880:IFb

[994] Anonymous. Improved fishway. Scientific American, 43(2):22, July 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n2/pdf/scientificamerican07101880-22.pdf.

Anonymous:1880:NCS

[995] Anonymous. Novel corn sled. Scientific American, 43(2):22–23, July 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n2/pdf/scientificamerican07101880-22a.pdf.

Anonymous:1880:SOK

[996] Anonymous. Sea otter, or kalan. Scientific American, 43(2):23, July 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n2/pdf/scientificamerican07101880-23.pdf.

Anonymous:1880:NSPc

[997] Anonymous. Nemertes, or striped polia. Scientific American, 43(2):23, July 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n2/pdf/scientificamerican07101880-23a.pdf.

Anonymous:1880:NHNj

[998] Anonymous. Natural history notes. Scientific American, 43(2):23–24, July 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n2/pdf/scientificamerican07101880-23b.pdf.

Anonymous:1880:IE

[999] Anonymous. Improved elevator. Scientific American, 43(2):24, July 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n2/pdf/scientificamerican07101880-24.pdf.

Anonymous:1880:SMa

[1000] Anonymous. The structure of matter. Scientific American, 43(2):25, July 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n2/pdf/scientificamerican07101880-25.pdf.

Anonymous:1880:NId

[1001] Anonymous. New inventions. Scientific American, 43(2):25–26, July 10, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n2/pdf/scientificamerican07101880-25a.pdf.

Anonymous:1880:IRIh

[1002] Anonymous. Inventions: Recent inventions. Scientific American, 43 (20):310–311, November 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n20/pdf/scientificamerican11131880-
310c.pdf.

Anonymous:1880:BPbq

[1003] Anonymous. Business and personal. Scientific American, 43 (20):314, November 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n20/pdf/scientificamerican11131880-
314.pdf.

Anonymous:1880:BNBw

[1004] Anonymous. Books: New books and publications. Scientific American, 43(20):314, November 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n20/pdf/scientificamerican11131880-
314a.pdf.

Anonymous:1880:NQbk

[1005] Anonymous. Notes and queries. Scientific American, 43(20): 314–315, November 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n20/pdf/scientificamerican11131880-
314b.pdf.

Anonymous:1880:IIbl

[1006] Anonymous. Index of inventions. Scientific American, 43(20): 315, November 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n20/pdf/scientificamerican11131880-
315a.pdf.

Anonymous:1880:PTa

[1007] Anonymous. The prospects of trade. Scientific American, 43 (20):304, November 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n20/pdf/scientificamerican11131880-
304.pdf.

Anonymous:1880:PPO

[1008] Anonymous. Proposed palm oil industry. Scientific American, 43 (20):304, November 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n20/pdf/scientificamerican11131880-
304a.pdf.

Anonymous:1880:EFU

[1009] Anonymous. The effect of forests upon rainfall. Scientific American, 43(20):304, November 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n20/pdf/scientificamerican11131880-
304b.pdf.

Anonymous:1880:BTM

[1010] Anonymous. The Brunton tunneling machine. Scientific American, 43(20):304–305, November 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n20/pdf/scientificamerican11131880-
304c.pdf.

Anonymous:1880:ICb

[1011] Anonymous. Infected cigars. Scientific American, 43(20):305, November 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n20/pdf/scientificamerican11131880-305.pdf.

Anonymous:1880:PTc

[1012] Anonymous. Police telephones. Scientific American, 43(20):305– 307, November 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n20/pdf/scientificamerican11131880-
305a.pdf.

Anonymous:1880:NNB

[1013] Anonymous. New nursing bottle. Scientific American, 43(20): 307, November 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n20/pdf/scientificamerican11131880-
307.pdf.

Anonymous:1880:AAB

[1014] Anonymous. Apparatus for adjusting balance wheels of watches. Scientific American, 43(20):307, November 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v43/n20/pdf/scientificamerican11131880-
307a.pdf.

Anonymous:1880:MIcj

[1015] Anonymous. Mechanical inventions. Scientific American, 43(20): 307–308, November 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n20/pdf/scientificamerican11131880-
307b.pdf.

Anonymous:1880:LPO

[1016] Anonymous. Lightning-proof oil tank. Scientific American, 43 (20):308, November 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n20/pdf/scientificamerican11131880-
308.pdf.

Anonymous:1880:MESa

[1017] Anonymous. Mr. Eads’ ship railway for the American Isthmus. Scientific American, 43(20):308–309, November 13, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v43/
n20/pdf/scientificamerican11131880-308a.pdf.

Anonymous:1880:MIck

[1018] Anonymous. Miscellaneous inventions. Scientific American, 43(20): 309–310, November 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n20/pdf/scientificamerican11131880-
309.pdf.

Anonymous:1880:IBS

[1019] Anonymous. Improved bottle stopper. Scientific American, 43 (20):310, November 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n20/pdf/scientificamerican11131880-
310.pdf.

Anonymous:1880:IBD

[1020] Anonymous. Improved bolt for double doors. Scientific American, 43(20):310, November 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n20/pdf/scientificamerican11131880-
310a.pdf.

Anonymous:1880:DSS

[1021] Anonymous. Deep sea-sounding apparatus. Scientific American, 43(20):310, November 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n20/pdf/scientificamerican11131880-
310b.pdf.

Anonymous:1880:HFb

[1022] Anonymous. The hunting falcon. Scientific American, 43(20): 311–312, November 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n20/pdf/scientificamerican11131880-
311.pdf.

Anonymous:1880:NR

[1023] Anonymous. The nutrition of roots. Scientific American, 43 (20):312, November 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n20/pdf/scientificamerican11131880-
312.pdf.

Anonymous:1880:IGBa

[1024] Anonymous. Improved gas burner. Scientific American, 43(20): 312, November 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n20/pdf/scientificamerican11131880-
312a.pdf.

Anonymous:1880:DRPg

[1025] Anonymous. Decisions relating to patents and trade marks, by the Commissioner of Patents, and more. Scientific American, 43 (20):313, November 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n20/pdf/scientificamerican11131880-
313.pdf.

Anonymous:1880:CRy

[1026] Anonymous. Communications received. Scientific American, 43 (20):315, November 13, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n20/pdf/scientificamerican11131880-
315.pdf.

Anonymous:1880:INIp

[1027] Anonymous. Inventions: New inventions. Scientific American, 43 (21):329–330, November 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n21/pdf/scientificamerican11201880-
329.pdf.

Anonymous:1880:BPbr

[1028] Anonymous. Business and personal. Scientific American, 43 (21):331, November 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n21/pdf/scientificamerican11201880-
331.pdf.

Anonymous:1880:NQbl

[1029] Anonymous. Notes and queries. Scientific American, 43(21): 331, November 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n21/pdf/scientificamerican11201880-
331a.pdf.

Anonymous:1880:IIbm

[1030] Anonymous. Index of inventions. Scientific American, 43(21): 331–332, November 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n21/pdf/scientificamerican11201880-
331b.pdf.

Anonymous:1880:AME

[1031] Anonymous. American mechanical engineers. Scientific American, 43(21):320, November 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n21/pdf/scientificamerican11201880-
320.pdf.

Anonymous:1880:EDA

[1032] Anonymous. Electric discharges from animals and men. Scientific American, 43(21):320, November 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v43/n21/pdf/scientificamerican11201880-
320a.pdf.

Anonymous:1880:NSAb

[1033] Anonymous. A new system of apprenticeship. Scientific American, 43(21):320, November 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n21/pdf/scientificamerican11201880-
320b.pdf.

Thurston:1880:ES

[1034] R. H. Thurston. The expansion of steam. Scientific American, 43(21):321, November 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n21/pdf/scientificamerican11201880-
321.pdf.

Anonymous:1880:LF

[1035] Anonymous. London fogs. Scientific American, 43(21):321, November 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n21/pdf/scientificamerican11201880-321a.pdf.

Anonymous:1880:BCF

[1036] Anonymous. The Butler colliery fire. Scientific American, 43 (21):321, November 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n21/pdf/scientificamerican11201880-
321b.pdf.

Anonymous:1880:AINba

[1037] Anonymous. American industries — no. 60. Scientific American, 43 (21):322–325, November 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n21/pdf/scientificamerican11201880-
322.pdf.

Anonymous:1880:SMb

[1038] Anonymous. The step of man. Scientific American, 43(21):325, November 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n21/pdf/scientificamerican11201880-325.pdf.

Anonymous:1880:CTAa

[1039] Anonymous. Center turning attachment for lathes. Scientific American, 43(21):326, November 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n21/pdf/scientificamerican11201880-
326.pdf.

Anonymous:1880:EDC

[1040] Anonymous. Ericsson’s duplex caloric pumping engine. Scientific American, 43(21):326–327, November 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v43/n21/pdf/scientificamerican11201880-
326a.pdf.

Anonymous:1880:GE

[1041] Anonymous. The golden eagle. Scientific American, 43(21):327– 328, November 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n21/pdf/scientificamerican11201880-
327.pdf.

Anonymous:1880:NPCb

[1042] Anonymous. Natural ponds for the cultivation of carp for private use. Scientific American, 43(21):328–329, November 20, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v43/
n21/pdf/scientificamerican11201880-328.pdf.

Anonymous:1880:AIw

[1043] Anonymous. Agricultural inventions. Scientific American, 43 (21):330, November 20, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n21/pdf/scientificamerican11201880-
330.pdf.

Anonymous:1880:BPbs

[1044] Anonymous. Business and personal. Scientific American, 43 (22):346, November 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n22/pdf/scientificamerican11271880-
346.pdf.

Anonymous:1880:NQbm

[1045] Anonymous. Notes and queries. Scientific American, 43(22): 346–347, November 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n22/pdf/scientificamerican11271880-
346a.pdf.

Anonymous:1880:IIbn

[1046] Anonymous. Index of inventions. Scientific American, 43(22): 347–350, November 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n22/pdf/scientificamerican11271880-
347.pdf.

Anonymous:1880:HVP

[1047] Anonymous. A huge vacuum pan. Scientific American, 43(22): 335–336, November 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n22/pdf/scientificamerican11271880-
335.pdf.

Anonymous:1880:ELI

[1048] Anonymous. Electric lighting by incandescence. Scientific American, 43(22):336, November 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n22/pdf/scientificamerican11271880-
336.pdf.

Anonymous:1880:PPL

[1049] Anonymous. Progress of patent law. Scientific American, 43(22): 336–337, November 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n22/pdf/scientificamerican11271880-
336a.pdf.

Anonymous:1880:Lb

[1050] Anonymous. The Livadia. Scientific American, 43(22):337, November 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n22/pdf/scientificamerican11271880-337.pdf.

Anonymous:1880:CES

[1051] Anonymous. Capt. Eads’ ship railway. Scientific American, 43 (22):337, November 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n22/pdf/scientificamerican11271880-
337a.pdf.

Anonymous:1880:NSAc

[1052] Anonymous. New stove attachment. Scientific American, 43 (22):338, November 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n22/pdf/scientificamerican11271880-
338.pdf.

Anonymous:1880:ICP

[1053] Anonymous. Improved cattle pen. Scientific American, 43(22): 338–339, November 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n22/pdf/scientificamerican11271880-
338a.pdf.

Anonymous:1880:NWS

[1054] Anonymous. Novel window screen. Scientific American, 43(22): 339, November 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n22/pdf/scientificamerican11271880-
339.pdf.

Anonymous:1880:NBb

[1055] Anonymous. A novel blind. Scientific American, 43(22):339, November 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n22/pdf/scientificamerican11271880-339a.pdf.

Anonymous:1880:ICc

[1056] Anonymous. An improved churn. Scientific American, 43(22): 339, November 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n22/pdf/scientificamerican11271880-
339b.pdf.

Anonymous:1880:MIcl

[1057] Anonymous. Miscellaneous inventions. Scientific American, 43(22): 339–340, November 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n22/pdf/scientificamerican11271880-
339c.pdf.

Anonymous:1880:FC

[1058] Anonymous. Filtering cisterns. Scientific American, 43(22):340, November 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n22/pdf/scientificamerican11271880-340.pdf.

Hovey:1880:HC

[1059] H. C. Hovey. Hydraulic cement. Scientific American, 43(22): 340–341, November 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n22/pdf/scientificamerican11271880-
340a.pdf.

Anonymous:1880:DRPh

[1060] Anonymous. Decisions relating to patents. Scientific American, 43 (22):341–342, November 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n22/pdf/scientificamerican11271880-
341.pdf.

Anonymous:1880:IT

[1061] Anonymous. An improved telephone. Scientific American, 43 (22):342, November 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n22/pdf/scientificamerican11271880-
342.pdf.

Anonymous:1880:NMPb

[1062] Anonymous. Novel method of precipitating rainfalls. Scientific American, 43(22):342, November 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n22/pdf/scientificamerican11271880-
342a.pdf.

Anonymous:1880:MIcm

[1063] Anonymous. Mechanical inventions. Scientific American, 43(22): 342, November 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n22/pdf/scientificamerican11271880-
342b.pdf.

Anonymous:1880:SPa

[1064] Anonymous. Stormy petrel. Scientific American, 43(22):343– 344, November 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n22/pdf/scientificamerican11271880-
343.pdf.

Anonymous:1880:FOH

[1065] Anonymous. The fan as an object of hygiene. Scientific American, 43(22):344, November 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n22/pdf/scientificamerican11271880-
344.pdf.

Anonymous:1880:EIx

[1066] Anonymous. Engineering inventions. Scientific American, 43(22): 344–345, November 27, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n22/pdf/scientificamerican11271880-
344a.pdf.

Anonymous:1880:Cx

[1067] Anonymous. Correspondence. Scientific American, 43(23):360–361, December 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n23/pdf/scientificamerican12041880-360.pdf.

Anonymous:1880:BPbt

[1068] Anonymous. Business and personal. Scientific American, 43 (23):362, December 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n23/pdf/scientificamerican12041880-
362b.pdf.

Anonymous:1880:BNBx

[1069] Anonymous. Books: New books and publications. Scientific American, 43(23):362, December 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n23/pdf/scientificamerican12041880-
362c.pdf.

Anonymous:1880:NQbn

[1070] Anonymous. Notes and queries. Scientific American, 43(23): 362–363, December 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n23/pdf/scientificamerican12041880-
362d.pdf.

Anonymous:1880:IIbo

[1071] Anonymous. Index of inventions. Scientific American, 43(23):363, December 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n23/pdf/scientificamerican12041880-363.pdf.

Anonymous:1880:AINbb

[1072] Anonymous. American industries — no. 61. Scientific American, 43 (23):351–352, December 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n23/pdf/scientificamerican12041880-
351.pdf.

Anonymous:1880:MMc

[1073] Anonymous. The magnet in medicine. Scientific American, 43 (23):352, December 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n23/pdf/scientificamerican12041880-
352.pdf.

Anonymous:1880:ATB

[1074] Anonymous. The Anthracite trial at the Brooklyn Navy Yard. Scientific American, 43(23):352–353, December 4, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v43/
n23/pdf/scientificamerican12041880-352a.pdf.

Anonymous:1880:PAT

[1075] Anonymous. Progress in American telegraphy. Scientific American, 43(23):353, December 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n23/pdf/scientificamerican12041880-
353.pdf.

Anonymous:1880:NASc

[1076] Anonymous. The National Academy of Science. Scientific American, 43(23):353–354, December 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n23/pdf/scientificamerican12041880-
353a.pdf.

Hopkins:1880:WAL

[1077] M. Hopkins. Wire apparatus for laboratory use. Scientific American, 43(23):354–355, December 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n23/pdf/scientificamerican12041880-
354.pdf.

Anonymous:1880:SBS

[1078] Anonymous. The steam boiler smith. Scientific American, 43 (23):355–356, December 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n23/pdf/scientificamerican12041880-
355.pdf.

Anonymous:1880:CJa

[1079] Anonymous. Changes on Jupiter. Scientific American, 43(23): 356–357, December 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n23/pdf/scientificamerican12041880-
356.pdf.

Anonymous:1880:FACa

[1080] Anonymous. Felt and asbestos covering for steam surfaces. Scientific American, 43(23):357, December 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v43/n23/pdf/scientificamerican12041880-
357.pdf.

Anonymous:1880:MIcn

[1081] Anonymous. Miscellaneous inventions. Scientific American, 43 (23):357–358, December 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n23/pdf/scientificamerican12041880-
357a.pdf.

Anonymous:1880:PM

[1082] Anonymous. Pulping machine. Scientific American, 43(23):358, December 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n23/pdf/scientificamerican12041880-358.pdf.

Anonymous:1880:ISCa

[1083] Anonymous. Improved stone crusher. Scientific American, 43(23): 358–359, December 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n23/pdf/scientificamerican12041880-
358a.pdf.

Anonymous:1880:G

[1084] Anonymous. The Gurami. Scientific American, 43(23):359, December 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n23/pdf/scientificamerican12041880-359.pdf.

Seiss:1880:GS

[1085] O. Few Seiss. The glutinous salamander. Scientific American, 43(23):359, December 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n23/pdf/scientificamerican12041880-
359a.pdf.

Anonymous:1880:CCd

[1086] Anonymous. Chandelier clock. Scientific American, 43(23):361, December 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n23/pdf/scientificamerican12041880-361.pdf.

Anonymous:1880:GP

[1087] Anonymous. A great paper. Scientific American, 43(23):362, December 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n23/pdf/scientificamerican12041880-362.pdf.

Anonymous:1880:SAHa

[1088] Anonymous. Science in aid of the housewife. Scientific American, 43(23):362, December 4, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n23/pdf/scientificamerican12041880-
362a.pdf.

Anonymous:1880:IRIi

[1089] Anonymous. Inventions: Recent inventions. Scientific American, 43 (24):371–372, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1880:Cy

[1090] Anonymous. Correspondence. Scientific American, 43(24):373, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1880:BPbu

[1091] Anonymous. Business and personal. Scientific American, 43 (24):378, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n24/pdf/scientificamerican12111880-
378.pdf.

Anonymous:1880:NQbo

[1092] Anonymous. Notes and queries. Scientific American, 43(24): 378–379, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n24/pdf/scientificamerican12111880-
378a.pdf.

Anonymous:1880:BNBy

[1093] Anonymous. Books: New books and publications. Scientific American, 43(24):379, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n24/pdf/scientificamerican12111880-
379a.pdf.

Anonymous:1880:IIbp

[1094] Anonymous. Index of inventions. Scientific American, 43(24): 379–380, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n24/pdf/scientificamerican12111880-
379b.pdf.

Anonymous:1880:PLb

[1095] Anonymous. On protection from lightning. Scientific American, 43(24):368, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n24/pdf/scientificamerican12111880-
368.pdf.

Anonymous:1880:ELWa

[1096] Anonymous. The electric light on Western river steamers. Scientific American, 43(24):368, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v43/n24/pdf/scientificamerican12111880-
368a.pdf.

Anonymous:1880:HT

[1097] Anonymous. The harmonic telegraph. Scientific American, 43 (24):368, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n24/pdf/scientificamerican12111880-
368b.pdf.

Anonymous:1880:REGa

[1098] Anonymous. Responsibility of employers in Germany. Scientific American, 43(24):368–369, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v43/n24/pdf/scientificamerican12111880-
368c.pdf.

Anonymous:1880:CMM

[1099] Anonymous. The corundum mines, Macon County, N.C. Scientific American, 43(24):369, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1880:HJ

[1100] Anonymous. Hot journals. Scientific American, 43(24):369, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1880:WFb

[1101] Anonymous. World’s Fair in 1883. Scientific American, 43(24):369, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1880:IWb

[1102] Anonymous. An invention wanted. Scientific American, 43(24):369, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1880:AMb

[1103] Anonymous. Amateur mechanics. Scientific American, 43(24):370, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1880:PHB

[1104] Anonymous. Professor Henry’s big magnet. Scientific American, 43(24): 370, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1880:ICCa

[1105] Anonymous. Improved car coupling. Scientific American, 43(24):371, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1880:NRI

[1106] Anonymous. New reference index. Scientific American, 43(24):371, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1880:SAS

[1107] Anonymous. The Stoner automatic scale. Scientific American, 43(24): 371, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1880:NASd

[1108] Anonymous. The National Academy of Sciences. Scientific American, 43(24):372–373, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1880:AINbc

[1109] Anonymous. American industries — no. 62. Scientific American, 43 (24):373, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1880:MIco

[1110] Anonymous. Mechanical inventions. Scientific American, 43(24): 373–374, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n24/pdf/scientificamerican12111880-
373b.pdf.

Anonymous:1880:NMW

[1111] Anonymous. New machine for washing bottles. Scientific American, 43(24):374, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n24/pdf/scientificamerican12111880-
374.pdf.

Anonymous:1880:IFW

[1112] Anonymous. Improvement in feeding and watering live stock on cars. Scientific American, 43(24):374, December 11, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v43/
n24/pdf/scientificamerican12111880-374a.pdf.

Beard:1880:BBD

[1113] Daniel C. Beard. Born blind and deaf. Scientific American, 43 (24):375–376, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n24/pdf/scientificamerican12111880-
375.pdf.

Anonymous:1880:DRPi

[1114] Anonymous. Decisions relating to patents. Scientific American, 43 (24):376–377, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n24/pdf/scientificamerican12111880-
376.pdf.

Anonymous:1880:CRz

[1115] Anonymous. Communications received. Scientific American, 43 (24):379, December 11, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n24/pdf/scientificamerican12111880-
379.pdf.

Anonymous:1880:IRIj

[1116] Anonymous. Inventions: Recent inventions. Scientific American, 43(25):387, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
387b.pdf.

Anonymous:1880:Cz

[1117] Anonymous. Correspondence. Scientific American, 43(25):393– 394, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
393b.pdf.

Anonymous:1880:BPbv

[1118] Anonymous. Business and personal. Scientific American, 43 (25):394, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
394.pdf.

Anonymous:1880:NQbp

[1119] Anonymous. Notes and queries. Scientific American, 43(25): 394–395, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
394a.pdf.

Anonymous:1880:IIbq

[1120] Anonymous. Index of inventions. Scientific American, 43(25): 395, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
395a.pdf.

Anonymous:1880:OMN

[1121] Anonymous. One more number. Scientific American, 43(25): 384, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
384.pdf.

Anonymous:1880:FCP

[1122] Anonymous. Fires-causes and prevention. Scientific American, 43 (25):384, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
384a.pdf.

Anonymous:1880:WL

[1123] Anonymous. What is light? Scientific American, 43(25):384– 385, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
384b.pdf.

Anonymous:1880:GSW

[1124] Anonymous. Glass spinning and weaving. Scientific American, 43 (25):385, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
385.pdf.

Anonymous:1880:REM

[1125] Anonymous. Remarkable eruption of Mauna Loa. Scientific American, 43(25):385, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
385a.pdf.

Anonymous:1880:NAEb

[1126] Anonymous. New air engine. Scientific American, 43(25):385, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n25/pdf/scientificamerican12181880-385b.pdf.

Anonymous:1880:ISN

[1127] Anonymous. Improved safety nut. Scientific American, 43(25): 386, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
386.pdf.

Anonymous:1880:TDC

[1128] Anonymous. Tool for driving and clinching nails. Scientific American, 43(25):386, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
386a.pdf.

Anonymous:1880:ABA

[1129] Anonymous. Automatic balance attachment for valves. Scientific American, 43(25):386, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
386b.pdf.

Anonymous:1880:ISV

[1130] Anonymous. Improved safety valve. Scientific American, 43(25): 386, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
386c.pdf.

Anonymous:1880:NHP

[1131] Anonymous. New hand power band saw. Scientific American, 43 (25):387, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
387.pdf.

Anonymous:1880:ARI

[1132] Anonymous. Apparatus for removing ice from railroads. Scientific American, 43(25):387, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
387a.pdf.

Anonymous:1880:NTI

[1133] Anonymous. New telegraph insulator. Scientific American, 43 (25):387, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
387c.pdf.

Anonymous:1880:BC

[1134] Anonymous. Business colleges. Scientific American, 43(25):388– 389, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
388.pdf.

Anonymous:1880:AMc

[1135] Anonymous. Amateur mechanics. Scientific American, 43(25): 390, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
390.pdf.

Anonymous:1880:MIcp

[1136] Anonymous. Miscellaneous inventions. Scientific American, 43(25): 390–391, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
390a.pdf.

Anonymous:1880:HB

[1137] Anonymous. The Hercules beetle. Scientific American, 43(25): 391, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
391.pdf.

Hogg:1880:VL

[1138] James Hogg. Variegation of leaves. Scientific American, 43(25): 392, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
392.pdf.

Anonymous:1880:DRPj

[1139] Anonymous. Decisions relating to patents. Scientific American, 43(25):393, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
393.pdf.

Anonymous:1880:CWF

[1140] Anonymous. Chinese women’s feet. Scientific American, 43(25): 393, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
393a.pdf.

Anonymous:1880:CRba

[1141] Anonymous. Communications received. Scientific American, 43 (25):395, December 18, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n25/pdf/scientificamerican12181880-
395.pdf.

Anonymous:1880:IRIk

[1142] Anonymous. Inventions: Recent inventions. Scientific American, 43(26):407, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n26/pdf/scientificamerican12251880-
407a.pdf.

Anonymous:1880:BPbw

[1143] Anonymous. Business and personal. Scientific American, 43 (26):408, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n26/pdf/scientificamerican12251880-
408.pdf.

Anonymous:1880:BNBz

[1144] Anonymous. Books: New books and publications. Scientific American, 43(26):408, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n26/pdf/scientificamerican12251880-
408a.pdf.

Anonymous:1880:NQbq

[1145] Anonymous. Notes and queries. Scientific American, 43(26): 408, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n26/pdf/scientificamerican12251880-
408b.pdf.

Anonymous:1880:IIbr

[1146] Anonymous. Index of inventions. Scientific American, 43(26): 409–413, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n26/pdf/scientificamerican12251880-
409.pdf.

Anonymous:1880:AINbd

[1147] Anonymous. American industries — no.63. Scientific American, 43 (26):399–400, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n26/pdf/scientificamerican12251880-
399.pdf.

Anonymous:1880:LN

[1148] Anonymous. The last number. Scientific American, 43(26):400, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n26/pdf/scientificamerican12251880-400.pdf.

Anonymous:1880:EHE

[1149] Anonymous. Eighteen hundred and eighty. Scientific American, 43(26):400, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n26/pdf/scientificamerican12251880-
400a.pdf.

Anonymous:1880:ELB

[1150] Anonymous. Electric lights in Broadway, New York. Scientific American, 43(26):400, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n26/pdf/scientificamerican12251880-
400b.pdf.

Anonymous:1880:SAb

[1151] Anonymous. The Scientific American for 1881. Scientific American, 43(26):401, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n26/pdf/scientificamerican12251880-
401.pdf.

Anonymous:1880:SNY

[1152] Anonymous. Site of the New York Fair of 1883. Scientific American, 43(26):401, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n26/pdf/scientificamerican12251880-
401a.pdf.

Anonymous:1880:CAM

[1153] Anonymous. Compressed air as a motive power. Scientific American, 43(26):401, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n26/pdf/scientificamerican12251880-
401b.pdf.

Anonymous:1880:PHRb

[1154] Anonymous. Progress of the Hudson River Tunnel. Scientific American, 43(26):401, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n26/pdf/scientificamerican12251880-
401c.pdf.

Anonymous:1880:SD

[1155] Anonymous. The spread of diphtheria. Scientific American, 43 (26):401, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n26/pdf/scientificamerican12251880-
401d.pdf.

Anonymous:1880:PHD

[1156] Anonymous. Petroleum for harbor defense. Scientific American, 43(26):401, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n26/pdf/scientificamerican12251880-
401e.pdf.

Anonymous:1880:WSW

[1157] Anonymous. Western side wheelers. Scientific American, 43 (26):402, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n26/pdf/scientificamerican12251880-
402.pdf.

Anonymous:1880:NPMc

[1158] Anonymous. New polishing machine. Scientific American, 43 (26):403, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n26/pdf/scientificamerican12251880-
403.pdf.

Anonymous:1880:IBL

[1159] Anonymous. Improved bomb lance. Scientific American, 43(26): 403, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n26/pdf/scientificamerican12251880-
403a.pdf.

Anonymous:1880:NCF

[1160] Anonymous. New cistern filter. Scientific American, 43(26):403, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n26/pdf/scientificamerican12251880-403b.pdf.

Anonymous:1880:HMH

[1161] Anonymous. Hand mirror holder. Scientific American, 43(26): 403, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n26/pdf/scientificamerican12251880-
403c.pdf.

Anonymous:1880:AIx

[1162] Anonymous. Agricultural inventions. Scientific American, 43(26): 403–404, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n26/pdf/scientificamerican12251880-
403d.pdf.

Anonymous:1880:MIcq

[1163] Anonymous. Miscellaneous inventions. Scientific American, 43(26): 405–406, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n26/pdf/scientificamerican12251880-
405.pdf.

Anonymous:1880:BIS

[1164] Anonymous. Burtons improved street lamp. Scientific American, 43(26):406, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n26/pdf/scientificamerican12251880-
406.pdf.

Anonymous:1880:EIy

[1165] Anonymous. Engineering inventions. Scientific American, 43(26): 406, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n26/pdf/scientificamerican12251880-
406a.pdf.

Anonymous:1880:SWS

[1166] Anonymous. Sand and water spouts. Scientific American, 43 (26):407, December 25, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n26/pdf/scientificamerican12251880-
407.pdf.

Anonymous:1880:Cbb

[1167] Anonymous. Correspondence. Scientific American, 43(3):40, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n3/pdf/scientificamerican07171880-40.pdf.

Anonymous:1880:BPbx

[1168] Anonymous. Business and personal. Scientific American, 43(3):43, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n3/pdf/scientificamerican07171880-43.pdf.

Anonymous:1880:NQbr

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n3/pdf/scientificamerican07171880-43a.pdf.

Anonymous:1880:Ig

[1170] Anonymous. Index. Scientific American, 43(3):44, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v43/
n3/pdf/scientificamerican07171880-44.pdf.

Anonymous:1880:GUP

[1171] Anonymous. Government use of patents. Scientific American, 43(3):32, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n3/pdf/scientificamerican07171880-32.pdf.

Anonymous:1880:AIOa

[1172] Anonymous. American industries — our illustrated series. Scientific American, 43(3):32, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n3/pdf/scientificamerican07171880-
32a.pdf.

Anonymous:1880:QFM

[1173] Anonymous. A quarantine fleet for the Mississippi. Scientific American, 43(3):32, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n3/pdf/scientificamerican07171880-
32b.pdf.

Anonymous:1880:IRM

[1174] Anonymous. Increase in the rubber manufacture. Scientific American, 43(3):32–33, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n3/pdf/scientificamerican07171880-
32c.pdf.

Anonymous:1880:PCG

[1175] Anonymous. A patent case of general interest. Scientific American, 43(3):33, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n3/pdf/scientificamerican07171880-
33.pdf.

Anonymous:1880:BOT

[1176] Anonymous. Burning of oil tanks by lightning. Scientific American, 43(3):33, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n3/pdf/scientificamerican07171880-
33a.pdf.

Anonymous:1880:SFN

[1177] Anonymous. Series forty-nine. Scientific American, 43(3):33, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n3/pdf/scientificamerican07171880-33b.pdf.

Anonymous:1880:BPAa

[1178] Anonymous. The boss puzzle abroad. Scientific American, 43(3):33, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n3/pdf/scientificamerican07171880-33c.pdf.

Anonymous:1880:DSA

[1179] Anonymous. Disastrous steamboat accident. Scientific American, 43(3):33, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n3/pdf/scientificamerican07171880-
33d.pdf.

Anonymous:1880:AINbe

[1180] Anonymous. American industries. — no.49. Scientific American, 43(3):34–35, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n3/pdf/scientificamerican07171880-
34.pdf.

Anonymous:1880:ISL

[1181] Anonymous. Improved street lamp. Scientific American, 43(3):35, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n3/pdf/scientificamerican07171880-35.pdf.

Anonymous:1880:ASCb

[1182] Anonymous. Automatic safety cylinder cock. Scientific American, 43(3):35, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n3/pdf/scientificamerican07171880-
35a.pdf.

Anonymous:1880:EIz

[1183] Anonymous. Engineering inventions. Scientific American, 43(3):35, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n3/pdf/scientificamerican07171880-35b.pdf.

Anonymous:1880:NW

[1184] Anonymous. A novel wheelbarrow. Scientific American, 43(3):35, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n3/pdf/scientificamerican07171880-35c.pdf.

Maxwell:1880:RSS

[1185] E. N. Maxwell. Repairing spiral springs. Scientific American, 43(3):36, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n3/pdf/scientificamerican07171880-36.pdf.

Anonymous:1880:NOS

[1186] Anonymous. New ore separator. Scientific American, 43(3):36, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n3/pdf/scientificamerican07171880-36a.pdf.

Anonymous:1880:ICd

[1187] Anonymous. Improved cartridge. Scientific American, 43(3):36, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n3/pdf/scientificamerican07171880-36b.pdf.

Anonymous:1880:ETT

[1188] Anonymous. Electro-thermic telephone. Scientific American, 43(3):37, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n3/pdf/scientificamerican07171880-37.pdf.

Anonymous:1880:NTE

[1189] Anonymous. New telephone experiments. Scientific American, 43(3):37, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n3/pdf/scientificamerican07171880-37a.pdf.

Anonymous:1880:NGC

[1190] Anonymous. Novel gate closer. Scientific American, 43(3):38, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n3/pdf/scientificamerican07171880-38.pdf.

Anonymous:1880:NCM

[1191] Anonymous. New car mover. Scientific American, 43(3):38, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n3/pdf/scientificamerican07171880-38a.pdf.

Anonymous:1880:MIcr

[1192] Anonymous. Miscellaneous inventions. Scientific American, 43(3):38, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n3/pdf/scientificamerican07171880-38b.pdf.

Anonymous:1880:TH

[1193] Anonymous. The Terebella and Hermella. Scientific American, 43(3):39, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n3/pdf/scientificamerican07171880-
39.pdf.

Roberts:1880:FP

[1194] A. W. Roberts. Fish parasites. Scientific American, 43(3):39–40, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n3/pdf/scientificamerican07171880-39a.pdf.

Anonymous:1880:SMc

[1195] Anonymous. The structure of matter. Scientific American, 43(3):41–42, July 17, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n3/pdf/scientificamerican07171880-41.pdf.

Anonymous:1880:CRbb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n3/pdf/scientificamerican07171880-43b.pdf.

Anonymous:1880:BPby

[1197] Anonymous. Business and personal. Scientific American, 43(4):59, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n4/pdf/scientificamerican07241880-59.pdf.

Anonymous:1880:BNBba

[1198] Anonymous. Books: New books and publications. Scientific American, 43(4):59, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n4/pdf/scientificamerican07241880-
59a.pdf.

Anonymous:1880:NQbs

[1199] Anonymous. Notes and queries. Scientific American, 43(4):59–60, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n4/pdf/scientificamerican07241880-59b.pdf.

Anonymous:1880:Ik

[1200] Anonymous. Index. Scientific American, 43(4):60, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v43/
n4/pdf/scientificamerican07241880-60.pdf.

Anonymous:1880:PNPv

[1201] Anonymous. Patent notes: Patents. Scientific American, 43(4):61, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n4/pdf/scientificamerican07241880-61.pdf.

Anonymous:1880:MMM

[1202] Anonymous. Matter as a mode of motion. Scientific American, 43(4):48, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n4/pdf/scientificamerican07241880-48.pdf.

Anonymous:1880:E

[1203] Anonymous. Explosives. Scientific American, 43(4):48–49, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n4/pdf/scientificamerican07241880-48a.pdf.

Anonymous:1880:CCP

[1204] Anonymous. Curious capillary phenomena. Scientific American, 43(4):49, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n4/pdf/scientificamerican07241880-
49.pdf.

Anonymous:1880:PHa

[1205] Anonymous. Panama hats. Scientific American, 43(4):49, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n4/pdf/scientificamerican07241880-49a.pdf.

Anonymous:1880:PWb

[1206] Anonymous. The protection of woodwork. Scientific American, 43(4):49, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n4/pdf/scientificamerican07241880-
49b.pdf.

Anonymous:1880:BSL

[1207] Anonymous. The Bartholdi Statue of Liberty. Scientific American, 43(4):49, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n4/pdf/scientificamerican07241880-
49c.pdf.

Anonymous:1880:MIcs

[1208] Anonymous. Mechanical inventions. Scientific American, 43(4):50, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n4/pdf/scientificamerican07241880-50.pdf.

Anonymous:1880:IRP

[1209] Anonymous. Improved rotary pump. Scientific American, 43(4):50, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n4/pdf/scientificamerican07241880-50a.pdf.

Anonymous:1880:IPH

[1210] Anonymous. Improved picture hanger. Scientific American, 43(4):50, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n4/pdf/scientificamerican07241880-50b.pdf.

Anonymous:1880:EIba

[1211] Anonymous. Engineering inventions. Scientific American, 43(4):50–51, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n4/pdf/scientificamerican07241880-50c.pdf.

Anonymous:1880:AMd

[1212] Anonymous. Amateur mechanics. Scientific American, 43(4):51, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n4/pdf/scientificamerican07241880-51.pdf.

Anonymous:1880:MIct

[1213] Anonymous. Miscellaneous inventions. Scientific American, 43 (4):51–52, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n4/pdf/scientificamerican07241880-
51a.pdf.

Anonymous:1880:FLE

[1214] Anonymous. Fonvielle & Lontin’s electrical motor. Scientific American, 43(4):52, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n4/pdf/scientificamerican07241880-
52.pdf.

Anonymous:1880:NIe

[1215] Anonymous. New inventions. Scientific American, 43(4):52, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n4/pdf/scientificamerican07241880-52a.pdf.

Anonymous:1880:AINbf

[1216] Anonymous. American industries. — no. 50. Scientific American, 43(4):53, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n4/pdf/scientificamerican07241880-
53.pdf.

Anonymous:1880:DHH

[1217] Anonymous. The demand for heavy horses. Scientific American, 43(4):53, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n4/pdf/scientificamerican07241880-
53a.pdf.

Anonymous:1880:HHR

[1218] Anonymous. The Holland hydrocarbon retort. Scientific American, 43(4):54, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n4/pdf/scientificamerican07241880-
54.pdf.

Anonymous:1880:SKF

[1219] Anonymous. Simple key fastener. Scientific American, 43(4):54–55, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n4/pdf/scientificamerican07241880-54a.pdf.

Anonymous:1880:AIy

[1220] Anonymous. Agricultural inventions. Scientific American, 43(4):55, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n4/pdf/scientificamerican07241880-55.pdf.

Anonymous:1880:MESb

[1221] Anonymous. The mouse-eating spider at the zoo. Scientific American, 43(4):55, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n4/pdf/scientificamerican07241880-
55a.pdf.

Anonymous:1880:JA

[1222] Anonymous. Japanese art. Scientific American, 43(4):55, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n4/pdf/scientificamerican07241880-55b.pdf.

Anonymous:1880:NHNk

[1223] Anonymous. Natural history notes. Scientific American, 43(4):55–56, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n4/pdf/scientificamerican07241880-55c.pdf.

Anonymous:1880:CDS

[1224] Anonymous. Capt. Dobbins’s self-righting life boat. Scientific American, 43(4):56–58, July 24, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n4/pdf/scientificamerican07241880-
56.pdf.

Anonymous:1880:INIq

[1225] Anonymous. Inventions: New inventions. Scientific American, 43(5):70, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n5/pdf/scientificamerican07311880-70b.pdf.

Anonymous:1880:BPbz

[1226] Anonymous. Business and personal. Scientific American, 43(5):75, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n5/pdf/scientificamerican07311880-75.pdf.

Anonymous:1880:BNBbb

[1227] Anonymous. Books: New books and publications. Scientific American, 43(5):75, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n5/pdf/scientificamerican07311880-
75a.pdf.

Anonymous:1880:NQbt

[1228] Anonymous. Notes and queries. Scientific American, 43(5):75–76, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1880:IIbs

[1229] Anonymous. Index of inventions. Scientific American, 43(5):76, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n5/pdf/scientificamerican07311880-76a.pdf.

Anonymous:1880:WFc

[1230] Anonymous. The World’s Fair of 1883. Scientific American, 43(5):64, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n5/pdf/scientificamerican07311880-64.pdf.

Anonymous:1880:RDRh

[1231] Anonymous. The removal of diamond reef New York Harbor. Scientific American, 43(5):64, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n5/pdf/scientificamerican07311880-
64a.pdf.

Riley:1880:NOA

[1232] C. V. Riley. Notes and observations on the army worm. Scientific American, 43(5):64–65, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n5/pdf/scientificamerican07311880-
64b.pdf.

Anonymous:1880:MOT

[1233] Anonymous. More oil tanks fired by lightning. Scientific American, 43(5):65, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n5/pdf/scientificamerican07311880-
65.pdf.

Anonymous:1880:ACSb

[1234] Anonymous. The American Chemical Society. Scientific American, 43(5):65, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n5/pdf/scientificamerican07311880-
65a.pdf.

Anonymous:1880:SIOa

[1235] Anonymous. Staten Island and oysters. Scientific American, 43(5):65, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n5/pdf/scientificamerican07311880-65b.pdf.

Anonymous:1880:NSMc

[1236] Anonymous. Novel steam motor. Scientific American, 43(5):66, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n5/pdf/scientificamerican07311880-66.pdf.

Anonymous:1880:Pj

[1237] Anonymous. The profilograph. Scientific American, 43(5):66, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n5/pdf/scientificamerican07311880-66a.pdf.

Anonymous:1880:MIcu

[1238] Anonymous. Mechanical inventions. Scientific American, 43(5):66–67, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n5/pdf/scientificamerican07311880-66b.pdf.

Anonymous:1880:NVd

[1239] Anonymous. A new velocipede. Scientific American, 43(5):67, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n5/pdf/scientificamerican07311880-67.pdf.

Anonymous:1880:ICe

[1240] Anonymous. Improved closets. Scientific American, 43(5):67, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n5/pdf/scientificamerican07311880-67a.pdf.

Anonymous:1880:MIcv

[1241] Anonymous. Miscellaneous inventions. Scientific American, 43 (5):67–68, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n5/pdf/scientificamerican07311880-
67b.pdf.

Anonymous:1880:SAc

[1242] Anonymous. The steamer Anthracite. Scientific American, 43(5):68, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n5/pdf/scientificamerican07311880-68.pdf.

Anonymous:1880:NSGb

[1243] Anonymous. Novel spy-glass. Scientific American, 43(5):68, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n5/pdf/scientificamerican07311880-68a.pdf.

Anonymous:1880:NTK

[1244] Anonymous. New table knife. Scientific American, 43(5):68, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n5/pdf/scientificamerican07311880-68b.pdf.

Anonymous:1880:MWR

[1245] Anonymous. The man who ran the first locomotive in America. Scientific American, 43(5):68, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n5/pdf/scientificamerican07311880-
68c.pdf.

Anonymous:1880:BT

[1246] Anonymous. Brick tea. Scientific American, 43(5):68, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n5/pdf/scientificamerican07311880-68d.pdf.

Anonymous:1880:AINbg

[1247] Anonymous. American industries — no. 51. Scientific American, 43(5):69, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n5/pdf/scientificamerican07311880-
69.pdf.

Anonymous:1880:AIz

[1248] Anonymous. Agricultural inventions. Scientific American, 43(5):69, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n5/pdf/scientificamerican07311880-69a.pdf.

Anonymous:1880:IMSb

[1249] Anonymous. Improved method of securing car wheels to the axles. Scientific American, 43(5):70, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v43/n5/pdf/scientificamerican07311880-
70.pdf.

Anonymous:1880:NMR

[1250] Anonymous. Novel method of running railroad cars. Scientific American, 43(5):70, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n5/pdf/scientificamerican07311880-
70a.pdf.

Roberts:1880:LC

[1251] A. W. Roberts. The leather carp. Scientific American, 43(5):71, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n5/pdf/scientificamerican07311880-71.pdf.

Holder:1880:MBF

[1252] C. F. Holder. Monster bones (fossils) from the Ancient Cretaceous Seas of Kansas and Colorado. Scientific American, 43(5):71–74, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n5/pdf/scientificamerican07311880-71a.pdf.

Anonymous:1880:CRbc

[1253] Anonymous. Communications received. Scientific American, 43(5):76, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n5/pdf/scientificamerican07311880-76.pdf.

Anonymous:1880:Aa

[1254] Anonymous. Advertisements. Scientific American, 43(5):76–78, July 31, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n5/pdf/scientificamerican07311880-76b.pdf.

Anonymous:1880:NQbu

[1255] Anonymous. Notes and queries. Scientific American, 43(6):91–94, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n6/pdf/scientificamerican08071880-91b.pdf.

Anonymous:1880:IIbt

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n6/pdf/scientificamerican08071880-91c.pdf.

Anonymous:1880:SAHb

[1257] Anonymous. Serious accident in the Hudson River Tunnel. Scientific American, 43(6):80, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n6/pdf/scientificamerican08071880-
80.pdf.

Anonymous:1880:HTS

[1258] Anonymous. How trees are struck by lightning. Scientific American, 43(6):80, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n6/pdf/scientificamerican08071880-
80a.pdf.

Anonymous:1880:AEO

[1259] Anonymous. Arrival of the Egyptian Obelisk at New York. Scientific American, 43(6):80, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n6/pdf/scientificamerican08071880-
80b.pdf.

Anonymous:1880:RME

[1260] Anonymous. The recent Milling Exhibition. Scientific American, 43(6):81, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n6/pdf/scientificamerican08071880-
81.pdf.

Anonymous:1880:AWCb

[1261] Anonymous. Artesian wells in California. Scientific American, 43(6):81, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n6/pdf/scientificamerican08071880-81a.pdf.

Anonymous:1880:DAT

[1262] Anonymous. Dental attachment for telephones. Scientific American, 43(6):82, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n6/pdf/scientificamerican08071880-
82.pdf.

Anonymous:1880:ISe

[1263] Anonymous. Improvement in sewers. Scientific American, 43(6):82, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n6/pdf/scientificamerican08071880-82a.pdf.

Anonymous:1880:NCO

[1264] Anonymous. Novel can opener. Scientific American, 43(6):82, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n6/pdf/scientificamerican08071880-82b.pdf.

Anonymous:1880:NIf

[1265] Anonymous. New inventions. Scientific American, 43(6):82–83, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n6/pdf/scientificamerican08071880-82c.pdf.

Anonymous:1880:IHd

[1266] Anonymous. An improved harvester. Scientific American, 43(6):83, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n6/pdf/scientificamerican08071880-83.pdf.

Anonymous:1880:NPP

[1267] Anonymous. A new punching press. Scientific American, 43(6):83–84, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n6/pdf/scientificamerican08071880-83a.pdf.

Anonymous:1880:AINbh

[1268] Anonymous. American industries — no. 52. Scientific American, 43(6):84–85, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n6/pdf/scientificamerican08071880-
84.pdf.

Anonymous:1880:NFG

[1269] Anonymous. Novel fruit gatherer. Scientific American, 43(6):85, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n6/pdf/scientificamerican08071880-85.pdf.

Anonymous:1880:PDH

[1270] Anonymous. Pneumatic drill-hole cleaner. Scientific American, 43(6):85, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n6/pdf/scientificamerican08071880-
85a.pdf.

Anonymous:1880:ISf

[1271] Anonymous. Improvement in steamships. Scientific American, 43(6):86, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n6/pdf/scientificamerican08071880-86.pdf.

Anonymous:1880:MIcw

[1272] Anonymous. Mechanical inventions. Scientific American, 43(6):86–87, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n6/pdf/scientificamerican08071880-86a.pdf.

Anonymous:1880:Ab

[1273] Anonymous. The Anthracite. Scientific American, 43(6):87–88, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n6/pdf/scientificamerican08071880-87.pdf.

Anonymous:1880:SIOb

[1274] Anonymous. Staten Island and oysters. Scientific American, 43(6):88, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n6/pdf/scientificamerican08071880-88.pdf.

Anonymous:1880:EIbb

[1275] Anonymous. Engineering inventions. Scientific American, 43(6):88–89, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n6/pdf/scientificamerican08071880-88a.pdf.

Anonymous:1880:MIcx

[1276] Anonymous. Miscellaneous inventions. Scientific American, 43 (6):89–90, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n6/pdf/scientificamerican08071880-
89.pdf.

Anonymous:1880:RDRi

[1277] Anonymous. Recent decisions relating to patents. Scientific American, 43(6):90, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n6/pdf/scientificamerican08071880-
90.pdf.

Anonymous:1880:BPca

[1278] Anonymous. Business and personal. Scientific American, 43(6):91, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n6/pdf/scientificamerican08071880-91.pdf.

Anonymous:1880:NBP

[1279] Anonymous. New books and publications. Scientific American, 43(6):91, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n6/pdf/scientificamerican08071880-
91a.pdf.

Anonymous:1880:Ae

[1280] Anonymous. Advertisements. Scientific American, 43(6):92–94, August 7, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n6/pdf/scientificamerican08071880-92.pdf.

Anonymous:1880:NQbv

[1281] Anonymous. Notes and queries. Scientific American, 43(07):107, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n7/pdf/scientificamerican08141880-107a.pdf.

Anonymous:1880:ICR

[1282] Anonymous. Index: Communications received. Scientific American, 43(07):107, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n7/pdf/scientificamerican08141880-
107b.pdf.

Anonymous:1880:IIbu 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n7/pdf/scientificamerican08141880-
107c.pdf.

Anonymous:1880:RPR

[1284] Anonymous. Railway progress and requirements, and more. Scientific American, 43(07):96, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n7/pdf/scientificamerican08141880-
96.pdf.

Anonymous:1880:EIbc

[1285] Anonymous. The evolution of ideas. Scientific American, 43(07):97, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n7/pdf/scientificamerican08141880-97.pdf.

Anonymous:1880:PVD

[1286] Anonymous. The prevention of virulent diseases. Scientific American, 43(07):97, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n7/pdf/scientificamerican08141880-
97a.pdf.

Anonymous:1880:NCL

[1287] Anonymous. Novel clothes line pulley. Scientific American, 43 (07):98, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n7/pdf/scientificamerican08141880-
98.pdf.

Anonymous:1880:NFBa

[1288] Anonymous. New folding boat. Scientific American, 43(07):98, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n7/pdf/scientificamerican08141880-98a.pdf.

Anonymous:1880:VTS

[1289] Anonymous. A visit to the Testing Station of the Industrial Academy, Berlin. Scientific American, 43(07):98, August 14, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v43/
n7/pdf/scientificamerican08141880-98b.pdf.

Anonymous:1880:EIbd

[1290] Anonymous. Engineering inventions. Scientific American, 43 (07):98, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n7/pdf/scientificamerican08141880-
98c.pdf.

Anonymous:1880:ISg

[1291] Anonymous. Improvement in stoves. Scientific American, 43 (07):99, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n7/pdf/scientificamerican08141880-
99.pdf.

Anonymous:1880:ICSb

[1292] Anonymous. Improved car starter. Scientific American, 43(07): 99–100, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n7/pdf/scientificamerican08141880-
99a.pdf.

Anonymous:1880:AINbi

[1293] Anonymous. American industries. — no. 54. Scientific American, 43(07):100, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n7/pdf/scientificamerican08141880-
100.pdf.

Anonymous:1880:AOH

[1294] Anonymous. Astronomical observations at high elevations. Scientific American, 43(07):100–101, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v43/n7/pdf/scientificamerican08141880-
100a.pdf.

Anonymous:1880:ICT

[1295] Anonymous. Improved cross tie. Scientific American, 43(07):101, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n7/pdf/scientificamerican08141880-101.pdf.

Anonymous:1880:AMS

[1296] Anonymous. American milling as seen by an Englishman. Scientific American, 43(07):101, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n7/pdf/scientificamerican08141880-
101a.pdf.

Anonymous:1880:MIcy

[1297] Anonymous. Mechanical inventions. Scientific American, 43(07): 101, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n7/pdf/scientificamerican08141880-
101b.pdf.

Anonymous:1880:ESS

[1298] Anonymous. The extraction of the salts of sodium and potassium from the mother lye of the Mediterranean salt pits. Scientific American, 43(07):102, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n7/pdf/scientificamerican08141880-
102.pdf.

Anonymous:1880:Ld

[1299] Anonymous. The Livadia. Scientific American, 43(07):102, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n7/pdf/scientificamerican08141880-102a.pdf.

Anonymous:1880:NIg

[1300] Anonymous. New inventions. Scientific American, 43(07):102, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n7/pdf/scientificamerican08141880-102b.pdf.

Roberts:1880:Sc

[1301] A. W. Roberts. The sturgeon. Scientific American, 43(07):103–104, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n7/pdf/scientificamerican08141880-103.pdf.

Anonymous:1880:FACb

[1302] Anonymous. Formation of American coal. Scientific American, 43(07):104, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n7/pdf/scientificamerican08141880-
104.pdf.

Anonymous:1880:NUC

[1303] Anonymous. New use for Cyprus reed stems, and more. Scientific American, 43(07):104–105, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n7/pdf/scientificamerican08141880-
104a.pdf.

Anonymous:1880:RDRj

[1304] Anonymous. Recent decisions relating to patents, and more. Scientific American, 43(07):106, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n7/pdf/scientificamerican08141880-
106.pdf.

Anonymous:1880:AIbb

[1305] Anonymous. Agricultural inventions. Scientific American, 43 (07):106, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n7/pdf/scientificamerican08141880-
106a.pdf.

Anonymous:1880:GHC

[1306] Anonymous. The geological history of the Catskills, and more. Scientific American, 43(07):106, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n7/pdf/scientificamerican08141880-
106b.pdf.

Anonymous:1880:BPcb

[1307] Anonymous. Business and personal. Scientific American, 43 (07):107, August 14, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n7/pdf/scientificamerican08141880-
107.pdf.

Anonymous:1880:IRIl

[1308] Anonymous. Inventions: Recent inventions. Scientific American, 43(8):121, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n8/pdf/scientificamerican08211880-
121.pdf.

Anonymous:1880:Cbc

[1309] Anonymous. Correspondence. Scientific American, 43(8):121–122, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n8/pdf/scientificamerican08211880-121b.pdf.

Anonymous:1880:BPcc

[1310] Anonymous. Business and personal. Scientific American, 43(8):123, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n8/pdf/scientificamerican08211880-123.pdf.

Anonymous:1880:NQbw

[1311] Anonymous. Notes and queries. Scientific American, 43(8):123, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n8/pdf/scientificamerican08211880-123a.pdf.

Anonymous:1880:IIbv

[1312] Anonymous. Index of inventions. Scientific American, 43(8): 123–124, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n8/pdf/scientificamerican08211880-
123c.pdf.

Anonymous:1880:PNPw

[1313] Anonymous. Patent notes: Patents. Scientific American, 43(8):125, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n8/pdf/scientificamerican08211880-125.pdf.

Anonymous:1880:DTF

[1314] Anonymous. Dr. Tanner’s fasting experiment. Scientific American, 43(8):112, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n8/pdf/scientificamerican08211880-
112.pdf.

Anonymous:1880:OPI

[1315] Anonymous. Our pottery industry. Scientific American, 43(8): 112–113, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n8/pdf/scientificamerican08211880-
112a.pdf.

Anonymous:1880:ARC

[1316] Anonymous. Another rain controller. Scientific American, 43 (8):113, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n8/pdf/scientificamerican08211880-
113.pdf.

Anonymous:1880:REGb

[1317] Anonymous. Remarkable explosions of gas. Scientific American, 43(8):113, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n8/pdf/scientificamerican08211880-
113a.pdf.

Anonymous:1880:THM

[1318] Anonymous. The Texas hydraulic mineral belt. Scientific American, 43(8):113, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n8/pdf/scientificamerican08211880-
113b.pdf.

Anonymous:1880:AWP

[1319] Anonymous. Arsenio in wall papers. Scientific American, 43(8):113, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n8/pdf/scientificamerican08211880-113c.pdf.

Anonymous:1880:MMd

[1320] Anonymous. The magnet in medicine. Scientific American, 43 (8):113–114, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n8/pdf/scientificamerican08211880-
113d.pdf.

Anonymous:1880:ISGa

[1321] Anonymous. Improved steam generator. Scientific American, 43 (8):114, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n8/pdf/scientificamerican08211880-
114.pdf.

Anonymous:1880:MIcz

[1322] Anonymous. Miscellaneous inventions. Scientific American, 43 (8):114–115, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n8/pdf/scientificamerican08211880-
114a.pdf.

Anonymous:1880:EIbe

[1323] Anonymous. Engineering inventions. Scientific American, 43(8): 115, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n8/pdf/scientificamerican08211880-
115.pdf.

Anonymous:1880:LT

[1324] Anonymous. Lifting tackle. Scientific American, 43(8):115, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n8/pdf/scientificamerican08211880-115a.pdf.

Anonymous:1880:IGBb

[1325] Anonymous. Improved galvanic battery. Scientific American, 43 (8):115, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n8/pdf/scientificamerican08211880-
115b.pdf.

Anonymous:1880:NIh

[1326] Anonymous. New inventions. Scientific American, 43(8):115–116, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n8/pdf/scientificamerican08211880-115c.pdf.

Anonymous:1880:ELWb

[1327] Anonymous. The electric lamps of W. Tchikoleff. Scientific American, 43(8):116, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n8/pdf/scientificamerican08211880-
116.pdf.

Anonymous:1880:DRPk

[1328] Anonymous. Decisions relating to patents. Scientific American, 43 (8):116–117, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n8/pdf/scientificamerican08211880-
116a.pdf.

Anonymous:1880:AINbj

[1329] Anonymous. American industries — no. 64. Scientific American, 43(8):117–118, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n8/pdf/scientificamerican08211880-
117.pdf.

Anonymous:1880:IWJ

[1330] Anonymous. Improved wagon jack. Scientific American, 43(8):118, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n8/pdf/scientificamerican08211880-118.pdf.

Anonymous:1880:NAC

[1331] Anonymous. New automatic clamping and pneumatic paper cutting machine. Scientific American, 43(8):118, August 21, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v43/
n8/pdf/scientificamerican08211880-118a.pdf.

Anonymous:1880:IHe

[1332] Anonymous. Improved hame. Scientific American, 43(8):118, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n8/pdf/scientificamerican08211880-118b.pdf.

Anonymous:1880:MIda

[1333] Anonymous. Mechanical inventions. Scientific American, 43(8): 118–119, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n8/pdf/scientificamerican08211880-
118c.pdf.

Lucas:1880:BFE

[1334] Frederic A. Lucas. The Berlin Fisheries Exhibition. Scientific American, 43(8):119, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n8/pdf/scientificamerican08211880-
119.pdf.

Anonymous:1880:NHNl

[1335] Anonymous. Natural history notes. Scientific American, 43(8): 119–120, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n8/pdf/scientificamerican08211880-
119a.pdf.

Anonymous:1880:AT

[1336] Anonymous. The Ailantus tree. Scientific American, 43(8):120, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n8/pdf/scientificamerican08211880-120.pdf.

Anonymous:1880:AIbc

[1337] Anonymous. Agricultural inventions. Scientific American, 43(8): 120–121, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n8/pdf/scientificamerican08211880-
120a.pdf.

VanderWeyde:1880:SAH

[1338] P. H. Vander Weyde. A safety appliance for the Hudson River Tunnel. Scientific American, 43(8):121, August 21, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v43/
n8/pdf/scientificamerican08211880-121a.pdf.

Anonymous:1880:ACP

[1339] Anonymous. The Acme cube pipe tongs. Scientific American, 43 (8):122–123, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n8/pdf/scientificamerican08211880-
122.pdf.

Anonymous:1880:CRbd

[1340] Anonymous. Communications received. Scientific American, 43 (8):123, August 21, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n8/pdf/scientificamerican08211880-
123b.pdf.

Anonymous:1880:INIr

[1341] Anonymous. Inventions: New inventions. Scientific American, 43 (9):131–132, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n9/pdf/scientificamerican08281880-
131b.pdf.

Anonymous:1880:IRIm

[1342] Anonymous. Inventions: Recent inventions. Scientific American, 43(9):138, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n9/pdf/scientificamerican08281880-
138b.pdf.

Anonymous:1880:BPcd

[1343] Anonymous. Business and personal. Scientific American, 43(9):139, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n9/pdf/scientificamerican08281880-139.pdf.

Anonymous:1880:BNBbc

[1344] Anonymous. Books: New books and publications. Scientific American, 43(9):139, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n9/pdf/scientificamerican08281880-
139a.pdf.

Anonymous:1880:NQbx

[1345] Anonymous. Notes and queries. Scientific American, 43(9):139, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n9/pdf/scientificamerican08281880-139b.pdf.

Anonymous:1880:IIbw

[1346] Anonymous. Index of inventions. Scientific American, 43(9):140, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n9/pdf/scientificamerican08281880-140.pdf.

Anonymous:1880:PNPx

[1347] Anonymous. Patent notes: Patents. Scientific American, 43(9):141, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n9/pdf/scientificamerican08281880-141.pdf.

Anonymous:1880:PIM

[1348] Anonymous. A Peruvian infernal machine. Scientific American, 43(9):127, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n9/pdf/scientificamerican08281880-
127.pdf.

Anonymous:1880:ELM

[1349] Anonymous. Electric light for marine use. Scientific American, 43(9):127, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n9/pdf/scientificamerican08281880-
127a.pdf.

Anonymous:1880:ESBb

[1350] Anonymous. Effect of starvation on the blood. Scientific American, 43(9):128, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n9/pdf/scientificamerican08281880-
128.pdf.

Anonymous:1880:CBCa

[1351] Anonymous. Color blindness in Connecticut. Scientific American, 43(9):128, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n9/pdf/scientificamerican08281880-
128a.pdf.

Anonymous:1880:ISI

[1352] Anonymous. Influx the source of invention. Scientific American, 43(9):129, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n9/pdf/scientificamerican08281880-
129.pdf.

Anonymous:1880:ACA

[1353] Anonymous. Artificial citric acid. Scientific American, 43(9):129, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n9/pdf/scientificamerican08281880-129a.pdf.

Anonymous:1880:GYS

[1354] Anonymous. A good year for statistics. Scientific American, 43(9):129, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n9/pdf/scientificamerican08281880-
129b.pdf.

Anonymous:1880:FPC

[1355] Anonymous. Fire in a Pennsylvania coal mine. Scientific American, 43(9):130, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n9/pdf/scientificamerican08281880-
130.pdf.

Anonymous:1880:CBS

[1356] Anonymous. Clutch for bow strings. Scientific American, 43(9):130, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n9/pdf/scientificamerican08281880-130a.pdf.

Anonymous:1880:MIdb

[1357] Anonymous. Mechanical inventions. Scientific American, 43(9):130, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n9/pdf/scientificamerican08281880-130b.pdf.

Anonymous:1880:LVF

[1358] Anonymous. A lens with variable focus. Scientific American, 43(9):131, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n9/pdf/scientificamerican08281880-
131.pdf.

Anonymous:1880:NCh

[1359] Anonymous. New coupler. Scientific American, 43(9):131, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n9/pdf/scientificamerican08281880-131a.pdf.

Eichler:1880:YRP

[1360] E. Eichler. Yellow and red phosphorus in the manufacture of matches. Scientific American, 43(9):132, August 28, 1880. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v43/
n9/pdf/scientificamerican08281880-132.pdf.

Anonymous:1880:SCc

[1361] Anonymous. A substitute for the crank. Scientific American, 43(9):133, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n9/pdf/scientificamerican08281880-
133.pdf.

Anonymous:1880:IBR

[1362] Anonymous. Improved book-rack. Scientific American, 43(9):133, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n9/pdf/scientificamerican08281880-133a.pdf.

Anonymous:1880:MIdc

[1363] Anonymous. Miscellaneous inventions. Scientific American, 43 (9):133, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n9/pdf/scientificamerican08281880-
133b.pdf.

Anonymous:1880:NAG

[1364] Anonymous. New air gun. Scientific American, 43(9):134, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n9/pdf/scientificamerican08281880-134.pdf.

Anonymous:1880:SPc

[1365] Anonymous. The steamer Pittsburg. Scientific American, 43 (9):134, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n9/pdf/scientificamerican08281880-
134a.pdf.

Anonymous:1880:GB

[1366] Anonymous. The glycerine barometer. Scientific American, 43 (9):134–135, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n9/pdf/scientificamerican08281880-
134b.pdf.

Anonymous:1880:TFS

[1367] Anonymous. Taguan flying squirrel. Scientific American, 43(9): 135–136, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n9/pdf/scientificamerican08281880-
135.pdf.

Anonymous:1880:CRHa

[1368] Anonymous. Clothing in its relation to health. Scientific American, 43(9):136–137, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n9/pdf/scientificamerican08281880-
136.pdf.

Anonymous:1880:EIbf

[1369] Anonymous. Engineering inventions. Scientific American, 43(9): 137, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n9/pdf/scientificamerican08281880-
137.pdf.

Anonymous:1880:FACc

[1370] Anonymous. Facts about cheese. Scientific American, 43(9):138, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n9/pdf/scientificamerican08281880-138.pdf.

Anonymous:1880:ACc

[1371] Anonymous. Artificial cold. Scientific American, 43(9):138, August 28, 1880. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v43/n9/pdf/scientificamerican08281880-138a.pdf.

Anonymous:1880:CRbe (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v43/n9/pdf/scientificamerican08281880-
139c.pdf.

Anonymous:1881:IRIa

[1373] Anonymous. Inventions: Recent inventions. Scientific American, 44(1):8–9, January 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n1/pdf/scientificamerican01011881-
8b.pdf.

Anonymous:1881:BPa

[1374] Anonymous. Business and personal. Scientific American, 44(1):10, January 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n1/pdf/scientificamerican01011881-10.pdf.

Anonymous:1881:BNBa

[1375] Anonymous. Books: New books and publications. Scientific American, 44(1):10, January 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n1/pdf/scientificamerican01011881-
10a.pdf.

Anonymous:1881:NQa

[1376] Anonymous. Notes and queries. Scientific American, 44(1):10–11, January 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n1/pdf/scientificamerican01011881-10b.pdf.

Anonymous:1881:IIa

[1377] Anonymous. Index of inventions. Scientific American, 44(1):11, January 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n1/pdf/scientificamerican01011881-11.pdf.

Anonymous:1881:PNPa

[1378] Anonymous. Patent notes: Patents. Scientific American, 44(1):12, January 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n1/pdf/scientificamerican01011881-12.pdf.

Anonymous:1881:BPb

[1379] Anonymous. Bell’s photophone. Scientific American, 44(1):1, January 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n1/pdf/scientificamerican01011881-1.pdf.

Anonymous:1881:PRS

[1380] Anonymous. Patent royalties on shoe machinery. Scientific American, 44(1):2, January 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n1/pdf/scientificamerican01011881-
2.pdf.

Anonymous:1881:PSH

[1381] Anonymous. The Prall system of heating. Scientific American, 44(1):2, January 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n1/pdf/scientificamerican01011881-2a.pdf.

Anonymous:1881:ROC

[1382] Anonymous. The restoration of our commercial and Naval marine. Scientific American, 44(1):2–3, January 1, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v44/
n1/pdf/scientificamerican01011881-2b.pdf.

Anonymous:1881:EBP

[1383] Anonymous. Electro-brass plating. Scientific American, 44(1):3, January 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n1/pdf/scientificamerican01011881-3.pdf.

Anonymous:1881:CTa

[1384] Anonymous. Cremation temple. Scientific American, 44(1):5, January 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n1/pdf/scientificamerican01011881-5.pdf.

Anonymous:1881:AGW

[1385] Anonymous. Automatic governor for waste water pipes. Scientific American, 44(1):5, January 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n1/pdf/scientificamerican01011881-
5a.pdf.

Anonymous:1881:EIa

[1386] Anonymous. Engineering inventions. Scientific American, 44(1):5, January 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n1/pdf/scientificamerican01011881-5b.pdf.

Anonymous:1881:GLE

[1387] Anonymous. The growth of large elms. Scientific American, 44(1):6, January 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n1/pdf/scientificamerican01011881-6.pdf.

Anonymous:1881:MIa

[1388] Anonymous. Mechanical inventions. Scientific American, 44(1):6, January 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n1/pdf/scientificamerican01011881-6a.pdf.

Anonymous:1881:DRPa

[1389] Anonymous. Decisions relating to patents, trade marks, etc. Scientific American, 44(1):7, January 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n1/pdf/scientificamerican01011881-
7.pdf.

Anonymous:1881:AIa

[1390] Anonymous. Agricultural inventions. Scientific American, 44(1):7–8, January 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n1/pdf/scientificamerican01011881-7a.pdf.

Anonymous:1881:NRE

[1391] Anonymous. Novel road engine. Scientific American, 44(1):8, January 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n1/pdf/scientificamerican01011881-8.pdf.

Anonymous:1881:IH

[1392] Anonymous. Improved hinge. Scientific American, 44(1):8, January 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n1/pdf/scientificamerican01011881-8a.pdf.

Roberts:1881:S

[1393] A. W. Roberts. Scallops. Scientific American, 44(1):9, January 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n1/pdf/scientificamerican01011881-9.pdf.

Anonymous:1881:INIa

[1394] Anonymous. Inventions: New inventions. Scientific American, 44(10):147, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n10/pdf/scientificamerican03051881-
147b.pdf.

Anonymous:1881:Ca

[1395] Anonymous. Correspondence. Scientific American, 44(10):148, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n10/pdf/scientificamerican03051881-148.pdf.

Anonymous:1881:IRIb

[1396] Anonymous. Inventions: Recent inventions. Scientific American, 44(10):152, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n10/pdf/scientificamerican03051881-
152.pdf.

Anonymous:1881:BPc

[1397] Anonymous. Business and personal. Scientific American, 44(10):154, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n10/pdf/scientificamerican03051881-154.pdf.

Anonymous:1881:BNBb

[1398] Anonymous. Books: New books and publications. Scientific American, 44(10):154, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n10/pdf/scientificamerican03051881-
154a.pdf.

Anonymous:1881:NQb

[1399] Anonymous. Notes and queries. Scientific American, 44(10):154–155, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n10/pdf/scientificamerican03051881-154b.pdf.

Anonymous:1881:IIb

[1400] Anonymous. Index of inventions. Scientific American, 44(10):155, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n10/pdf/scientificamerican03051881-155a.pdf.

Anonymous:1881:MFB

[1401] Anonymous. Mammoth fire boat. Scientific American, 44(10):143, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n10/pdf/scientificamerican03051881-143.pdf.

Anonymous:1881:TWC

[1402] Anonymous. Telegraph wires in cities. Scientific American, 44(10):144, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n10/pdf/scientificamerican03051881-144.pdf.

Anonymous:1881:AIR

[1403] Anonymous. Another important reissue decision by the Supreme Court. Scientific American, 44(10):144, March 5, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v44/
n10/pdf/scientificamerican03051881-144a.pdf.

Anonymous:1881:PTM

[1404] Anonymous. Physical training as a means of mental health. Scientific American, 44(10):144–145, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v44/n10/pdf/scientificamerican03051881-
144b.pdf.

Anonymous:1881:SEE

[1405] Anonymous. Steam engines for electric light machinery. Scientific American, 44(10):145, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n10/pdf/scientificamerican03051881-
145.pdf.

Anonymous:1881:NTRa

[1406] Anonymous. New theory in regard to lunar volcanoes. Scientific American, 44(10):145, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n10/pdf/scientificamerican03051881-
145a.pdf.

Anonymous:1881:WAC

[1407] Anonymous. White ants in court. Scientific American, 44(10):145, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n10/pdf/scientificamerican03051881-145b.pdf.

Anonymous:1881:HAC

[1408] Anonymous. The Humane Association’s Cattle Car Competition. Scientific American, 44(10):145, March 5, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v44/
n10/pdf/scientificamerican03051881-145c.pdf.

Anonymous:1881:TR

[1409] Anonymous. A telephone reissue. Scientific American, 44(10):145, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n10/pdf/scientificamerican03051881-145d.pdf.

Anonymous:1881:SWW

[1410] Anonymous. Self-winding watch. Scientific American, 44(10):146, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n10/pdf/scientificamerican03051881-146.pdf.

Anonymous:1881:SSa

[1411] Anonymous. Skate sailing. Scientific American, 44(10):146, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n10/pdf/scientificamerican03051881-146a.pdf.

Anonymous:1881:NDAa

[1412] Anonymous. New drop attachment for bottles. Scientific American, 44(10):146, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n10/pdf/scientificamerican03051881-
146b.pdf.

Anonymous:1881:MIb

[1413] Anonymous. Mechanical inventions. Scientific American, 44(10):146, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n10/pdf/scientificamerican03051881-146c.pdf.

Anonymous:1881:NHD

[1414] Anonymous. New hatcheling device. Scientific American, 44(10):147, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n10/pdf/scientificamerican03051881-147.pdf.

Anonymous:1881:IFH

[1415] Anonymous. Improved fog horn. Scientific American, 44(10):147, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n10/pdf/scientificamerican03051881-147a.pdf.

Anonymous:1881:DRPb

[1416] Anonymous. Decisions relating to patents. Scientific American, 44(10):148, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n10/pdf/scientificamerican03051881-
148a.pdf.

Anonymous:1881:AIb

[1417] Anonymous. Agricultural inventions. Scientific American, 44(10):149, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n10/pdf/scientificamerican03051881-149.pdf.

Anonymous:1881:PM

[1418] Anonymous. Pinking machine. Scientific American, 44(10):150, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n10/pdf/scientificamerican03051881-150.pdf.

Anonymous:1881:AMa

[1419] Anonymous. Amateur mechanics. Scientific American, 44(10):150–151, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n10/pdf/scientificamerican03051881-150a.pdf.

Anonymous:1881:O

[1420] Anonymous. The osprey. Scientific American, 44(10):151, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n10/pdf/scientificamerican03051881-151.pdf.

Anonymous:1881:DBH

[1421] Anonymous. A ducking battery and how to make it. Scientific American, 44(10):153, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n10/pdf/scientificamerican03051881-
153.pdf.

Anonymous:1881:MIc

[1422] Anonymous. Miscellaneous inventions. Scientific American, 44(10):153, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n10/pdf/scientificamerican03051881-153a.pdf.

Anonymous:1881:CRa

[1423] Anonymous. Communications received. Scientific American, 44 (10):155, March 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n10/pdf/scientificamerican03051881-
155.pdf.

Anonymous:1881:INIb

[1424] Anonymous. Inventions: New inventions. Scientific American, 44(11):162, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n11/pdf/scientificamerican03121881-
162.pdf.

Anonymous:1881:IRIc

[1425] Anonymous. Inventions: Recent inventions. Scientific American, 44(11):164, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n11/pdf/scientificamerican03121881-
164.pdf.

Anonymous:1881:Cb

[1426] Anonymous. Correspondence. Scientific American, 44(11):168–169, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n11/pdf/scientificamerican03121881-168.pdf.

Anonymous:1881:BPd

[1427] Anonymous. Business and personal. Scientific American, 44(11):170, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n11/pdf/scientificamerican03121881-170.pdf.

Anonymous:1881:BNBc

[1428] Anonymous. Books: New books and publications. Scientific American, 44(11):170, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n11/pdf/scientificamerican03121881-
170a.pdf.

Anonymous:1881:NQc

[1429] Anonymous. Notes and queries. Scientific American, 44(11):170–171, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n11/pdf/scientificamerican03121881-170b.pdf.

Anonymous:1881:IIc

[1430] Anonymous. Index of inventions. Scientific American, 44(11):171, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n11/pdf/scientificamerican03121881-171a.pdf.

Anonymous:1881:GDC

[1431] Anonymous. Garbage destructor and carbonizer. Scientific American, 44(11):159, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n11/pdf/scientificamerican03121881-
159.pdf.

Anonymous:1881:SVG

[1432] Anonymous. The sounds of vapors and gases — the photophone as an instrument of physical investigation. Scientific American, 44(11):160, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n11/pdf/scientificamerican03121881-160.pdf.

Anonymous:1881:MM

[1433] Anonymous. Midwinter malaria. Scientific American, 44(11):160, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n11/pdf/scientificamerican03121881-160a.pdf.

Anonymous:1881:RCSa

[1434] Anonymous. Rice culture in the Southwest. Scientific American, 44(11):160, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n11/pdf/scientificamerican03121881-
160b.pdf.

Anonymous:1881:CLF

[1435] Anonymous. Compressed lignite as fuel. Scientific American, 44(11):161, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n11/pdf/scientificamerican03121881-
161.pdf.

Anonymous:1881:SFC

[1436] Anonymous. Smokeless fuel from coal. Scientific American, 44(11):161, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n11/pdf/scientificamerican03121881-161a.pdf.

Anonymous:1881:SIU

[1437] Anonymous. The silk industry of the United States. Scientific American, 44(11):161, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n11/pdf/scientificamerican03121881-
161b.pdf.

Anonymous:1881:SOC

[1438] Anonymous. Suppression of one class of interferences. Scientific American, 44(11):161, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n11/pdf/scientificamerican03121881-
161c.pdf.

Anonymous:1881:ICa

[1439] Anonymous. Improvement in currycombs. Scientific American, 44(11):162, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n11/pdf/scientificamerican03121881-
162a.pdf.

Anonymous:1881:NTSa

[1440] Anonymous. New telescope of short focus. Scientific American, 44(11):162, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n11/pdf/scientificamerican03121881-
162b.pdf.

Anonymous:1881:CAV

[1441] Anonymous. Combined anvil and vise. Scientific American, 44(11):162, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n11/pdf/scientificamerican03121881-162c.pdf.

Anonymous:1881:SEa

[1442] Anonymous. Soldering by electricity. Scientific American, 44(11):163, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n11/pdf/scientificamerican03121881-163.pdf.

Anonymous:1881:APS

[1443] Anonymous. Apparatus for preparing starch for finishing linen and cotton goods. Scientific American, 44(11):163, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n11/pdf/scientificamerican03121881-163a.pdf.

Anonymous:1881:SEP

[1444] Anonymous. A simple experiment with polarized light. Scientific American, 44(11):163–164, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n11/pdf/scientificamerican03121881-
163b.pdf.

Anonymous:1881:RTR

[1445] Anonymous. Recording telephonic receiver. Scientific American, 44(11):165, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n11/pdf/scientificamerican03121881-
165.pdf.

Anonymous:1881:RDRa

[1446] Anonymous. Recent decisions relating to patents. Scientific American, 44(11):165–166, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n11/pdf/scientificamerican03121881-
165a.pdf.

Anonymous:1881:SPM

[1447] Anonymous. Slate pencil machine. Scientific American, 44(11):166, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n11/pdf/scientificamerican03121881-166.pdf.

Anonymous:1881:EGF

[1448] Anonymous. Engine with gas-fired boiler. Scientific American, 44(11):166, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n11/pdf/scientificamerican03121881-
166a.pdf.

Anonymous:1881:MId

[1449] Anonymous. Miscellaneous inventions. Scientific American, 44(11):166, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n11/pdf/scientificamerican03121881-166b.pdf.

Roberts:1881:WLS

[1450] A. W. Roberts. The winkle, or ladle-shell. Scientific American, 44(11):167–168, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n11/pdf/scientificamerican03121881-
167.pdf.

Anonymous:1881:ICP

[1451] Anonymous. Improved coffee pot. Scientific American, 44(11):169, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n11/pdf/scientificamerican03121881-169.pdf.

Anonymous:1881:CRb

[1452] Anonymous. Communications received. Scientific American, 44 (11):171, March 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n11/pdf/scientificamerican03121881-
171.pdf.

Anonymous:1881:INIc

[1453] Anonymous. Inventions: New inventions. Scientific American, 44 (12):181–182, March 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n12/pdf/scientificamerican03191881-
181a.pdf.

Anonymous:1881:BPe

[1454] Anonymous. Business and personal. Scientific American, 44(12):186, March 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n12/pdf/scientificamerican03191881-186.pdf.

Anonymous:1881:NQd

[1455] Anonymous. Notes and queries. Scientific American, 44(12):186–187, March 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n12/pdf/scientificamerican03191881-186a.pdf.

Anonymous:1881:IId

[1456] Anonymous. Index of inventions. Scientific American, 44(12):187, March 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n12/pdf/scientificamerican03191881-187a.pdf.

Anonymous:1881:LC

[1457] Anonymous. Lard cheese. Scientific American, 44(12):175, March 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n12/pdf/scientificamerican03191881-175.pdf.

Anonymous:1881:BM

[1458] Anonymous. Barrel machinery. Scientific American, 44(12):175, March 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n12/pdf/scientificamerican03191881-175a.pdf.

Anonymous:1881:ZL

[1459] Anonymous. The zodiacal light. Scientific American, 44(12):176, March 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n12/pdf/scientificamerican03191881-176.pdf.

Anonymous:1881:WBEa

[1460] Anonymous. Whose boilers explode. Scientific American, 44(12):176, March 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n12/pdf/scientificamerican03191881-176a.pdf.

Anonymous:1881:SWC

[1461] Anonymous. Storm warnings in court. Scientific American, 44(12):176, March 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n12/pdf/scientificamerican03191881-176b.pdf.

Anonymous:1881:AW

[1462] Anonymous. Air and water. Scientific American, 44(12):176, March 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n12/pdf/scientificamerican03191881-176c.pdf.

Anonymous:1881:ICCa

[1463] Anonymous. The industrial condition of Canada. Scientific American, 44(12):176–177, March 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n12/pdf/scientificamerican03191881-
176d.pdf.

Anonymous:1881:AD

[1464] Anonymous. Artificial daylight. Scientific American, 44(12):177, March 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n12/pdf/scientificamerican03191881-177.pdf.

Anonymous:1881:SAH

[1465] Anonymous. Sanitary arrangements in houses. Scientific American, 44(12):177, March 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n12/pdf/scientificamerican03191881-
177a.pdf.

Anonymous:1881:MIe

[1466] Anonymous. Miscellaneous inventions. Scientific American, 44 (12):178–179, March 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n12/pdf/scientificamerican03191881-
178.pdf.

Anonymous:1881:NFBa

[1467] Anonymous. New folding bath tub. Scientific American, 44(12):179, March 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n12/pdf/scientificamerican03191881-179.pdf.

Anonymous:1881:IRC

[1468] Anonymous. Improved rope-clamp. Scientific American, 44(12): 179–181, March 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n12/pdf/scientificamerican03191881-
179a.pdf.

Anonymous:1881:ISHa

[1469] Anonymous. Improved swivel-holder for fish-hooks. Scientific American, 44(12):181, March 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n12/pdf/scientificamerican03191881-
181.pdf.

Anonymous:1881:NS

[1470] Anonymous. Novel scissors. Scientific American, 44(12):182, March 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n12/pdf/scientificamerican03191881-182.pdf.

Hopkins:1881:NTT

[1471] Geo. M. Hopkins. New telephone transmitter. Scientific American, 44(12):182, March 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n12/pdf/scientificamerican03191881-
182a.pdf.

Lucas:1881:C

[1472] Frederio A. Lucas. The Cheiromeles. Scientific American, 44(12):183, March 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n12/pdf/scientificamerican03191881-183.pdf.

Anonymous:1881:HSa

[1473] Anonymous. The horned screamer. Scientific American, 44(12): 183–185, March 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n12/pdf/scientificamerican03191881-
183a.pdf.

Anonymous:1881:MIf

[1474] Anonymous. Mechanical inventions. Scientific American, 44(12):185, March 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n12/pdf/scientificamerican03191881-185.pdf.

Anonymous:1881:CRc

[1475] Anonymous. Communications received. Scientific American, 44 (12):187, March 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n12/pdf/scientificamerican03191881-
187.pdf.

Anonymous:1881:IRId

[1476] Anonymous. Inventions: Recent inventions. Scientific American, 44(13):198–199, March 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n13/pdf/scientificamerican03261881-
198a.pdf.

Anonymous:1881:INId

[1477] Anonymous. Inventions: New inventions. Scientific American, 44(13):201, March 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n13/pdf/scientificamerican03261881-
201.pdf.

Anonymous:1881:BPf

[1478] Anonymous. Business and personal. Scientific American, 44(13):202, March 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n13/pdf/scientificamerican03261881-202.pdf.

Anonymous:1881:NQe

[1479] Anonymous. Notes and queries. Scientific American, 44(13):202–203, March 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n13/pdf/scientificamerican03261881-202a.pdf.

Anonymous:1881:IIe

[1480] Anonymous. Index of inventions. Scientific American, 44(13):203, March 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n13/pdf/scientificamerican03261881-203a.pdf.

Anonymous:1881:PNPb

[1481] Anonymous. Patent notes: Patents. Scientific American, 44(13):203, March 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n13/pdf/scientificamerican03261881-203b.pdf.

Anonymous:1881:NTM

[1482] Anonymous. A new trade mark law. Scientific American, 44(13):192, March 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n13/pdf/scientificamerican03261881-192.pdf.

Anonymous:1881:ACM

[1483] Anonymous. The actual cost of making illuminating gas. Scientific American, 44(13):192, March 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n13/pdf/scientificamerican03261881-
192a.pdf.

Anonymous:1881:LB

[1484] Anonymous. Lager beer. Scientific American, 44(13):192–193, March 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n13/pdf/scientificamerican03261881-192b.pdf.

Anonymous:1881:MML

[1485] Anonymous. To make and maintain a lawn. Scientific American, 44(13):193, March 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n13/pdf/scientificamerican03261881-
193.pdf.

Anonymous:1881:IRa

[1486] Anonymous. An interesting region. Scientific American, 44(13):193, March 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n13/pdf/scientificamerican03261881-193a.pdf.

Anonymous:1881:NAGa

[1487] Anonymous. A new American gem. Scientific American, 44(13):193, March 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n13/pdf/scientificamerican03261881-193b.pdf.

Anonymous:1881:AINa

[1488] Anonymous. American industries — no. 68. Scientific American, 44(13):194–195, March 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n13/pdf/scientificamerican03261881-
194.pdf.

Anonymous:1881:NSGa

[1489] Anonymous. New swinging gate. Scientific American, 44(13):195, March 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n13/pdf/scientificamerican03261881-195.pdf.

Anonymous:1881:MIg

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n13/pdf/scientificamerican03261881-195a.pdf.

Anonymous:1881:HNM

[1491] Anonymous. Hiddenite — a new mineral. Scientific American, 44(13):196, March 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n13/pdf/scientificamerican03261881-
196.pdf.

Anonymous:1881:SFL

[1492] Anonymous. Singular flower-like forms of ice. Scientific American, 44(13):196, March 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n13/pdf/scientificamerican03261881-
196a.pdf.

Anonymous:1881:LDT

[1493] Anonymous. Long distance telephone system of Dr. Herz. Scientific American, 44(13):196, March 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n13/pdf/scientificamerican03261881-
196b.pdf.

Anonymous:1881:UWC

[1494] Anonymous. Utilization of the waste of cities. Scientific American, 44(13):197, March 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n13/pdf/scientificamerican03261881-
197.pdf.

Anonymous:1881:NM

[1495] Anonymous. A novel motor. Scientific American, 44(13):198, March 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n13/pdf/scientificamerican03261881-198.pdf.

Holder:1881:PMJ

[1496] C. F. Holder. The parasites of a monster jelly fish. Scientific American, 44(13):199, March 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n13/pdf/scientificamerican03261881-
199.pdf.

Anonymous:1881:NLL

[1497] Anonymous. New lime light. Scientific American, 44(13):199–201, March 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n13/pdf/scientificamerican03261881-199a.pdf.

Anonymous:1881:CRd

[1498] Anonymous. Communications received. Scientific American, 44 (13):203, March 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n13/pdf/scientificamerican03261881-
203.pdf.

Anonymous:1881:BPg

[1499] Anonymous. Business and personal. Scientific American, 44(14):218, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n14/pdf/scientificamerican04021881-218.pdf.

Anonymous:1881:NQf

[1500] Anonymous. Notes and queries. Scientific American, 44(14):218–219, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n14/pdf/scientificamerican04021881-218a.pdf.

Anonymous:1881:BNBd

[1501] Anonymous. Books: New books and publications. Scientific American, 44(14):219, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n14/pdf/scientificamerican04021881-
219a.pdf.

Anonymous:1881:IIf

[1502] Anonymous. Index of inventions. Scientific American, 44(14):219, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n14/pdf/scientificamerican04021881-219b.pdf.

Anonymous:1881:PNPc

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n14/pdf/scientificamerican04021881-219c.pdf.

Anonymous:1881:ODS

[1504] Anonymous. Our deep sea fisheries. Scientific American, 44(14):208, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n14/pdf/scientificamerican04021881-208.pdf.

Anonymous:1881:IUS

[1505] Anonymous. Index of United States patents. Scientific American, 44(14):208, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n14/pdf/scientificamerican04021881-
208a.pdf.

Anonymous:1881:BSM

[1506] Anonymous. The barge system on the Mississippi. Scientific American, 44(14):208, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n14/pdf/scientificamerican04021881-
208b.pdf.

Anonymous:1881:LCL

[1507] Anonymous. Large craft on the Lakes. Scientific American, 44 (14):208–209, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n14/pdf/scientificamerican04021881-
208c.pdf.

Anonymous:1881:WSC

[1508] Anonymous. Water supply of cincinnati. Scientific American, 44 (14):209, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n14/pdf/scientificamerican04021881-
209.pdf.

Anonymous:1881:RCSb

[1509] Anonymous. The Reese circular saw. Scientific American, 44(14):209, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n14/pdf/scientificamerican04021881-209a.pdf.

Reynier:1881:CEM

[1510] E. Reynier. Comparative experiments made with naked and metallized carbons. Scientific American, 44(14):210, April 2, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v44/
n14/pdf/scientificamerican04021881-210.pdf.

Hopkins:1881:TER

[1511] Geo. M. Hopkins. Tyndall’s experiment on radiant heat. Scientific American, 44(14):210, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n14/pdf/scientificamerican04021881-
210a.pdf.

Anonymous:1881:EFI

[1512] Anonymous. Electrical fire indicator of M. G. Dupré. Scientific American, 44(14):210, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n14/pdf/scientificamerican04021881-
210b.pdf.

Anonymous:1881:AINb

[1513] Anonymous. American industries. — no. 69. Scientific American, 44(14):211–212, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n14/pdf/scientificamerican04021881-
211.pdf.

Anonymous:1881:AIc

[1514] Anonymous. Agricultural inventions. Scientific American, 44 (14):212–213, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n14/pdf/scientificamerican04021881-
212.pdf.

Anonymous:1881:NHS

[1515] Anonymous. New handle for soldering irons. Scientific American, 44(14):213, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n14/pdf/scientificamerican04021881-
213.pdf.

Anonymous:1881:RDRb

[1516] Anonymous. Recent decisions relating to patents. Scientific American, 44(14):213, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n14/pdf/scientificamerican04021881-
213a.pdf.

Anonymous:1881:MIh

[1517] Anonymous. Mechanical inventions. Scientific American, 44(14):213, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n14/pdf/scientificamerican04021881-213b.pdf.

Anonymous:1881:ITT

[1518] Anonymous. Improvement in telephone and telegraph lines. Scientific American, 44(14):213, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n14/pdf/scientificamerican04021881-
213c.pdf.

Anonymous:1881:NSW

[1519] Anonymous. New safety whiffletree. Scientific American, 44(14):214, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n14/pdf/scientificamerican04021881-214.pdf.

Anonymous:1881:IHR

[1520] Anonymous. Improved hydraulic ram. Scientific American, 44(14):214, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n14/pdf/scientificamerican04021881-214a.pdf.

Anonymous:1881:NFK

[1521] Anonymous. Novel fire kindler. Scientific American, 44(14):214–215, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n14/pdf/scientificamerican04021881-214b.pdf.

Anonymous:1881:RBF

[1522] Anonymous. Russian beer flagon. Scientific American, 44(14):215, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n14/pdf/scientificamerican04021881-215.pdf.

Anonymous:1881:SFF

[1523] Anonymous. Scaly-finned fishes. Scientific American, 44(14):215–216, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n14/pdf/scientificamerican04021881-215a.pdf.

Anonymous:1881:MIi

[1524] Anonymous. Miscellaneous inventions. Scientific American, 44 (14):216–217, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n14/pdf/scientificamerican04021881-
216.pdf.

Anonymous:1881:CRe

[1525] Anonymous. Communications received. Scientific American, 44 (14):219, April 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n14/pdf/scientificamerican04021881-
219.pdf.

Anonymous:1881:Cc

[1526] Anonymous. Correspondence. Scientific American, 44(15):228, April 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n15/pdf/scientificamerican04091881-228.pdf.

Anonymous:1881:INIe

[1527] Anonymous. Inventions: New inventions. Scientific American, 44(15):230, April 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n15/pdf/scientificamerican04091881-
230b.pdf.

Anonymous:1881:BPh

[1528] Anonymous. Business and personal. Scientific American, 44(15):234, April 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n15/pdf/scientificamerican04091881-234.pdf.

Anonymous:1881:NQg

[1529] Anonymous. Notes and queries. Scientific American, 44(15):234–235, April 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n15/pdf/scientificamerican04091881-234a.pdf.

Anonymous:1881:BNBe

[1530] Anonymous. Books: New books and publications. Scientific American, 44(15):235, April 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n15/pdf/scientificamerican04091881-
235a.pdf.

Anonymous:1881:IIg

[1531] Anonymous. Index of inventions. Scientific American, 44(15):235, April 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n15/pdf/scientificamerican04091881-235b.pdf.

Anonymous:1881:PNPd

[1532] Anonymous. Patent notes: Patents. Scientific American, 44(15):235, April 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n15/pdf/scientificamerican04091881-235c.pdf.

Anonymous:1881:AINc

[1533] Anonymous. American industries — no. 70. Scientific American, 44(15):223, April 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n15/pdf/scientificamerican04091881-
223.pdf.

Anonymous:1881:RID

[1534] Anonymous. Recent infringement decisions. Scientific American, 44(15):224, April 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n15/pdf/scientificamerican04091881-
224.pdf.

Anonymous:1881:SSb

[1535] Anonymous. Sea sickness. Scientific American, 44(15):224–225, April 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n15/pdf/scientificamerican04091881-224a.pdf.

Anonymous:1881:RE

[1536] Anonymous. Report of the expert. Scientific American, 44(15):225, April 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n15/pdf/scientificamerican04091881-225.pdf.

Anonymous:1881:GAM

[1537] Anonymous. A gigantic artificial moon. Scientific American, 44(15):225, April 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n15/pdf/scientificamerican04091881-225a.pdf.

Anonymous:1881:MIj

[1538] Anonymous. Mechanical inventions. Scientific American, 44(15): 226–227, April 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n15/pdf/scientificamerican04091881-
226.pdf.

Anonymous:1881:SPE

[1539] Anonymous. Stationary and portable engines. Scientific American, 44(15):227, April 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n15/pdf/scientificamerican04091881-
227.pdf.

Anonymous:1881:MIk

[1540] Anonymous. Miscellaneous inventions. Scientific American, 44 (15):227–228, April 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n15/pdf/scientificamerican04091881-
227a.pdf.

Anonymous:1881:AId

[1541] Anonymous. Agricultural inventions. Scientific American, 44(15):228, April 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n15/pdf/scientificamerican04091881-228a.pdf.

Anonymous:1881:NGT

[1542] Anonymous. New grappling tool for oil wells. Scientific American, 44(15):229, April 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n15/pdf/scientificamerican04091881-
229.pdf.

Anonymous:1881:IASa

[1543] Anonymous. Improved ash sifter. Scientific American, 44(15):229, April 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n15/pdf/scientificamerican04091881-229a.pdf.

Anonymous:1881:IDK

[1544] Anonymous. Improved drying kiln. Scientific American, 44(15):230, April 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n15/pdf/scientificamerican04091881-230.pdf.

Anonymous:1881:IIh

[1545] Anonymous. Improved incubator. Scientific American, 44(15):230, April 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n15/pdf/scientificamerican04091881-230a.pdf.

Anonymous:1881:BCa

[1546] Anonymous. Branchiopod crustaceans. Scientific American, 44(15):231, April 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n15/pdf/scientificamerican04091881-231.pdf.

Anonymous:1881:SFa

[1547] Anonymous. The sturgeon family. Scientific American, 44(15):231–233, April 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n15/pdf/scientificamerican04091881-231a.pdf.

Anonymous:1881:NFBb

[1548] Anonymous. Novel fish basket. Scientific American, 44(15):233, April 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n15/pdf/scientificamerican04091881-233.pdf.

Anonymous:1881:CRf

[1549] Anonymous. Communications received. Scientific American, 44 (15):235, April 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n15/pdf/scientificamerican04091881-
235.pdf.

Anonymous:1881:Cd

[1550] Anonymous. Correspondence. Scientific American, 44(16):244–246, April 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n16/pdf/scientificamerican04161881-244a.pdf.

Anonymous:1881:IRIe

[1551] Anonymous. Inventions: Recent inventions. Scientific American, 44(16):249, April 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n16/pdf/scientificamerican04161881-
249b.pdf.

Anonymous:1881:BPi

[1552] Anonymous. Business and personal. Scientific American, 44(16):250, April 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n16/pdf/scientificamerican04161881-250.pdf.

Anonymous:1881:BNBf

[1553] Anonymous. Books: New books and publications. Scientific American, 44(16):250, April 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n16/pdf/scientificamerican04161881-
250a.pdf.

Anonymous:1881:NQh

[1554] Anonymous. Notes and queries. Scientific American, 44(16):250–251, April 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n16/pdf/scientificamerican04161881-250b.pdf.

Anonymous:1881:IIi

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n16/pdf/scientificamerican04161881-251.pdf.

Anonymous:1881:PNPe

[1556] Anonymous. Patent notes: Patents. Scientific American, 44(16):251, April 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n16/pdf/scientificamerican04161881-251a.pdf.

Anonymous:1881:UK

[1557] Anonymous. The utilization of knowledge. Scientific American, 44(16):240, April 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n16/pdf/scientificamerican04161881-
240.pdf.

Anonymous:1881:AFC

[1558] Anonymous. The American Fish Culturists’ Convention. Scientific American, 44(16):240, April 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n16/pdf/scientificamerican04161881-
240a.pdf.

Anonymous:1881:PB

[1559] Anonymous. Pigs and babies. Scientific American, 44(16):240–241, April 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n16/pdf/scientificamerican04161881-240b.pdf.

Anonymous:1881:CHG

[1560] Anonymous. Charges for handling grain in New York Harbor. Scientific American, 44(16):241, April 16, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v44/
n16/pdf/scientificamerican04161881-241.pdf.

Anonymous:1881:IEE

[1561] Anonymous. The International Exhibition of Electricity at Paris, 1881. Scientific American, 44(16):241, April 16, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v44/
n16/pdf/scientificamerican04161881-241a.pdf.

Anonymous:1881:IWU

[1562] Anonymous. An invention wanted for utilizing flax refuse. Scientific American, 44(16):241, April 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n16/pdf/scientificamerican04161881-
241b.pdf.

Anonymous:1881:NEF

[1563] Anonymous. New embroidering frame. Scientific American, 44(16):242, April 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n16/pdf/scientificamerican04161881-242.pdf.

Anonymous:1881:NPM

[1564] Anonymous. New perforating machine. Scientific American, 44(16):242, April 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n16/pdf/scientificamerican04161881-242a.pdf.

Anonymous:1881:AINd

[1565] Anonymous. American industries. — no. 71. Scientific American, 44(16):243–244, April 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n16/pdf/scientificamerican04161881-
243.pdf.

Anonymous:1881:RFD

[1566] Anonymous. The Reese fusion disk. Scientific American, 44(16):244, April 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n16/pdf/scientificamerican04161881-244.pdf.

Anonymous:1881:VHC

[1567] Anonymous. The velocipede hand car. Scientific American, 44(16):246, April 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n16/pdf/scientificamerican04161881-246.pdf.

Anonymous:1881:IRb

[1568] Anonymous. Improved refrigerator. Scientific American, 44(16):246, April 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n16/pdf/scientificamerican04161881-246a.pdf.

Anonymous:1881:MIl

[1569] Anonymous. Miscellaneous inventions. Scientific American, 44(16):246, April 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n16/pdf/scientificamerican04161881-246b.pdf.

Hopkins:1881:SHE

[1570] Geo M. Hopkins. Simplified Holtz electrical machine. Scientific American, 44(16):247–248, April 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n16/pdf/scientificamerican04161881-
247.pdf.

Anonymous:1881:MIm

[1571] Anonymous. Mechanical inventions. Scientific American, 44(16):249, April 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n16/pdf/scientificamerican04161881-249.pdf.

Anonymous:1881:LAT

[1572] Anonymous. Lavater’s apparatus for taking silhouettes. Scientific American, 44(16):249, April 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n16/pdf/scientificamerican04161881-
249a.pdf.

Anonymous:1881:IRIf

[1573] Anonymous. Inventions: Recent inventions. Scientific American, 44(17):258, April 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n17/pdf/scientificamerican04231881-
258a.pdf.

Anonymous:1881:INIf

[1574] Anonymous. Inventions: New inventions. Scientific American, 44(17):262, April 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n17/pdf/scientificamerican04231881-
262b.pdf.

Anonymous:1881:BPj

[1575] Anonymous. Business and personal. Scientific American, 44(17):266, April 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n17/pdf/scientificamerican04231881-266.pdf.

Anonymous:1881:NQi

[1576] Anonymous. Notes and queries. Scientific American, 44(17):266–267, April 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n17/pdf/scientificamerican04231881-266a.pdf.

Anonymous:1881:BNBg

[1577] Anonymous. Books: New books and publications. Scientific American, 44(17):267, April 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n17/pdf/scientificamerican04231881-
267.pdf.

Anonymous:1881:IIj

[1578] Anonymous. Index of inventions. Scientific American, 44(17):267, April 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n17/pdf/scientificamerican04231881-267a.pdf.

Anonymous:1881:PNPf

[1579] Anonymous. Patent notes: Patents. Scientific American, 44(17):267, April 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n17/pdf/scientificamerican04231881-267b.pdf.

Anonymous:1881:CPT

[1580] Anonymous. The Chicago Police Telephone and Patrol System. Scientific American, 44(17):255, April 23, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v44/
n17/pdf/scientificamerican04231881-255.pdf.

Anonymous:1881:LAP

[1581] Anonymous. Legislating against progress. Scientific American, 44(17):256, April 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n17/pdf/scientificamerican04231881-
256.pdf.

Anonymous:1881:OCS

[1582] Anonymous. Our city sewerage. Scientific American, 44(17):256, April 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n17/pdf/scientificamerican04231881-256a.pdf.

Anonymous:1881:PPRa

[1583] Anonymous. Permanent pictures on the retina. Scientific American, 44(17):257, April 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n17/pdf/scientificamerican04231881-
257.pdf.

Hovey:1881:PCa

[1584] H. C. Hovey. Pickett’s Cave. Scientific American, 44(17):257–258, April 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n17/pdf/scientificamerican04231881-257a.pdf.

Anonymous:1881:NCCa

[1585] Anonymous. New cooling cask. Scientific American, 44(17):258, April 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n17/pdf/scientificamerican04231881-258.pdf.

Anonymous:1881:RRB

[1586] Anonymous. Richard’s registering barometer. Scientific American, 44(17):259, April 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n17/pdf/scientificamerican04231881-
259.pdf.

Anonymous:1881:RT

[1587] Anonymous. Reversible tools. Scientific American, 44(17):259, April 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n17/pdf/scientificamerican04231881-259a.pdf.

Anonymous:1881:NGC

[1588] Anonymous. New game counter. Scientific American, 44(17):259, April 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n17/pdf/scientificamerican04231881-259b.pdf.

Anonymous:1881:EIb

[1589] Anonymous. Engineering inventions. Scientific American, 44(17):259, April 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n17/pdf/scientificamerican04231881-259c.pdf.

Anonymous:1881:CTCa

[1590] Anonymous. Cane with toilet combination. Scientific American, 44(17):259–261, April 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n17/pdf/scientificamerican04231881-
259d.pdf.

Anonymous:1881:MFU

[1591] Anonymous. A mechanical finger for use in the photographing of engravings. Scientific American, 44(17):261, April 23, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v44/
n17/pdf/scientificamerican04231881-261.pdf.

Anonymous:1881:MGB

[1592] Anonymous. Machine for grinding bands on goblets. Scientific American, 44(17):262, April 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n17/pdf/scientificamerican04231881-
262.pdf.

Hopkins:1881:BFW

[1593] Geo. M. Hopkins. Bursting of fly-wheels. Scientific American, 44(17):262, April 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n17/pdf/scientificamerican04231881-
262a.pdf.

Anonymous:1881:CIS

[1594] Anonymous. A curious inhabitant of the Sargasso Sea and its nest. Scientific American, 44(17):263, April 23, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v44/
n17/pdf/scientificamerican04231881-263.pdf.

Anonymous:1881:SVa

[1595] Anonymous. Sèvres vase. Scientific American, 44(17):263–264, April 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n17/pdf/scientificamerican04231881-263a.pdf.

Anonymous:1881:OP

[1596] Anonymous. Ozoniferous plants. Scientific American, 44(17):265, April 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n17/pdf/scientificamerican04231881-265.pdf.

Anonymous:1881:BTS

[1597] Anonymous. Blackford’s trout show. Scientific American, 44(17):265, April 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n17/pdf/scientificamerican04231881-265a.pdf.

Anonymous:1881:IRIg

[1598] Anonymous. Inventions: Recent inventions. Scientific American, 44(18):276, April 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n18/pdf/scientificamerican04301881-
276a.pdf.

Anonymous:1881:Ce

[1599] Anonymous. Correspondence. Scientific American, 44(18):280, April 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n18/pdf/scientificamerican04301881-280.pdf.

Anonymous:1881:BPk

[1600] Anonymous. Business and personal. Scientific American, 44(18):282, April 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n18/pdf/scientificamerican04301881-282.pdf.

Anonymous:1881:NQj

[1601] Anonymous. Notes and queries. Scientific American, 44(18):282–283, April 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n18/pdf/scientificamerican04301881-282a.pdf.

Anonymous:1881:IIk

[1602] Anonymous. Index of inventions. Scientific American, 44(18):283, April 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n18/pdf/scientificamerican04301881-283a.pdf.

Anonymous:1881:PNPg

[1603] Anonymous. Patent notes: Patents. Scientific American, 44(18):284, April 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n18/pdf/scientificamerican04301881-284.pdf.

Anonymous:1881:DL

[1604] Anonymous. The division of labor. Scientific American, 44(18):272, April 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n18/pdf/scientificamerican04301881-272.pdf.

Anonymous:1881:SAM

[1605] Anonymous. The study of animal motions. Scientific American, 44(18):272, April 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n18/pdf/scientificamerican04301881-
272a.pdf.

Anonymous:1881:ELAa

[1606] Anonymous. The electric light in Akron, Ohio. Scientific American, 44(18):272, April 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n18/pdf/scientificamerican04301881-
272b.pdf.

Hovey:1881:MAP

[1607] H. C. Hovey. More about Pickett’s Cave. Scientific American, 44(18):273, April 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n18/pdf/scientificamerican04301881-
273.pdf.

Anonymous:1881:STU

[1608] Anonymous. Standard time in the United States. Scientific American, 44(18):273–274, April 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n18/pdf/scientificamerican04301881-
273a.pdf.

Anonymous:1881:IWMa

[1609] Anonymous. Improved water meter. Scientific American, 44(18):274, April 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n18/pdf/scientificamerican04301881-274.pdf.

Anonymous:1881:IFP

[1610] Anonymous. Improved force pump. Scientific American, 44(18):274, April 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n18/pdf/scientificamerican04301881-274a.pdf.

Anonymous:1881:MIn

[1611] Anonymous. Mechanical inventions. Scientific American, 44(18):274, April 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n18/pdf/scientificamerican04301881-274b.pdf.

Anonymous:1881:AINe

[1612] Anonymous. American industries. — no. 72. Scientific American, 44(18):275–276, April 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n18/pdf/scientificamerican04301881-
275.pdf.

Anonymous:1881:HSb

[1613] Anonymous. Hanger for shafting. Scientific American, 44(18):276, April 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n18/pdf/scientificamerican04301881-276.pdf.

Anonymous:1881:RDRc

[1614] Anonymous. Recent decisions relating to patents. Scientific American, 44(18):277, April 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n18/pdf/scientificamerican04301881-
277.pdf.

Anonymous:1881:HCS

[1615] Anonymous. Hulling cotton seed. Scientific American, 44(18):278, April 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n18/pdf/scientificamerican04301881-278.pdf.

Anonymous:1881:MIo

[1616] Anonymous. Miscellaneous inventions. Scientific American, 44 (18):278–279, April 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n18/pdf/scientificamerican04301881-
278a.pdf.

Anonymous:1881:SBC

[1617] Anonymous. The stag beetle and champion beetle. Scientific American, 44(18):279–280, April 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n18/pdf/scientificamerican04301881-
279.pdf.

Anonymous:1881:CSa

[1618] Anonymous. The colors of the stars. Scientific American, 44(18):281, April 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n18/pdf/scientificamerican04301881-281.pdf.

Anonymous:1881:RDV

[1619] Anonymous. Richly decorated vase. Scientific American, 44(18):281, April 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n18/pdf/scientificamerican04301881-281a.pdf.

Anonymous:1881:CRg

[1620] Anonymous. Communications received. Scientific American, 44 (18):283, April 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n18/pdf/scientificamerican04301881-
283.pdf.

Anonymous:1881:Cf

[1621] Anonymous. Correspondence. Scientific American, 44(19):296, May 7, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n19/pdf/scientificamerican05071881-296a.pdf.

Anonymous:1881:BPl

[1622] Anonymous. Business and personal. Scientific American, 44(19):298, May 7, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n19/pdf/scientificamerican05071881-298.pdf.

Anonymous:1881:NQk

[1623] Anonymous. Notes and queries. Scientific American, 44(19):298–299, May 7, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n19/pdf/scientificamerican05071881-298a.pdf.

Anonymous:1881:BNBh

[1624] Anonymous. Books: New books and publications. Scientific American, 44(19):299, May 7, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n19/pdf/scientificamerican05071881-
299.pdf.

Anonymous:1881:IIl

[1625] Anonymous. Index of inventions. Scientific American, 44(19):299–300, May 7, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n19/pdf/scientificamerican05071881-299a.pdf.

Anonymous:1881:PNPh

[1626] Anonymous. Patent notes: Patents. Scientific American, 44(19):300, May 7, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n19/pdf/scientificamerican05071881-300.pdf.

Anonymous:1881:CW

[1627] Anonymous. The Chicago Waterworks. Scientific American, 44(19):287, May 7, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n19/pdf/scientificamerican05071881-287.pdf.

Anonymous:1881:PP

[1628] Anonymous. “Patent perfidy”. Scientific American, 44(19):288, May 7, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n19/pdf/scientificamerican05071881-288.pdf.

Anonymous:1881:OSO

[1629] Anonymous. Ozone and the sanitology of odors. Scientific American, 44(19):288–289, May 7, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n19/pdf/scientificamerican05071881-
288a.pdf.

Anonymous:1881:PWN

[1630] Anonymous. Public works in New York City. Scientific American, 44(19):289, May 7, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n19/pdf/scientificamerican05071881-
289.pdf.

Anonymous:1881:NASa

[1631] Anonymous. The National Academy of Sciences. Scientific American, 44(19):289, May 7, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n19/pdf/scientificamerican05071881-
289a.pdf.

Anonymous:1881:DGE

[1632] Anonymous. The date of the glacial era in Eastern North America. Scientific American, 44(19):289, May 7, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v44/
n19/pdf/scientificamerican05071881-289b.pdf.

Anonymous:1881:PSD

[1633] Anonymous. Photophonic and spectrophonic discoveries. Scientific American, 44(19):289, May 7, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n19/pdf/scientificamerican05071881-
289c.pdf.

Anonymous:1881:ICFa

[1634] Anonymous. Improvement in the construction of fences and posts. Scientific American, 44(19):291–292, May 7, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v44/
n19/pdf/scientificamerican05071881-291.pdf.

Anonymous:1881:RDRd

[1635] Anonymous. Recent decisions relating to patents. Scientific American, 44(19):292–293, May 7, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n19/pdf/scientificamerican05071881-
292.pdf.

Anonymous:1881:IO

[1636] Anonymous. Improved optometer. Scientific American, 44(19):294, May 7, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n19/pdf/scientificamerican05071881-294.pdf.

Anonymous:1881:IASb

[1637] Anonymous. Improvement in animal shears. Scientific American, 44(19):294, May 7, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n19/pdf/scientificamerican05071881-
294a.pdf.

Anonymous:1881:IWE

[1638] Anonymous. Improvement in water elevators. Scientific American, 44(19):294, May 7, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n19/pdf/scientificamerican05071881-
294b.pdf.

Anonymous:1881:MIp

[1639] Anonymous. Miscellaneous inventions. Scientific American, 44 (19):294–295, May 7, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n19/pdf/scientificamerican05071881-
294c.pdf.

Anonymous:1881:MRL

[1640] Anonymous. Manufacture of real lace by machinery. Scientific American, 44(19):295–296, May 7, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n19/pdf/scientificamerican05071881-
295.pdf.

Anonymous:1881:AIe

[1641] Anonymous. Agricultural inventions. Scientific American, 44(19):296, May 7, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n19/pdf/scientificamerican05071881-296.pdf.

Segur:1881:CUM

[1642] Henry Segur. Coffee. — its uses and medicinal qualities. Scientific American, 44(19):297, May 7, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n19/pdf/scientificamerican05071881-
297.pdf.

Anonymous:1881:Cg

[1643] Anonymous. Correspondence. Scientific American, 44(2):24, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n2/pdf/scientificamerican01081881-24.pdf.

Anonymous:1881:INIg

[1644] Anonymous. Inventions: New inventions. Scientific American, 44(2):26, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n2/pdf/scientificamerican01081881-26.pdf.

Anonymous:1881:BPm

[1645] Anonymous. Business and personal. Scientific American, 44(2):27, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n2/pdf/scientificamerican01081881-27.pdf.

Anonymous:1881:NQl

[1646] Anonymous. Notes and queries. Scientific American, 44(2):27, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n2/pdf/scientificamerican01081881-27a.pdf.

Anonymous:1881:BNBi

[1647] Anonymous. Books: New books and publications. Scientific American, 44(2):27, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n2/pdf/scientificamerican01081881-
27b.pdf.

Anonymous:1881:IIm

[1648] Anonymous. Index of inventions. Scientific American, 44(2):27–28, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n2/pdf/scientificamerican01081881-27c.pdf.

Anonymous:1881:PNPi

[1649] Anonymous. Patent notes: Patents. Scientific American, 44(2):28, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n2/pdf/scientificamerican01081881-28.pdf.

Anonymous:1881:ELB

[1650] Anonymous. The electric lights on Broadway, New York. Scientific American, 44(2):16, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n2/pdf/scientificamerican01081881-
16.pdf.

Anonymous:1881:GTP

[1651] Anonymous. Groove track pavement. Scientific American, 44(2):16, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n2/pdf/scientificamerican01081881-16a.pdf.

Anonymous:1881:UP

[1652] Anonymous. An unwise physician. Scientific American, 44(2):16, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n2/pdf/scientificamerican01081881-16b.pdf.

Anonymous:1881:IEF

[1653] Anonymous. International exchange of food fish. Scientific American, 44(2):16, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n2/pdf/scientificamerican01081881-
16c.pdf.

Anonymous:1881:TD

[1654] Anonymous. Temporary deafness. Scientific American, 44(2):16–17, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n2/pdf/scientificamerican01081881-16d.pdf.

Thurston:1881:ES

[1655] R. H. Thurston. The expansion of steam. Scientific American, 44(2):17–18, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n2/pdf/scientificamerican01081881-
17.pdf.

Anonymous:1881:BH

[1656] Anonymous. Brush holder. Scientific American, 44(2):18, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n2/pdf/scientificamerican01081881-18.pdf.

Anonymous:1881:ADM

[1657] Anonymous. Apparatus for demonstrating mechanical principles. Scientific American, 44(2):18, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v44/n2/pdf/scientificamerican01081881-
18a.pdf.

Anonymous:1881:EIc

[1658] Anonymous. Engineering inventions. Scientific American, 44(2):18–19, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n2/pdf/scientificamerican01081881-18b.pdf.

Anonymous:1881:TOS

[1659] Anonymous. Tidal outlets for sewers. Scientific American, 44(2):19, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n2/pdf/scientificamerican01081881-19.pdf.

Anonymous:1881:JPV

[1660] Anonymous. Jenkins’ packing and valves. Scientific American, 44(2):19, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n2/pdf/scientificamerican01081881-19a.pdf.

Anonymous:1881:AIf

[1661] Anonymous. Agricultural inventions. Scientific American, 44(2):19, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n2/pdf/scientificamerican01081881-19b.pdf.

Hoyet:1881:ACM

[1662] H. C. Hoyet. The Albion coal mines. Scientific American, 44(2):20–21, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n2/pdf/scientificamerican01081881-20.pdf.

Anonymous:1881:AINf

[1663] Anonymous. American industries — no. 64. Scientific American, 44(2):21, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n2/pdf/scientificamerican01081881-
21.pdf.

Anonymous:1881:DRPc

[1664] Anonymous. Decisions relating to patents. Scientific American, 44(2):21, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n2/pdf/scientificamerican01081881-
21a.pdf.

Anonymous:1881:IHB

[1665] Anonymous. Improved hand and bench vise. Scientific American, 44(2):22, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n2/pdf/scientificamerican01081881-
22.pdf.

Anonymous:1881:IAV

[1666] Anonymous. Improved awning and ventilator. Scientific American, 44(2):22–23, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n2/pdf/scientificamerican01081881-
22a.pdf.

Holder:1881:SPF

[1667] C. P. Holder. A strange parasitic fish. Scientific American, 44(2):23, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n2/pdf/scientificamerican01081881-23.pdf.

Anonymous:1881:SBH

[1668] Anonymous. The sword bill humming bird. Scientific American, 44(2):23–24, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n2/pdf/scientificamerican01081881-
23a.pdf.

Anonymous:1881:MIq

[1669] Anonymous. Miscellaneous inventions. Scientific American, 44 (2):24–26, January 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n2/pdf/scientificamerican01081881-
24a.pdf.

Anonymous:1881:Ch

[1670] Anonymous. Correspondence. Scientific American, 44(20):308, May 14, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n20/pdf/scientificamerican05141881-308a.pdf.

Anonymous:1881:INIh

[1671] Anonymous. Inventions: New inventions. Scientific American, 44 (20):309, May 14, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n20/pdf/scientificamerican05141881-
309.pdf.

Anonymous:1881:IRIh

[1672] Anonymous. Inventions: Recent inventions. Scientific American, 44(20):310, May 14, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n20/pdf/scientificamerican05141881-
310b.pdf.

Anonymous:1881:BPn

[1673] Anonymous. Business and personal. Scientific American, 44(20):315, May 14, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n20/pdf/scientificamerican05141881-315.pdf.

Anonymous:1881:NQm

[1674] Anonymous. Notes and queries. Scientific American, 44(20):315, May 14, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n20/pdf/scientificamerican05141881-315a.pdf.

Anonymous:1881:IIn

[1675] Anonymous. Index of invention. Scientific American, 44(20):315–318, May 14, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n20/pdf/scientificamerican05141881-315b.pdf.

Anonymous:1881:THW

[1676] Anonymous. Tornadoes, hailstorms, and waterspouts. Scientific American, 44(20):304, May 14, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n20/pdf/scientificamerican05141881-
304.pdf.

Anonymous:1881:HLP

[1677] Anonymous. Heat, light, and power without cost. Scientific American, 44(20):304, May 14, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n20/pdf/scientificamerican05141881-
304a.pdf.

Anonymous:1881:ANM

[1678] Anonymous. Another new motor. Scientific American, 44(20):305, May 14, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n20/pdf/scientificamerican05141881-305.pdf.

Anonymous:1881:KMDa

[1679] Anonymous. The Keely motor deception. Scientific American, 44(20):305, May 14, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n20/pdf/scientificamerican05141881-
305a.pdf.

Anonymous:1881:GPMa

[1680] Anonymous. The Gamgee perpetual motion. Scientific American, 44(20):305, May 14, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n20/pdf/scientificamerican05141881-
305b.pdf.

Anonymous:1881:DMB

[1681] Anonymous. Dose-measuring bottle. Scientific American, 44(20):306, May 14, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n20/pdf/scientificamerican05141881-306.pdf.

Anonymous:1881:IPH

[1682] Anonymous. Improved post hole digger. Scientific American, 44 (20):306, May 14, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n20/pdf/scientificamerican05141881-
306a.pdf.

Anonymous:1881:IDF

[1683] Anonymous. Improved door fastener. Scientific American, 44(20):306, May 14, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n20/pdf/scientificamerican05141881-306b.pdf.

Anonymous:1881:AINg

[1684] Anonymous. American industries — no. 73. Scientific American, 44(20):307–308, May 14, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n20/pdf/scientificamerican05141881-
307.pdf.

Anonymous:1881:RDRe

[1685] Anonymous. Recent decisions relating to patents. Scientific American, 44(20):308, May 14, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n20/pdf/scientificamerican05141881-
308.pdf.

Anonymous:1881:AME

[1686] Anonymous. Araometric method for the estimation of fat in milk. Scientific American, 44(20):309, May 14, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v44/
n20/pdf/scientificamerican05141881-309a.pdf.

Anonymous:1881:NHA

[1687] Anonymous. New hot-air bath chamber. Scientific American, 44 (20):310, May 14, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n20/pdf/scientificamerican05141881-
310.pdf.

Anonymous:1881:NRC

[1688] Anonymous. New road car. Scientific American, 44(20):310, May 14, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n20/pdf/scientificamerican05141881-310a.pdf.

Anonymous:1881:ELAb

[1689] Anonymous. The electric light in an art gallery. Scientific American, 44(20):311, May 14, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n20/pdf/scientificamerican05141881-
311.pdf.

Anonymous:1881:MIr

[1690] Anonymous. Miscellaneous inventions. Scientific American, 44(20):311, May 14, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n20/pdf/scientificamerican05141881-311a.pdf.

Anonymous:1881:KJ

[1691] Anonymous. Krugg jugs. Scientific American, 44(20):311, May 14, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n20/pdf/scientificamerican05141881-311b.pdf.

Anonymous:1881:AST

[1692] Anonymous. The alligator snapping turtle. Scientific American, 44(20):311–312, May 14, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n20/pdf/scientificamerican05141881-
311c.pdf.

Anonymous:1881:BNa

[1693] Anonymous. Botanical notes. Scientific American, 44(20):312–314, May 14, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n20/pdf/scientificamerican05141881-312.pdf.

Anonymous:1881:BPo

[1694] Anonymous. Business and personal. Scientific American, 44(21):330, May 21, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n21/pdf/scientificamerican05211881-330.pdf.

Anonymous:1881:BNBj

[1695] Anonymous. Books: New books and publications. Scientific American, 44(21):330, May 21, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n21/pdf/scientificamerican05211881-
330a.pdf.

Anonymous:1881:NQn

[1696] Anonymous. Notes and queries. Scientific American, 44(21):330–331, May 21, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n21/pdf/scientificamerican05211881-330b.pdf.

Anonymous:1881:IIo

[1697] Anonymous. Index of inventions. Scientific American, 44(21):331, May 21, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n21/pdf/scientificamerican05211881-331a.pdf.

Anonymous:1881:PGSa

[1698] Anonymous. Progress of the great suspension bridge between New York and Brooklyn. Scientific American, 44(21):319–322, May 21, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n21/pdf/scientificamerican05211881-319.pdf.

Anonymous:1881:HVF

[1699] Anonymous. Humanity versus fanaticism. Scientific American, 44 (21):320, May 21, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n21/pdf/scientificamerican05211881-
320.pdf.

Anonymous:1881:IAPa

[1700] Anonymous. Intolerance among physicians. Scientific American, 44(21):320, May 21, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n21/pdf/scientificamerican05211881-
320a.pdf.

Anonymous:1881:CFa

[1701] Anonymous. Corn as fuel. Scientific American, 44(21):320–321, May 21, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n21/pdf/scientificamerican05211881-320b.pdf.

Anonymous:1881:HGR

[1702] Anonymous. How to get rid of foundlings. Scientific American, 44(21):321, May 21, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n21/pdf/scientificamerican05211881-
321.pdf.

Anonymous:1881:EPC

[1703] Anonymous. Extinguishing petroleum with chloroform. Scientific American, 44(21):321, May 21, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n21/pdf/scientificamerican05211881-
321a.pdf.

Anonymous:1881:PGP

[1704] Anonymous. Plea for a government perpetual motion. Scientific American, 44(21):321, May 21, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n21/pdf/scientificamerican05211881-
321b.pdf.

Anonymous:1881:NOS

[1705] Anonymous. Novel oil separator. Scientific American, 44(21):322, May 21, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n21/pdf/scientificamerican05211881-322.pdf.

Anonymous:1881:MIs

[1706] Anonymous. Miscellaneous inventions. Scientific American, 44 (21):322–323, May 21, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n21/pdf/scientificamerican05211881-
322a.pdf.

Anonymous:1881:NCCb

[1707] Anonymous. Novel car cooler. Scientific American, 44(21):323–324, May 21, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n21/pdf/scientificamerican05211881-323.pdf.

Anonymous:1881:GPMb

[1708] Anonymous. The Gamgee perpetual motion. Scientific American, 44(21):324–325, May 21, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n21/pdf/scientificamerican05211881-
324.pdf.

Anonymous:1881:NWM

[1709] Anonymous. Novel washing machine. Scientific American, 44(21):325, May 21, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n21/pdf/scientificamerican05211881-325.pdf.

Anonymous:1881:EId

[1710] Anonymous. Engineering inventions. Scientific American, 44(21): 325–326, May 21, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n21/pdf/scientificamerican05211881-
325a.pdf.

Anonymous:1881:NC

[1711] Anonymous. Novel candlestick. Scientific American, 44(21):326, May 21, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n21/pdf/scientificamerican05211881-326.pdf.

Anonymous:1881:NTTa

[1712] Anonymous. New telephone transmitter. Scientific American, 44 (21):326, May 21, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n21/pdf/scientificamerican05211881-
326a.pdf.

Eitel:1881:SFP

[1713] John Eitel. A simple filter pump. Scientific American, 44(21):326–327, May 21, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n21/pdf/scientificamerican05211881-326b.pdf.

Anonymous:1881:LV

[1714] Anonymous. The lemur vari. Scientific American, 44(21):327–329, May 21, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n21/pdf/scientificamerican05211881-327.pdf.

Anonymous:1881:CRh

[1715] Anonymous. Communications received. Scientific American, 44 (21):331, May 21, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n21/pdf/scientificamerican05211881-
331.pdf.

Anonymous:1881:IRIi

[1716] Anonymous. Inventions: Recent inventions. Scientific American, 44(22):342–343, May 28, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n22/pdf/scientificamerican05281881-
342b.pdf.

Anonymous:1881:INIi

[1717] Anonymous. Inventions: New inventions. Scientific American, 44 (22):346, May 28, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n22/pdf/scientificamerican05281881-
346.pdf.

Anonymous:1881:BPp

[1718] Anonymous. Business and personal. Scientific American, 44(22):347, May 28, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n22/pdf/scientificamerican05281881-347.pdf.

Anonymous:1881:NQo

[1719] Anonymous. Notes and queries. Scientific American, 44(22):347, May 28, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n22/pdf/scientificamerican05281881-347a.pdf.

Anonymous:1881:IIp

[1720] Anonymous. Index of inventions. Scientific American, 44(22):347–348, May 28, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n22/pdf/scientificamerican05281881-347b.pdf.

Anonymous:1881:SBB

[1721] Anonymous. The second bridge between New York and Brooklyn. Scientific American, 44(22):335, May 28, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v44/n22/pdf/scientificamerican05281881-
335.pdf.

Anonymous:1881:CTb

[1722] Anonymous. Comet telegraphy. Scientific American, 44(22):336, May 28, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n22/pdf/scientificamerican05281881-336.pdf.

Anonymous:1881:OER

[1723] Anonymous. Opening of the electric railway in Berlin. Scientific American, 44(22):336, May 28, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n22/pdf/scientificamerican05281881-
336a.pdf.

Anonymous:1881:Ia

[1724] Anonymous. Iridium. Scientific American, 44(22):336, May 28, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n22/pdf/scientificamerican05281881-336b.pdf.

Anonymous:1881:GPMc

[1725] Anonymous. The Gamgee perpetual motion. Scientific American, 44(22):337, May 28, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n22/pdf/scientificamerican05281881-
337.pdf.

Anonymous:1881:IPa

[1726] Anonymous. Insectivorous plants. Scientific American, 44(22):337, May 28, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n22/pdf/scientificamerican05281881-337a.pdf.

Anonymous:1881:SA

[1727] Anonymous. Silk adulterations. Scientific American, 44(22):337–338, May 28, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n22/pdf/scientificamerican05281881-337b.pdf.

Anonymous:1881:NCB

[1728] Anonymous. New can and bottle opener. Scientific American, 44(22):338, May 28, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n22/pdf/scientificamerican05281881-
338.pdf.

Anonymous:1881:NMT

[1729] Anonymous. New mole trap. Scientific American, 44(22):338, May 28, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n22/pdf/scientificamerican05281881-338a.pdf.

Anonymous:1881:IFJ

[1730] Anonymous. Improved fruit jar. Scientific American, 44(22):338, May 28, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n22/pdf/scientificamerican05281881-338b.pdf.

Anonymous:1881:A

[1731] Anonymous. Assaying. Scientific American, 44(22):339–340, May 28, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n22/pdf/scientificamerican05281881-339.pdf.

Carey:1881:T

[1732] G. R. Carey. Thermophones. Scientific American, 44(22):341, May 28, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n22/pdf/scientificamerican05281881-341.pdf.

Anonymous:1881:MIt

[1733] Anonymous. Miscellaneous inventions. Scientific American, 44(22):341, May 28, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n22/pdf/scientificamerican05281881-341a.pdf.

Anonymous:1881:NFC

[1734] Anonymous. Novel folding cradle. Scientific American, 44(22):342, May 28, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n22/pdf/scientificamerican05281881-342.pdf.

Anonymous:1881:NBH

[1735] Anonymous. Novel book holder. Scientific American, 44(22):342, May 28, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n22/pdf/scientificamerican05281881-342a.pdf.

Anonymous:1881:FIa

[1736] Anonymous. Faience of india. Scientific American, 44(22):343, May 28, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n22/pdf/scientificamerican05281881-343.pdf.

Anonymous:1881:EC

[1737] Anonymous. Elegant chairs. Scientific American, 44(22):343, May 28, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n22/pdf/scientificamerican05281881-343a.pdf.

Anonymous:1881:MIu

[1738] Anonymous. Mechanical inventions. Scientific American, 44(22): 343–345, May 28, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n22/pdf/scientificamerican05281881-
343b.pdf.

Anonymous:1881:Ci

[1739] Anonymous. Correspondence. Scientific American, 44(23):353–354, June 4, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n23/pdf/scientificamerican06041881-353.pdf.

Anonymous:1881:IRIj

[1740] Anonymous. Inventions: Recent inventions. Scientific American, 44(23):355, June 4, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n23/pdf/scientificamerican06041881-
355b.pdf.

Anonymous:1881:INIj

[1741] Anonymous. Inventions: New inventions. Scientific American, 44(23):357–358, June 4, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n23/pdf/scientificamerican06041881-
357c.pdf.

Anonymous:1881:BPq

[1742] Anonymous. Business and personal. Scientific American, 44(23):362, June 4, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n23/pdf/scientificamerican06041881-362.pdf.

Anonymous:1881:NQp

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n23/pdf/scientificamerican06041881-362a.pdf.

Anonymous:1881:IIq

[1744] Anonymous. Index of inventions. Scientific American, 44(23):363, June 4, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n23/pdf/scientificamerican06041881-363a.pdf.

Anonymous:1881:PHR

[1745] Anonymous. Progress of the Hudson River Tunnel. Scientific American, 44(23):351–356, June 4, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n23/pdf/scientificamerican06041881-
351.pdf.

Anonymous:1881:PMD

[1746] Anonymous. Perpetual motion delusions. Scientific American, 44 (23):352, June 4, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n23/pdf/scientificamerican06041881-
352.pdf.

Anonymous:1881:PSa

[1747] Anonymous. Polarization of sound. Scientific American, 44(23):352, June 4, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n23/pdf/scientificamerican06041881-352a.pdf.

Anonymous:1881:EP

[1748] Anonymous. Electrical paper. Scientific American, 44(23):352, June 4, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n23/pdf/scientificamerican06041881-352b.pdf.

Anonymous:1881:ONW

[1749] Anonymous. Opening of the new workshop of the Stevens Institute of Technology. Scientific American, 44(23):354, June 4, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n23/pdf/scientificamerican06041881-354.pdf.

Anonymous:1881:NBC

[1750] Anonymous. New belt clasp. Scientific American, 44(23):355, June 4, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n23/pdf/scientificamerican06041881-355.pdf.

Anonymous:1881:THC

[1751] Anonymous. Telegraph hand car. Scientific American, 44(23):355, June 4, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n23/pdf/scientificamerican06041881-355a.pdf.

Anonymous:1881:EIe

[1752] Anonymous. Engineering inventions. Scientific American, 44(23):356, June 4, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n23/pdf/scientificamerican06041881-356.pdf.

Anonymous:1881:RDRf

[1753] Anonymous. Recent decisions relating to patents. Scientific American, 44(23):357, June 4, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n23/pdf/scientificamerican06041881-
357.pdf.

Anonymous:1881:NGBa

[1754] Anonymous. New gang boring machine. Scientific American, 44 (23):357, June 4, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n23/pdf/scientificamerican06041881-
357a.pdf.

Anonymous:1881:NSGb

[1755] Anonymous. New spring grass shears. Scientific American, 44(23):357, June 4, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n23/pdf/scientificamerican06041881-357b.pdf.

Anonymous:1881:FWH

[1756] Anonymous. Feed water heater and feed pump. Scientific American, 44(23):358, June 4, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n23/pdf/scientificamerican06041881-
358.pdf.

Anonymous:1881:NMSa

[1757] Anonymous. New method of spacing and lettering signs. Scientific American, 44(23):358, June 4, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n23/pdf/scientificamerican06041881-
358a.pdf.

Anonymous:1881:MIv

[1758] Anonymous. Mechanical inventions. Scientific American, 44(23):358, June 4, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n23/pdf/scientificamerican06041881-358b.pdf.

Anonymous:1881:JFS

[1759] Anonymous. John Fitch’s steamboat experiment on Collect Pond. Scientific American, 44(23):359, June 4, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v44/n23/pdf/scientificamerican06041881-
359.pdf.

Roberts:1881:LF

[1760] A. W. Roberts. Leech farming. Scientific American, 44(23):359–361, June 4, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n23/pdf/scientificamerican06041881-359a.pdf.

Anonymous:1881:CRi

[1761] Anonymous. Communications received. Scientific American, 44 (23):363, June 4, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n23/pdf/scientificamerican06041881-
363.pdf.

Anonymous:1881:Cj

[1762] Anonymous. Correspondence. Scientific American, 44(24):372, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n24/pdf/scientificamerican06111881-372.pdf.

Anonymous:1881:IRIk

[1763] Anonymous. Inventions: Recent inventions. Scientific American, 44(24):373, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n24/pdf/scientificamerican06111881-
373.pdf.

Anonymous:1881:BPr

[1764] Anonymous. Business and personal. Scientific American, 44(24):378, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n24/pdf/scientificamerican06111881-378.pdf.

Anonymous:1881:NQq

[1765] Anonymous. Notes and queries. Scientific American, 44(24):378–379, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n24/pdf/scientificamerican06111881-378a.pdf.

Anonymous:1881:IIr

[1766] Anonymous. Index of inventions. Scientific American, 44(24):379–380, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n24/pdf/scientificamerican06111881-379a.pdf.

Anonymous:1881:MFF

[1767] Anonymous. Machines for finishing fabrics. Scientific American, 44(24):367, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n24/pdf/scientificamerican06111881-
367.pdf.

Anonymous:1881:PGSb

[1768] Anonymous. The preservation of gas service-pipes. Scientific American, 44(24):368, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n24/pdf/scientificamerican06111881-
368.pdf.

Anonymous:1881:WLW

[1769] Anonymous. Where the letters are written. Scientific American, 44(24):368, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n24/pdf/scientificamerican06111881-
368a.pdf.

Anonymous:1881:TMN

[1770] Anonymous. Trade mark notes. Scientific American, 44(24):368–369, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n24/pdf/scientificamerican06111881-368b.pdf.

Anonymous:1881:DCRa

[1771] Anonymous. Deafness as a cause of railway risks. Scientific American, 44(24):369, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n24/pdf/scientificamerican06111881-
369.pdf.

Anonymous:1881:DD

[1772] Anonymous. Dangers of dentistry. Scientific American, 44(24):369, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n24/pdf/scientificamerican06111881-369a.pdf.

Hovey:1881:MFM

[1773] H. C. Hovey. A mushroom farm in mammoth cave. Scientific American, 44(24):369–370, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n24/pdf/scientificamerican06111881-
369b.pdf.

Anonymous:1881:NBM

[1774] Anonymous. New bending machine. Scientific American, 44(24):370, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n24/pdf/scientificamerican06111881-370.pdf.

Anonymous:1881:NPE

[1775] Anonymous. New pumping engine. Scientific American, 44(24):370, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n24/pdf/scientificamerican06111881-370a.pdf.

Anonymous:1881:MIw

[1776] Anonymous. Miscellaneous inventions. Scientific American, 44(24):370, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n24/pdf/scientificamerican06111881-370b.pdf.

Anonymous:1881:KMDb

[1777] Anonymous. The Keely motor deception. Scientific American, 44 (24):371–372, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n24/pdf/scientificamerican06111881-
371.pdf.

Anonymous:1881:NIF

[1778] Anonymous. New iron fence. Scientific American, 44(24):373–374, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n24/pdf/scientificamerican06111881-373a.pdf.

Anonymous:1881:ICR

[1779] Anonymous. Improved connecting rod. Scientific American, 44(24):374, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n24/pdf/scientificamerican06111881-374.pdf.

Anonymous:1881:NSP

[1780] Anonymous. New style power mortiser. Scientific American, 44(24):374, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n24/pdf/scientificamerican06111881-374a.pdf.

Anonymous:1881:EIf

[1781] Anonymous. Engineering inventions. Scientific American, 44(24): 374–375, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n24/pdf/scientificamerican06111881-
374b.pdf.

Anonymous:1881:ESP

[1782] Anonymous. English soft porcelain. Scientific American, 44(24):375, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n24/pdf/scientificamerican06111881-375.pdf.

Roberts:1881:L

[1783] A. W. Roberts. Lobsters. Scientific American, 44(24):375–376, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n24/pdf/scientificamerican06111881-375a.pdf.

Anonymous:1881:OBP

[1784] Anonymous. Oxygen from bleaching powder. Scientific American, 44(24):376, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n24/pdf/scientificamerican06111881-
376.pdf.

Anonymous:1881:AIg

[1785] Anonymous. Agricultural inventions. Scientific American, 44(24):377, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n24/pdf/scientificamerican06111881-377.pdf.

Anonymous:1881:CRj

[1786] Anonymous. Communications received. Scientific American, 44 (24):379, June 11, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n24/pdf/scientificamerican06111881-
379.pdf.

Anonymous:1881:IRIl

[1787] Anonymous. Inventions: Recent inventions. Scientific American, 44(25):387, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n25/pdf/scientificamerican06181881-
387b.pdf.

Anonymous:1881:INIk

[1788] Anonymous. Inventions: New inventions. Scientific American, 44(25):393, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n25/pdf/scientificamerican06181881-
393.pdf.

Anonymous:1881:BPs

[1789] Anonymous. Business and personal. Scientific American, 44(25):394, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n25/pdf/scientificamerican06181881-394.pdf.

Anonymous:1881:NQr

[1790] Anonymous. Notes and queries. Scientific American, 44(25):394–395, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n25/pdf/scientificamerican06181881-394a.pdf.

Anonymous:1881:BNBk

[1791] Anonymous. Books: New books and publications. Scientific American, 44(25):395, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n25/pdf/scientificamerican06181881-
395.pdf.

Anonymous:1881:IIs

[1792] Anonymous. Index of inventions. Scientific American, 44(25):395–396, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n25/pdf/scientificamerican06181881-395a.pdf.

Anonymous:1881:NTSb

[1793] Anonymous. A new telephone system. Scientific American, 44(25):384, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n25/pdf/scientificamerican06181881-384.pdf.

Anonymous:1881:MPT

[1794] Anonymous. Medical patents and trade marks. Scientific American, 44(25):384, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n25/pdf/scientificamerican06181881-
384a.pdf.

Anonymous:1881:TAL

[1795] Anonymous. Testing argentiferous lead ores. Scientific American, 44(25):384–385, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n25/pdf/scientificamerican06181881-
384b.pdf.

Hovey:1881:RDD

[1796] H. C. Hovey. Rural drainage and disease. Scientific American, 44(25):385, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n25/pdf/scientificamerican06181881-
385.pdf.

Anonymous:1881:VS

[1797] Anonymous. Vaccination in smallpox. Scientific American, 44 (25):385–386, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n25/pdf/scientificamerican06181881-
385a.pdf.

Roberts:1881:FSF

[1798] A. W. Roberts. Feast of strange fish. Scientific American, 44(25):386, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n25/pdf/scientificamerican06181881-386.pdf.

Anonymous:1881:AIh

[1799] Anonymous. Agricultural inventions. Scientific American, 44 (25):386–387, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n25/pdf/scientificamerican06181881-
386a.pdf.

Anonymous:1881:DTH

[1800] Anonymous. Double tongue for harvesters. Scientific American, 44(25):387, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n25/pdf/scientificamerican06181881-
387.pdf.

Anonymous:1881:NRH

[1801] Anonymous. Novel rein holder. Scientific American, 44(25):387, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n25/pdf/scientificamerican06181881-387a.pdf.

Dolbear:1881:NST

[1802] A. E. Dolbear. A new system of telephony. Scientific American, 44(25):388–389, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n25/pdf/scientificamerican06181881-
388.pdf.

Anonymous:1881:MIx

[1803] Anonymous. Mechanical inventions. Scientific American, 44(25):389, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n25/pdf/scientificamerican06181881-389.pdf.

Cuttriss:1881:PVV

[1804] C. Cuttriss. Photographing the vibrations of the voice. Scientific American, 44(25):389, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n25/pdf/scientificamerican06181881-
389a.pdf.

Anonymous:1881:SCA

[1805] Anonymous. Silk culture in America. Scientific American, 44(25):389, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n25/pdf/scientificamerican06181881-389b.pdf.

Anonymous:1881:EIg

[1806] Anonymous. Engineering inventions. Scientific American, 44(25):390, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n25/pdf/scientificamerican06181881-390.pdf.

Anonymous:1881:IWMb

[1807] Anonymous. Improved wind motor. Scientific American, 44(25):390, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n25/pdf/scientificamerican06181881-390a.pdf.

Anonymous:1881:NBP

[1808] Anonymous. New book protector. Scientific American, 44(25):390, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n25/pdf/scientificamerican06181881-390b.pdf.

Anonymous:1881:MIy

[1809] Anonymous. Miscellaneous inventions. Scientific American, 44 (25):390–391, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n25/pdf/scientificamerican06181881-
390c.pdf.

Anonymous:1881:LF

[1810] Anonymous. The lump fish. Scientific American, 44(25):391, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n25/pdf/scientificamerican06181881-391.pdf.

Anonymous:1881:ABW

[1811] Anonymous. Amphora of bronze and white metal. Scientific American, 44(25):391–392, June 18, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n25/pdf/scientificamerican06181881-
391a.pdf.

Anonymous:1881:IRIm

[1812] Anonymous. Inventions: Recent inventions. Scientific American, 44(26):402, June 25, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n26/pdf/scientificamerican06251881-
402.pdf.

Anonymous:1881:INIl

[1813] Anonymous. Inventions: New inventions. Scientific American, 44 (26):403–404, June 25, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n26/pdf/scientificamerican06251881-
403c.pdf.

Anonymous:1881:Ck

[1814] Anonymous. Correspondence. Scientific American, 44(26):405, June 25, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n26/pdf/scientificamerican06251881-405.pdf.

Anonymous:1881:BPt

[1815] Anonymous. Business and personal. Scientific American, 44(26):409, June 25, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n26/pdf/scientificamerican06251881-409.pdf.

Anonymous:1881:NQs

[1816] Anonymous. Notes and queries. Scientific American, 44(26):409, June 25, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n26/pdf/scientificamerican06251881-409a.pdf.

Anonymous:1881:IIt

[1817] Anonymous. Index of inventions. Scientific American, 44(26):409, June 25, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n26/pdf/scientificamerican06251881-409b.pdf.

Anonymous:1881:PNPj

[1818] Anonymous. Patent notes: Patents. Scientific American, 44(26):412, June 25, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n26/pdf/scientificamerican06251881-412.pdf.

Anonymous:1881:Sb

[1819] Anonymous. Siphons. Scientific American, 44(26):399, June 25, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n26/pdf/scientificamerican06251881-399.pdf.

Anonymous:1881:ASR

[1820] Anonymous. Attractive suburban residences. Scientific American, 44(26):399, June 25, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n26/pdf/scientificamerican06251881-
399a.pdf.

Anonymous:1881:PIa

[1821] Anonymous. Printing ink. Scientific American, 44(26):400, June 25, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n26/pdf/scientificamerican06251881-400.pdf.

Anonymous:1881:PHW

[1822] Anonymous. The problem of healthy water. Scientific American, 44(26):400–401, June 25, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n26/pdf/scientificamerican06251881-
400a.pdf.

Anonymous:1881:GS

[1823] Anonymous. George Stephenson. Scientific American, 44(26):401, June 25, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n26/pdf/scientificamerican06251881-401.pdf.

Anonymous:1881:TRB

[1824] Anonymous. Two recent boiler explosions. Scientific American, 44(26):401, June 25, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n26/pdf/scientificamerican06251881-
401a.pdf.

Anonymous:1881:CSE

[1825] Anonymous. Concentrating or storing up electricity. Scientific American, 44(26):401–402, June 25, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n26/pdf/scientificamerican06251881-
401b.pdf.

Anonymous:1881:NDW

[1826] Anonymous. Novel device for watering animals. Scientific American, 44(26):403, June 25, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n26/pdf/scientificamerican06251881-
403.pdf.

Hartwell:1881:STE

[1827] S. N. Hartwell. Steam tugboat explosion. Scientific American, 44(26):403, June 25, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n26/pdf/scientificamerican06251881-
403a.pdf.

Anonymous:1881:VC

[1828] Anonymous. Velocipede carriage. Scientific American, 44(26):403, June 25, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n26/pdf/scientificamerican06251881-403b.pdf.

Anonymous:1881:MIz

[1829] Anonymous. Miscellaneous inventions. Scientific American, 44(26):404, June 25, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n26/pdf/scientificamerican06251881-404.pdf.

Anonymous:1881:MIba

[1830] Anonymous. Mechanical inventions. Scientific American, 44(26):405, June 25, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n26/pdf/scientificamerican06251881-405a.pdf.

Anonymous:1881:NFG

[1831] Anonymous. New fastener for grain-car doors. Scientific American, 44(26):406, June 25, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n26/pdf/scientificamerican06251881-
406.pdf.

Anonymous:1881:SEb

[1832] Anonymous. Storing of electricity. Scientific American, 44(26):406–407, June 25, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n26/pdf/scientificamerican06251881-406a.pdf.

Anonymous:1881:SDF

[1833] Anonymous. The slender dragon fly. Scientific American, 44 (26):407–408, June 25, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n26/pdf/scientificamerican06251881-
407.pdf.

Anonymous:1881:NHNa

[1834] Anonymous. Natural history notes. Scientific American, 44(26):408, June 25, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n26/pdf/scientificamerican06251881-408.pdf.

Anonymous:1881:AIi

[1835] Anonymous. Agricultural inventions. Scientific American, 44(26):408, June 25, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n26/pdf/scientificamerican06251881-408a.pdf.

Anonymous:1881:IRIn

[1836] Anonymous. Inventions: Recent inventions. Scientific American, 44(3):36–37, January 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n3/pdf/scientificamerican01151881-
36.pdf.

Anonymous:1881:BPu

[1837] Anonymous. Business and personal. Scientific American, 44(3):43, January 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n3/pdf/scientificamerican01151881-43.pdf.

Anonymous:1881:NQt

[1838] Anonymous. Notes and queries. Scientific American, 44(3):43, January 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n3/pdf/scientificamerican01151881-43a.pdf.

Anonymous:1881:IIu

[1839] Anonymous. Index of inventions. Scientific American, 44(3):43–44, January 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n3/pdf/scientificamerican01151881-43b.pdf.

Anonymous:1881:PNPk

[1840] Anonymous. Patent notes: Patents. Scientific American, 44(3):44, January 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n3/pdf/scientificamerican01151881-44.pdf.

Anonymous:1881:PBB

[1841] Anonymous. Progress of the Brooklyn Bridge. Scientific American, 44(3):31–32, January 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n3/pdf/scientificamerican01151881-
31.pdf.

Anonymous:1881:PBS

[1842] Anonymous. Progress of the brush system of electric lighting. Scientific American, 44(3):32, January 15, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v44/
n3/pdf/scientificamerican01151881-32.pdf.

Anonymous:1881:AH

[1843] Anonymous. On aids to hearing. Scientific American, 44(3):32–33, January 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n3/pdf/scientificamerican01151881-32a.pdf.

Anonymous:1881:HF

[1844] Anonymous. Harvest figures. Scientific American, 44(3):33, January 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n3/pdf/scientificamerican01151881-33.pdf.

Anonymous:1881:PGI

[1845] Anonymous. Professor Gamgee’s ice machine. Scientific American, 44(3):33, January 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n3/pdf/scientificamerican01151881-
33a.pdf.

Anonymous:1881:EMa

[1846] Anonymous. Electro-metallurgy. Scientific American, 44(3):33, January 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n3/pdf/scientificamerican01151881-33b.pdf.

Anonymous:1881:PFW

[1847] Anonymous. Portable Fountain water closet. Scientific American, 44(3):34, January 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n3/pdf/scientificamerican01151881-
34.pdf.

Anonymous:1881:SSAa

[1848] Anonymous. Steamboats for South American rivers. Scientific American, 44(3):34, January 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n3/pdf/scientificamerican01151881-
34a.pdf.

Anonymous:1881:ITa

[1849] Anonymous. Improved tongs. Scientific American, 44(3):34–35, January 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n3/pdf/scientificamerican01151881-34b.pdf.

Anonymous:1881:IMG

[1850] Anonymous. Improvement in machine guns. Scientific American, 44(3):35, January 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n3/pdf/scientificamerican01151881-
35.pdf.

Hovey:1881:HC

[1851] H. C. Hovey. Howe’s Cave. Scientific American, 44(3):35–36, January 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n3/pdf/scientificamerican01151881-35a.pdf.

Anonymous:1881:FHD

[1852] Anonymous. The face in health and disease. Scientific American, 44(3):37, January 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n3/pdf/scientificamerican01151881-
37.pdf.

Anonymous:1881:CPA

[1853] Anonymous. The Chicago Police Alarm System. Scientific American, 44(3):37, January 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n3/pdf/scientificamerican01151881-
37a.pdf.

Anonymous:1881:EIh (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n3/pdf/scientificamerican01151881-
37b.pdf.

Rose:1881:MEC

[1855] Joshua Rose. A mysterious explosion and its causes. Scientific American, 44(3):38, January 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n3/pdf/scientificamerican01151881-
38.pdf.

Anonymous:1881:CTM

[1856] Anonymous. Curious time measurers. Scientific American, 44 (3):39–40, January 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n3/pdf/scientificamerican01151881-
39.pdf.

Anonymous:1881:MIbb

[1857] Anonymous. Mechanical inventions. Scientific American, 44(3):40, January 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n3/pdf/scientificamerican01151881-40.pdf.

Anonymous:1881:AIj

[1858] Anonymous. Agricultural inventions. Scientific American, 44 (3):41–42, January 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n3/pdf/scientificamerican01151881-
41.pdf.

Anonymous:1881:CMI

[1859] Anonymous. Cellar for manufacturing ice. Scientific American, 44(3):42, January 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n3/pdf/scientificamerican01151881-
42.pdf.

Anonymous:1881:HFK

[1860] Anonymous. Human footprints in Kentucky sandstone. Scientific American, 44(3):42, January 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n3/pdf/scientificamerican01151881-
42a.pdf.

Anonymous:1881:MIbc

[1861] Anonymous. Miscellaneous inventions. Scientific American, 44 (3):42, January 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n3/pdf/scientificamerican01151881-
42b.pdf.

Anonymous:1881:IRIo

[1862] Anonymous. Inventions: Recent inventions. Scientific American, 44(4):51–52, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n4/pdf/scientificamerican01221881-
51c.pdf.

Anonymous:1881:NIa

[1863] Anonymous. New inventions. Scientific American, 44(4):52, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n4/pdf/scientificamerican01221881-52.pdf.

Anonymous:1881:Cl

[1864] Anonymous. Correspondence. Scientific American, 44(4):52, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n4/pdf/scientificamerican01221881-52a.pdf.

Anonymous:1881:BPv

[1865] Anonymous. Business and personal. Scientific American, 44(4):58, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n4/pdf/scientificamerican01221881-58.pdf.

Anonymous:1881:NQu

[1866] Anonymous. Notes and queries. Scientific American, 44(4):58–59, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n4/pdf/scientificamerican01221881-58a.pdf.

Anonymous:1881:BNBl

[1867] Anonymous. Books: New books and publications. Scientific American, 44(4):59, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n4/pdf/scientificamerican01221881-
59a.pdf.

Anonymous:1881:IIv

[1868] Anonymous. Index of inventions. Scientific American, 44(4):59, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n4/pdf/scientificamerican01221881-59b.pdf.

Anonymous:1881:PNPl

[1869] Anonymous. Patent notes: Patents. Scientific American, 44(4):60, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n4/pdf/scientificamerican01221881-60.pdf.

Anonymous:1881:END

[1870] Anonymous. Edison’s new dynamo-electric machine. Scientific American, 44(4):47–48, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n4/pdf/scientificamerican01221881-
47.pdf.

Anonymous:1881:ISMa

[1871] Anonymous. Improvements in the silvering of mirrors. Scientific American, 44(4):48, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n4/pdf/scientificamerican01221881-
48.pdf.

Anonymous:1881:PST

[1872] Anonymous. Pharaoh’s serpents — are they dangerous toys? Scientific American, 44(4):48, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n4/pdf/scientificamerican01221881-
48a.pdf.

Anonymous:1881:IRR

[1873] Anonymous. Ice roads and railways on ice. Scientific American, 44(4):48–49, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n4/pdf/scientificamerican01221881-
48b.pdf.

Anonymous:1881:EIM

[1874] Anonymous. Electric illumination at Menlo Park. Scientific American, 44(4):49, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n4/pdf/scientificamerican01221881-
49.pdf.

Anonymous:1881:TSR

[1875] Anonymous. The Tehuantepec ship railway. Scientific American, 44(4):49, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n4/pdf/scientificamerican01221881-
49a.pdf.

Anonymous:1881:SBF

[1876] Anonymous. Should a baby be fat? Scientific American, 44(4):49, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n4/pdf/scientificamerican01221881-49b.pdf.

Anonymous:1881:ELGa

[1877] Anonymous. Electric light good for the eyes. Scientific American, 44(4):49–50, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n4/pdf/scientificamerican01221881-
49c.pdf.

Anonymous:1881:IBP

[1878] Anonymous. Improved bench-plane. Scientific American, 44(4):50, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n4/pdf/scientificamerican01221881-50.pdf.

Anonymous:1881:MRT

[1879] Anonymous. Machine for riveting the tubes of Galloway boilers. Scientific American, 44(4):50, January 22, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v44/
n4/pdf/scientificamerican01221881-50a.pdf.

Anonymous:1881:TLC

[1880] Anonymous. Tobacco-leaf cutting knife. Scientific American, 44 (4):50, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n4/pdf/scientificamerican01221881-
50b.pdf.

Anonymous:1881:IAU

[1881] Anonymous. Improved apparatus for upward percolation. Scientific American, 44(4):50–51, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v44/n4/pdf/scientificamerican01221881-
50c.pdf.

Anonymous:1881:NSB

[1882] Anonymous. Novel steam boiler. Scientific American, 44(4):51, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n4/pdf/scientificamerican01221881-51.pdf.

Anonymous:1881:MIbd

[1883] Anonymous. Mechanical inventions. Scientific American, 44(4):51, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n4/pdf/scientificamerican01221881-51a.pdf.

Anonymous:1881:BCE

[1884] Anonymous. Bottle cock for effervescing liquids. Scientific American, 44(4):51, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n4/pdf/scientificamerican01221881-
51b.pdf.

Bertin:1881:MM

[1885] ?. Bertin and ?. Dubosq. On magic mirrors. Scientific American, 44(4):53, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n4/pdf/scientificamerican01221881-
53.pdf.

Anonymous:1881:NRD

[1886] Anonymous. New radial drill. Scientific American, 44(4):54, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n4/pdf/scientificamerican01221881-54.pdf.

Anonymous:1881:ISFa

[1887] Anonymous. Improved steam fog horn. Scientific American, 44 (4):54, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n4/pdf/scientificamerican01221881-
54a.pdf.

Anonymous:1881:MIbe

[1888] Anonymous. Miscellaneous inventions. Scientific American, 44 (4):54, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n4/pdf/scientificamerican01221881-
54b.pdf.

Holder:1881:CFN

[1889] C. F. Holder. Curious fact in natural history. Scientific American, 44(4):55, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n4/pdf/scientificamerican01221881-
55.pdf.

Anonymous:1881:FWM

[1890] Anonymous. Fresh-water medusæ. Scientific American, 44(4):55, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n4/pdf/scientificamerican01221881-55a.pdf.

Anonymous:1881:NHNb

[1891] Anonymous. Natural history notes. Scientific American, 44(4): 55–56, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n4/pdf/scientificamerican01221881-
55b.pdf.

Anonymous:1881:TAS

[1892] Anonymous. Transactions of the American Society of Engineers. Scientific American, 44(4):56, January 22, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v44/
n4/pdf/scientificamerican01221881-56.pdf.

Anonymous:1881:EIi

[1893] Anonymous. Engineering inventions. Scientific American, 44(4): 56–57, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n4/pdf/scientificamerican01221881-
56a.pdf.

Anonymous:1881:DRPd

[1894] Anonymous. Decisions relating to patents. Scientific American, 44(4):57, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n4/pdf/scientificamerican01221881-
57.pdf.

Anonymous:1881:CRk

[1895] Anonymous. Communications received. Scientific American, 44 (4):59, January 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n4/pdf/scientificamerican01221881-
59.pdf.

Anonymous:1881:IRIp

[1896] Anonymous. Inventions: Recent inventions. Scientific American, 44(5):66, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n5/pdf/scientificamerican01291881-
66b.pdf.

Anonymous:1881:Cm

[1897] Anonymous. Correspondence. Scientific American, 44(5):68, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n5/pdf/scientificamerican01291881-68.pdf.

Anonymous:1881:INIm

[1898] Anonymous. Inventions: New inventions. Scientific American, 44(5):70, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n5/pdf/scientificamerican01291881-
70c.pdf.

Anonymous:1881:BPw

[1899] Anonymous. Business and personal. Scientific American, 44(5):74, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n5/pdf/scientificamerican01291881-74.pdf.

Anonymous:1881:NQv

[1900] Anonymous. Notes and queries. Scientific American, 44(5):74–75, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n5/pdf/scientificamerican01291881-74a.pdf.

Anonymous:1881:BNBm

[1901] Anonymous. Books: New books and publications. Scientific American, 44(5):75, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n5/pdf/scientificamerican01291881-
75.pdf.

Anonymous:1881:IIw

[1902] Anonymous. Index of inventions. Scientific American, 44(5):75, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n5/pdf/scientificamerican01291881-75a.pdf.

Anonymous:1881:PNPm

[1903] Anonymous. Patent notes: Patents. Scientific American, 44(5):76, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n5/pdf/scientificamerican01291881-76.pdf.

Anonymous:1881:IMR

[1904] Anonymous. An improved mountain railway system. Scientific American, 44(5):63, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n5/pdf/scientificamerican01291881-
63.pdf.

Anonymous:1881:TRC

[1905] Anonymous. The tendency of recent court decisions with regard to reissued patents. Scientific American, 44(5):64, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n5/pdf/scientificamerican01291881-64.pdf.

Anonymous:1881:ORS

[1906] Anonymous. Official report on the steamer Anthracite. Scientific American, 44(5):64–65, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n5/pdf/scientificamerican01291881-
64a.pdf.

Anonymous:1881:T

[1907] Anonymous. Thermometers. Scientific American, 44(5):65, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n5/pdf/scientificamerican01291881-65.pdf.

Anonymous:1881:HMC

[1908] Anonymous. Hydraulic mortars and cements. Scientific American, 44(5):65, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n5/pdf/scientificamerican01291881-
65a.pdf.

Anonymous:1881:GGP

[1909] Anonymous. General Grant as President of the World’s Fair Commission. Scientific American, 44(5):65, January 29, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v44/
n5/pdf/scientificamerican01291881-65b.pdf.

Anonymous:1881:SCS

[1910] Anonymous. Solar clouds and Sun spots. Scientific American, 44(5):65, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n5/pdf/scientificamerican01291881-
65c.pdf.

Anonymous:1881:NMMa

[1911] Anonymous. New mechanical movement. Scientific American, 44(5):66, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n5/pdf/scientificamerican01291881-
66.pdf.

Anonymous:1881:IWSa

[1912] Anonymous. Improved wood screw. Scientific American, 44(5):66, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n5/pdf/scientificamerican01291881-66a.pdf.

Anonymous:1881:SML

[1913] Anonymous. Simple magic lantern. Scientific American, 44(5):67, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n5/pdf/scientificamerican01291881-67.pdf.

Anonymous:1881:ICFb

[1914] Anonymous. Improved cattle feeder. Scientific American, 44(5):67, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n5/pdf/scientificamerican01291881-67a.pdf.

Anonymous:1881:NMSb

[1915] Anonymous. New milking stool. Scientific American, 44(5):67, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n5/pdf/scientificamerican01291881-67b.pdf.

Anonymous:1881:DT

[1916] Anonymous. Dynamo-telegraphy. Scientific American, 44(5):69, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n5/pdf/scientificamerican01291881-69.pdf.

Anonymous:1881:IHP

[1917] Anonymous. Improved hand press. Scientific American, 44(5):70, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n5/pdf/scientificamerican01291881-70.pdf.

Anonymous:1881:IOS

[1918] Anonymous. Improved oil-stove wick trimmer. Scientific American, 44(5):70, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n5/pdf/scientificamerican01291881-
70a.pdf.

Anonymous:1881:SMT

[1919] Anonymous. Spring-making tool. Scientific American, 44(5):70, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n5/pdf/scientificamerican01291881-70b.pdf.

Anonymous:1881:CHa

[1920] Anonymous. The common hornbill. Scientific American, 44(5):71, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n5/pdf/scientificamerican01291881-71.pdf.

Beard:1881:HTI

[1921] Daniel C. Beard. Human trees of india. Scientific American, 44 (5):71–72, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n5/pdf/scientificamerican01291881-
71a.pdf.

Anonymous:1881:CGa

[1922] Anonymous. A curious grave. Scientific American, 44(5):72, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n5/pdf/scientificamerican01291881-72.pdf.

Anonymous:1881:MIbf

[1923] Anonymous. Miscellaneous inventions. Scientific American, 44 (5):72–73, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n5/pdf/scientificamerican01291881-
72a.pdf.

Anonymous:1881:DRPe

[1924] Anonymous. Decisions relating to patents. Scientific American, 44(5):73, January 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n5/pdf/scientificamerican01291881-
73.pdf.

Anonymous:1881:IRIq

[1925] Anonymous. Inventions: Recent inventions. Scientific American, 44(6):83–84, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n6/pdf/scientificamerican02051881-
83b.pdf.

Anonymous:1881:Cn

[1926] Anonymous. Correspondence. Scientific American, 44(6):84, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n6/pdf/scientificamerican02051881-84a.pdf.

Anonymous:1881:INIn

[1927] Anonymous. Inventions: New inventions. Scientific American, 44(6):86, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n6/pdf/scientificamerican02051881-
86b.pdf.

Anonymous:1881:BPx

[1928] Anonymous. Business and personal. Scientific American, 44 (6):90, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n6/pdf/scientificamerican02051881-
90.pdf.

Anonymous:1881:NQw

[1929] Anonymous. Notes and queries. Scientific American, 44(6):90– 91, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n6/pdf/scientificamerican02051881-
90a.pdf.

Anonymous:1881:BNBn

[1930] Anonymous. Books: New books and publications. Scientific American, 44(6):91, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n6/pdf/scientificamerican02051881-
91a.pdf.

Anonymous:1881:IIx

[1931] Anonymous. Index of inventions. Scientific American, 44(6):91, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n6/pdf/scientificamerican02051881-91b.pdf.

Anonymous:1881:PNPn

[1932] Anonymous. Patent notes: Patents. Scientific American, 44 (6):92, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n6/pdf/scientificamerican02051881-
92.pdf.

Anonymous:1881:SI

[1933] Anonymous. Street illumination. Scientific American, 44(6):80, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n6/pdf/scientificamerican02051881-80.pdf.

Anonymous:1881:AQC

[1934] Anonymous. Aunrecognized qualities in charcoal. Scientific American, 44(6):80, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n6/pdf/scientificamerican02051881-
80a.pdf.

Anonymous:1881:ESN

[1935] Anonymous. Electric signals for the New York Elevated Railways. Scientific American, 44(6):80, February 5, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v44/
n6/pdf/scientificamerican02051881-80b.pdf.

Anonymous:1881:GSG

[1936] Anonymous. Gas in steel and glass making. Scientific American, 44(6):80–81, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n6/pdf/scientificamerican02051881-
80c.pdf.

Anonymous:1881:TCa

[1937] Anonymous. Trees in cities. Scientific American, 44(6):81, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n6/pdf/scientificamerican02051881-81.pdf.

Anonymous:1881:SDC

[1938] Anonymous. Singular discovery in connection with phosphorescence. Scientific American, 44(6):81, February 5, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v44/
n6/pdf/scientificamerican02051881-81a.pdf.

Anonymous:1881:DCP

[1939] Anonymous. A description of Caucasian petroleum. Scientific American, 44(6):81, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n6/pdf/scientificamerican02051881-
81b.pdf.

Anonymous:1881:EMb

[1940] Anonymous. Electro-metallurgy. Scientific American, 44(6):81– 82, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n6/pdf/scientificamerican02051881-
81c.pdf.

Anonymous:1881:ILM

[1941] Anonymous. Improved lock for museum cases. Scientific American, 44(6):82, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n6/pdf/scientificamerican02051881-
82.pdf.

Anonymous:1881:BCW

[1942] Anonymous. Butter cooler and water holder. Scientific American, 44(6):82, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n6/pdf/scientificamerican02051881-
82a.pdf.

Anonymous:1881:MIbg

[1943] Anonymous. Miscellaneous inventions. Scientific American, 44 (6):82–83, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n6/pdf/scientificamerican02051881-
82b.pdf.

Anonymous:1881:MCP

[1944] Anonymous. Many-color printing press. Scientific American, 44 (6):83, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n6/pdf/scientificamerican02051881-
83.pdf.

Boudet:1881:NM

[1945] M. Boudet. New microphone. Scientific American, 44(6):83, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n6/pdf/scientificamerican02051881-83a.pdf.

Anonymous:1881:DRPf

[1946] Anonymous. Decisions relating to patents. Scientific American, 44(6):84, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n6/pdf/scientificamerican02051881-
84.pdf.

Anonymous:1881:OCP

[1947] Anonymous. The obelisk in Central Park. Scientific American, 44(6):85, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n6/pdf/scientificamerican02051881-
85.pdf.

Anonymous:1881:IBC

[1948] Anonymous. Improved bench clamp. Scientific American, 44 (6):86, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n6/pdf/scientificamerican02051881-
86.pdf.

Anonymous:1881:IET

[1949] Anonymous. Improved extension table. Scientific American, 44 (6):86, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n6/pdf/scientificamerican02051881-
86a.pdf.

Anonymous:1881:SCb

[1950] Anonymous. The sea cat. Scientific American, 44(6):87, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n6/pdf/scientificamerican02051881-87.pdf.

Anonymous:1881:JBD

[1951] Anonymous. Jersey bull Diavolo. Scientific American, 44(6):87, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n6/pdf/scientificamerican02051881-87a.pdf.

Anonymous:1881:NHNc

[1952] Anonymous. Natural history notes. Scientific American, 44(6): 87–89, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n6/pdf/scientificamerican02051881-
87b.pdf.

Anonymous:1881:CRl

[1953] Anonymous. Communications received. Scientific American, 44 (6):91, February 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n6/pdf/scientificamerican02051881-
91.pdf.

Anonymous:1881:INIo

[1954] Anonymous. Inventions: New inventions. Scientific American, 44 (7):100–101, February 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n7/pdf/scientificamerican02121881-
100a.pdf.

Anonymous:1881:IRIr

[1955] Anonymous. Inventions: Recent inventions. Scientific American, 44(7):102, February 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n7/pdf/scientificamerican02121881-
102b.pdf.

Anonymous:1881:BPy

[1956] Anonymous. Business and personal. Scientific American, 44(7): 106, February 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n7/pdf/scientificamerican02121881-
106.pdf.

Anonymous:1881:NQx

[1957] Anonymous. Notes and queries. Scientific American, 44(7):106– 107, February 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n7/pdf/scientificamerican02121881-
106a.pdf.

Anonymous:1881:BNBo

[1958] Anonymous. Books: New books and publications. Scientific American, 44(7):107, February 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n7/pdf/scientificamerican02121881-
107.pdf.

Anonymous:1881:IIy

[1959] Anonymous. Index of inventions. Scientific American, 44(7):107, February 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n7/pdf/scientificamerican02121881-107a.pdf.

Anonymous:1881:PPL

[1960] Anonymous. Progress of patent law. Scientific American, 44 (7):96, February 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n7/pdf/scientificamerican02121881-
96.pdf.

Anonymous:1881:SPFa

[1961] Anonymous. The surplus patent funds. Scientific American, 44 (7):96, February 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n7/pdf/scientificamerican02121881-
96a.pdf.

Anonymous:1881:NHJ

[1962] Anonymous. The natural history of the Jews. Scientific American, 44(7):97, February 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n7/pdf/scientificamerican02121881-
97.pdf.

Anonymous:1881:DAC

[1963] Anonymous. Diagonal avenues in cities. Scientific American, 44(7):97, February 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n7/pdf/scientificamerican02121881-
97a.pdf.

Anonymous:1881:SGM

[1964] Anonymous. Subaqueous gold mining. Scientific American, 44 (7):97, February 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n7/pdf/scientificamerican02121881-
97b.pdf.

Anonymous:1881:TIS

[1965] Anonymous. The technological, industrial, and sanitary museum of New South Wales. Scientific American, 44(7):97, February 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n7/pdf/scientificamerican02121881-97c.pdf.

Rose:1881:AMB

[1966] Joshua Rose. Another “mysterious” boiler explosion. Scientific American, 44(7):98–99, February 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n7/pdf/scientificamerican02121881-
98.pdf.

Anonymous:1881:AINh

[1967] Anonymous. American industries — no. 65. Scientific American, 44(7):99–100, February 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n7/pdf/scientificamerican02121881-
99.pdf.

Anonymous:1881:DRPg

[1968] Anonymous. Decisions relating to patents. Scientific American, 44(7):100, February 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n7/pdf/scientificamerican02121881-
100.pdf.

Anonymous:1881:CSL

[1969] Anonymous. Continuous-slide lantern. Scientific American, 44 (7):101, February 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n7/pdf/scientificamerican02121881-
101.pdf.

Anonymous:1881:MIbh

[1970] Anonymous. Miscellaneous inventions. Scientific American, 44 (7):101, February 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n7/pdf/scientificamerican02121881-
101a.pdf.

Anonymous:1881:IAB

[1971] Anonymous. Improved air brake. Scientific American, 44(7):102, February 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n7/pdf/scientificamerican02121881-102.pdf.

Anonymous:1881:LSC

[1972] Anonymous. Locomotive steam crane. Scientific American, 44 (7):102, February 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n7/pdf/scientificamerican02121881-
102a.pdf.

Anonymous:1881:HC

[1973] Anonymous. Helmet crests. Scientific American, 44(7):103–104, February 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n7/pdf/scientificamerican02121881-103.pdf.

Anonymous:1881:NET

[1974] Anonymous. The Navies of Europe — ten years’ progress in ships of war. Scientific American, 44(7):104–105, February 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n7/pdf/scientificamerican02121881-104.pdf.

Anonymous:1881:AIk (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n7/pdf/scientificamerican02121881-
105.pdf.

Anonymous:1881:IRIs

[1976] Anonymous. Inventions: Recent inventions. Scientific American, 44 (8):114–115, February 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n8/pdf/scientificamerican02191881-
114.pdf.

Anonymous:1881:Co

[1977] Anonymous. Correspondence. Scientific American, 44(8):116– 117, February 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n8/pdf/scientificamerican02191881-
116a.pdf.

Anonymous:1881:BPz

[1978] Anonymous. Business and personal. Scientific American, 44(8): 122, February 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n8/pdf/scientificamerican02191881-
122.pdf.

Anonymous:1881:NQy

[1979] Anonymous. Notes and queries. Scientific American, 44(8):122– 123, February 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n8/pdf/scientificamerican02191881-
122a.pdf.

Anonymous:1881:IIz

[1980] Anonymous. Index of inventions. Scientific American, 44(8):123, February 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n8/pdf/scientificamerican02191881-123a.pdf.

Anonymous:1881:AINi

[1981] Anonymous. American industries — no. 66. Scientific American, 44(8):111, February 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n8/pdf/scientificamerican02191881-
111.pdf.

Anonymous:1881:RPW

[1982] Anonymous. Rights of patentees while in the employ of the government. Scientific American, 44(8):112, February 19, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v44/
n8/pdf/scientificamerican02191881-112.pdf.

Anonymous:1881:VR

[1983] Anonymous. Vaccination and revaccination. Scientific American, 44 (8):112–113, February 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n8/pdf/scientificamerican02191881-
112a.pdf.

Anonymous:1881:PBW

[1984] Anonymous. The petroleum basins of Wyoming. Scientific American, 44(8):113, February 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n8/pdf/scientificamerican02191881-
113.pdf.

Anonymous:1881:EUW

[1985] Anonymous. Experiments with underground wires. Scientific American, 44(8):113, February 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n8/pdf/scientificamerican02191881-
113a.pdf.

Anonymous:1881:TCS

[1986] Anonymous. A telephonic controversy settled. Scientific American, 44(8):114, February 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n8/pdf/scientificamerican02191881-
113b.pdf.

Anonymous:1881:ICb

[1987] Anonymous. Improved clevis. Scientific American, 44(8):115, February 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n8/pdf/scientificamerican02191881-115.pdf.

Anonymous:1881:HSL

[1988] Anonymous. The Huston self-leveling berth. Scientific American, 44(8):115, February 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n8/pdf/scientificamerican02191881-
115a.pdf.

Anonymous:1881:ISCa

[1989] Anonymous. Improved steam crane. Scientific American, 44(8): 115, February 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n8/pdf/scientificamerican02191881-
115b.pdf.

Anonymous:1881:SSG

[1990] Anonymous. A seat for shop girls. Scientific American, 44(8): 115, February 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n8/pdf/scientificamerican02191881-
115c.pdf.

Anonymous:1881:EMc

[1991] Anonymous. Electro-metallurgy. Scientific American, 44(8):116, February 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n8/pdf/scientificamerican02191881-116.pdf.

Anonymous:1881:IAR

[1992] Anonymous. Improved air refrigerating machine. Scientific American, 44(8):118, February 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n8/pdf/scientificamerican02191881-
118.pdf.

Bridwell:1881:ABS

[1993] H. L. Bridwell. American built steel steamer for the River Magdalena. Scientific American, 44(8):119, February 19, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v44/
n8/pdf/scientificamerican02191881-119.pdf.

Anonymous:1881:GWL

[1994] Anonymous. The gecko, or wall lizard. Scientific American, 44 (8):119–121, February 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n8/pdf/scientificamerican02191881-
119a.pdf.

Anonymous:1881:CRm

[1995] Anonymous. Communications received. Scientific American, 44 (8):123, February 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n8/pdf/scientificamerican02191881-
123.pdf.

Anonymous:1881:IRIt

[1996] Anonymous. Inventions: Recent inventions. Scientific American, 44 (9):131–132, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n9/pdf/scientificamerican02261881-
131d.pdf.

Anonymous:1881:INIp

[1997] Anonymous. Inventions: New inventions. Scientific American, 44(9):136, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n9/pdf/scientificamerican02261881-
136a.pdf.

Anonymous:1881:BPba

[1998] Anonymous. Business and personal. Scientific American, 44(9): 138, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n9/pdf/scientificamerican02261881-
138.pdf.

Anonymous:1881:NQz

[1999] Anonymous. Notes and queries. Scientific American, 44(9):138– 139, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n9/pdf/scientificamerican02261881-
138a.pdf.

Anonymous:1881:BNBp

[2000] Anonymous. Books: New books and publications. Scientific American, 44(9):139, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n9/pdf/scientificamerican02261881-
139a.pdf.

Anonymous:1881:IIba

[2001] Anonymous. Index of inventions. Scientific American, 44(9):139, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n9/pdf/scientificamerican02261881-139b.pdf.

Anonymous:1881:AINj

[2002] Anonymous. American industries. — no. 67. Scientific American, 44(9):127, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n9/pdf/scientificamerican02261881-
127.pdf.

Anonymous:1881:EIR

[2003] Anonymous. The encouragement of inventions — with a reservation. Scientific American, 44(9):128, February 26, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v44/
n9/pdf/scientificamerican02261881-128.pdf.

Anonymous:1881:GU

[2004] Anonymous. Is glucose unwholesome? Scientific American, 44 (9):128, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n9/pdf/scientificamerican02261881-
128a.pdf.

Anonymous:1881:WI

[2005] Anonymous. Writing ink. Scientific American, 44(9):128–129, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n9/pdf/scientificamerican02261881-128b.pdf.

Anonymous:1881:PDP

[2006] Anonymous. Proposed digest of patents. Scientific American, 44(9):129, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n9/pdf/scientificamerican02261881-
129.pdf.

Anonymous:1881:SSM

[2007] Anonymous. The Sun spot maximum. Scientific American, 44 (9):129, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n9/pdf/scientificamerican02261881-
129a.pdf.

Anonymous:1881:AND

[2008] Anonymous. Another new disinfectant. Scientific American, 44 (9):129, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n9/pdf/scientificamerican02261881-
129b.pdf.

Anonymous:1881:PHS

[2009] Anonymous. Pencil holder and scissors. Scientific American, 44 (9):131, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n9/pdf/scientificamerican02261881-
131.pdf.

Anonymous:1881:NIC

[2010] Anonymous. New ice crusher. Scientific American, 44(9):131, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n9/pdf/scientificamerican02261881-131a.pdf.

Anonymous:1881:NAM

[2011] Anonymous. New adjuster for mirror and picture frames. Scientific American, 44(9):131, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v44/n9/pdf/scientificamerican02261881-
131b.pdf.

Anonymous:1881:IHHa

[2012] Anonymous. Improved hand hoe. Scientific American, 44(9):131, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n9/pdf/scientificamerican02261881-131c.pdf.

Anonymous:1881:CSb

[2013] Anonymous. Cooking by steam. Scientific American, 44(9):132, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n9/pdf/scientificamerican02261881-132.pdf.

Anonymous:1881:CTCb

[2014] Anonymous. Convertible traveling cap and scarf. Scientific American, 44(9):132, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n9/pdf/scientificamerican02261881-
132a.pdf.

Anonymous:1881:IGC

[2015] Anonymous. International Geographical Congress. Scientific American, 44(9):132, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n9/pdf/scientificamerican02261881-
132b.pdf.

Anonymous:1881:EMd

[2016] Anonymous. Electro-metallurgy. Scientific American, 44(9):133, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n9/pdf/scientificamerican02261881-133.pdf.

Anonymous:1881:IWB

[2017] Anonymous. Improved wagon brake. Scientific American, 44 (9):134, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n9/pdf/scientificamerican02261881-
134.pdf.

Anonymous:1881:ADP

[2018] Anonymous. Apparatus for decorating pottery. Scientific American, 44(9):134, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n9/pdf/scientificamerican02261881-
134a.pdf.

Anonymous:1881:MIbi

[2019] Anonymous. Miscellaneous inventions. Scientific American, 44 (9):134, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n9/pdf/scientificamerican02261881-
134b.pdf.

Anonymous:1881:WB

[2020] Anonymous. Weaver birds. Scientific American, 44(9):135–136, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n9/pdf/scientificamerican02261881-135.pdf.

Anonymous:1881:BNc

[2021] Anonymous. Botanical notes. Scientific American, 44(9):136, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n9/pdf/scientificamerican02261881-136.pdf.

Anonymous:1881:MIbj

[2022] Anonymous. Mechanical inventions. Scientific American, 44(9): 137, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n9/pdf/scientificamerican02261881-
137.pdf.

Anonymous:1881:ELGb

[2023] Anonymous. Experiment on the Lake of Geneva to ascertain the velocity of sound in water. Scientific American, 44(9):137, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v44/n9/pdf/scientificamerican02261881-137a.pdf.

Anonymous:1881:GSN

[2024] Anonymous. Geological survey of New Jersey. Scientific American, 44(9):137, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n9/pdf/scientificamerican02261881-
137b.pdf.

Anonymous:1881:CRn

[2025] Anonymous. Communications received. Scientific American, 44 (9):139, February 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v44/n9/pdf/scientificamerican02261881-
139.pdf.

Anonymous:1881:INIq

[2026] Anonymous. Inventions: New inventions. Scientific American, 45(1):5– 6, July 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n1/pdf/scientificamerican07021881-5a.pdf.

Anonymous:1881:Cp

[2027] Anonymous. Correspondence. Scientific American, 45(1):8, July 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n1/pdf/scientificamerican07021881-8.pdf.

Anonymous:1881:BPbb

[2028] Anonymous. Business and personal. Scientific American, 45(1):11, July 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n1/pdf/scientificamerican07021881-11.pdf.

Anonymous:1881:NQba

[2029] Anonymous. Notes and queries. Scientific American, 45(1):11, July 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n1/pdf/scientificamerican07021881-11a.pdf.

Anonymous:1881:IIbb

[2030] Anonymous. Index of inventions. Scientific American, 45(1):11–12, July 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n1/pdf/scientificamerican07021881-11b.pdf.

Anonymous:1881:GCF

[2031] Anonymous. Garbage cremating furnace. Scientific American, 45(1):1– 2, July 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n1/pdf/scientificamerican07021881-1.pdf.

Anonymous:1881:BEN

[2032] Anonymous. Boiler explosion notes. Scientific American, 45(1):2, July 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n1/pdf/scientificamerican07021881-2.pdf.

Anonymous:1881:PE

[2033] Anonymous. The preservation of eggs. Scientific American, 45(1):3, July 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n1/pdf/scientificamerican07021881-3.pdf.

Anonymous:1881:LEO

[2034] Anonymous. Loan exhibition of the Ohio Mechanics’ Institute. Scientific American, 45(1):3, July 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n1/pdf/scientificamerican07021881-
3a.pdf.

Anonymous:1881:JRE

[2035] Anonymous. The Jeannette Relief Expedition. Scientific American, 45(1):3, July 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n1/pdf/scientificamerican07021881-
3b.pdf.

Anonymous:1881:NSO

[2036] Anonymous. New strainer for the outlet of tubs and basins. Scientific American, 45(1):4, July 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n1/pdf/scientificamerican07021881-
4.pdf.

Anonymous:1881:NCRa

[2037] Anonymous. Novel chart rack. Scientific American, 45(1):4, July 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n1/pdf/scientificamerican07021881-4a.pdf.

Anonymous:1881:EIj

[2038] Anonymous. Engineering inventions. Scientific American, 45(1):4, July 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n1/pdf/scientificamerican07021881-4b.pdf.

Anonymous:1881:IHC

[2039] Anonymous. An improvement in horse collars. Scientific American, 45(1):4, July 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n1/pdf/scientificamerican07021881-
4c.pdf.

Anonymous:1881:ISBa

[2040] Anonymous. Improvement in steam boiler furnaces. Scientific American, 45(1):5, July 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n1/pdf/scientificamerican07021881-
5.pdf.

Anonymous:1881:EIS

[2041] Anonymous. Electric induction by stress. Scientific American, 45(1):6, July 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n1/pdf/scientificamerican07021881-6.pdf.

Anonymous:1881:RI

[2042] Anonymous. Recent inventions. Scientific American, 45(1):6, July 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n1/pdf/scientificamerican07021881-6a.pdf.

Anonymous:1881:NRSa

[2043] Anonymous. New road scraper. Scientific American, 45(1):6–7, July 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n1/pdf/scientificamerican07021881-6b.pdf.

Anonymous:1881:PU

[2044] Anonymous. Porcelain urn. Scientific American, 45(1):7, July 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n1/pdf/scientificamerican07021881-7.pdf.

Anonymous:1881:AIl

[2045] Anonymous. Agricultural inventions. Scientific American, 45(1):8, July 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n1/pdf/scientificamerican07021881-8a.pdf.

Anonymous:1881:PPSa

[2046] Anonymous. The prospects and present state of photography in natural colors. Scientific American, 45(1):9, July 2, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v45/
n1/pdf/scientificamerican07021881-9.pdf.

Anonymous:1881:MIbk

[2047] Anonymous. Miscellaneous inventions. Scientific American, 45(1):9–10, July 2, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n1/pdf/scientificamerican07021881-9a.pdf.

Anonymous:1881:Cq

[2048] Anonymous. Correspondence. Scientific American, 45(11):164, September 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n11/pdf/scientificamerican09101881-164.pdf.

Anonymous:1881:BPbc

[2049] Anonymous. Business and personal. Scientific American, 45 (11):171, September 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n11/pdf/scientificamerican09101881-
171.pdf.

Anonymous:1881:NQbb

[2050] Anonymous. Notes and queries. Scientific American, 45(11): 171, September 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n11/pdf/scientificamerican09101881-
171a.pdf.

Anonymous:1881:IIbc

[2051] Anonymous. Index of inventions. Scientific American, 45(11): 171–172, September 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n11/pdf/scientificamerican09101881-
171b.pdf.

Anonymous:1881:PNPo 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n11/pdf/scientificamerican09101881-
173.pdf.

Anonymous:1881:PEE

[2053] Anonymous. The Paris Electrical Exhibition. Scientific American, 45(11):159, September 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n11/pdf/scientificamerican09101881-
159.pdf.

Anonymous:1881:BSSa

[2054] Anonymous. Boot and shoe sewing-machine patents. Scientific American, 45(11):160, September 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n11/pdf/scientificamerican09101881-
160.pdf.

Anonymous:1881:SVS

[2055] Anonymous. Steam versus sails in oyster culture. Scientific American, 45(11):160, September 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n11/pdf/scientificamerican09101881-
160a.pdf.

Anonymous:1881:FFa

[2056] Anonymous. The fishing fish. Scientific American, 45(11):160– 161, September 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n11/pdf/scientificamerican09101881-
160b.pdf.

Anonymous:1881:TFS

[2057] Anonymous. Testing full-sized bridge columns. Scientific American, 45(11):161, September 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n11/pdf/scientificamerican09101881-
161.pdf.

Anonymous:1881:ASA

[2058] Anonymous. The American Science Association. Scientific American, 45(11):161, September 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n11/pdf/scientificamerican09101881-
161a.pdf.

Anonymous:1881:SBNa

[2059] Anonymous. Steam-boiler notes. Scientific American, 45(11): 161–162, September 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n11/pdf/scientificamerican09101881-
161b.pdf.

Anonymous:1881:WBEb

[2060] Anonymous. Watertown boiler explosion. Scientific American, 45 (11):162–163, September 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n11/pdf/scientificamerican09101881-
162.pdf.

Anonymous:1881:IHHb

[2061] Anonymous. Improved horse hay-fork. Scientific American, 45 (11):163, September 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n11/pdf/scientificamerican09101881-
163.pdf.

Anonymous:1881:ICCb

[2062] Anonymous. Improvement in casting car wheels. Scientific American, 45(11):163–164, September 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n11/pdf/scientificamerican09101881-
163a.pdf.

Anonymous:1881:TPP

[2063] Anonymous. Tin-plating processes. Scientific American, 45(11): 165, September 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n11/pdf/scientificamerican09101881-
165.pdf.

Anonymous:1881:AIM

[2064] Anonymous. Artificial indigo in Meister, Lucius & Bruning’s exhibition of dyestuffs. Scientific American, 45(11):165, September 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n11/pdf/scientificamerican09101881-165a.pdf.

Anonymous:1881:IEG

[2065] Anonymous. Improved electric generator. Scientific American, 45 (11):166, September 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n11/pdf/scientificamerican09101881-
166.pdf.

Anonymous:1881:IFC

[2066] Anonymous. Improved freight car. Scientific American, 45(11): 166, September 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n11/pdf/scientificamerican09101881-
166a.pdf.

Anonymous:1881:MIbl

[2067] Anonymous. Mechanical inventions. Scientific American, 45(11): 166–167, September 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n11/pdf/scientificamerican09101881-
166b.pdf.

Roberts:1881:OE

[2068] A. W. Roberts. Operculums and eyestones. Scientific American, 45(11):167, September 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n11/pdf/scientificamerican09101881-
167.pdf.

Holder:1881:PC

[2069] C. F. Holder. Peculiarities of the cephalopoda. Scientific American, 45(11):167–169, September 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n11/pdf/scientificamerican09101881-
167a.pdf.

Anonymous:1881:MIbm

[2070] Anonymous. Miscellaneous inventions. Scientific American, 45 (11):170, September 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n11/pdf/scientificamerican09101881-
170.pdf.

Anonymous:1881:INIr

[2071] Anonymous. Inventions: New inventions. Scientific American, 45 (13):198–199, September 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n13/pdf/scientificamerican09241881-
198c.pdf.

Anonymous:1881:IRIu

[2072] Anonymous. Inventions: Recent inventions. Scientific American, 45(13):201, September 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n13/pdf/scientificamerican09241881-
201.pdf.

Anonymous:1881:BPbd

[2073] Anonymous. Business and personal. Scientific American, 45 (13):202, September 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n13/pdf/scientificamerican09241881-
202.pdf.

Anonymous:1881:NQbc

[2074] Anonymous. Notes and queries. Scientific American, 45(13): 202–203, September 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n13/pdf/scientificamerican09241881-
202a.pdf.

Anonymous:1881:IIbd

[2075] Anonymous. Index of inventions. Scientific American, 45(13): 203, September 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n13/pdf/scientificamerican09241881-
203.pdf.

Anonymous:1881:PNPp

[2076] Anonymous. Patent notes: Patents. Scientific American, 45 (13):204, September 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n13/pdf/scientificamerican09241881-
204.pdf.

Anonymous:1881:SH

[2077] Anonymous. Storage of heat. Scientific American, 45(13):192, September 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n13/pdf/scientificamerican09241881-192.pdf.

Anonymous:1881:OGI

[2078] Anonymous. Opening of Great Industrial Exhibitions, Boston. Scientific American, 45(13):192–193, September 24, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v45/
n13/pdf/scientificamerican09241881-192a.pdf.

Anonymous:1881:PPO

[2079] Anonymous. Promotions in the Patent Office. Scientific American, 45(13):193, September 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n13/pdf/scientificamerican09241881-
193.pdf.

Anonymous:1881:ROR

[2080] Anonymous. Reviving old rejected cases. Scientific American, 45 (13):193, September 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n13/pdf/scientificamerican09241881-
193a.pdf.

Anonymous:1881:SBNb

[2081] Anonymous. Steam-boiler notes. Scientific American, 45(13): 193, September 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n13/pdf/scientificamerican09241881-
193b.pdf.

Anonymous:1881:DHD

[2082] Anonymous. Dr. Heath’s discoveries in South America. Scientific American, 45(13):193, September 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n13/pdf/scientificamerican09241881-
193c.pdf.

Anonymous:1881:IPb

[2083] Anonymous. Instantaneous photography. Scientific American, 45 (13):194, September 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n13/pdf/scientificamerican09241881-
194.pdf.

Anonymous:1881:MIbn

[2084] Anonymous. Mechanical inventions. Scientific American, 45(13): 194, September 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n13/pdf/scientificamerican09241881-
194a.pdf.

Anonymous:1881:EFM

[2085] Anonymous. Electric flying machine. Scientific American, 45(13): 195–196, September 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n13/pdf/scientificamerican09241881-
195.pdf.

Anonymous:1881:CSW

[2086] Anonymous. De Combette’s solar watch. Scientific American, 45 (13):196, September 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n13/pdf/scientificamerican09241881-
196.pdf.

Anonymous:1881:SRT

[2087] Anonymous. Self-registering target. Scientific American, 45(13): 196, September 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n13/pdf/scientificamerican09241881-
196a.pdf.

Anonymous:1881:DFN

[2088] Anonymous. A deadly fly — new case of myiasis observed in the Argentine Republic. Scientific American, 45(13):196–197, September 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n13/pdf/scientificamerican09241881-196b.pdf.

Anonymous:1881:RFR

[2089] Anonymous. Removal of Flood Rock, New York City. Scientific American, 45(13):197–198, September 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v45/n13/pdf/scientificamerican09241881-
197.pdf.

Anonymous:1881:SAS

[2090] Anonymous. Spinning attachment for sewing machines. Scientific American, 45(13):198, September 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v45/n13/pdf/scientificamerican09241881-
198.pdf.

Anonymous:1881:IBa

[2091] Anonymous. An improved battery. Scientific American, 45(13): 198, September 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n13/pdf/scientificamerican09241881-
198a.pdf.

Anonymous:1881:ITb

[2092] Anonymous. Improvement in telephones. Scientific American, 45 (13):198, September 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n13/pdf/scientificamerican09241881-
198b.pdf.

Anonymous:1881:SU

[2093] Anonymous. Sevres urn. Scientific American, 45(13):199–201, September 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n13/pdf/scientificamerican09241881-199.pdf.

Anonymous:1881:IRIv

[2094] Anonymous. Inventions: Recent inventions. Scientific American, 45(14):210, October 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n14/pdf/scientificamerican10011881-
210b.pdf.

Anonymous:1881:INIs

[2095] Anonymous. Inventions: New inventions. Scientific American, 45(14):214, October 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n14/pdf/scientificamerican10011881-
214a.pdf.

Anonymous:1881:BPbe

[2096] Anonymous. Business and personal. Scientific American, 45(14):218, October 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n14/pdf/scientificamerican10011881-218.pdf.

Anonymous:1881:NQbd

[2097] Anonymous. Notes and queries. Scientific American, 45(14):218, October 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n14/pdf/scientificamerican10011881-218a.pdf.

Anonymous:1881:BNBq

[2098] Anonymous. Books: New books and publications. Scientific American, 45(14):219, October 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n14/pdf/scientificamerican10011881-
219.pdf.

Anonymous:1881:IIbe

[2099] Anonymous. Index of inventions. Scientific American, 45(14):219, October 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n14/pdf/scientificamerican10011881-219a.pdf.

Anonymous:1881:PNPq

[2100] Anonymous. Patent notes: Patents. Scientific American, 45(14):219, October 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n14/pdf/scientificamerican10011881-219b.pdf.

Anonymous:1881:FFF

[2101] Anonymous. Field and forest fires. Scientific American, 45(14):208, October 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n14/pdf/scientificamerican10011881-208.pdf.

Anonymous:1881:MN

[2102] Anonymous. Metallurgy of nickel. Scientific American, 45(14):208, October 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n14/pdf/scientificamerican10011881-208a.pdf.

Anonymous:1881:GVN

[2103] Anonymous. The germinal value of new truths. Scientific American, 45(14):208–209, October 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n14/pdf/scientificamerican10011881-
208b.pdf.

Anonymous:1881:BED

[2104] Anonymous. Boiler explosion on a dry dock. Scientific American, 45(14):209–210, October 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n14/pdf/scientificamerican10011881-
209.pdf.

Anonymous:1881:EMCa

[2105] Anonymous. An easily made chair. Scientific American, 45(14):210, October 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n14/pdf/scientificamerican10011881-210.pdf.

Anonymous:1881:IFS

[2106] Anonymous. Improved flower stand. Scientific American, 45(14):210, October 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n14/pdf/scientificamerican10011881-210a.pdf.

Anonymous:1881:CFO

[2107] Anonymous. Cotton and its future — an opportunity for invention. Scientific American, 45(14):211–212, October 1, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v45/
n14/pdf/scientificamerican10011881-211.pdf.

Anonymous:1881:PSS

[2108] Anonymous. Preventing ships from sinking. Scientific American, 45(14):212, October 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n14/pdf/scientificamerican10011881-
212.pdf.

Anonymous:1881:SBNc

[2109] Anonymous. Steam boiler notes. Scientific American, 45(14):213, October 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n14/pdf/scientificamerican10011881-213.pdf.

Anonymous:1881:AMb

[2110] Anonymous. Ajax metal. Scientific American, 45(14):213–214, October 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n14/pdf/scientificamerican10011881-213a.pdf.

Anonymous:1881:IDA

[2111] Anonymous. Improved double-acting steam pump. Scientific American, 45(14):214, October 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n14/pdf/scientificamerican10011881-
214.pdf.

Anonymous:1881:W

[2112] Anonymous. Woodpeckers. Scientific American, 45(14):215–216, October 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n14/pdf/scientificamerican10011881-215.pdf.

Anonymous:1881:MIbo

[2113] Anonymous. Mechanical inventions. Scientific American, 45(14): 216–218, October 1, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n14/pdf/scientificamerican10011881-
216.pdf.

Anonymous:1881:IRIw 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n15/pdf/scientificamerican10081881-
230a.pdf.

Anonymous:1881:NIb

[2115] Anonymous. New inventions. Scientific American, 45(15):231, October 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n15/pdf/scientificamerican10081881-231.pdf.

Anonymous:1881:BPbf

[2116] Anonymous. Business and personal. Scientific American, 45(15):234, October 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n15/pdf/scientificamerican10081881-234.pdf.

Anonymous:1881:NQbe

[2117] Anonymous. Notes and queries. Scientific American, 45(15):234–235, October 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n15/pdf/scientificamerican10081881-234a.pdf.

Anonymous:1881:BNBr

[2118] Anonymous. Books: New books and publications. Scientific American, 45(15):235, October 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n15/pdf/scientificamerican10081881-
235.pdf.

Anonymous:1881:IIbf

[2119] Anonymous. Index of inventions. Scientific American, 45(15):235, October 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n15/pdf/scientificamerican10081881-235a.pdf.

Anonymous:1881:PNPr

[2120] Anonymous. Patent notes: Patents. Scientific American, 45(15):236, October 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n15/pdf/scientificamerican10081881-236.pdf.

Anonymous:1881:SP

[2121] Anonymous. The storage of petroleum. Scientific American, 45(15):224, October 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n15/pdf/scientificamerican10081881-224.pdf.

Anonymous:1881:PGF

[2122] Anonymous. President Garfield’s fatal wound. Scientific American, 45(15):224, October 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n15/pdf/scientificamerican10081881-
224a.pdf.

Anonymous:1881:HSS

[2123] Anonymous. Heat, Sun storms, and yellow light. Scientific American, 45(15):224–225, October 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n15/pdf/scientificamerican10081881-
224b.pdf.

Hovey:1881:TMC

[2124] H. C. Hovey. The temperature of Mammoth Cave. Scientific American, 45(15):225, October 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n15/pdf/scientificamerican10081881-
225.pdf.

Anonymous:1881:MIbp

[2125] Anonymous. Mechanical inventions. Scientific American, 45(15):226, October 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n15/pdf/scientificamerican10081881-226.pdf.

Anonymous:1881:ISE

[2126] Anonymous. Improved steam engine piston. Scientific American, 45(15):226, October 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n15/pdf/scientificamerican10081881-
226a.pdf.

Anonymous:1881:CEL

[2127] Anonymous. Cut-off for electric lamps. Scientific American, 45 (15):226–227, October 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n15/pdf/scientificamerican10081881-
226b.pdf.

Anonymous:1881:LEH

[2128] Anonymous. Lime-extracting heater and filter. Scientific American, 45(15):227, October 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n15/pdf/scientificamerican10081881-
227.pdf.

Anonymous:1881:ICTa

[2129] Anonymous. Improved combination tool. Scientific American, 45(15):227, October 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n15/pdf/scientificamerican10081881-
227a.pdf.

Anonymous:1881:AINk

[2130] Anonymous. American industries — no. 76. Scientific American, 45(15):228, October 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n15/pdf/scientificamerican10081881-
228.pdf.

Anonymous:1881:PTU

[2131] Anonymous. Potatoes and their utilization. Scientific American, 45(15):229, October 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n15/pdf/scientificamerican10081881-
229.pdf.

Anonymous:1881:AIo

[2132] Anonymous. Agricultural inventions. Scientific American, 45(15):229, October 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n15/pdf/scientificamerican10081881-229a.pdf.

Anonymous:1881:FLa

[2133] Anonymous. The Fontaine locomotive. Scientific American, 45(15):230, October 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n15/pdf/scientificamerican10081881-230.pdf.

Anonymous:1881:SFb

[2134] Anonymous. Sleeping fish. Scientific American, 45(15):231, October 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n15/pdf/scientificamerican10081881-231a.pdf.

Anonymous:1881:EIk

[2135] Anonymous. Engineering inventions. Scientific American, 45(15): 232–233, October 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n15/pdf/scientificamerican10081881-
232.pdf.

Anonymous:1881:MIbq

[2136] Anonymous. Miscellaneous inventions. Scientific American, 45(15):233, October 8, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n15/pdf/scientificamerican10081881-233.pdf.

Anonymous:1881:INIt

[2137] Anonymous. Inventions: New inventions. Scientific American, 45(16):243, October 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n16/pdf/scientificamerican10151881-
243a.pdf.

Anonymous:1881:Cu

[2138] Anonymous. Correspondence. Scientific American, 45(16):244, October 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n16/pdf/scientificamerican10151881-244.pdf.

Anonymous:1881:IRIx

[2139] Anonymous. Inventions: Recent inventions. Scientific American, 45 (16):246–247, October 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n16/pdf/scientificamerican10151881-
246a.pdf.

Anonymous:1881:BPbg

[2140] Anonymous. Business and personal. Scientific American, 45 (16):250, October 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n16/pdf/scientificamerican10151881-
250.pdf.

Anonymous:1881:NQbf

[2141] Anonymous. Notes and queries. Scientific American, 45(16): 250–251, October 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n16/pdf/scientificamerican10151881-
250a.pdf.

Anonymous:1881:IIbg

[2142] Anonymous. Index of inventions. Scientific American, 45(16):251, October 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n16/pdf/scientificamerican10151881-251a.pdf.

Anonymous:1881:PNPs

[2143] Anonymous. Patent notes: Patents. Scientific American, 45 (16):252, October 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n16/pdf/scientificamerican10151881-
252.pdf.

Anonymous:1881:RCPa

[2144] Anonymous. The Raub central power locomotive. Scientific American, 45(16):239, October 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n16/pdf/scientificamerican10151881-
239.pdf.

Anonymous:1881:FIb

[2145] Anonymous. A field for invention. Scientific American, 45(16):240, October 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n16/pdf/scientificamerican10151881-240.pdf.

Anonymous:1881:MITa

[2146] Anonymous. The moral influence of the telegraph. Scientific American, 45(16):240, October 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n16/pdf/scientificamerican10151881-
240a.pdf.

Anonymous:1881:LMG

[2147] Anonymous. The lesson of Mr. Garfield’s youth. Scientific American, 45(16):240–241, October 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n16/pdf/scientificamerican10151881-
240b.pdf.

Anonymous:1881:HWM

[2148] Anonymous. The Holy Well at Mecca. Scientific American, 45 (16):241, October 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n16/pdf/scientificamerican10151881-
241.pdf.

Anonymous:1881:SBNd

[2149] Anonymous. Steam boiler notes. Scientific American, 45(16):241, October 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n16/pdf/scientificamerican10151881-241a.pdf.

Anonymous:1881:EMe

[2150] Anonymous. Electro-metallurgy. Scientific American, 45(16):241– 242, October 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n16/pdf/scientificamerican10151881-
241b.pdf.

Anonymous:1881:NCCc

[2151] Anonymous. New cloth-cutting machine. Scientific American, 45 (16):242, October 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n16/pdf/scientificamerican10151881-
242.pdf.

Anonymous:1881:NAR

[2152] Anonymous. The new Astronomer Royal. Scientific American, 45(16):242, October 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n16/pdf/scientificamerican10151881-
242a.pdf.

Anonymous:1881:MIbr

[2153] Anonymous. Miscellaneous inventions. Scientific American, 45 (16):242, October 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n16/pdf/scientificamerican10151881-
242b.pdf.

Anonymous:1881:PIJ

[2154] Anonymous. Patented improvements in jewelry. Scientific American, 45(16):243, October 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n16/pdf/scientificamerican10151881-
243.pdf.

Anonymous:1881:NTP

[2155] Anonymous. New ticket punch. Scientific American, 45(16): 243–244, October 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n16/pdf/scientificamerican10151881-
243b.pdf.

Anonymous:1881:SEN

[2156] Anonymous. Steam engine notes. Scientific American, 45(16):244, October 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n16/pdf/scientificamerican10151881-244a.pdf.

Anonymous:1881:NDE

[2157] Anonymous. A new dynamo-electric machine. Scientific American, 45(16):246, October 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n16/pdf/scientificamerican10151881-
246.pdf.

Anonymous:1881:ED

[2158] Anonymous. The eider ducks. Scientific American, 45(16):247–249, October 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n16/pdf/scientificamerican10151881-247.pdf.

Anonymous:1881:CRo

[2159] Anonymous. Communications received. Scientific American, 45 (16):251, October 15, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n16/pdf/scientificamerican10151881-
251.pdf.

Anonymous:1881:INIu

[2160] Anonymous. Inventions: New inventions. Scientific American, 45 (17):258–259, October 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n17/pdf/scientificamerican10221881-
258b.pdf.

Anonymous:1881:IRIy

[2161] Anonymous. Inventions: Recent inventions. Scientific American, 45 (17):262–263, October 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n17/pdf/scientificamerican10221881-
262c.pdf.

Anonymous:1881:BPbh

[2162] Anonymous. Business and personal. Scientific American, 45 (17):267, October 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n17/pdf/scientificamerican10221881-
267.pdf.

Anonymous:1881:NQbg

[2163] Anonymous. Notes and queries. Scientific American, 45(17):267, October 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n17/pdf/scientificamerican10221881-267a.pdf.

Anonymous:1881:IIbh

[2164] Anonymous. Index of inventions. Scientific American, 45(17): 267–268, October 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n17/pdf/scientificamerican10221881-
267b.pdf.

Anonymous:1881:PNPt

[2165] Anonymous. Patent notes: Patents. Scientific American, 45 (17):268, October 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n17/pdf/scientificamerican10221881-
268.pdf.

Anonymous:1881:AAT

[2166] Anonymous. American art in tools. Scientific American, 45(17):256, October 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n17/pdf/scientificamerican10221881-256.pdf.

Anonymous:1881:TCW

[2167] Anonymous. A tea cultivator wanted. Scientific American, 45 (17):256, October 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n17/pdf/scientificamerican10221881-
256a.pdf.

Anonymous:1881:SBNe

[2168] Anonymous. Steam boiler notes. Scientific American, 45(17):257, October 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n17/pdf/scientificamerican10221881-257.pdf.

Anonymous:1881:M

[2169] Anonymous. Mica. Scientific American, 45(17):257, October 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n17/pdf/scientificamerican10221881-257a.pdf.

Anonymous:1881:NGA

[2170] Anonymous. New graduating apparatus. Scientific American, 45 (17):258, October 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n17/pdf/scientificamerican10221881-
258.pdf.

Anonymous:1881:ITc

[2171] Anonymous. Improvement in tables. Scientific American, 45 (17):258, October 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n17/pdf/scientificamerican10221881-
258a.pdf.

Anonymous:1881:JPB

[2172] Anonymous. Jointed pitch-board for squares. Scientific American, 45(17):259, October 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n17/pdf/scientificamerican10221881-
259.pdf.

Anonymous:1881:IDP

[2173] Anonymous. Improved drawing pen. Scientific American, 45 (17):259, October 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n17/pdf/scientificamerican10221881-
259a.pdf.

Anonymous:1881:TJL

[2174] Anonymous. The telephone. — Judge Lowell’s opinions commented upon. — ingenious transformations. Scientific American, 45(17):260, October 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n17/pdf/scientificamerican10221881-
260.pdf.

Anonymous:1881:MIbs

[2175] Anonymous. Mechanical inventions. Scientific American, 45(17): 261, October 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n17/pdf/scientificamerican10221881-
261.pdf.

Anonymous:1881:NCCd

[2176] Anonymous. New car coupling. Scientific American, 45(17):262, October 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n17/pdf/scientificamerican10221881-262.pdf.

Anonymous:1881:IMM

[2177] Anonymous. Improved mortising machine. Scientific American, 45(17):262, October 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n17/pdf/scientificamerican10221881-
262a.pdf.

Anonymous:1881:ICCc

[2178] Anonymous. Improved curtain-cord tightener. Scientific American, 45(17):262, October 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n17/pdf/scientificamerican10221881-
262b.pdf.

Anonymous:1881:VBP

[2179] Anonymous. Velocipede boat in the Public Garden, Boston. Scientific American, 45(17):263–264, October 22, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v45/
n17/pdf/scientificamerican10221881-263.pdf.

Anonymous:1881:AINl

[2180] Anonymous. American industries. — no. 77. Scientific American, 45(17):264–266, October 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n17/pdf/scientificamerican10221881-
264.pdf.

Anonymous:1881:MIbt

[2181] Anonymous. Miscellaneous inventions. Scientific American, 45 (17):266, October 22, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n17/pdf/scientificamerican10221881-
266.pdf.

Anonymous:1881:IRIz

[2182] Anonymous. Inventions: Recent inventions. Scientific American, 45(18):274, October 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n18/pdf/scientificamerican10291881-
274.pdf.

Anonymous:1881:INIv

[2183] Anonymous. Inventions: New inventions. Scientific American, 45(18):278, October 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n18/pdf/scientificamerican10291881-
278.pdf.

Anonymous:1881:BPbi

[2184] Anonymous. Business and personal. Scientific American, 45 (18):283, October 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n18/pdf/scientificamerican10291881-
283.pdf.

Anonymous:1881:NQbh

[2185] Anonymous. Notes and queries. Scientific American, 45(18):283, October 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n18/pdf/scientificamerican10291881-283a.pdf.

Anonymous:1881:IIbi

[2186] Anonymous. Index of inventions. Scientific American, 45(18): 283–284, October 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n18/pdf/scientificamerican10291881-
283b.pdf.

Anonymous:1881:IPc

[2187] Anonymous. Inventions: Patents. Scientific American, 45(18):284, October 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n18/pdf/scientificamerican10291881-284.pdf.

Anonymous:1881:PRa

[2188] Anonymous. Patent reform. Scientific American, 45(18):272, October 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n18/pdf/scientificamerican10291881-272.pdf.

Anonymous:1881:HTa

[2189] Anonymous. Hamilton E. Towle. Scientific American, 45(18):272, October 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n18/pdf/scientificamerican10291881-272a.pdf.

Anonymous:1881:CPF

[2190] Anonymous. The cotton plant and its fiber. Scientific American, 45(18):273–274, October 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n18/pdf/scientificamerican10291881-
273.pdf.

Anonymous:1881:IIBj

[2191] Anonymous. Improved invalid bed. Scientific American, 45(18):274, October 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n18/pdf/scientificamerican10291881-274a.pdf.

Anonymous:1881:IMT

[2192] Anonymous. Improvement in the manufacture of tweezers. Scientific American, 45(18):274, October 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v45/n18/pdf/scientificamerican10291881-
274b.pdf.

Hovey:1881:NSC

[2193] H. C. Hovey. The new Scylla and Charybdis. Scientific American, 45(18):275–276, October 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n18/pdf/scientificamerican10291881-
275.pdf.

Anonymous:1881:AINm

[2194] Anonymous. American industries — no. 78. Scientific American, 45(18):276–277, October 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n18/pdf/scientificamerican10291881-
276.pdf.

Anonymous:1881:SBNf

[2195] Anonymous. Steam boiler notes. Scientific American, 45(18):277, October 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n18/pdf/scientificamerican10291881-277.pdf.

Anonymous:1881:IWC

[2196] Anonymous. Improved window chair. Scientific American, 45 (18):277, October 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n18/pdf/scientificamerican10291881-
277a.pdf.

Anonymous:1881:IG

[2197] Anonymous. Improved governor. Scientific American, 45(18):278, October 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n18/pdf/scientificamerican10291881-278a.pdf.

Anonymous:1881:NDAb

[2198] Anonymous. New diving apparatus. Scientific American, 45(18): 278–279, October 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n18/pdf/scientificamerican10291881-
278b.pdf.

Anonymous:1881:BSSb

[2199] Anonymous. Breakfast set of sevres porcelain. Scientific American, 45(18):279, October 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n18/pdf/scientificamerican10291881-
279.pdf.

Anonymous:1881:BNd

[2200] Anonymous. Botanical notes. Scientific American, 45(18):280, October 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n18/pdf/scientificamerican10291881-280.pdf.

Anonymous:1881:MIbu

[2201] Anonymous. Miscellaneous inventions. Scientific American, 45 (18):280–281, October 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n18/pdf/scientificamerican10291881-
280a.pdf.

Anonymous:1881:GEL

[2202] Anonymous. The gastroscope — an electric light for the human stomach. Scientific American, 45(18):281–282, October 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n18/pdf/scientificamerican10291881-281.pdf.

Anonymous:1881:AIp

[2203] Anonymous. Agricultural inventions. Scientific American, 45 (18):282, October 29, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n18/pdf/scientificamerican10291881-
282.pdf.

Anonymous:1881:INIw

[2204] Anonymous. Inventions: New inventions. Scientific American, 45 (19):291, November 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n19/pdf/scientificamerican11051881-
291.pdf.

Anonymous:1881:BPbj

[2205] Anonymous. Business and personal. Scientific American, 45 (19):298, November 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n19/pdf/scientificamerican11051881-
298.pdf.

Anonymous:1881:NQbi

[2206] Anonymous. Notes and queries. Scientific American, 45(19): 298–299, November 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n19/pdf/scientificamerican11051881-
298a.pdf.

Anonymous:1881:IIbk

[2207] Anonymous. Index of inventions. Scientific American, 45(19): 299–300, November 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n19/pdf/scientificamerican11051881-
299.pdf.

Anonymous:1881:PNPu

[2208] Anonymous. Patent notes: Patents. Scientific American, 45 (19):300, November 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n19/pdf/scientificamerican11051881-
300.pdf.

Anonymous:1881:OR

[2209] Anonymous. The Ohio River. Scientific American, 45(19):287, November 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n19/pdf/scientificamerican11051881-287.pdf.

Anonymous:1881:FLb

[2210] Anonymous. The Fontaine locomotive. Scientific American, 45 (19):288, November 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n19/pdf/scientificamerican11051881-
288.pdf.

Anonymous:1881:CBa

[2211] Anonymous. Casualties in blasting. Scientific American, 45(19): 288–289, November 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n19/pdf/scientificamerican11051881-
288a.pdf.

Anonymous:1881:ON

[2212] Anonymous. Oyster notes. Scientific American, 45(19):289, November 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n19/pdf/scientificamerican11051881-289.pdf.

Anonymous:1881:SBNg

[2213] Anonymous. Steam boiler notes. Scientific American, 45(19): 289, November 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n19/pdf/scientificamerican11051881-
289a.pdf.

Anonymous:1881:CAP

[2214] Anonymous. Combined anvil, punch, and shears. Scientific American, 45(19):290, November 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n19/pdf/scientificamerican11051881-
290.pdf.

Anonymous:1881:IK

[2215] Anonymous. Improved kaleidoscope. Scientific American, 45(19): 290, November 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n19/pdf/scientificamerican11051881-
290a.pdf.

Anonymous:1881:MIbv

[2216] Anonymous. Mechanical inventions. Scientific American, 45(19): 290, November 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n19/pdf/scientificamerican11051881-
290b.pdf.

Anonymous:1881:NDO

[2217] Anonymous. New device for operating the tool rests of lathes. Scientific American, 45(19):291, November 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v45/n19/pdf/scientificamerican11051881-
291a.pdf.

Anonymous:1881:NDAc

[2218] Anonymous. New draught appliance for stoves. Scientific American, 45(19):291, November 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n19/pdf/scientificamerican11051881-
291b.pdf.

Anonymous:1881:MIbw

[2219] Anonymous. Miscellaneous inventions. Scientific American, 45 (19):292, November 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n19/pdf/scientificamerican11051881-
292.pdf.

Anonymous:1881:TMD

[2220] Anonymous. Trade mark decisions by the Commissioner of Patents. Scientific American, 45(19):292–293, November 5, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v45/
n19/pdf/scientificamerican11051881-292a.pdf.

Anonymous:1881:SSC

[2221] Anonymous. Safety self-cocking arms. Scientific American, 45 (19):294, November 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n19/pdf/scientificamerican11051881-
294.pdf.

Anonymous:1881:NEC

[2222] Anonymous. New engine cut-off. Scientific American, 45(19): 294–295, November 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n19/pdf/scientificamerican11051881-
294a.pdf.

Anonymous:1881:ECD

[2223] Anonymous. Electric clock-dial mechanism. Scientific American, 45(19):295, November 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n19/pdf/scientificamerican11051881-
295.pdf.

Anonymous:1881:IHD

[2224] Anonymous. Improved hatchway doors. Scientific American, 45 (19):295–296, November 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n19/pdf/scientificamerican11051881-
295a.pdf.

Anonymous:1881:EIl

[2225] Anonymous. Engineering inventions. Scientific American, 45(19): 296, November 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n19/pdf/scientificamerican11051881-
296.pdf.

Anonymous:1881:Cv

[2226] Anonymous. The condor. Scientific American, 45(19):297, November 5, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n19/pdf/scientificamerican11051881-297.pdf.

Anonymous:1881:Cx

[2227] Anonymous. Correspondence. Scientific American, 45(2):24, July 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n2/pdf/scientificamerican07091881-24.pdf.

Anonymous:1881:INIx

[2228] Anonymous. Inventions: New inventions. Scientific American, 45(2):25, July 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n2/pdf/scientificamerican07091881-25a.pdf.

Anonymous:1881:BPbk

[2229] Anonymous. Business and personal. Scientific American, 45(2):26, July 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n2/pdf/scientificamerican07091881-26.pdf.

Anonymous:1881:NQbj

[2230] Anonymous. Notes and queries. Scientific American, 45(2):26–27, July 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n2/pdf/scientificamerican07091881-26a.pdf.

Anonymous:1881:BNBs

[2231] Anonymous. Books: New books and publications. Scientific American, 45(2):27, July 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n2/pdf/scientificamerican07091881-
27a.pdf.

Anonymous:1881:IIbl

[2232] Anonymous. Index of inventions. Scientific American, 45(2):27, July 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n2/pdf/scientificamerican07091881-27b.pdf.

Anonymous:1881:PNPv

[2233] Anonymous. Patent notes: Patents. Scientific American, 45(2):29, July 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n2/pdf/scientificamerican07091881-29.pdf.

Anonymous:1881:SLA

[2234] Anonymous. State laws affecting patents. Scientific American, 45(2):16, July 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n2/pdf/scientificamerican07091881-16.pdf.

Anonymous:1881:WG

[2235] Anonymous. Water glass. Scientific American, 45(2):16–17, July 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n2/pdf/scientificamerican07091881-16a.pdf.

Anonymous:1881:SCc

[2236] Anonymous. The Stephenson Centenary. Scientific American, 45(2):17, July 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n2/pdf/scientificamerican07091881-17.pdf.

Anonymous:1881:GCN

[2237] Anonymous. The Great Comet now in sight. Scientific American, 45(2): 17, July 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n2/pdf/scientificamerican07091881-17a.pdf.

Anonymous:1881:NCRb

[2238] Anonymous. Novel cattle ringer. Scientific American, 45(2):18, July 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n2/pdf/scientificamerican07091881-18.pdf.

Anonymous:1881:IFW

[2239] Anonymous. Improved feed-water regulator. Scientific American, 45(2): 18, July 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n2/pdf/scientificamerican07091881-18a.pdf.

Anonymous:1881:MIbx

[2240] Anonymous. Mechanical inventions. Scientific American, 45(2):18–19, July 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n2/pdf/scientificamerican07091881-18b.pdf.

Anonymous:1881:IWSb

[2241] Anonymous. Improved wagon spring. Scientific American, 45(2):19, July 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n2/pdf/scientificamerican07091881-19.pdf.

Anonymous:1881:IAPb

[2242] Anonymous. Improved air pump. Scientific American, 45(2):19, July 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n2/pdf/scientificamerican07091881-19a.pdf.

Hartwell:1881:EPC

[2243] S. N. Hartwell. Explosion of a plain cylinder boiler in Philadelphia. Scientific American, 45(2):20, July 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v45/n2/pdf/scientificamerican07091881-
20.pdf.

Anonymous:1881:RDRg

[2244] Anonymous. Recent decisions relating to patents. Scientific American, 45(2):20, July 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n2/pdf/scientificamerican07091881-
20a.pdf.

Anonymous:1881:IBE

[2245] Anonymous. An interesting boiler experiment. Scientific American, 45(2):21, July 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n2/pdf/scientificamerican07091881-
21.pdf.

Anonymous:1881:ICTb

[2246] Anonymous. Improved car truck. Scientific American, 45(2):22, July 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n2/pdf/scientificamerican07091881-22.pdf.

Hopkins:1881:EV

[2247] M. Hopkins. Expansion voltameter. Scientific American, 45(2):22–23, July 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n2/pdf/scientificamerican07091881-22a.pdf.

Anonymous:1881:EB

[2248] Anonymous. The electric boat. Scientific American, 45(2):23, July 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n2/pdf/scientificamerican07091881-23.pdf.

Anonymous:1881:FBS

[2249] Anonymous. The Faure battery-stored-up electricity. Scientific American, 45(2):23–24, July 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n2/pdf/scientificamerican07091881-
23a.pdf.

Anonymous:1881:PPSb

[2250] Anonymous. The prospects and present state of photo-graphy in natural colors. Scientific American, 45(2):25, July 9, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v45/
n2/pdf/scientificamerican07091881-25.pdf.

Anonymous:1881:CRp

[2251] Anonymous. Communications received. Scientific American, 45(2):27, July 9, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n2/pdf/scientificamerican07091881-27.pdf.

Anonymous:1881:IRIba

[2252] Anonymous. Inventions: Recent inventions. Scientific American, 45(20):307, November 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n20/pdf/scientificamerican11121881-
307c.pdf.

Anonymous:1881:Id

[2253] Anonymous. Inventions. Scientific American, 45(20):308, November 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n20/pdf/scientificamerican11121881-308.pdf.

Anonymous:1881:Cy

[2254] Anonymous. Correspondence. Scientific American, 45(20):309, November 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n20/pdf/scientificamerican11121881-309.pdf.

Anonymous:1881:INIy

[2255] Anonymous. Inventions: New inventions. Scientific American, 45 (20):310, November 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n20/pdf/scientificamerican11121881-
310b.pdf.

Anonymous:1881:BPbl

[2256] Anonymous. Business and personal. Scientific American, 45 (20):315, November 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n20/pdf/scientificamerican11121881-
315.pdf.

Anonymous:1881:NQbk

[2257] Anonymous. Notes and queries. Scientific American, 45(20): 315, November 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n20/pdf/scientificamerican11121881-
315a.pdf.

Anonymous:1881:IIbm

[2258] Anonymous. Index of inventions. Scientific American, 45(20): 315–316, November 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n20/pdf/scientificamerican11121881-
315c.pdf.

Anonymous:1881:PNPw

[2259] Anonymous. Patent notes: Patents. Scientific American, 45 (20):316, November 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n20/pdf/scientificamerican11121881-
316.pdf.

Anonymous:1881:OGC

[2260] Anonymous. Our grain crop and its commercial importance. Scientific American, 45(20):303, November 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v45/n20/pdf/scientificamerican11121881-
303.pdf.

Anonymous:1881:NLH

[2261] Anonymous. The new Lyman–Haskell cannon. Scientific American, 45(20):304, November 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n20/pdf/scientificamerican11121881-
304.pdf.

Anonymous:1881:SCF

[2262] Anonymous. The structure of cotton fiber in relation to dyeing. Scientific American, 45(20):305, November 12, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v45/
n20/pdf/scientificamerican11121881-305.pdf.

Anonymous:1881:RME

[2263] Anonymous. Reproduction of medals, etc. Scientific American, 45(20):305, November 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n20/pdf/scientificamerican11121881-
305a.pdf.

Anonymous:1881:PPRb

[2264] Anonymous. Portable pneumatic riveting machine. Scientific American, 45(20):306, November 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n20/pdf/scientificamerican11121881-
306.pdf.

Anonymous:1881:EIn

[2265] Anonymous. Engineering inventions. Scientific American, 45(20): 307, November 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n20/pdf/scientificamerican11121881-
307.pdf.

Anonymous:1881:NTa

[2266] Anonymous. New tools. Scientific American, 45(20):307, November 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n20/pdf/scientificamerican11121881-307a.pdf.

Anonymous:1881:NOR

[2267] Anonymous. New organ reed. Scientific American, 45(20):307, November 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n20/pdf/scientificamerican11121881-307b.pdf.

Anonymous:1881:AIq

[2268] Anonymous. Agricultural inventions. Scientific American, 45 (20):309, November 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n20/pdf/scientificamerican11121881-
309a.pdf.

Anonymous:1881:TWJ

[2269] Anonymous. Telegraph wire jointer. Scientific American, 45 (20):310, November 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n20/pdf/scientificamerican11121881-
310.pdf.

Anonymous:1881:PGB

[2270] Anonymous. Plate-glass beveling machine. Scientific American, 45 (20):310, November 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n20/pdf/scientificamerican11121881-
310a.pdf.

Anonymous:1881:MIby

[2271] Anonymous. Mechanical inventions. Scientific American, 45(20): 311, November 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n20/pdf/scientificamerican11121881-
311.pdf.

Anonymous:1881:NSC

[2272] Anonymous. New shot-case and distributer. Scientific American, 45(20):311, November 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n20/pdf/scientificamerican11121881-
311a.pdf.

Anonymous:1881:IV

[2273] Anonymous. Improved vise. Scientific American, 45(20):311– 312, November 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n20/pdf/scientificamerican11121881-
311b.pdf.

Anonymous:1881:MIbz

[2274] Anonymous. Miscellaneous inventions. Scientific American, 45 (20):312, November 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n20/pdf/scientificamerican11121881-
312.pdf.

Anonymous:1881:WTN

[2275] Anonymous. Wasps and their nests. Scientific American, 45(20): 313–314, November 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n20/pdf/scientificamerican11121881-
313.pdf.

Anonymous:1881:CRq

[2276] Anonymous. Communications received. Scientific American, 45 (20):315, November 12, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n20/pdf/scientificamerican11121881-
315b.pdf.

Anonymous:1881:INIz

[2277] Anonymous. Inventions: New inventions. Scientific American, 45 (21):325, November 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n21/pdf/scientificamerican11191881-
325.pdf.

Anonymous:1881:BPbm

[2278] Anonymous. Business and personal. Scientific American, 45 (21):330, November 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n21/pdf/scientificamerican11191881-
330.pdf.

Anonymous:1881:NQbl

[2279] Anonymous. Notes and queries. Scientific American, 45(21): 330–331, November 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n21/pdf/scientificamerican11191881-
330a.pdf.

Anonymous:1881:IIbn

[2280] Anonymous. Index of inventions. Scientific American, 45(21): 331, November 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n21/pdf/scientificamerican11191881-
331a.pdf.

Anonymous:1881:PNPx

[2281] Anonymous. Patent notes: Patents. Scientific American, 45 (21):332, November 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n21/pdf/scientificamerican11191881-
332.pdf.

Anonymous:1881:DLH

[2282] Anonymous. The distribution of light and heat in New York City. Scientific American, 45(21):319–320, November 19, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v45/
n21/pdf/scientificamerican11191881-319.pdf.

Anonymous:1881:MHV

[2283] Anonymous. The Madge and her victories. Scientific American, 45(21):320, November 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n21/pdf/scientificamerican11191881-
320.pdf.

Anonymous:1881:SGT

[2284] Anonymous. The St. Gothard Tunnel. Scientific American, 45 (21):320–321, November 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n21/pdf/scientificamerican11191881-
320a.pdf.

Anonymous:1881:IIbo

[2285] Anonymous. Is insanity increasing? Scientific American, 45 (21):321, November 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n21/pdf/scientificamerican11191881-
321.pdf.

Anonymous:1881:RIW

[2286] Anonymous. Railroad inventions wanted. Scientific American, 45 (21):321, November 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n21/pdf/scientificamerican11191881-
321a.pdf.

Anonymous:1881:NCA

[2287] Anonymous. New compressed air locomotive on the elevated railway, New York. Scientific American, 45(21):321, November 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n21/pdf/scientificamerican11191881-321b.pdf.

Anonymous:1881:APC

[2288] Anonymous. Allen’s portable compression riveter. Scientific American, 45(21):322, November 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n21/pdf/scientificamerican11191881-
322.pdf.

Anonymous:1881:MIca

[2289] Anonymous. Mechanical inventions. Scientific American, 45(21): 322–323, November 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n21/pdf/scientificamerican11191881-
322a.pdf.

Anonymous:1881:IDB

[2290] Anonymous. Improved dinner bucket. Scientific American, 45 (21):323, November 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n21/pdf/scientificamerican11191881-
323.pdf.

Anonymous:1881:TIP

[2291] Anonymous. Tile for illuminating purposes. Scientific American, 45 (21):323–324, November 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n21/pdf/scientificamerican11191881-
323a.pdf.

Anonymous:1881:SBNh

[2292] Anonymous. Steam boiler notes. Scientific American, 45(21): 324, November 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n21/pdf/scientificamerican11191881-
324.pdf.

Anonymous:1881:MIcb

[2293] Anonymous. Miscellaneous inventions. Scientific American, 45 (21):324, November 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n21/pdf/scientificamerican11191881-
324a.pdf.

Anonymous:1881:EIo

[2294] Anonymous. Engineering inventions. Scientific American, 45(21): 326, November 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n21/pdf/scientificamerican11191881-
326.pdf.

Anonymous:1881:IPK

[2295] Anonymous. Improved pocket knife. Scientific American, 45 (21):326, November 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n21/pdf/scientificamerican11191881-
326a.pdf.

Anonymous:1881:IBR

[2296] Anonymous. Improved book rest. Scientific American, 45(21): 326, November 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n21/pdf/scientificamerican11191881-
326b.pdf.

Anonymous:1881:CEC

[2297] Anonymous. Centrifugal extractor for cane juice. Scientific American, 45(21):326, November 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n21/pdf/scientificamerican11191881-
326c.pdf.

Anonymous:1881:EMf

[2298] Anonymous. Electrical magic. Scientific American, 45(21):327– 329, November 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n21/pdf/scientificamerican11191881-
327.pdf.

Seiss:1881:SA

[2299] C. Few Seiss. The spotted Amblystome. Scientific American, 45 (21):329, November 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n21/pdf/scientificamerican11191881-
329.pdf.

Anonymous:1881:CRr

[2300] Anonymous. Communications received. Scientific American, 45 (21):331, November 19, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n21/pdf/scientificamerican11191881-
331.pdf.

Anonymous:1881:IRIbb

[2301] Anonymous. Inventions: Recent inventions. Scientific American, 45 (22):339–340, November 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n22/pdf/scientificamerican11261881-
339.pdf.

Anonymous:1881:Cz

[2302] Anonymous. Correspondence. Scientific American, 45(22):340– 341, November 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n22/pdf/scientificamerican11261881-
340.pdf.

Anonymous:1881:INIba

[2303] Anonymous. Inventions: New inventions. Scientific American, 45 (22):345, November 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n22/pdf/scientificamerican11261881-
345.pdf.

Anonymous:1881:BPbn

[2304] Anonymous. Business and personal. Scientific American, 45 (22):346, November 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n22/pdf/scientificamerican11261881-
346.pdf.

Anonymous:1881:NQbm

[2305] Anonymous. Notes and queries. Scientific American, 45(22): 346, November 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n22/pdf/scientificamerican11261881-
346a.pdf.

Anonymous:1881:BNBt

[2306] Anonymous. Books: New books and publications. Scientific American, 45(22):346–347, November 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n22/pdf/scientificamerican11261881-
346b.pdf.

Anonymous:1881:IIbp

[2307] Anonymous. Index of inventions. Scientific American, 45(22): 347, November 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n22/pdf/scientificamerican11261881-
347.pdf.

Anonymous:1881:PNPy

[2308] Anonymous. Patent notes: Patents. Scientific American, 45 (22):348, November 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n22/pdf/scientificamerican11261881-
348.pdf.

Anonymous:1881:MWP

[2309] Anonymous. The manufacture of wall papers. Scientific American, 45(22):335, November 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n22/pdf/scientificamerican11261881-
335.pdf.

Anonymous:1881:FLc

[2310] Anonymous. The Fontaine locomotive. Scientific American, 45 (22):336, November 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n22/pdf/scientificamerican11261881-
336.pdf.

Anonymous:1881:PCI

[2311] Anonymous. The possibilities of the cotton industry. Scientific American, 45(22):336, November 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n22/pdf/scientificamerican11261881-
336a.pdf.

Anonymous:1881:CFS

[2312] Anonymous. The cause of failure of steel boiler plates. Scientific American, 45(22):336–337, November 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v45/n22/pdf/scientificamerican11261881-
336b.pdf.

Anonymous:1881:SBNi

[2313] Anonymous. Steam boiler notes. Scientific American, 45(22): 337, November 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n22/pdf/scientificamerican11261881-
337.pdf.

Anonymous:1881:RSW

[2314] Anonymous. A rain of spider webs. Scientific American, 45 (22):337, November 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n22/pdf/scientificamerican11261881-
337a.pdf.

Anonymous:1881:EGT

[2315] Anonymous. Experiments with the government testing machine. Scientific American, 45(22):337, November 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v45/n22/pdf/scientificamerican11261881-
337b.pdf.

Anonymous:1881:PMW

[2316] Anonymous. The phantasmoscope, or magic wheel. Scientific American, 45(22):338, November 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n22/pdf/scientificamerican11261881-
338.pdf.

Anonymous:1881:MIcc

[2317] Anonymous. Mechanical inventions. Scientific American, 45(22): 341, November 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n22/pdf/scientificamerican11261881-
341.pdf.

Anonymous:1881:ADL

[2318] Anonymous. Automatic detaching and lock hook. Scientific American, 45(22):342, November 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n22/pdf/scientificamerican11261881-
342.pdf.

Anonymous:1881:ICSa

[2319] Anonymous. Improved cotton stalk cutter and puller. Scientific American, 45(22):342, November 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n22/pdf/scientificamerican11261881-
342a.pdf.

Anonymous:1881:ICG

[2320] Anonymous. Improved cushioned gun stock. Scientific American, 45 (22):342–343, November 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n22/pdf/scientificamerican11261881-
342b.pdf.

Anonymous:1881:GEA

[2321] Anonymous. The Great Exhibition at Atlanta Ga. Scientific American, 45(22):343, November 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n22/pdf/scientificamerican11261881-
343.pdf.

Anonymous:1881:EIp

[2322] Anonymous. Engineering inventions. Scientific American, 45(22): 343, November 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n22/pdf/scientificamerican11261881-
343a.pdf.

Anonymous:1881:AIs

[2323] Anonymous. Agricultural inventions. Scientific American, 45 (22):344, November 26, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n22/pdf/scientificamerican11261881-
344.pdf.

Anonymous:1881:Cbb

[2324] Anonymous. Correspondence. Scientific American, 45(23):354, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n23/pdf/scientificamerican12031881-354d.pdf.

Anonymous:1881:IRIbc

[2325] Anonymous. Inventions: Recent inventions. Scientific American, 45(23):355, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n23/pdf/scientificamerican12031881-
355b.pdf.

Anonymous:1881:BPbo

[2326] Anonymous. Business and personal. Scientific American, 45 (23):362, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n23/pdf/scientificamerican12031881-
362.pdf.

Anonymous:1881:NQbn

[2327] Anonymous. Notes and queries. Scientific American, 45(23): 362–363, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n23/pdf/scientificamerican12031881-
362a.pdf.

Anonymous:1881:BNBu

[2328] Anonymous. Books: New books and publications. Scientific American, 45(23):363, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n23/pdf/scientificamerican12031881-
363a.pdf.

Anonymous:1881:IIbq

[2329] Anonymous. Index of inventions. Scientific American, 45(23):363, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n23/pdf/scientificamerican12031881-363b.pdf.

Anonymous:1881:AMc

[2330] Anonymous. The Acme of misinstruction. Scientific American, 45(23):352, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n23/pdf/scientificamerican12031881-
352.pdf.

Anonymous:1881:WFL

[2331] Anonymous. Water fuel on a locomotive. Scientific American, 45(23):352, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n23/pdf/scientificamerican12031881-
352a.pdf.

Anonymous:1881:CCL

[2332] Anonymous. Commander Cheyne’s lectures. Scientific American, 45(23):352, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n23/pdf/scientificamerican12031881-
352b.pdf.

Anonymous:1881:RAM

[2333] Anonymous. The relation of agriculture and manufactures to population. Scientific American, 45(23):352–353, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v45/
n23/pdf/scientificamerican12031881-352c.pdf.

Anonymous:1881:NASb

[2334] Anonymous. The National Academy of Science. Scientific American, 45(23):353, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n23/pdf/scientificamerican12031881-
353.pdf.

Anonymous:1881:HKL

[2335] Anonymous. How kid leather is prepared. Scientific American, 45(23):354, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n23/pdf/scientificamerican12031881-
354.pdf.

Anonymous:1881:EIq

[2336] Anonymous. Engineering inventions. Scientific American, 45(23): 354, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n23/pdf/scientificamerican12031881-
354a.pdf.

Anonymous:1881:IPd

[2337] Anonymous. Improved pick. Scientific American, 45(23):354, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n23/pdf/scientificamerican12031881-354b.pdf.

Anonymous:1881:NG

[2338] Anonymous. New galvanometer. Scientific American, 45(23):354, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n23/pdf/scientificamerican12031881-354c.pdf.

Anonymous:1881:CLR

[2339] Anonymous. Combined lock and reversible latch. Scientific American, 45(23):355, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n23/pdf/scientificamerican12031881-
355.pdf.

Anonymous:1881:IDc

[2340] Anonymous. Improved damper. Scientific American, 45(23):355, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n23/pdf/scientificamerican12031881-355a.pdf.

Hopkins:1881:EMI

[2341] Geo. M. Hopkins. Electricity by magnetic induction. Scientific American, 45(23):356, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n23/pdf/scientificamerican12031881-
356.pdf.

Anonymous:1881:MIcd

[2342] Anonymous. Miscellaneous inventions. Scientific American, 45 (23):357, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n23/pdf/scientificamerican12031881-
357.pdf.

Anonymous:1881:SBNj

[2343] Anonymous. Steam boiler notes. Scientific American, 45(23):357, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n23/pdf/scientificamerican12031881-357a.pdf.

Anonymous:1881:NTRb

[2344] Anonymous. Novel ticket reel and receptacle. Scientific American, 45(23):358, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n23/pdf/scientificamerican12031881-
358.pdf.

Anonymous:1881:NGBb

[2345] Anonymous. Novel grate-bar. Scientific American, 45(23):358–359, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n23/pdf/scientificamerican12031881-358a.pdf.

Anonymous:1881:ITR

[2346] Anonymous. Improved twisting reel. Scientific American, 45 (23):359, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n23/pdf/scientificamerican12031881-
359.pdf.

Anonymous:1881:MIce

[2347] Anonymous. Mechanical inventions. Scientific American, 45(23): 359–360, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n23/pdf/scientificamerican12031881-
359a.pdf.

Roberts:1881:GS

[2348] A. W. Roberts. The great sunfish. Scientific American, 45(23):361, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n23/pdf/scientificamerican12031881-361.pdf.

Anonymous:1881:CRs

[2349] Anonymous. Communications received. Scientific American, 45 (23):363, December 3, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n23/pdf/scientificamerican12031881-
363.pdf.

Anonymous:1881:Cbc

[2350] Anonymous. Correspondence. Scientific American, 45(24):372, December 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n24/pdf/scientificamerican12101881-372.pdf.

Anonymous:1881:IRIbd

[2351] Anonymous. Inventions: Recent inventions. Scientific American, 45 (24):373–374, December 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n24/pdf/scientificamerican12101881-
373.pdf.

Anonymous:1881:INIbb

[2352] Anonymous. Inventions: New inventions. Scientific American, 45 (24):377, December 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n24/pdf/scientificamerican12101881-
377.pdf.

Anonymous:1881:BPbp 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n24/pdf/scientificamerican12101881-
379.pdf.

Anonymous:1881:NQbo

[2354] Anonymous. Notes and queries. Scientific American, 45(24): 379, December 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n24/pdf/scientificamerican12101881-
379a.pdf.

Anonymous:1881:IIbr

[2355] Anonymous. Index of inventions. Scientific American, 45(24): 379, December 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n24/pdf/scientificamerican12101881-
379b.pdf.

Anonymous:1881:PNPz

[2356] Anonymous. Patent notes: Patents. Scientific American, 45 (24):380, December 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n24/pdf/scientificamerican12101881-
380.pdf.

Anonymous:1881:ENS

[2357] Anonymous. Edison’s new steam: Dynamo. Scientific American, 45(24):367, December 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n24/pdf/scientificamerican12101881-
367.pdf.

Anonymous:1881:CPS

[2358] Anonymous. Can a patent be sold by sheriff for debt like other property? Scientific American, 45(24):368, December 10, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v45/
n24/pdf/scientificamerican12101881-368.pdf.

Anonymous:1881:LP

[2359] Anonymous. Luminous paint. Scientific American, 45(24):368, December 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n24/pdf/scientificamerican12101881-368a.pdf.

Anonymous:1881:Sd

[2360] Anonymous. Smallpox. Scientific American, 45(24):369, December 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n24/pdf/scientificamerican12101881-369.pdf.

Anonymous:1881:SBNk

[2361] Anonymous. Steam boiler notes. Scientific American, 45(24): 369, December 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n24/pdf/scientificamerican12101881-
369a.pdf.

Anonymous:1881:RCPb

[2362] Anonymous. Resignation of the Commissioner of Patents. Scientific American, 45(24):369, December 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v45/n24/pdf/scientificamerican12101881-
369b.pdf.

Anonymous:1881:MPI

[2363] Anonymous. The muscular power of insects. Scientific American, 45 (24):370–371, December 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n24/pdf/scientificamerican12101881-
370.pdf.

Anonymous:1881:NBW

[2364] Anonymous. Novel bottle wrapper. Scientific American, 45(24): 371, December 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n24/pdf/scientificamerican12101881-
371.pdf.

Anonymous:1881:ISI

[2365] Anonymous. Improved soldering iron. Scientific American, 45 (24):371, December 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n24/pdf/scientificamerican12101881-
371a.pdf.

Anonymous:1881:MIcf

[2366] Anonymous. Miscellaneous inventions. Scientific American, 45(24): 371–372, December 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n24/pdf/scientificamerican12101881-
371b.pdf.

Anonymous:1881:NHG

[2367] Anonymous. New husking glove. Scientific American, 45(24): 374, December 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n24/pdf/scientificamerican12101881-
374.pdf.

Anonymous:1881:NTSc

[2368] Anonymous. Novel trotting sulky. Scientific American, 45(24): 374, December 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n24/pdf/scientificamerican12101881-
374a.pdf.

Anonymous:1881:MIcg

[2369] Anonymous. Mechanical inventions. Scientific American, 45(24): 374–375, December 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n24/pdf/scientificamerican12101881-
374b.pdf.

Anonymous:1881:AMd

[2370] Anonymous. Artistic mantelpiece. Scientific American, 45(24): 375, December 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n24/pdf/scientificamerican12101881-
375.pdf.

Anonymous:1881:EIt

[2371] Anonymous. Engineering inventions. Scientific American, 45(24): 376, December 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n24/pdf/scientificamerican12101881-
376.pdf.

Anonymous:1881:AIt

[2372] Anonymous. Agricultural inventions. Scientific American, 45 (24):378, December 10, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n24/pdf/scientificamerican12101881-
378.pdf.

Anonymous:1881:Cbd

[2373] Anonymous. Correspondence. Scientific American, 45(25):389, December 17, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n25/pdf/scientificamerican12171881-389.pdf.

Anonymous:1881:INIbc

[2374] Anonymous. Inventions: New inventions. Scientific American, 45 (25):391–392, December 17, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n25/pdf/scientificamerican12171881-
391a.pdf.

Anonymous:1881:BPbq

[2375] Anonymous. Business and personal. Scientific American, 45 (25):394, December 17, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n25/pdf/scientificamerican12171881-
394.pdf.

Anonymous:1881:NQbp

[2376] Anonymous. Notes and queries. Scientific American, 45(25): 394–395, December 17, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n25/pdf/scientificamerican12171881-
394a.pdf.

Anonymous:1881:IIbs

[2377] Anonymous. Index of inventions. Scientific American, 45(25): 395, December 17, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n25/pdf/scientificamerican12171881-
395.pdf.

Anonymous:1881:BNBv

[2378] Anonymous. Books: New books and publications. Scientific American, 45(25):395, December 17, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n25/pdf/scientificamerican12171881-
395a.pdf.

Anonymous:1881:AINn

[2379] Anonymous. American industries — no 79. Scientific American, 45(25):383, December 17, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n25/pdf/scientificamerican12171881-
383.pdf.

Anonymous:1881:SGE

[2380] Anonymous. Some great engineering project. Scientific American, 45(25):384, December 17, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n25/pdf/scientificamerican12171881-
384.pdf.

Anonymous:1881:RA

[2381] Anonymous. Railroad accidents. Scientific American, 45(25): 384, December 17, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n25/pdf/scientificamerican12171881-
384a.pdf.

Anonymous:1881:TBE

[2382] Anonymous. Terrible boiler explosion in Ohio. Scientific American, 45(25):384–385, December 17, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n25/pdf/scientificamerican12171881-
384b.pdf.

Anonymous:1881:MIch

[2383] Anonymous. Miscellaneous inventions. Scientific American, 45(25): 388–389, December 17, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n25/pdf/scientificamerican12171881-
388.pdf.

Anonymous:1881:SBNl

[2384] Anonymous. Steam boiler notes. Scientific American, 45(25): 390, December 17, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n25/pdf/scientificamerican12171881-
390.pdf.

Anonymous:1881:ISFb

[2385] Anonymous. Improved shutter fastener. Scientific American, 45 (25):390, December 17, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n25/pdf/scientificamerican12171881-
390a.pdf.

Anonymous:1881:PSL

[2386] Anonymous. Proposed storage of lightning. Scientific American, 45(25):390, December 17, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n25/pdf/scientificamerican12171881-
390b.pdf.

Anonymous:1881:AMf

[2387] Anonymous. Amateur mechanics. Scientific American, 45(25): 391, December 17, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n25/pdf/scientificamerican12171881-
391.pdf.

Anonymous:1881:MIci

[2388] Anonymous. Mechanical inventions. Scientific American, 45(25): 393, December 17, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n25/pdf/scientificamerican12171881-
393.pdf.

Anonymous:1881:Cbe

[2389] Anonymous. Correspondence. Scientific American, 45(26):405, December 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n26/pdf/scientificamerican12241881-405a.pdf.

Anonymous:1881:IRIbe

[2390] Anonymous. Inventions: Recent inventions. Scientific American, 45(26):408, December 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n26/pdf/scientificamerican12241881-
408.pdf.

Anonymous:1881:BPbr

[2391] Anonymous. Business and personal. Scientific American, 45 (26):410, December 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n26/pdf/scientificamerican12241881-
410.pdf.

Anonymous:1881:NQbq

[2392] Anonymous. Notes and queries. Scientific American, 45(26): 410–411, December 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n26/pdf/scientificamerican12241881-
410a.pdf.

Anonymous:1881:IIbt

[2393] Anonymous. Index of inventions. Scientific American, 45(26): 411, December 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n26/pdf/scientificamerican12241881-
411.pdf.

Anonymous:1881:PNS

[2394] Anonymous. Proposed new system of waterworks for Chicago. Scientific American, 45(26):399, December 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v45/n26/pdf/scientificamerican12241881-
399.pdf.

Anonymous:1881:GBF

[2395] Anonymous. A glance backward and forward. Scientific American, 45(26):400, December 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n26/pdf/scientificamerican12241881-
400.pdf.

Anonymous:1881:PIb

[2396] Anonymous. Patents as investments. Scientific American, 45(26): 400–401, December 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n26/pdf/scientificamerican12241881-
400a.pdf.

Anonymous:1881:VRB

[2397] Anonymous. Vibration of railway bridges. Scientific American, 45 (26):401, December 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n26/pdf/scientificamerican12241881-
401.pdf.

Anonymous:1881:NCD

[2398] Anonymous. Naval and coast defense. Scientific American, 45 (26):401–402, December 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n26/pdf/scientificamerican12241881-
401a.pdf.

Anonymous:1881:PIBa

[2399] Anonymous. Pig iron breaker. Scientific American, 45(26):402, December 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n26/pdf/scientificamerican12241881-402.pdf.

Anonymous:1881:ISCb

[2400] Anonymous. Improved shaft coupling. Scientific American, 45 (26):402, December 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n26/pdf/scientificamerican12241881-
402a.pdf.

Anonymous:1881:EBE

[2401] Anonymous. Experimental boiler explosion. Scientific American, 45(26):403, December 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n26/pdf/scientificamerican12241881-
403.pdf.

Anonymous:1881:MIcj

[2402] Anonymous. Mechanical inventions. Scientific American, 45(26): 403–404, December 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n26/pdf/scientificamerican12241881-
403a.pdf.

Anonymous:1881:DRPh

[2403] Anonymous. Decisions relating to patents. Scientific American, 45(26):404, December 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n26/pdf/scientificamerican12241881-
404.pdf.

Anonymous:1881:SBNm

[2404] Anonymous. Steam boiler notes. Scientific American, 45(26): 405, December 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n26/pdf/scientificamerican12241881-
405.pdf.

Anonymous:1881:NWW

[2405] Anonymous. New worm and worm wheel and gear cutter. Scientific American, 45(26):406, December 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v45/n26/pdf/scientificamerican12241881-
406.pdf.

Anonymous:1881:EIu

[2406] Anonymous. Engineering inventions. Scientific American, 45(26): 406, December 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n26/pdf/scientificamerican12241881-
406a.pdf.

Hopkins:1881:DEa

[2407] Geo. M. Hopkins. Dynamic electricity. Scientific American, 45 (26):407, December 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n26/pdf/scientificamerican12241881-
407.pdf.

Seiss:1881:FS

[2408] R. W. Seiss. The flying squirrel. Scientific American, 45(26): 409, December 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n26/pdf/scientificamerican12241881-
409.pdf.

Anonymous:1881:TW

[2409] Anonymous. The tape worm. Scientific American, 45(26):409, December 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n26/pdf/scientificamerican12241881-409a.pdf.

Anonymous:1881:MIck

[2410] Anonymous. Miscellaneous inventions. Scientific American, 45 (26):409, December 24, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n26/pdf/scientificamerican12241881-
409b.pdf.

Anonymous:1881:Cbf

[2411] Anonymous. Correspondence. Scientific American, 45(3):37–38, July 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n3/pdf/scientificamerican07161881-37.pdf.

Anonymous:1881:BPbs

[2412] Anonymous. Business and personal. Scientific American, 45(3):43, July 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n3/pdf/scientificamerican07161881-43.pdf.

Anonymous:1881:NQbr

[2413] Anonymous. Notes and queries. Scientific American, 45(3):43, July 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n3/pdf/scientificamerican07161881-43a.pdf.

Anonymous:1881:IIbu

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n3/pdf/scientificamerican07161881-43c.pdf.

Anonymous:1881:PNPba

[2415] Anonymous. Patent notes: Patents. Scientific American, 45(3):45, July 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n3/pdf/scientificamerican07161881-45.pdf.

Anonymous:1881:RNS

[2416] Anonymous. Renewal of Niagara Suspension Bridge. Scientific American, 45(3):31, July 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n3/pdf/scientificamerican07161881-
31.pdf.

Anonymous:1881:ITD

[2417] Anonymous. Important telephone decision. Scientific American, 45(3):32, July 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n3/pdf/scientificamerican07161881-
32.pdf.

Anonymous:1881:GC

[2418] Anonymous. The Great Comet of 1881. Scientific American, 45(3):32, July 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n3/pdf/scientificamerican07161881-32a.pdf.

Anonymous:1881:CAM

[2419] Anonymous. Chemical action in a magnetic field. Scientific American, 45(3):32, July 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n3/pdf/scientificamerican07161881-
32b.pdf.

Anonymous:1881:SCBa

[2420] Anonymous. The spread of the carpet beetle. Scientific American, 45(3):33, July 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n3/pdf/scientificamerican07161881-
33.pdf.

Anonymous:1881:GZ

[2421] Anonymous. Gamgee’s zeromotor. Scientific American, 45(3):33, July 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n3/pdf/scientificamerican07161881-33a.pdf.

Anonymous:1881:WW

[2422] Anonymous. Wood weaving. Scientific American, 45(3):33, July 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n3/pdf/scientificamerican07161881-33b.pdf.

Hartwell:1881:ERM

[2423] S. N. Hartwell. Explosion of a rolling mill boiler in Pottsville, Pennsylvania. Scientific American, 45(3):34, July 16, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v45/
n3/pdf/scientificamerican07161881-34.pdf.

Anonymous:1881:IHA

[2424] Anonymous. Improved hoisting apparatus. Scientific American, 45(3):38, July 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n3/pdf/scientificamerican07161881-
38.pdf.

Anonymous:1881:LGE

[2425] Anonymous. Lighting gas by electricity. Scientific American, 45 (3):38–39, July 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n3/pdf/scientificamerican07161881-
38a.pdf.

Anonymous:1881:PDE

[2426] Anonymous. Panel decorations for Eaton Hall. Scientific American, 45(3):39–40, July 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n3/pdf/scientificamerican07161881-
39.pdf.

Anonymous:1881:L

[2427] Anonymous. Lubricants. Scientific American, 45(3):41, July 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n3/pdf/scientificamerican07161881-41.pdf.

Anonymous:1881:CPb

[2428] Anonymous. Chemical paradoxes. Scientific American, 45(3):42, July 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n3/pdf/scientificamerican07161881-42.pdf.

Anonymous:1881:FFb

[2429] Anonymous. Firework formulæ. Scientific American, 45(3):42, July 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n3/pdf/scientificamerican07161881-42a.pdf.

Anonymous:1881:CRt

[2430] Anonymous. Communication received. Scientific American, 45(3):43, July 16, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n3/pdf/scientificamerican07161881-43b.pdf.

Anonymous:1881:IRIbf

[2431] Anonymous. Inventions: Recent inventions. Scientific American, 45(4):53, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n4/pdf/scientificamerican07231881-
53.pdf.

Anonymous:1881:INIbd

[2432] Anonymous. Inventions: New inventions. Scientific American, 45(4):57–58, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n4/pdf/scientificamerican07231881-
57.pdf.

Anonymous:1881:BPbt

[2433] Anonymous. Business and personal. Scientific American, 45(4):59, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n4/pdf/scientificamerican07231881-59.pdf.

Anonymous:1881:BNBw

[2434] Anonymous. Books: New books and publications. Scientific American, 45(4):59, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n4/pdf/scientificamerican07231881-
59a.pdf.

Anonymous:1881:NQbs

[2435] Anonymous. Notes and queries. Scientific American, 45(4):59, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n4/pdf/scientificamerican07231881-59b.pdf.

Anonymous:1881:IIbv

[2436] Anonymous. Index of inventions. Scientific American, 45(4):59–60, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n4/pdf/scientificamerican07231881-59c.pdf.

Anonymous:1881:PNPbb

[2437] Anonymous. Patent notes: Patents. Scientific American, 45(4):60, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n4/pdf/scientificamerican07231881-60.pdf.

Anonymous:1881:UAW

[2438] Anonymous. Union Arch, Washington Aqueduct. Scientific American, 45(4):47, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n4/pdf/scientificamerican07231881-
47.pdf.

Anonymous:1881:SCd

[2439] Anonymous. Separation of the Comet. Scientific American, 45(4):48, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n4/pdf/scientificamerican07231881-48.pdf.

Anonymous:1881:FC

[2440] Anonymous. Fulminating compounds. Scientific American, 45(4):48, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n4/pdf/scientificamerican07231881-48a.pdf.

Anonymous:1881:PRb

[2441] Anonymous. Poisonous refreshments. Scientific American, 45(4):48, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n4/pdf/scientificamerican07231881-48b.pdf.

Anonymous:1881:MC

[2442] Anonymous. The manufacture of celluloid. Scientific American, 45(4):48–49, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n4/pdf/scientificamerican07231881-
48c.pdf.

Anonymous:1881:SBNn

[2443] Anonymous. Steam boiler notes. Scientific American, 45(4):49, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n4/pdf/scientificamerican07231881-49.pdf.

Anonymous:1881:CCM

[2444] Anonymous. Centripetal and centrifugal motions in animals. Scientific American, 45(4):49, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n4/pdf/scientificamerican07231881-
49a.pdf.

Anonymous:1881:ISHb

[2445] Anonymous. Improved spring hinge. Scientific American, 45(4):50, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n4/pdf/scientificamerican07231881-50.pdf.

Anonymous:1881:NMMb

[2446] Anonymous. New mechanical movement. Scientific American, 45(4):50, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n4/pdf/scientificamerican07231881-50a.pdf.

Anonymous:1881:NGF

[2447] Anonymous. New glove fastener. Scientific American, 45(4):50, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n4/pdf/scientificamerican07231881-50b.pdf.

Anonymous:1881:MIcl

[2448] Anonymous. Mechanical inventions. Scientific American, 45(4):50, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n4/pdf/scientificamerican07231881-50c.pdf.

Anonymous:1881:AIu

[2449] Anonymous. Agricultural inventions. Scientific American, 45(4):51, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n4/pdf/scientificamerican07231881-51.pdf.

Anonymous:1881:NTTb

[2450] Anonymous. New tire tightener. Scientific American, 45(4):51, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n4/pdf/scientificamerican07231881-51a.pdf.

Hartwell:1881:CSB

[2451] S. N. Hartwell. Collapse of a steel boiler flue. Scientific American, 45(4):51–53, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n4/pdf/scientificamerican07231881-
51b.pdf.

Anonymous:1881:PSb

[2452] Anonymous. Phosphorescent substances. Scientific American, 45 (4):53–54, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n4/pdf/scientificamerican07231881-
53a.pdf.

Anonymous:1881:EMg

[2453] Anonymous. Electro-massage. Scientific American, 45(4):54, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n4/pdf/scientificamerican07231881-54.pdf.

Anonymous:1881:NRSb

[2454] Anonymous. New reflector for suspended lamp. Scientific American, 45(4):54, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n4/pdf/scientificamerican07231881-
54a.pdf.

Anonymous:1881:IBb

[2455] Anonymous. Improvement in buttons. Scientific American, 45(4):54, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n4/pdf/scientificamerican07231881-54b.pdf.

Anonymous:1881:MIcm

[2456] Anonymous. Miscellaneous inventions. Scientific American, 45(4):54, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n4/pdf/scientificamerican07231881-54c.pdf.

Anonymous:1881:VAB

[2457] Anonymous. Volumetric assay of bullion, alloys, etc. Scientific American, 45(4):55, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n4/pdf/scientificamerican07231881-
55.pdf.

Hopkins:1881:DEb

[2458] M. Hopkins. Dynamic electricity. Scientific American, 45(4):55–56, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n4/pdf/scientificamerican07231881-55a.pdf.

Anonymous:1881:EIv

[2459] Anonymous. Engineering inventions. Scientific American, 45(4):58, July 23, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n4/pdf/scientificamerican07231881-58.pdf.

Anonymous:1881:IRIbg

[2460] Anonymous. Inventions: Recent inventions. Scientific American, 45(5):66, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n5/pdf/scientificamerican07301881-
66b.pdf.

Anonymous:1881:Cbg

[2461] Anonymous. Correspondence. Scientific American, 45(5):69, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n5/pdf/scientificamerican07301881-69.pdf.

Anonymous:1881:BPbu

[2462] Anonymous. Business and personal. Scientific American, 45(5):74, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n5/pdf/scientificamerican07301881-74.pdf.

Anonymous:1881:NQbt

[2463] Anonymous. Notes and queries. Scientific American, 45(5):74, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n5/pdf/scientificamerican07301881-74a.pdf.

Anonymous:1881:BNBx

[2464] Anonymous. Books: New books and publications. Scientific American, 45(5):74–75, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n5/pdf/scientificamerican07301881-
74b.pdf.

Anonymous:1881:IIbw

[2465] Anonymous. Index of inventions. Scientific American, 45(5):75, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n5/pdf/scientificamerican07301881-75.pdf.

Anonymous:1881:PNPbc

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n5/pdf/scientificamerican07301881-77.pdf.

Anonymous:1881:NLM

[2467] Anonymous. New lasting machine. Scientific American, 45(5):63, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n5/pdf/scientificamerican07301881-63.pdf.

Anonymous:1881:DGF

[2468] Anonymous. Dr. Griscom’s fast. Scientific American, 45(5):64, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n5/pdf/scientificamerican07301881-64.pdf.

Anonymous:1881:AAR

[2469] Anonymous. An ancient aqueduct reopened. Scientific American, 45(5):64, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n5/pdf/scientificamerican07301881-
64a.pdf.

Anonymous:1881:IES

[2470] Anonymous. Inconveniences of electric street railways. Scientific American, 45(5):64, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n5/pdf/scientificamerican07301881-
64b.pdf.

Anonymous:1881:GBE

[2471] Anonymous. The Gaffney boiler explosion. Scientific American, 45(5):64, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n5/pdf/scientificamerican07301881-
64c.pdf.

Anonymous:1881:SCP

[2472] Anonymous. Stencil copying process. Scientific American, 45(5):65, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n5/pdf/scientificamerican07301881-65.pdf.

Anonymous:1881:ITCa

[2473] Anonymous. Interesting test of cast iron boiler heads. Scientific American, 45(5):65, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n5/pdf/scientificamerican07301881-
65a.pdf.

Anonymous:1881:LSL

[2474] Anonymous. A life-saving lesson in physics. Scientific American, 45(5):65, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n5/pdf/scientificamerican07301881-
65b.pdf.

Anonymous:1881:NAGb

[2475] Anonymous. New adjustable gun stock. Scientific American, 45(5):66, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n5/pdf/scientificamerican07301881-66.pdf.

Hopkins:1881:NUI

[2476] M. Hopkins. A new use for the induction balance. Scientific American, 45(5):66, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n5/pdf/scientificamerican07301881-
66a.pdf.

Anonymous:1881:ICSb

[2477] Anonymous. Improved canceling stamp. Scientific American, 45(5):67, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n5/pdf/scientificamerican07301881-67.pdf.

Anonymous:1881:NBS

[2478] Anonymous. New band saw machine. Scientific American, 45(5):67, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n5/pdf/scientificamerican07301881-67a.pdf.

Anonymous:1881:IDCa

[2479] Anonymous. Improved dividers and calipers. Scientific American, 45(5):67, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n5/pdf/scientificamerican07301881-
67b.pdf.

Anonymous:1881:NCE

[2480] Anonymous. Novel cork extractor. Scientific American, 45(5):67–68, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n5/pdf/scientificamerican07301881-67c.pdf.

Anonymous:1881:PTI

[2481] Anonymous. Preparation of tinned iron — tin plate. Scientific American, 45(5):68, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n5/pdf/scientificamerican07301881-
68.pdf.

Anonymous:1881:MB

[2482] Anonymous. Maiche’s battery. Scientific American, 45(5):68, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n5/pdf/scientificamerican07301881-68a.pdf.

Anonymous:1881:MIcn

[2483] Anonymous. Miscellaneous inventions. Scientific American, 45(5):69, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n5/pdf/scientificamerican07301881-69a.pdf.

Anonymous:1881:ISMb

[2484] Anonymous. Improved slotting machine. Scientific American, 45(5):70, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n5/pdf/scientificamerican07301881-70.pdf.

Anonymous:1881:IL

[2485] Anonymous. Improved lubricator. Scientific American, 45(5):70, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n5/pdf/scientificamerican07301881-70a.pdf.

Beard:1881:HAO

[2486] Daniel C. Beard. Home of the American osprey. Scientific American, 45(5):71–72, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n5/pdf/scientificamerican07301881-
71.pdf.

Anonymous:1881:QMW

[2487] Anonymous. Quartz and marl as wood fillers. Scientific American, 45(5):73, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n5/pdf/scientificamerican07301881-
73.pdf.

Anonymous:1881:MIco

[2488] Anonymous. Mechanical inventions. Scientific American, 45(5):73, July 30, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n5/pdf/scientificamerican07301881-73a.pdf.

Anonymous:1881:Cbh

[2489] Anonymous. Correspondence. Scientific American, 45(7):100, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n7/pdf/scientificamerican08131881-100a.pdf.

Anonymous:1881:INIbe

[2490] Anonymous. Inventions: New inventions. Scientific American, 45(7):102, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n7/pdf/scientificamerican08131881-
102.pdf.

Anonymous:1881:IRIbh

[2491] Anonymous. Inventions: Recent inventions. Scientific American, 45(7):105, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n7/pdf/scientificamerican08131881-
105.pdf.

Anonymous:1881:Ie

[2492] Anonymous. Inventions. Scientific American, 45(7):106, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n7/pdf/scientificamerican08131881-106.pdf.

Anonymous:1881:BPbv

[2493] Anonymous. Business and personal. Scientific American, 45(7):107, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n7/pdf/scientificamerican08131881-107.pdf.

Anonymous:1881:NQbu

[2494] Anonymous. Notes and queries. Scientific American, 45(7):107, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n7/pdf/scientificamerican08131881-107a.pdf.

Anonymous:1881:IIbx

[2495] Anonymous. Index of inventions. Scientific American, 45(7): 107–108, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n7/pdf/scientificamerican08131881-
107b.pdf.

Anonymous:1881:PNPbd

[2496] Anonymous. Patent notes: Patents. Scientific American, 45(7):109, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n7/pdf/scientificamerican08131881-109.pdf.

Anonymous:1881:IBL

[2497] Anonymous. Improved band log saw mills. Scientific American, 45(7):95, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n7/pdf/scientificamerican08131881-
95.pdf.

Anonymous:1881:VHA

[2498] Anonymous. Ventilation of halls of audience. Scientific American, 45(7):96, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n7/pdf/scientificamerican08131881-
96.pdf.

Anonymous:1881:BSR

[2499] Anonymous. Button-set riveting for boilers. Scientific American, 45(7):96, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n7/pdf/scientificamerican08131881-
96a.pdf.

Anonymous:1881:TSAa

[2500] Anonymous. The torpedo steam alarm. Scientific American, 45 (7):96–97, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n7/pdf/scientificamerican08131881-
96b.pdf.

Hartwell:1881:MLB

[2501] S. N. Hartwell. Mr. Lawson’s boiler explosion. Scientific American, 45(7):97, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n7/pdf/scientificamerican08131881-
97.pdf.

Hovey:1881:PCb

[2502] H. C. Hovey. Pioneer canning. Scientific American, 45(7):97, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n7/pdf/scientificamerican08131881-97a.pdf.

Anonymous:1881:GEP

[2503] Anonymous. Galvanization of an engine piston. Scientific American, 45(7):97, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n7/pdf/scientificamerican08131881-
97b.pdf.

Anonymous:1881:AIv

[2504] Anonymous. Agricultural inventions. Scientific American, 45(7):98, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n7/pdf/scientificamerican08131881-98.pdf.

Anonymous:1881:BMI

[2505] Anonymous. Bending machine for iron and steel. Scientific American, 45(7):98, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n7/pdf/scientificamerican08131881-
98a.pdf.

Anonymous:1881:EMCb

[2506] Anonymous. Extension magazines for coal stoves. Scientific American, 45(7):98, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n7/pdf/scientificamerican08131881-
98b.pdf.

Anonymous:1881:SBa

[2507] Anonymous. A simple balance. Scientific American, 45(7):99, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n7/pdf/scientificamerican08131881-99.pdf.

Anonymous:1881:VBB

[2508] Anonymous. Ventilated bath box for chemical and photographic purposes. Scientific American, 45(7):99, August 13, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v45/
n7/pdf/scientificamerican08131881-99a.pdf.

Anonymous:1881:CMa

[2509] Anonymous. Condensed milk. Scientific American, 45(7):99, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n7/pdf/scientificamerican08131881-99b.pdf.

Anonymous:1881:EAC

[2510] Anonymous. Electro-assay of copper ores. Scientific American, 45(7):99–100, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n7/pdf/scientificamerican08131881-
99c.pdf.

Anonymous:1881:NB

[2511] Anonymous. A novel bottle. Scientific American, 45(7):100, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n7/pdf/scientificamerican08131881-100.pdf.

Anonymous:1881:MIcp

[2512] Anonymous. Miscellaneous inventions. Scientific American, 45 (7):100–101, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n7/pdf/scientificamerican08131881-
100b.pdf.

Anonymous:1881:SEF

[2513] Anonymous. The secret of easy floating. Scientific American, 45(7):101, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n7/pdf/scientificamerican08131881-
101.pdf.

Anonymous:1881:NDP

[2514] Anonymous. New double plunges steam pump. Scientific American, 45(7):102, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n7/pdf/scientificamerican08131881-
102a.pdf.

Anonymous:1881:NFP

[2515] Anonymous. Novel fruit picker. Scientific American, 45(7):102, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n7/pdf/scientificamerican08131881-102b.pdf.

Anonymous:1881:EIw

[2516] Anonymous. Engineering inventions. Scientific American, 45(7): 103, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n7/pdf/scientificamerican08131881-
103.pdf.

Anonymous:1881:SVb

[2517] Anonymous. Sevres vase. Scientific American, 45(7):103, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n7/pdf/scientificamerican08131881-103a.pdf.

Anonymous:1881:SPFb

[2518] Anonymous. The Sèvres Porcelain Factory. Scientific American, 45(7):103–104, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n7/pdf/scientificamerican08131881-
103b.pdf.

Anonymous:1881:SBNo

[2519] Anonymous. Steam boiler notes. Scientific American, 45(7):104, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n7/pdf/scientificamerican08131881-104.pdf.

Anonymous:1881:MIcq

[2520] Anonymous. Mechanical inventions. Scientific American, 45(7):104, August 13, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n7/pdf/scientificamerican08131881-104a.pdf.

Anonymous:1881:INIbf

[2521] Anonymous. Inventions: New inventions. Scientific American, 45(8):114, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n8/pdf/scientificamerican08201881-
114.pdf.

Anonymous:1881:IRIbi

[2522] Anonymous. Inventions: Recent inventions. Scientific American, 45(8):117, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n8/pdf/scientificamerican08201881-
117a.pdf.

Anonymous:1881:BPbw

[2523] Anonymous. Business and personal. Scientific American, 45(8):123, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n8/pdf/scientificamerican08201881-123.pdf.

Anonymous:1881:NQbv

[2524] Anonymous. Notes and queries. Scientific American, 45(8):123, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n8/pdf/scientificamerican08201881-123a.pdf.

Anonymous:1881:IIby

[2525] Anonymous. Index of inventions. Scientific American, 45(8): 123–124, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n8/pdf/scientificamerican08201881-
123b.pdf.

Anonymous:1881:PNPbe

[2526] Anonymous. Patent notes: Patents. Scientific American, 45(8):125, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n8/pdf/scientificamerican08201881-125.pdf.

Anonymous:1881:NEK

[2527] Anonymous. Non-explosive kerosene. Scientific American, 45(8): 112, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n8/pdf/scientificamerican08201881-
112.pdf.

Anonymous:1881:STP

[2528] Anonymous. State taxation of patented articles. Scientific American, 45(8):112, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n8/pdf/scientificamerican08201881-
112a.pdf.

Anonymous:1881:BCb

[2529] Anonymous. Butter coloring. Scientific American, 45(8):112, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n8/pdf/scientificamerican08201881-112b.pdf.

Anonymous:1881:TCb

[2530] Anonymous. The tails of comets. Scientific American, 45(8): 112–113, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n8/pdf/scientificamerican08201881-
112c.pdf.

Anonymous:1881:SEc

[2531] Anonymous. A strange explosion. Scientific American, 45(8):113, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n8/pdf/scientificamerican08201881-113.pdf.

Anonymous:1881:DEE

[2532] Anonymous. Dust explosion in Ehret’s Brewery. Scientific American, 45(8):113, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n8/pdf/scientificamerican08201881-
113a.pdf.

Anonymous:1881:LOD

[2533] Anonymous. Latchinoff’s optical dynamometer. Scientific American, 45(8):114, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n8/pdf/scientificamerican08201881-
114a.pdf.

Anonymous:1881:EIx

[2534] Anonymous. Engineering inventions. Scientific American, 45(8): 114–115, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n8/pdf/scientificamerican08201881-
114b.pdf.

Anonymous:1881:NRF

[2535] Anonymous. New rail fastening. Scientific American, 45(8):115, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n8/pdf/scientificamerican08201881-115.pdf.

Anonymous:1881:PMS

[2536] Anonymous. Portable meteorological station. Scientific American, 45(8):115, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n8/pdf/scientificamerican08201881-
115a.pdf.

Anonymous:1881:SBNp

[2537] Anonymous. Steam-boiler notes. Scientific American, 45(8):115– 116, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n8/pdf/scientificamerican08201881-
115b.pdf.

Anonymous:1881:AP

[2538] Anonymous. Attachment for pants. Scientific American, 45(8):116, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n8/pdf/scientificamerican08201881-116.pdf.

Eilet:1881:RCR

[2539] C. T. Eilet. A remarkable case of retarded development. Scientific American, 45(8):116, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n8/pdf/scientificamerican08201881-
116a.pdf.

Anonymous:1881:AINo

[2540] Anonymous. American industries — no. 74. Scientific American, 45(8):117, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n8/pdf/scientificamerican08201881-
117.pdf.

Anonymous:1881:NTb

[2541] Anonymous. A new tool. Scientific American, 45(8):118, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n8/pdf/scientificamerican08201881-118.pdf.

Anonymous:1881:IIFa

[2542] Anonymous. An improvement in iron fences. Scientific American, 45(8):118, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n8/pdf/scientificamerican08201881-
118a.pdf.

Anonymous:1881:MIcr

[2543] Anonymous. Miscellaneous inventions. Scientific American, 45 (8):118, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n8/pdf/scientificamerican08201881-
118b.pdf.

Anonymous:1881:HTb

[2544] Anonymous. A handy tool. Scientific American, 45(8):118–119, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n8/pdf/scientificamerican08201881-118c.pdf.

Anonymous:1881:ES

[2545] Anonymous. Earth stars. Scientific American, 45(8):119, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n8/pdf/scientificamerican08201881-119.pdf.

Anonymous:1881:THD

[2546] Anonymous. A three hundred dollar turtle. Scientific American, 45(8):119–120, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n8/pdf/scientificamerican08201881-
119a.pdf.

Anonymous:1881:NHNd

[2547] Anonymous. Natural history notes. Scientific American, 45(8): 120–122, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n8/pdf/scientificamerican08201881-
120.pdf.

Anonymous:1881:DCRb

[2548] Anonymous. Decisions of the courts relating to patents. Scientific American, 45(8):122, August 20, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n8/pdf/scientificamerican08201881-
122.pdf.

Anonymous:1881:Cbi

[2549] Anonymous. Correspondence. Scientific American, 45(9):132, August 27, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n9/pdf/scientificamerican08271881-132.pdf.

Anonymous:1881:INIbg

[2550] Anonymous. Inventions: New inventions. Scientific American, 45(9):133, August 27, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n9/pdf/scientificamerican08271881-
133.pdf.

Anonymous:1881:BPbx

[2551] Anonymous. Business and personal. Scientific American, 45(9):139, August 27, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n9/pdf/scientificamerican08271881-139.pdf.

Anonymous:1881:BNBy

[2552] Anonymous. Books: New books and publications. Scientific American, 45(9):139, August 27, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n9/pdf/scientificamerican08271881-
139a.pdf.

Anonymous:1881:NQbw

[2553] Anonymous. Notes and queries. Scientific American, 45(9):139, August 27, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n9/pdf/scientificamerican08271881-139b.pdf.

Anonymous:1881:IIbz

[2554] Anonymous. Index of inventions. Scientific American, 45(9): 139–140, August 27, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n9/pdf/scientificamerican08271881-
139d.pdf.

Anonymous:1881:PNPbf

[2555] Anonymous. Patent notes: Patents. Scientific American, 45(9):141, August 27, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n9/pdf/scientificamerican08271881-141.pdf.

Anonymous:1881:WWP

[2556] Anonymous. Watertown, N.Y., and its water power. Scientific American, 45(9):127, August 27, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n9/pdf/scientificamerican08271881-
127.pdf.

Anonymous:1881:MS

[2557] Anonymous. “A militia for the sea”. Scientific American, 45(9):128, August 27, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n9/pdf/scientificamerican08271881-128.pdf.

Anonymous:1881:CMb

[2558] Anonymous. The cotton manufacture. Scientific American, 45 (9):128, August 27, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n9/pdf/scientificamerican08271881-
128a.pdf.

Anonymous:1881:CCa

[2559] Anonymous. Is consumption contagious? Scientific American, 45 (9):128–129, August 27, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n9/pdf/scientificamerican08271881-
128b.pdf.

Anonymous:1881:OPE

[2560] Anonymous. Opening of the Paris Electric Exhibition. Scientific American, 45(9):129, August 27, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n9/pdf/scientificamerican08271881-
129.pdf.

Anonymous:1881:SCe

[2561] Anonymous. Schaeberle’s Comet. Scientific American, 45(9):129, August 27, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n9/pdf/scientificamerican08271881-129a.pdf.

Hovey:1881:RIR

[2562] H. C. Hovey. A remarkable instance of retention of heat by the Earth. Scientific American, 45(9):129, August 27, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v45/
n9/pdf/scientificamerican08271881-129b.pdf.

Anonymous:1881:ISW

[2563] Anonymous. Improved shutter-worker. Scientific American, 45 (9):130, August 27, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n9/pdf/scientificamerican08271881-
130.pdf.

Anonymous:1881:IF

[2564] Anonymous. Improved filter. Scientific American, 45(9):130, August 27, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n9/pdf/scientificamerican08271881-130a.pdf.

Anonymous:1881:IDL

[2565] Anonymous. Improved door latch. Scientific American, 45(9):130, August 27, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n9/pdf/scientificamerican08271881-130b.pdf.

Anonymous:1881:MIcs

[2566] Anonymous. Mechanical inventions. Scientific American, 45(9): 130–131, August 27, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n9/pdf/scientificamerican08271881-
130c.pdf.

Anonymous:1881:TSAb

[2567] Anonymous. Temporary steering apparatus of H.M.S. Bacchante. Scientific American, 45(9):131, August 27, 1881. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v45/
n9/pdf/scientificamerican08271881-131.pdf.

Anonymous:1881:ISBb

[2568] Anonymous. Improved spring bed. Scientific American, 45(9):131, August 27, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v45/n9/pdf/scientificamerican08271881-131a.pdf.

Anonymous:1881:HMG

[2569] Anonymous. How to make good cider and to keep it. Scientific American, 45(9):133–134, August 27, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n9/pdf/scientificamerican08271881-
133a.pdf.

Anonymous:1881:NMD

[2570] Anonymous. Novel machine for demagnetizing watches. Scientific American, 45(9):134, August 27, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n9/pdf/scientificamerican08271881-
134.pdf.

Anonymous:1881:GOB

[2571] Anonymous. The guacharo, or oil bird. Scientific American, 45 (9):135–136, August 27, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n9/pdf/scientificamerican08271881-
135.pdf.

Anonymous:1881:DCRc

[2572] Anonymous. Decisions of the courts relating to patents. Scientific American, 45(9):137–138, August 27, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v45/n9/pdf/scientificamerican08271881-
137.pdf.

Anonymous:1881:MIct

[2573] Anonymous. Miscellaneous inventions. Scientific American, 45 (9):138, August 27, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n9/pdf/scientificamerican08271881-
138.pdf.

Anonymous:1881:CRu

[2574] Anonymous. Communications received. Scientific American, 45 (9):139, August 27, 1881. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v45/n9/pdf/scientificamerican08271881-
139c.pdf.

Anonymous:1882:IRIa

[2575] Anonymous. Inventions: Recent inventions. Scientific American, 46(1):4–5, January 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n1/pdf/scientificamerican01071882-
4b.pdf.

Anonymous:1882:INIa

[2576] Anonymous. Inventions: New inventions. Scientific American, 46(1):8–9, January 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n1/pdf/scientificamerican01071882-
8b.pdf.

Anonymous:1882:Ca

[2577] Anonymous. Correspondence. Scientific American, 46(1):10, January 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n1/pdf/scientificamerican01071882-10.pdf.

Anonymous:1882:BPa

[2578] Anonymous. Business and personal. Scientific American, 46(1):11, January 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n1/pdf/scientificamerican01071882-11.pdf.

Anonymous:1882:NQa

[2579] Anonymous. Notes and queries. Scientific American, 46(1):11, January 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n1/pdf/scientificamerican01071882-11a.pdf.

Anonymous:1882:IIa

[2580] Anonymous. Index of inventions. Scientific American, 46(1):11–12, January 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n1/pdf/scientificamerican01071882-11c.pdf.

Anonymous:1882:SE

[2581] Anonymous. Stone engraving. Scientific American, 46(1):1, January 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n1/pdf/scientificamerican01071882-1.pdf.

Anonymous:1882:NLA

[2582] Anonymous. New laws for analyzing food and drugs. Scientific American, 46(1):2, January 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n1/pdf/scientificamerican01071882-
2.pdf.

Anonymous:1882:ESF

[2583] Anonymous. English as the speech of the future. Scientific American, 46(1):2, January 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n1/pdf/scientificamerican01071882-
2a.pdf.

Anonymous:1882:LJ

[2584] Anonymous. The loss of the Jeannette. Scientific American, 46(1):2–3, January 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n1/pdf/scientificamerican01071882-2b.pdf.

Anonymous:1882:ICCa

[2585] Anonymous. Improved car coupling. Scientific American, 46(1):4, January 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n1/pdf/scientificamerican01071882-4.pdf.

Anonymous:1882:IFa

[2586] Anonymous. Improved filter. Scientific American, 46(1):4, January 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n1/pdf/scientificamerican01071882-4a.pdf.

Anonymous:1882:ICT

[2587] Anonymous. Improved car truck. Scientific American, 46(1):5, January 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n1/pdf/scientificamerican01071882-5.pdf.

Anonymous:1882:IPH

[2588] Anonymous. Improved power and hand presses. Scientific American, 46(1):5, January 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n1/pdf/scientificamerican01071882-
5a.pdf.

Anonymous:1882:MIa

[2589] Anonymous. Miscellaneous inventions. Scientific American, 46(1):6, January 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n1/pdf/scientificamerican01071882-6.pdf.

Anonymous:1882:NFSa

[2590] Anonymous. New fret sawing machine. Scientific American, 46(1):7, January 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n1/pdf/scientificamerican01071882-7.pdf.

Anonymous:1882:APG

[2591] Anonymous. The ætna pattern grate bar. Scientific American, 46(1):8, January 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n1/pdf/scientificamerican01071882-8.pdf.

Anonymous:1882:IPSa

[2592] Anonymous. Improved punch and shear. Scientific American, 46(1):8, January 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n1/pdf/scientificamerican01071882-8a.pdf.

Anonymous:1882:IFEa

[2593] Anonymous. Improved fire escape. Scientific American, 46(1):9, January 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n1/pdf/scientificamerican01071882-9.pdf.

Anonymous:1882:ITT

[2594] Anonymous. Improved thill tug. Scientific American, 46(1):9–10, January 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n1/pdf/scientificamerican01071882-9a.pdf.

Anonymous:1882:CRa

[2595] Anonymous. Communications received. Scientific American, 46(1):11, January 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n1/pdf/scientificamerican01071882-11b.pdf.

Anonymous:1882:IRIb

[2596] Anonymous. Inventions: Recent inventions. Scientific American, 46(10):147–148, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n10/pdf/scientificamerican03111882-
147a.pdf.

Anonymous:1882:INIb

[2597] Anonymous. Inventions: New inventions. Scientific American, 46 (10):150–151, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n10/pdf/scientificamerican03111882-
150b.pdf.

Anonymous:1882:BPb

[2598] Anonymous. Business and personal. Scientific American, 46(10):154, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n10/pdf/scientificamerican03111882-154.pdf.

Anonymous:1882:NQb

[2599] Anonymous. Notes and queries. Scientific American, 46(10):154, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n10/pdf/scientificamerican03111882-154a.pdf.

Anonymous:1882:BNBa

[2600] Anonymous. Books: New books and publications. Scientific American, 46(10):154–155, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n10/pdf/scientificamerican03111882-
154c.pdf.

Anonymous:1882:IIb

[2601] Anonymous. Index of inventions. Scientific American, 46(10):155–156, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n10/pdf/scientificamerican03111882-155.pdf.

Anonymous:1882:PNPa

[2602] Anonymous. Patent notes: Patents. Scientific American, 46(10):156, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n10/pdf/scientificamerican03111882-156.pdf.

Anonymous:1882:PFP

[2603] Anonymous. Preserving fence posts. Scientific American, 46(10):143, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n10/pdf/scientificamerican03111882-143.pdf.

Anonymous:1882:DPM

[2604] Anonymous. Duc’s patent mechanical atomizer. Scientific American, 46(10):143, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n10/pdf/scientificamerican03111882-
143a.pdf.

Anonymous:1882:IPPa

[2605] Anonymous. Invalidation of patents by public use. Scientific American, 46(10):144, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n10/pdf/scientificamerican03111882-
144.pdf.

Anonymous:1882:PC

[2606] Anonymous. The propagation of codfish. Scientific American, 46(10):144, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n10/pdf/scientificamerican03111882-
144a.pdf.

Anonymous:1882:SNP

[2607] Anonymous. The smoke nuisance. — problems for inventors. Scientific American, 46(10):144–145, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v46/n10/pdf/scientificamerican03111882-
144b.pdf.

Anonymous:1882:SBNa

[2608] Anonymous. Steam boiler notes. Scientific American, 46(10):145, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n10/pdf/scientificamerican03111882-145.pdf.

Riley:1882:CPM

[2609] C. V. Riley. Cotton picking by machinery. Scientific American, 46(10):145–146, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n10/pdf/scientificamerican03111882-
145a.pdf.

Anonymous:1882:MCM

[2610] Anonymous. M. Carpentier’s melograph. Scientific American, 46 (10):146–147, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n10/pdf/scientificamerican03111882-
146.pdf.

Anonymous:1882:IKB

[2611] Anonymous. Improved knockdown barrel. Scientific American, 46(10):147, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n10/pdf/scientificamerican03111882-
147.pdf.

Anonymous:1882:MIb

[2612] Anonymous. Mechanical inventions. Scientific American, 46(10): 148–149, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n10/pdf/scientificamerican03111882-
148.pdf.

Anonymous:1882:S

[2613] Anonymous. The skiagraph. Scientific American, 46(10):149–150, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n10/pdf/scientificamerican03111882-149.pdf.

Anonymous:1882:IV

[2614] Anonymous. Improved velocipede. Scientific American, 46(10):150, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n10/pdf/scientificamerican03111882-150.pdf.

Anonymous:1882:IBSa

[2615] Anonymous. Improved band sawing machine. Scientific American, 46(10):150–151, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n10/pdf/scientificamerican03111882-
150a.pdf.

Anonymous:1882:IBSb

[2616] Anonymous. Improved belt stretcher. Scientific American, 46(10):151, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n10/pdf/scientificamerican03111882-151.pdf.

Anonymous:1882:EHU

[2617] Anonymous. Effects of heat upon steel. Scientific American, 46 (10):151–152, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n10/pdf/scientificamerican03111882-
151a.pdf.

Anonymous:1882:AIa

[2618] Anonymous. Agricultural inventions. Scientific American, 46 (10):152–153, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n10/pdf/scientificamerican03111882-
152.pdf.

Anonymous:1882:DRPa

[2619] Anonymous. Decisions relating to patents. Scientific American, 46(10):153, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n10/pdf/scientificamerican03111882-
153.pdf.

Anonymous:1882:MIc

[2620] Anonymous. Miscellaneous inventions. Scientific American, 46(10):153, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n10/pdf/scientificamerican03111882-153a.pdf.

Anonymous:1882:CRb

[2621] Anonymous. Communications received. Scientific American, 46 (10):154, March 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n10/pdf/scientificamerican03111882-
154b.pdf.

Anonymous:1882:IRIc

[2622] Anonymous. Inventions: Recent inventions. Scientific American, 46(11):163, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n11/pdf/scientificamerican03181882-
163b.pdf.

Anonymous:1882:INIc

[2623] Anonymous. Inventions: New inventions. Scientific American, 46(11):166, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n11/pdf/scientificamerican03181882-
166.pdf.

Anonymous:1882:BPc

[2624] Anonymous. Business and personal. Scientific American, 46(11):171, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n11/pdf/scientificamerican03181882-171.pdf.

Anonymous:1882:NQc

[2625] Anonymous. Notes and queries. Scientific American, 46(11):171, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n11/pdf/scientificamerican03181882-171a.pdf.

Anonymous:1882:IIc

[2626] Anonymous. Index of inventions. Scientific American, 46(11):171–172, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n11/pdf/scientificamerican03181882-171c.pdf.

Anonymous:1882:PNPb

[2627] Anonymous. Patent notes: Patents. Scientific American, 46(11):172, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n11/pdf/scientificamerican03181882-172.pdf.

Anonymous:1882:PEL

[2628] Anonymous. Proposed electric light tower at New Orleans. Scientific American, 46(11):159, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n11/pdf/scientificamerican03181882-
159.pdf.

Anonymous:1882:PPLa

[2629] Anonymous. Proposed patent legislation. Scientific American, 46(11):160, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n11/pdf/scientificamerican03181882-
160.pdf.

Anonymous:1882:TB

[2630] Anonymous. The transformation of bacteria. Scientific American, 46(11):160, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n11/pdf/scientificamerican03181882-
160a.pdf.

Anonymous:1882:RCI

[2631] Anonymous. A railway-crossing invention wanted. Scientific American, 46(11):161–162, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n11/pdf/scientificamerican03181882-
161.pdf.

Anonymous:1882:NSSa

[2632] Anonymous. New side-saddle girth. Scientific American, 46(11):162, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n11/pdf/scientificamerican03181882-162.pdf.

Anonymous:1882:MId

[2633] Anonymous. Miscellaneous inventions. Scientific American, 46(11):162, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n11/pdf/scientificamerican03181882-162a.pdf.

Anonymous:1882:NSC

[2634] Anonymous. New safety car coupling. Scientific American, 46(11):163, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n11/pdf/scientificamerican03181882-163.pdf.

Anonymous:1882:INL

[2635] Anonymous. Improved nut lock. Scientific American, 46(11):163, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n11/pdf/scientificamerican03181882-163a.pdf.

Anonymous:1882:PHH

[2636] Anonymous. Pocket hanger for hats and garments. Scientific American, 46(11):163, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n11/pdf/scientificamerican03181882-
163c.pdf.

Hovey:1882:LCE

[2637] H. C. Hovey. The Luraz Caverns by electric light. Scientific American, 46(11):164, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n11/pdf/scientificamerican03181882-
164.pdf.

Anonymous:1882:MIe

[2638] Anonymous. Miscellaneous inventions. Scientific American, 46 (11):164–165, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n11/pdf/scientificamerican03181882-
164a.pdf.

Anonymous:1882:FMP

[2639] Anonymous. Failure of mats as a protection to the river front of New Orleans. — something new wanted. Scientific American, 46(11):165, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n11/pdf/scientificamerican03181882-165.pdf.

Anonymous:1882:NSA

[2640] Anonymous. New stock-alarm for locomotives. Scientific American, 46(11):166, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n11/pdf/scientificamerican03181882-
166a.pdf.

Anonymous:1882:NRH

[2641] Anonymous. Novel rein-holder. Scientific American, 46(11):166, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n11/pdf/scientificamerican03181882-166b.pdf.

Anonymous:1882:NTH

[2642] Anonymous. New tool holder for grindstones. Scientific American, 46(11):167, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n11/pdf/scientificamerican03181882-
167.pdf.

Anonymous:1882:FBH

[2643] Anonymous. Four-barreled hammerless gun. Scientific American, 46(11):167, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n11/pdf/scientificamerican03181882-
167a.pdf.

Anonymous:1882:IFEb

[2644] Anonymous. Improved fire escape. Scientific American, 46(11):167–168, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n11/pdf/scientificamerican03181882-167b.pdf.

Anonymous:1882:MIf

[2645] Anonymous. Mechanical inventions. Scientific American, 46(11):168, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n11/pdf/scientificamerican03181882-168.pdf.

Beard:1882:TSB

[2646] Daniel C. Beard. Two-story birds’ nests. Scientific American, 46(11):169, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n11/pdf/scientificamerican03181882-
169.pdf.

Anonymous:1882:EIa

[2647] Anonymous. Engineering inventions. Scientific American, 46(11):170, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n11/pdf/scientificamerican03181882-170.pdf.

Anonymous:1882:CRc

[2648] Anonymous. Communications received. Scientific American, 46 (11):171, March 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n11/pdf/scientificamerican03181882-
171b.pdf.

Anonymous:1882:Cb

[2649] Anonymous. Correspondence. Scientific American, 46(12):184, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n12/pdf/scientificamerican03251882-184.pdf.

Anonymous:1882:BPd

[2650] Anonymous. Business and personal. Scientific American, 46(12):187, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n12/pdf/scientificamerican03251882-187.pdf.

Anonymous:1882:NQd

[2651] Anonymous. Notes and queries. Scientific American, 46(12):187, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n12/pdf/scientificamerican03251882-187a.pdf.

Anonymous:1882:BNBb

[2652] Anonymous. Books: New books and publications. Scientific American, 46(12):187, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n12/pdf/scientificamerican03251882-
187b.pdf.

Anonymous:1882:IId

[2653] Anonymous. Index of inventions. Scientific American, 46(12):187–188, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n12/pdf/scientificamerican03251882-187c.pdf.

Anonymous:1882:PNPc

[2654] Anonymous. Patent notes: Patents. Scientific American, 46(12):188, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n12/pdf/scientificamerican03251882-188.pdf.

Anonymous:1882:ELS

[2655] Anonymous. Electric lighting for siege purposes. Scientific American, 46(12):175–176, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n12/pdf/scientificamerican03251882-
175.pdf.

Anonymous:1882:PAP

[2656] Anonymous. Proposed amendments of the Patent Laws. Scientific American, 46(12):176–177, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v46/n12/pdf/scientificamerican03251882-
176.pdf.

Anonymous:1882:CCa

[2657] Anonymous. The Chinese craze. Scientific American, 46(12):177, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n12/pdf/scientificamerican03251882-177.pdf.

Anonymous:1882:NRT

[2658] Anonymous. New railway track gauge and level. Scientific American, 46(12):178, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n12/pdf/scientificamerican03251882-
178.pdf.

Anonymous:1882:IGa

[2659] Anonymous. Improved gate. Scientific American, 46(12):178, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n12/pdf/scientificamerican03251882-178a.pdf.

Anonymous:1882:NHV

[2660] Anonymous. New hand vise. Scientific American, 46(12):179, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n12/pdf/scientificamerican03251882-179.pdf.

Anonymous:1882:IPPb

[2661] Anonymous. Improved pump piston. Scientific American, 46(12):179, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n12/pdf/scientificamerican03251882-179a.pdf.

Anonymous:1882:EIb

[2662] Anonymous. Engineering inventions. Scientific American, 46(12):179, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n12/pdf/scientificamerican03251882-179b.pdf.

Anonymous:1882:NDH

[2663] Anonymous. Novel door hanger. Scientific American, 46(12):179, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n12/pdf/scientificamerican03251882-179c.pdf.

Anonymous:1882:NSFa

[2664] Anonymous. New saw frame. Scientific American, 46(12):179, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n12/pdf/scientificamerican03251882-179d.pdf.

Deprez:1882:NFM

[2665] Marcel Deprez. On a new form of magneto-electric exploder. Scientific American, 46(12):180–181, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v46/n12/pdf/scientificamerican03251882-
180.pdf.

Anonymous:1882:SBNb

[2666] Anonymous. Steam boiler notes. Scientific American, 46(12):181–182, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n12/pdf/scientificamerican03251882-181.pdf.

Anonymous:1882:ISFa

[2667] Anonymous. Improved sash fastener. Scientific American, 46(12):182, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n12/pdf/scientificamerican03251882-182.pdf.

Anonymous:1882:CWH

[2668] Anonymous. A curiously worn hammer handle. Scientific American, 46(12):182, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n12/pdf/scientificamerican03251882-
182a.pdf.

Anonymous:1882:AIb

[2669] Anonymous. Agricultural inventions. Scientific American, 46(12):182, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n12/pdf/scientificamerican03251882-182b.pdf.

Anonymous:1882:NSBa

[2670] Anonymous. New show-box cover. Scientific American, 46(12):182–183, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n12/pdf/scientificamerican03251882-182c.pdf.

Anonymous:1882:FGT

[2671] Anonymous. Fresh grapes for the table. Scientific American, 46(12):183, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n12/pdf/scientificamerican03251882-183.pdf.

Anonymous:1882:SFC

[2672] Anonymous. Spongilla Fluviatilis — the cause of cucumber taste in water. Scientific American, 46(12):183, March 25, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v46/
n12/pdf/scientificamerican03251882-183a.pdf.

Anonymous:1882:MIg

[2673] Anonymous. Miscellaneous inventions. Scientific American, 46(12):183, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n12/pdf/scientificamerican03251882-183b.pdf.

Anonymous:1882:DRPb

[2674] Anonymous. Decisions relating to patents, trade marks etc. Scientific American, 46(12):184–185, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v46/n12/pdf/scientificamerican03251882-
184a.pdf.

Kroeh:1882:P

[2675] C. F. Kroeh. Parthenogenesis. Scientific American, 46(12):186, March 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n12/pdf/scientificamerican03251882-186.pdf.

Anonymous:1882:IRId

[2676] Anonymous. Inventions: Recent inventions. Scientific American, 46(13):194, April 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n13/pdf/scientificamerican04011882-
194.pdf.

Anonymous:1882:INId

[2677] Anonymous. Inventions: New inventions. Scientific American, 46(13):200, April 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n13/pdf/scientificamerican04011882-
200.pdf.

Anonymous:1882:BPe

[2678] Anonymous. Business and personal. Scientific American, 46(13):203, April 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n13/pdf/scientificamerican04011882-203.pdf.

Anonymous:1882:NQe

[2679] Anonymous. Notes and queries. Scientific American, 46(13):203, April 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n13/pdf/scientificamerican04011882-203a.pdf.

Anonymous:1882:IIe

[2680] Anonymous. Index of inventions. Scientific American, 46(13):203–204, April 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n13/pdf/scientificamerican04011882-203b.pdf.

Anonymous:1882:PNPd

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n13/pdf/scientificamerican04011882-204.pdf.

Anonymous:1882:ESPa

[2682] Anonymous. Engines of steamship Parisian. Scientific American, 46(13):191, April 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n13/pdf/scientificamerican04011882-
191.pdf.

Anonymous:1882:MM

[2683] Anonymous. A moral misapplied. Scientific American, 46(13):192, April 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n13/pdf/scientificamerican04011882-192.pdf.

Anonymous:1882:SGP

[2684] Anonymous. Specialists and general problems. Scientific American, 46(13):192, April 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n13/pdf/scientificamerican04011882-
192a.pdf.

Riley:1882:SCU

[2685] C. V. Riley. Silk culture in the United States. Scientific American, 46(13):193–194, April 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n13/pdf/scientificamerican04011882-
193.pdf.

Anonymous:1882:ISR

[2686] Anonymous. Improved stalk rake. Scientific American, 46(13):194, April 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n13/pdf/scientificamerican04011882-194a.pdf.

Anonymous:1882:IIPa

[2687] Anonymous. Improved insulator and protector for under-ground lines. Scientific American, 46(13):194, April 1, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v46/
n13/pdf/scientificamerican04011882-194b.pdf.

Anonymous:1882:MIh

[2688] Anonymous. Mechanical inventions. Scientific American, 46(13):194, April 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n13/pdf/scientificamerican04011882-194c.pdf.

Anonymous:1882:NTR

[2689] Anonymous. A new telephone receiver. Scientific American, 46(13):195, April 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n13/pdf/scientificamerican04011882-195.pdf.

Anonymous:1882:NMa

[2690] Anonymous. Novel motor. Scientific American, 46(13):195, April 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n13/pdf/scientificamerican04011882-195a.pdf.

Anonymous:1882:APAa

[2691] Anonymous. Aspects of the planets for April. Scientific American, 46(13):196, April 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n13/pdf/scientificamerican04011882-
196.pdf.

Anonymous:1882:MIi

[2692] Anonymous. Miscellaneous inventions. Scientific American, 46 (13):196–198, April 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n13/pdf/scientificamerican04011882-
196a.pdf.

Anonymous:1882:IJB

[2693] Anonymous. Improved journal box. Scientific American, 46(13):198, April 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n13/pdf/scientificamerican04011882-198.pdf.

Anonymous:1882:NFEa

[2694] Anonymous. A new fire escape. Scientific American, 46(13):198–199, April 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n13/pdf/scientificamerican04011882-198a.pdf.

Anonymous:1882:DSP

[2695] Anonymous. Dry steam portable engine. Scientific American, 46 (13):199–200, April 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n13/pdf/scientificamerican04011882-
199.pdf.

Anonymous:1882:MMB

[2696] Anonymous. Mound-making birds of Australasia. Scientific American, 46(13):201–202, April 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n13/pdf/scientificamerican04011882-
201.pdf.

Anonymous:1882:IRIe

[2697] Anonymous. Inventions: Recent inventions. Scientific American, 46(14):218, April 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n14/pdf/scientificamerican04081882-
218.pdf.

Anonymous:1882:BPf

[2698] Anonymous. Business and personal. Scientific American, 46(14):219, April 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n14/pdf/scientificamerican04081882-219.pdf.

Anonymous:1882:NQf

[2699] Anonymous. Notes and queries. Scientific American, 46(14):219, April 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n14/pdf/scientificamerican04081882-219a.pdf.

Anonymous:1882:IIf

[2700] Anonymous. Index of inventions. Scientific American, 46(14):219–220, April 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n14/pdf/scientificamerican04081882-219b.pdf.

Anonymous:1882:PNPe

[2701] Anonymous. Patent notes: Patents. Scientific American, 46(14):220, April 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n14/pdf/scientificamerican04081882-220.pdf.

Anonymous:1882:NMC

[2702] Anonymous. New machine for cleaning and separating the fibers of leaves. Scientific American, 46(14):207, April 8, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v46/
n14/pdf/scientificamerican04081882-207.pdf.

Anonymous:1882:SPG

[2703] Anonymous. The steering and propelling gear of the alarm. Scientific American, 46(14):208, April 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n14/pdf/scientificamerican04081882-
208.pdf.

Anonymous:1882:VSB

[2704] Anonymous. Virchow on soups and broths. Scientific American, 46(14):208–209, April 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n14/pdf/scientificamerican04081882-
208a.pdf.

Anonymous:1882:MIj

[2705] Anonymous. Mechanical inventions. Scientific American, 46(14):210, April 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n14/pdf/scientificamerican04081882-210.pdf.

Anonymous:1882:ISEa

[2706] Anonymous. Improved steam engine. Scientific American, 46(14):210, April 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n14/pdf/scientificamerican04081882-210a.pdf.

Anonymous:1882:ISa

[2707] Anonymous. Improved spindle. Scientific American, 46(14):210–211, April 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n14/pdf/scientificamerican04081882-210b.pdf.

Anonymous:1882:NPA

[2708] Anonymous. New pegging awl. Scientific American, 46(14):211, April 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n14/pdf/scientificamerican04081882-211.pdf.

Anonymous:1882:NSM

[2709] Anonymous. Novel striking mechanism for electric bells. Scientific American, 46(14):211, April 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n14/pdf/scientificamerican04081882-
211a.pdf.

Anonymous:1882:J

[2710] Anonymous. Jupiter. Scientific American, 46(14):212, April 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n14/pdf/scientificamerican04081882-212.pdf.

Anonymous:1882:RDRa

[2711] Anonymous. Recent decisions relating to patents. Scientific American, 46(14):213, April 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n14/pdf/scientificamerican04081882-
213.pdf.

Anonymous:1882:NSBb

[2712] Anonymous. New system of brick making. Scientific American, 46(14):214–215, April 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n14/pdf/scientificamerican04081882-
214.pdf.

Anonymous:1882:TY

[2713] Anonymous. Tapir and young. Scientific American, 46(14):215, April 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n14/pdf/scientificamerican04081882-215.pdf.

Anonymous:1882:NHNa

[2714] Anonymous. Natural history notes. Scientific American, 46(14): 216–217, April 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n14/pdf/scientificamerican04081882-
216.pdf.

Anonymous:1882:MIk

[2715] Anonymous. Miscellaneous inventions. Scientific American, 46 (14):217–218, April 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n14/pdf/scientificamerican04081882-
217.pdf.

Anonymous:1882:INIe

[2716] Anonymous. Inventions: New inventions. Scientific American, 46(15):227, April 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n15/pdf/scientificamerican04151882-
227.pdf.

Anonymous:1882:RIa

[2717] Anonymous. Recent inventions. Scientific American, 46(15):228, April 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n15/pdf/scientificamerican04151882-228.pdf.

Anonymous:1882:Cd

[2718] Anonymous. Correspondence. Scientific American, 46(15):228, April 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n15/pdf/scientificamerican04151882-228a.pdf.

Anonymous:1882:BPg

[2719] Anonymous. Business and personal. Scientific American, 46(15):234, April 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n15/pdf/scientificamerican04151882-234.pdf.

Anonymous:1882:NQg

[2720] Anonymous. Notes and queries. Scientific American, 46(15):234–235, April 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n15/pdf/scientificamerican04151882-234a.pdf.

Anonymous:1882:BNBc

[2721] Anonymous. Books: New books and publications. Scientific American, 46(15):235, April 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n15/pdf/scientificamerican04151882-
235.pdf.

Anonymous:1882:IIg

[2722] Anonymous. Index of inventions. Scientific American, 46(15):235–236, April 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n15/pdf/scientificamerican04151882-235a.pdf.

Anonymous:1882:PNPf

[2723] Anonymous. Patent notes: Patents. Scientific American, 46(15):236, April 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n15/pdf/scientificamerican04151882-236.pdf.

Anonymous:1882:NFI

[2724] Anonymous. A new field for invention. Scientific American, 46(15):224, April 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n15/pdf/scientificamerican04151882-224.pdf.

Anonymous:1882:Ce

[2725] Anonymous. Comet a 1882. Scientific American, 46(15):224, April 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n15/pdf/scientificamerican04151882-224a.pdf.

Anonymous:1882:FCA

[2726] Anonymous. Fish culture in America. Scientific American, 46 (15):224–225, April 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n15/pdf/scientificamerican04151882-
224b.pdf.

Anonymous:1882:LM

[2727] Anonymous. Locomotives and malaria. Scientific American, 46(15):225, April 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n15/pdf/scientificamerican04151882-225.pdf.

Anonymous:1882:FII

[2728] Anonymous. Foods for infants and invalids. Scientific American, 46(15):225–226, April 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n15/pdf/scientificamerican04151882-
225a.pdf.

Anonymous:1882:MIl

[2729] Anonymous. Miscellaneous inventions. Scientific American, 46(15):226, April 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n15/pdf/scientificamerican04151882-226.pdf.

Anonymous:1882:GCP

[2730] Anonymous. The gravity coal piers at Hoboken. Scientific American, 46(15):226–227, April 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n15/pdf/scientificamerican04151882-
226a.pdf.

Anonymous:1882:SBNc

[2731] Anonymous. Steam boiler notes. Scientific American, 46(15):229, April 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n15/pdf/scientificamerican04151882-229.pdf.

Anonymous:1882:EIc

[2732] Anonymous. Engineering inventions. Scientific American, 46(15):229, April 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n15/pdf/scientificamerican04151882-229a.pdf.

Anonymous:1882:IEB

[2733] Anonymous. Interesting experimental boiler explosion. Scientific American, 46(15):230–231, April 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n15/pdf/scientificamerican04151882-
230.pdf.

Anonymous:1882:PBI

[2734] Anonymous. Powerful bar iron shear. Scientific American, 46(15):231, April 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n15/pdf/scientificamerican04151882-231.pdf.

Anonymous:1882:IPPc

[2735] Anonymous. Improved plate planer. Scientific American, 46(15):231, April 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n15/pdf/scientificamerican04151882-231a.pdf.

Anonymous:1882:MIm

[2736] Anonymous. Mechanical inventions. Scientific American, 46(15): 231–232, April 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n15/pdf/scientificamerican04151882-
231b.pdf.

Anonymous:1882:AIc

[2737] Anonymous. Agricultural inventions. Scientific American, 46(15):233, April 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n15/pdf/scientificamerican04151882-233.pdf.

Anonymous:1882:INIf

[2738] Anonymous. Inventions: New inventions. Scientific American, 46 (16):242–243, April 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n16/pdf/scientificamerican04221882-
242b.pdf.

Anonymous:1882:Cf

[2739] Anonymous. Correspondence. Scientific American, 46(16):245, April 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n16/pdf/scientificamerican04221882-245.pdf.

Anonymous:1882:IRIf

[2740] Anonymous. Inventions: Recent inventions. Scientific American, 46(16):245, April 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n16/pdf/scientificamerican04221882-
245a.pdf.

Anonymous:1882:BPh

[2741] Anonymous. Business and personal. Scientific American, 46(16):251, April 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n16/pdf/scientificamerican04221882-251.pdf.

Anonymous:1882:NQh

[2742] Anonymous. Notes and queries. Scientific American, 46(16):251, April 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n16/pdf/scientificamerican04221882-251a.pdf.

Anonymous:1882:BNBd

[2743] Anonymous. Books: New books and publications. Scientific American, 46(16):251, April 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n16/pdf/scientificamerican04221882-
251b.pdf.

Anonymous:1882:IIh

[2744] Anonymous. Index of inventions. Scientific American, 46(16):251–252, April 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n16/pdf/scientificamerican04221882-251c.pdf.

Anonymous:1882:PNPg

[2745] Anonymous. Patent notes: Patents. Scientific American, 46(16):252, April 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n16/pdf/scientificamerican04221882-252.pdf.

Anonymous:1882:NHS

[2746] Anonymous. The new high-service pumping works, New York City. Scientific American, 46(16):239–240, April 22, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v46/
n16/pdf/scientificamerican04221882-239.pdf.

Anonymous:1882:UCN

[2747] Anonymous. Uniformity in chemical nomenclature. Scientific American, 46(16):240, April 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n16/pdf/scientificamerican04221882-
240.pdf.

Anonymous:1882:CCF

[2748] Anonymous. Carelessness as a cause of fire. Scientific American, 46(16):240–241, April 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n16/pdf/scientificamerican04221882-
240a.pdf.

Anonymous:1882:TC

[2749] Anonymous. Testing chocolate. Scientific American, 46(16):241, April 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n16/pdf/scientificamerican04221882-241.pdf.

Anonymous:1882:NGS

[2750] Anonymous. New gold separator. Scientific American, 46(16):242, April 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n16/pdf/scientificamerican04221882-242.pdf.

Anonymous:1882:IAG

[2751] Anonymous. Improved automatic gate. Scientific American, 46(16):242, April 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n16/pdf/scientificamerican04221882-242a.pdf.

Anonymous:1882:NBFa

[2752] Anonymous. Novel belt fastener. Scientific American, 46(16):243, April 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n16/pdf/scientificamerican04221882-243.pdf.

Anonymous:1882:CAM

[2753] Anonymous. Compressed-air machine used in the channel tunnel. Scientific American, 46(16):243–244, April 22, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v46/
n16/pdf/scientificamerican04221882-243a.pdf.

Anonymous:1882:RDRb

[2754] Anonymous. Recent decisions relating to patents. Scientific American, 46(16):244, April 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n16/pdf/scientificamerican04221882-
244.pdf.

Anonymous:1882:ADA

[2755] Anonymous. The attempt to destroy the Andre Monument. Scientific American, 46(16):245, April 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n16/pdf/scientificamerican04221882-
245b.pdf.

Anonymous:1882:PLE

[2756] Anonymous. Plow for laying electric cables. Scientific American, 46(16):246, April 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n16/pdf/scientificamerican04221882-
246.pdf.

Anonymous:1882:NPTa

[2757] Anonymous. New plant-trimming shears. Scientific American, 46(16):246, April 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n16/pdf/scientificamerican04221882-
246a.pdf.

Anonymous:1882:MIn

[2758] Anonymous. Miscellaneous inventions. Scientific American, 46(16):246, April 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n16/pdf/scientificamerican04221882-246b.pdf.

Holder:1882:RBW

[2759] J. B. Holder. A rediscovery — “the black whale”. Scientific American, 46(16):247–248, April 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n16/pdf/scientificamerican04221882-
247.pdf.

Anonymous:1882:MIo

[2760] Anonymous. Mechanical inventions. Scientific American, 46(16):248, April 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n16/pdf/scientificamerican04221882-248.pdf.

Anonymous:1882:EId

[2761] Anonymous. Engineering inventions. Scientific American, 46(16):249, April 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n16/pdf/scientificamerican04221882-249.pdf.

Anonymous:1882:AId

[2762] Anonymous. Agricultural inventions. Scientific American, 46(16):250, April 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n16/pdf/scientificamerican04221882-250.pdf.

Anonymous:1882:Cg

[2763] Anonymous. Correspondence. Scientific American, 46(17):260–261, April 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n17/pdf/scientificamerican04291882-260a.pdf.

Anonymous:1882:BPi

[2764] Anonymous. Business and personal. Scientific American, 46(17):267, April 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n17/pdf/scientificamerican04291882-267.pdf.

Anonymous:1882:NQi

[2765] Anonymous. Notes and queries. Scientific American, 46(17):267, April 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n17/pdf/scientificamerican04291882-267a.pdf.

Anonymous:1882:IIi

[2766] Anonymous. Index of inventions. Scientific American, 46(17):267–268, April 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n17/pdf/scientificamerican04291882-267c.pdf.

Anonymous:1882:PNPh

[2767] Anonymous. Patent notes: Patents. Scientific American, 46(17):268, April 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n17/pdf/scientificamerican04291882-268.pdf.

Anonymous:1882:BIB

[2768] Anonymous. The Blackwell’s Island Bridge. Scientific American, 46(17):255, April 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n17/pdf/scientificamerican04291882-
255.pdf.

Anonymous:1882:CDa

[2769] Anonymous. Charles Darwin. Scientific American, 46(17):256, April 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n17/pdf/scientificamerican04291882-256.pdf.

Anonymous:1882:LAS

[2770] Anonymous. The loss of the Arctic search steamer Rodgers. Scientific American, 46(17):256, April 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n17/pdf/scientificamerican04291882-
256a.pdf.

Anonymous:1882:APMa

[2771] Anonymous. Aspects of the planets for May. Scientific American, 46(17):257–258, April 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n17/pdf/scientificamerican04291882-
257.pdf.

Anonymous:1882:NEL

[2772] Anonymous. New electric log. Scientific American, 46(17):258, April 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n17/pdf/scientificamerican04291882-258.pdf.

Anonymous:1882:NS

[2773] Anonymous. A novel sled. Scientific American, 46(17):258, April 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n17/pdf/scientificamerican04291882-258a.pdf.

Anonymous:1882:NGD

[2774] Anonymous. New grain drier. Scientific American, 46(17):258, April 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n17/pdf/scientificamerican04291882-258b.pdf.

Anonymous:1882:NLB

[2775] Anonymous. New life boat plug. Scientific American, 46(17):259, April 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n17/pdf/scientificamerican04291882-259.pdf.

Anonymous:1882:NAT

[2776] Anonymous. New acoustic telephone. Scientific American, 46 (17):259–260, April 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n17/pdf/scientificamerican04291882-
259a.pdf.

Anonymous:1882:EIe

[2777] Anonymous. Engineering inventions. Scientific American, 46(17):260, April 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n17/pdf/scientificamerican04291882-260.pdf.

Anonymous:1882:MIp

[2778] Anonymous. Mechanical inventions. Scientific American, 46(17):261, April 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n17/pdf/scientificamerican04291882-261.pdf.

Anonymous:1882:NOM

[2779] Anonymous. A new ore mill. Scientific American, 46(17):262, April 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n17/pdf/scientificamerican04291882-262.pdf.

Anonymous:1882:MIq

[2780] Anonymous. Miscellaneous inventions. Scientific American, 46(17):262, April 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n17/pdf/scientificamerican04291882-262a.pdf.

Beard:1882:TF

[2781] Daniel C. Beard. The tile fish. Scientific American, 46(17):263, April 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n17/pdf/scientificamerican04291882-263.pdf.

Anonymous:1882:JLM

[2782] Anonymous. Jerboas, or leaping mice. Scientific American, 46 (17):263–265, April 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n17/pdf/scientificamerican04291882-
263a.pdf.

Anonymous:1882:MIr

[2783] Anonymous. Miscellaneous inventions. Scientific American, 46 (17):265–266, April 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n17/pdf/scientificamerican04291882-
265.pdf.

Anonymous:1882:CRd

[2784] Anonymous. Communications received. Scientific American, 46 (17):267, April 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n17/pdf/scientificamerican04291882-
267b.pdf.

Anonymous:1882:IRIg

[2785] Anonymous. Inventions: Recent inventions. Scientific American, 46(18):274, May 6, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n18/pdf/scientificamerican05061882-
274c.pdf.

Anonymous:1882:INIg

[2786] Anonymous. Inventions: New inventions. Scientific American, 46(18):278, May 6, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n18/pdf/scientificamerican05061882-
278b.pdf.

Anonymous:1882:BPj

[2787] Anonymous. Business and personal. Scientific American, 46(18):283, May 6, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n18/pdf/scientificamerican05061882-283.pdf.

Anonymous:1882:NQj

[2788] Anonymous. Notes and queries. Scientific American, 46(18):283, May 6, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n18/pdf/scientificamerican05061882-283a.pdf.

Anonymous:1882:IIj

[2789] Anonymous. Index of inventions. Scientific American, 46(18):283–284, May 6, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n18/pdf/scientificamerican05061882-283b.pdf.

Anonymous:1882:PNPi

[2790] Anonymous. Patent notes: Patents. Scientific American, 46(18):284, May 6, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n18/pdf/scientificamerican05061882-284.pdf.

Anonymous:1882:ESPb

[2791] Anonymous. Engines of the steamship Parisian. Scientific American, 46(18):271, May 6, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n18/pdf/scientificamerican05061882-
271.pdf.

Anonymous:1882:LHP

[2792] Anonymous. Light, heat, and power at little cost. Scientific American, 46(18):272, May 6, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n18/pdf/scientificamerican05061882-
272.pdf.

Anonymous:1882:PSIa

[2793] Anonymous. The protection of small inventions. Scientific American, 46(18):272, May 6, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n18/pdf/scientificamerican05061882-
272a.pdf.

Anonymous:1882:BBP

[2794] Anonymous. Bessemer’s bronze powder — how the public gains by granting patents for invention. Scientific American, 46(18):272–273, May 6, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n18/pdf/scientificamerican05061882-272b.pdf.

Anonymous:1882:MIs

[2795] Anonymous. Mechanical inventions. Scientific American, 46(18):274, May 6, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n18/pdf/scientificamerican05061882-274.pdf.

Anonymous:1882:IWMa

[2796] Anonymous. Improved washing machine. Scientific American, 46 (18):274, May 6, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n18/pdf/scientificamerican05061882-
274a.pdf.

Anonymous:1882:SCW

[2797] Anonymous. Sir Charles Wyville Thomson. Scientific American, 46(18):274, May 6, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n18/pdf/scientificamerican05061882-
274b.pdf.

Anonymous:1882:IGE

[2798] Anonymous. Improved gas engine. Scientific American, 46(18):275, May 6, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n18/pdf/scientificamerican05061882-275.pdf.

Anonymous:1882:SM

[2799] Anonymous. Sewage machine. Scientific American, 46(18):275, May 6, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n18/pdf/scientificamerican05061882-275a.pdf.

Anonymous:1882:EIf

[2800] Anonymous. Engineering inventions. Scientific American, 46(18): 275–278, May 6, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n18/pdf/scientificamerican05061882-
275b.pdf.

Anonymous:1882:IBF

[2801] Anonymous. Improved boiler and furnace. Scientific American, 46(18):278, May 6, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n18/pdf/scientificamerican05061882-
278.pdf.

Anonymous:1882:NSL

[2802] Anonymous. Novel ship’s log. Scientific American, 46(18):278, May 6, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n18/pdf/scientificamerican05061882-278a.pdf.

Anonymous:1882:SCB

[2803] Anonymous. Some curious bugs. Scientific American, 46(18):279, May 6, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n18/pdf/scientificamerican05061882-279.pdf.

Anonymous:1882:MIt

[2804] Anonymous. Miscellaneous inventions. Scientific American, 46(18):279, May 6, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n18/pdf/scientificamerican05061882-279a.pdf.

Anonymous:1882:SGH

[2805] Anonymous. Sugar — a glance at its history, production, and manufacture. Scientific American, 46(18):280–282, May 6, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v46/
n18/pdf/scientificamerican05061882-280.pdf.

Anonymous:1882:BPk

[2806] Anonymous. Business and personal. Scientific American, 46(19):303, May 13, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n19/pdf/scientificamerican05131882-303.pdf.

Anonymous:1882:BNBe

[2807] Anonymous. Books: New books and publications. Scientific American, 46(19):303, May 13, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n19/pdf/scientificamerican05131882-
303a.pdf.

Anonymous:1882:NQk

[2808] Anonymous. Notes and queries. Scientific American, 46(19):303, May 13, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n19/pdf/scientificamerican05131882-303b.pdf.

Anonymous:1882:IIk

[2809] Anonymous. Index of inventions. Scientific American, 46(19):303–304, May 13, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n19/pdf/scientificamerican05131882-303d.pdf.

Anonymous:1882:PNPj

[2810] Anonymous. Patent notes: Patents. Scientific American, 46(19):304, May 13, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n19/pdf/scientificamerican05131882-304.pdf.

Anonymous:1882:KV

[2811] Anonymous. The Kinzua Viaduct. Scientific American, 46(19):287–288, May 13, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n19/pdf/scientificamerican05131882-287.pdf.

Anonymous:1882:LD

[2812] Anonymous. The Libyan desert. Scientific American, 46(19):288–290, May 13, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n19/pdf/scientificamerican05131882-288.pdf.

Anonymous:1882:NSEa

[2813] Anonymous. A new steam engine. Scientific American, 46(19):290–291, May 13, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n19/pdf/scientificamerican05131882-290.pdf.

Eckart:1882:CEP

[2814] W. R. Eckart. Chronograph for engineering purposes. Scientific American, 46(19):291–293, May 13, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n19/pdf/scientificamerican05131882-
291.pdf.

Anonymous:1882:ICCb

[2815] Anonymous. Improved car coupling. Scientific American, 46(19):293, May 13, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n19/pdf/scientificamerican05131882-293.pdf.

Anonymous:1882:BHV

[2816] Anonymous. Barker’s heating and ventilating apparatus. Scientific American, 46(19):293, May 13, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n19/pdf/scientificamerican05131882-
293a.pdf.

Anonymous:1882:MIu

[2817] Anonymous. Miscellaneous inventions. Scientific American, 46(19):293, May 13, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n19/pdf/scientificamerican05131882-293b.pdf.

Anonymous:1882:ISSa

[2818] Anonymous. Improved sheep shears. Scientific American, 46(19):294, May 13, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n19/pdf/scientificamerican05131882-294.pdf.

Anonymous:1882:IFEc

[2819] Anonymous. Improved fruit evaporator. Scientific American, 46 (19):294–295, May 13, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n19/pdf/scientificamerican05131882-
294a.pdf.

Anonymous:1882:PHD

[2820] Anonymous. Peruke-horned doe. Scientific American, 46(19):295, May 13, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n19/pdf/scientificamerican05131882-295.pdf.

Anonymous:1882:ETW

[2821] Anonymous. The electric telegraph and the woodpeckers. Scientific American, 46(19):295–298, May 13, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v46/n19/pdf/scientificamerican05131882-
295a.pdf.

Anonymous:1882:EIg

[2822] Anonymous. Engineering invention. Scientific American, 46(19):299, May 13, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n19/pdf/scientificamerican05131882-299.pdf.

Anonymous:1882:MIv

[2823] Anonymous. Mechanical inventions. Scientific American, 46(19): 299–300, May 13, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n19/pdf/scientificamerican05131882-
299a.pdf.

Anonymous:1882:EIh

[2824] Anonymous. Electrical inventions. Scientific American, 46(19):300, May 13, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n19/pdf/scientificamerican05131882-300.pdf.

Anonymous:1882:MIw

[2825] Anonymous. Metallurgical invention. Scientific American, 46(19):300, May 13, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n19/pdf/scientificamerican05131882-300a.pdf.

Anonymous:1882:MIx

[2826] Anonymous. Miscellaneous inventions. Scientific American, 46 (19):300–302, May 13, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n19/pdf/scientificamerican05131882-
300b.pdf.

Anonymous:1882:CRe

[2827] Anonymous. Communications received. Scientific American, 46 (19):303, May 13, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n19/pdf/scientificamerican05131882-
303c.pdf.

Anonymous:1882:IRIh

[2828] Anonymous. Inventions: Recent inventions. Scientific American, 46(2):23–24, January 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n2/pdf/scientificamerican01141882-
23b.pdf.

Anonymous:1882:NIa

[2829] Anonymous. New inventions. Scientific American, 46(2):24–25, January 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n2/pdf/scientificamerican01141882-24.pdf.

Anonymous:1882:BPl

[2830] Anonymous. Business and personal. Scientific American, 46(2):26, January 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n2/pdf/scientificamerican01141882-26.pdf.

Anonymous:1882:NQl

[2831] Anonymous. Notes and queries. Scientific American, 46(2):26, January 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n2/pdf/scientificamerican01141882-26a.pdf.

Anonymous:1882:BNBf

[2832] Anonymous. Books: New books and publications. Scientific American, 46(2):26, January 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n2/pdf/scientificamerican01141882-
26c.pdf.

Anonymous:1882:IIl

[2833] Anonymous. Index of inventions. Scientific American, 46(2):26–28, January 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n2/pdf/scientificamerican01141882-26d.pdf.

Anonymous:1882:PNPk

[2834] Anonymous. Patent notes: Patents. Scientific American, 46(2):28, January 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n2/pdf/scientificamerican01141882-28.pdf.

Anonymous:1882:IPVa

[2835] Anonymous. Innocent patentee versus innocent purchaser. Scientific American, 46(2):16, January 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n2/pdf/scientificamerican01141882-
16.pdf.

Anonymous:1882:JWD

[2836] Anonymous. John William Draper. Scientific American, 46(2):16, January 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n2/pdf/scientificamerican01141882-16a.pdf.

Anonymous:1882:EUB

[2837] Anonymous. Explosion of an upright boiler. Scientific American, 46(2):17–18, January 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n2/pdf/scientificamerican01141882-
17.pdf.

Anonymous:1882:ECTa

[2838] Anonymous. Earth connections for telegraph wires and lightning rods. Scientific American, 46(2):18, January 14, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v46/
n2/pdf/scientificamerican01141882-18.pdf.

Anonymous:1882:PCB

[2839] Anonymous. Portable combined bath tub and heater. Scientific American, 46(2):18, January 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n2/pdf/scientificamerican01141882-
18a.pdf.

Anonymous:1882:NTC

[2840] Anonymous. New toilet combination. Scientific American, 46 (2):18, January 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n2/pdf/scientificamerican01141882-
18b.pdf.

Anonymous:1882:MIy

[2841] Anonymous. Miscellaneous inventions. Scientific American, 46 (2):18, January 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n2/pdf/scientificamerican01141882-
18c.pdf.

Anonymous:1882:AINa

[2842] Anonymous. American industries — no. 80. Scientific American, 46(2):19–22, January 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n2/pdf/scientificamerican01141882-
19.pdf.

Anonymous:1882:IUC

[2843] Anonymous. Improved universal chuck. Scientific American, 46 (2):22, January 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n2/pdf/scientificamerican01141882-
22.pdf.

Anonymous:1882:IK

[2844] Anonymous. Improvement in kettles. Scientific American, 46 (2):22–23, January 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n2/pdf/scientificamerican01141882-
22a.pdf.

Anonymous:1882:FSE

[2845] Anonymous. Finger-supporting and exercising device. Scientific American, 46(2):23, January 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n2/pdf/scientificamerican01141882-
23.pdf.

Anonymous:1882:ISb

[2846] Anonymous. Improved t-square. Scientific American, 46(2):23, January 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n2/pdf/scientificamerican01141882-23a.pdf.

Anonymous:1882:CRf

[2847] Anonymous. Communication received. Scientific American, 46 (2):26, January 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n2/pdf/scientificamerican01141882-
26b.pdf.

Anonymous:1882:BPm

[2848] Anonymous. Business and personal. Scientific American, 46(20):322, May 20, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n20/pdf/scientificamerican05201882-322.pdf.

Anonymous:1882:BNBg

[2849] Anonymous. Books: New books and publications. Scientific American, 46(20):322, May 20, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n20/pdf/scientificamerican05201882-
322a.pdf.

Anonymous:1882:NQm

[2850] Anonymous. Notes and queries. Scientific American, 46(20):322–323, May 20, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n20/pdf/scientificamerican05201882-322b.pdf.

Anonymous:1882:IIm

[2851] Anonymous. Index of inventions. Scientific American, 46(20):323–324, May 20, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n20/pdf/scientificamerican05201882-323.pdf.

Anonymous:1882:PNPl

[2852] Anonymous. Patent notes: Patents. Scientific American, 46(20):324, May 20, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n20/pdf/scientificamerican05201882-324.pdf.

Anonymous:1882:HLM

[2853] Anonymous. Hydraulic lift on the Morris Canal. Scientific American, 46(20):307, May 20, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n20/pdf/scientificamerican05201882-
307.pdf.

Anonymous:1882:KWC

[2854] Anonymous. Klinkerfues’ weather compass. Scientific American, 46(20):308, May 20, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n20/pdf/scientificamerican05201882-
308.pdf.

Anonymous:1882:SCG

[2855] Anonymous. Suffocation by coal gas. Scientific American, 46(20):308, May 20, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n20/pdf/scientificamerican05201882-308a.pdf.

Anonymous:1882:MC

[2856] Anonymous. Metallic cæsium. Scientific American, 46(20):308, May 20, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n20/pdf/scientificamerican05201882-308b.pdf.

Gissler:1882:PC

[2857] C. F. Gissler. The parasite of the clam. Scientific American, 46(20):308–309, May 20, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n20/pdf/scientificamerican05201882-
308c.pdf.

Anonymous:1882:DRPc

[2858] Anonymous. Decisions relating to patents. Scientific American, 46(20):309, May 20, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n20/pdf/scientificamerican05201882-
309.pdf.

Anonymous:1882:SEA

[2859] Anonymous. Stationary engine with adjustable cut-off valves. Scientific American, 46(20):310–311, May 20, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v46/n20/pdf/scientificamerican05201882-
310.pdf.

Anonymous:1882:IPa

[2860] Anonymous. Improved plow. Scientific American, 46(20):311–313, May 20, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n20/pdf/scientificamerican05201882-311.pdf.

Anonymous:1882:NCCa

[2861] Anonymous. New cotton chopper. Scientific American, 46(20):314, May 20, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n20/pdf/scientificamerican05201882-314.pdf.

Anonymous:1882:NDR

[2862] Anonymous. New disintegrating riffles. Scientific American, 46(20):314, May 20, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n20/pdf/scientificamerican05201882-314a.pdf.

Holder:1882:GAE

[2863] C. F. Holder. The great ant-eater and its young. Scientific American, 46(20):315–317, May 20, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n20/pdf/scientificamerican05201882-
315.pdf.

Anonymous:1882:EIi

[2864] Anonymous. Engineering inventions. Scientific American, 46(20):318, May 20, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n20/pdf/scientificamerican05201882-318.pdf.

Anonymous:1882:MIz

[2865] Anonymous. Mechanical inventions. Scientific American, 46(20): 318–319, May 20, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n20/pdf/scientificamerican05201882-
318a.pdf.

Anonymous:1882:EIj

[2866] Anonymous. Electrical invention. Scientific American, 46(20):319, May 20, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n20/pdf/scientificamerican05201882-319.pdf.

Anonymous:1882:AIe

[2867] Anonymous. Agricultural inventions. Scientific American, 46 (20):319–320, May 20, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n20/pdf/scientificamerican05201882-
319a.pdf.

Anonymous:1882:MIba

[2868] Anonymous. Miscellaneous inventions. Scientific American, 46 (20):320–321, May 20, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n20/pdf/scientificamerican05201882-
320.pdf.

Anonymous:1882:BPn

[2869] Anonymous. Business and personal. Scientific American, 46(21):339, May 27, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n21/pdf/scientificamerican05271882-339.pdf.

Anonymous:1882:NQn

[2870] Anonymous. Notes and queries. Scientific American, 46(21):339, May 27, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n21/pdf/scientificamerican05271882-339a.pdf.

Anonymous:1882:IIn

[2871] Anonymous. Index of inventions. Scientific American, 46(21):339–340, May 27, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n21/pdf/scientificamerican05271882-339b.pdf.

Anonymous:1882:PNPm

[2872] Anonymous. Patent notes: Patents. Scientific American, 46(21):340, May 27, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n21/pdf/scientificamerican05271882-340.pdf.

Anonymous:1882:PLA

[2873] Anonymous. Preparation of lactic acid. Scientific American, 46(21):327, May 27, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n21/pdf/scientificamerican05271882-327.pdf.

Anonymous:1882:LSA

[2874] Anonymous. Life-saving apparatus at the recent Naval and Submarine Exhibition. Scientific American, 46(21):327, May 27, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v46/
n21/pdf/scientificamerican05271882-327a.pdf.

Anonymous:1882:NPL

[2875] Anonymous. Nullification of the Patent Laws. Scientific American, 46(21):328, May 27, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n21/pdf/scientificamerican05271882-
328.pdf.

Anonymous:1882:APJa

[2876] Anonymous. Aspects of the planets for June. Scientific American, 46(21):328–329, May 27, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n21/pdf/scientificamerican05271882-
328a.pdf.

Anonymous:1882:CEa

[2877] Anonymous. Congressional engineering. Scientific American, 46 (21):329, May 27, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n21/pdf/scientificamerican05271882-
329.pdf.

Anonymous:1882:TBI

[2878] Anonymous. Torpedo boat for the Italian Government. Scientific American, 46(21):330, May 27, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n21/pdf/scientificamerican05271882-
330.pdf.

Anonymous:1882:NBC

[2879] Anonymous. New bottle cleaner. Scientific American, 46(21):330–331, May 27, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n21/pdf/scientificamerican05271882-330a.pdf.

Anonymous:1882:NWB

[2880] Anonymous. Novel window blind operator. Scientific American, 46(21):331, May 27, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n21/pdf/scientificamerican05271882-
331.pdf.

Anonymous:1882:NBA

[2881] Anonymous. Novel burglar alarm. Scientific American, 46(21):331, May 27, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n21/pdf/scientificamerican05271882-331a.pdf.

Anonymous:1882:NNLa

[2882] Anonymous. New nut lock. Scientific American, 46(21):331, May 27, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n21/pdf/scientificamerican05271882-331b.pdf.

Hopkins:1882:SSR

[2883] M. Hopkins. Simple sound recorder. Scientific American, 46(21):334, May 27, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n21/pdf/scientificamerican05271882-334.pdf.

Anonymous:1882:IRSa

[2884] Anonymous. Improved railroad switch. Scientific American, 46(21):334, May 27, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n21/pdf/scientificamerican05271882-334a.pdf.

Riley:1882:SMI

[2885] C. V. Riley. Successful management of the insects most destructive to the orange. Scientific American, 46(21):335–336, May 27, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n21/pdf/scientificamerican05271882-335.pdf.

Anonymous:1882:UAB

[2886] Anonymous. The use of ammonia in baking powders and its importance as a culinary agent. Scientific American, 46(21):336, May 27, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n21/pdf/scientificamerican05271882-336.pdf.

Anonymous:1882:EIk

[2887] Anonymous. Engineering inventions. Scientific American, 46(21):337, May 27, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n21/pdf/scientificamerican05271882-337.pdf.

Anonymous:1882:MIbb

[2888] Anonymous. Mechanical inventions. Scientific American, 46(21):337, May 27, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n21/pdf/scientificamerican05271882-337a.pdf.

Anonymous:1882:AIf

[2889] Anonymous. Agricultural invention. Scientific American, 46(21):337, May 27, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n21/pdf/scientificamerican05271882-337b.pdf.

Anonymous:1882:MIbc

[2890] Anonymous. Metallurgical invention. Scientific American, 46(21):337, May 27, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n21/pdf/scientificamerican05271882-337c.pdf.

Anonymous:1882:MIbd

[2891] Anonymous. Miscellaneous inventions. Scientific American, 46 (21):337–338, May 27, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n21/pdf/scientificamerican05271882-
337d.pdf.

Anonymous:1882:Ch

[2892] Anonymous. Correspondence. Scientific American, 46(22):349, June 3, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n22/pdf/scientificamerican06031882-349.pdf.

Anonymous:1882:BPo

[2893] Anonymous. Business and personal. Scientific American, 46(22):355, June 3, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n22/pdf/scientificamerican06031882-355.pdf.

Anonymous:1882:NQo

[2894] Anonymous. Notes and queries. Scientific American, 46(22):355, June 3, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n22/pdf/scientificamerican06031882-355a.pdf.

Anonymous:1882:IIo

[2895] Anonymous. Index of inventions. Scientific American, 46(22):355–356, June 3, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n22/pdf/scientificamerican06031882-355b.pdf.

Anonymous:1882:PNPn

[2896] Anonymous. Patent notes: Patents. Scientific American, 46(22):356, June 3, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n22/pdf/scientificamerican06031882-356.pdf.

Anonymous:1882:NOB

[2897] Anonymous. New ore boasting furnace. Scientific American, 46(22):343, June 3, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n22/pdf/scientificamerican06031882-343.pdf.

Anonymous:1882:TSE

[2898] Anonymous. The total solar eclipse. Scientific American, 46(22):344, June 3, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n22/pdf/scientificamerican06031882-344.pdf.

Anonymous:1882:IVM

[2899] Anonymous. Insatiate vampires in Michigan. Scientific American, 46(22):344, June 3, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n22/pdf/scientificamerican06031882-
344a.pdf.

Anonymous:1882:LC

[2900] Anonymous. Legislative confiscation. Scientific American, 46 (22):344–345, June 3, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n22/pdf/scientificamerican06031882-
344b.pdf.

Anonymous:1882:TP

[2901] Anonymous. Tubercle parasite. Scientific American, 46(22):345, June 3, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n22/pdf/scientificamerican06031882-345.pdf.

Anonymous:1882:DFA

[2902] Anonymous. Detective fire alarm box. Scientific American, 46(22):346, June 3, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n22/pdf/scientificamerican06031882-346.pdf.

Anonymous:1882:NT

[2903] Anonymous. New tracheotome. Scientific American, 46(22):346–347, June 3, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n22/pdf/scientificamerican06031882-346a.pdf.

Anonymous:1882:NIW

[2904] Anonymous. Novel indicator for weighing scales. Scientific American, 46(22):347, June 3, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n22/pdf/scientificamerican06031882-
347.pdf.

Anonymous:1882:NVB

[2905] Anonymous. New vapor burner. Scientific American, 46(22):347, June 3, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n22/pdf/scientificamerican06031882-347a.pdf.

Anonymous:1882:MR

[2906] Anonymous. The Mississippi River. Scientific American, 46(22):348, June 3, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n22/pdf/scientificamerican06031882-348.pdf.

Anonymous:1882:NPB

[2907] Anonymous. The new Patent Bill. Scientific American, 46(22):348, June 3, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n22/pdf/scientificamerican06031882-348a.pdf.

Anonymous:1882:RLB

[2908] Anonymous. The recent Lawson boiler experiment. Scientific American, 46(22):348, June 3, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n22/pdf/scientificamerican06031882-
348b.pdf.

Anonymous:1882:IPEa

[2909] Anonymous. Improved portable engines. Scientific American, 46 (22):350, June 3, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n22/pdf/scientificamerican06031882-
350.pdf.

Anonymous:1882:FJB

[2910] Anonymous. Filigree jewelry box. Scientific American, 46(22):351, June 3, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n22/pdf/scientificamerican06031882-351.pdf.

Anonymous:1882:WAC

[2911] Anonymous. Where agates come from. Scientific American, 46(22):352, June 3, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n22/pdf/scientificamerican06031882-352.pdf.

Anonymous:1882:EIl

[2912] Anonymous. Engineering inventions. Scientific American, 46(22):353, June 3, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n22/pdf/scientificamerican06031882-353.pdf.

Anonymous:1882:AIg

[2913] Anonymous. Agricultural invention. Scientific American, 46(22):353, June 3, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n22/pdf/scientificamerican06031882-353a.pdf.

Anonymous:1882:MIbe

[2914] Anonymous. Miscellaneous inventions. Scientific American, 46 (22):353–354, June 3, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n22/pdf/scientificamerican06031882-
353b.pdf.

Anonymous:1882:Ci

[2915] Anonymous. Correspondence. Scientific American, 46(23):364–366, June 10, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n23/pdf/scientificamerican06101882-364.pdf.

Anonymous:1882:BPp

[2916] Anonymous. Business and personal. Scientific American, 46(23):371, June 10, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n23/pdf/scientificamerican06101882-371.pdf.

Anonymous:1882:NQp

[2917] Anonymous. Notes and queries. Scientific American, 46(23):371, June 10, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n23/pdf/scientificamerican06101882-371a.pdf.

Anonymous:1882:IIq

[2918] Anonymous. Index of inventions. Scientific American, 46(23):371–372, June 10, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n23/pdf/scientificamerican06101882-371b.pdf.

Anonymous:1882:PNPo

[2919] Anonymous. Patent notes: Patents. Scientific American, 46(23):372, June 10, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n23/pdf/scientificamerican06101882-372.pdf.

Anonymous:1882:AINb

[2920] Anonymous. American industries — no, 83. Scientific American, 46(23):359, June 10, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n23/pdf/scientificamerican06101882-
359.pdf.

Anonymous:1882:VM

[2921] Anonymous. Vivisection and mortisection. Scientific American, 46(23):360, June 10, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n23/pdf/scientificamerican06101882-
360.pdf.

Anonymous:1882:HSR

[2922] Anonymous. How silk is reeled in France. Scientific American, 46(23):360–361, June 10, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n23/pdf/scientificamerican06101882-
360a.pdf.

Miller:1882:CG

[2923] J. E. Miller. Crystalline glass. Scientific American, 46(23):361, June 10, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n23/pdf/scientificamerican06101882-361.pdf.

Anonymous:1882:NSP

[2924] Anonymous. A new silvering process for iron and steel. Scientific American, 46(23):361, June 10, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n23/pdf/scientificamerican06101882-
361a.pdf.

Anonymous:1882:TSW

[2925] Anonymous. Temperature of steam and water as affected by pressure. Scientific American, 46(23):361, June 10, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v46/
n23/pdf/scientificamerican06101882-361b.pdf.

Anonymous:1882:MSC

[2926] Anonymous. Method or stopping crevasses. Scientific American, 46(23):366, June 10, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n23/pdf/scientificamerican06101882-
366.pdf.

Blinn:1882:SWC

[2927] F. G. Blinn. A simple way to convert a pillar and claw stand to an equatorial. Scientific American, 46(23):366, June 10, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v46/
n23/pdf/scientificamerican06101882-366a.pdf.

Anonymous:1882:KCA

[2928] Anonymous. Kabath’s corrugated accumulators. Scientific American, 46(23):367–368, June 10, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n23/pdf/scientificamerican06101882-
367.pdf.

Anonymous:1882:EIm

[2929] Anonymous. Engineering inventions. Scientific American, 46(23): 369–370, June 10, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n23/pdf/scientificamerican06101882-
369.pdf.

Anonymous:1882:MIbf

[2930] Anonymous. Mechanical inventions. Scientific American, 46(23):369, June 10, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n23/pdf/scientificamerican06101882-369a.pdf.

Anonymous:1882:AIh

[2931] Anonymous. Agricultural inventions. Scientific American, 46 (23):369–370, June 10, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n23/pdf/scientificamerican06101882-
369b.pdf.

Anonymous:1882:MIbg

[2932] Anonymous. Miscellaneous inventions. Scientific American, 46(23):370, June 10, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n23/pdf/scientificamerican06101882-370.pdf.

Anonymous:1882:Cj

[2933] Anonymous. Correspondence. Scientific American, 46(24):381, June 17, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n24/pdf/scientificamerican06171882-381.pdf.

Anonymous:1882:BPq

[2934] Anonymous. Business and personal. Scientific American, 46(24):386, June 17, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n24/pdf/scientificamerican06171882-386.pdf.

Anonymous:1882:NQq

[2935] Anonymous. Notes and queries. Scientific American, 46(24):386–387, June 17, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n24/pdf/scientificamerican06171882-386a.pdf.

Anonymous:1882:IIr

[2936] Anonymous. Index of inventions. Scientific American, 46(24):387, June 17, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n24/pdf/scientificamerican06171882-387.pdf.

Anonymous:1882:PNPp

[2937] Anonymous. Patent notes: Patents. Scientific American, 46(24):388, June 17, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n24/pdf/scientificamerican06171882-388.pdf.

Anonymous:1882:LA

[2938] Anonymous. Legislative atavism. Scientific American, 46(24):376, June 17, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n24/pdf/scientificamerican06171882-376.pdf.

Anonymous:1882:SS

[2939] Anonymous. The summer solstice. Scientific American, 46(24):376–377, June 17, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n24/pdf/scientificamerican06171882-376a.pdf.

Anonymous:1882:PS

[2940] Anonymous. The prevailing strikes. Scientific American, 46(24):377, June 17, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n24/pdf/scientificamerican06171882-377.pdf.

Anonymous:1882:MLI

[2941] Anonymous. Misrepresentation as a legislative influence. Scientific American, 46(24):377, June 17, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n24/pdf/scientificamerican06171882-
377a.pdf.

Hovey:1882:ERP

[2942] H. C. Hovey. East Rock Park New Haven. Scientific American, 46(24):378–379, June 17, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n24/pdf/scientificamerican06171882-
378.pdf.

Anonymous:1882:NVS

[2943] Anonymous. New ventilating system. Scientific American, 46 (24):379–380, June 17, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n24/pdf/scientificamerican06171882-
379.pdf.

Anonymous:1882:MIbh

[2944] Anonymous. Miscellaneous inventions. Scientific American, 46(24):380, June 17, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n24/pdf/scientificamerican06171882-380.pdf.

Anonymous:1882:L

[2945] Anonymous. Lock. Scientific American, 46(24):382, June 17, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n24/pdf/scientificamerican06171882-382.pdf.

Anonymous:1882:NOC

[2946] Anonymous. New oil cup. Scientific American, 46(24):382, June 17, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n24/pdf/scientificamerican06171882-382a.pdf.

Anonymous:1882:NFEb

[2947] Anonymous. Novel fire escape. Scientific American, 46(24):382, June 17, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n24/pdf/scientificamerican06171882-382b.pdf.

Anonymous:1882:FKB

[2948] Anonymous. The fox Kusu at the Berlin Aquarium. Scientific American, 46(24):383, June 17, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n24/pdf/scientificamerican06171882-
383.pdf.

Anonymous:1882:RIb

[2949] Anonymous. Recent inventions. Scientific American, 46(24):384–385, June 17, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n24/pdf/scientificamerican06171882-384.pdf.

Anonymous:1882:BPr

[2950] Anonymous. Business and personal. Scientific American, 46(25):400, June 24, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n25/pdf/scientificamerican06241882-400.pdf.

Anonymous:1882:BNBh

[2951] Anonymous. Books: New books and publications. Scientific American, 46(25):400, June 24, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n25/pdf/scientificamerican06241882-
400a.pdf.

Anonymous:1882:NQr

[2952] Anonymous. Notes and queries. Scientific American, 46(25):400–401, June 24, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n25/pdf/scientificamerican06241882-400b.pdf.

Anonymous:1882:CRg

[2953] Anonymous. Communications received. Scientific American, 46 (25):401, June 24, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n25/pdf/scientificamerican06241882-
401.pdf.

Anonymous:1882:IIs

[2954] Anonymous. Index of inventions. Scientific American, 46(25):401–402, June 24, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n25/pdf/scientificamerican06241882-401a.pdf.

Anonymous:1882:PNPq

[2955] Anonymous. Patent notes: Patents. Scientific American, 46(25):402, June 24, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n25/pdf/scientificamerican06241882-402.pdf.

Anonymous:1882:I

[2956] Anonymous. Index. Scientific American, 46(25):403–404, June 24, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n25/pdf/scientificamerican06241882-403.pdf.

Anonymous:1882:INE

[2957] Anonymous. Inauguration of the new Eddystone Lighthouse. Scientific American, 46(25):391, June 24, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n25/pdf/scientificamerican06241882-
391.pdf.

Anonymous:1882:NCF

[2958] Anonymous. Non-combustible fabrics. Scientific American, 46(25):392, June 24, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n25/pdf/scientificamerican06241882-392.pdf.

Anonymous:1882:DIR

[2959] Anonymous. Defective instruction in reading. Scientific American, 46(25):392, June 24, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n25/pdf/scientificamerican06241882-
392a.pdf.

Anonymous:1882:LSU

[2960] Anonymous. Live stock in the United States. Scientific American, 46(25):392–393, June 24, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n25/pdf/scientificamerican06241882-
392b.pdf.

Anonymous:1882:GGA

[2961] Anonymous. Gentilli’s glossograph. — an automatic short-hand apparatus. Scientific American, 46(25):394, June 24, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v46/
n25/pdf/scientificamerican06241882-394.pdf.

Anonymous:1882:NSSb

[2962] Anonymous. New suspended safe. Scientific American, 46(25):395, June 24, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n25/pdf/scientificamerican06241882-395.pdf.

Anonymous:1882:ICSa

[2963] Anonymous. Improvement in cooking stoves and ranges. Scientific American, 46(25):395, June 24, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n25/pdf/scientificamerican06241882-
395a.pdf.

Anonymous:1882:SHa

[2964] Anonymous. Spittoon holder. Scientific American, 46(25):395, June 24, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n25/pdf/scientificamerican06241882-395b.pdf.

Anonymous:1882:DRPd

[2965] Anonymous. Decisions relating to patents. Scientific American, 46(25):396, June 24, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n25/pdf/scientificamerican06241882-
396.pdf.

Anonymous:1882:SDE

[2966] Anonymous. Small dynamo-electric machine. Scientific American, 46(25):396, June 24, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n25/pdf/scientificamerican06241882-
396a.pdf.

Anonymous:1882:MIbi

[2967] Anonymous. Miscellaneous inventions. Scientific American, 46(25):397, June 24, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n25/pdf/scientificamerican06241882-397.pdf.

Anonymous:1882:MIbj

[2968] Anonymous. Mechanical inventions. Scientific American, 46(25):397, June 24, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n25/pdf/scientificamerican06241882-397a.pdf.

Anonymous:1882:PW

[2969] Anonymous. The paradox wheel. Scientific American, 46(25):397, June 24, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n25/pdf/scientificamerican06241882-397b.pdf.

Anonymous:1882:WPM

[2970] Anonymous. Wave power machine. Scientific American, 46(25):398, June 24, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n25/pdf/scientificamerican06241882-398.pdf.

Anonymous:1882:RIL

[2971] Anonymous. Regnard’s incandescent lamp. Scientific American, 46(25):398, June 24, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n25/pdf/scientificamerican06241882-
398a.pdf.

Anonymous:1882:DIM

[2972] Anonymous. The double induction motor. Scientific American, 46(25):399, June 24, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n25/pdf/scientificamerican06241882-
399.pdf.

Anonymous:1882:Ck

[2973] Anonymous. Correspondence. Scientific American, 46(3):40, January 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n3/pdf/scientificamerican01211882-40a.pdf.

Anonymous:1882:INIh

[2974] Anonymous. Inventions: New inventions. Scientific American, 46(3):41, January 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n3/pdf/scientificamerican01211882-
41.pdf.

Anonymous:1882:RIc

[2975] Anonymous. Recent inventions. Scientific American, 46(3):42, January 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n3/pdf/scientificamerican01211882-42.pdf.

Anonymous:1882:BPs

[2976] Anonymous. Business and personal. Scientific American, 46(3):43, January 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n3/pdf/scientificamerican01211882-43.pdf.

Anonymous:1882:NQs

[2977] Anonymous. Notes and queries. Scientific American, 46(3):43, January 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n3/pdf/scientificamerican01211882-43a.pdf.

Anonymous:1882:IIt

[2978] Anonymous. Index of inventions. Scientific American, 46(3):43–44, January 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n3/pdf/scientificamerican01211882-43b.pdf.

Anonymous:1882:UEG

[2979] Anonymous. The use and economy of gas engines. Scientific American, 46(3):32, January 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n3/pdf/scientificamerican01211882-
32.pdf.

Anonymous:1882:OLI

[2980] Anonymous. Our leather industry. Scientific American, 46(3):32, January 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n3/pdf/scientificamerican01211882-32a.pdf.

Anonymous:1882:RIW

[2981] Anonymous. Railway inventions wanted. Scientific American, 46 (3):33, January 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n3/pdf/scientificamerican01211882-
33.pdf.

Anonymous:1882:CPa

[2982] Anonymous. The “City of Puebla”. Scientific American, 46(3):33, January 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n3/pdf/scientificamerican01211882-33a.pdf.

Anonymous:1882:IRSb

[2983] Anonymous. Improved railroad spike. Scientific American, 46 (3):34, January 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n3/pdf/scientificamerican01211882-
34.pdf.

Anonymous:1882:IFF

[2984] Anonymous. Improved furnace feeder. Scientific American, 46 (3):34, January 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n3/pdf/scientificamerican01211882-
34a.pdf.

Anonymous:1882:ISCa

[2985] Anonymous. Improved sash-cord guide. Scientific American, 46 (3):34, January 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n3/pdf/scientificamerican01211882-
34b.pdf.

Anonymous:1882:AINc

[2986] Anonymous. American industries — no. 81. Scientific American, 46(3):35–38, January 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n3/pdf/scientificamerican01211882-
35.pdf.

Anonymous:1882:ICSb

[2987] Anonymous. Improved circular sawmill. Scientific American, 46 (3):38, January 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n3/pdf/scientificamerican01211882-
38.pdf.

Anonymous:1882:NAP

[2988] Anonymous. New automatic pencil. Scientific American, 46(3):38, January 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n3/pdf/scientificamerican01211882-38a.pdf.

Anonymous:1882:MIbk

[2989] Anonymous. Mechanical inventions. Scientific American, 46(3): 38–39, January 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n3/pdf/scientificamerican01211882-
38b.pdf.

Anonymous:1882:IQS

[2990] Anonymous. Improved quartz stamper. Scientific American, 46 (3):39, January 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n3/pdf/scientificamerican01211882-
39.pdf.

Anonymous:1882:EIn

[2991] Anonymous. Engineering inventions. Scientific American, 46(3): 40, January 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n3/pdf/scientificamerican01211882-
40.pdf.

Anonymous:1882:IRIi

[2992] Anonymous. Inventions: Recent inventions. Scientific American, 46(4):50–51, January 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n4/pdf/scientificamerican01281882-
50b.pdf.

Anonymous:1882:Cl

[2993] Anonymous. Correspondence. Scientific American, 46(4):52–53, January 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n4/pdf/scientificamerican01281882-52.pdf.

Anonymous:1882:INIi

[2994] Anonymous. Inventions: New inventions. Scientific American, 46 (4):53–54, January 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n4/pdf/scientificamerican01281882-
53.pdf.

Anonymous:1882:BPt

[2995] Anonymous. Business and personal. Scientific American, 46(4):58, January 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n4/pdf/scientificamerican01281882-58.pdf.

Anonymous:1882:BNBi

[2996] Anonymous. Books: New books and publications. Scientific American, 46(4):58, January 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n4/pdf/scientificamerican01281882-
58a.pdf.

Anonymous:1882:NQt

[2997] Anonymous. Notes and queries. Scientific American, 46(4):58–59, January 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n4/pdf/scientificamerican01281882-58b.pdf.

Anonymous:1882:IIu

[2998] Anonymous. Index of inventions. Scientific American, 46(4):59–60, January 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n4/pdf/scientificamerican01281882-59a.pdf.

Anonymous:1882:FCR

[2999] Anonymous. Fraudulent claims in reissued patents. Scientific American, 46(4):48, January 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n4/pdf/scientificamerican01281882-
48.pdf.

Anonymous:1882:LF

[3000] Anonymous. Losses by fire in 1881. Scientific American, 46(4):48, January 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n4/pdf/scientificamerican01281882-48a.pdf.

Anonymous:1882:LHG

[3001] Anonymous. Left-handed generosity. Scientific American, 46(4): 48, January 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n4/pdf/scientificamerican01281882-
48b.pdf.

Anonymous:1882:BE

[3002] Anonymous. Boiler, explosions in 1881. Scientific American, 46 (4):49, January 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n4/pdf/scientificamerican01281882-
49.pdf.

Anonymous:1882:ESS

[3003] Anonymous. Electrical sheep shears. Scientific American, 46 (4):49–50, January 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n4/pdf/scientificamerican01281882-
49a.pdf.

Anonymous:1882:ISd

[3004] Anonymous. Improvement in steamboats. Scientific American, 46(4):50, January 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n4/pdf/scientificamerican01281882-
50.pdf.

Anonymous:1882:ILL

[3005] Anonymous. Improvement in lighthouse lanterns. Scientific American, 46(4):50, January 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n4/pdf/scientificamerican01281882-
50a.pdf.

Anonymous:1882:MCC

[3006] Anonymous. The multichange cannon for coast and harbor defense. Scientific American, 46(4):51, January 28, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v46/
n4/pdf/scientificamerican01281882-51.pdf.

Anonymous:1882:HSE

[3007] Anonymous. Horizontal steam engine with automatic slide valve cut-off. Scientific American, 46(4):54, January 28, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v46/
n4/pdf/scientificamerican01281882-54.pdf.

Anonymous:1882:MIbl

[3008] Anonymous. Mechanical inventions. Scientific American, 46(4): 54–55, January 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n4/pdf/scientificamerican01281882-
54a.pdf.

Anonymous:1882:IIMa

[3009] Anonymous. Improved ice machine. Scientific American, 46(4):55, January 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n4/pdf/scientificamerican01281882-55.pdf.

Anonymous:1882:EIo

[3010] Anonymous. Engineering inventions. Scientific American, 46(4): 55, January 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n4/pdf/scientificamerican01281882-
55a.pdf.

Anonymous:1882:MIbm

[3011] Anonymous. Miscellaneous inventions. Scientific American, 46 (4):56–57, January 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n4/pdf/scientificamerican01281882-
56.pdf.

Anonymous:1882:AIi

[3012] Anonymous. Agricultural inventions. Scientific American, 46 (4):57, January 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n4/pdf/scientificamerican01281882-
57.pdf.

Holder:1882:PCO

[3013] C. F. Holder. A phosphorescent coral and other marine phosanimals. Scientific American, 46(4):57, January 28, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v46/
n4/pdf/scientificamerican01281882-57a.pdf.

Anonymous:1882:CRh

[3014] Anonymous. Communication received. Scientific American, 46 (4):59, January 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n4/pdf/scientificamerican01281882-
59.pdf.

Anonymous:1882:IRIj

[3015] Anonymous. Inventions: Recent inventions. Scientific American, 46(5):72–74, February 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n5/pdf/scientificamerican02041882-
72a.pdf.

Anonymous:1882:BPu

[3016] Anonymous. Business and personal. Scientific American, 46 (5):75, February 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n5/pdf/scientificamerican02041882-
75.pdf.

Anonymous:1882:NQu

[3017] Anonymous. Notes and queries. Scientific American, 46(5):75, February 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n5/pdf/scientificamerican02041882-75a.pdf.

Anonymous:1882:IIv

[3018] Anonymous. Index of inventions. Scientific American, 46(5): 75–76, February 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n5/pdf/scientificamerican02041882-
75c.pdf.

Anonymous:1882:HRT

[3019] Anonymous. The Hudson River Tunnel — New York end. Scientific American, 46(5):63, February 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n5/pdf/scientificamerican02041882-
63.pdf.

Anonymous:1882:CNR

[3020] Anonymous. Criminal neglect in railway management. Scientific American, 46(5):64, February 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n5/pdf/scientificamerican02041882-
64.pdf.

Anonymous:1882:SP

[3021] Anonymous. Solid petroleum. Scientific American, 46(5):64, February 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n5/pdf/scientificamerican02041882-64a.pdf.

Anonymous:1882:SCC

[3022] Anonymous. The Stock Car Competition. Scientific American, 46(5):64–65, February 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n5/pdf/scientificamerican02041882-
64b.pdf.

Anonymous:1882:PCS

[3023] Anonymous. Plating cotton with silk. Scientific American, 46 (5):65, February 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n5/pdf/scientificamerican02041882-
65.pdf.

Anonymous:1882:ME

[3024] Anonymous. Marine economy. Scientific American, 46(5):65, February 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n5/pdf/scientificamerican02041882-65a.pdf.

Anonymous:1882:ISP

[3025] Anonymous. Improved shovel plow. Scientific American, 46(5): 66, February 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n5/pdf/scientificamerican02041882-
66.pdf.

Anonymous:1882:IWF

[3026] Anonymous. Improved wire fence. Scientific American, 46(5):66, February 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n5/pdf/scientificamerican02041882-66a.pdf.

Anonymous:1882:EIp

[3027] Anonymous. Engineering inventions. Scientific American, 46(5): 66, February 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n5/pdf/scientificamerican02041882-
66b.pdf.

Anonymous:1882:KEA

[3028] Anonymous. Kohler’s extinguisher for Argand burners. Scientific American, 46(5):67, February 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n5/pdf/scientificamerican02041882-
67.pdf.

Anonymous:1882:IDM

[3029] Anonymous. Improved drilling machine. Scientific American, 46 (5):67, February 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n5/pdf/scientificamerican02041882-
67a.pdf.

Anonymous:1882:MIbn

[3030] Anonymous. Miscellaneous inventions. Scientific American, 46 (5):67, February 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n5/pdf/scientificamerican02041882-
67b.pdf.

Anonymous:1882:IH

[3031] Anonymous. Improved horseshoe. Scientific American, 46(5): 67–68, February 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n5/pdf/scientificamerican02041882-
67c.pdf.

Deprez:1882:RA

[3032] Marcel Deprez. Rheometric apparatus. Scientific American, 46 (5):68, February 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n5/pdf/scientificamerican02041882-
68.pdf.

Anonymous:1882:MIbo

[3033] Anonymous. Mechanical inventions. Scientific American, 46(5): 68–69, February 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n5/pdf/scientificamerican02041882-
68a.pdf.

Anonymous:1882:NCSa

[3034] Anonymous. New coin scale. Scientific American, 46(5):69, February 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n5/pdf/scientificamerican02041882-69.pdf.

Anonymous:1882:ETB

[3035] Anonymous. Explosion of two boilers at Pittsburg, PA. Scientific American, 46(5):70–72, February 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n5/pdf/scientificamerican02041882-
70.pdf.

Anonymous:1882:BN

[3036] Anonymous. Botanical notes. Scientific American, 46(5):72, February 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n5/pdf/scientificamerican02041882-72.pdf.

Anonymous:1882:CRi

[3037] Anonymous. Communication received. Scientific American, 46 (5):75, February 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n5/pdf/scientificamerican02041882-
75b.pdf.

Anonymous:1882:INIj

[3038] Anonymous. Inventions: New inventions. Scientific American, 46(6):90, February 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n6/pdf/scientificamerican02111882-
90.pdf.

Anonymous:1882:BPv

[3039] Anonymous. Business and personal. Scientific American, 46(6): 91, February 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n6/pdf/scientificamerican02111882-
91.pdf.

Anonymous:1882:NQv

[3040] Anonymous. Notes and queries. Scientific American, 46(6):91, February 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n6/pdf/scientificamerican02111882-91a.pdf.

Anonymous:1882:IIw

[3041] Anonymous. Index of inventions. Scientific American, 46(6):91– 92, February 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n6/pdf/scientificamerican02111882-
91b.pdf.

Anonymous:1882:PNPr

[3042] Anonymous. Patent notes: Patents. Scientific American, 46(6): 92, February 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n6/pdf/scientificamerican02111882-
92.pdf.

Anonymous:1882:CW

[3043] Anonymous. The City of Worcester. Scientific American, 46 (6):79, February 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n6/pdf/scientificamerican02111882-
79.pdf.

Anonymous:1882:ALH

[3044] Anonymous. Alexander Lyman Holley. Scientific American, 46 (6):80, February 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n6/pdf/scientificamerican02111882-
80.pdf.

Anonymous:1882:SQC

[3045] Anonymous. Some questions of coast defense. Scientific American, 46(6):80–81, February 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n6/pdf/scientificamerican02111882-
80a.pdf.

Anonymous:1882:BOO

[3046] Anonymous. Burning of the old and opening of the new Scientific American offices, New York. Scientific American, 46(6):81–82, February 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n6/pdf/scientificamerican02111882-81.pdf.

Anonymous:1882:IPb

[3047] Anonymous. Improved pulverizer. Scientific American, 46(6):82, February 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n6/pdf/scientificamerican02111882-82.pdf.

Anonymous:1882:MIbp

[3048] Anonymous. Miscellaneous inventions. Scientific American, 46 (6):83, February 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n6/pdf/scientificamerican02111882-
83.pdf.

Anonymous:1882:NEM

[3049] Anonymous. A new electric motor and its applications. Scientific American, 46(6):83, February 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n6/pdf/scientificamerican02111882-
83a.pdf.

Anonymous:1882:MPS

[3050] Anonymous. Multiple pressure sugar mill. Scientific American, 46(6):86, February 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n6/pdf/scientificamerican02111882-
86.pdf.

Anonymous:1882:PMB

[3051] Anonymous. Press for the manufacture of belts. Scientific American, 46(6):86, February 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n6/pdf/scientificamerican02111882-
86a.pdf.

Anonymous:1882:CO

[3052] Anonymous. The cork oak. Scientific American, 46(6):87–88, February 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n6/pdf/scientificamerican02111882-87.pdf.

Anonymous:1882:AIj

[3053] Anonymous. Agricultural inventions. Scientific American, 46 (6):88–90, February 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n6/pdf/scientificamerican02111882-
88.pdf.

Anonymous:1882:IRIk

[3054] Anonymous. Inventions: Recent inventions. Scientific American, 46(7):102, February 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n7/pdf/scientificamerican02181882-
102a.pdf.

Anonymous:1882:NIb

[3055] Anonymous. New inventions. Scientific American, 46(7):103, February 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n7/pdf/scientificamerican02181882-103.pdf.

Anonymous:1882:BPw

[3056] Anonymous. Business and personal. Scientific American, 46(7): 107, February 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n7/pdf/scientificamerican02181882-
107.pdf.

Anonymous:1882:NQw

[3057] Anonymous. Notes and queries. Scientific American, 46(7):107, February 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n7/pdf/scientificamerican02181882-107a.pdf.

Anonymous:1882:IIx

[3058] Anonymous. Index of inventions. Scientific American, 46(7): 107–108, February 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n7/pdf/scientificamerican02181882-
107b.pdf.

Anonymous:1882:PNPs

[3059] Anonymous. Patent notes: Patents. Scientific American, 46 (7):108, February 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n7/pdf/scientificamerican02181882-
108.pdf.

Anonymous:1882:AINd

[3060] Anonymous. American industries — no. 82. Scientific American, 46(7):95, February 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n7/pdf/scientificamerican02181882-
95.pdf.

Anonymous:1882:ACE

[3061] Anonymous. Another case of expanded reissue. Scientific American, 46(7):96, February 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n7/pdf/scientificamerican02181882-
96.pdf.

Anonymous:1882:SRJ

[3062] Anonymous. The scientific results of the Jeannette Expedition. Scientific American, 46(7):96, February 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v46/n7/pdf/scientificamerican02181882-
96a.pdf.

Anonymous:1882:RGP

[3063] Anonymous. The relation of the Government to patentees. Scientific American, 46(7):96, February 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n7/pdf/scientificamerican02181882-
96b.pdf.

Anonymous:1882:CS

[3064] Anonymous. Comet seeking. Scientific American, 46(7):96, February 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n7/pdf/scientificamerican02181882-96c.pdf.

Anonymous:1882:SBNd

[3065] Anonymous. Steam boiler notes. Scientific American, 46(7):97, February 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n7/pdf/scientificamerican02181882-97.pdf.

Anonymous:1882:TYP

[3066] Anonymous. Ten years’ progress in iron and steel production. Scientific American, 46(7):97–98, February 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v46/n7/pdf/scientificamerican02181882-
97a.pdf.

Anonymous:1882:AMT

[3067] Anonymous. Ader’s magnifying telephone. Scientific American, 46(7):98, February 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n7/pdf/scientificamerican02181882-
98.pdf.

Anonymous:1882:SSB

[3068] Anonymous. Supplying steam boilers with water. Scientific American, 46(7):98, February 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n7/pdf/scientificamerican02181882-
98a.pdf.

Anonymous:1882:MIbq

[3069] Anonymous. Mechanical inventions. Scientific American, 46(7): 98, February 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n7/pdf/scientificamerican02181882-
98b.pdf.

Anonymous:1882:AIk

[3070] Anonymous. Agricultural inventions. Scientific American, 46(7): 100–101, February 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n7/pdf/scientificamerican02181882-
100.pdf.

Anonymous:1882:IPEb

[3071] Anonymous. Improved portable engine. Scientific American, 46 (7):102, February 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n7/pdf/scientificamerican02181882-
102.pdf.

Anonymous:1882:GTN

[3072] Anonymous. Grebes and their nests. Scientific American, 46 (7):103, February 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n7/pdf/scientificamerican02181882-
103a.pdf.

Anonymous:1882:SSP

[3073] Anonymous. The symbolical signs of the planets. Scientific American, 46(7):103–104, February 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n7/pdf/scientificamerican02181882-
103b.pdf.

Anonymous:1882:MIbr

[3074] Anonymous. Miscellaneous inventions. Scientific American, 46(7): 104–106, February 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n7/pdf/scientificamerican02181882-
104.pdf.

Anonymous:1882:EIq

[3075] Anonymous. Engineering intentions. Scientific American, 46(7): 106, February 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n7/pdf/scientificamerican02181882-
106.pdf.

Anonymous:1882:IRIl

[3076] Anonymous. Inventions: Recent invention. Scientific American, 46 (8):115–117, February 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n8/pdf/scientificamerican02251882-
116.pdf.

Anonymous:1882:INIk

[3077] Anonymous. Inventions: New inventions. Scientific American, 46(8):121, February 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n8/pdf/scientificamerican02251882-
121.pdf.

Anonymous:1882:BPx

[3078] Anonymous. Business and personal. Scientific American, 46(8): 122, February 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n8/pdf/scientificamerican02251882-
122.pdf.

Anonymous:1882:NQx

[3079] Anonymous. Notes and queries. Scientific American, 46(8):122– 123, February 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n8/pdf/scientificamerican02251882-
122a.pdf.

Anonymous:1882:BNBj

[3080] Anonymous. Books: New books and publications. Scientific American, 46(8):123, February 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n8/pdf/scientificamerican02251882-
123.pdf.

Anonymous:1882:IIy

[3081] Anonymous. Index of inventions. Scientific American, 46(8): 123–124, February 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n8/pdf/scientificamerican02251882-
123a.pdf.

Anonymous:1882:PNPt

[3082] Anonymous. Patent notes: Patents. Scientific American, 46 (8):124, February 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n8/pdf/scientificamerican02251882-
124.pdf.

Anonymous:1882:FWM

[3083] Anonymous. Frager’s water meter. Scientific American, 46(8): 111, February 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n8/pdf/scientificamerican02251882-
111.pdf.

Anonymous:1882:WCM

[3084] Anonymous. The Wilson cooked meat patents. Scientific American, 46(8):112, February 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n8/pdf/scientificamerican02251882-
112.pdf.

Anonymous:1882:RCT

[3085] Anonymous. Railroad collisions and their preventives — room for invention. Scientific American, 46(8):112, February 25, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v46/
n8/pdf/scientificamerican02251882-112a.pdf.

Anonymous:1882:PRF

[3086] Anonymous. The Park Row Fire. Scientific American, 46(8): 112–113, February 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n8/pdf/scientificamerican02251882-
112b.pdf.

Anonymous:1882:APMb

[3087] Anonymous. Aspects of the planets for March. Scientific American, 46(8):113, February 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n8/pdf/scientificamerican02251882-
113.pdf.

Anonymous:1882:AEA

[3088] Anonymous. Apparatus for extracting ammonia from gas liquors. Scientific American, 46(8):114, February 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v46/n8/pdf/scientificamerican02251882-
114.pdf.

Anonymous:1882:IRF

[3089] Anonymous. Improvement in revolving fire arms. Scientific American, 46(8):115, February 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n8/pdf/scientificamerican02251882-
115.pdf.

Anonymous:1882:ITM

[3090] Anonymous. Improved testing machine. Scientific American, 46 (8):115–116, February 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n8/pdf/scientificamerican02251882-
115a.pdf.

Anonymous:1882:EIr

[3091] Anonymous. Engineering inventions. Scientific American, 46(8): 117, February 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n8/pdf/scientificamerican02251882-
117.pdf.

Anonymous:1882:IW

[3092] Anonymous. Improvement in wringers. Scientific American, 46 (8):118, February 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n8/pdf/scientificamerican02251882-
118.pdf.

Anonymous:1882:IFEd

[3093] Anonymous. Improved fire-extinguishing apparatus. Scientific American, 46(8):118, February 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n8/pdf/scientificamerican02251882-
118a.pdf.

Anonymous:1882:MIbs

[3094] Anonymous. Miscellaneous inventions. Scientific American, 46 (8):118, February 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n8/pdf/scientificamerican02251882-
118b.pdf.

Anonymous:1882:RTT

[3095] Anonymous. Revolving telegraph table. Scientific American, 46 (8):118, February 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n8/pdf/scientificamerican02251882-
118c.pdf.

Anonymous:1882:NPH

[3096] Anonymous. Novel passenger head rest. Scientific American, 46 (8):119, February 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n8/pdf/scientificamerican02251882-
119.pdf.

Anonymous:1882:FCJ

[3097] Anonymous. The force of a crocodile’s jaw. Scientific American, 46(8):119, February 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n8/pdf/scientificamerican02251882-
119a.pdf.

Anonymous:1882:ICSc

[3098] Anonymous. Improved carpet stretcher. Scientific American, 46 (8):119, February 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n8/pdf/scientificamerican02251882-
119b.pdf.

Anonymous:1882:MIbt

[3099] Anonymous. Mechanical inventions. Scientific American, 46(8): 119–120, February 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n8/pdf/scientificamerican02251882-
119c.pdf.

Anonymous:1882:HDH

[3100] Anonymous. How dwelling houses are poisoned. Scientific American, 46(8):120–121, February 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n8/pdf/scientificamerican02251882-
120.pdf.

Anonymous:1882:IRIm

[3101] Anonymous. Inventions: Recent inventions. Scientific American, 46(9):131–132, March 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n9/pdf/scientificamerican03041882-
131.pdf.

Anonymous:1882:BPy

[3102] Anonymous. Business and personal. Scientific American, 46(9):139, March 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n9/pdf/scientificamerican03041882-139.pdf.

Anonymous:1882:IIz

[3103] Anonymous. Index of inventions. Scientific American, 46(9):139–140, March 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n9/pdf/scientificamerican03041882-139a.pdf.

Anonymous:1882:PNPu

[3104] Anonymous. Patent notes: Patents. Scientific American, 46(9):140, March 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n9/pdf/scientificamerican03041882-140.pdf.

Anonymous:1882:NYW

[3105] Anonymous. The “New York World” newspaper. — a remarkable establishment. Scientific American, 46(9):127, March 4, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v46/
n9/pdf/scientificamerican03041882-127.pdf.

Anonymous:1882:FYP

[3106] Anonymous. Forty years in the Patent Office. Scientific American, 46(9):128, March 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n9/pdf/scientificamerican03041882-
128.pdf.

Anonymous:1882:PSCa

[3107] Anonymous. The profits of silk culture. Scientific American, 46(9):128, March 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n9/pdf/scientificamerican03041882-128a.pdf.

Anonymous:1882:NPO

[3108] Anonymous. Needs of the Patent Office. Scientific American, 46(9):128, March 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n9/pdf/scientificamerican03041882-128b.pdf.

Anonymous:1882:SBNe

[3109] Anonymous. Steam boiler notes. Scientific American, 46(9):129, March 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n9/pdf/scientificamerican03041882-129.pdf.

Anonymous:1882:MIbu

[3110] Anonymous. Miscellaneous inventions. Scientific American, 46(9):133, March 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n9/pdf/scientificamerican03041882-133.pdf.

Anonymous:1882:SSC

[3111] Anonymous. Safety shield for circular saws. Scientific American, 46(9):134, March 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n9/pdf/scientificamerican03041882-
134.pdf.

Anonymous:1882:EMB

[3112] Anonymous. Electro-magnetic brake. Scientific American, 46(9):134, March 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n9/pdf/scientificamerican03041882-134a.pdf.

Anonymous:1882:MIbv

[3113] Anonymous. Mechanical inventions. Scientific American, 46(9): 134–135, March 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n9/pdf/scientificamerican03041882-
134b.pdf.

Anonymous:1882:HB

[3114] Anonymous. The Hercules beetle. Scientific American, 46(9):135–136, March 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n9/pdf/scientificamerican03041882-135.pdf.

Anonymous:1882:AIl

[3115] Anonymous. Agricultural inventions. Scientific American, 46(9): 136–137, March 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v46/n9/pdf/scientificamerican03041882-
136.pdf.

Anonymous:1882:EIs

[3116] Anonymous. Engineering inventions. Scientific American, 46(9):138, March 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v46/n9/pdf/scientificamerican03041882-138.pdf.

Anonymous:1882:IRIn

[3117] Anonymous. Inventions: Recent inventions. Scientific American, 47(1): 10, July 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n1/pdf/scientificamerican07011882-10.pdf.

Anonymous:1882:BPz

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n1/pdf/scientificamerican07011882-11.pdf.

Anonymous:1882:NQy

[3119] Anonymous. Notes and queries. Scientific American, 47(1):11, July 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n1/pdf/scientificamerican07011882-11a.pdf.

Anonymous:1882:IIba

[3120] Anonymous. Index of inventions. Scientific American, 47(1):11–12, July 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n1/pdf/scientificamerican07011882-11b.pdf.

Anonymous:1882:PNPv

[3121] Anonymous. Patent notes: Patents. Scientific American, 47(1):12, July 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n1/pdf/scientificamerican07011882-12.pdf.

Anonymous:1882:GBS

[3122] Anonymous. The Great Bell for St. Paul’s. Scientific American, 47(1): 1, July 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n1/pdf/scientificamerican07011882-1.pdf.

Anonymous:1882:APJb

[3123] Anonymous. Aspects of the planets for July. Scientific American, 47(1):2–3, July 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n1/pdf/scientificamerican07011882-
2.pdf.

Anonymous:1882:TWM

[3124] Anonymous. Is there water on the Moon? Scientific American, 47(1): 3, July 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n1/pdf/scientificamerican07011882-3.pdf.

Anonymous:1882:NQAa

[3125] Anonymous. New quick adjusting vise. Scientific American, 47(1):4, July 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n1/pdf/scientificamerican07011882-4.pdf.

Anonymous:1882:NCSb

[3126] Anonymous. Novel carpet stretcher and carpet fastener. Scientific American, 47(1):4–5, July 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n1/pdf/scientificamerican07011882-
4a.pdf.

Anonymous:1882:NBPa

[3127] Anonymous. New baling press. Scientific American, 47(1):5, July 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n1/pdf/scientificamerican07011882-5.pdf.

Anonymous:1882:NMD

[3128] Anonymous. New mill dog. Scientific American, 47(1):5, July 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n1/pdf/scientificamerican07011882-5a.pdf.

Anonymous:1882:MIbw

[3129] Anonymous. Metallurgical invention. Scientific American, 47(1):6, July 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n1/pdf/scientificamerican07011882-6.pdf.

Anonymous:1882:MIbx

[3130] Anonymous. Miscellaneous invention. Scientific American, 47(1):6, July 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n1/pdf/scientificamerican07011882-6a.pdf.

Anonymous:1882:RDP

[3131] Anonymous. Recent discoveries in the Planet Mars. Scientific American, 47(1):7, July 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n1/pdf/scientificamerican07011882-
7.pdf.

Anonymous:1882:NNT

[3132] Anonymous. New nut tapping machine. Scientific American, 47(1):8–9, July 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n1/pdf/scientificamerican07011882-8.pdf.

Anonymous:1882:PHN

[3133] Anonymous. The pygmy hog of Nepaul. Scientific American, 47(1):9, July 1, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n1/pdf/scientificamerican07011882-9.pdf.

Anonymous:1882:Cm

[3134] Anonymous. Correspondence. Scientific American, 47(10):149, September 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n10/pdf/scientificamerican09021882-149.pdf.

Anonymous:1882:IRIo

[3135] Anonymous. Inventions: Recent inventions. Scientific American, 47(10):153, September 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n10/pdf/scientificamerican09021882-
153.pdf.

Anonymous:1882:BPba

[3136] Anonymous. Business and personal. Scientific American, 47 (10):155, September 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n10/pdf/scientificamerican09021882-
155.pdf.

Anonymous:1882:NQz

[3137] Anonymous. Notes and queries. Scientific American, 47(10): 155, September 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n10/pdf/scientificamerican09021882-
155a.pdf.

Anonymous:1882:IIbb 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n10/pdf/scientificamerican09021882-
155b.pdf.

Anonymous:1882:PNPw

[3139] Anonymous. Patent notes: Patents. Scientific American, 47 (10):156, September 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n10/pdf/scientificamerican09021882-
156.pdf.

Anonymous:1882:MSM

[3140] Anonymous. The Montreal scientific meetings. Scientific American, 47(10):144, September 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n10/pdf/scientificamerican09021882-
144.pdf.

Anonymous:1882:KMD

[3141] Anonymous. The Keely motor deception. Scientific American, 47 (10):144, September 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n10/pdf/scientificamerican09021882-
144a.pdf.

Anonymous:1882:APS

[3142] Anonymous. Aspects of the planets for September. Scientific American, 47(10):145, September 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n10/pdf/scientificamerican09021882-
145.pdf.

Anonymous:1882:LPM

[3143] Anonymous. Lyall’s positive motion loom. Scientific American, 47 (10):146–147, September 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n10/pdf/scientificamerican09021882-
146.pdf.

Anonymous:1882:BHR

[3144] Anonymous. Break in the Hudson River Tunnel. Scientific American, 47(10):147–148, September 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n10/pdf/scientificamerican09021882-
147.pdf.

Anonymous:1882:GPA

[3145] Anonymous. Geared pendulum air pump. Scientific American, 47(10):150, September 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n10/pdf/scientificamerican09021882-
150.pdf.

Anonymous:1882:IIPb

[3146] Anonymous. Improved iron planer. Scientific American, 47(10): 150, September 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n10/pdf/scientificamerican09021882-
150a.pdf.

Anonymous:1882:JSa

[3147] Anonymous. The Jabiru of Senegal. Scientific American, 47 (10):151, September 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n10/pdf/scientificamerican09021882-
151.pdf.

Anonymous:1882:DSM

[3148] Anonymous. The daet snake, or milk snake. Scientific American, 47 (10):151–152, September 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n10/pdf/scientificamerican09021882-
151a.pdf.

Anonymous:1882:TEB

[3149] Anonymous. Trout with elastic band. Scientific American, 47 (10):152, September 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n10/pdf/scientificamerican09021882-
152.pdf.

Anonymous:1882:LFS

[3150] Anonymous. London fire service. Scientific American, 47(10): 153, September 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n10/pdf/scientificamerican09021882-
153a.pdf.

Anonymous:1882:MIby

[3151] Anonymous. Mechanical inventions. Scientific American, 47(10): 154, September 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n10/pdf/scientificamerican09021882-
154.pdf.

Anonymous:1882:EIt

[3152] Anonymous. Electrical inventions. Scientific American, 47(10): 154, September 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n10/pdf/scientificamerican09021882-
154a.pdf.

Anonymous:1882:AIm

[3153] Anonymous. Agricultural inventions. Scientific American, 47 (10):154, September 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n10/pdf/scientificamerican09021882-
154b.pdf.

Anonymous:1882:MIbz

[3154] Anonymous. Miscellaneous inventions. Scientific American, 47 (10):154, September 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n10/pdf/scientificamerican09021882-
154c.pdf.

Anonymous:1882:EIu

[3155] Anonymous. Engineering invention. Scientific American, 47(10): 154, September 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n10/pdf/scientificamerican09021882-
154d.pdf.

Anonymous:1882:Cn

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n11/pdf/scientificamerican09091882-164.pdf.

Anonymous:1882:IRIp

[3157] Anonymous. Inventions: Recent inventions. Scientific American, 47(11):169, September 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n11/pdf/scientificamerican09091882-
169.pdf.

Anonymous:1882:NQba

[3158] Anonymous. Notes and queries. Scientific American, 47(11): 170, September 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n11/pdf/scientificamerican09091882-
171a.pdf.

Anonymous:1882:IIbc

[3159] Anonymous. Index of inventions. Scientific American, 47(11): 170, September 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n11/pdf/scientificamerican09091882-
171b.pdf.

Anonymous:1882:BPbb

[3160] Anonymous. Business and personal. Scientific American, 47 (11):171, September 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n11/pdf/scientificamerican09091882-
171.pdf.

Anonymous:1882:PNPx

[3161] Anonymous. Patent notes: Patents. Scientific American, 47 (11):172, September 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n11/pdf/scientificamerican09091882-
172.pdf.

Anonymous:1882:EY

[3162] Anonymous. An electrical yarn. Scientific American, 47(11): 160, September 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n11/pdf/scientificamerican09091882-
160.pdf.

Anonymous:1882:AIEa

[3163] Anonymous. American inventions in the Egyptian War. Scientific American, 47(11):160, September 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v47/n11/pdf/scientificamerican09091882-
160a.pdf.

Anonymous:1882:DLE

[3164] Anonymous. Duty of the locomotive engineer. Scientific American, 47(11):160, September 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n11/pdf/scientificamerican09091882-
160b.pdf.

Anonymous:1882:IIbd

[3165] Anonymous. Idlers as inventors. Scientific American, 47(11): 161, September 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n11/pdf/scientificamerican09091882-
161.pdf.

Anonymous:1882:CCE

[3166] Anonymous. Car couplings in England — a chance for American inventors. Scientific American, 47(11):161, September 9, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v47/
n11/pdf/scientificamerican09091882-161a.pdf.

Anonymous:1882:IFb

[3167] Anonymous. Improved fence. Scientific American, 47(11):162, September 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n11/pdf/scientificamerican09091882-162.pdf.

Anonymous:1882:VMC

[3168] Anonymous. Vernier and micrometer calipers. Scientific American, 47(11):162, September 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n11/pdf/scientificamerican09091882-
162a.pdf.

Anonymous:1882:NRS

[3169] Anonymous. New railroad signaling apparatus. Scientific American, 47(11):162, September 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n11/pdf/scientificamerican09091882-
162b.pdf.

Anonymous:1882:HGI

[3170] Anonymous. The history of a great invention — the Giffard injector. Scientific American, 47(11):163–164, September 9, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v47/
n11/pdf/scientificamerican09091882-163.pdf.

Anonymous:1882:AINe

[3171] Anonymous. American industries — no. 84. Scientific American, 47 (11):165–166, September 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n11/pdf/scientificamerican09091882-
165.pdf.

Anonymous:1882:EAL

[3172] Anonymous. Experiment with an air locomotive on the elevated railroad. Scientific American, 47(11):166, September 9, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v47/
n11/pdf/scientificamerican09091882-166.pdf.

Anonymous:1882:WFM

[3173] Anonymous. The white-footed mouse, or deer mouse. Scientific American, 47(11):167, September 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n11/pdf/scientificamerican09091882-
167.pdf.

Anonymous:1882:PM

[3174] Anonymous. The photographing of motion. Scientific American, 47 (11):167–168, September 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n11/pdf/scientificamerican09091882-
167a.pdf.

Anonymous:1882:DRPe

[3175] Anonymous. Decisions relating to patents, trade marks, etc. Scientific American, 47(11):168, September 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v47/n11/pdf/scientificamerican09091882-
168.pdf.

Anonymous:1882:EIv

[3176] Anonymous. Engineering inventions. Scientific American, 47(11): 170, September 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n11/pdf/scientificamerican09091882-
170.pdf.

Anonymous:1882:MIca

[3177] Anonymous. Mechanical inventions. Scientific American, 47(11): 170, September 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n11/pdf/scientificamerican09091882-
170a.pdf.

Anonymous:1882:EIw

[3178] Anonymous. Electrical inventions. Scientific American, 47(11): 170, September 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n11/pdf/scientificamerican09091882-
170b.pdf.

Anonymous:1882:AIo

[3179] Anonymous. Agricultural inventions. Scientific American, 47 (11):170, September 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n11/pdf/scientificamerican09091882-
170c.pdf.

Anonymous:1882:MIcb

[3180] Anonymous. Miscellaneous inventions. Scientific American, 47 (11):170, September 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n11/pdf/scientificamerican09091882-
170d.pdf.

Anonymous:1882:IRIq

[3181] Anonymous. Inventions: Recent inventions. Scientific American, 47(12):185, September 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n12/pdf/scientificamerican09161882-
185.pdf.

Anonymous:1882:NQbb

[3182] Anonymous. Notes and queries. Scientific American, 47(12): 186, September 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n12/pdf/scientificamerican09161882-
186f.pdf.

Anonymous:1882:BPbc

[3183] Anonymous. Business and personal. Scientific American, 47 (12):187, September 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n12/pdf/scientificamerican09161882-
187.pdf.

Anonymous:1882:IIbe

[3184] Anonymous. Index of inventions. Scientific American, 47(12): 187–188, September 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n12/pdf/scientificamerican09161882-
187a.pdf.

Anonymous:1882:PNPy

[3185] Anonymous. Patent notes: Patents. Scientific American, 47 (12):188, September 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n12/pdf/scientificamerican09161882-
188.pdf.

Anonymous:1882:IBL

[3186] Anonymous. Improved boat-lowering apparatus. Scientific American, 47(12):175–176, September 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n12/pdf/scientificamerican09161882-
175.pdf.

Anonymous:1882:ODI

[3187] Anonymous. One has-done-it better than three can’t-do-its. Scientific American, 47(12):176, September 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v47/n12/pdf/scientificamerican09161882-
176.pdf.

Anonymous:1882:M

[3188] Anonymous. Mica. Scientific American, 47(12):176, September 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n12/pdf/scientificamerican09161882-176a.pdf.

Anonymous:1882:GIE

[3189] Anonymous. General incandescent electric lighting in New York. Scientific American, 47(12):176–177, September 16, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v47/
n12/pdf/scientificamerican09161882-176b.pdf.

Anonymous:1882:NCCb

[3190] Anonymous. New combination clutch. Scientific American, 47 (12):178, September 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n12/pdf/scientificamerican09161882-
178.pdf.

Anonymous:1882:HHS

[3191] Anonymous. Hodson’s high speed rotary engine. Scientific American, 47(12):178–179, September 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n12/pdf/scientificamerican09161882-
178a.pdf.

Anonymous:1882:PBH

[3192] Anonymous. Paper-bag holder. Scientific American, 47(12): 179, September 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n12/pdf/scientificamerican09161882-
179.pdf.

Anonymous:1882:IR

[3193] Anonymous. Improvement in razors. Scientific American, 47 (12):179, September 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n12/pdf/scientificamerican09161882-
179a.pdf.

Anonymous:1882:NCSc

[3194] Anonymous. New cotton seed huller. Scientific American, 47 (12):179, September 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n12/pdf/scientificamerican09161882-
179b.pdf.

Hovet:1882:AAA

[3195] H. C. Hovet. The American Association for the Advancement of Science. Scientific American, 47(12):180–181, September 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n12/pdf/scientificamerican09161882-180.pdf.

Anonymous:1882:IAE

[3196] Anonymous. Improved agricultural engine. Scientific American, 47 (12):182, September 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n12/pdf/scientificamerican09161882-
182.pdf.

Anonymous:1882:FD

[3197] Anonymous. Ferments and diseases. Scientific American, 47 (12):183, September 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n12/pdf/scientificamerican09161882-
183.pdf.

Anonymous:1882:DRPf

[3198] Anonymous. Decisions relating to patents. Scientific American, 47 (12):184, September 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n12/pdf/scientificamerican09161882-
184.pdf.

Anonymous:1882:EIx

[3199] Anonymous. Engineering inventions. Scientific American, 47(12): 186, September 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n12/pdf/scientificamerican09161882-
186.pdf.

Anonymous:1882:MIcc

[3200] Anonymous. Mechanical inventions. Scientific American, 47(12): 186, September 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n12/pdf/scientificamerican09161882-
186a.pdf.

Anonymous:1882:AIp

[3201] Anonymous. Agricultural inventions. Scientific American, 47 (12):186, September 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n12/pdf/scientificamerican09161882-
186b.pdf.

Anonymous:1882:TIa

[3202] Anonymous.  Textile invention.  Scientific American, 47(12):

186, September 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n12/pdf/scientificamerican09161882-
186c.pdf.

Anonymous:1882:MIcd

[3203] Anonymous. Metallurgical invention. Scientific American, 47 (12):186, September 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n12/pdf/scientificamerican09161882-
186d.pdf.

Anonymous:1882:MIce

[3204] Anonymous. Miscellaneous inventions. Scientific American, 47 (12):186, September 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n12/pdf/scientificamerican09161882-
186e.pdf.

Anonymous:1882:Cq

[3205] Anonymous. Correspondence. Scientific American, 47(13):197, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n13/pdf/scientificamerican09231882-197.pdf.

Anonymous:1882:IRIr

[3206] Anonymous. Inventions: Recent inventions. Scientific American, 47(13):201, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
201.pdf.

Anonymous:1882:BPbd

[3207] Anonymous. Business and personal. Scientific American, 47 (13):203, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
203.pdf.

Anonymous:1882:BNBk

[3208] Anonymous. Books: New books and publications. Scientific American, 47(13):203, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
203a.pdf.

Anonymous:1882:NQbc

[3209] Anonymous. Notes and queries. Scientific American, 47(13): 203, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
203b.pdf.

Anonymous:1882:IIbf

[3210] Anonymous. Index of inventions. Scientific American, 47(13): 204, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
204.pdf.

Anonymous:1882:PNPz

[3211] Anonymous. Patent notes: Patents. Scientific American, 47 (13):205, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
205.pdf.

Anonymous:1882:RCP

[3212] Anonymous. Rosaline color patent, 250,247. Scientific American, 47(13):192, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
192.pdf.

Anonymous:1882:ECE

[3213] Anonymous. The Egyptian Campaign ended. Scientific American, 47(13):192, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
192a.pdf.

Anonymous:1882:TH

[3214] Anonymous. A telephonic hoax. Scientific American, 47(13): 192, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
192b.pdf.

Anonymous:1882:SHG

[3215] Anonymous. Steam for heating greenhouses. Scientific American, 47(13):192, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
192c.pdf.

Anonymous:1882:AIFa

[3216] Anonymous. The American Institute Fair. Scientific American, 47 (13):192, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
192d.pdf.

Anonymous:1882:PSF

[3217] Anonymous. Postage stamp frauds. Scientific American, 47(13): 192, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
192e.pdf.

Anonymous:1882:PMN

[3218] Anonymous. Pliocene man in Nevada. Scientific American, 47(13): 192–193, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
192f.pdf.

Anonymous:1882:BAa

[3219] Anonymous. The British Association. Scientific American, 47(13): 193–194, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
193.pdf.

Anonymous:1882:IWMb

[3220] Anonymous. Improved washing machine. Scientific American, 47 (13):194, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
194.pdf.

Anonymous:1882:ICCc

[3221] Anonymous. Improved corn crusher. Scientific American, 47 (13):194, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
194a.pdf.

Anonymous:1882:BSF

[3222] Anonymous. Bloomsdale seed farm. Scientific American, 47(13): 195–196, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
195.pdf.

Anonymous:1882:DRPg

[3223] Anonymous. Decisions relating to patents, etc. Scientific American, 47(13):196–197, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
196.pdf.

Anonymous:1882:ISBa

[3224] Anonymous. Improvement in steam boilers. Scientific American, 47(13):198, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
198.pdf.

Anonymous:1882:ISFb

[3225] Anonymous. Improved spinning frame. Scientific American, 47(13): 198–199, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
198a.pdf.

Anonymous:1882:NRW

[3226] Anonymous. The natural reddening of water. Scientific American, 47 (13):199–200, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
199.pdf.

Anonymous:1882:EIy

[3227] Anonymous. Engineering inventions. Scientific American, 47(13): 202, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
202.pdf.

Anonymous:1882:MIcf

[3228] Anonymous. Mechanical inventions. Scientific American, 47(13): 202, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
202a.pdf.

Anonymous:1882:AIq

[3229] Anonymous. Agricultural inventions. Scientific American, 47 (13):202, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
202b.pdf.

Anonymous:1882:TIb

[3230] Anonymous. Textile inventions. Scientific American, 47(13): 202, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
202c.pdf.

Anonymous:1882:MIcg

[3231] Anonymous. Metallurgical invention. Scientific American, 47 (13):202, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
202d.pdf.

Anonymous:1882:EIz

[3232] Anonymous. Electrical invention. Scientific American, 47(13): 202, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
202e.pdf.

Anonymous:1882:MIch

[3233] Anonymous. Miscellaneous inventions. Scientific American, 47 (13):202, September 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n13/pdf/scientificamerican09231882-
202f.pdf.

Anonymous:1882:Ct

[3234] Anonymous. Correspondence. Scientific American, 47(14):213, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n14/pdf/scientificamerican09301882-213.pdf.

Anonymous:1882:IRIs

[3235] Anonymous. Inventions: Recent inventions. Scientific American, 47(14):217, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
217.pdf.

Anonymous:1882:NQbd

[3236] Anonymous. Notes and queries. Scientific American, 47(14): 218, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
218c.pdf.

Anonymous:1882:BPbe

[3237] Anonymous. Business and personal. Scientific American, 47 (14):219, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
219.pdf.

Anonymous:1882:IIbg

[3238] Anonymous. Index of inventions. Scientific American, 47(14): 219–220, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
219a.pdf.

Anonymous:1882:PNPba

[3239] Anonymous. Patent notes: Patents. Scientific American, 47 (14):220, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
220.pdf.

Anonymous:1882:NPTb

[3240] Anonymous. The new Princeton Telescope. Scientific American, 47(14):207, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
207.pdf.

Anonymous:1882:CDb

[3241] Anonymous. Comets D and E 1882. Scientific American, 47 (14):208, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
208.pdf.

Anonymous:1882:ITB

[3242] Anonymous. Improved transportation of beef and beef cattle. Scientific American, 47(14):208, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
208a.pdf.

Anonymous:1882:NTA

[3243] Anonymous. National Telephone Association. Scientific American, 47(14):209, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
209.pdf.

Anonymous:1882:DII

[3244] Anonymous. Dairy interests of Italy. Scientific American, 47(14): 209–210, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
209a.pdf.

Anonymous:1882:IPC

[3245] Anonymous. Improved plow colter. Scientific American, 47(14): 210, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
210.pdf.

Anonymous:1882:CDC

[3246] Anonymous. Combined drier and cooler. Scientific American, 47 (14):210, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
210a.pdf.

Anonymous:1882:DPB

[3247] Anonymous. Drop press belt protector. Scientific American, 47 (14):210–211, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
210b.pdf.

Anonymous:1882:PRH

[3248] Anonymous. Portable roof for hay and grain ricks. Scientific American, 47(14):211, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
211.pdf.

Anonymous:1882:BPbf

[3249] Anonymous. Baling press. Scientific American, 47(14):211, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n14/pdf/scientificamerican09301882-211a.pdf.

Anonymous:1882:NMP

[3250] Anonymous. New millstone paint staff. Scientific American, 47 (14):211, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
211b.pdf.

Anonymous:1882:DRPh

[3251] Anonymous. Decisions relating to patents, trade marks, etc. Scientific American, 47(14):212, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
212.pdf.

Anonymous:1882:CRHa

[3252] Anonymous. Check rein holder. Scientific American, 47(14): 213, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
213a.pdf.

Anonymous:1882:ISSb

[3253] Anonymous. Improved squaring shear. Scientific American, 47 (14):214, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
214.pdf.

Anonymous:1882:CIG

[3254] Anonymous. Clamond’s incandescent gas lamp. Scientific American, 47(14):214, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
214a.pdf.

Anonymous:1882:G

[3255] Anonymous. Galvani. Scientific American, 47(14):215, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n14/pdf/scientificamerican09301882-215.pdf.

Anonymous:1882:RTR

[3256] Anonymous. Regnard’s temperature regulator. Scientific American, 47(14):215–216, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
215a.pdf.

Anonymous:1882:SV

[3257] Anonymous. The steam velocipede. Scientific American, 47(14): 216, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
216.pdf.

Hovey:1882:AC

[3258] H. C. Hovey. The Ausable Chasm. Scientific American, 47 (14):217, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
217a.pdf.

Anonymous:1882:EIba

[3259] Anonymous. Engineering inventions. Scientific American, 47(14): 218, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
218.pdf.

Anonymous:1882:MIci

[3260] Anonymous. Metallurgical invention. Scientific American, 47 (14):218, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
218a.pdf.

Anonymous:1882:MIcj

[3261] Anonymous. Miscellaneous inventions. Scientific American, 47 (14):218, September 30, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n14/pdf/scientificamerican09301882-
218b.pdf.

Anonymous:1882:Cu

[3262] Anonymous. Correspondence. Scientific American, 47(15):229, October 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n15/pdf/scientificamerican10071882-229.pdf.

Anonymous:1882:IRIt

[3263] Anonymous. Inventions: Recent inventions. Scientific American, 47(15):233, October 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n15/pdf/scientificamerican10071882-
233.pdf.

Anonymous:1882:BPbg

[3264] Anonymous. Business and personal. Scientific American, 47(15):235, October 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n15/pdf/scientificamerican10071882-235.pdf.

Anonymous:1882:NQbe

[3265] Anonymous. Notes and queries. Scientific American, 47(15):235, October 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n15/pdf/scientificamerican10071882-235a.pdf.

Anonymous:1882:IIbh

[3266] Anonymous. Index of inventions. Scientific American, 47(15):235–236, October 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n15/pdf/scientificamerican10071882-235b.pdf.

Anonymous:1882:PNPbb

[3267] Anonymous. Patent notes: Patents. Scientific American, 47(15):236, October 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n15/pdf/scientificamerican10071882-236.pdf.

Anonymous:1882:IIDa

[3268] Anonymous. The Italian ironclad Dandolo. Scientific American, 47(15):223, October 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n15/pdf/scientificamerican10071882-
223.pdf.

Anonymous:1882:BFE

[3269] Anonymous. Blue fire as an explosive. Scientific American, 47(15):224, October 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n15/pdf/scientificamerican10071882-224.pdf.

Anonymous:1882:CMP

[3270] Anonymous. The common mushroom and its poison. Scientific American, 47(15):224, October 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n15/pdf/scientificamerican10071882-
224a.pdf.

Anonymous:1882:MTM

[3271] Anonymous. Modification of the trade mark law. Scientific American, 47(15):224, October 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n15/pdf/scientificamerican10071882-
224b.pdf.

Anonymous:1882:MGT

[3272] Anonymous. Mr. Golding’s theory of Mississippi floods and their prevention. Scientific American, 47(15):225, October 7, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v47/
n15/pdf/scientificamerican10071882-225.pdf.

Anonymous:1882:TF

[3273] Anonymous. Tile fish. Scientific American, 47(15):225, October 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n15/pdf/scientificamerican10071882-225a.pdf.

Anonymous:1882:NOF

[3274] Anonymous. New old fashioned houses. Scientific American, 47 (15):226–227, October 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n15/pdf/scientificamerican10071882-
226.pdf.

Anonymous:1882:NAR

[3275] Anonymous. New alarm register. Scientific American, 47(15):227, October 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n15/pdf/scientificamerican10071882-227.pdf.

Anonymous:1882:IPW

[3276] Anonymous. Improved portable water closet. Scientific American, 47(15):227, October 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n15/pdf/scientificamerican10071882-
227a.pdf.

Anonymous:1882:APO

[3277] Anonymous. Aspects of the planets for October. Scientific American, 47(15):228, October 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n15/pdf/scientificamerican10071882-
228.pdf.

Anonymous:1882:NHNb

[3278] Anonymous. Natural history notes. Scientific American, 47(15):229, October 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n15/pdf/scientificamerican10071882-229a.pdf.

Anonymous:1882:CEP

[3279] Anonymous. Chronogbaph for engineering purposes, with the hipp escapement. Scientific American, 47(15):230, October 7, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v47/
n15/pdf/scientificamerican10071882-230.pdf.

Anonymous:1882:HH

[3280] Anonymous. The Heloderma Horridum. Scientific American, 47 (15):231–232, October 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n15/pdf/scientificamerican10071882-
231.pdf.

Anonymous:1882:NKR

[3281] Anonymous. New key ring. Scientific American, 47(15):232, October 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n15/pdf/scientificamerican10071882-232.pdf.

Anonymous:1882:EIbb

[3282] Anonymous. Engineering inventions. Scientific American, 47(15):234, October 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n15/pdf/scientificamerican10071882-234.pdf.

Anonymous:1882:MIck

[3283] Anonymous. Metallurgical inventions. Scientific American, 47(15):234, October 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n15/pdf/scientificamerican10071882-234a.pdf.

Anonymous:1882:EIbc

[3284] Anonymous. Electrical invention. Scientific American, 47(15):234, October 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n15/pdf/scientificamerican10071882-234b.pdf.

Anonymous:1882:TIc

[3285] Anonymous. Textile invention. Scientific American, 47(15):234, October 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n15/pdf/scientificamerican10071882-234c.pdf.

Anonymous:1882:AIr

[3286] Anonymous. Agricultural inventions. Scientific American, 47(15):234, October 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n15/pdf/scientificamerican10071882-234d.pdf.

Anonymous:1882:MIcl

[3287] Anonymous. Miscellaneous inventions. Scientific American, 47(15):234, October 7, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n15/pdf/scientificamerican10071882-234e.pdf.

Anonymous:1882:IRIu

[3288] Anonymous. Inventions: Recent inventions. Scientific American, 47(16):249, October 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n16/pdf/scientificamerican10141882-
249.pdf.

Anonymous:1882:BPbh

[3289] Anonymous. Business and personal. Scientific American, 47 (16):251, October 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n16/pdf/scientificamerican10141882-
251.pdf.

Anonymous:1882:NQbf

[3290] Anonymous. Notes and queries. Scientific American, 47(16):251, October 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n16/pdf/scientificamerican10141882-251a.pdf.

Anonymous:1882:IIbi

[3291] Anonymous. Index of inventions. Scientific American, 47(16): 251–252, October 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n16/pdf/scientificamerican10141882-
251b.pdf.

Anonymous:1882:PNPbc

[3292] Anonymous. Patent notes: Patents. Scientific American, 47 (16):252, October 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n16/pdf/scientificamerican10141882-
252.pdf.

Lentz:1882:FLT

[3293] G. Lentz. Fireless locomotives for tramways. Scientific American, 47(16):239–240, October 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n16/pdf/scientificamerican10141882-
239.pdf.

Anonymous:1882:JEL

[3294] Anonymous. The Jeannette Expedition — Lieutenant Danenhower’s lecture. Scientific American, 47(16):240, October 14, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v47/
n16/pdf/scientificamerican10141882-240.pdf.

Anonymous:1882:AEL

[3295] Anonymous. Accident with electric light wires. Scientific American, 47(16):240, October 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n16/pdf/scientificamerican10141882-
240a.pdf.

Anonymous:1882:FUW

[3296] Anonymous. Fireproof upper works for steamers. Scientific American, 47(16):240–241, October 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n16/pdf/scientificamerican10141882-
240b.pdf.

Anonymous:1882:GCa

[3297] Anonymous. The Great Comet of 1882. Scientific American, 47 (16):241, October 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n16/pdf/scientificamerican10141882-
241.pdf.

Anonymous:1882:GCE

[3298] Anonymous. The Great Comet in the East. Scientific American, 47(16):241, October 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n16/pdf/scientificamerican10141882-
241a.pdf.

Anonymous:1882:SFF

[3299] Anonymous. Storm, flood, and frost signals by telegraph and cannon. Scientific American, 47(16):241, October 14, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v47/
n16/pdf/scientificamerican10141882-241b.pdf.

Anonymous:1882:TRG

[3300] Anonymous. Tardily recognized genius. Scientific American, 47 (16):241, October 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n16/pdf/scientificamerican10141882-
241c.pdf.

Anonymous:1882:IOS

[3301] Anonymous. Improved ore separator. Scientific American, 47 (16):242–243, October 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n16/pdf/scientificamerican10141882-
242.pdf.

Anonymous:1882:SHb

[3302] Anonymous. Steam hammers. Scientific American, 47(16):243, October 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n16/pdf/scientificamerican10141882-243.pdf.

Anonymous:1882:DRPi

[3303] Anonymous. Decisions relating to patents. Scientific American, 47 (16):244–245, October 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n16/pdf/scientificamerican10141882-
244.pdf.

Anonymous:1882:GCb

[3304] Anonymous. The Great Comet of 1882. Scientific American, 47 (16):245, October 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n16/pdf/scientificamerican10141882-
245.pdf.

Anonymous:1882:ISEb

[3305] Anonymous. Improved stationary engine. Scientific American, 47(16):246, October 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n16/pdf/scientificamerican10141882-
246.pdf.

Anonymous:1882:NFF

[3306] Anonymous. The new food fish — (Sebastes Dactyloptera). Scientific American, 47(16):247, October 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v47/n16/pdf/scientificamerican10141882-
247.pdf.

Jago:1882:C

[3307] William Jago. Crystals. Scientific American, 47(16):248, October 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n16/pdf/scientificamerican10141882-248.pdf.

Anonymous:1882:EIbd

[3308] Anonymous. Engineering inventions. Scientific American, 47(16): 250, October 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n16/pdf/scientificamerican10141882-
250.pdf.

Anonymous:1882:MIcm

[3309] Anonymous. Mechanical inventions. Scientific American, 47(16): 250, October 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n16/pdf/scientificamerican10141882-
250a.pdf.

Anonymous:1882:EIbe

[3310] Anonymous. Electrical invention. Scientific American, 47(16):250, October 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n16/pdf/scientificamerican10141882-250b.pdf.

Anonymous:1882:TId

[3311] Anonymous. Textile invention. Scientific American, 47(16):250, October 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n16/pdf/scientificamerican10141882-250c.pdf.

Anonymous:1882:AIs

[3312] Anonymous. Agricultural inventions. Scientific American, 47 (16):250, October 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n16/pdf/scientificamerican10141882-
250d.pdf.

Anonymous:1882:MIcn

[3313] Anonymous. Miscellaneous inventions. Scientific American, 47 (16):250, October 14, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n16/pdf/scientificamerican10141882-
250e.pdf.

Anonymous:1882:IRIv

[3314] Anonymous. Inventions: Recent inventions. Scientific American, 47(17):264, October 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n17/pdf/scientificamerican10211882-
264.pdf.

Anonymous:1882:BPbi

[3315] Anonymous. Business and personal. Scientific American, 47 (17):266, October 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n17/pdf/scientificamerican10211882-
266.pdf.

Anonymous:1882:BNBl 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n17/pdf/scientificamerican10211882-
266a.pdf.

Anonymous:1882:NQbg

[3317] Anonymous. Notes and queries. Scientific American, 47(17):266, October 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n17/pdf/scientificamerican10211882-266b.pdf.

Anonymous:1882:IIbj

[3318] Anonymous. Index of inventions. Scientific American, 47(17): 266–268, October 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n17/pdf/scientificamerican10211882-
266d.pdf.

Anonymous:1882:PNPbd

[3319] Anonymous. Patent notes: Patents. Scientific American, 47 (17):268, October 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n17/pdf/scientificamerican10211882-
268.pdf.

Anonymous:1882:PAL

[3320] Anonymous. Projected American line of fast ocean steamships. Scientific American, 47(17):255–256, October 21, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v47/
n17/pdf/scientificamerican10211882-255.pdf.

Anonymous:1882:DAC

[3321] Anonymous. A dubious amendment of the copyright law. Scientific American, 47(17):256, October 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v47/n17/pdf/scientificamerican10211882-
256.pdf.

Anonymous:1882:NASa

[3322] Anonymous. New antiseptics for surgical use. Scientific American, 47(17):256–257, October 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n17/pdf/scientificamerican10211882-
256a.pdf.

Anonymous:1882:SMG

[3323] Anonymous. Sewing machines in Germany. Scientific American, 47(17):257, October 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n17/pdf/scientificamerican10211882-
257.pdf.

Anonymous:1882:NSFb

[3324] Anonymous. A new smelting furnace. Scientific American, 47 (17):258–259, October 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n17/pdf/scientificamerican10211882-
258.pdf.

Anonymous:1882:ICSd

[3325] Anonymous. Improved circular saw mill. Scientific American, 47 (17):259–261, October 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n17/pdf/scientificamerican10211882-
259.pdf.

Anonymous:1882:ACC

[3326] Anonymous. Apparatus for changing car trucks. Scientific American, 47(17):261, October 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n17/pdf/scientificamerican10211882-
261.pdf.

Anonymous:1882:ITC

[3327] Anonymous. Improved three-cylinder engine. Scientific American, 47(17):262, October 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n17/pdf/scientificamerican10211882-
262.pdf.

Anonymous:1882:NFSb

[3328] Anonymous. New four-sided moulder. Scientific American, 47 (17):262, October 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n17/pdf/scientificamerican10211882-
262a.pdf.

Anonymous:1882:PSCb

[3329] Anonymous. The prospects of silk culture in the United States. Scientific American, 47(17):263, October 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v47/n17/pdf/scientificamerican10211882-
263.pdf.

Anonymous:1882:EIbf

[3330] Anonymous. Engineering inventions. Scientific American, 47(17): 265, October 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n17/pdf/scientificamerican10211882-
265.pdf.

Anonymous:1882:EIbg

[3331] Anonymous. Electrical invention. Scientific American, 47(17):265, October 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n17/pdf/scientificamerican10211882-265a.pdf.

Anonymous:1882:AIt

[3332] Anonymous. Agricultural inventions. Scientific American, 47 (17):265, October 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n17/pdf/scientificamerican10211882-
265b.pdf.

Anonymous:1882:TIe

[3333] Anonymous. Textile invention. Scientific American, 47(17):265, October 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n17/pdf/scientificamerican10211882-265c.pdf.

Anonymous:1882:MIco

[3334] Anonymous. Miscellaneous inventions. Scientific American, 47 (17):265, October 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n17/pdf/scientificamerican10211882-
265d.pdf.

Anonymous:1882:CRj

[3335] Anonymous. Communications received. Scientific American, 47 (17):266, October 21, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n17/pdf/scientificamerican10211882-
266c.pdf.

Anonymous:1882:Cv

[3336] Anonymous. Correspondence. Scientific American, 47(18):276, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n18/pdf/scientificamerican10281882-276.pdf.

Anonymous:1882:IRIw

[3337] Anonymous. Inventions: Recent inventions. Scientific American, 47(18):280, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n18/pdf/scientificamerican10281882-
280.pdf.

Anonymous:1882:BPbj

[3338] Anonymous. Business and personal. Scientific American, 47 (18):282, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n18/pdf/scientificamerican10281882-
282.pdf.

Anonymous:1882:BNBm

[3339] Anonymous. Books: New books and publications. Scientific American, 47(18):282, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n18/pdf/scientificamerican10281882-
282a.pdf.

Anonymous:1882:NQbh

[3340] Anonymous. Notes and queries. Scientific American, 47(18): 282–283, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n18/pdf/scientificamerican10281882-
282b.pdf.

Anonymous:1882:IIbk

[3341] Anonymous. Index of inventions. Scientific American, 47(18): 283–284, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n18/pdf/scientificamerican10281882-
283.pdf.

Anonymous:1882:PNPbe

[3342] Anonymous. Patent notes: Patents. Scientific American, 47 (18):284, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n18/pdf/scientificamerican10281882-
284.pdf.

Anonymous:1882:MFT

[3343] Anonymous. Machine for fixing the threads of fabrics. Scientific American, 47(18):271–272, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n18/pdf/scientificamerican10281882-
271.pdf.

Anonymous:1882:PYW

[3344] Anonymous. The past year’s work in the Patent Office. Scientific American, 47(18):272, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n18/pdf/scientificamerican10281882-
272.pdf.

Anonymous:1882:GME

[3345] Anonymous. The Garfield Monument Exhibition. Scientific American, 47(18):272, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n18/pdf/scientificamerican10281882-
272a.pdf.

Anonymous:1882:QDO

[3346] Anonymous. Queer doings in oil. Scientific American, 47(18): 272–273, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n18/pdf/scientificamerican10281882-
272b.pdf.

Anonymous:1882:SWV

[3347] Anonymous. Speed in war vessels. Scientific American, 47(18):273, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n18/pdf/scientificamerican10281882-273.pdf.

Anonymous:1882:MRI

[3348] Anonymous. More railroad inventions wanted. Scientific American, 47(18):273, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n18/pdf/scientificamerican10281882-
273a.pdf.

Anonymous:1882:EWS

[3349] Anonymous. The educational want of the South. Scientific American, 47(18):273, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n18/pdf/scientificamerican10281882-
273b.pdf.

Anonymous:1882:CMN

[3350] Anonymous. Cheaper meat for New York. Scientific American, 47(18):273, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n18/pdf/scientificamerican10281882-
273c.pdf.

Anonymous:1882:PG

[3351] Anonymous. Plane guide. Scientific American, 47(18):274, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n18/pdf/scientificamerican10281882-274.pdf.

Anonymous:1882:IHA

[3352] Anonymous. Improved hot air engine. Scientific American, 47 (18):274, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n18/pdf/scientificamerican10281882-
274a.pdf.

Anonymous:1882:NBFb

[3353] Anonymous. New box fastener. Scientific American, 47(18):275, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n18/pdf/scientificamerican10281882-275.pdf.

Anonymous:1882:EFS

[3354] Anonymous. An English fourteen-story house. Scientific American, 47(18):275, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n18/pdf/scientificamerican10281882-
275a.pdf.

Anonymous:1882:APN

[3355] Anonymous. Aspects of the planets for November. Scientific American, 47(18):277, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n18/pdf/scientificamerican10281882-
277.pdf.

Anonymous:1882:JWM

[3356] Anonymous. Jacquet’s water meter. Scientific American, 47 (18):278, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n18/pdf/scientificamerican10281882-
278.pdf.

Anonymous:1882:FPE

[3357] Anonymous. A fuel that produces electricity. Scientific American, 47(18):279, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n18/pdf/scientificamerican10281882-
279.pdf.

Anonymous:1882:MFA

[3358] Anonymous. Meeting of the French Association for the Advancement of Science at La Rochelle. Scientific American, 47(18):279–280, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n18/pdf/scientificamerican10281882-279a.pdf.

Anonymous:1882:EIbh

[3359] Anonymous. Engineering inventions. Scientific American, 47(18): 281, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n18/pdf/scientificamerican10281882-
281.pdf.

Anonymous:1882:MIcp

[3360] Anonymous. Mechanical inventions. Scientific American, 47(18): 281, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n18/pdf/scientificamerican10281882-
281a.pdf.

Anonymous:1882:MIcq

[3361] Anonymous. Miscellaneous inventions. Scientific American, 47 (18):281, October 28, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n18/pdf/scientificamerican10281882-
281b.pdf.

Anonymous:1882:IRIx

[3362] Anonymous. Inventions: Recent inventions. Scientific American, 47(19):297, November 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n19/pdf/scientificamerican11041882-
297.pdf.

Anonymous:1882:BPbk

[3363] Anonymous. Business and personal. Scientific American, 47 (19):299, November 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n19/pdf/scientificamerican11041882-
299.pdf.

Anonymous:1882:NQbi

[3364] Anonymous. Notes and queries. Scientific American, 47(19): 299, November 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n19/pdf/scientificamerican11041882-
299a.pdf.

Anonymous:1882:IIbl

[3365] Anonymous. Index of inventions. Scientific American, 47(19): 299–300, November 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n19/pdf/scientificamerican11041882-
299b.pdf.

Anonymous:1882:PNPbf

[3366] Anonymous. Patent notes: Patents. Scientific American, 47 (19):300, November 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n19/pdf/scientificamerican11041882-
300.pdf.

Anonymous:1882:TGC

[3367] Anonymous. The transfer of gold coin in New York, progress of telegraphy, and more. Scientific American, 47(19):287, November 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n19/pdf/scientificamerican11041882-287.pdf.

Anonymous:1882:TTP

[3368] Anonymous. Twenty-ton portable steam crane. Scientific American, 47(19):287, November 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n19/pdf/scientificamerican11041882-
287a.pdf.

Anonymous:1882:WWN

[3369] Anonymous. Well water in New York City. Scientific American, 47(19):288, November 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n19/pdf/scientificamerican11041882-
288.pdf.

Anonymous:1882:FF

[3370] Anonymous. Fire without fuels. Scientific American, 47(19): 288, November 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n19/pdf/scientificamerican11041882-
288a.pdf.

Anonymous:1882:AES

[3371] Anonymous. Annular eclipse of the Sun. Scientific American, 47 (19):288, November 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n19/pdf/scientificamerican11041882-
288b.pdf.

Anonymous:1882:Cx

[3372] Anonymous. Corpulence. Scientific American, 47(19):289, November 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n19/pdf/scientificamerican11041882-289.pdf.

Anonymous:1882:WS

[3373] Anonymous. Wool shearing. Scientific American, 47(19):290, November 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n19/pdf/scientificamerican11041882-290.pdf.

Anonymous:1882:CDAa

[3374] Anonymous. Chatel’s direct acting steam cock. Scientific American, 47(19):291, November 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n19/pdf/scientificamerican11041882-
291.pdf.

Anonymous:1882:PST

[3375] Anonymous. Perreaux’s steam tricycle. Scientific American, 47 (19):291–293, November 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n19/pdf/scientificamerican11041882-
291a.pdf.

Anonymous:1882:ADD

[3376] Anonymous. Apparatus for determining the degree of moistuer of steam. Scientific American, 47(19):294, November 4, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v47/
n19/pdf/scientificamerican11041882-294.pdf.

Anonymous:1882:SAR

[3377] Anonymous. Sanitary arrangements in the residence of Mr. Cornelius Vanderbilt. Scientific American, 47(19):294, November 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n19/pdf/scientificamerican11041882-294a.pdf.

Anonymous:1882:TRS

[3378] Anonymous. Tricycle roller skate. Scientific American, 47(19): 295, November 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n19/pdf/scientificamerican11041882-
295.pdf.

Holder:1882:GHH

[3379] C. F. Holder. Giant herrings and herring pikes. Scientific American, 47(19):295–296, November 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n19/pdf/scientificamerican11041882-
295a.pdf.

Anonymous:1882:EIbi

[3380] Anonymous. Engineering inventions. Scientific American, 47(19): 298, November 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n19/pdf/scientificamerican11041882-
298.pdf.

Anonymous:1882:MIcr

[3381] Anonymous. Mechanical inventions. Scientific American, 47(19): 298, November 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n19/pdf/scientificamerican11041882-
298a.pdf.

Anonymous:1882:AIu

[3382] Anonymous. Agricultural inventions. Scientific American, 47 (19):298, November 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n19/pdf/scientificamerican11041882-
298b.pdf.

Anonymous:1882:MIcs

[3383] Anonymous. Miscellaneous inventions. Scientific American, 47 (19):298, November 4, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n19/pdf/scientificamerican11041882-
298c.pdf.

Anonymous:1882:IRIy

[3384] Anonymous. Inventions: Recent inventions. Scientific American, 47(2): 25, July 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n2/pdf/scientificamerican07081882-25.pdf.

Anonymous:1882:BPbl

[3385] Anonymous. Business and personal. Scientific American, 47(2):27, July 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n2/pdf/scientificamerican07081882-27.pdf.

Anonymous:1882:NQbj

[3386] Anonymous. Notes and queries. Scientific American, 47(2):27, July 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n2/pdf/scientificamerican07081882-27a.pdf.

Anonymous:1882:IIbm

[3387] Anonymous. Index of inventions. Scientific American, 47(2):27–28, July 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n2/pdf/scientificamerican07081882-27b.pdf.

Anonymous:1882:PNPbg

[3388] Anonymous. Patent notes: Patents. Scientific American, 47(2):28, July 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n2/pdf/scientificamerican07081882-28.pdf.

Anonymous:1882:SRM

[3389] Anonymous. The slaughter of railway men. Scientific American, 47(2): 16, July 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n2/pdf/scientificamerican07081882-16.pdf.

Anonymous:1882:EIbj

[3390] Anonymous. Employment for the idle. Scientific American, 47(2):16– 18, July 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n2/pdf/scientificamerican07081882-16a.pdf.

Anonymous:1882:BI

[3391] Anonymous. The binary injector. Scientific American, 47(2):18, July 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n2/pdf/scientificamerican07081882-18.pdf.

Anonymous:1882:RTD

[3392] Anonymous. Raffard’s transmission dynamometer. Scientific American, 47(2):18, July 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n2/pdf/scientificamerican07081882-
18a.pdf.

Anonymous:1882:JRN

[3393] Anonymous. John Roach, of New York. Scientific American, 47(2):19– 20, July 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n2/pdf/scientificamerican07081882-19.pdf.

Anonymous:1882:SFS

[3394] Anonymous. A simple form of storage or secondary battery. Scientific American, 47(2):20–21, July 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n2/pdf/scientificamerican07081882-
20.pdf.

Anonymous:1882:ABF

[3395] Anonymous. Automatic boiler feeder. Scientific American, 47(2):22, July 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n2/pdf/scientificamerican07081882-22.pdf.

Anonymous:1882:PMO

[3396] Anonymous. Press for the manufacture of oils. Scientific American, 47(2):22, July 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n2/pdf/scientificamerican07081882-
22a.pdf.

Anonymous:1882:EMA

[3397] Anonymous. Electricity as a motor for aerial navigation. Scientific American, 47(2):23–24, July 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n2/pdf/scientificamerican07081882-
23.pdf.

Anonymous:1882:MIct

[3398] Anonymous. Mechanical inventions. Scientific American, 47(2):26, July 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n2/pdf/scientificamerican07081882-26.pdf.

Anonymous:1882:MIcu

[3399] Anonymous. Miscellaneous inventions. Scientific American, 47(2):26, July 8, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n2/pdf/scientificamerican07081882-26a.pdf.

Anonymous:1882:Cy

[3400] Anonymous. Correspondence. Scientific American, 47(20):308, November 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n20/pdf/scientificamerican11111882-308.pdf.

Anonymous:1882:IRIz

[3401] Anonymous. Inventions: Recent inventions. Scientific American, 47(20):312, November 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n20/pdf/scientificamerican11111882-
312.pdf.

Anonymous:1882:BPbm

[3402] Anonymous. Business and personal. Scientific American, 47 (20):314, November 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n20/pdf/scientificamerican11111882-
314.pdf.

Anonymous:1882:BNBn

[3403] Anonymous. Books: New books and publications. Scientific American, 47(20):314, November 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n20/pdf/scientificamerican11111882-
314a.pdf.

Anonymous:1882:NQbk

[3404] Anonymous. Notes and queries. Scientific American, 47(20): 314–315, November 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n20/pdf/scientificamerican11111882-
314b.pdf.

Anonymous:1882:IIbn

[3405] Anonymous. Index of inventions. Scientific American, 47(20): 315–316, November 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n20/pdf/scientificamerican11111882-
315.pdf.

Anonymous:1882:PNPbh

[3406] Anonymous. Patent notes: Patents. Scientific American, 47 (20):316, November 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n20/pdf/scientificamerican11111882-
316.pdf.

Anonymous:1882:MDG

[3407] Anonymous. Machine for drawing and glossing silk in skeins. Scientific American, 47(20):303–304, November 11, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v47/
n20/pdf/scientificamerican11111882-303.pdf.

Anonymous:1882:IDC

[3408] Anonymous. An important decision by the Commissioner of Patents. Scientific American, 47(20):304–305, November 11, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v47/
n20/pdf/scientificamerican11111882-304.pdf.

Anonymous:1882:NMb

[3409] Anonymous. The November meteors. Scientific American, 47 (20):305, November 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n20/pdf/scientificamerican11111882-
305.pdf.

Anonymous:1882:MCS

[3410] Anonymous. Magneso-calcite for safe linings. Scientific American, 47(20):305, November 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n20/pdf/scientificamerican11111882-
305a.pdf.

Anonymous:1882:CDd

[3411] Anonymous. Charles Desnos. Scientific American, 47(20):305, November 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n20/pdf/scientificamerican11111882-305b.pdf.

Anonymous:1882:ELW

[3412] Anonymous. Evaporation of liquids. — Wahl’s improved vacuum pan. Scientific American, 47(20):306–307, November 11, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v47/
n20/pdf/scientificamerican11111882-306.pdf.

Anonymous:1882:NSEb

[3413] Anonymous. Naval search electric light. Scientific American, 47 (20):307, November 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n20/pdf/scientificamerican11111882-
307.pdf.

Anonymous:1882:NHNc

[3414] Anonymous. Natural history notes. Scientific American, 47(20): 309–310, November 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n20/pdf/scientificamerican11111882-
309.pdf.

Anonymous:1882:NBPb

[3415] Anonymous. New boat propelled by electricity. Scientific American, 47(20):310, November 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n20/pdf/scientificamerican11111882-
310.pdf.

Holder:1882:PRM

[3416] C. F. Holder. The Peripatus and its remarkable method of capturing prey. Scientific American, 47(20):311, November 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v47/
n20/pdf/scientificamerican11111882-311.pdf.

Anonymous:1882:DF

[3417] Anonymous. The dove flower. Scientific American, 47(20):311, November 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n20/pdf/scientificamerican11111882-311a.pdf.

Anonymous:1882:EIbk

[3418] Anonymous. Engineering inventions. Scientific American, 47(20): 313, November 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n20/pdf/scientificamerican11111882-
313.pdf.

Anonymous:1882:MIcv

[3419] Anonymous. Mechanical inventions. Scientific American, 47(20): 313, November 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n20/pdf/scientificamerican11111882-
313a.pdf.

Anonymous:1882:AIv

[3420] Anonymous. Agricultural inventions. Scientific American, 47 (20):313, November 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n20/pdf/scientificamerican11111882-
313b.pdf.

Anonymous:1882:EIbl

[3421] Anonymous. Electrical invention. Scientific American, 47(20): 313, November 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n20/pdf/scientificamerican11111882-
313c.pdf.

Anonymous:1882:MIcw

[3422] Anonymous. Miscellaneous inventions. Scientific American, 47 (20):313, November 11, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n20/pdf/scientificamerican11111882-
313d.pdf.

Anonymous:1882:Cz

[3423] Anonymous. Correspondence. Scientific American, 47(21):324– 325, November 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n21/pdf/scientificamerican11181882-
324.pdf.

Anonymous:1882:RPI

[3424] Anonymous. Recently patented inventions. Scientific American, 47(21):326, November 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n21/pdf/scientificamerican11181882-
326.pdf.

Anonymous:1882:IRIba

[3425] Anonymous. Inventions: Recent inventions. Scientific American, 47(21):329, November 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n21/pdf/scientificamerican11181882-
329.pdf.

Anonymous:1882:BPbn

[3426] Anonymous. Business and personal. Scientific American, 47 (21):331, November 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n21/pdf/scientificamerican11181882-
331.pdf.

Anonymous:1882:NQbl

[3427] Anonymous. Notes and queries. Scientific American, 47(21): 331, November 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n21/pdf/scientificamerican11181882-
331a.pdf.

Anonymous:1882:IIbo

[3428] Anonymous. Index of inventions. Scientific American, 47(21): 331–332, November 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n21/pdf/scientificamerican11181882-
331b.pdf.

Anonymous:1882:PNPbi

[3429] Anonymous. Patent notes: Patents. Scientific American, 47 (21):332, November 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n21/pdf/scientificamerican11181882-
332.pdf.

Anonymous:1882:NEA

[3430] Anonymous. A new era of alloys. Scientific American, 47(21): 319, November 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n21/pdf/scientificamerican11181882-
319.pdf.

Anonymous:1882:CTC

[3431] Anonymous. Crane for transferring cars. Scientific American, 47 (21):319, November 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n21/pdf/scientificamerican11181882-
319a.pdf.

Anonymous:1882:TVa

[3432] Anonymous. The Transit of Venus. Scientific American, 47 (21):320, November 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n21/pdf/scientificamerican11181882-
320.pdf.

Anonymous:1882:RVP

[3433] Anonymous. Rural views of patents and patent rights. Scientific American, 47(21):320–321, November 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v47/n21/pdf/scientificamerican11181882-
320a.pdf.

Hovey:1882:RPO

[3434] H. C. Hovey. Recent progress in oyster farming. Scientific American, 47(21):321, November 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n21/pdf/scientificamerican11181882-
321.pdf.

Anonymous:1882:NSV

[3435] Anonymous. New stop-valve for locomotive steam pipes. Scientific American, 47(21):322, November 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v47/n21/pdf/scientificamerican11181882-
322.pdf.

Anonymous:1882:MMS

[3436] Anonymous. Machine for making spikes. Scientific American, 47 (21):323, November 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n21/pdf/scientificamerican11181882-
323.pdf.

Anonymous:1882:SUM

[3437] Anonymous. Siemens’ unipolar machine. Scientific American, 47 (21):323, November 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n21/pdf/scientificamerican11181882-
323a.pdf.

Anonymous:1882:GDG

[3438] Anonymous. Garden destroyers. — gall flies of the oak — “Cynipidæ”. Scientific American, 47(21):327–328, November 18, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v47/
n21/pdf/scientificamerican11181882-327.pdf.

Anonymous:1882:TFF

[3439] Anonymous. Theater fires — fireproof fittings demanded. Scientific American, 47(21):329, November 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v47/n21/pdf/scientificamerican11181882-
329a.pdf.

Anonymous:1882:EIbm

[3440] Anonymous. Engineering inventions. Scientific American, 47(21): 330, November 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n21/pdf/scientificamerican11181882-
330.pdf.

Anonymous:1882:MIcx

[3441] Anonymous. Mechanical inventions. Scientific American, 47(21): 330, November 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n21/pdf/scientificamerican11181882-
330a.pdf.

Anonymous:1882:AIw

[3442] Anonymous. Agricultural inventions. Scientific American, 47 (21):330, November 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n21/pdf/scientificamerican11181882-
330b.pdf.

Anonymous:1882:MIcy

[3443] Anonymous. Miscellaneous inventions. Scientific American, 47 (21):330, November 18, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n21/pdf/scientificamerican11181882-
330c.pdf.

Anonymous:1882:IRIbb

[3444] Anonymous. Inventions: Recent inventions. Scientific American, 47(22):345, November 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n22/pdf/scientificamerican11251882-
345.pdf.

Anonymous:1882:IIbp

[3445] Anonymous. Index of inventions. Scientific American, 47(22): 346–347, November 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n22/pdf/scientificamerican11251882-
346d.pdf.

Anonymous:1882:NQbm

[3446] Anonymous. Notes and queries. Scientific American, 47(22): 347–348, November 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n22/pdf/scientificamerican11251882-
347.pdf.

Anonymous:1882:BNBo

[3447] Anonymous. Books: New books and publications. Scientific American, 47(22):348, November 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n22/pdf/scientificamerican11251882-
348.pdf.

Anonymous:1882:BPbo

[3448] Anonymous. Business and personal. Scientific American, 47 (22):348, November 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n22/pdf/scientificamerican11251882-
348a.pdf.

Anonymous:1882:PNPbj

[3449] Anonymous. Patent notes: Patents. Scientific American, 47 (22):348, November 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n22/pdf/scientificamerican11251882-
348b.pdf.

Anonymous:1882:TVb

[3450] Anonymous. Transits of Venus. Scientific American, 47(22): 336, November 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n22/pdf/scientificamerican11251882-
336.pdf.

Anonymous:1882:SBC

[3451] Anonymous. Shipwrecks on the British coasts. Scientific American, 47(22):336, November 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n22/pdf/scientificamerican11251882-
336a.pdf.

Anonymous:1882:RPA

[3452] Anonymous. The rabbit plague in Australia. — a big chance for a paying invention. Scientific American, 47(22):336, November 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n22/pdf/scientificamerican11251882-336b.pdf.

Anonymous:1882:APD

[3453] Anonymous. Aspects of the planets for December. Scientific American, 47(22):337, November 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n22/pdf/scientificamerican11251882-
337.pdf.

Anonymous:1882:ISCb

[3454] Anonymous. Improved straw cutter. Scientific American, 47 (22):338, November 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n22/pdf/scientificamerican11251882-
338.pdf.

Anonymous:1882:EAS

[3455] Anonymous. Electrical apparatus for stopping steam engines. Scientific American, 47(22):338, November 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v47/n22/pdf/scientificamerican11251882-
338a.pdf.

Anonymous:1882:IRC

[3456] Anonymous. Improved railway car. Scientific American, 47(22): 339, November 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n22/pdf/scientificamerican11251882-
339.pdf.

Anonymous:1882:MSS

[3457] Anonymous. Multiple-spindle slot drill. Scientific American, 47 (22):339–340, November 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n22/pdf/scientificamerican11251882-
339a.pdf.

Anonymous:1882:NNE

[3458] Anonymous. Novelties at the New England Institute Fair. Scientific American, 47(22):340, November 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v47/n22/pdf/scientificamerican11251882-
340.pdf.

Anonymous:1882:NASb

[3459] Anonymous. The National Academy of Sciences. Scientific American, 47(22):340–342, November 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n22/pdf/scientificamerican11251882-
340a.pdf.

Anonymous:1882:SRS

[3460] Anonymous. Self-registering ship’s compass. Scientific American, 47(22):342, November 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n22/pdf/scientificamerican11251882-
342.pdf.

Anonymous:1882:DP

[3461] Anonymous. Diaphragm pump. Scientific American, 47(22): 343, November 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n22/pdf/scientificamerican11251882-
343.pdf.

Anonymous:1882:JNE

[3462] Anonymous. Jablochkoff’s new electric motor, the “Ecliptic”. Scientific American, 47(22):343–344, November 25, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v47/
n22/pdf/scientificamerican11251882-343a.pdf.

Anonymous:1882:SCT

[3463] Anonymous. The sea cucumber’s tenant. Scientific American, 47 (22):344, November 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n22/pdf/scientificamerican11251882-
344.pdf.

Anonymous:1882:ICO

[3464] Anonymous. Improved can opener. Scientific American, 47(22): 345, November 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n22/pdf/scientificamerican11251882-
345a.pdf.

Anonymous:1882:EIbn

[3465] Anonymous. Engineering inventions. Scientific American, 47(22): 346, November 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n22/pdf/scientificamerican11251882-
346.pdf.

Anonymous:1882:MIcz

[3466] Anonymous. Mechanical inventions. Scientific American, 47(22): 346, November 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n22/pdf/scientificamerican11251882-
346a.pdf.

Anonymous:1882:AIx

[3467] Anonymous. Agricultural inventions. Scientific American, 47 (22):346, November 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n22/pdf/scientificamerican11251882-
346b.pdf.

Anonymous:1882:MIda

[3468] Anonymous. Miscellaneous inventions. Scientific American, 47 (22):346, November 25, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n22/pdf/scientificamerican11251882-
346c.pdf.

Anonymous:1882:Cba

[3469] Anonymous. Correspondence. Scientific American, 47(23):357, December 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n23/pdf/scientificamerican12021882-357.pdf.

Anonymous:1882:IRIbc

[3470] Anonymous. Inventions: Recent inventions. Scientific American, 47(23):361, December 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n23/pdf/scientificamerican12021882-
361.pdf.

Anonymous:1882:IIbq

[3471] Anonymous. Index of inventions. Scientific American, 47(23): 362–363, December 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n23/pdf/scientificamerican12021882-
362d.pdf.

Anonymous:1882:BNBp

[3472] Anonymous. Books: New books and publications. Scientific American, 47(23):363, December 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n23/pdf/scientificamerican12021882-
363.pdf.

Anonymous:1882:NQbn

[3473] Anonymous. Notes and queries. Scientific American, 47(23): 363–364, December 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n23/pdf/scientificamerican12021882-
363a.pdf.

Anonymous:1882:BPbp

[3474] Anonymous. Business and personal. Scientific American, 47 (23):364, December 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n23/pdf/scientificamerican12021882-
364a.pdf.

Anonymous:1882:PNPbk

[3475] Anonymous. Patent notes: Patents. Scientific American, 47 (23):364, December 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n23/pdf/scientificamerican12021882-
364b.pdf.

Anonymous:1882:ECTb

[3476] Anonymous. Explosives from coal tar. Scientific American, 47 (23):351, December 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n23/pdf/scientificamerican12021882-
351.pdf.

Anonymous:1882:IFEe

[3477] Anonymous. Improved floating elevator. Scientific American, 47 (23):351, December 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n23/pdf/scientificamerican12021882-
351a.pdf.

Anonymous:1882:HD

[3478] Anonymous. Henry Draper. Scientific American, 47(23):352, December 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n23/pdf/scientificamerican12021882-352.pdf.

Anonymous:1882:NIE

[3479] Anonymous. National and International Exhibitions in 1883. Scientific American, 47(23):352, December 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v47/n23/pdf/scientificamerican12021882-
352a.pdf.

Anonymous:1882:RRA

[3480] Anonymous. Romance and reality of animal motion. Scientific American, 47(23):352–353, December 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n23/pdf/scientificamerican12021882-
352b.pdf.

Anonymous:1882:ATP

[3481] Anonymous. Apparatus for testing pressure gauges. Scientific American, 47(23):355, December 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n23/pdf/scientificamerican12021882-
355.pdf.

Anonymous:1882:CTL

[3482] Anonymous. Catapult for throwing life lines. Scientific American, 47(23):355–356, December 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n23/pdf/scientificamerican12021882-
355a.pdf.

Anonymous:1882:NFEc

[3483] Anonymous. New fire escape. Scientific American, 47(23):358, December 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n23/pdf/scientificamerican12021882-358.pdf.

Anonymous:1882:NWW

[3484] Anonymous. New wood-working machine. Scientific American, 47(23):358, December 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n23/pdf/scientificamerican12021882-
358a.pdf.

Coryell:1882:GTW

[3485] John R. Coryell. The great tumble weed of the prairies. Scientific American, 47(23):359–360, December 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v47/n23/pdf/scientificamerican12021882-
359.pdf.

Anonymous:1882:EIbo

[3486] Anonymous. Engineering inventions. Scientific American, 47(23): 362, December 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n23/pdf/scientificamerican12021882-
362.pdf.

Anonymous:1882:MIdb

[3487] Anonymous. Mechanical inventions. Scientific American, 47(23): 362, December 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n23/pdf/scientificamerican12021882-
362a.pdf.

Anonymous:1882:AIy

[3488] Anonymous. Agricultural inventions. Scientific American, 47 (23):362, December 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n23/pdf/scientificamerican12021882-
362b.pdf.

Anonymous:1882:MIdc

[3489] Anonymous. Miscellaneous inventions. Scientific American, 47 (23):362, December 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n23/pdf/scientificamerican12021882-
362c.pdf.

Anonymous:1882:CRk

[3490] Anonymous. Communications received. Scientific American, 47 (23):364, December 2, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n23/pdf/scientificamerican12021882-
364.pdf.

Anonymous:1882:Cbb

[3491] Anonymous. Correspondence. Scientific American, 47(24):373, December 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n24/pdf/scientificamerican12091882-373.pdf.

Anonymous:1882:IRIbd

[3492] Anonymous. Inventions: Recent inventions. Scientific American, 47(24):377, December 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n24/pdf/scientificamerican12091882-
377.pdf.

Anonymous:1882:IIbr

[3493] Anonymous. Index of inventions. Scientific American, 47(24): 378–379, December 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n24/pdf/scientificamerican12091882-
378d.pdf.

Anonymous:1882:BNBq

[3494] Anonymous. Books: New books and publications. Scientific American, 47(24):379, December 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n24/pdf/scientificamerican12091882-
379.pdf.

Anonymous:1882:NQbo

[3495] Anonymous. Notes and queries. Scientific American, 47(24):379, December 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n24/pdf/scientificamerican12091882-379a.pdf.

Anonymous:1882:PNPbl

[3496] Anonymous. Patent notes: Patents. Scientific American, 47 (24):379, December 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n24/pdf/scientificamerican12091882-
379b.pdf.

Anonymous:1882:BPbq

[3497] Anonymous. Business and personal. Scientific American, 47 (24):380, December 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n24/pdf/scientificamerican12091882-
380.pdf.

Anonymous:1882:PSS

[3498] Anonymous. Progress of the street steam supply in New York City. Scientific American, 47(24):368, December 9, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v47/
n24/pdf/scientificamerican12091882-368.pdf.

Anonymous:1882:GLW

[3499] Anonymous. General and local weather service. Scientific American, 47(24):368, December 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n24/pdf/scientificamerican12091882-
368a.pdf.

Anonymous:1882:CSR

[3500] Anonymous. Convention of street railway officials. Scientific American, 47(24):368, December 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n24/pdf/scientificamerican12091882-
368b.pdf.

Anonymous:1882:HOF

[3501] Anonymous. How our farmers are helped by the patent system. Scientific American, 47(24):368–369, December 9, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v47/
n24/pdf/scientificamerican12091882-368c.pdf.

Anonymous:1882:SHF

[3502] Anonymous. Serious hurts that fail to kill. Scientific American, 47(24):369, December 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n24/pdf/scientificamerican12091882-
369.pdf.

Anonymous:1882:ASP

[3503] Anonymous. Alcohol sugar and paper pulp from wood. Scientific American, 47(24):369, December 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n24/pdf/scientificamerican12091882-
369a.pdf.

Mueller:1882:CWN

[3504] Alexander Mueller. Condensed whey. — a new industry and a new food product. Scientific American, 47(24):369–370, December 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n24/pdf/scientificamerican12091882-369b.pdf.

Anonymous:1882:WFA

[3505] Anonymous. Whitcomb’s fishing apparatus. Scientific American, 47(24):370, December 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n24/pdf/scientificamerican12091882-
370.pdf.

Anonymous:1882:HCF

[3506] Anonymous. Hirsch’s continuous furnace for flattening window glass. Scientific American, 47(24):370–371, December 9, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v47/
n24/pdf/scientificamerican12091882-370a.pdf.

Anonymous:1882:GUE

[3507] Anonymous. Gabert’s universal excavator. Scientific American, 47(24):371, December 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n24/pdf/scientificamerican12091882-
371.pdf.

Anonymous:1882:IPVb

[3508] Anonymous. Improved piston valve. Scientific American, 47 (24):371, December 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n24/pdf/scientificamerican12091882-
371a.pdf.

Anonymous:1882:NEN

[3509] Anonymous. Notable exhibits at the New England Fair. Scientific American, 47(24):372, December 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v47/n24/pdf/scientificamerican12091882-
372.pdf.

Anonymous:1882:MBF

[3510] Anonymous. Machine for breaking fabrics. Scientific American, 47(24):374, December 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n24/pdf/scientificamerican12091882-
374.pdf.

Smedley:1882:HMC

[3511] Victor Smedley. How to make a crab-apple bow. Scientific American, 47(24):375, December 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n24/pdf/scientificamerican12091882-
375.pdf.

Holder:1882:EVS

[3512] C. F. Holder. The economic value of sharks. Scientific American, 47(24):375–376, December 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n24/pdf/scientificamerican12091882-
375a.pdf.

Anonymous:1882:EIbp

[3513] Anonymous. Engineering inventions. Scientific American, 47(24): 378, December 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n24/pdf/scientificamerican12091882-
378.pdf.

Anonymous:1882:MIdd

[3514] Anonymous. Mechanical inventions. Scientific American, 47(24): 378, December 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n24/pdf/scientificamerican12091882-
378a.pdf.

Anonymous:1882:AIz

[3515] Anonymous. Agricultural inventions. Scientific American, 47 (24):378, December 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n24/pdf/scientificamerican12091882-
378b.pdf.

Anonymous:1882:MIde

[3516] Anonymous. Miscellaneous inventions. Scientific American, 47 (24):378, December 9, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n24/pdf/scientificamerican12091882-
378c.pdf.

Anonymous:1882:IRIbe

[3517] Anonymous. Inventions: Recent inventions. Scientific American, 47(25):393, December 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n25/pdf/scientificamerican12161882-
393.pdf.

Anonymous:1882:IIbs

[3518] Anonymous. Index of inventions. Scientific American, 47(25): 394–395, December 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n25/pdf/scientificamerican12161882-
394d.pdf.

Anonymous:1882:NQbp

[3519] Anonymous. Notes and queries. Scientific American, 47(25): 395–396, December 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n25/pdf/scientificamerican12161882-
395.pdf.

Anonymous:1882:BPbr

[3520] Anonymous. Business and personal. Scientific American, 47 (25):396, December 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n25/pdf/scientificamerican12161882-
396a.pdf.

Anonymous:1882:PNPbm

[3521] Anonymous. Patent notes: Patents. Scientific American, 47 (25):396, December 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n25/pdf/scientificamerican12161882-
396b.pdf.

Anonymous:1882:TVc

[3522] Anonymous. The Transit of Venus. Scientific American, 47 (25):384, December 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n25/pdf/scientificamerican12161882-
384.pdf.

Anonymous:1882:GSL

[3523] Anonymous. The great Statue of Liberty. Scientific American, 47(25):384, December 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n25/pdf/scientificamerican12161882-
384a.pdf.

Anonymous:1882:PCC

[3524] Anonymous. The proposed cotton centennial. Scientific American, 47 (25):384–385, December 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n25/pdf/scientificamerican12161882-
384b.pdf.

Anonymous:1882:TVS

[3525] Anonymous. The Transit of Venus as seen at the Seagrave Observatory. Scientific American, 47(25):385, December 16, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v47/
n25/pdf/scientificamerican12161882-385.pdf.

Anonymous:1882:PFR

[3526] Anonymous. A possible field for railway enterprise. Scientific American, 47(25):385, December 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n25/pdf/scientificamerican12161882-
385a.pdf.

Anonymous:1882:IME

[3527] Anonymous. Invention as a means of education. Scientific American, 47(25):385, December 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n25/pdf/scientificamerican12161882-
385b.pdf.

Anonymous:1882:TS

[3528] Anonymous. Tempering steel. Scientific American, 47(25):385, December 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n25/pdf/scientificamerican12161882-385c.pdf.

Anonymous:1882:IMR

[3529] Anonymous. Ice making and refrigeration. — the pictet process. Scientific American, 47(25):386–387, December 16, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v47/
n25/pdf/scientificamerican12161882-386.pdf.

Anonymous:1882:NRV

[3530] Anonymous. Novel road vehicle. Scientific American, 47(25): 387, December 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n25/pdf/scientificamerican12161882-
387.pdf.

Anonymous:1882:OMF

[3531] Anonymous. Oberstadt’s melting furnace with drying chamber. Scientific American, 47(25):387–388, December 16, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v47/
n25/pdf/scientificamerican12161882-387a.pdf.

Anonymous:1882:FEB

[3532] Anonymous. The first electric boat. Scientific American, 47(25): 388–389, December 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n25/pdf/scientificamerican12161882-
388.pdf.

Anonymous:1882:IWH

[3533] Anonymous. Improvement in watch hands. Scientific American, 47(25):390, December 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n25/pdf/scientificamerican12161882-
390.pdf.

Anonymous:1882:ISM

[3534] Anonymous. Improved saw mill. Scientific American, 47(25): 390, December 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n25/pdf/scientificamerican12161882-
390a.pdf.

Anonymous:1882:GHB

[3535] Anonymous. The great harlequin beetle of Cayenne. Scientific American, 47(25):391, December 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n25/pdf/scientificamerican12161882-
391.pdf.

Holder:1882:COA

[3536] C. F. Holder. Care of offspring in animals. Scientific American, 47(25):392, December 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n25/pdf/scientificamerican12161882-
392.pdf.

Anonymous:1882:EIbq

[3537] Anonymous. Engineering inventions. Scientific American, 47(25): 394, December 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n25/pdf/scientificamerican12161882-
394.pdf.

Anonymous:1882:MIdf

[3538] Anonymous. Mechanical inventions. Scientific American, 47(25): 394, December 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n25/pdf/scientificamerican12161882-
394a.pdf.

Anonymous:1882:AIba

[3539] Anonymous. Agricultural inventions. Scientific American, 47 (25):394, December 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n25/pdf/scientificamerican12161882-
394b.pdf.

Anonymous:1882:MIdg

[3540] Anonymous. Miscellaneous inventions. Scientific American, 47 (25):394, December 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n25/pdf/scientificamerican12161882-
394c.pdf.

Anonymous:1882:CRl

[3541] Anonymous. Communications received. Scientific American, 47 (25):396, December 16, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n25/pdf/scientificamerican12161882-
396.pdf.

Anonymous:1882:IIbt

[3542] Anonymous. Index of inventions. Scientific American, 47(26): 410–411, December 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n26/pdf/scientificamerican12231882-
410d.pdf.

Anonymous:1882:NQbq

[3543] Anonymous. Notes and queries. Scientific American, 47(26): 411, December 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n26/pdf/scientificamerican12231882-
411.pdf.

Anonymous:1882:PNPbn

[3544] Anonymous. Patent notes: Patents. Scientific American, 47 (26):411, December 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n26/pdf/scientificamerican12231882-
411a.pdf.

Anonymous:1882:BPbs

[3545] Anonymous. Business and personal. Scientific American, 47 (26):412, December 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n26/pdf/scientificamerican12231882-
412.pdf.

Anonymous:1882:ISSc

[3546] Anonymous. Improved steam street car. Scientific American, 47 (26):399, December 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n26/pdf/scientificamerican12231882-
399.pdf.

Anonymous:1882:NER

[3547] Anonymous. National Exposition of Railway Appliances. Scientific American, 47(26):400, December 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v47/n26/pdf/scientificamerican12231882-
400.pdf.

Anonymous:1882:PPLb

[3548] Anonymous. Proposed Patent Laws for Japan. Scientific American, 47(26):400, December 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n26/pdf/scientificamerican12231882-
400a.pdf.

Anonymous:1882:SCM

[3549] Anonymous. The structure of the cascade mountains. Scientific American, 47(26):400, December 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n26/pdf/scientificamerican12231882-
400b.pdf.

Anonymous:1882:WWI

[3550] Anonymous. Winter work for inventors. Scientific American, 47 (26):401, December 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n26/pdf/scientificamerican12231882-
401.pdf.

Anonymous:1882:SSS

[3551] Anonymous. The steam street supply in New York. Scientific American, 47(26):401, December 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n26/pdf/scientificamerican12231882-
401a.pdf.

Anonymous:1882:ACM

[3552] Anonymous. Ancient counterparts to modern inventions. Scientific American, 47(26):402–403, December 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v47/n26/pdf/scientificamerican12231882-
402.pdf.

Anonymous:1882:PHM

[3553] Anonymous. Portable hay and moss press. Scientific American, 47(26):403, December 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n26/pdf/scientificamerican12231882-
403.pdf.

Anonymous:1882:IVA

[3554] Anonymous. Improved vehicle axle. Scientific American, 47(26): 406, December 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n26/pdf/scientificamerican12231882-
406.pdf.

Anonymous:1882:PP

[3555] Anonymous. The projecting praxinoscope. Scientific American, 47 (26):406–407, December 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n26/pdf/scientificamerican12231882-
406a.pdf.

Anonymous:1882:MAC

[3556] Anonymous. Meteorological apparatus at the Comptoir D’Escompte, Paris. Scientific American, 47(26):407–408, December 23, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v47/
n26/pdf/scientificamerican12231882-407.pdf.

Anonymous:1882:JSb

[3557] Anonymous. Jumping seeds. Scientific American, 47(26):408– 409, December 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n26/pdf/scientificamerican12231882-
408.pdf.

Anonymous:1882:EIbr

[3558] Anonymous. Engineering inventions. Scientific American, 47(26): 410, December 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n26/pdf/scientificamerican12231882-
410.pdf.

Anonymous:1882:MIdh

[3559] Anonymous. Mechanical inventions. Scientific American, 47(26): 410, December 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n26/pdf/scientificamerican12231882-
410a.pdf.

Anonymous:1882:AIbb

[3560] Anonymous. Agricultural inventions. Scientific American, 47 (26):410, December 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n26/pdf/scientificamerican12231882-
410b.pdf.

Anonymous:1882:MIdi

[3561] Anonymous. Miscellaneous inventions. Scientific American, 47 (26):410, December 23, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n26/pdf/scientificamerican12231882-
410c.pdf.

Anonymous:1882:BPbt

[3562] Anonymous. Business and personal. Scientific American, 47(3):43, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n3/pdf/scientificamerican07151882-43.pdf.

Anonymous:1882:BNBr

[3563] Anonymous. Books: New books and publications. Scientific American, 47(3):43, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n3/pdf/scientificamerican07151882-
43a.pdf.

Anonymous:1882:NQbr

[3564] Anonymous. Notes and queries. Scientific American, 47(3):43, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n3/pdf/scientificamerican07151882-43b.pdf.

Anonymous:1882:IIbu

[3565] Anonymous. Index of inventions. Scientific American, 47(3):43–44, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n3/pdf/scientificamerican07151882-43c.pdf.

Anonymous:1882:PNPbo

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n3/pdf/scientificamerican07151882-45.pdf.

Anonymous:1882:IPSb

[3567] Anonymous. Innocent purchasers in the Senate. Scientific American, 47(3):32, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n3/pdf/scientificamerican07151882-
32.pdf.

Anonymous:1882:MPM

[3568] Anonymous. Markings on the Planet Mars. Scientific American, 47(3):32, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n3/pdf/scientificamerican07151882-
32a.pdf.

Anonymous:1882:AC

[3569] Anonymous. American crocodiles. Scientific American, 47(3):33–34, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n3/pdf/scientificamerican07151882-33.pdf.

Anonymous:1882:NFD

[3570] Anonymous. New fertilizer distributer. Scientific American, 47(3):34, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n3/pdf/scientificamerican07151882-34.pdf.

Anonymous:1882:NBS

[3571] Anonymous. Novel boot supporter. Scientific American, 47(3):34, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n3/pdf/scientificamerican07151882-34a.pdf.

Anonymous:1882:IGb

[3572] Anonymous. Improved gate. Scientific American, 47(3):35, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n3/pdf/scientificamerican07151882-35.pdf.

Anonymous:1882:GE

[3573] Anonymous. Gas exhauster. Scientific American, 47(3):35–36, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n3/pdf/scientificamerican07151882-35a.pdf.

Anonymous:1882:DRPj

[3574] Anonymous. Decisions relating to patents. Scientific American, 47(3):37, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n3/pdf/scientificamerican07151882-
37.pdf.

Anonymous:1882:MCGa

[3575] Anonymous. Moore county grit mill stones. Scientific American, 47(3):38, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n3/pdf/scientificamerican07151882-
38.pdf.

Anonymous:1882:ESN

[3576] Anonymous. The eclipse as seen on the Nile. Scientific American, 47(3):39, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n3/pdf/scientificamerican07151882-
39.pdf.

Anonymous:1882:PHR

[3577] Anonymous. Progress of the Hudson River Tunnel. Scientific American, 47(3):39–40, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n3/pdf/scientificamerican07151882-
39a.pdf.

Anonymous:1882:DBP

[3578] Anonymous. Drawing board paper holder. Scientific American, 47(3):40, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n3/pdf/scientificamerican07151882-
40.pdf.

Anonymous:1882:AIbc

[3579] Anonymous. Agricultural inventions. Scientific American, 47(3):41, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n3/pdf/scientificamerican07151882-41.pdf.

Anonymous:1882:MIdj

[3580] Anonymous. Miscellaneous inventions. Scientific American, 47(3):41, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n3/pdf/scientificamerican07151882-41a.pdf.

Anonymous:1882:EIbs

[3581] Anonymous. Engineering invention. Scientific American, 47(3):41, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n3/pdf/scientificamerican07151882-41b.pdf.

Anonymous:1882:MIdk

[3582] Anonymous. Mechanical inventions. Scientific American, 47(3):41, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n3/pdf/scientificamerican07151882-41c.pdf.

Anonymous:1882:EIbt

[3583] Anonymous. Engineering inventions. Scientific American, 47(3):42, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n3/pdf/scientificamerican07151882-42.pdf.

Anonymous:1882:EIbu

[3584] Anonymous. Electrical invention. Scientific American, 47(3):42, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n3/pdf/scientificamerican07151882-42a.pdf.

Anonymous:1882:MIdl

[3585] Anonymous. Mechanical inventions. Scientific American, 47(3):42, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n3/pdf/scientificamerican07151882-42b.pdf.

Anonymous:1882:AIbd

[3586] Anonymous. Agricultural inventions. Scientific American, 47(3):42, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n3/pdf/scientificamerican07151882-42c.pdf.

Anonymous:1882:MIdm

[3587] Anonymous. Miscellaneous inventions. Scientific American, 47(3):42, July 15, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n3/pdf/scientificamerican07151882-42d.pdf.

Anonymous:1882:Cbc

[3588] Anonymous. Correspondence. Scientific American, 47(4):53, July 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n4/pdf/scientificamerican07221882-53.pdf.

Anonymous:1882:IRIbf

[3589] Anonymous. Inventions: Recent inventions. Scientific American, 47(4):57, July 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n4/pdf/scientificamerican07221882-
57.pdf.

Anonymous:1882:BPbu

[3590] Anonymous. Business and personal. Scientific American, 47(4):59, July 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n4/pdf/scientificamerican07221882-59.pdf.

Anonymous:1882:BNBs

[3591] Anonymous. Books: New books and publications. Scientific American, 47(4):59, July 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n4/pdf/scientificamerican07221882-
59a.pdf.

Anonymous:1882:NQbs

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n4/pdf/scientificamerican07221882-59b.pdf.

Anonymous:1882:IIbv

[3593] Anonymous. Index of inventions. Scientific American, 47(4):59–60, July 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n4/pdf/scientificamerican07221882-59c.pdf.

Anonymous:1882:PNPbp

[3594] Anonymous. Patent notes: Patents. Scientific American, 47(4):61, July 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n4/pdf/scientificamerican07221882-61.pdf.

Anonymous:1882:NIC

[3595] Anonymous. New ice cutting machine. Scientific American, 47(4):47, July 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n4/pdf/scientificamerican07221882-47.pdf.

Anonymous:1882:BAb

[3596] Anonymous. The bombardment of Alexandria. Scientific American, 47(4):48, July 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n4/pdf/scientificamerican07221882-
48.pdf.

Anonymous:1882:CAa

[3597] Anonymous. The City of Alexandria. Scientific American, 47(4):48, July 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n4/pdf/scientificamerican07221882-48a.pdf.

Anonymous:1882:FAH

[3598] Anonymous. Firing alleged to be heard 1,000 miles. Scientific American, 47(4):48, July 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n4/pdf/scientificamerican07221882-
48b.pdf.

Anonymous:1882:RDO

[3599] Anonymous. Remarkable developments in oil. Scientific American, 47(4):48–50, July 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n4/pdf/scientificamerican07221882-
48c.pdf.

Anonymous:1882:TSD

[3600] Anonymous. The telephonic systems of Dr. Herz. Scientific American, 47(4):50–52, July 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n4/pdf/scientificamerican07221882-
50.pdf.

Anonymous:1882:NHNd

[3601] Anonymous. Natural history notes. Scientific American, 47(4):53, July 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n4/pdf/scientificamerican07221882-53a.pdf.

Anonymous:1882:BDG

[3602] Anonymous. Bourry’s dynamometric governor. Scientific American, 47(4):54–55, July 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n4/pdf/scientificamerican07221882-
54.pdf.

Anonymous:1882:BRP

[3603] Anonymous. Bass-reliefs by photography. Scientific American, 47(4):55, July 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n4/pdf/scientificamerican07221882-55.pdf.

Holder:1882:LS

[3604] C. F. Holder. A luminous shark. Scientific American, 47(4):55–56, July 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n4/pdf/scientificamerican07221882-55a.pdf.

Anonymous:1882:EIbv

[3605] Anonymous. Engineering invention. Scientific American, 47(4):58, July 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n4/pdf/scientificamerican07221882-58.pdf.

Anonymous:1882:MIdn

[3606] Anonymous. Mechanical inventions. Scientific American, 47(4):58, July 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n4/pdf/scientificamerican07221882-58a.pdf.

Anonymous:1882:MIdo

[3607] Anonymous. Mining invention. Scientific American, 47(4):58, July 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n4/pdf/scientificamerican07221882-58b.pdf.

Anonymous:1882:EIbw

[3608] Anonymous. Electrical invention. Scientific American, 47(4):58, July 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n4/pdf/scientificamerican07221882-58c.pdf.

Anonymous:1882:TIf

[3609] Anonymous. Textile invention. Scientific American, 47(4):58, July 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n4/pdf/scientificamerican07221882-58d.pdf.

Anonymous:1882:AIbe

[3610] Anonymous. Agricultural inventions. Scientific American, 47(4):58, July 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n4/pdf/scientificamerican07221882-58e.pdf.

Anonymous:1882:MIdp

[3611] Anonymous. Miscellaneous inventions. Scientific American, 47(4):58, July 22, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n4/pdf/scientificamerican07221882-58f.pdf.

Anonymous:1882:Cbd

[3612] Anonymous. Correspondence. Scientific American, 47(5):68–69, July 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n5/pdf/scientificamerican07291882-68.pdf.

Anonymous:1882:BPbv

[3613] Anonymous. Business and personal. Scientific American, 47(5):75, July 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n5/pdf/scientificamerican07291882-75.pdf.

Anonymous:1882:NQbt

[3614] Anonymous. Notes and queries. Scientific American, 47(5):75, July 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n5/pdf/scientificamerican07291882-75a.pdf.

Anonymous:1882:IIbw

[3615] Anonymous. Index of inventions. Scientific American, 47(5):75–76, July 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n5/pdf/scientificamerican07291882-75b.pdf.

Anonymous:1882:PNPbq

[3616] Anonymous. Patent notes: Patents. Scientific American, 47(5):76, July 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n5/pdf/scientificamerican07291882-76.pdf.

Anonymous:1882:NIF

[3617] Anonymous. A new and important field for research. Scientific American, 47(5):64, July 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n5/pdf/scientificamerican07291882-
64.pdf.

Anonymous:1882:H

[3618] Anonymous. The hektograph. Scientific American, 47(5):64, July 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n5/pdf/scientificamerican07291882-64a.pdf.

Anonymous:1882:APAb

[3619] Anonymous. Aspects of the planets for August. Scientific American, 47(5):65, July 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n5/pdf/scientificamerican07291882-
65.pdf.

Anonymous:1882:MO

[3620] Anonymous. The manufacture of opium. Scientific American, 47(5):66, July 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n5/pdf/scientificamerican07291882-66.pdf.

Anonymous:1882:CDM

[3621] Anonymous. Cutting drains by machinery. Scientific American, 47(5):66, July 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n5/pdf/scientificamerican07291882-
66a.pdf.

Anonymous:1882:ICB

[3622] Anonymous. Improved cutter bar for reapers and mowers. Scientific American, 47(5):67, July 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n5/pdf/scientificamerican07291882-
67.pdf.

Anonymous:1882:NAF

[3623] Anonymous. Novel attachment to fountains and bottle stoppers. Scientific American, 47(5):67, July 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n5/pdf/scientificamerican07291882-
67a.pdf.

Anonymous:1882:IPd

[3624] Anonymous. Improved pulverizer. Scientific American, 47(5):70, July 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n5/pdf/scientificamerican07291882-70.pdf.

Anonymous:1882:JSR

[3625] Anonymous. John Scott Russell. Scientific American, 47(5):71, July 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n5/pdf/scientificamerican07291882-71.pdf.

Cutter:1882:MEI

[3626] Ephraim Cutter. Microscopical examination of ice. Scientific American, 47(5):71–73, July 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n5/pdf/scientificamerican07291882-
71a.pdf.

Anonymous:1882:MIdq

[3627] Anonymous. Mechanical inventions. Scientific American, 47(5):73, July 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n5/pdf/scientificamerican07291882-73.pdf.

Anonymous:1882:MIdr

[3628] Anonymous. Miscellaneous inventions. Scientific American, 47(5):73, July 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n5/pdf/scientificamerican07291882-73a.pdf.

Anonymous:1882:EIbx

[3629] Anonymous. Engineering inventions. Scientific American, 47(5):74, July 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n5/pdf/scientificamerican07291882-74.pdf.

Anonymous:1882:MIds

[3630] Anonymous. Mechanical inventions. Scientific American, 47(5):74, July 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n5/pdf/scientificamerican07291882-74a.pdf.

Anonymous:1882:AIbf

[3631] Anonymous. Agricultural inventions. Scientific American, 47(5):74, July 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n5/pdf/scientificamerican07291882-74b.pdf.

Anonymous:1882:MIdt

[3632] Anonymous. Miscellaneous inventions. Scientific American, 47(5):74, July 29, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n5/pdf/scientificamerican07291882-74c.pdf.

Anonymous:1882:Cbe

[3633] Anonymous. Correspondence. Scientific American, 47(6):85, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n6/pdf/scientificamerican08051882-85.pdf.

Anonymous:1882:IRIbg

[3634] Anonymous. Inventions: Recent inventions. Scientific American, 47(6):89, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n6/pdf/scientificamerican08051882-
89.pdf.

Anonymous:1882:BPbw

[3635] Anonymous. Business and personal. Scientific American, 47(6):91, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n6/pdf/scientificamerican08051882-91.pdf.

Anonymous:1882:BNBt

[3636] Anonymous. Books: New books and publications. Scientific American, 47(6):91, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n6/pdf/scientificamerican08051882-
91a.pdf.

Anonymous:1882:NQbu

[3637] Anonymous. Notes and queries. Scientific American, 47(6):91, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n6/pdf/scientificamerican08051882-91b.pdf.

Anonymous:1882:IIbx

[3638] Anonymous. Index of inventions. Scientific American, 47(6):91–92, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n6/pdf/scientificamerican08051882-91c.pdf.

Anonymous:1882:PNPbr

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n6/pdf/scientificamerican08051882-92.pdf.

Anonymous:1882:DS

[3640] Anonymous. The decline of schooling. Scientific American, 47(6):80, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n6/pdf/scientificamerican08051882-80.pdf.

Anonymous:1882:MIBba

[3641] Anonymous. More innocent buyers needing protection. Scientific American, 47(6):80, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n6/pdf/scientificamerican08051882-
80a.pdf.

Anonymous:1882:LSC

[3642] Anonymous. A labor storm-center. Scientific American, 47(6):80, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n6/pdf/scientificamerican08051882-80b.pdf.

Anonymous:1882:FRE

[3643] Anonymous. Fire risks with electric lamps. Scientific American, 47(6):80–81, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n6/pdf/scientificamerican08051882-
80c.pdf.

Anonymous:1882:AMO

[3644] Anonymous. The absorption of metallic oxides by plants. Scientific American, 47(6):81, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n6/pdf/scientificamerican08051882-
81.pdf.

Anonymous:1882:AT

[3645] Anonymous. Accuracy in telegraphing. Scientific American, 47(6):81, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n6/pdf/scientificamerican08051882-81a.pdf.

Anonymous:1882:WBD

[3646] Anonymous. Why beef is dear. Scientific American, 47(6):81, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n6/pdf/scientificamerican08051882-81b.pdf.

Anonymous:1882:NSD

[3647] Anonymous. A new Sun dial. Scientific American, 47(6):82, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n6/pdf/scientificamerican08051882-82.pdf.

Anonymous:1882:DUP

[3648] Anonymous. Detachable umbrella and parasol covers. Scientific American, 47(6):82, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n6/pdf/scientificamerican08051882-
82a.pdf.

Anonymous:1882:TMP

[3649] Anonymous. Transmission of motive power by rarefied air. Scientific American, 47(6):83, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n6/pdf/scientificamerican08051882-
83.pdf.

Anonymous:1882:PER

[3650] Anonymous. Progress of the East River Bridge. Scientific American, 47(6):84, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n6/pdf/scientificamerican08051882-
84.pdf.

Anonymous:1882:HG

[3651] Anonymous. Henri Giffard. Scientific American, 47(6):84, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n6/pdf/scientificamerican08051882-84a.pdf.

Anonymous:1882:RCU

[3652] Anonymous. Rules in the cutting of uppers. Scientific American, 47(6):85, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n6/pdf/scientificamerican08051882-
85a.pdf.

Anonymous:1882:SA

[3653] Anonymous. The steamship Austral. Scientific American, 47(6):86, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n6/pdf/scientificamerican08051882-86.pdf.

Anonymous:1882:DTB

[3654] Anonymous. The direct transfusion of blood. Scientific American, 47(6):86–87, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n6/pdf/scientificamerican08051882-
86a.pdf.

Holder:1882:AT

[3655] C. F. Holder. The American tunny. Scientific American, 47(6):87–88, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n6/pdf/scientificamerican08051882-87.pdf.

Anonymous:1882:NFB

[3656] Anonymous. New form of bridge supports. Scientific American, 47(6):88, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n6/pdf/scientificamerican08051882-
88.pdf.

Anonymous:1882:EIby

[3657] Anonymous. Engineering inventions. Scientific American, 47(6):90, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n6/pdf/scientificamerican08051882-90.pdf.

Anonymous:1882:MIdu

[3658] Anonymous. Mechanical inventions. Scientific American, 47(6):90, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n6/pdf/scientificamerican08051882-90a.pdf.

Anonymous:1882:AIbg

[3659] Anonymous. Agricultural inventions. Scientific American, 47(6):90, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n6/pdf/scientificamerican08051882-90b.pdf.

Anonymous:1882:MIdv

[3660] Anonymous. Miscellaneous inventions. Scientific American, 47(6):90, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n6/pdf/scientificamerican08051882-90c.pdf.

Anonymous:1882:TIg

[3661] Anonymous. Textile invention. Scientific American, 47(6):90, August 5, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n6/pdf/scientificamerican08051882-90d.pdf.

Anonymous:1882:BPbx

[3662] Anonymous. Business and personal. Scientific American, 47(7):107, August 12, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n7/pdf/scientificamerican08121882-107.pdf.

Anonymous:1882:BNBu

[3663] Anonymous. Books: New books and publications. Scientific American, 47(7):107, August 12, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n7/pdf/scientificamerican08121882-
107a.pdf.

Anonymous:1882:NQbv

[3664] Anonymous. Notes and queries. Scientific American, 47(7):107, August 12, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n7/pdf/scientificamerican08121882-107b.pdf.

Anonymous:1882:IIby

[3665] Anonymous. Index of inventions. Scientific American, 47(7): 107–108, August 12, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n7/pdf/scientificamerican08121882-
107c.pdf.

Anonymous:1882:PNPbs

[3666] Anonymous. Patent notes: Patents. Scientific American, 47(7):108, August 12, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n7/pdf/scientificamerican08121882-108a.pdf.

Anonymous:1882:RWW

[3667] Anonymous. Rochester water works. Scientific American, 47(7):95, August 12, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n7/pdf/scientificamerican08121882-95.pdf.

Anonymous:1882:MD

[3668] Anonymous. Manual dexterity. Scientific American, 47(7):96, August 12, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n7/pdf/scientificamerican08121882-96.pdf.

Anonymous:1882:EHG

[3669] Anonymous. Early history of gas lighting. Scientific American, 47(7):96–98, August 12, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n7/pdf/scientificamerican08121882-
96a.pdf.

Anonymous:1882:BL

[3670] Anonymous. Blasting with lime. Scientific American, 47(7):98, August 12, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n7/pdf/scientificamerican08121882-98.pdf.

Anonymous:1882:NNLb

[3671] Anonymous. New nut lock. Scientific American, 47(7):98, August 12, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n7/pdf/scientificamerican08121882-98a.pdf.

Anonymous:1882:NV

[3672] Anonymous. New ventilator. Scientific American, 47(7):98, August 12, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n7/pdf/scientificamerican08121882-98b.pdf.

Anonymous:1882:FEW

[3673] Anonymous. French and English war ships compared. Scientific American, 47(7):99, August 12, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n7/pdf/scientificamerican08121882-
99.pdf.

Anonymous:1882:MCGb

[3674] Anonymous. Machine for coloring and grounding paper for paper hangings etc. Scientific American, 47(7):99, August 12, 1882. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v47/
n7/pdf/scientificamerican08121882-99a.pdf.

Anonymous:1882:CCb

[3675] Anonymous. A cheap cottage. Scientific American, 47(7):100–101, August 12, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n7/pdf/scientificamerican08121882-100.pdf.

Anonymous:1882:CPb

[3676] Anonymous. Contractors’ plant. Scientific American, 47(7):102, August 12, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n7/pdf/scientificamerican08121882-102.pdf.

Anonymous:1882:NSZ

[3677] Anonymous. A new species of zebra. Scientific American, 47 (7):103–104, August 12, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n7/pdf/scientificamerican08121882-
103.pdf.

Anonymous:1882:RId

[3678] Anonymous. Recent inventions. Scientific American, 47(7):105, August 12, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n7/pdf/scientificamerican08121882-105.pdf.

Anonymous:1882:EIbz (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n7/pdf/scientificamerican08121882-
106.pdf.

Anonymous:1882:MIdw

[3680] Anonymous. Mechanical inventions. Scientific American, 47(7):106, August 12, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n7/pdf/scientificamerican08121882-106a.pdf.

Anonymous:1882:AIbh

[3681] Anonymous. Agricultural inventions. Scientific American, 47(7): 106, August 12, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n7/pdf/scientificamerican08121882-
106b.pdf.

Anonymous:1882:MIdx

[3682] Anonymous. Miscellaneous inventions. Scientific American, 47 (7):106, August 12, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n7/pdf/scientificamerican08121882-
106c.pdf.

Anonymous:1882:EPI

[3683] Anonymous. English patents issued to Americans. Scientific American, 47(7):108, August 12, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n7/pdf/scientificamerican08121882-
108.pdf.

Anonymous:1882:Cbf

[3684] Anonymous. Correspondence. Scientific American, 47(8):117, August 19, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n8/pdf/scientificamerican08191882-117.pdf.

Anonymous:1882:IRIbh

[3685] Anonymous. Inventions: Recent inventions. Scientific American, 47(8):121, August 19, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n8/pdf/scientificamerican08191882-
121.pdf.

Anonymous:1882:BPby

[3686] Anonymous. Business and personal. Scientific American, 47(8):123, August 19, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n8/pdf/scientificamerican08191882-123.pdf.

Anonymous:1882:NQbw

[3687] Anonymous. Notes and queries. Scientific American, 47(8):123, August 19, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n8/pdf/scientificamerican08191882-123a.pdf.

Anonymous:1882:IIbz

[3688] Anonymous. Index of inventions. Scientific American, 47(8): 123–124, August 19, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n8/pdf/scientificamerican08191882-
123c.pdf.

Anonymous:1882:PNPbt

[3689] Anonymous. Patent notes: Patents. Scientific American, 47(8):124, August 19, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n8/pdf/scientificamerican08191882-124.pdf.

Anonymous:1882:LME

[3690] Anonymous. The Levitt–Muller electric lighting system and multiple circuit dynamo. Scientific American, 47(8):111, August 19, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n8/pdf/scientificamerican08191882-111.pdf.

Anonymous:1882:WE

[3691] Anonymous. The War in Egypt. Scientific American, 47(8):112, August 19, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n8/pdf/scientificamerican08191882-112.pdf.

Anonymous:1882:RA

[3692] Anonymous. A recent Aurora. Scientific American, 47(8):112, August 19, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n8/pdf/scientificamerican08191882-112a.pdf.

Anonymous:1882:FCC

[3693] Anonymous. Freight car couplers. Scientific American, 47(8): 112–113, August 19, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n8/pdf/scientificamerican08191882-
112b.pdf.

Anonymous:1882:NDS

[3694] Anonymous. Novel door securer. Scientific American, 47(8):115, August 19, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n8/pdf/scientificamerican08191882-115.pdf.

Anonymous:1882:IFC

[3695] Anonymous. Improved freight car. Scientific American, 47(8):115, August 19, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n8/pdf/scientificamerican08191882-115a.pdf.

Anonymous:1882:FPC

[3696] Anonymous. Frey’s plumber’s coupling. Scientific American, 47 (8):115–116, August 19, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n8/pdf/scientificamerican08191882-
115b.pdf.

Anonymous:1882:IGc

[3697] Anonymous. Improved gate. Scientific American, 47(8):118, August 19, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n8/pdf/scientificamerican08191882-118.pdf.

Anonymous:1882:SSH

[3698] Anonymous. Spring steam hammers. Scientific American, 47 (8):118, August 19, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n8/pdf/scientificamerican08191882-
118a.pdf.

Beard:1882:CNC

[3699] Daniel C. Beard. Compound nest of the C. Palustris and the A. Phoeniceus. Scientific American, 47(8):119–120, August 19, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n8/pdf/scientificamerican08191882-119.pdf.

Anonymous:1882:EIca

[3700] Anonymous. Engineering inventions. Scientific American, 47(8): 122, August 19, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n8/pdf/scientificamerican08191882-
122.pdf.

Anonymous:1882:MIdy

[3701] Anonymous. Mechanical inventions. Scientific American, 47(8):122, August 19, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n8/pdf/scientificamerican08191882-122a.pdf.

Anonymous:1882:EIcb

[3702] Anonymous. Electrical invention. Scientific American, 47(8):122, August 19, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n8/pdf/scientificamerican08191882-122b.pdf.

Anonymous:1882:AIbi

[3703] Anonymous. Agricultural inventions. Scientific American, 47(8): 122, August 19, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n8/pdf/scientificamerican08191882-
122c.pdf.

Anonymous:1882:MIdz

[3704] Anonymous. Miscellaneous inventions. Scientific American, 47 (8):122, August 19, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n8/pdf/scientificamerican08191882-
122d.pdf.

Anonymous:1882:CRm (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n8/pdf/scientificamerican08191882-
123b.pdf.

Anonymous:1882:IRIbi

[3706] Anonymous. Inventions: Recent inventions. Scientific American, 47(9):137, August 26, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n9/pdf/scientificamerican08261882-
137.pdf.

Anonymous:1882:BPbz

[3707] Anonymous. Business and personal. Scientific American, 47(9):139, August 26, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n9/pdf/scientificamerican08261882-139.pdf.

Anonymous:1882:BNBv

[3708] Anonymous. Books: New books and publications. Scientific American, 47(9):139, August 26, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n9/pdf/scientificamerican08261882-
139a.pdf.

Anonymous:1882:NQbx

[3709] Anonymous. Notes and queries. Scientific American, 47(9):139– 140, August 26, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n9/pdf/scientificamerican08261882-
139b.pdf.

Anonymous:1882:IIca

[3710] Anonymous. Index of inventions. Scientific American, 47(9):140, August 26, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n9/pdf/scientificamerican08261882-140a.pdf.

Anonymous:1882:PNPbu

[3711] Anonymous. Patent notes: Patents. Scientific American, 47(9):141, August 26, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n9/pdf/scientificamerican08261882-141.pdf.

Anonymous:1882:EEL

[3712] Anonymous. The Edison electric lighting station. Scientific American, 47(9):127, August 26, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n9/pdf/scientificamerican08261882-
127.pdf.

Anonymous:1882:PSL

[3713] Anonymous. Postage stamp losses. Scientific American, 47(9):128, August 26, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n9/pdf/scientificamerican08261882-128.pdf.

Anonymous:1882:DME

[3714] Anonymous. The Denver Mining Exposition. Scientific American, 47(9):128–129, August 26, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n9/pdf/scientificamerican08261882-
128a.pdf.

Anonymous:1882:PE

[3715] Anonymous. Prison electricity. Scientific American, 47(9):129, August 26, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n9/pdf/scientificamerican08261882-129.pdf.

Anonymous:1882:PSIb

[3716] Anonymous. Pictet’s small ice-machine. Scientific American, 47 (9):131, August 26, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n9/pdf/scientificamerican08261882-
131.pdf.

Anonymous:1882:HPM

[3717] Anonymous. Hydraulic punching machine. Scientific American, 47 (9):131–132, August 26, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n9/pdf/scientificamerican08261882-
131a.pdf.

Anonymous:1882:IBP

[3718] Anonymous. Improvement in bed-pans. Scientific American, 47 (9):132, August 26, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n9/pdf/scientificamerican08261882-
132.pdf.

Anonymous:1882:RDRc

[3719] Anonymous. Recent decisions relating to patents. Scientific American, 47(9):133, August 26, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n9/pdf/scientificamerican08261882-
133.pdf.

Anonymous:1882:ERW

[3720] Anonymous. The electric railway at West End, near Berlin. Scientific American, 47(9):134, August 26, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n9/pdf/scientificamerican08261882-
134.pdf.

Anonymous:1882:ABL

[3721] Anonymous. Ancient breech-loader. Scientific American, 47(9):135, August 26, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n9/pdf/scientificamerican08261882-135.pdf.

Anonymous:1882:HK

[3722] Anonymous. The herring king. Scientific American, 47(9):135–136, August 26, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n9/pdf/scientificamerican08261882-135a.pdf.

Anonymous:1882:WCS

[3723] Anonymous. The Wells Comet, as seen in Sydney. Scientific American, 47(9):136, August 26, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n9/pdf/scientificamerican08261882-
136.pdf.

Anonymous:1882:EIcc

[3724] Anonymous. Engineering inventions. Scientific American, 47(9): 138, August 26, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n9/pdf/scientificamerican08261882-
138.pdf.

Anonymous:1882:MIea

[3725] Anonymous. Mechanical inventions. Scientific American, 47(9):138, August 26, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v47/n9/pdf/scientificamerican08261882-138a.pdf.

Anonymous:1882:AIbj

[3726] Anonymous. Agricultural inventions. Scientific American, 47(9): 138, August 26, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n9/pdf/scientificamerican08261882-
138b.pdf.

Anonymous:1882:MIeb

[3727] Anonymous. Miscellaneous inventions. Scientific American, 47 (9):138, August 26, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n9/pdf/scientificamerican08261882-
138c.pdf.

Anonymous:1882:CRn

[3728] Anonymous. Communications received. Scientific American, 47 (9):140, August 26, 1882. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v47/n9/pdf/scientificamerican08261882-
140.pdf.

Anonymous:1883:IRIa

[3729] Anonymous. Inventions: Recent inventions. Scientific American, 48(1):7, January 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n1/pdf/scientificamerican01061883-
7.pdf.

Anonymous:1883:Ca

[3730] Anonymous. Correspondence. Scientific American, 48(1):7, January 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n1/pdf/scientificamerican01061883-7a.pdf.

Anonymous:1883:BPa

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n1/pdf/scientificamerican01061883-11.pdf.

Anonymous:1883:NQa

[3732] Anonymous. Notes and queries. Scientific American, 48(1):11, January 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n1/pdf/scientificamerican01061883-11a.pdf.

Anonymous:1883:PNPa

[3733] Anonymous. Patent notes: Patents. Scientific American, 48(1):12, January 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n1/pdf/scientificamerican01061883-12.pdf.

Anonymous:1883:GDE

[3734] Anonymous. Gordon’s dynamo electric machine. Scientific American, 48(1):1, January 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n1/pdf/scientificamerican01061883-
1.pdf.

Anonymous:1883:DBW

[3735] Anonymous. The death and burial of Woehler. Scientific American, 48(1):2, January 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n1/pdf/scientificamerican01061883-
2.pdf.

Anonymous:1883:RT

[3736] Anonymous. Railway transportation. Scientific American, 48(1):2, January 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n1/pdf/scientificamerican01061883-2a.pdf.

Anonymous:1883:DFA

[3737] Anonymous. Dangerous funeral appliances. Scientific American, 48(1):2, January 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n1/pdf/scientificamerican01061883-
2b.pdf.

Anonymous:1883:MLa

[3738] Anonymous. Machinery and labor. Scientific American, 48(1):2–3, January 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n1/pdf/scientificamerican01061883-2c.pdf.

Riley:1883:EPI

[3739] C. V. Riley. Emulsions of petroleum as insecticides. Scientific American, 48(1):3, January 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n1/pdf/scientificamerican01061883-
3.pdf.

Anonymous:1883:RMC

[3740] Anonymous. Rotary mould for casting pipes. Scientific American, 48(1):4, January 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n1/pdf/scientificamerican01061883-
4.pdf.

Anonymous:1883:IIM

[3741] Anonymous. Improved ice machine. Scientific American, 48(1):4, January 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n1/pdf/scientificamerican01061883-4a.pdf.

Anonymous:1883:IGB

[3742] Anonymous. Improved galvanic battery. Scientific American, 48(1):5, January 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n1/pdf/scientificamerican01061883-5.pdf.

Anonymous:1883:SFM

[3743] Anonymous. Saw filing machine. Scientific American, 48(1):5, January 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n1/pdf/scientificamerican01061883-5a.pdf.

Anonymous:1883:IWT

[3744] Anonymous. Improved wheel tire. Scientific American, 48(1):6, January 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n1/pdf/scientificamerican01061883-6.pdf.

Anonymous:1883:AAF

[3745] Anonymous. Allen’s apparatus for feeding blast and other furnaces. Scientific American, 48(1):8, January 6, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v48/
n1/pdf/scientificamerican01061883-8.pdf.

Anonymous:1883:NTE

[3746] Anonymous. New traction engine. Scientific American, 48(1):8–9, January 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n1/pdf/scientificamerican01061883-8a.pdf.

Anonymous:1883:DTL

[3747] Anonymous. Duck with three legs. Scientific American, 48(1):9, January 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n1/pdf/scientificamerican01061883-9.pdf.

Anonymous:1883:EIP

[3748] Anonymous. The elephant by instantaneous photography. Scientific American, 48(1):9, January 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n1/pdf/scientificamerican01061883-
9a.pdf.

Anonymous:1883:EIa

[3749] Anonymous. Engineering inventions. Scientific American, 48(1):10, January 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n1/pdf/scientificamerican01061883-10.pdf.

Anonymous:1883:MIa

[3750] Anonymous. Mechanical inventions. Scientific American, 48(1):10, January 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n1/pdf/scientificamerican01061883-10a.pdf.

Anonymous:1883:AIa

[3751] Anonymous. Agricultural inventions. Scientific American, 48(1):10, January 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n1/pdf/scientificamerican01061883-10b.pdf.

Anonymous:1883:MIb

[3752] Anonymous. Miscellaneous inventions. Scientific American, 48(1):10, January 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n1/pdf/scientificamerican01061883-10c.pdf.

Anonymous:1883:Cb

[3753] Anonymous. Correspondence. Scientific American, 48(10):149–150, March 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n10/pdf/scientificamerican03101883-149.pdf.

Anonymous:1883:RI

[3754] Anonymous. Recent invention. Scientific American, 48(10):153, March 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n10/pdf/scientificamerican03101883-153.pdf.

Anonymous:1883:BPb

[3755] Anonymous. Business and personal. Scientific American, 48 (10):154–155, March 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n10/pdf/scientificamerican03101883-
154c.pdf.

Anonymous:1883:NQb

[3756] Anonymous. Notes and queries. Scientific American, 48(10):155, March 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n10/pdf/scientificamerican03101883-155.pdf.

Anonymous:1883:IIa

[3757] Anonymous. Index of inventions. Scientific American, 48(10):155–156, March 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n10/pdf/scientificamerican03101883-155a.pdf.

Anonymous:1883:PFG

[3758] Anonymous. Paint for floors, a good old miller. Scientific American, 48(10):143, March 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n10/pdf/scientificamerican03101883-
143.pdf.

Anonymous:1883:PBE

[3759] Anonymous. Propulsion of balloons by electricity. Scientific American, 48(10):143–147, March 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n10/pdf/scientificamerican03101883-
143a.pdf.

Anonymous:1883:CBO

[3760] Anonymous. Compressed bran. — one thousand dollars reward for a new invention. Scientific American, 48(10):144, March 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n10/pdf/scientificamerican03101883-144.pdf.

Anonymous:1883:SE

[3761] Anonymous. Schooled but not educated. Scientific American, 48 (10):144–145, March 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n10/pdf/scientificamerican03101883-
144a.pdf.

Anonymous:1883:SHI

[3762] Anonymous. Silk and how it is dyed. Scientific American, 48(10):145, March 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n10/pdf/scientificamerican03101883-145.pdf.

Anonymous:1883:TAE

[3763] Anonymous. Thomson’s air extractor. Scientific American, 48(10):146, March 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n10/pdf/scientificamerican03101883-146.pdf.

Anonymous:1883:IFCa

[3764] Anonymous. Improved foundry cupola. Scientific American, 48(10):146, March 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n10/pdf/scientificamerican03101883-146a.pdf.

Anonymous:1883:IG

[3765] Anonymous. The isometric governor. Scientific American, 48 (10):147–148, March 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n10/pdf/scientificamerican03101883-
147.pdf.

Anonymous:1883:PEa

[3766] Anonymous. The “Peerless” engine. Scientific American, 48(10):150, March 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n10/pdf/scientificamerican03101883-150.pdf.

Anonymous:1883:TMS

[3767] Anonymous. Toy money safe. Scientific American, 48(10):150, March 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n10/pdf/scientificamerican03101883-150a.pdf.

Anonymous:1883:SSSa

[3768] Anonymous. Side show science. Scientific American, 48(10):151–152, March 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n10/pdf/scientificamerican03101883-151.pdf.

Anonymous:1883:MIc

[3769] Anonymous. Mechanical inventions. Scientific American, 48(10):154, March 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n10/pdf/scientificamerican03101883-154.pdf.

Anonymous:1883:AIb

[3770] Anonymous. Agricultural inventions. Scientific American, 48(10):154, March 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n10/pdf/scientificamerican03101883-154a.pdf.

Anonymous:1883:MId

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n10/pdf/scientificamerican03101883-154b.pdf.

Anonymous:1883:Cc

[3772] Anonymous. Correspondence. Scientific American, 48(11):165, March 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n11/pdf/scientificamerican03171883-165.pdf.

Anonymous:1883:IRIb

[3773] Anonymous. Inventions: Recent inventions. Scientific American, 48(11):169, March 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n11/pdf/scientificamerican03171883-
169.pdf.

Anonymous:1883:BPc

[3774] Anonymous. Business and personal. Scientific American, 48(11):170, March 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n11/pdf/scientificamerican03171883-170d.pdf.

Anonymous:1883:NQc

[3775] Anonymous. Notes and queries. Scientific American, 48(11):171, March 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n11/pdf/scientificamerican03171883-171.pdf.

Anonymous:1883:IIb

[3776] Anonymous. Index of inventions. Scientific American, 48(11):171–172, March 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n11/pdf/scientificamerican03171883-171b.pdf.

Anonymous:1883:SPM

[3777] Anonymous. Sale of patented machines by sheriff. Scientific American, 48(11):160, March 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n11/pdf/scientificamerican03171883-
160.pdf.

Anonymous:1883:ET

[3778] Anonymous. Expert testimony. Scientific American, 48(11):160, March 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n11/pdf/scientificamerican03171883-160a.pdf.

Anonymous:1883:RSJ

[3779] Anonymous. The Red Spot on Jupiter. Scientific American, 48(11):160, March 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n11/pdf/scientificamerican03171883-160b.pdf.

Anonymous:1883:SSFa

[3780] Anonymous. Safety at sea in a fog. Scientific American, 48(11):160, March 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n11/pdf/scientificamerican03171883-160c.pdf.

Anonymous:1883:GW

[3781] Anonymous. Is glucose wholesome? Scientific American, 48(11):161, March 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n11/pdf/scientificamerican03171883-161.pdf.

Anonymous:1883:CPL

[3782] Anonymous. Cheap postage at last, and postal notes in addition. Scientific American, 48(11):161, March 17, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v48/
n11/pdf/scientificamerican03171883-161a.pdf.

Anonymous:1883:LDT

[3783] Anonymous. Long distance telephoning. Scientific American, 48 (11):161, March 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n11/pdf/scientificamerican03171883-
161b.pdf.

Anonymous:1883:CRCa

[3784] Anonymous. Curiosities of the railway census. Scientific American, 48(11):161, March 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n11/pdf/scientificamerican03171883-
161c.pdf.

Anonymous:1883:IR

[3785] Anonymous. Improved refrigerator. Scientific American, 48(11):162, March 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n11/pdf/scientificamerican03171883-162.pdf.

Anonymous:1883:BRL

[3786] Anonymous. The De Bergen rail lifter. Scientific American, 48(11):162, March 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n11/pdf/scientificamerican03171883-162a.pdf.

Anonymous:1883:MES

[3787] Anonymous. The manufacture of engines, saw mills, and grain separators. Scientific American, 48(11):163–164, March 17, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v48/
n11/pdf/scientificamerican03171883-163.pdf.

Anonymous:1883:ICCa

[3788] Anonymous. Improved car coupling. Scientific American, 48(11):164, March 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n11/pdf/scientificamerican03171883-164.pdf.

Anonymous:1883:NHNa

[3789] Anonymous. Natural history notes. Scientific American, 48(11):165, March 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n11/pdf/scientificamerican03171883-165a.pdf.

Warnerke:1883:NL

[3790] Leon Warnerke. A new light. Scientific American, 48(11):166, March 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n11/pdf/scientificamerican03171883-166.pdf.

Anonymous:1883:PHB

[3791] Anonymous. Post hole boring machine. Scientific American, 48(11):166, March 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n11/pdf/scientificamerican03171883-166a.pdf.

Anonymous:1883:NDS

[3792] Anonymous. A new deep sea fish. — (Eurypharyna Pelecanoides). Scientific American, 48(11):167–168, March 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v48/n11/pdf/scientificamerican03171883-
167.pdf.

Anonymous:1883:EIb

[3793] Anonymous. Engineering inventions. Scientific American, 48(11):170, March 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n11/pdf/scientificamerican03171883-170.pdf.

Anonymous:1883:MIe

[3794] Anonymous. Mechanical inventions. Scientific American, 48(11):170, March 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n11/pdf/scientificamerican03171883-170a.pdf.

Anonymous:1883:AIc

[3795] Anonymous. Agricultural inventions. Scientific American, 48(11):170, March 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n11/pdf/scientificamerican03171883-170b.pdf.

Anonymous:1883:MIf

[3796] Anonymous. Miscellaneous inventions. Scientific American, 48(11):170, March 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n11/pdf/scientificamerican03171883-170c.pdf.

Anonymous:1883:CRa

[3797] Anonymous. Communications received. Scientific American, 48 (11):171, March 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n11/pdf/scientificamerican03171883-
171a.pdf.

Anonymous:1883:Cd

[3798] Anonymous. Correspondence. Scientific American, 48(12):181, March 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n12/pdf/scientificamerican03241883-181.pdf.

Anonymous:1883:IRIc

[3799] Anonymous. Inventions: Recent inventions. Scientific American, 48(12):185, March 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n12/pdf/scientificamerican03241883-
185.pdf.

Anonymous:1883:BPd

[3800] Anonymous. Business and personal. Scientific American, 48(12):187, March 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n12/pdf/scientificamerican03241883-187.pdf.

Anonymous:1883:NQd

[3801] Anonymous. Notes and queries. Scientific American, 48(12):187, March 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n12/pdf/scientificamerican03241883-187a.pdf.

Anonymous:1883:IIc

[3802] Anonymous. Index of inventions. Scientific American, 48(12):187–188, March 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n12/pdf/scientificamerican03241883-187b.pdf.

Anonymous:1883:AINa

[3803] Anonymous. American industries. — no. 87. Scientific American, 48(12):175–176, March 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n12/pdf/scientificamerican03241883-
175.pdf.

Anonymous:1883:SHF

[3804] Anonymous. Six hundred and fifty miles by telephone. Scientific American, 48(12):176–177, March 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n12/pdf/scientificamerican03241883-
176.pdf.

Anonymous:1883:ESE

[3805] Anonymous. Early steam engines and boilers. Scientific American, 48(12):177, March 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n12/pdf/scientificamerican03241883-
177.pdf.

Anonymous:1883:LT

[3806] Anonymous. The Leffel turbine. Scientific American, 48(12):177, March 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n12/pdf/scientificamerican03241883-177a.pdf.

Anonymous:1883:SPA

[3807] Anonymous. Sanitary precautions after floods. Scientific American, 48(12):177, March 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n12/pdf/scientificamerican03241883-
177b.pdf.

Anonymous:1883:IRE

[3808] Anonymous. Improved railway elevator. Scientific American, 48 (12):178, March 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n12/pdf/scientificamerican03241883-
178.pdf.

Anonymous:1883:NFW

[3809] Anonymous. New feed water heater and lime extractor. Scientific American, 48(12):179, March 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n12/pdf/scientificamerican03241883-
179.pdf.

Anonymous:1883:IFPa

[3810] Anonymous. Improved fence post. Scientific American, 48(12):179–180, March 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n12/pdf/scientificamerican03241883-179a.pdf.

Anonymous:1883:AASa

[3811] Anonymous. Abacus attachment for school slates. Scientific American, 48(12):180, March 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n12/pdf/scientificamerican03241883-
180.pdf.

Anonymous:1883:IVEa

[3812] Anonymous. Improved vertical engine. Scientific American, 48 (12):182–183, March 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n12/pdf/scientificamerican03241883-
182.pdf.

Anonymous:1883:IRC

[3813] Anonymous. Improved railroad crane. Scientific American, 48 (12):183–184, March 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n12/pdf/scientificamerican03241883-
183.pdf.

Anonymous:1883:MIg

[3814] Anonymous. Mechanical inventions. Scientific American, 48(12):186, March 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n12/pdf/scientificamerican03241883-186.pdf.

Anonymous:1883:AId

[3815] Anonymous. Agricultural inventions. Scientific American, 48(12):186, March 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n12/pdf/scientificamerican03241883-186a.pdf.

Anonymous:1883:MIh

[3816] Anonymous. Miscellaneous inventions. Scientific American, 48(12):186, March 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n12/pdf/scientificamerican03241883-186b.pdf.

Anonymous:1883:Ce

[3817] Anonymous. Correspondence. Scientific American, 48(13):197–198, March 31, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n13/pdf/scientificamerican03311883-197.pdf.

Anonymous:1883:IRId

[3818] Anonymous. Inventions: Recent inventions. Scientific American, 48(13):201, March 31, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n13/pdf/scientificamerican03311883-
201.pdf.

Anonymous:1883:BPe

[3819] Anonymous. Business and personal. Scientific American, 48(13):203, March 31, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n13/pdf/scientificamerican03311883-203.pdf.

Anonymous:1883:NQe

[3820] Anonymous. Notes and queries. Scientific American, 48(13):203, March 31, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n13/pdf/scientificamerican03311883-203a.pdf.

Anonymous:1883:IId

[3821] Anonymous. Index of inventions. Scientific American, 48(13):206–207, March 31, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n13/pdf/scientificamerican03311883-203b.pdf.

Anonymous:1883:IFa

[3822] Anonymous. Improved filters. Scientific American, 48(13):191–196, March 31, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n13/pdf/scientificamerican03311883-191.pdf.

Anonymous:1883:EMa

[3823] Anonymous. Education for mechanics. Scientific American, 48(13):192, March 31, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n13/pdf/scientificamerican03311883-192.pdf.

Anonymous:1883:AWa

[3824] Anonymous. Analysis of water. Scientific American, 48(13):192, March 31, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n13/pdf/scientificamerican03311883-192a.pdf.

Anonymous:1883:SWS

[3825] Anonymous. Sites for water supplies. Scientific American, 48(13):192, March 31, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n13/pdf/scientificamerican03311883-192b.pdf.

Anonymous:1883:APAa

[3826] Anonymous. Aspects of the planets for April. Scientific American, 48(13):193, March 31, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n13/pdf/scientificamerican03311883-
193.pdf.

Anonymous:1883:SACa

[3827] Anonymous. Some answers to correspondents. Scientific American, 48(13):193–194, March 31, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n13/pdf/scientificamerican03311883-
193a.pdf.

Anonymous:1883:NFA

[3828] Anonymous. New fanning apparatus. Scientific American, 48(13):194, March 31, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n13/pdf/scientificamerican03311883-194.pdf.

Anonymous:1883:NPG

[3829] Anonymous. New pile for Galvano Cautery. Scientific American, 48(13):194, March 31, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n13/pdf/scientificamerican03311883-
194a.pdf.

Anonymous:1883:BBS

[3830] Anonymous. Buggy bow spring. Scientific American, 48(13):198, March 31, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n13/pdf/scientificamerican03311883-198.pdf.

Anonymous:1883:IHb

[3831] Anonymous. Improved hat holder. Scientific American, 48(13):198, March 31, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n13/pdf/scientificamerican03311883-198a.pdf.

Anonymous:1883:NG

[3832] Anonymous. The Nordenfelt gun. Scientific American, 48(13):198, March 31, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n13/pdf/scientificamerican03311883-198b.pdf.

Anonymous:1883:AF

[3833] Anonymous. The archer fish. Scientific American, 48(13):199–200, March 31, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n13/pdf/scientificamerican03311883-199.pdf.

Anonymous:1883:EIc

[3834] Anonymous. Engineering inventions. Scientific American, 48(13):202, March 31, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n13/pdf/scientificamerican03311883-202.pdf.

Anonymous:1883:MIi

[3835] Anonymous. Mechanical inventions. Scientific American, 48(13):202, March 31, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n13/pdf/scientificamerican03311883-202a.pdf.

Anonymous:1883:AIe

[3836] Anonymous. Agricultural inventions. Scientific American, 48(13):202, March 31, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n13/pdf/scientificamerican03311883-202b.pdf.

Anonymous:1883:MIj

[3837] Anonymous. Miscellaneous inventions. Scientific American, 48(13):202, March 31, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n13/pdf/scientificamerican03311883-202c.pdf.

Anonymous:1883:Cf

[3838] Anonymous. Correspondence. Scientific American, 48(14):213, April 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n14/pdf/scientificamerican04071883-213.pdf.

Anonymous:1883:BNBa

[3839] Anonymous. Books: New books and publications. Scientific American, 48(14):216, April 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n14/pdf/scientificamerican04071883-
216a.pdf.

Anonymous:1883:IRIe

[3840] Anonymous. Inventions: Recent inventions. Scientific American, 48(14):217, April 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n14/pdf/scientificamerican04071883-
217.pdf.

Anonymous:1883:BPf

[3841] Anonymous. Business and personal. Scientific American, 48(14):219, April 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n14/pdf/scientificamerican04071883-219.pdf.

Anonymous:1883:BNBb

[3842] Anonymous. Books: New books and publications. Scientific American, 48(14):219, April 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n14/pdf/scientificamerican04071883-
219a.pdf.

Anonymous:1883:NQf

[3843] Anonymous. Notes and queries. Scientific American, 48(14):219, April 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n14/pdf/scientificamerican04071883-219b.pdf.

Anonymous:1883:IIe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n14/pdf/scientificamerican04071883-219c.pdf.

Anonymous:1883:GRT

[3845] Anonymous. The great Russian telescope. Scientific American, 48(14):207–208, April 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n14/pdf/scientificamerican04071883-
207.pdf.

Anonymous:1883:TSEa

[3846] Anonymous. The total solar eclipse of May 6. Scientific American, 48(14):208–209, April 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n14/pdf/scientificamerican04071883-
208.pdf.

Anonymous:1883:ENM

[3847] Anonymous. English nail makers and their wretchedness. Scientific American, 48(14):209, April 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n14/pdf/scientificamerican04071883-
209.pdf.

Anonymous:1883:HAB

[3848] Anonymous. Hot air for boiler furnaces. Scientific American, 48 (14):209, April 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n14/pdf/scientificamerican04071883-
209a.pdf.

Anonymous:1883:STa

[3849] Anonymous. Swearing by telephone. Scientific American, 48(14):209, April 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n14/pdf/scientificamerican04071883-209b.pdf.

Anonymous:1883:SSSb

[3850] Anonymous. Side show science. Scientific American, 48(14):210, April 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n14/pdf/scientificamerican04071883-210.pdf.

Anonymous:1883:OEG

[3851] Anonymous. Ornamental entrance gate. Scientific American, 48 (14):211–212, April 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n14/pdf/scientificamerican04071883-
211.pdf.

Anonymous:1883:NTC

[3852] Anonymous. A new tool cabinet. Scientific American, 48(14):212, April 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n14/pdf/scientificamerican04071883-212.pdf.

Anonymous:1883:HSF

[3853] Anonymous. Huttinger’s stovepipe fastener. Scientific American, 48(14):212, April 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n14/pdf/scientificamerican04071883-
212a.pdf.

Anonymous:1883:ATB

[3854] Anonymous. Apparatus for testing breadstuffs. Scientific American, 48(14):214–215, April 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n14/pdf/scientificamerican04071883-
214.pdf.

Anonymous:1883:MCa

[3855] Anonymous. The Masdevallia Chimera. Scientific American, 48 (14):215, April 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n14/pdf/scientificamerican04071883-
215.pdf.

Anonymous:1883:DRPa

[3856] Anonymous. Decisions relating to patents. Scientific American, 48(14):216, April 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n14/pdf/scientificamerican04071883-
216.pdf.

Anonymous:1883:EId

[3857] Anonymous. Engineering inventions. Scientific American, 48(14):218, April 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n14/pdf/scientificamerican04071883-218.pdf.

Anonymous:1883:MIk

[3858] Anonymous. Mechanical inventions. Scientific American, 48(14):218, April 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n14/pdf/scientificamerican04071883-218a.pdf.

Anonymous:1883:AIf

[3859] Anonymous. Agricultural inventions. Scientific American, 48(14):218, April 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n14/pdf/scientificamerican04071883-218b.pdf.

Anonymous:1883:MIl

[3860] Anonymous. Miscellaneous inventions. Scientific American, 48(14):218, April 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n14/pdf/scientificamerican04071883-218c.pdf.

Anonymous:1883:Cg

[3861] Anonymous. Correspondence. Scientific American, 48(15):229, April 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n15/pdf/scientificamerican04141883-229.pdf.

Anonymous:1883:IRIf

[3862] Anonymous. Inventions: Recent inventions. Scientific American, 48(15):232, April 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n15/pdf/scientificamerican04141883-
232.pdf.

Anonymous:1883:BPg

[3863] Anonymous. Business and personal. Scientific American, 48(15):234, April 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n15/pdf/scientificamerican04141883-234.pdf.

Anonymous:1883:BNBc

[3864] Anonymous. Books: New books and publications. Scientific American, 48(15):234, April 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n15/pdf/scientificamerican04141883-
234a.pdf.

Anonymous:1883:NQg

[3865] Anonymous. Notes and queries. Scientific American, 48(15):234–235, April 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n15/pdf/scientificamerican04141883-234b.pdf.

Anonymous:1883:IIf

[3866] Anonymous. Index of inventions. Scientific American, 48(15):235–236, April 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n15/pdf/scientificamerican04141883-235a.pdf.

Anonymous:1883:TPM

[3867] Anonymous. Transmission of power by means of electricity. Scientific American, 48(15):223, April 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n15/pdf/scientificamerican04141883-
223.pdf.

Anonymous:1883:PC

[3868] Anonymous. Peter Cooper. Scientific American, 48(15):224, April 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n15/pdf/scientificamerican04141883-224.pdf.

Anonymous:1883:MVM

[3869] Anonymous. Mosquitoes vs. malaria. Scientific American, 48 (15):224–225, April 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n15/pdf/scientificamerican04141883-
224a.pdf.

Anonymous:1883:EVS

[3870] Anonymous. The extermination of venomous serpents. Scientific American, 48(15):225–226, April 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n15/pdf/scientificamerican04141883-
225.pdf.

Anonymous:1883:MEG

[3871] Anonymous. Machine for examining goods. Scientific American, 48(15):226, April 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n15/pdf/scientificamerican04141883-
226.pdf.

Anonymous:1883:RAR

[3872] Anonymous. Robert’s automatic rain water separator. Scientific American, 48(15):226–227, April 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n15/pdf/scientificamerican04141883-
226a.pdf.

Anonymous:1883:ITa

[3873] Anonymous. Improved tricycles. Scientific American, 48(15):227, April 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n15/pdf/scientificamerican04141883-227.pdf.

Anonymous:1883:IFEa

[3874] Anonymous. Improved fire escape. Scientific American, 48(15):227–228, April 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n15/pdf/scientificamerican04141883-227a.pdf.

Riley:1883:JSG

[3875] C. V. Riley. Jumping seeds and galls. Scientific American, 48(15):228, April 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n15/pdf/scientificamerican04141883-228.pdf.

Anonymous:1883:DGP

[3876] Anonymous. Dowson’s gas producer. Scientific American, 48(15):230, April 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n15/pdf/scientificamerican04141883-230.pdf.

Anonymous:1883:CGM

[3877] Anonymous. Combined gas motor and refrigerator. Scientific American, 48(15):230–231, April 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n15/pdf/scientificamerican04141883-
230a.pdf.

Anonymous:1883:ELS

[3878] Anonymous. The electric light in surgical diagnosis. Scientific American, 48(15):231, April 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n15/pdf/scientificamerican04141883-
231.pdf.

Anonymous:1883:PFa

[3879] Anonymous. The paradise fish. Scientific American, 48(15):231–232, April 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n15/pdf/scientificamerican04141883-231a.pdf.

Anonymous:1883:EIe

[3880] Anonymous. Engineering inventions. Scientific American, 48(15):233, April 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n15/pdf/scientificamerican04141883-233.pdf.

Anonymous:1883:MIm

[3881] Anonymous. Mechanical inventions. Scientific American, 48(15):233, April 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n15/pdf/scientificamerican04141883-233a.pdf.

Anonymous:1883:AIg

[3882] Anonymous. Agricultural inventions. Scientific American, 48(15):233, April 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n15/pdf/scientificamerican04141883-233b.pdf.

Anonymous:1883:MIn

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n15/pdf/scientificamerican04141883-233c.pdf.

Anonymous:1883:CRb

[3884] Anonymous. Communications received. Scientific American, 48 (15):235, April 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n15/pdf/scientificamerican04141883-
235.pdf.

Anonymous:1883:Ch

[3885] Anonymous. Correspondence. Scientific American, 48(16):245–246, April 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n16/pdf/scientificamerican04211883-245.pdf.

Anonymous:1883:IRIg

[3886] Anonymous. Inventions: Recent inventions. Scientific American, 48(16):248, April 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n16/pdf/scientificamerican04211883-
248.pdf.

Anonymous:1883:BPh

[3887] Anonymous. Business and personal. Scientific American, 48(16):250, April 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n16/pdf/scientificamerican04211883-250.pdf.

Anonymous:1883:BNBd

[3888] Anonymous. Books: New books and publications. Scientific American, 48(16):250, April 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n16/pdf/scientificamerican04211883-
250a.pdf.

Anonymous:1883:NQh

[3889] Anonymous. Notes and queries. Scientific American, 48(16):250–251, April 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n16/pdf/scientificamerican04211883-250b.pdf.

Anonymous:1883:IIg

[3890] Anonymous. Index of inventions. Scientific American, 48(16):251–252, April 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n16/pdf/scientificamerican04211883-251.pdf.

Anonymous:1883:La

[3891] Anonymous. The Lepanto. Scientific American, 48(16):239, April 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n16/pdf/scientificamerican04211883-239.pdf.

Anonymous:1883:PAD

[3892] Anonymous. Progress of archeological discovery in Central America. Scientific American, 48(16):240, April 21, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v48/
n16/pdf/scientificamerican04211883-240.pdf.

Anonymous:1883:BA

[3893] Anonymous. The bursting of armatures. Scientific American, 48 (16):240, April 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n16/pdf/scientificamerican04211883-
240a.pdf.

Anonymous:1883:CSa

[3894] Anonymous. The clock of Strassburg. Scientific American, 48(16):241, April 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n16/pdf/scientificamerican04211883-241.pdf.

Anonymous:1883:IPa

[3895] Anonymous. Insects and plants. Scientific American, 48(16):241–242, April 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n16/pdf/scientificamerican04211883-241a.pdf.

Anonymous:1883:ISM

[3896] Anonymous. Improved sawing machine. Scientific American, 48 (16):242, April 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n16/pdf/scientificamerican04211883-
242.pdf.

Anonymous:1883:NLE

[3897] Anonymous. New lamp extinguisher. Scientific American, 48(16):242, April 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n16/pdf/scientificamerican04211883-242a.pdf.

Anonymous:1883:TBC

[3898] Anonymous. Torpedo boats in collision. Scientific American, 48 (16):243, April 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n16/pdf/scientificamerican04211883-
243.pdf.

Anonymous:1883:CAR

[3899] Anonymous. Chenot’s atmospheric rock drill. Scientific American, 48(16):243–244, April 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n16/pdf/scientificamerican04211883-
243a.pdf.

Anonymous:1883:FTD

[3900] Anonymous. The first telephone, as described by the inventor. Scientific American, 48(16):244, April 21, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v48/
n16/pdf/scientificamerican04211883-244.pdf.

Anonymous:1883:MMT

[3901] Anonymous. Machine for measuring textile fabrics. Scientific American, 48(16):246, April 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n16/pdf/scientificamerican04211883-
246.pdf.

Anonymous:1883:IFEb

[3902] Anonymous. Improved fire escape. Scientific American, 48(16):246, April 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n16/pdf/scientificamerican04211883-246a.pdf.

Anonymous:1883:EPA

[3903] Anonymous. Electro pulverizer and amalgamator. Scientific American, 48(16):246–247, April 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n16/pdf/scientificamerican04211883-
246b.pdf.

Anonymous:1883:HCa

[3904] Anonymous. A hairy child. Scientific American, 48(16):247, April 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n16/pdf/scientificamerican04211883-247.pdf.

Anonymous:1883:MIo

[3905] Anonymous. Mechanical inventions. Scientific American, 48(16):249, April 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n16/pdf/scientificamerican04211883-249.pdf.

Anonymous:1883:AIh

[3906] Anonymous. Agricultural inventions. Scientific American, 48(16):249, April 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n16/pdf/scientificamerican04211883-249a.pdf.

Anonymous:1883:MIp

[3907] Anonymous. Miscellaneous inventions. Scientific American, 48(16):249, April 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n16/pdf/scientificamerican04211883-249b.pdf.

Anonymous:1883:IRIh

[3908] Anonymous. Inventions: Recent inventions. Scientific American, 48(17):264, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n17/pdf/scientificamerican04281883-
264.pdf.

Anonymous:1883:BPi

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n17/pdf/scientificamerican04281883-266.pdf.

Anonymous:1883:BNBe

[3910] Anonymous. Books: New books and publications. Scientific American, 48(17):266, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n17/pdf/scientificamerican04281883-
266a.pdf.

Anonymous:1883:NQi

[3911] Anonymous. Notes and queries. Scientific American, 48(17):266–267, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n17/pdf/scientificamerican04281883-266b.pdf.

Anonymous:1883:IIh

[3912] Anonymous. Index of inventions. Scientific American, 48(17):267–268, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n17/pdf/scientificamerican04281883-267a.pdf.

Anonymous:1883:PNPb

[3913] Anonymous. Patent notes: Patents. Scientific American, 48(17):268, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n17/pdf/scientificamerican04281883-268.pdf.

Anonymous:1883:ORD

[3914] Anonymous. The Ottawa River Dam. Scientific American, 48 (17):255–256, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n17/pdf/scientificamerican04281883-
255.pdf.

Anonymous:1883:DGW

[3915] Anonymous. Does good work pay? Scientific American, 48(17):256, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n17/pdf/scientificamerican04281883-256.pdf.

Anonymous:1883:GBA

[3916] Anonymous. The great bridge almost completed. Scientific American, 48(17):256, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n17/pdf/scientificamerican04281883-
256a.pdf.

Anonymous:1883:RCS

[3917] Anonymous. Relative costs of street lighting by electricity and gas in New York. Scientific American, 48(17):256, April 28, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v48/
n17/pdf/scientificamerican04281883-256b.pdf.

Anonymous:1883:MDC

[3918] Anonymous. Making a drill chuck. Scientific American, 48(17):256, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n17/pdf/scientificamerican04281883-256c.pdf.

Anonymous:1883:APMa

[3919] Anonymous. Aspects of the planets for May. Scientific American, 48(17):257, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n17/pdf/scientificamerican04281883-
257.pdf.

Anonymous:1883:AAB

[3920] Anonymous. An artificial Aurora Borealis. Scientific American, 48(17):257, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n17/pdf/scientificamerican04281883-
257a.pdf.

Hovey:1883:SLI

[3921] H. C. Hovey. Seals in Long Island Sound. Scientific American, 48(17):257, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n17/pdf/scientificamerican04281883-
257b.pdf.

Anonymous:1883:BS

[3922] Anonymous. Bulkheads for steamships. Scientific American, 48 (17):258–259, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n17/pdf/scientificamerican04281883-
258.pdf.

Anonymous:1883:PFb

[3923] Anonymous. Piefke’s filter. Scientific American, 48(17):259–260, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n17/pdf/scientificamerican04281883-259.pdf.

Anonymous:1883:HRT

[3924] Anonymous. The Hudson River Tunnel. Scientific American, 48 (17):260–261, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n17/pdf/scientificamerican04281883-
260.pdf.

Anonymous:1883:MGTa

[3925] Anonymous. Mice-girdled trees. Scientific American, 48(17):261–262, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n17/pdf/scientificamerican04281883-261.pdf.

Anonymous:1883:ITH

[3926] Anonymous. Improvement in tunnel and house building. Scientific American, 48(17):262, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n17/pdf/scientificamerican04281883-
262.pdf.

Anonymous:1883:AFR

[3927] Anonymous. Automatic four-roller mill. Scientific American, 48 (17):262, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n17/pdf/scientificamerican04281883-
262a.pdf.

Anonymous:1883:MCb

[3928] Anonymous. Manufacture of charcoal. Scientific American, 48(17):263, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n17/pdf/scientificamerican04281883-263.pdf.

Anonymous:1883:PA

[3929] Anonymous. Physics without apparatus. Scientific American, 48 (17):263, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n17/pdf/scientificamerican04281883-
263a.pdf.

Anonymous:1883:EIf

[3930] Anonymous. Engineering inventions. Scientific American, 48(17):265, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n17/pdf/scientificamerican04281883-265.pdf.

Anonymous:1883:MIq

[3931] Anonymous. Mechanical inventions. Scientific American, 48(17):265, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n17/pdf/scientificamerican04281883-265a.pdf.

Anonymous:1883:AIi

[3932] Anonymous. Agricultural inventions. Scientific American, 48(17):265, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n17/pdf/scientificamerican04281883-265b.pdf.

Anonymous:1883:MIr

[3933] Anonymous. Miscellaneous inventions. Scientific American, 48(17):265, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n17/pdf/scientificamerican04281883-265c.pdf.

Anonymous:1883:CRd

[3934] Anonymous. Communications received. Scientific American, 48 (17):267, April 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n17/pdf/scientificamerican04281883-
267.pdf.

Anonymous:1883:NQj

[3935] Anonymous. Notes and queries. Scientific American, 48(2):27–28, January 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n2/pdf/scientificamerican01131883-27.pdf.

Anonymous:1883:BPj

[3936] Anonymous. Business and personal. Scientific American, 48(2):28, January 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n2/pdf/scientificamerican01131883-28.pdf.

Anonymous:1883:PNPc

[3937] Anonymous. Patent notes: Patents. Scientific American, 48(2):28, January 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n2/pdf/scientificamerican01131883-28a.pdf.

Anonymous:1883:GDD

[3938] Anonymous. The great dry docks at the Erie Basin. Scientific American, 48(2):15, January 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n2/pdf/scientificamerican01131883-
15.pdf.

Anonymous:1883:PSEa

[3939] Anonymous. Perils from suspended electric wires. Scientific American, 48(2):16, January 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n2/pdf/scientificamerican01131883-
16.pdf.

Anonymous:1883:TICa

[3940] Anonymous. Testing and improving cast iron. Scientific American, 48(2):16, January 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n2/pdf/scientificamerican01131883-
16a.pdf.

Anonymous:1883:MWI

[3941] Anonymous. More work for inventors. Scientific American, 48 (2):17, January 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n2/pdf/scientificamerican01131883-
17.pdf.

Anonymous:1883:SHS

[3942] Anonymous. Sawing hard steel with sand. Scientific American, 48(2):17–18, January 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n2/pdf/scientificamerican01131883-
17a.pdf.

Anonymous:1883:NHC

[3943] Anonymous. New hose coupling. Scientific American, 48(2):18–19, January 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n2/pdf/scientificamerican01131883-18.pdf.

Anonymous:1883:LSA

[3944] Anonymous. Log setting apparatus for sawmills. Scientific American, 48(2):19, January 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n2/pdf/scientificamerican01131883-
19.pdf.

Anonymous:1883:IVa

[3945] Anonymous. Improved vise. Scientific American, 48(2):19, January 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n2/pdf/scientificamerican01131883-19a.pdf.

Anonymous:1883:NPD

[3946] Anonymous. New potato digger. Scientific American, 48(2):22, January 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n2/pdf/scientificamerican01131883-22.pdf.

Anonymous:1883:LAHa

[3947] Anonymous. Laboratory apparatus for heating water. Scientific American, 48(2):22–23, January 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n2/pdf/scientificamerican01131883-
22a.pdf.

Anonymous:1883:OWJ

[3948] Anonymous. An ornamental water jar and basin. Scientific American, 48(2):23, January 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n2/pdf/scientificamerican01131883-
23.pdf.

Anonymous:1883:WPW

[3949] Anonymous. What is a pike? — what is a pickerel? Scientific American, 48(2):24–25, January 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n2/pdf/scientificamerican01131883-
24.pdf.

Anonymous:1883:EIg

[3950] Anonymous. Engineering inventions. Scientific American, 48(2): 26, January 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n2/pdf/scientificamerican01131883-
26.pdf.

Anonymous:1883:MIs

[3951] Anonymous. Mechanical inventions. Scientific American, 48(2):26, January 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n2/pdf/scientificamerican01131883-26a.pdf.

Anonymous:1883:AIj

[3952] Anonymous. Agricultural inventions. Scientific American, 48 (2):26, January 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n2/pdf/scientificamerican01131883-
26b.pdf.

Anonymous:1883:MIt

[3953] Anonymous. Miscellaneous inventions. Scientific American, 48 (2):26, January 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n2/pdf/scientificamerican01131883-
26c.pdf.

Anonymous:1883:IRIi

[3954] Anonymous. Inventions: Recent inventions. Scientific American, 48(21):328, May 26, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n21/pdf/scientificamerican05261883-
328.pdf.

Anonymous:1883:BPk

[3955] Anonymous. Business and personal. Scientific American, 48(21):330, May 26, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n21/pdf/scientificamerican05261883-330.pdf.

Anonymous:1883:BNBf

[3956] Anonymous. Books: New books and publications. Scientific American, 48(21):330, May 26, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n21/pdf/scientificamerican05261883-
330a.pdf.

Anonymous:1883:NQk

[3957] Anonymous. Notes and queries. Scientific American, 48(21):330–331, May 26, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n21/pdf/scientificamerican05261883-330b.pdf.

Anonymous:1883:IIi

[3958] Anonymous. Index of inventions. Scientific American, 48(21):331–332, May 26, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n21/pdf/scientificamerican05261883-331.pdf.

Anonymous:1883:PNPd

[3959] Anonymous. Patent notes: Patents. Scientific American, 48(21):332, May 26, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n21/pdf/scientificamerican05261883-332.pdf.

Anonymous:1883:OGB

[3960] Anonymous. Opening of the great bridge. Scientific American, 48(21):320, May 26, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n21/pdf/scientificamerican05261883-
320.pdf.

Anonymous:1883:PSL

[3961] Anonymous. Placing the Statue of Liberty. Scientific American, 48(21):320, May 26, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n21/pdf/scientificamerican05261883-
320a.pdf.

Anonymous:1883:DWP

[3962] Anonymous. The Drive Well Patent. Scientific American, 48(21):320, May 26, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n21/pdf/scientificamerican05261883-320b.pdf.

Anonymous:1883:UP

[3963] Anonymous. Uses of paper. Scientific American, 48(21):320, May 26, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n21/pdf/scientificamerican05261883-320c.pdf.

Anonymous:1883:BW

[3964] Anonymous. Belt widths. Scientific American, 48(21):320–321, May 26, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n21/pdf/scientificamerican05261883-320d.pdf.

Anonymous:1883:GPE

[3965] Anonymous. Goodridge’s plan for the erection of Bartholdi’s Statue. Scientific American, 48(21):322–323, May 26, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v48/
n21/pdf/scientificamerican05261883-322.pdf.

Anonymous:1883:LPM

[3966] Anonymous. De Labastide’s pocket mitrailleuse. Scientific American, 48(21):323, May 26, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n21/pdf/scientificamerican05261883-
323.pdf.

Anonymous:1883:NYB

[3967] Anonymous. The New York and Brooklyn Bridge. Scientific American, 48(21):324–326, May 26, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n21/pdf/scientificamerican05261883-
324.pdf.

Anonymous:1883:MBCa

[3968] Anonymous. Machine for branding cores and stoppers. Scientific American, 48(21):326, May 26, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n21/pdf/scientificamerican05261883-
326.pdf.

Anonymous:1883:PGa

[3969] Anonymous. The pouched gopher. Scientific American, 48(21):327, May 26, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n21/pdf/scientificamerican05261883-327.pdf.

Anonymous:1883:NHNb

[3970] Anonymous. Natural history notes. Scientific American, 48(21):327, May 26, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n21/pdf/scientificamerican05261883-327a.pdf.

Anonymous:1883:EIh

[3971] Anonymous. Engineering inventions. Scientific American, 48(21):329, May 26, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n21/pdf/scientificamerican05261883-329.pdf.

Anonymous:1883:MIu

[3972] Anonymous. Mechanical inventions. Scientific American, 48(21):329, May 26, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n21/pdf/scientificamerican05261883-329a.pdf.

Anonymous:1883:MIv

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n21/pdf/scientificamerican05261883-329b.pdf.

Anonymous:1883:Ci

[3974] Anonymous. Correspondence. Scientific American, 48(22):341, June 2, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n22/pdf/scientificamerican06021883-341.pdf.

Anonymous:1883:BPl

[3975] Anonymous. Business and personal. Scientific American, 48(22):346, June 2, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n22/pdf/scientificamerican06021883-346.pdf.

Anonymous:1883:BNBg

[3976] Anonymous. Books: New books and publications. Scientific American, 48(22):346, June 2, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n22/pdf/scientificamerican06021883-
346a.pdf.

Anonymous:1883:NQl

[3977] Anonymous. Notes and queries. Scientific American, 48(22):346–347, June 2, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n22/pdf/scientificamerican06021883-346b.pdf.

Anonymous:1883:IIj

[3978] Anonymous. Index of inventions. Scientific American, 48(22):347–348, June 2, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n22/pdf/scientificamerican06021883-347a.pdf.

Anonymous:1883:PNPe

[3979] Anonymous. Patent notes: Patents. Scientific American, 48(22):348, June 2, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n22/pdf/scientificamerican06021883-348.pdf.

Anonymous:1883:BO

[3980] Anonymous. The bridge opened. Scientific American, 48(22):336, June 2, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n22/pdf/scientificamerican06021883-336.pdf.

Anonymous:1883:DWM

[3981] Anonymous. Death of William Mason. Scientific American, 48(22):336, June 2, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n22/pdf/scientificamerican06021883-336a.pdf.

Anonymous:1883:COE

[3982] Anonymous. Consideration for old employes. Scientific American, 48(22):336, June 2, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n22/pdf/scientificamerican06021883-
336b.pdf.

Anonymous:1883:RWS

[3983] Anonymous. A remarkable wind storm. Scientific American, 48 (22):336–337, June 2, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n22/pdf/scientificamerican06021883-
336c.pdf.

Anonymous:1883:AT

[3984] Anonymous. The Athenian trireme. Scientific American, 48 (22):338–339, June 2, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n22/pdf/scientificamerican06021883-
338.pdf.

Anonymous:1883:ARP

[3985] Anonymous. Apparatus for regulating the pressure in water conduits. Scientific American, 48(22):339, June 2, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v48/
n22/pdf/scientificamerican06021883-339.pdf.

Anonymous:1883:NEA

[3986] Anonymous. New exercising apparatus. Scientific American, 48 (22):339, June 2, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n22/pdf/scientificamerican06021883-
339a.pdf.

Anonymous:1883:EBR

[3987] Anonymous. Engine and boiler rooms of the great bridge. Scientific American, 48(22):340, June 2, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n22/pdf/scientificamerican06021883-
340.pdf.

Anonymous:1883:DRPb

[3988] Anonymous. Decisions relating to patents, trade marks, etc. Scientific American, 48(22):341–342, June 2, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v48/n22/pdf/scientificamerican06021883-
341a.pdf.

Anonymous:1883:GWS

[3989] Anonymous. The great war ship Lepanto. Scientific American, 48(22):342, June 2, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n22/pdf/scientificamerican06021883-
342.pdf.

Anonymous:1883:PSM

[3990] Anonymous. Punching and shearing machine. Scientific American, 48(22):342, June 2, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n22/pdf/scientificamerican06021883-
342a.pdf.

Coryell:1883:GP

[3991] John R. Coryell. Gulling the pelican. Scientific American, 48 (22):343–344, June 2, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n22/pdf/scientificamerican06021883-
343.pdf.

Anonymous:1883:EIi

[3992] Anonymous. Engineering inventions. Scientific American, 48(22):345, June 2, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n22/pdf/scientificamerican06021883-345.pdf.

Anonymous:1883:MIw

[3993] Anonymous. Mechanical inventions. Scientific American, 48(22):345, June 2, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n22/pdf/scientificamerican06021883-345a.pdf.

Anonymous:1883:AIk

[3994] Anonymous. Agricultural inventions. Scientific American, 48(22):345, June 2, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n22/pdf/scientificamerican06021883-345b.pdf.

Anonymous:1883:MIx

[3995] Anonymous. Miscellaneous inventions. Scientific American, 48(22):345, June 2, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n22/pdf/scientificamerican06021883-345c.pdf.

Anonymous:1883:CRe

[3996] Anonymous. Communications received. Scientific American, 48 (22):347, June 2, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n22/pdf/scientificamerican06021883-
347.pdf.

Anonymous:1883:Cj

[3997] Anonymous. Correspondence. Scientific American, 48(23):357, June 9, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n23/pdf/scientificamerican06091883-357.pdf.

Anonymous:1883:IRIj

[3998] Anonymous. Inventions: Recent inventions. Scientific American, 48(23):360, June 9, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n23/pdf/scientificamerican06091883-
360.pdf.

Anonymous:1883:BPm

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n23/pdf/scientificamerican06091883-362.pdf.

Anonymous:1883:BNBh

[4000] Anonymous. Books: New books and publications. Scientific American, 48(23):362, June 9, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n23/pdf/scientificamerican06091883-
362a.pdf.

Anonymous:1883:NQm

[4001] Anonymous. Notes and queries. Scientific American, 48(23):362–363, June 9, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n23/pdf/scientificamerican06091883-362b.pdf.

Anonymous:1883:IIk

[4002] Anonymous. Index of inventions. Scientific American, 48(23):363–364, June 9, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n23/pdf/scientificamerican06091883-363.pdf.

Anonymous:1883:PNPf

[4003] Anonymous. Patent notes: Patents. Scientific American, 48(23):364, June 9, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n23/pdf/scientificamerican06091883-364.pdf.

Anonymous:1883:DFT

[4004] Anonymous. Drop forging of tool steel. Scientific American, 48(23):352, June 9, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n23/pdf/scientificamerican06091883-352.pdf.

Anonymous:1883:AGB

[4005] Anonymous. An accident on the great bridge. Scientific American, 48(23):353, June 9, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n23/pdf/scientificamerican06091883-
353.pdf.

Anonymous:1883:APJa

[4006] Anonymous. Aspects of the pianets for June. Scientific American, 48(23):353, June 9, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n23/pdf/scientificamerican06091883-
353a.pdf.

Anonymous:1883:NAM

[4007] Anonymous. A novel air motor. Scientific American, 48(23):354, June 9, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n23/pdf/scientificamerican06091883-354.pdf.

Anonymous:1883:IWH

[4008] Anonymous. Improved water heater. Scientific American, 48(23):354, June 9, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n23/pdf/scientificamerican06091883-354a.pdf.

Anonymous:1883:ADG

[4009] Anonymous. Atmospheric dust and germs. Scientific American, 48(23):355–356, June 9, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n23/pdf/scientificamerican06091883-
355.pdf.

Anonymous:1883:FBP

[4010] Anonymous. Fire and burglar proof safes. Scientific American, 48(23):356, June 9, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n23/pdf/scientificamerican06091883-
356.pdf.

Anonymous:1883:RDRa

[4011] Anonymous. Recent decisions relating to patents. Scientific American, 48(23):356–357, June 9, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n23/pdf/scientificamerican06091883-
356a.pdf.

Anonymous:1883:ELA

[4012] Anonymous. Electric lighting on the Arizona. Scientific American, 48(23):358, June 9, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n23/pdf/scientificamerican06091883-
358.pdf.

Anonymous:1883:HCb

[4013] Anonymous. The helmet cassowary. Scientific American, 48(23):359, June 9, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n23/pdf/scientificamerican06091883-359.pdf.

Anonymous:1883:EIj

[4014] Anonymous. Engineering inventions. Scientific American, 48(23):361, June 9, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n23/pdf/scientificamerican06091883-361.pdf.

Anonymous:1883:MIy

[4015] Anonymous. Mechanical inventions. Scientific American, 48(23):361, June 9, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n23/pdf/scientificamerican06091883-361a.pdf.

Anonymous:1883:AIl

[4016] Anonymous. Agricultural inventions. Scientific American, 48(23):361, June 9, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n23/pdf/scientificamerican06091883-361b.pdf.

Anonymous:1883:MIz

[4017] Anonymous. Miscellaneous inventions. Scientific American, 48(23):361, June 9, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n23/pdf/scientificamerican06091883-361c.pdf.

Anonymous:1883:Ck

[4018] Anonymous. Correspondence. Scientific American, 48(24):373, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n24/pdf/scientificamerican06161883-373.pdf.

Anonymous:1883:IRIk

[4019] Anonymous. Inventions: Recent inventions. Scientific American, 48(24):377, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n24/pdf/scientificamerican06161883-
377.pdf.

Anonymous:1883:BPn

[4020] Anonymous. Business and personal. Scientific American, 48(24):379, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n24/pdf/scientificamerican06161883-379.pdf.

Anonymous:1883:NQn

[4021] Anonymous. Notes and queries. Scientific American, 48(24):379, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n24/pdf/scientificamerican06161883-379a.pdf.

Anonymous:1883:IIl

[4022] Anonymous. Index of inventions. Scientific American, 48(24):379–380, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n24/pdf/scientificamerican06161883-379b.pdf.

Anonymous:1883:ICE

[4023] Anonymous. Improved caloric engine. Scientific American, 48 (24):367–368, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n24/pdf/scientificamerican06161883-
367.pdf.

Anonymous:1883:RP

[4024] Anonymous. The rebacking of paintings. Scientific American, 48 (24):368, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n24/pdf/scientificamerican06161883-
368.pdf.

Anonymous:1883:MYB

[4025] Anonymous. “Mind your business”. Scientific American, 48(24):368, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n24/pdf/scientificamerican06161883-368a.pdf.

Anonymous:1883:OBF

[4026] Anonymous. Overthrow of the barb fence patents. Scientific American, 48(24):368, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n24/pdf/scientificamerican06161883-
368b.pdf.

Anonymous:1883:ICa

[4027] Anonymous. The induced current. Scientific American, 48(24):368–369, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n24/pdf/scientificamerican06161883-368c.pdf.

Anonymous:1883:AASb

[4028] Anonymous. American Astronomical Society. Scientific American, 48(24):369, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n24/pdf/scientificamerican06161883-
369.pdf.

Anonymous:1883:FPL

[4029] Anonymous. The French Physical Laboratories. Scientific American, 48(24):369, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n24/pdf/scientificamerican06161883-
369a.pdf.

Anonymous:1883:RNV

[4030] Anonymous. Round noses vs. diamond shape. Scientific American, 48(24):369–370, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n24/pdf/scientificamerican06161883-
369b.pdf.

Anonymous:1883:CTa

[4031] Anonymous. The Columbia tricycle. Scientific American, 48(24):370, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n24/pdf/scientificamerican06161883-370.pdf.

Anonymous:1883:MCG

[4032] Anonymous. Mixer for carbonic gas and water. Scientific American, 48(24):371, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n24/pdf/scientificamerican06161883-
371.pdf.

Anonymous:1883:ISU

[4033] Anonymous. The influence of storms upon water traps. Scientific American, 48(24):371, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n24/pdf/scientificamerican06161883-
371a.pdf.

Anonymous:1883:PMS

[4034] Anonymous. Portable meat safe. Scientific American, 48(24):371–372, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n24/pdf/scientificamerican06161883-371b.pdf.

Anonymous:1883:RDRb

[4035] Anonymous. Recent decisions relating to patents. Scientific American, 48(24):373, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n24/pdf/scientificamerican06161883-
373a.pdf.

Anonymous:1883:MDH

[4036] Anonymous. Making the deaf to hear. Scientific American, 48 (24):373–374, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n24/pdf/scientificamerican06161883-
373b.pdf.

Anonymous:1883:QLW

[4037] Anonymous. A quadricycle for land and water. Scientific American, 48(24):374, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n24/pdf/scientificamerican06161883-
374.pdf.

Anonymous:1883:PGT

[4038] Anonymous. Pressure gauge testing apparatus. Scientific American, 48(24):374, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n24/pdf/scientificamerican06161883-
374a.pdf.

Anonymous:1883:GL

[4039] Anonymous. The giant locust. Scientific American, 48(24):375, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n24/pdf/scientificamerican06161883-375.pdf.

Anonymous:1883:FHC

[4040] Anonymous. Farm hand carts. Scientific American, 48(24):376, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n24/pdf/scientificamerican06161883-376.pdf.

Anonymous:1883:ERS

[4041] Anonymous. Electrical railway signaling. Scientific American, 48 (24):376, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n24/pdf/scientificamerican06161883-
376a.pdf.

Anonymous:1883:EIk

[4042] Anonymous. Engineering inventions. Scientific American, 48(24):378, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n24/pdf/scientificamerican06161883-378.pdf.

Anonymous:1883:MIba

[4043] Anonymous. Mechanical inventions. Scientific American, 48(24):378, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n24/pdf/scientificamerican06161883-378a.pdf.

Anonymous:1883:AIm

[4044] Anonymous. Agricultural inventions. Scientific American, 48(24):378, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n24/pdf/scientificamerican06161883-378b.pdf.

Anonymous:1883:MIbb

[4045] Anonymous. Miscellaneous inventions. Scientific American, 48(24):378, June 16, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n24/pdf/scientificamerican06161883-378c.pdf.

Anonymous:1883:IRIl

[4046] Anonymous. Inventions: Recent inventions. Scientific American, 48(25):393, June 23, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n25/pdf/scientificamerican06231883-
393.pdf.

Anonymous:1883:BPo

[4047] Anonymous. Business and personal. Scientific American, 48(25):395, June 23, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n25/pdf/scientificamerican06231883-395.pdf.

Anonymous:1883:BNBi

[4048] Anonymous. Books: New books and publications. Scientific American, 48(25):395, June 23, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n25/pdf/scientificamerican06231883-
395a.pdf.

Anonymous:1883:NQo

[4049] Anonymous. Notes and queries. Scientific American, 48(25):395, June 23, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n25/pdf/scientificamerican06231883-395b.pdf.

Anonymous:1883:IIn

[4050] Anonymous. Index of inventions. Scientific American, 48(25):395–396, June 23, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n25/pdf/scientificamerican06231883-395c.pdf.

Anonymous:1883:NER

[4051] Anonymous. The National Exposition of Railway Appliances at Chicago. Scientific American, 48(25):374, June 23, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v48/
n25/pdf/scientificamerican06231883-384.pdf.

Anonymous:1883:MRP

[4052] Anonymous. Marinoni’s rotary printing press. Scientific American, 48(25):383, June 23, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n25/pdf/scientificamerican06231883-
383.pdf.

Anonymous:1883:RIP

[4053] Anonymous. Rainfall on the Isthmus of Panama. Scientific American, 48(25):384, June 23, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n25/pdf/scientificamerican06231883-
384a.pdf.

Anonymous:1883:MV

[4054] Anonymous. Molecular vibrations. Scientific American, 48(25): 385–386, June 23, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n25/pdf/scientificamerican06231883-
385.pdf.

Anonymous:1883:CREa

[4055] Anonymous. The Chicago Railway Exposition. Scientific American, 48(25):386–387, June 23, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n25/pdf/scientificamerican06231883-
386.pdf.

Anonymous:1883:NFE

[4056] Anonymous. New fuel economizer. Scientific American, 48(25):387–388, June 23, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n25/pdf/scientificamerican06231883-387.pdf.

Anonymous:1883:DRPc

[4057] Anonymous. Decisions relating to patents. Scientific American, 48(25):389, June 23, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n25/pdf/scientificamerican06231883-
389.pdf.

Anonymous:1883:INA

[4058] Anonymous. Improved nut arbor. Scientific American, 48(25):390, June 23, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n25/pdf/scientificamerican06231883-390.pdf.

Anonymous:1883:NAS

[4059] Anonymous. New apparatus for stirring, mixing, and melting. Scientific American, 48(25):390, June 23, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v48/
n25/pdf/scientificamerican06231883-390a.pdf.

Anonymous:1883:NCCa

[4060] Anonymous. A new constant current pile. Scientific American, 48(25):391, June 23, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n25/pdf/scientificamerican06231883-
391.pdf.

Anonymous:1883:GDO

[4061] Anonymous. Garden destroyers. — the oak leaf roller moth. Scientific American, 48(25):391, June 23, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n25/pdf/scientificamerican06231883-
391a.pdf.

Anonymous:1883:IRRa

[4062] Anonymous. Improved railway rail chair. Scientific American, 48 (25):391–392, June 23, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n25/pdf/scientificamerican06231883-
391b.pdf.

Anonymous:1883:EIl

[4063] Anonymous. Engineering inventions. Scientific American, 48(25):394, June 23, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n25/pdf/scientificamerican06231883-394.pdf.

Anonymous:1883:MIbc

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n25/pdf/scientificamerican06231883-394a.pdf.

Anonymous:1883:AIo

[4065] Anonymous. Agricultural inventions. Scientific American, 48(25):394, June 23, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n25/pdf/scientificamerican06231883-394b.pdf.

Anonymous:1883:MIbd

[4066] Anonymous. Miscellaneous inventions. Scientific American, 48(25):394, June 23, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n25/pdf/scientificamerican06231883-394c.pdf.

Anonymous:1883:Cl

[4067] Anonymous. Correspondence. Scientific American, 48(26):404, June 30, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n26/pdf/scientificamerican06301883-404.pdf.

Anonymous:1883:IRIm

[4068] Anonymous. Inventions: Recent inventions. Scientific American, 48(26):408, June 30, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n26/pdf/scientificamerican06301883-
408.pdf.

Anonymous:1883:BPp

[4069] Anonymous. Business and personal. Scientific American, 48(26):409, June 30, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n26/pdf/scientificamerican06301883-409.pdf.

Anonymous:1883:BNBj

[4070] Anonymous. Books: New books and publications. Scientific American, 48(26):409, June 30, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n26/pdf/scientificamerican06301883-
409a.pdf.

Anonymous:1883:NQp

[4071] Anonymous. Notes and queries. Scientific American, 48(26):409, June 30, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n26/pdf/scientificamerican06301883-409b.pdf.

Anonymous:1883:IIo

[4072] Anonymous. Index of inventions. Scientific American, 48(26):410, June 30, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n26/pdf/scientificamerican06301883-410.pdf.

Anonymous:1883:Ia

[4073] Anonymous. Index. Scientific American, 48(26):411–412, June 30, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n26/pdf/scientificamerican06301883-411.pdf.

Anonymous:1883:MEL

[4074] Anonymous. Manufacture of electric lamps. Scientific American, 48(26):399, June 30, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n26/pdf/scientificamerican06301883-
399.pdf.

Anonymous:1883:IPO

[4075] Anonymous. An interesting Patent Office question. Scientific American, 48(26):400, June 30, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n26/pdf/scientificamerican06301883-
400.pdf.

Anonymous:1883:CFG

[4076] Anonymous. Calipers and fixed gauges. Scientific American, 48(26):400, June 30, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n26/pdf/scientificamerican06301883-400a.pdf.

Anonymous:1883:HBF

[4077] Anonymous. High buildings and fires. Scientific American, 48(26):400, June 30, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n26/pdf/scientificamerican06301883-400b.pdf.

Anonymous:1883:RRS

[4078] Anonymous. Reports of the recent solar eclipse. Scientific American, 48(26):401, June 30, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n26/pdf/scientificamerican06301883-
401.pdf.

Anonymous:1883:APJb

[4079] Anonymous. Aspects of the planets for July. Scientific American, 48(26):401, June 30, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n26/pdf/scientificamerican06301883-
401a.pdf.

Anonymous:1883:RA

[4080] Anonymous. Raising of the Austral. Scientific American, 48(26):402, June 30, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n26/pdf/scientificamerican06301883-402.pdf.

Anonymous:1883:SP

[4081] Anonymous. The steamer Pilgrim. Scientific American, 48(26):402–403, June 30, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n26/pdf/scientificamerican06301883-402a.pdf.

Anonymous:1883:RBG

[4082] Anonymous. A railway bracket girder. Scientific American, 48(26):403, June 30, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n26/pdf/scientificamerican06301883-403.pdf.

Anonymous:1883:IWJ

[4083] Anonymous. Improved wall for jails and treasure vaults. Scientific American, 48(26):403, June 30, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n26/pdf/scientificamerican06301883-
403a.pdf.

Anonymous:1883:TWP

[4084] Anonymous. Two wheeled portable engine. Scientific American, 48(26):403, June 30, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n26/pdf/scientificamerican06301883-
403b.pdf.

Anonymous:1883:DRPd

[4085] Anonymous. Decisions relating to patents. Scientific American, 48(26):404–406, June 30, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n26/pdf/scientificamerican06301883-
404a.pdf.

Anonymous:1883:ICPa

[4086] Anonymous. Improved cotton press. Scientific American, 48(26):406, June 30, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n26/pdf/scientificamerican06301883-406.pdf.

Anonymous:1883:ILa

[4087] Anonymous. Improvement in lanterns. Scientific American, 48 (26):406–407, June 30, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n26/pdf/scientificamerican06301883-
406a.pdf.

Anonymous:1883:S

[4088] Anonymous. The shoebill. Scientific American, 48(26):407, June 30, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n26/pdf/scientificamerican06301883-407.pdf.

Anonymous:1883:CNJ

[4089] Anonymous. The College of New Jersey. Scientific American, 48(26):408, June 30, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n26/pdf/scientificamerican06301883-
408a.pdf.

Anonymous:1883:Cm

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n3/pdf/scientificamerican01201883-37.pdf.

Anonymous:1883:BPq

[4091] Anonymous. Business and personal. Scientific American, 48(3):43, January 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n3/pdf/scientificamerican01201883-43.pdf.

Anonymous:1883:NQq

[4092] Anonymous. Notes and queries. Scientific American, 48(3):43, January 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n3/pdf/scientificamerican01201883-43a.pdf.

Anonymous:1883:SCSa

[4093] Anonymous. Steel and cast steel. Scientific American, 48(3):32, January 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n3/pdf/scientificamerican01201883-32.pdf.

Anonymous:1883:FF

[4094] Anonymous. Finical finish. Scientific American, 48(3):32–33, January 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n3/pdf/scientificamerican01201883-32a.pdf.

Anonymous:1883:OFI

[4095] Anonymous. One of the first ice machines. Scientific American, 48(3):34, January 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n3/pdf/scientificamerican01201883-
34.pdf.

Anonymous:1883:NAF

[4096] Anonymous. New apparatus for feeding shavings to furnaces. Scientific American, 48(3):34, January 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n3/pdf/scientificamerican01201883-
34a.pdf.

Anonymous:1883:NHNc

[4097] Anonymous. Novel hose nozzle. Scientific American, 48(3):35, January 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n3/pdf/scientificamerican01201883-35.pdf.

Anonymous:1883:AGF

[4098] Anonymous. Adjustable guard for fireplaces. Scientific American, 48(3):35, January 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n3/pdf/scientificamerican01201883-
35a.pdf.

Anonymous:1883:AINb

[4099] Anonymous. American industries — no.86. Scientific American, 48(3):36, January 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n3/pdf/scientificamerican01201883-
36.pdf.

Anonymous:1883:APP

[4100] Anonymous. The American paper pulley. Scientific American, 48(3):38, January 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n3/pdf/scientificamerican01201883-
38.pdf.

Anonymous:1883:IVEb

[4101] Anonymous. Improved vertical engine. Scientific American, 48 (3):38, January 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n3/pdf/scientificamerican01201883-
38a.pdf.

Anonymous:1883:GBM

[4102] Anonymous. A glass blowing machine. Scientific American, 48 (3):39–41, January 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n3/pdf/scientificamerican01201883-
39.pdf.

Anonymous:1883:NHNd

[4103] Anonymous. Natural history notes. Scientific American, 48(3):41, January 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n3/pdf/scientificamerican01201883-41.pdf.

Anonymous:1883:EIm

[4104] Anonymous. Engineering inventions. Scientific American, 48(3): 42, January 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n3/pdf/scientificamerican01201883-
42.pdf.

Anonymous:1883:MIbe

[4105] Anonymous. Mechanical inventions. Scientific American, 48(3):42, January 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n3/pdf/scientificamerican01201883-42a.pdf.

Anonymous:1883:AIp

[4106] Anonymous. Agricultural inventions. Scientific American, 48 (3):42, January 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n3/pdf/scientificamerican01201883-
42b.pdf.

Anonymous:1883:MIbf

[4107] Anonymous. Miscellaneous inventions. Scientific American, 48 (3):42, January 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n3/pdf/scientificamerican01201883-
42c.pdf.

Anonymous:1883:CRf

[4108] Anonymous. Communications received. Scientific American, 48 (3):43, January 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n3/pdf/scientificamerican01201883-
43b.pdf.

Anonymous:1883:IRIn

[4109] Anonymous. Inventions: Recent inventions. Scientific American, 48(4):57, January 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n4/pdf/scientificamerican01271883-
57.pdf.

Anonymous:1883:BPr

[4110] Anonymous. Business and personal. Scientific American, 48(4):59, January 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n4/pdf/scientificamerican01271883-59.pdf.

Anonymous:1883:NQr

[4111] Anonymous. Notes and queries. Scientific American, 48(4):59–60, January 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n4/pdf/scientificamerican01271883-59a.pdf.

Anonymous:1883:PNPg

[4112] Anonymous. Patent notes: Patents. Scientific American, 48(4):60, January 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n4/pdf/scientificamerican01271883-60.pdf.

Anonymous:1883:PMM

[4113] Anonymous. Perreaux’s micrometer machine. Scientific American, 48(4):47, January 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n4/pdf/scientificamerican01271883-
47.pdf.

Anonymous:1883:GIC

[4114] Anonymous. Giffard’s ice and cold air machine. Scientific American, 48(4):47–48, January 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n4/pdf/scientificamerican01271883-
47a.pdf.

Anonymous:1883:FCL

[4115] Anonymous. Fever caused by lemonade. Scientific American, 48(4):48, January 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n4/pdf/scientificamerican01271883-
48.pdf.

Anonymous:1883:SIHa

[4116] Anonymous. Shellac for imitations of horn, ivory, etc. Scientific American, 48(4):48, January 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n4/pdf/scientificamerican01271883-
48a.pdf.

Anonymous:1883:APF

[4117] Anonymous. Aspects of the planets for February. Scientific American, 48(4):49, January 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n4/pdf/scientificamerican01271883-
49.pdf.

Riley:1883:UAH

[4118] C. T. Riley. Utilization of ants in horticulture. Scientific American, 48(4):49, January 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n4/pdf/scientificamerican01271883-
49a.pdf.

Anonymous:1883:CU

[4119] Anonymous. Cut or uncut. Scientific American, 48(4):49–50, January 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n4/pdf/scientificamerican01271883-49b.pdf.

Anonymous:1883:NB

[4120] Anonymous. New bagholder. Scientific American, 48(4):50, January 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n4/pdf/scientificamerican01271883-50.pdf.

Anonymous:1883:NFF

[4121] Anonymous. New fireproof floor and ceiling. Scientific American, 48(4):50–51, January 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n4/pdf/scientificamerican01271883-
50a.pdf.

Anonymous:1883:NT

[4122] Anonymous. Novel Tellurian. Scientific American, 48(4):51, January 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n4/pdf/scientificamerican01271883-51.pdf.

Anonymous:1883:IWa

[4123] Anonymous. Improved wrench. Scientific American, 48(4):51, January 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n4/pdf/scientificamerican01271883-51a.pdf.

Anonymous:1883:IPU

[4124] Anonymous. Improvement in parasols and umbrellas. Scientific American, 48(4):51, January 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n4/pdf/scientificamerican01271883-
51b.pdf.

Vogel:1883:VCM

[4125] August Vogel. Vegetable coloring matters. Scientific American, 48(4):52–53, January 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n4/pdf/scientificamerican01271883-
52.pdf.

Anonymous:1883:MST

[4126] Anonymous. Machine for sharpening twist drills. Scientific American, 48(4):54–55, January 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n4/pdf/scientificamerican01271883-
54.pdf.

Anonymous:1883:JPR

[4127] Anonymous. The Jardin Des Plantes reproduced by photography. Scientific American, 48(4):55, January 27, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v48/
n4/pdf/scientificamerican01271883-55.pdf.

Anonymous:1883:PN

[4128] Anonymous. Photographic notes. Scientific American, 48(4):56, January 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n4/pdf/scientificamerican01271883-56.pdf.

Anonymous:1883:ACE

[4129] Anonymous. American civil engineers. Scientific American, 48 (4):57, January 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n4/pdf/scientificamerican01271883-
57a.pdf.

Anonymous:1883:EIn

[4130] Anonymous. Engineering inventions. Scientific American, 48(4): 58, January 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n4/pdf/scientificamerican01271883-
58.pdf.

Anonymous:1883:MIbg

[4131] Anonymous. Mechanical inventions. Scientific American, 48(4):58, January 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n4/pdf/scientificamerican01271883-58a.pdf.

Anonymous:1883:AIq

[4132] Anonymous. Agricultural inventions. Scientific American, 48 (4):58, January 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n4/pdf/scientificamerican01271883-
58b.pdf.

Anonymous:1883:MIbh

[4133] Anonymous. Miscellaneous inventions. Scientific American, 48 (4):58, January 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n4/pdf/scientificamerican01271883-
58c.pdf.

Anonymous:1883:IRIo

[4134] Anonymous. Inventions: Recent inventions. Scientific American, 48(5):73, February 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n5/pdf/scientificamerican02031883-
73.pdf.

Anonymous:1883:BPs

[4135] Anonymous. Business and personal. Scientific American, 48 (5):75, February 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n5/pdf/scientificamerican02031883-
75.pdf.

Anonymous:1883:BNBk

[4136] Anonymous. Books: New books and publications. Scientific American, 48(5):75, February 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n5/pdf/scientificamerican02031883-
75a.pdf.

Anonymous:1883:NQs

[4137] Anonymous. Notes and queries. Scientific American, 48(5):75– 76, February 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n5/pdf/scientificamerican02031883-
75b.pdf.

Anonymous:1883:PNPh

[4138] Anonymous. Patent notes: Patents. Scientific American, 48 (5):76, February 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n5/pdf/scientificamerican02031883-
76a.pdf.

Anonymous:1883:CPF

[4139] Anonymous. Clothes pins, a famous yacht builder. Scientific American, 48(5):63, February 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n5/pdf/scientificamerican02031883-
63.pdf.

Anonymous:1883:TMC

[4140] Anonymous. Taylor Manufacturing Co’s patent saw mill. Scientific American, 48(5):63, February 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n5/pdf/scientificamerican02031883-
63a.pdf.

Anonymous:1883:AASc

[4141] Anonymous. An American Astronomical Society. Scientific American, 48(5):64, February 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n5/pdf/scientificamerican02031883-
64.pdf.

Anonymous:1883:CSL

[4142] Anonymous. A cure for seasickness at last. Scientific American, 48(5):64, February 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n5/pdf/scientificamerican02031883-
64a.pdf.

Anonymous:1883:AWb

[4143] Anonymous. The Adirondack Wilderness. Scientific American, 48(5):64, February 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n5/pdf/scientificamerican02031883-
64b.pdf.

Anonymous:1883:SSFb

[4144] Anonymous. The starlit sky in February. Scientific American, 48(5):64–65, February 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n5/pdf/scientificamerican02031883-
64c.pdf.

Anonymous:1883:HPE

[4145] Anonymous. The Harding process of embossing on walls. Scientific American, 48(5):66, February 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n5/pdf/scientificamerican02031883-
66.pdf.

Anonymous:1883:ISa

[4146] Anonymous. Iron shutter. Scientific American, 48(5):67, February 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n5/pdf/scientificamerican02031883-67.pdf.

Anonymous:1883:ICPb

[4147] Anonymous. Improved cotton planter. Scientific American, 48 (5):67, February 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n5/pdf/scientificamerican02031883-
67a.pdf.

Anonymous:1883:NYA

[4148] Anonymous. New York Academy of Sciences. Scientific American, 48(5):68–69, February 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n5/pdf/scientificamerican02031883-
68.pdf.

Anonymous:1883:CCH

[4149] Anonymous. Carrier for coal, hay, etc. Scientific American, 48 (5):70, February 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n5/pdf/scientificamerican02031883-
70.pdf.

Anonymous:1883:NPC

[4150] Anonymous. New postal coin scale. Scientific American, 48 (5):71, February 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n5/pdf/scientificamerican02031883-
71.pdf.

Anonymous:1883:NIE

[4151] Anonymous. A new and interesting electrical machine. Scientific American, 48(5):71–72, February 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n5/pdf/scientificamerican02031883-
71a.pdf.

Anonymous:1883:EIo

[4152] Anonymous. Engineering inventions. Scientific American, 48(5): 74, February 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n5/pdf/scientificamerican02031883-
74.pdf.

Anonymous:1883:MIbi

[4153] Anonymous. Mechanical inventions. Scientific American, 48(5): 74, February 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n5/pdf/scientificamerican02031883-
74a.pdf.

Anonymous:1883:AIr

[4154] Anonymous. Agricultural inventions. Scientific American, 48 (5):74, February 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n5/pdf/scientificamerican02031883-
74b.pdf.

Anonymous:1883:MIbj

[4155] Anonymous. Miscellaneous inventions. Scientific American, 48 (5):74, February 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n5/pdf/scientificamerican02031883-
74c.pdf.

Anonymous:1883:CRg

[4156] Anonymous. Communications received. Scientific American, 48 (5):76, February 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n5/pdf/scientificamerican02031883-
76.pdf.

Anonymous:1883:Cn

[4157] Anonymous. Correspondence. Scientific American, 48(6):85, February 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n6/pdf/scientificamerican02101883-85.pdf.

Anonymous:1883:IRIp

[4158] Anonymous. Inventions: Recent inventions. Scientific American, 48(6):89, February 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n6/pdf/scientificamerican02101883-
89.pdf.

Anonymous:1883:BPt

[4159] Anonymous. Business and personal. Scientific American, 48(6): 91, February 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n6/pdf/scientificamerican02101883-
91.pdf.

Anonymous:1883:NQt

[4160] Anonymous. Notes and queries. Scientific American, 48(6):91, February 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n6/pdf/scientificamerican02101883-91a.pdf.

Anonymous:1883:IIp

[4161] Anonymous. Index of inventions. Scientific American, 48(6):91– 92, February 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n6/pdf/scientificamerican02101883-
91c.pdf.

Anonymous:1883:PNPi

[4162] Anonymous. Patent notes: Patents. Scientific American, 48(6): 92, February 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n6/pdf/scientificamerican02101883-
92.pdf.

Anonymous:1883:AES

[4163] Anonymous. Artillery experiments at Spezia. Scientific American, 48(6):79, February 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n6/pdf/scientificamerican02101883-
79.pdf.

Anonymous:1883:CIa

[4164] Anonymous. The care of the insane. Scientific American, 48 (6):80, February 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n6/pdf/scientificamerican02101883-
80.pdf.

Anonymous:1883:PME

[4165] Anonymous. Progress in medical ethics. Scientific American, 48(6):80, February 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n6/pdf/scientificamerican02101883-
80a.pdf.

Anonymous:1883:STW

[4166] Anonymous. A snake trap wanted. Scientific American, 48(6):80, February 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n6/pdf/scientificamerican02101883-80b.pdf.

Anonymous:1883:HSEa

[4167] Anonymous. How steam engine valves are finished. Scientific American, 48(6):81, February 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n6/pdf/scientificamerican02101883-
81.pdf.

Anonymous:1883:LIW

[4168] Anonymous. The Lambertville Iron Works automatic cut-off engine. Scientific American, 48(6):82–84, February 10, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v48/
n6/pdf/scientificamerican02101883-82.pdf.

Anonymous:1883:IBS

[4169] Anonymous. Improved band saw. Scientific American, 48(6):86, February 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n6/pdf/scientificamerican02101883-86.pdf.

Anonymous:1883:ICCb

[4170] Anonymous. Improved cotton chopper. Scientific American, 48 (6):86, February 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n6/pdf/scientificamerican02101883-
86a.pdf.

Anonymous:1883:ARG

[4171] Anonymous. Apparatus for removing grease, air, and impurities from feed water. Scientific American, 48(6):87, February 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n6/pdf/scientificamerican02101883-87.pdf.

Anonymous:1883:ANa

[4172] Anonymous. Aerial navigation. Scientific American, 48(6):87– 88, February 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n6/pdf/scientificamerican02101883-
87a.pdf.

Anonymous:1883:EIq

[4173] Anonymous. Engineering inventions. Scientific American, 48(6): 90, February 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n6/pdf/scientificamerican02101883-
90.pdf.

Anonymous:1883:MIbk

[4174] Anonymous. Mechanical inventions. Scientific American, 48(6): 90, February 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n6/pdf/scientificamerican02101883-
90a.pdf.

Anonymous:1883:AIs

[4175] Anonymous. Agricultural inventions. Scientific American, 48(6): 90, February 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n6/pdf/scientificamerican02101883-
90b.pdf.

Anonymous:1883:MIbl

[4176] Anonymous. Miscellaneous inventions. Scientific American, 48 (6):90, February 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n6/pdf/scientificamerican02101883-
90c.pdf.

Anonymous:1883:CRh

[4177] Anonymous. Communications received. Scientific American, 48 (6):91, February 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n6/pdf/scientificamerican02101883-
91b.pdf.

Anonymous:1883:Co

[4178] Anonymous. Correspondence. Scientific American, 48(7):101– 102, February 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n7/pdf/scientificamerican02171883-
101.pdf.

Anonymous:1883:IRIq

[4179] Anonymous. Inventions: Recent inventions. Scientific American, 48(7):105, February 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n7/pdf/scientificamerican02171883-
105.pdf.

Anonymous:1883:BPu

[4180] Anonymous. Business and personal. Scientific American, 48(7): 106–107, February 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n7/pdf/scientificamerican02171883-
106d.pdf.

Anonymous:1883:NQu

[4181] Anonymous. Notes and queries. Scientific American, 48(7):107, February 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n7/pdf/scientificamerican02171883-107.pdf.

Anonymous:1883:IIq

[4182] Anonymous. Index of inventions. Scientific American, 48(7): 107–108, February 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n7/pdf/scientificamerican02171883-
107a.pdf.

Anonymous:1883:PNPj

[4183] Anonymous. Patent notes: Patents. Scientific American, 48 (7):108, February 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n7/pdf/scientificamerican02171883-
108.pdf.

Anonymous:1883:VIE

[4184] Anonymous. Vienna International Electric Exhibition. Scientific American, 48(7):96, February 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n7/pdf/scientificamerican02171883-
96.pdf.

Anonymous:1883:PT

[4185] Anonymous. The problem of the telephone. Scientific American, 48(7):96, February 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n7/pdf/scientificamerican02171883-
96a.pdf.

Anonymous:1883:FMD

[4186] Anonymous. A four-mile deposit of iron ore. Scientific American, 48(7):96, February 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n7/pdf/scientificamerican02171883-
96b.pdf.

Anonymous:1883:WA

[4187] Anonymous. What is an artist? Scientific American, 48(7):97, February 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n7/pdf/scientificamerican02171883-97.pdf.

Anonymous:1883:SB

[4188] Anonymous. The Star of Bethlehem. Scientific American, 48 (7):97, February 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n7/pdf/scientificamerican02171883-
97a.pdf.

Anonymous:1883:PYP

[4189] Anonymous. The past year in the Patent Office. Scientific American, 48(7):97, February 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n7/pdf/scientificamerican02171883-
97b.pdf.

Anonymous:1883:ASS

[4190] Anonymous. An ammonia still for small gas works. Scientific American, 48(7):98, February 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n7/pdf/scientificamerican02171883-
98.pdf.

Anonymous:1883:EMI

[4191] Anonymous. Electric motors with inductors containing no iron. Scientific American, 48(7):98, February 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v48/n7/pdf/scientificamerican02171883-
98a.pdf.

Anonymous:1883:ISF

[4192] Anonymous. Improved sponge filters. Scientific American, 48 (7):98–99, February 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n7/pdf/scientificamerican02171883-
98b.pdf.

Anonymous:1883:TCL

[4193] Anonymous. Teera cotta lumber. Scientific American, 48(7):100, February 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n7/pdf/scientificamerican02171883-100.pdf.

Anonymous:1883:FSG

[4194] Anonymous. Flexible self-gauging faucet. Scientific American, 48(7):102, February 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n7/pdf/scientificamerican02171883-
102.pdf.

Anonymous:1883:NOO

[4195] Anonymous. The new Otto one-horse power silent gas engine. Scientific American, 48(7):102, February 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v48/n7/pdf/scientificamerican02171883-
102a.pdf.

Anonymous:1883:GGW

[4196] Anonymous. A glass globe from which three hundred watch crystals were cut. Scientific American, 48(7):103, February 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v48/
n7/pdf/scientificamerican02171883-103.pdf.

Anonymous:1883:ACP

[4197] Anonymous. The automaton chess player. Scientific American, 48 (7):103–104, February 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n7/pdf/scientificamerican02171883-
103a.pdf.

Anonymous:1883:EIr

[4198] Anonymous. Engineering inventions. Scientific American, 48(7): 106, February 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n7/pdf/scientificamerican02171883-
106.pdf.

Anonymous:1883:MIbm

[4199] Anonymous. Mechanical inventions. Scientific American, 48(7): 106, February 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n7/pdf/scientificamerican02171883-
106a.pdf.

Anonymous:1883:AIt

[4200] Anonymous. Agricultural inventions. Scientific American, 48(7): 106, February 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n7/pdf/scientificamerican02171883-
106b.pdf.

Anonymous:1883:MIbn

[4201] Anonymous. Miscellaneous inventions. Scientific American, 48 (7):106, February 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n7/pdf/scientificamerican02171883-
106c.pdf.

Anonymous:1883:Cp

[4202] Anonymous. Correspondence. Scientific American, 48(8):117, February 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n8/pdf/scientificamerican02241883-117.pdf.

Anonymous:1883:IRIr

[4203] Anonymous. Inventions: Recent inventions. Scientific American, 48(8):121, February 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n8/pdf/scientificamerican02241883-
121.pdf.

Anonymous:1883:BPv

[4204] Anonymous. Business and personal. Scientific American, 48(8): 122, February 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n8/pdf/scientificamerican02241883-
122c.pdf.

Anonymous:1883:NQv

[4205] Anonymous. Notes and queries. Scientific American, 48(8):123, February 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n8/pdf/scientificamerican02241883-123.pdf.

Anonymous:1883:IIr

[4206] Anonymous. Index of inventions. Scientific American, 48(8): 123–124, February 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n8/pdf/scientificamerican02241883-
123a.pdf.

Anonymous:1883:PNPk

[4207] Anonymous. Patent notes: Patents. Scientific American, 48 (8):124, February 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n8/pdf/scientificamerican02241883-
124.pdf.

Anonymous:1883:ACS

[4208] Anonymous. Apparatus for coaling ships. Scientific American, 48(8):111, February 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n8/pdf/scientificamerican02241883-
111.pdf.

Anonymous:1883:MECa

[4209] Anonymous. Machine for excavating the channel tunnel. Scientific American, 48(8):111, February 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v48/n8/pdf/scientificamerican02241883-
111a.pdf.

Anonymous:1883:PIW

[4210] Anonymous. Projects for increasing the water supply of New York City. Scientific American, 48(8):112, February 24, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v48/
n8/pdf/scientificamerican02241883-112.pdf.

Anonymous:1883:TTC

[4211] Anonymous. Torsion tests of cast steel. Scientific American, 48 (8):112, February 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n8/pdf/scientificamerican02241883-
112a.pdf.

Anonymous:1883:PMJ

[4212] Anonymous. Progress of music in Japan. Scientific American, 48(8):112, February 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n8/pdf/scientificamerican02241883-
112b.pdf.

Anonymous:1883:SFD

[4213] Anonymous. Shall failure to develop forfeit patent rights? Scientific American, 48(8):112–113, February 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v48/n8/pdf/scientificamerican02241883-
112c.pdf.

Stockwell:1883:ZW

[4214] G. A. Stockwell. The zebra wolf. Scientific American, 48(8):113, February 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n8/pdf/scientificamerican02241883-113.pdf.

Anonymous:1883:NKF

[4215] Anonymous. New key fastener. Scientific American, 48(8):114, February 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n8/pdf/scientificamerican02241883-114.pdf.

Anonymous:1883:ITC

[4216] Anonymous. Improved thill coupling. Scientific American, 48(8): 114–115, February 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n8/pdf/scientificamerican02241883-
114a.pdf.

Anonymous:1883:ISS

[4217] Anonymous. Improved seed sower. Scientific American, 48(8): 115, February 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n8/pdf/scientificamerican02241883-
115.pdf.

Hopkins:1883:LI

[4218] Geo. M. Hopkins. Lantern illumination. Scientific American, 48 (8):118, February 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n8/pdf/scientificamerican02241883-
118.pdf.

Anonymous:1883:NBT

[4219] Anonymous. New bag and twine holder. Scientific American, 48(8):118, February 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n8/pdf/scientificamerican02241883-
118a.pdf.

Anonymous:1883:AH

[4220] Anonymous. The æeolian harp. Scientific American, 48(8):119– 120, February 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n8/pdf/scientificamerican02241883-
119.pdf.

Anonymous:1883:EIs

[4221] Anonymous. Engineering inventions. Scientific American, 48(8): 122, February 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n8/pdf/scientificamerican02241883-
122.pdf.

Anonymous:1883:MIbo

[4222] Anonymous. Mechanical inventions. Scientific American, 48(8): 122, February 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n8/pdf/scientificamerican02241883-
122a.pdf.

Anonymous:1883:MIbp

[4223] Anonymous. Miscellaneous inventions. Scientific American, 48 (8):122, February 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n8/pdf/scientificamerican02241883-
122b.pdf.

Anonymous:1883:Cq

[4224] Anonymous. Correspondence. Scientific American, 48(9):133, March 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n9/pdf/scientificamerican03031883-133.pdf.

Anonymous:1883:IRIs

[4225] Anonymous. Inventions: Recent inventions. Scientific American, 48(9):137, March 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n9/pdf/scientificamerican03031883-
137.pdf.

Anonymous:1883:BPw

[4226] Anonymous. Business and personal. Scientific American, 48(9):138, March 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n9/pdf/scientificamerican03031883-138d.pdf.

Anonymous:1883:NQw

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n9/pdf/scientificamerican03031883-139.pdf.

Anonymous:1883:IIs

[4228] Anonymous. Index of inventions. Scientific American, 48(9):139–140, March 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n9/pdf/scientificamerican03031883-139a.pdf.

Anonymous:1883:PNPl

[4229] Anonymous. Patent notes: Patents. Scientific American, 48(9):140, March 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n9/pdf/scientificamerican03031883-140.pdf.

Anonymous:1883:MDP

[4230] Anonymous. Monster dredge for the Panama Canal. Scientific American, 48(9):127, March 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n9/pdf/scientificamerican03031883-
127.pdf.

Anonymous:1883:CRCb

[4231] Anonymous. Curiosities of the railway census. Scientific American, 48(9):128, March 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n9/pdf/scientificamerican03031883-
128.pdf.

Anonymous:1883:WBC

[4232] Anonymous. The wringer before Congress. Scientific American, 48(9):128, March 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n9/pdf/scientificamerican03031883-
128a.pdf.

Anonymous:1883:FEa

[4233] Anonymous. Florida exploration. Scientific American, 48(9):128, March 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n9/pdf/scientificamerican03031883-128b.pdf.

Anonymous:1883:APMb

[4234] Anonymous. Aspects of the planets for March. Scientific American, 48(9):129, March 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n9/pdf/scientificamerican03031883-
129.pdf.

Anonymous:1883:RF

[4235] Anonymous. Resharpening files. Scientific American, 48(9):129, March 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n9/pdf/scientificamerican03031883-129a.pdf.

Anonymous:1883:IWb

[4236] Anonymous. Improved windmill. Scientific American, 48(9):130, March 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n9/pdf/scientificamerican03031883-130.pdf.

Anonymous:1883:FFC

[4237] Anonymous. Fireproof floor and ceiling. Scientific American, 48 (9):130–131, March 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n9/pdf/scientificamerican03031883-
130a.pdf.

Anonymous:1883:ICH

[4238] Anonymous. Improvement in cutter heads. Scientific American, 48(9):131–132, March 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n9/pdf/scientificamerican03031883-
131.pdf.

Anonymous:1883:IMR

[4239] Anonymous. Improved method of raising cream from milk. Scientific American, 48(9):132, March 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n9/pdf/scientificamerican03031883-
132.pdf.

Anonymous:1883:UDP

[4240] Anonymous. Upright drill press used as a boring lathe. Scientific American, 48(9):134, March 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n9/pdf/scientificamerican03031883-
134.pdf.

Anonymous:1883:UEW

[4241] Anonymous. Underground electric wires. Scientific American, 48 (9):134, March 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n9/pdf/scientificamerican03031883-
134a.pdf.

Anonymous:1883:ICT

[4242] Anonymous. Improved car truck. Scientific American, 48(9):135–136, March 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n9/pdf/scientificamerican03031883-135.pdf.

Anonymous:1883:RTT

[4243] Anonymous. The Reis–Thompson telephone receiver. Scientific American, 48(9):136, March 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v48/n9/pdf/scientificamerican03031883-
136.pdf.

Anonymous:1883:EIt

[4244] Anonymous. Engineering inventions. Scientific American, 48(9):138, March 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n9/pdf/scientificamerican03031883-138.pdf.

Anonymous:1883:MIbq

[4245] Anonymous. Mechanical inventions. Scientific American, 48(9):138, March 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n9/pdf/scientificamerican03031883-138a.pdf.

Anonymous:1883:AIu

[4246] Anonymous. Agricultural inventions. Scientific American, 48(9):138, March 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n9/pdf/scientificamerican03031883-138b.pdf.

Anonymous:1883:MIbr

[4247] Anonymous. Miscellaneous inventions. Scientific American, 48(9):138, March 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v48/n9/pdf/scientificamerican03031883-138c.pdf.

Anonymous:1883:IRIt

[4248] Anonymous. Inventions: Recent inventions. Scientific American, 49(1): 8, July 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n1/pdf/scientificamerican07071883-8.pdf.

Anonymous:1883:Cv

[4249] Anonymous. Correspondence. Scientific American, 49(1):8, July 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n1/pdf/scientificamerican07071883-8a.pdf.

Anonymous:1883:BPx

[4250] Anonymous. Business and personal. Scientific American, 49(1):10, July 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n1/pdf/scientificamerican07071883-10.pdf.

Anonymous:1883:BNBl

[4251] Anonymous. Books: New books and publications. Scientific American, 49(1):10, July 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n1/pdf/scientificamerican07071883-
10a.pdf.

Anonymous:1883:NQx

[4252] Anonymous. Notes and queries. Scientific American, 49(1):10–11, July 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n1/pdf/scientificamerican07071883-10b.pdf.

Anonymous:1883:IIt

[4253] Anonymous. Index of inventions. Scientific American, 49(1):11–12, July 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n1/pdf/scientificamerican07071883-11a.pdf.

Anonymous:1883:PNPm

[4254] Anonymous. Patent notes: Patents. Scientific American, 49(1):12, July 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n1/pdf/scientificamerican07071883-12.pdf.

Anonymous:1883:RBa

[4255] Anonymous. The Rondout Bridge. Scientific American, 49(1):1–2, July 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n1/pdf/scientificamerican07071883-1.pdf.

Anonymous:1883:TSEb

[4256] Anonymous. The total solar eclipse of the 6th of May. Scientific American, 49(1):2–3, July 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n1/pdf/scientificamerican07071883-
2.pdf.

Anonymous:1883:HSTa

[4257] Anonymous. How screws are threaded. Scientific American, 49(1):3, July 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n1/pdf/scientificamerican07071883-3.pdf.

Anonymous:1883:DC

[4258] Anonymous. D’arrest’s Comet. Scientific American, 49(1):3–4, July 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n1/pdf/scientificamerican07071883-3a.pdf.

Anonymous:1883:CVD

[4259] Anonymous. Chemical vaporizer and deodorizer. Scientific American, 49(1):4, July 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n1/pdf/scientificamerican07071883-
4.pdf.

Anonymous:1883:IHP

[4260] Anonymous. Improved hop press. Scientific American, 49(1):4–5, July 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n1/pdf/scientificamerican07071883-4a.pdf.

Hervier:1883:RTB

[4261] L. Hervier. Radial tubular boiler. Scientific American, 49(1):5–6, July 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n1/pdf/scientificamerican07071883-5.pdf.

Anonymous:1883:BC

[4262] Anonymous. A beaver colony. Scientific American, 49(1):7, July 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n1/pdf/scientificamerican07071883-7.pdf.

Anonymous:1883:EIu

[4263] Anonymous. Engineering inventions. Scientific American, 49(1):9, July 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n1/pdf/scientificamerican07071883-9.pdf.

Anonymous:1883:MIbs

[4264] Anonymous. Mechanical inventions. Scientific American, 49(1):9, July 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n1/pdf/scientificamerican07071883-9a.pdf.

Anonymous:1883:AIv

[4265] Anonymous. Agricultural inventions. Scientific American, 49(1):9, July 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n1/pdf/scientificamerican07071883-9b.pdf.

Anonymous:1883:MIbt

[4266] Anonymous. Miscellaneous inventions. Scientific American, 49(1):9, July 7, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n1/pdf/scientificamerican07071883-9c.pdf.

Anonymous:1883:CRi

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n1/pdf/scientificamerican07071883-11.pdf.

Anonymous:1883:Cw

[4268] Anonymous. Correspondence. Scientific American, 49(10):149– 150, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
149.pdf.

Anonymous:1883:BPy

[4269] Anonymous. Business and personal. Scientific American, 49 (10):155, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
155.pdf.

Anonymous:1883:BNBm

[4270] Anonymous. Books: New books and publications. Scientific American, 49(10):155, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
155a.pdf.

Anonymous:1883:NQy

[4271] Anonymous. Notes and queries. Scientific American, 49(10): 155, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
155b.pdf.

Anonymous:1883:IIu

[4272] Anonymous. Index of inventions. Scientific American, 49(10): 155–156, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
155d.pdf.

Anonymous:1883:PNPn

[4273] Anonymous. Patent notes: Patents. Scientific American, 49 (10):157, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
157.pdf.

Anonymous:1883:MCBa

[4274] Anonymous. Mechanical criticisms, barbs in beef, and more. Scientific American, 49(10):143, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
143.pdf.

Anonymous:1883:MSPa

[4275] Anonymous. Motors for small power. Scientific American, 49 (10):143, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
143a.pdf.

Anonymous:1883:BDW

[4276] Anonymous. Boring deep wells by water pressure. Scientific American, 49(10):144, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
144.pdf.

Anonymous:1883:PR

[4277] Anonymous. The potato rot. Scientific American, 49(10):144, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n10/pdf/scientificamerican09081883-144a.pdf.

Anonymous:1883:CDa

[4278] Anonymous. Chemistry for digestion. Scientific American, 49(10): 144–145, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
144b.pdf.

Hovey:1883:AAAa

[4279] H. C. Hovey. The American Association for the Advancement of Science. Scientific American, 49(10):145, September 8, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v49/
n10/pdf/scientificamerican09081883-145.pdf.

Anonymous:1883:AWS

[4280] Anonymous. Automatic water still. Scientific American, 49(10): 146, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
146.pdf.

Anonymous:1883:IRP

[4281] Anonymous. Improved rotary plow. Scientific American, 49(10): 146, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
146a.pdf.

Anonymous:1883:TP

[4282] Anonymous.  Tree planter.  Scientific American, 49(10):146–

147, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
146b.pdf.

Anonymous:1883:LAHb

[4283] Anonymous. Luminous attachment for harness. Scientific American, 49(10):147, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
147.pdf.

Anonymous:1883:CSH

[4284] Anonymous. Circular saw for hot iron. Scientific American, 49 (10):147, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
147a.pdf.

Anonymous:1883:ETC

[4285] Anonymous. Ear trumpet and cane handle. Scientific American, 49(10):147, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
147b.pdf.

Anonymous:1883:US

[4286] Anonymous. The University of Strasbourg. Scientific American, 49(10):148, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
148.pdf.

Anonymous:1883:SWPa

[4287] Anonymous. “Storing wind power” revised and corrected. Scientific American, 49(10):148, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
148a.pdf.

Anonymous:1883:FEM

[4288] Anonymous. The first electric motor. Scientific American, 49(10): 150–151, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
150.pdf.

Anonymous:1883:HHSa

[4289] Anonymous. Hammer-headed shark. Scientific American, 49(10): 151–153, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
151.pdf.

Anonymous:1883:EIv

[4290] Anonymous. Engineering inventions. Scientific American, 49(10): 154, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
154.pdf.

Anonymous:1883:MIbu

[4291] Anonymous. Mechanical inventions. Scientific American, 49(10): 154, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
154a.pdf.

Anonymous:1883:AIw

[4292] Anonymous. Agricultural inventions. Scientific American, 49 (10):154, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
154b.pdf.

Anonymous:1883:MIbv

[4293] Anonymous. Miscellaneous inventions. Scientific American, 49 (10):154, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
154c.pdf.

Anonymous:1883:CRj

[4294] Anonymous. Communications received. Scientific American, 49 (10):155, September 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n10/pdf/scientificamerican09081883-
155c.pdf.

Anonymous:1883:Cx

[4295] Anonymous. Correspondence. Scientific American, 49(11):165, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n11/pdf/scientificamerican09151883-165.pdf.

Anonymous:1883:IRIu

[4296] Anonymous. Inventions: Recent inventions. Scientific American, 49(11):168, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
168.pdf.

Anonymous:1883:BPz

[4297] Anonymous. Business and personal. Scientific American, 49 (11):170, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
170.pdf.

Anonymous:1883:NQz

[4298] Anonymous. Notes and queries. Scientific American, 49(11): 170, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
170a.pdf.

Anonymous:1883:BNBn

[4299] Anonymous. Books: New books and publications. Scientific American, 49(11):170, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
170b.pdf.

Anonymous:1883:IIv

[4300] Anonymous. Index of inventions. Scientific American, 49(11): 170–171, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
170c.pdf.

Anonymous:1883:PNPo

[4301] Anonymous. Patent notes: Patents. Scientific American, 49 (11):172, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
172.pdf.

Anonymous:1883:BCS

[4302] Anonymous. Blake’s compound steam pumps. Scientific American, 49(11):159, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
159.pdf.

Anonymous:1883:OIB

[4303] Anonymous. The Observatory of the International Bureau of Weights and Measures. Scientific American, 49(11):159, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n11/pdf/scientificamerican09151883-159a.pdf.

Anonymous:1883:PL

[4304] Anonymous. Petroleum for light. Scientific American, 49(11): 160, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
160.pdf.

Anonymous:1883:EMb

[4305] Anonymous. Exports of manufactures. Scientific American, 49(11): 160–161, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
160a.pdf.

Anonymous:1883:ES

[4306] Anonymous. Economy of steam. Scientific American, 49(11): 161, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
161.pdf.

Anonymous:1883:PNBa

[4307] Anonymous. Passage of the new British Patent Law. Scientific American, 49(11):161, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
161a.pdf.

Anonymous:1883:SSSc

[4308] Anonymous. Scientific shows — sword swallowers. Scientific American, 49(11):162–163, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
162.pdf.

Anonymous:1883:RSW

[4309] Anonymous. Rocker seat worker. Scientific American, 49(11): 163, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
163.pdf.

Anonymous:1883:DPO

[4310] Anonymous. Device for protecting oil tanks from lightning. Scientific American, 49(11):163, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
163a.pdf.

Anonymous:1883:IFPb

[4311] Anonymous. Improved frying pan. Scientific American, 49(11): 163, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
163b.pdf.

Anonymous:1883:FFL

[4312] Anonymous. Force feed lubricator. Scientific American, 49(11): 163, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
163c.pdf.

Anonymous:1883:NUSa

[4313] Anonymous. New umbrella support. Scientific American, 49 (11):166, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
166.pdf.

Anonymous:1883:ADV

[4314] Anonymous. Apparatus for disintegrating vegetable substances. Scientific American, 49(11):166, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
166a.pdf.

Anonymous:1883:IPG

[4315] Anonymous. Improved platform gear for wagons. Scientific American, 49(11):166, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
166b.pdf.

Anonymous:1883:NBSa

[4316] Anonymous. New balanced slide valve. Scientific American, 49 (11):166, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
166c.pdf.

Anonymous:1883:DPC

[4317] Anonymous. The Demoiselle and Peacock crane. Scientific American, 49(11):167, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
167.pdf.

Anonymous:1883:EIw

[4318] Anonymous. Engineering inventions. Scientific American, 49(11): 169, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
169.pdf.

Anonymous:1883:MIbw

[4319] Anonymous. Mechanical inventions. Scientific American, 49(11): 169, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
169a.pdf.

Anonymous:1883:AIx

[4320] Anonymous. Agricultural inventions. Scientific American, 49 (11):169, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
169b.pdf.

Anonymous:1883:MIbx

[4321] Anonymous. Miscellaneous inventions. Scientific American, 49 (11):169, September 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n11/pdf/scientificamerican09151883-
169c.pdf.

Anonymous:1883:Cy

[4322] Anonymous. Correspondence. Scientific American, 49(12):181– 182, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
181.pdf.

Anonymous:1883:IRIv

[4323] Anonymous. Inventions: Recent inventions. Scientific American, 49(12):185, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
185.pdf.

Anonymous:1883:BPba

[4324] Anonymous. Business and personal. Scientific American, 49 (12):187, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
187.pdf.

Anonymous:1883:BNBo

[4325] Anonymous. Books: New books and publications. Scientific American, 49(12):187, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
187a.pdf.

Anonymous:1883:NQba

[4326] Anonymous. Notes and queries. Scientific American, 49(12): 187, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
187b.pdf.

Anonymous:1883:IIw

[4327] Anonymous. Index of inventions. Scientific American, 49(12): 187–188, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
187c.pdf.

Anonymous:1883:PNPp

[4328] Anonymous. Patent notes: Patents. Scientific American, 49 (12):188, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
188.pdf.

Anonymous:1883:SWPb

[4329] Anonymous. The storage of wind power. Scientific American, 49 (12):176, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
176.pdf.

Anonymous:1883:GBS

[4330] Anonymous. The great bridge as a spectacle. Scientific American, 49(12):176, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
176a.pdf.

Anonymous:1883:Fc

[4331] Anonymous. The foreman. Scientific American, 49(12):176, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n12/pdf/scientificamerican09221883-176b.pdf.

Anonymous:1883:MSM

[4332] Anonymous. Machine shop management. Scientific American, 49 (12):176, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
176c.pdf.

Anonymous:1883:SBI

[4333] Anonymous. A sham boiler inspection. Scientific American, 49 (12):177, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
177.pdf.

Anonymous:1883:MG

[4334] Anonymous. The mildew of the grape. Scientific American, 49 (12):177–178, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
177a.pdf.

Anonymous:1883:IHG

[4335] Anonymous. Improvement in heating grates. Scientific American, 49(12):178, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
178.pdf.

Anonymous:1883:ITF

[4336] Anonymous. Improved textile filter. Scientific American, 49 (12):178, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
178a.pdf.

Anonymous:1883:IBC

[4337] Anonymous. Improved billiard cushion. Scientific American, 49 (12):179, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
179.pdf.

Anonymous:1883:CLH

[4338] Anonymous. A convenient lever hand cart. Scientific American, 49(12):179, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
179a.pdf.

Anonymous:1883:IFCb

[4339] Anonymous. Improved feeder for cotton gins. Scientific American, 49(12):179, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
179b.pdf.

Anonymous:1883:NCB

[4340] Anonymous. Novel candlestick, broiler, etc. Scientific American, 49(12):179, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
179c.pdf.

Anonymous:1883:NBP

[4341] Anonymous. New bridge at Pittsburg, PA. Scientific American, 49(12):180, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
180.pdf.

Anonymous:1883:NHPa

[4342] Anonymous. New hand power rock drill. Scientific American, 49 (12):182, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
182.pdf.

Anonymous:1883:MDA

[4343] Anonymous. Metamorphosis of the deer’s antlers. Scientific American, 49(12):183, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
183.pdf.

Anonymous:1883:CGN

[4344] Anonymous. Completion of the Great Northern Pacific Railroad. Scientific American, 49(12):184, September 22, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v49/
n12/pdf/scientificamerican09221883-184.pdf.

Anonymous:1883:POR

[4345] Anonymous. The Patent Office Report for 1883. Scientific American, 49(12):184, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
184a.pdf.

Anonymous:1883:EIx

[4346] Anonymous. Engineering inventions. Scientific American, 49(12): 186, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
186.pdf.

Anonymous:1883:MIby

[4347] Anonymous. Mechanical inventions. Scientific American, 49(12): 186, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
186a.pdf.

Anonymous:1883:AIy

[4348] Anonymous. Agricultural inventions. Scientific American, 49 (12):186, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
186b.pdf.

Anonymous:1883:MIbz

[4349] Anonymous. Miscellaneous inventions. Scientific American, 49 (12):186, September 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n12/pdf/scientificamerican09221883-
186c.pdf.

Anonymous:1883:Cz

[4350] Anonymous. Correspondence. Scientific American, 49(13):197, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n13/pdf/scientificamerican09291883-197.pdf.

Anonymous:1883:IRIw

[4351] Anonymous. Inventions: Recent inventions. Scientific American, 49(13):201, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n13/pdf/scientificamerican09291883-
201.pdf.

Anonymous:1883:BPbb

[4352] Anonymous. Business and personal. Scientific American, 49 (13):203, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n13/pdf/scientificamerican09291883-
203.pdf.

Anonymous:1883:BNBp

[4353] Anonymous. Books: New books and publications. Scientific American, 49(13):203, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n13/pdf/scientificamerican09291883-
203a.pdf.

Anonymous:1883:NQbb

[4354] Anonymous. Notes and queries. Scientific American, 49(13): 203, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n13/pdf/scientificamerican09291883-
203b.pdf.

Anonymous:1883:IIx

[4355] Anonymous. Index of inventions. Scientific American, 49(13): 203–204, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n13/pdf/scientificamerican09291883-
203c.pdf.

Anonymous:1883:PNPq

[4356] Anonymous. Patent notes: Patents. Scientific American, 49 (13):204, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n13/pdf/scientificamerican09291883-
204.pdf.

Anonymous:1883:EDA

[4357] Anonymous. Excavating and dredging apparatus. Scientific American, 49(13):191–192, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n13/pdf/scientificamerican09291883-
191.pdf.

Anonymous:1883:LE

[4358] Anonymous. “Labor and education”. Scientific American, 49(13): 192–193, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n13/pdf/scientificamerican09291883-
192.pdf.

Anonymous:1883:ASA

[4359] Anonymous. Automatic safety appliances. Scientific American, 49 (13):193, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n13/pdf/scientificamerican09291883-
193.pdf.

Anonymous:1883:APO

[4360] Anonymous. Aspects of the planets for October. Scientific American, 49(13):193–194, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n13/pdf/scientificamerican09291883-
193a.pdf.

Anonymous:1883:IPb

[4361] Anonymous. Improved plow. Scientific American, 49(13):194, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n13/pdf/scientificamerican09291883-194.pdf.

Anonymous:1883:WUU

[4362] Anonymous. The Western Union underground systems. Scientific American, 49(13):194, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v49/n13/pdf/scientificamerican09291883-
194a.pdf.

Anonymous:1883:ISBa

[4363] Anonymous. Improved sash balance. Scientific American, 49 (13):195, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n13/pdf/scientificamerican09291883-
195.pdf.

Anonymous:1883:NSW

[4364] Anonymous. New Spanish war steamers. Scientific American, 49 (13):195, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n13/pdf/scientificamerican09291883-
195a.pdf.

Anonymous:1883:ISV

[4365] Anonymous. Improved safety valve. Scientific American, 49(13): 195–196, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n13/pdf/scientificamerican09291883-
195b.pdf.

Anonymous:1883:SIHb

[4366] Anonymous. Sad-iron holder. Scientific American, 49(13):198, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n13/pdf/scientificamerican09291883-198.pdf.

Anonymous:1883:IFEc

[4367] Anonymous. Improved fire escape. Scientific American, 49(13): 198, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n13/pdf/scientificamerican09291883-
198a.pdf.

Anonymous:1883:IDP

[4368] Anonymous. Improvement in dust-pans. Scientific American, 49 (13):198, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n13/pdf/scientificamerican09291883-
198b.pdf.

Anonymous:1883:IFb

[4369] Anonymous. Improved filter. Scientific American, 49(13):198– 199, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n13/pdf/scientificamerican09291883-
198c.pdf.

Anonymous:1883:PCC

[4370] Anonymous. Pyrethrum, or Chrysanthemum Corymbosum. Scientific American, 49(13):199–200, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v49/n13/pdf/scientificamerican09291883-
199.pdf.

Anonymous:1883:EIy

[4371] Anonymous. Engineering inventions. Scientific American, 49(13): 202, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n13/pdf/scientificamerican09291883-
202.pdf.

Anonymous:1883:MIca

[4372] Anonymous. Mechanical inventions. Scientific American, 49(13): 202, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n13/pdf/scientificamerican09291883-
202a.pdf.

Anonymous:1883:AIz

[4373] Anonymous. Agricultural inventions. Scientific American, 49 (13):202, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n13/pdf/scientificamerican09291883-
202b.pdf.

Anonymous:1883:MIcb

[4374] Anonymous. Miscellaneous inventions. Scientific American, 49 (13):202, September 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n13/pdf/scientificamerican09291883-
202c.pdf.

Anonymous:1883:Cba

[4375] Anonymous. Correspondence. Scientific American, 49(14):213, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n14/pdf/scientificamerican10061883-213.pdf.

Anonymous:1883:IRIx

[4376] Anonymous. Inventions: Recent invention. Scientific American, 49(14):217, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n14/pdf/scientificamerican10061883-
217.pdf.

Anonymous:1883:BNBq

[4377] Anonymous. Books: New books and publications. Scientific American, 49(14):218, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n14/pdf/scientificamerican10061883-
218d.pdf.

Anonymous:1883:BPbc

[4378] Anonymous. Business and personal. Scientific American, 49 (14):218–219, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n14/pdf/scientificamerican10061883-
218e.pdf.

Anonymous:1883:NQbc

[4379] Anonymous. Notes and queries. Scientific American, 49(14):219, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n14/pdf/scientificamerican10061883-219.pdf.

Anonymous:1883:IIy

[4380] Anonymous. Index of inventions. Scientific American, 49(14):219–220, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n14/pdf/scientificamerican10061883-219a.pdf.

Anonymous:1883:PNPr

[4381] Anonymous. Patent notes: Patents. Scientific American, 49(14):220, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n14/pdf/scientificamerican10061883-220.pdf.

Anonymous:1883:IDD

[4382] Anonymous. Improved depositing dry dock. Scientific American, 49(14):207, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n14/pdf/scientificamerican10061883-
207.pdf.

Anonymous:1883:FPa

[4383] Anonymous. Forging by pressure. Scientific American, 49(14):208, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n14/pdf/scientificamerican10061883-208.pdf.

Anonymous:1883:GMR

[4384] Anonymous. The Green Mountain Railway. Scientific American, 49(14):208, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n14/pdf/scientificamerican10061883-
208a.pdf.

Anonymous:1883:SIR

[4385] Anonymous. Self-imposed risks. Scientific American, 49(14):208, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n14/pdf/scientificamerican10061883-208b.pdf.

Anonymous:1883:CAA

[4386] Anonymous. Carbonic acid in the air. Scientific American, 49 (14):208–209, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n14/pdf/scientificamerican10061883-
208c.pdf.

Anonymous:1883:OSa

[4387] Anonymous. The oscillations of the sea. Scientific American, 49 (14):209–210, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n14/pdf/scientificamerican10061883-
209.pdf.

Anonymous:1883:IQF

[4388] Anonymous. Improved quilting frame. Scientific American, 49(14):210, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n14/pdf/scientificamerican10061883-210.pdf.

Anonymous:1883:SLR

[4389] Anonymous. Surveyor’s leveling rod. Scientific American, 49 (14):210–211, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n14/pdf/scientificamerican10061883-
210a.pdf.

Anonymous:1883:CAb

[4390] Anonymous. Car axle. Scientific American, 49(14):211, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n14/pdf/scientificamerican10061883-211.pdf.

Anonymous:1883:SSE

[4391] Anonymous. Safety stop for elevators. Scientific American, 49(14):211, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n14/pdf/scientificamerican10061883-211a.pdf.

Anonymous:1883:LB

[4392] Anonymous. Life boat. Scientific American, 49(14):211–212, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n14/pdf/scientificamerican10061883-211b.pdf.

Anonymous:1883:NYT

[4393] Anonymous. New York Terminus of the N.Y., W.S. & B.R. Scientific American, 49(14):212, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n14/pdf/scientificamerican10061883-
212.pdf.

Anonymous:1883:SFE

[4394] Anonymous. Sketches at the Fisheries Exhibition. Scientific American, 49(14):214–215, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n14/pdf/scientificamerican10061883-
214.pdf.

Anonymous:1883:LF

[4395] Anonymous. The lancelet fish. Scientific American, 49(14):215, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n14/pdf/scientificamerican10061883-215.pdf.

Anonymous:1883:SNM

[4396] Anonymous. The star nosed mole. Scientific American, 49(14):215–216, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n14/pdf/scientificamerican10061883-215a.pdf.

Anonymous:1883:ICb

[4397] Anonymous. Improved cultivator. Scientific American, 49(14):216, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n14/pdf/scientificamerican10061883-216.pdf.

Anonymous:1883:EIz

[4398] Anonymous. Engineering invention. Scientific American, 49(14):218, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n14/pdf/scientificamerican10061883-218.pdf.

Anonymous:1883:MIcc

[4399] Anonymous. Mechanical inventions. Scientific American, 49(14):218, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n14/pdf/scientificamerican10061883-218a.pdf.

Anonymous:1883:AIba

[4400] Anonymous. Agricultural inventions. Scientific American, 49(14):218, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n14/pdf/scientificamerican10061883-218b.pdf.

Anonymous:1883:MIcd

[4401] Anonymous. Miscellaneous inventions. Scientific American, 49(14):218, October 6, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n14/pdf/scientificamerican10061883-218c.pdf.

Anonymous:1883:Cbb

[4402] Anonymous. Correspondence. Scientific American, 49(15):229, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n15/pdf/scientificamerican10131883-229.pdf.

Anonymous:1883:BNBr

[4403] Anonymous. Books: New books and publications. Scientific American, 49(15):234, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n15/pdf/scientificamerican10131883-
234d.pdf.

Anonymous:1883:BPbd

[4404] Anonymous. Business and personal. Scientific American, 49(15): 234–235, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n15/pdf/scientificamerican10131883-
234e.pdf.

Anonymous:1883:NQbd

[4405] Anonymous. Notes and queries. Scientific American, 49(15):235, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n15/pdf/scientificamerican10131883-235.pdf.

Anonymous:1883:IIz

[4406] Anonymous. Index of inventions. Scientific American, 49(15): 235–236, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n15/pdf/scientificamerican10131883-
235a.pdf.

Anonymous:1883:PNPs

[4407] Anonymous. Patent notes: Patents. Scientific American, 49 (15):236, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n15/pdf/scientificamerican10131883-
236.pdf.

Anonymous:1883:EDE

[4408] Anonymous. Edison dynamo electric light machine. Scientific American, 49(15):223, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n15/pdf/scientificamerican10131883-
223.pdf.

Anonymous:1883:OG

[4409] Anonymous. Obach’s galvanometers. Scientific American, 49 (15):223, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n15/pdf/scientificamerican10131883-
223a.pdf.

Anonymous:1883:NHPb

[4410] Anonymous. Natural history in public schools. Scientific American, 49(15):224, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n15/pdf/scientificamerican10131883-
224.pdf.

Anonymous:1883:MSPb

[4411] Anonymous. The making of steel pens. Scientific American, 49 (15):224, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n15/pdf/scientificamerican10131883-
224a.pdf.

Hovey:1883:MC

[4412] H. C. Hovey. Marengo cavern. Scientific American, 49(15):225, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n15/pdf/scientificamerican10131883-225.pdf.

Anonymous:1883:PH

[4413] Anonymous. Petroleum for heat. Scientific American, 49(15): 225–226, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n15/pdf/scientificamerican10131883-
225a.pdf.

Anonymous:1883:ACC

[4414] Anonymous. Automatic circuit closer. Scientific American, 49 (15):226, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n15/pdf/scientificamerican10131883-
226.pdf.

Anonymous:1883:Cbc

[4415] Anonymous. Counterbalance. Scientific American, 49(15):226, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n15/pdf/scientificamerican10131883-226a.pdf.

Anonymous:1883:NSM

[4416] Anonymous. Novel sewing machine shuttle. Scientific American, 49(15):226, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n15/pdf/scientificamerican10131883-
226b.pdf.

Anonymous:1883:FA

[4417] Anonymous. Fire arm. Scientific American, 49(15):227, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n15/pdf/scientificamerican10131883-227.pdf.

Anonymous:1883:DRC

[4418] Anonymous. Deep rock cuts near New York. Scientific American, 49(15):227, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n15/pdf/scientificamerican10131883-
227a.pdf.

Anonymous:1883:WTS

[4419] Anonymous. Wagon tongue support. Scientific American, 49 (15):227, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n15/pdf/scientificamerican10131883-
227b.pdf.

Anonymous:1883:CCB

[4420] Anonymous. The cable cars of the Brooklyn Bridge. Scientific American, 49(15):230, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n15/pdf/scientificamerican10131883-
230.pdf.

Anonymous:1883:CPa

[4421] Anonymous. The crowned pigeon. Scientific American, 49(15): 231–232, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n15/pdf/scientificamerican10131883-
231.pdf.

Anonymous:1883:CST

[4422] Anonymous. The champion steam trap. Scientific American, 49 (15):232, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n15/pdf/scientificamerican10131883-
232.pdf.

Anonymous:1883:PSV

[4423] Anonymous. Pole and shaft for vehicles. Scientific American, 49 (15):232–233, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n15/pdf/scientificamerican10131883-
232a.pdf.

Anonymous:1883:EIba

[4424] Anonymous. Engineering inventions. Scientific American, 49(15): 234, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n15/pdf/scientificamerican10131883-
234.pdf.

Anonymous:1883:MIce

[4425] Anonymous. Mechanical inventions. Scientific American, 49(15): 234, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n15/pdf/scientificamerican10131883-
234a.pdf.

Anonymous:1883:AIbb

[4426] Anonymous. Agricultural inventions. Scientific American, 49 (15):234, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n15/pdf/scientificamerican10131883-
234b.pdf.

Anonymous:1883:MIcf

[4427] Anonymous. Miscellaneous inventions. Scientific American, 49 (15):234, October 13, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n15/pdf/scientificamerican10131883-
234c.pdf.

Anonymous:1883:Cbd

[4428] Anonymous. Correspondence. Scientific American, 49(16):245–246, October 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n16/pdf/scientificamerican10201883-245.pdf.

Anonymous:1883:BNBs

[4429] Anonymous. Books: New books and publications. Scientific American, 49(16):250, October 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n16/pdf/scientificamerican10201883-
250d.pdf.

Anonymous:1883:BPbe

[4430] Anonymous. Business and personal. Scientific American, 49(16): 250–251, October 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n16/pdf/scientificamerican10201883-
250e.pdf.

Anonymous:1883:NQbe

[4431] Anonymous. Notes and queries. Scientific American, 49(16):251, October 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n16/pdf/scientificamerican10201883-251.pdf.

Anonymous:1883:IIba

[4432] Anonymous. Index of inventions. Scientific American, 49(16): 251–252, October 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n16/pdf/scientificamerican10201883-
251a.pdf.

Anonymous:1883:PNPt

[4433] Anonymous. Patent notes: Patents. Scientific American, 49 (16):252, October 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n16/pdf/scientificamerican10201883-
252.pdf.

Anonymous:1883:NYO

[4434] Anonymous. A New York office building. Scientific American, 49(16):240, October 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n16/pdf/scientificamerican10201883-
240.pdf.

Anonymous:1883:RBb

[4435] Anonymous. Raising bread. Scientific American, 49(16):240–241, October 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n16/pdf/scientificamerican10201883-240a.pdf.

Anonymous:1883:SCa

[4436] Anonymous. Steel castings. Scientific American, 49(16):241, October 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n16/pdf/scientificamerican10201883-241.pdf.

Anonymous:1883:MLb

[4437] Anonymous. Miner’s lamp. Scientific American, 49(16):242, October 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n16/pdf/scientificamerican10201883-242.pdf.

Anonymous:1883:ISI

[4438] Anonymous. Improved station indicator. Scientific American, 49 (16):242, October 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n16/pdf/scientificamerican10201883-
242a.pdf.

Anonymous:1883:FEb

[4439] Anonymous. Fire escape. Scientific American, 49(16):242–243, October 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n16/pdf/scientificamerican10201883-242b.pdf.

Anonymous:1883:AFC

[4440] Anonymous. Automatic freight car brake. Scientific American, 49(16):243, October 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n16/pdf/scientificamerican10201883-
243.pdf.

Anonymous:1883:LBD

[4441] Anonymous. Lifting bridge for double track railway. Scientific American, 49(16):244, October 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n16/pdf/scientificamerican10201883-
244.pdf.

Anonymous:1883:SACb

[4442] Anonymous. Science in antiquity. — curio’s pivoted theaters. Scientific American, 49(16):246–247, October 20, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v49/
n16/pdf/scientificamerican10201883-246.pdf.

Anonymous:1883:JBI

[4443] Anonymous. Javanese boats at the International Colonial Exhibition at Amsterdam. Scientific American, 49(16):247, October 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n16/pdf/scientificamerican10201883-247.pdf.

Anonymous:1883:CGa

[4444] Anonymous. Caliper gauge. Scientific American, 49(16):248, October 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n16/pdf/scientificamerican10201883-248.pdf.

Anonymous:1883:PGb

[4445] Anonymous. Pipe grapple. Scientific American, 49(16):248, October 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n16/pdf/scientificamerican10201883-248a.pdf.

Anonymous:1883:CWR

[4446] Anonymous. Combined whip and rein holder. Scientific American, 49(16):248, October 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n16/pdf/scientificamerican10201883-
248b.pdf.

Anonymous:1883:PHC

[4447] Anonymous. Pipe and hose coupling. Scientific American, 49 (16):248–249, October 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n16/pdf/scientificamerican10201883-
248c.pdf.

Anonymous:1883:ER

[4448] Anonymous. Evaporator for registers. Scientific American, 49 (16):249, October 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n16/pdf/scientificamerican10201883-
249.pdf.

Anonymous:1883:EIbb

[4449] Anonymous. Engineering inventions. Scientific American, 49(16): 250, October 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n16/pdf/scientificamerican10201883-
250.pdf.

Anonymous:1883:MIcg

[4450] Anonymous. Mechanical inventions. Scientific American, 49(16): 250, October 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n16/pdf/scientificamerican10201883-
250a.pdf.

Anonymous:1883:AIbc

[4451] Anonymous. Agricultural inventions. Scientific American, 49 (16):250, October 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n16/pdf/scientificamerican10201883-
250b.pdf.

Anonymous:1883:MIch

[4452] Anonymous. Miscellaneous inventions. Scientific American, 49 (16):250, October 20, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n16/pdf/scientificamerican10201883-
250c.pdf.

Anonymous:1883:Cbe

[4453] Anonymous. Correspondence. Scientific American, 49(17):261–262, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n17/pdf/scientificamerican10271883-261.pdf.

Anonymous:1883:BNBt

[4454] Anonymous. Books: New books and publications. Scientific American, 49(17):266, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n17/pdf/scientificamerican10271883-
266d.pdf.

Anonymous:1883:BPbf

[4455] Anonymous. Business and personal. Scientific American, 49(17): 266–267, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n17/pdf/scientificamerican10271883-
266e.pdf.

Anonymous:1883:NQbf

[4456] Anonymous. Notes and queries. Scientific American, 49(17):267, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n17/pdf/scientificamerican10271883-267.pdf.

Anonymous:1883:IIbb

[4457] Anonymous. Index of inventions. Scientific American, 49(17): 267–268, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n17/pdf/scientificamerican10271883-
267a.pdf.

Anonymous:1883:PNPu

[4458] Anonymous. Patent notes: Patents. Scientific American, 49 (17):269, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n17/pdf/scientificamerican10271883-
269.pdf.

Anonymous:1883:NBSb

[4459] Anonymous. A new band sawmill. Scientific American, 49(17):255, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n17/pdf/scientificamerican10271883-255.pdf.

Anonymous:1883:STC

[4460] Anonymous. Successful treatment of cholera. Scientific American, 49(17):256, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n17/pdf/scientificamerican10271883-
256.pdf.

Anonymous:1883:FSB

[4461] Anonymous. Fusible safety boiler plugs. Scientific American, 49 (17):256, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n17/pdf/scientificamerican10271883-
256a.pdf.

Anonymous:1883:PIC

[4462] Anonymous. Preparing for the ice crop. Scientific American, 49 (17):256, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n17/pdf/scientificamerican10271883-
256b.pdf.

Anonymous:1883:WFG

[4463] Anonymous. Wages in France and Germany. Scientific American, 49(17):256, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n17/pdf/scientificamerican10271883-
256c.pdf.

Anonymous:1883:PGC

[4464] Anonymous. Petroleum geologically considered. Scientific American, 49(17):257, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n17/pdf/scientificamerican10271883-
257.pdf.

Anonymous:1883:T

[4465] Anonymous. Tonkin. Scientific American, 49(17):257–258, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n17/pdf/scientificamerican10271883-257a.pdf.

Anonymous:1883:BBF

[4466] Anonymous. Bed bottom fire escape. Scientific American, 49 (17):258, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n17/pdf/scientificamerican10271883-
258.pdf.

Anonymous:1883:MBCb

[4467] Anonymous. Mail bag catcher. Scientific American, 49(17):258, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n17/pdf/scientificamerican10271883-258a.pdf.

Anonymous:1883:FEAa

[4468] Anonymous. Fire extinguishing apparatus. Scientific American, 49 (17):258–259, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n17/pdf/scientificamerican10271883-
258b.pdf.

Anonymous:1883:NTS

[4469] Anonymous. New turret ship conqueror. Scientific American, 49 (17):259, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n17/pdf/scientificamerican10271883-
259.pdf.

Anonymous:1883:ISG

[4470] Anonymous. Improved sand guard for railways. Scientific American, 49(17):259–260, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n17/pdf/scientificamerican10271883-
259a.pdf.

Anonymous:1883:IHA

[4471] Anonymous. An improved hand air condenser. Scientific American, 49(17):260, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n17/pdf/scientificamerican10271883-
260.pdf.

Anonymous:1883:NGE

[4472] Anonymous. New gas engine pump. Scientific American, 49(17): 262–263, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n17/pdf/scientificamerican10271883-
262.pdf.

Anonymous:1883:SES

[4473] Anonymous. Sacred Egyptian scarabæus. Scientific American, 49(17):263, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n17/pdf/scientificamerican10271883-
263.pdf.

Anonymous:1883:NHNe

[4474] Anonymous. Natural history notes. Scientific American, 49(17): 263–264, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n17/pdf/scientificamerican10271883-
263a.pdf.

Anonymous:1883:ITb

[4475] Anonymous. Improved tricycle. Scientific American, 49(17):264, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n17/pdf/scientificamerican10271883-264.pdf.

Anonymous:1883:FBa

[4476] Anonymous. Folding basket. Scientific American, 49(17):264–265, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n17/pdf/scientificamerican10271883-264a.pdf.

Anonymous:1883:EIbc

[4477] Anonymous. Engineering inventions. Scientific American, 49(17): 266, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n17/pdf/scientificamerican10271883-
266.pdf.

Anonymous:1883:MIci

[4478] Anonymous. Mechanical inventions. Scientific American, 49(17): 266, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n17/pdf/scientificamerican10271883-
266a.pdf.

Anonymous:1883:AIbd

[4479] Anonymous. Agricultural inventions. Scientific American, 49 (17):266, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n17/pdf/scientificamerican10271883-
266b.pdf.

Anonymous:1883:MIcj

[4480] Anonymous. Miscellaneous inventions. Scientific American, 49 (17):266, October 27, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n17/pdf/scientificamerican10271883-
266c.pdf.

Anonymous:1883:Cbf

[4481] Anonymous. Correspondence. Scientific American, 49(18):277– 278, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n18/pdf/scientificamerican11031883-
277.pdf.

Anonymous:1883:BNBu

[4482] Anonymous. Books: New books and publications. Scientific American, 49(18):282, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n18/pdf/scientificamerican11031883-
282d.pdf.

Anonymous:1883:BPbg

[4483] Anonymous. Business and personal. Scientific American, 49(18): 282–283, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n18/pdf/scientificamerican11031883-
282e.pdf.

Anonymous:1883:NQbg (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n18/pdf/scientificamerican11031883-
283.pdf.

Anonymous:1883:IIbc

[4485] Anonymous. Index of inventions. Scientific American, 49(18): 283–284, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n18/pdf/scientificamerican11031883-
283a.pdf.

Anonymous:1883:PNPv

[4486] Anonymous. Patent notes: Patents. Scientific American, 49 (18):284, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n18/pdf/scientificamerican11031883-
284.pdf.

Anonymous:1883:TWI

[4487] Anonymous. Three wheeled inspection car. Scientific American, 49(18):271, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n18/pdf/scientificamerican11031883-
271.pdf.

Anonymous:1883:BAD

[4488] Anonymous. The Barney automatic dumping boat. Scientific American, 49(18):271, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n18/pdf/scientificamerican11031883-
271a.pdf.

Anonymous:1883:CIb

[4489] Anonymous. A convention of inventors. Scientific American, 49 (18):272, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n18/pdf/scientificamerican11031883-
272.pdf.

Anonymous:1883:OSb

[4490] Anonymous. The oil stone. Scientific American, 49(18):272, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n18/pdf/scientificamerican11031883-272a.pdf.

Anonymous:1883:MWP

[4491] Anonymous. Making wind power auxiliary. Scientific American, 49 (18):272–273, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n18/pdf/scientificamerican11031883-
272b.pdf.

Anonymous:1883:APN

[4492] Anonymous. Aspects of the planets for November. Scientific American, 49(18):273–274, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n18/pdf/scientificamerican11031883-
273.pdf.

Anonymous:1883:PGd

[4493] Anonymous. Plow gauge. Scientific American, 49(18):274, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n18/pdf/scientificamerican11031883-274.pdf.

Anonymous:1883:SG

[4494] Anonymous. Sheep gate. Scientific American, 49(18):274, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n18/pdf/scientificamerican11031883-274a.pdf.

Anonymous:1883:Fd

[4495] Anonymous. Fence. Scientific American, 49(18):274–275, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n18/pdf/scientificamerican11031883-274b.pdf.

Anonymous:1883:SV

[4496] Anonymous. Slide valve. Scientific American, 49(18):275, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n18/pdf/scientificamerican11031883-275.pdf.

Anonymous:1883:BSV

[4497] Anonymous. Balanced slide valve. Scientific American, 49(18): 275, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n18/pdf/scientificamerican11031883-
275a.pdf.

Anonymous:1883:CSb

[4498] Anonymous. Coasting without snow. Scientific American, 49 (18):276, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n18/pdf/scientificamerican11031883-
276.pdf.

Anonymous:1883:ASF

[4499] Anonymous. An Australian steam ferry boat. Scientific American, 49(18):278–279, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n18/pdf/scientificamerican11031883-
278.pdf.

Anonymous:1883:K

[4500] Anonymous. Krakatoa. Scientific American, 49(18):279, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n18/pdf/scientificamerican11031883-279.pdf.

Anonymous:1883:BNW

[4501] Anonymous. The bottle-nosed whale. Scientific American, 49 (18):279, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n18/pdf/scientificamerican11031883-
279a.pdf.

Anonymous:1883:AWAa

[4502] Anonymous. Artesian wells, Australia. Scientific American, 49 (18):279–281, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n18/pdf/scientificamerican11031883-
279b.pdf.

Anonymous:1883:CDb

[4503] Anonymous. Calipers and dividers. Scientific American, 49(18): 281, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n18/pdf/scientificamerican11031883-
281.pdf.

Anonymous:1883:CTb

[4504] Anonymous. Combination tool. Scientific American, 49(18): 281, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n18/pdf/scientificamerican11031883-
281a.pdf.

Anonymous:1883:EIbd

[4505] Anonymous. Engineering inventions. Scientific American, 49(18): 282, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n18/pdf/scientificamerican11031883-
282.pdf.

Anonymous:1883:MIck

[4506] Anonymous. Mechanical intentions. Scientific American, 49(18): 282, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n18/pdf/scientificamerican11031883-
282a.pdf.

Anonymous:1883:AIbe

[4507] Anonymous. Agricultural inventions. Scientific American, 49 (18):282, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n18/pdf/scientificamerican11031883-
282b.pdf.

Anonymous:1883:MIcl

[4508] Anonymous. Miscellaneous inventions. Scientific American, 49 (18):282, November 3, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n18/pdf/scientificamerican11031883-
282c.pdf.

Anonymous:1883:Cbg

[4509] Anonymous. Correspondence. Scientific American, 49(19):293– 294, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n19/pdf/scientificamerican11101883-
293.pdf.

Anonymous:1883:BNBv

[4510] Anonymous. Books: New books and publications. Scientific American, 49(19):297, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n19/pdf/scientificamerican11101883-
297d.pdf.

Anonymous:1883:NQbh

[4511] Anonymous. Notes and queries. Scientific American, 49(19): 297–299, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n19/pdf/scientificamerican11101883-
297e.pdf.

Anonymous:1883:BPbh

[4512] Anonymous. Business and personal. Scientific American, 49 (19):299, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n19/pdf/scientificamerican11101883-
299.pdf.

Anonymous:1883:IIbd

[4513] Anonymous. Index of inventions. Scientific American, 49(19): 299–300, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n19/pdf/scientificamerican11101883-
299a.pdf.

Anonymous:1883:PNPw

[4514] Anonymous. Patent notes: Patents. Scientific American, 49 (19):300, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n19/pdf/scientificamerican11101883-
300.pdf.

Anonymous:1883:TEL

[4515] Anonymous. The traveling electric light. Scientific American, 49 (19):287–288, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n19/pdf/scientificamerican11101883-
287.pdf.

Anonymous:1883:SSI

[4516] Anonymous. Substitution of steel for iron. Scientific American, 49(19):288, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n19/pdf/scientificamerican11101883-
288.pdf.

Anonymous:1883:NM

[4517] Anonymous. November meteors. Scientific American, 49(19): 288–289, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n19/pdf/scientificamerican11101883-
288a.pdf.

Anonymous:1883:OFC

[4518] Anonymous. Ostrich farming in this country. Scientific American, 49(19):289, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n19/pdf/scientificamerican11101883-
289.pdf.

Anonymous:1883:TF

[4519] Anonymous. Tobacco in France. Scientific American, 49(19): 289, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n19/pdf/scientificamerican11101883-
289a.pdf.

Anonymous:1883:DRPe

[4520] Anonymous. Decisions relating to patents. Scientific American, 49 (19):289–290, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n19/pdf/scientificamerican11101883-
289b.pdf.

Anonymous:1883:BB

[4521] Anonymous. Butter box. Scientific American, 49(19):290, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n19/pdf/scientificamerican11101883-290.pdf.

Anonymous:1883:EFA

[4522] Anonymous. Electric fire alarm. Scientific American, 49(19): 290, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n19/pdf/scientificamerican11101883-
290a.pdf.

Anonymous:1883:Cbh

[4523] Anonymous. Churn. Scientific American, 49(19):290, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n19/pdf/scientificamerican11101883-290b.pdf.

Anonymous:1883:EDM

[4524] Anonymous. The electro dynamic motor. Scientific American, 49 (19):291–292, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n19/pdf/scientificamerican11101883-
291.pdf.

Hiscox:1883:RED

[4525] G. D. Hiscox. The re-enforcement of deficient water supply in wells. Scientific American, 49(19):292, November 10, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v49/
n19/pdf/scientificamerican11101883-292.pdf.

Anonymous:1883:HPD

[4526] Anonymous. Hydraulic pile driver. Scientific American, 49(19): 294, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n19/pdf/scientificamerican11101883-
294.pdf.

Anonymous:1883:NRB

[4527] Anonymous. A novel Russian boat. Scientific American, 49(19): 294–295, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n19/pdf/scientificamerican11101883-
294a.pdf.

Anonymous:1883:RAC

[4528] Anonymous. Ringed adders creeping out from the eggs, in the Berlin Aquarium. Scientific American, 49(19):295–296, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n19/pdf/scientificamerican11101883-295.pdf.

Anonymous:1883:APW

[4529] Anonymous. Adjustable pipe wrench. Scientific American, 49 (19):296, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n19/pdf/scientificamerican11101883-
296.pdf.

Anonymous:1883:PFc

[4530] Anonymous. Patent fence. Scientific American, 49(19):296, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n19/pdf/scientificamerican11101883-296a.pdf.

Anonymous:1883:EIbe

[4531] Anonymous. Engineering inventions. Scientific American, 49(19): 297, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n19/pdf/scientificamerican11101883-
297.pdf.

Anonymous:1883:MIcm

[4532] Anonymous. Mechanical inventions. Scientific American, 49(19): 297, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n19/pdf/scientificamerican11101883-
297a.pdf.

Anonymous:1883:AIbf

[4533] Anonymous. Agricultural inventions. Scientific American, 49 (19):297, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n19/pdf/scientificamerican11101883-
297b.pdf.

Anonymous:1883:MIcn

[4534] Anonymous. Miscellaneous inventions. Scientific American, 49 (19):297, November 10, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n19/pdf/scientificamerican11101883-
297c.pdf.

Anonymous:1883:Cbi

[4535] Anonymous. Correspondence. Scientific American, 49(2):21–22, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n2/pdf/scientificamerican07141883-21.pdf.

Anonymous:1883:IRIy

[4536] Anonymous. Inventions: Recent inventions. Scientific American, 49(2):25, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n2/pdf/scientificamerican07141883-
25.pdf.

Anonymous:1883:BPbi

[4537] Anonymous. Business and personal. Scientific American, 49(2):27, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n2/pdf/scientificamerican07141883-27.pdf.

Anonymous:1883:BNBw

[4538] Anonymous. Books: New books and publications. Scientific American, 49(2):27, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n2/pdf/scientificamerican07141883-
27a.pdf.

Anonymous:1883:NQbi

[4539] Anonymous. Notes and queries. Scientific American, 49(2):27, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n2/pdf/scientificamerican07141883-27b.pdf.

Anonymous:1883:IIbe

[4540] Anonymous. Index of inventions. Scientific American, 49(2):27–28, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n2/pdf/scientificamerican07141883-27c.pdf.

Anonymous:1883:PNPx

[4541] Anonymous. Patent notes: Patents. Scientific American, 49(2):28, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n2/pdf/scientificamerican07141883-28.pdf.

Anonymous:1883:MSE

[4542] Anonymous. A modern steam engine. Scientific American, 49(2):15, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n2/pdf/scientificamerican07141883-15.pdf.

Anonymous:1883:ABL

[4543] Anonymous. Asquith’s brake lathe. Scientific American, 49(2):15, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n2/pdf/scientificamerican07141883-15a.pdf.

Anonymous:1883:WS

[4544] Anonymous. The workingman’s school. Scientific American, 49(2):16, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n2/pdf/scientificamerican07141883-16.pdf.

Anonymous:1883:LLS

[4545] Anonymous. Loss by lack of system. Scientific American, 49(2):16, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n2/pdf/scientificamerican07141883-16a.pdf.

Anonymous:1883:SSa

[4546] Anonymous. Science on sorghum. Scientific American, 49(2):16–17, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n2/pdf/scientificamerican07141883-16b.pdf.

Anonymous:1883:SWPc

[4547] Anonymous. The storage of wind power. Scientific American, 49(2):17, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n2/pdf/scientificamerican07141883-17.pdf.

Anonymous:1883:ABE

[4548] Anonymous. The Aurania’s broken engines. Scientific American, 49(2):17, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n2/pdf/scientificamerican07141883-
17a.pdf.

Anonymous:1883:FEL

[4549] Anonymous. The Fisheries Exhibition, London. Scientific American, 49(2):18, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n2/pdf/scientificamerican07141883-
18.pdf.

Anonymous:1883:DSF

[4550] Anonymous. Dredge for shell fish. Scientific American, 49(2):18–19, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n2/pdf/scientificamerican07141883-18a.pdf.

Anonymous:1883:CDT

[4551] Anonymous. Cloth drying and tentering machine. Scientific American, 49(2):19, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n2/pdf/scientificamerican07141883-
19.pdf.

Anonymous:1883:FEC

[4552] Anonymous. Fire escape cable. Scientific American, 49(2):19–20, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n2/pdf/scientificamerican07141883-19a.pdf.

Anonymous:1883:FBb

[4553] Anonymous. Fern basket. Scientific American, 49(2):20, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n2/pdf/scientificamerican07141883-20.pdf.

Anonymous:1883:NRS

[4554] Anonymous. New railroad signal. Scientific American, 49(2):22–23, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n2/pdf/scientificamerican07141883-22.pdf.

Anonymous:1883:CREb

[4555] Anonymous. Curious result of an earthquake. Scientific American, 49(2):23, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n2/pdf/scientificamerican07141883-
23.pdf.

Anonymous:1883:RS

[4556] Anonymous. A remarkable shark. Scientific American, 49(2):23–24, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n2/pdf/scientificamerican07141883-23a.pdf.

Anonymous:1883:EIbf

[4557] Anonymous. Engineering inventions. Scientific American, 49(2):26, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n2/pdf/scientificamerican07141883-26.pdf.

Anonymous:1883:MIco

[4558] Anonymous. Mechanical inventions. Scientific American, 49(2):26, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n2/pdf/scientificamerican07141883-26a.pdf.

Anonymous:1883:AIbg

[4559] Anonymous. Agricultural inventions. Scientific American, 49(2):26, July 14, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n2/pdf/scientificamerican07141883-26b.pdf.

Anonymous:1883:MIcp

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n2/pdf/scientificamerican07141883-26c.pdf.

Anonymous:1883:Cbj

[4561] Anonymous. Correspondence. Scientific American, 49(20):309, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n20/pdf/scientificamerican11171883-309.pdf.

Anonymous:1883:BPbj

[4562] Anonymous. Business and personal. Scientific American, 49 (20):314, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n20/pdf/scientificamerican11171883-
314g.pdf.

Anonymous:1883:NQbj

[4563] Anonymous. Notes and queries. Scientific American, 49(20): 315, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n20/pdf/scientificamerican11171883-
315.pdf.

Anonymous:1883:IIbf

[4564] Anonymous. Index of inventions. Scientific American, 49(20): 315–316, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n20/pdf/scientificamerican11171883-
315a.pdf.

Anonymous:1883:USS

[4565] Anonymous. United States screw steam cruiser Atlanta. Scientific American, 49(20):303–308, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v49/n20/pdf/scientificamerican11171883-
303.pdf.

Anonymous:1883:CCE

[4566] Anonymous. A Centennial Cotton Exhibition. Scientific American, 49(20):304, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n20/pdf/scientificamerican11171883-
304.pdf.

Anonymous:1883:AWW

[4567] Anonymous. Adulterations what the wholesale druggists say. Scientific American, 49(20):304, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v49/n20/pdf/scientificamerican11171883-
304a.pdf.

Anonymous:1883:BFa

[4568] Anonymous. Bearings and friction. Scientific American, 49(20): 304, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n20/pdf/scientificamerican11171883-
304b.pdf.

Anonymous:1883:CSO

[4569] Anonymous. Cotton seed oil. Scientific American, 49(20):304, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n20/pdf/scientificamerican11171883-304c.pdf.

Anonymous:1883:PSEb

[4570] Anonymous. Present steam engine practice. Scientific American, 49(20):304, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n20/pdf/scientificamerican11171883-
304d.pdf.

Anonymous:1883:SCSb

[4571] Anonymous. Some curiosities of steel. Scientific American, 49(20): 304–305, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n20/pdf/scientificamerican11171883-
304e.pdf.

Anonymous:1883:NTD

[4572] Anonymous. A new treatment for the dead. Scientific American, 49(20):305, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n20/pdf/scientificamerican11171883-
305.pdf.

Anonymous:1883:OOV

[4573] Anonymous. Osage orange vs. mulberry for the silkworm. Scientific American, 49(20):305, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v49/n20/pdf/scientificamerican11171883-
305a.pdf.

Anonymous:1883:RCa

[4574] Anonymous. Reappearance of the Comet of 1812. Scientific American, 49(20):305–306, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n20/pdf/scientificamerican11171883-
305b.pdf.

Anonymous:1883:OE

[4575] Anonymous. Oil extractor. Scientific American, 49(20):306, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n20/pdf/scientificamerican11171883-306.pdf.

Anonymous:1883:FBc

[4576] Anonymous. Folding boat. Scientific American, 49(20):306, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n20/pdf/scientificamerican11171883-306a.pdf.

Anonymous:1883:STb

[4577] Anonymous. Steam trap. Scientific American, 49(20):306, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n20/pdf/scientificamerican11171883-306b.pdf.

Anonymous:1883:CCa

[4578] Anonymous. Compression cock. Scientific American, 49(20): 307, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n20/pdf/scientificamerican11171883-
307.pdf.

Anonymous:1883:DCD

[4579] Anonymous. Dust and cinder deflector. Scientific American, 49 (20):307, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n20/pdf/scientificamerican11171883-
307a.pdf.

Anonymous:1883:FT

[4580] Anonymous. Folding table. Scientific American, 49(20):307, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n20/pdf/scientificamerican11171883-307b.pdf.

Anonymous:1883:Cbk

[4581] Anonymous. Clevis. Scientific American, 49(20):307, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n20/pdf/scientificamerican11171883-307c.pdf.

Anonymous:1883:ICS

[4582] Anonymous. Improved combined shears and punch. Scientific American, 49(20):310, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n20/pdf/scientificamerican11171883-
310.pdf.

Anonymous:1883:BMH

[4583] Anonymous. Bread making at home. Scientific American, 49(20): 311–312, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n20/pdf/scientificamerican11171883-
311.pdf.

Anonymous:1883:SR

[4584] Anonymous. Stock rest. Scientific American, 49(20):312, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n20/pdf/scientificamerican11171883-312.pdf.

Anonymous:1883:FST

[4585] Anonymous. Felloe and spoke tightener. Scientific American, 49 (20):312–313, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n20/pdf/scientificamerican11171883-
312a.pdf.

Anonymous:1883:EIbg

[4586] Anonymous. Engineering inventions. Scientific American, 49(20): 314, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n20/pdf/scientificamerican11171883-
314.pdf.

Anonymous:1883:MIcq

[4587] Anonymous. Mechanical inventions. Scientific American, 49(20): 314, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n20/pdf/scientificamerican11171883-
314a.pdf.

Anonymous:1883:AIbh

[4588] Anonymous. Agricultural inventions. Scientific American, 49 (20):314, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n20/pdf/scientificamerican11171883-
314b.pdf.

Anonymous:1883:MIcr

[4589] Anonymous. Miscellaneous inventions. Scientific American, 49 (20):314, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n20/pdf/scientificamerican11171883-
314c.pdf.

Anonymous:1883:Ib

[4590] Anonymous. Insurance. Scientific American, 49(20):314, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n20/pdf/scientificamerican11171883-314d.pdf.

Anonymous:1883:Fg

[4591] Anonymous. Financial. Scientific American, 49(20):314, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n20/pdf/scientificamerican11171883-314e.pdf.

Anonymous:1883:MNa

[4592] Anonymous. Manufacturing notes. Scientific American, 49(20): 314, November 17, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n20/pdf/scientificamerican11171883-
314f.pdf.

Anonymous:1883:Cbl

[4593] Anonymous. Correspondence. Scientific American, 49(21):325– 326, November 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n21/pdf/scientificamerican11241883-
325.pdf.

Anonymous:1883:BPbk

[4594] Anonymous. Business and personal. Scientific American, 49 (21):330, November 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n21/pdf/scientificamerican11241883-
330g.pdf.

Anonymous:1883:NQbk

[4595] Anonymous. Notes and queries. Scientific American, 49(21): 331, November 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n21/pdf/scientificamerican11241883-
331.pdf.

Anonymous:1883:IIbg

[4596] Anonymous. Index of inventions. Scientific American, 49(21): 331–332, November 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n21/pdf/scientificamerican11241883-
331a.pdf.

Anonymous:1883:CBNa

[4597] Anonymous. The coal bunkers at New Tacoma, W.T. Scientific American, 49(21):319, November 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n21/pdf/scientificamerican11241883-
319.pdf.

Anonymous:1883:THM

[4598] Anonymous. Trial of the Haskell multicharge gun. Scientific American, 49(21):320, November 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n21/pdf/scientificamerican11241883-
320.pdf.

Anonymous:1883:RSB

[4599] Anonymous. Repairing suspension bridge cables. Scientific American, 49(21):320, November 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n21/pdf/scientificamerican11241883-
320a.pdf.

Anonymous:1883:AIFa

[4600] Anonymous. The American Institute Fair. Scientific American, 49 (21):321–322, November 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n21/pdf/scientificamerican11241883-
321.pdf.

Anonymous:1883:CCS

[4601] Anonymous. Cross cut saw frame. Scientific American, 49(21): 322, November 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n21/pdf/scientificamerican11241883-
322.pdf.

Anonymous:1883:HA

[4602] Anonymous. Hollow auger. Scientific American, 49(21):322, November 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n21/pdf/scientificamerican11241883-322a.pdf.

Anonymous:1883:IVb

[4603] Anonymous. Improved vise. Scientific American, 49(21):322– 323, November 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n21/pdf/scientificamerican11241883-
322b.pdf.

Anonymous:1883:IGE

[4604] Anonymous. Improved grain elevator. Scientific American, 49 (21):323, November 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n21/pdf/scientificamerican11241883-
323.pdf.

Anonymous:1883:ATR

[4605] Anonymous. Adjustable telephone receiver. Scientific American, 49 (21):323–324, November 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n21/pdf/scientificamerican11241883-
323a.pdf.

Anonymous:1883:EMc

[4606] Anonymous. The electro-magnetophone. Scientific American, 49 (21):324, November 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n21/pdf/scientificamerican11241883-
324.pdf.

Anonymous:1883:TEB

[4607] Anonymous. Tissandier’s electric balloon. Scientific American, 49 (21):326–328, November 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n21/pdf/scientificamerican11241883-
326.pdf.

Anonymous:1883:HT

[4608] Anonymous. Hame tug. Scientific American, 49(21):328, November 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n21/pdf/scientificamerican11241883-328.pdf.

Anonymous:1883:TB

[4609] Anonymous. Trace buckle. Scientific American, 49(21):328, November 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n21/pdf/scientificamerican11241883-328a.pdf.

Anonymous:1883:HF

[4610] Anonymous. Hame fastener. Scientific American, 49(21):328– 329, November 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n21/pdf/scientificamerican11241883-
328b.pdf.

Anonymous:1883:EIbh (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n21/pdf/scientificamerican11241883-
330.pdf.

Anonymous:1883:MIcs

[4612] Anonymous. Mechanical inventions. Scientific American, 49(21): 330, November 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n21/pdf/scientificamerican11241883-
330a.pdf.

Anonymous:1883:AIbi

[4613] Anonymous. Agricultural inventions. Scientific American, 49 (21):330, November 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n21/pdf/scientificamerican11241883-
330b.pdf.

Anonymous:1883:MIct

[4614] Anonymous. Miscellaneous inventions. Scientific American, 49 (21):330, November 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n21/pdf/scientificamerican11241883-
330c.pdf.

Anonymous:1883:Id

[4615] Anonymous. Insurance. Scientific American, 49(21):330, November 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n21/pdf/scientificamerican11241883-330d.pdf.

Anonymous:1883:Fh

[4616] Anonymous. Financial. Scientific American, 49(21):330, November 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n21/pdf/scientificamerican11241883-330e.pdf.

Anonymous:1883:MNb

[4617] Anonymous. Manufacturing notes. Scientific American, 49(21): 330, November 24, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n21/pdf/scientificamerican11241883-
330f.pdf.

Anonymous:1883:BPbl

[4618] Anonymous. Business and personal. Scientific American, 49 (22):345, December 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n22/pdf/scientificamerican12011883-
345g.pdf.

Anonymous:1883:BNBx

[4619] Anonymous. Books: New books and publications. Scientific American, 49(22):346, December 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n22/pdf/scientificamerican12011883-
346.pdf.

Anonymous:1883:NQbl

[4620] Anonymous. Notes and queries. Scientific American, 49(22):346, December 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n22/pdf/scientificamerican12011883-346a.pdf.

Anonymous:1883:IIbh

[4621] Anonymous. Index of inventions. Scientific American, 49(22): 346–347, December 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n22/pdf/scientificamerican12011883-
346b.pdf.

Anonymous:1883:GCB

[4622] Anonymous. The great cantilever bridge over Niagara River. Scientific American, 49(22):335–342, December 1, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v49/
n22/pdf/scientificamerican12011883-335.pdf.

Anonymous:1883:UAD

[4623] Anonymous. The use of Argol in dyeing. Scientific American, 49(22):336, December 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n22/pdf/scientificamerican12011883-
336.pdf.

Anonymous:1883:APD

[4624] Anonymous. Aspects of the planets for December. Scientific American, 49(22):337, December 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n22/pdf/scientificamerican12011883-
337.pdf.

Anonymous:1883:BDI

[4625] Anonymous. Bacterial disease of the imported cabbage worm. Scientific American, 49(22):337–338, December 1, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v49/
n22/pdf/scientificamerican12011883-337a.pdf.

Anonymous:1883:Cbm

[4626] Anonymous. Cultivator. Scientific American, 49(22):338, December 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n22/pdf/scientificamerican12011883-338.pdf.

Anonymous:1883:BFb

[4627] Anonymous. Box fastener. Scientific American, 49(22):338, December 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n22/pdf/scientificamerican12011883-338a.pdf.

Anonymous:1883:BPbm

[4628] Anonymous. Barrel pump. Scientific American, 49(22):338, December 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n22/pdf/scientificamerican12011883-338b.pdf.

Anonymous:1883:SSb

[4629] Anonymous. Slotting shears. Scientific American, 49(22):338–339, December 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n22/pdf/scientificamerican12011883-338c.pdf.

Anonymous:1883:FEd

[4630] Anonymous. Fire escape. Scientific American, 49(22):339, December 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n22/pdf/scientificamerican12011883-339.pdf.

Anonymous:1883:NDF

[4631] Anonymous. A new disk fan or exhauster. Scientific American, 49(22):339, December 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n22/pdf/scientificamerican12011883-
339a.pdf.

Anonymous:1883:LEL

[4632] Anonymous. Locomotive electric lights. Scientific American, 49 (22):342, December 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n22/pdf/scientificamerican12011883-
342.pdf.

Anonymous:1883:CPb

[4633] Anonymous. The Canadian porcupine. Scientific American, 49 (22):343, December 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n22/pdf/scientificamerican12011883-
343.pdf.

Anonymous:1883:YM

[4634] Anonymous. Young marmosets. Scientific American, 49(22):343– 344, December 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n22/pdf/scientificamerican12011883-
343a.pdf.

Anonymous:1883:CFa

[4635] Anonymous. Carpet fastener. Scientific American, 49(22):344, December 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n22/pdf/scientificamerican12011883-344.pdf.

Anonymous:1883:EIbi

[4636] Anonymous. Engineering inventions. Scientific American, 49(22): 345, December 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n22/pdf/scientificamerican12011883-
345.pdf.

Anonymous:1883:MIcu

[4637] Anonymous. Mechanical inventions. Scientific American, 49(22): 345, December 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n22/pdf/scientificamerican12011883-
345a.pdf.

Anonymous:1883:AIbj

[4638] Anonymous. Agricultural inventions. Scientific American, 49 (22):345, December 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n22/pdf/scientificamerican12011883-
345b.pdf.

Anonymous:1883:MIcv

[4639] Anonymous. Miscellaneous inventions. Scientific American, 49 (22):345, December 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n22/pdf/scientificamerican12011883-
345c.pdf.

Anonymous:1883:Ie

[4640] Anonymous. Insurance. Scientific American, 49(22):345, December 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n22/pdf/scientificamerican12011883-345d.pdf.

Anonymous:1883:Fi

[4641] Anonymous. Financial. Scientific American, 49(22):345, December 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n22/pdf/scientificamerican12011883-345e.pdf.

Anonymous:1883:MNc

[4642] Anonymous. Manufacturing notes. Scientific American, 49(22):345, December 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n22/pdf/scientificamerican12011883-345f.pdf.

Anonymous:1883:Cbp

[4643] Anonymous. Correspondence. Scientific American, 49(23):357, December 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n23/pdf/scientificamerican12081883-357.pdf.

Anonymous:1883:BPbn

[4644] Anonymous. Business and personal. Scientific American, 49(23): 362–363, December 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n23/pdf/scientificamerican12081883-
362d.pdf.

Anonymous:1883:BNBy

[4645] Anonymous. Books: New books and publications. Scientific American, 49(23):363, December 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n23/pdf/scientificamerican12081883-
363.pdf.

Anonymous:1883:NQbm

[4646] Anonymous. Notes and queries. Scientific American, 49(23):363, December 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n23/pdf/scientificamerican12081883-363a.pdf.

Anonymous:1883:IIbi

[4647] Anonymous. Index of inventions. Scientific American, 49(23): 363–364, December 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n23/pdf/scientificamerican12081883-
363b.pdf.

Anonymous:1883:NYW

[4648] Anonymous. The New York, West Shore, and Buffalo Railway Terminus. Scientific American, 49(23):351, December 8, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v49/
n23/pdf/scientificamerican12081883-351.pdf.

Anonymous:1883:ME

[4649] Anonymous. Mechanical exactness. Scientific American, 49(23): 352, December 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n23/pdf/scientificamerican12081883-
352.pdf.

Anonymous:1883:SWPd

[4650] Anonymous. Storing wind power for small motors. Scientific American, 49(23):353–354, December 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n23/pdf/scientificamerican12081883-
353.pdf.

Anonymous:1883:CBFa

[4651] Anonymous. Corrugated boiler furnaces. Scientific American, 49 (23):354, December 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n23/pdf/scientificamerican12081883-
354.pdf.

Anonymous:1883:NFP

[4652] Anonymous. A novel force pump. Scientific American, 49(23): 354–355, December 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n23/pdf/scientificamerican12081883-
354a.pdf.

Anonymous:1883:DS

[4653] Anonymous. Door securer. Scientific American, 49(23):355, December 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n23/pdf/scientificamerican12081883-355.pdf.

Anonymous:1883:EW

[4654] Anonymous. Emery wheel. Scientific American, 49(23):355, December 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n23/pdf/scientificamerican12081883-355a.pdf.

Anonymous:1883:HRS

[4655] Anonymous. Hand rail for stairways. Scientific American, 49 (23):355–356, December 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n23/pdf/scientificamerican12081883-
355b.pdf.

Anonymous:1883:HBT

[4656] Anonymous. Hemens’ bell for the transmission of signals on railroads. Scientific American, 49(23):358, December 8, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v49/
n23/pdf/scientificamerican12081883-358.pdf.

Anonymous:1883:UFS

[4657] Anonymous. Useful furniture by Sheraton. Scientific American, 49 (23):359–360, December 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n23/pdf/scientificamerican12081883-
359.pdf.

Anonymous:1883:PWC

[4658] Anonymous. Paper weight, calendar, and pen back. Scientific American, 49(23):360–361, December 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n23/pdf/scientificamerican12081883-
360.pdf.

Anonymous:1883:EIbj

[4659] Anonymous. Engineering inventions. Scientific American, 49(23): 362, December 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n23/pdf/scientificamerican12081883-
362.pdf.

Anonymous:1883:MIcw

[4660] Anonymous. Mechanical inventions. Scientific American, 49(23): 362, December 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n23/pdf/scientificamerican12081883-
362a.pdf.

Anonymous:1883:AIbk

[4661] Anonymous. Agricultural inventions. Scientific American, 49 (23):362, December 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n23/pdf/scientificamerican12081883-
362b.pdf.

Anonymous:1883:MIcx

[4662] Anonymous. Miscellaneous inventions. Scientific American, 49 (23):362, December 8, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n23/pdf/scientificamerican12081883-
362c.pdf.

Anonymous:1883:Cbq

[4663] Anonymous. Correspondence. Scientific American, 49(24):373– 374, December 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n24/pdf/scientificamerican12151883-
373.pdf.

Anonymous:1883:BPbo

[4664] Anonymous. Business and personal. Scientific American, 49(24): 378–379, December 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n24/pdf/scientificamerican12151883-
378d.pdf.

Anonymous:1883:NQbn

[4665] Anonymous. Notes and queries. Scientific American, 49(24): 379, December 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n24/pdf/scientificamerican12151883-
379.pdf.

Anonymous:1883:IIbj

[4666] Anonymous. Index of inventions. Scientific American, 49(24): 379–380, December 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n24/pdf/scientificamerican12151883-
379a.pdf.

Anonymous:1883:EPO

[4667] Anonymous. The Equatorial of the Paris Observatory. Scientific American, 49(24):367–372, December 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v49/n24/pdf/scientificamerican12151883-
367.pdf.

Anonymous:1883:EEL

[4668] Anonymous. The Edison Electric Light Co. Scientific American, 49(24):368, December 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n24/pdf/scientificamerican12151883-
368.pdf.

Anonymous:1883:IHc

[4669] Anonymous. Injury by heat. Scientific American, 49(24):368, December 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n24/pdf/scientificamerican12151883-368a.pdf.

Anonymous:1883:LS

[4670] Anonymous. The life of stones. Scientific American, 49(24):368, December 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n24/pdf/scientificamerican12151883-368b.pdf.

Anonymous:1883:CEa

[4671] Anonymous. The cholera in Egypt. Scientific American, 49 (24):369, December 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n24/pdf/scientificamerican12151883-
369.pdf.

Anonymous:1883:A

[4672] Anonymous. Argols. Scientific American, 49(24):369, December 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n24/pdf/scientificamerican12151883-369a.pdf.

Anonymous:1883:MTP

[4673] Anonymous. Machine for trimming paper on four sides. Scientific American, 49(24):370, December 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n24/pdf/scientificamerican12151883-
370.pdf.

Anonymous:1883:SA

[4674] Anonymous. Spark arrester. Scientific American, 49(24):370– 371, December 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n24/pdf/scientificamerican12151883-
370a.pdf.

Anonymous:1883:H

[4675] Anonymous. Harrow. Scientific American, 49(24):371, December 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n24/pdf/scientificamerican12151883-371.pdf.

Anonymous:1883:FEe

[4676] Anonymous. Fire escape. Scientific American, 49(24):371, December 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n24/pdf/scientificamerican12151883-371a.pdf.

Anonymous:1883:P

[4677] Anonymous. Plow. Scientific American, 49(24):371, December 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n24/pdf/scientificamerican12151883-371b.pdf.

Anonymous:1883:SGI

[4678] Anonymous. Simple garden improvement for promoting winter growth. Scientific American, 49(24):372, December 15, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v49/
n24/pdf/scientificamerican12151883-372.pdf.

Anonymous:1883:MS

[4679] Anonymous. A massive scaffolding. Scientific American, 49(24): 374, December 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n24/pdf/scientificamerican12151883-
374.pdf.

Anonymous:1883:EE

[4680] Anonymous. Engraved eggs. Scientific American, 49(24):374– 375, December 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n24/pdf/scientificamerican12151883-
374a.pdf.

Anonymous:1883:GH

[4681] Anonymous. The giant heron. Scientific American, 49(24):375– 376, December 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n24/pdf/scientificamerican12151883-
375.pdf.

Anonymous:1883:BH

[4682] Anonymous. Belt holder. Scientific American, 49(24):376, December 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n24/pdf/scientificamerican12151883-376.pdf.

Anonymous:1883:GS

[4683] Anonymous. Gold separator. Scientific American, 49(24):376– 377, December 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n24/pdf/scientificamerican12151883-
376a.pdf.

Anonymous:1883:EIbk

[4684] Anonymous. Engineering inventions. Scientific American, 49(24): 378, December 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n24/pdf/scientificamerican12151883-
378.pdf.

Anonymous:1883:MIcy

[4685] Anonymous. Mechanical intentions. Scientific American, 49(24): 378, December 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n24/pdf/scientificamerican12151883-
378a.pdf.

Anonymous:1883:AIbl

[4686] Anonymous. Agricultural inventions. Scientific American, 49 (24):378, December 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n24/pdf/scientificamerican12151883-
378b.pdf.

Anonymous:1883:MIcz

[4687] Anonymous. Miscellaneous inventions. Scientific American, 49 (24):378, December 15, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n24/pdf/scientificamerican12151883-
378c.pdf.

Anonymous:1883:Cbr

[4688] Anonymous. Correspondence. Scientific American, 49(25):389, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n25/pdf/scientificamerican12221883-389.pdf.

Anonymous:1883:BNBz

[4689] Anonymous. Books: New books and publications. Scientific American, 49(25):393, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n25/pdf/scientificamerican12221883-
393d.pdf.

Anonymous:1883:NQbo

[4690] Anonymous. Notes and queries. Scientific American, 49(25): 394, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n25/pdf/scientificamerican12221883-
394.pdf.

Anonymous:1883:BPbp

[4691] Anonymous. Business and personal. Scientific American, 49(25): 394–395, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n25/pdf/scientificamerican12221883-
394a.pdf.

Anonymous:1883:IIbk

[4692] Anonymous. Index of inventions. Scientific American, 49(25): 395–396, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n25/pdf/scientificamerican12221883-
395.pdf.

Anonymous:1883:PNPy

[4693] Anonymous. Patent notes: Patents. Scientific American, 49 (25):396, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n25/pdf/scientificamerican12221883-
396.pdf.

Anonymous:1883:NUSb

[4694] Anonymous. The new United States cruisers. Scientific American, 49(25):383, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n25/pdf/scientificamerican12221883-
383.pdf.

Anonymous:1883:ICd

[4695] Anonymous. Irrigation in California. Scientific American, 49 (25):384, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n25/pdf/scientificamerican12221883-
384.pdf.

Anonymous:1883:ICL

[4696] Anonymous. Imitations of costly leather. Scientific American, 49 (25):384, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n25/pdf/scientificamerican12221883-
384a.pdf.

Anonymous:1883:CBNb

[4697] Anonymous. The chinch-bug in New York. Scientific American, 49(25):384, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n25/pdf/scientificamerican12221883-
384b.pdf.

Anonymous:1883:AFG

[4698] Anonymous. Adulterations of foods — glucose in sugar and in sirups. Scientific American, 49(25):384–385, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v49/
n25/pdf/scientificamerican12221883-384c.pdf.

Anonymous:1883:FCP

[4699] Anonymous. “Fish culture for profit”. Scientific American, 49 (25):385, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n25/pdf/scientificamerican12221883-
385.pdf.

Riley:1883:DLH

[4700] C. V. Riley. Dipterous larvæ in the human body. Scientific American, 49(25):385, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n25/pdf/scientificamerican12221883-
385a.pdf.

Anonymous:1883:ILb

[4701] Anonymous. Impoverishment of land. Scientific American, 49 (25):385, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n25/pdf/scientificamerican12221883-
385b.pdf.

Anonymous:1883:HBMa

[4702] Anonymous. The “Hercules” bone mill. Scientific American, 49 (25):386, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n25/pdf/scientificamerican12221883-
386.pdf.

Anonymous:1883:SHT

[4703] Anonymous. Socket for harrow teeth. Scientific American, 49 (25):386, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n25/pdf/scientificamerican12221883-
386a.pdf.

Anonymous:1883:WJ

[4704] Anonymous. Wagon jack. Scientific American, 49(25):387, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n25/pdf/scientificamerican12221883-387.pdf.

Anonymous:1883:IRF

[4705] Anonymous. Improved Ruelle furnace for revivifying bone black. Scientific American, 49(25):387, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v49/n25/pdf/scientificamerican12221883-
387a.pdf.

Anonymous:1883:IAC

[4706] Anonymous. Improved air compressor. Scientific American, 49 (25):387–388, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n25/pdf/scientificamerican12221883-
387b.pdf.

Anonymous:1883:BI

[4707] Anonymous. The Bernissart iguanodons. Scientific American, 49 (25):391–392, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n25/pdf/scientificamerican12221883-
391.pdf.

Anonymous:1883:FPb

[4708] Anonymous. Foot power. Scientific American, 49(25):392, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n25/pdf/scientificamerican12221883-392.pdf.

Anonymous:1883:EIbl

[4709] Anonymous. Engineering inventions. Scientific American, 49(25): 393, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n25/pdf/scientificamerican12221883-
393.pdf.

Anonymous:1883:MIda

[4710] Anonymous. Mechanical inventions. Scientific American, 49(25): 393, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n25/pdf/scientificamerican12221883-
393a.pdf.

Anonymous:1883:AIbm

[4711] Anonymous. Agricultural inventions. Scientific American, 49 (25):393, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n25/pdf/scientificamerican12221883-
393b.pdf.

Anonymous:1883:MIdb

[4712] Anonymous. Miscellaneous inventions. Scientific American, 49 (25):393, December 22, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n25/pdf/scientificamerican12221883-
393c.pdf.

Anonymous:1883:BPbq

[4713] Anonymous. Business and personal. Scientific American, 49(26): 408–409, December 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n26/pdf/scientificamerican12291883-
408d.pdf.

Anonymous:1883:NQbp

[4714] Anonymous. Notes and queries. Scientific American, 49(26): 409, December 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n26/pdf/scientificamerican12291883-
409.pdf.

Anonymous:1883:IIbl

[4715] Anonymous. Index of inventions. Scientific American, 49(26): 409–410, December 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n26/pdf/scientificamerican12291883-
409a.pdf.

Anonymous:1883:Ij

[4716] Anonymous. Index. Scientific American, 49(26):411–412, December 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n26/pdf/scientificamerican12291883-411.pdf.

Anonymous:1883:PNPz

[4717] Anonymous. Patent notes: Patents. Scientific American, 49 (26):413, December 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n26/pdf/scientificamerican12291883-
413.pdf.

Anonymous:1883:MGTb

[4718] Anonymous. The Marent Gulch trestle. Scientific American, 49 (26):399, December 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n26/pdf/scientificamerican12291883-
399.pdf.

Anonymous:1883:PEb

[4719] Anonymous. Pink eye. Scientific American, 49(26):400, December 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n26/pdf/scientificamerican12291883-400.pdf.

Anonymous:1883:LIC

[4720] Anonymous. Light iron castings. Scientific American, 49(26): 400, December 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n26/pdf/scientificamerican12291883-
400a.pdf.

Anonymous:1883:TNS

[4721] Anonymous. Toothing a nicking saw. Scientific American, 49 (26):400, December 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n26/pdf/scientificamerican12291883-
400b.pdf.

Anonymous:1883:BE

[4722] Anonymous. The Bottlers’ Exhibition. Scientific American, 49 (26):400, December 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n26/pdf/scientificamerican12291883-
400c.pdf.

Anonymous:1883:NMR

[4723] Anonymous. Need of a manufacturing reform. Scientific American, 49(26):400, December 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n26/pdf/scientificamerican12291883-
400d.pdf.

Anonymous:1883:APJc

[4724] Anonymous. Aspects of the planets for January. Scientific American, 49(26):401, December 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n26/pdf/scientificamerican12291883-
401.pdf.

Anonymous:1883:MCc

[4725] Anonymous. Miner’s candlestick. Scientific American, 49(26): 402, December 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n26/pdf/scientificamerican12291883-
402.pdf.

Anonymous:1883:EAS

[4726] Anonymous. Electric alarm for steam boilers, etc. Scientific American, 49(26):402, December 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n26/pdf/scientificamerican12291883-
402a.pdf.

Anonymous:1883:MT

[4727] Anonymous. Mechanical telephone. Scientific American, 49(26): 402, December 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n26/pdf/scientificamerican12291883-
402b.pdf.

Anonymous:1883:CCc

[4728] Anonymous. Car coupling. Scientific American, 49(26):402– 403, December 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n26/pdf/scientificamerican12291883-
402c.pdf.

Anonymous:1883:MSS

[4729] Anonymous. The Magnetic Station at the Saint Maur Park Observatory. Scientific American, 49(26):403–405, December 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v49/
n26/pdf/scientificamerican12291883-403.pdf.

Anonymous:1883:WSA

[4730] Anonymous. Window sash adjuster. Scientific American, 49(26): 405–406, December 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n26/pdf/scientificamerican12291883-
405.pdf.

Anonymous:1883:FSA

[4731] Anonymous. Fog signal apparatus. Scientific American, 49(26): 406, December 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n26/pdf/scientificamerican12291883-
406.pdf.

Anonymous:1883:NFS

[4732] Anonymous. The needle fish, sea horses, and rag fish. Scientific American, 49(26):407, December 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n26/pdf/scientificamerican12291883-
407.pdf.

Anonymous:1883:EIbm

[4733] Anonymous. Engineering inventions. Scientific American, 49(26): 408, December 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n26/pdf/scientificamerican12291883-
408.pdf.

Anonymous:1883:MIdc

[4734] Anonymous. Mechanical inventions. Scientific American, 49(26): 408, December 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n26/pdf/scientificamerican12291883-
408a.pdf.

Anonymous:1883:AIbn

[4735] Anonymous. Agricultural inventions. Scientific American, 49 (26):408, December 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n26/pdf/scientificamerican12291883-
408b.pdf.

Anonymous:1883:MIdd

[4736] Anonymous. Miscellaneous inventions. Scientific American, 49 (26):408, December 29, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n26/pdf/scientificamerican12291883-
408c.pdf.

Anonymous:1883:Cbs

[4737] Anonymous. Correspondence. Scientific American, 49(3):37, July 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n3/pdf/scientificamerican07211883-37.pdf.

Anonymous:1883:BPbr

[4738] Anonymous. Business and personal. Scientific American, 49(3):43, July 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n3/pdf/scientificamerican07211883-43.pdf.

Anonymous:1883:BNBba

[4739] Anonymous. Books: New books and publications. Scientific American, 49(3):43, July 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n3/pdf/scientificamerican07211883-
43a.pdf.

Anonymous:1883:NQbq

[4740] Anonymous. Notes and queries. Scientific American, 49(3):43, July 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n3/pdf/scientificamerican07211883-43b.pdf.

Anonymous:1883:IIbm

[4741] Anonymous. Index of inventions. Scientific American, 49(3):43–44, July 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n3/pdf/scientificamerican07211883-43c.pdf.

Anonymous:1883:PNPba

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n3/pdf/scientificamerican07211883-44.pdf.

Anonymous:1883:PHH

[4743] Anonymous. Poisoned by handling hides. Scientific American, 49(3):32, July 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n3/pdf/scientificamerican07211883-32.pdf.

Anonymous:1883:SCb

[4744] Anonymous. Spontaneous combustion. Scientific American, 49(3):32, July 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n3/pdf/scientificamerican07211883-32a.pdf.

Anonymous:1883:SCc

[4745] Anonymous. Science in coloring. Scientific American, 49(3):32, July 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n3/pdf/scientificamerican07211883-32b.pdf.

Anonymous:1883:Lc

[4746] Anonymous. The lathe. Scientific American, 49(3):32, July 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n3/pdf/scientificamerican07211883-32c.pdf.

Anonymous:1883:SPWa

[4747] Anonymous. Storing the power of the wind. Scientific American, 49(3):33, July 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n3/pdf/scientificamerican07211883-
33.pdf.

Anonymous:1883:IFEd

[4748] Anonymous. Improved fruit evaporator. Scientific American, 49(3):34, July 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n3/pdf/scientificamerican07211883-34.pdf.

Anonymous:1883:HSTb

[4749] Anonymous. Herrington’s system of transportation and delivery. Scientific American, 49(3):34–35, July 21, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v49/
n3/pdf/scientificamerican07211883-34a.pdf.

Anonymous:1883:NCM

[4750] Anonymous. New cigarette machine. Scientific American, 49(3):35, July 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n3/pdf/scientificamerican07211883-35.pdf.

Anonymous:1883:NC

[4751] Anonymous. A new crutch. Scientific American, 49(3):35, July 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n3/pdf/scientificamerican07211883-35a.pdf.

Anonymous:1883:PNBb

[4752] Anonymous. Proposed new bridge, London. Scientific American, 49(3):36, July 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n3/pdf/scientificamerican07211883-
36.pdf.

Anonymous:1883:DRPf

[4753] Anonymous. Decisions relating to patents. Scientific American, 49(3):37, July 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n3/pdf/scientificamerican07211883-
37a.pdf.

Anonymous:1883:IPK

[4754] Anonymous. Improvement in pocket knives. Scientific American, 49(3):38, July 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n3/pdf/scientificamerican07211883-
38.pdf.

Anonymous:1883:ATW

[4755] Anonymous. An aerial turbine wheel. Scientific American, 49(3):38, July 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n3/pdf/scientificamerican07211883-38a.pdf.

Anonymous:1883:DFS

[4756] Anonymous. Device for fitting stove pipes. Scientific American, 49(3):38–39, July 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n3/pdf/scientificamerican07211883-
38b.pdf.

Anonymous:1883:SMa

[4757] Anonymous. Silkworms and moths. Scientific American, 49(3):39–41, July 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n3/pdf/scientificamerican07211883-39.pdf.

Anonymous:1883:EIbn

[4758] Anonymous. Engineering inventions. Scientific American, 49(3):42, July 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n3/pdf/scientificamerican07211883-42.pdf.

Anonymous:1883:MIde

[4759] Anonymous. Mechanical inventions. Scientific American, 49(3):42, July 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n3/pdf/scientificamerican07211883-42a.pdf.

Anonymous:1883:AIbo

[4760] Anonymous. Agricultural inventions. Scientific American, 49(3):42, July 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n3/pdf/scientificamerican07211883-42b.pdf.

Anonymous:1883:MIdf

[4761] Anonymous. Miscellaneous inventions. Scientific American, 49(3):42, July 21, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n3/pdf/scientificamerican07211883-42c.pdf.

Anonymous:1883:Cbt

[4762] Anonymous. Correspondence. Scientific American, 49(4):53–54, July 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n4/pdf/scientificamerican07281883-53.pdf.

Anonymous:1883:IRIz

[4763] Anonymous. Inventions: Recent inventions. Scientific American, 49(4):57, July 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n4/pdf/scientificamerican07281883-
57.pdf.

Anonymous:1883:BPbs

[4764] Anonymous. Business and personal. Scientific American, 49(4):59, July 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n4/pdf/scientificamerican07281883-59.pdf.

Anonymous:1883:BNBbb

[4765] Anonymous. Books: New books and publications. Scientific American, 49(4):59, July 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n4/pdf/scientificamerican07281883-
59a.pdf.

Anonymous:1883:NQbr

[4766] Anonymous. Notes and queries. Scientific American, 49(4):59, July 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n4/pdf/scientificamerican07281883-59b.pdf.

Anonymous:1883:IIbn

[4767] Anonymous. Index of inventions. Scientific American, 49(4):59–60, July 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n4/pdf/scientificamerican07281883-59d.pdf.

Anonymous:1883:PNPbb

[4768] Anonymous. Patent notes: Patents. Scientific American, 49(4):60, July 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n4/pdf/scientificamerican07281883-60.pdf.

Anonymous:1883:PAM

[4769] Anonymous. Process and apparatus for the manufacture of grape sugar. Scientific American, 49(4):47–48, July 28, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v49/
n4/pdf/scientificamerican07281883-47.pdf.

Anonymous:1883:LCM

[4770] Anonymous. Lock-out of cigar makers. Scientific American, 49(4):48, July 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n4/pdf/scientificamerican07281883-48.pdf.

Anonymous:1883:TS

[4771] Anonymous. The telegraphers’ strike. Scientific American, 49(4):48, July 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n4/pdf/scientificamerican07281883-48a.pdf.

Anonymous:1883:EB

[4772] Anonymous. The examination of bridges. Scientific American, 49(4):48–49, July 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n4/pdf/scientificamerican07281883-
48b.pdf.

Anonymous:1883:HHSb

[4773] Anonymous. Heating and hardening of steel. Scientific American, 49(4):49, July 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n4/pdf/scientificamerican07281883-
49.pdf.

Anonymous:1883:APAb

[4774] Anonymous. Aspects of the planets for August. Scientific American, 49(4):49, July 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n4/pdf/scientificamerican07281883-
49a.pdf.

Anonymous:1883:SPWb

[4775] Anonymous. Storing the power of the wind. Scientific American, 49(4):49–50, July 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n4/pdf/scientificamerican07281883-
49b.pdf.

Anonymous:1883:SCO

[4776] Anonymous. Simple can opener. Scientific American, 49(4):50, July 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n4/pdf/scientificamerican07281883-50.pdf.

Anonymous:1883:SOC

[4777] Anonymous. Steam oven for cloth plates. Scientific American, 49(4):50–51, July 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n4/pdf/scientificamerican07281883-
50a.pdf.

Anonymous:1883:GCA

[4778] Anonymous. Gear cutting attachment for lathes. Scientific American, 49(4):51, July 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n4/pdf/scientificamerican07281883-
51.pdf.

Anonymous:1883:NCCb

[4779] Anonymous. New car coupler. Scientific American, 49(4):51, July 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n4/pdf/scientificamerican07281883-51a.pdf.

Hovey:1883:SF

[4780] H. C. Hovey. The sturgeon fishery. Scientific American, 49(4):52, July 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n4/pdf/scientificamerican07281883-52.pdf.

Anonymous:1883:IMI

[4781] Anonymous. Improvement in the manufacture of illuminating gas. Scientific American, 49(4):54–55, July 28, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v49/
n4/pdf/scientificamerican07281883-54.pdf.

Anonymous:1883:AP

[4782] Anonymous. The Argus pheasant. Scientific American, 49(4):55–56, July 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n4/pdf/scientificamerican07281883-55.pdf.

Anonymous:1883:EIbo

[4783] Anonymous. Engineering inventions. Scientific American, 49(4):58, July 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n4/pdf/scientificamerican07281883-58.pdf.

Anonymous:1883:MIdg

[4784] Anonymous. Mechanical inventions. Scientific American, 49(4):58, July 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n4/pdf/scientificamerican07281883-58a.pdf.

Anonymous:1883:AIbp

[4785] Anonymous. Agricultural inventions. Scientific American, 49(4):58, July 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n4/pdf/scientificamerican07281883-58b.pdf.

Anonymous:1883:MIdh

[4786] Anonymous. Miscellaneous inventions. Scientific American, 49(4):58, July 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n4/pdf/scientificamerican07281883-58c.pdf.

Anonymous:1883:CRk

[4787] Anonymous. Communications received. Scientific American, 49(4):59, July 28, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n4/pdf/scientificamerican07281883-59c.pdf.

Anonymous:1883:IRIba

[4788] Anonymous. Inventions: Recent inventions. Scientific American, 49(5):73, August 4, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n5/pdf/scientificamerican08041883-
73.pdf.

Anonymous:1883:BPbt

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n5/pdf/scientificamerican08041883-75.pdf.

Anonymous:1883:BNBbc

[4790] Anonymous. Books: New books and publications. Scientific American, 49(5):75, August 4, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n5/pdf/scientificamerican08041883-
75a.pdf.

Anonymous:1883:NQbs

[4791] Anonymous. Notes and queries. Scientific American, 49(5):75, August 4, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n5/pdf/scientificamerican08041883-75b.pdf.

Anonymous:1883:IIbo

[4792] Anonymous. Index of inventions. Scientific American, 49(5):75–76, August 4, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n5/pdf/scientificamerican08041883-75d.pdf.

Anonymous:1883:PNPbc

[4793] Anonymous. Patent notes: Patents. Scientific American, 49(5):76, August 4, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n5/pdf/scientificamerican08041883-76.pdf.

Anonymous:1883:MECb

[4794] Anonymous. The manufacture of electrical conductors. Scientific American, 49(5):63, August 4, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n5/pdf/scientificamerican08041883-
63.pdf.

Anonymous:1883:BLG

[4795] Anonymous. A brave life gone out. Scientific American, 49(5):64, August 4, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n5/pdf/scientificamerican08041883-64.pdf.

Anonymous:1883:CGL

[4796] Anonymous. Completion of the great Lyman–Haskell gun. Scientific American, 49(5):64, August 4, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n5/pdf/scientificamerican08041883-
64a.pdf.

Anonymous:1883:TICb

[4797] Anonymous. The telephone interference case decided. Scientific American, 49(5):64, August 4, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n5/pdf/scientificamerican08041883-
64b.pdf.

Anonymous:1883:OBM

[4798] Anonymous. Old building material. Scientific American, 49(5):64, August 4, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n5/pdf/scientificamerican08041883-64c.pdf.

Anonymous:1883:RPE

[4799] Anonymous. The relations of Patent experts to the Courts. Scientific American, 49(5):64–65, August 4, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n5/pdf/scientificamerican08041883-
64d.pdf.

Anonymous:1883:CEb

[4800] Anonymous. The caterpillar and its enemy. Scientific American, 49(5):65, August 4, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n5/pdf/scientificamerican08041883-
65.pdf.

Anonymous:1883:SMb

[4801] Anonymous. Sketching for mechanics. Scientific American, 49(5):65, August 4, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n5/pdf/scientificamerican08041883-65a.pdf.

Anonymous:1883:PFd

[4802] Anonymous. Picker fires. Scientific American, 49(5):65, August 4, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n5/pdf/scientificamerican08041883-65b.pdf.

Anonymous:1883:TTO

[4803] Anonymous. The telescope at the Trocadero Observatory. Scientific American, 49(5):67, August 4, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n5/pdf/scientificamerican08041883-
67.pdf.

Anonymous:1883:PBP

[4804] Anonymous. Parallels for balancing pulleys, etc. Scientific American, 49(5):67–70, August 4, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n5/pdf/scientificamerican08041883-
67a.pdf.

Anonymous:1883:IRRb

[4805] Anonymous. Improved reversing rail mill engines. Scientific American, 49(5):70, August 4, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n5/pdf/scientificamerican08041883-
70.pdf.

Anonymous:1883:TGS

[4806] Anonymous. The turbot, goldbutt, and sole. Scientific American, 49(5):71, August 4, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n5/pdf/scientificamerican08041883-
71.pdf.

Anonymous:1883:NHNf

[4807] Anonymous. Natural history notes. Scientific American, 49(5):71–72, August 4, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n5/pdf/scientificamerican08041883-71a.pdf.

Anonymous:1883:EIbp

[4808] Anonymous. Engineering inventions. Scientific American, 49(5):74, August 4, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n5/pdf/scientificamerican08041883-74.pdf.

Anonymous:1883:MIdi

[4809] Anonymous. Mechanical inventions. Scientific American, 49(5):74, August 4, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n5/pdf/scientificamerican08041883-74a.pdf.

Anonymous:1883:AIbq

[4810] Anonymous. Agricultural inventions. Scientific American, 49(5):74, August 4, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n5/pdf/scientificamerican08041883-74b.pdf.

Anonymous:1883:MIdj

[4811] Anonymous. Miscellaneous inventions. Scientific American, 49(5):74, August 4, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n5/pdf/scientificamerican08041883-74c.pdf.

Anonymous:1883:CRl

[4812] Anonymous. Communications received. Scientific American, 49(5):75, August 4, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n5/pdf/scientificamerican08041883-75c.pdf.

Anonymous:1883:Cbu

[4813] Anonymous. Correspondence. Scientific American, 49(6):85, August 11, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n6/pdf/scientificamerican08111883-85.pdf.

Anonymous:1883:BPbu

[4814] Anonymous. Business and personal. Scientific American, 49(6):90, August 11, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n6/pdf/scientificamerican08111883-90d.pdf.

Anonymous:1883:BNBbd

[4815] Anonymous. Books: New books and publications. Scientific American, 49(6):91, August 11, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n6/pdf/scientificamerican08111883-
91.pdf.

Anonymous:1883:NQbt

[4816] Anonymous. Notes and queries. Scientific American, 49(6):91, August 11, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n6/pdf/scientificamerican08111883-91a.pdf.

Anonymous:1883:IIbp

[4817] Anonymous. Index of inventions. Scientific American, 49(6):91–92, August 11, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n6/pdf/scientificamerican08111883-91b.pdf.

Anonymous:1883:PNPbd

[4818] Anonymous. Patent notes: Patents. Scientific American, 49(6):92, August 11, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n6/pdf/scientificamerican08111883-92.pdf.

Anonymous:1883:TCO

[4819] Anonymous. The Telephone Central Office System of Paris. Scientific American, 49(6):79–80, August 11, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v49/n6/pdf/scientificamerican08111883-
79.pdf.

Anonymous:1883:RCb

[4820] Anonymous. The revival of cherry. Scientific American, 49(6):80, August 11, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n6/pdf/scientificamerican08111883-80.pdf.

Anonymous:1883:EAU

[4821] Anonymous. Effect of alcohol upon digestion. Scientific American, 49(6):80, August 11, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n6/pdf/scientificamerican08111883-
80a.pdf.

Anonymous:1883:ANb

[4822] Anonymous. Aerial navigation. Scientific American, 49(6):80–81, August 11, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n6/pdf/scientificamerican08111883-80b.pdf.

Anonymous:1883:Cbv

[4823] Anonymous. Cholera. Scientific American, 49(6):81, August 11, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n6/pdf/scientificamerican08111883-81.pdf.

Anonymous:1883:CMa

[4824] Anonymous. China as a market. Scientific American, 49(6):81, August 11, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n6/pdf/scientificamerican08111883-81a.pdf.

Anonymous:1883:DGL

[4825] Anonymous. The detection of gas leakages. Scientific American, 49(6):82, August 11, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n6/pdf/scientificamerican08111883-
82.pdf.

Anonymous:1883:NNL

[4826] Anonymous. New nut lock. Scientific American, 49(6):82, August 11, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n6/pdf/scientificamerican08111883-82a.pdf.

Anonymous:1883:CTE

[4827] Anonymous. Combined tag and envelope. Scientific American, 49(6):82, August 11, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n6/pdf/scientificamerican08111883-
82b.pdf.

Anonymous:1883:BEL

[4828] Anonymous. The Brush Electric Light Works, Rochester, N.Y. Scientific American, 49(6):83–84, August 11, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v49/n6/pdf/scientificamerican08111883-
83.pdf.

Anonymous:1883:TBM

[4829] Anonymous. Tunnel boring machine. Scientific American, 49(6):86, August 11, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n6/pdf/scientificamerican08111883-86.pdf.

Anonymous:1883:GLD

[4830] Anonymous. Great loss in the diamond fields. Scientific American, 49(6):87–89, August 11, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n6/pdf/scientificamerican08111883-
87.pdf.

Anonymous:1883:EIbq

[4831] Anonymous. Engineering inventions. Scientific American, 49(6):90, August 11, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n6/pdf/scientificamerican08111883-90.pdf.

Anonymous:1883:MIdk

[4832] Anonymous. Mechanical inventions. Scientific American, 49(6):90, August 11, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n6/pdf/scientificamerican08111883-90a.pdf.

Anonymous:1883:AIbr

[4833] Anonymous. Agricultural inventions. Scientific American, 49(6):90, August 11, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n6/pdf/scientificamerican08111883-90b.pdf.

Anonymous:1883:MIdl

[4834] Anonymous. Miscellaneous inventions. Scientific American, 49 (6):90, August 11, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n6/pdf/scientificamerican08111883-
90c.pdf.

Anonymous:1883:BPbv

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n7/pdf/scientificamerican08181883-107.pdf.

Anonymous:1883:BNBbe

[4836] Anonymous. Books: New books and publications. Scientific American, 49(7):107, August 18, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n7/pdf/scientificamerican08181883-
107a.pdf.

Anonymous:1883:NQbu

[4837] Anonymous. Notes and queries. Scientific American, 49(7):107, August 18, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n7/pdf/scientificamerican08181883-107b.pdf.

Anonymous:1883:IIbq

[4838] Anonymous. Index of inventions. Scientific American, 49(7): 107–108, August 18, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n7/pdf/scientificamerican08181883-
107c.pdf.

Anonymous:1883:PNPbe

[4839] Anonymous. Patent notes: Patents. Scientific American, 49(7):109, August 18, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n7/pdf/scientificamerican08181883-109.pdf.

Anonymous:1883:SCd

[4840] Anonymous. The sugar canes. Scientific American, 49(7):96, August 18, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n7/pdf/scientificamerican08181883-96.pdf.

Anonymous:1883:ISBb

[4841] Anonymous. Increased speed between New York and Liverpool. Scientific American, 49(7):96, August 18, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v49/
n7/pdf/scientificamerican08181883-96a.pdf.

Anonymous:1883:NFB

[4842] Anonymous. New forms for boats. Scientific American, 49(7):97, August 18, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n7/pdf/scientificamerican08181883-97.pdf.

Anonymous:1883:FR

[4843] Anonymous. Faith remedies. Scientific American, 49(7):97, August 18, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n7/pdf/scientificamerican08181883-97a.pdf.

Anonymous:1883:SMc

[4844] Anonymous. Solar machinery. Scientific American, 49(7):97, August 18, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n7/pdf/scientificamerican08181883-97b.pdf.

Anonymous:1883:IFH

[4845] Anonymous. Improved fire-hose nozzle. Scientific American, 49 (7):98, August 18, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n7/pdf/scientificamerican08181883-
98.pdf.

Anonymous:1883:ISc

[4846] Anonymous. Improved sleigh. Scientific American, 49(7):98, August 18, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n7/pdf/scientificamerican08181883-98a.pdf.

Anonymous:1883:ITK

[4847] Anonymous. Improvement in tea kettles. Scientific American, 49(7):99, August 18, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n7/pdf/scientificamerican08181883-
99.pdf.

Anonymous:1883:CSN

[4848] Anonymous. Chair seating needle. Scientific American, 49(7):99, August 18, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n7/pdf/scientificamerican08181883-99a.pdf.

Anonymous:1883:HSEb

[4849] Anonymous. Headless shell extractor. Scientific American, 49 (7):99, August 18, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n7/pdf/scientificamerican08181883-
99b.pdf.

Anonymous:1883:HSB

[4850] Anonymous. Hartford steam boiler inspection and insurance company. Scientific American, 49(7):100–101, August 18, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v49/
n7/pdf/scientificamerican08181883-100.pdf.

Anonymous:1883:IAM

[4851] Anonymous. Improved adding machine. Scientific American, 49 (7):102, August 18, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n7/pdf/scientificamerican08181883-
102.pdf.

Anonymous:1883:NDE

[4852] Anonymous. New dynamo electric machine. Scientific American, 49(7):102, August 18, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n7/pdf/scientificamerican08181883-
102a.pdf.

Anonymous:1883:GHL

[4853] Anonymous. Gas heated laundry iron. Scientific American, 49 (7):102, August 18, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n7/pdf/scientificamerican08181883-
102b.pdf.

Anonymous:1883:AK

[4854] Anonymous. The apteryx, or kiwi. Scientific American, 49(7): 103–105, August 18, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n7/pdf/scientificamerican08181883-
103.pdf.

Anonymous:1883:EIbr

[4855] Anonymous. Engineering inventions. Scientific American, 49(7): 106, August 18, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n7/pdf/scientificamerican08181883-
106.pdf.

Anonymous:1883:MIdm

[4856] Anonymous. Mechanical inventions. Scientific American, 49(7):106, August 18, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n7/pdf/scientificamerican08181883-106a.pdf.

Anonymous:1883:AIbs

[4857] Anonymous. Agricultural inventions. Scientific American, 49(7): 106, August 18, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n7/pdf/scientificamerican08181883-
106b.pdf.

Anonymous:1883:MIdn

[4858] Anonymous. Miscellaneous inventions. Scientific American, 49 (7):106, August 18, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n7/pdf/scientificamerican08181883-
106c.pdf.

Anonymous:1883:Cbw

[4859] Anonymous. Correspondence. Scientific American, 49(8):117, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n8/pdf/scientificamerican08251883-117.pdf.

Anonymous:1883:IRIbb

[4860] Anonymous. Inventions: Recent inventions. Scientific American, 49(8):121, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n8/pdf/scientificamerican08251883-
121.pdf.

Anonymous:1883:BPbw

[4861] Anonymous. Business and personal. Scientific American, 49(8):123, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n8/pdf/scientificamerican08251883-123.pdf.

Anonymous:1883:BNBbf

[4862] Anonymous. Books: New books and publications. Scientific American, 49(8):123, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n8/pdf/scientificamerican08251883-
123a.pdf.

Anonymous:1883:NQbv

[4863] Anonymous. Notes and queries. Scientific American, 49(8):123, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n8/pdf/scientificamerican08251883-123b.pdf.

Anonymous:1883:IIbr

[4864] Anonymous. Index of inventions. Scientific American, 49(8): 123–124, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n8/pdf/scientificamerican08251883-
123c.pdf.

Anonymous:1883:PNPbf

[4865] Anonymous. Patent notes: Patents. Scientific American, 49(8):125, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n8/pdf/scientificamerican08251883-125.pdf.

Anonymous:1883:HBMb

[4866] Anonymous. Hooping barrels by machinery. Scientific American, 49(8):111, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n8/pdf/scientificamerican08251883-
111.pdf.

Anonymous:1883:WSN

[4867] Anonymous. Water supply for New York City. Scientific American, 49(8):112, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n8/pdf/scientificamerican08251883-
112.pdf.

Anonymous:1883:CFb

[4868] Anonymous. The Carson footprints. Scientific American, 49(8):112, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n8/pdf/scientificamerican08251883-112a.pdf.

Anonymous:1883:HWF

[4869] Anonymous. Hard wood flooring. Scientific American, 49(8):112, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n8/pdf/scientificamerican08251883-112b.pdf.

Anonymous:1883:HGC

[4870] Anonymous. How granite columns are polished. Scientific American, 49(8):112–113, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n8/pdf/scientificamerican08251883-
112c.pdf.

Anonymous:1883:PRN

[4871] Anonymous. Palestine as a route for a new ship canal. Scientific American, 49(8):113, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n8/pdf/scientificamerican08251883-
113.pdf.

Coryell:1883:CJ

[4872] John R. Coryell. The Chinese junk. Scientific American, 49 (8):113–114, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n8/pdf/scientificamerican08251883-
113a.pdf.

Anonymous:1883:NLT

[4873] Anonymous. New lighthouse at Tampico. Scientific American, 49 (8):114–115, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n8/pdf/scientificamerican08251883-
114.pdf.

Anonymous:1883:SWT

[4874] Anonymous. Spring wheel for traction engines. Scientific American, 49(8):115, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n8/pdf/scientificamerican08251883-
115.pdf.

Anonymous:1883:MPE

[4875] Anonymous. Method of preserving ensilage in silos. Scientific American, 49(8):115–116, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n8/pdf/scientificamerican08251883-
115a.pdf.

Anonymous:1883:PRJ

[4876] Anonymous. From Paris to Rio Janeiro in 42 minutes and 11 seconds. Scientific American, 49(8):116, August 25, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v49/
n8/pdf/scientificamerican08251883-116.pdf.

Anonymous:1883:NCL

[4877] Anonymous. New combination lock. Scientific American, 49(8):118, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n8/pdf/scientificamerican08251883-118.pdf.

Anonymous:1883:NET

[4878] Anonymous. New expansion trap. Scientific American, 49(8): 118–119, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n8/pdf/scientificamerican08251883-
118a.pdf.

Anonymous:1883:JBC

[4879] Anonymous. A jardiniere, bird cage, and aquarium combined. Scientific American, 49(8):119, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n8/pdf/scientificamerican08251883-
119.pdf.

Anonymous:1883:LPH

[4880] Anonymous. Lightning prints on the human body. Scientific American, 49(8):119–120, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n8/pdf/scientificamerican08251883-
119a.pdf.

Anonymous:1883:EIbs

[4881] Anonymous. Engineering inventions. Scientific American, 49(8): 122, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n8/pdf/scientificamerican08251883-
122.pdf.

Anonymous:1883:MIdo

[4882] Anonymous. Mechanical inventions. Scientific American, 49(8):122, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n8/pdf/scientificamerican08251883-122a.pdf.

Anonymous:1883:AIbt

[4883] Anonymous. Agricultural inventions. Scientific American, 49(8): 122, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n8/pdf/scientificamerican08251883-
122b.pdf.

Anonymous:1883:MIdp

[4884] Anonymous. Miscellaneous inventions. Scientific American, 49 (8):122, August 25, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n8/pdf/scientificamerican08251883-
122c.pdf.

Anonymous:1883:Cbx

[4885] Anonymous. Correspondence. Scientific American, 49(9):133– 134, September 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n9/pdf/scientificamerican09011883-
133.pdf.

Anonymous:1883:IRIbc

[4886] Anonymous. Inventions: Recent inventions. Scientific American, 49 (9):136–137, September 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n9/pdf/scientificamerican09011883-
136.pdf.

Anonymous:1883:BPbx

[4887] Anonymous. Business and personal. Scientific American, 49(9): 138, September 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n9/pdf/scientificamerican09011883-
138.pdf.

Anonymous:1883:BNBbg

[4888] Anonymous. Books: New books and publications. Scientific American, 49(9):138, September 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n9/pdf/scientificamerican09011883-
138a.pdf.

Anonymous:1883:NQbw

[4889] Anonymous. Notes and queries. Scientific American, 49(9):138– 139, September 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n9/pdf/scientificamerican09011883-
138b.pdf.

Anonymous:1883:IIbs

[4890] Anonymous. Index of inventions. Scientific American, 49(9): 139–140, September 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n9/pdf/scientificamerican09011883-
139a.pdf.

Anonymous:1883:PNPbg

[4891] Anonymous. Patent notes: Patents. Scientific American, 49 (9):140, September 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n9/pdf/scientificamerican09011883-
140.pdf.

Anonymous:1883:GAC

[4892] Anonymous. Gerard’s alternating current machine. Scientific American, 49(9):127, September 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n9/pdf/scientificamerican09011883-
127.pdf.

Anonymous:1883:SFO

[4893] Anonymous. Sugar and fodder, in one. Scientific American, 49 (9):128, September 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n9/pdf/scientificamerican09011883-
128.pdf.

Anonymous:1883:PON

[4894] Anonymous. Petroleum — the old in the new. Scientific American, 49(9):128–129, September 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n9/pdf/scientificamerican09011883-
128a.pdf.

Hovey:1883:AAAb

[4895] H. C. Hovey. The American Association for the Advancement of Science. Scientific American, 49(9):129–130, September 1, 1883. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v49/
n9/pdf/scientificamerican09011883-129.pdf.

Anonymous:1883:NBM

[4896] Anonymous. New boring machine. Scientific American, 49(9): 130, September 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n9/pdf/scientificamerican09011883-
130.pdf.

Anonymous:1883:VS

[4897] Anonymous. Velocipede sleigh. Scientific American, 49(9):130, September 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n9/pdf/scientificamerican09011883-130a.pdf.

Anonymous:1883:CTW

[4898] Anonymous. Combined trunk and wardrobe. Scientific American, 49(9):130, September 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n9/pdf/scientificamerican09011883-
130b.pdf.

Anonymous:1883:WRT

[4899] Anonymous. Wire rope towage. Scientific American, 49(9):131, September 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v49/n9/pdf/scientificamerican09011883-131.pdf.

Anonymous:1883:APS

[4900] Anonymous. Aspects of the planets for September. Scientific American, 49(9):132–133, September 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n9/pdf/scientificamerican09011883-
132.pdf.

Anonymous:1883:CDAa

[4901] Anonymous. Compound double acting horizontal engine. Scientific American, 49(9):134, September 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v49/n9/pdf/scientificamerican09011883-
134.pdf.

Anonymous:1883:EDF

[4902] Anonymous. Ephemera, or day flies. Scientific American, 49 (9):135, September 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n9/pdf/scientificamerican09011883-
135.pdf.

Anonymous:1883:EIbt

[4903] Anonymous. Engineering inventions. Scientific American, 49(9): 137, September 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n9/pdf/scientificamerican09011883-
137.pdf.

Anonymous:1883:MIdq

[4904] Anonymous. Mechanical inventions. Scientific American, 49(9): 137, September 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n9/pdf/scientificamerican09011883-
137a.pdf.

Anonymous:1883:AIbu

[4905] Anonymous. Agricultural inventions. Scientific American, 49(9): 137, September 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n9/pdf/scientificamerican09011883-
137b.pdf.

Anonymous:1883:MIdr

[4906] Anonymous. Miscellaneous inventions. Scientific American, 49 (9):137, September 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n9/pdf/scientificamerican09011883-
137c.pdf.

Anonymous:1883:CRm

[4907] Anonymous. Communications received. Scientific American, 49 (9):139, September 1, 1883. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v49/n9/pdf/scientificamerican09011883-
139.pdf.

Anonymous:1884:Ca

[4908] Anonymous. Correspondence. Scientific American, 50(1):7, January 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n1/pdf/scientificamerican01051884-7.pdf.

Anonymous:1884:BPa

[4909] Anonymous. Business and personal. Scientific American, 50(1):10, January 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n1/pdf/scientificamerican01051884-10d.pdf.

Anonymous:1884:NQa

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n1/pdf/scientificamerican01051884-11.pdf.

Anonymous:1884:BNBa

[4911] Anonymous. Books: New books and publications. Scientific American, 50(1):11, January 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n1/pdf/scientificamerican01051884-
11a.pdf.

Anonymous:1884:IIa

[4912] Anonymous. Index of inventions. Scientific American, 50(1):11–12, January 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n1/pdf/scientificamerican01051884-11b.pdf.

Anonymous:1884:PNPa

[4913] Anonymous. Patent notes: Patents. Scientific American, 50(1):13, January 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n1/pdf/scientificamerican01051884-13.pdf.

Anonymous:1884:NLH

[4914] Anonymous. New lighthouse, Hudson River. Scientific American, 50(1):1, January 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n1/pdf/scientificamerican01051884-
1.pdf.

Anonymous:1884:CIA

[4915] Anonymous. Converting iron articles into steel. Scientific American, 50(1):2, January 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n1/pdf/scientificamerican01051884-
2.pdf.

Anonymous:1884:HWD

[4916] Anonymous. Heavy work with the diamond drill. Scientific American, 50(1):2, January 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n1/pdf/scientificamerican01051884-
2a.pdf.

Anonymous:1884:DWG

[4917] Anonymous. Does not want to be grammared out of his invention. Scientific American, 50(1):2, January 5, 1884. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v50/
n1/pdf/scientificamerican01051884-2b.pdf.

Anonymous:1884:RMa

[4918] Anonymous. Robert Macfarlane. Scientific American, 50(1):2, January 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n1/pdf/scientificamerican01051884-2c.pdf.

Anonymous:1884:PPV

[4919] Anonymous. Pneumatic propulsion of vessels. Scientific American, 50(1):3–4, January 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n1/pdf/scientificamerican01051884-
3.pdf.

Anonymous:1884:SPa

[4920] Anonymous. Shovel plow. Scientific American, 50(1):4, January 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n1/pdf/scientificamerican01051884-4.pdf.

Anonymous:1884:DFS

[4921] Anonymous. Device for fastening saw-clamps, etc, to workbenches. Scientific American, 50(1):4, January 5, 1884. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v50/
n1/pdf/scientificamerican01051884-4a.pdf.

Anonymous:1884:FD

[4922] Anonymous. Field drag. Scientific American, 50(1):4, January 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n1/pdf/scientificamerican01051884-4b.pdf.

Anonymous:1884:CSa

[4923] Anonymous. Cooking steamer. Scientific American, 50(1):5, January 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n1/pdf/scientificamerican01051884-5.pdf.

Anonymous:1884:RCS

[4924] Anonymous. Regulator and condenser for steam pumps. Scientific American, 50(1):5–6, January 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n1/pdf/scientificamerican01051884-
5a.pdf.

Anonymous:1884:GT

[4925] Anonymous. Game table. Scientific American, 50(1):6, January 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n1/pdf/scientificamerican01051884-6.pdf.

Anonymous:1884:RRMa

[4926] Anonymous. Railway riveting machine. Scientific American, 50(1):8–9, January 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n1/pdf/scientificamerican01051884-8.pdf.

Anonymous:1884:EIa

[4927] Anonymous. Engineering inventions. Scientific American, 50(1):10, January 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n1/pdf/scientificamerican01051884-10.pdf.

Anonymous:1884:MIa

[4928] Anonymous. Mechanical inventions. Scientific American, 50(1):10, January 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n1/pdf/scientificamerican01051884-10a.pdf.

Anonymous:1884:AIa

[4929] Anonymous. Agricultural inventions. Scientific American, 50(1):10, January 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n1/pdf/scientificamerican01051884-10b.pdf.

Anonymous:1884:MIb

[4930] Anonymous. Miscellaneous inventions. Scientific American, 50(1):10, January 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n1/pdf/scientificamerican01051884-10c.pdf.

Anonymous:1884:Cb

[4931] Anonymous. Correspondence. Scientific American, 50(10):149–150, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n10/pdf/scientificamerican03081884-149.pdf.

Anonymous:1884:BNBb

[4932] Anonymous. Books: New books and publications. Scientific American, 50(10):154, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n10/pdf/scientificamerican03081884-
154d.pdf.

Anonymous:1884:BPb

[4933] Anonymous. Business and personal. Scientific American, 50(10):154, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n10/pdf/scientificamerican03081884-154e.pdf.

Anonymous:1884:NQb

[4934] Anonymous. Notes and queries. Scientific American, 50(10):154–155, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n10/pdf/scientificamerican03081884-154f.pdf.

Anonymous:1884:IIb

[4935] Anonymous. Index of inventions. Scientific American, 50(10):155–156, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n10/pdf/scientificamerican03081884-155.pdf.

Anonymous:1884:PNPb

[4936] Anonymous. Patent notes: Patents. Scientific American, 50(10):157, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n10/pdf/scientificamerican03081884-157.pdf.

Anonymous:1884:GLW

[4937] Anonymous. Gigantic lifting wheels for the copper mines. Scientific American, 50(10):143–144, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v50/n10/pdf/scientificamerican03081884-
143.pdf.

Anonymous:1884:PSE

[4938] Anonymous. Present steam engine practice. Scientific American, 50(10):144–145, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n10/pdf/scientificamerican03081884-
144.pdf.

Anonymous:1884:FSS

[4939] Anonymous. Forming spiral springs. Scientific American, 50(10):145, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n10/pdf/scientificamerican03081884-145.pdf.

Anonymous:1884:HPBa

[4940] Anonymous. Hostile patent bills now before Congress. Scientific American, 50(10):145, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n10/pdf/scientificamerican03081884-
145a.pdf.

Anonymous:1884:NSb

[4941] Anonymous. The Nile and the Soudan. Scientific American, 50(10):145, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n10/pdf/scientificamerican03081884-145b.pdf.

Anonymous:1884:DRPa

[4942] Anonymous. Decisions relating to patents. Scientific American, 50(10):145, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n10/pdf/scientificamerican03081884-
145c.pdf.

Anonymous:1884:USH

[4943] Anonymous. Uniformity of shop heat. Scientific American, 50 (10):145–146, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n10/pdf/scientificamerican03081884-
145d.pdf.

Anonymous:1884:ISCa

[4944] Anonymous. Improved steam cooker. Scientific American, 50(10):146, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n10/pdf/scientificamerican03081884-146.pdf.

Anonymous:1884:TLP

[4945] Anonymous. Trunk life preserver. Scientific American, 50(10):146, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n10/pdf/scientificamerican03081884-146a.pdf.

Anonymous:1884:IHCa

[4946] Anonymous. Improved hose coupling. Scientific American, 50 (10):146–147, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n10/pdf/scientificamerican03081884-
146b.pdf.

Anonymous:1884:INL

[4947] Anonymous. Improved nut lock. Scientific American, 50(10):147, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n10/pdf/scientificamerican03081884-147.pdf.

Anonymous:1884:BDS

[4948] Anonymous. Boring, drilling, and surfacing machine. Scientific American, 50(10):147–148, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n10/pdf/scientificamerican03081884-
147a.pdf.

Anonymous:1884:DTCa

[4949] Anonymous. Dredger on the Tancarville Canal. Scientific American, 50(10):150, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n10/pdf/scientificamerican03081884-
150.pdf.

Anonymous:1884:NV

[4950] Anonymous. A novel velocipede. Scientific American, 50(10):150, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n10/pdf/scientificamerican03081884-150a.pdf.

Anonymous:1884:MPa

[4951] Anonymous. Max Von Pettenkofer. Scientific American, 50(10): 151–152, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n10/pdf/scientificamerican03081884-
151.pdf.

Anonymous:1884:RBE

[4952] Anonymous. A recent boiler explosion at Cincinnati, Ohio. Scientific American, 50(10):152–153, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v50/n10/pdf/scientificamerican03081884-
152.pdf.

Anonymous:1884:EIb

[4953] Anonymous. Engineering inventions. Scientific American, 50(10):154, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n10/pdf/scientificamerican03081884-154.pdf.

Anonymous:1884:MIc

[4954] Anonymous. Mechanical inventions. Scientific American, 50(10):154, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n10/pdf/scientificamerican03081884-154a.pdf.

Anonymous:1884:AIb

[4955] Anonymous. Agricultural inventions. Scientific American, 50(10):154, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n10/pdf/scientificamerican03081884-154b.pdf.

Anonymous:1884:MId

[4956] Anonymous. Miscellaneous inventions. Scientific American, 50(10):154, March 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n10/pdf/scientificamerican03081884-154c.pdf.

Anonymous:1884:Cc

[4957] Anonymous. Correspondence. Scientific American, 50(11):165, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n11/pdf/scientificamerican03151884-165.pdf.

Anonymous:1884:BPc

[4958] Anonymous. Business and personal. Scientific American, 50(11):170, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n11/pdf/scientificamerican03151884-170d.pdf.

Anonymous:1884:BNBc

[4959] Anonymous. Books: New books and publications. Scientific American, 50(11):170, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n11/pdf/scientificamerican03151884-
170f.pdf.

Anonymous:1884:NQc

[4960] Anonymous. Notes and queries. Scientific American, 50(11):170–171, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n11/pdf/scientificamerican03151884-170g.pdf.

Anonymous:1884:IIc

[4961] Anonymous. Index of inventions. Scientific American, 50(11):171–172, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n11/pdf/scientificamerican03151884-171.pdf.

Anonymous:1884:PNPc

[4962] Anonymous. Patent notes: Patents. Scientific American, 50(11):173, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n11/pdf/scientificamerican03151884-173.pdf.

Hovey:1884:VFS

[4963] H. C. Hovey. The viaduct at the Falls of St. Anthony. Scientific American, 50(11):159–160, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n11/pdf/scientificamerican03151884-
159.pdf.

Anonymous:1884:AHJ

[4964] Anonymous. America a huge “jelly fish”. Scientific American, 50(11):160, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n11/pdf/scientificamerican03151884-
160.pdf.

Anonymous:1884:VIW

[4965] Anonymous. Value of inventions to the world. Scientific American, 50(11):160, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n11/pdf/scientificamerican03151884-
160a.pdf.

Anonymous:1884:PE

[4966] Anonymous. Progress in exactness. Scientific American, 50(11): 160–161, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n11/pdf/scientificamerican03151884-
160b.pdf.

Anonymous:1884:DRPb

[4967] Anonymous. Decisions relating to patents. Scientific American, 50(11):161, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n11/pdf/scientificamerican03151884-
161.pdf.

Anonymous:1884:PPB

[4968] Anonymous. Prospects of the Patent Bills. Scientific American, 50(11):161, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n11/pdf/scientificamerican03151884-
161a.pdf.

Anonymous:1884:PPa

[4969] Anonymous. Patents and politics. Scientific American, 50(11):161, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n11/pdf/scientificamerican03151884-161b.pdf.

Anonymous:1884:TTM

[4970] Anonymous. The triple thermic motor. Scientific American, 50(11):161, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n11/pdf/scientificamerican03151884-161c.pdf.

Anonymous:1884:IKS

[4971] Anonymous. Improved knife scourer. Scientific American, 50(11):162, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n11/pdf/scientificamerican03151884-162.pdf.

Anonymous:1884:RHP

[4972] Anonymous. The Remington horse power fire engine. Scientific American, 50(11):162, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n11/pdf/scientificamerican03151884-
162a.pdf.

Anonymous:1884:ILL

[4973] Anonymous. Improved lock and latch. Scientific American, 50(11):162, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n11/pdf/scientificamerican03151884-162b.pdf.

Anonymous:1884:EDN

[4974] Anonymous. The electric diadems of the new ballet “La Farandole”. Scientific American, 50(11):163, March 15, 1884. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v50/
n11/pdf/scientificamerican03151884-163.pdf.

Anonymous:1884:DBS

[4975] Anonymous. Dujour and Bianchi’s single axle windlass. Scientific American, 50(11):163–164, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n11/pdf/scientificamerican03151884-
163a.pdf.

Anonymous:1884:PBa

[4976] Anonymous. Petroleum at Baku. Scientific American, 50(11):166–167, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n11/pdf/scientificamerican03151884-166.pdf.

Gbatacap:1884:O

[4977] L. P. Gbatacap. The ornithorhynchus. Scientific American, 50 (11):167–168, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n11/pdf/scientificamerican03151884-
167.pdf.

Anonymous:1884:IRS

[4978] Anonymous. Improved rotating sample stand. Scientific American, 50(11):168, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n11/pdf/scientificamerican03151884-
168.pdf.

Anonymous:1884:IPa

[4979] Anonymous. Improved pail. Scientific American, 50(11):168, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n11/pdf/scientificamerican03151884-168a.pdf.

Anonymous:1884:IML

[4980] Anonymous. Improved music leaf turner. Scientific American, 50 (11):168–169, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n11/pdf/scientificamerican03151884-
168b.pdf.

Anonymous:1884:EIc

[4981] Anonymous. Engineering inventions. Scientific American, 50(11):170, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n11/pdf/scientificamerican03151884-170.pdf.

Anonymous:1884:MIe

[4982] Anonymous. Mechanical inventions. Scientific American, 50(11):170, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n11/pdf/scientificamerican03151884-170a.pdf.

Anonymous:1884:AIc

[4983] Anonymous. Agricultural inventions. Scientific American, 50(11):170, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n11/pdf/scientificamerican03151884-170b.pdf.

Anonymous:1884:MIf

[4984] Anonymous. Miscellaneous inventions. Scientific American, 50(11):170, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n11/pdf/scientificamerican03151884-170c.pdf.

Anonymous:1884:Sb

[4985] Anonymous. Special. Scientific American, 50(11):170, March 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n11/pdf/scientificamerican03151884-170e.pdf.

Anonymous:1884:Cd

[4986] Anonymous. Correspondence. Scientific American, 50(12):181–182, March 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n12/pdf/scientificamerican03221884-181.pdf.

Anonymous:1884:BPd

[4987] Anonymous. Business and personal. Scientific American, 50(12):186, March 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n12/pdf/scientificamerican03221884-186d.pdf.

Anonymous:1884:BNBd

[4988] Anonymous. Books: New books and publications. Scientific American, 50(12):186, March 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n12/pdf/scientificamerican03221884-
186e.pdf.

Anonymous:1884:NQd

[4989] Anonymous. Notes and queries. Scientific American, 50(12):186–187, March 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n12/pdf/scientificamerican03221884-186f.pdf.

Anonymous:1884:IId

[4990] Anonymous. Index of inventions. Scientific American, 50(12):187–188, March 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n12/pdf/scientificamerican03221884-187.pdf.

Anonymous:1884:PNPd

[4991] Anonymous. Patent notes: Patents. Scientific American, 50(12):189, March 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n12/pdf/scientificamerican03221884-189.pdf.

Anonymous:1884:NEL

[4992] Anonymous. The new electric light at Hell Gate Lighthouse. Scientific American, 50(12):175, March 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n12/pdf/scientificamerican03221884-
175.pdf.

Anonymous:1884:PAP

[4993] Anonymous. The plot against patents. Scientific American, 50(12):176, March 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n12/pdf/scientificamerican03221884-176.pdf.

Anonymous:1884:CRN

[4994] Anonymous. Cable railways for New York City. Scientific American, 50(12):176, March 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n12/pdf/scientificamerican03221884-
176a.pdf.

Anonymous:1884:MFa

[4995] Anonymous. Malarial fevers. Scientific American, 50(12):176–177, March 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n12/pdf/scientificamerican03221884-176b.pdf.

Anonymous:1884:HPBb

[4996] Anonymous. The hostile Patent Bills. Scientific American, 50(12):177, March 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n12/pdf/scientificamerican03221884-177.pdf.

Anonymous:1884:WLD

[4997] Anonymous. Water line defense and gun shields for cruisers. Scientific American, 50(12):177, March 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n12/pdf/scientificamerican03221884-
177a.pdf.

Anonymous:1884:MRT

[4998] Anonymous. Metallic railroad tie. Scientific American, 50(12):178, March 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n12/pdf/scientificamerican03221884-178.pdf.

Anonymous:1884:IAT

[4999] Anonymous. Improved acoustic telephone. Scientific American, 50(12):179, March 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n12/pdf/scientificamerican03221884-
179.pdf.

Anonymous:1884:ICCa

[5000] Anonymous. Improved car coupling. Scientific American, 50(12):179, March 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n12/pdf/scientificamerican03221884-179a.pdf.

Anonymous:1884:ILB

[5001] Anonymous. Improvement in the Leclanche battery. Scientific American, 50(12):179–180, March 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n12/pdf/scientificamerican03221884-
179b.pdf.

Anonymous:1884:IHT

[5002] Anonymous. Improved hand truck. Scientific American, 50(12):180, March 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n12/pdf/scientificamerican03221884-180.pdf.

Anonymous:1884:ACC

[5003] Anonymous. Another curious case of freezing. Scientific American, 50(12):182, March 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n12/pdf/scientificamerican03221884-
182.pdf.

Anonymous:1884:OTW

[5004] Anonymous. Oil on troubled waters. Scientific American, 50(12):182, March 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n12/pdf/scientificamerican03221884-182a.pdf.

Anonymous:1884:TTP

[5005] Anonymous. The twelve thread plume bird. Scientific American, 50(12):183–185, March 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n12/pdf/scientificamerican03221884-
183.pdf.

Anonymous:1884:EId

[5006] Anonymous. Engineering inventions. Scientific American, 50(12):186, March 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n12/pdf/scientificamerican03221884-186.pdf.

Anonymous:1884:MIg

[5007] Anonymous. Mechanical inventions. Scientific American, 50(12):186, March 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n12/pdf/scientificamerican03221884-186a.pdf.

Anonymous:1884:AId

[5008] Anonymous. Agricultural inventions. Scientific American, 50(12):186, March 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n12/pdf/scientificamerican03221884-186b.pdf.

Anonymous:1884:MIh

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n12/pdf/scientificamerican03221884-186c.pdf.

Anonymous:1884:Ce

[5010] Anonymous. Correspondence. Scientific American, 50(13):197, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n13/pdf/scientificamerican03291884-197.pdf.

Anonymous:1884:BPe

[5011] Anonymous. Business and personal. Scientific American, 50 (13):201–202, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n13/pdf/scientificamerican03291884-
201e.pdf.

Anonymous:1884:BNBe

[5012] Anonymous. Books: New books and publications. Scientific American, 50(13):202, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n13/pdf/scientificamerican03291884-
202.pdf.

Anonymous:1884:NQe

[5013] Anonymous. Notes and queries. Scientific American, 50(13):202–203, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n13/pdf/scientificamerican03291884-202a.pdf.

Anonymous:1884:IIe

[5014] Anonymous. Index of inventions. Scientific American, 50(13):203–204, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n13/pdf/scientificamerican03291884-203.pdf.

Anonymous:1884:PNPe

[5015] Anonymous. Patent notes: Patents. Scientific American, 50(13):205, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n13/pdf/scientificamerican03291884-205.pdf.

Anonymous:1884:TES

[5016] Anonymous. Traction engine with spring wheels. Scientific American, 50(13):191–192, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n13/pdf/scientificamerican03291884-
191.pdf.

Anonymous:1884:PBBa

[5017] Anonymous. The Patent Bills before Congress. Scientific American, 50(13):192, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n13/pdf/scientificamerican03291884-
192.pdf.

Anonymous:1884:ICPa

[5018] Anonymous. Items concerning the Patent Bills. Scientific American, 50(13):193, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n13/pdf/scientificamerican03291884-
193.pdf.

Anonymous:1884:FPa

[5019] Anonymous. Finishing by pressure. Scientific American, 50(13):193, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n13/pdf/scientificamerican03291884-193a.pdf.

Anonymous:1884:LSE

[5020] Anonymous. A long straight edge. Scientific American, 50(13):193, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n13/pdf/scientificamerican03291884-193b.pdf.

Anonymous:1884:SCFa

[5021] Anonymous. Sash-cord fastener and window sash. Scientific American, 50(13):194, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n13/pdf/scientificamerican03291884-
194.pdf.

Anonymous:1884:ILa

[5022] Anonymous. An improved lock. Scientific American, 50(13):194, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n13/pdf/scientificamerican03291884-194a.pdf.

Anonymous:1884:IRDa

[5023] Anonymous. Improved ratchet drill. Scientific American, 50(13):194, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n13/pdf/scientificamerican03291884-194b.pdf.

Anonymous:1884:ITS

[5024] Anonymous. Improved truck for sugar refineries. Scientific American, 50(13):194, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n13/pdf/scientificamerican03291884-
194c.pdf.

Anonymous:1884:SAa

[5025] Anonymous. Science in antiquity. Scientific American, 50(13):195, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n13/pdf/scientificamerican03291884-195.pdf.

Anonymous:1884:APA

[5026] Anonymous. Aspects of the planets for April. Scientific American, 50(13):196, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n13/pdf/scientificamerican03291884-
196.pdf.

Anonymous:1884:IER

[5027] Anonymous. An improved electric railway signal. Scientific American, 50(13):198, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n13/pdf/scientificamerican03291884-
198.pdf.

Anonymous:1884:NWE

[5028] Anonymous. A new wind engine. Scientific American, 50(13):198, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n13/pdf/scientificamerican03291884-198a.pdf.

Anonymous:1884:CWa

[5029] Anonymous. A combination wardrobe. Scientific American, 50(13):199, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n13/pdf/scientificamerican03291884-199.pdf.

Anonymous:1884:BBH

[5030] Anonymous. The boat bill heron. Scientific American, 50(13):199–200, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n13/pdf/scientificamerican03291884-199a.pdf.

Anonymous:1884:EIe

[5031] Anonymous. Engineering inventions. Scientific American, 50(13):201, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n13/pdf/scientificamerican03291884-201.pdf.

Anonymous:1884:MIi

[5032] Anonymous. Mechanical inventions. Scientific American, 50(13):201, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n13/pdf/scientificamerican03291884-201a.pdf.

Anonymous:1884:AIe

[5033] Anonymous. Agricultural inventions. Scientific American, 50(13):201, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n13/pdf/scientificamerican03291884-201b.pdf.

Anonymous:1884:MIj

[5034] Anonymous. Miscellaneous inventions. Scientific American, 50(13):201, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n13/pdf/scientificamerican03291884-201c.pdf.

Anonymous:1884:Sc

[5035] Anonymous. Special. Scientific American, 50(13):201, March 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n13/pdf/scientificamerican03291884-201d.pdf.

Anonymous:1884:Cf

[5036] Anonymous. Correspondence. Scientific American, 50(14):213, April 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n14/pdf/scientificamerican04051884-213.pdf.

Anonymous:1884:BNBf

[5037] Anonymous. Books: New books and publications. Scientific American, 50(14):217, April 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n14/pdf/scientificamerican04051884-
217d.pdf.

Anonymous:1884:BPf

[5038] Anonymous. Business and personal. Scientific American, 50 (14):217–218, April 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n14/pdf/scientificamerican04051884-
217e.pdf.

Anonymous:1884:NQf

[5039] Anonymous. Notes and queries. Scientific American, 50(14):218–219, April 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n14/pdf/scientificamerican04051884-218.pdf.

Anonymous:1884:IIf

[5040] Anonymous. Index of inventions. Scientific American, 50(14):219–220, April 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n14/pdf/scientificamerican04051884-219.pdf.

Anonymous:1884:PNPf

[5041] Anonymous. Patent notes: Patents. Scientific American, 50(14):221, April 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n14/pdf/scientificamerican04051884-221.pdf.

Anonymous:1884:MTS

[5042] Anonymous. Movable theater stages. Scientific American, 50 (14):207–208, April 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n14/pdf/scientificamerican04051884-
207.pdf.

Anonymous:1884:PDB

[5043] Anonymous. Probable defeat of the bad Patent Bills. Scientific American, 50(14):208–209, April 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n14/pdf/scientificamerican04051884-
208.pdf.

Anonymous:1884:ELH

[5044] Anonymous. Extracts from the letter of the Hon. Benjamin Butterworth, Commissioner of Patents, to the Cincinnati Convention of Inventors. Scientific American, 50(14):209–210, April 5, 1884. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v50/
n14/pdf/scientificamerican04051884-209.pdf.

Anonymous:1884:IR

[5045] Anonymous. Improved refrigerator. Scientific American, 50(14):210, April 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n14/pdf/scientificamerican04051884-210.pdf.

Anonymous:1884:IEEa

[5046] Anonymous. The International Electrical Exhibition. Scientific American, 50(14):210, April 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n14/pdf/scientificamerican04051884-
210a.pdf.

Anonymous:1884:ISCb

[5047] Anonymous. Improved ship cleaner. Scientific American, 50 (14):210–211, April 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n14/pdf/scientificamerican04051884-
210b.pdf.

Anonymous:1884:RDH

[5048] Anonymous. The Remington drop hammer. Scientific American, 50(14):211, April 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n14/pdf/scientificamerican04051884-
211.pdf.

Anonymous:1884:ISE

[5049] Anonymous. Improved steam engine. Scientific American, 50 (14):211–212, April 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n14/pdf/scientificamerican04051884-
211a.pdf.

Anonymous:1884:ISHa

[5050] Anonymous. Improved steel horse collar. Scientific American, 50 (14):212, April 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n14/pdf/scientificamerican04051884-
212.pdf.

Anonymous:1884:NDG

[5051] Anonymous. The new dynamite gun. Scientific American, 50 (14):214–215, April 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n14/pdf/scientificamerican04051884-
214.pdf.

Anonymous:1884:JJO

[5052] Anonymous. Jade and jade objects from Central America. Scientific American, 50(14):215–216, April 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v50/n14/pdf/scientificamerican04051884-
215.pdf.

Anonymous:1884:EIf

[5053] Anonymous. Engineering inventions. Scientific American, 50(14):217, April 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n14/pdf/scientificamerican04051884-217.pdf.

Anonymous:1884:MIk

[5054] Anonymous. Mechanical inventions. Scientific American, 50(14):217, April 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n14/pdf/scientificamerican04051884-217a.pdf.

Anonymous:1884:AIf

[5055] Anonymous. Agricultural inventions. Scientific American, 50(14):217, April 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n14/pdf/scientificamerican04051884-217b.pdf.

Anonymous:1884:MIl

[5056] Anonymous. Miscellaneous inventions. Scientific American, 50(14):217, April 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n14/pdf/scientificamerican04051884-217c.pdf.

Anonymous:1884:Cg

[5057] Anonymous. Correspondence. Scientific American, 50(15):229–230, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n15/pdf/scientificamerican04121884-229.pdf.

Anonymous:1884:BNBg

[5058] Anonymous. Books: New books and publications. Scientific American, 50(15):233, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n15/pdf/scientificamerican04121884-
233d.pdf.

Anonymous:1884:BPg

[5059] Anonymous. Business and personal. Scientific American, 50(15):233, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n15/pdf/scientificamerican04121884-233f.pdf.

Anonymous:1884:NQg

[5060] Anonymous. Notes and queries. Scientific American, 50(15):233–235, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n15/pdf/scientificamerican04121884-233g.pdf.

Anonymous:1884:IIg

[5061] Anonymous. Index of inventions. Scientific American, 50(15):235–236, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n15/pdf/scientificamerican04121884-235.pdf.

Anonymous:1884:PNPg

[5062] Anonymous. Patent notes: Patents. Scientific American, 50(15):237, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n15/pdf/scientificamerican04121884-237.pdf.

Anonymous:1884:IBF

[5063] Anonymous. Increasing the bridge facilities. Scientific American, 50(15):223–224, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n15/pdf/scientificamerican04121884-
223.pdf.

Anonymous:1884:PCa

[5064] Anonymous. Patents in Congress. Scientific American, 50(15):224–225, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n15/pdf/scientificamerican04121884-224.pdf.

Anonymous:1884:MFb

[5065] Anonymous. Malarial fevers. Scientific American, 50(15):225, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n15/pdf/scientificamerican04121884-225.pdf.

Anonymous:1884:NSA

[5066] Anonymous. New Scientific American offices. Scientific American, 50(15):225, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n15/pdf/scientificamerican04121884-
225a.pdf.

Anonymous:1884:US

[5067] Anonymous. Uneven shrinking. Scientific American, 50(15):225, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n15/pdf/scientificamerican04121884-225b.pdf.

Anonymous:1884:RBT

[5068] Anonymous. Removable bottle trap. Scientific American, 50(15):226, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n15/pdf/scientificamerican04121884-226.pdf.

Anonymous:1884:CPa

[5069] Anonymous. Corn planter. Scientific American, 50(15):226, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n15/pdf/scientificamerican04121884-226a.pdf.

Anonymous:1884:IFPa

[5070] Anonymous. Improved fence post. Scientific American, 50(15):226, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n15/pdf/scientificamerican04121884-226b.pdf.

Anonymous:1884:IVa

[5071] Anonymous. Improved valve. Scientific American, 50(15):226–227, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n15/pdf/scientificamerican04121884-226c.pdf.

Anonymous:1884:Pa

[5072] Anonymous. Panolastite. Scientific American, 50(15):227–228, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n15/pdf/scientificamerican04121884-227.pdf.

Anonymous:1884:IWM

[5073] Anonymous. Improved warping machine. Scientific American, 50 (15):230–231, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n15/pdf/scientificamerican04121884-
230.pdf.

Anonymous:1884:CFU

[5074] Anonymous. Central France under the clouds. Scientific American, 50(15):231–232, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n15/pdf/scientificamerican04121884-
231.pdf.

Anonymous:1884:ICa

[5075] Anonymous. An improved crutch. Scientific American, 50(15):232, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n15/pdf/scientificamerican04121884-232.pdf.

Anonymous:1884:WBF

[5076] Anonymous. Window bead fastener. Scientific American, 50(15):232, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n15/pdf/scientificamerican04121884-232a.pdf.

Anonymous:1884:RASa

[5077] Anonymous. Ripping attachment for scissors. Scientific American, 50(15):232, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n15/pdf/scientificamerican04121884-
232b.pdf.

Anonymous:1884:EIg

[5078] Anonymous. Engineering inventions. Scientific American, 50(15):233, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n15/pdf/scientificamerican04121884-233.pdf.

Anonymous:1884:MIm

[5079] Anonymous. Mechanical inventions. Scientific American, 50(15):233, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n15/pdf/scientificamerican04121884-233a.pdf.

Anonymous:1884:AIg

[5080] Anonymous. Agricultural inventions. Scientific American, 50(15):233, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n15/pdf/scientificamerican04121884-233b.pdf.

Anonymous:1884:MIn

[5081] Anonymous. Miscellaneous inventions. Scientific American, 50(15):233, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n15/pdf/scientificamerican04121884-233c.pdf.

Anonymous:1884:Sd

[5082] Anonymous. Special. Scientific American, 50(15):233, April 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n15/pdf/scientificamerican04121884-233e.pdf.

Anonymous:1884:Ch

[5083] Anonymous. Correspondence. Scientific American, 50(16):245–246, April 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n16/pdf/scientificamerican04191884-245.pdf.

Anonymous:1884:BPh

[5084] Anonymous. Business and personal. Scientific American, 50(16):249, April 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n16/pdf/scientificamerican04191884-249c.pdf.

Anonymous:1884:BNBh

[5085] Anonymous. Books: New books and publications. Scientific American, 50(16):249, April 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n16/pdf/scientificamerican04191884-
249d.pdf.

Anonymous:1884:NQh

[5086] Anonymous. Notes and queries. Scientific American, 50(16):249–251, April 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n16/pdf/scientificamerican04191884-249e.pdf.

Anonymous:1884:IIh

[5087] Anonymous. Index of inventions. Scientific American, 50(16):251–252, April 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n16/pdf/scientificamerican04191884-251.pdf.

Anonymous:1884:PNPh

[5088] Anonymous. Patent notes: Patents. Scientific American, 50(16):253, April 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n16/pdf/scientificamerican04191884-253.pdf.

Anonymous:1884:DPLa

[5089] Anonymous. Dr. Le Plongeon’s latest and most important discoveries among the ruined cities of Yucatan. Scientific American, 50 (16):239–240, April 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n16/pdf/scientificamerican04191884-
239.pdf.

Anonymous:1884:DAPa

[5090] Anonymous. Dr. Augustus Le Plongeon’s latest and most remarkable discoveries in Yucatan. Scientific American, 50(16):240, April 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n16/pdf/scientificamerican04191884-240.pdf.

Anonymous:1884:PM

[5091] Anonymous. Protests from manufacturers. Scientific American, 50 (16):240–241, April 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n16/pdf/scientificamerican04191884-
240a.pdf.

Anonymous:1884:IMMa

[5092] Anonymous. Improvements in men and machines. Scientific American, 50(16):241, April 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n16/pdf/scientificamerican04191884-
241.pdf.

Anonymous:1884:ASB

[5093] Anonymous. American standard for boilers. Scientific American, 50(16):241, April 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n16/pdf/scientificamerican04191884-
241a.pdf.

Anonymous:1884:SI

[5094] Anonymous. Suggestions for inventors. Scientific American, 50(16):241, April 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n16/pdf/scientificamerican04191884-241b.pdf.

Anonymous:1884:CCa

[5095] Anonymous. The cheapness of cost. Scientific American, 50 (16):241–242, April 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n16/pdf/scientificamerican04191884-
241c.pdf.

Anonymous:1884:NLC

[5096] Anonymous. New lathe chuck. Scientific American, 50(16):242, April 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n16/pdf/scientificamerican04191884-242.pdf.

Anonymous:1884:ITB

[5097] Anonymous. Improved torpedo boat. Scientific American, 50(16):242, April 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n16/pdf/scientificamerican04191884-242a.pdf.

Anonymous:1884:MBP

[5098] Anonymous. Memorandum book and pad. Scientific American, 50(16):242–243, April 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n16/pdf/scientificamerican04191884-
242b.pdf.

Anonymous:1884:ATa

[5099] Anonymous. The atmospheric turbine. Scientific American, 50(16):243, April 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n16/pdf/scientificamerican04191884-243.pdf.

Anonymous:1884:STa

[5100] Anonymous. Saw tooth. Scientific American, 50(16):243–245, April 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n16/pdf/scientificamerican04191884-243a.pdf.

Anonymous:1884:SSS

[5101] Anonymous. The sectional steamer Le Stanley. Scientific American, 50(16):246–248, April 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n16/pdf/scientificamerican04191884-
246.pdf.

Anonymous:1884:ATb

[5102] Anonymous. Animal trap. Scientific American, 50(16):248, April 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n16/pdf/scientificamerican04191884-248.pdf.

Anonymous:1884:FEa

[5103] Anonymous. Fire escape. Scientific American, 50(16):248–249, April 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n16/pdf/scientificamerican04191884-248a.pdf.

Anonymous:1884:EIh

[5104] Anonymous. Engineering inventions. Scientific American, 50(16):249, April 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n16/pdf/scientificamerican04191884-249.pdf.

Anonymous:1884:MIo

[5105] Anonymous. Mechanical inventions. Scientific American, 50(16):249, April 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n16/pdf/scientificamerican04191884-249a.pdf.

Anonymous:1884:MIp

[5106] Anonymous. Miscellaneous inventions. Scientific American, 50(16):249, April 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n16/pdf/scientificamerican04191884-249b.pdf.

Anonymous:1884:BPi

[5107] Anonymous. Business and personal. Scientific American, 50 (17):266–267, April 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n17/pdf/scientificamerican04261884-
266e.pdf.

Anonymous:1884:BNBi

[5108] Anonymous. Books: New books and publications. Scientific American, 50(17):267, April 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n17/pdf/scientificamerican04261884-
267.pdf.

Anonymous:1884:NQi

[5109] Anonymous. Notes and queries. Scientific American, 50(17):267, April 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n17/pdf/scientificamerican04261884-267a.pdf.

Anonymous:1884:IIi

[5110] Anonymous. Index of inventions. Scientific American, 50(17):267–268, April 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n17/pdf/scientificamerican04261884-267b.pdf.

Anonymous:1884:PNPi

[5111] Anonymous. Patent notes: Patents. Scientific American, 50(17):269, April 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n17/pdf/scientificamerican04261884-269.pdf.

Anonymous:1884:BCC

[5112] Anonymous. The bridge over Catskill Creek. Scientific American, 50(17):255–256, April 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n17/pdf/scientificamerican04261884-
255.pdf.

Anonymous:1884:SOI

[5113] Anonymous. Separation of oil from iron chips. Scientific American, 50(17):256, April 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n17/pdf/scientificamerican04261884-
256.pdf.

Anonymous:1884:PIF

[5114] Anonymous. Progress in iron founding. Scientific American, 50(17):256, April 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n17/pdf/scientificamerican04261884-256a.pdf.

Anonymous:1884:SCO

[5115] Anonymous. Saving by co-operation. Scientific American, 50(17):257, April 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n17/pdf/scientificamerican04261884-257.pdf.

Anonymous:1884:NFEa

[5116] Anonymous. A new fire escape. Scientific American, 50(17):258, April 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n17/pdf/scientificamerican04261884-258.pdf.

Anonymous:1884:IHCb

[5117] Anonymous. Improved hose coupling. Scientific American, 50(17):258, April 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n17/pdf/scientificamerican04261884-258a.pdf.

Anonymous:1884:BPEa

[5118] Anonymous. Binding post for electric conductors. Scientific American, 50(17):258, April 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n17/pdf/scientificamerican04261884-
258b.pdf.

Anonymous:1884:NRC

[5119] Anonymous. New reclining chair. Scientific American, 50(17):258–259, April 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n17/pdf/scientificamerican04261884-258c.pdf.

Anonymous:1884:MW

[5120] Anonymous. Musical wagon. Scientific American, 50(17):259, April 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n17/pdf/scientificamerican04261884-259.pdf.

Anonymous:1884:OGP

[5121] Anonymous. Oil gas plant for Ailsa Craig Lighthouse. Scientific American, 50(17):259–262, April 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n17/pdf/scientificamerican04261884-
259a.pdf.

Anonymous:1884:DPLb

[5122] Anonymous. Dr. Le Plongeon’s latest and most important discoveries among the ruined cities of Yucatan. Scientific American, 50 (17):262–264, April 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n17/pdf/scientificamerican04261884-
262.pdf.

Anonymous:1884:BPj

[5123] Anonymous. Butter package. Scientific American, 50(17):264, April 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n17/pdf/scientificamerican04261884-264.pdf.

Anonymous:1884:BCT

[5124] Anonymous. Billiard cue tip. Scientific American, 50(17):264, April 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n17/pdf/scientificamerican04261884-264a.pdf.

Anonymous:1884:NGT

[5125] Anonymous. New game table. Scientific American, 50(17):264–265, April 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n17/pdf/scientificamerican04261884-264b.pdf.

Anonymous:1884:EIi

[5126] Anonymous. Engineering inventions. Scientific American, 50(17):266, April 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n17/pdf/scientificamerican04261884-266.pdf.

Anonymous:1884:MIq

[5127] Anonymous. Mechanical inventions. Scientific American, 50(17):266, April 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n17/pdf/scientificamerican04261884-266a.pdf.

Anonymous:1884:AIh

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n17/pdf/scientificamerican04261884-266b.pdf.

Anonymous:1884:MIr

[5129] Anonymous. Miscellaneous inventions. Scientific American, 50(17):266, April 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n17/pdf/scientificamerican04261884-266c.pdf.

Anonymous:1884:Se

[5130] Anonymous. Special. Scientific American, 50(17):266, April 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n17/pdf/scientificamerican04261884-266d.pdf.

Anonymous:1884:Ci

[5131] Anonymous. Correspondence. Scientific American, 50(18):277–279, May 3, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n18/pdf/scientificamerican05031884-277.pdf.

Anonymous:1884:BPk

[5132] Anonymous. Business and personal. Scientific American, 50(18):282, May 3, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n18/pdf/scientificamerican05031884-282d.pdf.

Anonymous:1884:NQj

[5133] Anonymous. Notes and queries. Scientific American, 50(18):282–283, May 3, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n18/pdf/scientificamerican05031884-282e.pdf.

Anonymous:1884:IIj

[5134] Anonymous. Index of inventions. Scientific American, 50(18):283–284, May 3, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n18/pdf/scientificamerican05031884-283.pdf.

Anonymous:1884:PNPj

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n18/pdf/scientificamerican05031884-285.pdf.

Anonymous:1884:PIW

[5136] Anonymous. Plan to increase the water supply of New York. Scientific American, 50(18):271, May 3, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n18/pdf/scientificamerican05031884-
271.pdf.

Anonymous:1884:IPU

[5137] Anonymous. The International Patent Union. Scientific American, 50(18):272, May 3, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n18/pdf/scientificamerican05031884-
272.pdf.

Anonymous:1884:SCP

[5138] Anonymous. Status of the Congressional Patent Bills. Scientific American, 50(18):272, May 3, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n18/pdf/scientificamerican05031884-
272a.pdf.

Anonymous:1884:HIa

[5139] Anonymous. Hints to inventors. Scientific American, 50(18):272, May 3, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n18/pdf/scientificamerican05031884-272b.pdf.

Anonymous:1884:APMa

[5140] Anonymous. Aspects of the planets for May. Scientific American, 50(18):273, May 3, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n18/pdf/scientificamerican05031884-
273.pdf.

Anonymous:1884:NFEb

[5141] Anonymous. A new fire escape. Scientific American, 50(18):274, May 3, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n18/pdf/scientificamerican05031884-274.pdf.

Anonymous:1884:GBO

[5142] Anonymous. Guide bar oiler. Scientific American, 50(18):274, May 3, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n18/pdf/scientificamerican05031884-274a.pdf.

Anonymous:1884:HBC

[5143] Anonymous. Head for barrels, casks, etc. Scientific American, 50(18):274–275, May 3, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n18/pdf/scientificamerican05031884-
274b.pdf.

Anonymous:1884:TPM

[5144] Anonymous. Taking a photograph by the magnesium light. Scientific American, 50(18):275, May 3, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n18/pdf/scientificamerican05031884-
275.pdf.

Anonymous:1884:ADF

[5145] Anonymous. Apparatus for drying fruit. Scientific American, 50 (18):275, May 3, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n18/pdf/scientificamerican05031884-
275a.pdf.

Anonymous:1884:AEM

[5146] Anonymous. Apparatus for enlarging microscopic slides. Scientific American, 50(18):276, May 3, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n18/pdf/scientificamerican05031884-
276.pdf.

Anonymous:1884:ICPb

[5147] Anonymous. An industrial city — Pullman, Ill. Scientific American, 50(18):279–280, May 3, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n18/pdf/scientificamerican05031884-
279.pdf.

Anonymous:1884:NOEa

[5148] Anonymous. New ore elevator. Scientific American, 50(18):280, May 3, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n18/pdf/scientificamerican05031884-280.pdf.

Anonymous:1884:RLG

[5149] Anonymous. To the rescue of Lieut Greely. Scientific American, 50(18):281, May 3, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n18/pdf/scientificamerican05031884-
281.pdf.

Anonymous:1884:EIj

[5150] Anonymous. Engineering inventions. Scientific American, 50(18):282, May 3, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n18/pdf/scientificamerican05031884-282.pdf.

Anonymous:1884:MIs

[5151] Anonymous. Mechanical inventions. Scientific American, 50(18):282, May 3, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n18/pdf/scientificamerican05031884-282a.pdf.

Anonymous:1884:AIi

[5152] Anonymous. Agricultural inventions. Scientific American, 50(18):282, May 3, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n18/pdf/scientificamerican05031884-282b.pdf.

Anonymous:1884:MIt

[5153] Anonymous. Miscellaneous inventions. Scientific American, 50(18):282, May 3, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n18/pdf/scientificamerican05031884-282c.pdf.

Anonymous:1884:Cj

[5154] Anonymous. Correspondence. Scientific American, 50(19):293–294, May 10, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n19/pdf/scientificamerican05101884-293.pdf.

Anonymous:1884:BPl

[5155] Anonymous. Business and personal. Scientific American, 50(19):297, May 10, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n19/pdf/scientificamerican05101884-297d.pdf.

Anonymous:1884:NQk

[5156] Anonymous. Notes and queries. Scientific American, 50(19):297–299, May 10, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n19/pdf/scientificamerican05101884-297e.pdf.

Anonymous:1884:IIk

[5157] Anonymous. Index of inventions. Scientific American, 50(19):299–300, May 10, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n19/pdf/scientificamerican05101884-299.pdf.

Anonymous:1884:PNPk

[5158] Anonymous. Patent notes: Patents. Scientific American, 50(19):301, May 10, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n19/pdf/scientificamerican05101884-301.pdf.

Anonymous:1884:PFG

[5159] Anonymous. Patents to foreigners — Government purchase of patents. Scientific American, 50(19):288, May 10, 1884. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v50/
n19/pdf/scientificamerican05101884-288.pdf.

Anonymous:1884:SOW

[5160] Anonymous. A suggestion as to overhead wires. Scientific American, 50(19):288, May 10, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n19/pdf/scientificamerican05101884-
288a.pdf.

Anonymous:1884:MSSa

[5161] Anonymous. The maximum of Sun spots. Scientific American, 50(19):288–290, May 10, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n19/pdf/scientificamerican05101884-
288b.pdf.

Anonymous:1884:IBD

[5162] Anonymous. Improved breast drill. Scientific American, 50(19):290, May 10, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n19/pdf/scientificamerican05101884-290.pdf.

Anonymous:1884:NDC

[5163] Anonymous. Novel door check. Scientific American, 50(19):290, May 10, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n19/pdf/scientificamerican05101884-290a.pdf.

Anonymous:1884:FSL

[5164] Anonymous. Folding step ladder. Scientific American, 50(19):290, May 10, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n19/pdf/scientificamerican05101884-290b.pdf.

Anonymous:1884:ITH

[5165] Anonymous. An improved tool handle. Scientific American, 50 (19):290–291, May 10, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n19/pdf/scientificamerican05101884-
290c.pdf.

Anonymous:1884:IFG

[5166] Anonymous. Improved furnace grate. Scientific American, 50(19):291, May 10, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n19/pdf/scientificamerican05101884-291.pdf.

Anonymous:1884:NBP

[5167] Anonymous. New buildings of prominence in New York City. Scientific American, 50(19):292–293, May 10, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v50/n19/pdf/scientificamerican05101884-
292.pdf.

Anonymous:1884:DPLc

[5168] Anonymous. Dr. Le Plongeon’s latest and most important discoveries among the ruined cities of Yucatan. Scientific American, 50(19):294–295, May 10, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n19/pdf/scientificamerican05101884-
294.pdf.

Anonymous:1884:DTCb

[5169] Anonymous. Dynamite thrown from cannon by powder. Scientific American, 50(19):295, May 10, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n19/pdf/scientificamerican05101884-
295.pdf.

Anonymous:1884:BGB

[5170] Anonymous. The bursting of gun barrels. Scientific American, 50(19):295–296, May 10, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n19/pdf/scientificamerican05101884-
295a.pdf.

Anonymous:1884:IPL

[5171] Anonymous. Improved permutation lock. Scientific American, 50 (19):296, May 10, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n19/pdf/scientificamerican05101884-
296.pdf.

Anonymous:1884:IFPb

[5172] Anonymous. An improved fence post. Scientific American, 50(19):296, May 10, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n19/pdf/scientificamerican05101884-296a.pdf.

Anonymous:1884:EIk

[5173] Anonymous. Engineering inventions. Scientific American, 50(19):297, May 10, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n19/pdf/scientificamerican05101884-297.pdf.

Anonymous:1884:MIu

[5174] Anonymous. Mechanical inventions. Scientific American, 50(19):297, May 10, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n19/pdf/scientificamerican05101884-297a.pdf.

Anonymous:1884:AIj

[5175] Anonymous. Agricultural inventions. Scientific American, 50(19):297, May 10, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n19/pdf/scientificamerican05101884-297b.pdf.

Anonymous:1884:MIv

[5176] Anonymous. Miscellaneous inventions. Scientific American, 50(19):297, May 10, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n19/pdf/scientificamerican05101884-297c.pdf.

Anonymous:1884:Ck

[5177] Anonymous. Correspondence. Scientific American, 50(2):21–22, January 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n2/pdf/scientificamerican01121884-21.pdf.

Anonymous:1884:BPm

[5178] Anonymous. Business and personal. Scientific American, 50 (2):26–27, January 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n2/pdf/scientificamerican01121884-
26d.pdf.

Anonymous:1884:BNBj

[5179] Anonymous. Books: New books and publications. Scientific American, 50(2):27, January 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n2/pdf/scientificamerican01121884-
27.pdf.

Anonymous:1884:NQl

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n2/pdf/scientificamerican01121884-27a.pdf.

Anonymous:1884:IIl

[5181] Anonymous. Index of inventions. Scientific American, 50(2):27–28, January 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n2/pdf/scientificamerican01121884-27b.pdf.

Anonymous:1884:PNPl

[5182] Anonymous. Patent notes: Patents. Scientific American, 50 (2):27–28, January 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n2/pdf/scientificamerican01121884-
29.pdf.

Anonymous:1884:QWP

[5183] Anonymous. Quantity of water per horse power. Scientific American, 50(2):16, January 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n2/pdf/scientificamerican01121884-
16.pdf.

Anonymous:1884:AB

[5184] Anonymous. Alcohol from bread. Scientific American, 50(2):16, January 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n2/pdf/scientificamerican01121884-16a.pdf.

Anonymous:1884:DSa

[5185] Anonymous. The demand for skill. Scientific American, 50(2):16, January 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n2/pdf/scientificamerican01121884-16b.pdf.

Anonymous:1884:PBCa

[5186] Anonymous. The Pons–Brooks Comet. Scientific American, 50 (2):16–17, January 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n2/pdf/scientificamerican01121884-
16c.pdf.

Anonymous:1884:DBG

[5187] Anonymous. Drilling and boring gun barrels. Scientific American, 50(2):17–18, January 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n2/pdf/scientificamerican01121884-
17.pdf.

Anonymous:1884:GMa

[5188] Anonymous. Grinding mills. Scientific American, 50(2):18, January 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n2/pdf/scientificamerican01121884-18.pdf.

Anonymous:1884:EB

[5189] Anonymous. Elevator bucket. Scientific American, 50(2):18–19, January 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n2/pdf/scientificamerican01121884-18a.pdf.

Anonymous:1884:CCb

[5190] Anonymous. Car coupling. Scientific American, 50(2):19, January 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n2/pdf/scientificamerican01121884-19.pdf.

Anonymous:1884:BH

[5191] Anonymous. Book holder. Scientific American, 50(2):19, January 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n2/pdf/scientificamerican01121884-19a.pdf.

Anonymous:1884:LCC

[5192] Anonymous. Locomotive cow catcher. Scientific American, 50 (2):19–20, January 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n2/pdf/scientificamerican01121884-
19b.pdf.

Anonymous:1884:AINa

[5193] Anonymous. American industries — no. 88. Scientific American, 50(2):20–21, January 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n2/pdf/scientificamerican01121884-
20.pdf.

Anonymous:1884:WTA

[5194] Anonymous. Wire tram across the Teremakau, N.Z. Scientific American, 50(2):22–23, January 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n2/pdf/scientificamerican01121884-
22.pdf.

Anonymous:1884:TD

[5195] Anonymous. The Thibet dog. Scientific American, 50(2):23–25, January 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n2/pdf/scientificamerican01121884-23.pdf.

Anonymous:1884:Ba

[5196] Anonymous. Buckle. Scientific American, 50(2):25, January 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n2/pdf/scientificamerican01121884-25.pdf.

Anonymous:1884:LN

[5197] Anonymous. Locking nut. Scientific American, 50(2):25–26, January 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n2/pdf/scientificamerican01121884-25a.pdf.

Anonymous:1884:EIl

[5198] Anonymous. Engineering inventions. Scientific American, 50(2): 26, January 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n2/pdf/scientificamerican01121884-
26.pdf.

Anonymous:1884:AIk

[5199] Anonymous. Agricultural inventions. Scientific American, 50 (2):26, January 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n2/pdf/scientificamerican01121884-
26a.pdf.

Anonymous:1884:MIw

[5200] Anonymous. Mechanical inventions. Scientific American, 50(2):26, January 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n2/pdf/scientificamerican01121884-26b.pdf.

Anonymous:1884:MIx

[5201] Anonymous. Miscellaneous inventions. Scientific American, 50 (2):26, January 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n2/pdf/scientificamerican01121884-
26c.pdf.

Anonymous:1884:Cl

[5202] Anonymous. Correspondence. Scientific American, 50(20):309, May 17, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n20/pdf/scientificamerican05171884-309.pdf.

Anonymous:1884:BNBk

[5203] Anonymous. Books: New books and publications. Scientific American, 50(20):314, May 17, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n20/pdf/scientificamerican05171884-
314d.pdf.

Anonymous:1884:BPn

[5204] Anonymous. Business and personal. Scientific American, 50(20):314, May 17, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n20/pdf/scientificamerican05171884-314e.pdf.

Anonymous:1884:NQm

[5205] Anonymous. Notes and queries. Scientific American, 50(20):315, May 17, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n20/pdf/scientificamerican05171884-315.pdf.

Anonymous:1884:IIm

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n20/pdf/scientificamerican05171884-315a.pdf.

Anonymous:1884:PNPm

[5207] Anonymous. Patent notes: Patents. Scientific American, 50(20):317, May 17, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n20/pdf/scientificamerican05171884-317.pdf.

Anonymous:1884:ECR

[5208] Anonymous. Elevated city railways cause eye troubles. Scientific American, 50(20):304, May 17, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n20/pdf/scientificamerican05171884-
304.pdf.

Anonymous:1884:LS

[5209] Anonymous. The little shop. Scientific American, 50(20):304–305, May 17, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n20/pdf/scientificamerican05171884-304a.pdf.

Anonymous:1884:CRW

[5210] Anonymous. Collecting and removing waste from spinning machines. Scientific American, 50(20):306, May 17, 1884. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v50/
n20/pdf/scientificamerican05171884-306.pdf.

Anonymous:1884:IWa

[5211] Anonymous. Improved wrench. Scientific American, 50(20):306, May 17, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n20/pdf/scientificamerican05171884-306a.pdf.

Anonymous:1884:CTa

[5212] Anonymous. Car truck. Scientific American, 50(20):306, May 17, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n20/pdf/scientificamerican05171884-306b.pdf.

Anonymous:1884:NLa

[5213] Anonymous. A new lock. Scientific American, 50(20):306–307, May 17, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n20/pdf/scientificamerican05171884-306c.pdf.

Anonymous:1884:BPo

[5214] Anonymous. Bench plane. Scientific American, 50(20):307, May 17, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n20/pdf/scientificamerican05171884-307.pdf.

Anonymous:1884:PRAa

[5215] Anonymous. A photographic revolver for amateurs. Scientific American, 50(20):307–308, May 17, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n20/pdf/scientificamerican05171884-
307a.pdf.

Anonymous:1884:AINb

[5216] Anonymous. American industries — no. 89. Scientific American, 50(20):308, May 17, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n20/pdf/scientificamerican05171884-
308.pdf.

Anonymous:1884:Ta

[5217] Anonymous. Tombstone. Scientific American, 50(20):309–310, May 17, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n20/pdf/scientificamerican05171884-309a.pdf.

Anonymous:1884:ESM

[5218] Anonymous. Ericsson’s Sun motor. Scientific American, 50(20):310, May 17, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n20/pdf/scientificamerican05171884-310.pdf.

Anonymous:1884:BHP

[5219] Anonymous. A beautiful house plant. Scientific American, 50(20):310, May 17, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n20/pdf/scientificamerican05171884-310a.pdf.

Anonymous:1884:DPLd

[5220] Anonymous. Dr. Le Plongeon’s latest and most important discoveries among the ruined cities of Yucatan. Scientific American, 50(20):311–312, May 17, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n20/pdf/scientificamerican05171884-
311.pdf.

Anonymous:1884:FAM

[5221] Anonymous. Fountain attachment for marking brushes. Scientific American, 50(20):312–313, May 17, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v50/n20/pdf/scientificamerican05171884-
312.pdf.

Anonymous:1884:EIm

[5222] Anonymous. Engineering inventions. Scientific American, 50(20):314, May 17, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n20/pdf/scientificamerican05171884-314.pdf.

Anonymous:1884:MIy

[5223] Anonymous. Mechanical inventions. Scientific American, 50(20):314, May 17, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n20/pdf/scientificamerican05171884-314a.pdf.

Anonymous:1884:AIl

[5224] Anonymous. Agricultural inventions. Scientific American, 50(20):314, May 17, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n20/pdf/scientificamerican05171884-314b.pdf.

Anonymous:1884:MIz

[5225] Anonymous. Miscellaneous inventions. Scientific American, 50(20):314, May 17, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n20/pdf/scientificamerican05171884-314c.pdf.

Anonymous:1884:Cm

[5226] Anonymous. Correspondence. Scientific American, 50(21):325, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n21/pdf/scientificamerican05241884-325.pdf.

Anonymous:1884:BPp

[5227] Anonymous. Business and personal. Scientific American, 50(21):330, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n21/pdf/scientificamerican05241884-330d.pdf.

Anonymous:1884:NQn

[5228] Anonymous. Notes and queries. Scientific American, 50(21):330–331, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n21/pdf/scientificamerican05241884-330e.pdf.

Anonymous:1884:IIn

[5229] Anonymous. Index of inventions. Scientific American, 50(21):331–332, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n21/pdf/scientificamerican05241884-331.pdf.

Anonymous:1884:NCR

[5230] Anonymous. A new cable railway system. Scientific American, 50(21):319, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n21/pdf/scientificamerican05241884-
319.pdf.

Anonymous:1884:STP

[5231] Anonymous. Steam at a thousand pounds pressure. Scientific American, 50(21):320, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n21/pdf/scientificamerican05241884-
320.pdf.

Anonymous:1884:ELB

[5232] Anonymous. Elasticity of leather belts. Scientific American, 50(21):320, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n21/pdf/scientificamerican05241884-320a.pdf.

Anonymous:1884:N

[5233] Anonymous. Nails. Scientific American, 50(21):320–321, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n21/pdf/scientificamerican05241884-320b.pdf.

Anonymous:1884:DRPc

[5234] Anonymous. Decisions relating to patents. Scientific American, 50(21):321–322, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n21/pdf/scientificamerican05241884-
321.pdf.

Anonymous:1884:G

[5235] Anonymous. Gravimotor. Scientific American, 50(21):322, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n21/pdf/scientificamerican05241884-322.pdf.

Anonymous:1884:FS

[5236] Anonymous. Folding staircase. Scientific American, 50(21):322, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n21/pdf/scientificamerican05241884-322a.pdf.

Anonymous:1884:ILJ

[5237] Anonymous. Improved lifting jack. Scientific American, 50(21): 322–323, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n21/pdf/scientificamerican05241884-
322b.pdf.

Anonymous:1884:ZHM

[5238] Anonymous. Zschiesche’s hydraulic motor. Scientific American, 50(21):323, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n21/pdf/scientificamerican05241884-
323.pdf.

Anonymous:1884:DPO

[5239] Anonymous. Design for a parlor organ. Scientific American, 50 (21):323–324, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n21/pdf/scientificamerican05241884-
323a.pdf.

Anonymous:1884:BSL

[5240] Anonymous. Bartholdi’s Statue of Liberty. Scientific American, 50(21):323, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n21/pdf/scientificamerican05241884-
323b.pdf.

Anonymous:1884:DEA

[5241] Anonymous. The doctrine of evolution, an antique Roman mosaic from Carthage. Scientific American, 50(21):324, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n21/pdf/scientificamerican05241884-324.pdf.

Anonymous:1884:TBP

[5242] Anonymous. Thompson’s bleaching process. Scientific American, 50(21):326–327, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n21/pdf/scientificamerican05241884-
326.pdf.

Anonymous:1884:DPLe

[5243] Anonymous. Dr. Le Plongeon’s latest and most important discoveries among the ruined cities of Yucatan. Scientific American, 50(21):327–328, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n21/pdf/scientificamerican05241884-
327.pdf.

Anonymous:1884:BFC

[5244] Anonymous. Boiler flue cleaner. Scientific American, 50(21):328, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n21/pdf/scientificamerican05241884-328.pdf.

Anonymous:1884:IFB

[5245] Anonymous. Improved form of bell. Scientific American, 50(21):328, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n21/pdf/scientificamerican05241884-328a.pdf.

Anonymous:1884:NSCa

[5246] Anonymous. A new ship cleaner. Scientific American, 50(21):328–329, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n21/pdf/scientificamerican05241884-328b.pdf.

Anonymous:1884:EIn

[5247] Anonymous. Engineering inventions. Scientific American, 50(21):330, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n21/pdf/scientificamerican05241884-330.pdf.

Anonymous:1884:MIba

[5248] Anonymous. Mechanical inventions. Scientific American, 50(21):330, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n21/pdf/scientificamerican05241884-330a.pdf.

Anonymous:1884:AIm

[5249] Anonymous. Agricultural inventions. Scientific American, 50(21):330, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n21/pdf/scientificamerican05241884-330b.pdf.

Anonymous:1884:MIbb

[5250] Anonymous. Miscellaneous inventions. Scientific American, 50(21):330, May 24, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n21/pdf/scientificamerican05241884-330c.pdf.

Anonymous:1884:BNBl

[5251] Anonymous. Books: New books and publications. Scientific American, 50(22):345, May 31, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n22/pdf/scientificamerican05311884-
345d.pdf.

Anonymous:1884:BPq

[5252] Anonymous. Business and personal. Scientific American, 50(22):345, May 31, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n22/pdf/scientificamerican05311884-345e.pdf.

Anonymous:1884:NQo

[5253] Anonymous. Notes and queries. Scientific American, 50(22):346–347, May 31, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n22/pdf/scientificamerican05311884-346.pdf.

Anonymous:1884:IIo

[5254] Anonymous. Index of inventions. Scientific American, 50(22):347–348, May 31, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n22/pdf/scientificamerican05311884-347.pdf.

Anonymous:1884:PNPn

[5255] Anonymous. Patent notes: Patents. Scientific American, 50(22):349, May 31, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n22/pdf/scientificamerican05311884-349.pdf.

Anonymous:1884:AC

[5256] Anonymous. American chains. Scientific American, 50(22):336, May 31, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n22/pdf/scientificamerican05311884-336.pdf.

Anonymous:1884:APJa

[5257] Anonymous. Aspects of the planets for June. Scientific American, 50(22):337, May 31, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n22/pdf/scientificamerican05311884-
337.pdf.

Anonymous:1884:CDa

[5258] Anonymous. Chest of drawers. Scientific American, 50(22):338, May 31, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n22/pdf/scientificamerican05311884-338.pdf.

Anonymous:1884:ID

[5259] Anonymous. Improved desk. Scientific American, 50(22):338, May 31, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n22/pdf/scientificamerican05311884-338a.pdf.

Anonymous:1884:CSb

[5260] Anonymous. Car starter. Scientific American, 50(22):338, May 31, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n22/pdf/scientificamerican05311884-338b.pdf.

Anonymous:1884:SPb

[5261] Anonymous. Stump puller. Scientific American, 50(22):338, May 31, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n22/pdf/scientificamerican05311884-338c.pdf.

Anonymous:1884:HCa

[5262] Anonymous. Horse cleaner. Scientific American, 50(22):338, May 31, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n22/pdf/scientificamerican05311884-338d.pdf.

Anonymous:1884:AINc

[5263] Anonymous. American industries. — no. 90. Scientific American, 50(22):339–341, May 31, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n22/pdf/scientificamerican05311884-
339.pdf.

Anonymous:1884:DFD

[5264] Anonymous. Dolman’s fire dampers. Scientific American, 50(22):341, May 31, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n22/pdf/scientificamerican05311884-341.pdf.

Anonymous:1884:ISPa

[5265] Anonymous. Improved sewer pipe trap. Scientific American, 50(22):341, May 31, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n22/pdf/scientificamerican05311884-341a.pdf.

Anonymous:1884:FCD

[5266] Anonymous. Freight car door locks. Scientific American, 50(22):342, May 31, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n22/pdf/scientificamerican05311884-342.pdf.

Anonymous:1884:PIB

[5267] Anonymous. Petroleum industry of Baku. Scientific American, 50(22):342–343, May 31, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n22/pdf/scientificamerican05311884-
342a.pdf.

Anonymous:1884:DPH

[5268] Anonymous. Dr. Le Plongeon and his party. Scientific American, 50(22):343–344, May 31, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n22/pdf/scientificamerican05311884-
343.pdf.

Anonymous:1884:FGP

[5269] Anonymous. Feed guide for printing presses. Scientific American, 50(22):344, May 31, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n22/pdf/scientificamerican05311884-
344.pdf.

Anonymous:1884:WBa

[5270] Anonymous. Wagon body. Scientific American, 50(22):344, May 31, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n22/pdf/scientificamerican05311884-344a.pdf.

Anonymous:1884:WCa

[5271] Anonymous. Water closet. Scientific American, 50(22):344, May 31, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n22/pdf/scientificamerican05311884-344b.pdf.

Anonymous:1884:EIo

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n22/pdf/scientificamerican05311884-345.pdf.

Anonymous:1884:MIbc

[5273] Anonymous. Mechanical inventions. Scientific American, 50(22):345, May 31, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n22/pdf/scientificamerican05311884-345a.pdf.

Anonymous:1884:AIo

[5274] Anonymous. Agricultural inventions. Scientific American, 50(22):345, May 31, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n22/pdf/scientificamerican05311884-345b.pdf.

Anonymous:1884:MIbd

[5275] Anonymous. Miscellaneous inventions. Scientific American, 50(22):345, May 31, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n22/pdf/scientificamerican05311884-345c.pdf.

Anonymous:1884:Cn

[5276] Anonymous. Correspondence. Scientific American, 50(23):357, June 7, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n23/pdf/scientificamerican06071884-357.pdf.

Anonymous:1884:BPr

[5277] Anonymous. Business and personal. Scientific American, 50(23):362, June 7, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n23/pdf/scientificamerican06071884-362c.pdf.

Anonymous:1884:BNBm

[5278] Anonymous. Books: New books and publications. Scientific American, 50(23):362, June 7, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n23/pdf/scientificamerican06071884-
362d.pdf.

Anonymous:1884:NQp

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n23/pdf/scientificamerican06071884-362e.pdf.

Anonymous:1884:IIp

[5280] Anonymous. Index of inventions. Scientific American, 50(23):363–364, June 7, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n23/pdf/scientificamerican06071884-363.pdf.

Anonymous:1884:PNPo

[5281] Anonymous. Patent notes: Patents. Scientific American, 50(23):365, June 7, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n23/pdf/scientificamerican06071884-365.pdf.

Anonymous:1884:WIT

[5282] Anonymous. Wrought iron tube for cable railways. Scientific American, 50(23):351–352, June 7, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n23/pdf/scientificamerican06071884-
351.pdf.

Anonymous:1884:SMa

[5283] Anonymous. The state of manufacturing. Scientific American, 50(23):352, June 7, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n23/pdf/scientificamerican06071884-
352.pdf.

Anonymous:1884:PHN

[5284] Anonymous. Porcelain headed nails. Scientific American, 50(23):352, June 7, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n23/pdf/scientificamerican06071884-352a.pdf.

Anonymous:1884:SHC

[5285] Anonymous. Softening and hardening cast iron. Scientific American, 50(23):352–354, June 7, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n23/pdf/scientificamerican06071884-
352b.pdf.

Anonymous:1884:FEC

[5286] Anonymous. Folding egg case. Scientific American, 50(23):354, June 7, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n23/pdf/scientificamerican06071884-354.pdf.

Anonymous:1884:SPC

[5287] Anonymous. Stove pipe connection. Scientific American, 50(23):354, June 7, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n23/pdf/scientificamerican06071884-354a.pdf.

Anonymous:1884:EDK

[5288] Anonymous. Electric door keepers. Scientific American, 50(23):354, June 7, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n23/pdf/scientificamerican06071884-354b.pdf.

Anonymous:1884:CBa

[5289] Anonymous. Canal boat. Scientific American, 50(23):354, June 7, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n23/pdf/scientificamerican06071884-354c.pdf.

Anonymous:1884:DHG

[5290] Anonymous. Device for holding and guiding the fingers in writing. Scientific American, 50(23):355, June 7, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v50/n23/pdf/scientificamerican06071884-
355.pdf.

Anonymous:1884:PFL

[5291] Anonymous. Photographing a flash of lightning. Scientific American, 50(23):355, June 7, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n23/pdf/scientificamerican06071884-
355a.pdf.

Anonymous:1884:CVG

[5292] Anonymous. Cut-off valve gear. Scientific American, 50(23):355–356, June 7, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n23/pdf/scientificamerican06071884-355b.pdf.

Anonymous:1884:ISHb

[5293] Anonymous. Improved steam hammer. Scientific American, 50(23):358, June 7, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n23/pdf/scientificamerican06071884-358.pdf.

Anonymous:1884:IAF

[5294] Anonymous. The isatis, or Arctic fox. Scientific American, 50(23):359, June 7, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n23/pdf/scientificamerican06071884-359.pdf.

Anonymous:1884:WA

[5295] Anonymous. A whalebone arch. Scientific American, 50(23):359–360, June 7, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n23/pdf/scientificamerican06071884-359a.pdf.

Anonymous:1884:CHa

[5296] Anonymous. Cuff holder. Scientific American, 50(23):360, June 7, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n23/pdf/scientificamerican06071884-360.pdf.

Anonymous:1884:WMH

[5297] Anonymous. Watch maker’s hand vise and ring bender. Scientific American, 50(23):360–361, June 7, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n23/pdf/scientificamerican06071884-
360a.pdf.

Anonymous:1884:EIp

[5298] Anonymous. Engineering inventions. Scientific American, 50(23):362, June 7, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n23/pdf/scientificamerican06071884-362.pdf.

Anonymous:1884:MIbe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n23/pdf/scientificamerican06071884-362a.pdf.

Anonymous:1884:MIbf

[5300] Anonymous. Miscellaneous inventions. Scientific American, 50(23):362, June 7, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n23/pdf/scientificamerican06071884-362b.pdf.

Anonymous:1884:Co

[5301] Anonymous. Correspondence. Scientific American, 50(24):373–374, June 14, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n24/pdf/scientificamerican06141884-373.pdf.

Anonymous:1884:BPs

[5302] Anonymous. Business and personal. Scientific American, 50 (24):378–379, June 14, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n24/pdf/scientificamerican06141884-
378c.pdf.

Anonymous:1884:NQq

[5303] Anonymous. Notes and queries. Scientific American, 50(24):379, June 14, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n24/pdf/scientificamerican06141884-379.pdf.

Anonymous:1884:IIq

[5304] Anonymous. Index of inventions. Scientific American, 50(24):379–380, June 14, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n24/pdf/scientificamerican06141884-379a.pdf.

Anonymous:1884:BSQ

[5305] Anonymous. The brown stone quarries of Connecticut. Scientific American, 50(24):368, June 14, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n24/pdf/scientificamerican06141884-
368.pdf.

Anonymous:1884:GMW

[5306] Anonymous. Gauges for mechanical work. Scientific American, 50(24):368, June 14, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n24/pdf/scientificamerican06141884-
368a.pdf.

Anonymous:1884:ONS

[5307] Anonymous. Our new supplement catalogue. Scientific American, 50(24):368, June 14, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n24/pdf/scientificamerican06141884-
368b.pdf.

Anonymous:1884:SLI

[5308] Anonymous. A suggested lathe improvement. Scientific American, 50(24):368, June 14, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n24/pdf/scientificamerican06141884-
368c.pdf.

Anonymous:1884:RSE

[5309] Anonymous. A remarkable straight edge. Scientific American, 50 (24):368–370, June 14, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n24/pdf/scientificamerican06141884-
368d.pdf.

Anonymous:1884:MWB

[5310] Anonymous. Machine for working butter. Scientific American, 50(24):370, June 14, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n24/pdf/scientificamerican06141884-
370.pdf.

Anonymous:1884:ISA

[5311] Anonymous. Improved saw arbor. Scientific American, 50(24):370, June 14, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n24/pdf/scientificamerican06141884-370a.pdf.

Anonymous:1884:CTCa

[5312] Anonymous. Car track cleaner. Scientific American, 50(24):370, June 14, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n24/pdf/scientificamerican06141884-370b.pdf.

Anonymous:1884:ADA

[5313] Anonymous. The Allen dense air ice machine. Scientific American, 50(24):371, June 14, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n24/pdf/scientificamerican06141884-
371.pdf.

Anonymous:1884:SWa

[5314] Anonymous. Sail wagon. Scientific American, 50(24):371, June 14, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n24/pdf/scientificamerican06141884-371a.pdf.

Anonymous:1884:AINd

[5315] Anonymous. American industries — no. 91. Scientific American, 50(24):372–373, June 14, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n24/pdf/scientificamerican06141884-
372.pdf.

Anonymous:1884:PET

[5316] Anonymous. Portable electric testing apparatus. Scientific American, 50(24):374–375, June 14, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n24/pdf/scientificamerican06141884-
374.pdf.

Anonymous:1884:NHN

[5317] Anonymous. Natural history notes. Scientific American, 50(24):375, June 14, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n24/pdf/scientificamerican06141884-375.pdf.

Anonymous:1884:TIP

[5318] Anonymous. The tiger by instantaneous photography. Scientific American, 50(24):375–376, June 14, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n24/pdf/scientificamerican06141884-
375a.pdf.

Anonymous:1884:OUT

[5319] Anonymous. An old universal tool. Scientific American, 50 (24):376–377, June 14, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n24/pdf/scientificamerican06141884-
376.pdf.

Anonymous:1884:ICB

[5320] Anonymous. Improved claw bar. Scientific American, 50(24):377, June 14, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n24/pdf/scientificamerican06141884-377.pdf.

Anonymous:1884:WCb

[5321] Anonymous. Water cooler. Scientific American, 50(24):377, June 14, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n24/pdf/scientificamerican06141884-377a.pdf.

Anonymous:1884:EIq

[5322] Anonymous. Engineering inventions. Scientific American, 50(24):378, June 14, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n24/pdf/scientificamerican06141884-378.pdf.

Anonymous:1884:MIbg

[5323] Anonymous. Mechanical inventions. Scientific American, 50(24):378, June 14, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n24/pdf/scientificamerican06141884-378a.pdf.

Anonymous:1884:MIbh

[5324] Anonymous. Miscellaneous inventions. Scientific American, 50(24):378, June 14, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n24/pdf/scientificamerican06141884-378b.pdf.

Anonymous:1884:Cq

[5325] Anonymous. Correspondence. Scientific American, 50(25):389, June 21, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n25/pdf/scientificamerican06211884-389.pdf.

Anonymous:1884:BNBn

[5326] Anonymous. Books: New books and publications. Scientific American, 50(25):394, June 21, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n25/pdf/scientificamerican06211884-
394d.pdf.

Anonymous:1884:BPt

[5327] Anonymous. Business and personal. Scientific American, 50(25):394, June 21, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n25/pdf/scientificamerican06211884-394f.pdf.

Anonymous:1884:NQr

[5328] Anonymous. Notes and queries. Scientific American, 50(25):394–395, June 21, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n25/pdf/scientificamerican06211884-394g.pdf.

Anonymous:1884:IIr

[5329] Anonymous. Index of inventions. Scientific American, 50(25):395–396, June 21, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n25/pdf/scientificamerican06211884-395.pdf.

Anonymous:1884:PNPp

[5330] Anonymous. Patent notes: Patents. Scientific American, 50(25):397, June 21, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n25/pdf/scientificamerican06211884-397.pdf.

Anonymous:1884:CSC

[5331] Anonymous. Curiosities of screw cutting. Scientific American, 50(25):384, June 21, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n25/pdf/scientificamerican06211884-
384.pdf.

Anonymous:1884:CCc

[5332] Anonymous. Chewing the cud. Scientific American, 50(25):384–385, June 21, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n25/pdf/scientificamerican06211884-384a.pdf.

Anonymous:1884:AINe

[5333] Anonymous. American industries — no. 92. Scientific American, 50(25):386, June 21, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n25/pdf/scientificamerican06211884-
386.pdf.

Anonymous:1884:FEb

[5334] Anonymous. Fire escape. Scientific American, 50(25):387, June 21, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n25/pdf/scientificamerican06211884-387.pdf.

Anonymous:1884:LM

[5335] Anonymous. Lubricator and measure. Scientific American, 50(25):387, June 21, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n25/pdf/scientificamerican06211884-387a.pdf.

Anonymous:1884:CAB

[5336] Anonymous. Car axle box. Scientific American, 50(25):387, June 21, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n25/pdf/scientificamerican06211884-387b.pdf.

Anonymous:1884:MEG

[5337] Anonymous. Marine engine governor. Scientific American, 50(25):387, June 21, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n25/pdf/scientificamerican06211884-387c.pdf.

Anonymous:1884:WWC

[5338] Anonymous. Wire wrapping or covering machine. Scientific American, 50(25):388, June 21, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n25/pdf/scientificamerican06211884-
388.pdf.

Anonymous:1884:LCA

[5339] Anonymous. Liquid carbonic acid — its industrial production and uses. Scientific American, 50(25):390, June 21, 1884. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v50/
n25/pdf/scientificamerican06211884-390.pdf.

Anonymous:1884:DSF

[5340] Anonymous. Deep sea fishes. Scientific American, 50(25):391–392, June 21, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n25/pdf/scientificamerican06211884-391.pdf.

Anonymous:1884:CHb

[5341] Anonymous. Clothes horse. Scientific American, 50(25):392, June 21, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n25/pdf/scientificamerican06211884-392.pdf.

Anonymous:1884:IFA

[5342] Anonymous. Improved fan attachment. Scientific American, 50 (25):392–393, June 21, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n25/pdf/scientificamerican06211884-
392a.pdf.

Anonymous:1884:SAb

[5343] Anonymous. The steamer America. Scientific American, 50(25):393, June 21, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n25/pdf/scientificamerican06211884-393.pdf.

Anonymous:1884:EIr

[5344] Anonymous. Engineering inventions. Scientific American, 50(25):394, June 21, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n25/pdf/scientificamerican06211884-394.pdf.

Anonymous:1884:MIbi

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n25/pdf/scientificamerican06211884-394a.pdf.

Anonymous:1884:AIp

[5346] Anonymous. Agricultural inventions. Scientific American, 50(25):394, June 21, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n25/pdf/scientificamerican06211884-394b.pdf.

Anonymous:1884:MIbj

[5347] Anonymous. Miscellaneous inventions. Scientific American, 50(25):394, June 21, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n25/pdf/scientificamerican06211884-394c.pdf.

Anonymous:1884:Sf

[5348] Anonymous. Special. Scientific American, 50(25):394, June 21, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n25/pdf/scientificamerican06211884-394e.pdf.

Anonymous:1884:Cr

[5349] Anonymous. Correspondence. Scientific American, 50(26):405, June 28, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n26/pdf/scientificamerican06281884-405.pdf.

Anonymous:1884:BPu

[5350] Anonymous. Business and personal. Scientific American, 50(26):408, June 28, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n26/pdf/scientificamerican06281884-408d.pdf.

Anonymous:1884:BNBo

[5351] Anonymous. Books: New books and publications. Scientific American, 50(26):408, June 28, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n26/pdf/scientificamerican06281884-
408e.pdf.

Anonymous:1884:NQs

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n26/pdf/scientificamerican06281884-408f.pdf.

Anonymous:1884:IIs

[5353] Anonymous. Index of inventions. Scientific American, 50(26):409–412, June 28, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n26/pdf/scientificamerican06281884-409.pdf.

Anonymous:1884:CLD

[5354] Anonymous. Conditions of life in the depths of the sea. Scientific American, 50(26):400, June 28, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n26/pdf/scientificamerican06281884-
400.pdf.

Anonymous:1884:BF

[5355] Anonymous. From Broadway to the Ferries. Scientific American, 50(26):400, June 28, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n26/pdf/scientificamerican06281884-
400a.pdf.

Anonymous:1884:APJb

[5356] Anonymous. Aspects of the planets for July. Scientific American, 50(26):401, June 28, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n26/pdf/scientificamerican06281884-
401.pdf.

Anonymous:1884:HPBc

[5357] Anonymous. How pile buttons are made. Scientific American, 50(26):401, June 28, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n26/pdf/scientificamerican06281884-
401a.pdf.

Anonymous:1884:NSG

[5358] Anonymous. New spring gearing for vehicles. Scientific American, 50(26):402, June 28, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n26/pdf/scientificamerican06281884-
402.pdf.

Anonymous:1884:WRG

[5359] Anonymous. Wagon running gear. Scientific American, 50(26):402, June 28, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n26/pdf/scientificamerican06281884-402a.pdf.

Anonymous:1884:HAD

[5360] Anonymous. Hub attaching device. Scientific American, 50(26):402, June 28, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n26/pdf/scientificamerican06281884-402b.pdf.

Anonymous:1884:SSa

[5361] Anonymous. Shaft support. Scientific American, 50(26):403, June 28, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n26/pdf/scientificamerican06281884-403.pdf.

Anonymous:1884:FPb

[5362] Anonymous. Foot power. Scientific American, 50(26):403, June 28, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n26/pdf/scientificamerican06281884-403a.pdf.

Anonymous:1884:IJB

[5363] Anonymous. Improved journal bearing. Scientific American, 50 (26):403, June 28, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n26/pdf/scientificamerican06281884-
403b.pdf.

Anonymous:1884:CPH

[5364] Anonymous. Combined plow and harrow. Scientific American, 50 (26):403–404, June 28, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n26/pdf/scientificamerican06281884-
403c.pdf.

Anonymous:1884:TB

[5365] Anonymous. Trace buckle. Scientific American, 50(26):404, June 28, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n26/pdf/scientificamerican06281884-404.pdf.

Anonymous:1884:UPF

[5366] Anonymous. Umbrella and parasol frame. Scientific American, 50(26):404, June 28, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n26/pdf/scientificamerican06281884-
404a.pdf.

Anonymous:1884:NTa

[5367] Anonymous. A novel toy. Scientific American, 50(26):405, June 28, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n26/pdf/scientificamerican06281884-405a.pdf.

Anonymous:1884:PCb

[5368] Anonymous. The Panama Canal. Scientific American, 50(26):406, June 28, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n26/pdf/scientificamerican06281884-406.pdf.

Anonymous:1884:GPMa

[5369] Anonymous. The gorilla at the Paris Museum of Natural History. Scientific American, 50(26):407, June 28, 1884. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v50/
n26/pdf/scientificamerican06281884-407.pdf.

Anonymous:1884:SES

[5370] Anonymous. Street elevator at Stockholm. Scientific American, 50(26):407, June 28, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n26/pdf/scientificamerican06281884-
407a.pdf.

Anonymous:1884:EIs

[5371] Anonymous. Engineering inventions. Scientific American, 50(26):408, June 28, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n26/pdf/scientificamerican06281884-408.pdf.

Anonymous:1884:MIbk

[5372] Anonymous. Mechanical inventions. Scientific American, 50(26):408, June 28, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n26/pdf/scientificamerican06281884-408a.pdf.

Anonymous:1884:AIq

[5373] Anonymous. Agricultural inventions. Scientific American, 50(26):408, June 28, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n26/pdf/scientificamerican06281884-408b.pdf.

Anonymous:1884:MIbl

[5374] Anonymous. Miscellaneous inventions. Scientific American, 50(26):408, June 28, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n26/pdf/scientificamerican06281884-408c.pdf.

Anonymous:1884:Cu

[5375] Anonymous. Correspondence. Scientific American, 50(3):41, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n3/pdf/scientificamerican01191884-41.pdf.

Anonymous:1884:BPv

[5376] Anonymous. Business and personal. Scientific American, 50 (3):42–43, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n3/pdf/scientificamerican01191884-
42e.pdf.

Anonymous:1884:NQt

[5377] Anonymous. Notes and queries. Scientific American, 50(3):43, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n3/pdf/scientificamerican01191884-43.pdf.

Anonymous:1884:IIt

[5378] Anonymous. Index of inventions. Scientific American, 50(3):43–44, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n3/pdf/scientificamerican01191884-43a.pdf.

Anonymous:1884:PNPq

[5379] Anonymous. Patent notes: Patents. Scientific American, 50(3):45, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n3/pdf/scientificamerican01191884-45.pdf.

Anonymous:1884:AET

[5380] Anonymous. Application of electricity to tunneling. Scientific American, 50(3):31–32, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n3/pdf/scientificamerican01191884-
31.pdf.

Anonymous:1884:NPLa

[5381] Anonymous. The nullification of the Patent Laws. Scientific American, 50(3):32, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n3/pdf/scientificamerican01191884-
32.pdf.

Anonymous:1884:MCa

[5382] Anonymous. Magnetism in cities. Scientific American, 50(3):33, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n3/pdf/scientificamerican01191884-33.pdf.

Anonymous:1884:BII

[5383] Anonymous. A blow at the industries of indiana. Scientific American, 50(3):33, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n3/pdf/scientificamerican01191884-
33a.pdf.

Anonymous:1884:WCW

[5384] Anonymous. Working cold wrought iron. Scientific American, 50(3):33, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n3/pdf/scientificamerican01191884-
33b.pdf.

Anonymous:1884:RASb

[5385] Anonymous. Restoration of American shipbuilding. Scientific American, 50(3):33, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n3/pdf/scientificamerican01191884-
33c.pdf.

Anonymous:1884:PAB

[5386] Anonymous. Process of and apparatus for blocking ice. Scientific American, 50(3):34, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n3/pdf/scientificamerican01191884-
34.pdf.

Anonymous:1884:Bb

[5387] Anonymous. Blotter. Scientific American, 50(3):34, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n3/pdf/scientificamerican01191884-34a.pdf.

Anonymous:1884:IDM

[5388] Anonymous. Improved double mould board plow. Scientific American, 50(3):35, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n3/pdf/scientificamerican01191884-
35.pdf.

Anonymous:1884:HR

[5389] Anonymous. Hose reel. Scientific American, 50(3):35, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n3/pdf/scientificamerican01191884-35a.pdf.

Anonymous:1884:Pd

[5390] Anonymous. Plow. Scientific American, 50(3):35, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n3/pdf/scientificamerican01191884-35b.pdf.

Anonymous:1884:EL

[5391] Anonymous. Engine lubricator. Scientific American, 50(3):35–37, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n3/pdf/scientificamerican01191884-35c.pdf.

Anonymous:1884:SLP

[5392] Anonymous. The strength of lead pipes. Scientific American, 50(3):38, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n3/pdf/scientificamerican01191884-
38.pdf.

Anonymous:1884:SET

[5393] Anonymous. Soemmering’s electric telegraph. Scientific American, 50(3):38–39, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n3/pdf/scientificamerican01191884-
38a.pdf.

Anonymous:1884:CKa

[5394] Anonymous. The chimpanzee and Koolokamba. Scientific American, 50(3):39–40, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n3/pdf/scientificamerican01191884-
39.pdf.

Anonymous:1884:NLb

[5395] Anonymous. Nut lock. Scientific American, 50(3):40, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n3/pdf/scientificamerican01191884-40.pdf.

Anonymous:1884:EIt

[5396] Anonymous. Engineering inventions. Scientific American, 50(3): 42, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n3/pdf/scientificamerican01191884-
42.pdf.

Anonymous:1884:MIbm

[5397] Anonymous. Mechanical inventions. Scientific American, 50(3):42, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n3/pdf/scientificamerican01191884-42a.pdf.

Anonymous:1884:AIr

[5398] Anonymous. Agricultural inventions. Scientific American, 50 (3):42, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n3/pdf/scientificamerican01191884-
42b.pdf.

Anonymous:1884:MIbn

[5399] Anonymous. Miscellaneous inventions. Scientific American, 50 (3):42, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n3/pdf/scientificamerican01191884-
42c.pdf.

Anonymous:1884:I

[5400] Anonymous. Insurance. Scientific American, 50(3):42, January 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n3/pdf/scientificamerican01191884-42d.pdf.

Anonymous:1884:Cv

[5401] Anonymous. Correspondence. Scientific American, 50(4):53–54, January 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n4/pdf/scientificamerican01261884-53.pdf.

Anonymous:1884:BNBp

[5402] Anonymous. Books: New books and publications. Scientific American, 50(4):58, January 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n4/pdf/scientificamerican01261884-
58d.pdf.

Anonymous:1884:BPw

[5403] Anonymous. Business and personal. Scientific American, 50 (4):58–59, January 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n4/pdf/scientificamerican01261884-
58e.pdf.

Anonymous:1884:NQu

[5404] Anonymous. Notes and queries. Scientific American, 50(4):59, January 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n4/pdf/scientificamerican01261884-59.pdf.

Anonymous:1884:IIu

[5405] Anonymous. Index of inventions. Scientific American, 50(4):59–60, January 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n4/pdf/scientificamerican01261884-59a.pdf.

Anonymous:1884:PNPr

[5406] Anonymous. Patent notes: Patents. Scientific American, 50(4):61, January 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n4/pdf/scientificamerican01261884-61.pdf.

Anonymous:1884:RBB

[5407] Anonymous. Railroad building on the bank of the Hudson. Scientific American, 50(4):47–48, January 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v50/n4/pdf/scientificamerican01261884-
47.pdf.

Anonymous:1884:BNP

[5408] Anonymous. Bills for the nullification of patents. Scientific American, 50(4):48, January 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n4/pdf/scientificamerican01261884-
48.pdf.

Anonymous:1884:ACH

[5409] Anonymous. Annealing and case hardening. Scientific American, 50(4):48, January 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n4/pdf/scientificamerican01261884-
48a.pdf.

Anonymous:1884:PAC

[5410] Anonymous. Proposed amendment of the copyright law. Scientific American, 50(4):48, January 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n4/pdf/scientificamerican01261884-
48b.pdf.

Anonymous:1884:OWP

[5411] Anonymous. Is our water power failing? Scientific American, 50(4):48–49, January 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n4/pdf/scientificamerican01261884-
48c.pdf.

Anonymous:1884:BFA

[5412] Anonymous. Brass finishing by acids. Scientific American, 50 (4):49–50, January 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n4/pdf/scientificamerican01261884-
49.pdf.

Anonymous:1884:MPS

[5413] Anonymous. Metallic plastering surface. Scientific American, 50 (4):50, January 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n4/pdf/scientificamerican01261884-
50.pdf.

Anonymous:1884:SBC

[5414] Anonymous. Show box cover. Scientific American, 50(4):50, January 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n4/pdf/scientificamerican01261884-50a.pdf.

Anonymous:1884:MCb

[5415] Anonymous. Milk cooler. Scientific American, 50(4):50–51, January 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n4/pdf/scientificamerican01261884-50b.pdf.

Anonymous:1884:SWS

[5416] Anonymous. Sliding window shutter. Scientific American, 50 (4):51, January 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n4/pdf/scientificamerican01261884-
51.pdf.

Anonymous:1884:LHH

[5417] Anonymous. Laporte’s hydraulic hay press. Scientific American, 50(4):51–53, January 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n4/pdf/scientificamerican01261884-
51a.pdf.

Anonymous:1884:BTS

[5418] Anonymous. Bradford technical school. Scientific American, 50 (4):54–55, January 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n4/pdf/scientificamerican01261884-
54.pdf.

Anonymous:1884:DP

[5419] Anonymous. The diamond pheasant. Scientific American, 50 (4):55–56, January 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n4/pdf/scientificamerican01261884-
55.pdf.

Anonymous:1884:CSd

[5420] Anonymous. Car seat. Scientific American, 50(4):56–58, January 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n4/pdf/scientificamerican01261884-56.pdf.

Anonymous:1884:EIu

[5421] Anonymous. Engineering inventions. Scientific American, 50(4): 58, January 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n4/pdf/scientificamerican01261884-
58.pdf.

Anonymous:1884:MIbo

[5422] Anonymous. Mechanical inventions. Scientific American, 50(4):58, January 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n4/pdf/scientificamerican01261884-58a.pdf.

Anonymous:1884:AIs

[5423] Anonymous. Agricultural inventions. Scientific American, 50 (4):58, January 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n4/pdf/scientificamerican01261884-
58b.pdf.

Anonymous:1884:MIbp

[5424] Anonymous. Miscellaneous inventions. Scientific American, 50 (4):58, January 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n4/pdf/scientificamerican01261884-
58c.pdf.

Anonymous:1884:Cx

[5425] Anonymous. Correspondence. Scientific American, 50(5):69, February 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n5/pdf/scientificamerican02021884-69.pdf.

Anonymous:1884:BNBq

[5426] Anonymous. Books: New books and publications. Scientific American, 50(5):74, February 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n5/pdf/scientificamerican02021884-
74d.pdf.

Anonymous:1884:BPx

[5427] Anonymous. Business and personal. Scientific American, 50(5): 74–75, February 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n5/pdf/scientificamerican02021884-
74e.pdf.

Anonymous:1884:NQv

[5428] Anonymous. Notes and queries. Scientific American, 50(5):75, February 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n5/pdf/scientificamerican02021884-75.pdf.

Anonymous:1884:IIv

[5429] Anonymous. Index of inventions. Scientific American, 50(5): 75–76, February 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n5/pdf/scientificamerican02021884-
75a.pdf.

Anonymous:1884:PNPs

[5430] Anonymous. Patent notes: Patents. Scientific American, 50 (5):77, February 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n5/pdf/scientificamerican02021884-
77.pdf.

Anonymous:1884:ICS

[5431] Anonymous. Improved crank shaft lathe. Scientific American, 50(5):63, February 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n5/pdf/scientificamerican02021884-
63.pdf.

Anonymous:1884:ER

[5432] Anonymous. The earthquake recorder. Scientific American, 50 (5):63–64, February 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n5/pdf/scientificamerican02021884-
63a.pdf.

Anonymous:1884:NPLb

[5433] Anonymous. Nullification of the Patent Laws. Scientific American, 50(5):64, February 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n5/pdf/scientificamerican02021884-
64.pdf.

Anonymous:1884:AMS

[5434] Anonymous. Another movement for a Swiss Patent Law. Scientific American, 50(5):64–65, February 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v50/n5/pdf/scientificamerican02021884-
64a.pdf.

Anonymous:1884:BAS

[5435] Anonymous. Boracic acid; its sources. Scientific American, 50 (5):65, February 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n5/pdf/scientificamerican02021884-
65.pdf.

Anonymous:1884:PVL

[5436] Anonymous. Proportions for vessels of light draught. Scientific American, 50(5):65, February 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n5/pdf/scientificamerican02021884-
65a.pdf.

Anonymous:1884:CPb

[5437] Anonymous. Care of patterns. Scientific American, 50(5):65, February 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n5/pdf/scientificamerican02021884-65b.pdf.

Anonymous:1884:LL

[5438] Anonymous. Locking latch. Scientific American, 50(5):66, February 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n5/pdf/scientificamerican02021884-66.pdf.

Anonymous:1884:AMC

[5439] Anonymous. Automatic machine for cutting paper and cardboard. Scientific American, 50(5):66, February 2, 1884. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v50/
n5/pdf/scientificamerican02021884-66a.pdf.

Dodson:1884:PBC

[5440] C. Marion Dodson. The propeller, bone — a contribution to science. Scientific American, 50(5):67, February 2, 1884. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v50/
n5/pdf/scientificamerican02021884-67.pdf.

Anonymous:1884:CKB

[5441] Anonymous. Combined knife brick box and grinder. Scientific American, 50(5):67, February 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n5/pdf/scientificamerican02021884-
67a.pdf.

Anonymous:1884:APF

[5442] Anonymous. Aspects of the planets for February. Scientific American, 50(5):68, February 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n5/pdf/scientificamerican02021884-
68.pdf.

Anonymous:1884:BFN

[5443] Anonymous. Bee farming in New South Wales. Scientific American, 50(5):70–71, February 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n5/pdf/scientificamerican02021884-
70.pdf.

Anonymous:1884:EJ

[5444] Anonymous. Electric jewelry. Scientific American, 50(5):71–72, February 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n5/pdf/scientificamerican02021884-71.pdf.

Anonymous:1884:SG

[5445] Anonymous. Scratch gauge. Scientific American, 50(5):72, February 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n5/pdf/scientificamerican02021884-72.pdf.

Anonymous:1884:PT

[5446] Anonymous. Pipe tongs. Scientific American, 50(5):72–73, February 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n5/pdf/scientificamerican02021884-72a.pdf.

Anonymous:1884:EIv

[5447] Anonymous. Engineering inventions. Scientific American, 50(5): 74, February 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n5/pdf/scientificamerican02021884-
74.pdf.

Anonymous:1884:MIbq

[5448] Anonymous. Mechanical inventions. Scientific American, 50(5): 74, February 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n5/pdf/scientificamerican02021884-
74a.pdf.

Anonymous:1884:AIt

[5449] Anonymous. Agricultural inventions. Scientific American, 50 (5):74, February 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n5/pdf/scientificamerican02021884-
74b.pdf.

Anonymous:1884:MIbr

[5450] Anonymous. Miscellaneous inventions. Scientific American, 50 (5):74, February 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n5/pdf/scientificamerican02021884-
74c.pdf.

Anonymous:1884:Cy

[5451] Anonymous. Correspondence. Scientific American, 50(6):85, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n6/pdf/scientificamerican02091884-85.pdf.

Anonymous:1884:BPy

[5452] Anonymous. Business and personal. Scientific American, 50 (6):90, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n6/pdf/scientificamerican02091884-
90d.pdf.

Anonymous:1884:BNBr

[5453] Anonymous. Books: New books and publications. Scientific American, 50(6):90, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n6/pdf/scientificamerican02091884-
90e.pdf.

Anonymous:1884:NQw

[5454] Anonymous. Notes and queries. Scientific American, 50(6):90– 91, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n6/pdf/scientificamerican02091884-
90f.pdf.

Anonymous:1884:IIw

[5455] Anonymous. Index of inventions. Scientific American, 50(6):92, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n6/pdf/scientificamerican02091884-92.pdf.

Anonymous:1884:PNPt

[5456] Anonymous. Patent notes: Patents. Scientific American, 50 (6):93, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n6/pdf/scientificamerican02091884-
93.pdf.

Anonymous:1884:IOI

[5457] Anonymous. An inquiry into the origin of invention. Scientific American, 50(6):80, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n6/pdf/scientificamerican02091884-
80.pdf.

Anonymous:1884:SST

[5458] Anonymous. The standard screw threads. Scientific American, 50(6):80, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n6/pdf/scientificamerican02091884-
80a.pdf.

Anonymous:1884:DHL

[5459] Anonymous. To drill holes in line. Scientific American, 50(6):80, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n6/pdf/scientificamerican02091884-80b.pdf.

Anonymous:1884:MSSb

[5460] Anonymous. Manufacture of silver spoons. Scientific American, 50(6):80, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n6/pdf/scientificamerican02091884-
80c.pdf.

Anonymous:1884:SWb

[5461] Anonymous. Setting-up with the wrench. Scientific American, 50(6):80, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n6/pdf/scientificamerican02091884-
80d.pdf.

Anonymous:1884:Tc

[5462] Anonymous. Trichinosis. Scientific American, 50(6):81, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n6/pdf/scientificamerican02091884-81.pdf.

Anonymous:1884:PCE

[5463] Anonymous. The Patent Committee’s error. Scientific American, 50(6):81, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n6/pdf/scientificamerican02091884-
81a.pdf.

Anonymous:1884:MAR

[5464] Anonymous. Modern and ancient rainfall. Scientific American, 50(6):81, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n6/pdf/scientificamerican02091884-
81b.pdf.

Anonymous:1884:IDI

[5465] Anonymous. Improved dam for irrigating-ditches. Scientific American, 50(6):82, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n6/pdf/scientificamerican02091884-
82.pdf.

Anonymous:1884:SRE

[5466] Anonymous. Single rail elevated railway and train. Scientific American, 50(6):82, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n6/pdf/scientificamerican02091884-
82a.pdf.

Anonymous:1884:IG

[5467] Anonymous. Improved gate. Scientific American, 50(6):82–83, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n6/pdf/scientificamerican02091884-82b.pdf.

Anonymous:1884:CVC

[5468] Anonymous. Constructing, ventilating, and cooling cellars. Scientific American, 50(6):83, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n6/pdf/scientificamerican02091884-
83.pdf.

Anonymous:1884:CPS

[5469] Anonymous. Crank pin for steam engines. Scientific American, 50(6):83–84, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n6/pdf/scientificamerican02091884-
83a.pdf.

Smedley:1884:AFH

[5470] Victor Smedley. The automatic flute and how to make it. Scientific American, 50(6):84–85, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v50/n6/pdf/scientificamerican02091884-
84.pdf.

Anonymous:1884:SPE

[5471] Anonymous. Simple photo-enlarging apparatus. Scientific American, 50(6):86, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n6/pdf/scientificamerican02091884-
86.pdf.

Anonymous:1884:IPC

[5472] Anonymous. Isolating pavilions for contagious diseases. Scientific American, 50(6):87, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n6/pdf/scientificamerican02091884-
87.pdf.

Anonymous:1884:IY

[5473] Anonymous. Ice yachting. Scientific American, 50(6):88, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n6/pdf/scientificamerican02091884-88.pdf.

Anonymous:1884:IBW

[5474] Anonymous. Improved burial windlass. Scientific American, 50 (6):88–89, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n6/pdf/scientificamerican02091884-
88a.pdf.

Anonymous:1884:EIw

[5475] Anonymous. Engineering inventions. Scientific American, 50(6): 90, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n6/pdf/scientificamerican02091884-
90.pdf.

Anonymous:1884:MIbs

[5476] Anonymous. Mechanical inventions. Scientific American, 50(6): 90, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n6/pdf/scientificamerican02091884-
90a.pdf.

Anonymous:1884:AIu

[5477] Anonymous. Agricultural inventions. Scientific American, 50 (6):90, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n6/pdf/scientificamerican02091884-
90b.pdf.

Anonymous:1884:MIbt

[5478] Anonymous. Miscellaneous inventions. Scientific American, 50 (6):90, February 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n6/pdf/scientificamerican02091884-
90c.pdf.

Anonymous:1884:Cz

[5479] Anonymous. Correspondence. Scientific American, 50(7):101– 102, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n7/pdf/scientificamerican02161884-
101.pdf.

Anonymous:1884:BNBs

[5480] Anonymous. Books: New books and publications. Scientific American, 50(7):106, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n7/pdf/scientificamerican02161884-
106d.pdf.

Anonymous:1884:BPz

[5481] Anonymous. Business and personal. Scientific American, 50(7): 107, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n7/pdf/scientificamerican02161884-
107.pdf.

Anonymous:1884:NQx

[5482] Anonymous. Notes and queries. Scientific American, 50(7):107, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n7/pdf/scientificamerican02161884-107a.pdf.

Anonymous:1884:IIx

[5483] Anonymous. Index of inventions. Scientific American, 50(7): 107–108, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n7/pdf/scientificamerican02161884-
107b.pdf.

Anonymous:1884:PNPu

[5484] Anonymous. Patent notes: Patents. Scientific American, 50 (7):109, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n7/pdf/scientificamerican02161884-
109.pdf.

Anonymous:1884:PI

[5485] Anonymous. The position of inventors. Scientific American, 50 (7):96, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n7/pdf/scientificamerican02161884-
96.pdf.

Anonymous:1884:PBCb

[5486] Anonymous. The Pons–Brooks Comet. Scientific American, 50 (7):96, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n7/pdf/scientificamerican02161884-
96a.pdf.

Anonymous:1884:PWW

[5487] Anonymous. Peculiar wood working. Scientific American, 50 (7):96–97, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n7/pdf/scientificamerican02161884-
96b.pdf.

Anonymous:1884:MPb

[5488] Anonymous. A menace to prosperity. Scientific American, 50 (7):97, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n7/pdf/scientificamerican02161884-
97.pdf.

Anonymous:1884:FF

[5489] Anonymous. The February floods. Scientific American, 50(7):97, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n7/pdf/scientificamerican02161884-97a.pdf.

Anonymous:1884:RPS

[5490] Anonymous. A remarkable phenomenon seen at Sulphur Springs, Ohio. Scientific American, 50(7):97, February 16, 1884. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v50/
n7/pdf/scientificamerican02161884-97b.pdf.

Anonymous:1884:POW

[5491] Anonymous. Patent Office work of 1883. Scientific American, 50(7):97, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n7/pdf/scientificamerican02161884-
97c.pdf.

Anonymous:1884:DLW

[5492] Anonymous. Device for lubricating wheels and pulleys. Scientific American, 50(7):98, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n7/pdf/scientificamerican02161884-
98.pdf.

Anonymous:1884:CSP

[5493] Anonymous. A Canadian snow plow. Scientific American, 50 (7):98, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n7/pdf/scientificamerican02161884-
98a.pdf.

Anonymous:1884:MM

[5494] Anonymous. Mechanical movement. Scientific American, 50(7): 98–99, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n7/pdf/scientificamerican02161884-
98b.pdf.

Anonymous:1884:Pf

[5495] Anonymous. “percentograph”. Scientific American, 50(7):99, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n7/pdf/scientificamerican02161884-99.pdf.

Anonymous:1884:FEd

[5496] Anonymous. Fire escape. Scientific American, 50(7):99, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n7/pdf/scientificamerican02161884-99a.pdf.

Anonymous:1884:FSFa

[5497] Anonymous. The floating steam fire engines of New York. Scientific American, 50(7):100, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v50/n7/pdf/scientificamerican02161884-
100.pdf.

Anonymous:1884:TAT

[5498] Anonymous. Terry’s aquatic tricycle. Scientific American, 50(7): 102–103, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n7/pdf/scientificamerican02161884-
102.pdf.

Anonymous:1884:Pg

[5499] Anonymous. Periophthalmus. Scientific American, 50(7):103, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n7/pdf/scientificamerican02161884-103.pdf.

Anonymous:1884:CCF

[5500] Anonymous. A curious case of freezing. Scientific American, 50 (7):103, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n7/pdf/scientificamerican02161884-
103a.pdf.

Anonymous:1884:SSV

[5501] Anonymous. Side spring for vehicles. Scientific American, 50 (7):104, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n7/pdf/scientificamerican02161884-
104.pdf.

Hiscox:1884:FC

[5502] G. D. Hiscox. Filtering cisterns. Scientific American, 50(7): 104–105, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n7/pdf/scientificamerican02161884-
104a.pdf.

Anonymous:1884:EIx

[5503] Anonymous. Engineering inventions. Scientific American, 50(7): 106, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n7/pdf/scientificamerican02161884-
106.pdf.

Anonymous:1884:MIbu

[5504] Anonymous. Mechanical inventions. Scientific American, 50(7): 106, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n7/pdf/scientificamerican02161884-
106a.pdf.

Anonymous:1884:AIv (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n7/pdf/scientificamerican02161884-
106b.pdf.

Anonymous:1884:MIbv

[5506] Anonymous. Miscellaneous inventions. Scientific American, 50 (7):106, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n7/pdf/scientificamerican02161884-
106c.pdf.

Anonymous:1884:Sh

[5507] Anonymous. Special. Scientific American, 50(7):106–107, February 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n7/pdf/scientificamerican02161884-106e.pdf.

Anonymous:1884:Cbb

[5508] Anonymous. Correspondence. Scientific American, 50(8):117– 118, February 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n8/pdf/scientificamerican02231884-
117.pdf.

Anonymous:1884:BPba

[5509] Anonymous. Business and personal. Scientific American, 50(8): 122–123, February 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n8/pdf/scientificamerican02231884-
122d.pdf.

Anonymous:1884:BNBt

[5510] Anonymous. Books: New books and publications. Scientific American, 50(8):123, February 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n8/pdf/scientificamerican02231884-
123.pdf.

Anonymous:1884:NQy

[5511] Anonymous. Notes and queries. Scientific American, 50(8):123, February 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n8/pdf/scientificamerican02231884-123a.pdf.

Anonymous:1884:IIy

[5512] Anonymous. Index of inventions. Scientific American, 50(8): 123–124, February 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n8/pdf/scientificamerican02231884-
123b.pdf.

Anonymous:1884:PNPv

[5513] Anonymous. Patent notes: Patents. Scientific American, 50 (8):125, February 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n8/pdf/scientificamerican02231884-
125.pdf.

Anonymous:1884:IWP

[5514] Anonymous. It was the Pons–Brooks Comet. Scientific American, 50(8):112, February 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n8/pdf/scientificamerican02231884-
112.pdf.

Anonymous:1884:REA

[5515] Anonymous. Relief expeditions for the American Colony at Lady Franklin Bay. Scientific American, 50(8):112–113, February 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n8/pdf/scientificamerican02231884-112a.pdf.

Anonymous:1884:ICCb

[5516] Anonymous. Improved car coupling. Scientific American, 50 (8):114, February 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n8/pdf/scientificamerican02231884-
114.pdf.

Anonymous:1884:SLS

[5517] Anonymous. Steam log setting apparatus for saw mills. Scientific American, 50(8):114, February 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n8/pdf/scientificamerican02231884-
114a.pdf.

Anonymous:1884:ICCc

[5518] Anonymous. Improved car coupling. Scientific American, 50(8): 114–115, February 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n8/pdf/scientificamerican02231884-
114b.pdf.

Anonymous:1884:FWR

[5519] Anonymous. Feed water regulator and alarm. Scientific American, 50(8):115, February 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n8/pdf/scientificamerican02231884-
115.pdf.

Anonymous:1884:STM

[5520] Anonymous. Speaking tube mouth piece. Scientific American, 50(8):115, February 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n8/pdf/scientificamerican02231884-
115a.pdf.

Anonymous:1884:EJE

[5521] Anonymous. End of the Jeannette Expedition. Scientific American, 50(8):116, February 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n8/pdf/scientificamerican02231884-
116.pdf.

Anonymous:1884:DTW

[5522] Anonymous. The duplex tide wheel. Scientific American, 50 (8):118, February 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n8/pdf/scientificamerican02231884-
118.pdf.

Anonymous:1884:EAa

[5523] Anonymous. The exploration of the atmosphere. Scientific American, 50(8):119–120, February 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n8/pdf/scientificamerican02231884-
119.pdf.

Anonymous:1884:IDE

[5524] Anonymous. Improved draught equalizer. Scientific American, 50(8):120, February 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n8/pdf/scientificamerican02231884-
120.pdf.

Anonymous:1884:IT

[5525] Anonymous. Improved truss. Scientific American, 50(8):120– 121, February 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n8/pdf/scientificamerican02231884-
120a.pdf.

Anonymous:1884:EIy

[5526] Anonymous. Engineering inventions. Scientific American, 50(8): 122, February 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n8/pdf/scientificamerican02231884-
122.pdf.

Anonymous:1884:MIbw

[5527] Anonymous. Mechanical inventions. Scientific American, 50(8): 122, February 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n8/pdf/scientificamerican02231884-
122a.pdf.

Anonymous:1884:AIw

[5528] Anonymous. Agricultural inventions. Scientific American, 50(8): 122, February 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n8/pdf/scientificamerican02231884-
122b.pdf.

Anonymous:1884:MIbx

[5529] Anonymous. Miscellaneous inventions. Scientific American, 50 (8):122, February 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n8/pdf/scientificamerican02231884-
122c.pdf.

Anonymous:1884:Cbc

[5530] Anonymous. Correspondence. Scientific American, 50(9):133–134, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n9/pdf/scientificamerican03011884-133.pdf.

Anonymous:1884:BPbb

[5531] Anonymous. Business and personal. Scientific American, 50(9):138, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n9/pdf/scientificamerican03011884-138d.pdf.

Anonymous:1884:BNBu

[5532] Anonymous. Books: New books and publications. Scientific American, 50(9):139, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n9/pdf/scientificamerican03011884-
139.pdf.

Anonymous:1884:NQz

[5533] Anonymous. Notes and queries. Scientific American, 50(9):139, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n9/pdf/scientificamerican03011884-139a.pdf.

Anonymous:1884:IIz

[5534] Anonymous. Index of inventions. Scientific American, 50(9):139–140, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n9/pdf/scientificamerican03011884-139b.pdf.

Anonymous:1884:PNPw

[5535] Anonymous. Patent notes: Patents. Scientific American, 50(9):141, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n9/pdf/scientificamerican03011884-141.pdf.

Anonymous:1884:HFM

[5536] Anonymous. The Hyatt filters — manufactured by the Newark Filtering Co., Newark, N.J. Scientific American, 50(9):127, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n9/pdf/scientificamerican03011884-127.pdf.

Anonymous:1884:RBP

[5537] Anonymous. Robbery as a basis of property right. Scientific American, 50(9):128, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n9/pdf/scientificamerican03011884-
128.pdf.

Anonymous:1884:LPI

[5538] Anonymous. Loose pulleys, or idlers. Scientific American, 50(9):128, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n9/pdf/scientificamerican03011884-128a.pdf.

Anonymous:1884:HKM

[5539] Anonymous. How knurls are made. Scientific American, 50(9):128, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n9/pdf/scientificamerican03011884-128b.pdf.

Anonymous:1884:TLJ

[5540] Anonymous. Tetanus-lock jaw. Scientific American, 50(9):128–129, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n9/pdf/scientificamerican03011884-128c.pdf.

Anonymous:1884:PBCc

[5541] Anonymous. Patent Bills in Congress. Scientific American, 50(9):129, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n9/pdf/scientificamerican03011884-129.pdf.

Anonymous:1884:FDE

[5542] Anonymous. Fire damp explosions. Scientific American, 50(9):129, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n9/pdf/scientificamerican03011884-129a.pdf.

Anonymous:1884:FSFb

[5543] Anonymous. First steam fire engine. Scientific American, 50(9):129, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n9/pdf/scientificamerican03011884-129b.pdf.

Anonymous:1884:BC

[5544] Anonymous. Bee culture. Scientific American, 50(9):129, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n9/pdf/scientificamerican03011884-129c.pdf.

Anonymous:1884:IRH

[5545] Anonymous. Improved revolving harrow. Scientific American, 50 (9):131, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n9/pdf/scientificamerican03011884-
131.pdf.

Anonymous:1884:ISS

[5546] Anonymous. Improved safety switch guard. Scientific American, 50(9):131, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n9/pdf/scientificamerican03011884-
131a.pdf.

Anonymous:1884:IFPc

[5547] Anonymous. Improved foot power hammer. Scientific American, 50(9):131, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n9/pdf/scientificamerican03011884-
131b.pdf.

Anonymous:1884:APMb

[5548] Anonymous. Aspects of the planets for March. Scientific American, 50(9):132, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n9/pdf/scientificamerican03011884-
132.pdf.

Anonymous:1884:IRDb

[5549] Anonymous. Improved rapid dumping cart. Scientific American, 50(9):134, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n9/pdf/scientificamerican03011884-
134.pdf.

Anonymous:1884:PBCd

[5550] Anonymous. The Pons–Brooks Comet. Scientific American, 50 (9):134–135, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n9/pdf/scientificamerican03011884-
134a.pdf.

Anonymous:1884:YPZ

[5551] Anonymous. Young panthers in the Zoological Garden at Dusseldorf, Germany. Scientific American, 50(9):135–136, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n9/pdf/scientificamerican03011884-135.pdf.

Anonymous:1884:NAV

[5552] Anonymous. A new attachment for vises. Scientific American, 50(9):136–137, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v50/n9/pdf/scientificamerican03011884-
136.pdf.

Anonymous:1884:EIz

[5553] Anonymous. Engineering inventions. Scientific American, 50(9):138, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n9/pdf/scientificamerican03011884-138.pdf.

Anonymous:1884:MIby

[5554] Anonymous. Mechanical inventions. Scientific American, 50(9):138, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n9/pdf/scientificamerican03011884-138a.pdf.

Anonymous:1884:AIx

[5555] Anonymous. Agricultural inventions. Scientific American, 50(9):138, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n9/pdf/scientificamerican03011884-138b.pdf.

Anonymous:1884:MIbz

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n9/pdf/scientificamerican03011884-138c.pdf.

Anonymous:1884:Sj

[5557] Anonymous. Special. Scientific American, 50(9):138–139, March 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v50/n9/pdf/scientificamerican03011884-138e.pdf.

Anonymous:1884:Cbd

[5558] Anonymous. Correspondence. Scientific American, 51(1):7, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n1/pdf/scientificamerican07051884-7a.pdf.

Anonymous:1884:BPbc

[5559] Anonymous. Business and personal. Scientific American, 51(1):10, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n1/pdf/scientificamerican07051884-10d.pdf.

Anonymous:1884:BNBv

[5560] Anonymous. Books: New books and publications. Scientific American, 51(1):10, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n1/pdf/scientificamerican07051884-
10e.pdf.

Anonymous:1884:NQba

[5561] Anonymous. Notes and queries. Scientific American, 51(1):10–11, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n1/pdf/scientificamerican07051884-10f.pdf.

Anonymous:1884:IIba

[5562] Anonymous. Index of inventions. Scientific American, 51(1):11–12, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1884:GSP

[5563] Anonymous. The Gaskill steam pumping engines. Scientific American, 51(1):1, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n1/pdf/scientificamerican07051884-
1.pdf.

Anonymous:1884:MDN

[5564] Anonymous. A mechanical dictionary needed. Scientific American, 51(1):2, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n1/pdf/scientificamerican07051884-
2.pdf.

Anonymous:1884:UIS

[5565] Anonymous. The union of iron and steel. Scientific American, 51(1):2, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n1/pdf/scientificamerican07051884-2a.pdf.

Anonymous:1884:NGI

[5566] Anonymous. Natural gas as an industrial factor. Scientific American, 51(1):2, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n1/pdf/scientificamerican07051884-
2b.pdf.

Anonymous:1884:HGB

[5567] Anonymous. How globes are built. Scientific American, 51(1):2, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n1/pdf/scientificamerican07051884-2c.pdf.

Anonymous:1884:ERI

[5568] Anonymous. The equitable rights of inventors before Congress. Scientific American, 51(1):3, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n1/pdf/scientificamerican07051884-
3.pdf.

Anonymous:1884:HCP

[5569] Anonymous. Hand corn planter. Scientific American, 51(1):4, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n1/pdf/scientificamerican07051884-4.pdf.

Anonymous:1884:STb

[5570] Anonymous. Steam trap. Scientific American, 51(1):4, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n1/pdf/scientificamerican07051884-4a.pdf.

Anonymous:1884:RWT

[5571] Anonymous. Rain water trap. Scientific American, 51(1):4, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n1/pdf/scientificamerican07051884-4b.pdf.

Anonymous:1884:WBW

[5572] Anonymous. Well bucket windlass. Scientific American, 51(1):4–5, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n1/pdf/scientificamerican07051884-4c.pdf.

Anonymous:1884:FMP

[5573] Anonymous. Feeder for mills, purifiers, etc. Scientific American, 51(1): 5, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n1/pdf/scientificamerican07051884-5.pdf.

Anonymous:1884:SEF

[5574] Anonymous. Shell extractor for fire arms. Scientific American, 51(1):5, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n1/pdf/scientificamerican07051884-5a.pdf.

Anonymous:1884:MBa

[5575] Anonymous. Miter box. Scientific American, 51(1):5, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n1/pdf/scientificamerican07051884-5b.pdf.

Anonymous:1884:IPW

[5576] Anonymous. Improved propeller wheel. Scientific American, 51(1):7, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n1/pdf/scientificamerican07051884-7.pdf.

Anonymous:1884:CAL

[5577] Anonymous. Cailletet’s apparatus for the liquefaction of gases. Scientific American, 51(1):8, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n1/pdf/scientificamerican07051884-
8.pdf.

Anonymous:1884:NAW

[5578] Anonymous. Novel arrangement of water-tight compartments. Scientific American, 51(1):8, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n1/pdf/scientificamerican07051884-
8a.pdf.

Anonymous:1884:MAa

[5579] Anonymous. The Mexican axolotl. Scientific American, 51(1):9, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n1/pdf/scientificamerican07051884-9.pdf.

Anonymous:1884:EIba

[5580] Anonymous. Engineering inventions. Scientific American, 51(1):10, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n1/pdf/scientificamerican07051884-10.pdf.

Anonymous:1884:MIca

[5581] Anonymous. Mechanical inventions. Scientific American, 51(1):10, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n1/pdf/scientificamerican07051884-10a.pdf.

Anonymous:1884:AIy

[5582] Anonymous. Agricultural inventions. Scientific American, 51(1):10, July 5, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n1/pdf/scientificamerican07051884-10b.pdf.

Anonymous:1884:MIcb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n1/pdf/scientificamerican07051884-10c.pdf.

Anonymous:1884:Cbe

[5584] Anonymous. Correspondence. Scientific American, 51(10):149, September 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n10/pdf/scientificamerican09061884-149.pdf.

Anonymous:1884:BPbd

[5585] Anonymous. Business and personal. Scientific American, 51(10): 154–155, September 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n10/pdf/scientificamerican09061884-
154d.pdf.

Anonymous:1884:NQbb

[5586] Anonymous. Notes and queries. Scientific American, 51(10): 155, September 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n10/pdf/scientificamerican09061884-
155.pdf.

Anonymous:1884:IIbb

[5587] Anonymous. Index of inventions. Scientific American, 51(10): 155–156, September 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n10/pdf/scientificamerican09061884-
155a.pdf.

Anonymous:1884:PNPx

[5588] Anonymous. Patent notes: Patents. Scientific American, 51 (10):157, September 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n10/pdf/scientificamerican09061884-
157.pdf.

Anonymous:1884:RPA

[5589] Anonymous. Ragona’s Portable Astronomical, Magnetic, and Meteorological Observatory. Scientific American, 51(10):143–144, September 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n10/pdf/scientificamerican09061884-143.pdf.

Anonymous:1884:RPF

[5590] Anonymous. Rufus Porter, founder of the Scientific American. Scientific American, 51(10):144–145, September 6, 1884. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v51/
n10/pdf/scientificamerican09061884-144.pdf.

Anonymous:1884:ME

[5591] Anonymous. Mechanics in education. Scientific American, 51 (10):145, September 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n10/pdf/scientificamerican09061884-
145.pdf.

Anonymous:1884:PPN

[5592] Anonymous. The problematic Planet Neith. Scientific American, 51 (10):145–146, September 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n10/pdf/scientificamerican09061884-
145a.pdf.

Anonymous:1884:GCD

[5593] Anonymous. Grain crusher and disintegrator. Scientific American, 51(10):146, September 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n10/pdf/scientificamerican09061884-
146.pdf.

Anonymous:1884:SOL

[5594] Anonymous. Self-oiling loose pulley. Scientific American, 51 (10):146, September 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n10/pdf/scientificamerican09061884-
146a.pdf.

Anonymous:1884:DGL

[5595] Anonymous. Door and gate latch. Scientific American, 51(10): 146, September 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n10/pdf/scientificamerican09061884-
146b.pdf.

Anonymous:1884:TR

[5596] Anonymous.  Tinsmith’s roller.  Scientific American, 51(10):

147, September 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n10/pdf/scientificamerican09061884-
147.pdf.

Anonymous:1884:SCEa

[5597] Anonymous. Safety catch for elevators. Scientific American, 51 (10):147, September 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n10/pdf/scientificamerican09061884-
147a.pdf.

Anonymous:1884:ISHc

[5598] Anonymous. Improved steam heater. Scientific American, 51(10): 147–148, September 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n10/pdf/scientificamerican09061884-
147b.pdf.

Anonymous:1884:Cbf

[5599] Anonymous. The cholera in 1884. Scientific American, 51(10): 150–152, September 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n10/pdf/scientificamerican09061884-
150.pdf.

Anonymous:1884:EWa

[5600] Anonymous. An English wolf. Scientific American, 51(10):152, September 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n10/pdf/scientificamerican09061884-152.pdf.

Anonymous:1884:FPC

[5601] Anonymous. Fumigating passengers for cholera. Scientific American, 51(10):152–153, September 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n10/pdf/scientificamerican09061884-
152a.pdf.

Anonymous:1884:EIbb

[5602] Anonymous. Engineering inventions. Scientific American, 51(10): 154, September 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n10/pdf/scientificamerican09061884-
154.pdf.

Anonymous:1884:MIcc

[5603] Anonymous. Mechanical inventions. Scientific American, 51(10): 154, September 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n10/pdf/scientificamerican09061884-
154a.pdf.

Anonymous:1884:AIz

[5604] Anonymous. Agricultural inventions. Scientific American, 51 (10):154, September 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n10/pdf/scientificamerican09061884-
154b.pdf.

Anonymous:1884:MIcd

[5605] Anonymous. Miscellaneous inventions. Scientific American, 51 (10):154, September 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n10/pdf/scientificamerican09061884-
154c.pdf.

Anonymous:1884:Cbg

[5606] Anonymous. Correspondence. Scientific American, 51(11):165, September 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n11/pdf/scientificamerican09131884-165.pdf.

Anonymous:1884:BPbe

[5607] Anonymous. Business and personal. Scientific American, 51 (11):169, September 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n11/pdf/scientificamerican09131884-
169d.pdf.

Anonymous:1884:NQbc

[5608] Anonymous. Notes and queries. Scientific American, 51(11): 169–171, September 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n11/pdf/scientificamerican09131884-
169e.pdf.

Anonymous:1884:IIbc

[5609] Anonymous. Index of inventions. Scientific American, 51(11): 171–172, September 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n11/pdf/scientificamerican09131884-
171.pdf.

Anonymous:1884:PNPy

[5610] Anonymous. Patent notes: Patents. Scientific American, 51 (11):173, September 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n11/pdf/scientificamerican09131884-
173.pdf.

Anonymous:1884:SSE

[5611] Anonymous. The Shipman steam engine. Scientific American, 51 (11):159, September 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n11/pdf/scientificamerican09131884-
159.pdf.

Anonymous:1884:IEEb

[5612] Anonymous. The International Electrical Exposition. Scientific American, 51(11):160, September 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n11/pdf/scientificamerican09131884-
160.pdf.

Anonymous:1884:EWb

[5613] Anonymous. Emery wheels. Scientific American, 51(11):160, September 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n11/pdf/scientificamerican09131884-160a.pdf.

Anonymous:1884:WPO

[5614] Anonymous. White pine ornamentation. Scientific American, 51 (11):160–162, September 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n11/pdf/scientificamerican09131884-
160b.pdf.

Anonymous:1884:DMV

[5615] Anonymous. Driving mechanism for velocipedes. Scientific American, 51(11):162, September 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n11/pdf/scientificamerican09131884-
162.pdf.

Anonymous:1884:TDT

[5616] Anonymous. Tightening device for trunk straps. Scientific American, 51(11):162, September 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n11/pdf/scientificamerican09131884-
162a.pdf.

Anonymous:1884:SCC

[5617] Anonymous. Sulky cornstalk cutter. Scientific American, 51 (11):162, September 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n11/pdf/scientificamerican09131884-
162b.pdf.

Anonymous:1884:AHL

[5618] Anonymous. Apparatus for hydrating lime. Scientific American, 51(11):163, September 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n11/pdf/scientificamerican09131884-
163.pdf.

Anonymous:1884:CCP

[5619] Anonymous. Combined calendar, paper weight, etc. Scientific American, 51(11):163, September 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n11/pdf/scientificamerican09131884-
163a.pdf.

Anonymous:1884:RHB

[5620] Anonymous. Radiators for heating buildings. Scientific American, 51(11):165, September 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n11/pdf/scientificamerican09131884-
165a.pdf.

Anonymous:1884:HMT

[5621] Anonymous. A home-made turntable. Scientific American, 51 (11):165, September 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n11/pdf/scientificamerican09131884-
165b.pdf.

Anonymous:1884:VG

[5622] Anonymous. The Victorian gunboats. Scientific American, 51 (11):166, September 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n11/pdf/scientificamerican09131884-
166.pdf.

Anonymous:1884:PKL

[5623] Anonymous. Phases of kangaroo life. Scientific American, 51(11): 167–168, September 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n11/pdf/scientificamerican09131884-
167.pdf.

Anonymous:1884:EIbc

[5624] Anonymous. Engineering inventions. Scientific American, 51(11): 169, September 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n11/pdf/scientificamerican09131884-
169.pdf.

Anonymous:1884:MIce

[5625] Anonymous. Mechanical inventions. Scientific American, 51(11): 169, September 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n11/pdf/scientificamerican09131884-
169a.pdf.

Anonymous:1884:AIba

[5626] Anonymous. Agricultural inventions. Scientific American, 51 (11):169, September 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n11/pdf/scientificamerican09131884-
169b.pdf.

Anonymous:1884:MIcf

[5627] Anonymous. Miscellaneous inventions. Scientific American, 51 (11):169, September 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n11/pdf/scientificamerican09131884-
169c.pdf.

Anonymous:1884:Cbh

[5628] Anonymous. Correspondence. Scientific American, 51(12):181, September 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n12/pdf/scientificamerican09201884-181.pdf.

Anonymous:1884:BPbf

[5629] Anonymous. Business and personal. Scientific American, 51 (12):187, September 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n12/pdf/scientificamerican09201884-
187d.pdf.

Anonymous:1884:NQbd

[5630] Anonymous. Notes and queries. Scientific American, 51(12): 187, September 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n12/pdf/scientificamerican09201884-
187e.pdf.

Anonymous:1884:IIbd

[5631] Anonymous. Index of inventions. Scientific American, 51(12): 187–188, September 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n12/pdf/scientificamerican09201884-
187f.pdf.

Anonymous:1884:PNPz

[5632] Anonymous. Patent notes: Patents. Scientific American, 51 (12):189, September 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n12/pdf/scientificamerican09201884-
189.pdf.

Anonymous:1884:MAb

[5633] Anonymous. More about the A.A.A.S. Scientific American, 51 (12):176–177, September 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n12/pdf/scientificamerican09201884-
176.pdf.

Anonymous:1884:AS

[5634] Anonymous. For the advancement of science. Scientific American, 51(12):177, September 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n12/pdf/scientificamerican09201884-
177.pdf.

Anonymous:1884:MPW

[5635] Anonymous. Magnetization of pocket watches at the Electrical Exhibition. Scientific American, 51(12):177, September 20, 1884. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v51/
n12/pdf/scientificamerican09201884-177a.pdf.

Anonymous:1884:DBPa

[5636] Anonymous. A double baking pan. Scientific American, 51 (12):178, September 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n12/pdf/scientificamerican09201884-
178.pdf.

Anonymous:1884:CWb

[5637] Anonymous. Current wheel. Scientific American, 51(12):178, September 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n12/pdf/scientificamerican09201884-178a.pdf.

Anonymous:1884:THD

[5638] Anonymous. Trace holder and detacher. Scientific American, 51 (12):178, September 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n12/pdf/scientificamerican09201884-
178b.pdf.

Anonymous:1884:CTCb

[5639] Anonymous. Combination tool for carpenters. Scientific American, 51(12):178, September 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n12/pdf/scientificamerican09201884-
178c.pdf.

Anonymous:1884:IMMb

[5640] Anonymous. Improved mechanical movement. Scientific American, 51(12):179, September 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n12/pdf/scientificamerican09201884-
179.pdf.

Anonymous:1884:IMT

[5641] Anonymous. Improved machine for tonguing and grooving boards. Scientific American, 51(12):179, September 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v51/n12/pdf/scientificamerican09201884-
179a.pdf.

Anonymous:1884:SBa

[5642] Anonymous. Steam boiler. Scientific American, 51(12):179, September 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n12/pdf/scientificamerican09201884-179b.pdf.

Anonymous:1884:Te

[5643] Anonymous. The telephone. Scientific American, 51(12):180– 181, September 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n12/pdf/scientificamerican09201884-
180.pdf.

Anonymous:1884:TPC

[5644] Anonymous. Transmitting power from a central station. Scientific American, 51(12):182, September 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v51/n12/pdf/scientificamerican09201884-
182.pdf.

Anonymous:1884:NCN

[5645] Anonymous. The Northeast Coast of New Guinea. Scientific American, 51(12):183–184, September 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n12/pdf/scientificamerican09201884-
183.pdf.

Anonymous:1884:DRPd

[5646] Anonymous. Decisions relating to patents, trade marks, etc. Scientific American, 51(12):185–186, September 20, 1884. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v51/
n12/pdf/scientificamerican09201884-185.pdf.

Anonymous:1884:EIbd

[5647] Anonymous. Engineering inventions. Scientific American, 51(12): 187, September 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n12/pdf/scientificamerican09201884-
187.pdf.

Anonymous:1884:MIcg

[5648] Anonymous. Mechanical inventions. Scientific American, 51(12): 187, September 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n12/pdf/scientificamerican09201884-
187a.pdf.

Anonymous:1884:AIbb

[5649] Anonymous. Agricultural inventions. Scientific American, 51 (12):187, September 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n12/pdf/scientificamerican09201884-
187b.pdf.

Anonymous:1884:MIch

[5650] Anonymous. Miscellaneous inventions. Scientific American, 51 (12):187, September 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n12/pdf/scientificamerican09201884-
187c.pdf.

Anonymous:1884:Cbi

[5651] Anonymous. Correspondence. Scientific American, 51(13):197, September 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n13/pdf/scientificamerican09271884-197.pdf.

Anonymous:1884:BPbg

[5652] Anonymous. Business and personal. Scientific American, 51 (13):202, September 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n13/pdf/scientificamerican09271884-
202c.pdf.

Anonymous:1884:NQbe

[5653] Anonymous. Notes and queries. Scientific American, 51(13): 202–203, September 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n13/pdf/scientificamerican09271884-
202d.pdf.

Anonymous:1884:IIbe

[5654] Anonymous. Index of inventions. Scientific American, 51(13): 203–204, September 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n13/pdf/scientificamerican09271884-
203.pdf.

Anonymous:1884:SRG

[5655] Anonymous. The Siemens regenerative gas lamp. Scientific American, 51(13):191, September 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n13/pdf/scientificamerican09271884-
191.pdf.

Anonymous:1884:SCIa

[5656] Anonymous. Soldering on cast iron. Scientific American, 51 (13):192, September 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n13/pdf/scientificamerican09271884-
192.pdf.

Anonymous:1884:RHW

[5657] Anonymous. Raw hide wheels. Scientific American, 51(13):192– 193, September 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n13/pdf/scientificamerican09271884-
192a.pdf.

Anonymous:1884:WDSa

[5658] Anonymous. A watering device for stock car. Scientific American, 51(13):194, September 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n13/pdf/scientificamerican09271884-
194.pdf.

Anonymous:1884:IDK

[5659] Anonymous. An improved door knob. Scientific American, 51 (13):194, September 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n13/pdf/scientificamerican09271884-
194a.pdf.

Anonymous:1884:IFEa

[5660] Anonymous. An improved fire escape. Scientific American, 51 (13):194, September 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n13/pdf/scientificamerican09271884-
194b.pdf.

Anonymous:1884:RTH

[5661] Anonymous. Ratchet tool handle. Scientific American, 51(13): 195, September 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n13/pdf/scientificamerican09271884-
195.pdf.

Anonymous:1884:Ha

[5662] Anonymous. Horseshoes. Scientific American, 51(13):195, September 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n13/pdf/scientificamerican09271884-195a.pdf.

Anonymous:1884:ICF

[5663] Anonymous. An improved clock frame. Scientific American, 51 (13):195–196, September 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n13/pdf/scientificamerican09271884-
195b.pdf.

Anonymous:1884:BHB

[5664] Anonymous. Boiler for heating buildings. Scientific American, 51 (13):196, September 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n13/pdf/scientificamerican09271884-
196.pdf.

Anonymous:1884:NSCb

[5665] Anonymous. A new steam carriage. Scientific American, 51 (13):198, September 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n13/pdf/scientificamerican09271884-
198.pdf.

Anonymous:1884:IHR

[5666] Anonymous. Improved Hotchkiss rapid six pounder gun. Scientific American, 51(13):198–199, September 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v51/n13/pdf/scientificamerican09271884-
198a.pdf.

Anonymous:1884:MRK

[5667] Anonymous. Messrs. Renard and Krebs’ electric balloon. Scientific American, 51(13):199–200, September 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v51/n13/pdf/scientificamerican09271884-
199.pdf.

Anonymous:1884:DRPe

[5668] Anonymous. Decisions relating to patents. Scientific American, 51 (13):201, September 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n13/pdf/scientificamerican09271884-
201.pdf.

Anonymous:1884:EIbe

[5669] Anonymous. Engineering inventions. Scientific American, 51(13): 202, September 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n13/pdf/scientificamerican09271884-
202.pdf.

Anonymous:1884:AIbc

[5670] Anonymous. Agricultural inventions. Scientific American, 51 (13):202, September 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n13/pdf/scientificamerican09271884-
202a.pdf.

Anonymous:1884:MIci

[5671] Anonymous. Miscellaneous inventions. Scientific American, 51 (13):202, September 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n13/pdf/scientificamerican09271884-
202b.pdf.

Anonymous:1884:BNBw

[5672] Anonymous. Books: New books and publications. Scientific American, 51(14):218, October 4, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n14/pdf/scientificamerican10041884-
218c.pdf.

Anonymous:1884:BPbh

[5673] Anonymous. Business and personal. Scientific American, 51(14):218, October 4, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n14/pdf/scientificamerican10041884-218d.pdf.

Anonymous:1884:NQbf

[5674] Anonymous. Notes and queries. Scientific American, 51(14):218–219, October 4, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n14/pdf/scientificamerican10041884-218e.pdf.

Anonymous:1884:IIbf

[5675] Anonymous. Index of inventions. Scientific American, 51(14):219–220, October 4, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n14/pdf/scientificamerican10041884-219.pdf.

Anonymous:1884:BSM

[5676] Anonymous. Band saw mills for logs. Scientific American, 51 (14):207–208, October 4, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n14/pdf/scientificamerican10041884-
207.pdf.

Anonymous:1884:GTS

[5677] Anonymous. The great telephone suit. Scientific American, 51 (14):208–209, October 4, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n14/pdf/scientificamerican10041884-
208.pdf.

Anonymous:1884:GMb

[5678] Anonymous. Grinding materials. Scientific American, 51(14):209, October 4, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n14/pdf/scientificamerican10041884-209.pdf.

Anonymous:1884:APO

[5679] Anonymous. Aspects of the planets for October. Scientific American, 51(14):209, October 4, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n14/pdf/scientificamerican10041884-
209a.pdf.

Anonymous:1884:IA

[5680] Anonymous. An improved axle. Scientific American, 51(14):210, October 4, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n14/pdf/scientificamerican10041884-210.pdf.

Anonymous:1884:ICd

[5681] Anonymous. An improved cultivator. Scientific American, 51(14):210, October 4, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n14/pdf/scientificamerican10041884-210a.pdf.

Anonymous:1884:IF

[5682] Anonymous. Improved faucet. Scientific American, 51(14):210, October 4, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n14/pdf/scientificamerican10041884-210b.pdf.

Anonymous:1884:DC

[5683] Anonymous. Dumping car. Scientific American, 51(14):210, October 4, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n14/pdf/scientificamerican10041884-210c.pdf.

Anonymous:1884:CTb

[5684] Anonymous. Combination tool. Scientific American, 51(14):211, October 4, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n14/pdf/scientificamerican10041884-211.pdf.

Anonymous:1884:SFB

[5685] Anonymous. Seat and foot board for row boats. Scientific American, 51(14):211, October 4, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n14/pdf/scientificamerican10041884-
211a.pdf.

Anonymous:1884:IRCa

[5686] Anonymous. Improved “rapid” cupola. Scientific American, 51 (14):211–212, October 4, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n14/pdf/scientificamerican10041884-
211b.pdf.

Anonymous:1884:SASa

[5687] Anonymous. Self-acting Speing leg brace. Scientific American, 51 (14):212–213, October 4, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n14/pdf/scientificamerican10041884-
212.pdf.

Anonymous:1884:MRP

[5688] Anonymous. The Metropolitan Railway of Paris. Scientific American, 51(14):214, October 4, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n14/pdf/scientificamerican10041884-
214.pdf.

Anonymous:1884:PBd

[5689] Anonymous. Paraffine brush. Scientific American, 51(14):214, October 4, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n14/pdf/scientificamerican10041884-214a.pdf.

Anonymous:1884:EC

[5690] Anonymous. The edible crab. Scientific American, 51(14):215, October 4, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n14/pdf/scientificamerican10041884-215.pdf.

Anonymous:1884:SMb

[5691] Anonymous. Saddle men. Scientific American, 51(14):215–216, October 4, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n14/pdf/scientificamerican10041884-215a.pdf.

Anonymous:1884:O

[5692] Anonymous. The ocarina. Scientific American, 51(14):216–217, October 4, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n14/pdf/scientificamerican10041884-216.pdf.

Anonymous:1884:EIbf

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n14/pdf/scientificamerican10041884-218.pdf.

Anonymous:1884:MIcj

[5694] Anonymous. Mechanical inventions. Scientific American, 51(14):218, October 4, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n14/pdf/scientificamerican10041884-218a.pdf.

Anonymous:1884:MIck

[5695] Anonymous. Miscellaneous inventions. Scientific American, 51(14):218, October 4, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n14/pdf/scientificamerican10041884-218b.pdf.

Anonymous:1884:Cbj

[5696] Anonymous. Correspondence. Scientific American, 51(15):228, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n15/pdf/scientificamerican10111884-228.pdf.

Anonymous:1884:BNBx

[5697] Anonymous. Books: New books and publications. Scientific American, 51(15):234, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n15/pdf/scientificamerican10111884-
234c.pdf.

Anonymous:1884:BPbi

[5698] Anonymous. Business and personal. Scientific American, 51 (15):234, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n15/pdf/scientificamerican10111884-
234d.pdf.

Anonymous:1884:NQbg

[5699] Anonymous. Notes and queries. Scientific American, 51(15): 234–235, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n15/pdf/scientificamerican10111884-
234e.pdf.

Anonymous:1884:IIbg

[5700] Anonymous. Index of inventions. Scientific American, 51(15): 235–236, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n15/pdf/scientificamerican10111884-
235.pdf.

Anonymous:1884:LLS

[5701] Anonymous. Leveling and lining shafts. Scientific American, 51 (15):224, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n15/pdf/scientificamerican10111884-
224.pdf.

Anonymous:1884:TMP

[5702] Anonymous. Turning metals to patterns. Scientific American, 51 (15):224–225, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n15/pdf/scientificamerican10111884-
224a.pdf.

Anonymous:1884:RSa

[5703] Anonymous. Rein support. Scientific American, 51(15):226, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n15/pdf/scientificamerican10111884-226.pdf.

Anonymous:1884:CCd

[5704] Anonymous. Car coupling. Scientific American, 51(15):226, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n15/pdf/scientificamerican10111884-226a.pdf.

Anonymous:1884:PFH

[5705] Anonymous. Paper file or holder. Scientific American, 51(15):226, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n15/pdf/scientificamerican10111884-226b.pdf.

Anonymous:1884:TP

[5706] Anonymous. Tree protector. Scientific American, 51(15):226, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n15/pdf/scientificamerican10111884-226c.pdf.

Anonymous:1884:CGa

[5707] Anonymous. Car gate. Scientific American, 51(15):226, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n15/pdf/scientificamerican10111884-226d.pdf.

Anonymous:1884:ERF

[5708] Anonymous. The electric railway at Frankfort, Germany. Scientific American, 51(15):227, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v51/n15/pdf/scientificamerican10111884-
227.pdf.

Anonymous:1884:WSa

[5709] Anonymous. Wagon seat. Scientific American, 51(15):228, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n15/pdf/scientificamerican10111884-228a.pdf.

Richter:1884:FRC

[5710] J. E. Richter. The Fish River Caves, Near Sydney, Australia. Scientific American, 51(15):229, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v51/n15/pdf/scientificamerican10111884-
229.pdf.

Anonymous:1884:KMD

[5711] Anonymous. The Keely motor deception. Scientific American, 51 (15):230–231, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n15/pdf/scientificamerican10111884-
230.pdf.

Hovey:1884:AL

[5712] H. C. Hovey. The ant lion. Scientific American, 51(15):231–232, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n15/pdf/scientificamerican10111884-231.pdf.

Anonymous:1884:GPMb

[5713] Anonymous. Gas pressure moderator. Scientific American, 51 (15):232–233, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n15/pdf/scientificamerican10111884-
232.pdf.

Anonymous:1884:EIbg

[5714] Anonymous. Engineering inventions. Scientific American, 51(15): 234, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n15/pdf/scientificamerican10111884-
234.pdf.

Anonymous:1884:AIbd

[5715] Anonymous. Agricultural inventions. Scientific American, 51 (15):234, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n15/pdf/scientificamerican10111884-
234a.pdf.

Anonymous:1884:MIcl

[5716] Anonymous. Miscellaneous inventions. Scientific American, 51 (15):234, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n15/pdf/scientificamerican10111884-
234b.pdf.

Anonymous:1884:Cbk

[5717] Anonymous. Correspondence. Scientific American, 51(16):245, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n16/pdf/scientificamerican10181884-245.pdf.

Anonymous:1884:BNBy

[5718] Anonymous. Books: New books and publications. Scientific American, 51(16):250, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n16/pdf/scientificamerican10181884-
250c.pdf.

Anonymous:1884:BPbj

[5719] Anonymous. Business and personal. Scientific American, 51 (16):250, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n16/pdf/scientificamerican10181884-
250d.pdf.

Anonymous:1884:NQbh

[5720] Anonymous. Notes and queries. Scientific American, 51(16):251, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n16/pdf/scientificamerican10181884-251.pdf.

Anonymous:1884:IIbh

[5721] Anonymous. Index of inventions. Scientific American, 51(16): 251–252, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n16/pdf/scientificamerican10181884-
251a.pdf.

Anonymous:1884:EEP

[5722] Anonymous. The Edison Exhibit at the Philadelphia Electrical Exhibition. Scientific American, 51(16):239–246, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v51/
n16/pdf/scientificamerican10181884-239.pdf.

Anonymous:1884:RDC

[5723] Anonymous. Recent decisions of the Commissioner of Patents concerning trade marks and labels. Scientific American, 51(16):240, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n16/pdf/scientificamerican10181884-240.pdf.

Anonymous:1884:SSb

[5724] Anonymous. Shop savings. Scientific American, 51(16):240, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n16/pdf/scientificamerican10181884-240a.pdf.

Anonymous:1884:PRAb

[5725] Anonymous. A project for the reorganization of the Army. Scientific American, 51(16):241, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v51/n16/pdf/scientificamerican10181884-
241.pdf.

Anonymous:1884:SBb

[5726] Anonymous. Shafts and belts. Scientific American, 51(16):241, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n16/pdf/scientificamerican10181884-241a.pdf.

Anonymous:1884:PGI

[5727] Anonymous. Poisoning from galvanized iron. Scientific American, 51(16):241, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n16/pdf/scientificamerican10181884-
241b.pdf.

Anonymous:1884:SDT

[5728] Anonymous. Setting diamonds in tools. Scientific American, 51 (16):242, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n16/pdf/scientificamerican10181884-
242.pdf.

Anonymous:1884:IPb

[5729] Anonymous. Improved press. Scientific American, 51(16):242, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n16/pdf/scientificamerican10181884-242a.pdf.

Anonymous:1884:LR

[5730] Anonymous. Life raft. Scientific American, 51(16):242, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n16/pdf/scientificamerican10181884-242b.pdf.

Anonymous:1884:IPO

[5731] Anonymous. The industrial production of oxygen and nitrogen. Scientific American, 51(16):243–244, October 11, 1884. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v51/
n16/pdf/scientificamerican10181884-243.pdf.

Anonymous:1884:USE

[5732] Anonymous. The United States Electric Light Co.’s exhibit at the Philadelphia Electrical Exhibition. Scientific American, 51(16):246, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n16/pdf/scientificamerican10181884-246.pdf.

Anonymous:1884:BUC

[5733] Anonymous. Botanizing upon a coin. Scientific American, 51 (16):247, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n16/pdf/scientificamerican10181884-
247.pdf.

Anonymous:1884:YC

[5734] Anonymous. Young chimpanzees. Scientific American, 51(16): 247–248, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n16/pdf/scientificamerican10181884-
247a.pdf.

Anonymous:1884:IVb

[5735] Anonymous. Improved vise. Scientific American, 51(16):248, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n16/pdf/scientificamerican10181884-248.pdf.

Anonymous:1884:CPW

[5736] Anonymous. Combined paper weight and pencil sharpener. Scientific American, 51(16):248, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v51/n16/pdf/scientificamerican10181884-
248a.pdf.

Anonymous:1884:DRPf

[5737] Anonymous. Decisions relating to patents. Scientific American, 51 (16):248–249, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n16/pdf/scientificamerican10181884-
248b.pdf.

Anonymous:1884:MIcm

[5738] Anonymous. Mechanical inventions. Scientific American, 51(16): 250, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n16/pdf/scientificamerican10181884-
250.pdf.

Anonymous:1884:AIbe

[5739] Anonymous. Agricultural inventions. Scientific American, 51 (16):250, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n16/pdf/scientificamerican10181884-
250a.pdf.

Anonymous:1884:MIcn

[5740] Anonymous. Miscellaneous inventions. Scientific American, 51 (16):250, October 11, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n16/pdf/scientificamerican10181884-
250b.pdf.

Anonymous:1884:Cbl

[5741] Anonymous. Correspondence. Scientific American, 51(17):264, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n17/pdf/scientificamerican10251884-264.pdf.

Anonymous:1884:BNBz

[5742] Anonymous. Books: New books and publications. Scientific American, 51(17):266, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n17/pdf/scientificamerican10251884-
266d.pdf.

Anonymous:1884:BPbk

[5743] Anonymous. Business and personal. Scientific American, 51 (17):266, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n17/pdf/scientificamerican10251884-
266e.pdf.

Anonymous:1884:NQbi

[5744] Anonymous. Notes and queries. Scientific American, 51(17): 266–267, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n17/pdf/scientificamerican10251884-
266f.pdf.

Anonymous:1884:IIbi

[5745] Anonymous. Index of inventions. Scientific American, 51(17):268, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n17/pdf/scientificamerican10251884-268.pdf.

Anonymous:1884:THE

[5746] Anonymous. The Thomson–Houston Exhibit at the International Electrical Exposition, Philadelphia. Scientific American, 51(17): 255–256, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n17/pdf/scientificamerican10251884-
255.pdf.

Anonymous:1884:PIN

[5747] Anonymous. A prize for inventors. — need of passenger and freight car brakes. Scientific American, 51(17):256–257, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n17/pdf/scientificamerican10251884-256.pdf.

Anonymous:1884:HGQ

[5748] Anonymous. The heavy gun question. Scientific American, 51 (17):257, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n17/pdf/scientificamerican10251884-
257.pdf.

Anonymous:1884:TRM

[5749] Anonymous. Two remarkable meteors. Scientific American, 51 (17):257, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n17/pdf/scientificamerican10251884-
257a.pdf.

Anonymous:1884:ISCc

[5750] Anonymous. An improved screw cutting lathe. Scientific American, 51(17):258, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n17/pdf/scientificamerican10251884-
258.pdf.

Anonymous:1884:MBb

[5751] Anonymous. Mail bag. Scientific American, 51(17):258, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n17/pdf/scientificamerican10251884-258a.pdf.

Anonymous:1884:CCe

[5752] Anonymous. Car coupling. Scientific American, 51(17):258, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n17/pdf/scientificamerican10251884-258b.pdf.

Anonymous:1884:JEC

[5753] Anonymous. Joint for electric conductors. Scientific American, 51(17):258, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n17/pdf/scientificamerican10251884-
258c.pdf.

Anonymous:1884:EML

[5754] Anonymous. Electric mouth lamp and laryngoscope. Scientific American, 51(17):259, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n17/pdf/scientificamerican10251884-
259.pdf.

Anonymous:1884:HM

[5755] Anonymous. Hoisting machine. Scientific American, 51(17):259– 260, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n17/pdf/scientificamerican10251884-
259a.pdf.

Anonymous:1884:TDN

[5756] Anonymous. Treatment of deformities of the nose. Scientific American, 51(17):260–261, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n17/pdf/scientificamerican10251884-
260.pdf.

Anonymous:1884:MGC

[5757] Anonymous. Making gas from crude petroleum. Scientific American, 51(17):262, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n17/pdf/scientificamerican10251884-
262.pdf.

Anonymous:1884:GRT

[5758] Anonymous. Glass rolling table. Scientific American, 51(17):262, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n17/pdf/scientificamerican10251884-262a.pdf.

Anonymous:1884:AHC

[5759] Anonymous. Ancient hydraulic clocks. Scientific American, 51 (17):263, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n17/pdf/scientificamerican10251884-
263.pdf.

Anonymous:1884:RWS

[5760] Anonymous. Reversible wall scraper. Scientific American, 51 (17):264, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n17/pdf/scientificamerican10251884-
264a.pdf.

Anonymous:1884:LC

[5761] Anonymous. Lunar chart. Scientific American, 51(17):264–265, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n17/pdf/scientificamerican10251884-264b.pdf.

Anonymous:1884:EIbh

[5762] Anonymous. Engineering inventions. Scientific American, 51(17): 266, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n17/pdf/scientificamerican10251884-
266.pdf.

Anonymous:1884:MIco

[5763] Anonymous. Mechanical inventions. Scientific American, 51(17): 266, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n17/pdf/scientificamerican10251884-
266a.pdf.

Anonymous:1884:AIbf

[5764] Anonymous. Agricultural inventions. Scientific American, 51 (17):266, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n17/pdf/scientificamerican10251884-
266b.pdf.

Anonymous:1884:MIcp

[5765] Anonymous. Miscellaneous inventions. Scientific American, 51 (17):266, October 25, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n17/pdf/scientificamerican10251884-
266c.pdf.

Anonymous:1884:BNBba

[5766] Anonymous. Books: New books and publications. Scientific American, 51(18):283, November 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n18/pdf/scientificamerican11011884-
283c.pdf.

Anonymous:1884:BPbl

[5767] Anonymous. Business and personal. Scientific American, 51 (18):283, November 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n18/pdf/scientificamerican11011884-
283d.pdf.

Anonymous:1884:NQbj

[5768] Anonymous. Notes and queries. Scientific American, 51(18): 283, November 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n18/pdf/scientificamerican11011884-
283e.pdf.

Anonymous:1884:IIbj

[5769] Anonymous. Index of inventions. Scientific American, 51(18): 283–284, November 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n18/pdf/scientificamerican11011884-
283f.pdf.

Anonymous:1884:PNPba

[5770] Anonymous. Patent notes: Patents. Scientific American, 51 (18):285, November 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n18/pdf/scientificamerican11011884-
285.pdf.

Anonymous:1884:EBEa

[5771] Anonymous. The exhibit of the Brush Electric Company at the International Electrical Exposition Philadelphia. Scientific American, 51(18):271, November 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n18/pdf/scientificamerican11011884-
271.pdf.

Anonymous:1884:CSe

[5772] Anonymous. Changeable speeds. Scientific American, 51(18): 272, November 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n18/pdf/scientificamerican11011884-
272.pdf.

Anonymous:1884:ATc

[5773] Anonymous. Automatic torpedoes. Scientific American, 51(18): 272, November 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n18/pdf/scientificamerican11011884-
272a.pdf.

Anonymous:1884:BPCa

[5774] Anonymous. Baldness — its prevention and cure. Scientific American, 51(18):272, November 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n18/pdf/scientificamerican11011884-
272b.pdf.

Anonymous:1884:UI

[5775] Anonymous. Upsetting of iron. Scientific American, 51(18):273, November 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n18/pdf/scientificamerican11011884-273.pdf.

Anonymous:1884:APN

[5776] Anonymous. Aspects of the planets in November. Scientific American, 51(18):273, November 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n18/pdf/scientificamerican11011884-
273a.pdf.

Anonymous:1884:LPB

[5777] Anonymous. Low pressure boiler. Scientific American, 51(18): 274, November 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n18/pdf/scientificamerican11011884-
274.pdf.

Anonymous:1884:SRH

[5778] Anonymous. Shade roller hanger. Scientific American, 51(18): 274, November 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n18/pdf/scientificamerican11011884-
274a.pdf.

Anonymous:1884:IES

[5779] Anonymous. Improved explosive shell. Scientific American, 51 (18):274, November 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n18/pdf/scientificamerican11011884-
274b.pdf.

Anonymous:1884:SCD

[5780] Anonymous. Street car driver’s change box. Scientific American, 51(18):275, November 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n18/pdf/scientificamerican11011884-
275.pdf.

Anonymous:1884:ICH

[5781] Anonymous. Improved colter hanger. Scientific American, 51 (18):275, November 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n18/pdf/scientificamerican11011884-
275a.pdf.

Anonymous:1884:ICCd

[5782] Anonymous. Improved coal crane. Scientific American, 51(18): 275–276, November 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n18/pdf/scientificamerican11011884-
275b.pdf.

Anonymous:1884:IGS

[5783] Anonymous. An improved gun sight. Scientific American, 51 (18):276, November 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n18/pdf/scientificamerican11011884-
276.pdf.

Anonymous:1884:SCa

[5784] Anonymous. Steam catamaran. Scientific American, 51(18): 278, November 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n18/pdf/scientificamerican11011884-
278.pdf.

Anonymous:1884:TWF

[5785] Anonymous. Twelve-wheel freight locomotive. Scientific American, 51(18):278–279, November 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n18/pdf/scientificamerican11011884-
278a.pdf.

Anonymous:1884:LOA

[5786] Anonymous. The living organisms of the atmosphere. Scientific American, 51(18):279–282, November 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n18/pdf/scientificamerican11011884-
279.pdf.

Anonymous:1884:EIbi

[5787] Anonymous. Engineering inventions. Scientific American, 51(18): 283, November 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n18/pdf/scientificamerican11011884-
283.pdf.

Anonymous:1884:AIbg

[5788] Anonymous. Agricultural inventions. Scientific American, 51 (18):283, November 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n18/pdf/scientificamerican11011884-
283a.pdf.

Anonymous:1884:MIcq

[5789] Anonymous. Miscellaneous inventions. Scientific American, 51 (18):283, November 1, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n18/pdf/scientificamerican11011884-
283b.pdf.

Anonymous:1884:Cbm

[5790] Anonymous. Correspondence. Scientific American, 51(19):293, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n19/pdf/scientificamerican11081884-293.pdf.

Anonymous:1884:BNBbb

[5791] Anonymous. Books: New books and publications. Scientific American, 51(19):298, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n19/pdf/scientificamerican11081884-
298c.pdf.

Anonymous:1884:BPbm

[5792] Anonymous. Business and personal. Scientific American, 51 (19):298, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n19/pdf/scientificamerican11081884-
298d.pdf.

Anonymous:1884:NQbk

[5793] Anonymous. Notes and queries. Scientific American, 51(19): 298–299, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n19/pdf/scientificamerican11081884-
298e.pdf.

Anonymous:1884:IIbk

[5794] Anonymous. Index of inventions. Scientific American, 51(19): 299–300, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n19/pdf/scientificamerican11081884-
299.pdf.

Anonymous:1884:PNPbb

[5795] Anonymous. Patent notes: Patents. Scientific American, 51 (19):301, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n19/pdf/scientificamerican11081884-
301.pdf.

Anonymous:1884:EWE

[5796] Anonymous. The Edward Weston Exhibit at the International Electrical Exposition, Philadelphia. Scientific American, 51(19):287–288, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n19/pdf/scientificamerican11081884-287.pdf.

Anonymous:1884:CSI

[5797] Anonymous. Composite steel and iron. Scientific American, 51 (19):288, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n19/pdf/scientificamerican11081884-
288.pdf.

Anonymous:1884:UGP

[5798] Anonymous. Uses of gas pipe. Scientific American, 51(19):288, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n19/pdf/scientificamerican11081884-288a.pdf.

Anonymous:1884:DBW

[5799] Anonymous. Drilling and boring wood. Scientific American, 51 (19):288, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n19/pdf/scientificamerican11081884-
288b.pdf.

Anonymous:1884:EE

[5800] Anonymous. Employer and employed. Scientific American, 51 (19):289, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n19/pdf/scientificamerican11081884-
289.pdf.

Anonymous:1884:CIC

[5801] Anonymous. Cast iron cutlery. Scientific American, 51(19):289, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n19/pdf/scientificamerican11081884-289a.pdf.

Anonymous:1884:RMb

[5802] Anonymous. Retenoning machine. Scientific American, 51(19): 290, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n19/pdf/scientificamerican11081884-
290.pdf.

Anonymous:1884:LHG

[5803] Anonymous. Loop for hanging garments. Scientific American, 51 (19):290, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n19/pdf/scientificamerican11081884-
290a.pdf.

Anonymous:1884:CLR

[5804] Anonymous. Charger for loading revolvers. Scientific American, 51(19):290, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n19/pdf/scientificamerican11081884-
290b.pdf.

Anonymous:1884:WMD

[5805] Anonymous. Wire measuring device. Scientific American, 51 (19):291, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n19/pdf/scientificamerican11081884-
291.pdf.

Anonymous:1884:RRMb

[5806] Anonymous. Reversing rail mill engines. Scientific American, 51 (19):291–292, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n19/pdf/scientificamerican11081884-
291a.pdf.

Anonymous:1884:CRa

[5807] Anonymous. Clothes reel. Scientific American, 51(19):292, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n19/pdf/scientificamerican11081884-292.pdf.

Anonymous:1884:SHa

[5808] Anonymous. Shade hanger. Scientific American, 51(19):292, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n19/pdf/scientificamerican11081884-292a.pdf.

Anonymous:1884:NVP

[5809] Anonymous. A new vacuum pump, and its application in the manufacture of ice. Scientific American, 51(19):294–295, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n19/pdf/scientificamerican11081884-294.pdf.

Anonymous:1884:LP

[5810] Anonymous. Louis Pasteur. Scientific American, 51(19):295, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n19/pdf/scientificamerican11081884-295.pdf.

Coryell:1884:SD

[5811] John R. Coryell. A sheep destroyer. Scientific American, 51(19): 296–297, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n19/pdf/scientificamerican11081884-
296.pdf.

Anonymous:1884:EIbj

[5812] Anonymous. Engineering inventions. Scientific American, 51(19): 298, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n19/pdf/scientificamerican11081884-
298.pdf.

Anonymous:1884:AIbh

[5813] Anonymous. Agricultural inventions. Scientific American, 51 (19):298, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n19/pdf/scientificamerican11081884-
298a.pdf.

Anonymous:1884:MIcr

[5814] Anonymous. Miscellaneous inventions. Scientific American, 51 (19):298, November 8, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n19/pdf/scientificamerican11081884-
298b.pdf.

Anonymous:1884:Cbn

[5815] Anonymous. Correspondence. Scientific American, 51(2):21–22, July 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n2/pdf/scientificamerican07121884-21.pdf.

Anonymous:1884:BNBbc

[5816] Anonymous. Books: New books and publications. Scientific American, 51(2):26, July 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n2/pdf/scientificamerican07121884-
26d.pdf.

Anonymous:1884:NQbl

[5817] Anonymous. Notes and queries. Scientific American, 51(2):26–27, July 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n2/pdf/scientificamerican07121884-26e.pdf.

Anonymous:1884:IIbl

[5818] Anonymous. Index of inventions. Scientific American, 51(2):27–28, July 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n2/pdf/scientificamerican07121884-27.pdf.

Anonymous:1884:HLB

[5819] Anonymous. Hydraulic lift bridge. Scientific American, 51(2):15–16, July 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n2/pdf/scientificamerican07121884-15.pdf.

Anonymous:1884:CSf

[5820] Anonymous. Contraction of steel. Scientific American, 51(2):16, July 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n2/pdf/scientificamerican07121884-16.pdf.

Anonymous:1884:CLS

[5821] Anonymous. Cutting leading screws. Scientific American, 51(2):16–17, July 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n2/pdf/scientificamerican07121884-16a.pdf.

Anonymous:1884:PLa

[5822] Anonymous. Pneumatic lock. Scientific American, 51(2):18, July 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n2/pdf/scientificamerican07121884-18.pdf.

Anonymous:1884:BSV

[5823] Anonymous. Balanced slide valve. Scientific American, 51(2):18, July 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n2/pdf/scientificamerican07121884-18a.pdf.

Anonymous:1884:ITCa

[5824] Anonymous. Improved thill coupling. Scientific American, 51(2):18, July 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n2/pdf/scientificamerican07121884-18b.pdf.

Anonymous:1884:DTM

[5825] Anonymous. Device for transmitting motion. Scientific American, 51(2):18–19, July 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n2/pdf/scientificamerican07121884-
18c.pdf.

Anonymous:1884:DSB

[5826] Anonymous. Device for securing box covers. Scientific American, 51(2):19, July 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n2/pdf/scientificamerican07121884-
19.pdf.

Anonymous:1884:HHGa

[5827] Anonymous. The Harden hand-grenade fire extinguisher. Scientific American, 51(2):19, July 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n2/pdf/scientificamerican07121884-
19a.pdf.

Anonymous:1884:IJL

[5828] Anonymous. Index for journals, letter files, etc. Scientific American, 51(2):19–20, July 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n2/pdf/scientificamerican07121884-
19b.pdf.

Anonymous:1884:SCb

[5829] Anonymous. Shaft coupling. Scientific American, 51(2):20, July 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n2/pdf/scientificamerican07121884-20.pdf.

Anonymous:1884:DRPg

[5830] Anonymous. Decisions relating to patents trade marks etc. Scientific American, 51(2):20, July 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n2/pdf/scientificamerican07121884-
20a.pdf.

Anonymous:1884:MGI

[5831] Anonymous. Machine for grinding indigo and ultramarine. Scientific American, 51(2):22, July 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n2/pdf/scientificamerican07121884-
22.pdf.

Anonymous:1884:NAC

[5832] Anonymous. New automatic cut-off agricultural engine. Scientific American, 51(2):23, July 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n2/pdf/scientificamerican07121884-
23.pdf.

Brown-Sequard:1884:AAD

[5833] C. B. Brown-Sequard. Attitudes after death. Scientific American, 51(2):23–24, July 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n2/pdf/scientificamerican07121884-
23a.pdf.

Anonymous:1884:TS

[5834] Anonymous. Telephone support. Scientific American, 51(2):24–25, July 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n2/pdf/scientificamerican07121884-24.pdf.

Anonymous:1884:EIbk

[5835] Anonymous. Engineering inventions. Scientific American, 51(2):26, July 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n2/pdf/scientificamerican07121884-26.pdf.

Anonymous:1884:MIcs

[5836] Anonymous. Mechanical inventions. Scientific American, 51(2):26, July 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n2/pdf/scientificamerican07121884-26a.pdf.

Anonymous:1884:AIbi

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n2/pdf/scientificamerican07121884-26b.pdf.

Anonymous:1884:MIct

[5838] Anonymous. Miscellaneous inventions. Scientific American, 51(2):26, July 12, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n2/pdf/scientificamerican07121884-26c.pdf.

Anonymous:1884:Cbo

[5839] Anonymous. Correspondence. Scientific American, 51(20):312– 313, November 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n20/pdf/scientificamerican11151884-
312.pdf.

Anonymous:1884:BPbn

[5840] Anonymous. Business and personal. Scientific American, 51 (20):314, November 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n20/pdf/scientificamerican11151884-
314c.pdf.

Anonymous:1884:NQbm

[5841] Anonymous. Notes and queries. Scientific American, 51(20): 314–315, November 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n20/pdf/scientificamerican11151884-
314d.pdf.

Anonymous:1884:IIbm

[5842] Anonymous. Index of inventions. Scientific American, 51(20): 315–316, November 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n20/pdf/scientificamerican11151884-
315.pdf.

Anonymous:1884:PNPbc

[5843] Anonymous. Patent notes: Patents. Scientific American, 51 (20):317, November 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n20/pdf/scientificamerican11151884-
317.pdf.

Anonymous:1884:ASY

[5844] Anonymous. American steam yachts. Scientific American, 51(20): 303–304, November 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n20/pdf/scientificamerican11151884-
303.pdf.

Anonymous:1884:RSb

[5845] Anonymous. The return screw. Scientific American, 51(20):304, November 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n20/pdf/scientificamerican11151884-304.pdf.

Anonymous:1884:AIAa

[5846] Anonymous. The ancient interior African sea. Scientific American, 51(20):304–305, November 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n20/pdf/scientificamerican11151884-
304a.pdf.

Anonymous:1884:Hb

[5847] Anonymous. Horseradish. Scientific American, 51(20):305, November 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n20/pdf/scientificamerican11151884-305.pdf.

Anonymous:1884:EHF

[5848] Anonymous. An electric hair and flesh brush. Scientific American, 51(20):306, November 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n20/pdf/scientificamerican11151884-
306.pdf.

Anonymous:1884:IMC

[5849] Anonymous. Improved measuring canister. Scientific American, 51 (20):306, November 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n20/pdf/scientificamerican11151884-
306a.pdf.

Anonymous:1884:IVT

[5850] Anonymous. Improved vehicle top. Scientific American, 51(20): 306, November 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n20/pdf/scientificamerican11151884-
306b.pdf.

Anonymous:1884:ICe

[5851] Anonymous. An improved churn. Scientific American, 51(20): 306, November 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n20/pdf/scientificamerican11151884-
306c.pdf.

Anonymous:1884:CPM

[5852] Anonymous. Cane planting machine. Scientific American, 51 (20):307, November 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n20/pdf/scientificamerican11151884-
307.pdf.

Anonymous:1884:IDP

[5853] Anonymous. Imperial Dom Pedro II. Bridge. Scientific American, 51 (20):307–308, November 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n20/pdf/scientificamerican11151884-
307a.pdf.

Anonymous:1884:PBBb

[5854] Anonymous. Packing box for bottles. Scientific American, 51(20): 308–309, November 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n20/pdf/scientificamerican11151884-
308.pdf.

Anonymous:1884:ETP

[5855] Anonymous. An electric torchlight procession. Scientific American, 51(20):310–311, November 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n20/pdf/scientificamerican11151884-
310.pdf.

Anonymous:1884:TNO

[5856] Anonymous. Two new optical illusions. Scientific American, 51 (20):311–312, November 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n20/pdf/scientificamerican11151884-
311.pdf.

Anonymous:1884:EIbl

[5857] Anonymous. Engineering inventions. Scientific American, 51(20): 314, November 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n20/pdf/scientificamerican11151884-
314.pdf.

Anonymous:1884:AIbj

[5858] Anonymous. Agricultural inventions. Scientific American, 51 (20):314, November 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n20/pdf/scientificamerican11151884-
314a.pdf.

Anonymous:1884:MIcu

[5859] Anonymous. Miscellaneous inventions. Scientific American, 51 (20):314, November 15, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n20/pdf/scientificamerican11151884-
314b.pdf.

Anonymous:1884:Cbp

[5860] Anonymous. Correspondence. Scientific American, 51(21):325, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n21/pdf/scientificamerican11221884-325.pdf.

Anonymous:1884:BNBbd

[5861] Anonymous. Books: New books and publications. Scientific American, 51(21):329, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n21/pdf/scientificamerican11221884-
329d.pdf.

Anonymous:1884:BPbo

[5862] Anonymous. Business and personal. Scientific American, 51 (21):329, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n21/pdf/scientificamerican11221884-
329f.pdf.

Anonymous:1884:NQbn

[5863] Anonymous. Notes and queries. Scientific American, 51(21): 329–331, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n21/pdf/scientificamerican11221884-
329g.pdf.

Anonymous:1884:IIbn

[5864] Anonymous. Index of inventions. Scientific American, 51(21): 331, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n21/pdf/scientificamerican11221884-
331.pdf.

Anonymous:1884:PNPbd

[5865] Anonymous. Patent notes: Patents. Scientific American, 51 (21):332, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n21/pdf/scientificamerican11221884-
332.pdf.

Anonymous:1884:CWS

[5866] Anonymous. The Chilian war ship Esmeralda. Scientific American, 51(21):319–320, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n21/pdf/scientificamerican11221884-
319.pdf.

Anonymous:1884:GR

[5867] Anonymous. Grinding reamers. Scientific American, 51(21): 320, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n21/pdf/scientificamerican11221884-
320.pdf.

Anonymous:1884:WDSb

[5868] Anonymous. What the doctors say about bicycle riding. Scientific American, 51(21):320, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v51/n21/pdf/scientificamerican11221884-
320a.pdf.

Anonymous:1884:ESR

[5869] Anonymous. The Eads ship railway. Scientific American, 51(21): 320–321, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n21/pdf/scientificamerican11221884-
320b.pdf.

Anonymous:1884:IBSa

[5870] Anonymous. Improved bench stop. Scientific American, 51(21): 322, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n21/pdf/scientificamerican11221884-
322.pdf.

Anonymous:1884:FTC

[5871] Anonymous. Folding top crib. Scientific American, 51(21):322, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n21/pdf/scientificamerican11221884-322a.pdf.

Anonymous:1884:AL

[5872] Anonymous. Axle lubricator. Scientific American, 51(21):322, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n21/pdf/scientificamerican11221884-322b.pdf.

Anonymous:1884:IRV

[5873] Anonymous. Improved road vehicle. Scientific American, 51 (21):322, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n21/pdf/scientificamerican11221884-
322c.pdf.

Anonymous:1884:IWS

[5874] Anonymous. Improved water softening apparatus. Scientific American, 51(21):323–324, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n21/pdf/scientificamerican11221884-
323.pdf.

Anonymous:1884:CMa

[5875] Anonymous. Capsule machine. Scientific American, 51(21):324, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n21/pdf/scientificamerican11221884-324.pdf.

Anonymous:1884:IAB

[5876] Anonymous. Improved auger bit. Scientific American, 51(21): 324, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n21/pdf/scientificamerican11221884-
324a.pdf.

Anonymous:1884:TVV

[5877] Anonymous. Telescopic views of Venus and Jupiter. Scientific American, 51(21):325, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n21/pdf/scientificamerican11221884-
325a.pdf.

Anonymous:1884:OGE

[5878] Anonymous. Otto gas engines at the Philadelphia Electrical Exhibition. Scientific American, 51(21):326, November 22, 1884. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v51/
n21/pdf/scientificamerican11221884-326.pdf.

Anonymous:1884:NGS

[5879] Anonymous. New grip system for electric railroads. Scientific American, 51(21):326–327, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n21/pdf/scientificamerican11221884-
326a.pdf.

Anonymous:1884:CMb

[5880] Anonymous. Cholera Morbus. Scientific American, 51(21):327– 328, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n21/pdf/scientificamerican11221884-
327.pdf.

Anonymous:1884:HHGb

[5881] Anonymous. The Harden hand grenade fire extinguisher. Scientific American, 51(21):328, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v51/n21/pdf/scientificamerican11221884-
328.pdf.

Anonymous:1884:EIbm

[5882] Anonymous. Engineering inventions. Scientific American, 51(21): 329, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n21/pdf/scientificamerican11221884-
329.pdf.

Anonymous:1884:MIcv

[5883] Anonymous. Mechanical inventions. Scientific American, 51(21): 329, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n21/pdf/scientificamerican11221884-
329a.pdf.

Anonymous:1884:AIbk

[5884] Anonymous. Agricultural inventions. Scientific American, 51 (21):329, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n21/pdf/scientificamerican11221884-
329b.pdf.

Anonymous:1884:MIcw

[5885] Anonymous. Miscellaneous inventions. Scientific American, 51 (21):329, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n21/pdf/scientificamerican11221884-
329c.pdf.

Anonymous:1884:Sk

[5886] Anonymous. Special. Scientific American, 51(21):329, November 22, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n21/pdf/scientificamerican11221884-329e.pdf.

Anonymous:1884:Cbq

[5887] Anonymous. Correspondence. Scientific American, 51(22):341, November 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n22/pdf/scientificamerican11291884-341.pdf.

Anonymous:1884:BNBbe

[5888] Anonymous. Books: New books and publications. Scientific American, 51(22):346, November 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n22/pdf/scientificamerican11291884-
346c.pdf.

Anonymous:1884:BPbp

[5889] Anonymous. Business and personal. Scientific American, 51 (22):346, November 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n22/pdf/scientificamerican11291884-
346d.pdf.

Anonymous:1884:NQbo

[5890] Anonymous. Notes and queries. Scientific American, 51(22): 346–347, November 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n22/pdf/scientificamerican11291884-
346e.pdf.

Anonymous:1884:IIbo

[5891] Anonymous. Index of inventions. Scientific American, 51(22): 347–348, November 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n22/pdf/scientificamerican11291884-
347.pdf.

Anonymous:1884:PNPbe

[5892] Anonymous. Patent notes: Patents. Scientific American, 51 (22):349, November 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n22/pdf/scientificamerican11291884-
349.pdf.

Anonymous:1884:PPc

[5893] Anonymous. Property in patents. Scientific American, 51(22): 336, November 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n22/pdf/scientificamerican11291884-
336.pdf.

Anonymous:1884:WCc

[5894] Anonymous. Wire cutting. Scientific American, 51(22):336, November 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n22/pdf/scientificamerican11291884-336a.pdf.

Anonymous:1884:APD

[5895] Anonymous. Aspects of the planets for December. Scientific American, 51(22):337, November 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n22/pdf/scientificamerican11291884-
337.pdf.

Anonymous:1884:NSV

[5896] Anonymous. New system of ventilating vessels. Scientific American, 51(22):338, November 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n22/pdf/scientificamerican11291884-
338.pdf.

Anonymous:1884:HPS

[5897] Anonymous. Hand power for sewing machines. Scientific American, 51(22):338, November 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n22/pdf/scientificamerican11291884-
338a.pdf.

Anonymous:1884:CDV

[5898] Anonymous. Cellar drain and ventilator. Scientific American, 51 (22):338, November 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n22/pdf/scientificamerican11291884-
338b.pdf.

Anonymous:1884:CTH

[5899] Anonymous. Combination tool holder. Scientific American, 51 (22):338, November 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n22/pdf/scientificamerican11291884-
338c.pdf.

Anonymous:1884:JCS

[5900] Anonymous. Jacquelin and Chevre’s steam excavator. Scientific American, 51(22):339–340, November 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v51/n22/pdf/scientificamerican11291884-
339.pdf.

Anonymous:1884:CSL

[5901] Anonymous. The “Champion” six-lever rim night latch. Scientific American, 51(22):340, November 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v51/n22/pdf/scientificamerican11291884-
340.pdf.

Anonymous:1884:ISW

[5902] Anonymous. Improved steam wheel. Scientific American, 51 (22):341, November 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n22/pdf/scientificamerican11291884-
341a.pdf.

Tatin:1884:AN

[5903] Victor Tatin. Aerial navigation. Scientific American, 51(22): 342–343, November 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n22/pdf/scientificamerican11291884-
342.pdf.

Anonymous:1884:D

[5904] Anonymous. Dinosaurs. Scientific American, 51(22):343–344, November 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n22/pdf/scientificamerican11291884-343.pdf.

Anonymous:1884:ISO

[5905] Anonymous. Improved stove and other pipes. Scientific American, 51 (22):344–345, November 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n22/pdf/scientificamerican11291884-
344.pdf.

Anonymous:1884:DRPh

[5906] Anonymous. Decisions relating to patents. Scientific American, 51(22):345, November 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n22/pdf/scientificamerican11291884-
345.pdf.

Anonymous:1884:EIbn

[5907] Anonymous. Engineering inventions. Scientific American, 51(22): 346, November 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n22/pdf/scientificamerican11291884-
346.pdf.

Anonymous:1884:AIbl

[5908] Anonymous. Agricultural inventions. Scientific American, 51 (22):346, November 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n22/pdf/scientificamerican11291884-
346a.pdf.

Anonymous:1884:MIcx

[5909] Anonymous. Miscellaneous inventions. Scientific American, 51 (22):346, November 29, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n22/pdf/scientificamerican11291884-
346b.pdf.

Anonymous:1884:Cbr

[5910] Anonymous. Correspondence. Scientific American, 51(23):357, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n23/pdf/scientificamerican12061884-357.pdf.

Anonymous:1884:BNBbf

[5911] Anonymous. Books: New books and publications. Scientific American, 51(23):362, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n23/pdf/scientificamerican12061884-
362c.pdf.

Anonymous:1884:BPbq

[5912] Anonymous. Business and personal. Scientific American, 51 (23):362, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n23/pdf/scientificamerican12061884-
362e.pdf.

Anonymous:1884:NQbp

[5913] Anonymous. Notes and queries. Scientific American, 51(23):387, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n23/pdf/scientificamerican12061884-387a.pdf.

Anonymous:1884:IIbp

[5914] Anonymous. Index of inventions. Scientific American, 51(23): 387–388, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n23/pdf/scientificamerican12061884-
387b.pdf.

Anonymous:1884:PNPbf

[5915] Anonymous. Patent notes: Patents. Scientific American, 51 (23):389, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n23/pdf/scientificamerican12061884-
389.pdf.

Anonymous:1884:IVP

[5916] Anonymous. An immense vacuum pan for sugar making. Scientific American, 51(23):351, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v51/n23/pdf/scientificamerican12061884-
351.pdf.

Anonymous:1884:SOR

[5917] Anonymous. Specially for our readers. Scientific American, 51 (23):352, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n23/pdf/scientificamerican12061884-
352.pdf.

Anonymous:1884:EFV

[5918] Anonymous. The efficiency of fluid in vapor engines. Scientific American, 51(23):352, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n23/pdf/scientificamerican12061884-
352a.pdf.

Anonymous:1884:SU

[5919] Anonymous. Steel uniformity. Scientific American, 51(23):352, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n23/pdf/scientificamerican12061884-352b.pdf.

Anonymous:1884:MED

[5920] Anonymous. Methods of estimating distances. Scientific American, 51(23):353–354, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n23/pdf/scientificamerican12061884-
353.pdf.

Anonymous:1884:CBs

[5921] Anonymous. Claw bar. Scientific American, 51(23):354, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n23/pdf/scientificamerican12061884-354.pdf.

Anonymous:1884:FMR

[5922] Anonymous. Feed mechanism for roller mills. Scientific American, 51(23):354, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n23/pdf/scientificamerican12061884-
354a.pdf.

Anonymous:1884:RSM

[5923] Anonymous. Rope serving machine. Scientific American, 51 (23):354, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n23/pdf/scientificamerican12061884-
354b.pdf.

Anonymous:1884:IPd

[5924] Anonymous. An improved plow. Scientific American, 51(23):354, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n23/pdf/scientificamerican12061884-354c.pdf.

Anonymous:1884:IUP

[5925] Anonymous. Improvements in upright pianos. Scientific American, 51(23):355–356, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n23/pdf/scientificamerican12061884-
355.pdf.

Anonymous:1884:NFEc

[5926] Anonymous. A novel fire escape. Scientific American, 51(23):356, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n23/pdf/scientificamerican12061884-356.pdf.

Anonymous:1884:NET

[5927] Anonymous. A new elevated tramway without rails. Scientific American, 51(23):358, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n23/pdf/scientificamerican12061884-
358.pdf.

Anonymous:1884:CBt

[5928] Anonymous. The carbon blast. Scientific American, 51(23):359, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n23/pdf/scientificamerican12061884-359.pdf.

Anonymous:1884:CFa

[5929] Anonymous. The Colladon Fountain. Scientific American, 51(23): 359–360, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n23/pdf/scientificamerican12061884-
359a.pdf.

Anonymous:1884:DOG

[5930] Anonymous. Dr. Otto Gmelin’s cupola. Scientific American, 51 (23):360–361, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n23/pdf/scientificamerican12061884-
360.pdf.

Anonymous:1884:EIbo

[5931] Anonymous. Engineering inventions. Scientific American, 51(23): 362, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n23/pdf/scientificamerican12061884-
362.pdf.

Anonymous:1884:AIbm

[5932] Anonymous. Agricultural inventions. Scientific American, 51 (23):362, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n23/pdf/scientificamerican12061884-
362a.pdf.

Anonymous:1884:MIcy

[5933] Anonymous. Miscellaneous inventions. Scientific American, 51 (23):362, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n23/pdf/scientificamerican12061884-
362b.pdf.

Anonymous:1884:Sl

[5934] Anonymous. Special. Scientific American, 51(23):362, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n23/pdf/scientificamerican12061884-362d.pdf.

Anonymous:1884:SAC

[5935] Anonymous. Scientific American catalogue. Scientific American, 51 (23):363–386, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n23/pdf/scientificamerican12061884-
363.pdf.

Anonymous:1884:SASb

[5936] Anonymous. The Scientific American supplement. Scientific American, 51(23):387, December 6, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n23/pdf/scientificamerican12061884-
387.pdf.

Anonymous:1884:BNBbg

[5937] Anonymous. Books: New books and publications. Scientific American, 51(24):402, December 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n24/pdf/scientificamerican12131884-
402c.pdf.

Anonymous:1884:BPbr

[5938] Anonymous. Business and personal. Scientific American, 51 (24):402, December 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n24/pdf/scientificamerican12131884-
402d.pdf.

Anonymous:1884:NQbq

[5939] Anonymous. Notes and queries. Scientific American, 51(24): 402–404, December 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n24/pdf/scientificamerican12131884-
402e.pdf.

Anonymous:1884:IIbq

[5940] Anonymous. Index of inventions. Scientific American, 51(24): 404, December 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n24/pdf/scientificamerican12131884-
404.pdf.

Anonymous:1884:PNPbg

[5941] Anonymous. Patent notes: Patents. Scientific American, 51 (24):405, December 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n24/pdf/scientificamerican12131884-
405.pdf.

Anonymous:1884:MMG

[5942] Anonymous. The Maxim machine gun. Scientific American, 51 (24):391–398, December 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n24/pdf/scientificamerican12131884-
391.pdf.

Anonymous:1884:BMa

[5943] Anonymous. Belt material. Scientific American, 51(24):392, December 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n24/pdf/scientificamerican12131884-392.pdf.

Anonymous:1884:GL

[5944] Anonymous. Gold leaf. Scientific American, 51(24):392, December 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n24/pdf/scientificamerican12131884-392a.pdf.

Anonymous:1884:CIb

[5945] Anonymous. Criminal ignorance. Scientific American, 51(24): 392, December 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n24/pdf/scientificamerican12131884-
392b.pdf.

Anonymous:1884:PGa

[5946] Anonymous. Pulleys and gears. Scientific American, 51(24): 392–393, December 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n24/pdf/scientificamerican12131884-
392c.pdf.

Anonymous:1884:BMb

[5947] Anonymous. A basic metal. Scientific American, 51(24):393, December 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n24/pdf/scientificamerican12131884-393.pdf.

Anonymous:1884:ESC

[5948] Anonymous. Experiments on superheating as a cause of boiler explosion. Scientific American, 51(24):394–397, December 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v51/
n24/pdf/scientificamerican12131884-394.pdf.

Anonymous:1884:PTL

[5949] Anonymous. The pneumatic telegraph line at Paris. Scientific American, 51(24):395, December 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n24/pdf/scientificamerican12131884-
395.pdf.

Anonymous:1884:IRCb

[5950] Anonymous. Improved refrigerating cars. Scientific American, 51 (24):395–396, December 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n24/pdf/scientificamerican12131884-
395a.pdf.

Anonymous:1884:PDa

[5951] Anonymous. Patent drawers. Scientific American, 51(24):396, December 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n24/pdf/scientificamerican12131884-396.pdf.

Anonymous:1884:CMS

[5952] Anonymous. Composite mica sheet. Scientific American, 51 (24):396, December 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n24/pdf/scientificamerican12131884-
396a.pdf.

Anonymous:1884:SSFa

[5953] Anonymous. Some specimens of fine cattle. Scientific American, 51 (24):399–401, December 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n24/pdf/scientificamerican12131884-
399.pdf.

Anonymous:1884:EIbp

[5954] Anonymous. Engineering inventions. Scientific American, 51(24): 402, December 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n24/pdf/scientificamerican12131884-
402.pdf.

Anonymous:1884:AIbn

[5955] Anonymous. Agricultural inventions. Scientific American, 51 (24):402, December 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n24/pdf/scientificamerican12131884-
402a.pdf.

Anonymous:1884:MIcz

[5956] Anonymous. Miscellaneous inventions. Scientific American, 51 (24):402, December 13, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n24/pdf/scientificamerican12131884-
402b.pdf.

Anonymous:1884:Cbu

[5957] Anonymous. Correspondence. Scientific American, 51(25):413, December 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n25/pdf/scientificamerican12201884-413.pdf.

Anonymous:1884:BPbs

[5958] Anonymous. Business and personal. Scientific American, 51(25): 418–419, December 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n25/pdf/scientificamerican12201884-
418d.pdf.

Anonymous:1884:NQbr

[5959] Anonymous. Notes and queries. Scientific American, 51(25): 419, December 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n25/pdf/scientificamerican12201884-
419.pdf.

Anonymous:1884:IIbr

[5960] Anonymous. Index of inventions. Scientific American, 51(25): 419–420, December 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n25/pdf/scientificamerican12201884-
419a.pdf.

Anonymous:1884:PNPbh

[5961] Anonymous. Patent notes: Patents. Scientific American, 51 (25):420, December 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n25/pdf/scientificamerican12201884-
420.pdf.

Anonymous:1884:DBD

[5962] Anonymous. The decision in the Bell–Drawbaugh Telephone Suit. Scientific American, 51(25):408, December 20, 1884. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v51/
n25/pdf/scientificamerican12201884-408.pdf.

Anonymous:1884:HIb

[5963] Anonymous. Hints to inventors. Scientific American, 51(25): 408–409, December 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n25/pdf/scientificamerican12201884-
408a.pdf.

Anonymous:1884:DBPb

[5964] Anonymous. The delay of business in the Patent Office. Scientific American, 51(25):409, December 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n25/pdf/scientificamerican12201884-
409.pdf.

Anonymous:1884:AGC

[5965] Anonymous. Automatic game counter. Scientific American, 51 (25):410, December 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n25/pdf/scientificamerican12201884-
410.pdf.

Anonymous:1884:NTb

[5966] Anonymous. A novel table. Scientific American, 51(25):410, December 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n25/pdf/scientificamerican12201884-410a.pdf.

Anonymous:1884:AAB

[5967] Anonymous. Automatic alarm for bearings. Scientific American, 51(25):410, December 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n25/pdf/scientificamerican12201884-
410b.pdf.

Anonymous:1884:CPc

[5968] Anonymous. Casing for pipes. Scientific American, 51(25):410– 411, December 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n25/pdf/scientificamerican12201884-
410c.pdf.

Anonymous:1884:DSH

[5969] Anonymous. Double steam hammer. Scientific American, 51 (25):411, December 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n25/pdf/scientificamerican12201884-
411.pdf.

Anonymous:1884:FRD

[5970] Anonymous. Frankel’s registering dynamometer. Scientific American, 51(25):411–412, December 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n25/pdf/scientificamerican12201884-
411a.pdf.

Anonymous:1884:NOEb

[5971] Anonymous. The New Orleans Exposition. Scientific American, 51(25):412, December 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n25/pdf/scientificamerican12201884-
412.pdf.

Anonymous:1884:NGE

[5972] Anonymous. A new gas engine. Scientific American, 51(25):414, December 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n25/pdf/scientificamerican12201884-414.pdf.

Anonymous:1884:ISPb

[5973] Anonymous. Improved silo press. Scientific American, 51(25): 414, December 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n25/pdf/scientificamerican12201884-
414a.pdf.

Anonymous:1884:CCBa

[5974] Anonymous. The Camel Corps of the British Army. Scientific American, 51(25):415–416, December 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v51/n25/pdf/scientificamerican12201884-
415.pdf.

Anonymous:1884:RDR

[5975] Anonymous. Recent decisions relating to patents. Scientific American, 51(25):417, December 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n25/pdf/scientificamerican12201884-
417.pdf.

Anonymous:1884:ISI

[5976] Anonymous. An improved speed indicator. Scientific American, 51(25):417, December 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n25/pdf/scientificamerican12201884-
417a.pdf.

Anonymous:1884:EIbq

[5977] Anonymous. Engineering inventions. Scientific American, 51(25): 418, December 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n25/pdf/scientificamerican12201884-
418.pdf.

Anonymous:1884:AIbo

[5978] Anonymous. Agricultural inventions. Scientific American, 51 (25):418, December 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n25/pdf/scientificamerican12201884-
418a.pdf.

Anonymous:1884:MIda

[5979] Anonymous. Miscellaneous inventions. Scientific American, 51 (25):418, December 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n25/pdf/scientificamerican12201884-
418b.pdf.

Anonymous:1884:Sn

[5980] Anonymous. Special. Scientific American, 51(25):418, December 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n25/pdf/scientificamerican12201884-418c.pdf.

Anonymous:1884:FPd

[5981] Anonymous. Foreign patents. Scientific American, 51(25):421, December 20, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n25/pdf/scientificamerican12201884-421.pdf.

Anonymous:1884:BPbt

[5982] Anonymous. Business and personal. Scientific American, 51 (26):432, December 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n26/pdf/scientificamerican12271884-
432c.pdf.

Anonymous:1884:NQbs

[5983] Anonymous. Notes and queries. Scientific American, 51(26): 432–433, December 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n26/pdf/scientificamerican12271884-
432d.pdf.

Anonymous:1884:IIbs

[5984] Anonymous. Index of inventions. Scientific American, 51(26): 433–434, December 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n26/pdf/scientificamerican12271884-
433.pdf.

Anonymous:1884:PNPbi

[5985] Anonymous. Patent notes: Patents. Scientific American, 51 (26):434, December 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n26/pdf/scientificamerican12271884-
434.pdf.

Anonymous:1884:HBa

[5986] Anonymous. How is business? Scientific American, 51(26):424, December 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n26/pdf/scientificamerican12271884-424.pdf.

Anonymous:1884:IWb

[5987] Anonymous. Improved workmanship. Scientific American, 51 (26):424, December 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n26/pdf/scientificamerican12271884-
424a.pdf.

Anonymous:1884:APJc

[5988] Anonymous. Aspects of the planets for January. Scientific American, 51(26):425, December 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n26/pdf/scientificamerican12271884-
425.pdf.

Anonymous:1884:NSR

[5989] Anonymous. A new safety regulator for elevators. Scientific American, 51(26):426, December 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n26/pdf/scientificamerican12271884-
426.pdf.

Anonymous:1884:NMP

[5990] Anonymous. Novel method of propelling vessels. Scientific American, 51(26):426, December 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n26/pdf/scientificamerican12271884-
426a.pdf.

Anonymous:1884:CHc

[5991] Anonymous. A crow hut. Scientific American, 51(26):426, December 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n26/pdf/scientificamerican12271884-426b.pdf.

Anonymous:1884:WR

[5992] Anonymous. Watch regulator. Scientific American, 51(26):426, December 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n26/pdf/scientificamerican12271884-426c.pdf.

Anonymous:1884:BDV

[5993] Anonymous. Balancing device for vessels. Scientific American, 51 (26):426, December 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n26/pdf/scientificamerican12271884-
426d.pdf.

Anonymous:1884:ISR

[5994] Anonymous. The Interoceanic Ship Railway. Scientific American, 51 (26):428–431, December 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n26/pdf/scientificamerican12271884-
428.pdf.

Anonymous:1884:EIbr

[5995] Anonymous. Engineering inventions. Scientific American, 51(26): 432, December 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n26/pdf/scientificamerican12271884-
432.pdf.

Anonymous:1884:MIdb

[5996] Anonymous. Mechanical inventions. Scientific American, 51(26): 432, December 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n26/pdf/scientificamerican12271884-
432a.pdf.

Anonymous:1884:MIdc

[5997] Anonymous. Miscellaneous inventions. Scientific American, 51 (26):432, December 27, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n26/pdf/scientificamerican12271884-
432b.pdf.

Anonymous:1884:BPbu

[5998] Anonymous. Business and personal. Scientific American, 51(3):41, July 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n3/pdf/scientificamerican07191884-41d.pdf.

Anonymous:1884:BNBbh

[5999] Anonymous. Books: New books and publications. Scientific American, 51(3):41, July 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n3/pdf/scientificamerican07191884-
41e.pdf.

Anonymous:1884:NQbt

[6000] Anonymous. Notes and queries. Scientific American, 51(3):41–43, July 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n3/pdf/scientificamerican07191884-41f.pdf.

Anonymous:1884:IIbt

[6001] Anonymous. Index of inventions. Scientific American, 51(3):43–44, July 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n3/pdf/scientificamerican07191884-43.pdf.

Anonymous:1884:IUD

[6002] Anonymous. Improved upright drying machine. Scientific American, 51(3):31, July 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n3/pdf/scientificamerican07191884-
31.pdf.

Anonymous:1884:SSFb

[6003] Anonymous. Scraping surfaces to fit. Scientific American, 51(3):32, July 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n3/pdf/scientificamerican07191884-32.pdf.

Anonymous:1884:HCC

[6004] Anonymous. Handles for cold chisels. Scientific American, 51(3):32–33, July 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n3/pdf/scientificamerican07191884-32a.pdf.

Anonymous:1884:WPC

[6005] Anonymous. Well pipe cutter. Scientific American, 51(3):34, July 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n3/pdf/scientificamerican07191884-34.pdf.

Anonymous:1884:SCg

[6006] Anonymous. Sail clamp. Scientific American, 51(3):34, July 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n3/pdf/scientificamerican07191884-34a.pdf.

Anonymous:1884:WEG

[6007] Anonymous. Wagon end gate. Scientific American, 51(3):34, July 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n3/pdf/scientificamerican07191884-34b.pdf.

Anonymous:1884:DAPb

[6008] Anonymous. Double acting pump. Scientific American, 51(3):34, July 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n3/pdf/scientificamerican07191884-34c.pdf.

Anonymous:1884:PCD

[6009] Anonymous. Phantom curve. — Denver and Rio Grande Railway. Scientific American, 51(3):35, July 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n3/pdf/scientificamerican07191884-
35.pdf.

Anonymous:1884:EBEb

[6010] Anonymous. An extraordinary boiler explosion at Orléans, France. Scientific American, 51(3):35–36, July 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v51/n3/pdf/scientificamerican07191884-
35a.pdf.

Anonymous:1884:SCEb

[6011] Anonymous. Safety catch for elevators. Scientific American, 51 (3):36–37, July 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n3/pdf/scientificamerican07191884-
36.pdf.

Anonymous:1884:DRPi

[6012] Anonymous. Decisions relating to patents. Scientific American, 51(3):37, July 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n3/pdf/scientificamerican07191884-
37.pdf.

Anonymous:1884:WSb

[6013] Anonymous. Wireworms and skipjacks. Scientific American, 51 (3):37–38, July 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n3/pdf/scientificamerican07191884-
37a.pdf.

Anonymous:1884:SCEc

[6014] Anonymous. Shanks’ compound engine for small vessels. Scientific American, 51(3):38–39, July 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n3/pdf/scientificamerican07191884-
38.pdf.

Anonymous:1884:EAb

[6015] Anonymous. Electricity without apparatus. Scientific American, 51(3):39, July 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n3/pdf/scientificamerican07191884-
39.pdf.

Anonymous:1884:TSC

[6016] Anonymous. The tarantula of Southern California. Scientific American, 51(3):39–40, July 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n3/pdf/scientificamerican07191884-
39a.pdf.

Anonymous:1884:ITCb

[6017] Anonymous. Improved thill coupling. Scientific American, 51(3):40, July 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n3/pdf/scientificamerican07191884-40.pdf.

Anonymous:1884:EIbs

[6018] Anonymous. Engineering inventions. Scientific American, 51(3):41, July 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n3/pdf/scientificamerican07191884-41.pdf.

Anonymous:1884:MIdd

[6019] Anonymous. Mechanical inventions. Scientific American, 51(3):41, July 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n3/pdf/scientificamerican07191884-41a.pdf.

Anonymous:1884:AIbp

[6020] Anonymous. Agricultural inventions. Scientific American, 51(3):41, July 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n3/pdf/scientificamerican07191884-41b.pdf.

Anonymous:1884:MIde

[6021] Anonymous. Miscellaneous inventions. Scientific American, 51(3):41, July 19, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n3/pdf/scientificamerican07191884-41c.pdf.

Anonymous:1884:Cbv

[6022] Anonymous. Correspondence. Scientific American, 51(4):53, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n4/pdf/scientificamerican07261884-53.pdf.

Anonymous:1884:BPbv

[6023] Anonymous. Business and personal. Scientific American, 51(4):57–58, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n4/pdf/scientificamerican07261884-57d.pdf.

Anonymous:1884:BNBbi

[6024] Anonymous. Books: New books and publications. Scientific American, 51(4):58, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n4/pdf/scientificamerican07261884-
58.pdf.

Anonymous:1884:NQbu

[6025] Anonymous. Notes and queries. Scientific American, 51(4):58–59, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n4/pdf/scientificamerican07261884-58a.pdf.

Anonymous:1884:IIbu

[6026] Anonymous. Index of inventions. Scientific American, 51(4):59–60, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n4/pdf/scientificamerican07261884-59.pdf.

Anonymous:1884:PNPbj

[6027] Anonymous. Patent notes: Patents. Scientific American, 51(4):61, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n4/pdf/scientificamerican07261884-61.pdf.

Anonymous:1884:CHCa

[6028] Anonymous. Casting a huge cannon. Scientific American, 51(4):47–48, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n4/pdf/scientificamerican07261884-47.pdf.

Anonymous:1884:CPO

[6029] Anonymous. Congress and the Patent Office. Scientific American, 51(4):48, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n4/pdf/scientificamerican07261884-
48.pdf.

Anonymous:1884:DF

[6030] Anonymous. Driving by friction. Scientific American, 51(4):48–49, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n4/pdf/scientificamerican07261884-48a.pdf.

Anonymous:1884:VS

[6031] Anonymous. Variations of speed. Scientific American, 51(4):49, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n4/pdf/scientificamerican07261884-49.pdf.

Anonymous:1884:GE

[6032] Anonymous. The Greely Expedition. Scientific American, 51(4):49, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n4/pdf/scientificamerican07261884-49a.pdf.

Anonymous:1884:LT

[6033] Anonymous. Labels and trademarks. Scientific American, 51(4):49, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n4/pdf/scientificamerican07261884-49b.pdf.

Anonymous:1884:HP

[6034] Anonymous. Horse power. Scientific American, 51(4):50, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n4/pdf/scientificamerican07261884-50.pdf.

Anonymous:1884:IOC

[6035] Anonymous. Improved oil cup. Scientific American, 51(4):50, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n4/pdf/scientificamerican07261884-50a.pdf.

Anonymous:1884:CSAa

[6036] Anonymous. Compound and self-acting plug valve for wash basins. Scientific American, 51(4):50, July 26, 1884. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v51/
n4/pdf/scientificamerican07261884-50b.pdf.

Anonymous:1884:MFc

[6037] Anonymous. Metallurgic furnace. Scientific American, 51(4):51, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n4/pdf/scientificamerican07261884-51.pdf.

Anonymous:1884:IE

[6038] Anonymous. Improved elevator. Scientific American, 51(4):51, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n4/pdf/scientificamerican07261884-51a.pdf.

Anonymous:1884:BMc

[6039] Anonymous. Bending machine. Scientific American, 51(4):51–52, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n4/pdf/scientificamerican07261884-51b.pdf.

Anonymous:1884:STc

[6040] Anonymous. Screw tap. Scientific American, 51(4):52, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n4/pdf/scientificamerican07261884-52.pdf.

Anonymous:1884:WBb

[6041] Anonymous. Well bucket. Scientific American, 51(4):52, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n4/pdf/scientificamerican07261884-52a.pdf.

Anonymous:1884:CPT

[6042] Anonymous. Conical press for treating beet pulp. Scientific American, 51(4):54, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n4/pdf/scientificamerican07261884-
54.pdf.

Anonymous:1884:PME

[6043] Anonymous. A practical method of estimating distances. Scientific American, 51(4):55–56, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n4/pdf/scientificamerican07261884-
55.pdf.

Anonymous:1884:AWL

[6044] Anonymous. Automatic wave lubricating life buoy. Scientific American, 51(4):56, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n4/pdf/scientificamerican07261884-
56.pdf.

Anonymous:1884:EIbt

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n4/pdf/scientificamerican07261884-57.pdf.

Anonymous:1884:MIdf

[6046] Anonymous. Mechanical inventions. Scientific American, 51(4):57, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n4/pdf/scientificamerican07261884-57a.pdf.

Anonymous:1884:AIbq

[6047] Anonymous. Agricultural inventions. Scientific American, 51(4):57, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n4/pdf/scientificamerican07261884-57b.pdf.

Anonymous:1884:MIdg

[6048] Anonymous. Miscellaneous inventions. Scientific American, 51(4):57, July 26, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n4/pdf/scientificamerican07261884-57c.pdf.

Anonymous:1884:Cbw

[6049] Anonymous. Correspondence. Scientific American, 51(5):69, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-69.pdf.

Anonymous:1884:BNBbj

[6050] Anonymous. Books: New books and publications. Scientific American, 51(5):73, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n5/pdf/scientificamerican08021884-
73d.pdf.

Anonymous:1884:BPbw

[6051] Anonymous. Business and personal. Scientific American, 51(5):73, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-73e.pdf.

Anonymous:1884:NQbv

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-73f.pdf.

Anonymous:1884:IIbv

[6053] Anonymous. Index of inventions. Scientific American, 51(5):75–76, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-75.pdf.

Anonymous:1884:PNPbk

[6054] Anonymous. Patent notes: Patents. Scientific American, 51(5):77, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-77.pdf.

Anonymous:1884:NSCc

[6055] Anonymous. A new steam catamaran. Scientific American, 51(5):63, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-63.pdf.

Anonymous:1884:FPe

[6056] Anonymous. Finishing by pressure. Scientific American, 51(5):64, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-64.pdf.

Anonymous:1884:BPPa

[6057] Anonymous. The Barff process for preserving iron. Scientific American, 51(5):64, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n5/pdf/scientificamerican08021884-
64a.pdf.

Anonymous:1884:PMT

[6058] Anonymous. Portable machine tools. Scientific American, 51(5):64, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-64b.pdf.

Anonymous:1884:PCc

[6059] Anonymous. Poisonous coffee. Scientific American, 51(5):64, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-64c.pdf.

Anonymous:1884:PCf

[6060] Anonymous. Poisonous coffee. Scientific American, 51(5):65, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-65.pdf.

Anonymous:1884:WSc

[6061] Anonymous. Weather strip. Scientific American, 51(5):66, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-66.pdf.

Anonymous:1884:RBS

[6062] Anonymous. Revolving book stand. Scientific American, 51(5):66, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-66a.pdf.

Anonymous:1884:HCb

[6063] Anonymous. Hay carrier. Scientific American, 51(5):66, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-66b.pdf.

Anonymous:1884:SHP

[6064] Anonymous. Side hill plow. Scientific American, 51(5):66, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-66c.pdf.

Anonymous:1884:SCFb

[6065] Anonymous. Shade and curtain fixture. Scientific American, 51(5):67, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-67.pdf.

Anonymous:1884:CCG

[6066] Anonymous. Combination cellar grate. Scientific American, 51(5):67, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-67a.pdf.

Anonymous:1884:FFF

[6067] Anonymous. Folding fireplace fender. Scientific American, 51(5):67–68, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-67b.pdf.

Anonymous:1884:IST

[6068] Anonymous. Insertible saw tooth. Scientific American, 51(5):68, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-68.pdf.

Anonymous:1884:IO

[6069] Anonymous. Improved oar. Scientific American, 51(5):68, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-68a.pdf.

Anonymous:1884:DRPj

[6070] Anonymous. Decisions relating to patents. Scientific American, 51(5):69–70, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n5/pdf/scientificamerican08021884-
69a.pdf.

Anonymous:1884:RCM

[6071] Anonymous. Rock cutting machine. Scientific American, 51(5):70, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-70.pdf.

Anonymous:1884:DSc

[6072] Anonymous. Drawbridge signal. Scientific American, 51(5):70, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-70a.pdf.

Anonymous:1884:SLA

[6073] Anonymous. The sun lamp applied to photography. Scientific American, 51(5):71–72, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n5/pdf/scientificamerican08021884-
71.pdf.

Anonymous:1884:EIbu

[6074] Anonymous. Engineering inventions. Scientific American, 51(5):73, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-73.pdf.

Anonymous:1884:MIdh

[6075] Anonymous. Mechanical inventions. Scientific American, 51(5):73, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-73a.pdf.

Anonymous:1884:AIbr

[6076] Anonymous. Agricultural inventions. Scientific American, 51(5):73, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-73b.pdf.

Anonymous:1884:MIdi

[6077] Anonymous. Miscellaneous inventions. Scientific American, 51(5):73, August 2, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n5/pdf/scientificamerican08021884-73c.pdf.

Anonymous:1884:Cbx

[6078] Anonymous. Correspondence. Scientific American, 51(6):90, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n6/pdf/scientificamerican08091884-90.pdf.

Anonymous:1884:BNBbk

[6079] Anonymous. Books: New books and publications. Scientific American, 51(6):91, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n6/pdf/scientificamerican08091884-
91d.pdf.

Anonymous:1884:BPbx

[6080] Anonymous. Business and personal. Scientific American, 51(6):91, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n6/pdf/scientificamerican08091884-91e.pdf.

Anonymous:1884:NQbw

[6081] Anonymous. Notes and queries. Scientific American, 51(6):91, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n6/pdf/scientificamerican08091884-91f.pdf.

Anonymous:1884:IIbw

[6082] Anonymous. Index of inventions. Scientific American, 51(6):92, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n6/pdf/scientificamerican08091884-92.pdf.

Anonymous:1884:PNPbl

[6083] Anonymous. Patent notes: Patents. Scientific American, 51(6):93, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n6/pdf/scientificamerican08091884-93.pdf.

Anonymous:1884:LSC

[6084] Anonymous. Labels as subjects of copyright. Scientific American, 51(6):80, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n6/pdf/scientificamerican08091884-
80.pdf.

Anonymous:1884:NTc

[6085] Anonymous. Now is the time. Scientific American, 51(6):80, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n6/pdf/scientificamerican08091884-80a.pdf.

Anonymous:1884:ASM

[6086] Anonymous. Anomalies of the sewing-machine business. Scientific American, 51(6):80–81, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n6/pdf/scientificamerican08091884-
80b.pdf.

Anonymous:1884:LF

[6087] Anonymous. Lathe feeds. Scientific American, 51(6):81, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n6/pdf/scientificamerican08091884-81.pdf.

Anonymous:1884:MCAa

[6088] Anonymous. Muscular contraction after death. Scientific American, 51(6):81, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n6/pdf/scientificamerican08091884-
81a.pdf.

Anonymous:1884:FCC

[6089] Anonymous. Forms of cold chisels. Scientific American, 51(6):81–82, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n6/pdf/scientificamerican08091884-81b.pdf.

Anonymous:1884:CRC

[6090] Anonymous. Check row corn planter. Scientific American, 51(6):82, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n6/pdf/scientificamerican08091884-82.pdf.

Anonymous:1884:STV

[6091] Anonymous. Slide trombone valve for cornets, etc. Scientific American, 51(6):82, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n6/pdf/scientificamerican08091884-
82a.pdf.

Anonymous:1884:DSd

[6092] Anonymous. Drag saw. Scientific American, 51(6):82–83, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n6/pdf/scientificamerican08091884-82b.pdf.

Anonymous:1884:CCT

[6093] Anonymous. Combined curbstone and telegraph wire conduit. Scientific American, 51(6):83, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n6/pdf/scientificamerican08091884-
83.pdf.

Anonymous:1884:ML

[6094] Anonymous. Manure loader. Scientific American, 51(6):83–84, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n6/pdf/scientificamerican08091884-83a.pdf.

Anonymous:1884:IFEb

[6095] Anonymous. Improved fire escape. Scientific American, 51(6):84, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n6/pdf/scientificamerican08091884-84.pdf.

Anonymous:1884:TF

[6096] Anonymous. Trace fastener. Scientific American, 51(6):84, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n6/pdf/scientificamerican08091884-84a.pdf.

Anonymous:1884:NYF

[6097] Anonymous. The New York Fire Department. Scientific American, 51(6):85–86, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n6/pdf/scientificamerican08091884-
85.pdf.

Anonymous:1884:IFS

[6098] Anonymous. Improved friction screw press. Scientific American, 51(6):86, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n6/pdf/scientificamerican08091884-
86.pdf.

Anonymous:1884:NSW

[6099] Anonymous. The new stern-wheel gunboats of the French Navy. Scientific American, 51(6):87–89, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v51/n6/pdf/scientificamerican08091884-
87.pdf.

Anonymous:1884:EIbv

[6100] Anonymous. Engineering inventions. Scientific American, 51(6):91, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n6/pdf/scientificamerican08091884-91.pdf.

Anonymous:1884:MIdj

[6101] Anonymous. Mechanical inventions. Scientific American, 51(6):91, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n6/pdf/scientificamerican08091884-91a.pdf.

Anonymous:1884:AIbs

[6102] Anonymous. Agricultural inventions. Scientific American, 51(6):91, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n6/pdf/scientificamerican08091884-91b.pdf.

Anonymous:1884:MIdk

[6103] Anonymous. Miscellaneous inventions. Scientific American, 51(6):91, August 9, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n6/pdf/scientificamerican08091884-91c.pdf.

Anonymous:1884:BPby

[6104] Anonymous. Business and personal. Scientific American, 51(7):107, August 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n7/pdf/scientificamerican08161884-107d.pdf.

Anonymous:1884:NQbx

[6105] Anonymous. Notes and queries. Scientific American, 51(7):107, August 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n7/pdf/scientificamerican08161884-107e.pdf.

Anonymous:1884:IIbx

[6106] Anonymous. Index of inventions. Scientific American, 51(7): 107–108, August 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n7/pdf/scientificamerican08161884-
107f.pdf.

Anonymous:1884:PNPbm

[6107] Anonymous. Patent notes: Patents. Scientific American, 51(7):109, August 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n7/pdf/scientificamerican08161884-109.pdf.

Hovey:1884:AWW

[6108] H. C. Hovey. The American Watch Works. Scientific American, 51(7):95, August 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n7/pdf/scientificamerican08161884-
95.pdf.

Anonymous:1884:GEH

[6109] Anonymous. The Greely explorers at home. Scientific American, 51(7):96, August 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n7/pdf/scientificamerican08161884-
96.pdf.

Anonymous:1884:SBH

[6110] Anonymous. Screw bolt heads. Scientific American, 51(7):96, August 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n7/pdf/scientificamerican08161884-96a.pdf.

Anonymous:1884:SSc

[6111] Anonymous. Straightening shafts. Scientific American, 51(7):96, August 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n7/pdf/scientificamerican08161884-96b.pdf.

Anonymous:1884:STd

[6112] Anonymous. Steel tests. Scientific American, 51(7):96–98, August 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n7/pdf/scientificamerican08161884-96c.pdf.

Anonymous:1884:PBe

[6113] Anonymous. Portable barn. Scientific American, 51(7):98, August 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n7/pdf/scientificamerican08161884-98.pdf.

Anonymous:1884:SSW

[6114] Anonymous. Spring seat for wagons. Scientific American, 51(7):98, August 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n7/pdf/scientificamerican08161884-98a.pdf.

Anonymous:1884:LGC

[6115] Anonymous. Lumber grader’s combination tool. Scientific American, 51(7):98–99, August 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n7/pdf/scientificamerican08161884-
98b.pdf.

Anonymous:1884:MTM

[6116] Anonymous. Means for transmitting motion. Scientific American, 51(7):99, August 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n7/pdf/scientificamerican08161884-
99.pdf.

Anonymous:1884:HAF

[6117] Anonymous. Hot air furnace. Scientific American, 51(7):99–100, August 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n7/pdf/scientificamerican08161884-99a.pdf.

Anonymous:1884:DOC

[6118] Anonymous. Dredging at Oakland, Cal. Scientific American, 51(7):101, August 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n7/pdf/scientificamerican08161884-
101.pdf.

Anonymous:1884:EPV

[6119] Anonymous. Electro-pneumatic valve. Scientific American, 51 (7):105, August 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n7/pdf/scientificamerican08161884-
105.pdf.

Anonymous:1884:HOC

[6120] Anonymous. How our carp are destroyed. Scientific American, 51 (7):105–106, August 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n7/pdf/scientificamerican08161884-
105a.pdf.

Anonymous:1884:RFB

[6121] Anonymous. A Russian flower basket. Scientific American, 51 (7):106, August 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n7/pdf/scientificamerican08161884-
106.pdf.

Anonymous:1884:AFE

[6122] Anonymous. An automatic fire extinguisher. Scientific American, 51(7):106, August 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n7/pdf/scientificamerican08161884-
106a.pdf.

Anonymous:1884:CSg

[6123] Anonymous. Cuspidor stand. Scientific American, 51(7):106, August 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n7/pdf/scientificamerican08161884-106b.pdf.

Anonymous:1884:EIbw

[6124] Anonymous. Engineering inventions. Scientific American, 51(7): 107, August 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n7/pdf/scientificamerican08161884-
107.pdf.

Anonymous:1884:MIdl

[6125] Anonymous. Mechanical inventions. Scientific American, 51(7):107, August 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n7/pdf/scientificamerican08161884-107a.pdf.

Anonymous:1884:AIbt (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n7/pdf/scientificamerican08161884-
107b.pdf.

Anonymous:1884:MIdm

[6127] Anonymous. Miscellaneous inventions. Scientific American, 51 (7):107, August 16, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n7/pdf/scientificamerican08161884-
107c.pdf.

Anonymous:1884:Cby

[6128] Anonymous. Correspondence. Scientific American, 51(8):117, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n8/pdf/scientificamerican08231884-117.pdf.

Anonymous:1884:BPbz

[6129] Anonymous. Business and personal. Scientific American, 51(8):122, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n8/pdf/scientificamerican08231884-122d.pdf.

Anonymous:1884:NQby

[6130] Anonymous. Notes and queries. Scientific American, 51(8):122– 123, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n8/pdf/scientificamerican08231884-
122e.pdf.

Anonymous:1884:IIby

[6131] Anonymous. Index of inventions. Scientific American, 51(8): 123–124, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n8/pdf/scientificamerican08231884-
123.pdf.

Anonymous:1884:PNPbn

[6132] Anonymous. Patent notes: Patents. Scientific American, 51(8):125, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n8/pdf/scientificamerican08231884-125.pdf.

Anonymous:1884:WCCa

[6133] Anonymous. Work on the Corinth Canal. Scientific American, 51(8):111, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n8/pdf/scientificamerican08231884-
111.pdf.

Anonymous:1884:EAc

[6134] Anonymous. The earthquake of August 10. Scientific American, 51(8):112–113, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n8/pdf/scientificamerican08231884-
112.pdf.

Anonymous:1884:SCIb

[6135] Anonymous. Smelting and casting of iron. Scientific American, 51(8):113, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n8/pdf/scientificamerican08231884-
113.pdf.

Anonymous:1884:CEa

[6136] Anonymous. Cooling by evaporation. Scientific American, 51 (8):113, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n8/pdf/scientificamerican08231884-
113a.pdf.

Anonymous:1884:EHS

[6137] Anonymous. The effect of hardening on steel. Scientific American, 51(8):113, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n8/pdf/scientificamerican08231884-
113b.pdf.

Anonymous:1884:BBB

[6138] Anonymous. Back band buckle. Scientific American, 51(8):114, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n8/pdf/scientificamerican08231884-114.pdf.

Anonymous:1884:SPP

[6139] Anonymous. Semicircular pointed pens. Scientific American, 51 (8):114, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n8/pdf/scientificamerican08231884-
114a.pdf.

Anonymous:1884:GPG

[6140] Anonymous. Geographical puzzles and games. Scientific American, 51(8):114–115, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n8/pdf/scientificamerican08231884-
114b.pdf.

Anonymous:1884:PLb

[6141] Anonymous. Pole ladder. Scientific American, 51(8):115, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n8/pdf/scientificamerican08231884-115.pdf.

Anonymous:1884:ILc

[6142] Anonymous. An improved locomotive. Scientific American, 51 (8):115–116, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n8/pdf/scientificamerican08231884-
115a.pdf.

Anonymous:1884:GP

[6143] Anonymous. The Greindl pump. Scientific American, 51(8):118, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n8/pdf/scientificamerican08231884-118.pdf.

Anonymous:1884:NSd

[6144] Anonymous. A novel sled. Scientific American, 51(8):118, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n8/pdf/scientificamerican08231884-118a.pdf.

Anonymous:1884:IBSb

[6145] Anonymous. Improved buck saw. Scientific American, 51(8):118, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n8/pdf/scientificamerican08231884-118b.pdf.

Anonymous:1884:FHP

[6146] Anonymous. Fifty horse power engine and boiler. Scientific American, 51(8):119, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n8/pdf/scientificamerican08231884-
119.pdf.

Anonymous:1884:NWN

[6147] Anonymous. A nurse that will never be careless. Scientific American, 51(8):120, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n8/pdf/scientificamerican08231884-
120.pdf.

Anonymous:1884:PMA

[6148] Anonymous. Post mortem attitudes. Scientific American, 51 (8):120–121, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n8/pdf/scientificamerican08231884-
120a.pdf.

Anonymous:1884:COI

[6149] Anonymous. Chemical obstructions in iron water pipes. Scientific American, 51(8):121, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n8/pdf/scientificamerican08231884-
121.pdf.

Anonymous:1884:NTd

[6150] Anonymous. A novel toy. Scientific American, 51(8):121, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n8/pdf/scientificamerican08231884-121a.pdf.

Anonymous:1884:EIbx

[6151] Anonymous. Engineering inventions. Scientific American, 51(8): 122, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n8/pdf/scientificamerican08231884-
122.pdf.

Anonymous:1884:MIdn

[6152] Anonymous. Mechanical inventions. Scientific American, 51(8):122, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n8/pdf/scientificamerican08231884-122a.pdf.

Anonymous:1884:AIbu

[6153] Anonymous. Agricultural inventions. Scientific American, 51(8): 122, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n8/pdf/scientificamerican08231884-
122b.pdf.

Anonymous:1884:MIdo

[6154] Anonymous. Miscellaneous inventions. Scientific American, 51 (8):122, August 23, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n8/pdf/scientificamerican08231884-
122c.pdf.

Anonymous:1884:Cbz

[6155] Anonymous. Correspondence. Scientific American, 51(9):133, August 30, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n9/pdf/scientificamerican08301884-133.pdf.

Anonymous:1884:BPcb

[6156] Anonymous. Business and personal. Scientific American, 51(9):138, August 30, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n9/pdf/scientificamerican08301884-138d.pdf.

Anonymous:1884:NQbz

[6157] Anonymous. Notes and queries. Scientific American, 51(9):138– 139, August 30, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n9/pdf/scientificamerican08301884-
138e.pdf.

Anonymous:1884:IIbz 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n9/pdf/scientificamerican08301884-
139.pdf.

Anonymous:1884:PNPbo

[6159] Anonymous. Patent notes: Patents. Scientific American, 51(9):141, August 30, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n9/pdf/scientificamerican08301884-141.pdf.

Anonymous:1884:FTL

[6160] Anonymous. Fourteen ton locomotive crane. Scientific American, 51(9):127, August 30, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n9/pdf/scientificamerican08301884-
127.pdf.

Anonymous:1884:BWC

[6161] Anonymous. Balancing wheels and cylinders. Scientific American, 51(9):128, August 30, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n9/pdf/scientificamerican08301884-
128.pdf.

Anonymous:1884:APS

[6162] Anonymous. Aspects of the planets in September. Scientific American, 51(9):129–130, August 30, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n9/pdf/scientificamerican08301884-
129.pdf.

Anonymous:1884:CBPa

[6163] Anonymous. Cotton baling press. Scientific American, 51(9):130, August 30, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n9/pdf/scientificamerican08301884-130.pdf.

Anonymous:1884:PCCa

[6164] Anonymous. Paper and card cutter. Scientific American, 51(9):130, August 30, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n9/pdf/scientificamerican08301884-130a.pdf.

Anonymous:1884:AP

[6165] Anonymous. An aerial propeller. Scientific American, 51(9): 130–131, August 30, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n9/pdf/scientificamerican08301884-
130b.pdf.

Anonymous:1884:PLBa

[6166] Anonymous. Paper or letter box. Scientific American, 51(9):131, August 30, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n9/pdf/scientificamerican08301884-131.pdf.

Anonymous:1884:BTH

[6167] Anonymous. Bag and twine holder. Scientific American, 51(9):131, August 30, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n9/pdf/scientificamerican08301884-131a.pdf.

Anonymous:1884:SEM

[6168] Anonymous. Stamp and envelope moistener. Scientific American, 51(9):131–132, August 30, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n9/pdf/scientificamerican08301884-
131b.pdf.

Anonymous:1884:DM

[6169] Anonymous. Delta metal. Scientific American, 51(9):134, August 30, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n9/pdf/scientificamerican08301884-134.pdf.

Anonymous:1884:CCh

[6170] Anonymous. Car coupling. Scientific American, 51(9):134, August 30, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n9/pdf/scientificamerican08301884-134a.pdf.

Anonymous:1884:CEP

[6171] Anonymous. Couvreux’s excavator with pivoting bucket frame. Scientific American, 51(9):135–137, August 30, 1884. CODEN SCAMAC.

ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v51/n9/pdf/scientificamerican08301884-
135.pdf.

Anonymous:1884:EIby

[6172] Anonymous. Engineering inventions. Scientific American, 51(9): 138, August 30, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n9/pdf/scientificamerican08301884-
138.pdf.

Anonymous:1884:MIdp

[6173] Anonymous. Mechanical inventions. Scientific American, 51(9):138, August 30, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v51/n9/pdf/scientificamerican08301884-138a.pdf.

Anonymous:1884:AIbv

[6174] Anonymous. Agricultural inventions. Scientific American, 51(9): 138, August 30, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n9/pdf/scientificamerican08301884-
138b.pdf.

Anonymous:1884:MIdq

[6175] Anonymous. Miscellaneous inventions. Scientific American, 51 (9):138, August 30, 1884. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v51/n9/pdf/scientificamerican08301884-
138c.pdf.

Anonymous:1885:BPa

[6176] Anonymous. Business and personal. Scientific American, 52(1):11, January 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n1/pdf/scientificamerican01031885-11a.pdf.

Anonymous:1885:NQa

[6177] Anonymous. Notes and queries. Scientific American, 52(1):11, January 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n1/pdf/scientificamerican01031885-11b.pdf.

Anonymous:1885:IIa

[6178] Anonymous. Index of inventions. Scientific American, 52(1):11–12, January 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n1/pdf/scientificamerican01031885-11c.pdf.

Anonymous:1885:PNPa

[6179] Anonymous. Patent notes: Patents. Scientific American, 52(1):13, January 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n1/pdf/scientificamerican01031885-13.pdf.

Anonymous:1885:NOEa

[6180] Anonymous. The New Orleans Exposition, the Panama Canal and its climatic results, and more. Scientific American, 52(1):2–3, January 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n1/pdf/scientificamerican01031885-2.pdf.

Anonymous:1885:FLF

[6181] Anonymous. A flexible ladder fire escape. Scientific American, 52(1):4, January 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n1/pdf/scientificamerican01031885-4.pdf.

Anonymous:1885:LI

[6182] Anonymous. Leveling instrument. Scientific American, 52(1):4, January 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n1/pdf/scientificamerican01031885-4a.pdf.

Anonymous:1885:GFC

[6183] Anonymous. Gravity friction clutch. Scientific American, 52(1):4, January 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n1/pdf/scientificamerican01031885-4b.pdf.

Anonymous:1885:DCD

[6184] Anonymous. Dust cover for drawers. Scientific American, 52(1):4, January 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n1/pdf/scientificamerican01031885-4c.pdf.

Anonymous:1885:WCM

[6185] Anonymous. A “Workingman’s” College, Melbourne. Scientific American, 52(1):5, January 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n1/pdf/scientificamerican01031885-
5.pdf.

Anonymous:1885:AIN

[6186] Anonymous. American industries — no. 93. Scientific American, 52(1):6–7, January 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n1/pdf/scientificamerican01031885-
6.pdf.

Anonymous:1885:SWRa

[6187] Anonymous. The Superga wire rope railway near Turin. Scientific American, 52(1):8, January 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n1/pdf/scientificamerican01031885-
8.pdf.

Beard:1885:AAW

[6188] Daniel Carter Beard. Architecture of the American white-footed mouse (deer mouse). Scientific American, 52(1):9, January 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n1/pdf/scientificamerican01031885-9.pdf.

Anonymous:1885:DRPa

[6189] Anonymous. Decisions relating to patents. Scientific American, 52(1):10, January 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n1/pdf/scientificamerican01031885-
10.pdf.

Anonymous:1885:MIa

[6190] Anonymous. Miscellaneous inventions. Scientific American, 52(1):11, January 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n1/pdf/scientificamerican01031885-11.pdf.

Anonymous:1885:Ca

[6191] Anonymous. Correspondence. Scientific American, 52(10):149, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n10/pdf/scientificamerican03071885-149.pdf.

Anonymous:1885:BNBa

[6192] Anonymous. Books: New books and publications. Scientific American, 52(10):153, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n10/pdf/scientificamerican03071885-
153c.pdf.

Anonymous:1885:BPb

[6193] Anonymous. Business and personal. Scientific American, 52(10):153, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n10/pdf/scientificamerican03071885-153d.pdf.

Anonymous:1885:NQb

[6194] Anonymous. Notes and queries. Scientific American, 52(10):153–155, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n10/pdf/scientificamerican03071885-153e.pdf.

Anonymous:1885:IIb

[6195] Anonymous. Index of inventions. Scientific American, 52(10):155–156, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n10/pdf/scientificamerican03071885-155.pdf.

Anonymous:1885:PNPb

[6196] Anonymous. Patent notes: Patents. Scientific American, 52(10):157, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n10/pdf/scientificamerican03071885-157.pdf.

Anonymous:1885:M

[6197] Anonymous. El Mahdi. Scientific American, 52(10):144, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n10/pdf/scientificamerican03071885-144.pdf.

Anonymous:1885:MCa

[6198] Anonymous. The mechanic’s capital. Scientific American, 52(10):144, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n10/pdf/scientificamerican03071885-144a.pdf.

Anonymous:1885:WP

[6199] Anonymous. Webbed pulleys. Scientific American, 52(10):144, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n10/pdf/scientificamerican03071885-144b.pdf.

Anonymous:1885:TLI

[6200] Anonymous. A thrifty life insurance company. Scientific American, 52(10):144, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n10/pdf/scientificamerican03071885-
144c.pdf.

Anonymous:1885:CWM

[6201] Anonymous. Completion of Washington’s Monument. Scientific American, 52(10):145, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n10/pdf/scientificamerican03071885-
145.pdf.

Anonymous:1885:AR

[6202] Anonymous. The Alaska’s rudder. Scientific American, 52(10):145, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n10/pdf/scientificamerican03071885-145a.pdf.

Anonymous:1885:IST

[6203] Anonymous. Improved shaft tugs. Scientific American, 52(10):146, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n10/pdf/scientificamerican03071885-146.pdf.

Anonymous:1885:FPa

[6204] Anonymous. Frying pan. Scientific American, 52(10):146, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n10/pdf/scientificamerican03071885-146a.pdf.

Anonymous:1885:SBT

[6205] Anonymous. Support for buggy tops. Scientific American, 52(10):146, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n10/pdf/scientificamerican03071885-146b.pdf.

Anonymous:1885:SIa

[6206] Anonymous. Stove implement. Scientific American, 52(10):146, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n10/pdf/scientificamerican03071885-146c.pdf.

Anonymous:1885:BD

[6207] Anonymous. Bridge over the Dnieper. Scientific American, 52(10):147, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n10/pdf/scientificamerican03071885-147.pdf.

Anonymous:1885:RGS

[6208] Anonymous. The rudder of a great steamship. Scientific American, 52(10):148, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n10/pdf/scientificamerican03071885-
148.pdf.

Anonymous:1885:ELc

[6209] Anonymous. Extinguisher for lamps. Scientific American, 52(10):148, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n10/pdf/scientificamerican03071885-148a.pdf.

Anonymous:1885:GEM

[6210] Anonymous. A great electric machine. Scientific American, 52(10):150, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n10/pdf/scientificamerican03071885-150.pdf.

Anonymous:1885:IEE

[6211] Anonymous. Interesting electrical experiments. Scientific American, 52(10):150–151, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n10/pdf/scientificamerican03071885-
150a.pdf.

Anonymous:1885:A

[6212] Anonymous. Æchmeas. Scientific American, 52(10):151–152, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n10/pdf/scientificamerican03071885-151.pdf.

Anonymous:1885:EIa

[6213] Anonymous. Engineering inventions. Scientific American, 52(10):153, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n10/pdf/scientificamerican03071885-153.pdf.

Anonymous:1885:AIa

[6214] Anonymous. Agricultural inventions. Scientific American, 52(10):153, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n10/pdf/scientificamerican03071885-153a.pdf.

Anonymous:1885:MIb

[6215] Anonymous. Miscellaneous inventions. Scientific American, 52(10):153, March 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n10/pdf/scientificamerican03071885-153b.pdf.

Anonymous:1885:Cb

[6216] Anonymous. Correspondence. Scientific American, 52(11):165, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n11/pdf/scientificamerican03141885-165.pdf.

Anonymous:1885:BNBb

[6217] Anonymous. Books: New books and publications. Scientific American, 52(11):169, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n11/pdf/scientificamerican03141885-
169d.pdf.

Anonymous:1885:BPc

[6218] Anonymous. Business and personal. Scientific American, 52(11):169, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n11/pdf/scientificamerican03141885-169f.pdf.

Anonymous:1885:NQc

[6219] Anonymous. Notes and queries. Scientific American, 52(11):170–171, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n11/pdf/scientificamerican03141885-170.pdf.

Anonymous:1885:IIc

[6220] Anonymous. Index of inventions. Scientific American, 52(11):171–172, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n11/pdf/scientificamerican03141885-171.pdf.

Anonymous:1885:PNPc

[6221] Anonymous. Patent notes: Patents. Scientific American, 52(11):173, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n11/pdf/scientificamerican03141885-173.pdf.

Anonymous:1885:NAC

[6222] Anonymous. A novel arrangement of cog wheels. Scientific American, 52(11):159, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n11/pdf/scientificamerican03141885-
159.pdf.

Anonymous:1885:LSA

[6223] Anonymous. Light-ships across the Atlantic. Scientific American, 52(11):160, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n11/pdf/scientificamerican03141885-
160.pdf.

Anonymous:1885:MT

[6224] Anonymous. Making a tube. Scientific American, 52(11):160, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n11/pdf/scientificamerican03141885-160a.pdf.

Anonymous:1885:HL

[6225] Anonymous. Horace lord. Scientific American, 52(11):160, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n11/pdf/scientificamerican03141885-160b.pdf.

Anonymous:1885:HSE

[6226] Anonymous. How shall the Erie Canal be improved? Scientific American, 52(11):160–161, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n11/pdf/scientificamerican03141885-
160c.pdf.

Anonymous:1885:DDP

[6227] Anonymous. Doctor Dowling on pneumonia. Scientific American, 52(11):161, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n11/pdf/scientificamerican03141885-
161.pdf.

Anonymous:1885:ISSa

[6228] Anonymous. An improved saw swage. Scientific American, 52(11):162, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n11/pdf/scientificamerican03141885-162.pdf.

Anonymous:1885:ISSb

[6229] Anonymous. Improved scroll saw. Scientific American, 52(11):162, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n11/pdf/scientificamerican03141885-162a.pdf.

Anonymous:1885:STF

[6230] Anonymous. A sectional tile fence post. Scientific American, 52(11):162, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n11/pdf/scientificamerican03141885-162b.pdf.

Anonymous:1885:ILa

[6231] Anonymous. Improved lath. Scientific American, 52(11):162, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n11/pdf/scientificamerican03141885-162c.pdf.

Anonymous:1885:RAP

[6232] Anonymous. A refrigerator and air pressure apparatus. Scientific American, 52(11):163, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n11/pdf/scientificamerican03141885-
163.pdf.

Anonymous:1885:PPa

[6233] Anonymous. The pneumatic pulverizer. Scientific American, 52 (11):163, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n11/pdf/scientificamerican03141885-
163a.pdf.

Anonymous:1885:DAC

[6234] Anonymous. A double automatic car coupler. Scientific American, 52(11):163, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n11/pdf/scientificamerican03141885-
163b.pdf.

Anonymous:1885:SFS

[6235] Anonymous. A suction fitting for steam heating apparatus. Scientific American, 52(11):164, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n11/pdf/scientificamerican03141885-
164.pdf.

Anonymous:1885:ELB

[6236] Anonymous. Explosion of a lard boiler. Scientific American, 52(11):166, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n11/pdf/scientificamerican03141885-166.pdf.

Anonymous:1885:ES

[6237] Anonymous. Experiments with the siphon. Scientific American, 52(11):166, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n11/pdf/scientificamerican03141885-
166a.pdf.

Anonymous:1885:FCB

[6238] Anonymous. Four-coupled bogie engine, Midland Railway. Scientific American, 52(11):166–167, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v52/n11/pdf/scientificamerican03141885-
166b.pdf.

Anonymous:1885:ECS

[6239] Anonymous. The Egyptian Camel Service. Scientific American, 52 (11):167–168, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n11/pdf/scientificamerican03141885-
167.pdf.

Anonymous:1885:EIb

[6240] Anonymous. Engineering inventions. Scientific American, 52(11):169, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n11/pdf/scientificamerican03141885-169.pdf.

Anonymous:1885:MIc

[6241] Anonymous. Mechanical inventions. Scientific American, 52(11):169, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n11/pdf/scientificamerican03141885-169a.pdf.

Anonymous:1885:AIb

[6242] Anonymous. Agricultural inventions. Scientific American, 52(11):169, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n11/pdf/scientificamerican03141885-169b.pdf.

Anonymous:1885:MId

[6243] Anonymous. Miscellaneous inventions. Scientific American, 52(11):169, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n11/pdf/scientificamerican03141885-169c.pdf.

Anonymous:1885:Sb

[6244] Anonymous. Special. Scientific American, 52(11):169, March 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n11/pdf/scientificamerican03141885-169e.pdf.

Anonymous:1885:BPd

[6245] Anonymous. Business and personal. Scientific American, 52(12):186, March 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n12/pdf/scientificamerican03211885-186c.pdf.

Anonymous:1885:NQd

[6246] Anonymous. Notes and queries. Scientific American, 52(12):186–187, March 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n12/pdf/scientificamerican03211885-186d.pdf.

Anonymous:1885:IId

[6247] Anonymous. Index of inventions. Scientific American, 52(12):187–188, March 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n12/pdf/scientificamerican03211885-187.pdf.

Anonymous:1885:PNPd

[6248] Anonymous. Patent notes: Patents. Scientific American, 52(12):189, March 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n12/pdf/scientificamerican03211885-189.pdf.

Anonymous:1885:EIR

[6249] Anonymous. Engines of the Italian Ram Etna. Scientific American, 52(12):175, March 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n12/pdf/scientificamerican03211885-
175.pdf.

Anonymous:1885:PK

[6250] Anonymous. Personal knowledge. Scientific American, 52(12):176, March 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n12/pdf/scientificamerican03211885-176.pdf.

Anonymous:1885:CAa

[6251] Anonymous. Comets and asteroids of 1884. Scientific American, 52(12):176, March 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n12/pdf/scientificamerican03211885-
176a.pdf.

Anonymous:1885:SAa

[6252] Anonymous. Steering the Alaska. Scientific American, 52(12):177, March 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n12/pdf/scientificamerican03211885-177.pdf.

Anonymous:1885:NSC

[6253] Anonymous. A new safety coupler. Scientific American, 52(12):178, March 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n12/pdf/scientificamerican03211885-178.pdf.

Anonymous:1885:IFC

[6254] Anonymous. An improved friction clutch. Scientific American, 52(12):178, March 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n12/pdf/scientificamerican03211885-
178a.pdf.

Anonymous:1885:SGBa

[6255] Anonymous. Safety gas burner. Scientific American, 52(12):178, March 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n12/pdf/scientificamerican03211885-178b.pdf.

Anonymous:1885:IMa

[6256] Anonymous. The illustrator’s microscope. Scientific American, 52(12):178, March 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n12/pdf/scientificamerican03211885-
178c.pdf.

Anonymous:1885:PL

[6257] Anonymous. Photographing the larynx. Scientific American, 52 (12):179–180, March 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n12/pdf/scientificamerican03211885-
179.pdf.

Anonymous:1885:ICFa

[6258] Anonymous. An improved cuff fastener. Scientific American, 52(12):180, March 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n12/pdf/scientificamerican03211885-
180.pdf.

Anonymous:1885:ISCa

[6259] Anonymous. Improved sleeve or collar button. Scientific American, 52(12):180, March 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n12/pdf/scientificamerican03211885-
180a.pdf.

Anonymous:1885:IRP

[6260] Anonymous. Improvement in running-part of baby carriage. Scientific American, 52(12):180, March 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n12/pdf/scientificamerican03211885-
180b.pdf.

Anonymous:1885:ACC

[6261] Anonymous. Automatic cut-off for cisterns. Scientific American, 52(12):180–183, March 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n12/pdf/scientificamerican03211885-
180c.pdf.

Anonymous:1885:BEC

[6262] Anonymous. A boiler explosion at Cincinnati, Ohio. Scientific American, 52(12):183, March 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n12/pdf/scientificamerican03211885-
183.pdf.

Anonymous:1885:PSMa

[6263] Anonymous. Practical studies of man’s locomotion. Scientific American, 52(12):183–185, March 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n12/pdf/scientificamerican03211885-
183a.pdf.

Anonymous:1885:EIc

[6264] Anonymous. Engineering inventions. Scientific American, 52(12):186, March 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n12/pdf/scientificamerican03211885-186.pdf.

Anonymous:1885:AIc

[6265] Anonymous. Agricultural inventions. Scientific American, 52(12):186, March 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n12/pdf/scientificamerican03211885-186a.pdf.

Anonymous:1885:MIe

[6266] Anonymous. Miscellaneous inventions. Scientific American, 52(12):186, March 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n12/pdf/scientificamerican03211885-186b.pdf.

Anonymous:1885:BPe

[6267] Anonymous. Business and personal. Scientific American, 52(13):202, March 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n13/pdf/scientificamerican03281885-202c.pdf.

Anonymous:1885:NQe

[6268] Anonymous. Notes and queries. Scientific American, 52(13):202–203, March 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n13/pdf/scientificamerican03281885-202d.pdf.

Anonymous:1885:IIe

[6269] Anonymous. Index of inventions. Scientific American, 52(13):203–204, March 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n13/pdf/scientificamerican03281885-203.pdf.

Anonymous:1885:PNPe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n13/pdf/scientificamerican03281885-206.pdf.

Anonymous:1885:NCPa

[6271] Anonymous. A new Commissioner of Patents. Scientific American, 52(13):192, March 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n13/pdf/scientificamerican03281885-
192.pdf.

Anonymous:1885:POE

[6272] Anonymous. Patent Office examinations of novelty of inventions. Scientific American, 52(13):192, March 28, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v52/
n13/pdf/scientificamerican03281885-192a.pdf.

Anonymous:1885:WPMa

[6273] Anonymous. The working power of man. Scientific American, 52 (13):192–193, March 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n13/pdf/scientificamerican03281885-
192b.pdf.

Anonymous:1885:APAa

[6274] Anonymous. Aspects of the planets for April. Scientific American, 52(13):193, March 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n13/pdf/scientificamerican03281885-
193.pdf.

Anonymous:1885:SEG

[6275] Anonymous. Steam engine governor. Scientific American, 52(13):194, March 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n13/pdf/scientificamerican03281885-194.pdf.

Anonymous:1885:ICCa

[6276] Anonymous. Improved car coupling. Scientific American, 52(13):194, March 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n13/pdf/scientificamerican03281885-194a.pdf.

Anonymous:1885:IBWa

[6277] Anonymous. Improved barbed wire fence. Scientific American, 52(13):194, March 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n13/pdf/scientificamerican03281885-
194b.pdf.

Anonymous:1885:ITT

[6278] Anonymous. Improved tire tightener. Scientific American, 52(13):194, March 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n13/pdf/scientificamerican03281885-194c.pdf.

Anonymous:1885:FRBa

[6279] Anonymous. Friction roller bearing. Scientific American, 52 (13):194–195, March 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n13/pdf/scientificamerican03281885-
194d.pdf.

Anonymous:1885:AED

[6280] Anonymous. Artificial ear drum. Scientific American, 52(13):195, March 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n13/pdf/scientificamerican03281885-195.pdf.

Anonymous:1885:TWM

[6281] Anonymous. Tchebichef’s walking machine. Scientific American, 52(13):195–196, March 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n13/pdf/scientificamerican03281885-
195a.pdf.

Anonymous:1885:NYT

[6282] Anonymous. The New York trade schools. Scientific American, 52 (13):196–197, March 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n13/pdf/scientificamerican03281885-
196.pdf.

Anonymous:1885:HPB

[6283] Anonymous. Hand propeller for boats. Scientific American, 52(13):197, March 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n13/pdf/scientificamerican03281885-197.pdf.

Anonymous:1885:MSL

[6284] Anonymous. Machine for sewing linings in hat bodies. Scientific American, 52(13):198, March 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n13/pdf/scientificamerican03281885-
198.pdf.

Anonymous:1885:IHSa

[6285] Anonymous. Improved hydraulic shearing machine. Scientific American, 52(13):198–199, March 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n13/pdf/scientificamerican03281885-
198a.pdf.

Seiss:1885:MBV

[6286] C. Few Seiss. Migration of blowing vipers. Scientific American, 52(13):199, March 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n13/pdf/scientificamerican03281885-
199.pdf.

Anonymous:1885:MPb

[6287] Anonymous. Mother-of-pearl. Scientific American, 52(13):199–200, March 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n13/pdf/scientificamerican03281885-199a.pdf.

Anonymous:1885:CTT

[6288] Anonymous. The Columbia two track tricycle. Scientific American, 52(13):200, March 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n13/pdf/scientificamerican03281885-
200.pdf.

Anonymous:1885:RAa

[6289] Anonymous. The rudder of the Alaska. Scientific American, 52(13):201, March 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n13/pdf/scientificamerican03281885-201.pdf.

Anonymous:1885:EId

[6290] Anonymous. Engineering inventions. Scientific American, 52(13):202, March 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n13/pdf/scientificamerican03281885-202.pdf.

Anonymous:1885:AId

[6291] Anonymous. Agricultural inventions. Scientific American, 52(13):202, March 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n13/pdf/scientificamerican03281885-202a.pdf.

Anonymous:1885:MIf

[6292] Anonymous. Miscellaneous inventions. Scientific American, 52(13):202, March 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n13/pdf/scientificamerican03281885-202b.pdf.

Anonymous:1885:Cc

[6293] Anonymous. Correspondence. Scientific American, 52(14):213, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n14/pdf/scientificamerican04041885-213.pdf.

Anonymous:1885:BPf

[6294] Anonymous. Business and personal. Scientific American, 52(14):218, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n14/pdf/scientificamerican04041885-218d.pdf.

Anonymous:1885:NQf

[6295] Anonymous. Notes and queries. Scientific American, 52(14):219, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n14/pdf/scientificamerican04041885-219.pdf.

Anonymous:1885:IIf

[6296] Anonymous. Index of inventions. Scientific American, 52(14):219–220, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n14/pdf/scientificamerican04041885-219a.pdf.

Anonymous:1885:PNPf

[6297] Anonymous. Patent notes: Patents. Scientific American, 52(14):222, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n14/pdf/scientificamerican04041885-222.pdf.

Anonymous:1885:FB

[6298] Anonymous. The Forth Bridge. Scientific American, 52(14):207–208, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n14/pdf/scientificamerican04041885-207.pdf.

Anonymous:1885:CT

[6299] Anonymous. Concerning telescopes. Scientific American, 52(14): 208–209, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n14/pdf/scientificamerican04041885-
208.pdf.

Anonymous:1885:PPP

[6300] Anonymous. Progress of pneumatic postal tubes. Scientific American, 52(14):209, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n14/pdf/scientificamerican04041885-
209.pdf.

Anonymous:1885:ATS

[6301] Anonymous. Apparatus for testing sealed cans. Scientific American, 52(14):210, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n14/pdf/scientificamerican04041885-
210.pdf.

Anonymous:1885:CCa

[6302] Anonymous. Car coupling. Scientific American, 52(14):210, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n14/pdf/scientificamerican04041885-210a.pdf.

Anonymous:1885:ISLa

[6303] Anonymous. Improved seal lock. Scientific American, 52(14):210, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n14/pdf/scientificamerican04041885-210b.pdf.

Anonymous:1885:IDO

[6304] Anonymous. Improved door opener. Scientific American, 52(14):210, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n14/pdf/scientificamerican04041885-210c.pdf.

Anonymous:1885:SWRb

[6305] Anonymous. A steam wheel, or rotary engine. Scientific American, 52(14):211, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n14/pdf/scientificamerican04041885-
211.pdf.

Anonymous:1885:SGBb

[6306] Anonymous. Safety gas burner. Scientific American, 52(14):211, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n14/pdf/scientificamerican04041885-211a.pdf.

Anonymous:1885:CGD

[6307] Anonymous. Combined grain drier and cooler. Scientific American, 52(14):211–212, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n14/pdf/scientificamerican04041885-
211b.pdf.

Anonymous:1885:SSF

[6308] Anonymous. The sinking of shafts by freezing. Scientific American, 52(14):212, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n14/pdf/scientificamerican04041885-
212.pdf.

Anonymous:1885:ALP

[6309] Anonymous. Automatic loose pulley oiler. Scientific American, 52(14):212, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n14/pdf/scientificamerican04041885-
212a.pdf.

Anonymous:1885:CCU

[6310] Anonymous. Cover for cooking utensils. Scientific American, 52 (14):212–213, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n14/pdf/scientificamerican04041885-
212b.pdf.

Anonymous:1885:IEP

[6311] Anonymous. Imitation of electro-physiological apparatus and electric tuning-forks by hydrodynamic tuning-forks. Scientific American, 52(14):214, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n14/pdf/scientificamerican04041885-
214.pdf.

Anonymous:1885:IHLa

[6312] Anonymous. The improved hand loom Eugenie. Scientific American, 52(14):215, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n14/pdf/scientificamerican04041885-
215.pdf.

Anonymous:1885:IHLb

[6313] Anonymous. The improved hand loom Penelope. Scientific American, 52(14):215, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n14/pdf/scientificamerican04041885-
215a.pdf.

Anonymous:1885:RAb

[6314] Anonymous. The rudder of the Alaska. Scientific American, 52(14):216, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n14/pdf/scientificamerican04041885-216.pdf.

Anonymous:1885:VR

[6315] Anonymous. The Victoria Regia. Scientific American, 52(14):217, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n14/pdf/scientificamerican04041885-217.pdf.

Anonymous:1885:EIe

[6316] Anonymous. Engineering inventions. Scientific American, 52(14):218, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n14/pdf/scientificamerican04041885-218.pdf.

Anonymous:1885:AIe

[6317] Anonymous. Agricultural inventions. Scientific American, 52(14):218, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n14/pdf/scientificamerican04041885-218a.pdf.

Anonymous:1885:MIg

[6318] Anonymous. Miscellaneous inventions. Scientific American, 52(14):218, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n14/pdf/scientificamerican04041885-218b.pdf.

Anonymous:1885:Sc

[6319] Anonymous. Special. Scientific American, 52(14):218, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n14/pdf/scientificamerican04041885-218c.pdf.

Anonymous:1885:FPb

[6320] Anonymous. Foreign patents. Scientific American, 52(14):221, April 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n14/pdf/scientificamerican04041885-221.pdf.

Anonymous:1885:Cd

[6321] Anonymous. Correspondence. Scientific American, 52(15):232, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n15/pdf/scientificamerican04111885-232.pdf.

Anonymous:1885:BNBc

[6322] Anonymous. Books: New books and publications. Scientific American, 52(15):233, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n15/pdf/scientificamerican04111885-
233c.pdf.

Anonymous:1885:BPg

[6323] Anonymous. Business and personal. Scientific American, 52(15):233, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n15/pdf/scientificamerican04111885-233d.pdf.

Anonymous:1885:NQg

[6324] Anonymous. Notes and queries. Scientific American, 52(15):234–235, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n15/pdf/scientificamerican04111885-234.pdf.

Anonymous:1885:IIg

[6325] Anonymous. Index of inventions. Scientific American, 52(15):235–236, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n15/pdf/scientificamerican04111885-235.pdf.

Anonymous:1885:PNPg

[6326] Anonymous. Patent notes: Patents. Scientific American, 52(15):238, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n15/pdf/scientificamerican04111885-238.pdf.

Anonymous:1885:NOEb

[6327] Anonymous. The New Orleans Exposition. Scientific American, 52(15):223, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n15/pdf/scientificamerican04111885-
223.pdf.

Anonymous:1885:ILP

[6328] Anonymous. Improved loose pulleys. Scientific American, 52(15):224, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n15/pdf/scientificamerican04111885-224.pdf.

Anonymous:1885:REV

[6329] Anonymous. Recent experiments in vivisection. Scientific American, 52(15):224, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n15/pdf/scientificamerican04111885-
224a.pdf.

Anonymous:1885:AMa

[6330] Anonymous. Artistic mechanics. Scientific American, 52(15):224, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n15/pdf/scientificamerican04111885-224b.pdf.

Anonymous:1885:JAT

[6331] Anonymous. The Jack of all trades. Scientific American, 52 (15):224–225, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n15/pdf/scientificamerican04111885-
224c.pdf.

Anonymous:1885:WGR

[6332] Anonymous. Water gas ruled out of Massachusetts. Scientific American, 52(15):225, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n15/pdf/scientificamerican04111885-
225.pdf.

Anonymous:1885:PWP

[6333] Anonymous. Protector for water pipes against freezing. Scientific American, 52(15):226, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n15/pdf/scientificamerican04111885-
226.pdf.

Anonymous:1885:ITH

[6334] Anonymous. Improved thill holder. Scientific American, 52(15):226, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n15/pdf/scientificamerican04111885-226a.pdf.

Anonymous:1885:IDC

[6335] Anonymous. Improved dividers and calipers. Scientific American, 52(15):226, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n15/pdf/scientificamerican04111885-
226b.pdf.

Anonymous:1885:RVH

[6336] Anonymous. Reversing valve for hoists, elevators, etc. Scientific American, 52(15):227, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n15/pdf/scientificamerican04111885-
227.pdf.

Anonymous:1885:CO

[6337] Anonymous. The hat as a camera obscura. Scientific American, 52(15):227–228, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n15/pdf/scientificamerican04111885-
227a.pdf.

Anonymous:1885:BPS

[6338] Anonymous. Back for piano stools. Scientific American, 52(15):228, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n15/pdf/scientificamerican04111885-228.pdf.

Anonymous:1885:WFP

[6339] Anonymous. Wire flower pot holder. Scientific American, 52(15):228, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n15/pdf/scientificamerican04111885-228a.pdf.

Anonymous:1885:SCD

[6340] Anonymous. The “Sun” column, designed for lighting entire Paris. Scientific American, 52(15):230, April 11, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v52/
n15/pdf/scientificamerican04111885-230.pdf.

Anonymous:1885:CBa

[6341] Anonymous. A cheap battery. Scientific American, 52(15):230, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n15/pdf/scientificamerican04111885-230a.pdf.

Anonymous:1885:ICV

[6342] Anonymous. Improved construction of vessels. Scientific American, 52(15):230–231, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n15/pdf/scientificamerican04111885-
230b.pdf.

Anonymous:1885:JCM

[6343] Anonymous. Jersey cow Mollie Garfield. Scientific American, 52(15):231, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n15/pdf/scientificamerican04111885-
231.pdf.

Anonymous:1885:FL

[6344] Anonymous. A fight for life. Scientific American, 52(15):231, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n15/pdf/scientificamerican04111885-231a.pdf.

Anonymous:1885:RAc

[6345] Anonymous. The rudder of the Alaska. Scientific American, 52(15):232, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n15/pdf/scientificamerican04111885-232a.pdf.

Anonymous:1885:EIf

[6346] Anonymous. Engineering inventions. Scientific American, 52(15):233, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n15/pdf/scientificamerican04111885-233.pdf.

Anonymous:1885:AIf

[6347] Anonymous. Agricultural inventions. Scientific American, 52(15):233, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n15/pdf/scientificamerican04111885-233a.pdf.

Anonymous:1885:MIh

[6348] Anonymous. Miscellaneous inventions. Scientific American, 52(15):233, April 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n15/pdf/scientificamerican04111885-233b.pdf.

Anonymous:1885:Ce

[6349] Anonymous. Correspondence. Scientific American, 52(16):245–246, April 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n16/pdf/scientificamerican04181885-245a.pdf.

Anonymous:1885:BNBd

[6350] Anonymous. Books: New books and publications. Scientific American, 52(16):249, April 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n16/pdf/scientificamerican04181885-
249c.pdf.

Anonymous:1885:BPh

[6351] Anonymous. Business and personal. Scientific American, 52(16):249, April 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n16/pdf/scientificamerican04181885-249e.pdf.

Anonymous:1885:NQh

[6352] Anonymous. Notes and queries. Scientific American, 52(16):250–251, April 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n16/pdf/scientificamerican04181885-250.pdf.

Anonymous:1885:IIh

[6353] Anonymous. Index of inventions. Scientific American, 52(16):251–252, April 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n16/pdf/scientificamerican04181885-251.pdf.

Anonymous:1885:PNPh

[6354] Anonymous. Patent notes: Patents. Scientific American, 52(16):254, April 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n16/pdf/scientificamerican04181885-254.pdf.

Anonymous:1885:NOEc

[6355] Anonymous. The New Orleans Exposition. Scientific American, 52(16):239, April 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n16/pdf/scientificamerican04181885-
239.pdf.

Anonymous:1885:WBE

[6356] Anonymous. War between England and Russia. Scientific American, 52(16):240, April 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n16/pdf/scientificamerican04181885-
240.pdf.

Anonymous:1885:LF

[6357] Anonymous. Lathe fits. Scientific American, 52(16):240–241, April 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n16/pdf/scientificamerican04181885-240a.pdf.

Anonymous:1885:MRTa

[6358] Anonymous. Metallic railway tie. Scientific American, 52(16):242, April 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n16/pdf/scientificamerican04181885-242.pdf.

Anonymous:1885:NA

[6359] Anonymous. A novel anchor. Scientific American, 52(16):242, April 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n16/pdf/scientificamerican04181885-242a.pdf.

Anonymous:1885:ICCb

[6360] Anonymous. Improved car coupling. Scientific American, 52(16):242, April 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n16/pdf/scientificamerican04181885-242b.pdf.

Anonymous:1885:FCa

[6361] Anonymous. Fowl coop. Scientific American, 52(16):242–243, April 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n16/pdf/scientificamerican04181885-242c.pdf.

Anonymous:1885:SFRa

[6362] Anonymous. Self-feeding rip saw table. Scientific American, 52(16):243, April 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n16/pdf/scientificamerican04181885-243.pdf.

Anonymous:1885:EDF

[6363] Anonymous. Evaporator for drying fruit, etc. Scientific American, 52(16):243–244, April 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n16/pdf/scientificamerican04181885-
243a.pdf.

Anonymous:1885:CCC

[6364] Anonymous. Combined cigar-cutter and match box. Scientific American, 52(16):245, April 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n16/pdf/scientificamerican04181885-
245.pdf.

Anonymous:1885:COC

[6365] Anonymous. Compound optometer for correcting errors of refraction. Scientific American, 52(16):246–247, April 18, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v52/
n16/pdf/scientificamerican04181885-246.pdf.

Anonymous:1885:P

[6366] Anonymous. Planchette. Scientific American, 52(16):247, April 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n16/pdf/scientificamerican04181885-247.pdf.

Anonymous:1885:BSa

[6367] Anonymous. The bird spider. Scientific American, 52(16):247–248, April 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n16/pdf/scientificamerican04181885-247a.pdf.

Anonymous:1885:NTC

[6368] Anonymous. New top for cooking stoves and ranges. Scientific American, 52(16):248, April 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n16/pdf/scientificamerican04181885-
248.pdf.

Anonymous:1885:ISa

[6369] Anonymous. Improved stove. Scientific American, 52(16):248, April 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n16/pdf/scientificamerican04181885-248a.pdf.

Anonymous:1885:EIg

[6370] Anonymous. Engineering inventions. Scientific American, 52(16):249, April 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n16/pdf/scientificamerican04181885-249.pdf.

Anonymous:1885:AIg

[6371] Anonymous. Agricultural inventions. Scientific American, 52(16):249, April 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n16/pdf/scientificamerican04181885-249a.pdf.

Anonymous:1885:MIi

[6372] Anonymous. Miscellaneous inventions. Scientific American, 52(16):249, April 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n16/pdf/scientificamerican04181885-249b.pdf.

Anonymous:1885:Sd

[6373] Anonymous. Special. Scientific American, 52(16):249, April 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n16/pdf/scientificamerican04181885-249d.pdf.

Anonymous:1885:BPi

[6374] Anonymous. Business and personal. Scientific American, 52(17):266, April 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n17/pdf/scientificamerican04251885-266.pdf.

Anonymous:1885:NQi

[6375] Anonymous. Notes and queries. Scientific American, 52(17):266–267, April 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n17/pdf/scientificamerican04251885-266a.pdf.

Anonymous:1885:IIi

[6376] Anonymous. Index of inventions. Scientific American, 52(17):267–268, April 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n17/pdf/scientificamerican04251885-267.pdf.

Anonymous:1885:PNPi

[6377] Anonymous. Patent notes: Patents. Scientific American, 52(17):270, April 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n17/pdf/scientificamerican04251885-270.pdf.

Anonymous:1885:CSH

[6378] Anonymous. Casting a statue of heroic size. Scientific American, 52(17):255, April 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n17/pdf/scientificamerican04251885-
255.pdf.

Anonymous:1885:CPO

[6379] Anonymous. Changes in the Patent Office. Scientific American, 52(17):256, April 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n17/pdf/scientificamerican04251885-
256.pdf.

Anonymous:1885:Ha

[6380] Anonymous. Herat. Scientific American, 52(17):256, April 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n17/pdf/scientificamerican04251885-256a.pdf.

Anonymous:1885:DWW

[6381] Anonymous. Do we work too much? Scientific American, 52 (17):256–257, April 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n17/pdf/scientificamerican04251885-
256b.pdf.

Anonymous:1885:AWBa

[6382] Anonymous. Adjustable window bead fastener. Scientific American, 52(17):258, April 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n17/pdf/scientificamerican04251885-
258.pdf.

Anonymous:1885:DBS

[6383] Anonymous. Double bolt sash lock. Scientific American, 52(17):258, April 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n17/pdf/scientificamerican04251885-258a.pdf.

Anonymous:1885:ISb

[6384] Anonymous. Improved sled. Scientific American, 52(17):258, April 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n17/pdf/scientificamerican04251885-258b.pdf.

Anonymous:1885:IMR

[6385] Anonymous. Improved mine railway. Scientific American, 52(17):258, April 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n17/pdf/scientificamerican04251885-258c.pdf.

Anonymous:1885:DAW

[6386] Anonymous. Double-acting water wheel. Scientific American, 52 (17):259, April 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n17/pdf/scientificamerican04251885-
259.pdf.

Anonymous:1885:RAd

[6387] Anonymous. Rowing apparatus. Scientific American, 52(17):259, April 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n17/pdf/scientificamerican04251885-259a.pdf.

Anonymous:1885:CIP

[6388] Anonymous. Cold iron punch. Scientific American, 52(17):259, April 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n17/pdf/scientificamerican04251885-259b.pdf.

Anonymous:1885:ESM

[6389] Anonymous. An electrical standard for measuring light. Scientific American, 52(17):261–262, April 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v52/n17/pdf/scientificamerican04251885-
261.pdf.

Holder:1885:MN

[6390] C. P. Holder. Monumental nests. Scientific American, 52(17):263, April 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n17/pdf/scientificamerican04251885-263.pdf.

Anonymous:1885:DOL

[6391] Anonymous. Dr. Ochorowicz’s loud-speaking telephone. Scientific American, 52(17):263, April 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n17/pdf/scientificamerican04251885-
263a.pdf.

Anonymous:1885:POT

[6392] Anonymous. Pruning ornamental trees. Scientific American, 52 (17):264, April 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n17/pdf/scientificamerican04251885-
264.pdf.

Anonymous:1885:EIh

[6393] Anonymous. Engineering inventions. Scientific American, 52(17):265, April 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n17/pdf/scientificamerican04251885-265.pdf.

Anonymous:1885:AIh

[6394] Anonymous. Agricultural inventions. Scientific American, 52(17):265, April 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n17/pdf/scientificamerican04251885-265a.pdf.

Anonymous:1885:MIj

[6395] Anonymous. Miscellaneous inventions. Scientific American, 52(17):265, April 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n17/pdf/scientificamerican04251885-265b.pdf.

Anonymous:1885:BPj

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n18/pdf/scientificamerican05021885-281e.pdf.

Anonymous:1885:BNBe

[6397] Anonymous. Books: New books and publications. Scientific American, 52(18):281, May 2, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n18/pdf/scientificamerican05021885-
281f.pdf.

Anonymous:1885:NQj

[6398] Anonymous. Notes and queries. Scientific American, 52(18):282, May 2, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n18/pdf/scientificamerican05021885-282.pdf.

Anonymous:1885:IIj

[6399] Anonymous. Index of inventions. Scientific American, 52(18):282–284, May 2, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n18/pdf/scientificamerican05021885-282a.pdf.

Anonymous:1885:PNPj

[6400] Anonymous. Patent notes: Patents. Scientific American, 52(18):286, May 2, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n18/pdf/scientificamerican05021885-286.pdf.

Anonymous:1885:EPE

[6401] Anonymous. Exhibition of pumping engines by C. H. De Lamater & Co. at the New Orleans Exposition. Scientific American, 52(18):271, May 2, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n18/pdf/scientificamerican05021885-271.pdf.

Anonymous:1885:TML

[6402] Anonymous. Trade marks and label registration. Scientific American, 52(18):272, May 2, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n18/pdf/scientificamerican05021885-
272.pdf.

Anonymous:1885:RCa

[6403] Anonymous. Rust cement. Scientific American, 52(18):272, May 2, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n18/pdf/scientificamerican05021885-272a.pdf.

Anonymous:1885:VDC

[6404] Anonymous. Views of the District of Columbia judges on trademark and label registration. Scientific American, 52(18):273, May 2, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n18/pdf/scientificamerican05021885-273.pdf.

Anonymous:1885:AOC

[6405] Anonymous. Automatic oiler for crank pins. Scientific American, 52(18):274, May 2, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n18/pdf/scientificamerican05021885-
274.pdf.

Anonymous:1885:HPA

[6406] Anonymous. Hand power apparatus for hoisting and conveying broken ice, etc. Scientific American, 52(18):274, May 2, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v52/
n18/pdf/scientificamerican05021885-274a.pdf.

Anonymous:1885:CPT

[6407] Anonymous. Combined potato and tree planter. Scientific American, 52(18):274–275, May 2, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n18/pdf/scientificamerican05021885-
274b.pdf.

Anonymous:1885:VCP

[6408] Anonymous. The “Vulcan” cushioned power hammer. Scientific American, 52(18):275, May 2, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n18/pdf/scientificamerican05021885-
275.pdf.

Anonymous:1885:DPG

[6409] Anonymous. Double plunger geared pump. Scientific American, 52(18):275–276, May 2, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n18/pdf/scientificamerican05021885-
275a.pdf.

Anonymous:1885:CDP

[6410] Anonymous. Combined detachable pocket and cap. Scientific American, 52(18):276, May 2, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n18/pdf/scientificamerican05021885-
276.pdf.

Anonymous:1885:HDO

[6411] Anonymous. Heat distributer for oil stoves. Scientific American, 52(18):276, May 2, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n18/pdf/scientificamerican05021885-
276a.pdf.

Anonymous:1885:GDMa

[6412] Anonymous. The great drilling machines of the Forth Bridge. Scientific American, 52(18):278–279, May 2, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v52/n18/pdf/scientificamerican05021885-
278.pdf.

Anonymous:1885:APMa

[6413] Anonymous. Aspects of the planets for May. Scientific American, 52(18):280, May 2, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n18/pdf/scientificamerican05021885-
280.pdf.

Anonymous:1885:EIi

[6414] Anonymous. Engineering inventions. Scientific American, 52(18):281, May 2, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n18/pdf/scientificamerican05021885-281.pdf.

Anonymous:1885:MIk

[6415] Anonymous. Mechanical inventions. Scientific American, 52(18):281, May 2, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n18/pdf/scientificamerican05021885-281a.pdf.

Anonymous:1885:AIi

[6416] Anonymous. Agricultural inventions. Scientific American, 52(18):281, May 2, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n18/pdf/scientificamerican05021885-281b.pdf.

Anonymous:1885:MIl

[6417] Anonymous. Miscellaneous inventions. Scientific American, 52(18):281, May 2, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n18/pdf/scientificamerican05021885-281c.pdf.

Anonymous:1885:Se

[6418] Anonymous. Special. Scientific American, 52(18):281, May 2, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n18/pdf/scientificamerican05021885-281d.pdf.

Anonymous:1885:Cf

[6419] Anonymous. Correspondence. Scientific American, 52(19):293, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n19/pdf/scientificamerican05091885-293.pdf.

Anonymous:1885:BNBf

[6420] Anonymous. Books: New books and publications. Scientific American, 52(19):297, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n19/pdf/scientificamerican05091885-
297d.pdf.

Anonymous:1885:BPk

[6421] Anonymous. Business and personal. Scientific American, 52(19):298, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n19/pdf/scientificamerican05091885-298.pdf.

Anonymous:1885:NQk

[6422] Anonymous. Notes and queries. Scientific American, 52(19):298–299, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n19/pdf/scientificamerican05091885-298a.pdf.

Anonymous:1885:IIk

[6423] Anonymous. Index of inventions. Scientific American, 52(19):299–300, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n19/pdf/scientificamerican05091885-299.pdf.

Anonymous:1885:PNPk

[6424] Anonymous. Patent notes: Patents. Scientific American, 52(19):302, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n19/pdf/scientificamerican05091885-302.pdf.

Anonymous:1885:ICPa

[6425] Anonymous. Improved cotton picker. Scientific American, 52 (19):287–288, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n19/pdf/scientificamerican05091885-
287.pdf.

Anonymous:1885:SCPa

[6426] Anonymous. Square cotter pins. Scientific American, 52(19):288, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n19/pdf/scientificamerican05091885-288.pdf.

Anonymous:1885:SGa

[6427] Anonymous. Spring gauges. Scientific American, 52(19):288, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n19/pdf/scientificamerican05091885-288a.pdf.

Anonymous:1885:CNW

[6428] Anonymous. Cut nails and wire nails. Scientific American, 52 (19):288–289, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n19/pdf/scientificamerican05091885-
288b.pdf.

Kendall:1885:DDD

[6429] E. Dwight Kendall. Dirt, disease and disinfection. Scientific American, 52(19):289, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n19/pdf/scientificamerican05091885-
289.pdf.

Anonymous:1885:HMSa

[6430] Anonymous. How much shall the doctor be paid? Scientific American, 52(19):289, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n19/pdf/scientificamerican05091885-
289a.pdf.

Anonymous:1885:IAa

[6431] Anonymous. Improved anemometer. Scientific American, 52(19):290, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n19/pdf/scientificamerican05091885-290.pdf.

Anonymous:1885:FCc

[6432] Anonymous. Folding chair. Scientific American, 52(19):290, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n19/pdf/scientificamerican05091885-290a.pdf.

Anonymous:1885:ADa

[6433] Anonymous. An automatic disinfecter. Scientific American, 52(19):290, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n19/pdf/scientificamerican05091885-290b.pdf.

Anonymous:1885:FTB

[6434] Anonymous. Fish trap and bucket. Scientific American, 52(19):290, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n19/pdf/scientificamerican05091885-290c.pdf.

Anonymous:1885:BCa

[6435] Anonymous. Boiler cleaner. Scientific American, 52(19):291, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n19/pdf/scientificamerican05091885-291.pdf.

Anonymous:1885:NAM

[6436] Anonymous. A new aerial machine. Scientific American, 52(19):291, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n19/pdf/scientificamerican05091885-291a.pdf.

Anonymous:1885:TL

[6437] Anonymous. Trunk lock. Scientific American, 52(19):291–292, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n19/pdf/scientificamerican05091885-291b.pdf.

Anonymous:1885:NTb

[6438] Anonymous. A novel toy. Scientific American, 52(19):292, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n19/pdf/scientificamerican05091885-292.pdf.

Anonymous:1885:HH

[6439] Anonymous. Hobby horse. Scientific American, 52(19):292, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n19/pdf/scientificamerican05091885-292a.pdf.

Anonymous:1885:ACO

[6440] Anonymous. Apparatus for cleaning oils. Scientific American, 52(19):294, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n19/pdf/scientificamerican05091885-
294.pdf.

Anonymous:1885:PPC

[6441] Anonymous. A portable photographic camera. Scientific American, 52(19):294–295, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n19/pdf/scientificamerican05091885-
294a.pdf.

Anonymous:1885:NPG

[6442] Anonymous. The New Paleontological Gallery of the Paris Museum. Scientific American, 52(19):296, May 9, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v52/
n19/pdf/scientificamerican05091885-295.pdf.

Anonymous:1885:EIj

[6443] Anonymous. Engineering inventions. Scientific American, 52(19):297, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n19/pdf/scientificamerican05091885-297.pdf.

Anonymous:1885:MIm

[6444] Anonymous. Mechanical inventions. Scientific American, 52(19):297, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n19/pdf/scientificamerican05091885-297a.pdf.

Anonymous:1885:AIj

[6445] Anonymous. Agricultural inventions. Scientific American, 52(19):297, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n19/pdf/scientificamerican05091885-297b.pdf.

Anonymous:1885:MIn

[6446] Anonymous. Miscellaneous inventions. Scientific American, 52(19):297, May 9, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n19/pdf/scientificamerican05091885-297c.pdf.

Anonymous:1885:Cg

[6447] Anonymous. Correspondence. Scientific American, 52(2):24–26, January 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n2/pdf/scientificamerican01101885-24.pdf.

Anonymous:1885:BPl

[6448] Anonymous. Business and personal. Scientific American, 52(2):27, January 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n2/pdf/scientificamerican01101885-27d.pdf.

Anonymous:1885:NQl

[6449] Anonymous. Notes and queries. Scientific American, 52(2):27–28, January 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n2/pdf/scientificamerican01101885-27e.pdf.

Anonymous:1885:IIl

[6450] Anonymous. Index of inventions. Scientific American, 52(2):28, January 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n2/pdf/scientificamerican01101885-28.pdf.

Anonymous:1885:PNPl

[6451] Anonymous. Patent notes: Patents. Scientific American, 52(2):29, January 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n2/pdf/scientificamerican01101885-29.pdf.

Anonymous:1885:YMC

[6452] Anonymous. The young mechanics capital — how to increase it. Scientific American, 52(2):16–17, January 10, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v52/
n2/pdf/scientificamerican01101885-16.pdf.

Anonymous:1885:CCb

[6453] Anonymous. Cut-off for cisterns. Scientific American, 52(2):18, January 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n2/pdf/scientificamerican01101885-18.pdf.

Anonymous:1885:DBF

[6454] Anonymous. Draw bar for freight cars. Scientific American, 52 (2):18, January 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n2/pdf/scientificamerican01101885-
18a.pdf.

Anonymous:1885:ISV

[6455] Anonymous. Improved safety valve. Scientific American, 52(2):18, January 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n2/pdf/scientificamerican01101885-18b.pdf.

Anonymous:1885:FCS

[6456] Anonymous. Freight car skid. Scientific American, 52(2):18, January 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n2/pdf/scientificamerican01101885-18c.pdf.

Anonymous:1885:HMSb

[6457] Anonymous. Henry Morton Stanley. Scientific American, 52(2):19, January 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n2/pdf/scientificamerican01101885-19.pdf.

Anonymous:1885:SAG

[6458] Anonymous. Strengthening the abutment of a great bridge. Scientific American, 52(2):20–22, January 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v52/n2/pdf/scientificamerican01101885-
20.pdf.

Anonymous:1885:GAa

[6459] Anonymous. Goupil’s aeroplane. Scientific American, 52(2):22, January 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n2/pdf/scientificamerican01101885-22.pdf.

Anonymous:1885:OTF

[6460] Anonymous. A one thousand foot tower. Scientific American, 52(2):23, January 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n2/pdf/scientificamerican01101885-
23.pdf.

Anonymous:1885:MMC

[6461] Anonymous. Machine for making compressed yeast. Scientific American, 52(2):23, January 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n2/pdf/scientificamerican01101885-
23a.pdf.

Anonymous:1885:EIk

[6462] Anonymous. Engineering inventions. Scientific American, 52(2): 27, January 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n2/pdf/scientificamerican01101885-
27.pdf.

Anonymous:1885:MIo

[6463] Anonymous. Mechanical inventions. Scientific American, 52(2):27, January 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n2/pdf/scientificamerican01101885-27a.pdf.

Anonymous:1885:AIk

[6464] Anonymous. Agricultural inventions. Scientific American, 52 (2):27, January 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n2/pdf/scientificamerican01101885-
27b.pdf.

Anonymous:1885:MIp

[6465] Anonymous. Miscellaneous inventions. Scientific American, 52 (2):27, January 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n2/pdf/scientificamerican01101885-
27c.pdf.

Anonymous:1885:BPm

[6466] Anonymous. Business and personal. Scientific American, 52(20):313, May 16, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n20/pdf/scientificamerican05161885-313c.pdf.

Anonymous:1885:BNBg

[6467] Anonymous. Books: New books and publications. Scientific American, 52(20):314, May 16, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n20/pdf/scientificamerican05161885-
314.pdf.

Anonymous:1885:NQm

[6468] Anonymous. Notes and queries. Scientific American, 52(20):314–315, May 16, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n20/pdf/scientificamerican05161885-314a.pdf.

Anonymous:1885:IIm

[6469] Anonymous. Index of inventions. Scientific American, 52(20):315–316, May 16, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n20/pdf/scientificamerican05161885-315.pdf.

Anonymous:1885:PNPm

[6470] Anonymous. Patent notes: Patents. Scientific American, 52(20):318, May 16, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n20/pdf/scientificamerican05161885-318.pdf.

Anonymous:1885:SPL

[6471] Anonymous. Standard passenger locomotive. Scientific American, 52(20):303, May 16, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n20/pdf/scientificamerican05161885-
303.pdf.

Anonymous:1885:TSO

[6472] Anonymous. Torpedo swimmers — old and new. Scientific American, 52(20):304, May 16, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n20/pdf/scientificamerican05161885-
304.pdf.

Anonymous:1885:RMa

[6473] Anonymous. Relative measurements. Scientific American, 52(20):304, May 16, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n20/pdf/scientificamerican05161885-304a.pdf.

Anonymous:1885:EA

[6474] Anonymous. Elecampane as an antiseptic. Scientific American, 52(20):305, May 16, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n20/pdf/scientificamerican05161885-
305.pdf.

Anonymous:1885:FAC

[6475] Anonymous. Facts about cholera. Scientific American, 52(20):305, May 16, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n20/pdf/scientificamerican05161885-305a.pdf.

Anonymous:1885:SCFa

[6476] Anonymous. Sash cord fastener. Scientific American, 52(20):306, May 16, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n20/pdf/scientificamerican05161885-306.pdf.

Anonymous:1885:CSa

[6477] Anonymous. Carpet stretcher. Scientific American, 52(20):306, May 16, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n20/pdf/scientificamerican05161885-306a.pdf.

Anonymous:1885:CRG

[6478] Anonymous. Cable railway grip. Scientific American, 52(20):306, May 16, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n20/pdf/scientificamerican05161885-306b.pdf.

Anonymous:1885:WFT

[6479] Anonymous. Wire fence tool. Scientific American, 52(20):306, May 16, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n20/pdf/scientificamerican05161885-306c.pdf.

Anonymous:1885:CRA

[6480] Anonymous. Course reading attachments for transits. Scientific American, 52(20):307, May 16, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n20/pdf/scientificamerican05161885-
307.pdf.

Anonymous:1885:TEA

[6481] Anonymous. Triple effect apparatus. Scientific American, 52(20):307, May 16, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n20/pdf/scientificamerican05161885-307a.pdf.

Anonymous:1885:RSA

[6482] Anonymous. Railroad signaling apparatus. Scientific American, 52(20):307–308, May 16, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n20/pdf/scientificamerican05161885-
307b.pdf.

Anonymous:1885:Ch

[6483] Anonymous. The cyclostyle. Scientific American, 52(20):308–310, May 16, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n20/pdf/scientificamerican05161885-308.pdf.

Anonymous:1885:NCS

[6484] Anonymous. The new Cunard steamship Etruria. Scientific American, 52(20):310, May 16, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n20/pdf/scientificamerican05161885-
310.pdf.

Anonymous:1885:IGE

[6485] Anonymous. An improved gas engine. Scientific American, 52(20):310, May 16, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n20/pdf/scientificamerican05161885-310a.pdf.

Anonymous:1885:WFC

[6486] Anonymous. Whale fishery off the coast of Norway. Scientific American, 52(20):311–312, May 16, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n20/pdf/scientificamerican05161885-
311.pdf.

Anonymous:1885:EIl

[6487] Anonymous. Engineering inventions. Scientific American, 52(20):313, May 16, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n20/pdf/scientificamerican05161885-313.pdf.

Anonymous:1885:AIl

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n20/pdf/scientificamerican05161885-313a.pdf.

Anonymous:1885:MIq

[6489] Anonymous. Miscellaneous inventions. Scientific American, 52(20):313, May 16, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n20/pdf/scientificamerican05161885-313b.pdf.

Anonymous:1885:Ci

[6490] Anonymous. Correspondence. Scientific American, 52(21):325, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n21/pdf/scientificamerican05231885-325b.pdf.

Anonymous:1885:BNBh

[6491] Anonymous. Books: New books and publications. Scientific American, 52(21):329, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n21/pdf/scientificamerican05231885-
329c.pdf.

Anonymous:1885:BPn

[6492] Anonymous. Business and personal. Scientific American, 52(21):329, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n21/pdf/scientificamerican05231885-329d.pdf.

Anonymous:1885:NQn

[6493] Anonymous. Notes and queries. Scientific American, 52(21):330–331, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n21/pdf/scientificamerican05231885-330.pdf.

Anonymous:1885:IIn

[6494] Anonymous. Index of inventions. Scientific American, 52(21):331–332, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n21/pdf/scientificamerican05231885-331.pdf.

Anonymous:1885:PNPn

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n21/pdf/scientificamerican05231885-334.pdf.

Anonymous:1885:CAS

[6496] Anonymous. Compressed air for small motors. Scientific American, 52(21):320, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n21/pdf/scientificamerican05231885-
320.pdf.

Anonymous:1885:GCI

[6497] Anonymous. Grain in cast iron. Scientific American, 52(21):320, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n21/pdf/scientificamerican05231885-320a.pdf.

Anonymous:1885:TM

[6498] Anonymous. The teeth of mills. Scientific American, 52(21):320, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n21/pdf/scientificamerican05231885-320b.pdf.

Anonymous:1885:EAL

[6499] Anonymous. Expected advent of the locust. Scientific American, 52(21):320–321, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n21/pdf/scientificamerican05231885-
320c.pdf.

Anonymous:1885:AF

[6500] Anonymous. Automatic fan. Scientific American, 52(21):322, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n21/pdf/scientificamerican05231885-322.pdf.

Anonymous:1885:ADG

[6501] Anonymous. Automatic drawbridge gate. Scientific American, 52 (21):322, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n21/pdf/scientificamerican05231885-
322a.pdf.

Anonymous:1885:MRTb

[6502] Anonymous. Metallic railroad tie. Scientific American, 52(21):322, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n21/pdf/scientificamerican05231885-322b.pdf.

Anonymous:1885:NSH

[6503] Anonymous. A novel soap holder. Scientific American, 52(21):322, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n21/pdf/scientificamerican05231885-322c.pdf.

Anonymous:1885:ISHa

[6504] Anonymous. Improved steam hammer. Scientific American, 52(21):323, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n21/pdf/scientificamerican05231885-323.pdf.

Anonymous:1885:MSS

[6505] Anonymous. A mammoth steam snow shovel. Scientific American, 52(21):323–324, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n21/pdf/scientificamerican05231885-
323a.pdf.

Anonymous:1885:PE

[6506] Anonymous. Pendulum escapement. Scientific American, 52(21):324, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n21/pdf/scientificamerican05231885-324.pdf.

Anonymous:1885:UBF

[6507] Anonymous. Unexplored branch of the fire escape problem. Scientific American, 52(21):325, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n21/pdf/scientificamerican05231885-
325.pdf.

Anonymous:1885:BFa

[6508] Anonymous. Belt fastener. Scientific American, 52(21):325, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n21/pdf/scientificamerican05231885-325a.pdf.

Anonymous:1885:LSM

[6509] Anonymous. Labor-saving machines for restaurants. Scientific American, 52(21):326, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n21/pdf/scientificamerican05231885-
326.pdf.

Anonymous:1885:BEF

[6510] Anonymous. Boiler explosion at Franklin, Ind. Scientific American, 52(21):327, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n21/pdf/scientificamerican05231885-
327.pdf.

Anonymous:1885:TFR

[6511] Anonymous. Tarpon fishing with rod and reel. Scientific American, 52(21):327–328, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n21/pdf/scientificamerican05231885-
327a.pdf.

Anonymous:1885:NBR

[6512] Anonymous. A new bag rack. Scientific American, 52(21):328, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n21/pdf/scientificamerican05231885-328.pdf.

Anonymous:1885:EIm

[6513] Anonymous. Engineering inventions. Scientific American, 52(21):329, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n21/pdf/scientificamerican05231885-329.pdf.

Anonymous:1885:AIm

[6514] Anonymous. Agricultural inventions. Scientific American, 52(21):329, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n21/pdf/scientificamerican05231885-329a.pdf.

Anonymous:1885:MIr

[6515] Anonymous. Miscellaneous inventions. Scientific American, 52(21):329, May 23, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n21/pdf/scientificamerican05231885-329b.pdf.

Anonymous:1885:BNBi

[6516] Anonymous. Books: New books and publications. Scientific American, 52(22):345, May 30, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n22/pdf/scientificamerican05301885-
345d.pdf.

Anonymous:1885:BPo

[6517] Anonymous. Business and personal. Scientific American, 52(22):346, May 30, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n22/pdf/scientificamerican05301885-346.pdf.

Anonymous:1885:NQo

[6518] Anonymous. Notes and queries. Scientific American, 52(22):346–347, May 30, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n22/pdf/scientificamerican05301885-346a.pdf.

Anonymous:1885:IIo

[6519] Anonymous. Index of inventions. Scientific American, 52(22):347–348, May 30, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n22/pdf/scientificamerican05301885-347.pdf.

Anonymous:1885:PNPo

[6520] Anonymous. Patent notes: Patents. Scientific American, 52(22):350, May 30, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n22/pdf/scientificamerican05301885-350.pdf.

Anonymous:1885:PBS

[6521] Anonymous. Proposed bridge over the St. Lawrence River. Scientific American, 52(22):335, May 30, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n22/pdf/scientificamerican05301885-
335.pdf.

Anonymous:1885:HOM

[6522] Anonymous. The higher organization of manufactories. Scientific American, 52(22):336, May 30, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n22/pdf/scientificamerican05301885-
336.pdf.

Anonymous:1885:APJa

[6523] Anonymous. Aspects of the planets for June. Scientific American, 52(22):337, May 30, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n22/pdf/scientificamerican05301885-
337.pdf.

Anonymous:1885:NIF

[6524] Anonymous. A new and improved filter. Scientific American, 52 (22):338, May 30, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n22/pdf/scientificamerican05301885-
338.pdf.

Anonymous:1885:IBP

[6525] Anonymous. Improved baling press. Scientific American, 52(22):338, May 30, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n22/pdf/scientificamerican05301885-338a.pdf.

Anonymous:1885:WCS

[6526] Anonymous. Wire carrier and stretcher. Scientific American, 52 (22):338, May 30, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n22/pdf/scientificamerican05301885-
338b.pdf.

Anonymous:1885:EHL

[6527] Anonymous. Electric high and low water detecter. Scientific American, 52(22):338, May 30, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n22/pdf/scientificamerican05301885-
338c.pdf.

Anonymous:1885:TNA

[6528] Anonymous. The Taylor new automatic stationary engine. Scientific American, 52(22):339, May 30, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n22/pdf/scientificamerican05301885-
339.pdf.

Anonymous:1885:IP

[6529] Anonymous. Improved plow. Scientific American, 52(22):339, May 30, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n22/pdf/scientificamerican05301885-339a.pdf.

Anonymous:1885:DCH

[6530] Anonymous. Door check or holder. Scientific American, 52(22):341, May 30, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n22/pdf/scientificamerican05301885-341.pdf.

Anonymous:1885:GSF

[6531] Anonymous. Glove or shoe fastener. Scientific American, 52(22):341, May 30, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n22/pdf/scientificamerican05301885-341a.pdf.

Anonymous:1885:MPE

[6532] Anonymous. Mather’s portable engine. Scientific American, 52(22):342, May 30, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n22/pdf/scientificamerican05301885-342.pdf.

Anonymous:1885:MBC

[6533] Anonymous. Mechanical boiler cleaner and water purifier. Scientific American, 52(22):342, May 30, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n22/pdf/scientificamerican05301885-
342a.pdf.

Anonymous:1885:AL

[6534] Anonymous. The art of leaping. Scientific American, 52(22):343–344, May 30, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n22/pdf/scientificamerican05301885-343.pdf.

Anonymous:1885:EIn

[6535] Anonymous. Engineering inventions. Scientific American, 52(22):345, May 30, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n22/pdf/scientificamerican05301885-345.pdf.

Anonymous:1885:MIs

[6536] Anonymous. Mechanical inventions. Scientific American, 52(22):345, May 30, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n22/pdf/scientificamerican05301885-345a.pdf.

Anonymous:1885:AIo

[6537] Anonymous. Agricultural inventions. Scientific American, 52(22):345, May 30, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n22/pdf/scientificamerican05301885-345b.pdf.

Anonymous:1885:MIt

[6538] Anonymous. Miscellaneous inventions. Scientific American, 52(22):345, May 30, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n22/pdf/scientificamerican05301885-345c.pdf.

Anonymous:1885:Cj

[6539] Anonymous. Correspondence. Scientific American, 52(23):357, June 6, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n23/pdf/scientificamerican06061885-357.pdf.

Anonymous:1885:BPp

[6540] Anonymous. Business and personal. Scientific American, 52(23):362, June 6, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n23/pdf/scientificamerican06061885-362c.pdf.

Anonymous:1885:NQp

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n23/pdf/scientificamerican06061885-362d.pdf.

Anonymous:1885:IIp

[6542] Anonymous. Index of inventions. Scientific American, 52(23):363–364, June 6, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n23/pdf/scientificamerican06061885-363.pdf.

Anonymous:1885:PNPp

[6543] Anonymous. Patent notes: Patents. Scientific American, 52(23):366, June 6, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n23/pdf/scientificamerican06061885-366.pdf.

Anonymous:1885:NWS

[6544] Anonymous. New war ships of the British Navy. Scientific American, 52(23):351, June 6, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n23/pdf/scientificamerican06061885-
351.pdf.

Anonymous:1885:CSF

[6545] Anonymous. Are costly ships and forts necessary to an effective defense? Scientific American, 52(23):352–353, June 6, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v52/
n23/pdf/scientificamerican06061885-352.pdf.

Anonymous:1885:NFP

[6546] Anonymous. The Niagara Falls Park. Scientific American, 52(23):353, June 6, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n23/pdf/scientificamerican06061885-353.pdf.

Anonymous:1885:CDa

[6547] Anonymous. Comets due in 1885. Scientific American, 52(23):353, June 6, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n23/pdf/scientificamerican06061885-353a.pdf.

Anonymous:1885:PM

[6548] Anonymous. The practical mechanic. Scientific American, 52(23):353, June 6, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n23/pdf/scientificamerican06061885-353b.pdf.

Anonymous:1885:MRTc

[6549] Anonymous. Metallic railway tie. Scientific American, 52(23):354, June 6, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n23/pdf/scientificamerican06061885-354.pdf.

Anonymous:1885:ICGa

[6550] Anonymous. An improved cornice gutter. Scientific American, 52(23):354, June 6, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n23/pdf/scientificamerican06061885-
354a.pdf.

Anonymous:1885:IFa

[6551] Anonymous. Improved fence. Scientific American, 52(23):354, June 6, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n23/pdf/scientificamerican06061885-354b.pdf.

Anonymous:1885:CCd

[6552] Anonymous. Car coupling. Scientific American, 52(23):354–355, June 6, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n23/pdf/scientificamerican06061885-354c.pdf.

Anonymous:1885:MSP

[6553] Anonymous. Machine for shelling peas. Scientific American, 52(23):355, June 6, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n23/pdf/scientificamerican06061885-355.pdf.

Anonymous:1885:CBSa

[6554] Anonymous. Combined bee swarmer and hiver. Scientific American, 52(23):355–356, June 6, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n23/pdf/scientificamerican06061885-
355a.pdf.

Anonymous:1885:REN

[6555] Anonymous. Rack for exhibiting notions. Scientific American, 52(23):356, June 6, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n23/pdf/scientificamerican06061885-
356.pdf.

Anonymous:1885:CIT

[6556] Anonymous. Cox’s improved table knives. Scientific American, 52(23):356, June 6, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n23/pdf/scientificamerican06061885-
356a.pdf.

Anonymous:1885:BHH

[6557] Anonymous. Button hook holder for shoes. Scientific American, 52(23):356, June 6, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n23/pdf/scientificamerican06061885-
356b.pdf.

Anonymous:1885:IICa

[6558] Anonymous. Improved ice creepers. Scientific American, 52(23):356, June 6, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n23/pdf/scientificamerican06061885-356c.pdf.

Anonymous:1885:PWC

[6559] Anonymous. Protector for water conductors. Scientific American, 52(23):357, June 6, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n23/pdf/scientificamerican06061885-
357a.pdf.

Anonymous:1885:DRPb

[6560] Anonymous. Decisions relating to patents. Scientific American, 52(23):357–358, June 6, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n23/pdf/scientificamerican06061885-
357b.pdf.

Anonymous:1885:PSC

[6561] Anonymous. Power shears for cutting sheet metal. Scientific American, 52(23):358, June 6, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n23/pdf/scientificamerican06061885-
358.pdf.

Anonymous:1885:WNM

[6562] Anonymous. The white-nosed monkey at the Berlin Aquarium. Scientific American, 52(23):359–361, June 6, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v52/
n23/pdf/scientificamerican06061885-359.pdf.

Anonymous:1885:EIo

[6563] Anonymous. Engineering inventions. Scientific American, 52(23):362, June 6, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n23/pdf/scientificamerican06061885-362.pdf.

Anonymous:1885:AIp

[6564] Anonymous. Agricultural inventions. Scientific American, 52(23):362, June 6, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n23/pdf/scientificamerican06061885-362a.pdf.

Anonymous:1885:MIu

[6565] Anonymous. Miscellaneous inventions. Scientific American, 52(23):362, June 6, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n23/pdf/scientificamerican06061885-362b.pdf.

Anonymous:1885:Ck

[6566] Anonymous. Correspondence. Scientific American, 52(25):389–390, June 20, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n25/pdf/scientificamerican06201885-389.pdf.

Anonymous:1885:BNBj

[6567] Anonymous. Books: New books and publications. Scientific American, 52(25):393, June 20, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n25/pdf/scientificamerican06201885-
393d.pdf.

Anonymous:1885:BPq

[6568] Anonymous. Business and personal. Scientific American, 52(25):394, June 20, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n25/pdf/scientificamerican06201885-394.pdf.

Anonymous:1885:NQq

[6569] Anonymous. Notes and queries. Scientific American, 52(25):394–395, June 20, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n25/pdf/scientificamerican06201885-394a.pdf.

Anonymous:1885:IIq

[6570] Anonymous. Index of inventions. Scientific American, 52(25):395–396, June 20, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n25/pdf/scientificamerican06201885-395.pdf.

Anonymous:1885:PNPq

[6571] Anonymous. Patent notes: Patents. Scientific American, 52(25):398, June 20, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n25/pdf/scientificamerican06201885-398.pdf.

Anonymous:1885:FRSa

[6572] Anonymous. The fastest river steamer beaten by a steam yacht. Scientific American, 52(25):383–384, June 20, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v52/n25/pdf/scientificamerican06201885-
383.pdf.

Anonymous:1885:ECG

[6573] Anonymous. The English cutter Genesta. Scientific American, 52(25):384, June 20, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n25/pdf/scientificamerican06201885-
384.pdf.

Anonymous:1885:NUC

[6574] Anonymous. The new U.S. cruiser Dolphin. Scientific American, 52(25):384–385, June 20, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n25/pdf/scientificamerican06201885-
384a.pdf.

Anonymous:1885:IDF

[6575] Anonymous. Improved double fireplace. Scientific American, 52 (25):386, June 20, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n25/pdf/scientificamerican06201885-
386.pdf.

Anonymous:1885:AGC

[6576] Anonymous. Attachment for gas cocks. Scientific American, 52(25):386, June 20, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n25/pdf/scientificamerican06201885-386a.pdf.

Anonymous:1885:CCe

[6577] Anonymous. Car coupling. Scientific American, 52(25):386, June 20, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n25/pdf/scientificamerican06201885-386b.pdf.

Anonymous:1885:IBSa

[6578] Anonymous. Improved brake shoe. Scientific American, 52(25):386, June 20, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n25/pdf/scientificamerican06201885-386c.pdf.

Anonymous:1885:LVH

[6579] Anonymous. The laws of vision and the harmony of colors. Scientific American, 52(25):387–388, June 20, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v52/n25/pdf/scientificamerican06201885-
387.pdf.

Anonymous:1885:UST

[6580] Anonymous. The “Unique” single tube injector. Scientific American, 52(25):388, June 20, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n25/pdf/scientificamerican06201885-
388.pdf.

Anonymous:1885:ITEa

[6581] Anonymous. Improved traction engine. Scientific American, 52(25):390, June 20, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n25/pdf/scientificamerican06201885-390.pdf.

Anonymous:1885:SVA

[6582] Anonymous. Safety valve and alarm for steam boilers. Scientific American, 52(25):390, June 20, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n25/pdf/scientificamerican06201885-
390a.pdf.

Sloane:1885:ESB

[6583] T. O’conor Sloane. Experiments with soap bubbles and films. Scientific American, 52(25):391–392, June 20, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v52/n25/pdf/scientificamerican06201885-
391.pdf.

Anonymous:1885:EIp

[6584] Anonymous. Engineering inventions. Scientific American, 52(25):393, June 20, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n25/pdf/scientificamerican06201885-393.pdf.

Anonymous:1885:MIv

[6585] Anonymous. Mechanical inventions. Scientific American, 52(25):393, June 20, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n25/pdf/scientificamerican06201885-393a.pdf.

Anonymous:1885:AIq

[6586] Anonymous. Agricultural inventions. Scientific American, 52(25):393, June 20, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n25/pdf/scientificamerican06201885-393b.pdf.

Anonymous:1885:MIw

[6587] Anonymous. Miscellaneous inventions. Scientific American, 52(25):393, June 20, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n25/pdf/scientificamerican06201885-393c.pdf.

Anonymous:1885:BPr

[6588] Anonymous. Business and personal. Scientific American, 52(26):408, June 27, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n26/pdf/scientificamerican06271885-408d.pdf.

Anonymous:1885:NQr

[6589] Anonymous. Notes and queries. Scientific American, 52(26):408–409, June 27, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n26/pdf/scientificamerican06271885-408e.pdf.

Anonymous:1885:IIr

[6590] Anonymous. Index of inventions. Scientific American, 52(26):409–410, June 27, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n26/pdf/scientificamerican06271885-409.pdf.

Anonymous:1885:I

[6591] Anonymous. Index. Scientific American, 52(26):411–412, June 27, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n26/pdf/scientificamerican06271885-411.pdf.

Anonymous:1885:PNPr

[6592] Anonymous. Patent notes: Patents. Scientific American, 52(26):414, June 27, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n26/pdf/scientificamerican06271885-414.pdf.

Anonymous:1885:TPS

[6593] Anonymous. The thermo phote, or self-registering photometer. Scientific American, 52(26):399, June 27, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v52/
n26/pdf/scientificamerican06271885-399.pdf.

Anonymous:1885:ASL

[6594] Anonymous. Arrival of the Statue of Liberty. Scientific American, 52(26):400, June 27, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n26/pdf/scientificamerican06271885-
400.pdf.

Anonymous:1885:JNO

[6595] Anonymous. Japan at the New Orleans Exposition. Scientific American, 52(26):400–401, June 27, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n26/pdf/scientificamerican06271885-
400a.pdf.

Anonymous:1885:PAS

[6596] Anonymous. Pitman attachment for sewing machines. Scientific American, 52(26):402, June 27, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n26/pdf/scientificamerican06271885-
402.pdf.

Anonymous:1885:EFS

[6597] Anonymous. An electric “front-sight” for firearms. Scientific American, 52(26):402, June 27, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n26/pdf/scientificamerican06271885-
402a.pdf.

Anonymous:1885:SDS

[6598] Anonymous. School desk and seat. Scientific American, 52(26):402, June 27, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n26/pdf/scientificamerican06271885-402b.pdf.

Anonymous:1885:RTM

[6599] Anonymous. Railway tamping machine. Scientific American, 52 (26):402, June 27, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n26/pdf/scientificamerican06271885-
402c.pdf.

Anonymous:1885:IUG

[6600] Anonymous. Improvements in upright grand pianos. Scientific American, 52(26):403–404, June 27, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n26/pdf/scientificamerican06271885-
403.pdf.

Anonymous:1885:SPB

[6601] Anonymous. Simple push button for electric bells. Scientific American, 52(26):404–405, June 27, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n26/pdf/scientificamerican06271885-
404.pdf.

Anonymous:1885:EBS

[6602] Anonymous. The engine and boiler of the Stiletto. Scientific American, 52(26):406, June 27, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n26/pdf/scientificamerican06271885-
406.pdf.

Anonymous:1885:NTP

[6603] Anonymous. The needle and thread plant. Scientific American, 52(26):407, June 27, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n26/pdf/scientificamerican06271885-
407.pdf.

Anonymous:1885:EIq

[6604] Anonymous. Engineering inventions. Scientific American, 52(26):408, June 27, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n26/pdf/scientificamerican06271885-408.pdf.

Anonymous:1885:MIx

[6605] Anonymous. Mechanical inventions. Scientific American, 52(26):408, June 27, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n26/pdf/scientificamerican06271885-408a.pdf.

Anonymous:1885:AIr

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n26/pdf/scientificamerican06271885-408b.pdf.

Anonymous:1885:MIy

[6607] Anonymous. Miscellaneous inventions. Scientific American, 52(26):408, June 27, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n26/pdf/scientificamerican06271885-408c.pdf.

Anonymous:1885:Cl

[6608] Anonymous. Correspondence. Scientific American, 52(3):37, January 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n3/pdf/scientificamerican01171885-37.pdf.

Anonymous:1885:BNBk

[6609] Anonymous. Books: New books and publications. Scientific American, 52(3):42, January 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n3/pdf/scientificamerican01171885-
42c.pdf.

Anonymous:1885:BPt

[6610] Anonymous. Business and personal. Scientific American, 52(3):42, January 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n3/pdf/scientificamerican01171885-42d.pdf.

Anonymous:1885:NQs

[6611] Anonymous. Notes and queries. Scientific American, 52(3):42–43, January 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n3/pdf/scientificamerican01171885-42e.pdf.

Anonymous:1885:IIs

[6612] Anonymous. Index of inventions. Scientific American, 52(3):43–44, January 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n3/pdf/scientificamerican01171885-43.pdf.

Anonymous:1885:PNPs

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n3/pdf/scientificamerican01171885-45.pdf.

Anonymous:1885:OTR

[6614] Anonymous. Oil of turpentine and resin. Scientific American, 52(3):31, January 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n3/pdf/scientificamerican01171885-
31.pdf.

Anonymous:1885:ITG

[6615] Anonymous. Improved tool grinder and press. Scientific American, 52(3):31, January 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n3/pdf/scientificamerican01171885-
31a.pdf.

Anonymous:1885:SMJ

[6616] Anonymous. Success of the Mississippi jetties. Scientific American, 52(3):32–34, January 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n3/pdf/scientificamerican01171885-
32.pdf.

Anonymous:1885:IPW

[6617] Anonymous. Improved pipe wrench. Scientific American, 52(3):34, January 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n3/pdf/scientificamerican01171885-34.pdf.

Anonymous:1885:DSa

[6618] Anonymous. Door spring. Scientific American, 52(3):34, January 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n3/pdf/scientificamerican01171885-34a.pdf.

Anonymous:1885:NRF

[6619] Anonymous. A novel rail fastening. Scientific American, 52 (3):34–35, January 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n3/pdf/scientificamerican01171885-
34b.pdf.

Anonymous:1885:IPRa

[6620] Anonymous. Ice planing and ridging machine. Scientific American, 52(3):35, January 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n3/pdf/scientificamerican01171885-
35.pdf.

Anonymous:1885:VB

[6621] Anonymous. Vehicle brake. Scientific American, 52(3):35, January 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n3/pdf/scientificamerican01171885-35a.pdf.

Anonymous:1885:BB

[6622] Anonymous. Back band. Scientific American, 52(3):36, January 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n3/pdf/scientificamerican01171885-36.pdf.

Anonymous:1885:NRB

[6623] Anonymous. New roadway bridge. Scientific American, 52(3):38, January 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n3/pdf/scientificamerican01171885-38.pdf.

Anonymous:1885:SBL

[6624] Anonymous. Steam bell for locomotives. Scientific American, 52(3):38, January 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n3/pdf/scientificamerican01171885-
38a.pdf.

Anonymous:1885:SNA

[6625] Anonymous. A study of the Navies of Antiquity. Scientific American, 52(3):39–41, January 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n3/pdf/scientificamerican01171885-
39.pdf.

Anonymous:1885:EIs (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n3/pdf/scientificamerican01171885-
42.pdf.

Anonymous:1885:AIs

[6627] Anonymous. Agricultural inventions. Scientific American, 52 (3):42, January 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n3/pdf/scientificamerican01171885-
42a.pdf.

Anonymous:1885:MIz

[6628] Anonymous. Miscellaneous inventions. Scientific American, 52 (3):42, January 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n3/pdf/scientificamerican01171885-
42b.pdf.

Anonymous:1885:Cm

[6629] Anonymous. Correspondence. Scientific American, 52(4):53–54, January 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n4/pdf/scientificamerican01241885-53.pdf.

Anonymous:1885:BNBl

[6630] Anonymous. Books: New books and publications. Scientific American, 52(4):58, January 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n4/pdf/scientificamerican01241885-
58c.pdf.

Anonymous:1885:BPu

[6631] Anonymous. Business and personal. Scientific American, 52(4):58, January 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n4/pdf/scientificamerican01241885-58d.pdf.

Anonymous:1885:NQt

[6632] Anonymous. Notes and queries. Scientific American, 52(4):58–59, January 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n4/pdf/scientificamerican01241885-58e.pdf.

Anonymous:1885:IIt

[6633] Anonymous. Index of inventions. Scientific American, 52(4):59–60, January 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n4/pdf/scientificamerican01241885-59.pdf.

Anonymous:1885:PNPt

[6634] Anonymous. Patent notes: Patents. Scientific American, 52(4):61, January 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n4/pdf/scientificamerican01241885-61.pdf.

Anonymous:1885:BMF

[6635] Anonymous. Bridge with movable floor system. Scientific American, 52(4):47, January 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n4/pdf/scientificamerican01241885-
47.pdf.

Anonymous:1885:NOEd

[6636] Anonymous. The New Orleans Exposition. Scientific American, 52(4):48, January 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n4/pdf/scientificamerican01241885-
48.pdf.

Anonymous:1885:TF

[6637] Anonymous. The taper fit. Scientific American, 52(4):48, January 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n4/pdf/scientificamerican01241885-48a.pdf.

Anonymous:1885:WCa

[6638] Anonymous. Wood carving. Scientific American, 52(4):48, January 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n4/pdf/scientificamerican01241885-48b.pdf.

Anonymous:1885:MCb

[6639] Anonymous. Making cherries. Scientific American, 52(4):48–49, January 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n4/pdf/scientificamerican01241885-48c.pdf.

Anonymous:1885:APF

[6640] Anonymous. Aspects of the planets for February. Scientific American, 52(4):49, January 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n4/pdf/scientificamerican01241885-
49.pdf.

Anonymous:1885:ISLb

[6641] Anonymous. Improved seat lock. Scientific American, 52(4):50, January 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n4/pdf/scientificamerican01241885-50.pdf.

Anonymous:1885:DSC

[6642] Anonymous. Design for spring clip or tongs. Scientific American, 52(4):50, January 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n4/pdf/scientificamerican01241885-
50a.pdf.

Anonymous:1885:CBM

[6643] Anonymous. Check billiard marker. Scientific American, 52(4):50, January 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n4/pdf/scientificamerican01241885-50b.pdf.

Anonymous:1885:EH

[6644] Anonymous. Egg holder. Scientific American, 52(4):51, January 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n4/pdf/scientificamerican01241885-51.pdf.

Anonymous:1885:ASR

[6645] Anonymous. Automatic signal for railroad crossings. Scientific American, 52(4):51, January 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n4/pdf/scientificamerican01241885-
51a.pdf.

Anonymous:1885:NH

[6646] Anonymous. A novel horseshoe. Scientific American, 52(4):51–52, January 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n4/pdf/scientificamerican01241885-51b.pdf.

Anonymous:1885:Hb

[6647] Anonymous. Hat holder. Scientific American, 52(4):52, January 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n4/pdf/scientificamerican01241885-52.pdf.

Anonymous:1885:ID

[6648] Anonymous. Improved dredger. Scientific American, 52(4):54, January 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n4/pdf/scientificamerican01241885-54.pdf.

Anonymous:1885:MAD

[6649] Anonymous. Magical apparition of a drawing on white paper. Scientific American, 52(4):54, January 24, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v52/
n4/pdf/scientificamerican01241885-54a.pdf.

vonMwell:1885:RT

[6650] F. von Mwell. Rhododendron Toverenæ. Scientific American, 52 (4):55–56, January 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n4/pdf/scientificamerican01241885-
55.pdf.

Anonymous:1885:MFT

[6651] Anonymous. Machine for framing timber. Scientific American, 52(4):56–57, January 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n4/pdf/scientificamerican01241885-
56.pdf.

Anonymous:1885:EIt (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n4/pdf/scientificamerican01241885-
58.pdf.

Anonymous:1885:AIt

[6653] Anonymous. Agricultural inventions. Scientific American, 52 (4):58, January 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n4/pdf/scientificamerican01241885-
58a.pdf.

Anonymous:1885:MIba

[6654] Anonymous. Miscellaneous inventions. Scientific American, 52 (4):58, January 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n4/pdf/scientificamerican01241885-
58b.pdf.

Anonymous:1885:Cn

[6655] Anonymous. Correspondence. Scientific American, 52(5):69–70, January 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n5/pdf/scientificamerican01311885-69.pdf.

Anonymous:1885:BPv

[6656] Anonymous. Business and personal. Scientific American, 52(5):73, January 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n5/pdf/scientificamerican01311885-73c.pdf.

Anonymous:1885:NQu

[6657] Anonymous. Notes and queries. Scientific American, 52(5):73–75, January 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n5/pdf/scientificamerican01311885-73d.pdf.

Anonymous:1885:IIu

[6658] Anonymous. Index of inventions. Scientific American, 52(5):75–76, January 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n5/pdf/scientificamerican01311885-75.pdf.

Anonymous:1885:PNPu

[6659] Anonymous. Patent notes: Patents. Scientific American, 52(5):77, January 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n5/pdf/scientificamerican01311885-77.pdf.

Anonymous:1885:MAF

[6660] Anonymous. The manufacture of agglomerated fuel. Scientific American, 52(5):63–64, January 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n5/pdf/scientificamerican01311885-
63.pdf.

Anonymous:1885:IMb

[6661] Anonymous. Improved methods. Scientific American, 52(5):64, January 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n5/pdf/scientificamerican01311885-64.pdf.

Anonymous:1885:DSb

[6662] Anonymous. Dropping and striking up. Scientific American, 52 (5):64–65, January 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n5/pdf/scientificamerican01311885-
64a.pdf.

Anonymous:1885:BT

[6663] Anonymous. Butter tub. Scientific American, 52(5):66, January 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n5/pdf/scientificamerican01311885-66.pdf.

Anonymous:1885:BFb

[6664] Anonymous. Belt fastener. Scientific American, 52(5):66, January 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n5/pdf/scientificamerican01311885-66a.pdf.

Anonymous:1885:ISW

[6665] Anonymous. Improved shutter worker. Scientific American, 52 (5):66, January 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n5/pdf/scientificamerican01311885-
66b.pdf.

Anonymous:1885:BA

[6666] Anonymous. Burglar alarm. Scientific American, 52(5):66, January 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n5/pdf/scientificamerican01311885-66c.pdf.

Anonymous:1885:SL

[6667] Anonymous. Spirit level. Scientific American, 52(5):67, January 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n5/pdf/scientificamerican01311885-67.pdf.

Anonymous:1885:IQP

[6668] Anonymous. Improved quadruple pump. Scientific American, 52 (5):67, January 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n5/pdf/scientificamerican01311885-
67a.pdf.

Anonymous:1885:CGW

[6669] Anonymous. Combined guard and weeder. Scientific American, 52(5):67–68, January 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n5/pdf/scientificamerican01311885-
67b.pdf.

Anonymous:1885:DSBa

[6670] Anonymous. Detachable shirt bosom. Scientific American, 52 (5):68, January 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n5/pdf/scientificamerican01311885-
68.pdf.

Anonymous:1885:SMA

[6671] Anonymous. A simple mode of ascertaining the revolutions of a shaft. Scientific American, 52(5):70, January 31, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v52/
n5/pdf/scientificamerican01311885-70.pdf.

Anonymous:1885:WPMb

[6672] Anonymous. Weight power machine. Scientific American, 52(5):70, January 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n5/pdf/scientificamerican01311885-70a.pdf.

Anonymous:1885:MW

[6673] Anonymous. The Mason wasp. Scientific American, 52(5):71–72, January 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n5/pdf/scientificamerican01311885-71.pdf.

Anonymous:1885:EIu

[6674] Anonymous. Engineering inventions. Scientific American, 52(5): 73, January 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n5/pdf/scientificamerican01311885-
73.pdf.

Anonymous:1885:AIu

[6675] Anonymous. Agricultural inventions. Scientific American, 52 (5):73, January 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n5/pdf/scientificamerican01311885-
73a.pdf.

Anonymous:1885:MIbb

[6676] Anonymous. Miscellaneous inventions. Scientific American, 52 (5):73, January 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n5/pdf/scientificamerican01311885-
73b.pdf.

Anonymous:1885:BNBm

[6677] Anonymous. Books: New books and publications. Scientific American, 52(6):89, February 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n6/pdf/scientificamerican02071885-
89c.pdf.

Anonymous:1885:BPw (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n6/pdf/scientificamerican02071885-
89d.pdf.

Anonymous:1885:NQv

[6679] Anonymous. Notes and queries. Scientific American, 52(6):89– 91, February 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n6/pdf/scientificamerican02071885-
89e.pdf.

Anonymous:1885:IIv

[6680] Anonymous. Index of inventions. Scientific American, 52(6): 91–92, February 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n6/pdf/scientificamerican02071885-
91.pdf.

Anonymous:1885:PNPv

[6681] Anonymous. Patent notes: Patents. Scientific American, 52 (6):93, February 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n6/pdf/scientificamerican02071885-
93.pdf.

Anonymous:1885:AB

[6682] Anonymous. The Attock Bridge. Scientific American, 52(6): 79–80, February 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n6/pdf/scientificamerican02071885-
79.pdf.

Anonymous:1885:EC

[6683] Anonymous. Encke’s Comet. Scientific American, 52(6):80–81, February 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n6/pdf/scientificamerican02071885-80.pdf.

Anonymous:1885:SDa

[6684] Anonymous. Shop driving. Scientific American, 52(6):81, February 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n6/pdf/scientificamerican02071885-81.pdf.

Anonymous:1885:AMb

[6685] Anonymous. Asbestos in mechanics. Scientific American, 52 (6):81, February 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n6/pdf/scientificamerican02071885-
81a.pdf.

Anonymous:1885:BEP

[6686] Anonymous. Binding eyelet for papers, magazines, etc. Scientific American, 52(6):82, February 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n6/pdf/scientificamerican02071885-
82.pdf.

Anonymous:1885:ASC

[6687] Anonymous. Anchor stake for check row wires. Scientific American, 52(6):82, February 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n6/pdf/scientificamerican02071885-
82a.pdf.

Anonymous:1885:ICCc

[6688] Anonymous. Improved car coupling. Scientific American, 52 (6):82, February 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n6/pdf/scientificamerican02071885-
82b.pdf.

Anonymous:1885:PHA

[6689] Anonymous. Post hole auger. Scientific American, 52(6):82, February 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n6/pdf/scientificamerican02071885-82c.pdf.

Anonymous:1885:NEM

[6690] Anonymous. A new electro magnet. Scientific American, 52 (6):83, February 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n6/pdf/scientificamerican02071885-
83.pdf.

Anonymous:1885:AFE

[6691] Anonymous. Automatic fire escape. Scientific American, 52(6): 83, February 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n6/pdf/scientificamerican02071885-
83a.pdf.

Anonymous:1885:AW

[6692] Anonymous. An Aztec warrior. Scientific American, 52(6):83– 84, February 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n6/pdf/scientificamerican02071885-
83b.pdf.

Anonymous:1885:ISSc

[6693] Anonymous. Improved shaft support. Scientific American, 52 (6):84, February 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n6/pdf/scientificamerican02071885-
84.pdf.

Anonymous:1885:ADP

[6694] Anonymous. Automatic double-pointed pen holder. Scientific American, 52(6):84–85, February 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n6/pdf/scientificamerican02071885-
84a.pdf.

Anonymous:1885:RTP

[6695] Anonymous. Registering thermometers and pressure gauges. Scientific American, 52(6):86, February 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n6/pdf/scientificamerican02071885-
86.pdf.

Anonymous:1885:HAL

[6696] Anonymous. Hydraulic accumulators for loading and unloading cars. Scientific American, 52(6):86–88, February 7, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v52/
n6/pdf/scientificamerican02071885-86a.pdf.

Anonymous:1885:Hc

[6697] Anonymous. The hypnoscope. Scientific American, 52(6):88, February 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n6/pdf/scientificamerican02071885-88.pdf.

Anonymous:1885:EIv

[6698] Anonymous. Engineering inventions. Scientific American, 52(6): 89, February 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n6/pdf/scientificamerican02071885-
89.pdf.

Anonymous:1885:AIv

[6699] Anonymous. Agricultural inventions. Scientific American, 52 (6):89, February 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n6/pdf/scientificamerican02071885-
89a.pdf.

Anonymous:1885:MIbc

[6700] Anonymous. Miscellaneous inventions. Scientific American, 52 (6):89, February 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n6/pdf/scientificamerican02071885-
89b.pdf.

Anonymous:1885:Cp

[6701] Anonymous. Correspondence. Scientific American, 52(7):101– 102, February 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n7/pdf/scientificamerican02141885-
101.pdf.

Anonymous:1885:BPx

[6702] Anonymous. Business and personal. Scientific American, 52(7): 107, February 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n7/pdf/scientificamerican02141885-
107d.pdf.

Anonymous:1885:NQw

[6703] Anonymous. Notes and queries. Scientific American, 52(7):107, February 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n7/pdf/scientificamerican02141885-107e.pdf.

Anonymous:1885:IIw

[6704] Anonymous. Index of inventions. Scientific American, 52(7): 107–108, February 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n7/pdf/scientificamerican02141885-
107f.pdf.

Anonymous:1885:PNPw

[6705] Anonymous. Patent notes: Patents. Scientific American, 52 (7):109, February 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n7/pdf/scientificamerican02141885-
109.pdf.

Anonymous:1885:SR

[6706] Anonymous. Shop remedies. Scientific American, 52(7):96, February 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n7/pdf/scientificamerican02141885-96.pdf.

Anonymous:1885:CIF

[6707] Anonymous. Cast iron finish. Scientific American, 52(7):96, February 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n7/pdf/scientificamerican02141885-96a.pdf.

Anonymous:1885:ICb

[6708] Anonymous. Inside calipering. Scientific American, 52(7):96, February 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n7/pdf/scientificamerican02141885-96b.pdf.

Anonymous:1885:TS

[6709] Anonymous. Treatment of steel. Scientific American, 52(7):96– 97, February 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n7/pdf/scientificamerican02141885-
96c.pdf.

Anonymous:1885:IAb

[6710] Anonymous. Invention as an art. Scientific American, 52(7):97, February 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n7/pdf/scientificamerican02141885-97.pdf.

Anonymous:1885:ELE

[6711] Anonymous. Has electric light any effect on the growth of plants? Scientific American, 52(7):97–98, February 14, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v52/
n7/pdf/scientificamerican02141885-97a.pdf.

Anonymous:1885:ITa

[6712] Anonymous. Improved tricycle. Scientific American, 52(7):98, February 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n7/pdf/scientificamerican02141885-98.pdf.

Anonymous:1885:ISLc

[6713] Anonymous. Improved spirit level. Scientific American, 52(7):98, February 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n7/pdf/scientificamerican02141885-98a.pdf.

Anonymous:1885:IHa

[6714] Anonymous. Improved harrow. Scientific American, 52(7):98, February 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n7/pdf/scientificamerican02141885-98b.pdf.

Anonymous:1885:NTA

[6715] Anonymous. Navigable trains of air ships. Scientific American, 52(7):99, February 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n7/pdf/scientificamerican02141885-
99.pdf.

Anonymous:1885:RFG

[6716] Anonymous. A remarkable fungoid growth. Scientific American, 52(7):99, February 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n7/pdf/scientificamerican02141885-
99a.pdf.

Anonymous:1885:PSW

[6717] Anonymous. The pneumatic system of the Western Union. Scientific American, 52(7):100, February 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v52/n7/pdf/scientificamerican02141885-
100.pdf.

Anonymous:1885:IRM

[6718] Anonymous. Improved rail mill engines. Scientific American, 52 (7):102–103, February 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n7/pdf/scientificamerican02141885-
102.pdf.

Anonymous:1885:AV

[6719] Anonymous. An aquatic velocipede. Scientific American, 52(7): 103–104, February 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n7/pdf/scientificamerican02141885-
103.pdf.

Anonymous:1885:DRPc

[6720] Anonymous. Decisions relating to patents. Scientific American, 52 (7):104–106, February 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n7/pdf/scientificamerican02141885-
104.pdf.

Anonymous:1885:EIw

[6721] Anonymous. Engineering inventions. Scientific American, 52(7): 107, February 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n7/pdf/scientificamerican02141885-
107.pdf.

Anonymous:1885:AIw

[6722] Anonymous. Agricultural inventions. Scientific American, 52(7): 107, February 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n7/pdf/scientificamerican02141885-
107a.pdf.

Anonymous:1885:MIbd

[6723] Anonymous. Miscellaneous inventions. Scientific American, 52 (7):107, February 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n7/pdf/scientificamerican02141885-
107b.pdf.

Anonymous:1885:Sf

[6724] Anonymous. Special. Scientific American, 52(7):107, February 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n7/pdf/scientificamerican02141885-107c.pdf.

Anonymous:1885:Cq

[6725] Anonymous. Correspondence. Scientific American, 52(8):116, February 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n8/pdf/scientificamerican02211885-116b.pdf.

Anonymous:1885:BPy

[6726] Anonymous. Business and personal. Scientific American, 52(8): 122, February 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n8/pdf/scientificamerican02211885-
122.pdf.

Anonymous:1885:BNBn

[6727] Anonymous. Books: New books and publications. Scientific American, 52(8):122, February 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n8/pdf/scientificamerican02211885-
122a.pdf.

Anonymous:1885:NQx (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n8/pdf/scientificamerican02211885-
122b.pdf.

Anonymous:1885:IIx

[6729] Anonymous. Index of inventions. Scientific American, 52(8): 123–124, February 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n8/pdf/scientificamerican02211885-
123.pdf.

Anonymous:1885:PNPx

[6730] Anonymous. Patent notes: Patents. Scientific American, 52 (8):125, February 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n8/pdf/scientificamerican02211885-
125.pdf.

Anonymous:1885:MPP

[6731] Anonymous. The motive power of the Philadelphia Cable Railway. Scientific American, 52(8):111, February 21, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v52/
n8/pdf/scientificamerican02211885-111.pdf.

Anonymous:1885:CIM

[6732] Anonymous. Cast iron malleable. Scientific American, 52(8):112, February 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n8/pdf/scientificamerican02211885-112.pdf.

Anonymous:1885:NRI

[6733] Anonymous. A new remedy for the imported cabbage worm. Scientific American, 52(8):112, February 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v52/n8/pdf/scientificamerican02211885-
112a.pdf.

Kendall:1885:CNC

[6734] E. Dwight Kendall. The chemical nature of cotton. Scientific American, 52(8):113, February 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n8/pdf/scientificamerican02211885-
113.pdf.

Anonymous:1885:IMBa

[6735] Anonymous. Improved miter box. Scientific American, 52(8): 114, February 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n8/pdf/scientificamerican02211885-
114.pdf.

Anonymous:1885:RCE

[6736] Anonymous. Regenerator for compound engines. Scientific American, 52(8):114, February 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n8/pdf/scientificamerican02211885-
114a.pdf.

Anonymous:1885:ABC

[6737] Anonymous. Apparatus for building concrete walls. Scientific American, 52(8):114, February 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n8/pdf/scientificamerican02211885-
114b.pdf.

Anonymous:1885:IPC

[6738] Anonymous. Improved printer’s chase. Scientific American, 52 (8):115, February 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n8/pdf/scientificamerican02211885-
115.pdf.

Anonymous:1885:ISE

[6739] Anonymous. An improved safety engine. Scientific American, 52(8):115, February 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n8/pdf/scientificamerican02211885-
115a.pdf.

Anonymous:1885:THT

[6740] Anonymous. Trace or hame tug buckle. Scientific American, 52 (8):115–116, February 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n8/pdf/scientificamerican02211885-
115b.pdf.

Anonymous:1885:ASB

[6741] Anonymous. Adjustable single-bell chime whistle. Scientific American, 52(8):116, February 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n8/pdf/scientificamerican02211885-
116.pdf.

Anonymous:1885:PPb

[6742] Anonymous. Plant protector. Scientific American, 52(8):116, February 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n8/pdf/scientificamerican02211885-116a.pdf.

Anonymous:1885:RPE

[6743] Anonymous. Recent progress in electricity — the Phelps system of telegraphing from a railway train while in motion. Scientific American, 52(8):118–120, February 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n8/pdf/scientificamerican02211885-
118.pdf.

Anonymous:1885:EIx

[6744] Anonymous. Engineering inventions. Scientific American, 52(8): 121, February 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n8/pdf/scientificamerican02211885-
121.pdf.

Anonymous:1885:MIbe

[6745] Anonymous. Mechanical inventions. Scientific American, 52(8): 121, February 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n8/pdf/scientificamerican02211885-
121a.pdf.

Anonymous:1885:AIx (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n8/pdf/scientificamerican02211885-
121b.pdf.

Anonymous:1885:MIbf

[6747] Anonymous. Miscellaneous inventions. Scientific American, 52 (8):121, February 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n8/pdf/scientificamerican02211885-
121c.pdf.

Anonymous:1885:Cr

[6748] Anonymous. Correspondence. Scientific American, 52(9):133, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n9/pdf/scientificamerican02281885-133.pdf.

Anonymous:1885:BNBo

[6749] Anonymous. Books: New books and publications. Scientific American, 52(9):138, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n9/pdf/scientificamerican02281885-
138d.pdf.

Anonymous:1885:BPz

[6750] Anonymous. Business and personal. Scientific American, 52(9): 138–139, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n9/pdf/scientificamerican02281885-
138f.pdf.

Anonymous:1885:NQy

[6751] Anonymous. Notes and queries. Scientific American, 52(9):139, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n9/pdf/scientificamerican02281885-139.pdf.

Anonymous:1885:IIy

[6752] Anonymous. Index of inventions. Scientific American, 52(9): 139–140, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n9/pdf/scientificamerican02281885-
139a.pdf.

Anonymous:1885:PNPy

[6753] Anonymous. Patent notes: Patents. Scientific American, 52 (9):141, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n9/pdf/scientificamerican02281885-
141.pdf.

Anonymous:1885:ILS

[6754] Anonymous. Improved locomotive steam crane. Scientific American, 52(9):127, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n9/pdf/scientificamerican02281885-
127.pdf.

Anonymous:1885:CPa

[6755] Anonymous. The Commissionership of Patents. Scientific American, 52(9):128, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n9/pdf/scientificamerican02281885-
128.pdf.

Anonymous:1885:BM

[6756] Anonymous. Built-up mills. Scientific American, 52(9):128, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n9/pdf/scientificamerican02281885-128a.pdf.

Anonymous:1885:TRa

[6757] Anonymous. A trade as a refuge. Scientific American, 52(9):128, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n9/pdf/scientificamerican02281885-128b.pdf.

Anonymous:1885:APMb

[6758] Anonymous. Aspects of the planets for March. Scientific American, 52(9):129, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n9/pdf/scientificamerican02281885-
129.pdf.

Anonymous:1885:RTH

[6759] Anonymous. Railroad ticket holder. Scientific American, 52(9): 130, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n9/pdf/scientificamerican02281885-
130.pdf.

Anonymous:1885:CRb

[6760] Anonymous. Car replacer. Scientific American, 52(9):130, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n9/pdf/scientificamerican02281885-130a.pdf.

Anonymous:1885:KSB

[6761] Anonymous. Keel for submarine boats. Scientific American, 52 (9):130, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n9/pdf/scientificamerican02281885-
130b.pdf.

Anonymous:1885:INLa

[6762] Anonymous. Improved nut lock. Scientific American, 52(9):130, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n9/pdf/scientificamerican02281885-130c.pdf.

Anonymous:1885:ISCb

[6763] Anonymous. Improved shipping case. Scientific American, 52 (9):130, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n9/pdf/scientificamerican02281885-
130d.pdf.

Anonymous:1885:IBWb

[6764] Anonymous. Improved back-washing machine. Scientific American, 52(9):131, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n9/pdf/scientificamerican02281885-
131.pdf.

Anonymous:1885:UEA

[6765] Anonymous. Unloading and elevating apparatus. Scientific American, 52(9):131–132, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n9/pdf/scientificamerican02281885-
131a.pdf.

Anonymous:1885:ICFb

[6766] Anonymous. Improved car fare box. Scientific American, 52 (9):132, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n9/pdf/scientificamerican02281885-
132.pdf.

Anonymous:1885:TAN

[6767] Anonymous. Torpedoes of the Austrian Navy. Scientific American, 52(9):134, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n9/pdf/scientificamerican02281885-
134.pdf.

Anonymous:1885:EEM

[6768] Anonymous. An early electro-magnetic locomotive. Scientific American, 52(9):134, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n9/pdf/scientificamerican02281885-
134a.pdf.

Anonymous:1885:FRSb

[6769] Anonymous. Fishways on the river sire. Scientific American, 52 (9):135–137, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n9/pdf/scientificamerican02281885-
135.pdf.

Anonymous:1885:EIy

[6770] Anonymous. Engineering inventions. Scientific American, 52(9): 138, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n9/pdf/scientificamerican02281885-
138.pdf.

Anonymous:1885:MIbg

[6771] Anonymous. Mechanical inventions. Scientific American, 52(9): 138, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n9/pdf/scientificamerican02281885-
138a.pdf.

Anonymous:1885:AIy

[6772] Anonymous. Agricultural inventions. Scientific American, 52(9): 138, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n9/pdf/scientificamerican02281885-
138b.pdf.

Anonymous:1885:MIbh

[6773] Anonymous. Miscellaneous inventions. Scientific American, 52 (9):138, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v52/n9/pdf/scientificamerican02281885-
138c.pdf.

Anonymous:1885:Sh

[6774] Anonymous. Special. Scientific American, 52(9):138, February 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v52/n9/pdf/scientificamerican02281885-138e.pdf.

Anonymous:1885:BNBp

[6775] Anonymous. Books: New books and publications. Scientific American, 53(1):10, July 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n1/pdf/scientificamerican07041885-
10c.pdf.

Anonymous:1885:BPba

[6776] Anonymous. Business and personal. Scientific American, 53(1):10, July 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n1/pdf/scientificamerican07041885-10d.pdf.

Anonymous:1885:NQz

[6777] Anonymous. Notes and queries. Scientific American, 53(1):10–11, July 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n1/pdf/scientificamerican07041885-10e.pdf.

Anonymous:1885:IIz

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n1/pdf/scientificamerican07041885-11.pdf.

Anonymous:1885:PNPz

[6779] Anonymous. Patent notes: Patents. Scientific American, 53(1):14, July 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n1/pdf/scientificamerican07041885-14.pdf.

Anonymous:1885:GFG

[6780] Anonymous. The great French gun. Scientific American, 53(1):1, July 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n1/pdf/scientificamerican07041885-1.pdf.

Anonymous:1885:EG

[6781] Anonymous. An electric gun. Scientific American, 53(1):2, July 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n1/pdf/scientificamerican07041885-2.pdf.

Anonymous:1885:SWP

[6782] Anonymous. Shall water pipes and gas pipes be connected with lightning rods? Scientific American, 53(1):2–3, July 4, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v53/
n1/pdf/scientificamerican07041885-2a.pdf.

Anonymous:1885:APJb

[6783] Anonymous. Aspects of the planets for July. Scientific American, 53(1): 3, July 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n1/pdf/scientificamerican07041885-3.pdf.

Anonymous:1885:ICR

[6784] Anonymous. Improved car replacer. Scientific American, 53(1):4, July 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n1/pdf/scientificamerican07041885-4.pdf.

Anonymous:1885:DHV

[6785] Anonymous. Detachable handle for valises. Scientific American, 53(1): 4, July 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n1/pdf/scientificamerican07041885-4a.pdf.

Anonymous:1885:SFB

[6786] Anonymous. Spring frame for beds. Scientific American, 53(1):4, July 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n1/pdf/scientificamerican07041885-4b.pdf.

Anonymous:1885:RMb

[6787] Anonymous. Rolling mill. Scientific American, 53(1):4, July 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n1/pdf/scientificamerican07041885-4c.pdf.

Anonymous:1885:TKS

[6788] Anonymous. Telegraph key and sounder. Scientific American, 53(1):5, July 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n1/pdf/scientificamerican07041885-5.pdf.

Anonymous:1885:ISM

[6789] Anonymous. Improved shaping machine. Scientific American, 53(1):5, July 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n1/pdf/scientificamerican07041885-5a.pdf.

Anonymous:1885:OLT

[6790] Anonymous. Object lens for telescopes and cameras. Scientific American, 53(1):5, July 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n1/pdf/scientificamerican07041885-
5b.pdf.

Anonymous:1885:MFH

[6791] Anonymous. Machine for forming heels for ladies’ shoes. Scientific American, 53(1):6–7, July 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n1/pdf/scientificamerican07041885-
6.pdf.

Anonymous:1885:PCI

[6792] Anonymous. Photographic camera for instantaneous views. Scientific American, 53(1):8, July 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n1/pdf/scientificamerican07041885-
8.pdf.

Anonymous:1885:HNP

[6793] Anonymous. Hoffmann’s new plan for a cemetery. Scientific American, 53(1):8–9, July 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n1/pdf/scientificamerican07041885-
8a.pdf.

Anonymous:1885:TC

[6794] Anonymous. Thoroughbred cattle. Scientific American, 53(1):9, July 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n1/pdf/scientificamerican07041885-9.pdf.

Anonymous:1885:EIz

[6795] Anonymous. Engineering inventions. Scientific American, 53(1):10, July 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n1/pdf/scientificamerican07041885-10.pdf.

Anonymous:1885:AIz

[6796] Anonymous. Agricultural inventions. Scientific American, 53(1):10, July 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n1/pdf/scientificamerican07041885-10a.pdf.

Anonymous:1885:MIbi

[6797] Anonymous. Miscellaneous inventions. Scientific American, 53(1):10, July 4, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n1/pdf/scientificamerican07041885-10b.pdf.

Anonymous:1885:Cu

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n10/pdf/scientificamerican09051885-149.pdf.

Anonymous:1885:BNBq

[6799] Anonymous. Books: New books and publications. Scientific American, 53(10):154, September 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n10/pdf/scientificamerican09051885-
154c.pdf.

Anonymous:1885:BPbb

[6800] Anonymous. Business and personal. Scientific American, 53 (10):154, September 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n10/pdf/scientificamerican09051885-
154d.pdf.

Anonymous:1885:NQba

[6801] Anonymous. Notes and queries. Scientific American, 53(10): 155, September 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n10/pdf/scientificamerican09051885-
155.pdf.

Anonymous:1885:IIba

[6802] Anonymous. Index of inventions. Scientific American, 53(10): 155–156, September 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n10/pdf/scientificamerican09051885-
155a.pdf.

Anonymous:1885:PNPba

[6803] Anonymous. Patent notes: Patents. Scientific American, 53 (10):158, September 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n10/pdf/scientificamerican09051885-
158.pdf.

Anonymous:1885:OT

[6804] Anonymous. Optical telegraphy. Scientific American, 53(10): 143, September 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n10/pdf/scientificamerican09051885-
143.pdf.

Anonymous:1885:MWC

[6805] Anonymous. The Metropolitan Water Company’s proposal. Scientific American, 53(10):144, September 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v53/n10/pdf/scientificamerican09051885-
144.pdf.

Anonymous:1885:SO

[6806] Anonymous. Speed on the ocean. Scientific American, 53(10): 144, September 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n10/pdf/scientificamerican09051885-
144a.pdf.

Anonymous:1885:L

[6807] Anonymous. A. S. Lyman. Scientific American, 53(10):145, September 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n10/pdf/scientificamerican09051885-145.pdf.

Anonymous:1885:ILW

[6808] Anonymous. An improved lever wrench. Scientific American, 53 (10):146, September 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n10/pdf/scientificamerican09051885-
146.pdf.

Anonymous:1885:PSRa

[6809] Anonymous. Performance sheets and recording work done by locomotives. Scientific American, 53(10):146, September 5, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v53/
n10/pdf/scientificamerican09051885-146a.pdf.

Anonymous:1885:CWA

[6810] Anonymous. A car wheel and axle. Scientific American, 53 (10):146, September 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n10/pdf/scientificamerican09051885-
146b.pdf.

Anonymous:1885:ICL

[6811] Anonymous. An improved cover for large books. Scientific American, 53(10):146, September 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n10/pdf/scientificamerican09051885-
146c.pdf.

Anonymous:1885:SCPb

[6812] Anonymous. A sod cutter and pulverizer. Scientific American, 53(10):147, September 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n10/pdf/scientificamerican09051885-
147.pdf.

Anonymous:1885:MAI

[6813] Anonymous. A moist air injector for furnaces and forges. Scientific American, 53(10):147, September 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v53/n10/pdf/scientificamerican09051885-
147a.pdf.

Anonymous:1885:IHSb

[6814] Anonymous. An improved hand snow plow. Scientific American, 53(10):147, September 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n10/pdf/scientificamerican09051885-
147b.pdf.

Anonymous:1885:VCS

[6815] Anonymous. A vibrating cylinder steam valve. Scientific American, 53(10):147, September 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n10/pdf/scientificamerican09051885-
147c.pdf.

Anonymous:1885:IBSb

[6816] Anonymous. Improved button or stud. Scientific American, 53 (10):149, September 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n10/pdf/scientificamerican09051885-
149a.pdf.

Anonymous:1885:ITEb

[6817] Anonymous. Improved traction engine and crane. Scientific American, 53(10):150, September 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n10/pdf/scientificamerican09051885-
150.pdf.

Anonymous:1885:CPM

[6818] Anonymous. Composite portraits of members of the National Academy of Sciences. Scientific American, 53(10):151–152, September 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n10/pdf/scientificamerican09051885-151.pdf.

Anonymous:1885:HMM

[6819] Anonymous. How to make a magic three-sided fan. Scientific American, 53(10):152–153, September 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n10/pdf/scientificamerican09051885-
152.pdf.

Anonymous:1885:EIba

[6820] Anonymous. Engineering inventions. Scientific American, 53(10): 154, September 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n10/pdf/scientificamerican09051885-
154.pdf.

Anonymous:1885:AIba

[6821] Anonymous. Agricultural inventions. Scientific American, 53 (10):154, September 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n10/pdf/scientificamerican09051885-
154a.pdf.

Anonymous:1885:MIbj

[6822] Anonymous. Miscellaneous inventions. Scientific American, 53 (10):154, September 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n10/pdf/scientificamerican09051885-
154b.pdf.

Anonymous:1885:BPbc

[6823] Anonymous. Business and personal. Scientific American, 53 (11):170, September 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n11/pdf/scientificamerican09121885-
170c.pdf.

Anonymous:1885:NQbb

[6824] Anonymous. Notes and queries. Scientific American, 53(11): 170–171, September 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n11/pdf/scientificamerican09121885-
170d.pdf.

Anonymous:1885:IIbb

[6825] Anonymous. Index of inventions. Scientific American, 53(11): 171–172, September 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n11/pdf/scientificamerican09121885-
171.pdf.

Anonymous:1885:PNPbb

[6826] Anonymous. Patent notes: Patents. Scientific American, 53 (11):174, September 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n11/pdf/scientificamerican09121885-
174.pdf.

Anonymous:1885:IYR

[6827] Anonymous. The International Yacht Race. Scientific American, 53 (11):159–160, September 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n11/pdf/scientificamerican09121885-
159.pdf.

Anonymous:1885:NYC

[6828] Anonymous. New York cable road. Scientific American, 53 (11):160, September 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n11/pdf/scientificamerican09121885-
160.pdf.

Anonymous:1885:SWa

[6829] Anonymous. Submarine warfare. Scientific American, 53(11): 160–161, September 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n11/pdf/scientificamerican09121885-
160a.pdf.

Anonymous:1885:HGW

[6830] Anonymous. A heavy gun that will not burst. Scientific American, 53(11):161, September 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n11/pdf/scientificamerican09121885-
161.pdf.

Anonymous:1885:DSd

[6831] Anonymous. The decline of the sailor. Scientific American, 53 (11):161, September 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n11/pdf/scientificamerican09121885-
161a.pdf.

Anonymous:1885:ANa

[6832] Anonymous. Astronomical notes. Scientific American, 53(11): 161, September 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n11/pdf/scientificamerican09121885-
161b.pdf.

Anonymous:1885:DB

[6833] Anonymous. Drenching bit. Scientific American, 53(11):162, September 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n11/pdf/scientificamerican09121885-162.pdf.

Anonymous:1885:USB

[6834] Anonymous. Upright steam boiler. Scientific American, 53(11): 162, September 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n11/pdf/scientificamerican09121885-
162a.pdf.

Anonymous:1885:ITS

[6835] Anonymous. An improved tunnel strainer. Scientific American, 53 (11):162, September 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n11/pdf/scientificamerican09121885-
162b.pdf.

Anonymous:1885:IEL

[6836] Anonymous. Improved extensible ladder. Scientific American, 53 (11):162, September 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n11/pdf/scientificamerican09121885-
162c.pdf.

Anonymous:1885:IAC

[6837] Anonymous. Independent air and circulating pumping engine for the United States cruiser Chicago. Scientific American, 53(11):163, September 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n11/pdf/scientificamerican09121885-163.pdf.

Anonymous:1885:NMG

[6838] Anonymous. New method of gripping and crossing cables. Scientific American, 53(11):163–164, September 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v53/n11/pdf/scientificamerican09121885-
163a.pdf.

Anonymous:1885:FS

[6839] Anonymous. Flower stand. Scientific American, 53(11):164– 166, September 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n11/pdf/scientificamerican09121885-
164.pdf.

Anonymous:1885:IGG

[6840] Anonymous. The improved Gatling gun. Scientific American, 53 (11):166–169, September 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n11/pdf/scientificamerican09121885-
166.pdf.

Anonymous:1885:GC

[6841] Anonymous. Greenhouses and conservatories. Scientific American, 53(11):169, September 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n11/pdf/scientificamerican09121885-
169.pdf.

Anonymous:1885:EIbb

[6842] Anonymous. Engineering inventions. Scientific American, 53(11): 170, September 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n11/pdf/scientificamerican09121885-
170.pdf.

Anonymous:1885:AIbb

[6843] Anonymous. Agricultural inventions. Scientific American, 53 (11):170, September 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n11/pdf/scientificamerican09121885-
170a.pdf.

Anonymous:1885:MIbk

[6844] Anonymous. Miscellaneous inventions. Scientific American, 53 (11):170, September 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n11/pdf/scientificamerican09121885-
170b.pdf.

Anonymous:1885:Cv

[6845] Anonymous. Correspondence. Scientific American, 53(12):181, September 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n12/pdf/scientificamerican09191885-181a.pdf.

Anonymous:1885:BNBr

[6846] Anonymous. Books: New books and publications. Scientific American, 53(12):186, September 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n12/pdf/scientificamerican09191885-
186b.pdf.

Anonymous:1885:BPbd

[6847] Anonymous. Business and personal. Scientific American, 53(12): 186–187, September 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n12/pdf/scientificamerican09191885-
186c.pdf.

Anonymous:1885:NQbc

[6848] Anonymous. Notes and queries. Scientific American, 53(12): 187, September 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n12/pdf/scientificamerican09191885-
187.pdf.

Anonymous:1885:IIbc

[6849] Anonymous. Index of inventions. Scientific American, 53(12): 187–188, September 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n12/pdf/scientificamerican09191885-
187a.pdf.

Anonymous:1885:PNPbc

[6850] Anonymous. Patent notes: Patents. Scientific American, 53 (12):190, September 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n12/pdf/scientificamerican09191885-
190.pdf.

Anonymous:1885:SCa

[6851] Anonymous. Spontaneous combustion. Scientific American, 53 (12):175, September 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n12/pdf/scientificamerican09191885-
175.pdf.

Anonymous:1885:GTG

[6852] Anonymous. The great tangent galvanometer of the Cornell University. Scientific American, 53(12):175, September 19, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v53/
n12/pdf/scientificamerican09191885-175a.pdf.

Anonymous:1885:CNa

[6853] Anonymous. Cylindrical nuts. Scientific American, 53(12):176, September 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n12/pdf/scientificamerican09191885-176.pdf.

Anonymous:1885:RIN

[6854] Anonymous. Railway improvements needed. Scientific American, 53(12):176, September 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n12/pdf/scientificamerican09191885-
176a.pdf.

Anonymous:1885:SWb

[6855] Anonymous. As to the sinking of the wires. Scientific American, 53(12):176, September 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n12/pdf/scientificamerican09191885-
176b.pdf.

Anonymous:1885:SIb

[6856] Anonymous. Shop independence. Scientific American, 53(12): 176–177, September 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n12/pdf/scientificamerican09191885-
176c.pdf.

Anonymous:1885:DV

[6857] Anonymous. Death Valley. Scientific American, 53(12):177, September 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n12/pdf/scientificamerican09191885-177.pdf.

Anonymous:1885:IBF

[6858] Anonymous. Improved boiler feeder. Scientific American, 53 (12):178, September 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n12/pdf/scientificamerican09191885-
178.pdf.

Anonymous:1885:PT

[6859] Anonymous. A photo-tricycle. Scientific American, 53(12):178, September 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n12/pdf/scientificamerican09191885-178a.pdf.

Anonymous:1885:ISBa

[6860] Anonymous. Improved steam boiler. Scientific American, 53 (12):178, September 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n12/pdf/scientificamerican09191885-
178b.pdf.

Anonymous:1885:SCMa

[6861] Anonymous. Safety check for musical boxes. Scientific American, 53(12):179, September 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n12/pdf/scientificamerican09191885-
179.pdf.

Anonymous:1885:PAR

[6862] Anonymous. Petroleum and its application to the running of locomotives. Scientific American, 53(12):179–180, September 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n12/pdf/scientificamerican09191885-179a.pdf.

Anonymous:1885:PNa

[6863] Anonymous. Photographic notes. Scientific American, 53(12): 180, September 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n12/pdf/scientificamerican09191885-
180.pdf.

Anonymous:1885:WNS

[6864] Anonymous. The wonderful new star in Andromeda. Scientific American, 53(12):181, September 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n12/pdf/scientificamerican09191885-
181.pdf.

Anonymous:1885:IBM

[6865] Anonymous. Improved blooming mill. Scientific American, 53 (12):182, September 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n12/pdf/scientificamerican09191885-
182.pdf.

Holder:1885:SO

[6866] C. F. Holder. Salt in ornamentation. Scientific American, 53(12): 183–185, September 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n12/pdf/scientificamerican09191885-
183.pdf.

Hopkins:1885:MMAa

[6867] Geo. M. Hopkins. The microscope in the mechanic arts. Scientific American, 53(12):185, September 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n12/pdf/scientificamerican09191885-
185.pdf.

Anonymous:1885:AIbc

[6868] Anonymous. Agricultural inventions. Scientific American, 53 (12):186, September 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n12/pdf/scientificamerican09191885-
186.pdf.

Anonymous:1885:MIbl

[6869] Anonymous. Miscellaneous inventions. Scientific American, 53 (12):186, September 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n12/pdf/scientificamerican09191885-
186a.pdf.

Anonymous:1885:BNBs

[6870] Anonymous. Books: New books and publications. Scientific American, 53(13):202, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n13/pdf/scientificamerican09261885-
202d.pdf.

Anonymous:1885:BPbe

[6871] Anonymous. Business and personal. Scientific American, 53 (13):202, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n13/pdf/scientificamerican09261885-
202e.pdf.

Anonymous:1885:NQbd

[6872] Anonymous. Notes and queries. Scientific American, 53(13): 203, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n13/pdf/scientificamerican09261885-
203.pdf.

Anonymous:1885:IIbd

[6873] Anonymous. Index of inventions. Scientific American, 53(13): 203–204, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n13/pdf/scientificamerican09261885-
203a.pdf.

Anonymous:1885:PNPbd

[6874] Anonymous. Patent notes: Patents. Scientific American, 53 (13):206, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n13/pdf/scientificamerican09261885-
206.pdf.

Anonymous:1885:VH

[6875] Anonymous. Valves and hydrants. Scientific American, 53(13): 191, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n13/pdf/scientificamerican09261885-
191.pdf.

Anonymous:1885:BTT

[6876] Anonymous. The Berdan twin torpedo. Scientific American, 53 (13):192, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n13/pdf/scientificamerican09261885-
192.pdf.

Anonymous:1885:AKC

[6877] Anonymous. America to keep the Cup. Scientific American, 53 (13):192, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n13/pdf/scientificamerican09261885-
192a.pdf.

Anonymous:1885:EBa

[6878] Anonymous. Examinations of beer. Scientific American, 53(13): 192–193, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n13/pdf/scientificamerican09261885-
192b.pdf.

Anonymous:1885:ISd

[6879] Anonymous. An improved skate. Scientific American, 53(13): 194, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n13/pdf/scientificamerican09261885-
194.pdf.

Anonymous:1885:RGa

[6880] Anonymous. Railroad gate. Scientific American, 53(13):194, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n13/pdf/scientificamerican09261885-194a.pdf.

Anonymous:1885:IHb

[6881] Anonymous. Improved horseshoe. Scientific American, 53(13): 194, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n13/pdf/scientificamerican09261885-
194b.pdf.

Anonymous:1885:VS

[6882] Anonymous. Vehicle seat. Scientific American, 53(13):194, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n13/pdf/scientificamerican09261885-194c.pdf.

Anonymous:1885:UCH

[6883] Anonymous. Upright cushioned helve hammer. Scientific American, 53(13):195, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n13/pdf/scientificamerican09261885-
195.pdf.

Anonymous:1885:WSV

[6884] Anonymous. Window screen and ventilator. Scientific American, 53(13):195, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n13/pdf/scientificamerican09261885-
195a.pdf.

Anonymous:1885:NTd

[6885] Anonymous. A new torpedo. Scientific American, 53(13):196, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n13/pdf/scientificamerican09261885-196.pdf.

Anonymous:1885:CMP

[6886] Anonymous. Combined milk pan and cheese mould. Scientific American, 53(13):196, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n13/pdf/scientificamerican09261885-
196a.pdf.

Anonymous:1885:FNM

[6887] Anonymous. A few new microscope accessories. Scientific American, 53(13):197, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n13/pdf/scientificamerican09261885-
197.pdf.

Anonymous:1885:PNb

[6888] Anonymous. Photographic notes. Scientific American, 53(13): 197, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n13/pdf/scientificamerican09261885-
197a.pdf.

Anonymous:1885:ISAa

[6889] Anonymous. Improved steering apparatus for rudderless ships. Scientific American, 53(13):197, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v53/n13/pdf/scientificamerican09261885-
197b.pdf.

Anonymous:1885:NC

[6890] Anonymous. A novel clock. Scientific American, 53(13):197, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n13/pdf/scientificamerican09261885-197c.pdf.

Anonymous:1885:IGD

[6891] Anonymous. An improved grain drier. Scientific American, 53 (13):199, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n13/pdf/scientificamerican09261885-
199.pdf.

Anonymous:1885:ELA

[6892] Anonymous. Electric light adapted for use of naturalists, chemists, etc. Scientific American, 53(13):200, September 26, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v53/
n13/pdf/scientificamerican09261885-200.pdf.

Anonymous:1885:EAS

[6893] Anonymous. Electric apparatus for submarine lighting. Scientific American, 53(13):200–201, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v53/n13/pdf/scientificamerican09261885-
200a.pdf.

Anonymous:1885:ELG

[6894] Anonymous. Electric lighting of the Grevin Museum. Scientific American, 53(13):201, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n13/pdf/scientificamerican09261885-
201.pdf.

Anonymous:1885:EIbc

[6895] Anonymous. Engineering inventions. Scientific American, 53(13): 202, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n13/pdf/scientificamerican09261885-
202.pdf.

Anonymous:1885:MIbm

[6896] Anonymous. Mechanical inventions. Scientific American, 53(13): 202, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n13/pdf/scientificamerican09261885-
202a.pdf.

Anonymous:1885:AIbd

[6897] Anonymous. Agricultural inventions. Scientific American, 53 (13):202, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n13/pdf/scientificamerican09261885-
202b.pdf.

Anonymous:1885:MIbn

[6898] Anonymous. Miscellaneous inventions. Scientific American, 53 (13):202, September 26, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n13/pdf/scientificamerican09261885-
202c.pdf.

Anonymous:1885:Cw

[6899] Anonymous. Correspondence. Scientific American, 53(14):213, October 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n14/pdf/scientificamerican10031885-213.pdf.

Anonymous:1885:BPbf

[6900] Anonymous. Business and personal. Scientific American, 53(14):217, October 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n14/pdf/scientificamerican10031885-217c.pdf.

Anonymous:1885:BNBt

[6901] Anonymous. Books: New books and publications. Scientific American, 53(14):217, October 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n14/pdf/scientificamerican10031885-
217d.pdf.

Anonymous:1885:NQbe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n14/pdf/scientificamerican10031885-218.pdf.

Anonymous:1885:IIbe

[6903] Anonymous. Index of inventions. Scientific American, 53(14):219–220, October 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n14/pdf/scientificamerican10031885-219.pdf.

Anonymous:1885:PNPbe

[6904] Anonymous. Patent notes: Patents. Scientific American, 53(14):222, October 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n14/pdf/scientificamerican10031885-222.pdf.

Anonymous:1885:FDD

[6905] Anonymous. Floating depositing dock and hydraulic grid dock. Scientific American, 53(14):207–208, October 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v53/n14/pdf/scientificamerican10031885-
207.pdf.

Anonymous:1885:NPT

[6906] Anonymous. A new phase in the telephone litigations. Scientific American, 53(14):208, October 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n14/pdf/scientificamerican10031885-
208.pdf.

Anonymous:1885:SAb

[6907] Anonymous. Stiletto and Atalanta. Scientific American, 53(14): 208–209, October 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n14/pdf/scientificamerican10031885-
208a.pdf.

Anonymous:1885:POa

[6908] Anonymous. Pneumonia and ozone. Scientific American, 53(14):209, October 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n14/pdf/scientificamerican10031885-209.pdf.

Anonymous:1885:APO

[6909] Anonymous. Aspects of the planets for October. Scientific American, 53(14):209, October 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n14/pdf/scientificamerican10031885-
209a.pdf.

Anonymous:1885:THH

[6910] Anonymous. Tricycle hobby horse. Scientific American, 53(14):210, October 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n14/pdf/scientificamerican10031885-210.pdf.

Anonymous:1885:ALE

[6911] Anonymous. Automatic lamp extinguisher. Scientific American, 53(14):210, October 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n14/pdf/scientificamerican10031885-
210a.pdf.

Anonymous:1885:IGB

[6912] Anonymous. Improvement in grate bars. Scientific American, 53(14):210, October 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n14/pdf/scientificamerican10031885-
210b.pdf.

Anonymous:1885:SFM

[6913] Anonymous. Spring friction mirror pivot. Scientific American, 53(14):210, October 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n14/pdf/scientificamerican10031885-
210c.pdf.

Anonymous:1885:SAR

[6914] Anonymous. Stop attachment for roller skates. Scientific American, 53(14):210, October 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n14/pdf/scientificamerican10031885-
210d.pdf.

Anonymous:1885:SET

[6915] Anonymous. Supporting and end thrust anti-friction pads for shafts. Scientific American, 53(14):211, October 3, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v53/
n14/pdf/scientificamerican10031885-211.pdf.

Anonymous:1885:IAT

[6916] Anonymous. Improved apparatus for training horses. Scientific American, 53(14):211–212, October 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n14/pdf/scientificamerican10031885-
211a.pdf.

Anonymous:1885:USH

[6917] Anonymous. The utilization of solar heat for the elevation of water. Scientific American, 53(14):214, October 3, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v53/
n14/pdf/scientificamerican10031885-214.pdf.

Anonymous:1885:FTS

[6918] Anonymous. Fifteen ton steam forge cranes. Scientific American, 53(14):214, October 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n14/pdf/scientificamerican10031885-
214a.pdf.

Anonymous:1885:NAE

[6919] Anonymous. New applications of electricity. Scientific American, 53(14):215, October 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n14/pdf/scientificamerican10031885-
215.pdf.

Holder:1885:CAF

[6920] C. F. Holder. Constructive ability of fishes. Scientific American, 53(14):215–216, October 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n14/pdf/scientificamerican10031885-
215a.pdf.

Anonymous:1885:PNc

[6921] Anonymous. Photographic notes. Scientific American, 53(14):216, October 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n14/pdf/scientificamerican10031885-216.pdf.

Anonymous:1885:EIbd

[6922] Anonymous. Engineering inventions. Scientific American, 53(14):217, October 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n14/pdf/scientificamerican10031885-217.pdf.

Anonymous:1885:AIbe

[6923] Anonymous. Agricultural inventions. Scientific American, 53(14):217, October 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n14/pdf/scientificamerican10031885-217a.pdf.

Anonymous:1885:MIbo

[6924] Anonymous. Miscellaneous inventions. Scientific American, 53(14):217, October 3, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n14/pdf/scientificamerican10031885-217b.pdf.

Anonymous:1885:BPbg

[6925] Anonymous. Business and personal. Scientific American, 53 (15):234, October 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n15/pdf/scientificamerican10101885-
234c.pdf.

Anonymous:1885:NQbf

[6926] Anonymous. Notes and queries. Scientific American, 53(15): 234–235, October 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n15/pdf/scientificamerican10101885-
234d.pdf.

Anonymous:1885:IIbf

[6927] Anonymous. Index of inventions. Scientific American, 53(15): 235–236, October 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n15/pdf/scientificamerican10101885-
235a.pdf.

Anonymous:1885:PNPbf

[6928] Anonymous. Patent notes: Patents. Scientific American, 53 (15):238, October 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n15/pdf/scientificamerican10101885-
238.pdf.

Anonymous:1885:PA

[6929] Anonymous. The Pennsylvania aerolite. Scientific American, 53 (15):224, October 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n15/pdf/scientificamerican10101885-
224.pdf.

Anonymous:1885:GCP

[6930] Anonymous. Graphical comparison of perforation formulæ. Scientific American, 53(15):224, October 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v53/n15/pdf/scientificamerican10101885-
224a.pdf.

Anonymous:1885:CFE

[6931] Anonymous. Curious features of the electric lighting business. Scientific American, 53(15):224, October 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v53/n15/pdf/scientificamerican10101885-
224b.pdf.

Anonymous:1885:SMb

[6932] Anonymous. Student mechanics. Scientific American, 53(15): 224–225, October 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n15/pdf/scientificamerican10101885-
224c.pdf.

Proctor:1885:NSW

[6933] Richard A. Proctor. A new star within a star cloud. Scientific American, 53(15):225, October 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n15/pdf/scientificamerican10101885-
225.pdf.

Anonymous:1885:AWBb

[6934] Anonymous. An automatic wagon brake. Scientific American, 53(15):226, October 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n15/pdf/scientificamerican10101885-
226.pdf.

Anonymous:1885:ISPa

[6935] Anonymous. An improved snow plow. Scientific American, 53 (15):226, October 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n15/pdf/scientificamerican10101885-
226a.pdf.

Anonymous:1885:IHc

[6936] Anonymous. An improved horseshoe. Scientific American, 53 (15):226, October 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n15/pdf/scientificamerican10101885-
226b.pdf.

Anonymous:1885:BFR

[6937] Anonymous. Blowing up Flood Rock, Hell Gate. Scientific American, 53(15):227–228, October 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n15/pdf/scientificamerican10101885-
227.pdf.

Anonymous:1885:IWa

[6938] Anonymous. An improved wrench. Scientific American, 53(15): 228–229, October 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n15/pdf/scientificamerican10101885-
228.pdf.

Anonymous:1885:ISG

[6939] Anonymous. Improvement in small gas engines. Scientific American, 53(15):230, October 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n15/pdf/scientificamerican10101885-
230.pdf.

Anonymous:1885:CEU

[6940] Anonymous. Chair elevator for use at fires. Scientific American, 53(15):231, October 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n15/pdf/scientificamerican10101885-
231.pdf.

Anonymous:1885:CCT

[6941] Anonymous. The car coupler trials at Buffalo, N.Y. Scientific American, 53(15):232–233, October 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n15/pdf/scientificamerican10101885-
232.pdf.

Anonymous:1885:EIbe

[6942] Anonymous. Engineering inventions. Scientific American, 53(15): 234, October 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n15/pdf/scientificamerican10101885-
234.pdf.

Anonymous:1885:AIbf

[6943] Anonymous. Agricultural inventions. Scientific American, 53 (15):234, October 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n15/pdf/scientificamerican10101885-
234a.pdf.

Anonymous:1885:MIbp

[6944] Anonymous. Miscellaneous inventions. Scientific American, 53 (15):234, October 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n15/pdf/scientificamerican10101885-
234b.pdf.

Anonymous:1885:CRc

[6945] Anonymous. Communications received. Scientific American, 53 (15):235, October 10, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n15/pdf/scientificamerican10101885-
235.pdf.

Anonymous:1885:Cx

[6946] Anonymous. Correspondence. Scientific American, 53(16):245, October 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n16/pdf/scientificamerican10171885-245.pdf.

Anonymous:1885:BPbh

[6947] Anonymous. Business and personal. Scientific American, 53(16): 250–251, October 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n16/pdf/scientificamerican10171885-
250c.pdf.

Anonymous:1885:NQbg

[6948] Anonymous. Notes and queries. Scientific American, 53(16):251, October 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n16/pdf/scientificamerican10171885-251.pdf.

Anonymous:1885:IIbg

[6949] Anonymous. Index of inventions. Scientific American, 53(16): 251–252, October 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n16/pdf/scientificamerican10171885-
251a.pdf.

Anonymous:1885:PNPbg

[6950] Anonymous. Patent notes: Patents. Scientific American, 53 (16):254, October 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n16/pdf/scientificamerican10171885-
254.pdf.

Anonymous:1885:MRI

[6951] Anonymous. Mississippi River improvements. Scientific American, 53(16):240, October 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n16/pdf/scientificamerican10171885-
240.pdf.

Anonymous:1885:NME

[6952] Anonymous. A necklace of mummy eyes. Scientific American, 53 (16):240–241, October 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n16/pdf/scientificamerican10171885-
240a.pdf.

Anonymous:1885:HPSa

[6953] Anonymous. A hand-protecting sad iron holder. Scientific American, 53(16):242, October 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n16/pdf/scientificamerican10171885-
242.pdf.

Anonymous:1885:IG

[6954] Anonymous. An improved gate. Scientific American, 53(16):242, October 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n16/pdf/scientificamerican10171885-242a.pdf.

Anonymous:1885:OSM

[6955] Anonymous. Ore separating machine. Scientific American, 53 (16):242, October 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n16/pdf/scientificamerican10171885-
242b.pdf.

Anonymous:1885:CCP

[6956] Anonymous. Combined caliper, protractor, and bevel. Scientific American, 53(16):242, October 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n16/pdf/scientificamerican10171885-
242c.pdf.

Anonymous:1885:IHF

[6957] Anonymous. An improved hame fastener. Scientific American, 53(16):242, October 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n16/pdf/scientificamerican10171885-
242d.pdf.

Anonymous:1885:FRBb

[6958] Anonymous. Flood rock blown up. Scientific American, 53(16): 243–244, October 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n16/pdf/scientificamerican10171885-
243.pdf.

Anonymous:1885:IEBa

[6959] Anonymous. Improved expansion bolt. Scientific American, 53 (16):244, October 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n16/pdf/scientificamerican10171885-
244.pdf.

Anonymous:1885:SGP

[6960] Anonymous. A substitute for glass in photography. Scientific American, 53(16):246–247, October 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n16/pdf/scientificamerican10171885-
246.pdf.

Anonymous:1885:SCC

[6961] Anonymous. Sibley College, Cornell University. Scientific American, 53(16):247–248, October 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n16/pdf/scientificamerican10171885-
247.pdf.

Anonymous:1885:MB

[6962] Anonymous. Machines for bakeries. Scientific American, 53(16): 249, October 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n16/pdf/scientificamerican10171885-
249.pdf.

Anonymous:1885:EIbf

[6963] Anonymous. Engineering inventions. Scientific American, 53(16): 250, October 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n16/pdf/scientificamerican10171885-
250.pdf.

Anonymous:1885:AIbg

[6964] Anonymous. Agricultural inventions. Scientific American, 53 (16):250, October 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n16/pdf/scientificamerican10171885-
250a.pdf.

Anonymous:1885:MIbq

[6965] Anonymous. Miscellaneous inventions. Scientific American, 53 (16):250, October 17, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n16/pdf/scientificamerican10171885-
250b.pdf.

Anonymous:1885:Cy

[6966] Anonymous. Correspondence. Scientific American, 53(17):264, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n17/pdf/scientificamerican10241885-264.pdf.

Anonymous:1885:BNBu

[6967] Anonymous. Books: New books and publications. Scientific American, 53(17):266, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n17/pdf/scientificamerican10241885-
266b.pdf.

Anonymous:1885:BPbi

[6968] Anonymous. Business and personal. Scientific American, 53(17): 266–267, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n17/pdf/scientificamerican10241885-
266c.pdf.

Anonymous:1885:NQbh

[6969] Anonymous. Notes and queries. Scientific American, 53(17):267, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n17/pdf/scientificamerican10241885-267.pdf.

Anonymous:1885:IIbh 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n17/pdf/scientificamerican10241885-
267a.pdf.

Anonymous:1885:PNPbh

[6971] Anonymous. Patent notes: Patents. Scientific American, 53 (17):270, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n17/pdf/scientificamerican10241885-
270.pdf.

Anonymous:1885:EFR

[6972] Anonymous. Effects of the flood rock explosion. Scientific American, 53(17):256, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n17/pdf/scientificamerican10241885-
256.pdf.

Anonymous:1885:NDG

[6973] Anonymous. A new dynamite gun. Scientific American, 53(17):256, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n17/pdf/scientificamerican10241885-256a.pdf.

Anonymous:1885:BCI

[6974] Anonymous. Bending cast iron. Scientific American, 53(17): 256–257, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n17/pdf/scientificamerican10241885-
256b.pdf.

Hopkins:1885:MEP

[6975] Geo. M. Hopkins. Microscopical examination of the phenomenon of colors of thin plates. Scientific American, 53(17):257, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n17/pdf/scientificamerican10241885-257.pdf.

Anonymous:1885:IFEa

[6976] Anonymous. Improved fire escape. Scientific American, 53(17):258, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n17/pdf/scientificamerican10241885-258.pdf.

Anonymous:1885:ICPb

[6977] Anonymous. Improved corn planter. Scientific American, 53 (17):258, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n17/pdf/scientificamerican10241885-
258a.pdf.

Anonymous:1885:IFB

[6978] Anonymous. Improved flour bolt. Scientific American, 53(17):258, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n17/pdf/scientificamerican10241885-258b.pdf.

Anonymous:1885:GDMb

[6979] Anonymous. Gas detecter for mines. Scientific American, 53 (17):258–259, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n17/pdf/scientificamerican10241885-
258c.pdf.

Anonymous:1885:OEC

[6980] Anonymous. Oiler for eccentrics, cranks, etc. Scientific American, 53(17):259, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n17/pdf/scientificamerican10241885-
259.pdf.

Anonymous:1885:RF

[6981] Anonymous. Repeating firearm. Scientific American, 53(17):259, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n17/pdf/scientificamerican10241885-259a.pdf.

Anonymous:1885:IEG

[6982] Anonymous. Improved engine governor. Scientific American, 53 (17):259, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n17/pdf/scientificamerican10241885-
259b.pdf.

Anonymous:1885:RWT

[6983] Anonymous. Railway water tank. Scientific American, 53(17): 259–260, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n17/pdf/scientificamerican10241885-
259c.pdf.

Anonymous:1885:FA

[6984] Anonymous. Fireplace attachment. Scientific American, 53(17): 260, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n17/pdf/scientificamerican10241885-
260.pdf.

Anonymous:1885:TT

[6985] Anonymous. Tire tightener. Scientific American, 53(17):260, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n17/pdf/scientificamerican10241885-260a.pdf.

Anonymous:1885:HBa

[6986] Anonymous. Hook buckle. Scientific American, 53(17):260, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n17/pdf/scientificamerican10241885-260b.pdf.

Anonymous:1885:ISSd

[6987] Anonymous. Improved saw set. Scientific American, 53(17):260, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n17/pdf/scientificamerican10241885-260c.pdf.

Anonymous:1885:NTBa

[6988] Anonymous. The new Tower Bridge London. Scientific American, 53(17):261, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n17/pdf/scientificamerican10241885-
261.pdf.

Anonymous:1885:PNd

[6989] Anonymous. Photographic notes. Scientific American, 53(17):261, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n17/pdf/scientificamerican10241885-261a.pdf.

Anonymous:1885:ABM

[6990] Anonymous. Aillot’s brick machine. Scientific American, 53(17): 262–263, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n17/pdf/scientificamerican10241885-
262.pdf.

Anonymous:1885:LMS

[6991] Anonymous. Locks of the Manchester Ship Canal. Scientific American, 53(17):263, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n17/pdf/scientificamerican10241885-
263.pdf.

Anonymous:1885:NS

[6992] Anonymous. A new steamboat. Scientific American, 53(17):263, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n17/pdf/scientificamerican10241885-263a.pdf.

Anonymous:1885:KTL

[6993] Anonymous. Keeping track of locomotives. Scientific American, 53(17):264, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n17/pdf/scientificamerican10241885-
264a.pdf.

Anonymous:1885:GBP

[6994] Anonymous. The great “boa” of the Paris Museum. Scientific American, 53(17):265, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n17/pdf/scientificamerican10241885-
265.pdf.

Anonymous:1885:AIbh

[6995] Anonymous. Agricultural inventions. Scientific American, 53 (17):266, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n17/pdf/scientificamerican10241885-
266.pdf.

Anonymous:1885:MIbr

[6996] Anonymous. Miscellaneous inventions. Scientific American, 53 (17):266, October 24, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n17/pdf/scientificamerican10241885-
266a.pdf.

Anonymous:1885:Cz

[6997] Anonymous. Correspondence. Scientific American, 53(18):277, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n18/pdf/scientificamerican10311885-277.pdf.

Anonymous:1885:BPbj

[6998] Anonymous. Business and personal. Scientific American, 53 (18):282, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n18/pdf/scientificamerican10311885-
282c.pdf.

Anonymous:1885:NQbi

[6999] Anonymous. Notes and queries. Scientific American, 53(18): 282–283, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n18/pdf/scientificamerican10311885-
282d.pdf.

Anonymous:1885:IIbi

[7000] Anonymous. Index of inventions. Scientific American, 53(18): 283–284, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n18/pdf/scientificamerican10311885-
283.pdf.

Anonymous:1885:PNPbi

[7001] Anonymous. Patent notes: Patents. Scientific American, 53 (18):286, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n18/pdf/scientificamerican10311885-
286.pdf.

Anonymous:1885:PDG

[7002] Anonymous. The pneumatic dynamite gun. Scientific American, 53(18):271, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n18/pdf/scientificamerican10311885-
271.pdf.

Anonymous:1885:DAS

[7003] Anonymous. Decline of American seamanship. Scientific American, 53(18):272, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n18/pdf/scientificamerican10311885-
272.pdf.

Anonymous:1885:ICPc

[7004] Anonymous. Instinctive claim to property. Scientific American, 53(18):272, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n18/pdf/scientificamerican10311885-
272a.pdf.

Anonymous:1885:APN

[7005] Anonymous. Aspects of the planets for November. Scientific American, 53(18):273, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n18/pdf/scientificamerican10311885-
273.pdf.

Anonymous:1885:IFd

[7006] Anonymous. An improved fence. Scientific American, 53(18):274, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n18/pdf/scientificamerican10311885-274.pdf.

Anonymous:1885:ELd

[7007] Anonymous. An electric laryngoscope. Scientific American, 53 (18):274, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n18/pdf/scientificamerican10311885-
274a.pdf.

Anonymous:1885:FEa

[7008] Anonymous. Fire escape. Scientific American, 53(18):274, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n18/pdf/scientificamerican10311885-274b.pdf.

Anonymous:1885:SMc

[7009] Anonymous. Stamp moistener. Scientific American, 53(18):274, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n18/pdf/scientificamerican10311885-274c.pdf.

Anonymous:1885:ST

[7010] Anonymous. Steam trap. Scientific American, 53(18):275, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n18/pdf/scientificamerican10311885-275.pdf.

Anonymous:1885:ISPb

[7011] Anonymous. Improved seed planter. Scientific American, 53 (18):275, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n18/pdf/scientificamerican10311885-
275a.pdf.

Anonymous:1885:WWa

[7012] Anonymous. Water wheel. Scientific American, 53(18):275, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n18/pdf/scientificamerican10311885-275b.pdf.

Anonymous:1885:PNe

[7013] Anonymous. Photographic notes. Scientific American, 53(18):276, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n18/pdf/scientificamerican10311885-276.pdf.

Anonymous:1885:CSP

[7014] Anonymous. Combined shafts and pole for vehicles. Scientific American, 53(18):276, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n18/pdf/scientificamerican10311885-
276a.pdf.

Anonymous:1885:IPE

[7015] Anonymous. Improved pumping engine. Scientific American, 53 (18):278, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n18/pdf/scientificamerican10311885-
278.pdf.

Anonymous:1885:HFM

[7016] Anonymous. Hydraulic flanging machine. Scientific American, 53(18):279, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n18/pdf/scientificamerican10311885-
279.pdf.

Anonymous:1885:SCb

[7017] Anonymous. Steam catamaran. Scientific American, 53(18):279– 280, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n18/pdf/scientificamerican10311885-
279a.pdf.

Anonymous:1885:NIT

[7018] Anonymous. A new individual telephone call. Scientific American, 53(18):280, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n18/pdf/scientificamerican10311885-
280.pdf.

Anonymous:1885:CTA

[7019] Anonymous. Can the temperature of the Atlantic states be changed? Scientific American, 53(18):280–281, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v53/
n18/pdf/scientificamerican10311885-280a.pdf.

Anonymous:1885:EIbg

[7020] Anonymous. Engineering inventions. Scientific American, 53(18): 282, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n18/pdf/scientificamerican10311885-
282.pdf.

Anonymous:1885:AIbi

[7021] Anonymous. Agricultural inventions. Scientific American, 53 (18):282, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n18/pdf/scientificamerican10311885-
282a.pdf.

Anonymous:1885:MIbs

[7022] Anonymous. Miscellaneous inventions. Scientific American, 53 (18):282, October 31, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n18/pdf/scientificamerican10311885-
282b.pdf.

Anonymous:1885:Cbb

[7023] Anonymous. Correspondence. Scientific American, 53(19):296, November 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n19/pdf/scientificamerican11071885-296.pdf.

Anonymous:1885:BNBv

[7024] Anonymous. Books: New books and publications. Scientific American, 53(19):297, November 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n19/pdf/scientificamerican11071885-
297c.pdf.

Anonymous:1885:BPbk

[7025] Anonymous. Business and personal. Scientific American, 53 (19):297, November 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n19/pdf/scientificamerican11071885-
297d.pdf.

Anonymous:1885:NQbj

[7026] Anonymous. Notes and queries. Scientific American, 53(19): 298, November 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n19/pdf/scientificamerican11071885-
298.pdf.

Anonymous:1885:IIbj

[7027] Anonymous. Index of inventions. Scientific American, 53(19): 298–299, November 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n19/pdf/scientificamerican11071885-
298a.pdf.

Anonymous:1885:PNPbj

[7028] Anonymous. Patent notes: Patents. Scientific American, 53 (19):302, November 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n19/pdf/scientificamerican11071885-
302.pdf.

Anonymous:1885:GIB

[7029] Anonymous. Government investigation of the bell patents. Scientific American, 53(19):288, November 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v53/n19/pdf/scientificamerican11071885-
288.pdf.

Anonymous:1885:DMS

[7030] Anonymous. “The Dollar Medical Shop”. Scientific American, 53 (19):288–289, November 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n19/pdf/scientificamerican11071885-
288a.pdf.

Proctor:1885:NSO

[7031] Richard A. Proctor. Night sky — October and November. Scientific American, 53(19):289–290, November 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v53/n19/pdf/scientificamerican11071885-
289.pdf.

Anonymous:1885:IKG

[7032] Anonymous. An improved keyhole guard. Scientific American, 53(19):290, November 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n19/pdf/scientificamerican11071885-
290.pdf.

Anonymous:1885:ICS

[7033] Anonymous. An improved carpet stretcher. Scientific American, 53(19):290, November 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n19/pdf/scientificamerican11071885-
290a.pdf.

Anonymous:1885:FEb

[7034] Anonymous. Fire escape. Scientific American, 53(19):290, November 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n19/pdf/scientificamerican11071885-290b.pdf.

Anonymous:1885:ERT

[7035] Anonymous. Egg register and tester. Scientific American, 53 (19):290, November 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n19/pdf/scientificamerican11071885-
290c.pdf.

Anonymous:1885:BHT

[7036] Anonymous. Bag holder for trucks. Scientific American, 53 (19):290, November 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n19/pdf/scientificamerican11071885-
290d.pdf.

Anonymous:1885:SOW

[7037] Anonymous. Some old Warwick houses. Scientific American, 53 (19):291, November 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n19/pdf/scientificamerican11071885-
291.pdf.

Anonymous:1885:RDC

[7038] Anonymous. The relation of domestic chimneys to the construction of roofs. Scientific American, 53(19):291–292, November 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n19/pdf/scientificamerican11071885-291a.pdf.

Anonymous:1885:ICCd

[7039] Anonymous. Improved chimney cap. Scientific American, 53 (19):292, November 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n19/pdf/scientificamerican11071885-
292.pdf.

Anonymous:1885:PNf

[7040] Anonymous. Photographic notes. Scientific American, 53(19): 292–293, November 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n19/pdf/scientificamerican11071885-
292a.pdf.

Anonymous:1885:NCA

[7041] Anonymous. The new Croton Aqueduct. Scientific American, 53 (19):293–294, November 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n19/pdf/scientificamerican11071885-
293.pdf.

Anonymous:1885:NLM

[7042] Anonymous. A new link movement for reversible engines. Scientific American, 53(19):294, November 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v53/n19/pdf/scientificamerican11071885-
294.pdf.

Anonymous:1885:NSB

[7043] Anonymous. The Nordenfelt submarine boat. Scientific American, 53(19):295, November 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n19/pdf/scientificamerican11071885-
295.pdf.

Anonymous:1885:EIbh

[7044] Anonymous. Engineering inventions. Scientific American, 53(19): 297, November 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n19/pdf/scientificamerican11071885-
297.pdf.

Anonymous:1885:AIbj

[7045] Anonymous. Agricultural inventions. Scientific American, 53 (19):297, November 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n19/pdf/scientificamerican11071885-
297a.pdf.

Anonymous:1885:MIbt

[7046] Anonymous. Miscellaneous inventions. Scientific American, 53 (19):297, November 7, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n19/pdf/scientificamerican11071885-
297b.pdf.

Anonymous:1885:Cbc

[7047] Anonymous. Correspondence. Scientific American, 53(2):21, July 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n2/pdf/scientificamerican07111885-21.pdf.

Anonymous:1885:BPbl

[7048] Anonymous. Business and personal. Scientific American, 53(2):26, July 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n2/pdf/scientificamerican07111885-26.pdf.

Anonymous:1885:NQbk

[7049] Anonymous. Notes and queries. Scientific American, 53(2):26–27, July 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n2/pdf/scientificamerican07111885-26a.pdf.

Anonymous:1885:IIbk

[7050] Anonymous. Index of inventions. Scientific American, 53(2):27–28, July 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n2/pdf/scientificamerican07111885-27.pdf.

Anonymous:1885:PNPbk

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n2/pdf/scientificamerican07111885-30.pdf.

Anonymous:1885:CEC

[7052] Anonymous. The colossal elephant of Coney Island. Scientific American, 53(2):15, July 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n2/pdf/scientificamerican07111885-
15.pdf.

Anonymous:1885:CGI

[7053] Anonymous. Cotton ginning improvements needed. Scientific American, 53(2):16, July 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n2/pdf/scientificamerican07111885-
16.pdf.

Anonymous:1885:EFa

[7054] Anonymous. Edging by forging. Scientific American, 53(2):16, July 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n2/pdf/scientificamerican07111885-16a.pdf.

Anonymous:1885:PFJ

[7055] Anonymous. Professor Fleeming Jenkin, LL.D., F.R.S. Scientific American, 53(2):16–17, July 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n2/pdf/scientificamerican07111885-
16b.pdf.

Anonymous:1885:LSF

[7056] Anonymous. Life-saving fire appliances in New York. Scientific American, 53(2):17, July 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n2/pdf/scientificamerican07111885-
17.pdf.

Anonymous:1885:AC

[7057] Anonymous. An accelerating cartridge. Scientific American, 53(2):17, July 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n2/pdf/scientificamerican07111885-17a.pdf.

Anonymous:1885:RS

[7058] Anonymous. Railroad spike. Scientific American, 53(2):18, July 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n2/pdf/scientificamerican07111885-18.pdf.

Anonymous:1885:ILJ

[7059] Anonymous. Improved lifting jack. Scientific American, 53(2):18, July 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n2/pdf/scientificamerican07111885-18a.pdf.

Anonymous:1885:RGb

[7060] Anonymous. Railroad gate. Scientific American, 53(2):18, July 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n2/pdf/scientificamerican07111885-18b.pdf.

Anonymous:1885:BSV

[7061] Anonymous. Balanced slide valve. Scientific American, 53(2):18, July 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n2/pdf/scientificamerican07111885-18c.pdf.

Anonymous:1885:EBb

[7062] Anonymous. The Exposition at Budapesth. Scientific American, 53(2):19–20, July 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n2/pdf/scientificamerican07111885-
19.pdf.

Anonymous:1885:PCR

[7063] Anonymous. Pie and cake rack. Scientific American, 53(2):20, July 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n2/pdf/scientificamerican07111885-20.pdf.

Anonymous:1885:ISHb

[7064] Anonymous. Improved snap hook. Scientific American, 53(2):20, July 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n2/pdf/scientificamerican07111885-20a.pdf.

Anonymous:1885:CCf

[7065] Anonymous. Chimney cowl. Scientific American, 53(2):20, July 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n2/pdf/scientificamerican07111885-20b.pdf.

Anonymous:1885:HMB

[7066] Anonymous. H.M.S. Benbow. Scientific American, 53(2):22, July 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n2/pdf/scientificamerican07111885-22.pdf.

Anonymous:1885:SRM

[7067] Anonymous. Sail rigged merry-go-round. Scientific American, 53(2):22, July 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n2/pdf/scientificamerican07111885-22a.pdf.

Holder:1885:PFN

[7068] C. F. Holder. The paradise fish and its nest. Scientific American, 53(2):23–24, July 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n2/pdf/scientificamerican07111885-
23.pdf.

Anonymous:1885:EIbi

[7069] Anonymous. Engineering inventions. Scientific American, 53(2):25, July 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n2/pdf/scientificamerican07111885-25.pdf.

Anonymous:1885:AIbk

[7070] Anonymous. Agricultural inventions. Scientific American, 53(2):25, July 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n2/pdf/scientificamerican07111885-25a.pdf.

Anonymous:1885:MIbu

[7071] Anonymous. Miscellaneous inventions. Scientific American, 53(2):25, July 11, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n2/pdf/scientificamerican07111885-25b.pdf.

Anonymous:1885:Cbd

[7072] Anonymous. Correspondence. Scientific American, 53(20):309– 310, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n20/pdf/scientificamerican11141885-
309a.pdf.

Anonymous:1885:BPbm

[7073] Anonymous. Business and personal. Scientific American, 53 (20):314, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n20/pdf/scientificamerican11141885-
314c.pdf.

Anonymous:1885:NQbl

[7074] Anonymous. Notes and queries. Scientific American, 53(20): 314–315, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n20/pdf/scientificamerican11141885-
314d.pdf.

Anonymous:1885:IIbl

[7075] Anonymous. Index of inventions. Scientific American, 53(20): 315–316, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n20/pdf/scientificamerican11141885-
315.pdf.

Anonymous:1885:PNPbl

[7076] Anonymous. Patent notes: Patents. Scientific American, 53 (20):318, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n20/pdf/scientificamerican11141885-
318.pdf.

Anonymous:1885:WWb

[7077] Anonymous. The waste of whistling. Scientific American, 53 (20):304, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n20/pdf/scientificamerican11141885-
304.pdf.

Anonymous:1885:ANb

[7078] Anonymous. Astronomical notes. Scientific American, 53(20): 304–305, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n20/pdf/scientificamerican11141885-
304a.pdf.

Anonymous:1885:PNg

[7079] Anonymous. Photographic notes. Scientific American, 53(20): 305, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n20/pdf/scientificamerican11141885-
305.pdf.

Anonymous:1885:IOC

[7080] Anonymous. Improved oil cup. Scientific American, 53(20):306, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n20/pdf/scientificamerican11141885-306.pdf.

Anonymous:1885:PWR

[7081] Anonymous. Portable wire reel. Scientific American, 53(20): 306, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n20/pdf/scientificamerican11141885-
306a.pdf.

Anonymous:1885:IAd

[7082] Anonymous. Improved amalgamator. Scientific American, 53 (20):306, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n20/pdf/scientificamerican11141885-
306b.pdf.

Anonymous:1885:LH

[7083] Anonymous. Label holder. Scientific American, 53(20):306, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n20/pdf/scientificamerican11141885-306c.pdf.

Anonymous:1885:FPAa

[7084] Anonymous. A fence post of artificial stone and wood. Scientific American, 53(20):306, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n20/pdf/scientificamerican11141885-
306d.pdf.

Anonymous:1885:BCH

[7085] Anonymous. Book and copy holder. Scientific American, 53(20): 306–307, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n20/pdf/scientificamerican11141885-
306e.pdf.

Anonymous:1885:IPRb

[7086] Anonymous. Improved plate rolling machine. Scientific American, 53(20):307, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n20/pdf/scientificamerican11141885-
307.pdf.

Anonymous:1885:IRS

[7087] Anonymous. Improved railroad spike. Scientific American, 53 (20):307, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n20/pdf/scientificamerican11141885-
307a.pdf.

Anonymous:1885:BSb

[7088] Anonymous. Bicycle saddle. Scientific American, 53(20):307, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n20/pdf/scientificamerican11141885-307b.pdf.

Anonymous:1885:WHM

[7089] Anonymous. A wrist-held memorandum pad. Scientific American, 53(20):308, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n20/pdf/scientificamerican11141885-
308.pdf.

Anonymous:1885:SS

[7090] Anonymous. Sail skating. Scientific American, 53(20):308, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n20/pdf/scientificamerican11141885-308a.pdf.

Anonymous:1885:IPB

[7091] Anonymous. Improved pocket for billiard and pool tables. Scientific American, 53(20):308, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v53/n20/pdf/scientificamerican11141885-
308b.pdf.

Anonymous:1885:WBa

[7092] Anonymous. White bronze. Scientific American, 53(20):309, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n20/pdf/scientificamerican11141885-309.pdf.

Anonymous:1885:ART

[7093] Anonymous. Apparatus for the recovery of tar and ammonia from blast furnaces. Scientific American, 53(20):310–311, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n20/pdf/scientificamerican11141885-310.pdf.

Anonymous:1885:IDM

[7094] Anonymous. Improved drilling machine. Scientific American, 53 (20):311, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n20/pdf/scientificamerican11141885-
311.pdf.

Anonymous:1885:HPM

[7095] Anonymous. How to pack a mule. — the diamond hitch. Scientific American, 53(20):312, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v53/n20/pdf/scientificamerican11141885-
312.pdf.

Anonymous:1885:PTS

[7096] Anonymous. The Pendulum Titan shown at the Inventions Exhibition, London. Scientific American, 53(20):312, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n20/pdf/scientificamerican11141885-312a.pdf.

Anonymous:1885:PTC

[7097] Anonymous. The prehensile tailed coendou. Scientific American, 53(20):313, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n20/pdf/scientificamerican11141885-
313.pdf.

Anonymous:1885:EIbj

[7098] Anonymous. Engineering inventions. Scientific American, 53(20): 314, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n20/pdf/scientificamerican11141885-
314.pdf.

Anonymous:1885:AIbl

[7099] Anonymous. Agricultural inventions. Scientific American, 53 (20):314, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n20/pdf/scientificamerican11141885-
314a.pdf.

Anonymous:1885:MIbv

[7100] Anonymous. Miscellaneous inventions. Scientific American, 53 (20):314, November 14, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n20/pdf/scientificamerican11141885-
314b.pdf.

Anonymous:1885:Cbe

[7101] Anonymous. Correspondence. Scientific American, 53(21):325, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n21/pdf/scientificamerican11211885-325.pdf.

Anonymous:1885:BNBw

[7102] Anonymous. Books: New books and publications. Scientific American, 53(21):330, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n21/pdf/scientificamerican11211885-
330c.pdf.

Anonymous:1885:BPbn

[7103] Anonymous. Business and personal. Scientific American, 53(21): 330–331, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n21/pdf/scientificamerican11211885-
330d.pdf.

Anonymous:1885:NQbm

[7104] Anonymous. Notes and queries. Scientific American, 53(21): 331, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n21/pdf/scientificamerican11211885-
331.pdf.

Anonymous:1885:IIbm

[7105] Anonymous. Index of inventions. Scientific American, 53(21): 331–332, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n21/pdf/scientificamerican11211885-
331a.pdf.

Anonymous:1885:PNPbm

[7106] Anonymous. Patent notes: Patents. Scientific American, 53 (21):334, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n21/pdf/scientificamerican11211885-
334.pdf.

Anonymous:1885:EMN

[7107] Anonymous. Electric motor on the New York Elevated Railroad. Scientific American, 53(21):319, November 21, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v53/
n21/pdf/scientificamerican11211885-319.pdf.

Anonymous:1885:NAS

[7108] Anonymous. The National Academy of Sciences. Scientific American, 53(21):320–321, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n21/pdf/scientificamerican11211885-
320.pdf.

Anonymous:1885:PNh

[7109] Anonymous. Photographic notes. Scientific American, 53(21): 321, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n21/pdf/scientificamerican11211885-
321.pdf.

Anonymous:1885:PSMb

[7110] Anonymous. Purifying and separating middlings. Scientific American, 53(21):322, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n21/pdf/scientificamerican11211885-
322.pdf.

Anonymous:1885:CCAa

[7111] Anonymous. Cotton chopper attachment for cultivators. Scientific American, 53(21):322, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v53/n21/pdf/scientificamerican11211885-
322a.pdf.

Anonymous:1885:CNb

[7112] Anonymous. Cutting nippers. Scientific American, 53(21):322, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n21/pdf/scientificamerican11211885-322b.pdf.

Anonymous:1885:RCb

[7113] Anonymous. Revolving cultivator. Scientific American, 53(21): 322, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n21/pdf/scientificamerican11211885-
322c.pdf.

Anonymous:1885:ITb

[7114] Anonymous. Improved trunk. Scientific American, 53(21):322– 323, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n21/pdf/scientificamerican11211885-
322d.pdf.

Anonymous:1885:ICCe

[7115] Anonymous. Improved car coupler. Scientific American, 53(21): 323, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n21/pdf/scientificamerican11211885-
323.pdf.

Anonymous:1885:RRJ

[7116] Anonymous. Railroad rail joint. Scientific American, 53(21): 323, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n21/pdf/scientificamerican11211885-
323a.pdf.

Anonymous:1885:DCa

[7117] Anonymous. Dumping car. Scientific American, 53(21):323, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n21/pdf/scientificamerican11211885-323b.pdf.

Anonymous:1885:MCc

[7118] Anonymous. The “Monarch” churn. Scientific American, 53(21): 323–324, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n21/pdf/scientificamerican11211885-
323c.pdf.

Hopkins:1885:MEC

[7119] George M. Hopkins. Microscopic examination of ciliated organisms by intermittent light. Scientific American, 53(21):324, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n21/pdf/scientificamerican11211885-324.pdf.

Lalande:1885:NAM

[7120] F. De Lalande. A new apparatus for measuring electric currents. Scientific American, 53(21):325, November 21, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v53/
n21/pdf/scientificamerican11211885-325a.pdf.

Anonymous:1885:OTL

[7121] Anonymous. Opening of a Telpher line. Scientific American, 53 (21):326–327, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n21/pdf/scientificamerican11211885-
326.pdf.

Anonymous:1885:HBG

[7122] Anonymous. Hub-boxing gauge. Scientific American, 53(21): 327, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n21/pdf/scientificamerican11211885-
327.pdf.

Anonymous:1885:ISLd

[7123] Anonymous. Improved split link. Scientific American, 53(21): 327, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n21/pdf/scientificamerican11211885-
327a.pdf.

Anonymous:1885:BCb

[7124] Anonymous. Buttonhole cutter. Scientific American, 53(21): 327, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n21/pdf/scientificamerican11211885-
327b.pdf.

Anonymous:1885:HCM

[7125] Anonymous. Hair cutting machine. Scientific American, 53(21): 327–328, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n21/pdf/scientificamerican11211885-
327c.pdf.

Anonymous:1885:CBAa

[7126] Anonymous. Captive balloons for Army use. Scientific American, 53 (21):328–329, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n21/pdf/scientificamerican11211885-
328.pdf.

Anonymous:1885:EIbk

[7127] Anonymous. Engineering inventions. Scientific American, 53(21): 330, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n21/pdf/scientificamerican11211885-
330.pdf.

Anonymous:1885:AIbm

[7128] Anonymous. Agricultural inventions. Scientific American, 53 (21):330, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n21/pdf/scientificamerican11211885-
330a.pdf.

Anonymous:1885:MIbw

[7129] Anonymous. Miscellaneous inventions. Scientific American, 53 (21):330, November 21, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n21/pdf/scientificamerican11211885-
330b.pdf.

Anonymous:1885:BNBx

[7130] Anonymous. Books: New books and publications. Scientific American, 53(22):346, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n22/pdf/scientificamerican11281885-
346d.pdf.

Anonymous:1885:BPbo

[7131] Anonymous. Business and personal. Scientific American, 53(22): 346–347, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n22/pdf/scientificamerican11281885-
346f.pdf.

Anonymous:1885:NQbn

[7132] Anonymous. Notes and queries. Scientific American, 53(22): 347, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n22/pdf/scientificamerican11281885-
347.pdf.

Anonymous:1885:IIbn

[7133] Anonymous. Index of inventions. Scientific American, 53(22): 347–348, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n22/pdf/scientificamerican11281885-
347a.pdf.

Anonymous:1885:PNPbn

[7134] Anonymous. Patent notes: Patents. Scientific American, 53 (22):350, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n22/pdf/scientificamerican11281885-
350.pdf.

Anonymous:1885:ORT

[7135] Anonymous. The original Reis telephone. Scientific American, 53 (22):335, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n22/pdf/scientificamerican11281885-
335.pdf.

Anonymous:1885:PTL

[7136] Anonymous. Progress of the telephone litigation. Scientific American, 53(22):336, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n22/pdf/scientificamerican11281885-
336.pdf.

Anonymous:1885:PRP

[7137] Anonymous. Is the Panama route practicable? Scientific American, 53(22):336, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n22/pdf/scientificamerican11281885-
336a.pdf.

Anonymous:1885:SCMb

[7138] Anonymous. Sugar cane mills. Scientific American, 53(22):336, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n22/pdf/scientificamerican11281885-336b.pdf.

Anonymous:1885:SFC

[7139] Anonymous. Shell fisheries of Connecticut. Scientific American, 53 (22):336–337, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n22/pdf/scientificamerican11281885-
336c.pdf.

Anonymous:1885:WBb

[7140] Anonymous. White bronze. Scientific American, 53(22):337, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n22/pdf/scientificamerican11281885-337.pdf.

Anonymous:1885:PNi

[7141] Anonymous. Photographic notes. Scientific American, 53(22): 337, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n22/pdf/scientificamerican11281885-
337a.pdf.

Anonymous:1885:ISBb

[7142] Anonymous. Improved sample box. Scientific American, 53(22): 338, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n22/pdf/scientificamerican11281885-
338.pdf.

Anonymous:1885:WM

[7143] Anonymous. Washing machine. Scientific American, 53(22): 338, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n22/pdf/scientificamerican11281885-
338a.pdf.

Anonymous:1885:CAM

[7144] Anonymous. Cutting apparatus for mowers and reapers. Scientific American, 53(22):338, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v53/n22/pdf/scientificamerican11281885-
338b.pdf.

Anonymous:1885:ITW

[7145] Anonymous. Improved two wheeled vehicle. Scientific American, 53(22):338, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n22/pdf/scientificamerican11281885-
338c.pdf.

Anonymous:1885:SJS

[7146] Anonymous. The St. John Steel Cantilever Bridge. Scientific American, 53(22):339–341, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n22/pdf/scientificamerican11281885-
339.pdf.

Anonymous:1885:DWD

[7147] Anonymous. Donald and Waddell’s dry dock. Scientific American, 53(22):343, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n22/pdf/scientificamerican11281885-
343.pdf.

Anonymous:1885:IGM

[7148] Anonymous. Improved grinding machine. Scientific American, 53 (22):343–344, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n22/pdf/scientificamerican11281885-
343a.pdf.

Battle:1885:IWB

[7149] H. B. Battle. An improved wash-bottle for chemical laboratories. Scientific American, 53(22):344, November 28, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v53/
n22/pdf/scientificamerican11281885-344.pdf.

Anonymous:1885:CAH

[7150] Anonymous. A convenient article for hospitals and sick rooms. Scientific American, 53(22):344, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v53/n22/pdf/scientificamerican11281885-
344a.pdf.

Anonymous:1885:INW

[7151] Anonymous. Improved “Noble” wool comber. Scientific American, 53(22):345, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n22/pdf/scientificamerican11281885-
345.pdf.

Anonymous:1885:EIbl

[7152] Anonymous. Engineering inventions. Scientific American, 53(22): 346, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n22/pdf/scientificamerican11281885-
346.pdf.

Anonymous:1885:MIbx

[7153] Anonymous. Mechanical inventions. Scientific American, 53(22): 346, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n22/pdf/scientificamerican11281885-
346a.pdf.

Anonymous:1885:AIbn

[7154] Anonymous. Agricultural inventions. Scientific American, 53 (22):346, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n22/pdf/scientificamerican11281885-
346b.pdf.

Anonymous:1885:MIby

[7155] Anonymous. Miscellaneous inventions. Scientific American, 53 (22):346, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n22/pdf/scientificamerican11281885-
346c.pdf.

Anonymous:1885:Sj

[7156] Anonymous. Special. Scientific American, 53(22):346, November 28, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n22/pdf/scientificamerican11281885-346e.pdf.

Anonymous:1885:Cbf

[7157] Anonymous. Correspondence. Scientific American, 53(23):357, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n23/pdf/scientificamerican12051885-357b.pdf.

Anonymous:1885:BPbp

[7158] Anonymous. Business and personal. Scientific American, 53 (23):361, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n23/pdf/scientificamerican12051885-
361c.pdf.

Anonymous:1885:BNBy

[7159] Anonymous. Books: New books and publications. Scientific American, 53(23):361, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n23/pdf/scientificamerican12051885-
361d.pdf.

Anonymous:1885:NQbo

[7160] Anonymous. Notes and queries. Scientific American, 53(23): 361–362, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n23/pdf/scientificamerican12051885-
361e.pdf.

Anonymous:1885:IIbo

[7161] Anonymous. Index of inventions. Scientific American, 53(23): 362–363, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n23/pdf/scientificamerican12051885-
362a.pdf.

Anonymous:1885:GAb

[7162] Anonymous. The greyhounds of the Atlantic. Scientific American, 53(23):352, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n23/pdf/scientificamerican12051885-
352.pdf.

Anonymous:1885:MPR

[7163] Anonymous. M. Pasteur’s researches in the treatment of hydrophobia. Scientific American, 53(23):352, December 5, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v53/
n23/pdf/scientificamerican12051885-352a.pdf.

Anonymous:1885:WBc

[7164] Anonymous. War balloons. Scientific American, 53(23):352, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n23/pdf/scientificamerican12051885-352b.pdf.

Proctor:1885:NSN

[7165] Richard A. Proctor. Night sky — November and December. Scientific American, 53(23):353, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v53/n23/pdf/scientificamerican12051885-
353.pdf.

Anonymous:1885:IMBb

[7166] Anonymous. Improved method of burning limestone. Scientific American, 53(23):354, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n23/pdf/scientificamerican12051885-
354.pdf.

Anonymous:1885:IWb

[7167] Anonymous. An improved window. Scientific American, 53(23):354, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n23/pdf/scientificamerican12051885-354a.pdf.

Anonymous:1885:MBA

[7168] Anonymous. Machine for beveling angle bars, etc. Scientific American, 53(23):355, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n23/pdf/scientificamerican12051885-
355.pdf.

Anonymous:1885:SGc

[7169] Anonymous. The sleeper’s guard. Scientific American, 53(23):355, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n23/pdf/scientificamerican12051885-355a.pdf.

Anonymous:1885:ITL

[7170] Anonymous. An improved turning lathe. Scientific American, 53 (23):355–356, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n23/pdf/scientificamerican12051885-
355b.pdf.

Anonymous:1885:BFC

[7171] Anonymous. Boiler flue cleaner. Scientific American, 53(23):356, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n23/pdf/scientificamerican12051885-356.pdf.

Anonymous:1885:CRS

[7172] Anonymous. Clamp for roof scaffolds, etc. Scientific American, 53(23):356, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n23/pdf/scientificamerican12051885-
356a.pdf.

Anonymous:1885:PSS

[7173] Anonymous. Portable sash securer. Scientific American, 53(23):356, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n23/pdf/scientificamerican12051885-356b.pdf.

Anonymous:1885:EOW

[7174] Anonymous. An easily operated washing machine. Scientific American, 53(23):356, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n23/pdf/scientificamerican12051885-
356c.pdf.

Anonymous:1885:RCI

[7175] Anonymous. Repairing the Cooper Institute. Scientific American, 53(23):357, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n23/pdf/scientificamerican12051885-
357.pdf.

Anonymous:1885:PNj

[7176] Anonymous. Photographic notes. Scientific American, 53(23):357, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n23/pdf/scientificamerican12051885-357a.pdf.

Anonymous:1885:POb

[7177] Anonymous. The preservation of the Obelisk. Scientific American, 53(23):358, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n23/pdf/scientificamerican12051885-
358.pdf.

Anonymous:1885:MGC

[7178] Anonymous. Moncrieff gun carriages for Russian iron clads. Scientific American, 53(23):359–360, December 5, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v53/
n23/pdf/scientificamerican12051885-359.pdf.

Anonymous:1885:SKN

[7179] Anonymous. The Silver King in Northern waters. Scientific American, 53(23):360, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n23/pdf/scientificamerican12051885-
360.pdf.

Anonymous:1885:EIbm

[7180] Anonymous. Engineering inventions. Scientific American, 53(23): 361, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n23/pdf/scientificamerican12051885-
361.pdf.

Anonymous:1885:MIbz

[7181] Anonymous. Miscellaneous inventions. Scientific American, 53 (23):361, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n23/pdf/scientificamerican12051885-
361a.pdf.

Anonymous:1885:Sk

[7182] Anonymous. Special. Scientific American, 53(23):361, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n23/pdf/scientificamerican12051885-361b.pdf.

Anonymous:1885:CRd

[7183] Anonymous. Communication received. Scientific American, 53 (23):362, December 5, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n23/pdf/scientificamerican12051885-
362.pdf.

Anonymous:1885:Cbg

[7184] Anonymous. Correspondence. Scientific American, 53(24):376, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n24/pdf/scientificamerican12121885-376.pdf.

Anonymous:1885:BPbq

[7185] Anonymous. Business and personal. Scientific American, 53 (24):378, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n24/pdf/scientificamerican12121885-
378.pdf.

Anonymous:1885:NQbp

[7186] Anonymous. Notes and queries. Scientific American, 53(24): 378–379, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n24/pdf/scientificamerican12121885-
378a.pdf.

Anonymous:1885:IIbp

[7187] Anonymous. Index of inventions. Scientific American, 53(24): 379–380, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n24/pdf/scientificamerican12121885-
379.pdf.

Anonymous:1885:PNPbo 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n24/pdf/scientificamerican12121885-
382.pdf.

Anonymous:1885:HMW

[7189] Anonymous. How to make winter evenings profitable. Scientific American, 53(24):368, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n24/pdf/scientificamerican12121885-
368.pdf.

Anonymous:1885:NMS

[7190] Anonymous. The November meteoric showers. Scientific American, 53(24):368, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n24/pdf/scientificamerican12121885-
368a.pdf.

Anonymous:1885:PNk

[7191] Anonymous. Photographic notes. Scientific American, 53(24): 368, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n24/pdf/scientificamerican12121885-
368b.pdf.

Anonymous:1885:APD

[7192] Anonymous. Aspects of the planets for December. Scientific American, 53(24):369, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n24/pdf/scientificamerican12121885-
369.pdf.

Anonymous:1885:GDMc

[7193] Anonymous. Grading and ditching machine. Scientific American, 53(24):370, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n24/pdf/scientificamerican12121885-
370.pdf.

Anonymous:1885:IFEb

[7194] Anonymous. Improved fire escape. Scientific American, 53(24): 370, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n24/pdf/scientificamerican12121885-
370a.pdf.

Anonymous:1885:IMT

[7195] Anonymous. Improved metal tongs. Scientific American, 53 (24):370, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n24/pdf/scientificamerican12121885-
370b.pdf.

Anonymous:1885:GW

[7196] Anonymous. Gate for waterwheels. Scientific American, 53(24): 370, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n24/pdf/scientificamerican12121885-
370c.pdf.

Anonymous:1885:SCFb

[7197] Anonymous. Smoke consuming furnace. Scientific American, 53 (24):371, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n24/pdf/scientificamerican12121885-
371.pdf.

Anonymous:1885:CCg

[7198] Anonymous. The contractor’s crane. Scientific American, 53(24): 371–372, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n24/pdf/scientificamerican12121885-
371a.pdf.

Anonymous:1885:IHd

[7199] Anonymous. An improved horseshoe. Scientific American, 53 (24):372, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n24/pdf/scientificamerican12121885-
372.pdf.

Anonymous:1885:IMBc

[7200] Anonymous. Improved mail bag and lock. Scientific American, 53(24):372, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n24/pdf/scientificamerican12121885-
372a.pdf.

Anonymous:1885:SCe

[7201] Anonymous. Stock and collar. Scientific American, 53(24):372, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n24/pdf/scientificamerican12121885-372b.pdf.

Anonymous:1885:DCb

[7202] Anonymous. Door check. Scientific American, 53(24):372, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n24/pdf/scientificamerican12121885-372c.pdf.

Anonymous:1885:KAE

[7203] Anonymous. The Knickerbocker Avenue extension sewer, Brooklyn, N.Y. Scientific American, 53(24):373, December 12, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v53/
n24/pdf/scientificamerican12121885-373.pdf.

Anonymous:1885:ICA

[7204] Anonymous. Improved cork attachment. Scientific American, 53 (24):373, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n24/pdf/scientificamerican12121885-
373a.pdf.

Anonymous:1885:BWM

[7205] Anonymous. Balanced water motor. Scientific American, 53(24): 373–374, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n24/pdf/scientificamerican12121885-
373b.pdf.

Anonymous:1885:NDM

[7206] Anonymous. A new domestic motor. Scientific American, 53 (24):374, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n24/pdf/scientificamerican12121885-
374.pdf.

Holder:1885:CCC

[7207] C. F. Holder. Candelabra cactus and California woodpecker. Scientific American, 53(24):375, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v53/n24/pdf/scientificamerican12121885-
375.pdf.

Anonymous:1885:EIbn

[7208] Anonymous. Engineering inventions. Scientific American, 53(24): 377, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n24/pdf/scientificamerican12121885-
377.pdf.

Anonymous:1885:AIbo

[7209] Anonymous. Agricultural inventions. Scientific American, 53 (24):377, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n24/pdf/scientificamerican12121885-
377a.pdf.

Anonymous:1885:MIca

[7210] Anonymous. Miscellaneous inventions. Scientific American, 53 (24):377, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n24/pdf/scientificamerican12121885-
377b.pdf.

Anonymous:1885:Sn

[7211] Anonymous. Special. Scientific American, 53(24):377, December 12, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n24/pdf/scientificamerican12121885-377c.pdf.

Anonymous:1885:Cbh

[7212] Anonymous. Correspondence. Scientific American, 53(25):389, December 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n25/pdf/scientificamerican12191885-389.pdf.

Anonymous:1885:BPbr 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n25/pdf/scientificamerican12191885-
394d.pdf.

Anonymous:1885:BNBz

[7214] Anonymous. Books: New books and publications. Scientific American, 53(25):394, December 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n25/pdf/scientificamerican12191885-
394e.pdf.

Anonymous:1885:NQbq

[7215] Anonymous. Notes and queries. Scientific American, 53(25): 395, December 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n25/pdf/scientificamerican12191885-
395.pdf.

Anonymous:1885:IIbq

[7216] Anonymous. Index of inventions. Scientific American, 53(25): 395–396, December 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n25/pdf/scientificamerican12191885-
395a.pdf.

Anonymous:1885:PNPbp

[7217] Anonymous. Patent notes: Patents. Scientific American, 53 (25):398, December 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n25/pdf/scientificamerican12191885-
398.pdf.

Anonymous:1885:SMF

[7218] Anonymous. The St. Mary’s Falls Canal. Scientific American, 53 (25):383, December 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n25/pdf/scientificamerican12191885-
383.pdf.

Anonymous:1885:RNC

[7219] Anonymous. Revival of the Nicaragua Canal Project. Scientific American, 53(25):384, December 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n25/pdf/scientificamerican12191885-
384.pdf.

Anonymous:1885:PRO

[7220] Anonymous. The power required to operate the ship railway. Scientific American, 53(25):385, December 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v53/n25/pdf/scientificamerican12191885-
385.pdf.

Anonymous:1885:PNl

[7221] Anonymous. Photographic notes. Scientific American, 53(25): 385, December 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n25/pdf/scientificamerican12191885-
385a.pdf.

Anonymous:1885:PHS

[7222] Anonymous. Pruning and house shears. Scientific American, 53 (25):388, December 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n25/pdf/scientificamerican12191885-
388.pdf.

Anonymous:1885:IFH

[7223] Anonymous. Improved fish hook. Scientific American, 53(25): 388, December 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n25/pdf/scientificamerican12191885-
388a.pdf.

Anonymous:1885:LIM

[7224] Anonymous. Luscher’s ice machine. Scientific American, 53(25): 388, December 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n25/pdf/scientificamerican12191885-
388b.pdf.

Anonymous:1885:TP

[7225] Anonymous. Tobacco pipe. Scientific American, 53(25):388, December 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n25/pdf/scientificamerican12191885-388c.pdf.

Anonymous:1885:PBH

[7226] Anonymous. Paper bag holder. Scientific American, 53(25):388, December 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n25/pdf/scientificamerican12191885-388d.pdf.

Anonymous:1885:BFd

[7227] Anonymous. Button fastener. Scientific American, 53(25):388, December 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n25/pdf/scientificamerican12191885-388e.pdf.

Anonymous:1885:ISBc

[7228] Anonymous. Improved sand blast machine. Scientific American, 53(25):390, December 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n25/pdf/scientificamerican12191885-
390.pdf.

Hopkins:1885:C

[7229] M. Hopkins. The cycloidotrope. Scientific American, 53(25): 390, December 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n25/pdf/scientificamerican12191885-
390a.pdf.

Anonymous:1885:MPA

[7230] Anonymous. M. Pasteur’s American patients. Scientific American, 53 (25):391–393, December 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n25/pdf/scientificamerican12191885-
391.pdf.

Anonymous:1885:BIP

[7231] Anonymous. Blake’s “improved 1884” pipe hanger. Scientific American, 53(25):393, December 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n25/pdf/scientificamerican12191885-
393.pdf.

Anonymous:1885:EIbo (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n25/pdf/scientificamerican12191885-
394.pdf.

Anonymous:1885:AIbp

[7233] Anonymous. Agricultural inventions. Scientific American, 53 (25):394, December 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n25/pdf/scientificamerican12191885-
394a.pdf.

Anonymous:1885:MIcb

[7234] Anonymous. Miscellaneous inventions. Scientific American, 53 (25):394, December 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n25/pdf/scientificamerican12191885-
394b.pdf.

Anonymous:1885:Sp

[7235] Anonymous. Special. Scientific American, 53(25):394, December 19, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n25/pdf/scientificamerican12191885-394c.pdf.

Anonymous:1885:BPbs

[7236] Anonymous. Business and personal. Scientific American, 53(3):41, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n3/pdf/scientificamerican07181885-41c.pdf.

Anonymous:1885:BNBba

[7237] Anonymous. Books: New books and publications. Scientific American, 53(3):42, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n3/pdf/scientificamerican07181885-
42.pdf.

Anonymous:1885:NQbr

[7238] Anonymous. Notes and queries. Scientific American, 53(3):42–43, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n3/pdf/scientificamerican07181885-42a.pdf.

Anonymous:1885:IIbr

[7239] Anonymous. Index of inventions. Scientific American, 53(3):43–44, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n3/pdf/scientificamerican07181885-43.pdf.

Anonymous:1885:PNPbq

[7240] Anonymous. Patent notes: Patents. Scientific American, 53(3):46, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n3/pdf/scientificamerican07181885-46.pdf.

Anonymous:1885:LBW

[7241] Anonymous. The latest British war ships. Scientific American, 53(3):31, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n3/pdf/scientificamerican07181885-31.pdf.

Anonymous:1885:ATD

[7242] Anonymous. Another telephone decision. Scientific American, 53(3):32, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n3/pdf/scientificamerican07181885-32.pdf.

Anonymous:1885:HKB

[7243] Anonymous. House knowledge for boys. Scientific American, 53(3):32, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n3/pdf/scientificamerican07181885-32a.pdf.

Anonymous:1885:NCPb

[7244] Anonymous. The normal concert pitch. Scientific American, 53(3):32, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n3/pdf/scientificamerican07181885-32b.pdf.

Anonymous:1885:SLD

[7245] Anonymous. Single lipped drills. Scientific American, 53(3):32–33, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n3/pdf/scientificamerican07181885-32c.pdf.

Anonymous:1885:ARC

[7246] Anonymous. Automatic rocking cradle. Scientific American, 53(3):34, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n3/pdf/scientificamerican07181885-34.pdf.

Anonymous:1885:DFA

[7247] Anonymous. Device for feeding air to furnaces. Scientific American, 53(3):34, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n3/pdf/scientificamerican07181885-
34a.pdf.

Anonymous:1885:IFf

[7248] Anonymous. Improved funnel. Scientific American, 53(3):34, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n3/pdf/scientificamerican07181885-34b.pdf.

Anonymous:1885:ICGb

[7249] Anonymous. Improved cable grip. Scientific American, 53(3):34, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n3/pdf/scientificamerican07181885-34c.pdf.

Anonymous:1885:LOa

[7250] Anonymous. The liquefaction of oxygen. Scientific American, 53(3):35, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n3/pdf/scientificamerican07181885-35.pdf.

Anonymous:1885:TAM

[7251] Anonymous. Telephone apparatus for military purposes. Scientific American, 53(3):35, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n3/pdf/scientificamerican07181885-
35a.pdf.

Anonymous:1885:LPL

[7252] Anonymous. Loose pulley lubricator. Scientific American, 53(3):36, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n3/pdf/scientificamerican07181885-36.pdf.

Anonymous:1885:FFC

[7253] Anonymous. Fastening for freight car doors. Scientific American, 53(3):36, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n3/pdf/scientificamerican07181885-
36a.pdf.

Anonymous:1885:SFRb

[7254] Anonymous. The smallest firing revolver. Scientific American, 53(3):36, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n3/pdf/scientificamerican07181885-36b.pdf.

Anonymous:1885:TRb

[7255] Anonymous. Towel rack. Scientific American, 53(3):36–37, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n3/pdf/scientificamerican07181885-36c.pdf.

Anonymous:1885:EFb

[7256] Anonymous. Electric fan. Scientific American, 53(3):37–38, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n3/pdf/scientificamerican07181885-37.pdf.

Anonymous:1885:GBD

[7257] Anonymous. The great blowers of the Denain and Anzin Forges. Scientific American, 53(3):38, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n3/pdf/scientificamerican07181885-
38.pdf.

Anonymous:1885:GF

[7258] Anonymous. Gigantic flowers. Scientific American, 53(3):39–40, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n3/pdf/scientificamerican07181885-39.pdf.

Anonymous:1885:EIbp

[7259] Anonymous. Engineering inventions. Scientific American, 53(3):41, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n3/pdf/scientificamerican07181885-41.pdf.

Anonymous:1885:AIbq

[7260] Anonymous. Agricultural inventions. Scientific American, 53(3):41, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n3/pdf/scientificamerican07181885-41a.pdf.

Anonymous:1885:MIcc

[7261] Anonymous. Miscellaneous inventions. Scientific American, 53(3):41, July 18, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n3/pdf/scientificamerican07181885-41b.pdf.

Anonymous:1885:BPbt

[7262] Anonymous. Business and personal. Scientific American, 53(4):58–59, July 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n4/pdf/scientificamerican07251885-58c.pdf.

Anonymous:1885:NQbs

[7263] Anonymous. Notes and queries. Scientific American, 53(4):59, July 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n4/pdf/scientificamerican07251885-59.pdf.

Anonymous:1885:IIbs

[7264] Anonymous. Index of inventions. Scientific American, 53(4):59–60, July 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n4/pdf/scientificamerican07251885-59a.pdf.

Anonymous:1885:PNPbr

[7265] Anonymous. Patent notes: Patents. Scientific American, 53(4):62, July 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n4/pdf/scientificamerican07251885-62.pdf.

Anonymous:1885:LB

[7266] Anonymous. The lead bath. Scientific American, 53(4):48–49, July 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n4/pdf/scientificamerican07251885-48.pdf.

Anonymous:1885:LAE

[7267] Anonymous. Lippmann’s apparatus for electric measurements. Scientific American, 53(4):50, July 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n4/pdf/scientificamerican07251885-
50.pdf.

Anonymous:1885:ILV

[7268] Anonymous. Improved liming vat. Scientific American, 53(4):50, July 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n4/pdf/scientificamerican07251885-50a.pdf.

Anonymous:1885:IRC

[7269] Anonymous. Improved rail chair. Scientific American, 53(4):51, July 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n4/pdf/scientificamerican07251885-51.pdf.

Anonymous:1885:IBC

[7270] Anonymous. Improved boiler cleanser. Scientific American, 53(4):51, July 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n4/pdf/scientificamerican07251885-51a.pdf.

Anonymous:1885:ILb

[7271] Anonymous. Improved lubricator. Scientific American, 53(4):51–52, July 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n4/pdf/scientificamerican07251885-51b.pdf.

Anonymous:1885:RH

[7272] Anonymous. Rein holder. Scientific American, 53(4):52, July 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n4/pdf/scientificamerican07251885-52.pdf.

Anonymous:1885:FOR

[7273] Anonymous. Final operations for the removal of flood rock. Scientific American, 53(4):53–54, July 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n4/pdf/scientificamerican07251885-
53.pdf.

Hopkins:1885:MMAb

[7274] M. Hopkins. The microscope in the mechanic arts. Scientific American, 53(4):54, July 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n4/pdf/scientificamerican07251885-
54.pdf.

Anonymous:1885:PCH

[7275] Anonymous. A portable covered hammock. Scientific American, 53(4):55, July 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n4/pdf/scientificamerican07251885-
55.pdf.

Anonymous:1885:BSP

[7276] Anonymous. Baudre’s silex piano. Scientific American, 53(4):55–56, July 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n4/pdf/scientificamerican07251885-55a.pdf.

Ebard:1885:IHI

[7277] Daniel C. Ebard. The iguana and how it is caught. Scientific American, 53(4):57, July 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n4/pdf/scientificamerican07251885-
57.pdf.

Anonymous:1885:EIbq

[7278] Anonymous. Engineering inventions. Scientific American, 53(4):58, July 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n4/pdf/scientificamerican07251885-58.pdf.

Anonymous:1885:AIbr

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n4/pdf/scientificamerican07251885-58a.pdf.

Anonymous:1885:MIcd

[7280] Anonymous. Miscellaneous inventions. Scientific American, 53(4):58, July 25, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n4/pdf/scientificamerican07251885-58b.pdf.

Anonymous:1885:Cbi

[7281] Anonymous. Correspondence. Scientific American, 53(5):69, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n5/pdf/scientificamerican08011885-69.pdf.

Anonymous:1885:BNBbb

[7282] Anonymous. Books: New books and publications. Scientific American, 53(5):74, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n5/pdf/scientificamerican08011885-
74c.pdf.

Anonymous:1885:BPbu

[7283] Anonymous. Business and personal. Scientific American, 53(5):74, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n5/pdf/scientificamerican08011885-74d.pdf.

Anonymous:1885:NQbt

[7284] Anonymous. Notes and queries. Scientific American, 53(5):74–75, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n5/pdf/scientificamerican08011885-74e.pdf.

Anonymous:1885:IIbt

[7285] Anonymous. Index of inventions. Scientific American, 53(5):75–76, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n5/pdf/scientificamerican08011885-75.pdf.

Anonymous:1885:PNPbs

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n5/pdf/scientificamerican08011885-78.pdf.

Anonymous:1885:PSRb

[7287] Anonymous. Plans of the suburban rapid transit co. Scientific American, 53(5):63–64, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n5/pdf/scientificamerican08011885-
63.pdf.

Anonymous:1885:GG

[7288] Anonymous. General Grant. Scientific American, 53(5):64, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n5/pdf/scientificamerican08011885-64.pdf.

Anonymous:1885:PHW

[7289] Anonymous. Peter H. Watson. Scientific American, 53(5):64–65, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n5/pdf/scientificamerican08011885-64a.pdf.

Anonymous:1885:PVT

[7290] Anonymous. Panama vs Tehuantepec. Scientific American, 53(5):65, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n5/pdf/scientificamerican08011885-65.pdf.

Anonymous:1885:PGS

[7291] Anonymous. Pressure at great sea depths. Scientific American, 53(5):65, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n5/pdf/scientificamerican08011885-
65a.pdf.

Anonymous:1885:LOb

[7292] Anonymous. The legal ohm. Scientific American, 53(5):66, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n5/pdf/scientificamerican08011885-66.pdf.

Anonymous:1885:ILT

[7293] Anonymous. Improved lawn tennis post. Scientific American, 53(5):66, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n5/pdf/scientificamerican08011885-66a.pdf.

Anonymous:1885:IEBb

[7294] Anonymous. Improved egg beater. Scientific American, 53(5):67, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n5/pdf/scientificamerican08011885-67.pdf.

Anonymous:1885:ICd

[7295] Anonymous. Improved caisson. Scientific American, 53(5):67, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n5/pdf/scientificamerican08011885-67a.pdf.

Anonymous:1885:BDH

[7296] Anonymous. Bundle dropper for harvesters. Scientific American, 53(5):67, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n5/pdf/scientificamerican08011885-
67b.pdf.

Anonymous:1885:RSV

[7297] Anonymous. Rotary steam valve. Scientific American, 53(5):67–68, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n5/pdf/scientificamerican08011885-67c.pdf.

Anonymous:1885:POW

[7298] Anonymous. Pump for oil wells. Scientific American, 53(5):68, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n5/pdf/scientificamerican08011885-68.pdf.

Anonymous:1885:CBSb

[7299] Anonymous. Check for blind slats. Scientific American, 53(5):68, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n5/pdf/scientificamerican08011885-68a.pdf.

Anonymous:1885:BHA

[7300] Anonymous. A button hole attachment for sewing machines. Scientific American, 53(5):68, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n5/pdf/scientificamerican08011885-
68b.pdf.

Anonymous:1885:BPTa

[7301] Anonymous. Baillehache’s printing telegraph. Scientific American, 53(5):70, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n5/pdf/scientificamerican08011885-
70.pdf.

Anonymous:1885:MTS

[7302] Anonymous. Maclaine’s twin screw propellers. Scientific American, 53(5):71, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n5/pdf/scientificamerican08011885-
71.pdf.

Anonymous:1885:ADb

[7303] Anonymous. The American Dicentras. Scientific American, 53 (5):71–72, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n5/pdf/scientificamerican08011885-
71a.pdf.

Anonymous:1885:HP

[7304] Anonymous. The Houses of Parliament. Scientific American, 53(5):73, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n5/pdf/scientificamerican08011885-73.pdf.

Anonymous:1885:EIbr

[7305] Anonymous. Engineering inventions. Scientific American, 53(5):74, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n5/pdf/scientificamerican08011885-74.pdf.

Anonymous:1885:AIbs

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n5/pdf/scientificamerican08011885-74a.pdf.

Anonymous:1885:MIce

[7307] Anonymous. Miscellaneous inventions. Scientific American, 53(5):74, August 1, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n5/pdf/scientificamerican08011885-74b.pdf.

Anonymous:1885:BNBbc

[7308] Anonymous. Books: New books and publications. Scientific American, 53(6):90, August 8, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n6/pdf/scientificamerican08081885-
90c.pdf.

Anonymous:1885:BPbv

[7309] Anonymous. Business and personal. Scientific American, 53(6):90, August 8, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n6/pdf/scientificamerican08081885-90d.pdf.

Anonymous:1885:NQbu

[7310] Anonymous. Notes and queries. Scientific American, 53(6):91, August 8, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n6/pdf/scientificamerican08081885-91.pdf.

Anonymous:1885:IIbu

[7311] Anonymous. Index of inventions. Scientific American, 53(6):91–92, August 8, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n6/pdf/scientificamerican08081885-91a.pdf.

Anonymous:1885:PNPbt

[7312] Anonymous. Patent notes: Patents. Scientific American, 53(6):94, August 8, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n6/pdf/scientificamerican08081885-94.pdf.

Anonymous:1885:RPB

[7313] Anonymous. Riverside Park — the burial place of Gen Grant. Scientific American, 53(6):80, August 8, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n6/pdf/scientificamerican08081885-
80.pdf.

Anonymous:1885:CGC

[7314] Anonymous. Charts for great circle sailing. Scientific American, 53(6):80, August 8, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n6/pdf/scientificamerican08081885-
80a.pdf.

Anonymous:1885:BFe

[7315] Anonymous. Bronze frames. Scientific American, 53(6):80–81, August 8, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n6/pdf/scientificamerican08081885-80b.pdf.

Anonymous:1885:COI

[7316] Anonymous. A curious optical illusion. Scientific American, 53(6):82, August 8, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n6/pdf/scientificamerican08081885-82.pdf.

Anonymous:1885:WES

[7317] Anonymous. Weber’s electric siren. Scientific American, 53(6):82, August 8, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n6/pdf/scientificamerican08081885-82a.pdf.

Anonymous:1885:DHS

[7318] Anonymous. Device for handling stamps. Scientific American, 53(6):83, August 8, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n6/pdf/scientificamerican08081885-83.pdf.

Anonymous:1885:BIA

[7319] Anonymous. Burgess’ improved atomizer. Scientific American, 53(6):83, August 8, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n6/pdf/scientificamerican08081885-
83a.pdf.

Anonymous:1885:IDB

[7320] Anonymous. Improved door bolt. Scientific American, 53(6):83, August 8, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n6/pdf/scientificamerican08081885-83b.pdf.

Anonymous:1885:ML

[7321] Anonymous. Magic lantern. Scientific American, 53(6):83–84, August 8, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n6/pdf/scientificamerican08081885-83c.pdf.

Anonymous:1885:ILc

[7322] Anonymous. Improved ladder. Scientific American, 53(6):84, August 8, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n6/pdf/scientificamerican08081885-84.pdf.

Anonymous:1885:DRPd

[7323] Anonymous. Decisions relating to patents. Scientific American, 53(6):84, August 8, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n6/pdf/scientificamerican08081885-
84a.pdf.

Anonymous:1885:APAb

[7324] Anonymous. Aspects of the planets for August. Scientific American, 53(6):85, August 8, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n6/pdf/scientificamerican08081885-
85.pdf.

Anonymous:1885:SPF

[7325] Anonymous. Surgery for piano-forte players. Scientific American, 53(6):86, August 8, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n6/pdf/scientificamerican08081885-
86.pdf.

Anonymous:1885:FHH

[7326] Anonymous. Fifty horse horizontal engine. Scientific American, 53(6):86, August 8, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n6/pdf/scientificamerican08081885-
86a.pdf.

Anonymous:1885:RWN

[7327] Anonymous. The right whale of the North Atlantic. Scientific American, 53(6):87–89, August 8, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n6/pdf/scientificamerican08081885-
87.pdf.

Anonymous:1885:EIbs

[7328] Anonymous. Engineering inventions. Scientific American, 53(6):90, August 8, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n6/pdf/scientificamerican08081885-90.pdf.

Anonymous:1885:AIbt

[7329] Anonymous. Agricultural inventions. Scientific American, 53(6):90, August 8, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n6/pdf/scientificamerican08081885-90a.pdf.

Anonymous:1885:MIcf

[7330] Anonymous. Miscellaneous inventions. Scientific American, 53(6):90, August 8, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n6/pdf/scientificamerican08081885-90b.pdf.

Anonymous:1885:Cbj

[7331] Anonymous. Correspondence. Scientific American, 53(7):101, August 15, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n7/pdf/scientificamerican08151885-101.pdf.

Anonymous:1885:BPbw

[7332] Anonymous. Business and personal. Scientific American, 53(7):106, August 15, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n7/pdf/scientificamerican08151885-106c.pdf.

Anonymous:1885:NQbv

[7333] Anonymous. Notes and queries. Scientific American, 53(7):107, August 15, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n7/pdf/scientificamerican08151885-107.pdf.

Anonymous:1885:IIbv

[7334] Anonymous. Index of inventions. Scientific American, 53(7): 107–108, August 15, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n7/pdf/scientificamerican08151885-
107a.pdf.

Anonymous:1885:PNPbu

[7335] Anonymous. Patent notes: Patents. Scientific American, 53(7):110, August 15, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n7/pdf/scientificamerican08151885-110.pdf.

Anonymous:1885:GAD

[7336] Anonymous. The great American dredges on the Panama Canal. Scientific American, 53(7):95, August 15, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v53/n7/pdf/scientificamerican08151885-
95.pdf.

Anonymous:1885:NGW

[7337] Anonymous. Natural gas wells in Ohio. Scientific American, 53 (7):96, August 15, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n7/pdf/scientificamerican08151885-
96.pdf.

Anonymous:1885:SOC

[7338] Anonymous. Shall our canals be maintained? Scientific American, 53(7):96–97, August 15, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n7/pdf/scientificamerican08151885-
96a.pdf.

Anonymous:1885:WCb

[7339] Anonymous. Wedge cutter. Scientific American, 53(7):98, August 15, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n7/pdf/scientificamerican08151885-98.pdf.

Anonymous:1885:AH

[7340] Anonymous. An adjustable harrow. Scientific American, 53(7):98, August 15, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n7/pdf/scientificamerican08151885-98a.pdf.

Anonymous:1885:MCS

[7341] Anonymous. Machine for cutting scrub, cane, etc. Scientific American, 53(7):98, August 15, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n7/pdf/scientificamerican08151885-
98b.pdf.

Anonymous:1885:EGF

[7342] Anonymous. End gate fastening. Scientific American, 53(7):99, August 15, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n7/pdf/scientificamerican08151885-99.pdf.

Anonymous:1885:SFL

[7343] Anonymous. Sand feeder for locomotives. Scientific American, 53(7):99, August 15, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n7/pdf/scientificamerican08151885-
99a.pdf.

Anonymous:1885:IMM

[7344] Anonymous. An improved mechanical movement. Scientific American, 53(7):99, August 15, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n7/pdf/scientificamerican08151885-
99b.pdf.

Anonymous:1885:HAF

[7345] Anonymous. Hot air furnace. Scientific American, 53(7):99–100, August 15, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n7/pdf/scientificamerican08151885-99c.pdf.

Anonymous:1885:IFP

[7346] Anonymous. Improved fence post for wire fences. Scientific American, 53(7):100, August 15, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n7/pdf/scientificamerican08151885-
100.pdf.

Anonymous:1885:GB

[7347] Anonymous. A guard for bedsteads. Scientific American, 53(7):100, August 15, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n7/pdf/scientificamerican08151885-100a.pdf.

Anonymous:1885:MCd

[7348] Anonymous. A muzzle for cattle. Scientific American, 53(7):100, August 15, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n7/pdf/scientificamerican08151885-100b.pdf.

Anonymous:1885:FDT

[7349] Anonymous. The French dispatch torpedo boat La Bombe. Scientific American, 53(7):103–104, August 15, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v53/n7/pdf/scientificamerican08151885-
103.pdf.

Anonymous:1885:EMR

[7350] Anonymous. An English military railway. Scientific American, 53(7):105, August 15, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n7/pdf/scientificamerican08151885-
105.pdf.

Anonymous:1885:EIbt

[7351] Anonymous. Engineering inventions. Scientific American, 53(7): 106, August 15, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n7/pdf/scientificamerican08151885-
106.pdf.

Anonymous:1885:AIbu (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n7/pdf/scientificamerican08151885-
106a.pdf.

Anonymous:1885:MIcg

[7353] Anonymous. Miscellaneous inventions. Scientific American, 53 (7):106, August 15, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n7/pdf/scientificamerican08151885-
106b.pdf.

Anonymous:1885:Cbk

[7354] Anonymous. Correspondence. Scientific American, 53(8):117, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n8/pdf/scientificamerican08221885-117.pdf.

Anonymous:1885:BNBbd

[7355] Anonymous. Books: New books and publications. Scientific American, 53(8):122, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n8/pdf/scientificamerican08221885-
122c.pdf.

Anonymous:1885:BPbx

[7356] Anonymous. Business and personal. Scientific American, 53(8): 122–123, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n8/pdf/scientificamerican08221885-
122d.pdf.

Anonymous:1885:NQbw

[7357] Anonymous. Notes and queries. Scientific American, 53(8):123, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n8/pdf/scientificamerican08221885-123.pdf.

Anonymous:1885:IIbw

[7358] Anonymous. Index of inventions. Scientific American, 53(8): 123–126, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n8/pdf/scientificamerican08221885-
123a.pdf.

Anonymous:1885:PNPbv

[7359] Anonymous. Patent notes: Patents. Scientific American, 53(8):126, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n8/pdf/scientificamerican08221885-126.pdf.

Anonymous:1885:PC

[7360] Anonymous. The Panama Canal. Scientific American, 53(8): 111–112, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n8/pdf/scientificamerican08221885-
111.pdf.

Anonymous:1885:MPc

[7361] Anonymous. Manufacturing prospects. Scientific American, 53 (8):112, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n8/pdf/scientificamerican08221885-
112.pdf.

Anonymous:1885:THC

[7362] Anonymous. The triangulation and hydrography of the coast. Scientific American, 53(8):112–113, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v53/n8/pdf/scientificamerican08221885-
112a.pdf.

Anonymous:1885:RT

[7363] Anonymous. The Reis telephones. Scientific American, 53(8):113, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n8/pdf/scientificamerican08221885-113.pdf.

Anonymous:1885:CCh

[7364] Anonymous. Car coupling. Scientific American, 53(8):114, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n8/pdf/scientificamerican08221885-114.pdf.

Anonymous:1885:WSF

[7365] Anonymous. Window sash and frame. Scientific American, 53 (8):114, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n8/pdf/scientificamerican08221885-
114a.pdf.

Anonymous:1885:Sq

[7366] Anonymous. The skiagraph. Scientific American, 53(8):114, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n8/pdf/scientificamerican08221885-114b.pdf.

Anonymous:1885:SCFc

[7367] Anonymous. Smoke-consuming furnace. Scientific American, 53 (8):114–115, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n8/pdf/scientificamerican08221885-
114c.pdf.

Anonymous:1885:CBHa

[7368] Anonymous. Combined bag holder and truck. Scientific American, 53(8):115, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n8/pdf/scientificamerican08221885-
115.pdf.

Anonymous:1885:FEc

[7369] Anonymous. Fire escape. Scientific American, 53(8):115, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n8/pdf/scientificamerican08221885-115a.pdf.

Anonymous:1885:HAa

[7370] Anonymous. Harness attachment. Scientific American, 53(8):115, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n8/pdf/scientificamerican08221885-115b.pdf.

Anonymous:1885:SAV

[7371] Anonymous. Steam actuated valve. Scientific American, 53(8):115, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n8/pdf/scientificamerican08221885-115c.pdf.

Anonymous:1885:IEBc

[7372] Anonymous. Improved egg beater. Scientific American, 53(8):116, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n8/pdf/scientificamerican08221885-116.pdf.

Anonymous:1885:IKF

[7373] Anonymous. Improved knockdown furniture. Scientific American, 53(8):116, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n8/pdf/scientificamerican08221885-
116a.pdf.

Anonymous:1885:INLb

[7374] Anonymous. Improved nut lock. Scientific American, 53(8):116, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n8/pdf/scientificamerican08221885-116b.pdf.

Anonymous:1885:TVS

[7375] Anonymous. “Trains” of vessels on the Suez Canal. Scientific American, 53(8):118, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n8/pdf/scientificamerican08221885-
118.pdf.

Anonymous:1885:NCL

[7376] Anonymous. New Chemical Laboratory, Cambridge University. Scientific American, 53(8):119–121, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v53/n8/pdf/scientificamerican08221885-
119.pdf.

Anonymous:1885:BH

[7377] Anonymous. A Belgian horse. Scientific American, 53(8):121, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n8/pdf/scientificamerican08221885-121.pdf.

Anonymous:1885:EIbu (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n8/pdf/scientificamerican08221885-
122.pdf.

Anonymous:1885:AIbv

[7379] Anonymous. Agricultural inventions. Scientific American, 53(8): 122, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n8/pdf/scientificamerican08221885-
122a.pdf.

Anonymous:1885:MIch

[7380] Anonymous. Miscellaneous inventions. Scientific American, 53 (8):122, August 22, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n8/pdf/scientificamerican08221885-
122b.pdf.

Anonymous:1885:Cbl

[7381] Anonymous. Correspondence. Scientific American, 53(9):133, August 29, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n9/pdf/scientificamerican08291885-133.pdf.

Anonymous:1885:BNBbe

[7382] Anonymous. Books: New books and publications. Scientific American, 53(9):138, August 29, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n9/pdf/scientificamerican08291885-
138c.pdf.

Anonymous:1885:BPby

[7383] Anonymous. Business and personal. Scientific American, 53(9):138, August 29, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n9/pdf/scientificamerican08291885-138d.pdf.

Anonymous:1885:NQbx

[7384] Anonymous. Notes and queries. Scientific American, 53(9):138– 139, August 29, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n9/pdf/scientificamerican08291885-
138e.pdf.

Anonymous:1885:IIbx

[7385] Anonymous. Index of inventions. Scientific American, 53(9): 139–140, August 29, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n9/pdf/scientificamerican08291885-
139.pdf.

Anonymous:1885:PNPbw

[7386] Anonymous. Patent notes: Patents. Scientific American, 53(9):142, August 29, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n9/pdf/scientificamerican08291885-142.pdf.

Anonymous:1885:AT

[7387] Anonymous. The autographic telegraph. Scientific American, 53 (9):127, August 29, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n9/pdf/scientificamerican08291885-
127.pdf.

Anonymous:1885:MNE

[7388] Anonymous. Meeting of the National Electric Light Association. Scientific American, 53(9):128, August 29, 1885. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v53/
n9/pdf/scientificamerican08291885-128.pdf.

Anonymous:1885:CSb

[7389] Anonymous. Car seats. Scientific American, 53(9):128–129, August 29, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n9/pdf/scientificamerican08291885-128a.pdf.

Anonymous:1885:APS

[7390] Anonymous. Aspects of the planets for September. Scientific American, 53(9):129, August 29, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n9/pdf/scientificamerican08291885-
129.pdf.

Anonymous:1885:CVa

[7391] Anonymous. Compound valve. Scientific American, 53(9):130, August 29, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n9/pdf/scientificamerican08291885-130.pdf.

Anonymous:1885:BCc

[7392] Anonymous. Boiler cleaner. Scientific American, 53(9):130, August 29, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n9/pdf/scientificamerican08291885-130a.pdf.

Anonymous:1885:HTF

[7393] Anonymous. Hose tower and fire escape. Scientific American, 53(9):130, August 29, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n9/pdf/scientificamerican08291885-
130b.pdf.

Anonymous:1885:RCc

[7394] Anonymous. Rocking chair. Scientific American, 53(9):130, August 29, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n9/pdf/scientificamerican08291885-130c.pdf.

Anonymous:1885:IWP

[7395] Anonymous. Improved wheel or pulley. Scientific American, 53 (9):131, August 29, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n9/pdf/scientificamerican08291885-
131.pdf.

Anonymous:1885:HPSb

[7396] Anonymous. A hand power sawing machine. Scientific American, 53(9):131, August 29, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n9/pdf/scientificamerican08291885-
131a.pdf.

Anonymous:1885:CBn

[7397] Anonymous. A car brake. Scientific American, 53(9):131, August 29, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n9/pdf/scientificamerican08291885-131b.pdf.

Anonymous:1885:AMc

[7398] Anonymous. Adding machine. Scientific American, 53(9):132, August 29, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n9/pdf/scientificamerican08291885-132.pdf.

Anonymous:1885:BTW

[7399] Anonymous. Burying telephone wires in the sewers of Paris. Scientific American, 53(9):134–136, August 29, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v53/n9/pdf/scientificamerican08291885-
134.pdf.

Anonymous:1885:CPS

[7400] Anonymous. Combined plow and seeder. Scientific American, 53(9):136, August 29, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n9/pdf/scientificamerican08291885-
136.pdf.

Anonymous:1885:NWM

[7401] Anonymous. A new washing machine. Scientific American, 53 (9):136, August 29, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n9/pdf/scientificamerican08291885-
136a.pdf.

Holder:1885:SWF

[7402] C. F. Holder. Some wonderful fliers. Scientific American, 53(9):137, August 29, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v53/n9/pdf/scientificamerican08291885-137.pdf.

Anonymous:1885:EIbv

[7403] Anonymous. Engineering inventions. Scientific American, 53(9): 138, August 29, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n9/pdf/scientificamerican08291885-
138.pdf.

Anonymous:1885:AIbw (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n9/pdf/scientificamerican08291885-
138a.pdf.

Anonymous:1885:MIci

[7405] Anonymous. Miscellaneous inventions. Scientific American, 53 (9):138, August 29, 1885. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v53/n9/pdf/scientificamerican08291885-
138b.pdf.

Anonymous:1886:Ca

[7406] Anonymous. Correspondence. Scientific American, 54(1):7, January 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n1/pdf/scientificamerican01021886-7.pdf.

Anonymous:1886:BNBa

[7407] Anonymous. Books: New books and publications. Scientific American, 54(1):10, January 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n1/pdf/scientificamerican01021886-
10c.pdf.

Anonymous:1886:BPa

[7408] Anonymous. Business and personal. Scientific American, 54(1):10, January 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n1/pdf/scientificamerican01021886-10d.pdf.

Anonymous:1886:NQa

[7409] Anonymous. Notes and queries. Scientific American, 54(1):10–11, January 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n1/pdf/scientificamerican01021886-10e.pdf.

Anonymous:1886:IIa

[7410] Anonymous. Index of inventions. Scientific American, 54(1):11–13, January 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n1/pdf/scientificamerican01021886-11.pdf.

Anonymous:1886:PNPa

[7411] Anonymous. Patent notes: Patents. Scientific American, 54(1):14, January 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n1/pdf/scientificamerican01021886-14.pdf.

Anonymous:1886:ICa

[7412] Anonymous. Improve the channelways. Scientific American, 54(1):2, January 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n1/pdf/scientificamerican01021886-2.pdf.

Anonymous:1886:BTCa

[7413] Anonymous. The Bell Telephone Case. Scientific American, 54(1):2, January 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n1/pdf/scientificamerican01021886-2a.pdf.

Proctor:1886:NSD

[7414] Richard A. Proctor. Night sky — December and January. Scientific American, 54(1):3, January 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n1/pdf/scientificamerican01021886-
3.pdf.

Anonymous:1886:FEL

[7415] Anonymous. Fire escape ladder. Scientific American, 54(1):4, January 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n1/pdf/scientificamerican01021886-4.pdf.

Anonymous:1886:ISa

[7416] Anonymous. Improved square. Scientific American, 54(1):4, January 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n1/pdf/scientificamerican01021886-4a.pdf.

Anonymous:1886:FFP

[7417] Anonymous. Farm fence post. Scientific American, 54(1):4, January 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n1/pdf/scientificamerican01021886-4b.pdf.

Anonymous:1886:SRI

[7418] Anonymous. A separator to remove impurities from wheat. Scientific American, 54(1):4, January 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n1/pdf/scientificamerican01021886-
4c.pdf.

Anonymous:1886:ADE

[7419] Anonymous. Automatic discharger for evaporating pans. Scientific American, 54(1):4, January 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n1/pdf/scientificamerican01021886-
4d.pdf.

Anonymous:1886:IM

[7420] Anonymous. Improved motor. Scientific American, 54(1):5, January 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n1/pdf/scientificamerican01021886-5.pdf.

Anonymous:1886:JIM

[7421] Anonymous. The Jarman ice machine. Scientific American, 54(1):5, January 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n1/pdf/scientificamerican01021886-5a.pdf.

Anonymous:1886:IGM

[7422] Anonymous. Improved grinding mill. Scientific American, 54(1):5, January 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n1/pdf/scientificamerican01021886-5b.pdf.

Anonymous:1886:HMO

[7423] Anonymous. Hydraulic machinery for operating a drawbridge. Scientific American, 54(1):6, January 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n1/pdf/scientificamerican01021886-
6.pdf.

Anonymous:1886:VFM

[7424] Anonymous. Vertical flour mill engines. Scientific American, 54(1):8–9, January 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n1/pdf/scientificamerican01021886-8.pdf.

Anonymous:1886:ESa

[7425] Anonymous. Electricity at the Salpetriere. Scientific American, 54(1):9, January 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n1/pdf/scientificamerican01021886-9.pdf.

Anonymous:1886:EIa

[7426] Anonymous. Engineering inventions. Scientific American, 54(1):10, January 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n1/pdf/scientificamerican01021886-10.pdf.

Anonymous:1886:AIa

[7427] Anonymous. Agricultural inventions. Scientific American, 54(1):10, January 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n1/pdf/scientificamerican01021886-10a.pdf.

Anonymous:1886:MIa

[7428] Anonymous. Miscellaneous inventions. Scientific American, 54(1):10, January 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n1/pdf/scientificamerican01021886-10b.pdf.

Anonymous:1886:BNBb

[7429] Anonymous. Books: New books and publications. Scientific American, 54(10):153, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n10/pdf/scientificamerican03061886-
153c.pdf.

Anonymous:1886:BPb

[7430] Anonymous. Business and personal. Scientific American, 54(10):154, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n10/pdf/scientificamerican03061886-154a.pdf.

Anonymous:1886:NQb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n10/pdf/scientificamerican03061886-154b.pdf.

Anonymous:1886:IIb

[7432] Anonymous. Index of inventions. Scientific American, 54(10):155–156, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n10/pdf/scientificamerican03061886-155.pdf.

Anonymous:1886:PNPb

[7433] Anonymous. Patent notes: Patents. Scientific American, 54(10):158, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n10/pdf/scientificamerican03061886-158.pdf.

Anonymous:1886:PHR

[7434] Anonymous. The proposed Harlem River Bridge. Scientific American, 54(10):143–144, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n10/pdf/scientificamerican03061886-
143.pdf.

Anonymous:1886:RSFa

[7435] Anonymous. A remarkable snow flow. Scientific American, 54(10):144, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n10/pdf/scientificamerican03061886-144.pdf.

Anonymous:1886:PFP

[7436] Anonymous. Paint, finish, and polish. Scientific American, 54 (10):144–145, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n10/pdf/scientificamerican03061886-
144a.pdf.

Proctor:1886:NSF

[7437] Richard A. Proctor. Night sky — February and March. Scientific American, 54(10):145, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n10/pdf/scientificamerican03061886-
145.pdf.

Anonymous:1886:ICH

[7438] Anonymous. Improved counting-house stamp. Scientific American, 54(10):146, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n10/pdf/scientificamerican03061886-
146.pdf.

Anonymous:1886:ICb

[7439] Anonymous. An improved churn. Scientific American, 54(10):146, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n10/pdf/scientificamerican03061886-146a.pdf.

Anonymous:1886:APT

[7440] Anonymous. An attachment for poles of two-horse wagons. Scientific American, 54(10):146, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n10/pdf/scientificamerican03061886-
146b.pdf.

Anonymous:1886:NDM

[7441] Anonymous. A novel device for milking cows. Scientific American, 54(10):146, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n10/pdf/scientificamerican03061886-
146c.pdf.

Anonymous:1886:SVW

[7442] Anonymous. Sanitary valve for wash basins, etc. Scientific American, 54(10):146, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n10/pdf/scientificamerican03061886-
146d.pdf.

Anonymous:1886:DQF

[7443] Anonymous. Davis’ quilting frame. Scientific American, 54(10):147, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n10/pdf/scientificamerican03061886-147.pdf.

Anonymous:1886:HWS

[7444] Anonymous. The helicoidal or wire stone saw. Scientific American, 54(10):147–149, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n10/pdf/scientificamerican03061886-
147a.pdf.

Anonymous:1886:TSRa

[7445] Anonymous. Thurston’s standard railroad oil testing machine. Scientific American, 54(10):149, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n10/pdf/scientificamerican03061886-
149.pdf.

Anonymous:1886:CWM

[7446] Anonymous. Caligraph writing machines. Scientific American, 54 (10):150–151, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n10/pdf/scientificamerican03061886-
150.pdf.

Anonymous:1886:PSCa

[7447] Anonymous. Palmers’ steam carriage. Scientific American, 54(10):151, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n10/pdf/scientificamerican03061886-151.pdf.

Seiss:1886:OC

[7448] C. Few Seiss. The oyster crab. Scientific American, 54(10):151, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n10/pdf/scientificamerican03061886-151a.pdf.

Anonymous:1886:NHNa

[7449] Anonymous. Natural history notes. Scientific American, 54(10): 151–152, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n10/pdf/scientificamerican03061886-
151b.pdf.

Anonymous:1886:THN

[7450] Anonymous. Tree houses in New Guinea. Scientific American, 54(10):152, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n10/pdf/scientificamerican03061886-
152.pdf.

Anonymous:1886:RDRa

[7451] Anonymous. Recent decisions relating to patents. Scientific American, 54(10):152, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n10/pdf/scientificamerican03061886-
152a.pdf.

Anonymous:1886:EIb

[7452] Anonymous. Engineering inventions. Scientific American, 54(10):153, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n10/pdf/scientificamerican03061886-153.pdf.

Anonymous:1886:AIb

[7453] Anonymous. Agricultural inventions. Scientific American, 54(10):153, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n10/pdf/scientificamerican03061886-153a.pdf.

Anonymous:1886:MIb

[7454] Anonymous. Miscellaneous inventions. Scientific American, 54(10):153, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n10/pdf/scientificamerican03061886-153b.pdf.

Anonymous:1886:Ia

[7455] Anonymous. Insurance. Scientific American, 54(10):154, March 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n10/pdf/scientificamerican03061886-154.pdf.

Anonymous:1886:Cb

[7456] Anonymous. Correspondence. Scientific American, 54(11):165–166, March 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n11/pdf/scientificamerican03131886-165.pdf.

Anonymous:1886:BNBc

[7457] Anonymous. Books: New books and publications. Scientific American, 54(11):169, March 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n11/pdf/scientificamerican03131886-
169b.pdf.

Anonymous:1886:BPc

[7458] Anonymous. Business and personal. Scientific American, 54(11):169, March 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n11/pdf/scientificamerican03131886-169c.pdf.

Anonymous:1886:NQc

[7459] Anonymous. Notes and queries. Scientific American, 54(11):170–171, March 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n11/pdf/scientificamerican03131886-170.pdf.

Anonymous:1886:IIc

[7460] Anonymous. Index of inventions. Scientific American, 54(11):171–172, March 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n11/pdf/scientificamerican03131886-171.pdf.

Anonymous:1886:PNPc

[7461] Anonymous. Patent notes: Patents. Scientific American, 54(11):174, March 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n11/pdf/scientificamerican03131886-174.pdf.

Anonymous:1886:MT

[7462] Anonymous. The Mersey Tunnel. Scientific American, 54(11):159, March 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n11/pdf/scientificamerican03131886-159.pdf.

Anonymous:1886:UI

[7463] Anonymous. Unsuccessful inventors. Scientific American, 54(11):160, March 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n11/pdf/scientificamerican03131886-160.pdf.

Anonymous:1886:HCS

[7464] Anonymous. Hardening cast steel. Scientific American, 54(11):160–161, March 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n11/pdf/scientificamerican03131886-160a.pdf.

Anonymous:1886:DRPa

[7465] Anonymous. Decisions relating to patents. Scientific American, 54(11):161, March 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n11/pdf/scientificamerican03131886-
161.pdf.

Anonymous:1886:IDa

[7466] Anonymous. Improved desk. Scientific American, 54(11):162, March 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n11/pdf/scientificamerican03131886-162.pdf.

Anonymous:1886:HP

[7467] Anonymous. Hop press. Scientific American, 54(11):162, March 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n11/pdf/scientificamerican03131886-162a.pdf.

Anonymous:1886:ICAa

[7468] Anonymous. Improved cuff adjuster. Scientific American, 54(11):162, March 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n11/pdf/scientificamerican03131886-162b.pdf.

Anonymous:1886:PHC

[7469] Anonymous. Parasol holder for children’s carriages. Scientific American, 54(11):162, March 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n11/pdf/scientificamerican03131886-
162c.pdf.

Anonymous:1886:BMa

[7470] Anonymous. Brushing machine. Scientific American, 54(11):162, March 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n11/pdf/scientificamerican03131886-162d.pdf.

Anonymous:1886:JC

[7471] Anonymous. Jar cover. Scientific American, 54(11):163, March 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n11/pdf/scientificamerican03131886-163.pdf.

Anonymous:1886:GCG

[7472] Anonymous. Glass cutting gauge. Scientific American, 54(11):163, March 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n11/pdf/scientificamerican03131886-163a.pdf.

Anonymous:1886:ESWa

[7473] Anonymous. An English stern wheel steamer for the Nile. Scientific American, 54(11):163–164, March 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v54/n11/pdf/scientificamerican03131886-
163b.pdf.

Anonymous:1886:TIB

[7474] Anonymous. Ten inch breech loading Naval gun. Scientific American, 54(11):166, March 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n11/pdf/scientificamerican03131886-
166.pdf.

Anonymous:1886:CG

[7475] Anonymous. The Cathedral of Granada. Scientific American, 54 (11):166–167, March 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n11/pdf/scientificamerican03131886-
166a.pdf.

Anonymous:1886:PNa

[7476] Anonymous. Photographic notes. Scientific American, 54(11):168, March 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n11/pdf/scientificamerican03131886-168.pdf.

Anonymous:1886:EIc

[7477] Anonymous. Engineering inventions. Scientific American, 54(11):169, March 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n11/pdf/scientificamerican03131886-169.pdf.

Anonymous:1886:MIc

[7478] Anonymous. Miscellaneous inventions. Scientific American, 54(11):169, March 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n11/pdf/scientificamerican03131886-169a.pdf.

Anonymous:1886:Cc

[7479] Anonymous. Correspondence. Scientific American, 54(12):181–182, March 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n12/pdf/scientificamerican03201886-181.pdf.

Anonymous:1886:BNBd

[7480] Anonymous. Books: New books and publications. Scientific American, 54(12):186, March 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n12/pdf/scientificamerican03201886-
186c.pdf.

Anonymous:1886:BPd

[7481] Anonymous. Business and personal. Scientific American, 54 (12):186–187, March 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n12/pdf/scientificamerican03201886-
186d.pdf.

Anonymous:1886:NQd

[7482] Anonymous. Notes and queries. Scientific American, 54(12):187, March 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n12/pdf/scientificamerican03201886-187.pdf.

Anonymous:1886:IId

[7483] Anonymous. Index of inventions. Scientific American, 54(12):187–188, March 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n12/pdf/scientificamerican03201886-187a.pdf.

Anonymous:1886:PNPd

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n12/pdf/scientificamerican03201886-190.pdf.

Anonymous:1886:MWG

[7485] Anonymous. The manufacture of window glass with natural gas. Scientific American, 54(12):175, March 20, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v54/
n12/pdf/scientificamerican03201886-175.pdf.

Anonymous:1886:TMI

[7486] Anonymous. Too much idle capital. Scientific American, 54(12):176, March 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n12/pdf/scientificamerican03201886-176.pdf.

Anonymous:1886:PAE

[7487] Anonymous. Proposed American Exhibition in 1889. Scientific American, 54(12):176, March 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n12/pdf/scientificamerican03201886-
176a.pdf.

Anonymous:1886:PNb

[7488] Anonymous. Photographic notes. Scientific American, 54(12):177, March 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n12/pdf/scientificamerican03201886-177.pdf.

Anonymous:1886:STa

[7489] Anonymous. Saw tooth. Scientific American, 54(12):178, March 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n12/pdf/scientificamerican03201886-178.pdf.

Anonymous:1886:ISPa

[7490] Anonymous. Improved snow plow. Scientific American, 54(12):178, March 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n12/pdf/scientificamerican03201886-178a.pdf.

Anonymous:1886:HCa

[7491] Anonymous. Handle for cans. Scientific American, 54(12):178, March 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n12/pdf/scientificamerican03201886-178b.pdf.

Anonymous:1886:AWW

[7492] Anonymous. Automatic water works. Scientific American, 54(12):178, March 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n12/pdf/scientificamerican03201886-178c.pdf.

Anonymous:1886:PI

[7493] Anonymous. Pruning implement. Scientific American, 54(12):178, March 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n12/pdf/scientificamerican03201886-178d.pdf.

Anonymous:1886:TBC

[7494] Anonymous. Testing bayonets and cavalry swords. Scientific American, 54(12):179, March 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n12/pdf/scientificamerican03201886-
179.pdf.

Anonymous:1886:SCSa

[7495] Anonymous. Screw cutting and self-acting sliding gap lathe. Scientific American, 54(12):179–180, March 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v54/n12/pdf/scientificamerican03201886-
179a.pdf.

Hopkins:1886:MOV

[7496] Geo.M. Hopkins. Microscopic observation of vibrating rods. Scientific American, 54(12):180, March 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n12/pdf/scientificamerican03201886-
180.pdf.

Anonymous:1886:DLT

[7497] Anonymous. The Diamond learner’s telegraph instrument. Scientific American, 54(12):180, March 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n12/pdf/scientificamerican03201886-
180a.pdf.

Anonymous:1886:DSG

[7498] Anonymous. Design for a summer garden house. Scientific American, 54(12):182, March 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n12/pdf/scientificamerican03201886-
182.pdf.

Anonymous:1886:BSC

[7499] Anonymous. Bollee’s steam carriage. Scientific American, 54(12):185, March 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n12/pdf/scientificamerican03201886-185.pdf.

Anonymous:1886:EId

[7500] Anonymous. Engineering inventions. Scientific American, 54(12):186, March 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n12/pdf/scientificamerican03201886-186.pdf.

Anonymous:1886:AIc

[7501] Anonymous. Agricultural inventions. Scientific American, 54(12):186, March 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n12/pdf/scientificamerican03201886-186a.pdf.

Anonymous:1886:MId

[7502] Anonymous. Miscellaneous inventions. Scientific American, 54(12):186, March 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n12/pdf/scientificamerican03201886-186b.pdf.

Anonymous:1886:Cd

[7503] Anonymous. Correspondence. Scientific American, 54(13):197, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n13/pdf/scientificamerican03271886-197.pdf.

Anonymous:1886:BNBe

[7504] Anonymous. Books: New books and publications. Scientific American, 54(13):202, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n13/pdf/scientificamerican03271886-
202c.pdf.

Anonymous:1886:BPe

[7505] Anonymous. Business and personal. Scientific American, 54 (13):202–203, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n13/pdf/scientificamerican03271886-
202d.pdf.

Anonymous:1886:NQe

[7506] Anonymous. Notes and queries. Scientific American, 54(13):203, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n13/pdf/scientificamerican03271886-203.pdf.

Anonymous:1886:IIe

[7507] Anonymous. Index of inventions. Scientific American, 54(13):203–204, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n13/pdf/scientificamerican03271886-203a.pdf.

Anonymous:1886:PNPe

[7508] Anonymous. Patent notes: Patents. Scientific American, 54(13):206, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n13/pdf/scientificamerican03271886-206.pdf.

Anonymous:1886:CGSa

[7509] Anonymous. The Clapp–Griffiths steel process. Scientific American, 54(13):191, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n13/pdf/scientificamerican03271886-
191.pdf.

Anonymous:1886:LCS

[7510] Anonymous. Loss of the Cunard steamer Oregon. Scientific American, 54(13):192, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n13/pdf/scientificamerican03271886-
192.pdf.

Anonymous:1886:SO

[7511] Anonymous. As to the sinking of the Oregon. Scientific American, 54(13):192–193, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n13/pdf/scientificamerican03271886-
192a.pdf.

Anonymous:1886:PDL

[7512] Anonymous. Practical directions for lightning rods. Scientific American, 54(13):193, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n13/pdf/scientificamerican03271886-
193.pdf.

Anonymous:1886:WWI

[7513] Anonymous. Waterproof writing ink and paper. Scientific American, 54(13):193, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n13/pdf/scientificamerican03271886-
193a.pdf.

Anonymous:1886:PNc

[7514] Anonymous. Photographic notes. Scientific American, 54(13):193, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n13/pdf/scientificamerican03271886-193b.pdf.

Anonymous:1886:WMa

[7515] Anonymous. Washing machine. Scientific American, 54(13):194, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n13/pdf/scientificamerican03271886-194.pdf.

Anonymous:1886:ISPb

[7516] Anonymous. Improved snow plow. Scientific American, 54(13):194, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n13/pdf/scientificamerican03271886-194a.pdf.

Anonymous:1886:SCa

[7517] Anonymous. Smoke conductor. Scientific American, 54(13):194, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n13/pdf/scientificamerican03271886-194b.pdf.

Anonymous:1886:CCa

[7518] Anonymous. Car coupling. Scientific American, 54(13):194, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n13/pdf/scientificamerican03271886-194c.pdf.

Anonymous:1886:IVB

[7519] Anonymous. An improved “Venetian” blind. Scientific American, 54(13):194, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n13/pdf/scientificamerican03271886-
194d.pdf.

Anonymous:1886:MCa

[7520] Anonymous. A mechanical counter. Scientific American, 54(13): 195–197, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n13/pdf/scientificamerican03271886-
195.pdf.

Anonymous:1886:IAEa

[7521] Anonymous. An improved automatic engine. Scientific American, 54(13):198, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n13/pdf/scientificamerican03271886-
198.pdf.

Anonymous:1886:PCL

[7522] Anonymous. Pipe cutting lathe. Scientific American, 54(13):198, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n13/pdf/scientificamerican03271886-198a.pdf.

Anonymous:1886:EIE

[7523] Anonymous. The Edinburgh International Exhibition. Scientific American, 54(13):199–200, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n13/pdf/scientificamerican03271886-
199.pdf.

Anonymous:1886:NE

[7524] Anonymous. A novel ellipsograph. Scientific American, 54(13):200, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n13/pdf/scientificamerican03271886-200.pdf.

Anonymous:1886:AVF

[7525] Anonymous. Armor for vessels, forts, etc. Scientific American, 54(13):201, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n13/pdf/scientificamerican03271886-
201.pdf.

Anonymous:1886:EIf

[7526] Anonymous. Engineering inventions. Scientific American, 54(13):202, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n13/pdf/scientificamerican03271886-202.pdf.

Anonymous:1886:AId

[7527] Anonymous. Agricultural inventions. Scientific American, 54(13):202, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n13/pdf/scientificamerican03271886-202a.pdf.

Anonymous:1886:MIe

[7528] Anonymous. Miscellaneous inventions. Scientific American, 54(13):202, March 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n13/pdf/scientificamerican03271886-202b.pdf.

Anonymous:1886:Ce

[7529] Anonymous. Correspondence. Scientific American, 54(14):213–214, April 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n14/pdf/scientificamerican04031886-213.pdf.

Anonymous:1886:BPf

[7530] Anonymous. Business and personal. Scientific American, 54(14):218, April 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n14/pdf/scientificamerican04031886-218c.pdf.

Anonymous:1886:NQf

[7531] Anonymous. Notes and queries. Scientific American, 54(14):218–219, April 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n14/pdf/scientificamerican04031886-218d.pdf.

Anonymous:1886:IIf

[7532] Anonymous. Index of inventions. Scientific American, 54(14):219–220, April 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n14/pdf/scientificamerican04031886-219a.pdf.

Anonymous:1886:PNPf

[7533] Anonymous. Patent notes: Patents. Scientific American, 54(14):222, April 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n14/pdf/scientificamerican04031886-222.pdf.

Anonymous:1886:DID

[7534] Anonymous. The Davis Island Dam on the Ohio River. Scientific American, 54(14):207, April 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n14/pdf/scientificamerican04031886-
207.pdf.

Anonymous:1886:PBB

[7535] Anonymous. The Patent Bills before Congress. Scientific American, 54(14):208, April 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n14/pdf/scientificamerican04031886-
208.pdf.

Anonymous:1886:RBP

[7536] Anonymous. Roots’ blowing and pumping machinery. Scientific American, 54(14):208–209, April 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n14/pdf/scientificamerican04031886-
208a.pdf.

Proctor:1886:NSMa

[7537] Richard A. Proctor. Night sky — March and April. Scientific American, 54(14):209, April 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n14/pdf/scientificamerican04031886-
209.pdf.

Anonymous:1886:ITa

[7538] Anonymous. Ironing table. Scientific American, 54(14):210, April 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n14/pdf/scientificamerican04031886-210.pdf.

Anonymous:1886:CCb

[7539] Anonymous. Car coupling. Scientific American, 54(14):210, April 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n14/pdf/scientificamerican04031886-210a.pdf.

Anonymous:1886:ICc

[7540] Anonymous. An impromptu hen coop. Scientific American, 54(14):210, April 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n14/pdf/scientificamerican04031886-210b.pdf.

Anonymous:1886:ISIa

[7541] Anonymous. Improved sad iron. Scientific American, 54(14):210, April 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n14/pdf/scientificamerican04031886-210c.pdf.

Anonymous:1886:HMA

[7542] Anonymous. Hill marking attachment for corn planters. Scientific American, 54(14):210–211, April 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v54/n14/pdf/scientificamerican04031886-
210d.pdf.

Anonymous:1886:CBB

[7543] Anonymous. A compact boat boiler. Scientific American, 54 (14):211–212, April 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n14/pdf/scientificamerican04031886-
211.pdf.

Anonymous:1886:PNd

[7544] Anonymous. Photographic notes. Scientific American, 54(14):212, April 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n14/pdf/scientificamerican04031886-212.pdf.

Anonymous:1886:AAD

[7545] Anonymous. Artistic architectural designs. Scientific American, 54(14):215–216, April 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n14/pdf/scientificamerican04031886-
215.pdf.

Anonymous:1886:RIB

[7546] Anonymous. Roots’ improvements in blowers and pumps. Scientific American, 54(14):217, April 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n14/pdf/scientificamerican04031886-
217.pdf.

Anonymous:1886:EIg

[7547] Anonymous. Engineering inventions. Scientific American, 54(14):218, April 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n14/pdf/scientificamerican04031886-218.pdf.

Anonymous:1886:AIe

[7548] Anonymous. Agricultural inventions. Scientific American, 54(14):218, April 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n14/pdf/scientificamerican04031886-218a.pdf.

Anonymous:1886:MIf

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n14/pdf/scientificamerican04031886-218b.pdf.

Anonymous:1886:Ib

[7550] Anonymous. To inventors. Scientific American, 54(14):219, April 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n14/pdf/scientificamerican04031886-219.pdf.

Anonymous:1886:Cf

[7551] Anonymous. Correspondence. Scientific American, 54(15):229–230, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n15/pdf/scientificamerican04101886-229.pdf.

Anonymous:1886:BNBf

[7552] Anonymous. Books: New books and publications. Scientific American, 54(15):234, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n15/pdf/scientificamerican04101886-
234c.pdf.

Anonymous:1886:BPg

[7553] Anonymous. Business and personal. Scientific American, 54 (15):234–235, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n15/pdf/scientificamerican04101886-
234e.pdf.

Anonymous:1886:NQg

[7554] Anonymous. Notes and queries. Scientific American, 54(15):235, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n15/pdf/scientificamerican04101886-235.pdf.

Anonymous:1886:IIg

[7555] Anonymous. Index of inventions. Scientific American, 54(15):235–236, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n15/pdf/scientificamerican04101886-235b.pdf.

Anonymous:1886:PNPg

[7556] Anonymous. Patent notes: Patents. Scientific American, 54(15):238, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n15/pdf/scientificamerican04101886-238.pdf.

Anonymous:1886:DWS

[7557] Anonymous. The driven well system of the Brooklyn Water Works. Scientific American, 54(15):223, April 10, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v54/
n15/pdf/scientificamerican04101886-223.pdf.

Anonymous:1886:TSRb

[7558] Anonymous. The Tehuantepec ship railway. Scientific American, 54(15):224, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n15/pdf/scientificamerican04101886-
224.pdf.

Anonymous:1886:LSS

[7559] Anonymous. Licenses for small steamboats. Scientific American, 54(15):224, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n15/pdf/scientificamerican04101886-
224a.pdf.

Anonymous:1886:SCSb

[7560] Anonymous. Shop common sense. Scientific American, 54(15):224–225, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n15/pdf/scientificamerican04101886-224b.pdf.

Anonymous:1886:DRPb

[7561] Anonymous. Decisions relating to patents. Scientific American, 54(15):225, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n15/pdf/scientificamerican04101886-
225.pdf.

Anonymous:1886:MVD

[7562] Anonymous. Mixing and vaporizing device for inhalers. Scientific American, 54(15):226, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n15/pdf/scientificamerican04101886-
226.pdf.

Anonymous:1886:G

[7563] Anonymous. Gasometer. Scientific American, 54(15):226, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n15/pdf/scientificamerican04101886-226a.pdf.

Anonymous:1886:TCa

[7564] Anonymous. Time calculator. Scientific American, 54(15):226, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n15/pdf/scientificamerican04101886-226b.pdf.

Anonymous:1886:CSBa

[7565] Anonymous. Combined spring bed and fire escape. Scientific American, 54(15):226, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n15/pdf/scientificamerican04101886-
226c.pdf.

Anonymous:1886:NBH

[7566] Anonymous. A novel barrel head. Scientific American, 54(15):226–227, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n15/pdf/scientificamerican04101886-226d.pdf.

Anonymous:1886:SDU

[7567] Anonymous. Steam disinfecter for use in hospitals. Scientific American, 54(15):227, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n15/pdf/scientificamerican04101886-
227.pdf.

Anonymous:1886:CKL

[7568] Anonymous. The “Champion” keyless locks. Scientific American, 54(15):227–228, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n15/pdf/scientificamerican04101886-
227a.pdf.

Anonymous:1886:AP

[7569] Anonymous. Astronomical photography. Scientific American, 54 (15):230–231, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n15/pdf/scientificamerican04101886-
230.pdf.

Anonymous:1886:RSSa

[7570] Anonymous. A railway safety switch. Scientific American, 54 (15):231–233, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n15/pdf/scientificamerican04101886-
231.pdf.

Anonymous:1886:NGW

[7571] Anonymous. A natural gab well. Scientific American, 54(15):233, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n15/pdf/scientificamerican04101886-233.pdf.

Anonymous:1886:EIh

[7572] Anonymous. Engineering inventions. Scientific American, 54(15):234, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n15/pdf/scientificamerican04101886-234.pdf.

Anonymous:1886:AIf

[7573] Anonymous. Agricultural inventions. Scientific American, 54(15):234, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n15/pdf/scientificamerican04101886-234a.pdf.

Anonymous:1886:MIg

[7574] Anonymous. Miscellaneous inventions. Scientific American, 54(15):234, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n15/pdf/scientificamerican04101886-234b.pdf.

Anonymous:1886:Sa

[7575] Anonymous. Special. Scientific American, 54(15):234, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n15/pdf/scientificamerican04101886-234d.pdf.

Anonymous:1886:Ie

[7576] Anonymous. To inventors. Scientific American, 54(15):235, April 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n15/pdf/scientificamerican04101886-235a.pdf.

Anonymous:1886:Ch

[7577] Anonymous. Correspondence. Scientific American, 54(16):245–246, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n16/pdf/scientificamerican04171886-245.pdf.

Anonymous:1886:BNBg

[7578] Anonymous. Books: New books and publications. Scientific American, 54(16):250, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n16/pdf/scientificamerican04171886-
250c.pdf.

Anonymous:1886:BPh

[7579] Anonymous. Business and personal. Scientific American, 54 (16):250–251, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n16/pdf/scientificamerican04171886-
250e.pdf.

Anonymous:1886:NQh

[7580] Anonymous. Notes and queries. Scientific American, 54(16):251, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n16/pdf/scientificamerican04171886-251.pdf.

Anonymous:1886:IIh

[7581] Anonymous. Index of inventions. Scientific American, 54(16):251–252, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n16/pdf/scientificamerican04171886-251a.pdf.

Anonymous:1886:PNPh

[7582] Anonymous. Patent notes: Patents. Scientific American, 54(16):254, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n16/pdf/scientificamerican04171886-254.pdf.

Gratacap:1886:GHA

[7583] L. P. Gratacap. The Geological Hall of the American Museum of Natural History. Scientific American, 54(16):239, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n16/pdf/scientificamerican04171886-239.pdf.

Anonymous:1886:CSa

[7584] Anonymous. The coast survey. Scientific American, 54(16):240, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n16/pdf/scientificamerican04171886-240.pdf.

Anonymous:1886:RC

[7585] Anonymous. Railways in China. Scientific American, 54(16):240–241, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n16/pdf/scientificamerican04171886-240a.pdf.

Anonymous:1886:MKC

[7586] Anonymous. Mechanics’ keying clamp. Scientific American, 54(16):242, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n16/pdf/scientificamerican04171886-242.pdf.

Anonymous:1886:RRC

[7587] Anonymous. Reclining rocking chair. Scientific American, 54(16):242, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n16/pdf/scientificamerican04171886-242a.pdf.

Anonymous:1886:CK

[7588] Anonymous. Convertible keg. Scientific American, 54(16):242, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n16/pdf/scientificamerican04171886-242b.pdf.

Anonymous:1886:EGW

[7589] Anonymous. End gate for wagons. Scientific American, 54(16):242, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n16/pdf/scientificamerican04171886-242c.pdf.

Anonymous:1886:IHS

[7590] Anonymous. Improved heating stove. Scientific American, 54(16):243, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n16/pdf/scientificamerican04171886-243.pdf.

Anonymous:1886:EBG

[7591] Anonymous. An engine and boiler for good service. Scientific American, 54(16):243, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n16/pdf/scientificamerican04171886-
243a.pdf.

Anonymous:1886:PCa

[7592] Anonymous. A positive cyclometer. Scientific American, 54(16): 243–244, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n16/pdf/scientificamerican04171886-
243b.pdf.

Anonymous:1886:ACD

[7593] Anonymous. Automatic clothes drier. Scientific American, 54(16):244, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n16/pdf/scientificamerican04171886-244.pdf.

Anonymous:1886:LLM

[7594] Anonymous. Leather-link machinery belt. Scientific American, 54(16):244, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n16/pdf/scientificamerican04171886-
244a.pdf.

Anonymous:1886:FDL

[7595] Anonymous. Furniture and decorations of a library. Scientific American, 54(16):246, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n16/pdf/scientificamerican04171886-
246.pdf.

Anonymous:1886:MGB

[7596] Anonymous. Manufacture of gelatino–bromide dry plates. Scientific American, 54(16):247–248, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v54/n16/pdf/scientificamerican04171886-
247.pdf.

Anonymous:1886:PCb

[7597] Anonymous. Pygmy cattle. Scientific American, 54(16):249, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n16/pdf/scientificamerican04171886-249.pdf.

Anonymous:1886:EIi

[7598] Anonymous. Engineering inventions. Scientific American, 54(16):250, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n16/pdf/scientificamerican04171886-250.pdf.

Anonymous:1886:AIg

[7599] Anonymous. Agricultural inventions. Scientific American, 54(16):250, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n16/pdf/scientificamerican04171886-250a.pdf.

Anonymous:1886:MIh

[7600] Anonymous. Miscellaneous inventions. Scientific American, 54(16):250, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n16/pdf/scientificamerican04171886-250b.pdf.

Anonymous:1886:Sb

[7601] Anonymous. Special. Scientific American, 54(16):250, April 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n16/pdf/scientificamerican04171886-250d.pdf.

Anonymous:1886:Ci

[7602] Anonymous. Correspondence. Scientific American, 54(17):261, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n17/pdf/scientificamerican04241886-261.pdf.

Anonymous:1886:BPi

[7603] Anonymous. Business and personal. Scientific American, 54(17):266, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n17/pdf/scientificamerican04241886-266a.pdf.

Anonymous:1886:NQi

[7604] Anonymous. Notes and queries. Scientific American, 54(17):266–267, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n17/pdf/scientificamerican04241886-266b.pdf.

Anonymous:1886:IIi

[7605] Anonymous. Index of inventions. Scientific American, 54(17):267–268, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n17/pdf/scientificamerican04241886-267.pdf.

Anonymous:1886:PNPi

[7606] Anonymous. Patent notes: Patents. Scientific American, 54(17):270, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n17/pdf/scientificamerican04241886-270.pdf.

Anonymous:1886:NFM

[7607] Anonymous. A novel feature in mining. Scientific American, 54(17):255, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n17/pdf/scientificamerican04241886-255.pdf.

Anonymous:1886:SCM

[7608] Anonymous. Smoke condensed by means of electricity. Scientific American, 54(17):255, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n17/pdf/scientificamerican04241886-
255a.pdf.

Anonymous:1886:SSO

[7609] Anonymous. The sinking of the steamer Oregon. Scientific American, 54(17):256, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n17/pdf/scientificamerican04241886-
256.pdf.

Anonymous:1886:ISCa

[7610] Anonymous. Inter-state commerce. Scientific American, 54(17):256, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n17/pdf/scientificamerican04241886-256a.pdf.

Anonymous:1886:SSa

[7611] Anonymous. Signals at sea. Scientific American, 54(17):256–257, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n17/pdf/scientificamerican04241886-256b.pdf.

Anonymous:1886:CGSb

[7612] Anonymous. The Clapp–Griffiths steel process. Scientific American, 54(17):257, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n17/pdf/scientificamerican04241886-
257.pdf.

Anonymous:1886:PNe

[7613] Anonymous. Photographic notes. Scientific American, 54(17):257, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n17/pdf/scientificamerican04241886-257a.pdf.

Anonymous:1886:IOC

[7614] Anonymous. Improved oil cup. Scientific American, 54(17):258, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n17/pdf/scientificamerican04241886-258.pdf.

Anonymous:1886:SFa

[7615] Anonymous. Sash fastener. Scientific American, 54(17):258, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n17/pdf/scientificamerican04241886-258a.pdf.

Anonymous:1886:BPj

[7616] Anonymous. Base for posts. Scientific American, 54(17):258, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n17/pdf/scientificamerican04241886-258b.pdf.

Anonymous:1886:SBa

[7617] Anonymous. Spring balance. Scientific American, 54(17):258–259, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n17/pdf/scientificamerican04241886-258c.pdf.

Anonymous:1886:BWS

[7618] Anonymous. Bevel wheel shaping and dividing machine. Scientific American, 54(17):259, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n17/pdf/scientificamerican04241886-
259.pdf.

Anonymous:1886:ACB

[7619] Anonymous. An adjustable crib and bedstead. Scientific American, 54(17):259, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n17/pdf/scientificamerican04241886-
259a.pdf.

Anonymous:1886:SSb

[7620] Anonymous. Spiked skid. Scientific American, 54(17):259–260, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n17/pdf/scientificamerican04241886-259b.pdf.

Anonymous:1886:CTS

[7621] Anonymous. Combined truss and supporter. Scientific American, 54(17):260, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n17/pdf/scientificamerican04241886-
260.pdf.

Anonymous:1886:SFM

[7622] Anonymous. Screw forging machine. Scientific American, 54 (17):262–263, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n17/pdf/scientificamerican04241886-
262.pdf.

Anonymous:1886:IAR

[7623] Anonymous. An Italian artist’s residence — Sixteenth Century. Scientific American, 54(17):263, April 24, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v54/
n17/pdf/scientificamerican04241886-263.pdf.

Anonymous:1886:AG

[7624] Anonymous. Art in the garden. Scientific American, 54(17):263, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n17/pdf/scientificamerican04241886-263a.pdf.

Anonymous:1886:WGSa

[7625] Anonymous. A window grille of the seventeenth century. Scientific American, 54(17):263, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n17/pdf/scientificamerican04241886-
263b.pdf.

Seiss:1886:TT

[7626] C. Few Seiss. Tree toads. Scientific American, 54(17):264, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n17/pdf/scientificamerican04241886-264.pdf.

Anonymous:1886:EIj

[7627] Anonymous. Engineering inventions. Scientific American, 54(17):265, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n17/pdf/scientificamerican04241886-265.pdf.

Anonymous:1886:AIh

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n17/pdf/scientificamerican04241886-265a.pdf.

Anonymous:1886:MIi

[7629] Anonymous. Miscellaneous inventions. Scientific American, 54(17):265, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n17/pdf/scientificamerican04241886-265b.pdf.

Anonymous:1886:Sd

[7630] Anonymous. Special. Scientific American, 54(17):266, April 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n17/pdf/scientificamerican04241886-266.pdf.

Anonymous:1886:Cj

[7631] Anonymous. Correspondence. Scientific American, 54(18):277, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n18/pdf/scientificamerican05011886-277.pdf.

Anonymous:1886:BNBh

[7632] Anonymous. Books: New books and publications. Scientific American, 54(18):282, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n18/pdf/scientificamerican05011886-
282b.pdf.

Anonymous:1886:BPk

[7633] Anonymous. Business and personal. Scientific American, 54 (18):282–283, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n18/pdf/scientificamerican05011886-
282d.pdf.

Anonymous:1886:NQj

[7634] Anonymous. Notes and queries. Scientific American, 54(18):283, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n18/pdf/scientificamerican05011886-283.pdf.

Anonymous:1886:IIj

[7635] Anonymous. Index of inventions. Scientific American, 54(18):283–284, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n18/pdf/scientificamerican05011886-283a.pdf.

Anonymous:1886:PNPj

[7636] Anonymous. Patent notes: Patents. Scientific American, 54(18):286, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n18/pdf/scientificamerican05011886-286.pdf.

Anonymous:1886:NFAa

[7637] Anonymous. The new French Atlantic steamship La Bourgogne. Scientific American, 54(18):271, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v54/n18/pdf/scientificamerican05011886-
271.pdf.

Anonymous:1886:IWI

[7638] Anonymous. Inventions wanted in India. Scientific American, 54(18):272, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n18/pdf/scientificamerican05011886-
272.pdf.

Anonymous:1886:AMN

[7639] Anonymous. The American Museum of Natural History. Scientific American, 54(18):272, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n18/pdf/scientificamerican05011886-
272a.pdf.

Anonymous:1886:NAS

[7640] Anonymous. National Academy of Sciences — annual meeting at Washington. Scientific American, 54(18):272–273, May 1, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v54/
n18/pdf/scientificamerican05011886-272b.pdf.

Proctor:1886:NSAa

[7641] Richard A. Proctor. Night sky — April and May. Scientific American, 54(18):273, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n18/pdf/scientificamerican05011886-
273.pdf.

Anonymous:1886:CSF

[7642] Anonymous. A crate for shipping fruits, vegetables, etc. Scientific American, 54(18):274, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n18/pdf/scientificamerican05011886-
274.pdf.

Anonymous:1886:TCLa

[7643] Anonymous. Twick’s combined land roller and clod crusher. Scientific American, 54(18):274, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n18/pdf/scientificamerican05011886-
274a.pdf.

Anonymous:1886:SLW

[7644] Anonymous. A spring for lumber wagons. Scientific American, 54(18):274, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n18/pdf/scientificamerican05011886-
274b.pdf.

Anonymous:1886:AAO

[7645] Anonymous. An adjustable automatic oil cup. Scientific American, 54(18):274, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n18/pdf/scientificamerican05011886-
274c.pdf.

Anonymous:1886:SFE

[7646] Anonymous. A simple fire escape. Scientific American, 54(18):274, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n18/pdf/scientificamerican05011886-274d.pdf.

Anonymous:1886:ISB

[7647] Anonymous. Improved steam boiler. Scientific American, 54(18):275, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n18/pdf/scientificamerican05011886-275.pdf.

Anonymous:1886:DLA

[7648] Anonymous. A double-link, automatic car coupling. Scientific American, 54(18):275, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n18/pdf/scientificamerican05011886-
275a.pdf.

Anonymous:1886:CTB

[7649] Anonymous. Construction of torpedo boats. Scientific American, 54(18):275–276, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n18/pdf/scientificamerican05011886-
275b.pdf.

Anonymous:1886:SMT

[7650] Anonymous. Simple match trick. Scientific American, 54(18):276, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n18/pdf/scientificamerican05011886-276.pdf.

Anonymous:1886:HMP

[7651] Anonymous. A home-made printing press. Scientific American, 54(18):276, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n18/pdf/scientificamerican05011886-
276a.pdf.

Anonymous:1886:DBC

[7652] Anonymous. A detachable billiard cue tip. Scientific American, 54(18):277–278, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n18/pdf/scientificamerican05011886-
277a.pdf.

Anonymous:1886:ICMa

[7653] Anonymous. Improved concrete making machine. Scientific American, 54(18):278, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n18/pdf/scientificamerican05011886-
278.pdf.

Anonymous:1886:SCE

[7654] Anonymous. A small condensing engine and boiler. Scientific American, 54(18):278, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n18/pdf/scientificamerican05011886-
278a.pdf.

Anonymous:1886:OCM

[7655] Anonymous. Othniel Charles Marsh. Scientific American, 54 (18):279–281, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n18/pdf/scientificamerican05011886-
279.pdf.

Anonymous:1886:NMF

[7656] Anonymous. Novel mode of feeding lambs. Scientific American, 54(18):281, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n18/pdf/scientificamerican05011886-
281.pdf.

Anonymous:1886:EIk

[7657] Anonymous. Engineering inventions. Scientific American, 54(18):282, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n18/pdf/scientificamerican05011886-282.pdf.

Anonymous:1886:MIj

[7658] Anonymous. Miscellaneous inventions. Scientific American, 54(18):282, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n18/pdf/scientificamerican05011886-282a.pdf.

Anonymous:1886:Se

[7659] Anonymous. Special. Scientific American, 54(18):282, May 1, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n18/pdf/scientificamerican05011886-282c.pdf.

Anonymous:1886:BNBi

[7660] Anonymous. Books: New books and publications. Scientific American, 54(19):297, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n19/pdf/scientificamerican05081886-
297c.pdf.

Anonymous:1886:BPl

[7661] Anonymous. Business and personal. Scientific American, 54 (19):297–298, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n19/pdf/scientificamerican05081886-
297e.pdf.

Anonymous:1886:NQk

[7662] Anonymous. Notes and queries. Scientific American, 54(19):298–299, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n19/pdf/scientificamerican05081886-298.pdf.

Anonymous:1886:IIk

[7663] Anonymous. Index of inventions. Scientific American, 54(19):299–300, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n19/pdf/scientificamerican05081886-299.pdf.

Anonymous:1886:PNPk

[7664] Anonymous. Patent notes: Patents. Scientific American, 54(19):302, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n19/pdf/scientificamerican05081886-302.pdf.

Anonymous:1886:DFW

[7665] Anonymous. The deepest foundation in the world. Scientific American, 54(19):287, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n19/pdf/scientificamerican05081886-
287.pdf.

Anonymous:1886:ONI

[7666] Anonymous. Opposition to notable inventions. Scientific American, 54(19):288, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n19/pdf/scientificamerican05081886-
288.pdf.

Anonymous:1886:CPCa

[7667] Anonymous. Condition of the Panama Canal. Scientific American, 54(19):288–289, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n19/pdf/scientificamerican05081886-
288a.pdf.

Anonymous:1886:MWU

[7668] Anonymous. Mineral waters of the United States. Scientific American, 54(19):289, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n19/pdf/scientificamerican05081886-
289.pdf.

Anonymous:1886:CFB

[7669] Anonymous. Comets Fabry and Barnard. Scientific American, 54(19):289, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n19/pdf/scientificamerican05081886-
289a.pdf.

Anonymous:1886:PNf

[7670] Anonymous. Photographic notes. Scientific American, 54(19):289, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n19/pdf/scientificamerican05081886-289b.pdf.

Anonymous:1886:RTW

[7671] Anonymous. The Remington type writer. Scientific American, 54(19):289, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n19/pdf/scientificamerican05081886-
289c.pdf.

Anonymous:1886:WC

[7672] Anonymous. Wagon coupling. Scientific American, 54(19):290, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n19/pdf/scientificamerican05081886-290.pdf.

Anonymous:1886:LAS

[7673] Anonymous. Leveling apparatus for steam boilers. Scientific American, 54(19):290, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n19/pdf/scientificamerican05081886-
290a.pdf.

Anonymous:1886:CPa

[7674] Anonymous. Copying press. Scientific American, 54(19):290, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n19/pdf/scientificamerican05081886-290b.pdf.

Anonymous:1886:PTO

[7675] Anonymous. Perforating type and overlay. Scientific American, 54(19):290, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n19/pdf/scientificamerican05081886-
290c.pdf.

Anonymous:1886:IPa

[7676] Anonymous. An improved pulverizer. Scientific American, 54(19):291, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n19/pdf/scientificamerican05081886-291.pdf.

Anonymous:1886:ACL

[7677] Anonymous. An automatic cut-out and lightning arrester. Scientific American, 54(19):291, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n19/pdf/scientificamerican05081886-
291a.pdf.

Anonymous:1886:IRV

[7678] Anonymous. Improved rotary ventilator. Scientific American, 54 (19):291, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n19/pdf/scientificamerican05081886-
291b.pdf.

Anonymous:1886:LAA

[7679] Anonymous. Lightning arrester and automatic cut-out. Scientific American, 54(19):291–292, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n19/pdf/scientificamerican05081886-
291c.pdf.

Anonymous:1886:BBH

[7680] Anonymous. Bale and box hook. Scientific American, 54(19):293, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n19/pdf/scientificamerican05081886-293.pdf.

Anonymous:1886:DST

[7681] Anonymous. Device for separating teeth. Scientific American, 54(19):293–294, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n19/pdf/scientificamerican05081886-
293a.pdf.

Anonymous:1886:CCS

[7682] Anonymous. A combined crusher and stamp mill. Scientific American, 54(19):294–295, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n19/pdf/scientificamerican05081886-
294.pdf.

Anonymous:1886:NCM

[7683] Anonymous. New club mouse of the st. louis jockey club. Scientific American, 54(19):295–296, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v54/n19/pdf/scientificamerican05081886-
295.pdf.

Anonymous:1886:EIl

[7684] Anonymous. Engineering inventions. Scientific American, 54(19):297, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n19/pdf/scientificamerican05081886-297.pdf.

Anonymous:1886:AIi

[7685] Anonymous. Agricultural inventions. Scientific American, 54(19):297, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n19/pdf/scientificamerican05081886-297a.pdf.

Anonymous:1886:MIk

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n19/pdf/scientificamerican05081886-297b.pdf.

Anonymous:1886:Sg

[7687] Anonymous. Special. Scientific American, 54(19):297, May 8, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n19/pdf/scientificamerican05081886-297d.pdf.

Anonymous:1886:Cl

[7688] Anonymous. Correspondence. Scientific American, 54(2):21, January 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n2/pdf/scientificamerican01091886-21.pdf.

Anonymous:1886:BPm

[7689] Anonymous. Business and personal. Scientific American, 54(2):26, January 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n2/pdf/scientificamerican01091886-26.pdf.

Anonymous:1886:BNBj

[7690] Anonymous. Books: New books and publications. Scientific American, 54(2):26, January 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n2/pdf/scientificamerican01091886-
26a.pdf.

Anonymous:1886:NQl

[7691] Anonymous. Notes and queries. Scientific American, 54(2):26–27, January 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n2/pdf/scientificamerican01091886-26b.pdf.

Anonymous:1886:IIl

[7692] Anonymous. Index of inventions. Scientific American, 54(2):27–29, January 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n2/pdf/scientificamerican01091886-27.pdf.

Anonymous:1886:PNPl

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n2/pdf/scientificamerican01091886-30.pdf.

Anonymous:1886:NTV

[7694] Anonymous. The new Tay Viaduct. Scientific American, 54(2):15, January 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n2/pdf/scientificamerican01091886-15.pdf.

Anonymous:1886:ELC

[7695] Anonymous. Electric light on captive balloons. Scientific American, 54(2):16, January 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n2/pdf/scientificamerican01091886-
16.pdf.

Anonymous:1886:PNg

[7696] Anonymous. Photographic notes. Scientific American, 54(2):17–18, January 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n2/pdf/scientificamerican01091886-17.pdf.

Anonymous:1886:VS

[7697] Anonymous. Vehicle spring. Scientific American, 54(2):18, January 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n2/pdf/scientificamerican01091886-18.pdf.

Anonymous:1886:IPb

[7698] Anonymous. Improved plow. Scientific American, 54(2):18, January 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n2/pdf/scientificamerican01091886-18a.pdf.

Anonymous:1886:IDD

[7699] Anonymous. Improved drier for damp grain, glucose, starch, etc. Scientific American, 54(2):18, January 9, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v54/
n2/pdf/scientificamerican01091886-18b.pdf.

Anonymous:1886:VW

[7700] Anonymous. Vehicle wheel. Scientific American, 54(2):18, January 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n2/pdf/scientificamerican01091886-18c.pdf.

Anonymous:1886:ANF

[7701] Anonymous. An automatic and non-freezing stock watering trough. Scientific American, 54(2):18–19, January 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v54/n2/pdf/scientificamerican01091886-
18d.pdf.

Anonymous:1886:IDb

[7702] Anonymous. Insect destroyer. Scientific American, 54(2):19, January 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n2/pdf/scientificamerican01091886-19.pdf.

Anonymous:1886:IDW

[7703] Anonymous. An improved dumping wagon. Scientific American, 54(2):19, January 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n2/pdf/scientificamerican01091886-
19a.pdf.

Anonymous:1886:HWL

[7704] Anonymous. Hammer for welding locomotive frames. Scientific American, 54(2):19, January 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n2/pdf/scientificamerican01091886-
19b.pdf.

Anonymous:1886:PSP

[7705] Anonymous. Packing for steam pipe joint. Scientific American, 54(2):19–20, January 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n2/pdf/scientificamerican01091886-
19c.pdf.

Anonymous:1886:IPBa

[7706] Anonymous. Improved plate bending machine. Scientific American, 54(2):22, January 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n2/pdf/scientificamerican01091886-
22.pdf.

Anonymous:1886:NSC

[7707] Anonymous. A new steam carriage. Scientific American, 54(2):22–23, January 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n2/pdf/scientificamerican01091886-22a.pdf.

Anonymous:1886:LAG

[7708] Anonymous. The Layton Art Gallery, Milwaukee, Wisconsin. Scientific American, 54(2):23, January 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n2/pdf/scientificamerican01091886-
23.pdf.

Holder:1886:WW

[7709] J. B. Holder. Whales and whaling. Scientific American, 54(2):24, January 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n2/pdf/scientificamerican01091886-24.pdf.

Anonymous:1886:EIm

[7710] Anonymous. Engineering inventions. Scientific American, 54(2):25, January 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n2/pdf/scientificamerican01091886-25.pdf.

Anonymous:1886:MIl

[7711] Anonymous. Mechanical inventions. Scientific American, 54(2):25, January 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n2/pdf/scientificamerican01091886-25a.pdf.

Anonymous:1886:AIj

[7712] Anonymous. Agricultural inventions. Scientific American, 54(2):25, January 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n2/pdf/scientificamerican01091886-25b.pdf.

Anonymous:1886:MIm

[7713] Anonymous. Miscellaneous inventions. Scientific American, 54(2):25, January 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n2/pdf/scientificamerican01091886-25c.pdf.

Anonymous:1886:Cm

[7714] Anonymous. Correspondence. Scientific American, 54(20):309–310, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n20/pdf/scientificamerican05151886-309a.pdf.

Anonymous:1886:BNBk

[7715] Anonymous. Books: New books and publications. Scientific American, 54(20):314, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n20/pdf/scientificamerican05151886-
314b.pdf.

Anonymous:1886:BPn

[7716] Anonymous. Business and personal. Scientific American, 54 (20):314–315, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n20/pdf/scientificamerican05151886-
314c.pdf.

Anonymous:1886:NQm

[7717] Anonymous. Notes and queries. Scientific American, 54(20):315, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n20/pdf/scientificamerican05151886-315.pdf.

Anonymous:1886:IIm

[7718] Anonymous. Index of inventions. Scientific American, 54(20):315–316, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n20/pdf/scientificamerican05151886-315a.pdf.

Anonymous:1886:PNPm

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n20/pdf/scientificamerican05151886-318.pdf.

Hovey:1886:BEW

[7720] H. C. Hovey. The Brush Electric Works at Cleveland, Ohio. Scientific American, 54(20):303, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n20/pdf/scientificamerican05151886-
303.pdf.

Anonymous:1886:SIP

[7721] Anonymous. State interference with patents. Scientific American, 54(20):304, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n20/pdf/scientificamerican05151886-
304.pdf.

Anonymous:1886:PEP

[7722] Anonymous. The proposed extension of patents. Scientific American, 54(20):304–305, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n20/pdf/scientificamerican05151886-
304a.pdf.

Anonymous:1886:DTN

[7723] Anonymous. Discovery of two new comets. Scientific American, 54(20):305, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n20/pdf/scientificamerican05151886-
305.pdf.

Anonymous:1886:ACc

[7724] Anonymous. The Aladdin cooker. Scientific American, 54(20):305, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n20/pdf/scientificamerican05151886-305a.pdf.

Anonymous:1886:CCVa

[7725] Anonymous. Combined chimney and ventilator. Scientific American, 54(20):306, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n20/pdf/scientificamerican05151886-
306.pdf.

Anonymous:1886:WRB

[7726] Anonymous. White rubber back square piston packing. Scientific American, 54(20):306, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n20/pdf/scientificamerican05151886-
306a.pdf.

Anonymous:1886:CLP

[7727] Anonymous. A compound lever pinch bar for moving cars. Scientific American, 54(20):306, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n20/pdf/scientificamerican05151886-
306b.pdf.

Anonymous:1886:PGa

[7728] Anonymous. Pressure gauge. Scientific American, 54(20):306, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n20/pdf/scientificamerican05151886-306c.pdf.

Anonymous:1886:LCC

[7729] Anonymous. Listed corn cultivator. Scientific American, 54(20): 306–307, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n20/pdf/scientificamerican05151886-
306d.pdf.

Anonymous:1886:PAC

[7730] Anonymous. Powerful automatic car coupler. Scientific American, 54(20):307, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n20/pdf/scientificamerican05151886-
307.pdf.

Anonymous:1886:NSE

[7731] Anonymous. A novel steam engine. Scientific American, 54(20):307, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n20/pdf/scientificamerican05151886-307a.pdf.

Anonymous:1886:CSI

[7732] Anonymous. Centrifugal speed indicator. Scientific American, 54 (20):309, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n20/pdf/scientificamerican05151886-
309.pdf.

Anonymous:1886:SEA

[7733] Anonymous. A singular effect of the action of frost. Scientific American, 54(20):310, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n20/pdf/scientificamerican05151886-
310.pdf.

Atkinson:1886:AC

[7734] Edward Atkinson. The Aladdin cooker. Scientific American, 54(20):310, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n20/pdf/scientificamerican05151886-310a.pdf.

Sloane:1886:HCW

[7735] T. O’conor Sloane. How to carry water in a sieve. Scientific American, 54(20):311–312, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n20/pdf/scientificamerican05151886-
311.pdf.

Anonymous:1886:WIR

[7736] Anonymous. Who invented the ratchet brace? Scientific American, 54(20):312, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n20/pdf/scientificamerican05151886-
312.pdf.

Anonymous:1886:APBa

[7737] Anonymous. Argus peacocks in the Berlin Zoological Garden. Scientific American, 54(20):313, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n20/pdf/scientificamerican05151886-
313.pdf.

Anonymous:1886:EIn

[7738] Anonymous. Engineering inventions. Scientific American, 54(20):314, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n20/pdf/scientificamerican05151886-314.pdf.

Anonymous:1886:MIn

[7739] Anonymous. Miscellaneous inventions. Scientific American, 54(20):314, May 15, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n20/pdf/scientificamerican05151886-314a.pdf.

Anonymous:1886:Cn

[7740] Anonymous. Correspondence. Scientific American, 54(21):325, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n21/pdf/scientificamerican05221886-325.pdf.

Anonymous:1886:BNBl

[7741] Anonymous. Books: New books and publications. Scientific American, 54(21):330, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n21/pdf/scientificamerican05221886-
330d.pdf.

Anonymous:1886:BPo

[7742] Anonymous. Business and personal. Scientific American, 54(21):330, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n21/pdf/scientificamerican05221886-330e.pdf.

Anonymous:1886:NQn

[7743] Anonymous. Notes and queries. Scientific American, 54(21):330–331, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n21/pdf/scientificamerican05221886-330f.pdf.

Anonymous:1886:IIn

[7744] Anonymous. Index of inventions. Scientific American, 54(21):331–332, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n21/pdf/scientificamerican05221886-331.pdf.

Anonymous:1886:PNPn

[7745] Anonymous. Patent notes: Patents. Scientific American, 54(21):334, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n21/pdf/scientificamerican05221886-334.pdf.

Anonymous:1886:NRS

[7746] Anonymous. The new racing sloop yacht Atlantic. Scientific American, 54(21):319–320, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n21/pdf/scientificamerican05221886-
319.pdf.

Anonymous:1886:URN

[7747] Anonymous. An underground railway for New York City. Scientific American, 54(21):320, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n21/pdf/scientificamerican05221886-
320.pdf.

Anonymous:1886:PIE

[7748] Anonymous. Progress of industrial electricity. Scientific American, 54(21):320–321, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n21/pdf/scientificamerican05221886-
320a.pdf.

Anonymous:1886:ITC

[7749] Anonymous. Improved tube cutter. Scientific American, 54(21):322, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n21/pdf/scientificamerican05221886-322.pdf.

Anonymous:1886:HS

[7750] Anonymous. Hay stacker. Scientific American, 54(21):322, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n21/pdf/scientificamerican05221886-322a.pdf.

Anonymous:1886:ICMb

[7751] Anonymous. Ice cutting machine. Scientific American, 54(21):322, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n21/pdf/scientificamerican05221886-322b.pdf.

Anonymous:1886:PF

[7752] Anonymous. Plumber’s furnace. Scientific American, 54(21):322, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n21/pdf/scientificamerican05221886-322c.pdf.

Anonymous:1886:FMS

[7753] Anonymous. Feed mechanism for shingle machines. Scientific American, 54(21):322, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n21/pdf/scientificamerican05221886-
322d.pdf.

Anonymous:1886:CCc

[7754] Anonymous. Car coupling. Scientific American, 54(21):323, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n21/pdf/scientificamerican05221886-323.pdf.

Anonymous:1886:IWS

[7755] Anonymous. Improved window sash. Scientific American, 54(21):323, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n21/pdf/scientificamerican05221886-323a.pdf.

Anonymous:1886:SMF

[7756] Anonymous. Simple means of fastening wagon tires. Scientific American, 54(21):323, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n21/pdf/scientificamerican05221886-
323b.pdf.

Anonymous:1886:HCP

[7757] Anonymous. Hay and cotton press. Scientific American, 54(21):323, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n21/pdf/scientificamerican05221886-323c.pdf.

Anonymous:1886:SCF

[7758] Anonymous. A smoke consuming furnace. Scientific American, 54(21):323–324, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n21/pdf/scientificamerican05221886-
323d.pdf.

Anonymous:1886:TEE

[7759] Anonymous. Triple expansion engines. Scientific American, 54(21):326, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n21/pdf/scientificamerican05221886-326.pdf.

Anonymous:1886:CMW

[7760] Anonymous. Carriage makers’ work in the last century. Scientific American, 54(21):327, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n21/pdf/scientificamerican05221886-
327.pdf.

Meigs:1886:SC

[7761] M. Meigs. A Scythian chariot. Scientific American, 54(21):327–328, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n21/pdf/scientificamerican05221886-327a.pdf.

Anonymous:1886:CCd

[7762] Anonymous. Concrete construction. Scientific American, 54(21):329, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n21/pdf/scientificamerican05221886-329.pdf.

Anonymous:1886:LGP

[7763] Anonymous. The laying out of garden paths. Scientific American, 54(21):329, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n21/pdf/scientificamerican05221886-
329a.pdf.

Anonymous:1886:EIo

[7764] Anonymous. Engineering inventions. Scientific American, 54(21):330, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n21/pdf/scientificamerican05221886-330.pdf.

Anonymous:1886:MIo

[7765] Anonymous. Mechanical inventions. Scientific American, 54(21):330, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n21/pdf/scientificamerican05221886-330a.pdf.

Anonymous:1886:AIk

[7766] Anonymous. Agricultural inventions. Scientific American, 54(21):330, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n21/pdf/scientificamerican05221886-330b.pdf.

Anonymous:1886:MIp

[7767] Anonymous. Miscellaneous inventions. Scientific American, 54(21):330, May 22, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n21/pdf/scientificamerican05221886-330c.pdf.

Anonymous:1886:Co

[7768] Anonymous. Correspondence. Scientific American, 54(22):341, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n22/pdf/scientificamerican05291886-341.pdf.

Anonymous:1886:BNBm

[7769] Anonymous. Books: New books and publications. Scientific American, 54(22):346, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n22/pdf/scientificamerican05291886-
346c.pdf.

Anonymous:1886:BPp

[7770] Anonymous. Business and personal. Scientific American, 54(22):346, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n22/pdf/scientificamerican05291886-346d.pdf.

Anonymous:1886:NQo

[7771] Anonymous. Notes and queries. Scientific American, 54(22):347, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n22/pdf/scientificamerican05291886-347.pdf.

Anonymous:1886:IIo

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n22/pdf/scientificamerican05291886-347a.pdf.

Anonymous:1886:PNPo

[7773] Anonymous. Patent notes: Patents. Scientific American, 54(22):350, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n22/pdf/scientificamerican05291886-350.pdf.

Anonymous:1886:ET

[7774] Anonymous. Early telephones. Scientific American, 54(22):335–336, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n22/pdf/scientificamerican05291886-335.pdf.

Anonymous:1886:ORP

[7775] Anonymous. Official report on the Panama Canal. Scientific American, 54(22):336, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n22/pdf/scientificamerican05291886-
336.pdf.

Anonymous:1886:GH

[7776] Anonymous. Galveston Harbor. Scientific American, 54(22):336–337, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n22/pdf/scientificamerican05291886-336a.pdf.

Anonymous:1886:DWC

[7777] Anonymous. Deep water canal transportation. Scientific American, 54(22):337, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n22/pdf/scientificamerican05291886-
337.pdf.

Anonymous:1886:OD

[7778] Anonymous. The Oregon Disaster. Scientific American, 54(22): 337–338, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n22/pdf/scientificamerican05291886-
337a.pdf.

Anonymous:1886:IRH

[7779] Anonymous. An improved reciprocating hand tool. Scientific American, 54(22):338, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n22/pdf/scientificamerican05291886-
338.pdf.

Anonymous:1886:IFGa

[7780] Anonymous. Improved farm gate. Scientific American, 54(22):338, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n22/pdf/scientificamerican05291886-338a.pdf.

Anonymous:1886:SMD

[7781] Anonymous. Stop motion for doubling machines. Scientific American, 54(22):338, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n22/pdf/scientificamerican05291886-
338b.pdf.

Anonymous:1886:CNS

[7782] Anonymous. Combined newspaper stand and file. Scientific American, 54(22):338–339, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n22/pdf/scientificamerican05291886-
338c.pdf.

Anonymous:1886:IHP

[7783] Anonymous. Improved horse power. Scientific American, 54(22):339, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n22/pdf/scientificamerican05291886-339.pdf.

Anonymous:1886:SGa

[7784] Anonymous. Stove grate. Scientific American, 54(22):339, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n22/pdf/scientificamerican05291886-339a.pdf.

Anonymous:1886:ITb

[7785] Anonymous. Improved tympan. Scientific American, 54(22):339, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n22/pdf/scientificamerican05291886-339b.pdf.

Anonymous:1886:LS

[7786] Anonymous. Loom shuttle. Scientific American, 54(22):339–340, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n22/pdf/scientificamerican05291886-339c.pdf.

Anonymous:1886:DP

[7787] Anonymous. Dust pan. Scientific American, 54(22):340, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n22/pdf/scientificamerican05291886-340.pdf.

Anonymous:1886:ILP

[7788] Anonymous. An improved life preserver. Scientific American, 54 (22):342, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n22/pdf/scientificamerican05291886-
342.pdf.

Anonymous:1886:IAEb

[7789] Anonymous. Improved air engine. Scientific American, 54(22):342–343, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n22/pdf/scientificamerican05291886-342a.pdf.

Anonymous:1886:LBZ

[7790] Anonymous. The lemurs in the Berlin Zoological Garden. Scientific American, 54(22):343–344, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v54/n22/pdf/scientificamerican05291886-
343.pdf.

Sloane:1886:D

[7791] T. O’Conor Sloane. Dilatancy. Scientific American, 54(22):345, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n22/pdf/scientificamerican05291886-345.pdf.

Anonymous:1886:EIp

[7792] Anonymous. Engineering inventions. Scientific American, 54(22):346, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n22/pdf/scientificamerican05291886-346.pdf.

Anonymous:1886:AIl

[7793] Anonymous. Agricultural inventions. Scientific American, 54(22):346, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n22/pdf/scientificamerican05291886-346a.pdf.

Anonymous:1886:MIq

[7794] Anonymous. Miscellaneous inventions. Scientific American, 54(22):346, May 29, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n22/pdf/scientificamerican05291886-346b.pdf.

Anonymous:1886:BPq

[7795] Anonymous. Business and personal. Scientific American, 54(23):363, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n23/pdf/scientificamerican06051886-363c.pdf.

Anonymous:1886:NQp

[7796] Anonymous. Notes and queries. Scientific American, 54(23):363, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n23/pdf/scientificamerican06051886-363d.pdf.

Anonymous:1886:IIp

[7797] Anonymous. Index of inventions. Scientific American, 54(23):363–364, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n23/pdf/scientificamerican06051886-363e.pdf.

Anonymous:1886:PNPp

[7798] Anonymous. Patent notes: Patents. Scientific American, 54(23):366, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n23/pdf/scientificamerican06051886-366.pdf.

Anonymous:1886:SSH

[7799] Anonymous. Stone steps at High Bridge, New York City. Scientific American, 54(23):351, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n23/pdf/scientificamerican06051886-
351.pdf.

Anonymous:1886:APN

[7800] Anonymous. Another patent nullification bill. Scientific American, 54(23):352, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n23/pdf/scientificamerican06051886-
352.pdf.

Anonymous:1886:GF

[7801] Anonymous. Gloucester fishers. Scientific American, 54(23):352–353, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n23/pdf/scientificamerican06051886-352a.pdf.

Proctor:1886:NSMb

[7802] Richard A. Proctor. Night sky — May and June. Scientific American, 54(23):353, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n23/pdf/scientificamerican06051886-
353.pdf.

Anonymous:1886:IPM

[7803] Anonymous. Improved planing machine. Scientific American, 54 (23):354, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n23/pdf/scientificamerican06051886-
354.pdf.

Anonymous:1886:IFCa

[7804] Anonymous. Improved freight car door lock. Scientific American, 54(23):355, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n23/pdf/scientificamerican06051886-
355.pdf.

Anonymous:1886:CSM

[7805] Anonymous. Circular saw mill. Scientific American, 54(23):355, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n23/pdf/scientificamerican06051886-355a.pdf.

Anonymous:1886:HL

[7806] Anonymous. Hay loader. Scientific American, 54(23):355, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n23/pdf/scientificamerican06051886-355b.pdf.

Anonymous:1886:ECE

[7807] Anonymous. Extensible cane, etc. Scientific American, 54(23):355, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n23/pdf/scientificamerican06051886-355c.pdf.

Anonymous:1886:FBF

[7808] Anonymous. Furnace for burning fine coal. Scientific American, 54(23):355, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n23/pdf/scientificamerican06051886-
355d.pdf.

Anonymous:1886:HDa

[7809] Anonymous. Horse detacher. Scientific American, 54(23):356, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n23/pdf/scientificamerican06051886-356.pdf.

Anonymous:1886:IPAa

[7810] Anonymous. Improved pneumatic action for organs. Scientific American, 54(23):356, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n23/pdf/scientificamerican06051886-
356a.pdf.

Anonymous:1886:FFa

[7811] Anonymous. The fluttering fly. Scientific American, 54(23):356, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n23/pdf/scientificamerican06051886-356b.pdf.

Anonymous:1886:PNh

[7812] Anonymous. Photographic notes. Scientific American, 54(23):357–358, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n23/pdf/scientificamerican06051886-357.pdf.

Anonymous:1886:EME

[7813] Anonymous. Eruption of Mount Etna. Scientific American, 54 (23):358–359, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n23/pdf/scientificamerican06051886-
358.pdf.

Anonymous:1886:ESb

[7814] Anonymous. The Egyptian Sphinx. Scientific American, 54(23): 359–360, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n23/pdf/scientificamerican06051886-
359.pdf.

Rood:1886:CC

[7815] Ogden N. Rood and Thomas Ewing. Color contrast. Scientific American, 54(23):361, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n23/pdf/scientificamerican06051886-
361.pdf.

Sloane:1886:CVa

[7816] T. O’Conor Sloane. Chemical vegetation. Scientific American, 54(23):361–362, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n23/pdf/scientificamerican06051886-
361a.pdf.

Anonymous:1886:EIq

[7817] Anonymous. Engineering inventions. Scientific American, 54(23):363, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n23/pdf/scientificamerican06051886-363.pdf.

Anonymous:1886:MIr

[7818] Anonymous. Miscellaneous inventions. Scientific American, 54(23):363, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n23/pdf/scientificamerican06051886-363a.pdf.

Anonymous:1886:Sh

[7819] Anonymous. Special. Scientific American, 54(23):363, June 5, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n23/pdf/scientificamerican06051886-363b.pdf.

Anonymous:1886:Cq

[7820] Anonymous. Correspondence. Scientific American, 54(25):389, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n25/pdf/scientificamerican06191886-389.pdf.

Anonymous:1886:BNBn

[7821] Anonymous. Books: New books and publications. Scientific American, 54(25):394, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n25/pdf/scientificamerican06191886-
394b.pdf.

Anonymous:1886:BPr

[7822] Anonymous. Business and personal. Scientific American, 54(25):394, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n25/pdf/scientificamerican06191886-394c.pdf.

Anonymous:1886:NQq

[7823] Anonymous. Notes and queries. Scientific American, 54(25):394–395, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n25/pdf/scientificamerican06191886-394d.pdf.

Anonymous:1886:IIq

[7824] Anonymous. Index of inventions. Scientific American, 54(25):395–396, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n25/pdf/scientificamerican06191886-395.pdf.

Anonymous:1886:PNPq

[7825] Anonymous. Patent notes: Patents. Scientific American, 54(25):398, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n25/pdf/scientificamerican06191886-398.pdf.

Anonymous:1886:UAS

[7826] Anonymous. Universal aiming stand. Scientific American, 54(25):383, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n25/pdf/scientificamerican06191886-383.pdf.

Anonymous:1886:D

[7827] Anonymous. The Duguesclin. Scientific American, 54(25):383, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n25/pdf/scientificamerican06191886-383a.pdf.

Anonymous:1886:NOT

[7828] Anonymous. The New Orleans Telephone Decision. Scientific American, 54(25):384, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n25/pdf/scientificamerican06191886-
384.pdf.

Anonymous:1886:DCR

[7829] Anonymous. The death of Col Richard M. Hoe. Scientific American, 54(25):384, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n25/pdf/scientificamerican06191886-
384a.pdf.

Anonymous:1886:PNi

[7830] Anonymous. Photographic notes. Scientific American, 54(25):385, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n25/pdf/scientificamerican06191886-385.pdf.

Anonymous:1886:PSW

[7831] Anonymous. Plug for stationary wash basins, etc. Scientific American, 54(25):386, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n25/pdf/scientificamerican06191886-
386.pdf.

Anonymous:1886:IFCb

[7832] Anonymous. Improved flooring clamp. Scientific American, 54(25):386, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n25/pdf/scientificamerican06191886-386a.pdf.

Anonymous:1886:IWE

[7833] Anonymous. Improved water elevator. Scientific American, 54(25):386, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n25/pdf/scientificamerican06191886-386b.pdf.

Anonymous:1886:GV

[7834] Anonymous. Governor valve. Scientific American, 54(25):386, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n25/pdf/scientificamerican06191886-386c.pdf.

Anonymous:1886:TCb

[7835] Anonymous. Track clearer. Scientific American, 54(25):386, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n25/pdf/scientificamerican06191886-386d.pdf.

Anonymous:1886:SBb

[7836] Anonymous. Shingling bracket. Scientific American, 54(25):386, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n25/pdf/scientificamerican06191886-386e.pdf.

Anonymous:1886:VPG

[7837] Anonymous. Vehicle platform gear. Scientific American, 54(25):387, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n25/pdf/scientificamerican06191886-387.pdf.

Anonymous:1886:VCM

[7838] Anonymous. Vertical chucking machine. Scientific American, 54 (25):387, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n25/pdf/scientificamerican06191886-
387a.pdf.

Anonymous:1886:HDb

[7839] Anonymous. Horse detacher. Scientific American, 54(25):387–388, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n25/pdf/scientificamerican06191886-387b.pdf.

Anonymous:1886:CPL

[7840] Anonymous. Central power locomotive. Scientific American, 54 (25):388, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n25/pdf/scientificamerican06191886-
388.pdf.

Anonymous:1886:SBD

[7841] Anonymous. Safety bulkhead doors. Scientific American, 54(25):388, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n25/pdf/scientificamerican06191886-388a.pdf.

Anonymous:1886:NDA

[7842] Anonymous. New dyeing apparatus. Scientific American, 54(25):390, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n25/pdf/scientificamerican06191886-390.pdf.

Anonymous:1886:IST

[7843] Anonymous. Improved sewer tunneling machine. Scientific American, 54(25):390, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n25/pdf/scientificamerican06191886-
390a.pdf.

Anonymous:1886:WJW

[7844] Anonymous. The workshop of James Watt. Scientific American, 54(25):391, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n25/pdf/scientificamerican06191886-
391.pdf.

Anonymous:1886:RDRb

[7845] Anonymous. Recent decisions relating to patents. Scientific American, 54(25):392, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n25/pdf/scientificamerican06191886-
392.pdf.

Sloane:1886:EL

[7846] T. O’conor Sloane. Experiments with Lycopodium. Scientific American, 54(25):393, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n25/pdf/scientificamerican06191886-
393.pdf.

Anonymous:1886:BNJ

[7847] Anonymous. Birds’ nests in Japanese houses. Scientific American, 54(25):393, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n25/pdf/scientificamerican06191886-
393a.pdf.

Anonymous:1886:EIr

[7848] Anonymous. Engineering inventions. Scientific American, 54(25):394, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n25/pdf/scientificamerican06191886-394.pdf.

Anonymous:1886:MIs

[7849] Anonymous. Miscellaneous inventions. Scientific American, 54(25):394, June 19, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n25/pdf/scientificamerican06191886-394a.pdf.

Anonymous:1886:Cr

[7850] Anonymous. Correspondence. Scientific American, 54(26):405, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n26/pdf/scientificamerican06261886-405.pdf.

Anonymous:1886:BNBo 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n26/pdf/scientificamerican06261886-
409b.pdf.

Anonymous:1886:BPs

[7852] Anonymous. Business and personal. Scientific American, 54(26):409, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n26/pdf/scientificamerican06261886-409c.pdf.

Anonymous:1886:NQr

[7853] Anonymous. Notes and queries. Scientific American, 54(26):409, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n26/pdf/scientificamerican06261886-409d.pdf.

Anonymous:1886:IIr

[7854] Anonymous. Index of inventions. Scientific American, 54(26):409–410, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n26/pdf/scientificamerican06261886-409f.pdf.

Anonymous:1886:If

[7855] Anonymous. Index. Scientific American, 54(26):411–412, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n26/pdf/scientificamerican06261886-411.pdf.

Anonymous:1886:PNPr

[7856] Anonymous. Patent notes: Patents. Scientific American, 54(26):414, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n26/pdf/scientificamerican06261886-414.pdf.

Anonymous:1886:DPB

[7857] Anonymous. Ducker’s portable barrack and field hospital. Scientific American, 54(26):399, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n26/pdf/scientificamerican06261886-
399.pdf.

Anonymous:1886:BPLa

[7858] Anonymous. The benefits of the Patent Laws. Scientific American, 54(26):400, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n26/pdf/scientificamerican06261886-
400.pdf.

Anonymous:1886:CBP

[7859] Anonymous. Car builders perplexed about car couplers. Scientific American, 54(26):400–401, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v54/n26/pdf/scientificamerican06261886-
400a.pdf.

Anonymous:1886:TBV

[7860] Anonymous. A transcontinental balloon voyage. Scientific American, 54(26):401, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n26/pdf/scientificamerican06261886-
401.pdf.

Anonymous:1886:ELI

[7861] Anonymous. Electric lighting as an investment. Scientific American, 54(26):401, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n26/pdf/scientificamerican06261886-
401a.pdf.

Anonymous:1886:PNj

[7862] Anonymous. Photographic notes. Scientific American, 54(26):401–402, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n26/pdf/scientificamerican06261886-401b.pdf.

Anonymous:1886:IHT

[7863] Anonymous. Improved hand truck. Scientific American, 54(26):402, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n26/pdf/scientificamerican06261886-402.pdf.

Anonymous:1886:ITG

[7864] Anonymous. Improved tool grinder. Scientific American, 54(26):402, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n26/pdf/scientificamerican06261886-402a.pdf.

Anonymous:1886:NHT

[7865] Anonymous. A new hand truck. Scientific American, 54(26):402, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n26/pdf/scientificamerican06261886-402b.pdf.

Anonymous:1886:MSV

[7866] Anonymous. Muffler for steam valves. Scientific American, 54(26):402, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n26/pdf/scientificamerican06261886-402c.pdf.

Anonymous:1886:FSPa

[7867] Anonymous. Feeding stand for poultry. Scientific American, 54(26):402, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n26/pdf/scientificamerican06261886-402d.pdf.

Anonymous:1886:NPP

[7868] Anonymous. New power punching presses. Scientific American, 54(26):403, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n26/pdf/scientificamerican06261886-
403.pdf.

Anonymous:1886:BFa

[7869] Anonymous. Boiler feeder. Scientific American, 54(26):403, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n26/pdf/scientificamerican06261886-403a.pdf.

Sloane:1886:SSW

[7870] T. O’conor Sloane. The spheroidal state of water. Scientific American, 54(26):404, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n26/pdf/scientificamerican06261886-
404.pdf.

Anonymous:1886:WW

[7871] Anonymous. Wood wool. Scientific American, 54(26):406, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n26/pdf/scientificamerican06261886-406.pdf.

Anonymous:1886:PGb

[7872] Anonymous. Prismatic gunpowder. Scientific American, 54(26): 406–407, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n26/pdf/scientificamerican06261886-
406a.pdf.

Anonymous:1886:WBa

[7873] Anonymous. The water buck. Scientific American, 54(26):407, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n26/pdf/scientificamerican06261886-407.pdf.

Anonymous:1886:HEa

[7874] Anonymous. A hairy elephant. Scientific American, 54(26):408, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n26/pdf/scientificamerican06261886-408.pdf.

Anonymous:1886:EIs

[7875] Anonymous. Engineering inventions. Scientific American, 54(26):409, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n26/pdf/scientificamerican06261886-409.pdf.

Anonymous:1886:MIt

[7876] Anonymous. Miscellaneous inventions. Scientific American, 54(26):409, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n26/pdf/scientificamerican06261886-409a.pdf.

Anonymous:1886:Ig

[7877] Anonymous. To inventors. Scientific American, 54(26):409, June 26, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n26/pdf/scientificamerican06261886-409e.pdf.

Anonymous:1886:Ct

[7878] Anonymous. Correspondence. Scientific American, 54(3):37–38, January 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n3/pdf/scientificamerican01161886-37.pdf.

Anonymous:1886:BNBp

[7879] Anonymous. Books: New books and publications. Scientific American, 54(3):42, January 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n3/pdf/scientificamerican01161886-
42c.pdf.

Anonymous:1886:BPt

[7880] Anonymous. Business and personal. Scientific American, 54 (3):42–43, January 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n3/pdf/scientificamerican01161886-
42d.pdf.

Anonymous:1886:NQs

[7881] Anonymous. Notes and queries. Scientific American, 54(3):43, January 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n3/pdf/scientificamerican01161886-43.pdf.

Anonymous:1886:IIs

[7882] Anonymous. Index of inventions. Scientific American, 54(3):43–45, January 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n3/pdf/scientificamerican01161886-43a.pdf.

Anonymous:1886:PNPs

[7883] Anonymous. Patent notes: Patents. Scientific American, 54(3):46, January 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n3/pdf/scientificamerican01161886-46.pdf.

Anonymous:1886:RRAa

[7884] Anonymous. A remarkable railroad accident. Scientific American, 54(3):31–32, January 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n3/pdf/scientificamerican01161886-
31.pdf.

Anonymous:1886:MCb

[7885] Anonymous. Mind cures. Scientific American, 54(3):32–33, January 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n3/pdf/scientificamerican01161886-32.pdf.

Anonymous:1886:FTE

[7886] Anonymous. The fish torpedo and its enemy. Scientific American, 54(3):33, January 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n3/pdf/scientificamerican01161886-
33.pdf.

Anonymous:1886:FSPb

[7887] Anonymous. Fire from steam pipes. Scientific American, 54(3):33, January 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n3/pdf/scientificamerican01161886-33a.pdf.

Anonymous:1886:PNk

[7888] Anonymous. Photographic notes. Scientific American, 54(3):33, January 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n3/pdf/scientificamerican01161886-33b.pdf.

Anonymous:1886:IAG

[7889] Anonymous. Improved automatic gas burner. Scientific American, 54(3):34, January 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n3/pdf/scientificamerican01161886-
34.pdf.

Anonymous:1886:IWa

[7890] Anonymous. Improved wrench. Scientific American, 54(3):34, January 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n3/pdf/scientificamerican01161886-34a.pdf.

Anonymous:1886:RGV

[7891] Anonymous. Running gear for vehicles. Scientific American, 54 (3):34, January 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n3/pdf/scientificamerican01161886-
34b.pdf.

Anonymous:1886:IMP

[7892] Anonymous. An improved middlings purifier. Scientific American, 54(3):34–35, January 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n3/pdf/scientificamerican01161886-
34c.pdf.

Anonymous:1886:AWM

[7893] Anonymous. Aerated water machinery and auxiliary apparatus. Scientific American, 54(3):35–37, January 16, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v54/
n3/pdf/scientificamerican01161886-35.pdf.

Anonymous:1886:NCa

[7894] Anonymous. A novel clock. Scientific American, 54(3):38–39, January 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n3/pdf/scientificamerican01161886-38.pdf.

Anonymous:1886:STCa

[7895] Anonymous. Sixty ton crane. Scientific American, 54(3):39–41, January 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n3/pdf/scientificamerican01161886-39.pdf.

Anonymous:1886:Ba

[7896] Anonymous. Burmah. Scientific American, 54(3):41–42, January 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n3/pdf/scientificamerican01161886-41.pdf.

Anonymous:1886:EIt (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n3/pdf/scientificamerican01161886-
42.pdf.

Anonymous:1886:AIm

[7898] Anonymous. Agricultural inventions. Scientific American, 54 (3):42, January 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n3/pdf/scientificamerican01161886-
42a.pdf.

Anonymous:1886:MIu

[7899] Anonymous. Miscellaneous inventions. Scientific American, 54 (3):42, January 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n3/pdf/scientificamerican01161886-
42b.pdf.

Anonymous:1886:Cu

[7900] Anonymous. Correspondence. Scientific American, 54(4):53–54, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n4/pdf/scientificamerican01231886-53.pdf.

Anonymous:1886:BNBq

[7901] Anonymous. Books: New books and publications. Scientific American, 54(4):58, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n4/pdf/scientificamerican01231886-
58c.pdf.

Anonymous:1886:BPu

[7902] Anonymous. Business and personal. Scientific American, 54(4):58, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n4/pdf/scientificamerican01231886-58d.pdf.

Anonymous:1886:NQt

[7903] Anonymous. Notes and queries. Scientific American, 54(4):58–59, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n4/pdf/scientificamerican01231886-58e.pdf.

Anonymous:1886:IIt

[7904] Anonymous. Index of inventions. Scientific American, 54(4):59–60, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n4/pdf/scientificamerican01231886-59.pdf.

Anonymous:1886:PNPt

[7905] Anonymous. Patent notes: Patents. Scientific American, 54(4):62, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n4/pdf/scientificamerican01231886-62.pdf.

Anonymous:1886:IFD

[7906] Anonymous. Indicating fire damp in coal mines. Scientific American, 54(4):47, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n4/pdf/scientificamerican01231886-
47.pdf.

Anonymous:1886:NIS

[7907] Anonymous. A new invention to secure safety in coal mines. Scientific American, 54(4):48, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n4/pdf/scientificamerican01231886-
48.pdf.

Anonymous:1886:NFAb

[7908] Anonymous. Is nerve force an agent in coloration. Scientific American, 54(4):48, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n4/pdf/scientificamerican01231886-
48a.pdf.

Anonymous:1886:RSSb

[7909] Anonymous. Removal of snow from streets by steam. Scientific American, 54(4):48, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n4/pdf/scientificamerican01231886-
48b.pdf.

Anonymous:1886:PNl

[7910] Anonymous. Photographic notes. Scientific American, 54(4):49, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n4/pdf/scientificamerican01231886-49.pdf.

Anonymous:1886:IPBb

[7911] Anonymous. Improved packing box. Scientific American, 54(4):50, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n4/pdf/scientificamerican01231886-50.pdf.

Anonymous:1886:ASD

[7912] Anonymous. Automatic switch for dovetailing machines. Scientific American, 54(4):50, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n4/pdf/scientificamerican01231886-
50a.pdf.

Anonymous:1886:FTP

[7913] Anonymous. Fastening for tied packages. Scientific American, 54(4):50, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n4/pdf/scientificamerican01231886-
50b.pdf.

Anonymous:1886:MPP

[7914] Anonymous. Metallic piston packing. Scientific American, 54 (4):50–51, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n4/pdf/scientificamerican01231886-
50c.pdf.

Anonymous:1886:MDC

[7915] Anonymous. Movable dressing closet for sleeping cars. Scientific American, 54(4):51, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n4/pdf/scientificamerican01231886-
51.pdf.

Anonymous:1886:SHS

[7916] Anonymous. Siemens and Halske’s soot indicator. Scientific American, 54(4):51, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n4/pdf/scientificamerican01231886-
51a.pdf.

Anonymous:1886:IMM

[7917] Anonymous. Improved milling machine. Scientific American, 54 (4):54, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n4/pdf/scientificamerican01231886-
54.pdf.

Anonymous:1886:ESP

[7918] Anonymous. The Empire steam pump. Scientific American, 54 (4):54–55, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n4/pdf/scientificamerican01231886-
54a.pdf.

Anonymous:1886:CIF

[7919] Anonymous. A curious ice formation. Scientific American, 54 (4):55, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n4/pdf/scientificamerican01231886-
55.pdf.

Coryell:1886:CR

[7920] John R. Coryell. The California roadrunner. Scientific American, 54(4):55–57, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n4/pdf/scientificamerican01231886-
55a.pdf.

Anonymous:1886:PSM

[7921] Anonymous. Parenty’s smoking machine. Scientific American, 54(4):57, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n4/pdf/scientificamerican01231886-
57.pdf.

Anonymous:1886:EIu

[7922] Anonymous. Engineering inventions. Scientific American, 54(4): 58, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n4/pdf/scientificamerican01231886-
58.pdf.

Anonymous:1886:AIn

[7923] Anonymous. Agricultural inventions. Scientific American, 54 (4):58, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n4/pdf/scientificamerican01231886-
58a.pdf.

Anonymous:1886:MIv

[7924] Anonymous. Miscellaneous inventions. Scientific American, 54 (4):58, January 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n4/pdf/scientificamerican01231886-
58b.pdf.

Anonymous:1886:Cv

[7925] Anonymous. Correspondence. Scientific American, 54(5):69, January 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n5/pdf/scientificamerican01301886-69.pdf.

Anonymous:1886:BNBr

[7926] Anonymous. Books: New books and publications. Scientific American, 54(5):74, January 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n5/pdf/scientificamerican01301886-
74c.pdf.

Anonymous:1886:BPv

[7927] Anonymous. Business and personal. Scientific American, 54(5):74, January 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n5/pdf/scientificamerican01301886-74d.pdf.

Anonymous:1886:NQu

[7928] Anonymous. Notes and queries. Scientific American, 54(5):74–75, January 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n5/pdf/scientificamerican01301886-74e.pdf.

Anonymous:1886:IIu

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n5/pdf/scientificamerican01301886-75.pdf.

Anonymous:1886:PNPu

[7930] Anonymous. Patent notes: Patents. Scientific American, 54(5):78, January 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n5/pdf/scientificamerican01301886-78.pdf.

Anonymous:1886:NYC

[7931] Anonymous. The New York Cable Railway. Scientific American, 54(5):63, January 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n5/pdf/scientificamerican01301886-
63.pdf.

Anonymous:1886:IDF

[7932] Anonymous. The Interior Department favors a suit against the Bell Telephone Patent. Scientific American, 54(5):64, January 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n5/pdf/scientificamerican01301886-64.pdf.

Anonymous:1886:SCL

[7933] Anonymous. Sibley College Lectures, Cornell University. Scientific American, 54(5):64, January 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n5/pdf/scientificamerican01301886-
64a.pdf.

Anonymous:1886:SHM

[7934] Anonymous. Steam heat for melting snow in streets. Scientific American, 54(5):64, January 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n5/pdf/scientificamerican01301886-
64b.pdf.

Anonymous:1886:PNm

[7935] Anonymous. Photographic notes. Scientific American, 54(5): 65–66, January 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n5/pdf/scientificamerican01301886-
65.pdf.

Anonymous:1886:FCF

[7936] Anonymous. Feeding coal to furnaces. Scientific American, 54 (5):66, January 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n5/pdf/scientificamerican01301886-
66.pdf.

Anonymous:1886:MA

[7937] Anonymous. The micro-audiphone. Scientific American, 54(5):66, January 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n5/pdf/scientificamerican01301886-66a.pdf.

Anonymous:1886:AMI

[7938] Anonymous. Apparatus for making illuminating gas. Scientific American, 54(5):66, January 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n5/pdf/scientificamerican01301886-
66b.pdf.

Anonymous:1886:CMS

[7939] Anonymous. Combined milling and slotting machine. Scientific American, 54(5):67–68, January 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n5/pdf/scientificamerican01301886-
67.pdf.

Anonymous:1886:MSF

[7940] Anonymous. Making steam by friction. Scientific American, 54 (5):70–71, January 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n5/pdf/scientificamerican01301886-
70.pdf.

Anonymous:1886:TFa

[7941] Anonymous. The Termes Flavipes. Scientific American, 54(5):71, January 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n5/pdf/scientificamerican01301886-71.pdf.

Anonymous:1886:FB

[7942] Anonymous. Feathered butchers. Scientific American, 54(5):71–73, January 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n5/pdf/scientificamerican01301886-71a.pdf.

Anonymous:1886:WF

[7943] Anonymous. Whittled fans. Scientific American, 54(5):73, January 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n5/pdf/scientificamerican01301886-73.pdf.

Anonymous:1886:EIv

[7944] Anonymous. Engineering inventions. Scientific American, 54(5): 74, January 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n5/pdf/scientificamerican01301886-
74.pdf.

Anonymous:1886:AIo

[7945] Anonymous. Agricultural inventions. Scientific American, 54 (5):74, January 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n5/pdf/scientificamerican01301886-
74a.pdf.

Anonymous:1886:MIw

[7946] Anonymous. Miscellaneous inventions. Scientific American, 54 (5):74, January 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n5/pdf/scientificamerican01301886-
74b.pdf.

Anonymous:1886:Cw

[7947] Anonymous. Correspondence. Scientific American, 54(6):88, February 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n6/pdf/scientificamerican02061886-88.pdf.

Anonymous:1886:BNBs

[7948] Anonymous. Books: New books and publications. Scientific American, 54(6):90, February 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n6/pdf/scientificamerican02061886-
90c.pdf.

Anonymous:1886:BPw

[7949] Anonymous. Business and personal. Scientific American, 54 (6):90, February 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n6/pdf/scientificamerican02061886-
90d.pdf.

Anonymous:1886:NQv

[7950] Anonymous. Notes and queries. Scientific American, 54(6):90– 91, February 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n6/pdf/scientificamerican02061886-
90e.pdf.

Anonymous:1886:IIv

[7951] Anonymous. Index of inventions. Scientific American, 54(6): 91–92, February 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n6/pdf/scientificamerican02061886-
91.pdf.

Anonymous:1886:PNPv

[7952] Anonymous. Patent notes: Patents. Scientific American, 54 (6):94, February 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n6/pdf/scientificamerican02061886-
94.pdf.

Anonymous:1886:SCW

[7953] Anonymous. Stoppage of the Cleveland Water Tunnel by ice spicules. Scientific American, 54(6):80, February 6, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v54/
n6/pdf/scientificamerican02061886-80.pdf.

Anonymous:1886:APS

[7954] Anonymous. American precious stones. Scientific American, 54 (6):80–81, February 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n6/pdf/scientificamerican02061886-
80a.pdf.

Proctor:1886:NSJa

[7955] Richard A. Proctor. Night sky — January and February. Scientific American, 54(6):81, February 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n6/pdf/scientificamerican02061886-
81.pdf.

Anonymous:1886:HEb

[7956] Anonymous. Honey extractor. Scientific American, 54(6):82, February 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n6/pdf/scientificamerican02061886-82.pdf.

Anonymous:1886:EP

[7957] Anonymous. Expanding pulley. Scientific American, 54(6):82, February 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n6/pdf/scientificamerican02061886-82a.pdf.

Anonymous:1886:STb

[7958] Anonymous. Spring trap. Scientific American, 54(6):82, February 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n6/pdf/scientificamerican02061886-82b.pdf.

Anonymous:1886:SCb

[7959] Anonymous. Strainer and cut-off. Scientific American, 54(6):82, February 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n6/pdf/scientificamerican02061886-82c.pdf.

Anonymous:1886:MCBa

[7960] Anonymous. Machine for carding buttons by electricity. Scientific American, 54(6):83, February 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n6/pdf/scientificamerican02061886-
83.pdf.

Anonymous:1886:STT

[7961] Anonymous. A simple traveling top. Scientific American, 54 (6):83–84, February 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n6/pdf/scientificamerican02061886-
83a.pdf.

Anonymous:1886:SCc

[7962] Anonymous. Saucepan and cover. Scientific American, 54(6):84, February 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n6/pdf/scientificamerican02061886-84.pdf.

Anonymous:1886:ISMa

[7963] Anonymous. Improvement in stone and marble sawing machines. Scientific American, 54(6):84, February 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v54/n6/pdf/scientificamerican02061886-
84a.pdf.

Anonymous:1886:PNn

[7964] Anonymous. Photographic notes. Scientific American, 54(6):84, February 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n6/pdf/scientificamerican02061886-84b.pdf.

Anonymous:1886:IFH

[7965] Anonymous. Improved fish hook. Scientific American, 54(6):84, February 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n6/pdf/scientificamerican02061886-84c.pdf.

Anonymous:1886:BBT

[7966] Anonymous. The Big Bend Tunnel in Butte County, California. Scientific American, 54(6):85, February 6, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v54/
n6/pdf/scientificamerican02061886-85.pdf.

Anonymous:1886:CEA

[7967] Anonymous. Condensing engine, Antwerp Exhibition. Scientific American, 54(6):86–87, February 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n6/pdf/scientificamerican02061886-
86.pdf.

Anonymous:1886:DOT

[7968] Anonymous. Design for an observatory tower. Scientific American, 54(6):87, February 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n6/pdf/scientificamerican02061886-
87.pdf.

Anonymous:1886:NMM

[7969] Anonymous. A new mass of meteoric iron. Scientific American, 54(6):89, February 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n6/pdf/scientificamerican02061886-
89.pdf.

Anonymous:1886:EIw

[7970] Anonymous. Engineering inventions. Scientific American, 54(6): 90, February 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n6/pdf/scientificamerican02061886-
90.pdf.

Anonymous:1886:AIp

[7971] Anonymous. Agricultural inventions. Scientific American, 54 (6):90, February 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n6/pdf/scientificamerican02061886-
90a.pdf.

Anonymous:1886:MIx

[7972] Anonymous. Miscellaneous inventions. Scientific American, 54 (6):90, February 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n6/pdf/scientificamerican02061886-
90b.pdf.

Anonymous:1886:Cx

[7973] Anonymous. Correspondence. Scientific American, 54(7):101, February 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n7/pdf/scientificamerican02131886-101.pdf.

Anonymous:1886:BNBt

[7974] Anonymous. Books: New books and publications. Scientific American, 54(7):106, February 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n7/pdf/scientificamerican02131886-
106c.pdf.

Anonymous:1886:BPx

[7975] Anonymous. Business and personal. Scientific American, 54(7): 106, February 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n7/pdf/scientificamerican02131886-
106d.pdf.

Anonymous:1886:NQw

[7976] Anonymous. Notes and queries. Scientific American, 54(7):106– 107, February 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n7/pdf/scientificamerican02131886-
106e.pdf.

Anonymous:1886:IIw

[7977] Anonymous. Index of inventions. Scientific American, 54(7): 107–108, February 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n7/pdf/scientificamerican02131886-
107.pdf.

Anonymous:1886:PNPw

[7978] Anonymous. Patent notes: Patents. Scientific American, 54 (7):109, February 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n7/pdf/scientificamerican02131886-
109.pdf.

Anonymous:1886:MSa

[7979] Anonymous. The Marvin safe. Scientific American, 54(7):95, February 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n7/pdf/scientificamerican02131886-95.pdf.

Anonymous:1886:OAR

[7980] Anonymous. Is our average rainfall diminishing? Scientific American, 54(7):96, February 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n7/pdf/scientificamerican02131886-
96.pdf.

Anonymous:1886:CAF

[7981] Anonymous. Carbonic acid as a fire extinguisher at sea. Scientific American, 54(7):96–97, February 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n7/pdf/scientificamerican02131886-
96a.pdf.

Anonymous:1886:PNo

[7982] Anonymous. Photographic notes. Scientific American, 54(7):97, February 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n7/pdf/scientificamerican02131886-97.pdf.

Anonymous:1886:IFE

[7983] Anonymous. Improved fire escape. Scientific American, 54(7):98, February 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n7/pdf/scientificamerican02131886-98.pdf.

Anonymous:1886:IBa

[7984] Anonymous. Ironing board. Scientific American, 54(7):98, February 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n7/pdf/scientificamerican02131886-98a.pdf.

Anonymous:1886:IBb

[7985] Anonymous. Improved bellows. Scientific American, 54(7):98, February 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n7/pdf/scientificamerican02131886-98b.pdf.

Anonymous:1886:CDa

[7986] Anonymous. Clothes drier. Scientific American, 54(7):98, February 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n7/pdf/scientificamerican02131886-98c.pdf.

Anonymous:1886:PL

[7987] Anonymous. The principle of the loom. Scientific American, 54(7):99, February 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n7/pdf/scientificamerican02131886-
99.pdf.

Anonymous:1886:Bb

[7988] Anonymous. Brickwork. Scientific American, 54(7):99, February 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n7/pdf/scientificamerican02131886-99a.pdf.

Anonymous:1886:IEM

[7989] Anonymous. Improvements in electro magnets. Scientific American, 54(7):102, February 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n7/pdf/scientificamerican02131886-
102.pdf.

Seiss:1886:BRC

[7990] C. Few Seiss. The bullfrog (Rana Catesbiana, Shaw). Scientific American, 54(7):103, February 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n7/pdf/scientificamerican02131886-
103.pdf.

Anonymous:1886:OOP

[7991] Anonymous. Old organs and pianofortes. Scientific American, 54 (7):103–104, February 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n7/pdf/scientificamerican02131886-
103a.pdf.

Anonymous:1886:CSH

[7992] Anonymous. Cowen steam hose. Scientific American, 54(7):104, February 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n7/pdf/scientificamerican02131886-104.pdf.

Anonymous:1886:ITBa

[7993] Anonymous. Improved treadle for bicycles. Scientific American, 54(7):105, February 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n7/pdf/scientificamerican02131886-
105.pdf.

Anonymous:1886:EIx

[7994] Anonymous. Engineering inventions. Scientific American, 54(7): 106, February 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n7/pdf/scientificamerican02131886-
106.pdf.

Anonymous:1886:AIq

[7995] Anonymous. Agricultural inventions. Scientific American, 54(7): 106, February 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n7/pdf/scientificamerican02131886-
106a.pdf.

Anonymous:1886:MIy

[7996] Anonymous. Miscellaneous inventions. Scientific American, 54 (7):106, February 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n7/pdf/scientificamerican02131886-
106b.pdf.

Anonymous:1886:Cy

[7997] Anonymous. Correspondence. Scientific American, 54(8):117, February 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n8/pdf/scientificamerican02201886-117.pdf.

Anonymous:1886:BPy

[7998] Anonymous. Business and personal. Scientific American, 54(8): 121, February 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n8/pdf/scientificamerican02201886-
121c.pdf.

Anonymous:1886:NQx

[7999] Anonymous. Notes and queries. Scientific American, 54(8):121– 123, February 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n8/pdf/scientificamerican02201886-
121d.pdf.

Anonymous:1886:IIx

[8000] Anonymous. Index of inventions. Scientific American, 54(8): 123–125, February 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n8/pdf/scientificamerican02201886-
123.pdf.

Anonymous:1886:PNPx

[8001] Anonymous. Patent notes: Patents. Scientific American, 54 (8):126, February 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n8/pdf/scientificamerican02201886-
126.pdf.

Anonymous:1886:HIC

[8002] Anonymous. The Hoboken Inclined Cable Railway. Scientific American, 54(8):111–112, February 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n8/pdf/scientificamerican02201886-
111.pdf.

Anonymous:1886:SOP

[8003] Anonymous. Seat and origin of pain. Scientific American, 54 (8):112, February 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n8/pdf/scientificamerican02201886-
112.pdf.

Anonymous:1886:EPDa

[8004] Anonymous. The exhaustion of the public domain. Scientific American, 54(8):112, February 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n8/pdf/scientificamerican02201886-
112a.pdf.

Anonymous:1886:AIS

[8005] Anonymous. Aluminum iron and steel. Scientific American, 54 (8):113, February 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n8/pdf/scientificamerican02201886-
113.pdf.

Hovey:1886:IPS

[8006] H. C. Hovey. The Ice Palace at St. Paul. Scientific American, 54(8):113, February 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n8/pdf/scientificamerican02201886-
113a.pdf.

Anonymous:1886:PNq

[8007] Anonymous. Photographic notes. Scientific American, 54(8): 113, February 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n8/pdf/scientificamerican02201886-
113b.pdf.

Anonymous:1886:LDa

[8008] Anonymous. Lumber drier. Scientific American, 54(8):114, February 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n8/pdf/scientificamerican02201886-114.pdf.

Anonymous:1886:CPCb

[8009] Anonymous. Combined plow, cultivator, and harrow. Scientific American, 54(8):114, February 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n8/pdf/scientificamerican02201886-
114a.pdf.

Anonymous:1886:ADB

[8010] Anonymous. Adjustable double bedstead. Scientific American, 54(8):114, February 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n8/pdf/scientificamerican02201886-
114b.pdf.

Anonymous:1886:TGC

[8011] Anonymous. Trellis for garden crops. Scientific American, 54 (8):114, February 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n8/pdf/scientificamerican02201886-
114c.pdf.

Anonymous:1886:ISCb

[8012] Anonymous. Improved stone channeling machine. Scientific American, 54(8):115–116, February 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n8/pdf/scientificamerican02201886-
115.pdf.

Anonymous:1886:IFL

[8013] Anonymous. An improved farm locomotive. Scientific American, 54(8):118, February 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n8/pdf/scientificamerican02201886-
118.pdf.

Anonymous:1886:FPP

[8014] Anonymous. Frame plate planing machine. Scientific American, 54(8):118, February 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n8/pdf/scientificamerican02201886-
118a.pdf.

Anonymous:1886:ESR

[8015] Anonymous. The Edison system of railway telegraphy. Scientific American, 54(8):119–120, February 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n8/pdf/scientificamerican02201886-
119.pdf.

Anonymous:1886:BL

[8016] Anonymous. Bicycle leg. Scientific American, 54(8):120, February 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n8/pdf/scientificamerican02201886-120.pdf.

Anonymous:1886:IBSa

[8017] Anonymous. Improved bottle and stopper. Scientific American, 54(8):120, February 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n8/pdf/scientificamerican02201886-
120a.pdf.

Anonymous:1886:EIy

[8018] Anonymous. Engineering inventions. Scientific American, 54(8): 121, February 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n8/pdf/scientificamerican02201886-
121.pdf.

Anonymous:1886:AIr

[8019] Anonymous. Agricultural inventions. Scientific American, 54(8): 121, February 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n8/pdf/scientificamerican02201886-
121a.pdf.

Anonymous:1886:MIz

[8020] Anonymous. Miscellaneous inventions. Scientific American, 54 (8):121, February 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n8/pdf/scientificamerican02201886-
121b.pdf.

Anonymous:1886:Cz

[8021] Anonymous. Correspondence. Scientific American, 54(9):137, February 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n9/pdf/scientificamerican02271886-137.pdf.

Anonymous:1886:BNBu

[8022] Anonymous. Books: New books and publications. Scientific American, 54(9):138, February 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n9/pdf/scientificamerican02271886-
138c.pdf.

Anonymous:1886:BPz

[8023] Anonymous. Business and personal. Scientific American, 54(9): 138, February 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n9/pdf/scientificamerican02271886-
138d.pdf.

Anonymous:1886:NQy

[8024] Anonymous. Notes and queries. Scientific American, 54(9):138– 139, February 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n9/pdf/scientificamerican02271886-
138e.pdf.

Anonymous:1886:IIy

[8025] Anonymous. Index of inventions. Scientific American, 54(9): 139–140, February 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n9/pdf/scientificamerican02271886-
139.pdf.

Anonymous:1886:PNPy

[8026] Anonymous. Patent notes: Patents. Scientific American, 54 (9):142, February 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n9/pdf/scientificamerican02271886-
142.pdf.

Anonymous:1886:UNG

[8027] Anonymous. The use of natural gas at Pittsburg. Scientific American, 54(9):127, February 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n9/pdf/scientificamerican02271886-
127.pdf.

Anonymous:1886:GCR

[8028] Anonymous. The La Guayra and Caracas Railway. — a reclamation. Scientific American, 54(9):128, February 27, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v54/
n9/pdf/scientificamerican02271886-128.pdf.

Anonymous:1886:TCc

[8029] Anonymous. The telephone controversy. Scientific American, 54 (9):128, February 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n9/pdf/scientificamerican02271886-
128a.pdf.

Anonymous:1886:PNr

[8030] Anonymous. Photographic notes. Scientific American, 54(9): 129, February 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n9/pdf/scientificamerican02271886-
129.pdf.

Anonymous:1886:WSM

[8031] Anonymous. Wood sawing machine. Scientific American, 54(9): 130, February 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n9/pdf/scientificamerican02271886-
130.pdf.

Anonymous:1886:IFB

[8032] Anonymous. An improved folding bed. Scientific American, 54 (9):130, February 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n9/pdf/scientificamerican02271886-
130a.pdf.

Anonymous:1886:IBSb

[8033] Anonymous. An improved bob-sleigh. Scientific American, 54 (9):130, February 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n9/pdf/scientificamerican02271886-
130b.pdf.

Anonymous:1886:BTS

[8034] Anonymous. Buggy top support. Scientific American, 54(9):130, February 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n9/pdf/scientificamerican02271886-130c.pdf.

Anonymous:1886:NGS

[8035] Anonymous. A nitro-glycerine shell. Scientific American, 54(9): 131, February 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n9/pdf/scientificamerican02271886-
131.pdf.

Anonymous:1886:DET

[8036] Anonymous. Distribution of energy by transformers. Scientific American, 54(9):131–132, February 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n9/pdf/scientificamerican02271886-
131a.pdf.

Anonymous:1886:CLW

[8037] Anonymous. The Clyde Locomotive Works. Scientific American, 54 (9):134–136, February 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n9/pdf/scientificamerican02271886-
134.pdf.

Anonymous:1886:ADR

[8038] Anonymous. Automatic damper regulator. Scientific American, 54(9):136, February 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n9/pdf/scientificamerican02271886-
136.pdf.

Anonymous:1886:ISd

[8039] Anonymous. An improved saw. Scientific American, 54(9):136, February 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v54/n9/pdf/scientificamerican02271886-136a.pdf.

Anonymous:1886:NSP

[8040] Anonymous. A new screw propeller. Scientific American, 54 (9):136, February 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n9/pdf/scientificamerican02271886-
136b.pdf.

Anonymous:1886:EIz

[8041] Anonymous. Engineering inventions. Scientific American, 54(9): 138, February 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n9/pdf/scientificamerican02271886-
138.pdf.

Anonymous:1886:AIt

[8042] Anonymous. Agricultural inventions. Scientific American, 54(9): 138, February 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n9/pdf/scientificamerican02271886-
138a.pdf.

Anonymous:1886:MIba

[8043] Anonymous. Miscellaneous inventions. Scientific American, 54 (9):138, February 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v54/n9/pdf/scientificamerican02271886-
138b.pdf.

Anonymous:1886:Cba

[8044] Anonymous. Correspondence. Scientific American, 55(1):7, July 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n1/pdf/scientificamerican07031886-7.pdf.

Anonymous:1886:BNBv

[8045] Anonymous. Books: New books and publications. Scientific American, 55(1):10, July 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n1/pdf/scientificamerican07031886-
10b.pdf.

Anonymous:1886:BPba

[8046] Anonymous. Business and personal. Scientific American, 55(1):10, July 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n1/pdf/scientificamerican07031886-10c.pdf.

Anonymous:1886:NQz

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n1/pdf/scientificamerican07031886-11.pdf.

Anonymous:1886:IIz

[8048] Anonymous. Index of inventions. Scientific American, 55(1):11–12, July 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n1/pdf/scientificamerican07031886-11b.pdf.

Anonymous:1886:PNPz

[8049] Anonymous. Patent notes: Patents. Scientific American, 55(1):14, July 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n1/pdf/scientificamerican07031886-14.pdf.

Hopkins:1886:SEP

[8050] Geo. M. Hopkins. Simple experiments in polarized light. Scientific American, 55(1):1, July 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n1/pdf/scientificamerican07031886-
1.pdf.

Anonymous:1886:IGI

[8051] Anonymous. The illuminating gas industry. Scientific American, 55(1): 2, July 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n1/pdf/scientificamerican07031886-2.pdf.

Anonymous:1886:PNs

[8052] Anonymous. Photographic notes. Scientific American, 55(1):2, July 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n1/pdf/scientificamerican07031886-2a.pdf.

Anonymous:1886:CAL

[8053] Anonymous. Compressed air for life boats. Scientific American, 55(1): 3, July 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n1/pdf/scientificamerican07031886-3.pdf.

Proctor:1886:NSJb

[8054] Richard A. Proctor. Night sky — June and July. Scientific American, 55(1):3, July 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n1/pdf/scientificamerican07031886-
3a.pdf.

Anonymous:1886:MH

[8055] Anonymous. Music holder. Scientific American, 55(1):4, July 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n1/pdf/scientificamerican07031886-4.pdf.

Anonymous:1886:LE

[8056] Anonymous. Liquid emptier. Scientific American, 55(1):4, July 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n1/pdf/scientificamerican07031886-4a.pdf.

Anonymous:1886:ISIb

[8057] Anonymous. Improved sad iron. Scientific American, 55(1):4, July 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n1/pdf/scientificamerican07031886-4b.pdf.

Anonymous:1886:SKP

[8058] Anonymous. Scraping knife for painters. Scientific American, 55(1):4, July 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n1/pdf/scientificamerican07031886-4c.pdf.

Anonymous:1886:ICSa

[8059] Anonymous. Improved car seat. Scientific American, 55(1):4, July 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n1/pdf/scientificamerican07031886-4d.pdf.

Anonymous:1886:ASE

[8060] Anonymous. Automatic stop for engines. Scientific American, 55(1):5, July 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n1/pdf/scientificamerican07031886-5.pdf.

Anonymous:1886:IWM

[8061] Anonymous. Improved woodworking machine. Scientific American, 55(1):5, July 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n1/pdf/scientificamerican07031886-
5a.pdf.

Anonymous:1886:SWD

[8062] Anonymous. Steel wire door mat. Scientific American, 55(1):5, July 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n1/pdf/scientificamerican07031886-5b.pdf.

Anonymous:1886:DRPc

[8063] Anonymous. Decisions relating to patents. Scientific American, 55(1): 7, July 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n1/pdf/scientificamerican07031886-7a.pdf.

Anonymous:1886:IBF

[8064] Anonymous. Improved boring and facing machine. Scientific American, 55(1):8, July 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n1/pdf/scientificamerican07031886-
8.pdf.

Anonymous:1886:TCLb

[8065] Anonymous. Triple calla lily. Scientific American, 55(1):9, July 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n1/pdf/scientificamerican07031886-9.pdf.

Sloane:1886:EAP

[8066] T. O’Connor Sloane. Experiments in atmospheric pressure and condensation of steam. Scientific American, 55(1):9, July 3, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v55/
n1/pdf/scientificamerican07031886-9a.pdf.

Anonymous:1886:EIba

[8067] Anonymous. Engineering inventions. Scientific American, 55(1):10, July 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n1/pdf/scientificamerican07031886-10.pdf.

Anonymous:1886:MIbb

[8068] Anonymous. Miscellaneous inventions. Scientific American, 55(1):10, July 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n1/pdf/scientificamerican07031886-10a.pdf.

Anonymous:1886:Ih

[8069] Anonymous. To inventors. Scientific American, 55(1):11, July 3, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n1/pdf/scientificamerican07031886-11a.pdf.

Anonymous:1886:BPbb

[8070] Anonymous. Business and personal. Scientific American, 55 (10):155, September 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n10/pdf/scientificamerican09041886-
155d.pdf.

Anonymous:1886:IIba

[8071] Anonymous. Index of inventions. Scientific American, 55(10): 155–156, September 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n10/pdf/scientificamerican09041886-
155f.pdf.

Anonymous:1886:PNPba

[8072] Anonymous. Patent notes: Patents. Scientific American, 55 (10):158, September 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n10/pdf/scientificamerican09041886-
158.pdf.

Anonymous:1886:WSN

[8073] Anonymous. The water supply of New York City. Scientific American, 55(10):144, September 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n10/pdf/scientificamerican09041886-
144.pdf.

Anonymous:1886:SGb

[8074] Anonymous. Scoring of grindstones. Scientific American, 55 (10):144, September 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n10/pdf/scientificamerican09041886-
144a.pdf.

Anonymous:1886:PNt

[8075] Anonymous. Photographic notes. Scientific American, 55(10): 144–145, September 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n10/pdf/scientificamerican09041886-
144b.pdf.

Anonymous:1886:BCG

[8076] Anonymous. Birmingham corporation gas holders. Scientific American, 55(10):146–147, September 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n10/pdf/scientificamerican09041886-
146.pdf.

Anonymous:1886:MCE

[8077] Anonymous. A milk cooler that excludes the air. Scientific American, 55(10):147, September 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n10/pdf/scientificamerican09041886-
147.pdf.

Anonymous:1886:OCW

[8078] Anonymous. An oil can with wick tube besides a nozzle. Scientific American, 55(10):147, September 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n10/pdf/scientificamerican09041886-
147a.pdf.

Anonymous:1886:TPP

[8079] Anonymous. A traveling post and pile driver. Scientific American, 55(10):147, September 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n10/pdf/scientificamerican09041886-
147b.pdf.

Anonymous:1886:AGP

[8080] Anonymous. Automatic guard and platform doors for railway cars. Scientific American, 55(10):148, September 4, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v55/
n10/pdf/scientificamerican09041886-148.pdf.

Anonymous:1886:FJC

[8081] Anonymous. A fruit jar cover fitted with air pump. Scientific American, 55(10):148, September 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n10/pdf/scientificamerican09041886-
148a.pdf.

Anonymous:1886:SCd

[8082] Anonymous. Surface condenser. Scientific American, 55(10): 148, September 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n10/pdf/scientificamerican09041886-
148b.pdf.

Anonymous:1886:WHD

[8083] Anonymous. Worrell’s “Hercules” drier. Scientific American, 55 (10):149, September 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n10/pdf/scientificamerican09041886-
149.pdf.

Anonymous:1886:MSW

[8084] Anonymous. Method of setting wagon tires without heating. Scientific American, 55(10):149, September 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n10/pdf/scientificamerican09041886-
149a.pdf.

Anonymous:1886:FNM

[8085] Anonymous. French Naval maneuvers off the coast of Corsica. Scientific American, 55(10):150–151, September 4, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v55/
n10/pdf/scientificamerican09041886-150.pdf.

Sloane:1886:ESC

[8086] T. O’Conor Sloane. Experiments in sound. — Chladni’s plates. Scientific American, 55(10):151, September 4, 1886. CODEN SCAMAC.

ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v55/n10/pdf/scientificamerican09041886-
151.pdf.

Anonymous:1886:DRPd

[8087] Anonymous. Decisions relating to patents. Scientific American, 55(10):152, September 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n10/pdf/scientificamerican09041886-
152.pdf.

Anonymous:1886:TSG

[8088] Anonymous. Temporary steering gear. Scientific American, 55 (10):152–153, September 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n10/pdf/scientificamerican09041886-
152a.pdf.

Anonymous:1886:WSG

[8089] Anonymous. A walking stick gun. Scientific American, 55(10): 153–154, September 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n10/pdf/scientificamerican09041886-
153.pdf.

Anonymous:1886:EIbb

[8090] Anonymous. Engineering inventions. Scientific American, 55(10): 155, September 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n10/pdf/scientificamerican09041886-
155.pdf.

Anonymous:1886:MIbc

[8091] Anonymous. Mechanical invention. Scientific American, 55(10): 155, September 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n10/pdf/scientificamerican09041886-
155a.pdf.

Anonymous:1886:AIu

[8092] Anonymous. Agricultural invention. Scientific American, 55(10): 155, September 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n10/pdf/scientificamerican09041886-
155b.pdf.

Anonymous:1886:MIbd

[8093] Anonymous. Miscellaneous inventions. Scientific American, 55 (10):155, September 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n10/pdf/scientificamerican09041886-
155c.pdf.

Anonymous:1886:Ij

[8094] Anonymous. To inventors. Scientific American, 55(10):155, September 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n10/pdf/scientificamerican09041886-155e.pdf.

Anonymous:1886:BPbc

[8095] Anonymous. Business and personal. Scientific American, 55 (11):171, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n11/pdf/scientificamerican09111886-
171b.pdf.

Anonymous:1886:NQba

[8096] Anonymous. Notes and queries. Scientific American, 55(11): 171, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n11/pdf/scientificamerican09111886-
171c.pdf.

Anonymous:1886:IIbb

[8097] Anonymous. Index of inventions. Scientific American, 55(11): 171–172, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n11/pdf/scientificamerican09111886-
171e.pdf.

Anonymous:1886:Pa

[8098] Anonymous. Patents. Scientific American, 55(11):174, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n11/pdf/scientificamerican09111886-174.pdf.

Anonymous:1886:M

[8099] Anonymous. Macadamsfins. Scientific American, 55(11):159, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n11/pdf/scientificamerican09111886-159.pdf.

Anonymous:1886:PAL

[8100] Anonymous. Practical application of Leidenfrost’s drops. Scientific American, 55(11):160, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n11/pdf/scientificamerican09111886-
160.pdf.

Anonymous:1886:W

[8101] Anonymous. Out of work. Scientific American, 55(11):160, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n11/pdf/scientificamerican09111886-160a.pdf.

Proctor:1886:NSAb

[8102] Richard A. Proctor. Night sky — August and September. Scientific American, 55(11):161, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n11/pdf/scientificamerican09111886-
161.pdf.

Anonymous:1886:SFb

[8103] Anonymous. Sash fastener. Scientific American, 55(11):162, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n11/pdf/scientificamerican09111886-162.pdf.

Anonymous:1886:IWb

[8104] Anonymous. Improved wrench. Scientific American, 55(11): 162, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n11/pdf/scientificamerican09111886-
162a.pdf.

Anonymous:1886:FSa

[8105] Anonymous. File stand. Scientific American, 55(11):162, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n11/pdf/scientificamerican09111886-162b.pdf.

Anonymous:1886:DAa

[8106] Anonymous. Drawing apparatus. Scientific American, 55(11): 162, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n11/pdf/scientificamerican09111886-
162c.pdf.

Anonymous:1886:WWP

[8107] Anonymous. The Water Works of Paris. Scientific American, 55 (11):163–164, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n11/pdf/scientificamerican09111886-
163.pdf.

Anonymous:1886:IDC

[8108] Anonymous. Improved dumping cart. Scientific American, 55 (11):164, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n11/pdf/scientificamerican09111886-
164.pdf.

Anonymous:1886:CFa

[8109] Anonymous. Curtain fastener. Scientific American, 55(11):164, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n11/pdf/scientificamerican09111886-164a.pdf.

Anonymous:1886:SBH

[8110] Anonymous. The Star blacksmith’s hand blower. Scientific American, 55(11):166, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n11/pdf/scientificamerican09111886-
166.pdf.

Anonymous:1886:QCG

[8111] Anonymous. The “Quaker City” grinding mills for corn and cobs, grain, bones, etc. Scientific American, 55(11):166, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n11/pdf/scientificamerican09111886-166a.pdf.

Anonymous:1886:IGS

[8112] Anonymous. Improved gas stove. Scientific American, 55(11): 166, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n11/pdf/scientificamerican09111886-
166b.pdf.

Anonymous:1886:SBN

[8113] Anonymous. The shadow bird and its nest. Scientific American, 55(11):167, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n11/pdf/scientificamerican09111886-
167.pdf.

Anonymous:1886:IPD

[8114] Anonymous. Improved post driver. Scientific American, 55(11): 168, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n11/pdf/scientificamerican09111886-
168.pdf.

Anonymous:1886:LJ

[8115] Anonymous. Lifting jack. Scientific American, 55(11):168, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n11/pdf/scientificamerican09111886-168a.pdf.

Chadwick:1886:BSB

[8116] Henry Chadwick. Baseball — scientific batting — no. 2. Scientific American, 55(11):168, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n11/pdf/scientificamerican09111886-
168b.pdf.

Anonymous:1886:MRI

[8117] Anonymous. Mummy of Rameses II. Scientific American, 55 (11):169, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n11/pdf/scientificamerican09111886-
169.pdf.

Anonymous:1886:PFE

[8118] Anonymous. Proposed fire engine elevators for use in the New York Engine Houses. Scientific American, 55(11):170, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n11/pdf/scientificamerican09111886-170.pdf.

Anonymous:1886:PNu

[8119] Anonymous. Photographic notes. Scientific American, 55(11): 170, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n11/pdf/scientificamerican09111886-
170a.pdf.

Anonymous:1886:EIbc

[8120] Anonymous. Engineering inventions. Scientific American, 55(11): 171, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n11/pdf/scientificamerican09111886-
171.pdf.

Anonymous:1886:MIbe

[8121] Anonymous. Miscellaneous inventions. Scientific American, 55 (11):171, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n11/pdf/scientificamerican09111886-
171a.pdf.

Anonymous:1886:Ik

[8122] Anonymous. To inventors. Scientific American, 55(11):171, September 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n11/pdf/scientificamerican09111886-171d.pdf.

Anonymous:1886:Cbc

[8123] Anonymous. Correspondence. Scientific American, 55(12):181, September 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n12/pdf/scientificamerican09181886-181.pdf.

Anonymous:1886:BPbd

[8124] Anonymous. Business and personal. Scientific American, 55 (12):186, September 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n12/pdf/scientificamerican09181886-
186c.pdf.

Anonymous:1886:NQbb

[8125] Anonymous. Notes and queries. Scientific American, 55(12): 187, September 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n12/pdf/scientificamerican09181886-
187.pdf.

Anonymous:1886:IIbc

[8126] Anonymous. Index of inventions. Scientific American, 55(12): 187–188, September 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n12/pdf/scientificamerican09181886-
187b.pdf.

Anonymous:1886:PNPbb

[8127] Anonymous. Patent notes: Patents. Scientific American, 55 (12):190, September 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n12/pdf/scientificamerican09181886-
190.pdf.

Anonymous:1886:GE

[8128] Anonymous. The great earthquake. Scientific American, 55(12): 176, September 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n12/pdf/scientificamerican09181886-
176.pdf.

Anonymous:1886:RMa

[8129] Anonymous. Rare monkeys. Scientific American, 55(12):176– 177, September 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n12/pdf/scientificamerican09181886-
176a.pdf.

Anonymous:1886:SLa

[8130] Anonymous. Spring lock. Scientific American, 55(12):178, September 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n12/pdf/scientificamerican09181886-178.pdf.

Anonymous:1886:LBC

[8131] Anonymous. Locomotive boiler check valve. Scientific American, 55(12):178, September 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n12/pdf/scientificamerican09181886-
178a.pdf.

Anonymous:1886:HCb

[8132] Anonymous. Horse collar. Scientific American, 55(12):178, September 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n12/pdf/scientificamerican09181886-178b.pdf.

Anonymous:1886:SEb

[8133] Anonymous. Steam engine. Scientific American, 55(12):178, September 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n12/pdf/scientificamerican09181886-178c.pdf.

Anonymous:1886:IEW

[8134] Anonymous. Insulator for electric wires. Scientific American, 55 (12):178, September 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n12/pdf/scientificamerican09181886-
178d.pdf.

Anonymous:1886:MEC

[8135] Anonymous. Michel Eugene Chevreul. Scientific American, 55(12): 179–180, September 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n12/pdf/scientificamerican09181886-
179.pdf.

Anonymous:1886:SFc

[8136] Anonymous. Stopper fastener. Scientific American, 55(12):180, September 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n12/pdf/scientificamerican09181886-180.pdf.

Anonymous:1886:ISe

[8137] Anonymous. Improved shovel. Scientific American, 55(12):180, September 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n12/pdf/scientificamerican09181886-180a.pdf.

Anonymous:1886:GM

[8138] Anonymous. The Galatea and Mayflower. Scientific American, 55 (12):181–182, September 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n12/pdf/scientificamerican09181886-
181a.pdf.

Anonymous:1886:AAS

[8139] Anonymous. The Acme automatic safety engine. Scientific American, 55(12):183, September 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n12/pdf/scientificamerican09181886-
183.pdf.

Anonymous:1886:NLD

[8140] Anonymous. A new long distance telephone system. Scientific American, 55(12):183–184, September 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n12/pdf/scientificamerican09181886-
183a.pdf.

Hopkins:1886:IAP

[8141] Geo. M. Hopkins. An inexpensive air pump. Scientific American, 55 (12):185–186, September 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n12/pdf/scientificamerican09181886-
185.pdf.

Anonymous:1886:EIbd

[8142] Anonymous. Engineering inventions. Scientific American, 55(12): 186, September 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n12/pdf/scientificamerican09181886-
186.pdf.

Anonymous:1886:AIv

[8143] Anonymous. Agricultural inventions. Scientific American, 55 (12):186, September 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n12/pdf/scientificamerican09181886-
186a.pdf.

Anonymous:1886:MIbf

[8144] Anonymous. Miscellaneous inventions. Scientific American, 55 (12):186, September 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n12/pdf/scientificamerican09181886-
186b.pdf.

Anonymous:1886:Il

[8145] Anonymous. To inventors. Scientific American, 55(12):187, September 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n12/pdf/scientificamerican09181886-187a.pdf.

Anonymous:1886:Cbd

[8146] Anonymous. Correspondence. Scientific American, 55(13):201– 202, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n13/pdf/scientificamerican09251886-
201.pdf.

Anonymous:1886:BPbe

[8147] Anonymous. Business and personal. Scientific American, 55 (13):203, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n13/pdf/scientificamerican09251886-
203c.pdf.

Anonymous:1886:NQbc

[8148] Anonymous. Notes and queries. Scientific American, 55(13): 203, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n13/pdf/scientificamerican09251886-
203d.pdf.

Anonymous:1886:IIbd

[8149] Anonymous. Index of inventions. Scientific American, 55(13): 203–204, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n13/pdf/scientificamerican09251886-
203f.pdf.

Anonymous:1886:PNPbc

[8150] Anonymous. Patent notes: Patents. Scientific American, 55 (13):206, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n13/pdf/scientificamerican09251886-
206.pdf.

Anonymous:1886:NYW

[8151] Anonymous. The New York Water Supply. Scientific American, 55(13):192, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n13/pdf/scientificamerican09251886-
192.pdf.

Anonymous:1886:AWC

[8152] Anonymous. Arrangement of wires in cities. Scientific American, 55 (13):192–193, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n13/pdf/scientificamerican09251886-
192a.pdf.

Anonymous:1886:IGH

[8153] Anonymous. Improved gate hinge. Scientific American, 55(13): 194, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n13/pdf/scientificamerican09251886-
194.pdf.

Anonymous:1886:FFL

[8154] Anonymous. Filling fork for loom stop motions. Scientific American, 55(13):194, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n13/pdf/scientificamerican09251886-
194a.pdf.

Anonymous:1886:ICd

[8155] Anonymous. Improved churn. Scientific American, 55(13):194, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n13/pdf/scientificamerican09251886-194b.pdf.

Anonymous:1886:FPa

[8156] Anonymous. Force pump. Scientific American, 55(13):194, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n13/pdf/scientificamerican09251886-194c.pdf.

Anonymous:1886:FPb

[8157] Anonymous. Fruit picker. Scientific American, 55(13):194, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n13/pdf/scientificamerican09251886-194d.pdf.

Anonymous:1886:GI

[8158] Anonymous. Garden implement. Scientific American, 55(13): 194, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n13/pdf/scientificamerican09251886-
194e.pdf.

Anonymous:1886:SMS

[8159] Anonymous. Ship model shaping machine. Scientific American, 55 (13):195–196, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n13/pdf/scientificamerican09251886-
195.pdf.

Anonymous:1886:IBc

[8160] Anonymous.  Improved buckle.  Scientific American, 55(13):

196, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n13/pdf/scientificamerican09251886-
196.pdf.

Anonymous:1886:WAG

[8161] Anonymous. Wardrobe attachment and garment hanger. Scientific American, 55(13):196, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n13/pdf/scientificamerican09251886-
196a.pdf.

Anonymous:1886:BH

[8162] Anonymous. Bouquet holder. Scientific American, 55(13):196, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n13/pdf/scientificamerican09251886-196b.pdf.

Anonymous:1886:MMG

[8163] Anonymous. Making mammoth guns. Scientific American, 55 (13):197, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n13/pdf/scientificamerican09251886-
197.pdf.

Anonymous:1886:NGB

[8164] Anonymous. New German-built steamers. Scientific American, 55 (13):198, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n13/pdf/scientificamerican09251886-
198.pdf.

Hopkins:1886:FES

[8165] Geo. M. Hopkins. A few experiments with the simple air pump. Scientific American, 55(13):199–200, September 25, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v55/
n13/pdf/scientificamerican09251886-199.pdf.

Seiss:1886:M

[8166] C. Few Seiss. The mink. Scientific American, 55(13):200, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n13/pdf/scientificamerican09251886-200.pdf.

Anonymous:1886:NES

[8167] Anonymous. A new experiment in static electricity. Scientific American, 55(13):202, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n13/pdf/scientificamerican09251886-
202.pdf.

Anonymous:1886:EIbe

[8168] Anonymous. Engineering inventions. Scientific American, 55(13): 203, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n13/pdf/scientificamerican09251886-
203.pdf.

Anonymous:1886:AIw

[8169] Anonymous. Agricultural inventions. Scientific American, 55 (13):203, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n13/pdf/scientificamerican09251886-
203a.pdf.

Anonymous:1886:MIbg

[8170] Anonymous. Miscellaneous inventions. Scientific American, 55 (13):203, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n13/pdf/scientificamerican09251886-
203b.pdf.

Anonymous:1886:In

[8171] Anonymous. To inventors. Scientific American, 55(13):203, September 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n13/pdf/scientificamerican09251886-203e.pdf.

Anonymous:1886:Cbe

[8172] Anonymous. Correspondence. Scientific American, 55(14):216, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n14/pdf/scientificamerican10021886-216.pdf.

Anonymous:1886:BPbf

[8173] Anonymous. Business and personal. Scientific American, 55(14):218, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n14/pdf/scientificamerican10021886-218c.pdf.

Anonymous:1886:BNBw

[8174] Anonymous. Books: New books and publications. Scientific American, 55(14):219, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n14/pdf/scientificamerican10021886-
219.pdf.

Anonymous:1886:NQbd

[8175] Anonymous. Notes and queries. Scientific American, 55(14):219, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n14/pdf/scientificamerican10021886-219a.pdf.

Anonymous:1886:IIbe

[8176] Anonymous. Index of inventions. Scientific American, 55(14):219–220, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n14/pdf/scientificamerican10021886-219b.pdf.

Anonymous:1886:PNPbd

[8177] Anonymous. Patent notes: Patents. Scientific American, 55(14):222, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n14/pdf/scientificamerican10021886-222.pdf.

Hopkins:1886:DW

[8178] M. Hopkins. The demagnetization of watches. Scientific American, 55(14):207, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n14/pdf/scientificamerican10021886-
207.pdf.

Anonymous:1886:LDT

[8179] Anonymous. Long distance telephony. Scientific American, 55(14):208, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n14/pdf/scientificamerican10021886-208.pdf.

Proctor:1886:NSS

[8180] Richard A. Proctor. Night sky — September and October. Scientific American, 55(14):209, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n14/pdf/scientificamerican10021886-
209.pdf.

Anonymous:1886:ETH

[8181] Anonymous. Eaves-trough hanger. Scientific American, 55(14):210, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n14/pdf/scientificamerican10021886-210.pdf.

Anonymous:1886:DSa

[8182] Anonymous. Drag saw. Scientific American, 55(14):210, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n14/pdf/scientificamerican10021886-210a.pdf.

Anonymous:1886:CCe

[8183] Anonymous. Car coupling. Scientific American, 55(14):210, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n14/pdf/scientificamerican10021886-210b.pdf.

Anonymous:1886:IPc

[8184] Anonymous. Improved padlock. Scientific American, 55(14):210, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n14/pdf/scientificamerican10021886-210c.pdf.

Anonymous:1886:ICSb

[8185] Anonymous. Improved car starter. Scientific American, 55(14):211, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n14/pdf/scientificamerican10021886-211.pdf.

Anonymous:1886:PH

[8186] Anonymous. Pendulum hoist. Scientific American, 55(14):211, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n14/pdf/scientificamerican10021886-211a.pdf.

Anonymous:1886:ASS

[8187] Anonymous. The Allard spiral screw driver. Scientific American, 55(14):211, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n14/pdf/scientificamerican10021886-
211b.pdf.

Anonymous:1886:CAMa

[8188] Anonymous. Cutting apparatus for mowers and reapers. Scientific American, 55(14):211, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n14/pdf/scientificamerican10021886-
211c.pdf.

Anonymous:1886:CTa

[8189] Anonymous. Combination tool. Scientific American, 55(14):211, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n14/pdf/scientificamerican10021886-211d.pdf.

Anonymous:1886:NO

[8190] Anonymous. New Orleans. Scientific American, 55(14):212–213, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n14/pdf/scientificamerican10021886-212.pdf.

Sloane:1886:EPSa

[8191] T. O’Conor Sloane. Experiments in pneumatics with a steam vacuum. Scientific American, 55(14):214, October 2, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v55/
n14/pdf/scientificamerican10021886-214.pdf.

Anonymous:1886:HET

[8192] Anonymous. The Hamburg electric tramway. Scientific American, 55(14):215, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n14/pdf/scientificamerican10021886-
215.pdf.

Anonymous:1886:PNv

[8193] Anonymous. Photographic notes. Scientific American, 55(14):216, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n14/pdf/scientificamerican10021886-216a.pdf.

Anonymous:1886:DRPe

[8194] Anonymous. Decision relating to patents. Scientific American, 55(14):216, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n14/pdf/scientificamerican10021886-
216b.pdf.

Anonymous:1886:TMD

[8195] Anonymous. Trade mark decision. Scientific American, 55(14):216, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n14/pdf/scientificamerican10021886-216c.pdf.

Anonymous:1886:TSB

[8196] Anonymous. A tropical stag beetle. Scientific American, 55(14):217, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n14/pdf/scientificamerican10021886-217.pdf.

Anonymous:1886:EIbf

[8197] Anonymous. Engineering inventions. Scientific American, 55(14):218, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n14/pdf/scientificamerican10021886-218.pdf.

Anonymous:1886:AIx

[8198] Anonymous. Agricultural invention. Scientific American, 55(14):218, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n14/pdf/scientificamerican10021886-218a.pdf.

Anonymous:1886:MIbh

[8199] Anonymous. Miscellaneous inventions. Scientific American, 55(14):218, October 2, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n14/pdf/scientificamerican10021886-218b.pdf.

Anonymous:1886:Cbf

[8200] Anonymous. Correspondence. Scientific American, 55(15):229, October 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n15/pdf/scientificamerican10091886-229.pdf.

Anonymous:1886:BPbg

[8201] Anonymous. Business and personal. Scientific American, 55(15):235, October 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n15/pdf/scientificamerican10091886-235d.pdf.

Anonymous:1886:BNBx

[8202] Anonymous. Books: New books and publications. Scientific American, 55(15):235, October 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n15/pdf/scientificamerican10091886-
235e.pdf.

Anonymous:1886:NQbe

[8203] Anonymous. Notes and queries. Scientific American, 55(15):235, October 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n15/pdf/scientificamerican10091886-235f.pdf.

Anonymous:1886:IIbf

[8204] Anonymous. Index of inventions. Scientific American, 55(15):235–236, October 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n15/pdf/scientificamerican10091886-235g.pdf.

Anonymous:1886:PNPbe

[8205] Anonymous. Patent notes: Patents. Scientific American, 55(15):238, October 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n15/pdf/scientificamerican10091886-238.pdf.

Anonymous:1886:ESC

[8206] Anonymous. The electric subway conduit in New York. Scientific American, 55(15):223, October 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n15/pdf/scientificamerican10091886-
223.pdf.

Anonymous:1886:RN

[8207] Anonymous. “The remedies of nature”. Scientific American, 55 (15):224–225, October 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n15/pdf/scientificamerican10091886-
224.pdf.

Anonymous:1886:IRS

[8208] Anonymous. Improved railway signal. Scientific American, 55(15):226, October 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n15/pdf/scientificamerican10091886-226.pdf.

Anonymous:1886:IEG

[8209] Anonymous. Improved emery grinder. Scientific American, 55(15):226, October 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n15/pdf/scientificamerican10091886-226a.pdf.

Anonymous:1886:ICe

[8210] Anonymous. An improved churn. Scientific American, 55(15):227, October 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n15/pdf/scientificamerican10091886-227.pdf.

Anonymous:1886:ICf

[8211] Anonymous. Improved carillon. Scientific American, 55(15):227, October 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n15/pdf/scientificamerican10091886-227a.pdf.

Anonymous:1886:BCa

[8212] Anonymous. Boiler cleaner. Scientific American, 55(15):227, October 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n15/pdf/scientificamerican10091886-227b.pdf.

Sloane:1886:EPSb

[8213] T. O’conor Sloane. Experiments in pneumatics with a steam vacuum. Scientific American, 55(15):228, October 9, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v55/
n15/pdf/scientificamerican10091886-228.pdf.

Anonymous:1886:PNw

[8214] Anonymous. Photographic notes. Scientific American, 55(15):229, October 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n15/pdf/scientificamerican10091886-229a.pdf.

Hopkins:1886:ST

[8215] M. Hopkins. A scientific top. Scientific American, 55(15):230–231, October 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n15/pdf/scientificamerican10091886-230.pdf.

Anonymous:1886:TPEa

[8216] Anonymous. The transmission of power by electricity between Creil and Paris. Scientific American, 55(15):232, October 9, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v55/
n15/pdf/scientificamerican10091886-232.pdf.

Anonymous:1886:BBa

[8217] Anonymous. A ball of birds. Scientific American, 55(15):233–234, October 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n15/pdf/scientificamerican10091886-233.pdf.

Anonymous:1886:EIbg

[8218] Anonymous. Engineering inventions. Scientific American, 55(15):235, October 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n15/pdf/scientificamerican10091886-235.pdf.

Anonymous:1886:MIbi

[8219] Anonymous. Mechanical invention. Scientific American, 55(15):235, October 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n15/pdf/scientificamerican10091886-235a.pdf.

Anonymous:1886:AIy

[8220] Anonymous. Agricultural inventions. Scientific American, 55(15):235, October 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n15/pdf/scientificamerican10091886-235b.pdf.

Anonymous:1886:MIbj

[8221] Anonymous. Miscellaneous inventions. Scientific American, 55(15):235, October 9, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n15/pdf/scientificamerican10091886-235c.pdf.

Anonymous:1886:BPbh

[8222] Anonymous. Business and personal. Scientific American, 55 (16):251, October 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n16/pdf/scientificamerican10161886-
251c.pdf.

Anonymous:1886:BNBy

[8223] Anonymous. Books: New books and publications. Scientific American, 55(16):251, October 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n16/pdf/scientificamerican10161886-
251d.pdf.

Anonymous:1886:NQbf

[8224] Anonymous. Notes and queries. Scientific American, 55(16):251, October 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n16/pdf/scientificamerican10161886-251e.pdf.

Anonymous:1886:IIbg

[8225] Anonymous. Index of inventions. Scientific American, 55(16): 251–252, October 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n16/pdf/scientificamerican10161886-
251g.pdf.

Anonymous:1886:PNPbf

[8226] Anonymous. Patent notes: Patents. Scientific American, 55 (16):254, October 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n16/pdf/scientificamerican10161886-
254.pdf.

Hopkins:1886:EST

[8227] M. Hopkins. Experiments with the scientific top. Scientific American, 55(16):239, October 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n16/pdf/scientificamerican10161886-
239.pdf.

Anonymous:1886:FMM

[8228] Anonymous. The “Fere” monster mortar. Scientific American, 55(16):240, October 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n16/pdf/scientificamerican10161886-
240.pdf.

Anonymous:1886:PNx

[8229] Anonymous. Photographic notes. Scientific American, 55(16):241, October 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n16/pdf/scientificamerican10161886-241.pdf.

Anonymous:1886:IOL

[8230] Anonymous. Instrument for obtaining lengths and bevels of rafters. Scientific American, 55(16):242, October 16, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v55/
n16/pdf/scientificamerican10161886-242.pdf.

Anonymous:1886:OC

[8231] Anonymous. Ore concentrator. Scientific American, 55(16):242, October 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n16/pdf/scientificamerican10161886-242a.pdf.

Anonymous:1886:RTS

[8232] Anonymous. Railway track system. Scientific American, 55(16): 242, October 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n16/pdf/scientificamerican10161886-
242b.pdf.

Anonymous:1886:RTa

[8233] Anonymous. Riveting tool. Scientific American, 55(16):242, October 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n16/pdf/scientificamerican10161886-242c.pdf.

Anonymous:1886:ITd

[8234] Anonymous. Improved toboggan. Scientific American, 55(16):243, October 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n16/pdf/scientificamerican10161886-243.pdf.

Anonymous:1886:GC

[8235] Anonymous. Game carrier. Scientific American, 55(16):243, October 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n16/pdf/scientificamerican10161886-243a.pdf.

Anonymous:1886:EPS

[8236] Anonymous. Engines of the paddle steamer Ozone. Scientific American, 55(16):243–244, October 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n16/pdf/scientificamerican10161886-
243b.pdf.

Anonymous:1886:SML

[8237] Anonymous. A stop motion for looms. Scientific American, 55 (16):244–245, October 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n16/pdf/scientificamerican10161886-
244.pdf.

Anonymous:1886:CTW

[8238] Anonymous. A cheap type writer. Scientific American, 55(16): 246–247, October 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n16/pdf/scientificamerican10161886-
246.pdf.

Sloane:1886:EPSc

[8239] T. O’conor Sloane. Experiments in pneumatics with a steam vacuum. Scientific American, 55(16):247–248, October 16, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v55/
n16/pdf/scientificamerican10161886-247.pdf.

Anonymous:1886:SSc

[8240] Anonymous. The “Saddle Sulky”. Scientific American, 55(16):248, October 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n16/pdf/scientificamerican10161886-248.pdf.

Anonymous:1886:IVSa

[8241] Anonymous. Improved vibrating stethoscope. Scientific American, 55(16):248, October 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n16/pdf/scientificamerican10161886-
248a.pdf.

Anonymous:1886:BTCb

[8242] Anonymous. A balloon trip from Cherbourg to London. Scientific American, 55(16):249–250, October 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n16/pdf/scientificamerican10161886-
249.pdf.

Anonymous:1886:NHNb

[8243] Anonymous. Natural history notes. Scientific American, 55(16): 250, October 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n16/pdf/scientificamerican10161886-
250.pdf.

Anonymous:1886:EIbh

[8244] Anonymous. Engineering inventions. Scientific American, 55(16): 251, October 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n16/pdf/scientificamerican10161886-
251.pdf.

Anonymous:1886:AIz

[8245] Anonymous. Agricultural invention. Scientific American, 55(16): 251, October 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n16/pdf/scientificamerican10161886-
251a.pdf.

Anonymous:1886:MIbk

[8246] Anonymous. Miscellaneous inventions. Scientific American, 55 (16):251, October 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n16/pdf/scientificamerican10161886-
251b.pdf.

Anonymous:1886:Io

[8247] Anonymous. To inventors. Scientific American, 55(16):251, October 16, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n16/pdf/scientificamerican10161886-251f.pdf.

Anonymous:1886:Cbg

[8248] Anonymous. Correspondence. Scientific American, 55(17):261, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n17/pdf/scientificamerican10231886-261.pdf.

Anonymous:1886:BPbi

[8249] Anonymous. Business and personal. Scientific American, 55 (17):267, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n17/pdf/scientificamerican10231886-
267.pdf.

Anonymous:1886:NQbg

[8250] Anonymous. Notes and queries. Scientific American, 55(17):267, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n17/pdf/scientificamerican10231886-267a.pdf.

Anonymous:1886:IIbh

[8251] Anonymous. Index of inventions. Scientific American, 55(17): 267–268, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n17/pdf/scientificamerican10231886-
267c.pdf.

Anonymous:1886:PNPbg

[8252] Anonymous. Patent notes: Patents. Scientific American, 55 (17):270, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n17/pdf/scientificamerican10231886-
270.pdf.

Anonymous:1886:HDS

[8253] Anonymous. Hilgard’s design for a steel arch bridge at Minneapolis. Scientific American, 55(17):255–256, October 23, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v55/
n17/pdf/scientificamerican10231886-255.pdf.

Anonymous:1886:T

[8254] Anonymous. Twine. Scientific American, 55(17):256, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n17/pdf/scientificamerican10231886-256.pdf.

Anonymous:1886:PNy

[8255] Anonymous. Photographic notes. Scientific American, 55(17):257, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n17/pdf/scientificamerican10231886-257.pdf.

Anonymous:1886:SLb

[8256] Anonymous. Seal lock. Scientific American, 55(17):258, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n17/pdf/scientificamerican10231886-258.pdf.

Anonymous:1886:MSS

[8257] Anonymous. Metallic spring for suspenders. Scientific American, 55(17):258, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n17/pdf/scientificamerican10231886-
258a.pdf.

Anonymous:1886:IPE

[8258] Anonymous. Improved pail ear. Scientific American, 55(17):258, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n17/pdf/scientificamerican10231886-258b.pdf.

Anonymous:1886:CCf

[8259] Anonymous. Car coupling. Scientific American, 55(17):258, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n17/pdf/scientificamerican10231886-258c.pdf.

Anonymous:1886:CCg

[8260] Anonymous. Car coupling. Scientific American, 55(17):258, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n17/pdf/scientificamerican10231886-258d.pdf.

Anonymous:1886:SCg

[8261] Anonymous. Soldering case. Scientific American, 55(17):258–259, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n17/pdf/scientificamerican10231886-258e.pdf.

Anonymous:1886:FI

[8262] Anonymous. Floods in India. Scientific American, 55(17):259, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n17/pdf/scientificamerican10231886-259.pdf.

Anonymous:1886:TFb

[8263] Anonymous. Tunneling by freezing. Scientific American, 55(17): 259–260, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n17/pdf/scientificamerican10231886-
259a.pdf.

Hopkins:1886:PLa

[8264] M. Hopkins. Polarized light. Scientific American, 55(17):262, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n17/pdf/scientificamerican10231886-262.pdf.

Gratacap:1886:Q

[8265] L. P. Gratacap. The Quadrumana. Scientific American, 55(17): 263–264, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n17/pdf/scientificamerican10231886-
263.pdf.

Anonymous:1886:WTV

[8266] Anonymous. The water tower at Victoria, Texas. Scientific American, 55(17):264, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n17/pdf/scientificamerican10231886-
264.pdf.

Anonymous:1886:DRPf

[8267] Anonymous. Decision relating to patents. Scientific American, 55(17):264, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n17/pdf/scientificamerican10231886-
264a.pdf.

Sloane:1886:TPM

[8268] T O’conor Sloane. Torricelli’s Principle and mercury fountain. Scientific American, 55(17):265, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n17/pdf/scientificamerican10231886-
265.pdf.

Anonymous:1886:EIbi

[8269] Anonymous. Engineering inventions. Scientific American, 55(17): 266, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n17/pdf/scientificamerican10231886-
266.pdf.

Anonymous:1886:MIbl

[8270] Anonymous. Mechanical invention. Scientific American, 55(17): 266, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n17/pdf/scientificamerican10231886-
266a.pdf.

Anonymous:1886:AIba

[8271] Anonymous. Agricultural inventions. Scientific American, 55 (17):266, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n17/pdf/scientificamerican10231886-
266b.pdf.

Anonymous:1886:MIbm

[8272] Anonymous. Miscellaneous inventions. Scientific American, 55 (17):266–267, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n17/pdf/scientificamerican10231886-
266c.pdf.

Anonymous:1886:Iq

[8273] Anonymous. To inventors. Scientific American, 55(17):267, October 23, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n17/pdf/scientificamerican10231886-267b.pdf.

Anonymous:1886:Cbh

[8274] Anonymous. Correspondence. Scientific American, 55(18):277–278, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n18/pdf/scientificamerican10301886-277.pdf.

Anonymous:1886:BPbj

[8275] Anonymous. Business and personal. Scientific American, 55(18): 282–283, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n18/pdf/scientificamerican10301886-
282d.pdf.

Anonymous:1886:NQbh

[8276] Anonymous. Notes and queries. Scientific American, 55(18):283, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n18/pdf/scientificamerican10301886-283.pdf.

Anonymous:1886:IIbi

[8277] Anonymous. Index of inventions. Scientific American, 55(18): 283–284, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n18/pdf/scientificamerican10301886-
283b.pdf.

Anonymous:1886:PNPbh

[8278] Anonymous. Patent notes: Patents. Scientific American, 55 (18):286, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n18/pdf/scientificamerican10301886-
286.pdf.

Anonymous:1886:HMB

[8279] Anonymous. H.M.S. Benbow. Scientific American, 55(18):271–272, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n18/pdf/scientificamerican10301886-271.pdf.

Anonymous:1886:SLF

[8280] Anonymous. The Statue of Liberty finishes. Scientific American, 55(18):272–273, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n18/pdf/scientificamerican10301886-
272.pdf.

Anonymous:1886:BCb

[8281] Anonymous. Barrel cover. Scientific American, 55(18):274, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n18/pdf/scientificamerican10301886-274.pdf.

Anonymous:1886:CCT

[8282] Anonymous. Construction cars for track laying. Scientific American, 55(18):274, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n18/pdf/scientificamerican10301886-
274a.pdf.

Anonymous:1886:CPb

[8283] Anonymous. Cotton planter. Scientific American, 55(18):274, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n18/pdf/scientificamerican10301886-274b.pdf.

Anonymous:1886:EB

[8284] Anonymous. Egg beater. Scientific American, 55(18):274, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n18/pdf/scientificamerican10301886-274c.pdf.

Anonymous:1886:WMb

[8285] Anonymous. Washing machine. Scientific American, 55(18):274, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n18/pdf/scientificamerican10301886-274d.pdf.

Anonymous:1886:ACV

[8286] Anonymous. Apparatus for carrying vessels over obstructions. Scientific American, 55(18):275, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n18/pdf/scientificamerican10301886-
275.pdf.

Anonymous:1886:SCV

[8287] Anonymous. Swinging centreboard for vessels. Scientific American, 55(18):275, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n18/pdf/scientificamerican10301886-
275a.pdf.

Anonymous:1886:SMI

[8288] Anonymous. A simple method of insulating underground and other wires. Scientific American, 55(18):275, October 30, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v55/
n18/pdf/scientificamerican10301886-275b.pdf.

Anonymous:1886:MSb

[8289] Anonymous. Mushrooms on shelves. Scientific American, 55(18): 275–276, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n18/pdf/scientificamerican10301886-
275c.pdf.

Hovey:1886:USW

[8290] E. O. Hovey. A unique system of water works. Scientific American, 55(18):276, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n18/pdf/scientificamerican10301886-
276.pdf.

Harding:1886:DPB

[8291] Ed Harding. Double parabolic bridge, submitted in competition, for crossing the Harlem River at 18lst Street, City of New York. Scientific American, 55(18):278, October 30, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v55/
n18/pdf/scientificamerican10301886-278.pdf.

Anonymous:1886:HWR

[8292] Anonymous. Heating water rapidly. Scientific American, 55(18): 279, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n18/pdf/scientificamerican10301886-
279.pdf.

Anonymous:1886:EPW

[8293] Anonymous. Egyptian petroleum wells. Scientific American, 55 (18):279, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n18/pdf/scientificamerican10301886-
279a.pdf.

Anonymous:1886:BEC

[8294] Anonymous. A boiler explosion at Charlotte, N.C. Scientific American, 55(18):279, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n18/pdf/scientificamerican10301886-
279b.pdf.

Anonymous:1886:AAP

[8295] Anonymous. The Aquatic Arena at Paris. Scientific American, 55(18):280, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n18/pdf/scientificamerican10301886-
280.pdf.

Sloane:1886:CVb

[8296] T. O’conor Sloane. Capillary vacua. Scientific American, 55 (18):281, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n18/pdf/scientificamerican10301886-
281.pdf.

Anonymous:1886:EIbj

[8297] Anonymous. Engineering inventions. Scientific American, 55(18): 282, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n18/pdf/scientificamerican10301886-
282.pdf.

Anonymous:1886:AIbb

[8298] Anonymous. Agricultural inventions. Scientific American, 55 (18):282, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n18/pdf/scientificamerican10301886-
282a.pdf.

Anonymous:1886:MIbn

[8299] Anonymous. Miscellaneous inventions. Scientific American, 55 (18):282, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n18/pdf/scientificamerican10301886-
282b.pdf.

Anonymous:1886:Si

[8300] Anonymous. Special. Scientific American, 55(18):282, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n18/pdf/scientificamerican10301886-282c.pdf.

Anonymous:1886:Ir

[8301] Anonymous. To inventors. Scientific American, 55(18):283, October 30, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n18/pdf/scientificamerican10301886-283a.pdf.

Anonymous:1886:BNBz

[8302] Anonymous. Books: New books and publications. Scientific American, 55(19):298, November 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n19/pdf/scientificamerican11061886-
298d.pdf.

Anonymous:1886:BPbk

[8303] Anonymous. Business and personal. Scientific American, 55 (19):298, November 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n19/pdf/scientificamerican11061886-
298e.pdf.

Anonymous:1886:NQbi

[8304] Anonymous. Notes and queries. Scientific American, 55(19): 298–299, November 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n19/pdf/scientificamerican11061886-
298f.pdf.

Anonymous:1886:IIbj

[8305] Anonymous. Index of inventions. Scientific American, 55(19): 299–300, November 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n19/pdf/scientificamerican11061886-
299.pdf.

Anonymous:1886:PNPbi

[8306] Anonymous. Patent notes: Patents. Scientific American, 55 (19):302, November 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n19/pdf/scientificamerican11061886-
302.pdf.

Anonymous:1886:ISL

[8307] Anonymous. The inauguration of the Statue of Liberty. Scientific American, 55(19):288, November 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n19/pdf/scientificamerican11061886-
288.pdf.

Anonymous:1886:SDa

[8308] Anonymous. The science of drinking. Scientific American, 55 (19):288–289, November 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n19/pdf/scientificamerican11061886-
288a.pdf.

Anonymous:1886:DRPg

[8309] Anonymous. Decision relating to patents. Scientific American, 55(19):289, November 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n19/pdf/scientificamerican11061886-
289.pdf.

Anonymous:1886:DG

[8310] Anonymous. Drawbridge gate. Scientific American, 55(19):290, November 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n19/pdf/scientificamerican11061886-290.pdf.

Anonymous:1886:STd

[8311] Anonymous. Scraping tool. Scientific American, 55(19):290, November 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n19/pdf/scientificamerican11061886-290a.pdf.

Anonymous:1886:STCb

[8312] Anonymous. Steam traveling crane. Scientific American, 55(19): 290, November 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n19/pdf/scientificamerican11061886-
290b.pdf.

Anonymous:1886:FSB

[8313] Anonymous. Furnace for steam boilers. Scientific American, 55 (19):290–291, November 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n19/pdf/scientificamerican11061886-
290c.pdf.

Anonymous:1886:NFC

[8314] Anonymous. New French cruiser Tonnant. Scientific American, 55(19):291, November 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n19/pdf/scientificamerican11061886-
291.pdf.

Anonymous:1886:ISHa

[8315] Anonymous. Improved steam heating boiler. Scientific American, 55(19):291, November 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n19/pdf/scientificamerican11061886-
291a.pdf.

Anonymous:1886:ALH

[8316] Anonymous. On the airing and lighting of houses. Scientific American, 55(19):292–294, November 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n19/pdf/scientificamerican11061886-
292.pdf.

Anonymous:1886:DD

[8317] Anonymous. Double dredger. Scientific American, 55(19):294, November 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n19/pdf/scientificamerican11061886-294.pdf.

Anonymous:1886:TLM

[8318] Anonymous. The tarsiers and lorises of the Malay Archipelago. Scientific American, 55(19):295–296, November 6, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v55/
n19/pdf/scientificamerican11061886-295.pdf.

Anonymous:1886:Cbi

[8319] Anonymous. The cyclorama. Scientific American, 55(19):296, November 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n19/pdf/scientificamerican11061886-296.pdf.

Sloane:1886:CH

[8320] T. O’conor Sloane. Capillarity and hydrostatics. Scientific American, 55(19):297, November 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n19/pdf/scientificamerican11061886-
297.pdf.

Anonymous:1886:EIbk

[8321] Anonymous. Engineering inventions. Scientific American, 55(19): 298, November 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n19/pdf/scientificamerican11061886-
298.pdf.

Anonymous:1886:MIbo

[8322] Anonymous. Mechanical invention. Scientific American, 55(19): 298, November 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n19/pdf/scientificamerican11061886-
298a.pdf.

Anonymous:1886:AIbc

[8323] Anonymous. Agricultural inventions. Scientific American, 55 (19):298, November 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n19/pdf/scientificamerican11061886-
298b.pdf.

Anonymous:1886:MIbp

[8324] Anonymous. Miscellaneous inventions. Scientific American, 55 (19):298, November 6, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n19/pdf/scientificamerican11061886-
298c.pdf.

Anonymous:1886:BPbl

[8325] Anonymous. Business and personal. Scientific American, 55(2):26, July 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n2/pdf/scientificamerican07101886-26b.pdf.

Anonymous:1886:NQbj

[8326] Anonymous. Notes and queries. Scientific American, 55(2):26–27, July 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n2/pdf/scientificamerican07101886-26c.pdf.

Anonymous:1886:IIbk

[8327] Anonymous. Index of inventions. Scientific American, 55(2):27–28, July 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n2/pdf/scientificamerican07101886-27.pdf.

Anonymous:1886:PNPbj

[8328] Anonymous. Patent notes: Patents. Scientific American, 55(2):30, July 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n2/pdf/scientificamerican07101886-30.pdf.

Anonymous:1886:FSL

[8329] Anonymous. The fastest steam launch. Scientific American, 55(2):16, July 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n2/pdf/scientificamerican07101886-16.pdf.

Anonymous:1886:PDG

[8330] Anonymous. The pneumatic dynamite gun. Scientific American, 55(2):16, July 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n2/pdf/scientificamerican07101886-
16a.pdf.

Anonymous:1886:PNz

[8331] Anonymous. Photographic notes. Scientific American, 55(2):17–18, July 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n2/pdf/scientificamerican07101886-17.pdf.

Anonymous:1886:FPc

[8332] Anonymous. Fruit pulper. Scientific American, 55(2):18, July 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n2/pdf/scientificamerican07101886-18.pdf.

Anonymous:1886:OMR

[8333] Anonymous. Operating mechanism for railway signals. Scientific American, 55(2):18, July 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n2/pdf/scientificamerican07101886-
18a.pdf.

Anonymous:1886:AM

[8334] Anonymous. Adding machine. Scientific American, 55(2):18, July 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n2/pdf/scientificamerican07101886-18b.pdf.

Anonymous:1886:BS

[8335] Anonymous. Boiler sweeper. Scientific American, 55(2):18, July 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n2/pdf/scientificamerican07101886-18c.pdf.

Anonymous:1886:IPAb

[8336] Anonymous. Improved pumping apparatus. Scientific American, 55(2):18–19, July 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n2/pdf/scientificamerican07101886-
18d.pdf.

Hopkins:1886:PLb

[8337] Geo. M. Hopkins. Polarized light. Scientific American, 55(2):19, July 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n2/pdf/scientificamerican07101886-19.pdf.

Anonymous:1886:AGW

[8338] Anonymous. Automatic grain weighing and registering machine. Scientific American, 55(2):19–20, July 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n2/pdf/scientificamerican07101886-
19a.pdf.

Anonymous:1886:MIEa

[8339] Anonymous. The Minneapolis Industrial Exposition. Scientific American, 55(2):20, July 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n2/pdf/scientificamerican07101886-
20.pdf.

Anonymous:1886:MER

[8340] Anonymous. The Meigs elevated railway. Scientific American, 55(2):21, July 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n2/pdf/scientificamerican07101886-21.pdf.

Anonymous:1886:NYH

[8341] Anonymous. New York Harbor. Scientific American, 55(2):22–23, July 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n2/pdf/scientificamerican07101886-22.pdf.

Anonymous:1886:HTW

[8342] Anonymous. The Hall type writer. Scientific American, 55(2):24–25, July 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n2/pdf/scientificamerican07101886-24.pdf.

Anonymous:1886:MSE

[8343] Anonymous. Mennig’s steam engine. Scientific American, 55(2):25, July 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n2/pdf/scientificamerican07101886-25.pdf.

Anonymous:1886:EIbl

[8344] Anonymous. Engineering inventions. Scientific American, 55(2):26, July 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n2/pdf/scientificamerican07101886-26.pdf.

Anonymous:1886:MIbq

[8345] Anonymous. Miscellaneous inventions. Scientific American, 55(2):26, July 10, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n2/pdf/scientificamerican07101886-26a.pdf.

Anonymous:1886:Cbj

[8346] Anonymous. Correspondence. Scientific American, 55(20):309, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n20/pdf/scientificamerican11131886-309.pdf.

Anonymous:1886:BNBba

[8347] Anonymous. Books: New books and publications. Scientific American, 55(20):314, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n20/pdf/scientificamerican11131886-
314d.pdf.

Anonymous:1886:BPbm

[8348] Anonymous. Business and personal. Scientific American, 55 (20):315, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n20/pdf/scientificamerican11131886-
315.pdf.

Anonymous:1886:NQbk (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n20/pdf/scientificamerican11131886-
315a.pdf.

Anonymous:1886:IIbl

[8350] Anonymous. Index of inventions. Scientific American, 55(20): 315–317, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n20/pdf/scientificamerican11131886-
315b.pdf.

Anonymous:1886:PNPbk

[8351] Anonymous. Patent notes: Patents. Scientific American, 55 (20):318, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n20/pdf/scientificamerican11131886-
318.pdf.

Anonymous:1886:RHT

[8352] Anonymous. Royal E. House’s telephone of 1868. Scientific American, 55(20):303, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n20/pdf/scientificamerican11131886-
303.pdf.

Anonymous:1886:BTP

[8353] Anonymous. The Bell Telephone Patent probably broken. Scientific American, 55(20):304, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n20/pdf/scientificamerican11131886-
304.pdf.

Anonymous:1886:RBE

[8354] Anonymous. The recent boiler explosion at Charlotte, N.C. Scientific American, 55(20):304, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n20/pdf/scientificamerican11131886-
304a.pdf.

Anonymous:1886:RRAb

[8355] Anonymous. A remarkable railway accident. Scientific American, 55 (20):304–305, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n20/pdf/scientificamerican11131886-
304b.pdf.

Anonymous:1886:NNa

[8356] Anonymous. The new Navy. Scientific American, 55(20):305, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n20/pdf/scientificamerican11131886-305.pdf.

Hale:1886:DCW

[8357] Wm. H. Hale. The dynamo Colossus at work. Scientific American, 55(20):305, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n20/pdf/scientificamerican11131886-
305a.pdf.

Anonymous:1886:FSc

[8358] Anonymous. Frogless switch. Scientific American, 55(20):306, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n20/pdf/scientificamerican11131886-306.pdf.

Anonymous:1886:CCh

[8359] Anonymous. Cylinder cock. Scientific American, 55(20):306, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n20/pdf/scientificamerican11131886-306a.pdf.

Anonymous:1886:OS

[8360] Anonymous. Ore separator. Scientific American, 55(20):306, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n20/pdf/scientificamerican11131886-306b.pdf.

Anonymous:1886:RF

[8361] Anonymous. Railroad fence. Scientific American, 55(20):306, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n20/pdf/scientificamerican11131886-306c.pdf.

Anonymous:1886:CAMb

[8362] Anonymous. Cutting apparatus for mowers and reapers. Scientific American, 55(20):306, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n20/pdf/scientificamerican11131886-
306d.pdf.

Anonymous:1886:ISf

[8363] Anonymous. Improved sleigh. Scientific American, 55(20):307, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n20/pdf/scientificamerican11131886-307.pdf.

Anonymous:1886:FCS

[8364] Anonymous. Flour chest and sifter. Scientific American, 55 (20):307, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n20/pdf/scientificamerican11131886-
307a.pdf.

Anonymous:1886:IPf

[8365] Anonymous. Improved propeller. Scientific American, 55(20): 307, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n20/pdf/scientificamerican11131886-
307b.pdf.

Anonymous:1886:RTb

[8366] Anonymous. Revolving target. Scientific American, 55(20):307, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n20/pdf/scientificamerican11131886-307c.pdf.

Anonymous:1886:EC

[8367] Anonymous. Electric clock. Scientific American, 55(20):307, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n20/pdf/scientificamerican11131886-307d.pdf.

Anonymous:1886:MP

[8368] Anonymous. Measuring pump. Scientific American, 55(20):307– 308, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n20/pdf/scientificamerican11131886-
307e.pdf.

J:1886:DRP

[8369] Blodgett J. Decisions relating to patents. Scientific American, 55 (20):309, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n20/pdf/scientificamerican11131886-
309a.pdf.

Anonymous:1886:GK

[8370] Anonymous. A gigantic kite. Scientific American, 55(20):310, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n20/pdf/scientificamerican11131886-310.pdf.

Anonymous:1886:NTK

[8371] Anonymous. The new tunnel, Königstrasse, Berlin. Scientific American, 55(20):310–311, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n20/pdf/scientificamerican11131886-
310a.pdf.

Sloane:1886:EEF

[8372] T. O’Conor Sloane. Experiments in equilibrium of fluids. Scientific American, 55(20):311–312, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v55/n20/pdf/scientificamerican11131886-
311.pdf.

Anonymous:1886:RBa

[8373] Anonymous. Rock borers. Scientific American, 55(20):312, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n20/pdf/scientificamerican11131886-312.pdf.

Anonymous:1886:SMa

[8374] Anonymous. A sumptuous mantel. Scientific American, 55(20): 313, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n20/pdf/scientificamerican11131886-
313.pdf.

Anonymous:1886:NHNc

[8375] Anonymous. Natural history notes. Scientific American, 55(20): 313, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n20/pdf/scientificamerican11131886-
313a.pdf.

Anonymous:1886:EIbm

[8376] Anonymous. Engineering inventions. Scientific American, 55(20): 314, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n20/pdf/scientificamerican11131886-
314.pdf.

Anonymous:1886:MIbr

[8377] Anonymous. Mechanical invention. Scientific American, 55(20): 314, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n20/pdf/scientificamerican11131886-
314a.pdf.

Anonymous:1886:AIbd

[8378] Anonymous. Agricultural invention. Scientific American, 55(20): 314, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n20/pdf/scientificamerican11131886-
314b.pdf.

Anonymous:1886:MIbs

[8379] Anonymous. Miscellaneous inventions. Scientific American, 55 (20):314, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n20/pdf/scientificamerican11131886-
314c.pdf.

Anonymous:1886:Sj

[8380] Anonymous. Special. Scientific American, 55(20):314, November 13, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n20/pdf/scientificamerican11131886-314e.pdf.

Anonymous:1886:BNBbb

[8381] Anonymous. Books: New books and publications. Scientific American, 55(21):331, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n21/pdf/scientificamerican11201886-
331c.pdf.

Anonymous:1886:BPbn

[8382] Anonymous. Business and personal. Scientific American, 55 (21):331, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n21/pdf/scientificamerican11201886-
331d.pdf.

Anonymous:1886:NQbl

[8383] Anonymous. Notes and queries. Scientific American, 55(21): 331, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n21/pdf/scientificamerican11201886-
331e.pdf.

Anonymous:1886:IIbm

[8384] Anonymous. Index of inventions. Scientific American, 55(21): 331–332, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n21/pdf/scientificamerican11201886-
331g.pdf.

Anonymous:1886:PNPbl

[8385] Anonymous. Patent notes: Patents. Scientific American, 55 (21):334, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n21/pdf/scientificamerican11201886-
334.pdf.

Anonymous:1886:GSC

[8386] Anonymous. The Government suit for the cancellation of the Bell Telephone Patents. Scientific American, 55(21):320, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n21/pdf/scientificamerican11201886-320.pdf.

Anonymous:1886:SLNa

[8387] Anonymous. The Statue of Liberty, New York. Scientific American, 55(21):320, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n21/pdf/scientificamerican11201886-
320a.pdf.

Anonymous:1886:EUW

[8388] Anonymous. Effect of the use of white paper upon the eyes. Scientific American, 55(21):320, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n21/pdf/scientificamerican11201886-
320b.pdf.

Anonymous:1886:NNb

[8389] Anonymous. The new North. Scientific American, 55(21):321, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n21/pdf/scientificamerican11201886-321.pdf.

Anonymous:1886:ARS

[8390] Anonymous. The Abendroth & Root safety boiler. Scientific American, 55(21):322, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n21/pdf/scientificamerican11201886-
322.pdf.

Anonymous:1886:CCi

[8391] Anonymous. Car coupling. Scientific American, 55(21):323, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n21/pdf/scientificamerican11201886-323.pdf.

Anonymous:1886:CBk

[8392] Anonymous. Car brake. Scientific American, 55(21):323, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n21/pdf/scientificamerican11201886-323a.pdf.

Anonymous:1886:ACA

[8393] Anonymous. Apparatus for closing apertures in the hulls of vessels. Scientific American, 55(21):323, November 20, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v55/
n21/pdf/scientificamerican11201886-323b.pdf.

Anonymous:1886:SGc

[8394] Anonymous. Steam generator. Scientific American, 55(21):323, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n21/pdf/scientificamerican11201886-323c.pdf.

Anonymous:1886:CVa

[8395] Anonymous. Centerboard for vessels. Scientific American, 55 (21):324, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n21/pdf/scientificamerican11201886-
324.pdf.

Anonymous:1886:MCM

[8396] Anonymous. A medicine chest with many conveniences. Scientific American, 55(21):324, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n21/pdf/scientificamerican11201886-
324a.pdf.

Anonymous:1886:US

[8397] Anonymous. Umbrella support. Scientific American, 55(21): 324, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n21/pdf/scientificamerican11201886-
324b.pdf.

Anonymous:1886:FC

[8398] Anonymous. Fly catcher. Scientific American, 55(21):324–325, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n21/pdf/scientificamerican11201886-324c.pdf.

Anonymous:1886:HWG

[8399] Anonymous. Handling the water for a great city. Scientific American, 55(21):326, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n21/pdf/scientificamerican11201886-
326.pdf.

Anonymous:1886:EES

[8400] Anonymous. Effect of the earthquake on the South Carolina Railway. Scientific American, 55(21):327, November 20, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v55/
n21/pdf/scientificamerican11201886-327.pdf.

Anonymous:1886:AAL

[8401] Anonymous. Automatic attachment for locomotives. Scientific American, 55(21):327–328, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n21/pdf/scientificamerican11201886-
327a.pdf.

Anonymous:1886:GAP

[8402] Anonymous. Gauge attachment for printing presses. Scientific American, 55(21):328, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n21/pdf/scientificamerican11201886-
328.pdf.

Sloane:1886:C

[8403] T. O’Conor Sloane. The cycloid. Scientific American, 55(21): 328, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n21/pdf/scientificamerican11201886-
328a.pdf.

Anonymous:1886:IJM

[8404] Anonymous. Improved Jacquard machine. Scientific American, 55 (21):328, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n21/pdf/scientificamerican11201886-
328b.pdf.

Anonymous:1886:FMA

[8405] Anonymous. Fishes of the Malay Archipelago. Scientific American, 55(21):329–330, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n21/pdf/scientificamerican11201886-
329.pdf.

Anonymous:1886:EIbn (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n21/pdf/scientificamerican11201886-
331.pdf.

Anonymous:1886:AIbe

[8407] Anonymous. Agricultural inventions. Scientific American, 55 (21):331, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n21/pdf/scientificamerican11201886-
331a.pdf.

Anonymous:1886:MIbt

[8408] Anonymous. Miscellaneous inventions. Scientific American, 55 (21):331, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n21/pdf/scientificamerican11201886-
331b.pdf.

Anonymous:1886:Iu

[8409] Anonymous. To inventors. Scientific American, 55(21):331, November 20, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n21/pdf/scientificamerican11201886-331f.pdf.

Anonymous:1886:Cbl

[8410] Anonymous. Correspondence. Scientific American, 55(22):341– 342, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n22/pdf/scientificamerican11271886-
341.pdf.

Anonymous:1886:BNBbc

[8411] Anonymous. Books: New books and publications. Scientific American, 55(22):346, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n22/pdf/scientificamerican11271886-
346c.pdf.

Anonymous:1886:BPbo

[8412] Anonymous. Business and personal. Scientific American, 55 (22):347, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n22/pdf/scientificamerican11271886-
347.pdf.

Anonymous:1886:NQbm

[8413] Anonymous. Notes and queries. Scientific American, 55(22): 347, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n22/pdf/scientificamerican11271886-
347a.pdf.

Anonymous:1886:IIbn

[8414] Anonymous. Index of inventions. Scientific American, 55(22): 347–348, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n22/pdf/scientificamerican11271886-
347b.pdf.

Anonymous:1886:Pb

[8415] Anonymous. Patents. Scientific American, 55(22):350, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n22/pdf/scientificamerican11271886-350.pdf.

Hopkins:1886:G

[8416] George M. Hopkins. The gyroscope. Scientific American, 55 (22):335, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n22/pdf/scientificamerican11271886-
335.pdf.

Anonymous:1886:DC

[8417] Anonymous. The dynamite cruiser. Scientific American, 55(22): 336, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n22/pdf/scientificamerican11271886-
336.pdf.

Anonymous:1886:TVR

[8418] Anonymous. Torpedoes vs. rams. Scientific American, 55(22): 336, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n22/pdf/scientificamerican11271886-
336a.pdf.

Anonymous:1886:WH

[8419] Anonymous. Work and habits. Scientific American, 55(22):336– 337, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n22/pdf/scientificamerican11271886-
336b.pdf.

Anonymous:1886:IBSc

[8420] Anonymous. Improved balanced steam engine. Scientific American, 55(22):338, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n22/pdf/scientificamerican11271886-
338.pdf.

Anonymous:1886:IWH

[8421] Anonymous. Improved wheel hub. Scientific American, 55(22): 338, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n22/pdf/scientificamerican11271886-
338a.pdf.

Anonymous:1886:IVSb

[8422] Anonymous. Improved vehicle seat. Scientific American, 55(22): 338, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n22/pdf/scientificamerican11271886-
338b.pdf.

Anonymous:1886:DE

[8423] Anonymous. Draught equalizer. Scientific American, 55(22):338– 339, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n22/pdf/scientificamerican11271886-
338c.pdf.

Anonymous:1886:ADI

[8424] Anonymous. The African diamond industry. Scientific American, 55(22):339, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n22/pdf/scientificamerican11271886-
339.pdf.

Anonymous:1886:ICCa

[8425] Anonymous. Improved car coupling. Scientific American, 55(22): 339–340, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n22/pdf/scientificamerican11271886-
339a.pdf.

Anonymous:1886:NFP

[8426] Anonymous. A novel flower pot. Scientific American, 55(22): 340, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n22/pdf/scientificamerican11271886-
340.pdf.

Anonymous:1886:BMb

[8427] Anonymous. Brick making. Scientific American, 55(22):343, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n22/pdf/scientificamerican11271886-343.pdf.

Anonymous:1886:SKM

[8428] Anonymous. The South Kensington Museum. Scientific American, 55 (22):343–344, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n22/pdf/scientificamerican11271886-
343a.pdf.

Anonymous:1886:RTR

[8429] Anonymous. The repairing of tanks and reservoirs. Scientific American, 55(22):344–345, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n22/pdf/scientificamerican11271886-
344.pdf.

Sloane:1886:CS

[8430] T. O’Conor Sloane. The capillary siphon. Scientific American, 55 (22):345, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n22/pdf/scientificamerican11271886-
345.pdf.

Anonymous:1886:EIbo

[8431] Anonymous. Engineering inventions. Scientific American, 55(22): 346, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n22/pdf/scientificamerican11271886-
346.pdf.

Anonymous:1886:AIbf

[8432] Anonymous. Agricultural inventions. Scientific American, 55 (22):346, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n22/pdf/scientificamerican11271886-
346a.pdf.

Anonymous:1886:MIbu

[8433] Anonymous. Miscellaneous inventions. Scientific American, 55 (22):346, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n22/pdf/scientificamerican11271886-
346b.pdf.

Anonymous:1886:Sk

[8434] Anonymous. Special. Scientific American, 55(22):346, November 27, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n22/pdf/scientificamerican11271886-346d.pdf.

Anonymous:1886:Cbm

[8435] Anonymous. Correspondence. Scientific American, 55(23):357, December 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n23/pdf/scientificamerican12041886-357.pdf.

Anonymous:1886:BNBbd

[8436] Anonymous. Books: New books and publications. Scientific American, 55(23):362, December 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n23/pdf/scientificamerican12041886-
362d.pdf.

Anonymous:1886:BPbp 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n23/pdf/scientificamerican12041886-
362e.pdf.

Anonymous:1886:NQbn

[8438] Anonymous. Notes and queries. Scientific American, 55(23):363, December 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n23/pdf/scientificamerican12041886-363.pdf.

Anonymous:1886:IIbo

[8439] Anonymous. Index of inventions. Scientific American, 55(23): 363–364, December 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n23/pdf/scientificamerican12041886-
363a.pdf.

Anonymous:1886:PNPbm

[8440] Anonymous. Patent notes: Patents. Scientific American, 55 (23):366, December 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n23/pdf/scientificamerican12041886-
366.pdf.

Anonymous:1886:GGM

[8441] Anonymous. The great gas meter of the Consolidated Gas Company of New York. Scientific American, 55(23):351, December 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n23/pdf/scientificamerican12041886-351.pdf.

Anonymous:1886:NNc

[8442] Anonymous. The Navy’s needs. Scientific American, 55(23):352, December 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n23/pdf/scientificamerican12041886-352.pdf.

Anonymous:1886:IBL

[8443] Anonymous. Incidents in bird life at the park. Scientific American, 55(23):352–353, December 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n23/pdf/scientificamerican12041886-
352a.pdf.

Anonymous:1886:SMPa

[8444] Anonymous. The submarine monitor Peacemaker. Scientific American, 55(23):354, December 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n23/pdf/scientificamerican12041886-
354.pdf.

Anonymous:1886:SAV

[8445] Anonymous. Skeg attachment for vessels. Scientific American, 55(23):355, December 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n23/pdf/scientificamerican12041886-
355.pdf.

Anonymous:1886:IGC

[8446] Anonymous. Improved grain conveyer. Scientific American, 55 (23):355, December 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n23/pdf/scientificamerican12041886-
355a.pdf.

Anonymous:1886:SFL

[8447] Anonymous. Safety fastening for lace pins, etc. Scientific American, 55(23):355, December 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n23/pdf/scientificamerican12041886-
355b.pdf.

Anonymous:1886:SPM

[8448] Anonymous. Sandpapering and polishing machine. Scientific American, 55(23):355, December 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n23/pdf/scientificamerican12041886-
355c.pdf.

Sloane:1886:CSH

[8449] T. O’conor Sloane. The capillary siphon — Hero’s engine. Scientific American, 55(23):356, December 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n23/pdf/scientificamerican12041886-
356.pdf.

Anonymous:1886:DRPh

[8450] Anonymous. Decisions relating to patents. Scientific American, 55(23):357, December 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n23/pdf/scientificamerican12041886-
357a.pdf.

Hopkins:1886:DG

[8451] M. Hopkins. Deprez’s galvanometer. Scientific American, 55(23): 358–359, December 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n23/pdf/scientificamerican12041886-
358.pdf.

Anonymous:1886:NSW

[8452] Anonymous. The new Spanish war steamer Reina Regente. Scientific American, 55(23):359, December 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n23/pdf/scientificamerican12041886-
359.pdf.

Anonymous:1886:SAI

[8453] Anonymous. Some ancient inventions. Scientific American, 55 (23):360, December 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n23/pdf/scientificamerican12041886-
360.pdf.

Anonymous:1886:EUG

[8454] Anonymous. Expulsion of unburned gunpowder from cannons. Scientific American, 55(23):361, December 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n23/pdf/scientificamerican12041886-
361.pdf.

Anonymous:1886:EIbp

[8455] Anonymous. Engineering inventions. Scientific American, 55(23): 362, December 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n23/pdf/scientificamerican12041886-
362.pdf.

Anonymous:1886:MIbv

[8456] Anonymous. Mechanical invention. Scientific American, 55(23): 362, December 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n23/pdf/scientificamerican12041886-
362a.pdf.

Anonymous:1886:AIbg

[8457] Anonymous. Agricultural inventions. Scientific American, 55 (23):362, December 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n23/pdf/scientificamerican12041886-
362b.pdf.

Anonymous:1886:MIbw

[8458] Anonymous. Miscellaneous inventions. Scientific American, 55 (23):362, December 4, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n23/pdf/scientificamerican12041886-
362c.pdf.

Anonymous:1886:BPbq

[8459] Anonymous. Business and personal. Scientific American, 55(24): 378–379, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
378e.pdf.

Anonymous:1886:NQbo

[8460] Anonymous. Notes and queries. Scientific American, 55(24): 379, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
379.pdf.

Anonymous:1886:IIbp

[8461] Anonymous. Index of inventions. Scientific American, 55(24): 379–380, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
379b.pdf.

Anonymous:1886:PNPbn

[8462] Anonymous. Patent notes: Patents. Scientific American, 55 (24):382, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
382.pdf.

Anonymous:1886:TEY

[8463] Anonymous. Torpedoes eighty years ago. Scientific American, 55 (24):368–369, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
368.pdf.

Anonymous:1886:CCM

[8464] Anonymous. A convenient and certain mode for tempering steel. Scientific American, 55(24):369, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
369.pdf.

Seiss:1886:ST

[8465] C. Few Seiss. The snapping turtle. Scientific American, 55 (24):370, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
370.pdf.

Sloane:1886:WG

[8466] T. O’Conor Sloane. Weighing a gas. Scientific American, 55 (24):370, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
370a.pdf.

Anonymous:1886:IFCc

[8467] Anonymous. Improved fence clamp. Scientific American, 55 (24):371, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
371.pdf.

Anonymous:1886:WGSb

[8468] Anonymous. Water gauge for steam generators. Scientific American, 55(24):371, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
371a.pdf.

Anonymous:1886:RSC

[8469] Anonymous. Revolving shot case. Scientific American, 55(24): 371, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
371b.pdf.

Anonymous:1886:SST

[8470] Anonymous. Sheep shearing table. Scientific American, 55(24): 371, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
371c.pdf.

Anonymous:1886:FWR

[8471] Anonymous. Feed water regulator. Scientific American, 55(24): 371, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
371d.pdf.

Anonymous:1886:ISSa

[8472] Anonymous. Improved saw swage. Scientific American, 55(24): 372, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
372.pdf.

Anonymous:1886:CMF

[8473] Anonymous. Clock movement frame. Scientific American, 55 (24):372, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
372a.pdf.

Anonymous:1886:ADS

[8474] Anonymous. Automatic danger signal. Scientific American, 55 (24):372, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
372b.pdf.

Anonymous:1886:ILF

[8475] Anonymous. Improved lock for firearms. Scientific American, 55 (24):372, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
372c.pdf.

Anonymous:1886:MML

[8476] Anonymous. The Miller Manual Labor School of Albemarle, Va. Scientific American, 55(24):373, December 11, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v55/
n24/pdf/scientificamerican12111886-373.pdf.

Anonymous:1886:IWBa

[8477] Anonymous. Improved wheel boring and facing machine. Scientific American, 55(24):374, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
374.pdf.

Anonymous:1886:NSM

[8478] Anonymous. A new system of manufacturing metal tubes. Scientific American, 55(24):374, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
374a.pdf.

Anonymous:1886:CBN

[8479] Anonymous. Collision between the Netherland steamer Waesland and a large whale. Scientific American, 55(24):375–376, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n24/pdf/scientificamerican12111886-375.pdf.

Anonymous:1886:IAC

[8480] Anonymous. Improved automatic cable grip. Scientific American, 55(24):376, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
376.pdf.

Anonymous:1886:LTS

[8481] Anonymous. Lathe for turning spirals. Scientific American, 55 (24):376, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
376a.pdf.

Anonymous:1886:CSBb

[8482] Anonymous. The colossal statues of Bamian. Scientific American, 55(24):377, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
377.pdf.

Anonymous:1886:EIbq

[8483] Anonymous. Engineering inventions. Scientific American, 55(24): 378, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
378.pdf.

Anonymous:1886:MIbx

[8484] Anonymous. Mechanical invention. Scientific American, 55(24): 378, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
378a.pdf.

Anonymous:1886:AIbh

[8485] Anonymous. Agricultural inventions. Scientific American, 55 (24):378, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
378b.pdf.

Anonymous:1886:MIby

[8486] Anonymous. Miscellaneous inventions. Scientific American, 55 (24):378, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n24/pdf/scientificamerican12111886-
378c.pdf.

Anonymous:1886:Sn

[8487] Anonymous. Special. Scientific American, 55(24):378, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n24/pdf/scientificamerican12111886-378d.pdf.

Anonymous:1886:Iv

[8488] Anonymous. To inventors. Scientific American, 55(24):379, December 11, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n24/pdf/scientificamerican12111886-379a.pdf.

Anonymous:1886:Cbo

[8489] Anonymous. Correspondence. Scientific American, 55(25):389, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n25/pdf/scientificamerican12181886-389.pdf.

Anonymous:1886:BPbr

[8490] Anonymous. Business and personal. Scientific American, 55 (25):395, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n25/pdf/scientificamerican12181886-
395c.pdf.

Anonymous:1886:NQbp

[8491] Anonymous. Notes and queries. Scientific American, 55(25): 395, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n25/pdf/scientificamerican12181886-
395d.pdf.

Anonymous:1886:IIbq 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n25/pdf/scientificamerican12181886-
395f.pdf.

Anonymous:1886:Pd

[8493] Anonymous. Patents. Scientific American, 55(25):398, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n25/pdf/scientificamerican12181886-398.pdf.

Anonymous:1886:FTW

[8494] Anonymous. The first type writing machine. Scientific American, 55(25):383, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n25/pdf/scientificamerican12181886-
383.pdf.

Anonymous:1886:PMP

[8495] Anonymous. The President’s message and the Patent Office. Scientific American, 55(25):384, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n25/pdf/scientificamerican12181886-
384.pdf.

Anonymous:1886:PCU

[8496] Anonymous. Progress of Cornell University. Scientific American, 55(25):384, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n25/pdf/scientificamerican12181886-
384a.pdf.

Anonymous:1886:WBW

[8497] Anonymous. Will the Brooklyn wires be buried? Scientific American, 55(25):384, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n25/pdf/scientificamerican12181886-
384b.pdf.

Anonymous:1886:PNba (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n25/pdf/scientificamerican12181886-
385.pdf.

Anonymous:1886:DRPi

[8499] Anonymous. Decisions relating to patents. Scientific American, 55(25):385, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n25/pdf/scientificamerican12181886-
385a.pdf.

Anonymous:1886:WPK

[8500] Anonymous. Watch pendant key and safety attachment for watches. Scientific American, 55(25):386, December 18, 1886. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.nature.com/scientificamerican/journal/v55/n25/pdf/scientificamerican12181886-
386.pdf.

Anonymous:1886:SSd

[8501] Anonymous. Safety stirrup. Scientific American, 55(25):386, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n25/pdf/scientificamerican12181886-386a.pdf.

Anonymous:1886:HR

[8502] Anonymous. Hotel register. Scientific American, 55(25):386, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n25/pdf/scientificamerican12181886-386b.pdf.

Anonymous:1886:IFGb

[8503] Anonymous. Improved furnace grate. Scientific American, 55 (25):386, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n25/pdf/scientificamerican12181886-
386c.pdf.

Anonymous:1886:NFV

[8504] Anonymous. A novel form of vessel. Scientific American, 55 (25):386, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n25/pdf/scientificamerican12181886-
386d.pdf.

Anonymous:1886:V

[8505] Anonymous. Vehicle. Scientific American, 55(25):387, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n25/pdf/scientificamerican12181886-387.pdf.

Anonymous:1886:NTB

[8506] Anonymous. New torpedo boats. Scientific American, 55(25): 387–388, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n25/pdf/scientificamerican12181886-
387a.pdf.

Anonymous:1886:TR

[8507] Anonymous. Time register. Scientific American, 55(25):388, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n25/pdf/scientificamerican12181886-388.pdf.

Anonymous:1886:PFS

[8508] Anonymous. Petroleum as a fuel for steam tricycles. Scientific American, 55(25):390, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n25/pdf/scientificamerican12181886-
390.pdf.

Anonymous:1886:NCb

[8509] Anonymous. A novel clock. Scientific American, 55(25):390, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n25/pdf/scientificamerican12181886-390a.pdf.

Anonymous:1886:GEE

[8510] Anonymous. The Gymnotus Electricus, or electric eel. Scientific American, 55(25):391–392, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n25/pdf/scientificamerican12181886-
391.pdf.

Hopkins:1886:CT

[8511] Geo. M. Hopkins. Centrifugal top. Scientific American, 55(25): 392, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n25/pdf/scientificamerican12181886-
392.pdf.

Anonymous:1886:GBC

[8512] Anonymous. A grizzly bear and its cubs. — “the bath”. Scientific American, 55(25):393–394, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n25/pdf/scientificamerican12181886-
393.pdf.

Anonymous:1886:EIbr

[8513] Anonymous. Engineering invention. Scientific American, 55(25): 395, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n25/pdf/scientificamerican12181886-
395.pdf.

Anonymous:1886:AIbi

[8514] Anonymous. Agricultural inventions. Scientific American, 55 (25):395, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n25/pdf/scientificamerican12181886-
395a.pdf.

Anonymous:1886:MIbz

[8515] Anonymous. Miscellaneous inventions. Scientific American, 55 (25):395, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n25/pdf/scientificamerican12181886-
395b.pdf.

Anonymous:1886:Ix

[8516] Anonymous. To inventors. Scientific American, 55(25):395, December 18, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n25/pdf/scientificamerican12181886-395e.pdf.

Anonymous:1886:BNBbe

[8517] Anonymous. Books: New books and publications. Scientific American, 55(26):408, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n26/pdf/scientificamerican12251886-
408d.pdf.

Anonymous:1886:BPbs

[8518] Anonymous. Business and personal. Scientific American, 55 (26):408, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n26/pdf/scientificamerican12251886-
408f.pdf.

Anonymous:1886:NQbq

[8519] Anonymous. Notes and queries. Scientific American, 55(26): 408–409, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n26/pdf/scientificamerican12251886-
408g.pdf.

Anonymous:1886:IIbr

[8520] Anonymous. Index of inventions. Scientific American, 55(26): 409–410, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n26/pdf/scientificamerican12251886-
409a.pdf.

Anonymous:1886:Iy

[8521] Anonymous. Index. Scientific American, 55(26):411–412, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n26/pdf/scientificamerican12251886-411.pdf.

Anonymous:1886:PNPbo

[8522] Anonymous. Patent notes: Patents. Scientific American, 55 (26):414, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n26/pdf/scientificamerican12251886-
414.pdf.

Anonymous:1886:RSSc

[8523] Anonymous. Removal of snow from streets. Scientific American, 55(26):400, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n26/pdf/scientificamerican12251886-
400.pdf.

Anonymous:1886:GRD

[8524] Anonymous. A good rule — doubt must be resolved in favor of the inventor. Scientific American, 55(26):400–401, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n26/pdf/scientificamerican12251886-400a.pdf.

Anonymous:1886:ACe

[8525] Anonymous. The atmosphere of caves. Scientific American, 55 (26):401, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n26/pdf/scientificamerican12251886-
401.pdf.

Anonymous:1886:DUS

[8526] Anonymous. Direct use of steam as motive power of aerostats. Scientific American, 55(26):401, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n26/pdf/scientificamerican12251886-
401a.pdf.

Anonymous:1886:POS

[8527] Anonymous. Propulsion for ocean steamers by the steam jet. Scientific American, 55(26):401, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n26/pdf/scientificamerican12251886-
401b.pdf.

Anonymous:1886:IHa

[8528] Anonymous. Improved harrow. Scientific American, 55(26):402, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n26/pdf/scientificamerican12251886-402.pdf.

Anonymous:1886:ICCb

[8529] Anonymous. Improved car coupling. Scientific American, 55 (26):402, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n26/pdf/scientificamerican12251886-
402a.pdf.

Anonymous:1886:IHb

[8530] Anonymous. Improved harrow. Scientific American, 55(26):402, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n26/pdf/scientificamerican12251886-402b.pdf.

Anonymous:1886:GB

[8531] Anonymous. The glycerine barometer. Scientific American, 55(26): 403–404, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n26/pdf/scientificamerican12251886-
403.pdf.

Anonymous:1886:CHB

[8532] Anonymous. Combined hook and buckle. Scientific American, 55 (26):404, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n26/pdf/scientificamerican12251886-
404.pdf.

Anonymous:1886:CCj

[8533] Anonymous. Cotton chopper. Scientific American, 55(26):404, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n26/pdf/scientificamerican12251886-404a.pdf.

Anonymous:1886:TTa

[8534] Anonymous. Telephone transmitter. Scientific American, 55(26): 404, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n26/pdf/scientificamerican12251886-
404b.pdf.

Anonymous:1886:ISMb

[8535] Anonymous. Improved sugar machinery. Scientific American, 55 (26):404, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n26/pdf/scientificamerican12251886-
404c.pdf.

Sloane:1886:MB

[8536] T. O’conor Sloane. Mercury bubbles. Scientific American, 55(26): 404–405, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n26/pdf/scientificamerican12251886-
404d.pdf.

Anonymous:1886:FGW

[8537] Anonymous. Fall of a great water tower. Scientific American, 55 (26):405, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n26/pdf/scientificamerican12251886-
405.pdf.

Anonymous:1886:RTG

[8538] Anonymous. Recent types of the Gatling gun. Scientific American, 55(26):406, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n26/pdf/scientificamerican12251886-
406.pdf.

Ayres:1886:TAM

[8539] W. O. Ayres. Trichechus Americanus — manatee. Scientific American, 55(26):407, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n26/pdf/scientificamerican12251886-
407.pdf.

Anonymous:1886:FFb

[8540] Anonymous. The fishing frog. Scientific American, 55(26):407, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n26/pdf/scientificamerican12251886-407a.pdf.

Anonymous:1886:EIbs

[8541] Anonymous. Engineering inventions. Scientific American, 55(26): 408, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n26/pdf/scientificamerican12251886-
408.pdf.

Anonymous:1886:MIca

[8542] Anonymous. Mechanical invention. Scientific American, 55(26): 408, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n26/pdf/scientificamerican12251886-
408a.pdf.

Anonymous:1886:AIbj

[8543] Anonymous. Agricultural inventions. Scientific American, 55 (26):408, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n26/pdf/scientificamerican12251886-
408b.pdf.

Anonymous:1886:MIcb

[8544] Anonymous. Miscellaneous inventions. Scientific American, 55 (26):408, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n26/pdf/scientificamerican12251886-
408c.pdf.

Anonymous:1886:Sp

[8545] Anonymous. Special. Scientific American, 55(26):408, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n26/pdf/scientificamerican12251886-408e.pdf.

Anonymous:1886:Iz

[8546] Anonymous. To inventors. Scientific American, 55(26):409, December 25, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n26/pdf/scientificamerican12251886-409.pdf.

Anonymous:1886:Cbq

[8547] Anonymous. Correspondence. Scientific American, 55(3):37, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n3/pdf/scientificamerican07171886-37.pdf.

Anonymous:1886:BNBbf

[8548] Anonymous. Books: New books and publications. Scientific American, 55(3):42, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n3/pdf/scientificamerican07171886-
42c.pdf.

Anonymous:1886:BPbt

[8549] Anonymous. Business and personal. Scientific American, 55(3):42, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n3/pdf/scientificamerican07171886-42d.pdf.

Anonymous:1886:NQbr

[8550] Anonymous. Notes and queries. Scientific American, 55(3):42–43, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n3/pdf/scientificamerican07171886-42e.pdf.

Anonymous:1886:IIbs

[8551] Anonymous. Index of inventions. Scientific American, 55(3):43–44, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n3/pdf/scientificamerican07171886-43.pdf.

Anonymous:1886:PNPbp

[8552] Anonymous. Patent notes: Patents. Scientific American, 55(3):46, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n3/pdf/scientificamerican07171886-46.pdf.

Anonymous:1886:TPF

[8553] Anonymous. Transformation of physical forces. Scientific American, 55(3):31, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n3/pdf/scientificamerican07171886-
31.pdf.

Anonymous:1886:EHS

[8554] Anonymous. Estrade’s high speed locomotive. Scientific American, 55(3):31, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n3/pdf/scientificamerican07171886-
31a.pdf.

Anonymous:1886:RPP

[8555] Anonymous. Recently proposed patent legislation. Scientific American, 55(3):32, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n3/pdf/scientificamerican07171886-
32.pdf.

Anonymous:1886:PSCb

[8556] Anonymous. Progress of silk culture in this country. Scientific American, 55(3):32, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n3/pdf/scientificamerican07171886-
32a.pdf.

Anonymous:1886:PPE

[8557] Anonymous. Prof Pickering’s early experiments in telegraphing sound. Scientific American, 55(3):32, July 17, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v55/
n3/pdf/scientificamerican07171886-32b.pdf.

Anonymous:1886:NGP

[8558] Anonymous. The nation’s great problem, how to cool a cellar, and more. Scientific American, 55(3):32–34, July 17, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v55/
n3/pdf/scientificamerican07171886-32c.pdf.

Anonymous:1886:IHc

[8559] Anonymous. Improved hinge. Scientific American, 55(3):34, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n3/pdf/scientificamerican07171886-34.pdf.

Anonymous:1886:CCk

[8560] Anonymous. Car coupling. Scientific American, 55(3):34, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n3/pdf/scientificamerican07171886-34a.pdf.

Anonymous:1886:DSb

[8561] Anonymous. Door securer. Scientific American, 55(3):34, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n3/pdf/scientificamerican07171886-34b.pdf.

Anonymous:1886:CCl

[8562] Anonymous. Car coupling. Scientific American, 55(3):34, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n3/pdf/scientificamerican07171886-34c.pdf.

Anonymous:1886:SMb

[8563] Anonymous. Shaving mirror. Scientific American, 55(3):34, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n3/pdf/scientificamerican07171886-34d.pdf.

Anonymous:1886:PPR

[8564] Anonymous. Pneumatic pressure relief governor. Scientific American, 55(3):35, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n3/pdf/scientificamerican07171886-
35.pdf.

Anonymous:1886:WLI

[8565] Anonymous. Water level indicator. Scientific American, 55(3):35–36, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n3/pdf/scientificamerican07171886-35a.pdf.

Sloane:1886:CBH

[8566] T. O’Conor Sloane. Collision balls and harmonic impact. Scientific American, 55(3):38, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n3/pdf/scientificamerican07171886-
38.pdf.

Hopkins:1886:PLN

[8567] Geo. M. Hopkins. Polarized light. — Norremberg doubler and the study of thin films. Scientific American, 55(3):38–39, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n3/pdf/scientificamerican07171886-38a.pdf.

Anonymous:1886:WEA

[8568] Anonymous. Winding engine for Australia. Scientific American, 55(3):39, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n3/pdf/scientificamerican07171886-
39.pdf.

Anonymous:1886:TEA

[8569] Anonymous. Thermometer with electric alarm. Scientific American, 55(3):39–40, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n3/pdf/scientificamerican07171886-
39a.pdf.

Anonymous:1886:DTP

[8570] Anonymous. Drain tile protector. Scientific American, 55(3):40, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n3/pdf/scientificamerican07171886-40.pdf.

Anonymous:1886:RCT

[8571] Anonymous. Railway cross ties. Scientific American, 55(3):40, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n3/pdf/scientificamerican07171886-40a.pdf.

Anonymous:1886:CCn

[8572] Anonymous. Child’s carriage. Scientific American, 55(3):40–41, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n3/pdf/scientificamerican07171886-40b.pdf.

Anonymous:1886:RSM

[8573] Anonymous. Rare specimens of marmosets. Scientific American, 55(3):41, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n3/pdf/scientificamerican07171886-
41.pdf.

Anonymous:1886:EIbt

[8574] Anonymous. Engineering inventions. Scientific American, 55(3):42, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n3/pdf/scientificamerican07171886-42.pdf.

Anonymous:1886:AIbk

[8575] Anonymous. Agricultural inventions. Scientific American, 55(3):42, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n3/pdf/scientificamerican07171886-42a.pdf.

Anonymous:1886:MIcc

[8576] Anonymous. Miscellaneous inventions. Scientific American, 55(3):42, July 17, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n3/pdf/scientificamerican07171886-42b.pdf.

Anonymous:1886:Cbr

[8577] Anonymous. Correspondence. Scientific American, 55(4):53, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n4/pdf/scientificamerican07241886-53.pdf.

Anonymous:1886:BPbu

[8578] Anonymous. Business and personal. Scientific American, 55(4):59, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n4/pdf/scientificamerican07241886-59c.pdf.

Anonymous:1886:NQbs

[8579] Anonymous. Notes and queries. Scientific American, 55(4):59, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n4/pdf/scientificamerican07241886-59d.pdf.

Anonymous:1886:IIbt

[8580] Anonymous. Index of inventions. Scientific American, 55(4):59–60, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n4/pdf/scientificamerican07241886-59f.pdf.

Anonymous:1886:PNPbq

[8581] Anonymous. Patent notes: Patents. Scientific American, 55(4):62, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n4/pdf/scientificamerican07241886-62.pdf.

Anonymous:1886:PSY

[8582] Anonymous. Propulsion of the Secor yacht Eureka by the explosion of petroleum vapor. Scientific American, 55(4):47, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n4/pdf/scientificamerican07241886-47.pdf.

Anonymous:1886:TBG

[8583] Anonymous. Torpedo boats in the gale. Scientific American, 55(4):48, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n4/pdf/scientificamerican07241886-48.pdf.

Anonymous:1886:BSP

[8584] Anonymous. The basic steel process. Scientific American, 55(4):48–49, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n4/pdf/scientificamerican07241886-48a.pdf.

Anonymous:1886:ESWb

[8585] Anonymous. Efficiency of small water motors. Scientific American, 55(4):49, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n4/pdf/scientificamerican07241886-
49.pdf.

Anonymous:1886:IWc

[8586] Anonymous. Improved windmill. Scientific American, 55(4):50, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n4/pdf/scientificamerican07241886-50.pdf.

Anonymous:1886:IWd

[8587] Anonymous. Improved whiffletree. Scientific American, 55(4):50, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n4/pdf/scientificamerican07241886-50a.pdf.

Anonymous:1886:IDAa

[8588] Anonymous. Improved desk attachment. Scientific American, 55(4):50, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n4/pdf/scientificamerican07241886-50b.pdf.

Anonymous:1886:WPM

[8589] Anonymous. Wire painting machine. Scientific American, 55(4):50, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n4/pdf/scientificamerican07241886-50c.pdf.

Anonymous:1886:RTc

[8590] Anonymous. Railway track. Scientific American, 55(4):50, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n4/pdf/scientificamerican07241886-50d.pdf.

Anonymous:1886:MFG

[8591] Anonymous. Making fine guns for sporting purposes. Scientific American, 55(4):51, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n4/pdf/scientificamerican07241886-
51.pdf.

Anonymous:1886:AFB

[8592] Anonymous. Anti-friction bearing. Scientific American, 55(4):51, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n4/pdf/scientificamerican07241886-51a.pdf.

Anonymous:1886:BFb

[8593] Anonymous. Banjo frame. Scientific American, 55(4):51, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n4/pdf/scientificamerican07241886-51b.pdf.

Anonymous:1886:PNbb

[8594] Anonymous. Photographic notes. Scientific American, 55(4):52, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n4/pdf/scientificamerican07241886-52.pdf.

Anonymous:1886:POW

[8595] Anonymous. The petroleum oil wells of Baku. Scientific American, 55(4):54, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n4/pdf/scientificamerican07241886-
54.pdf.

Hopkins:1886:PLc

[8596] Geo. M. Hopkins. Polarized light. Scientific American, 55(4):55, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n4/pdf/scientificamerican07241886-55.pdf.

Anonymous:1886:DRPj

[8597] Anonymous. Decisions relating to patents. Scientific American, 55(4):56, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n4/pdf/scientificamerican07241886-
56.pdf.

Anonymous:1886:VEN

[8598] Anonymous. Volcanic eruptions in New Zealand. Scientific American, 55(4):57, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n4/pdf/scientificamerican07241886-
57.pdf.

Sloane:1886:WPF

[8599] T. O’Conor Sloane. Wedge paradox and falling and projected ball. Scientific American, 55(4):57–58, July 24, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v55/
n4/pdf/scientificamerican07241886-57a.pdf.

Anonymous:1886:EIbu

[8600] Anonymous. Engineering inventions. Scientific American, 55(4):59, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n4/pdf/scientificamerican07241886-59.pdf.

Anonymous:1886:AIbl

[8601] Anonymous. Agricultural inventions. Scientific American, 55(4):59, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n4/pdf/scientificamerican07241886-59a.pdf.

Anonymous:1886:MIcd

[8602] Anonymous. Miscellaneous inventions. Scientific American, 55(4):59, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n4/pdf/scientificamerican07241886-59b.pdf.

Anonymous:1886:Ibd

[8603] Anonymous. To inventors. Scientific American, 55(4):59, July 24, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n4/pdf/scientificamerican07241886-59e.pdf.

Anonymous:1886:Cbs

[8604] Anonymous. Correspondence. Scientific American, 55(5):69, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n5/pdf/scientificamerican07311886-69a.pdf.

Anonymous:1886:BNBbg

[8605] Anonymous. Books: New books and publications. Scientific American, 55(5):74–75, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n5/pdf/scientificamerican07311886-
74b.pdf.

Anonymous:1886:BPbv

[8606] Anonymous. Business and personal. Scientific American, 55(5):75, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n5/pdf/scientificamerican07311886-75.pdf.

Anonymous:1886:NQbt

[8607] Anonymous. Notes and queries. Scientific American, 55(5):75, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n5/pdf/scientificamerican07311886-75a.pdf.

Anonymous:1886:IIbu

[8608] Anonymous. Index of inventions. Scientific American, 55(5):75–76, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n5/pdf/scientificamerican07311886-75c.pdf.

Anonymous:1886:PNPbr

[8609] Anonymous. Patent notes: Patents. Scientific American, 55(5):78, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n5/pdf/scientificamerican07311886-78.pdf.

Anonymous:1886:NYE

[8610] Anonymous. New York electrical subways. Scientific American, 55(5):64, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n5/pdf/scientificamerican07311886-
64.pdf.

Anonymous:1886:STAa

[8611] Anonymous. Ship transit across the Isthmus. Scientific American, 55(5):64, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n5/pdf/scientificamerican07311886-
64a.pdf.

Anonymous:1886:Bd

[8612] Anonymous. Baseball. Scientific American, 55(5):64–65, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n5/pdf/scientificamerican07311886-64b.pdf.

Anonymous:1886:DPV

[8613] Anonymous. Discharge pipe for vaults. Scientific American, 55(5):66, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n5/pdf/scientificamerican07311886-66.pdf.

Anonymous:1886:CRC

[8614] Anonymous. Check row corn planter. Scientific American, 55(5):66, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n5/pdf/scientificamerican07311886-66a.pdf.

Anonymous:1886:WS

[8615] Anonymous. Watchcase spring. Scientific American, 55(5):66, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n5/pdf/scientificamerican07311886-66b.pdf.

Anonymous:1886:MW

[8616] Anonymous. Measuring wheel. Scientific American, 55(5):66, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n5/pdf/scientificamerican07311886-66c.pdf.

Anonymous:1886:DCSa

[8617] Anonymous. Drawer check and support. Scientific American, 55(5):66, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n5/pdf/scientificamerican07311886-66d.pdf.

Anonymous:1886:CCVb

[8618] Anonymous. Clip coupling for vehicles. Scientific American, 55(5):67, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n5/pdf/scientificamerican07311886-67.pdf.

Anonymous:1886:FWT

[8619] Anonymous. Fence wire tightener. Scientific American, 55(5):67, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n5/pdf/scientificamerican07311886-67a.pdf.

Anonymous:1886:LFA

[8620] Anonymous. Lamp filling attachment for oil cans. Scientific American, 55(5):67, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n5/pdf/scientificamerican07311886-
67b.pdf.

Anonymous:1886:IQF

[8621] Anonymous. Improved quilting frame. Scientific American, 55 (5):67–68, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n5/pdf/scientificamerican07311886-
67c.pdf.

Anonymous:1886:FDA

[8622] Anonymous. Fireworks with dramatic accessories. Scientific American, 55(5):69, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n5/pdf/scientificamerican07311886-
69.pdf.

Hopkins:1886:PLd

[8623] Geo. M. Hopkins. Polarized light. Scientific American, 55(5):70, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n5/pdf/scientificamerican07311886-70.pdf.

Chadwick:1886:APB

[8624] Henry Chadwick. The art of pitching in baseball. Scientific American, 55(5):71–72, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n5/pdf/scientificamerican07311886-
71.pdf.

Anonymous:1886:PNbc

[8625] Anonymous. Photographic notes. Scientific American, 55(5):73, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n5/pdf/scientificamerican07311886-73.pdf.

Anonymous:1886:DAb

[8626] Anonymous. Driving apparatus. Scientific American, 55(5):73, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n5/pdf/scientificamerican07311886-73a.pdf.

Anonymous:1886:H

[8627] Anonymous. Hornbills. Scientific American, 55(5):73, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n5/pdf/scientificamerican07311886-73b.pdf.

Anonymous:1886:EIbv

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n5/pdf/scientificamerican07311886-74.pdf.

Anonymous:1886:MIce

[8629] Anonymous. Miscellaneous inventions. Scientific American, 55(5):74, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n5/pdf/scientificamerican07311886-74a.pdf.

Anonymous:1886:Ibe

[8630] Anonymous. To inventors. Scientific American, 55(5):75, July 31, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n5/pdf/scientificamerican07311886-75b.pdf.

Anonymous:1886:Cbt

[8631] Anonymous. Correspondence. Scientific American, 55(6):85, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n6/pdf/scientificamerican08071886-85a.pdf.

Anonymous:1886:BNBbh

[8632] Anonymous. Books: New books and publications. Scientific American, 55(6):90, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n6/pdf/scientificamerican08071886-
90c.pdf.

Anonymous:1886:BPbw

[8633] Anonymous. Business and personal. Scientific American, 55(6):90, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n6/pdf/scientificamerican08071886-90d.pdf.

Anonymous:1886:NQbu

[8634] Anonymous. Notes and queries. Scientific American, 55(6):91, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n6/pdf/scientificamerican08071886-91.pdf.

Anonymous:1886:IIbv

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n6/pdf/scientificamerican08071886-91b.pdf.

Anonymous:1886:PNPbs

[8636] Anonymous. Patent notes: Patents. Scientific American, 55(6):94, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n6/pdf/scientificamerican08071886-94.pdf.

Anonymous:1886:AAA

[8637] Anonymous. The American Association for the Advancement of Science. Scientific American, 55(6):80, August 7, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v55/
n6/pdf/scientificamerican08071886-80.pdf.

Anonymous:1886:TPEb

[8638] Anonymous. Transmission of power by electricity. Scientific American, 55(6):80, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n6/pdf/scientificamerican08071886-
80a.pdf.

Anonymous:1886:CNC

[8639] Anonymous. The character of the new cruisers. Scientific American, 55(6):80, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n6/pdf/scientificamerican08071886-
80b.pdf.

Anonymous:1886:OB

[8640] Anonymous. Our boys. Scientific American, 55(6):80–81, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n6/pdf/scientificamerican08071886-80c.pdf.

Proctor:1886:NSJc

[8641] Richard A. Proctor. Night sky — July and August. Scientific American, 55(6):81, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n6/pdf/scientificamerican08071886-
81.pdf.

Anonymous:1886:LB

[8642] Anonymous. Lamp bracket. Scientific American, 55(6):82, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n6/pdf/scientificamerican08071886-82.pdf.

Anonymous:1886:ISFa

[8643] Anonymous. Improved sash fastener. Scientific American, 55(6):82, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n6/pdf/scientificamerican08071886-82a.pdf.

Anonymous:1886:ISHb

[8644] Anonymous. Improved snap hook. Scientific American, 55(6):82, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n6/pdf/scientificamerican08071886-82b.pdf.

Anonymous:1886:HFN

[8645] Anonymous. A horseshoe to fit the natural foot of a horse. Scientific American, 55(6):82, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n6/pdf/scientificamerican08071886-
82c.pdf.

Anonymous:1886:ICg

[8646] Anonymous. Improved churn. Scientific American, 55(6):82–83, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n6/pdf/scientificamerican08071886-82d.pdf.

Anonymous:1886:CAE

[8647] Anonymous. Curious accretion of emery wheel dust. Scientific American, 55(6):83, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n6/pdf/scientificamerican08071886-
83.pdf.

Anonymous:1886:NEO

[8648] Anonymous. New electric organ movement. Scientific American, 55(6):83–84, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n6/pdf/scientificamerican08071886-
83a.pdf.

Anonymous:1886:IPW

[8649] Anonymous. Improved pipe wrench. Scientific American, 55(6):84, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n6/pdf/scientificamerican08071886-84.pdf.

Anonymous:1886:MPN

[8650] Anonymous. Morningside Park, New York City. Scientific American, 55(6):85, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n6/pdf/scientificamerican08071886-
85.pdf.

Anonymous:1886:PNbd

[8651] Anonymous. Photographic notes. Scientific American, 55(6):85, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n6/pdf/scientificamerican08071886-85b.pdf.

Anonymous:1886:DRPk

[8652] Anonymous. Decisions relating to patents. Scientific American, 55(6):85, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n6/pdf/scientificamerican08071886-
85c.pdf.

Anonymous:1886:SCQ

[8653] Anonymous. Six cylinder quadruple expansion engines. Scientific American, 55(6):86–87, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n6/pdf/scientificamerican08071886-
86.pdf.

Anonymous:1886:NST

[8654] Anonymous. A new submarine torpedo boat. Scientific American, 55(6):88, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n6/pdf/scientificamerican08071886-
88.pdf.

Sloane:1886:CF

[8655] T. O’conor Sloane. Centrifugal force. Scientific American, 55(6):89, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n6/pdf/scientificamerican08071886-89.pdf.

Anonymous:1886:EIbw

[8656] Anonymous. Engineering inventions. Scientific American, 55(6):90, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n6/pdf/scientificamerican08071886-90.pdf.

Anonymous:1886:AIbm

[8657] Anonymous. Agricultural inventions. Scientific American, 55(6):90, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n6/pdf/scientificamerican08071886-90a.pdf.

Anonymous:1886:MIcf

[8658] Anonymous. Miscellaneous inventions. Scientific American, 55(6):90, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n6/pdf/scientificamerican08071886-90b.pdf.

Anonymous:1886:Ibg

[8659] Anonymous. To inventor. Scientific American, 55(6):91, August 7, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n6/pdf/scientificamerican08071886-91a.pdf.

Anonymous:1886:Cbu

[8660] Anonymous. Correspondence. Scientific American, 55(7):106, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n7/pdf/scientificamerican08141886-106.pdf.

Anonymous:1886:BPbx

[8661] Anonymous. Business and personal. Scientific American, 55(7):107, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n7/pdf/scientificamerican08141886-107b.pdf.

Anonymous:1886:BNBbi

[8662] Anonymous. Books: New books and publications. Scientific American, 55(7):107, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n7/pdf/scientificamerican08141886-
107c.pdf.

Anonymous:1886:NQbv

[8663] Anonymous. Notes and queries. Scientific American, 55(7):107, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n7/pdf/scientificamerican08141886-107d.pdf.

Anonymous:1886:IIbw

[8664] Anonymous. Index of inventions. Scientific American, 55(7): 107–108, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n7/pdf/scientificamerican08141886-
107f.pdf.

Anonymous:1886:PNPbt

[8665] Anonymous. Patent notes: Patents. Scientific American, 55(7):110, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n7/pdf/scientificamerican08141886-110.pdf.

Anonymous:1886:PIS

[8666] Anonymous. Production of iron and steel in America. Scientific American, 55(7):96, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n7/pdf/scientificamerican08141886-
96.pdf.

Anonymous:1886:HOC

[8667] Anonymous. Heavy ordnance for coast defense. Scientific American, 55(7):96, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n7/pdf/scientificamerican08141886-
96a.pdf.

Anonymous:1886:DGM

[8668] Anonymous. Disease germs in milk. Scientific American, 55 (7):96–97, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n7/pdf/scientificamerican08141886-
96b.pdf.

Anonymous:1886:ICi

[8669] Anonymous. Improved cultivator. Scientific American, 55(7):98, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n7/pdf/scientificamerican08141886-98.pdf.

Anonymous:1886:RMb

[8670] Anonymous. Rubbing machine. Scientific American, 55(7):98, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n7/pdf/scientificamerican08141886-98a.pdf.

Anonymous:1886:PSV

[8671] Anonymous. Platform spring for vehicles. Scientific American, 55(7):98, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n7/pdf/scientificamerican08141886-
98b.pdf.

Anonymous:1886:RE

[8672] Anonymous. Rotary engine. Scientific American, 55(7):98, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n7/pdf/scientificamerican08141886-98c.pdf.

Anonymous:1886:SSe

[8673] Anonymous. Switch stand. Scientific American, 55(7):98, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n7/pdf/scientificamerican08141886-98d.pdf.

Anonymous:1886:ISHc

[8674] Anonymous. Improved steam heater. Scientific American, 55(7):99, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n7/pdf/scientificamerican08141886-99.pdf.

Anonymous:1886:LDb

[8675] Anonymous. Levesque’s diplograph. Scientific American, 55(7):99, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n7/pdf/scientificamerican08141886-99a.pdf.

Anonymous:1886:CBCa

[8676] Anonymous. Cracker box cover. Scientific American, 55(7):99–100, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n7/pdf/scientificamerican08141886-99b.pdf.

Anonymous:1886:SLNb

[8677] Anonymous. The Statue of Liberty nearing completion. Scientific American, 55(7):100–101, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n7/pdf/scientificamerican08141886-
100.pdf.

Hovey:1886:GLJ

[8678] H. C. Hovey. The great log jam in the St. Croix River. Scientific American, 55(7):101–102, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n7/pdf/scientificamerican08141886-
101.pdf.

Hopkins:1886:FFM

[8679] Geo. M. Hopkins. The formation and fixation of magnetic curves. Scientific American, 55(7):102, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n7/pdf/scientificamerican08141886-
102.pdf.

Anonymous:1886:OFR

[8680] Anonymous. The occurrence and fabrication of rock crystal. Scientific American, 55(7):103–104, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v55/n7/pdf/scientificamerican08141886-
103.pdf.

Anonymous:1886:WLC

[8681] Anonymous. A wrench with lift cams. Scientific American, 55 (7):104, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n7/pdf/scientificamerican08141886-
104.pdf.

Anonymous:1886:HBL

[8682] Anonymous. A hammerless breech-loading gun. Scientific American, 55(7):104, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n7/pdf/scientificamerican08141886-
104a.pdf.

Anonymous:1886:DCL

[8683] Anonymous. Device for controlling locks on railway cars. Scientific American, 55(7):104, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n7/pdf/scientificamerican08141886-
104b.pdf.

Anonymous:1886:RWF

[8684] Anonymous. Rain water filter. Scientific American, 55(7):104, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n7/pdf/scientificamerican08141886-104c.pdf.

Sloane:1886:ESa

[8685] T. O’conor Sloane. Experiments in sound. Scientific American, 55(7):105, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n7/pdf/scientificamerican08141886-
105.pdf.

Anonymous:1886:DCI

[8686] Anonymous. Disintegration of cast iron by coal gas flame. Scientific American, 55(7):105, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n7/pdf/scientificamerican08141886-
105a.pdf.

Anonymous:1886:PNbe

[8687] Anonymous. Photographic notes. Scientific American, 55(7):106, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n7/pdf/scientificamerican08141886-106a.pdf.

Anonymous:1886:EIbx

[8688] Anonymous. Engineering inventions. Scientific American, 55(7): 107, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n7/pdf/scientificamerican08141886-
107.pdf.

Anonymous:1886:MIcg

[8689] Anonymous. Miscellaneous inventions. Scientific American, 55 (7):107, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n7/pdf/scientificamerican08141886-
107a.pdf.

Anonymous:1886:Ibh

[8690] Anonymous. To inventors. Scientific American, 55(7):107, August 14, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n7/pdf/scientificamerican08141886-107e.pdf.

Anonymous:1886:Cbv

[8691] Anonymous. Correspondence. Scientific American, 55(8):117, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n8/pdf/scientificamerican08211886-117.pdf.

Anonymous:1886:BNBbj

[8692] Anonymous. Books: New books and publications. Scientific American, 55(8):123, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n8/pdf/scientificamerican08211886-
123c.pdf.

Anonymous:1886:BPby

[8693] Anonymous. Business and personal. Scientific American, 55(8):123, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n8/pdf/scientificamerican08211886-123d.pdf.

Anonymous:1886:NQbw

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n8/pdf/scientificamerican08211886-123e.pdf.

Anonymous:1886:IIbx

[8695] Anonymous. Index of inventions. Scientific American, 55(8): 123–124, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n8/pdf/scientificamerican08211886-
123f.pdf.

Anonymous:1886:PNPbu

[8696] Anonymous. Patent notes: Patents. Scientific American, 55(8):126, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n8/pdf/scientificamerican08211886-126.pdf.

Anonymous:1886:HDW

[8697] Anonymous. How deep wells are made. Scientific American, 55(8):111, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n8/pdf/scientificamerican08211886-
111.pdf.

Anonymous:1886:CNY

[8698] Anonymous. The channelways of New York. Scientific American, 55(8):112–113, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n8/pdf/scientificamerican08211886-
112.pdf.

Anonymous:1886:PNbf

[8699] Anonymous. Photographic notes. Scientific American, 55(8):113, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n8/pdf/scientificamerican08211886-113.pdf.

Anonymous:1886:LBS

[8700] Anonymous. Lee board for small vessels. Scientific American, 55(8):114, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n8/pdf/scientificamerican08211886-
114.pdf.

Anonymous:1886:DOW

[8701] Anonymous. Device for opening window sashes. Scientific American, 55(8):114, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n8/pdf/scientificamerican08211886-
114a.pdf.

Anonymous:1886:PP

[8702] Anonymous. Plant protector. Scientific American, 55(8):114, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n8/pdf/scientificamerican08211886-114b.pdf.

Anonymous:1886:WBb

[8703] Anonymous. Window blind. Scientific American, 55(8):114, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n8/pdf/scientificamerican08211886-114c.pdf.

Anonymous:1886:SBc

[8704] Anonymous. Sash balance. Scientific American, 55(8):114, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n8/pdf/scientificamerican08211886-114d.pdf.

Anonymous:1886:RBb

[8705] Anonymous. Rotary brush. Scientific American, 55(8):115, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n8/pdf/scientificamerican08211886-115.pdf.

Anonymous:1886:IDe

[8706] Anonymous. Improved derrick. Scientific American, 55(8):115, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n8/pdf/scientificamerican08211886-115a.pdf.

Anonymous:1886:IVa

[8707] Anonymous. Improved vulcanizer. Scientific American, 55(8):115, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n8/pdf/scientificamerican08211886-115b.pdf.

Anonymous:1886:RSFb

[8708] Anonymous. Refrigerator storehouse for fruits, etc. Scientific American, 55(8):115, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n8/pdf/scientificamerican08211886-
115c.pdf.

Anonymous:1886:DMC

[8709] Anonymous. Detachable milk can refrigerator. Scientific American, 55(8):118, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n8/pdf/scientificamerican08211886-
118.pdf.

Anonymous:1886:WT

[8710] Anonymous. Writing tablet. Scientific American, 55(8):118, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n8/pdf/scientificamerican08211886-118a.pdf.

Anonymous:1886:PGc

[8711] Anonymous. Protective garment. Scientific American, 55(8):118, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n8/pdf/scientificamerican08211886-118b.pdf.

Anonymous:1886:DCSb

[8712] Anonymous. Drawer check and support. Scientific American, 55 (8):118–119, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n8/pdf/scientificamerican08211886-
118c.pdf.

Sloane:1886:ESb

[8713] T. O’conor Sloane. Experiments in sound. Scientific American, 55(8):119, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n8/pdf/scientificamerican08211886-
119.pdf.

Anonymous:1886:Cbw

[8714] Anonymous. Chicago. Scientific American, 55(8):120–121, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n8/pdf/scientificamerican08211886-120.pdf.

Coryell:1886:ZM

[8715] John R. Coryell. The Zarabatana of the Macoushies. Scientific American, 55(8):121–122, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n8/pdf/scientificamerican08211886-
121.pdf.

LaRue:1886:DRP

[8716] ?. La Rue. Decisions relating to patents. Scientific American, 55(8):122, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n8/pdf/scientificamerican08211886-
122.pdf.

Anonymous:1886:EIby

[8717] Anonymous. Engineering invention. Scientific American, 55(8):123, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n8/pdf/scientificamerican08211886-123.pdf.

Anonymous:1886:MIch

[8718] Anonymous. Mechanical invention. Scientific American, 55(8):123, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n8/pdf/scientificamerican08211886-123a.pdf.

Anonymous:1886:MIci

[8719] Anonymous. Miscellaneous inventions. Scientific American, 55 (8):123, August 21, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n8/pdf/scientificamerican08211886-
123b.pdf.

Anonymous:1886:Cbx

[8720] Anonymous. Correspondence. Scientific American, 55(9):133, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n9/pdf/scientificamerican08281886-133.pdf.

Anonymous:1886:BNBbk

[8721] Anonymous. Books: New books and publications. Scientific American, 55(9):138, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n9/pdf/scientificamerican08281886-
138c.pdf.

Anonymous:1886:BPbz

[8722] Anonymous. Business and personal. Scientific American, 55(9):138, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n9/pdf/scientificamerican08281886-138d.pdf.

Anonymous:1886:NQbx

[8723] Anonymous. Notes and queries. Scientific American, 55(9):139, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n9/pdf/scientificamerican08281886-139.pdf.

Anonymous:1886:IIby

[8724] Anonymous. Index of inventions. Scientific American, 55(9): 139–142, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n9/pdf/scientificamerican08281886-
139b.pdf.

Anonymous:1886:PNPbv

[8725] Anonymous. Patent notes: Patents. Scientific American, 55(9):142, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n9/pdf/scientificamerican08281886-142.pdf.

Hovey:1886:BCD

[8726] H. C. Hovey. Brush’s colossal dynamo. Scientific American, 55 (9):127, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n9/pdf/scientificamerican08281886-
127.pdf.

Anonymous:1886:EPDb

[8727] Anonymous. The effect of patents in the development of the gas industry. Scientific American, 55(9):128, August 28, 1886. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v55/
n9/pdf/scientificamerican08281886-128.pdf.

Anonymous:1886:TTb

[8728] Anonymous. The Tilden Trust. Scientific American, 55(9):128–129, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n9/pdf/scientificamerican08281886-128a.pdf.

Anonymous:1886:APBb

[8729] Anonymous. The art of pitching in baseball. Scientific American, 55(9):129, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n9/pdf/scientificamerican08281886-
129.pdf.

Anonymous:1886:GBT

[8730] Anonymous. Good and bad taxi dermal art. Scientific American, 55(9):129, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n9/pdf/scientificamerican08281886-
129a.pdf.

Anonymous:1886:PNbg

[8731] Anonymous. Photographic notes. Scientific American, 55(9):129, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n9/pdf/scientificamerican08281886-129b.pdf.

Anonymous:1886:CSc

[8732] Anonymous. Carriage screen. Scientific American, 55(9):130, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n9/pdf/scientificamerican08281886-130.pdf.

Anonymous:1886:FE

[8733] Anonymous. Fire escape. Scientific American, 55(9):130, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n9/pdf/scientificamerican08281886-130a.pdf.

Anonymous:1886:ISSb

[8734] Anonymous. Improved single spring for vehicle. Scientific American, 55(9):130, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n9/pdf/scientificamerican08281886-
130b.pdf.

Anonymous:1886:SMPb

[8735] Anonymous. Sheet metal picket fence. Scientific American, 55 (9):130, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n9/pdf/scientificamerican08281886-
130c.pdf.

Anonymous:1886:CRa

[8736] Anonymous. Coal receptacle. Scientific American, 55(9):131, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n9/pdf/scientificamerican08281886-131.pdf.

Anonymous:1886:AS

[8737] Anonymous. Ash sifter. Scientific American, 55(9):131, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n9/pdf/scientificamerican08281886-131a.pdf.

Anonymous:1886:HRC

[8738] Anonymous. Head rest for chairs. Scientific American, 55(9):131, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n9/pdf/scientificamerican08281886-131b.pdf.

Anonymous:1886:CCo

[8739] Anonymous. Chimney cowl. Scientific American, 55(9):131–132, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n9/pdf/scientificamerican08281886-131c.pdf.

Anonymous:1886:AAT

[8740] Anonymous. Adhesive address tag for travelers. Scientific American, 55(9):132, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n9/pdf/scientificamerican08281886-
132.pdf.

Anonymous:1886:SPF

[8741] Anonymous. The Star portable forges. Scientific American, 55 (9):132, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n9/pdf/scientificamerican08281886-
132a.pdf.

Anonymous:1886:DAc

[8742] Anonymous. Disinfecting apparatus. Scientific American, 55(9): 134–135, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n9/pdf/scientificamerican08281886-
134.pdf.

Anonymous:1886:ERC

[8743] Anonymous. Estrade’s railway cars. Scientific American, 55(9):135, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n9/pdf/scientificamerican08281886-135.pdf.

Anonymous:1886:VDT

[8744] Anonymous. The value of a desert tree. Scientific American, 55 (9):135–136, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n9/pdf/scientificamerican08281886-
135a.pdf.

Anonymous:1886:HMH

[8745] Anonymous. A home-made humming top. Scientific American, 55(9):137, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n9/pdf/scientificamerican08281886-
137.pdf.

Anonymous:1886:BDM

[8746] Anonymous. Boiler drilling machine. Scientific American, 55 (9):137, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n9/pdf/scientificamerican08281886-
137a.pdf.

Anonymous:1886:EIbz

[8747] Anonymous. Engineering inventions. Scientific American, 55(9): 138, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n9/pdf/scientificamerican08281886-
138.pdf.

Anonymous:1886:AIbn

[8748] Anonymous. Agricultural inventions. Scientific American, 55(9): 138, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n9/pdf/scientificamerican08281886-
138a.pdf.

Anonymous:1886:MIcj

[8749] Anonymous. Miscellaneous inventions. Scientific American, 55 (9):138, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v55/n9/pdf/scientificamerican08281886-
138b.pdf.

Anonymous:1886:Ibi

[8750] Anonymous. To inventors. Scientific American, 55(9):139, August 28, 1886. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v55/n9/pdf/scientificamerican08281886-139a.pdf.

Anonymous:1887:Ca

[8751] Anonymous. Correspondence. Scientific American, 56(1):7, January 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n1/pdf/scientificamerican01011887-7.pdf.

Anonymous:1887:BPa

[8752] Anonymous. Business and personal. Scientific American, 56(1):11, January 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n1/pdf/scientificamerican01011887-11c.pdf.

Anonymous:1887:IIa

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n1/pdf/scientificamerican01011887-11e.pdf.

Anonymous:1887:NFAa

[8754] Anonymous. A novel form of aerial vessel. Scientific American, 56(1):1, January 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n1/pdf/scientificamerican01011887-1.pdf.

Anonymous:1887:HSCa

[8755] Anonymous. Hiring ships to carry the flag. Scientific American, 56(1):2, January 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n1/pdf/scientificamerican01011887-2.pdf.

Anonymous:1887:DSPa

[8756] Anonymous. Direct steam propulsion. Scientific American, 56(1):3, January 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n1/pdf/scientificamerican01011887-3.pdf.

Anonymous:1887:SMU

[8757] Anonymous. Snow melting and underground conduits. Scientific American, 56(1):4, January 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n1/pdf/scientificamerican01011887-
4.pdf.

Anonymous:1887:CCa

[8758] Anonymous. Car coupling. Scientific American, 56(1):5, January 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n1/pdf/scientificamerican01011887-5.pdf.

Anonymous:1887:CCb

[8759] Anonymous. Car coupling. Scientific American, 56(1):5, January 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n1/pdf/scientificamerican01011887-5a.pdf.

Anonymous:1887:CWB

[8760] Anonymous. Convertible wire basket. Scientific American, 56(1):5, January 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n1/pdf/scientificamerican01011887-5b.pdf.

Anonymous:1887:TM

[8761] Anonymous. Tape measure. Scientific American, 56(1):5, January 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n1/pdf/scientificamerican01011887-5c.pdf.

Anonymous:1887:OT

[8762] Anonymous. Outlining tool. Scientific American, 56(1):5–6, January 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n1/pdf/scientificamerican01011887-5d.pdf.

Anonymous:1887:PSW

[8763] Anonymous. Pendant stem for watches. Scientific American, 56(1):6, January 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n1/pdf/scientificamerican01011887-6.pdf.

Anonymous:1887:FS

[8764] Anonymous. Flexible scraper. Scientific American, 56(1):6, January 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n1/pdf/scientificamerican01011887-6a.pdf.

Anonymous:1887:ICa

[8765] Anonymous. Improved cultivator. Scientific American, 56(1):6, January 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n1/pdf/scientificamerican01011887-6b.pdf.

Anonymous:1887:LBP

[8766] Anonymous. The Leikoes Breakwater, Portugal. Scientific American, 56(1):8, January 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n1/pdf/scientificamerican01011887-
8.pdf.

Anonymous:1887:RPA

[8767] Anonymous. Rolling platforms and armor-clad batteries. Scientific American, 56(1):8–9, January 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n1/pdf/scientificamerican01011887-
8a.pdf.

Anonymous:1887:CGT

[8768] Anonymous. Curious growth of tree. Scientific American, 56(1):9, January 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n1/pdf/scientificamerican01011887-9.pdf.

Anonymous:1887:GE

[8769] Anonymous. The Great Eastern. Scientific American, 56(1):9–10, January 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n1/pdf/scientificamerican01011887-9a.pdf.

Anonymous:1887:EIa

[8770] Anonymous. Engineering inventions. Scientific American, 56(1):11, January 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n1/pdf/scientificamerican01011887-11.pdf.

Anonymous:1887:AIa

[8771] Anonymous. Agricultural inventions. Scientific American, 56(1):11, January 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n1/pdf/scientificamerican01011887-11a.pdf.

Anonymous:1887:MIa

[8772] Anonymous. Miscellaneous inventions. Scientific American, 56(1):11, January 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n1/pdf/scientificamerican01011887-11b.pdf.

Anonymous:1887:Ia

[8773] Anonymous. To inventors. Scientific American, 56(1):11, January 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n1/pdf/scientificamerican01011887-11d.pdf.

Anonymous:1887:Cb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n10/pdf/scientificamerican03051887-149.pdf.

Anonymous:1887:BPb

[8775] Anonymous. Business and personal. Scientific American, 56(10):154, March 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n10/pdf/scientificamerican03051887-154d.pdf.

Anonymous:1887:NQa

[8776] Anonymous. Notes and queries. Scientific American, 56(10):155, March 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n10/pdf/scientificamerican03051887-155.pdf.

Anonymous:1887:IIb

[8777] Anonymous. Index of inventions. Scientific American, 56(10):155–157, March 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n10/pdf/scientificamerican03051887-155a.pdf.

Anonymous:1887:PNPa

[8778] Anonymous. Patent notes: Patents. Scientific American, 56(10):158, March 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n10/pdf/scientificamerican03051887-158.pdf.

Anonymous:1887:LTM

[8779] Anonymous. Latest type of the Maxim machine gun. Scientific American, 56(10):143, March 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n10/pdf/scientificamerican03051887-
143.pdf.

Anonymous:1887:CD

[8780] Anonymous. Coast defenses. Scientific American, 56(10):144–145, March 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n10/pdf/scientificamerican03051887-144.pdf.

Anonymous:1887:NLT

[8781] Anonymous. A new laboratory tromp. Scientific American, 56(10):146, March 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n10/pdf/scientificamerican03051887-146.pdf.

Anonymous:1887:RPb

[8782] Anonymous. Rein protector. Scientific American, 56(10):147, March 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n10/pdf/scientificamerican03051887-147.pdf.

Anonymous:1887:RBP

[8783] Anonymous. Rafting boom pin. Scientific American, 56(10):147, March 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n10/pdf/scientificamerican03051887-147a.pdf.

Anonymous:1887:BFa

[8784] Anonymous. Binding file. Scientific American, 56(10):147, March 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n10/pdf/scientificamerican03051887-147b.pdf.

Anonymous:1887:CTb

[8785] Anonymous. Child’s tray. Scientific American, 56(10):147–148, March 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n10/pdf/scientificamerican03051887-147c.pdf.

Anonymous:1887:PHS

[8786] Anonymous. Piston head for steam engines. Scientific American, 56(10):148, March 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n10/pdf/scientificamerican03051887-
148.pdf.

Anonymous:1887:EE

[8787] Anonymous. The earthquake in Europe. Scientific American, 56 (10):150–151, March 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n10/pdf/scientificamerican03051887-
150.pdf.

Anonymous:1887:SFO

[8788] Anonymous. Some fine old carriages. Scientific American, 56 (10):151–152, March 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n10/pdf/scientificamerican03051887-
151.pdf.

Anonymous:1887:PNa

[8789] Anonymous. Photographic notes. Scientific American, 56(10):152, March 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n10/pdf/scientificamerican03051887-152.pdf.

Beard:1887:FMK

[8790] Daniel C. Beard. A Florida “mule killer” (Thelyphonus Giganteus). Scientific American, 56(10):153, March 5, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v56/
n10/pdf/scientificamerican03051887-153.pdf.

Anonymous:1887:CBT

[8791] Anonymous. A cheese box toboggan. Scientific American, 56(10):153, March 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n10/pdf/scientificamerican03051887-153a.pdf.

Anonymous:1887:EIb

[8792] Anonymous. Engineering inventions. Scientific American, 56(10):154, March 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n10/pdf/scientificamerican03051887-154.pdf.

Anonymous:1887:MIb

[8793] Anonymous. Mechanical invention. Scientific American, 56(10):154, March 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n10/pdf/scientificamerican03051887-154a.pdf.

Anonymous:1887:AIb

[8794] Anonymous. Agricultural invention. Scientific American, 56(10):154, March 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n10/pdf/scientificamerican03051887-154b.pdf.

Anonymous:1887:MIc

[8795] Anonymous. Miscellaneous inventions. Scientific American, 56(10):154, March 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n10/pdf/scientificamerican03051887-154c.pdf.

Anonymous:1887:Ce

[8796] Anonymous. Correspondence. Scientific American, 56(11):165, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n11/pdf/scientificamerican03121887-165a.pdf.

Anonymous:1887:BNBa

[8797] Anonymous. Books: New books and publications. Scientific American, 56(11):170, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n11/pdf/scientificamerican03121887-
170c.pdf.

Anonymous:1887:BPc

[8798] Anonymous. Business and personal. Scientific American, 56(11):170, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n11/pdf/scientificamerican03121887-170d.pdf.

Anonymous:1887:NQb

[8799] Anonymous. Notes and queries. Scientific American, 56(11):170–171, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n11/pdf/scientificamerican03121887-170e.pdf.

Anonymous:1887:IIc

[8800] Anonymous. Index of inventions. Scientific American, 56(11):171–172, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n11/pdf/scientificamerican03121887-171a.pdf.

Anonymous:1887:PNPb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n11/pdf/scientificamerican03121887-174.pdf.

Anonymous:1887:NNM

[8802] Anonymous. New Naval and military industries. Scientific American, 56(11):160, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n11/pdf/scientificamerican03121887-
160.pdf.

Anonymous:1887:WIa

[8803] Anonymous. War and invention. Scientific American, 56(11):160, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n11/pdf/scientificamerican03121887-160a.pdf.

Anonymous:1887:CCE

[8804] Anonymous. Celebration of the centennial of the enactment of the Patent Laws. Scientific American, 56(11):160–161, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n11/pdf/scientificamerican03121887-160b.pdf.

Anonymous:1887:DRA

[8805] Anonymous. Derailing and replacing attachment for street cars. Scientific American, 56(11):162, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v56/n11/pdf/scientificamerican03121887-
162.pdf.

Anonymous:1887:DSPb

[8806] Anonymous. Damper for stove pipes, chimneys, etc. Scientific American, 56(11):162, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n11/pdf/scientificamerican03121887-
162a.pdf.

Anonymous:1887:AFE

[8807] Anonymous. Automatic fire extinguisher. Scientific American, 56(11):162, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n11/pdf/scientificamerican03121887-
162b.pdf.

Anonymous:1887:ICR

[8808] Anonymous. Improved coal or rock drill. Scientific American, 56(11):162, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n11/pdf/scientificamerican03121887-
162c.pdf.

Anonymous:1887:IB

[8809] Anonymous. Improved bevel. Scientific American, 56(11):163, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n11/pdf/scientificamerican03121887-163.pdf.

Anonymous:1887:IEGa

[8810] Anonymous. Improvement in eye glasses. Scientific American, 56(11):163, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n11/pdf/scientificamerican03121887-
163a.pdf.

Anonymous:1887:PT

[8811] Anonymous. Plumber’s trap. Scientific American, 56(11):163, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n11/pdf/scientificamerican03121887-163b.pdf.

Anonymous:1887:ICTa

[8812] Anonymous. Improved cask or tub. Scientific American, 56(11):163, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n11/pdf/scientificamerican03121887-163c.pdf.

Anonymous:1887:IVP

[8813] Anonymous. Improved vehicle pole. Scientific American, 56(11):163, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n11/pdf/scientificamerican03121887-163d.pdf.

Anonymous:1887:CTC

[8814] Anonymous. Covering for traction cables. Scientific American, 56(11):163, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n11/pdf/scientificamerican03121887-
163e.pdf.

Anonymous:1887:DRPa

[8815] Anonymous. Decision relating to patents. Scientific American, 56(11):164, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n11/pdf/scientificamerican03121887-
164.pdf.

Anonymous:1887:IGL

[8816] Anonymous. Improved gate latch. Scientific American, 56(11):164, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n11/pdf/scientificamerican03121887-164a.pdf.

Anonymous:1887:CCNa

[8817] Anonymous. The College of the City of New York — the technical course. Scientific American, 56(11):165, March 12, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v56/
n11/pdf/scientificamerican03121887-165.pdf.

Anonymous:1887:SSB

[8818] Anonymous. Sectional steel boat for Mr. Stanley. Scientific American, 56(11):166, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n11/pdf/scientificamerican03121887-
166.pdf.

Anonymous:1887:CAF

[8819] Anonymous. Convertible anchor and freight hoisting apparatus for vessels. Scientific American, 56(11):166, March 12, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v56/
n11/pdf/scientificamerican03121887-166a.pdf.

Anonymous:1887:EWS

[8820] Anonymous. Ernst Werner Siemens. Scientific American, 56(11):167, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n11/pdf/scientificamerican03121887-167.pdf.

Anonymous:1887:MD

[8821] Anonymous. Mechanical dyeing. Scientific American, 56(11):167–168, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n11/pdf/scientificamerican03121887-167a.pdf.

Anonymous:1887:E

[8822] Anonymous. The Epiornis. Scientific American, 56(11):168, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n11/pdf/scientificamerican03121887-168.pdf.

Anonymous:1887:NTB

[8823] Anonymous. New torpedo boat for the Spanish Government. Scientific American, 56(11):169, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n11/pdf/scientificamerican03121887-
169.pdf.

Anonymous:1887:EIc

[8824] Anonymous. Engineering invention. Scientific American, 56(11):170, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n11/pdf/scientificamerican03121887-170.pdf.

Anonymous:1887:AIc

[8825] Anonymous. Agricultural inventions. Scientific American, 56(11):170, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n11/pdf/scientificamerican03121887-170a.pdf.

Anonymous:1887:MId

[8826] Anonymous. Miscellaneous inventions. Scientific American, 56(11):170, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n11/pdf/scientificamerican03121887-170b.pdf.

Anonymous:1887:Id

[8827] Anonymous. To inventors. Scientific American, 56(11):171, March 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n11/pdf/scientificamerican03121887-171.pdf.

Anonymous:1887:Cf

[8828] Anonymous. Correspondence. Scientific American, 56(12):181, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n12/pdf/scientificamerican03191887-181.pdf.

Anonymous:1887:BNBb

[8829] Anonymous. Books: New books and publications. Scientific American, 56(12):187, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n12/pdf/scientificamerican03191887-
187b.pdf.

Anonymous:1887:BPd

[8830] Anonymous. Business and personal. Scientific American, 56(12):187, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n12/pdf/scientificamerican03191887-187c.pdf.

Anonymous:1887:NQc

[8831] Anonymous. Notes and queries. Scientific American, 56(12):187, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n12/pdf/scientificamerican03191887-187d.pdf.

Anonymous:1887:IId

[8832] Anonymous. Index of inventions. Scientific American, 56(12):187–188, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n12/pdf/scientificamerican03191887-187e.pdf.

Anonymous:1887:PNPc

[8833] Anonymous. Patent notes: Patents. Scientific American, 56(12):190, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n12/pdf/scientificamerican03191887-190.pdf.

Anonymous:1887:SYD

[8834] Anonymous. A snow yacht on a Dakota prairie. Scientific American, 56(12):175, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n12/pdf/scientificamerican03191887-
175.pdf.

Anonymous:1887:SBA

[8835] Anonymous. The sale of the British Admiralty plans. Scientific American, 56(12):176, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n12/pdf/scientificamerican03191887-
176.pdf.

Anonymous:1887:JBEa

[8836] Anonymous. James Buchanan Eads, C.E. Scientific American, 56(12):176, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n12/pdf/scientificamerican03191887-
176a.pdf.

Anonymous:1887:STC

[8837] Anonymous. The state tax on commercial travelers unconstitutional. Scientific American, 56(12):176, March 19, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v56/
n12/pdf/scientificamerican03191887-176b.pdf.

Anonymous:1887:HWB

[8838] Anonymous. Henry Ward Beecher. Scientific American, 56(12):177, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n12/pdf/scientificamerican03191887-177.pdf.

Anonymous:1887:AN

[8839] Anonymous. Astronomical notes. Scientific American, 56(12):177, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n12/pdf/scientificamerican03191887-177a.pdf.

Anonymous:1887:MP

[8840] Anonymous. The Morgue at Paris. Scientific American, 56(12):178–179, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n12/pdf/scientificamerican03191887-178.pdf.

Anonymous:1887:EGA

[8841] Anonymous. Electrical gas alarm. Scientific American, 56(12):179, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n12/pdf/scientificamerican03191887-179.pdf.

Anonymous:1887:ACS

[8842] Anonymous. Attachment for cooking stoves. Scientific American, 56(12):179, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n12/pdf/scientificamerican03191887-
179a.pdf.

Anonymous:1887:FWH

[8843] Anonymous. Feed water heater. Scientific American, 56(12):179, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n12/pdf/scientificamerican03191887-179b.pdf.

Anonymous:1887:DPF

[8844] Anonymous. Device for placing fusible plugs in boilers. Scientific American, 56(12):179–180, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n12/pdf/scientificamerican03191887-
179c.pdf.

Anonymous:1887:RNE

[8845] Anonymous. Roburite — a new explosive. Scientific American, 56(12):180, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n12/pdf/scientificamerican03191887-
180.pdf.

Anonymous:1887:GAM

[8846] Anonymous. The German Army magazine gun. Scientific American, 56(12):182–183, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n12/pdf/scientificamerican03191887-
182.pdf.

Anonymous:1887:SHN

[8847] Anonymous. Suburban house of a New York artist. Scientific American, 56(12):183–184, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n12/pdf/scientificamerican03191887-
183.pdf.

Anonymous:1887:SES

[8848] Anonymous. Self-extinguishing stove for railway cars. Scientific American, 56(12):184–185, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n12/pdf/scientificamerican03191887-
184.pdf.

Anonymous:1887:ILH

[8849] Anonymous. Improved label holder. Scientific American, 56(12):185, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n12/pdf/scientificamerican03191887-185.pdf.

Anonymous:1887:MPI

[8850] Anonymous. The multum-in-parvo iron. Scientific American, 56 (12):185, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n12/pdf/scientificamerican03191887-
185a.pdf.

Anonymous:1887:IPe

[8851] Anonymous. Improved plow. Scientific American, 56(12):185–186, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n12/pdf/scientificamerican03191887-185b.pdf.

Anonymous:1887:EId

[8852] Anonymous. Engineering inventions. Scientific American, 56(12):187, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n12/pdf/scientificamerican03191887-187.pdf.

Anonymous:1887:MIe

[8853] Anonymous. Miscellaneous inventions. Scientific American, 56(12):187, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n12/pdf/scientificamerican03191887-187a.pdf.

Anonymous:1887:PBM

[8854] Anonymous. Patents in Brazil and Mexico. Scientific American, 56(12):189, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n12/pdf/scientificamerican03191887-
189.pdf.

Anonymous:1887:FPa

[8855] Anonymous. Foreign patents. Scientific American, 56(12):189, March 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n12/pdf/scientificamerican03191887-189a.pdf.

Anonymous:1887:Cg

[8856] Anonymous. Correspondence. Scientific American, 56(13):197, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n13/pdf/scientificamerican03261887-197.pdf.

Anonymous:1887:BNBc

[8857] Anonymous. Books: New books and publications. Scientific American, 56(13):202, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n13/pdf/scientificamerican03261887-
202c.pdf.

Anonymous:1887:BPe

[8858] Anonymous. Business and personal. Scientific American, 56(13):202, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n13/pdf/scientificamerican03261887-202d.pdf.

Anonymous:1887:NQd

[8859] Anonymous. Notes and queries. Scientific American, 56(13):203, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n13/pdf/scientificamerican03261887-203.pdf.

Anonymous:1887:IIe

[8860] Anonymous. Index of inventions. Scientific American, 56(13):203–204, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n13/pdf/scientificamerican03261887-203b.pdf.

Anonymous:1887:CAPa

[8861] Anonymous. Cruiser armed with pneumatic dynamite guns. Scientific American, 56(13):191, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n13/pdf/scientificamerican03261887-
191.pdf.

Anonymous:1887:SBN

[8862] Anonymous. The submarine boat Nautilus. Scientific American, 56(13):191–192, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n13/pdf/scientificamerican03261887-
191a.pdf.

Anonymous:1887:DNY

[8863] Anonymous. The defense of New York Harbor within thirty days. Scientific American, 56(13):192, March 26, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v56/
n13/pdf/scientificamerican03261887-192.pdf.

Anonymous:1887:STa

[8864] Anonymous. Sea telephony. Scientific American, 56(13):192–193, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n13/pdf/scientificamerican03261887-192a.pdf.

Anonymous:1887:SAA

[8865] Anonymous. The Scientific American — architects’ and builders’ edition. Scientific American, 56(13):193, March 26, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v56/
n13/pdf/scientificamerican03261887-193.pdf.

Anonymous:1887:SPCa

[8866] Anonymous. Safety platform for cars. Scientific American, 56(13):194, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n13/pdf/scientificamerican03261887-194.pdf.

Anonymous:1887:SER

[8867] Anonymous. Symington’s early road locomotives. Scientific American, 56(13):194, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n13/pdf/scientificamerican03261887-
194a.pdf.

Anonymous:1887:CR

[8868] Anonymous. Clothes rack. Scientific American, 56(13):194, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n13/pdf/scientificamerican03261887-194b.pdf.

Anonymous:1887:ISBa

[8869] Anonymous. Improved stuffing box. Scientific American, 56(13):194, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n13/pdf/scientificamerican03261887-194c.pdf.

Anonymous:1887:AH

[8870] Anonymous. Adjustable harrow. Scientific American, 56(13):195, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n13/pdf/scientificamerican03261887-195.pdf.

Anonymous:1887:IM

[8871] Anonymous. Improved mower. Scientific American, 56(13):195, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n13/pdf/scientificamerican03261887-195a.pdf.

Anonymous:1887:ICPa

[8872] Anonymous. Improved corn planter. Scientific American, 56(13):195, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n13/pdf/scientificamerican03261887-195b.pdf.

Anonymous:1887:MSa

[8873] Anonymous. Milking stool. Scientific American, 56(13):195, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n13/pdf/scientificamerican03261887-195c.pdf.

Anonymous:1887:HBS

[8874] Anonymous. Heels for boots and shoes. Scientific American, 56(13):196, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n13/pdf/scientificamerican03261887-196.pdf.

Anonymous:1887:NHI

[8875] Anonymous. Notebook holder and indicator. Scientific American, 56(13):196, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n13/pdf/scientificamerican03261887-
196a.pdf.

Anonymous:1887:IGFa

[8876] Anonymous. Improved grate-fire blower. Scientific American, 56 (13):196, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n13/pdf/scientificamerican03261887-
196b.pdf.

Anonymous:1887:STb

[8877] Anonymous. Science in toys. Scientific American, 56(13):198–199, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n13/pdf/scientificamerican03261887-198.pdf.

Anonymous:1887:RA

[8878] Anonymous. Rougerie’s anemogene. Scientific American, 56(13): 200–201, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n13/pdf/scientificamerican03261887-
200.pdf.

Anonymous:1887:PIa

[8879] Anonymous. The principle of inertia. Scientific American, 56(13):201, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n13/pdf/scientificamerican03261887-201.pdf.

Anonymous:1887:ER

[8880] Anonymous. Earthquake on the Riviera. Scientific American, 56(13):201, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n13/pdf/scientificamerican03261887-
201a.pdf.

Anonymous:1887:EIe

[8881] Anonymous. Engineering inventions. Scientific American, 56(13):202, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n13/pdf/scientificamerican03261887-202.pdf.

Anonymous:1887:AId

[8882] Anonymous. Agricultural invention. Scientific American, 56(13):202, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n13/pdf/scientificamerican03261887-202a.pdf.

Anonymous:1887:MIf

[8883] Anonymous. Miscellaneous inventions. Scientific American, 56(13):202, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n13/pdf/scientificamerican03261887-202b.pdf.

Anonymous:1887:Ie

[8884] Anonymous. To inventors. Scientific American, 56(13):203, March 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n13/pdf/scientificamerican03261887-203a.pdf.

Anonymous:1887:Ch

[8885] Anonymous. Correspondence. Scientific American, 56(14):213, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n14/pdf/scientificamerican04021887-213a.pdf.

Anonymous:1887:BPf

[8886] Anonymous. Business and personal. Scientific American, 56(14):218, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n14/pdf/scientificamerican04021887-218c.pdf.

Anonymous:1887:NQe

[8887] Anonymous. Notes and queries. Scientific American, 56(14):218–219, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n14/pdf/scientificamerican04021887-218d.pdf.

Anonymous:1887:IIf

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n14/pdf/scientificamerican04021887-219a.pdf.

Anonymous:1887:PNPd

[8889] Anonymous. Patent notes: Patents. Scientific American, 56(14):222, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n14/pdf/scientificamerican04021887-222.pdf.

Anonymous:1887:SFSa

[8890] Anonymous. Sham fighting shams. Scientific American, 56(14):208, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n14/pdf/scientificamerican04021887-208.pdf.

Anonymous:1887:NGN

[8891] Anonymous. The new guns for the Navy. Scientific American, 56(14):208–209, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n14/pdf/scientificamerican04021887-
208a.pdf.

Anonymous:1887:IAR

[8892] Anonymous. Improved account register. Scientific American, 56 (14):210, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n14/pdf/scientificamerican04021887-
210.pdf.

Anonymous:1887:INL

[8893] Anonymous. Improved number and letter plate. Scientific American, 56(14):210, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n14/pdf/scientificamerican04021887-
210a.pdf.

Anonymous:1887:ETR

[8894] Anonymous. Extension table for railway cars. Scientific American, 56(14):210, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n14/pdf/scientificamerican04021887-
210b.pdf.

Anonymous:1887:HSCb

[8895] Anonymous. Hair spring collet. Scientific American, 56(14):210, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n14/pdf/scientificamerican04021887-210c.pdf.

Anonymous:1887:BAa

[8896] Anonymous. Bottle attachment. Scientific American, 56(14):210, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n14/pdf/scientificamerican04021887-210d.pdf.

Anonymous:1887:ISa

[8897] Anonymous. Improved stove. Scientific American, 56(14):211, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n14/pdf/scientificamerican04021887-211.pdf.

Anonymous:1887:SIC

[8898] Anonymous. Station indicator for cars. Scientific American, 56(14):211, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n14/pdf/scientificamerican04021887-211a.pdf.

Anonymous:1887:ICCa

[8899] Anonymous. Improved car coupling. Scientific American, 56(14):211, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n14/pdf/scientificamerican04021887-211b.pdf.

Anonymous:1887:BC

[8900] Anonymous. Boiler cleaner. Scientific American, 56(14):211, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n14/pdf/scientificamerican04021887-211c.pdf.

Anonymous:1887:WBF

[8901] Anonymous. Wash boiler fountain. Scientific American, 56(14):211, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n14/pdf/scientificamerican04021887-211d.pdf.

Anonymous:1887:DRPb

[8902] Anonymous. Decisions relating to patents. Scientific American, 56(14):212, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n14/pdf/scientificamerican04021887-
212.pdf.

Anonymous:1887:ISHa

[8903] Anonymous. Improved shaft holder. Scientific American, 56(14):212, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n14/pdf/scientificamerican04021887-212a.pdf.

Anonymous:1887:IBM

[8904] Anonymous. Improved boring machine. Scientific American, 56 (14):212, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n14/pdf/scientificamerican04021887-
212b.pdf.

Anonymous:1887:YRA

[8905] Anonymous. Yacht race across the Atlantic. Scientific American, 56(14):213, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n14/pdf/scientificamerican04021887-
213.pdf.

Anonymous:1887:JEB

[8906] Anonymous. The Jensen electric bells. Scientific American, 56(14):214, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n14/pdf/scientificamerican04021887-214.pdf.

Anonymous:1887:MOF

[8907] Anonymous. Moulding an old fashioned knob. Scientific American, 56(14):214, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n14/pdf/scientificamerican04021887-
214a.pdf.

Anonymous:1887:PLA

[8908] Anonymous. Plague of locusts in Australia. Scientific American, 56(14):215, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n14/pdf/scientificamerican04021887-
215.pdf.

Anonymous:1887:TRB

[8909] Anonymous. Three related breeds of dogs. Scientific American, 56(14):215–216, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n14/pdf/scientificamerican04021887-
215a.pdf.

Anonymous:1887:BFR

[8910] Anonymous. Boiler feed regulator. Scientific American, 56(14):216, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n14/pdf/scientificamerican04021887-216.pdf.

Anonymous:1887:PCW

[8911] Anonymous. The Panama Canal works. Scientific American, 56 (14):217, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n14/pdf/scientificamerican04021887-
217.pdf.

Anonymous:1887:HAF

[8912] Anonymous. Hot air furnace. Scientific American, 56(14):217, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n14/pdf/scientificamerican04021887-217a.pdf.

Anonymous:1887:EIf

[8913] Anonymous. Engineering inventions. Scientific American, 56(14):218, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n14/pdf/scientificamerican04021887-218.pdf.

Anonymous:1887:AIe

[8914] Anonymous. Agricultural inventions. Scientific American, 56(14):218, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n14/pdf/scientificamerican04021887-218a.pdf.

Anonymous:1887:MIg

[8915] Anonymous. Miscellaneous inventions. Scientific American, 56(14):218, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n14/pdf/scientificamerican04021887-218b.pdf.

Anonymous:1887:If

[8916] Anonymous. To inventors. Scientific American, 56(14):219, April 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n14/pdf/scientificamerican04021887-219.pdf.

Anonymous:1887:Ci

[8917] Anonymous. Correspondence. Scientific American, 56(15):229–230, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n15/pdf/scientificamerican04091887-229.pdf.

Anonymous:1887:BNBd

[8918] Anonymous. Books: New books and publications. Scientific American, 56(15):234, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n15/pdf/scientificamerican04091887-
234c.pdf.

Anonymous:1887:BPg

[8919] Anonymous. Business and personal. Scientific American, 56(15):234, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n15/pdf/scientificamerican04091887-234d.pdf.

Anonymous:1887:NQf

[8920] Anonymous. Notes and queries. Scientific American, 56(15):235, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n15/pdf/scientificamerican04091887-235.pdf.

Anonymous:1887:IIg

[8921] Anonymous. Index of inventions. Scientific American, 56(15):235–236, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n15/pdf/scientificamerican04091887-235b.pdf.

Anonymous:1887:PNPe

[8922] Anonymous. Patent notes: Patents. Scientific American, 56(15):238, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n15/pdf/scientificamerican04091887-238.pdf.

Anonymous:1887:HMM

[8923] Anonymous. H.M.S. Mersey. Scientific American, 56(15):223, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n15/pdf/scientificamerican04091887-223.pdf.

Anonymous:1887:CLE

[8924] Anonymous. Car lighting by electricity. Scientific American, 56(15):224, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n15/pdf/scientificamerican04091887-224.pdf.

Anonymous:1887:HRC

[8925] Anonymous. Heating of railway cars by steam. Scientific American, 56(15):224–225, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n15/pdf/scientificamerican04091887-
224a.pdf.

Anonymous:1887:ICMa

[8926] Anonymous. Improved cake mixer. Scientific American, 56(15):226, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n15/pdf/scientificamerican04091887-226.pdf.

Anonymous:1887:SBL

[8927] Anonymous. Swivel bogie locomotive. Scientific American, 56(15):226, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n15/pdf/scientificamerican04091887-226a.pdf.

Anonymous:1887:SH

[8928] Anonymous. Spool holder. Scientific American, 56(15):226, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n15/pdf/scientificamerican04091887-226b.pdf.

Anonymous:1887:IFF

[8929] Anonymous. Improved flood fence. Scientific American, 56(15):226, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n15/pdf/scientificamerican04091887-226c.pdf.

Anonymous:1887:CCNb

[8930] Anonymous. Combined collar and necktie. Scientific American, 56(15):227, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n15/pdf/scientificamerican04091887-
227.pdf.

Anonymous:1887:CPD

[8931] Anonymous. Crosshead pushing device. Scientific American, 56 (15):227, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n15/pdf/scientificamerican04091887-
227a.pdf.

Anonymous:1887:ICCb

[8932] Anonymous. Improved car coupling. Scientific American, 56(15):227, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n15/pdf/scientificamerican04091887-227b.pdf.

Anonymous:1887:IDCa

[8933] Anonymous. Improved dumping cart. Scientific American, 56(15):227, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n15/pdf/scientificamerican04091887-227c.pdf.

Anonymous:1887:HT

[8934] Anonymous. Hat holder for travelers. Scientific American, 56(15):228, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n15/pdf/scientificamerican04091887-228.pdf.

Anonymous:1887:NWV

[8935] Anonymous. A novel weather vane. Scientific American, 56 (15):228–229, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n15/pdf/scientificamerican04091887-
228a.pdf.

Anonymous:1887:GSB

[8936] Anonymous. Gold stamper battery. Scientific American, 56(15): 230–231, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n15/pdf/scientificamerican04091887-
230.pdf.

Anonymous:1887:NMP

[8937] Anonymous. Novel method of protecting vessels against torpedo attacks. Scientific American, 56(15):231, April 9, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v56/
n15/pdf/scientificamerican04091887-231.pdf.

Anonymous:1887:SHR

[8938] Anonymous. Self-heating reversible sad iron. Scientific American, 56(15):232, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n15/pdf/scientificamerican04091887-
232.pdf.

Anonymous:1887:FBMa

[8939] Anonymous. A flying Batrachian of Malaisia. Scientific American, 56(15):233, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n15/pdf/scientificamerican04091887-
233.pdf.

Anonymous:1887:EIg

[8940] Anonymous. Engineering inventions. Scientific American, 56(15):234, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n15/pdf/scientificamerican04091887-234.pdf.

Anonymous:1887:AIf

[8941] Anonymous. Agricultural inventions. Scientific American, 56(15):234, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n15/pdf/scientificamerican04091887-234a.pdf.

Anonymous:1887:MIh

[8942] Anonymous. Miscellaneous inventions. Scientific American, 56(15):234, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n15/pdf/scientificamerican04091887-234b.pdf.

Anonymous:1887:Ig

[8943] Anonymous. To inventors. Scientific American, 56(15):235, April 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n15/pdf/scientificamerican04091887-235a.pdf.

Anonymous:1887:BNBe

[8944] Anonymous. Books: New books and publications. Scientific American, 56(16):250, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n16/pdf/scientificamerican04161887-
250c.pdf.

Anonymous:1887:BPh

[8945] Anonymous. Business and personal. Scientific American, 56(16):250, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n16/pdf/scientificamerican04161887-250d.pdf.

Anonymous:1887:NQg

[8946] Anonymous. Notes and queries. Scientific American, 56(16):251, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n16/pdf/scientificamerican04161887-251.pdf.

Anonymous:1887:IIh

[8947] Anonymous. Index of inventions. Scientific American, 56(16):251–252, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n16/pdf/scientificamerican04161887-251b.pdf.

Anonymous:1887:PNPf

[8948] Anonymous. Patent notes: Patents. Scientific American, 56(16):254, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n16/pdf/scientificamerican04161887-254.pdf.

Anonymous:1887:NCP

[8949] Anonymous. The new Commissioner of Patents. Scientific American, 56(16):240, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n16/pdf/scientificamerican04161887-
240.pdf.

Anonymous:1887:JM

[8950] Anonymous. Judge Montgomery. Scientific American, 56(16):240, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n16/pdf/scientificamerican04161887-240a.pdf.

Hovey:1887:NGI

[8951] H. C. Hovey. Natural gas in Indiana. Scientific American, 56 (16):240–241, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n16/pdf/scientificamerican04161887-
240b.pdf.

Anonymous:1887:WID

[8952] Anonymous. Warner’s improved dry plate holder. Scientific American, 56(16):242, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n16/pdf/scientificamerican04161887-
242.pdf.

Anonymous:1887:CUCa

[8953] Anonymous. Conduit for underground conductors. Scientific American, 56(16):242, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n16/pdf/scientificamerican04161887-
242a.pdf.

Anonymous:1887:FWR

[8954] Anonymous. Feed water regulator. Scientific American, 56(16):242–243, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n16/pdf/scientificamerican04161887-242b.pdf.

Anonymous:1887:NPE

[8955] Anonymous. A novel picture exhibitor. Scientific American, 56(16):243, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n16/pdf/scientificamerican04161887-243.pdf.

Anonymous:1887:TLM

[8956] Anonymous. Take-up and let-off mechanism for looms. Scientific American, 56(16):243, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n16/pdf/scientificamerican04161887-
243a.pdf.

Anonymous:1887:ICCc

[8957] Anonymous. Improved car coupling. Scientific American, 56(16):243, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n16/pdf/scientificamerican04161887-243b.pdf.

Anonymous:1887:ISCa

[8958] Anonymous. Improved saddletree and check-hook. Scientific American, 56(16):243, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n16/pdf/scientificamerican04161887-
243c.pdf.

Anonymous:1887:SFN

[8959] Anonymous. Sinking the foundation for the new Harlem River Bridge. Scientific American, 56(16):244, April 16, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v56/
n16/pdf/scientificamerican04161887-244.pdf.

Anonymous:1887:NTG

[8960] Anonymous. The new 110-ton gun. Scientific American, 56(16):245, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n16/pdf/scientificamerican04161887-245.pdf.

Anonymous:1887:CTS

[8961] Anonymous. Combination tool for squaring, leveling, etc. Scientific American, 56(16):245, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n16/pdf/scientificamerican04161887-
245a.pdf.

Anonymous:1887:NGB

[8962] Anonymous. A novel game board. Scientific American, 56(16):245–246, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n16/pdf/scientificamerican04161887-245b.pdf.

Anonymous:1887:STd

[8963] Anonymous. Science in toys. Scientific American, 56(16):246–248, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n16/pdf/scientificamerican04161887-246.pdf.

Anonymous:1887:DH

[8964] Anonymous. The decoration of a home. Scientific American, 56(16):248, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n16/pdf/scientificamerican04161887-248.pdf.

Anonymous:1887:BED

[8965] Anonymous. Braun’s electro-dynamic air ship. Scientific American, 56(16):249, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n16/pdf/scientificamerican04161887-
249.pdf.

Anonymous:1887:EIh

[8966] Anonymous. Engineering inventions. Scientific American, 56(16):250, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n16/pdf/scientificamerican04161887-250.pdf.

Anonymous:1887:AIg

[8967] Anonymous. Agricultural inventions. Scientific American, 56(16):250, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n16/pdf/scientificamerican04161887-250a.pdf.

Anonymous:1887:MIi

[8968] Anonymous. Miscellaneous inventions. Scientific American, 56(16):250, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n16/pdf/scientificamerican04161887-250b.pdf.

Anonymous:1887:Ih

[8969] Anonymous. To inventors. Scientific American, 56(16):251, April 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n16/pdf/scientificamerican04161887-251a.pdf.

Anonymous:1887:Cj

[8970] Anonymous. Correspondence. Scientific American, 56(17):261, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n17/pdf/scientificamerican04231887-261.pdf.

Anonymous:1887:BNBf

[8971] Anonymous. Books: New books and publications. Scientific American, 56(17):266, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n17/pdf/scientificamerican04231887-
266c.pdf.

Anonymous:1887:BPi

[8972] Anonymous. Business and personal. Scientific American, 56(17):266, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n17/pdf/scientificamerican04231887-266d.pdf.

Anonymous:1887:NQh

[8973] Anonymous. Notes and queries. Scientific American, 56(17):266–267, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n17/pdf/scientificamerican04231887-266e.pdf.

Anonymous:1887:IIi

[8974] Anonymous. Index of inventions. Scientific American, 56(17):267–268, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n17/pdf/scientificamerican04231887-267a.pdf.

Anonymous:1887:PNPg

[8975] Anonymous. Patent notes: Patents. Scientific American, 56(17):270, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n17/pdf/scientificamerican04231887-270.pdf.

Hopkins:1887:ELDa

[8976] M. Hopkins. Eight light dynamo. Scientific American, 56(17):255, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n17/pdf/scientificamerican04231887-255.pdf.

Anonymous:1887:NPP

[8977] Anonymous. New, process for the protection of iron. Scientific American, 56(17):256, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n17/pdf/scientificamerican04231887-
256.pdf.

Anonymous:1887:SLC

[8978] Anonymous. Sea lions in Central Park. Scientific American, 56(17):256, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n17/pdf/scientificamerican04231887-256a.pdf.

Anonymous:1887:DDa

[8979] Anonymous. Daniel davis. Scientific American, 56(17):256, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n17/pdf/scientificamerican04231887-256b.pdf.

Anonymous:1887:CWa

[8980] Anonymous. The celestial world. Scientific American, 56(17):256–257, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n17/pdf/scientificamerican04231887-256c.pdf.

Anonymous:1887:DUW

[8981] Anonymous. The designs for U.S. war vessels. Scientific American, 56(17):257, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n17/pdf/scientificamerican04231887-
257.pdf.

Anonymous:1887:IWR

[8982] Anonymous. Improved wagon rack. Scientific American, 56(17):258, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n17/pdf/scientificamerican04231887-258.pdf.

Anonymous:1887:ISd

[8983] Anonymous. Improved stove. Scientific American, 56(17):258, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n17/pdf/scientificamerican04231887-258a.pdf.

Anonymous:1887:DRPc

[8984] Anonymous. Decisions relating to patents. Scientific American, 56(17):258, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n17/pdf/scientificamerican04231887-
258b.pdf.

Anonymous:1887:IPL

[8985] Anonymous. Improved pan lifter. Scientific American, 56(17):258, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n17/pdf/scientificamerican04231887-258c.pdf.

Anonymous:1887:IPHa

[8986] Anonymous. Improved pen holder tip. Scientific American, 56(17):258, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n17/pdf/scientificamerican04231887-258d.pdf.

Anonymous:1887:CTP

[8987] Anonymous. Combined tablet, paper weight, and blotter. Scientific American, 56(17):258, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n17/pdf/scientificamerican04231887-
258e.pdf.

Anonymous:1887:CHE

[8988] Anonymous. Connection hanger for electrotyping apparatus. Scientific American, 56(17):259, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n17/pdf/scientificamerican04231887-
259.pdf.

Anonymous:1887:ICCd

[8989] Anonymous. Improved cash carrier. Scientific American, 56(17):259, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n17/pdf/scientificamerican04231887-259a.pdf.

Anonymous:1887:LTP

[8990] Anonymous. Lewis tent pole and ventilator. Scientific American, 56(17):259, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n17/pdf/scientificamerican04231887-
259b.pdf.

Anonymous:1887:DAC

[8991] Anonymous. Double air chamber furring tile. Scientific American, 56(17):259, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n17/pdf/scientificamerican04231887-
259c.pdf.

Anonymous:1887:SSS

[8992] Anonymous. Safety set screw for collars and pulleys. Scientific American, 56(17):259–260, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n17/pdf/scientificamerican04231887-
259d.pdf.

Anonymous:1887:JBEb

[8993] Anonymous. James Buchanan Eads. Scientific American, 56 (17):263–264, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n17/pdf/scientificamerican04231887-
263.pdf.

Anonymous:1887:ICSa

[8994] Anonymous. Improved compound steam engine. Scientific American, 56(17):265, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n17/pdf/scientificamerican04231887-
265.pdf.

Anonymous:1887:IIPa

[8995] Anonymous. Improved iron planer. Scientific American, 56(17):265, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n17/pdf/scientificamerican04231887-265a.pdf.

Anonymous:1887:EIi

[8996] Anonymous. Engineering inventions. Scientific American, 56(17):266, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n17/pdf/scientificamerican04231887-266.pdf.

Anonymous:1887:AIh

[8997] Anonymous. Agricultural inventions. Scientific American, 56(17):266, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n17/pdf/scientificamerican04231887-266a.pdf.

Anonymous:1887:MIj

[8998] Anonymous. Miscellaneous inventions. Scientific American, 56(17):266, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n17/pdf/scientificamerican04231887-266b.pdf.

Anonymous:1887:Ij

[8999] Anonymous. To inventors. Scientific American, 56(17):267, April 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n17/pdf/scientificamerican04231887-267.pdf.

Anonymous:1887:Ck

[9000] Anonymous. Correspondence. Scientific American, 56(18):277, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n18/pdf/scientificamerican04301887-277.pdf.

Anonymous:1887:BPj

[9001] Anonymous. Business and personal. Scientific American, 56(18):282, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n18/pdf/scientificamerican04301887-282b.pdf.

Anonymous:1887:BNBg

[9002] Anonymous. Books: New books and publications. Scientific American, 56(18):282, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n18/pdf/scientificamerican04301887-
282c.pdf.

Anonymous:1887:NQi

[9003] Anonymous. Notes and queries. Scientific American, 56(18):282–283, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n18/pdf/scientificamerican04301887-282d.pdf.

Anonymous:1887:IIj

[9004] Anonymous. Index of inventions. Scientific American, 56(18):283–284, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n18/pdf/scientificamerican04301887-283a.pdf.

Anonymous:1887:PNPh

[9005] Anonymous. Patent notes: Patents. Scientific American, 56 (18):283–284, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n18/pdf/scientificamerican04301887-
286.pdf.

Anonymous:1887:FAT

[9006] Anonymous. The first American type writer. Scientific American, 56(18):271–276, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n18/pdf/scientificamerican04301887-
271.pdf.

Anonymous:1887:USJ

[9007] Anonymous. The United States joins the industrial union. Scientific American, 56(18):272, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n18/pdf/scientificamerican04301887-
272.pdf.

Anonymous:1887:AEP

[9008] Anonymous. Ant-eater and penguin. Scientific American, 56 (18):272–273, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n18/pdf/scientificamerican04301887-
272a.pdf.

Anonymous:1887:ICBa

[9009] Anonymous. Improved cutter bar for mowers and reapers. Scientific American, 56(18):274, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n18/pdf/scientificamerican04301887-
274.pdf.

Anonymous:1887:ICWa

[9010] Anonymous. Improved condenser for wool, cotton, cotton waste, etc. Scientific American, 56(18):274, April 30, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v56/
n18/pdf/scientificamerican04301887-274a.pdf.

Anonymous:1887:KAC

[9011] Anonymous. The Keystone automatic car coupler. Scientific American, 56(18):274, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n18/pdf/scientificamerican04301887-
274b.pdf.

Anonymous:1887:CLL

[9012] Anonymous. Combined latch and lock. Scientific American, 56 (18):274–275, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n18/pdf/scientificamerican04301887-
274c.pdf.

Anonymous:1887:IFa

[9013] Anonymous. Improved firearm. Scientific American, 56(18):275, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n18/pdf/scientificamerican04301887-275.pdf.

Anonymous:1887:DSE

[9014] Anonymous. Device for securing envelopes, etc. Scientific American, 56(18):275, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n18/pdf/scientificamerican04301887-
275a.pdf.

Anonymous:1887:ISPa

[9015] Anonymous. Improved stove pipe collar. Scientific American, 56(18):275, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n18/pdf/scientificamerican04301887-
275b.pdf.

Anonymous:1887:IBJ

[9016] Anonymous. Improved butter jar. Scientific American, 56(18):275, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n18/pdf/scientificamerican04301887-275c.pdf.

Anonymous:1887:CCM

[9017] Anonymous. Combined cigarette and match box. Scientific American, 56(18):275, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n18/pdf/scientificamerican04301887-
275d.pdf.

Anonymous:1887:IOCa

[9018] Anonymous. Improved ore crusher and amalgamator. Scientific American, 56(18):276, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n18/pdf/scientificamerican04301887-
276.pdf.

Anonymous:1887:IPS

[9019] Anonymous. Improved peach stoner. Scientific American, 56(18):276, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n18/pdf/scientificamerican04301887-276a.pdf.

Hopkins:1887:ELDb

[9020] George M. Hopkins. Eight light dynamo. Scientific American, 56 (18):278–279, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n18/pdf/scientificamerican04301887-
278.pdf.

Anonymous:1887:EP

[9021] Anonymous. Exercises in prestidigitation. Scientific American, 56(18):279, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n18/pdf/scientificamerican04301887-
279.pdf.

Anonymous:1887:PBS

[9022] Anonymous. The polar bear and seals. Scientific American, 56 (18):279–280, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n18/pdf/scientificamerican04301887-
279a.pdf.

Anonymous:1887:CUCb

[9023] Anonymous. Conduit for underground conductors. Scientific American, 56(18):280, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n18/pdf/scientificamerican04301887-
280.pdf.

Sloane:1887:ADS

[9024] T. O’Conor Sloane. Apparatus for determining the specific gravity of liquids. Scientific American, 56(18):281, April 30, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v56/
n18/pdf/scientificamerican04301887-281.pdf.

Anonymous:1887:ECR

[9025] Anonymous. Elevated cable road with suspended car. Scientific American, 56(18):281, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n18/pdf/scientificamerican04301887-
281a.pdf.

Anonymous:1887:EIj

[9026] Anonymous. Engineering inventions. Scientific American, 56(18):282, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n18/pdf/scientificamerican04301887-282.pdf.

Anonymous:1887:MIk

[9027] Anonymous. Miscellaneous inventions. Scientific American, 56(18):282, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n18/pdf/scientificamerican04301887-282a.pdf.

Anonymous:1887:Ik

[9028] Anonymous. To inventors. Scientific American, 56(18):283, April 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n18/pdf/scientificamerican04301887-283.pdf.

Anonymous:1887:BPk

[9029] Anonymous. Business and personal. Scientific American, 56(19):299, May 7, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n19/pdf/scientificamerican05071887-299d.pdf.

Anonymous:1887:IIk

[9030] Anonymous. Index of inventions. Scientific American, 56(19):299–300, May 7, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n19/pdf/scientificamerican05071887-299e.pdf.

Anonymous:1887:PNPi

[9031] Anonymous. Patent notes: Patents. Scientific American, 56(19):302, May 7, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n19/pdf/scientificamerican05071887-302.pdf.

Anonymous:1887:JSL

[9032] Anonymous. The Julien system of lighting railroad cars. Scientific American, 56(19):287, May 7, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n19/pdf/scientificamerican05071887-
287.pdf.

Anonymous:1887:FEW

[9033] Anonymous. French export wine. Scientific American, 56(19):288, May 7, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n19/pdf/scientificamerican05071887-288.pdf.

Anonymous:1887:PCD

[9034] Anonymous. Panama Canal difficulties. Scientific American, 56 (19):288–289, May 7, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n19/pdf/scientificamerican05071887-
288a.pdf.

Anonymous:1887:IDSa

[9035] Anonymous. Improved drag saw. Scientific American, 56(19):290, May 7, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n19/pdf/scientificamerican05071887-290.pdf.

Anonymous:1887:LNM

[9036] Anonymous. Lovell’s nail making machine. Scientific American, 56(19):290, May 7, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n19/pdf/scientificamerican05071887-
290a.pdf.

Anonymous:1887:IVBa

[9037] Anonymous. Improved vehicle brake. Scientific American, 56(19):290, May 7, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n19/pdf/scientificamerican05071887-290b.pdf.

Anonymous:1887:ICPb

[9038] Anonymous. Improved coffee pot. Scientific American, 56(19):291, May 7, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n19/pdf/scientificamerican05071887-291.pdf.

Anonymous:1887:IOa

[9039] Anonymous. Improved oiler. Scientific American, 56(19):291, May 7, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n19/pdf/scientificamerican05071887-291a.pdf.

Anonymous:1887:FCP

[9040] Anonymous. Friction clutch pulley. Scientific American, 56(19):291, May 7, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n19/pdf/scientificamerican05071887-291b.pdf.

Anonymous:1887:SFC

[9041] Anonymous. A simple floor clamp. Scientific American, 56(19):291, May 7, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n19/pdf/scientificamerican05071887-291c.pdf.

Anonymous:1887:ALC

[9042] Anonymous. Adjustable lambrequin, curtain, and shade support. Scientific American, 56(19):291–292, May 7, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v56/
n19/pdf/scientificamerican05071887-291d.pdf.

Anonymous:1887:CEa

[9043] Anonymous. A convenient escritoire. Scientific American, 56(19):292, May 7, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n19/pdf/scientificamerican05071887-292.pdf.

Anonymous:1887:SDE

[9044] Anonymous. Small dynamo electric machine. Scientific American, 56(19):292, May 7, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n19/pdf/scientificamerican05071887-
292a.pdf.

Sloane:1887:AIL

[9045] T. O’Conor Sloane. Apparatus for illustrating the law of the reflection of light. Scientific American, 56(19):293, May 7, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v56/
n19/pdf/scientificamerican05071887-293.pdf.

Anonymous:1887:NBDa

[9046] Anonymous. New bench drill. Scientific American, 56(19):293, May 7, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n19/pdf/scientificamerican05071887-293a.pdf.

Hopkins:1887:MN

[9047] George M. Hopkins. Microscopic notes. Scientific American, 56(19):295, May 7, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n19/pdf/scientificamerican05071887-295.pdf.

Anonymous:1887:NEE

[9048] Anonymous. A novel electrical experiment. Scientific American, 56(19):295–296, May 7, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n19/pdf/scientificamerican05071887-
295a.pdf.

Anonymous:1887:HMS

[9049] Anonymous. How to make statuary in paper. Scientific American, 56(19):297–298, May 7, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n19/pdf/scientificamerican05071887-
297.pdf.

Anonymous:1887:EIk

[9050] Anonymous. Engineering inventions. Scientific American, 56(19):299, May 7, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n19/pdf/scientificamerican05071887-299.pdf.

Anonymous:1887:MIl

[9051] Anonymous. Mechanical invention. Scientific American, 56(19):299, May 7, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n19/pdf/scientificamerican05071887-299a.pdf.

Anonymous:1887:AIi

[9052] Anonymous. Agricultural inventions. Scientific American, 56(19):299, May 7, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n19/pdf/scientificamerican05071887-299b.pdf.

Anonymous:1887:MIm

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n19/pdf/scientificamerican05071887-299c.pdf.

Anonymous:1887:BNBh

[9054] Anonymous. Books: New books and publications. Scientific American, 56(2):26, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n2/pdf/scientificamerican01081887-
26d.pdf.

Anonymous:1887:BPl

[9055] Anonymous. Business and personal. Scientific American, 56(2):27, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n2/pdf/scientificamerican01081887-27.pdf.

Anonymous:1887:NQj

[9056] Anonymous. Notes and queries. Scientific American, 56(2):27, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n2/pdf/scientificamerican01081887-27a.pdf.

Anonymous:1887:IIl

[9057] Anonymous. Index of inventions. Scientific American, 56(2):27–28, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n2/pdf/scientificamerican01081887-27c.pdf.

Anonymous:1887:PNPj

[9058] Anonymous. Patent notes: Patents. Scientific American, 56(2):30, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n2/pdf/scientificamerican01081887-30.pdf.

Anonymous:1887:AAU

[9059] Anonymous. The annual attack upon our Patent Laws. Scientific American, 56(2):16, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n2/pdf/scientificamerican01081887-
16.pdf.

Anonymous:1887:MCa

[9060] Anonymous. The multicharge cannon. Scientific American, 56(2):16, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n2/pdf/scientificamerican01081887-16a.pdf.

Anonymous:1887:RPN

[9061] Anonymous. Revival of patent nullification. Scientific American, 56(2):16–17, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n2/pdf/scientificamerican01081887-
16b.pdf.

Anonymous:1887:OBS

[9062] Anonymous. Our building stone supply. Scientific American, 56(2):18, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n2/pdf/scientificamerican01081887-18.pdf.

Anonymous:1887:DEI

[9063] Anonymous. Dynamo electric igniting machine. Scientific American, 56(2):19, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n2/pdf/scientificamerican01081887-
19.pdf.

Anonymous:1887:IHD

[9064] Anonymous. Improved hopper dredger. Scientific American, 56(2):19, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n2/pdf/scientificamerican01081887-19a.pdf.

Anonymous:1887:WUP

[9065] Anonymous. Waves utilized to pump water. Scientific American, 56(2):19, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n2/pdf/scientificamerican01081887-
19b.pdf.

Anonymous:1887:SLH

[9066] Anonymous. Safety loop for harness checks. Scientific American, 56(2):20, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n2/pdf/scientificamerican01081887-
20.pdf.

Anonymous:1887:FD

[9067] Anonymous. Fertilizer distributer. Scientific American, 56(2):20, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n2/pdf/scientificamerican01081887-20a.pdf.

Anonymous:1887:ADB

[9068] Anonymous. Ax with detachable blade. Scientific American, 56(2):20, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n2/pdf/scientificamerican01081887-20b.pdf.

Anonymous:1887:RR

[9069] Anonymous. Repairing rails. Scientific American, 56(2):20, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n2/pdf/scientificamerican01081887-20c.pdf.

Anonymous:1887:NSF

[9070] Anonymous. New steam fire engine. Scientific American, 56(2):22, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n2/pdf/scientificamerican01081887-22.pdf.

Anonymous:1887:GBZ

[9071] Anonymous. The gibbon in the Berlin Zoological Garden. Scientific American, 56(2):23, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n2/pdf/scientificamerican01081887-
23.pdf.

Anonymous:1887:NPMa

[9072] Anonymous. New processes in the manufacture of beet sugar. Scientific American, 56(2):23–24, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v56/n2/pdf/scientificamerican01081887-
23a.pdf.

Anonymous:1887:STe

[9073] Anonymous. Science in toys. Scientific American, 56(2):25, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n2/pdf/scientificamerican01081887-25.pdf.

Anonymous:1887:EIl

[9074] Anonymous. Engineering inventions. Scientific American, 56(2):26, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n2/pdf/scientificamerican01081887-26.pdf.

Anonymous:1887:MIn

[9075] Anonymous. Mechanical invention. Scientific American, 56(2):26, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n2/pdf/scientificamerican01081887-26a.pdf.

Anonymous:1887:AIj

[9076] Anonymous. Agricultural invention. Scientific American, 56(2):26, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n2/pdf/scientificamerican01081887-26b.pdf.

Anonymous:1887:MIo

[9077] Anonymous. Miscellaneous inventions. Scientific American, 56(2):26, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n2/pdf/scientificamerican01081887-26c.pdf.

Anonymous:1887:Il

[9078] Anonymous. To inventors. Scientific American, 56(2):27, January 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n2/pdf/scientificamerican01081887-27b.pdf.

Anonymous:1887:Cl

[9079] Anonymous. Correspondence. Scientific American, 56(20):309, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n20/pdf/scientificamerican05141887-309a.pdf.

Anonymous:1887:BNBi

[9080] Anonymous. Books: New books and publications. Scientific American, 56(20):314, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n20/pdf/scientificamerican05141887-
314c.pdf.

Anonymous:1887:BPm

[9081] Anonymous. Business and personal. Scientific American, 56(20):314, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n20/pdf/scientificamerican05141887-314d.pdf.

Anonymous:1887:NQk

[9082] Anonymous. Notes and queries. Scientific American, 56(20):315, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n20/pdf/scientificamerican05141887-315.pdf.

Anonymous:1887:IIm

[9083] Anonymous. Index of inventions. Scientific American, 56(20):315, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n20/pdf/scientificamerican05141887-315b.pdf.

Anonymous:1887:PNPk

[9084] Anonymous. Patent notes: Patents. Scientific American, 56(20):318, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n20/pdf/scientificamerican05141887-318.pdf.

Anonymous:1887:CBN

[9085] Anonymous. Colossus, of the British Navy. Scientific American, 56(20):303, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n20/pdf/scientificamerican05141887-
303.pdf.

Anonymous:1887:EA

[9086] Anonymous. Electric accumulators. Scientific American, 56(20):304, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n20/pdf/scientificamerican05141887-304.pdf.

Anonymous:1887:HGS

[9087] Anonymous. Headway of great ships. Scientific American, 56 (20):304–305, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n20/pdf/scientificamerican05141887-
304a.pdf.

Anonymous:1887:VG

[9088] Anonymous. Vegetable grater. Scientific American, 56(20):306, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n20/pdf/scientificamerican05141887-306.pdf.

Anonymous:1887:DOA

[9089] Anonymous. Device for operating awnings. Scientific American, 56(20):306, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n20/pdf/scientificamerican05141887-
306a.pdf.

Anonymous:1887:IHP

[9090] Anonymous. Improved harrow and pulverizer. Scientific American, 56(20):306, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n20/pdf/scientificamerican05141887-
306b.pdf.

Anonymous:1887:IHS

[9091] Anonymous. Improved horizontal stationary engine. Scientific American, 56(20):306, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n20/pdf/scientificamerican05141887-
306c.pdf.

Anonymous:1887:WB

[9092] Anonymous. Watch for the blind. Scientific American, 56(20):306, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n20/pdf/scientificamerican05141887-306d.pdf.

Anonymous:1887:ILPa

[9093] Anonymous. Improved lamp post. Scientific American, 56(20):307, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n20/pdf/scientificamerican05141887-307.pdf.

Anonymous:1887:FPM

[9094] Anonymous. Folding and pasting machines. Scientific American, 56(20):307, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n20/pdf/scientificamerican05141887-
307a.pdf.

Anonymous:1887:CAGa

[9095] Anonymous. Clearing attachment for grain drills. Scientific American, 56(20):307–309, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n20/pdf/scientificamerican05141887-
307b.pdf.

Anonymous:1887:PMMa

[9096] Anonymous. A perpetual motion machine — will it work? Scientific American, 56(20):309, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n20/pdf/scientificamerican05141887-
309.pdf.

Anonymous:1887:VTP

[9097] Anonymous. Violin tuning peg. Scientific American, 56(20):309, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n20/pdf/scientificamerican05141887-309b.pdf.

Anonymous:1887:STf

[9098] Anonymous. Steno-telegraphy. Scientific American, 56(20):310–311, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n20/pdf/scientificamerican05141887-310.pdf.

Anonymous:1887:PBZ

[9099] Anonymous. The porcupine in the Berlin Zoological Garden. Scientific American, 56(20):311–312, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v56/n20/pdf/scientificamerican05141887-
311.pdf.

Anonymous:1887:SGa

[9100] Anonymous. A simple galvanometer. Scientific American, 56(20):312, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n20/pdf/scientificamerican05141887-312.pdf.

Anonymous:1887:EM

[9101] Anonymous. Experiment in magnetism. Scientific American, 56 (20):312, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n20/pdf/scientificamerican05141887-
312a.pdf.

Anonymous:1887:MWB

[9102] Anonymous. The mechanical working of ballast quarries. Scientific American, 56(20):312–313, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v56/n20/pdf/scientificamerican05141887-
312b.pdf.

Anonymous:1887:EIm

[9103] Anonymous. Engineering inventions. Scientific American, 56(20):314, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n20/pdf/scientificamerican05141887-314.pdf.

Anonymous:1887:AIk

[9104] Anonymous. Agricultural inventions. Scientific American, 56(20):314, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n20/pdf/scientificamerican05141887-314a.pdf.

Anonymous:1887:MIp

[9105] Anonymous. Miscellaneous inventions. Scientific American, 56(20):314, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n20/pdf/scientificamerican05141887-314b.pdf.

Anonymous:1887:In

[9106] Anonymous. To inventors. Scientific American, 56(20):315, May 14, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n20/pdf/scientificamerican05141887-315a.pdf.

Anonymous:1887:BPn

[9107] Anonymous. Business and personal. Scientific American, 56 (21):330–331, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n21/pdf/scientificamerican05211887-
330d.pdf.

Anonymous:1887:BNBj

[9108] Anonymous. Books: New books and publications. Scientific American, 56(21):331, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n21/pdf/scientificamerican05211887-
331.pdf.

Anonymous:1887:NQl

[9109] Anonymous. Notes and queries. Scientific American, 56(21):331, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n21/pdf/scientificamerican05211887-331a.pdf.

Anonymous:1887:IIn

[9110] Anonymous. Index of inventions. Scientific American, 56(21):331–332, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n21/pdf/scientificamerican05211887-331c.pdf.

Anonymous:1887:PNPl

[9111] Anonymous. Patent notes: Patents. Scientific American, 56(21):334, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n21/pdf/scientificamerican05211887-334.pdf.

Anonymous:1887:MGM

[9112] Anonymous. The magazine gun as a military arm. Scientific American, 56(21):320, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n21/pdf/scientificamerican05211887-
320.pdf.

Anonymous:1887:CWc

[9113] Anonymous. The celestial world. Scientific American, 56(21):320–321, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n21/pdf/scientificamerican05211887-320a.pdf.

Anonymous:1887:SCA

[9114] Anonymous. Steam cruiser Atlanta, U.S.N. Scientific American, 56(21):321, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n21/pdf/scientificamerican05211887-
321.pdf.

Anonymous:1887:ISCb

[9115] Anonymous. Improved shuttle cushioner for power looks. Scientific American, 56(21):322, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n21/pdf/scientificamerican05211887-
322.pdf.

Anonymous:1887:BSM

[9116] Anonymous. Beet sugar manufacturing processes. Scientific American, 56(21):322, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n21/pdf/scientificamerican05211887-
322a.pdf.

Anonymous:1887:DDH

[9117] Anonymous. Device for detaching houses from vehicles. Scientific American, 56(21):323, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n21/pdf/scientificamerican05211887-
323.pdf.

Anonymous:1887:NSO

[9118] Anonymous. A night sight for ordnance. Scientific American, 56 (21):323, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n21/pdf/scientificamerican05211887-
323a.pdf.

Anonymous:1887:ICSb

[9119] Anonymous. Improved camera shutter. Scientific American, 56(21):323, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n21/pdf/scientificamerican05211887-323b.pdf.

Anonymous:1887:CAPb

[9120] Anonymous. Carrier for ash pans. Scientific American, 56(21):323, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n21/pdf/scientificamerican05211887-323c.pdf.

Anonymous:1887:HRT

[9121] Anonymous. The Hudson River Tunnel. Scientific American, 56 (21):324–325, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n21/pdf/scientificamerican05211887-
324.pdf.

Anonymous:1887:NAC

[9122] Anonymous. New armored cruiser for the Spanish Navy. Scientific American, 56(21):325, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n21/pdf/scientificamerican05211887-
325.pdf.

Anonymous:1887:SCM

[9123] Anonymous. A self-cleaning metal valve. Scientific American, 56 (21):325, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n21/pdf/scientificamerican05211887-
325a.pdf.

Anonymous:1887:LC

[9124] Anonymous. A large casting. Scientific American, 56(21):326, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n21/pdf/scientificamerican05211887-326.pdf.

Anonymous:1887:BTBa

[9125] Anonymous. The Berdan torpedo boat. Scientific American, 56(21):326, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n21/pdf/scientificamerican05211887-326a.pdf.

Benjamin:1887:JDD

[9126] Marcus Benjamin. James dwight dana. Scientific American, 56 (21):327–328, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n21/pdf/scientificamerican05211887-
327.pdf.

Anonymous:1887:STg

[9127] Anonymous. Science in toys. Scientific American, 56(21):328, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n21/pdf/scientificamerican05211887-328.pdf.

Hopkins:1887:PL

[9128] Geo. M. Hopkins. Polarized light. Scientific American, 56(21):329, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n21/pdf/scientificamerican05211887-329.pdf.

Anonymous:1887:EIn

[9129] Anonymous. Engineering inventions. Scientific American, 56(21):330, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n21/pdf/scientificamerican05211887-330.pdf.

Anonymous:1887:AIl

[9130] Anonymous. Agricultural invention. Scientific American, 56(21):330, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n21/pdf/scientificamerican05211887-330a.pdf.

Anonymous:1887:MIq

[9131] Anonymous. Miscellaneous inventions. Scientific American, 56(21):330, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n21/pdf/scientificamerican05211887-330b.pdf.

Anonymous:1887:Sa

[9132] Anonymous. Special. Scientific American, 56(21):330, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n21/pdf/scientificamerican05211887-330c.pdf.

Anonymous:1887:Iq

[9133] Anonymous. To inventors. Scientific American, 56(21):331, May 21, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n21/pdf/scientificamerican05211887-331b.pdf.

Anonymous:1887:Cm

[9134] Anonymous. Correspondence. Scientific American, 56(22):341, May 28, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n22/pdf/scientificamerican05281887-341.pdf.

Anonymous:1887:BPo

[9135] Anonymous. Business and personal. Scientific American, 56(22):347, May 28, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n22/pdf/scientificamerican05281887-347.pdf.

Anonymous:1887:BNBk

[9136] Anonymous. Books: New books and publications. Scientific American, 56(22):347, May 28, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n22/pdf/scientificamerican05281887-
347a.pdf.

Anonymous:1887:NQm

[9137] Anonymous. Notes and queries. Scientific American, 56(22):347, May 28, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n22/pdf/scientificamerican05281887-347b.pdf.

Anonymous:1887:IIo

[9138] Anonymous. Index of inventions. Scientific American, 56(22):347–348, May 28, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n22/pdf/scientificamerican05281887-347d.pdf.

Anonymous:1887:PNPm

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n22/pdf/scientificamerican05281887-350.pdf.

Anonymous:1887:NBDb

[9140] Anonymous. Newark bay draw bridge. Scientific American, 56(22):335, May 28, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n22/pdf/scientificamerican05281887-335.pdf.

Anonymous:1887:PSTa

[9141] Anonymous. The protection of shade trees against insects. Scientific American, 56(22):336, May 28, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n22/pdf/scientificamerican05281887-
336.pdf.

Anonymous:1887:PCS

[9142] Anonymous. Present condition of shipping and sailors. Scientific American, 56(22):336–337, May 28, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n22/pdf/scientificamerican05281887-
336a.pdf.

Anonymous:1887:IOL

[9143] Anonymous. Improved oar lock. Scientific American, 56(22):338, May 28, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n22/pdf/scientificamerican05281887-338.pdf.

Anonymous:1887:IPCa

[9144] Anonymous. Improved printer’s chase. Scientific American, 56(22):338, May 28, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n22/pdf/scientificamerican05281887-338a.pdf.

Anonymous:1887:ICCe

[9145] Anonymous. Improved car coupling. Scientific American, 56(22):338, May 28, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n22/pdf/scientificamerican05281887-338b.pdf.

Anonymous:1887:ICL

[9146] Anonymous. Improved combination lock. Scientific American, 56 (22):338, May 28, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n22/pdf/scientificamerican05281887-
338c.pdf.

Anonymous:1887:RJB

[9147] Anonymous. Rounding jack for hat brims. Scientific American, 56(22):338–339, May 28, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n22/pdf/scientificamerican05281887-
338d.pdf.

Anonymous:1887:AP

[9148] Anonymous. Animal poke. Scientific American, 56(22):339, May 28, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n22/pdf/scientificamerican05281887-339.pdf.

Anonymous:1887:CFD

[9149] Anonymous. Combined fertilizer dropper and hill former for tobacco plants. Scientific American, 56(22):339, May 28, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v56/
n22/pdf/scientificamerican05281887-339a.pdf.

Sloane:1887:ML

[9150] T. O’conor Sloane. The momentum of liquids. Scientific American, 56(22):339, May 28, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n22/pdf/scientificamerican05281887-
339b.pdf.

Anonymous:1887:WM

[9151] Anonymous. Water motor. Scientific American, 56(22):339, May 28, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n22/pdf/scientificamerican05281887-339c.pdf.

Anonymous:1887:APJ

[9152] Anonymous. Attraction produced by jets of steam or air. Scientific American, 56(22):342, May 28, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n22/pdf/scientificamerican05281887-
342.pdf.

Anonymous:1887:ISR

[9153] Anonymous. Improved steam road roller. Scientific American, 56 (22):342, May 28, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n22/pdf/scientificamerican05281887-
342a.pdf.

Anonymous:1887:THC

[9154] Anonymous. The two headed cow. Scientific American, 56(22):343–344, May 28, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n22/pdf/scientificamerican05281887-343.pdf.

Hopkins:1887:AES

[9155] Geo. M. Hopkins. Autographs of the electric spark. Scientific American, 56(22):345, May 28, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n22/pdf/scientificamerican05281887-
345.pdf.

Anonymous:1887:EIo

[9156] Anonymous. Engineering inventions. Scientific American, 56(22):346, May 28, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n22/pdf/scientificamerican05281887-346.pdf.

Anonymous:1887:MIr

[9157] Anonymous. Miscellaneous inventions. Scientific American, 56(22):346, May 28, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n22/pdf/scientificamerican05281887-346a.pdf.

Anonymous:1887:Ir

[9158] Anonymous. To inventors. Scientific American, 56(22):347, May 28, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n22/pdf/scientificamerican05281887-347c.pdf.

Anonymous:1887:Cn

[9159] Anonymous. Correspondence. Scientific American, 56(23):357, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n23/pdf/scientificamerican06041887-357.pdf.

Anonymous:1887:BPp

[9160] Anonymous. Business and personal. Scientific American, 56(23):362, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n23/pdf/scientificamerican06041887-362d.pdf.

Anonymous:1887:BNBl

[9161] Anonymous. Books: New books and publications. Scientific American, 56(23):362, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n23/pdf/scientificamerican06041887-
362e.pdf.

Anonymous:1887:NQn

[9162] Anonymous. Notes and queries. Scientific American, 56(23):363, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n23/pdf/scientificamerican06041887-363.pdf.

Anonymous:1887:IIq

[9163] Anonymous. Index of inventions. Scientific American, 56(23):363–364, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n23/pdf/scientificamerican06041887-363a.pdf.

Anonymous:1887:PNPn

[9164] Anonymous. Patent notes: Patents. Scientific American, 56(23):366, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n23/pdf/scientificamerican06041887-366.pdf.

Anonymous:1887:NYW

[9165] Anonymous. New York Water Supply. Scientific American, 56(23):351, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n23/pdf/scientificamerican06041887-351.pdf.

Anonymous:1887:DWPa

[9166] Anonymous. The driven well patent sustained. Scientific American, 56(23):352–353, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n23/pdf/scientificamerican06041887-
352.pdf.

Anonymous:1887:IGa

[9167] Anonymous. Improved gate. Scientific American, 56(23):354, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n23/pdf/scientificamerican06041887-354.pdf.

Anonymous:1887:DLC

[9168] Anonymous. Device for lowering caskets. Scientific American, 56(23):354, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n23/pdf/scientificamerican06041887-
354a.pdf.

Anonymous:1887:PSTb

[9169] Anonymous. Portable switch table and car replacer. Scientific American, 56(23):354, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n23/pdf/scientificamerican06041887-
354b.pdf.

Anonymous:1887:IHR

[9170] Anonymous. Improved hay rake. Scientific American, 56(23):354, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n23/pdf/scientificamerican06041887-354c.pdf.

Anonymous:1887:FPS

[9171] Anonymous. A frost proof steam trap for car heating. Scientific American, 56(23):354, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n23/pdf/scientificamerican06041887-
354d.pdf.

Anonymous:1887:AC

[9172] Anonymous. Alphabet case. Scientific American, 56(23):354, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n23/pdf/scientificamerican06041887-354e.pdf.

Anonymous:1887:DA

[9173] Anonymous. Dental apparatus. Scientific American, 56(23):355, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n23/pdf/scientificamerican06041887-355.pdf.

Anonymous:1887:ICCf

[9174] Anonymous. Improved car coupling. Scientific American, 56(23):355, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n23/pdf/scientificamerican06041887-355a.pdf.

Anonymous:1887:SHF

[9175] Anonymous. Stock or hay frame. Scientific American, 56(23):355, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n23/pdf/scientificamerican06041887-355b.pdf.

Anonymous:1887:ICF

[9176] Anonymous. Improved curtain fixture. Scientific American, 56(23):355, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n23/pdf/scientificamerican06041887-355c.pdf.

Anonymous:1887:ISFa

[9177] Anonymous. Improved sash fastener. Scientific American, 56 (23):355–356, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n23/pdf/scientificamerican06041887-
355d.pdf.

Anonymous:1887:HIM

[9178] Anonymous. Hand ironing machine. Scientific American, 56(23):356, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n23/pdf/scientificamerican06041887-356.pdf.

Anonymous:1887:DAP

[9179] Anonymous. Double-acting pump. Scientific American, 56(23):356, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n23/pdf/scientificamerican06041887-356a.pdf.

Anonymous:1887:IW

[9180] Anonymous. Improved wrench. Scientific American, 56(23):356, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n23/pdf/scientificamerican06041887-356b.pdf.

Anonymous:1887:PNb

[9181] Anonymous. Photographic notes. Scientific American, 56(23):357, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n23/pdf/scientificamerican06041887-357a.pdf.

Anonymous:1887:DRPd

[9182] Anonymous. Decisions relating to patents. Scientific American, 56(23):357, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n23/pdf/scientificamerican06041887-
357b.pdf.

Anonymous:1887:PP

[9183] Anonymous. The pitcher plant. Scientific American, 56(23):359, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n23/pdf/scientificamerican06041887-359.pdf.

Seiss:1887:PSC

[9184] C. Few Seiss. The pine snake — so-called horn snake. Scientific American, 56(23):359–360, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n23/pdf/scientificamerican06041887-
359a.pdf.

Anonymous:1887:CBCa

[9185] Anonymous. Collision between the Celtic and Britannic. Scientific American, 56(23):361, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n23/pdf/scientificamerican06041887-
361.pdf.

Anonymous:1887:EIp

[9186] Anonymous. Engineering inventions. Scientific American, 56(23):362, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n23/pdf/scientificamerican06041887-362.pdf.

Anonymous:1887:AIm

[9187] Anonymous. Agricultural intention. Scientific American, 56(23):362, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n23/pdf/scientificamerican06041887-362a.pdf.

Anonymous:1887:MIs

[9188] Anonymous. Miscellaneous inventions. Scientific American, 56(23):362, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n23/pdf/scientificamerican06041887-362b.pdf.

Anonymous:1887:Sb

[9189] Anonymous. Special. Scientific American, 56(23):362, June 4, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n23/pdf/scientificamerican06041887-362c.pdf.

Anonymous:1887:Co

[9190] Anonymous. Correspondence. Scientific American, 56(24):373, June 11, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n24/pdf/scientificamerican06111887-373.pdf.

Anonymous:1887:BPq

[9191] Anonymous. Business and personal. Scientific American, 56(24):379, June 11, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n24/pdf/scientificamerican06111887-379c.pdf.

Anonymous:1887:IIr

[9192] Anonymous. Index of inventions. Scientific American, 56(24):379–380, June 11, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n24/pdf/scientificamerican06111887-379d.pdf.

Anonymous:1887:PNPo

[9193] Anonymous. Patent notes: Patents. Scientific American, 56(24):382, June 11, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n24/pdf/scientificamerican06111887-382.pdf.

Anonymous:1887:SLU

[9194] Anonymous. Search light on the U.S. cruiser Atlanta. Scientific American, 56(24):367–372, June 11, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n24/pdf/scientificamerican06111887-
372.pdf.

Anonymous:1887:CAN

[9195] Anonymous. Celebration of the anniversary of natural gas in Findlay, O. Scientific American, 56(24):368, June 11, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v56/
n24/pdf/scientificamerican06111887-368.pdf.

Anonymous:1887:DHN

[9196] Anonymous. The defense of the harbor of New York. Scientific American, 56(24):368, June 11, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n24/pdf/scientificamerican06111887-
368a.pdf.

Anonymous:1887:PNc

[9197] Anonymous. Photographic notes. Scientific American, 56(24):369, June 11, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n24/pdf/scientificamerican06111887-369.pdf.

Anonymous:1887:SP

[9198] Anonymous. Smoking pipe. Scientific American, 56(24):370, June 11, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n24/pdf/scientificamerican06111887-370.pdf.

Sloane:1887:SDB

[9199] T. O’Conor Sloane. Standard Daniell’s battery. Scientific American, 56(24):370, June 11, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n24/pdf/scientificamerican06111887-
370a.pdf.

Anonymous:1887:FG

[9200] Anonymous. Fruit gatherer. Scientific American, 56(24):370, June 11, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n24/pdf/scientificamerican06111887-370b.pdf.

Anonymous:1887:IDSb

[9201] Anonymous. Improved door securer. Scientific American, 56(24):370, June 11, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n24/pdf/scientificamerican06111887-370c.pdf.

Anonymous:1887:ICCg

[9202] Anonymous. Improved camera clamp and tripod head. Scientific American, 56(24):371, June 11, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n24/pdf/scientificamerican06111887-
371.pdf.

Anonymous:1887:NMM

[9203] Anonymous. The new military multicycle. Scientific American, 56(24):371, June 11, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n24/pdf/scientificamerican06111887-
371a.pdf.

Anonymous:1887:NT

[9204] Anonymous. A novel trunk. Scientific American, 56(24):371, June 11, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n24/pdf/scientificamerican06111887-371b.pdf.

Anonymous:1887:HGO

[9205] Anonymous. How to graft old trees. Scientific American, 56(24):373, June 11, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n24/pdf/scientificamerican06111887-373a.pdf.

Hopkins:1887:IAL

[9206] Geo. M. Hopkins. Inexpensive arc lamp. Scientific American, 56(24):374, June 11, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n24/pdf/scientificamerican06111887-
374.pdf.

Anonymous:1887:AM

[9207] Anonymous. Amateur mechanics. Scientific American, 56(24):374, June 11, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n24/pdf/scientificamerican06111887-374a.pdf.

Anonymous:1887:PLS

[9208] Anonymous. A pair of live salamanders. Scientific American, 56 (24):375–376, June 11, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n24/pdf/scientificamerican06111887-
375.pdf.

Anonymous:1887:PIE

[9209] Anonymous. The Paris International Exhibition. Scientific American, 56(24):377–378, June 11, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n24/pdf/scientificamerican06111887-
377.pdf.

Anonymous:1887:MIt

[9210] Anonymous. Miscellaneous inventions. Scientific American, 56(24):378, June 11, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n24/pdf/scientificamerican06111887-379b.pdf.

Anonymous:1887:EIq

[9211] Anonymous. Engineering invention. Scientific American, 56(24):379, June 11, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n24/pdf/scientificamerican06111887-379.pdf.

Anonymous:1887:AIn

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n24/pdf/scientificamerican06111887-379a.pdf.

Anonymous:1887:Cp

[9213] Anonymous. Correspondence. Scientific American, 56(25):389, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n25/pdf/scientificamerican06181887-389.pdf.

Anonymous:1887:BPr

[9214] Anonymous. Business and personal. Scientific American, 56 (25):394–395, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n25/pdf/scientificamerican06181887-
394d.pdf.

Anonymous:1887:NQo

[9215] Anonymous. Notes and queries. Scientific American, 56(25):395, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n25/pdf/scientificamerican06181887-395.pdf.

Anonymous:1887:IIs

[9216] Anonymous. Index of inventions. Scientific American, 56(25):395–396, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n25/pdf/scientificamerican06181887-395b.pdf.

Anonymous:1887:PNPp

[9217] Anonymous. Patent notes: Patents. Scientific American, 56(25):398, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n25/pdf/scientificamerican06181887-398.pdf.

Benjamin:1887:AA

[9218] Marcus Benjamin. Alexander Agassiz. Scientific American, 56 (25):383–389, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n25/pdf/scientificamerican06181887-
383.pdf.

Anonymous:1887:CTA

[9219] Anonymous. Coffee and tea adulteration. Scientific American, 56(25):384, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n25/pdf/scientificamerican06181887-
384.pdf.

Anonymous:1887:CWd

[9220] Anonymous. The celestial world. Scientific American, 56(25):384, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n25/pdf/scientificamerican06181887-384a.pdf.

Anonymous:1887:PNd

[9221] Anonymous. Photographic notes. Scientific American, 56(25):385, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n25/pdf/scientificamerican06181887-385.pdf.

Anonymous:1887:CRHa

[9222] Anonymous. Combined range and heater. Scientific American, 56(25):386, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n25/pdf/scientificamerican06181887-
386.pdf.

Anonymous:1887:IMF

[9223] Anonymous. Improved means of furling topsails. Scientific American, 56(25):386, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n25/pdf/scientificamerican06181887-
386a.pdf.

Anonymous:1887:CUCc

[9224] Anonymous. A combined uniform and civilian’s button. Scientific American, 56(25):386, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n25/pdf/scientificamerican06181887-
386b.pdf.

Anonymous:1887:SNR

[9225] Anonymous. A secure and non-rattling thill coupling. Scientific American, 56(25):386, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n25/pdf/scientificamerican06181887-
386c.pdf.

Anonymous:1887:TBS

[9226] Anonymous. A timber bunk for sleigh or other vehicle. Scientific American, 56(25):386, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n25/pdf/scientificamerican06181887-
386d.pdf.

Anonymous:1887:IPCb

[9227] Anonymous. Improved package for carrying eggs. Scientific American, 56(25):386, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n25/pdf/scientificamerican06181887-
386e.pdf.

Anonymous:1887:ISG

[9228] Anonymous. An improved sliding gate. Scientific American, 56(25):387, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n25/pdf/scientificamerican06181887-387.pdf.

Anonymous:1887:IBF

[9229] Anonymous. Improved Backus furnaces. Scientific American, 56 (25):387, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n25/pdf/scientificamerican06181887-
387a.pdf.

Anonymous:1887:ICWb

[9230] Anonymous. An improved clothes washer. Scientific American, 56(25):387, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n25/pdf/scientificamerican06181887-
387b.pdf.

Anonymous:1887:IDMa

[9231] Anonymous. An improved ditch making machine. Scientific American, 56(25):388, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n25/pdf/scientificamerican06181887-
388.pdf.

Anonymous:1887:IPB

[9232] Anonymous. An improved portable bed. Scientific American, 56 (25):388, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n25/pdf/scientificamerican06181887-
388a.pdf.

Anonymous:1887:WSB

[9233] Anonymous. A weather strip for the bottoms of doors. Scientific American, 56(25):388, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n25/pdf/scientificamerican06181887-
388b.pdf.

Anonymous:1887:NSB

[9234] Anonymous. The Nordenfelt submarine boat at Constantinople. Scientific American, 56(25):390, June 18, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v56/
n25/pdf/scientificamerican06181887-390.pdf.

Anonymous:1887:RPc

[9235] Anonymous. The royal palm. Scientific American, 56(25):391–392, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n25/pdf/scientificamerican06181887-391.pdf.

Anonymous:1887:STh

[9236] Anonymous. Science in toys. Scientific American, 56(25):393, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n25/pdf/scientificamerican06181887-393.pdf.

Anonymous:1887:CJa

[9237] Anonymous. The Crown Jewels. Scientific American, 56(25):393, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n25/pdf/scientificamerican06181887-393a.pdf.

Anonymous:1887:EIr

[9238] Anonymous. Engineering inventions. Scientific American, 56(25):394, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n25/pdf/scientificamerican06181887-394.pdf.

Anonymous:1887:AIo

[9239] Anonymous. Agricultural inventions. Scientific American, 56(25):394, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n25/pdf/scientificamerican06181887-394a.pdf.

Anonymous:1887:MIu

[9240] Anonymous. Miscellaneous inventions. Scientific American, 56(25):394, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n25/pdf/scientificamerican06181887-394b.pdf.

Anonymous:1887:Sc

[9241] Anonymous. Special. Scientific American, 56(25):394, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n25/pdf/scientificamerican06181887-394c.pdf.

Anonymous:1887:It

[9242] Anonymous. To inventors. Scientific American, 56(25):395, June 18, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n25/pdf/scientificamerican06181887-395a.pdf.

Anonymous:1887:Cq

[9243] Anonymous. Correspondence. Scientific American, 56(26):405, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n26/pdf/scientificamerican06251887-405.pdf.

Anonymous:1887:BPs

[9244] Anonymous. Business and personal. Scientific American, 56(26):409, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n26/pdf/scientificamerican06251887-409c.pdf.

Anonymous:1887:BNBm

[9245] Anonymous. Books: New books and publications. Scientific American, 56(26):409, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n26/pdf/scientificamerican06251887-
409d.pdf.

Anonymous:1887:IIt

[9246] Anonymous. Index of inventions. Scientific American, 56(26):409–410, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n26/pdf/scientificamerican06251887-409e.pdf.

Anonymous:1887:Iu

[9247] Anonymous. Index. Scientific American, 56(26):411–412, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n26/pdf/scientificamerican06251887-411.pdf.

Anonymous:1887:PNPq

[9248] Anonymous. Patent notes: Patents. Scientific American, 56(26):414, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n26/pdf/scientificamerican06251887-414.pdf.

Anonymous:1887:CME

[9249] Anonymous. Compound mill engine. Scientific American, 56(26):399, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n26/pdf/scientificamerican06251887-399.pdf.

Anonymous:1887:ISe

[9250] Anonymous. Investigating spiritualism. Scientific American, 56 (26):400–401, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n26/pdf/scientificamerican06251887-
400.pdf.

Anonymous:1887:NWH

[9251] Anonymous. A new way to hang vehicles. Scientific American, 56(26):402, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n26/pdf/scientificamerican06251887-
402.pdf.

Anonymous:1887:CSCa

[9252] Anonymous. Corn and sugar cane cutter. Scientific American, 56(26):402, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n26/pdf/scientificamerican06251887-
402a.pdf.

Anonymous:1887:DFH

[9253] Anonymous. A device to facilitate the heating of sad irons. Scientific American, 56(26):402, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n26/pdf/scientificamerican06251887-
402b.pdf.

Anonymous:1887:IIBa

[9254] Anonymous. An improved ironing board. Scientific American, 56(26):402, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n26/pdf/scientificamerican06251887-
402c.pdf.

Anonymous:1887:IFN

[9255] Anonymous. Improved fly net for horses. Scientific American, 56(26):402, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n26/pdf/scientificamerican06251887-
402d.pdf.

Anonymous:1887:CRHb

[9256] Anonymous. A clevis to relieve horses from strain. Scientific American, 56(26):403, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n26/pdf/scientificamerican06251887-
403.pdf.

Anonymous:1887:ICPc

[9257] Anonymous. An improved cotton picker. Scientific American, 56 (26):403, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n26/pdf/scientificamerican06251887-
403a.pdf.

Anonymous:1887:ISFb

[9258] Anonymous. An improved saw filing machine. Scientific American, 56(26):403, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n26/pdf/scientificamerican06251887-
403b.pdf.

Anonymous:1887:SSD

[9259] Anonymous. A seed sowing device. Scientific American, 56(26):403, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n26/pdf/scientificamerican06251887-403c.pdf.

Anonymous:1887:ITCa

[9260] Anonymous. An improved thill coupling. Scientific American, 56(26):404, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n26/pdf/scientificamerican06251887-
404.pdf.

Anonymous:1887:IV

[9261] Anonymous. An improved ventilator. Scientific American, 56(26):404, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n26/pdf/scientificamerican06251887-404a.pdf.

Anonymous:1887:RB

[9262] Anonymous. The Rudge bicyclette. Scientific American, 56(26):404, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n26/pdf/scientificamerican06251887-404b.pdf.

Anonymous:1887:TSE

[9263] Anonymous. The total solar eclipse of August 19, 1887. Scientific American, 56(26):404–405, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n26/pdf/scientificamerican06251887-
404c.pdf.

Anonymous:1887:USR

[9264] Anonymous. An unsinkable, self-righting lifeboat. Scientific American, 56(26):406, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n26/pdf/scientificamerican06251887-
406.pdf.

Anonymous:1887:TWB

[9265] Anonymous. The twelve-wired bird of paradise. Scientific American, 56(26):407, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n26/pdf/scientificamerican06251887-
407.pdf.

Michels:1887:CCP

[9266] John Michels. Condition of the Crown Prince of Germany. Scientific American, 56(26):407–408, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v56/n26/pdf/scientificamerican06251887-
407a.pdf.

Anonymous:1887:EIs

[9267] Anonymous. Engineering inventions. Scientific American, 56(26):409, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n26/pdf/scientificamerican06251887-409.pdf.

Anonymous:1887:AIp

[9268] Anonymous. Agricultural inventions. Scientific American, 56(26):409, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n26/pdf/scientificamerican06251887-409a.pdf.

Anonymous:1887:MIv

[9269] Anonymous. Miscellaneous inventions. Scientific American, 56(26):409, June 25, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n26/pdf/scientificamerican06251887-409b.pdf.

Anonymous:1887:Cs

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n3/pdf/scientificamerican01151887-37.pdf.

Anonymous:1887:BPt

[9271] Anonymous. Business and personal. Scientific American, 56 (3):42–43, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n3/pdf/scientificamerican01151887-
42c.pdf.

Anonymous:1887:BNBn

[9272] Anonymous. Books: New books and publications. Scientific American, 56(3):43, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n3/pdf/scientificamerican01151887-
43.pdf.

Anonymous:1887:NQp

[9273] Anonymous. Notes and queries. Scientific American, 56(3):43, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n3/pdf/scientificamerican01151887-43a.pdf.

Anonymous:1887:IIu

[9274] Anonymous. Index of inventions. Scientific American, 56(3):43–44, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n3/pdf/scientificamerican01151887-43c.pdf.

Anonymous:1887:PNPr

[9275] Anonymous. Patent notes: Patents. Scientific American, 56(3):46, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n3/pdf/scientificamerican01151887-46.pdf.

Anonymous:1887:IOH

[9276] Anonymous. Improved one hundred ton floating crane. Scientific American, 56(3):31, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n3/pdf/scientificamerican01151887-
31.pdf.

Anonymous:1887:WIb

[9277] Anonymous. War and invention. Scientific American, 56(3):32, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n3/pdf/scientificamerican01151887-32.pdf.

Anonymous:1887:PS

[9278] Anonymous. Promotion by seniority. Scientific American, 56 (3):32–33, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n3/pdf/scientificamerican01151887-
32a.pdf.

Anonymous:1887:MWA

[9279] Anonymous. Man and the wild animal. Scientific American, 56(3):33, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n3/pdf/scientificamerican01151887-
33.pdf.

Anonymous:1887:PNe

[9280] Anonymous. Photographic notes. Scientific American, 56(3):33, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n3/pdf/scientificamerican01151887-33a.pdf.

Anonymous:1887:ISBb

[9281] Anonymous. Improved steam boiler. Scientific American, 56(3):34, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n3/pdf/scientificamerican01151887-34.pdf.

Anonymous:1887:LPS

[9282] Anonymous. Large punching and shearing machine. Scientific American, 56(3):34, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n3/pdf/scientificamerican01151887-
34a.pdf.

Anonymous:1887:WGS

[9283] Anonymous. Water gauge for steam boiler. Scientific American, 56(3):34, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n3/pdf/scientificamerican01151887-
34b.pdf.

Anonymous:1887:PFA

[9284] Anonymous. Picture frame attachment. Scientific American, 56 (3):35, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n3/pdf/scientificamerican01151887-
35.pdf.

Anonymous:1887:CPP

[9285] Anonymous. Combined pulvebizer and plow. Scientific American, 56(3):35, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n3/pdf/scientificamerican01151887-
35a.pdf.

Anonymous:1887:ISDa

[9286] Anonymous. Improved spring and dust guard for watch cases. Scientific American, 56(3):35, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n3/pdf/scientificamerican01151887-
35b.pdf.

Anonymous:1887:FSL

[9287] Anonymous. Folding step ladder. Scientific American, 56(3):35, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n3/pdf/scientificamerican01151887-35c.pdf.

Anonymous:1887:IOCb

[9288] Anonymous. Improved oil cup. Scientific American, 56(3):35–36, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n3/pdf/scientificamerican01151887-35d.pdf.

Anonymous:1887:CST

[9289] Anonymous. Comfort and style too. Scientific American, 56(3):36, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n3/pdf/scientificamerican01151887-36.pdf.

Anonymous:1887:Pa

[9290] Anonymous. A puzzle. Scientific American, 56(3):36, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n3/pdf/scientificamerican01151887-36a.pdf.

Sloane:1887:MFE

[9291] T. O’conor Sloane. Miking of fluids — expansion by heat. Scientific American, 56(3):38, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n3/pdf/scientificamerican01151887-
38.pdf.

Anonymous:1887:PMP

[9292] Anonymous. Photography of a moving pendulum. Scientific American, 56(3):38–39, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n3/pdf/scientificamerican01151887-
38a.pdf.

Anonymous:1887:PAE

[9293] Anonymous. The porcupine ant eater. Scientific American, 56 (3):39–40, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n3/pdf/scientificamerican01151887-
39.pdf.

Anonymous:1887:STi

[9294] Anonymous. Science in toys. Scientific American, 56(3):40, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n3/pdf/scientificamerican01151887-40.pdf.

Anonymous:1887:IBD

[9295] Anonymous. Incandescent burner of dr. auer. Scientific American, 56(3):41, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n3/pdf/scientificamerican01151887-
41.pdf.

Anonymous:1887:JED

[9296] Anonymous. Joseph Echteler’s design for a Grant Monument. Scientific American, 56(3):41, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n3/pdf/scientificamerican01151887-
41a.pdf.

Anonymous:1887:EIt

[9297] Anonymous. Engineer inventions. Scientific American, 56(3):42, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n3/pdf/scientificamerican01151887-42.pdf.

Anonymous:1887:AIq

[9298] Anonymous. Agricultural invention. Scientific American, 56(3):42, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n3/pdf/scientificamerican01151887-42a.pdf.

Anonymous:1887:MIw

[9299] Anonymous. Miscellaneous inventions. Scientific American, 56 (3):42, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n3/pdf/scientificamerican01151887-
42b.pdf.

Anonymous:1887:Ix

[9300] Anonymous. To inventors. Scientific American, 56(3):43, January 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n3/pdf/scientificamerican01151887-43b.pdf.

Anonymous:1887:Cu

[9301] Anonymous. Correspondence. Scientific American, 56(4):53, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n4/pdf/scientificamerican01221887-53.pdf.

Anonymous:1887:BPu

[9302] Anonymous. Business and personal. Scientific American, 56(4):59, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n4/pdf/scientificamerican01221887-59d.pdf.

Anonymous:1887:NQq

[9303] Anonymous. Notes and queries. Scientific American, 56(4):59, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n4/pdf/scientificamerican01221887-59e.pdf.

Anonymous:1887:IIv

[9304] Anonymous. Index of inventions. Scientific American, 56(4):59–60, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n4/pdf/scientificamerican01221887-59g.pdf.

Anonymous:1887:PNPs

[9305] Anonymous. Patent notes: Patents. Scientific American, 56(4):62, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n4/pdf/scientificamerican01221887-62.pdf.

Hopkins:1887:RD

[9306] Geo. M. Hopkins. Reminiscences of Daguerreotypy. Scientific American, 56(4):47, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n4/pdf/scientificamerican01221887-
47.pdf.

Anonymous:1887:JR

[9307] Anonymous. John Roach. Scientific American, 56(4):48, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n4/pdf/scientificamerican01221887-48.pdf.

Anonymous:1887:IWA

[9308] Anonymous. The inventor’s work in agriculture. Scientific American, 56(4):48–49, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n4/pdf/scientificamerican01221887-
48a.pdf.

Anonymous:1887:IMPa

[9309] Anonymous. Improved metallic printing block. Scientific American, 56(4):50, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n4/pdf/scientificamerican01221887-
50.pdf.

Anonymous:1887:IGE

[9310] Anonymous. Improved gas engine. Scientific American, 56(4):50, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n4/pdf/scientificamerican01221887-50a.pdf.

Anonymous:1887:IDA

[9311] Anonymous. Improved door attachment. Scientific American, 56 (4):50, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n4/pdf/scientificamerican01221887-
50b.pdf.

Anonymous:1887:IASa

[9312] Anonymous. Improved automatic step ladder. Scientific American, 56(4):51, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n4/pdf/scientificamerican01221887-
51.pdf.

Anonymous:1887:IDCb

[9313] Anonymous. Improved door check. Scientific American, 56(4):51, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n4/pdf/scientificamerican01221887-51a.pdf.

Anonymous:1887:IWC

[9314] Anonymous. Improved wheel cultivator. Scientific American, 56 (4):51, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n4/pdf/scientificamerican01221887-
51b.pdf.

Anonymous:1887:CCc

[9315] Anonymous. Cheese cutter. Scientific American, 56(4):51, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n4/pdf/scientificamerican01221887-51c.pdf.

Anonymous:1887:CAGb

[9316] Anonymous. Compressed air grease cup. Scientific American, 56(4):51, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n4/pdf/scientificamerican01221887-
51d.pdf.

Anonymous:1887:ICd

[9317] Anonymous. Illuminated clock. Scientific American, 56(4):52, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n4/pdf/scientificamerican01221887-52.pdf.

Anonymous:1887:SMR

[9318] Anonymous. The Saint Malo rolling bridge. Scientific American, 56(4):54, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n4/pdf/scientificamerican01221887-
54.pdf.

Anonymous:1887:SCb

[9319] Anonymous. Steam carriages. Scientific American, 56(4):55, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n4/pdf/scientificamerican01221887-55.pdf.

Anonymous:1887:SPI

[9320] Anonymous. The St. Paul Ice Palace. Scientific American, 56 (4):55, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n4/pdf/scientificamerican01221887-
55a.pdf.

Anonymous:1887:SIN

[9321] Anonymous. Sir Isaac Newton’s steam carriage. Scientific American, 56(4):55, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n4/pdf/scientificamerican01221887-
55b.pdf.

Anonymous:1887:STj

[9322] Anonymous. Science in toys. Scientific American, 56(4):56, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n4/pdf/scientificamerican01221887-56.pdf.

Anonymous:1887:IWN

[9323] Anonymous. Improved wire netting machine. Scientific American, 56(4):57–58, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n4/pdf/scientificamerican01221887-
57.pdf.

Anonymous:1887:EIu

[9324] Anonymous. Engineering inventions. Scientific American, 56(4): 59, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n4/pdf/scientificamerican01221887-
59.pdf.

Anonymous:1887:MIx

[9325] Anonymous. Mechanical inventions. Scientific American, 56(4):59, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n4/pdf/scientificamerican01221887-59a.pdf.

Anonymous:1887:AIr

[9326] Anonymous. Agricultural inventions. Scientific American, 56 (4):59, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n4/pdf/scientificamerican01221887-
59b.pdf.

Anonymous:1887:MIy

[9327] Anonymous. Miscellaneous inventions. Scientific American, 56 (4):59, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n4/pdf/scientificamerican01221887-
59c.pdf.

Anonymous:1887:Iy

[9328] Anonymous. To inventors. Scientific American, 56(4):59, January 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n4/pdf/scientificamerican01221887-59f.pdf.

Anonymous:1887:BPv

[9329] Anonymous. Business and personal. Scientific American, 56(5):74, January 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n5/pdf/scientificamerican01291887-74d.pdf.

Anonymous:1887:NQr

[9330] Anonymous. Notes and queries. Scientific American, 56(5):75, January 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n5/pdf/scientificamerican01291887-75.pdf.

Anonymous:1887:PNPt

[9331] Anonymous. Patent notes: Patents. Scientific American, 56(5):78, January 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n5/pdf/scientificamerican01291887-78.pdf.

Anonymous:1887:WWH

[9332] Anonymous. The water wheels of Hamath. Scientific American, 56(5):63–64, January 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n5/pdf/scientificamerican01291887-
63.pdf.

Anonymous:1887:DOP

[9333] Anonymous. Defeat of the ornoxious Patent Bill. Scientific American, 56(5):64, January 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n5/pdf/scientificamerican01291887-
64.pdf.

Anonymous:1887:TUL

[9334] Anonymous. Two useful lives. Scientific American, 56(5):64–65, January 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n5/pdf/scientificamerican01291887-64a.pdf.

Anonymous:1887:CBA

[9335] Anonymous. Car brake attachment. Scientific American, 56(5):66, January 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n5/pdf/scientificamerican01291887-66.pdf.

Anonymous:1887:AEO

[9336] Anonymous. Apparatus for extracting oil from seeds. Scientific American, 56(5):66, January 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n5/pdf/scientificamerican01291887-
66a.pdf.

Anonymous:1887:CSBa

[9337] Anonymous. Car starter and brake. Scientific American, 56 (5):66–67, January 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n5/pdf/scientificamerican01291887-
66b.pdf.

Anonymous:1887:CPa

[9338] Anonymous. Combination plow. Scientific American, 56(5):67, January 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n5/pdf/scientificamerican01291887-67.pdf.

Anonymous:1887:CTW

[9339] Anonymous. Combination tool wrench. Scientific American, 56 (5):67, January 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n5/pdf/scientificamerican01291887-
67a.pdf.

Anonymous:1887:CBP

[9340] Anonymous. Cultivator beam and point. Scientific American, 56(5):67, January 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n5/pdf/scientificamerican01291887-
67b.pdf.

Anonymous:1887:ITP

[9341] Anonymous. Improved tobacco planter. Scientific American, 56 (5):67–69, January 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n5/pdf/scientificamerican01291887-
67c.pdf.

Anonymous:1887:LCG

[9342] Anonymous. Lathe center grinding machine. Scientific American, 56(5):70, January 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n5/pdf/scientificamerican01291887-
70.pdf.

Anonymous:1887:ETP

[9343] Anonymous. Electrical transmission of power to a distance. Scientific American, 56(5):70–71, January 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v56/n5/pdf/scientificamerican01291887-
70a.pdf.

Sloane:1887:ESL

[9344] T. O’conor Sloane. Expansion of solids, liquids, and gases. Scientific American, 56(5):71–72, January 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v56/n5/pdf/scientificamerican01291887-
71.pdf.

Anonymous:1887:RPd

[9345] Anonymous. The rhinoceros in the Park. Scientific American, 56(5):72, January 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n5/pdf/scientificamerican01291887-
72.pdf.

Anonymous:1887:STk

[9346] Anonymous. Science in toys. Scientific American, 56(5):72, January 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n5/pdf/scientificamerican01291887-72a.pdf.

Anonymous:1887:ICMb

[9347] Anonymous. Inverted cylinder mill engines. Scientific American, 56(5):73, January 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n5/pdf/scientificamerican01291887-
73.pdf.

Anonymous:1887:EIv

[9348] Anonymous. Engineering inventions. Scientific American, 56(5): 74, January 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n5/pdf/scientificamerican01291887-
74.pdf.

Anonymous:1887:MIz

[9349] Anonymous. Mechanical invention. Scientific American, 56(5):74, January 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n5/pdf/scientificamerican01291887-74a.pdf.

Anonymous:1887:AIs

[9350] Anonymous. Agricultural inventions. Scientific American, 56 (5):74, January 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n5/pdf/scientificamerican01291887-
74b.pdf.

Anonymous:1887:MIba

[9351] Anonymous. Miscellaneous inventions. Scientific American, 56 (5):74, January 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n5/pdf/scientificamerican01291887-
74c.pdf.

Anonymous:1887:IIw

[9352] Anonymous. Index of inventions. Scientific American, 56(5):75–76, January 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n5/pdf/scientificamerican01291887-75a.pdf.

Anonymous:1887:Cv

[9353] Anonymous. Correspondence. Scientific American, 56(6):85–86, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n6/pdf/scientificamerican02051887-85.pdf.

Anonymous:1887:BPw (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n6/pdf/scientificamerican02051887-
91d.pdf.

Anonymous:1887:NQs

[9355] Anonymous. Notes and queries. Scientific American, 56(6):91, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n6/pdf/scientificamerican02051887-91e.pdf.

Anonymous:1887:PNPu

[9356] Anonymous. Patent notes: Patents. Scientific American, 56 (6):94, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n6/pdf/scientificamerican02051887-
94.pdf.

Anonymous:1887:BHP

[9357] Anonymous. Bridging the Hudson at Poughkeepsie. Scientific American, 56(6):79, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n6/pdf/scientificamerican02051887-
79.pdf.

Anonymous:1887:BTS

[9358] Anonymous. The Bell Telephone Suits. Scientific American, 56 (6):80, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n6/pdf/scientificamerican02051887-
80.pdf.

Anonymous:1887:SJW

[9359] Anonymous. Sir Joseph Whitworth. Scientific American, 56 (6):80, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n6/pdf/scientificamerican02051887-
80a.pdf.

Anonymous:1887:PAJ

[9360] Anonymous. The pet ambition of John Roach. Scientific American, 56(6):81, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n6/pdf/scientificamerican02051887-
81.pdf.

Anonymous:1887:LAL

[9361] Anonymous. Lathe for amateurs and light shop work. Scientific American, 56(6):82, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n6/pdf/scientificamerican02051887-
82.pdf.

Anonymous:1887:EMBa

[9362] Anonymous. An electrical music baton. Scientific American, 56 (6):82, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n6/pdf/scientificamerican02051887-
82a.pdf.

Anonymous:1887:DTE

[9363] Anonymous. Device for trimming the ends of tubes. Scientific American, 56(6):82–83, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n6/pdf/scientificamerican02051887-
82b.pdf.

Anonymous:1887:BBH

[9364] Anonymous. Back band for harness. Scientific American, 56 (6):83, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n6/pdf/scientificamerican02051887-
83.pdf.

Anonymous:1887:BCS

[9365] Anonymous. Balanced cooking stove for ships. Scientific American, 56(6):83, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n6/pdf/scientificamerican02051887-
83a.pdf.

Anonymous:1887:IDRa

[9366] Anonymous. The Ingersoll duplex rock drilling machine. Scientific American, 56(6):83, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n6/pdf/scientificamerican02051887-
83b.pdf.

Anonymous:1887:IGP

[9367] Anonymous. Improved gas pliers. Scientific American, 56(6):83, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n6/pdf/scientificamerican02051887-83c.pdf.

Anonymous:1887:ILPb

[9368] Anonymous. Improved lead pipe reel. Scientific American, 56 (6):83–84, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n6/pdf/scientificamerican02051887-
83d.pdf.

Anonymous:1887:ISW

[9369] Anonymous. Improved spring wheel traction engine. Scientific American, 56(6):86, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n6/pdf/scientificamerican02051887-
86.pdf.

Holder:1887:SNF

[9370] C. F. Holder. Some new fishes. Scientific American, 56(6):87, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n6/pdf/scientificamerican02051887-87.pdf.

Anonymous:1887:SBb

[9371] Anonymous. Swing bicycle. Scientific American, 56(6):87–88, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n6/pdf/scientificamerican02051887-87a.pdf.

Anonymous:1887:ICBb

[9372] Anonymous. Improved car brake. Scientific American, 56(6):88, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n6/pdf/scientificamerican02051887-88.pdf.

Anonymous:1887:IRR

[9373] Anonymous. Improved railway rail joint. Scientific American, 56(6):88, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n6/pdf/scientificamerican02051887-
88a.pdf.

Anonymous:1887:STl

[9374] Anonymous. Science in toys. Scientific American, 56(6):89–90, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n6/pdf/scientificamerican02051887-89.pdf.

Anonymous:1887:EIw

[9375] Anonymous. Engineering inventions. Scientific American, 56(6): 91, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n6/pdf/scientificamerican02051887-
91.pdf.

Anonymous:1887:MIbb

[9376] Anonymous. Mechanical inventions. Scientific American, 56(6): 91, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n6/pdf/scientificamerican02051887-
91a.pdf.

Anonymous:1887:AIt

[9377] Anonymous. Agricultural inventions. Scientific American, 56 (6):91, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n6/pdf/scientificamerican02051887-
91b.pdf.

Anonymous:1887:MIbc

[9378] Anonymous. Miscellaneous inventions. Scientific American, 56 (6):91, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n6/pdf/scientificamerican02051887-
91c.pdf.

Anonymous:1887:Iz

[9379] Anonymous. To inventors. Scientific American, 56(6):91, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n6/pdf/scientificamerican02051887-91f.pdf.

Anonymous:1887:IIx

[9380] Anonymous. Index of inventions. Scientific American, 56(6): 91–92, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n6/pdf/scientificamerican02051887-
91g.pdf.

Anonymous:1887:FPb

[9381] Anonymous. Foreign patents. Scientific American, 56(6):93, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n6/pdf/scientificamerican02051887-93.pdf.

Anonymous:1887:PSA

[9382] Anonymous. Prospectus of the Scientific American for 1887. Scientific American, 56(6):94, February 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n6/pdf/scientificamerican02051887-
94a.pdf.

Anonymous:1887:Cx

[9383] Anonymous. Correspondence. Scientific American, 56(7):101– 102, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n7/pdf/scientificamerican02121887-
101.pdf.

Anonymous:1887:BPx

[9384] Anonymous. Business and personal. Scientific American, 56(7): 107, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n7/pdf/scientificamerican02121887-
107d.pdf.

Anonymous:1887:NQt

[9385] Anonymous. Notes and queries. Scientific American, 56(7):107, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n7/pdf/scientificamerican02121887-107e.pdf.

Anonymous:1887:IIy

[9386] Anonymous. Index of inventions. Scientific American, 56(7): 107–108, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n7/pdf/scientificamerican02121887-
107g.pdf.

Anonymous:1887:PNPv

[9387] Anonymous. Patent notes: Patents. Scientific American, 56 (7):110, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n7/pdf/scientificamerican02121887-
110.pdf.

Anonymous:1887:CWS

[9388] Anonymous. City water supply from driven wells. Scientific American, 56(7):95, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n7/pdf/scientificamerican02121887-
95.pdf.

Anonymous:1887:CDU

[9389] Anonymous. The coast defenses of the United States. Scientific American, 56(7):96, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n7/pdf/scientificamerican02121887-
96.pdf.

Anonymous:1887:PNf

[9390] Anonymous. Photographic notes. Scientific American, 56(7):97, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n7/pdf/scientificamerican02121887-97.pdf.

Anonymous:1887:ICSc

[9391] Anonymous. Improved cattle stanchion. Scientific American, 56 (7):98, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n7/pdf/scientificamerican02121887-
98.pdf.

Anonymous:1887:EML

[9392] Anonymous. Exhaust mechanism for locomotives. Scientific American, 56(7):98, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n7/pdf/scientificamerican02121887-
98a.pdf.

Anonymous:1887:IRSa

[9393] Anonymous. Improved railroad signal. Scientific American, 56 (7):98, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n7/pdf/scientificamerican02121887-
98b.pdf.

Anonymous:1887:TS

[9394] Anonymous. Trousers stretcher. Scientific American, 56(7):98, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n7/pdf/scientificamerican02121887-98c.pdf.

Anonymous:1887:EPR

[9395] Anonymous. Experiment with paper rings. Scientific American, 56(7):99, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n7/pdf/scientificamerican02121887-
99.pdf.

Anonymous:1887:NMC

[9396] Anonymous. A new method of constructing propeller screws. Scientific American, 56(7):99, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n7/pdf/scientificamerican02121887-
99a.pdf.

Matthews:1887:HNI

[9397] W. Matthews. How the Navajo Indians make blankets. Scientific American, 56(7):99–100, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n7/pdf/scientificamerican02121887-
99b.pdf.

Anonymous:1887:ISDb

[9398] Anonymous. Improved steam digger. Scientific American, 56 (7):102, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n7/pdf/scientificamerican02121887-
102.pdf.

Smith:1887:HLD

[9399] T. Berry Smith. How large does it appear? Scientific American, 56(7):102, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n7/pdf/scientificamerican02121887-
102a.pdf.

Anonymous:1887:AB

[9400] Anonymous. The Australian bustard. Scientific American, 56 (7):103, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n7/pdf/scientificamerican02121887-
103.pdf.

Anonymous:1887:BFb

[9401] Anonymous. The built up file. Scientific American, 56(7):104, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n7/pdf/scientificamerican02121887-104.pdf.

Anonymous:1887:ITa

[9402] Anonymous. Improved tether. Scientific American, 56(7):104, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n7/pdf/scientificamerican02121887-104a.pdf.

Anonymous:1887:STm

[9403] Anonymous. Science in toys. Scientific American, 56(7):105–106, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n7/pdf/scientificamerican02121887-105.pdf.

Anonymous:1887:EIx (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n7/pdf/scientificamerican02121887-
107.pdf.

Anonymous:1887:MIbd

[9405] Anonymous. Mechanical inventions. Scientific American, 56(7): 107, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n7/pdf/scientificamerican02121887-
107a.pdf.

Anonymous:1887:AIu

[9406] Anonymous. Agricultural inventions. Scientific American, 56(7): 107, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n7/pdf/scientificamerican02121887-
107b.pdf.

Anonymous:1887:MIbe

[9407] Anonymous. Miscellaneous inventions. Scientific American, 56 (7):107, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n7/pdf/scientificamerican02121887-
107c.pdf.

Anonymous:1887:Iba

[9408] Anonymous. To inventors. Scientific American, 56(7):107, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n7/pdf/scientificamerican02121887-107f.pdf.

Anonymous:1887:FPc

[9409] Anonymous. Foreign patents. Scientific American, 56(7):109, February 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n7/pdf/scientificamerican02121887-109.pdf.

Anonymous:1887:Cy

[9410] Anonymous. Correspondence. Scientific American, 56(8):117, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n8/pdf/scientificamerican02191887-117.pdf.

Anonymous:1887:BNBo

[9411] Anonymous. Books: New books and publications. Scientific American, 56(8):122, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n8/pdf/scientificamerican02191887-
122c.pdf.

Anonymous:1887:BPy

[9412] Anonymous. Business and personal. Scientific American, 56(8): 122, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n8/pdf/scientificamerican02191887-
122d.pdf.

Anonymous:1887:NQu

[9413] Anonymous. Notes and queries. Scientific American, 56(8):123, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n8/pdf/scientificamerican02191887-123.pdf.

Anonymous:1887:IIz

[9414] Anonymous. Index of inventions. Scientific American, 56(8): 123–124, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n8/pdf/scientificamerican02191887-
123b.pdf.

Anonymous:1887:PNPw

[9415] Anonymous. Patent notes: Patents. Scientific American, 56 (8):126, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n8/pdf/scientificamerican02191887-
126.pdf.

Anonymous:1887:RC

[9416] Anonymous. A remarkable collision. Scientific American, 56(8): 111, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n8/pdf/scientificamerican02191887-
111.pdf.

Anonymous:1887:GWS

[9417] Anonymous. Great war ships and forts. Scientific American, 56 (8):112, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n8/pdf/scientificamerican02191887-
112.pdf.

Anonymous:1887:EMV

[9418] Anonymous. Electro-motor vs. cable traction. Scientific American, 56(8):112–113, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n8/pdf/scientificamerican02191887-
112a.pdf.

Anonymous:1887:BTBb

[9419] Anonymous. The Bell Telephone before the Supreme Court. Scientific American, 56(8):113, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v56/n8/pdf/scientificamerican02191887-
113.pdf.

Anonymous:1887:DDS

[9420] Anonymous. Double-deck stock car. Scientific American, 56(8): 114, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n8/pdf/scientificamerican02191887-
114.pdf.

Anonymous:1887:ILBa

[9421] Anonymous. Improved letter box. Scientific American, 56(8): 114, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n8/pdf/scientificamerican02191887-
114a.pdf.

Anonymous:1887:IWS

[9422] Anonymous. Improved wagon stake. Scientific American, 56(8): 114, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n8/pdf/scientificamerican02191887-
114b.pdf.

Anonymous:1887:ITL

[9423] Anonymous. Improved transom lifter. Scientific American, 56 (8):114, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n8/pdf/scientificamerican02191887-
114c.pdf.

Anonymous:1887:CTd

[9424] Anonymous. Combination tool. Scientific American, 56(8):114, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n8/pdf/scientificamerican02191887-114d.pdf.

Anonymous:1887:LR

[9425] Anonymous. The Lartigue Railway. Scientific American, 56(8): 115, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n8/pdf/scientificamerican02191887-
115.pdf.

Anonymous:1887:AVN

[9426] Anonymous. The Alva — Vanderbilt’s new Yacht. Scientific American, 56(8):115–116, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n8/pdf/scientificamerican02191887-
115a.pdf.

Anonymous:1887:ITb

[9427] Anonymous. Improved tent. Scientific American, 56(8):116, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n8/pdf/scientificamerican02191887-116.pdf.

Anonymous:1887:ISS

[9428] Anonymous. Improved sheep shearing table. Scientific American, 56(8):116, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n8/pdf/scientificamerican02191887-
116a.pdf.

Anonymous:1887:GHS

[9429] Anonymous. Garment and hose supporter. Scientific American, 56(8):116, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n8/pdf/scientificamerican02191887-
116b.pdf.

Anonymous:1887:Pb

[9430] Anonymous. A puzzle. Scientific American, 56(8):117–118, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n8/pdf/scientificamerican02191887-117a.pdf.

Anonymous:1887:LMT

[9431] Anonymous. Londe’s method of timing photographic exposures. Scientific American, 56(8):118, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v56/n8/pdf/scientificamerican02191887-
118.pdf.

Anonymous:1887:MDM

[9432] Anonymous. Multiple drilling machine. Scientific American, 56 (8):119, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n8/pdf/scientificamerican02191887-
119.pdf.

Anonymous:1887:STn

[9433] Anonymous. Science in toys. Scientific American, 56(8):120–121, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n8/pdf/scientificamerican02191887-120.pdf.

Anonymous:1887:DC

[9434] Anonymous. Devon cattle. Scientific American, 56(8):121, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n8/pdf/scientificamerican02191887-121.pdf.

Anonymous:1887:EIy

[9435] Anonymous. Engineering inventions. Scientific American, 56(8): 122, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n8/pdf/scientificamerican02191887-
122.pdf.

Anonymous:1887:MIbf

[9436] Anonymous. Mechanical inventions. Scientific American, 56(8): 122, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n8/pdf/scientificamerican02191887-
122a.pdf.

Anonymous:1887:MIbg

[9437] Anonymous. Miscellaneous inventions. Scientific American, 56 (8):122, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n8/pdf/scientificamerican02191887-
122b.pdf.

Anonymous:1887:Ibb

[9438] Anonymous. To inventors. Scientific American, 56(8):123, February 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n8/pdf/scientificamerican02191887-123a.pdf.

Anonymous:1887:Cz

[9439] Anonymous. Correspondence. Scientific American, 56(9):133, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n9/pdf/scientificamerican02261887-133.pdf.

Anonymous:1887:BNBp

[9440] Anonymous. Books: New books and publications. Scientific American, 56(9):138, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
138c.pdf.

Anonymous:1887:BPz

[9441] Anonymous. Business and personal. Scientific American, 56(9): 138, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
138d.pdf.

Anonymous:1887:NQv (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
138e.pdf.

Anonymous:1887:IIbb

[9443] Anonymous. Index of inventions. Scientific American, 56(9): 139–140, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
139a.pdf.

Anonymous:1887:PNPx

[9444] Anonymous. Patent notes: Patents. Scientific American, 56 (9):142, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
142.pdf.

Anonymous:1887:ICE

[9445] Anonymous. Improved caloric engine. Scientific American, 56 (9):127, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
127.pdf.

Anonymous:1887:NAD

[9446] Anonymous. New amendment of the design patent law. Scientific American, 56(9):128, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
128.pdf.

Anonymous:1887:SGR

[9447] Anonymous. Are steel guns really superior? Scientific American, 56(9):128, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
128a.pdf.

Anonymous:1887:CHa

[9448] Anonymous. The cocaine habit. Scientific American, 56(9):128– 129, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
128b.pdf.

Anonymous:1887:EWG

[9449] Anonymous. Emery wheels for gumming saws. Scientific American, 56(9):130, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
130.pdf.

Anonymous:1887:FSO

[9450] Anonymous. The frilled shark — the oldest living type of vertebrates. Scientific American, 56(9):130, February 26, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v56/
n9/pdf/scientificamerican02261887-130a.pdf.

Anonymous:1887:CBI

[9451] Anonymous. Combined bench and ironing board. Scientific American, 56(9):131, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
131.pdf.

Anonymous:1887:IBSa

[9452] Anonymous. Improved blind stop. Scientific American, 56(9): 131, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
131a.pdf.

Anonymous:1887:SDC

[9453] Anonymous. A simple device for crimping boots or shoes. Scientific American, 56(9):131, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
131b.pdf.

Anonymous:1887:PH

[9454] Anonymous. Plasterer’s hawk. Scientific American, 56(9):131, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n9/pdf/scientificamerican02261887-131c.pdf.

Anonymous:1887:AWM

[9455] Anonymous. Adjustable wood measuring rack. Scientific American, 56(9):131–132, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
131d.pdf.

Anonymous:1887:PIb

[9456] Anonymous. Penberthy injector. Scientific American, 56(9):132, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n9/pdf/scientificamerican02261887-132.pdf.

Anonymous:1887:PLH

[9457] Anonymous. Protector for ladies’ hats. Scientific American, 56 (9):132, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
132a.pdf.

Anonymous:1887:ISPb

[9458] Anonymous. Improved stump puller. Scientific American, 56 (9):132, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
132b.pdf.

Anonymous:1887:CBCb

[9459] Anonymous. Combined brush and comb cleaner. Scientific American, 56(9):132, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
132c.pdf.

Hasbrouck:1887:OWB

[9460] B. M. Hasbrouck. Our winter birds. Scientific American, 56 (9):133, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
133a.pdf.

Hopkins:1887:TB

[9461] Geo. M. Hopkins. Thermoscopic balance. Scientific American, 56(9):134, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
134.pdf.

Anonymous:1887:IMD

[9462] Anonymous. Improved marine dredger. Scientific American, 56 (9):134, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
134a.pdf.

Anonymous:1887:NL

[9463] Anonymous. The Northern Lights. Scientific American, 56(9): 135–136, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
135.pdf.

Anonymous:1887:ETI

[9464] Anonymous. Effect of a torpedo on an ironclad. Scientific American, 56(9):136, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
136.pdf.

Anonymous:1887:TBA

[9465] Anonymous. Torpedo boat armed with pneumatic dynamite guns. Scientific American, 56(9):137, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
137.pdf.

Sloane:1887:CB

[9466] T. O’conor Sloane. Chevreul’s black. Scientific American, 56 (9):137, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
137a.pdf.

Anonymous:1887:EIz

[9467] Anonymous. Engineering inventions. Scientific American, 56(9): 138, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
138.pdf.

Anonymous:1887:AIv

[9468] Anonymous. Agricultural inventions. Scientific American, 56(9): 138, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
138a.pdf.

Anonymous:1887:MIbh

[9469] Anonymous. Miscellaneous inventions. Scientific American, 56 (9):138, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v56/n9/pdf/scientificamerican02261887-
138b.pdf.

Anonymous:1887:Ibc

[9470] Anonymous. To inventors. Scientific American, 56(9):139, February 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v56/n9/pdf/scientificamerican02261887-139.pdf.

Anonymous:1887:Cbb

[9471] Anonymous. Correspondence. Scientific American, 57(1):7, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n1/pdf/scientificamerican07021887-7.pdf.

Anonymous:1887:BNBq

[9472] Anonymous. Books: New books and publications. Scientific American, 57(1):10, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n1/pdf/scientificamerican07021887-
10c.pdf.

Anonymous:1887:BPba

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n1/pdf/scientificamerican07021887-10e.pdf.

Anonymous:1887:NQw

[9474] Anonymous. Notes and queries. Scientific American, 57(1):11, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n1/pdf/scientificamerican07021887-11.pdf.

Anonymous:1887:IIbc

[9475] Anonymous. Index of inventions. Scientific American, 57(1):11–12, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n1/pdf/scientificamerican07021887-11a.pdf.

Anonymous:1887:PNPy

[9476] Anonymous. Patent notes: Patents. Scientific American, 57(1):14, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n1/pdf/scientificamerican07021887-14.pdf.

Anonymous:1887:CCd

[9477] Anonymous. Corsets and consumption. Scientific American, 57(1):2, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n1/pdf/scientificamerican07021887-2.pdf.

Anonymous:1887:OSN

[9478] Anonymous. Our slow new Navy. Scientific American, 57(1):2, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n1/pdf/scientificamerican07021887-2a.pdf.

Anonymous:1887:PNg

[9479] Anonymous. Photographic notes. Scientific American, 57(1):3, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n1/pdf/scientificamerican07021887-3.pdf.

Anonymous:1887:IPP

[9480] Anonymous. An improved Pitman or pump rod. Scientific American, 57(1):4, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n1/pdf/scientificamerican07021887-
4.pdf.

Anonymous:1887:IUPa

[9481] Anonymous. An improved universal pantograph. Scientific American, 57(1):4, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n1/pdf/scientificamerican07021887-
4a.pdf.

Anonymous:1887:RSV

[9482] Anonymous. A revolving self-ventilating churn. Scientific American, 57(1):4, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n1/pdf/scientificamerican07021887-
4b.pdf.

Anonymous:1887:ICCh

[9483] Anonymous. An improved car coupling. Scientific American, 57(1):4, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n1/pdf/scientificamerican07021887-4c.pdf.

Anonymous:1887:CCW

[9484] Anonymous. A combined carpet and weather strip. Scientific American, 57(1):4, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n1/pdf/scientificamerican07021887-
4d.pdf.

Anonymous:1887:EAA

[9485] Anonymous. An easily adjusted awning for doors or windows. Scientific American, 57(1):5, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n1/pdf/scientificamerican07021887-
5.pdf.

Anonymous:1887:IJa

[9486] Anonymous. An improved jack. Scientific American, 57(1):5, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n1/pdf/scientificamerican07021887-5a.pdf.

Anonymous:1887:IFS

[9487] Anonymous. An improved feeder for stoves. Scientific American, 57(1): 5, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n1/pdf/scientificamerican07021887-5b.pdf.

Anonymous:1887:MLG

[9488] Anonymous. Manufacture of large guns. Scientific American, 57(1):5–6, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n1/pdf/scientificamerican07021887-5c.pdf.

Anonymous:1887:NBR

[9489] Anonymous. The new British racing yacht Thistle. Scientific American, 57(1):6, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n1/pdf/scientificamerican07021887-
6.pdf.

Anonymous:1887:WCR

[9490] Anonymous. A water cooled refrigerator. Scientific American, 57(1):7, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n1/pdf/scientificamerican07021887-7a.pdf.

Anonymous:1887:IFPa

[9491] Anonymous. Improved filter press pump. Scientific American, 57(1):8, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n1/pdf/scientificamerican07021887-8.pdf.

Anonymous:1887:DBC

[9492] Anonymous. Dutch belted cattle. Scientific American, 57(1):8, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n1/pdf/scientificamerican07021887-8a.pdf.

Anonymous:1887:FF

[9493] Anonymous. Flying foxes. Scientific American, 57(1):9, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n1/pdf/scientificamerican07021887-9.pdf.

Anonymous:1887:EIba

[9494] Anonymous. Engineering inventions. Scientific American, 57(1):10, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n1/pdf/scientificamerican07021887-10.pdf.

Anonymous:1887:AIw

[9495] Anonymous. Agricultural inventions. Scientific American, 57(1):10, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n1/pdf/scientificamerican07021887-10a.pdf.

Anonymous:1887:MIbi

[9496] Anonymous. Miscellaneous inventions. Scientific American, 57(1):10, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n1/pdf/scientificamerican07021887-10b.pdf.

Anonymous:1887:Sd

[9497] Anonymous. Special. Scientific American, 57(1):10, July 2, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v57/
n1/pdf/scientificamerican07021887-10d.pdf.

Anonymous:1887:BPbb

[9498] Anonymous. Business and personal. Scientific American, 57 (10):154, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
154c.pdf.

Anonymous:1887:NQx

[9499] Anonymous. Notes and queries. Scientific American, 57(10): 155, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
155.pdf.

Anonymous:1887:IIbd

[9500] Anonymous. Index of inventions. Scientific American, 57(10): 155–156, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
155b.pdf.

Anonymous:1887:PNPz

[9501] Anonymous. Patent notes: Patents. Scientific American, 57 (10):158, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
158.pdf.

Anonymous:1887:GBC

[9502] Anonymous. A great bronze casting. Scientific American, 57 (10):143, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
143.pdf.

Anonymous:1887:ISTa

[9503] Anonymous. Isthmian ship transit. Scientific American, 57(10): 144, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
144.pdf.

Anonymous:1887:DGS

[9504] Anonymous. Duplication of government scientific work. Scientific American, 57(10):144–145, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
144a.pdf.

Anonymous:1887:CWe

[9505] Anonymous. The celestial world. Scientific American, 57(10): 145, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
145.pdf.

Anonymous:1887:CDS

[9506] Anonymous. A convenient device in sharpening saws. Scientific American, 57(10):146, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
146.pdf.

Anonymous:1887:IMAa

[9507] Anonymous. Improved mortar attachments. Scientific American, 57(10):146, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
146a.pdf.

Anonymous:1887:ICg

[9508] Anonymous. An improved cultivator. Scientific American, 57 (10):146, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
146b.pdf.

Anonymous:1887:FCB

[9509] Anonymous. A folding center board for small boats. Scientific American, 57(10):146, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
146c.pdf.

Anonymous:1887:SIH

[9510] Anonymous. A simple and inexpensive hat hook. Scientific American, 57(10):146, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
146d.pdf.

Anonymous:1887:ICSd

[9511] Anonymous. An improved canopy-supporting frame. Scientific American, 57(10):147, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
147.pdf.

Anonymous:1887:IFPb

[9512] Anonymous. An improved finder for photographic cameras. Scientific American, 57(10):147, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
147a.pdf.

Anonymous:1887:IABa

[9513] Anonymous. An improved adjustable book holder. Scientific American, 57(10):147, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
147b.pdf.

Anonymous:1887:IAG

[9514] Anonymous. An improved axle gauge. Scientific American, 57 (10):147, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
147c.pdf.

Anonymous:1887:IMAb

[9515] Anonymous. An improved mop attachment. Scientific American, 57(10):147, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
147d.pdf.

Sloane:1887:FB

[9516] T. O’conor Sloane. The floating balance. Scientific American, 57 (10):149, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
149.pdf.

Anonymous:1887:FAP

[9517] Anonymous. The Fire Apparatus of Paris. Scientific American, 57(10):150, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
150.pdf.

Anonymous:1887:APM

[9518] Anonymous. An alleged perpetual motion. Scientific American, 57(10):151, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
151.pdf.

Benjamin:1887:SFB

[9519] Marous Benjamin. Spencer Fullerton Baird. Scientific American, 57 (10):151–152, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
151a.pdf.

Seiss:1887:SLG

[9520] C. Few Seiss. The shake lizard, glass snake, or joint snake. Scientific American, 57(10):152, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
152.pdf.

Hopkins:1887:CV

[9521] M. Hopkins. Composition of vibrations. Scientific American, 57 (10):153, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
153.pdf.

Anonymous:1887:EIbb

[9522] Anonymous. Engineering inventions. Scientific American, 57(10): 154, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
154.pdf.

Anonymous:1887:AIx

[9523] Anonymous. Agricultural invention. Scientific American, 57(10): 154, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
154a.pdf.

Anonymous:1887:MIbj

[9524] Anonymous. Miscellaneous inventions. Scientific American, 57 (10):154, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n10/pdf/scientificamerican09031887-
154b.pdf.

Anonymous:1887:Ibe

[9525] Anonymous. To inventors. Scientific American, 57(10):155, September 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n10/pdf/scientificamerican09031887-155a.pdf.

Anonymous:1887:Cbd

[9526] Anonymous. Correspondence. Scientific American, 57(11):165, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n11/pdf/scientificamerican09101887-165.pdf.

Anonymous:1887:BPbc

[9527] Anonymous. Business and personal. Scientific American, 57 (11):171, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n11/pdf/scientificamerican09101887-
171d.pdf.

Anonymous:1887:BNBr

[9528] Anonymous. Books: New books and publications. Scientific American, 57(11):171, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n11/pdf/scientificamerican09101887-
171e.pdf.

Anonymous:1887:IIbe

[9529] Anonymous. Index of inventions. Scientific American, 57(11): 171–172, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n11/pdf/scientificamerican09101887-
171g.pdf.

Anonymous:1887:PNPba

[9530] Anonymous. Patent notes: Patents. Scientific American, 57 (11):174, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n11/pdf/scientificamerican09101887-
174.pdf.

Anonymous:1887:BSL

[9531] Anonymous. The Barrett system of lighting railroad cars by electricity. Scientific American, 57(11):159–160, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v57/
n11/pdf/scientificamerican09101887-159.pdf.

Anonymous:1887:RI

[9532] Anonymous. Restricted immigration. Scientific American, 57 (11):160, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n11/pdf/scientificamerican09101887-
160.pdf.

Anonymous:1887:BNM

[9533] Anonymous. British Naval maneuvers. Scientific American, 57(11): 160–162, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n11/pdf/scientificamerican09101887-
160a.pdf.

Anonymous:1887:IOT

[9534] Anonymous. An improved overhead twine holder. Scientific American, 57(11):162, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n11/pdf/scientificamerican09101887-
162.pdf.

Anonymous:1887:BTT

[9535] Anonymous. The brush turkey (Talegallus Lathami). Scientific American, 57(11):162, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n11/pdf/scientificamerican09101887-
162a.pdf.

Anonymous:1887:DHJ

[9536] Anonymous. A device for heating a jet of air. Scientific American, 57(11):162, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n11/pdf/scientificamerican09101887-
162b.pdf.

Anonymous:1887:SHS

[9537] Anonymous. A spring holder for securing vines. Scientific American, 57(11):163, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n11/pdf/scientificamerican09101887-
163.pdf.

Anonymous:1887:ISBc

[9538] Anonymous. An improved straw burning stove. Scientific American, 57(11):163, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n11/pdf/scientificamerican09101887-
163a.pdf.

Anonymous:1887:ICSe

[9539] Anonymous. An improved car starter. Scientific American, 57 (11):163, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n11/pdf/scientificamerican09101887-
163b.pdf.

Anonymous:1887:IMCa

[9540] Anonymous. An improved machine for coloring paper. Scientific American, 57(11):163, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n11/pdf/scientificamerican09101887-
163c.pdf.

Anonymous:1887:INS

[9541] Anonymous. An improved nail set. Scientific American, 57 (11):163, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n11/pdf/scientificamerican09101887-
163d.pdf.

Anonymous:1887:ILJ

[9542] Anonymous. Improved locomotive journal box lubricator. Scientific American, 57(11):163–164, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v57/n11/pdf/scientificamerican09101887-
163e.pdf.

Anonymous:1887:PNh

[9543] Anonymous. Photographic notes. Scientific American, 57(11): 165, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n11/pdf/scientificamerican09101887-
165a.pdf.

Hopkins:1887:SPE

[9544] M. Hopkins. Simple process of engraving glass and metals. Scientific American, 57(11):166, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v57/n11/pdf/scientificamerican09101887-
166.pdf.

Anonymous:1887:YT

[9545] Anonymous. The yacht Thistle. Scientific American, 57(11): 167, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n11/pdf/scientificamerican09101887-
167.pdf.

Sloane:1887:AIS

[9546] T. O’conor Sloane. Apparatus for illustrating the specific gravity of liquids. Scientific American, 57(11):167–169, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n11/pdf/scientificamerican09101887-167a.pdf.

Edison:1887:MBB

[9547] Thomas A. Edison. On a magnetic bridge or balance for measuring magnetic conductivity. Scientific American, 57(11):169–170, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n11/pdf/scientificamerican09101887-169.pdf.

Anonymous:1887:EIbc

[9548] Anonymous. Engineering inventions. Scientific American, 57(11): 171, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n11/pdf/scientificamerican09101887-
171.pdf.

Anonymous:1887:AIy

[9549] Anonymous. Agricultural inventions. Scientific American, 57 (11):171, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n11/pdf/scientificamerican09101887-
171a.pdf.

Anonymous:1887:MIbk

[9550] Anonymous. Miscellaneous inventions. Scientific American, 57 (11):171, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n11/pdf/scientificamerican09101887-
171b.pdf.

Anonymous:1887:Se

[9551] Anonymous. Special. Scientific American, 57(11):171, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n11/pdf/scientificamerican09101887-171c.pdf.

Anonymous:1887:Ibg

[9552] Anonymous. To inventors. Scientific American, 57(11):171, September 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n11/pdf/scientificamerican09101887-171f.pdf.

Anonymous:1887:Cbe

[9553] Anonymous. Correspondence. Scientific American, 57(12):181, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n12/pdf/scientificamerican09171887-181.pdf.

Anonymous:1887:BPbd

[9554] Anonymous. Business and personal. Scientific American, 57 (12):186, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
186c.pdf.

Anonymous:1887:BNBs

[9555] Anonymous. Books: New books and publications. Scientific American, 57(12):186–187, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
186d.pdf.

Anonymous:1887:NQy

[9556] Anonymous. Notes and queries. Scientific American, 57(12): 187, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
187.pdf.

Anonymous:1887:IIbf

[9557] Anonymous. Index of inventions. Scientific American, 57(12): 187–188, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
187a.pdf.

Anonymous:1887:PNPbb

[9558] Anonymous. Patent notes: Patents. Scientific American, 57 (12):190, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
190.pdf.

Anonymous:1887:EMBb

[9559] Anonymous. Electro-motor batteries. Scientific American, 57(12): 176–177, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
176.pdf.

Anonymous:1887:ARC

[9560] Anonymous. An automatic railway crossing gate. Scientific American, 57(12):178, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
178.pdf.

Anonymous:1887:IDCc

[9561] Anonymous. An improved dumping car. Scientific American, 57 (12):178, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
178a.pdf.

Anonymous:1887:IER

[9562] Anonymous. An improved elevated railway gate. Scientific American, 57(12):178, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
178b.pdf.

Anonymous:1887:ICCi

[9563] Anonymous. An improved car coupling. Scientific American, 57 (12):178, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
178c.pdf.

Anonymous:1887:ISCc

[9564] Anonymous. An improved shaft coupling. Scientific American, 57 (12):178, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
178d.pdf.

Anonymous:1887:PTCa

[9565] Anonymous. A pattern tracing and cutting device. Scientific American, 57(12):179, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
179.pdf.

Anonymous:1887:IPTa

[9566] Anonymous. An improved pipe tongs or wrench. Scientific American, 57(12):179, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
179a.pdf.

Anonymous:1887:IBCa

[9567] Anonymous. An improved bell cord coupling. Scientific American, 57(12):179, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
179b.pdf.

Anonymous:1887:ISFc

[9568] Anonymous. An improved saw filing machine. Scientific American, 57(12):179, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
179c.pdf.

Anonymous:1887:IABb

[9569] Anonymous. Improved attachment for blinding horses. Scientific American, 57(12):179, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
179d.pdf.

Anonymous:1887:IBH

[9570] Anonymous. An improved bit for horses. Scientific American, 57 (12):179, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
179e.pdf.

Anonymous:1887:ONNa

[9571] Anonymous. Our new Navy — the cruiser Chicago. Scientific American, 57(12):180, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
180.pdf.

Anonymous:1887:OBC

[9572] Anonymous. The Olbers–Brooks Comet. Scientific American, 57 (12):181, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
181a.pdf.

Anonymous:1887:FML

[9573] Anonymous. Fireless mining locomotive. Scientific American, 57 (12):182, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
182.pdf.

Anonymous:1887:LDS

[9574] Anonymous. Light draught steam launches. Scientific American, 57(12):182, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
182a.pdf.

Anonymous:1887:SY

[9575] Anonymous. A Singhalese yacht. Scientific American, 57(12): 183–184, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
183.pdf.

Anonymous:1887:IIPb

[9576] Anonymous. An inexpensive index plate. Scientific American, 57 (12):184–185, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
184.pdf.

Anonymous:1887:EIbd

[9577] Anonymous. Engineering inventions. Scientific American, 57(12): 186, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
186.pdf.

Anonymous:1887:AIz

[9578] Anonymous. Agricultural invention. Scientific American, 57(12): 186, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
186a.pdf.

Anonymous:1887:MIbl

[9579] Anonymous. Miscellaneous inventions. Scientific American, 57 (12):186, September 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n12/pdf/scientificamerican09171887-
186b.pdf.

Anonymous:1887:Cbf

[9580] Anonymous. Correspondence. Scientific American, 57(13):197, September 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n13/pdf/scientificamerican09241887-197.pdf.

Anonymous:1887:BPbe

[9581] Anonymous. Business and personal. Scientific American, 57 (13):202, September 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n13/pdf/scientificamerican09241887-
202d.pdf.

Anonymous:1887:NQz

[9582] Anonymous. Notes and queries. Scientific American, 57(13): 202–203, September 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n13/pdf/scientificamerican09241887-
202e.pdf.

Anonymous:1887:IIbg

[9583] Anonymous. Index of inventions. Scientific American, 57(13): 203–204, September 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n13/pdf/scientificamerican09241887-
203a.pdf.

Anonymous:1887:PNPbc

[9584] Anonymous. Patent notes: Patents. Scientific American, 57 (13):206, September 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n13/pdf/scientificamerican09241887-
206.pdf.

Anonymous:1887:ROC

[9585] Anonymous. The return of Olbers’ Comet of 1815. Scientific American, 57(13):192, September 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n13/pdf/scientificamerican09241887-
192.pdf.

Anonymous:1887:SGb

[9586] Anonymous. For star gazers. Scientific American, 57(13):193, September 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n13/pdf/scientificamerican09241887-193.pdf.

Anonymous:1887:LPB

[9587] Anonymous. A low pressure boiler for steam heating. Scientific American, 57(13):194, September 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n13/pdf/scientificamerican09241887-
194.pdf.

Anonymous:1887:IPAa

[9588] Anonymous. An improved pulverizing attachment for seeders. Scientific American, 57(13):194, September 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v57/n13/pdf/scientificamerican09241887-
194a.pdf.

Anonymous:1887:IASb

[9589] Anonymous. An improved adjustable school seat. Scientific American, 57(13):194, September 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n13/pdf/scientificamerican09241887-
194b.pdf.

Anonymous:1887:CHR

[9590] Anonymous. A convenient holder for ribbons, edgings, etc. Scientific American, 57(13):194, September 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v57/n13/pdf/scientificamerican09241887-
194c.pdf.

Anonymous:1887:RRC

[9591] Anonymous. A ready-reckoning coin package. Scientific American, 57(13):194, September 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n13/pdf/scientificamerican09241887-
194d.pdf.

Anonymous:1887:SWD

[9592] Anonymous. Splendid work done by rubber belts. Scientific American, 57(13):195–196, September 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n13/pdf/scientificamerican09241887-
195.pdf.

Anonymous:1887:ACE

[9593] Anonymous. The Alvan Clark establishment. Scientific American, 57 (13):198–200, September 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n13/pdf/scientificamerican09241887-
198.pdf.

Anonymous:1887:TEE

[9594] Anonymous. Triple expansion engines. Scientific American, 57 (13):201, September 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n13/pdf/scientificamerican09241887-
201.pdf.

Anonymous:1887:EIbe

[9595] Anonymous. Engineering inventions. Scientific American, 57(13): 202, September 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n13/pdf/scientificamerican09241887-
202.pdf.

Anonymous:1887:AIba

[9596] Anonymous. Agricultural invention. Scientific American, 57(13): 202, September 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n13/pdf/scientificamerican09241887-
202a.pdf.

Anonymous:1887:MIbm

[9597] Anonymous. Miscellaneous inventions. Scientific American, 57 (13):202, September 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n13/pdf/scientificamerican09241887-
202b.pdf.

Anonymous:1887:Sf

[9598] Anonymous. Special. Scientific American, 57(13):202, September 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n13/pdf/scientificamerican09241887-202c.pdf.

Anonymous:1887:Ibi

[9599] Anonymous. To inventors. Scientific American, 57(13):203, September 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n13/pdf/scientificamerican09241887-203.pdf.

Anonymous:1887:Cbg

[9600] Anonymous. Correspondence. Scientific American, 57(14):213, October 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n14/pdf/scientificamerican10011887-213.pdf.

Anonymous:1887:BPbf

[9601] Anonymous. Business and personal. Scientific American, 57(14):219, October 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n14/pdf/scientificamerican10011887-219c.pdf.

Anonymous:1887:IIbh

[9602] Anonymous. Index of inventions. Scientific American, 57(14):219–220, October 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n14/pdf/scientificamerican10011887-219e.pdf.

Anonymous:1887:PNPbd

[9603] Anonymous. Patent notes: Patents. Scientific American, 57(14):222, October 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n14/pdf/scientificamerican10011887-222.pdf.

Anonymous:1887:CSCb

[9604] Anonymous. The centennial of the signing of the Constitution of the United States. Scientific American, 57(14):208, October 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n14/pdf/scientificamerican10011887-208.pdf.

Anonymous:1887:PPO

[9605] Anonymous. Positions of the planets in October. Scientific American, 57(14):208–209, October 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n14/pdf/scientificamerican10011887-
208a.pdf.

Anonymous:1887:IPD

[9606] Anonymous. An improved potato digger. Scientific American, 57(14):210, October 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n14/pdf/scientificamerican10011887-
210.pdf.

Anonymous:1887:VPB

[9607] Anonymous. Vertical plate bending rolls. Scientific American, 57(14):210, October 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n14/pdf/scientificamerican10011887-
210a.pdf.

Anonymous:1887:IOCc

[9608] Anonymous. An improved oil can holder. Scientific American, 57(14):211, October 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n14/pdf/scientificamerican10011887-
211.pdf.

Anonymous:1887:ISL

[9609] Anonymous. An improved seal lock. Scientific American, 57(14):211, October 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n14/pdf/scientificamerican10011887-211a.pdf.

Anonymous:1887:IMS

[9610] Anonymous. An improved metallic shingle. Scientific American, 57(14):211, October 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n14/pdf/scientificamerican10011887-
211b.pdf.

Anonymous:1887:IGFb

[9611] Anonymous. An improved glove fastener. Scientific American, 57(14):211, October 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n14/pdf/scientificamerican10011887-
211c.pdf.

Anonymous:1887:DRK

[9612] Anonymous. A device to retain keys in locks. Scientific American, 57(14):211–212, October 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n14/pdf/scientificamerican10011887-
211d.pdf.

Anonymous:1887:PTP

[9613] Anonymous. Practical trial of the pneumatic dynamite gun. Scientific American, 57(14):214, October 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n14/pdf/scientificamerican10011887-
214.pdf.

Anonymous:1887:BPbg

[9614] Anonymous. Birds of paradise. Scientific American, 57(14):215, October 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n14/pdf/scientificamerican10011887-215.pdf.

Anonymous:1887:KGG

[9615] Anonymous. Krupp’s great gun for the Italian Navy. Scientific American, 57(14):216–217, October 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n14/pdf/scientificamerican10011887-
216.pdf.

Anonymous:1887:FBMb

[9616] Anonymous. Fine Belgian metal work. Scientific American, 57(14):217, October 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n14/pdf/scientificamerican10011887-217.pdf.

Anonymous:1887:AWS

[9617] Anonymous. Artistic work in silver and bronze. Scientific American, 57(14):217–218, October 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n14/pdf/scientificamerican10011887-
217a.pdf.

Anonymous:1887:EIbf

[9618] Anonymous. Engineering inventions. Scientific American, 57(14):219, October 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n14/pdf/scientificamerican10011887-219.pdf.

Anonymous:1887:AIbb

[9619] Anonymous. Agricultural invention. Scientific American, 57(14):219, October 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n14/pdf/scientificamerican10011887-219a.pdf.

Anonymous:1887:MIbn

[9620] Anonymous. Miscellaneous inventions. Scientific American, 57(14):219, October 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n14/pdf/scientificamerican10011887-219b.pdf.

Anonymous:1887:Ibk

[9621] Anonymous. To inventors. Scientific American, 57(14):219, October 1, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n14/pdf/scientificamerican10011887-219d.pdf.

Anonymous:1887:Cbh

[9622] Anonymous. Correspondence. Scientific American, 57(15):229, October 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n15/pdf/scientificamerican10081887-229.pdf.

Anonymous:1887:BNBt

[9623] Anonymous. Books: New books and publications. Scientific American, 57(15):235, October 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n15/pdf/scientificamerican10081887-
235c.pdf.

Anonymous:1887:BPbh

[9624] Anonymous. Business and personal. Scientific American, 57(15):235, October 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n15/pdf/scientificamerican10081887-235e.pdf.

Anonymous:1887:NQba

[9625] Anonymous. Notes and queries. Scientific American, 57(15):235, October 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n15/pdf/scientificamerican10081887-235f.pdf.

Anonymous:1887:IIbi

[9626] Anonymous. Index of inventions. Scientific American, 57(15):235–236, October 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n15/pdf/scientificamerican10081887-235h.pdf.

Anonymous:1887:PNPbe

[9627] Anonymous. Patent notes: Patents. Scientific American, 57(15):238, October 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n15/pdf/scientificamerican10081887-238.pdf.

Anonymous:1887:TEA

[9628] Anonymous. The total eclipse of August 19. Scientific American, 57(15):224, October 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n15/pdf/scientificamerican10081887-
224.pdf.

Anonymous:1887:PNi

[9629] Anonymous. Photographic notes. Scientific American, 57(15):225, October 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n15/pdf/scientificamerican10081887-225.pdf.

Anonymous:1887:ILI

[9630] Anonymous. Improved lamp and incombustible burner. Scientific American, 57(15):226, October 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n15/pdf/scientificamerican10081887-
226.pdf.

Anonymous:1887:DKV

[9631] Anonymous. A device to keep vise jaws parallel. Scientific American, 57(15):226, October 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n15/pdf/scientificamerican10081887-
226a.pdf.

Anonymous:1887:SPM

[9632] Anonymous. A safety package for money. Scientific American, 57(15):226, October 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n15/pdf/scientificamerican10081887-
226b.pdf.

Anonymous:1887:IBSb

[9633] Anonymous. An improved belt shifting device. Scientific American, 57(15):226, October 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n15/pdf/scientificamerican10081887-
226c.pdf.

Anonymous:1887:RAC

[9634] Anonymous. The race for the America’s Cup. Scientific American, 57(15):227, October 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n15/pdf/scientificamerican10081887-
227.pdf.

Anonymous:1887:SRM

[9635] Anonymous. A self-registering measure for lumber. Scientific American, 57(15):228, October 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n15/pdf/scientificamerican10081887-
228.pdf.

Anonymous:1887:DRPe

[9636] Anonymous. Decisions relating to patents. Scientific American, 57(15):229, October 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n15/pdf/scientificamerican10081887-
229a.pdf.

Anonymous:1887:EED

[9637] Anonymous. Experiments on the effects of detonating gun cotton at the U.S. Torpedo Station, Newport, R.I. Scientific American, 57(15):230–231, October 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n15/pdf/scientificamerican10081887-
230.pdf.

Anonymous:1887:GFB

[9638] Anonymous. George Frederic Barker. Scientific American, 57 (15):231–232, October 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n15/pdf/scientificamerican10081887-
231.pdf.

Anonymous:1887:SGH

[9639] Anonymous. The staircase of a great house. Scientific American, 57(15):233, October 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n15/pdf/scientificamerican10081887-
233.pdf.

Forbes:1887:NEC

[9640] G. Forbes. A new electric current meter. Scientific American, 57 (15):233–234, October 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n15/pdf/scientificamerican10081887-
233a.pdf.

Anonymous:1887:EIbg

[9641] Anonymous. Engineering inventions. Scientific American, 57(15):235, October 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n15/pdf/scientificamerican10081887-235.pdf.

Anonymous:1887:AIbc

[9642] Anonymous. Agricultural inventions. Scientific American, 57(15):235, October 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n15/pdf/scientificamerican10081887-235a.pdf.

Anonymous:1887:MIbo

[9643] Anonymous. Miscellaneous inventions. Scientific American, 57(15):235, October 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n15/pdf/scientificamerican10081887-235b.pdf.

Anonymous:1887:Si

[9644] Anonymous. Special. Scientific American, 57(15):235, October 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n15/pdf/scientificamerican10081887-235d.pdf.

Anonymous:1887:Ibl

[9645] Anonymous. To inventors. Scientific American, 57(15):235, October 8, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n15/pdf/scientificamerican10081887-235g.pdf.

Anonymous:1887:BPbi

[9646] Anonymous. Business and personal. Scientific American, 57 (16):250, October 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n16/pdf/scientificamerican10151887-
250c.pdf.

Anonymous:1887:BNBu

[9647] Anonymous. Books: New books and publications. Scientific American, 57(16):250–251, October 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n16/pdf/scientificamerican10151887-
250d.pdf.

Anonymous:1887:NQbb

[9648] Anonymous. Notes and queries. Scientific American, 57(16):251, October 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n16/pdf/scientificamerican10151887-251.pdf.

Anonymous:1887:IIbj

[9649] Anonymous. Index of inventions. Scientific American, 57(16): 251–252, October 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n16/pdf/scientificamerican10151887-
251b.pdf.

Anonymous:1887:PNPbf

[9650] Anonymous. Patent notes: Patents. Scientific American, 57 (16):254, October 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n16/pdf/scientificamerican10151887-
254.pdf.

Anonymous:1887:HOHa

[9651] Anonymous. Harvard Observatory and the Henry Draper Memorial. Scientific American, 57(16):239, October 15, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v57/
n16/pdf/scientificamerican10151887-239.pdf.

Anonymous:1887:RMB

[9652] Anonymous. Railroad men’s building. Scientific American, 57 (16):240, October 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n16/pdf/scientificamerican10151887-
240.pdf.

Anonymous:1887:ETT

[9653] Anonymous. Exhibition of train telegraphy. Scientific American, 57(16):240, October 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n16/pdf/scientificamerican10151887-
240a.pdf.

Anonymous:1887:MR

[9654] Anonymous. Magazine rifles. Scientific American, 57(16):240–241, October 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n16/pdf/scientificamerican10151887-240b.pdf.

Anonymous:1887:PNj

[9655] Anonymous. Photographic notes. Scientific American, 57(16):241, October 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n16/pdf/scientificamerican10151887-241.pdf.

Anonymous:1887:VGO

[9656] Anonymous. Valve gear for oscillating engines. Scientific American, 57(16):242, October 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n16/pdf/scientificamerican10151887-
242.pdf.

Anonymous:1887:APS

[9657] Anonymous. An adjustable pitch screw propeller. Scientific American, 57(16):243, October 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n16/pdf/scientificamerican10151887-
243.pdf.

Anonymous:1887:IHA

[9658] Anonymous. An improved household ash sifter. Scientific American, 57(16):243, October 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n16/pdf/scientificamerican10151887-
243a.pdf.

Anonymous:1887:ISAa

[9659] Anonymous. An improved signaling apparatus. Scientific American, 57(16):243, October 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n16/pdf/scientificamerican10151887-
243b.pdf.

Anonymous:1887:ICH

[9660] Anonymous. An improved cutter head. Scientific American, 57 (16):243, October 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n16/pdf/scientificamerican10151887-
243c.pdf.

Anonymous:1887:IFW

[9661] Anonymous. An improved fifth wheel. Scientific American, 57 (16):243, October 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n16/pdf/scientificamerican10151887-
243d.pdf.

Anonymous:1887:IEGb

[9662] Anonymous. International Exhibition at Glasgow. Scientific American, 57(16):244–245, October 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n16/pdf/scientificamerican10151887-
244.pdf.

Anonymous:1887:MMP

[9663] Anonymous. Machines for making packing boxes. Scientific American, 57(16):246, October 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n16/pdf/scientificamerican10151887-
246.pdf.

Anonymous:1887:RWM

[9664] Anonymous. The regal walnut moth. Scientific American, 57 (16):249, October 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n16/pdf/scientificamerican10151887-
249.pdf.

Anonymous:1887:EIbh

[9665] Anonymous. Engineering inventions. Scientific American, 57(16): 250, October 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n16/pdf/scientificamerican10151887-
250.pdf.

Anonymous:1887:AIbd

[9666] Anonymous. Agricultural invention. Scientific American, 57(16): 250, October 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n16/pdf/scientificamerican10151887-
250a.pdf.

Anonymous:1887:MIbp

[9667] Anonymous. Miscellaneous inventions. Scientific American, 57 (16):250, October 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n16/pdf/scientificamerican10151887-
250b.pdf.

Anonymous:1887:Ibn

[9668] Anonymous. To inventors. Scientific American, 57(16):251, October 15, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n16/pdf/scientificamerican10151887-251a.pdf.

Anonymous:1887:Cbj

[9669] Anonymous. Correspondence. Scientific American, 57(17):261–262, October 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n17/pdf/scientificamerican10221887-261.pdf.

Anonymous:1887:BPbj

[9670] Anonymous. Business and personal. Scientific American, 57(17): 266–267, October 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n17/pdf/scientificamerican10221887-
266d.pdf.

Anonymous:1887:NQbc

[9671] Anonymous. Notes and queries. Scientific American, 57(17):267, October 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n17/pdf/scientificamerican10221887-267.pdf.

Anonymous:1887:BNBv

[9672] Anonymous. Books: New books and publications. Scientific American, 57(17):267, October 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n17/pdf/scientificamerican10221887-
267a.pdf.

Anonymous:1887:IIbk

[9673] Anonymous. Index of inventions. Scientific American, 57(17): 267–268, October 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n17/pdf/scientificamerican10221887-
267c.pdf.

Anonymous:1887:PNPbg

[9674] Anonymous. Patent notes: Patents. Scientific American, 57 (17):270, October 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n17/pdf/scientificamerican10221887-
270.pdf.

Anonymous:1887:TBF

[9675] Anonymous. A torpedo boat flotilla attack the Atlanta. Scientific American, 57(17):255, October 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n17/pdf/scientificamerican10221887-
255.pdf.

Anonymous:1887:NKA

[9676] Anonymous. New knowledge about car couplers. Scientific American, 57(17):256, October 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n17/pdf/scientificamerican10221887-
256.pdf.

Anonymous:1887:AIEa

[9677] Anonymous. The American Institute Exhibition. Scientific American, 57(17):256, October 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n17/pdf/scientificamerican10221887-
256a.pdf.

Anonymous:1887:PNk

[9678] Anonymous. Photographic notes. Scientific American, 57(17):257, October 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n17/pdf/scientificamerican10221887-257.pdf.

Anonymous:1887:HMT

[9679] Anonymous. H.M.S. Trafalgar. Scientific American, 57(17):259– 260, October 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n17/pdf/scientificamerican10221887-
259.pdf.

Anonymous:1887:WC

[9680] Anonymous. A watch camera. Scientific American, 57(17):260, October 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n17/pdf/scientificamerican10221887-260.pdf.

Anonymous:1887:ED

[9681] Anonymous. Engines of the Dogali. Scientific American, 57(17): 262–265, October 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n17/pdf/scientificamerican10221887-
262.pdf.

Anonymous:1887:CYa

[9682] Anonymous. Chevreul at 101 years. Scientific American, 57 (17):265, October 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n17/pdf/scientificamerican10221887-
265.pdf.

Anonymous:1887:EIbi

[9683] Anonymous. Engineering inventions. Scientific American, 57(17): 266, October 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n17/pdf/scientificamerican10221887-
266.pdf.

Anonymous:1887:AIbe

[9684] Anonymous. Agricultural inventions. Scientific American, 57 (17):266, October 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n17/pdf/scientificamerican10221887-
266a.pdf.

Anonymous:1887:MIbq

[9685] Anonymous. Miscellaneous inventions. Scientific American, 57 (17):266, October 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n17/pdf/scientificamerican10221887-
266b.pdf.

Anonymous:1887:Sj

[9686] Anonymous. Special. Scientific American, 57(17):266, October 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n17/pdf/scientificamerican10221887-266c.pdf.

Anonymous:1887:Ibo

[9687] Anonymous. To inventors. Scientific American, 57(17):267, October 22, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n17/pdf/scientificamerican10221887-267b.pdf.

Anonymous:1887:Cbk

[9688] Anonymous. Correspondence. Scientific American, 57(18):277, October 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n18/pdf/scientificamerican10291887-277.pdf.

Anonymous:1887:BPbk

[9689] Anonymous. Business and personal. Scientific American, 57 (18):283, October 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n18/pdf/scientificamerican10291887-
283c.pdf.

Anonymous:1887:NQbd

[9690] Anonymous. Notes and queries. Scientific American, 57(18):283, October 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n18/pdf/scientificamerican10291887-283d.pdf.

Anonymous:1887:IIbl

[9691] Anonymous. Index of inventions. Scientific American, 57(18): 283–284, October 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n18/pdf/scientificamerican10291887-
283f.pdf.

Anonymous:1887:PNPbh

[9692] Anonymous. Patent notes: Patents. Scientific American, 57 (18):286, October 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n18/pdf/scientificamerican10291887-
286.pdf.

Anonymous:1887:HMP

[9693] Anonymous. Hastening the military pioneer. Scientific American, 57(18):272, October 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n18/pdf/scientificamerican10291887-
272.pdf.

Anonymous:1887:CCf

[9694] Anonymous. The crowned crane. Scientific American, 57(18):272, October 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n18/pdf/scientificamerican10291887-272a.pdf.

Anonymous:1887:PPN

[9695] Anonymous. Positions of the planets in November. Scientific American, 57(18):272, October 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n18/pdf/scientificamerican10291887-
272b.pdf.

Anonymous:1887:PNl

[9696] Anonymous. Photographic notes. Scientific American, 57(18):273, October 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n18/pdf/scientificamerican10291887-273.pdf.

Anonymous:1887:SHA

[9697] Anonymous. A spoon holder attachment for bottles, etc. Scientific American, 57(18):274, October 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v57/n18/pdf/scientificamerican10291887-
274.pdf.

Anonymous:1887:IMCb

[9698] Anonymous. Improved manhole cover for sewers, street vaults, etc. Scientific American, 57(18):274, October 29, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v57/
n18/pdf/scientificamerican10291887-274a.pdf.

Anonymous:1887:MRS

[9699] Anonymous. A method of running street cars. Scientific American, 57(18):274, October 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n18/pdf/scientificamerican10291887-
274b.pdf.

Anonymous:1887:ICBc

[9700] Anonymous. An improved car brake and starter. Scientific American, 57(18):274, October 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n18/pdf/scientificamerican10291887-
274c.pdf.

Anonymous:1887:DCW

[9701] Anonymous. A display case for window shades, etc. Scientific American, 57(18):274–275, October 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n18/pdf/scientificamerican10291887-
274d.pdf.

Anonymous:1887:IATa

[9702] Anonymous. Improved adjustable tongs. Scientific American, 57 (18):275, October 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n18/pdf/scientificamerican10291887-
275.pdf.

Anonymous:1887:SOL

[9703] Anonymous. A switch operated by the locomotive. Scientific American, 57(18):275, October 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n18/pdf/scientificamerican10291887-
275a.pdf.

Anonymous:1887:ICPd

[9704] Anonymous. An improved corn planter. Scientific American, 57 (18):275, October 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n18/pdf/scientificamerican10291887-
275b.pdf.

Anonymous:1887:NAA

[9705] Anonymous. New all-around crane. Scientific American, 57(18): 275–276, October 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n18/pdf/scientificamerican10291887-
275c.pdf.

Anonymous:1887:HOHb

[9706] Anonymous. Harvard Observatory and the Henry Draper Memorial. Scientific American, 57(18):278–280, October 29, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v57/
n18/pdf/scientificamerican10291887-278.pdf.

Anonymous:1887:NCN

[9707] Anonymous. The new Chinese Naval squadron. Scientific American, 57(18):281–282, October 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n18/pdf/scientificamerican10291887-
281.pdf.

Anonymous:1887:EIbj

[9708] Anonymous. Engineering inventions. Scientific American, 57(18): 283, October 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n18/pdf/scientificamerican10291887-
283.pdf.

Anonymous:1887:AIbf

[9709] Anonymous. Agricultural inventions. Scientific American, 57 (18):283, October 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n18/pdf/scientificamerican10291887-
283a.pdf.

Anonymous:1887:MIbr

[9710] Anonymous. Miscellaneous inventions. Scientific American, 57 (18):283, October 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n18/pdf/scientificamerican10291887-
283b.pdf.

Anonymous:1887:Ibp

[9711] Anonymous. To inventors. Scientific American, 57(18):283, October 29, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n18/pdf/scientificamerican10291887-283e.pdf.

Anonymous:1887:BPbl

[9712] Anonymous. Business and personal. Scientific American, 57 (19):298, November 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n19/pdf/scientificamerican11051887-
298d.pdf.

Anonymous:1887:BNBw

[9713] Anonymous. Books: New books and publications. Scientific American, 57(19):299, November 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n19/pdf/scientificamerican11051887-
299.pdf.

Anonymous:1887:NQbe

[9714] Anonymous. Notes and queries. Scientific American, 57(19): 299, November 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n19/pdf/scientificamerican11051887-
299a.pdf.

Anonymous:1887:IIbm

[9715] Anonymous. Index of inventions. Scientific American, 57(19): 299–300, November 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n19/pdf/scientificamerican11051887-
299c.pdf.

Anonymous:1887:PNPbi

[9716] Anonymous. Patent notes: Patents. Scientific American, 57 (19):302, November 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n19/pdf/scientificamerican11051887-
302.pdf.

Anonymous:1887:RLM

[9717] Anonymous. The research laboratory of Mr. Edward Weston. Scientific American, 57(19):287, November 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v57/n19/pdf/scientificamerican11051887-
287.pdf.

Anonymous:1887:DW

[9718] Anonymous. Distillation of wood. Scientific American, 57(19): 288, November 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n19/pdf/scientificamerican11051887-
288.pdf.

Anonymous:1887:CSa

[9719] Anonymous. Coaling at sea. Scientific American, 57(19):288– 289, November 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n19/pdf/scientificamerican11051887-
288a.pdf.

Anonymous:1887:ASR

[9720] Anonymous. An adjustable support for a receiving telephone. Scientific American, 57(19):291, November 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v57/n19/pdf/scientificamerican11051887-
291.pdf.

Anonymous:1887:ILPc

[9721] Anonymous. An improved life preserver. Scientific American, 57 (19):291, November 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n19/pdf/scientificamerican11051887-
291a.pdf.

Anonymous:1887:FEC

[9722] Anonymous. A fire extinguisher for car stoves. Scientific American, 57(19):291, November 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n19/pdf/scientificamerican11051887-
291b.pdf.

Anonymous:1887:IRP

[9723] Anonymous. An improved railway plow and excavator. Scientific American, 57(19):291, November 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n19/pdf/scientificamerican11051887-
291c.pdf.

Anonymous:1887:IPR

[9724] Anonymous. An improved portable railway switch. Scientific American, 57(19):291, November 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n19/pdf/scientificamerican11051887-
291d.pdf.

Anonymous:1887:ICPe

[9725] Anonymous. An improved cotton planter. Scientific American, 57(19):292, November 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n19/pdf/scientificamerican11051887-
292.pdf.

Anonymous:1887:IACa

[9726] Anonymous. An improved automatic car coupler. Scientific American, 57(19):292–294, November 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n19/pdf/scientificamerican11051887-
292a.pdf.

Anonymous:1887:RLC

[9727] Anonymous. Raising liquids by compressed air. Scientific American, 57(19):294, November 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n19/pdf/scientificamerican11051887-
294.pdf.

Holder:1887:AME

[9728] J. B. Holder. The American mammoth (Elephas Primigenius). Scientific American, 57(19):295, November 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v57/n19/pdf/scientificamerican11051887-
295.pdf.

Sloane:1887:EII

[9729] T. O’conor Sloane. Experimental illustration of the injector. Scientific American, 57(19):295–296, November 5, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v57/
n19/pdf/scientificamerican11051887-295a.pdf.

Anonymous:1887:TCB

[9730] Anonymous. Talcott’s combination belt hooks. Scientific American, 57(19):296, November 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n19/pdf/scientificamerican11051887-
296.pdf.

Anonymous:1887:SCC

[9731] Anonymous. The statue of Christopher Columbus at Barcelona, Spain. Scientific American, 57(19):297, November 5, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v57/
n19/pdf/scientificamerican11051887-297.pdf.

Anonymous:1887:EIbk

[9732] Anonymous. Engineering inventions. Scientific American, 57(19): 298, November 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n19/pdf/scientificamerican11051887-
298.pdf.

Anonymous:1887:AIbg

[9733] Anonymous. Agricultural inventions. Scientific American, 57 (19):298, November 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n19/pdf/scientificamerican11051887-
298a.pdf.

Anonymous:1887:MIbs

[9734] Anonymous. Miscellaneous inventions. Scientific American, 57 (19):298, November 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n19/pdf/scientificamerican11051887-
298b.pdf.

Anonymous:1887:Sk

[9735] Anonymous. Special. Scientific American, 57(19):298, November 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n19/pdf/scientificamerican11051887-298c.pdf.

Anonymous:1887:Ibq

[9736] Anonymous. To inventors. Scientific American, 57(19):299, November 5, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n19/pdf/scientificamerican11051887-299b.pdf.

Anonymous:1887:BPbm

[9737] Anonymous. Business and personal. Scientific American, 57(2):26, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n2/pdf/scientificamerican07091887-26b.pdf.

Anonymous:1887:BNBx

[9738] Anonymous. Books: New books and publications. Scientific American, 57(2):26–27, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n2/pdf/scientificamerican07091887-
26c.pdf.

Anonymous:1887:NQbf

[9739] Anonymous. Notes and queries. Scientific American, 57(2):27, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n2/pdf/scientificamerican07091887-27.pdf.

Anonymous:1887:IIbn

[9740] Anonymous. Index of inventions. Scientific American, 57(2):27–28, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n2/pdf/scientificamerican07091887-27b.pdf.

Anonymous:1887:PNPbj

[9741] Anonymous. Patent notes: Patents. Scientific American, 57(2):30, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n2/pdf/scientificamerican07091887-30.pdf.

Anonymous:1887:RSS

[9742] Anonymous. Raising a sunken steamship. Scientific American, 57(2):15, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n2/pdf/scientificamerican07091887-15.pdf.

Anonymous:1887:AUS

[9743] Anonymous. Accession of the United States to the industrial union. Scientific American, 57(2):16, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v57/n2/pdf/scientificamerican07091887-
16.pdf.

Anonymous:1887:USM

[9744] Anonymous. Use of the siren in marine signaling. Scientific American, 57(2):16, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n2/pdf/scientificamerican07091887-
16a.pdf.

Anonymous:1887:EDP

[9745] Anonymous. The economical disposition of primary batteries. Scientific American, 57(2):16–17, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n2/pdf/scientificamerican07091887-
16b.pdf.

Anonymous:1887:MFR

[9746] Anonymous. A mill to free rice from its hulls. Scientific American, 57(2): 18, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n2/pdf/scientificamerican07091887-18.pdf.

Anonymous:1887:SHI

[9747] Anonymous. A steam heated ironing machine. Scientific American, 57(2):18, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n2/pdf/scientificamerican07091887-
18a.pdf.

Anonymous:1887:AWT

[9748] Anonymous. An automatic wood turning lathe. Scientific American, 57(2):18, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n2/pdf/scientificamerican07091887-
18b.pdf.

Anonymous:1887:MTC

[9749] Anonymous. A miner’s tool to cut fuses, set caps, etc. Scientific American, 57(2):18, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n2/pdf/scientificamerican07091887-
18c.pdf.

Anonymous:1887:DFS

[9750] Anonymous. A device to feed the seed in cotton planting. Scientific American, 57(2):18, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n2/pdf/scientificamerican07091887-
18d.pdf.

Anonymous:1887:PFS

[9751] Anonymous. The premium folding stool and chair. Scientific American, 57(2):19, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n2/pdf/scientificamerican07091887-
19.pdf.

Anonymous:1887:PDS

[9752] Anonymous. A pen for double and single line ruling. Scientific American, 57(2):19, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n2/pdf/scientificamerican07091887-
19a.pdf.

Anonymous:1887:DCS

[9753] Anonymous. A dustless carpet sweeper. Scientific American, 57(2):19, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n2/pdf/scientificamerican07091887-19b.pdf.

Anonymous:1887:WPA

[9754] Anonymous. A wall protecting attachment for furniture. Scientific American, 57(2):19, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n2/pdf/scientificamerican07091887-
19c.pdf.

Anonymous:1887:IKH

[9755] Anonymous. An improved key hole guide and escutcheon. Scientific American, 57(2):19, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n2/pdf/scientificamerican07091887-
19d.pdf.

Anonymous:1887:IWG

[9756] Anonymous. An improved water gauge float. Scientific American, 57(2):19–20, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n2/pdf/scientificamerican07091887-
19e.pdf.

Anonymous:1887:NST

[9757] Anonymous. A new style of thermometer. Scientific American, 57(2): 20, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n2/pdf/scientificamerican07091887-20.pdf.

Mdougall:1887:LBF

[9758] James F. M’dougall. Larvæ of bot fly in horse’s stomach. Scientific American, 57(2):20–21, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n2/pdf/scientificamerican07091887-
20a.pdf.

Anonymous:1887:NFAb

[9759] Anonymous. New French armorclad cruiser Le Terrible. Scientific American, 57(2):22, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n2/pdf/scientificamerican07091887-
22.pdf.

Anonymous:1887:IACb

[9760] Anonymous. Improved automatic cut off engine. Scientific American, 57(2):24, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n2/pdf/scientificamerican07091887-
24.pdf.

Anonymous:1887:SPP

[9761] Anonymous. Simple photographic and photo-micrographic apparatus. Scientific American, 57(2):25, July 9, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v57/
n2/pdf/scientificamerican07091887-25.pdf.

Anonymous:1887:EIbl

[9762] Anonymous. Engineering invention. Scientific American, 57(2):26, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n2/pdf/scientificamerican07091887-26.pdf.

Anonymous:1887:MIbt

[9763] Anonymous. Miscellaneous inventions. Scientific American, 57(2):26, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n2/pdf/scientificamerican07091887-26a.pdf.

Anonymous:1887:Ibr

[9764] Anonymous. To inventors. Scientific American, 57(2):27, July 9, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n2/pdf/scientificamerican07091887-27a.pdf.

Anonymous:1887:BPbn

[9765] Anonymous. Business and personal. Scientific American, 57 (20):314, November 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n20/pdf/scientificamerican11121887-
314c.pdf.

Anonymous:1887:BNBy

[9766] Anonymous. Books: New books and publications. Scientific American, 57(20):314, November 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n20/pdf/scientificamerican11121887-
314d.pdf.

Anonymous:1887:NQbg (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n20/pdf/scientificamerican11121887-
315.pdf.

Anonymous:1887:IIbo

[9768] Anonymous. Index of inventions. Scientific American, 57(20): 315–316, November 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n20/pdf/scientificamerican11121887-
315b.pdf.

Anonymous:1887:PNPbk

[9769] Anonymous. Patent notes: Patents. Scientific American, 57 (20):318, November 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n20/pdf/scientificamerican11121887-
318.pdf.

Anonymous:1887:WSS

[9770] Anonymous. The Weiss system of street car motors. Scientific American, 57(20):303, November 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n20/pdf/scientificamerican11121887-
303.pdf.

Anonymous:1887:RCF

[9771] Anonymous. The rich coal fields of Colorado. Scientific American, 57(20):304, November 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n20/pdf/scientificamerican11121887-
304.pdf.

Anonymous:1887:BIA

[9772] Anonymous. Books for the insane in asylums. Scientific American, 57(20):304, November 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n20/pdf/scientificamerican11121887-
304a.pdf.

Anonymous:1887:AIEb

[9773] Anonymous. The American Institute Exhibition. Scientific American, 57(20):304–305, November 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n20/pdf/scientificamerican11121887-
304b.pdf.

Holden:1887:IES

[9774] H. J. Holden. Interesting experiments with simple apparatus. Scientific American, 57(20):306, November 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v57/n20/pdf/scientificamerican11121887-
306.pdf.

Anonymous:1887:IDLa

[9775] Anonymous. An improved door latch and lock. Scientific American, 57(20):307, November 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n20/pdf/scientificamerican11121887-
307.pdf.

Anonymous:1887:ICMc

[9776] Anonymous. An improved cloth measuring device. Scientific American, 57(20):307, November 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n20/pdf/scientificamerican11121887-
307a.pdf.

Anonymous:1887:IDRb

[9777] Anonymous. An improved drill for renewing worn out valves and valve seats. Scientific American, 57(20):307, November 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n20/pdf/scientificamerican11121887-307b.pdf.

Anonymous:1887:ITS

[9778] Anonymous. An improved telescoping stovepipe joint. Scientific American, 57(20):307, November 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n20/pdf/scientificamerican11121887-
307c.pdf.

Anonymous:1887:IRT

[9779] Anonymous. An improved railway tie. Scientific American, 57 (20):307, November 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n20/pdf/scientificamerican11121887-
307d.pdf.

Anonymous:1887:ICCj

[9780] Anonymous. An improved car coupling. Scientific American, 57 (20):308–309, November 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n20/pdf/scientificamerican11121887-
308.pdf.

Anonymous:1887:HMC

[9781] Anonymous. How to make a camera bellows. Scientific American, 57 (20):310–311, November 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n20/pdf/scientificamerican11121887-
310.pdf.

Anonymous:1887:SPCb

[9782] Anonymous. Sargent’s palm. (ChamæPhenix Sargentii). Scientific American, 57(20):311–312, November 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v57/n20/pdf/scientificamerican11121887-
311.pdf.

Anonymous:1887:IHM

[9783] Anonymous. Improved Hotchkiss and mountain guns. Scientific American, 57(20):313, November 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n20/pdf/scientificamerican11121887-
313.pdf.

Anonymous:1887:EIbm

[9784] Anonymous. Engineering inventions. Scientific American, 57(20): 314, November 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n20/pdf/scientificamerican11121887-
314.pdf.

Anonymous:1887:AIbh

[9785] Anonymous. Agricultural inventions. Scientific American, 57 (20):314, November 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n20/pdf/scientificamerican11121887-
314a.pdf.

Anonymous:1887:MIbu

[9786] Anonymous. Miscellaneous inventions. Scientific American, 57 (20):314, November 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n20/pdf/scientificamerican11121887-
314b.pdf.

Anonymous:1887:Ibt

[9787] Anonymous. To inventors. Scientific American, 57(20):315, November 12, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n20/pdf/scientificamerican11121887-315a.pdf.

Anonymous:1887:Cbl

[9788] Anonymous. Correspondence. Scientific American, 57(21):325, November 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n21/pdf/scientificamerican11191887-325.pdf.

Anonymous:1887:BPbo

[9789] Anonymous. Business and personal. Scientific American, 57 (21):331, November 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n21/pdf/scientificamerican11191887-
331d.pdf.

Anonymous:1887:IIbp

[9790] Anonymous. Index of inventions. Scientific American, 57(21): 331–332, November 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n21/pdf/scientificamerican11191887-
331f.pdf.

Anonymous:1887:PNPbl

[9791] Anonymous. Patent notes: Patents. Scientific American, 57 (21):334, November 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n21/pdf/scientificamerican11191887-
334.pdf.

Anonymous:1887:IUPb

[9792] Anonymous. International union for the protection of industrial property. Scientific American, 57(21):320, November 19, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v57/
n21/pdf/scientificamerican11191887-320.pdf.

Anonymous:1887:GNM

[9793] Anonymous. The Geneva non-magnetic watch. Scientific American, 57(21):320, November 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n21/pdf/scientificamerican11191887-
320a.pdf.

Anonymous:1887:PNm

[9794] Anonymous. Photographic notes. Scientific American, 57(21): 321, November 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n21/pdf/scientificamerican11191887-
321.pdf.

Anonymous:1887:ICGa

[9795] Anonymous. Improved construction of grain elevators. Scientific American, 57(21):322, November 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n21/pdf/scientificamerican11191887-
322.pdf.

Anonymous:1887:IOJ

[9796] Anonymous. An improved ore jigger. Scientific American, 57 (21):322, November 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n21/pdf/scientificamerican11191887-
322a.pdf.

Anonymous:1887:IVT

[9797] Anonymous. An improved violin tuning peg. Scientific American, 57(21):322, November 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n21/pdf/scientificamerican11191887-
322b.pdf.

Anonymous:1887:IVE

[9798] Anonymous. An improved vermin exterminator. Scientific American, 57(21):322, November 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n21/pdf/scientificamerican11191887-
322c.pdf.

Anonymous:1887:ISPc

[9799] Anonymous. An improved shoemaker’s pegging jack. Scientific American, 57(21):323, November 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n21/pdf/scientificamerican11191887-
323.pdf.

Anonymous:1887:CAL

[9800] Anonymous. Compressed air locomotive for underground haulage. Scientific American, 57(21):323, November 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v57/n21/pdf/scientificamerican11191887-
323a.pdf.

Anonymous:1887:IRSb

[9801] Anonymous. An improved railroad signal. Scientific American, 57 (21):323, November 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n21/pdf/scientificamerican11191887-
323b.pdf.

Anonymous:1887:ONNb

[9802] Anonymous. Our new Navy. Scientific American, 57(21):324, November 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n21/pdf/scientificamerican11191887-324.pdf.

Anonymous:1887:NJB

[9803] Anonymous. The New Jersey ballot box. Scientific American, 57 (21):326, November 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n21/pdf/scientificamerican11191887-
326.pdf.

Sloane:1887:OGV

[9804] T. O’conor Sloane. Otto Von Guericke’s vacuum gun. Scientific American, 57(21):326–327, November 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v57/n21/pdf/scientificamerican11191887-
326a.pdf.

Anonymous:1887:NIC

[9805] Anonymous. New isochronal clock. Scientific American, 57(21): 327–328, November 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n21/pdf/scientificamerican11191887-
327.pdf.

Anonymous:1887:VKK

[9806] Anonymous. The volcano of Kluchi or Kluchefskaya, Kamschatka. Scientific American, 57(21):329, November 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v57/n21/pdf/scientificamerican11191887-
329.pdf.

Anonymous:1887:PSI

[9807] Anonymous. Pneumatic speed indicator. Scientific American, 57 (21):329–330, November 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n21/pdf/scientificamerican11191887-
329a.pdf.

Anonymous:1887:EIbn

[9808] Anonymous. Engineering inventions. Scientific American, 57(21): 331, November 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n21/pdf/scientificamerican11191887-
331.pdf.

Anonymous:1887:AIbi

[9809] Anonymous. Agricultural inventions. Scientific American, 57 (21):331, November 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n21/pdf/scientificamerican11191887-
331a.pdf.

Anonymous:1887:MIbv

[9810] Anonymous. Miscellaneous inventions. Scientific American, 57 (21):331, November 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n21/pdf/scientificamerican11191887-
331b.pdf.

Anonymous:1887:Sl

[9811] Anonymous. Special. Scientific American, 57(21):331, November 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n21/pdf/scientificamerican11191887-331c.pdf.

Anonymous:1887:Ibu

[9812] Anonymous. To inventors. Scientific American, 57(21):331, November 19, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n21/pdf/scientificamerican11191887-331e.pdf.

Anonymous:1887:Cbm

[9813] Anonymous. Correspondence. Scientific American, 57(22):341– 342, November 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n22/pdf/scientificamerican11261887-
341.pdf.

Anonymous:1887:BPbp

[9814] Anonymous. Business and personal. Scientific American, 57 (22):347, November 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n22/pdf/scientificamerican11261887-
347c.pdf.

Anonymous:1887:IIbq

[9815] Anonymous. Index of inventions. Scientific American, 57(22): 347–348, November 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n22/pdf/scientificamerican11261887-
347e.pdf.

Anonymous:1887:PNPbm

[9816] Anonymous. Patent notes: Patents. Scientific American, 57 (22):350, November 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n22/pdf/scientificamerican11261887-
350.pdf.

Anonymous:1887:EW

[9817] Anonymous. Electric welding. Scientific American, 57(22):335, November 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n22/pdf/scientificamerican11261887-335.pdf.

Anonymous:1887:SRF

[9818] Anonymous. The Siemens Regenerative Furnace Patent in the Supreme Court. Scientific American, 57(22):336, November 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n22/pdf/scientificamerican11261887-336.pdf.

Anonymous:1887:DWPb

[9819] Anonymous. The driven well patent invalid. Scientific American, 57(22):336, November 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n22/pdf/scientificamerican11261887-
336a.pdf.

Anonymous:1887:RPe

[9820] Anonymous. Roadside planting. Scientific American, 57(22): 336, November 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n22/pdf/scientificamerican11261887-
336b.pdf.

Asher:1887:NTS

[9821] J. Asher. New thermometer scale. Scientific American, 57(22): 337–338, November 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n22/pdf/scientificamerican11261887-
337.pdf.

Anonymous:1887:RVH

[9822] Anonymous. The “Robinson” Victoria hansom. Scientific American, 57(22):338, November 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n22/pdf/scientificamerican11261887-
338.pdf.

Anonymous:1887:ICBd

[9823] Anonymous. An improved clothes back. Scientific American, 57 (22):339, November 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n22/pdf/scientificamerican11261887-
339.pdf.

Anonymous:1887:IHCa

[9824] Anonymous. An improved hay cocking machine. Scientific American, 57(22):339, November 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n22/pdf/scientificamerican11261887-
339a.pdf.

Anonymous:1887:ITH

[9825] Anonymous. An improved trunk handle. Scientific American, 57 (22):339, November 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n22/pdf/scientificamerican11261887-
339b.pdf.

Anonymous:1887:ICi

[9826] Anonymous. An improved cultivator. Scientific American, 57 (22):339, November 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n22/pdf/scientificamerican11261887-
339c.pdf.

Anonymous:1887:SAS

[9827] Anonymous. A safety attachment for stem winding watches. Scientific American, 57(22):339–340, November 26, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v57/
n22/pdf/scientificamerican11261887-339d.pdf.

Anonymous:1887:ICGb

[9828] Anonymous. Improved cutter grinder. Scientific American, 57 (22):342, November 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n22/pdf/scientificamerican11261887-
342.pdf.

Hopkins:1887:I

[9829] Geo. M. Hopkins. Irradiation. Scientific American, 57(22):342, November 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n22/pdf/scientificamerican11261887-342a.pdf.

Anonymous:1887:AAa

[9830] Anonymous. Aquariums for amateurs. Scientific American, 57(22): 343–344, November 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n22/pdf/scientificamerican11261887-
343.pdf.

Anonymous:1887:WPI

[9831] Anonymous. The L. W. Pond iron planing machine. Scientific American, 57(22):345–346, November 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v57/n22/pdf/scientificamerican11261887-
345.pdf.

Anonymous:1887:EIbo

[9832] Anonymous. Engineering inventions. Scientific American, 57(22): 347, November 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n22/pdf/scientificamerican11261887-
347.pdf.

Anonymous:1887:AIbj

[9833] Anonymous. Agricultural inventions. Scientific American, 57 (22):347, November 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n22/pdf/scientificamerican11261887-
347a.pdf.

Anonymous:1887:MIbw

[9834] Anonymous. Miscellaneous inventions. Scientific American, 57 (22):347, November 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n22/pdf/scientificamerican11261887-
347b.pdf.

Anonymous:1887:Ibv

[9835] Anonymous. To inventors. Scientific American, 57(22):347, November 26, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n22/pdf/scientificamerican11261887-347d.pdf.

Anonymous:1887:Cbo

[9836] Anonymous. Correspondence. Scientific American, 57(23):357, December 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n23/pdf/scientificamerican12031887-357.pdf.

Anonymous:1887:BPbq

[9837] Anonymous. Business and personal. Scientific American, 57(23): 362–363, December 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n23/pdf/scientificamerican12031887-
362c.pdf.

Anonymous:1887:BNBz

[9838] Anonymous. Books: New books and publications. Scientific American, 57(23):363, December 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n23/pdf/scientificamerican12031887-
363.pdf.

Anonymous:1887:NQbh

[9839] Anonymous. Notes and queries. Scientific American, 57(23):363, December 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n23/pdf/scientificamerican12031887-363a.pdf.

Anonymous:1887:IIbr

[9840] Anonymous. Index of inventions. Scientific American, 57(23): 363–364, December 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n23/pdf/scientificamerican12031887-
363c.pdf.

Anonymous:1887:PNPbn 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n23/pdf/scientificamerican12031887-
366.pdf.

Anonymous:1887:BSS

[9842] Anonymous. The Boston sewer system and main drainage works. Scientific American, 57(23):351, December 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v57/n23/pdf/scientificamerican12031887-
351.pdf.

Anonymous:1887:DDW

[9843] Anonymous. The decision in the driven well suit. Scientific American, 57(23):352, December 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n23/pdf/scientificamerican12031887-
352.pdf.

Anonymous:1887:WSW

[9844] Anonymous. War ships that are weak and slow. Scientific American, 57(23):352, December 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n23/pdf/scientificamerican12031887-
352a.pdf.

Anonymous:1887:PPD

[9845] Anonymous. Position of the planets in December. Scientific American, 57(23):352–354, December 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n23/pdf/scientificamerican12031887-
352b.pdf.

Anonymous:1887:APF

[9846] Anonymous. Apparatus of the Paris Fire Department. Scientific American, 57(23):354, December 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n23/pdf/scientificamerican12031887-
354.pdf.

Anonymous:1887:CSBb

[9847] Anonymous. A convenient sketching board or book rest. Scientific American, 57(23):355, December 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v57/n23/pdf/scientificamerican12031887-
355.pdf.

Anonymous:1887:ICCk

[9848] Anonymous. An improved car coupling. Scientific American, 57 (23):355, December 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n23/pdf/scientificamerican12031887-
355a.pdf.

Anonymous:1887:ICSf

[9849] Anonymous. An improved cotton stalk cutter. Scientific American, 57(23):355, December 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n23/pdf/scientificamerican12031887-
355b.pdf.

Anonymous:1887:ICTb

[9850] Anonymous. An improved combination tool. Scientific American, 57(23):355, December 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n23/pdf/scientificamerican12031887-
355c.pdf.

Anonymous:1887:ACB

[9851] Anonymous. An anti-clogging boiler flue plug. Scientific American, 57(23):355–356, December 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n23/pdf/scientificamerican12031887-
355d.pdf.

Anonymous:1887:ICBe

[9852] Anonymous. Improved construction of buildings. Scientific American, 57(23):356, December 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n23/pdf/scientificamerican12031887-
356.pdf.

Anonymous:1887:MG

[9853] Anonymous. Madagascar geese. Scientific American, 57(23):361, December 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n23/pdf/scientificamerican12031887-361.pdf.

Anonymous:1887:JGT

[9854] Anonymous. Jay Gould’s tomb. Scientific American, 57(23):361, December 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n23/pdf/scientificamerican12031887-361a.pdf.

Anonymous:1887:EIbp

[9855] Anonymous. Engineering inventions. Scientific American, 57(23): 362, December 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n23/pdf/scientificamerican12031887-
362.pdf.

Anonymous:1887:MIbx

[9856] Anonymous. Miscellaneous inventions. Scientific American, 57 (23):362, December 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n23/pdf/scientificamerican12031887-
362a.pdf.

Anonymous:1887:Sm

[9857] Anonymous. Special. Scientific American, 57(23):362, December 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n23/pdf/scientificamerican12031887-362b.pdf.

Anonymous:1887:Ibw

[9858] Anonymous. To inventors. Scientific American, 57(23):363, December 3, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n23/pdf/scientificamerican12031887-363b.pdf.

Anonymous:1887:Cbq

[9859] Anonymous. Correspondence. Scientific American, 57(24):373, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n24/pdf/scientificamerican12101887-373.pdf.

Anonymous:1887:BPbr

[9860] Anonymous. Business and personal. Scientific American, 57 (24):378, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
378c.pdf.

Anonymous:1887:BNBba

[9861] Anonymous. Books: New books and publications. Scientific American, 57(24):378, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
378d.pdf.

Anonymous:1887:NQbi

[9862] Anonymous. Notes and queries. Scientific American, 57(24): 378–379, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
378e.pdf.

Anonymous:1887:IIbs

[9863] Anonymous. Index of inventions. Scientific American, 57(24): 379–380, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
379a.pdf.

Anonymous:1887:PNPbo

[9864] Anonymous. Patent notes: Patents. Scientific American, 57 (24):382, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
382.pdf.

Anonymous:1887:LBT

[9865] Anonymous. Lifting bridge at Tarante. Scientific American, 57 (24):367, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
367.pdf.

Anonymous:1887:IPAb

[9866] Anonymous. The illuminating power of arc lamps. Scientific American, 57(24):368, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
368.pdf.

Anonymous:1887:NPMb

[9867] Anonymous. A newly patented mode of preserving live fish. Scientific American, 57(24):368, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
368a.pdf.

Anonymous:1887:SRN

[9868] Anonymous. Survey of the route for the Nicaragua Canal. Scientific American, 57(24):368–369, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
368b.pdf.

Anonymous:1887:CWf

[9869] Anonymous. The celestial world. Scientific American, 57(24): 369, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
369.pdf.

Anonymous:1887:DRPf

[9870] Anonymous. Decisions relating to patents. Scientific American, 57(24):369, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
369a.pdf.

Anonymous:1887:IAP

[9871] Anonymous. An improved anchor for posts. Scientific American, 57(24):370, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
370.pdf.

Anonymous:1887:ISBd

[9872] Anonymous. An improved separable button or stud. Scientific American, 57(24):370, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
370a.pdf.

Anonymous:1887:SCF

[9873] Anonymous. A simple cover fastening for jars. Scientific American, 57(24):370, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
370b.pdf.

Anonymous:1887:ITd

[9874] Anonymous. An improved toaster. Scientific American, 57(24): 370, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
370c.pdf.

Anonymous:1887:PTCb

[9875] Anonymous. A protector for trunk corners. Scientific American, 57(24):370, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
370d.pdf.

Anonymous:1887:IVBb

[9876] Anonymous. An improved vehicle body support. Scientific American, 57(24):370, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
370e.pdf.

Anonymous:1887:IDMb

[9877] Anonymous. An improved discount measuring glass and bank note examiner. Scientific American, 57(24):371, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n24/pdf/scientificamerican12101887-371.pdf.

Anonymous:1887:ICSg

[9878] Anonymous. An improved carpet stretcher. Scientific American, 57(24):371, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
371a.pdf.

Anonymous:1887:SPT

[9879] Anonymous. Securing pictures to tombstones. Scientific American, 57(24):371, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
371b.pdf.

Anonymous:1887:IFPc

[9880] Anonymous. An improved fence post. Scientific American, 57 (24):371, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
371c.pdf.

Anonymous:1887:HGP

[9881] Anonymous. A heater for giant powder and other explosives. Scientific American, 57(24):371–372, December 10, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v57/
n24/pdf/scientificamerican12101887-371d.pdf.

Anonymous:1887:ICPf

[9882] Anonymous. An improved cotton planter. Scientific American, 57 (24):372, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
372.pdf.

Anonymous:1887:PNn

[9883] Anonymous. Photographic notes. Scientific American, 57(24): 373, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
373a.pdf.

Anonymous:1887:SMF

[9884] Anonymous. Separating the Manila fiber. Scientific American, 57 (24):374, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
374.pdf.

Anonymous:1887:FDA

[9885] Anonymous. French disinfecting apparatus. Scientific American, 57(24):374, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
374a.pdf.

Anonymous:1887:BZG

[9886] Anonymous. Birds in the Zoological Garden at Berlin. Scientific American, 57(24):375–376, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
375.pdf.

B:1887:JPC

[9887] M. B. Josiah Parsons Cooke. Scientific American, 57(24):377, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n24/pdf/scientificamerican12101887-377.pdf.

Anonymous:1887:EIbq

[9888] Anonymous. Engineering invention. Scientific American, 57(24): 378, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
378.pdf.

Anonymous:1887:AIbk

[9889] Anonymous. Agricultural invention. Scientific American, 57(24): 378, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
378a.pdf.

Anonymous:1887:MIby

[9890] Anonymous. Miscellaneous inventions. Scientific American, 57 (24):378, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n24/pdf/scientificamerican12101887-
378b.pdf.

Anonymous:1887:Ibx

[9891] Anonymous. To inventors. Scientific American, 57(24):379, December 10, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n24/pdf/scientificamerican12101887-379.pdf.

Anonymous:1887:Cbr

[9892] Anonymous. Correspondence. Scientific American, 57(25):389, December 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n25/pdf/scientificamerican12171887-389.pdf.

Anonymous:1887:BPbs

[9893] Anonymous. Business and personal. Scientific American, 57(25): 394–395, December 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n25/pdf/scientificamerican12171887-
394d.pdf.

Anonymous:1887:BNBbb

[9894] Anonymous. Books: New books and publications. Scientific American, 57(25):395, December 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n25/pdf/scientificamerican12171887-
395.pdf.

Anonymous:1887:NQbj

[9895] Anonymous. Notes and queries. Scientific American, 57(25): 395, December 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n25/pdf/scientificamerican12171887-
395a.pdf.

Anonymous:1887:IIbt 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n25/pdf/scientificamerican12171887-
395c.pdf.

Anonymous:1887:PNPbp

[9897] Anonymous. Patent notes: Patents. Scientific American, 57 (25):398, December 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n25/pdf/scientificamerican12171887-
398.pdf.

Anonymous:1887:AWE

[9898] Anonymous. Artesian well engineering. Scientific American, 57 (25):383, December 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n25/pdf/scientificamerican12171887-
383.pdf.

Anonymous:1887:OES

[9899] Anonymous. One effect of sham battles. Scientific American, 57 (25):384, December 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n25/pdf/scientificamerican12171887-
384.pdf.

Anonymous:1887:CPC

[9900] Anonymous. The change in the Panama Canal. Scientific American, 57(25):384, December 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n25/pdf/scientificamerican12171887-
384a.pdf.

Anonymous:1887:DRPg

[9901] Anonymous. Decisions relating to patents. Scientific American, 57(25):385, December 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n25/pdf/scientificamerican12171887-
385.pdf.

Anonymous:1887:PNo (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n25/pdf/scientificamerican12171887-
385a.pdf.

Anonymous:1887:MCb

[9903] Anonymous. Mechanical cooperage. Scientific American, 57(25): 386, December 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n25/pdf/scientificamerican12171887-
386.pdf.

Anonymous:1887:IMRa

[9904] Anonymous. An improved metallic railroad tie. Scientific American, 57(25):387, December 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n25/pdf/scientificamerican12171887-
387.pdf.

Anonymous:1887:CLR

[9905] Anonymous. A combined level and rule. Scientific American, 57 (25):387, December 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n25/pdf/scientificamerican12171887-
387a.pdf.

Anonymous:1887:IVBc

[9906] Anonymous. An improved vehicle brake. Scientific American, 57 (25):387, December 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n25/pdf/scientificamerican12171887-
387b.pdf.

Anonymous:1887:ISHb

[9907] Anonymous. An improved sash holder. Scientific American, 57 (25):387, December 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n25/pdf/scientificamerican12171887-
387c.pdf.

Anonymous:1887:IHb

[9908] Anonymous. An improved hinge. Scientific American, 57(25): 387–388, December 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n25/pdf/scientificamerican12171887-
387d.pdf.

Anonymous:1887:CDD

[9909] Anonymous. A cultivator, duster, and digger. Scientific American, 57(25):388, December 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n25/pdf/scientificamerican12171887-
388.pdf.

Sloane:1887:SFB

[9910] T. O’Conor Sloane. Single fluid battery. Scientific American, 57 (25):390, December 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n25/pdf/scientificamerican12171887-
390.pdf.

Anonymous:1887:BIE

[9911] Anonymous. The Brussels International Exhibition, 1888. Scientific American, 57(25):390–391, December 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v57/n25/pdf/scientificamerican12171887-
390a.pdf.

Anonymous:1887:AWR

[9912] Anonymous. An automatically working railroad gate. Scientific American, 57(25):391, December 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n25/pdf/scientificamerican12171887-
391.pdf.

Anonymous:1887:DFM

[9913] Anonymous. The distillation of fruits and manufacture of brandy. Scientific American, 57(25):391–393, December 17, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v57/
n25/pdf/scientificamerican12171887-391a.pdf.

Anonymous:1887:EIbr (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n25/pdf/scientificamerican12171887-
394.pdf.

Anonymous:1887:AIbl

[9915] Anonymous. Agricultural inventions. Scientific American, 57 (25):394, December 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n25/pdf/scientificamerican12171887-
394a.pdf.

Anonymous:1887:MIbz

[9916] Anonymous. Miscellaneous inventions. Scientific American, 57 (25):394, December 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n25/pdf/scientificamerican12171887-
394b.pdf.

Anonymous:1887:Sn

[9917] Anonymous. Special. Scientific American, 57(25):394, December 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n25/pdf/scientificamerican12171887-394c.pdf.

Anonymous:1887:Iby

[9918] Anonymous. To inventors. Scientific American, 57(25):395, December 17, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n25/pdf/scientificamerican12171887-395b.pdf.

Anonymous:1887:Cbs

[9919] Anonymous. Correspondence. Scientific American, 57(26):405, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n26/pdf/scientificamerican12241887-405.pdf.

Anonymous:1887:BPbt

[9920] Anonymous. Business and personal. Scientific American, 57 (26):410, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
410c.pdf.

Anonymous:1887:BNBbc

[9921] Anonymous. Books: New books and publications. Scientific American, 57(26):410–411, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
410d.pdf.

Anonymous:1887:NQbk

[9922] Anonymous. Notes and queries. Scientific American, 57(26): 411, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
411.pdf.

Anonymous:1887:IIbu

[9923] Anonymous. Index of inventions. Scientific American, 57(26): 411–412, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
411b.pdf.

Anonymous:1887:PNPbq

[9924] Anonymous. Patent notes: Patents. Scientific American, 57 (26):414, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
414.pdf.

Anonymous:1887:WAT

[9925] Anonymous. The Watertown Arsenal testing machine. Scientific American, 57(26):399–408, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
399.pdf.

Anonymous:1887:ESR

[9926] Anonymous. The Eads ship railroad. Scientific American, 57 (26):400, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
400.pdf.

Anonymous:1887:GRL

[9927] Anonymous. A great raft of lumber. Scientific American, 57 (26):400, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
400a.pdf.

Anonymous:1887:PIAa

[9928] Anonymous. Patent “innocents” again in Congress. Scientific American, 57(26):400–401, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
400b.pdf.

Anonymous:1887:IDCd

[9929] Anonymous. An improved dump cart. Scientific American, 57 (26):402, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
402.pdf.

Anonymous:1887:TWV

[9930] Anonymous. The transportation of war vessels by railway. Scientific American, 57(26):402, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
402a.pdf.

Anonymous:1887:IJH

[9931] Anonymous. An improved jib hank. Scientific American, 57(26): 402–403, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
402b.pdf.

Anonymous:1887:ISCd

[9932] Anonymous. An improved safety car heater. Scientific American, 57(26):403, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
403.pdf.

Anonymous:1887:IGb

[9933] Anonymous. An improved gate. Scientific American, 57(26): 403, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
403a.pdf.

Anonymous:1887:IAA

[9934] Anonymous. An improvement in attaching ax helves. Scientific American, 57(26):403, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
403b.pdf.

Anonymous:1887:IDB

[9935] Anonymous. An improved door bolt. Scientific American, 57 (26):403, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
403c.pdf.

Anonymous:1887:MFB

[9936] Anonymous. A machine for forming boot or shoe heels. Scientific American, 57(26):403, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
403d.pdf.

Anonymous:1887:IFC

[9937] Anonymous. An improved fireproof car heater. Scientific American, 57(26):403–404, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
403e.pdf.

Anonymous:1887:LCT

[9938] Anonymous. A lamp for Christmas trees. Scientific American, 57 (26):404, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
404.pdf.

Anonymous:1887:ASE

[9939] Anonymous. Automatic safety elevator. Scientific American, 57 (26):404, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
404a.pdf.

Anonymous:1887:IRO

[9940] Anonymous. Improved rocker for ordinary chairs. Scientific American, 57(26):404, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
404b.pdf.

Anonymous:1887:CDB

[9941] Anonymous. Centrifugal dyeing and bleaching apparatus. Scientific American, 57(26):406–407, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
406.pdf.

Hopkins:1887:SES

[9942] Geo. M. Hopkins. Scientific experiments with simple apparatus. Scientific American, 57(26):409, December 24, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v57/
n26/pdf/scientificamerican12241887-409.pdf.

Anonymous:1887:EIbs

[9943] Anonymous. Engineering inventions. Scientific American, 57(26): 410, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
410.pdf.

Anonymous:1887:AIbm

[9944] Anonymous. Agricultural invention. Scientific American, 57(26): 410, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
410a.pdf.

Anonymous:1887:MIca

[9945] Anonymous. Miscellaneous inventions. Scientific American, 57 (26):410, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n26/pdf/scientificamerican12241887-
410b.pdf.

Anonymous:1887:Ibz

[9946] Anonymous. To inventors. Scientific American, 57(26):411, December 24, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n26/pdf/scientificamerican12241887-411a.pdf.

Anonymous:1887:Cbu

[9947] Anonymous. Correspondence. Scientific American, 57(3):37, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n3/pdf/scientificamerican07161887-37.pdf.

Anonymous:1887:BPbu

[9948] Anonymous. Business and personal. Scientific American, 57(3):42–43, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n3/pdf/scientificamerican07161887-42e.pdf.

Anonymous:1887:BNBbd

[9949] Anonymous. Books: New books and publications. Scientific American, 57(3):43, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n3/pdf/scientificamerican07161887-
43.pdf.

Anonymous:1887:NQbl

[9950] Anonymous. Notes and queries. Scientific American, 57(3):43, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n3/pdf/scientificamerican07161887-43a.pdf.

Anonymous:1887:IIbv

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n3/pdf/scientificamerican07161887-43c.pdf.

Anonymous:1887:PNPbr

[9952] Anonymous. Patent notes: Patents. Scientific American, 57(3):46, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n3/pdf/scientificamerican07161887-46.pdf.

Anonymous:1887:WPC

[9953] Anonymous. Work on the Panama Canal likely to stop. Scientific American, 57(3):32, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n3/pdf/scientificamerican07161887-
32.pdf.

Anonymous:1887:HWU

[9954] Anonymous. How was the Unebi’kan lost? Scientific American, 57(3):32, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n3/pdf/scientificamerican07161887-
32a.pdf.

Anonymous:1887:IDPa

[9955] Anonymous. Infection from dairy products. Scientific American, 57(3):32–33, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n3/pdf/scientificamerican07161887-
32b.pdf.

Anonymous:1887:MMA

[9956] Anonymous. A mowing machine attachment. Scientific American, 57(3):34, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n3/pdf/scientificamerican07161887-
34.pdf.

Anonymous:1887:IBSc

[9957] Anonymous. An improved boiler for steam heaters. Scientific American, 57(3):34, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n3/pdf/scientificamerican07161887-
34a.pdf.

Anonymous:1887:IATb

[9958] Anonymous. An improved adjustable trammel. Scientific American, 57(3):34, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n3/pdf/scientificamerican07161887-
34b.pdf.

Anonymous:1887:DOS

[9959] Anonymous. Device for operating sewing machines by hand. Scientific American, 57(3):34, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n3/pdf/scientificamerican07161887-
34c.pdf.

Anonymous:1887:QSM

[9960] Anonymous. A quick steam making boiler for fire engines. Scientific American, 57(3):34, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n3/pdf/scientificamerican07161887-
34d.pdf.

Anonymous:1887:IFIa

[9961] Anonymous. An improved folding ironing table. Scientific American, 57(3):35, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n3/pdf/scientificamerican07161887-
35.pdf.

Anonymous:1887:IECa

[9962] Anonymous. An improved egg carrier. Scientific American, 57(3):35, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n3/pdf/scientificamerican07161887-35a.pdf.

Anonymous:1887:RTS

[9963] Anonymous. A roller toboggan slide. Scientific American, 57(3):35, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n3/pdf/scientificamerican07161887-35b.pdf.

Anonymous:1887:IFb

[9964] Anonymous. An improved filter. Scientific American, 57(3):35, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n3/pdf/scientificamerican07161887-35c.pdf.

Anonymous:1887:RCM

[9965] Anonymous. A recording compass and marine indicator. Scientific American, 57(3):35–36, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n3/pdf/scientificamerican07161887-
35d.pdf.

Anonymous:1887:ONNc

[9966] Anonymous. Our new Navy. Scientific American, 57(3):36, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n3/pdf/scientificamerican07161887-36.pdf.

Anonymous:1887:STo

[9967] Anonymous. Science in toys. Scientific American, 57(3):38, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n3/pdf/scientificamerican07161887-38.pdf.

Anonymous:1887:BM

[9968] Anonymous. Beetling machine. Scientific American, 57(3):38–39, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n3/pdf/scientificamerican07161887-38a.pdf.

Benjamin:1887:CFC

[9969] Marcus Benjamin. Charles Frederick Chandler. Scientific American, 57(3):39–41, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n3/pdf/scientificamerican07161887-
39.pdf.

Anonymous:1887:RTK

[9970] Anonymous. The racquet tailed kingfisher. Scientific American, 57(3):41, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n3/pdf/scientificamerican07161887-
41.pdf.

Anonymous:1887:EIbt

[9971] Anonymous. Engineering invention. Scientific American, 57(3):42, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n3/pdf/scientificamerican07161887-42.pdf.

Anonymous:1887:MIcb

[9972] Anonymous. Mechanical inventions. Scientific American, 57(3):42, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n3/pdf/scientificamerican07161887-42a.pdf.

Anonymous:1887:AIbn

[9973] Anonymous. Agricultural inventions. Scientific American, 57(3):42, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n3/pdf/scientificamerican07161887-42b.pdf.

Anonymous:1887:MIcc

[9974] Anonymous. Miscellaneous inventions. Scientific American, 57(3):42, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n3/pdf/scientificamerican07161887-42c.pdf.

Anonymous:1887:So

[9975] Anonymous. Special. Scientific American, 57(3):42, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n3/pdf/scientificamerican07161887-42d.pdf.

Anonymous:1887:Icj

[9976] Anonymous. To inventors. Scientific American, 57(3):43, July 16, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n3/pdf/scientificamerican07161887-43b.pdf.

Anonymous:1887:BPbv

[9977] Anonymous. Business and personal. Scientific American, 57(4):58, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n4/pdf/scientificamerican07231887-58d.pdf.

Anonymous:1887:BNBbe

[9978] Anonymous. Books: New books and publications. Scientific American, 57(4):58, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n4/pdf/scientificamerican07231887-
58e.pdf.

Anonymous:1887:NQbm

[9979] Anonymous. Notes and queries. Scientific American, 57(4):59, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n4/pdf/scientificamerican07231887-59.pdf.

Anonymous:1887:IIbw

[9980] Anonymous. Index of inventions. Scientific American, 57(4):59–60, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n4/pdf/scientificamerican07231887-59b.pdf.

Anonymous:1887:PNPbs

[9981] Anonymous. Patent notes: Patents. Scientific American, 57(4):62, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n4/pdf/scientificamerican07231887-62.pdf.

Anonymous:1887:ITCb

[9982] Anonymous. Improved type composing and distributing machine. Scientific American, 57(4):47, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v57/n4/pdf/scientificamerican07231887-
47.pdf.

Anonymous:1887:AK

[9983] Anonymous. Alfred krupp. Scientific American, 57(4):48, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n4/pdf/scientificamerican07231887-48.pdf.

Anonymous:1887:NFB

[9984] Anonymous. New fast boats. Scientific American, 57(4):48, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n4/pdf/scientificamerican07231887-48a.pdf.

Anonymous:1887:PNq

[9985] Anonymous. Photographic notes. Scientific American, 57(4):49, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n4/pdf/scientificamerican07231887-49.pdf.

Anonymous:1887:IFIb

[9986] Anonymous. An improved fastening for ice plow teeth. Scientific American, 57(4):50, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n4/pdf/scientificamerican07231887-
50.pdf.

Anonymous:1887:SAM

[9987] Anonymous. A stand adjustable for many uses. Scientific American, 57(4):50, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n4/pdf/scientificamerican07231887-
50a.pdf.

Anonymous:1887:SLD

[9988] Anonymous. A self-loading dirt cart. Scientific American, 57(4):50, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n4/pdf/scientificamerican07231887-50b.pdf.

Anonymous:1887:FBS

[9989] Anonymous. A folding back for solar printing frames. Scientific American, 57(4):50, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n4/pdf/scientificamerican07231887-
50c.pdf.

Anonymous:1887:IVL

[9990] Anonymous. An improved vehicle lubricator. Scientific American, 57(4):50, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n4/pdf/scientificamerican07231887-
50d.pdf.

Anonymous:1887:ILW

[9991] Anonymous. An improved lubricating wagon axle. Scientific American, 57(4):51, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n4/pdf/scientificamerican07231887-
51.pdf.

Anonymous:1887:IPHb

[9992] Anonymous. An improved pad for horse collars. Scientific American, 57(4):51, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n4/pdf/scientificamerican07231887-
51a.pdf.

Anonymous:1887:NCC

[9993] Anonymous. A nutshell cutter for confectioners, bakers, etc. Scientific American, 57(4):51, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n4/pdf/scientificamerican07231887-
51b.pdf.

Anonymous:1887:ICWc

[9994] Anonymous. An improved calf weaner. Scientific American, 57(4):51, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n4/pdf/scientificamerican07231887-51c.pdf.

Anonymous:1887:ICCl

[9995] Anonymous. An improved car coupling. Scientific American, 57(4):51, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n4/pdf/scientificamerican07231887-51d.pdf.

Anonymous:1887:RMP

[9996] Anonymous. The railroad on Mt. Pilatus. Scientific American, 57(4):55, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n4/pdf/scientificamerican07231887-55.pdf.

Anonymous:1887:NLS

[9997] Anonymous. Novel lantern slide. Scientific American, 57(4):55, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n4/pdf/scientificamerican07231887-55a.pdf.

Anonymous:1887:SFB

[9998] Anonymous. Suspension foot bridge. Scientific American, 57(4):55–56, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n4/pdf/scientificamerican07231887-55b.pdf.

Anonymous:1887:DEM

[9999] Anonymous. The Dinoceras, an extinct mammal. Scientific American, 57(4):57, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n4/pdf/scientificamerican07231887-
57.pdf.

Anonymous:1887:EIbu

[10000] Anonymous. Engineering inventions. Scientific American, 57(4):58, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n4/pdf/scientificamerican07231887-58.pdf.

Anonymous:1887:MIcd

[10001] Anonymous. Mechanical invention. Scientific American, 57(4):58, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n4/pdf/scientificamerican07231887-58a.pdf.

Anonymous:1887:AIbo

[10002] Anonymous. Agricultural inventions. Scientific American, 57(4):58, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n4/pdf/scientificamerican07231887-58b.pdf.

Anonymous:1887:MIce

[10003] Anonymous. Miscellaneous inventions. Scientific American, 57(4):58, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n4/pdf/scientificamerican07231887-58c.pdf.

Anonymous:1887:Ick

[10004] Anonymous. To inventors. Scientific American, 57(4):59, July 23, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n4/pdf/scientificamerican07231887-59a.pdf.

Anonymous:1887:Cbv

[10005] Anonymous. Correspondence. Scientific American, 57(5):69, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n5/pdf/scientificamerican07301887-69.pdf.

Anonymous:1887:BPbw

[10006] Anonymous. Business and personal. Scientific American, 57(5):75, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n5/pdf/scientificamerican07301887-75a.pdf.

Anonymous:1887:BNBbf

[10007] Anonymous. Books: New books and publications. Scientific American, 57(5):75, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n5/pdf/scientificamerican07301887-
75b.pdf.

Anonymous:1887:NQbn

[10008] Anonymous. Notes and queries. Scientific American, 57(5):75, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n5/pdf/scientificamerican07301887-75c.pdf.

Anonymous:1887:IIbx

[10009] Anonymous. Index of inventions. Scientific American, 57(5):75–76, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n5/pdf/scientificamerican07301887-75e.pdf.

Anonymous:1887:PNPbt

[10010] Anonymous. Patent notes: Patents. Scientific American, 57(5):78, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n5/pdf/scientificamerican07301887-78.pdf.

Anonymous:1887:IYR

[10011] Anonymous. The International Yacht Race. Scientific American, 57(5):64, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n5/pdf/scientificamerican07301887-
64.pdf.

Anonymous:1887:PPA

[10012] Anonymous. Positions of the planets in August. Scientific American, 57(5):64–65, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n5/pdf/scientificamerican07301887-
64a.pdf.

Anonymous:1887:IVF

[10013] Anonymous. An improved ventilating fan. Scientific American, 57(5):66, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n5/pdf/scientificamerican07301887-
66.pdf.

Anonymous:1887:IWP

[10014] Anonymous. An improved wear plate for vehicles. Scientific American, 57(5):66, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n5/pdf/scientificamerican07301887-
66a.pdf.

Anonymous:1887:IPF

[10015] Anonymous. Improved portable forges. Scientific American, 57(5):66, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n5/pdf/scientificamerican07301887-66b.pdf.

Anonymous:1887:IPTb

[10016] Anonymous. An improved pipe for tobacco smokers. Scientific American, 57(5):66, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n5/pdf/scientificamerican07301887-
66c.pdf.

Anonymous:1887:IDPb

[10017] Anonymous. An improved door plate and bell pull. Scientific American, 57(5):66, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n5/pdf/scientificamerican07301887-
66d.pdf.

Anonymous:1887:IACc

[10018] Anonymous. An improved automatic car brake. Scientific American, 57(5):67, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n5/pdf/scientificamerican07301887-
67.pdf.

Anonymous:1887:CFC

[10019] Anonymous. A combined feed cooker and scalder. Scientific American, 57(5):67, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n5/pdf/scientificamerican07301887-
67a.pdf.

Anonymous:1887:ICCm

[10020] Anonymous. An improved car coupling. Scientific American, 57(5):67, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n5/pdf/scientificamerican07301887-67b.pdf.

Anonymous:1887:NHF

[10021] Anonymous. The new headquarters of the Fire Department of the City of New York. Scientific American, 57(5):68, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n5/pdf/scientificamerican07301887-68.pdf.

Anonymous:1887:PMMb

[10022] Anonymous. A perpetual motion machine. Scientific American, 57(5):69, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n5/pdf/scientificamerican07301887-
69a.pdf.

Anonymous:1887:RBE

[10023] Anonymous. A remarkable boiler explosion. Scientific American, 57(5):70, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n5/pdf/scientificamerican07301887-
70.pdf.

Anonymous:1887:HC

[10024] Anonymous. Holstein cattle. Scientific American, 57(5):71, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n5/pdf/scientificamerican07301887-71.pdf.

Anonymous:1887:TGS

[10025] Anonymous. The two gray seals in the Zoological Garden in Berlin. Scientific American, 57(5):71–72, July 30, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v57/
n5/pdf/scientificamerican07301887-71a.pdf.

Anonymous:1887:BAb

[10026] Anonymous. Bicycle ambulance. Scientific American, 57(5):73, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n5/pdf/scientificamerican07301887-73.pdf.

Anonymous:1887:NNS

[10027] Anonymous. The new Nordenfelt submarine boat. Scientific American, 57(5):73, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n5/pdf/scientificamerican07301887-
73a.pdf.

Anonymous:1887:EIbv

[10028] Anonymous. Engineering inventions. Scientific American, 57(5):74, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n5/pdf/scientificamerican07301887-74.pdf.

Anonymous:1887:AIbp

[10029] Anonymous. Agricultural inventions. Scientific American, 57(5):74, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n5/pdf/scientificamerican07301887-74a.pdf.

Anonymous:1887:MIcf

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n5/pdf/scientificamerican07301887-74b.pdf.

Anonymous:1887:Sq

[10031] Anonymous. Special. Scientific American, 57(5):75, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n5/pdf/scientificamerican07301887-75.pdf.

Anonymous:1887:Icn

[10032] Anonymous. To inventors. Scientific American, 57(5):75, July 30, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n5/pdf/scientificamerican07301887-75d.pdf.

Anonymous:1887:Cbw

[10033] Anonymous. Correspondence. Scientific American, 57(6):85, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n6/pdf/scientificamerican08061887-85a.pdf.

Anonymous:1887:BPbx

[10034] Anonymous. Business and personal. Scientific American, 57(6):91, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n6/pdf/scientificamerican08061887-91c.pdf.

Anonymous:1887:BNBbg

[10035] Anonymous. Books: New books and publications. Scientific American, 57(6):91, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n6/pdf/scientificamerican08061887-
91d.pdf.

Anonymous:1887:IIby

[10036] Anonymous. Index of inventions. Scientific American, 57(6):91–92, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n6/pdf/scientificamerican08061887-91e.pdf.

Anonymous:1887:PNPbu

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n6/pdf/scientificamerican08061887-94.pdf.

Anonymous:1887:ELC

[10038] Anonymous. The Electric Lighting Convention. Scientific American, 57(6):80, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n6/pdf/scientificamerican08061887-
80.pdf.

Anonymous:1887:RRR

[10039] Anonymous. The recent remarkable rainfall. Scientific American, 57(6):81, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n6/pdf/scientificamerican08061887-
81.pdf.

Anonymous:1887:ILC

[10040] Anonymous. An improved lubricator for crank pins, etc. Scientific American, 57(6):82, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n6/pdf/scientificamerican08061887-
82.pdf.

Anonymous:1887:IFE

[10041] Anonymous. An improved fire escape. Scientific American, 57(6):82, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n6/pdf/scientificamerican08061887-82a.pdf.

Anonymous:1887:ISTb

[10042] Anonymous. An improved steam trap. Scientific American, 57(6):82, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n6/pdf/scientificamerican08061887-82b.pdf.

Anonymous:1887:IRSc

[10043] Anonymous. An improved reversible screw driver. Scientific American, 57(6):82, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n6/pdf/scientificamerican08061887-
82c.pdf.

Anonymous:1887:APP

[10044] Anonymous. An automatic perforator for printing presses. Scientific American, 57(6):82, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n6/pdf/scientificamerican08061887-
82d.pdf.

Anonymous:1887:IGBa

[10045] Anonymous. An improved gas burner. Scientific American, 57(6):82–83, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n6/pdf/scientificamerican08061887-82e.pdf.

Anonymous:1887:IIBz

[10046] Anonymous. An improved ironing board. Scientific American, 57(6):83, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n6/pdf/scientificamerican08061887-83.pdf.

Anonymous:1887:IDCe

[10047] Anonymous. An improved device for closing gates. Scientific American, 57(6):83, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n6/pdf/scientificamerican08061887-
83a.pdf.

Anonymous:1887:FCC

[10048] Anonymous. A floor cramp for carpenters, joiners, etc. Scientific American, 57(6):83, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n6/pdf/scientificamerican08061887-
83b.pdf.

Anonymous:1887:ISI

[10049] Anonymous. An improved surgical instrument holder. Scientific American, 57(6):83, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n6/pdf/scientificamerican08061887-
83c.pdf.

Anonymous:1887:IMRb

[10050] Anonymous. An improved method of raising bread, etc. Scientific American, 57(6):83, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n6/pdf/scientificamerican08061887-
83d.pdf.

Anonymous:1887:ISBe

[10051] Anonymous. An improved scaffold bracket. Scientific American, 57(6):83, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n6/pdf/scientificamerican08061887-
83e.pdf.

Anonymous:1887:IRL

[10052] Anonymous. An improved rope ladder. Scientific American, 57 (6):83–84, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n6/pdf/scientificamerican08061887-
83f.pdf.

Anonymous:1887:IFG

[10053] Anonymous. An improved fire grenade. Scientific American, 57(6):84, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n6/pdf/scientificamerican08061887-84.pdf.

Anonymous:1887:IDLb

[10054] Anonymous. An improved door latch and handle. Scientific American, 57(6):84, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n6/pdf/scientificamerican08061887-
84a.pdf.

Anonymous:1887:ILBb

[10055] Anonymous. An improved letter and bill file. Scientific American, 57(6):84, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n6/pdf/scientificamerican08061887-
84b.pdf.

Anonymous:1887:AAb

[10056] Anonymous. The armament of the Atlanta. Scientific American, 57(6):85, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n6/pdf/scientificamerican08061887-
85.pdf.

Anonymous:1887:ELE

[10057] Anonymous. Electric light engines at the Manchester Exhibition. Scientific American, 57(6):86, August 6, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v57/
n6/pdf/scientificamerican08061887-86.pdf.

Anonymous:1887:MSb

[10058] Anonymous. A Malaisian spider. Scientific American, 57(6):87, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n6/pdf/scientificamerican08061887-87.pdf.

Anonymous:1887:DDb

[10059] Anonymous. Disinfection and disinfectants. Scientific American, 57(6):88, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n6/pdf/scientificamerican08061887-
88.pdf.

Anonymous:1887:IIMa

[10060] Anonymous. Improved ice making machinery. Scientific American, 57(6):89–90, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n6/pdf/scientificamerican08061887-
89.pdf.

Anonymous:1887:EIbw

[10061] Anonymous. Engineering inventions. Scientific American, 57(6):91, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n6/pdf/scientificamerican08061887-91.pdf.

Anonymous:1887:AIbq

[10062] Anonymous. Agricultural inventions. Scientific American, 57(6):91, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n6/pdf/scientificamerican08061887-91a.pdf.

Anonymous:1887:MIcg

[10063] Anonymous. Miscellaneous inventions. Scientific American, 57(6):91, August 6, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n6/pdf/scientificamerican08061887-91b.pdf.

Anonymous:1887:Cbx

[10064] Anonymous. Correspondence. Scientific American, 57(7):101, August 13, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n7/pdf/scientificamerican08131887-101.pdf.

Anonymous:1887:BPby

[10065] Anonymous. Business and personal. Scientific American, 57(7):106, August 13, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n7/pdf/scientificamerican08131887-106d.pdf.

Anonymous:1887:BNBbh

[10066] Anonymous. Books: New books and publications. Scientific American, 57(7):106–107, August 13, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n7/pdf/scientificamerican08131887-
106e.pdf.

Anonymous:1887:NQbo

[10067] Anonymous. Notes and queries. Scientific American, 57(7):107, August 13, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n7/pdf/scientificamerican08131887-107.pdf.

Anonymous:1887:IIca

[10068] Anonymous. Index of inventions. Scientific American, 57(7): 107–108, August 13, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n7/pdf/scientificamerican08131887-
107b.pdf.

Anonymous:1887:PNPbv

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n7/pdf/scientificamerican08131887-110.pdf.

Anonymous:1887:MAA

[10070] Anonymous. Meeting of the American Association for the Advancement of Science, new york, 1887. Scientific American, 57(7):95, August 13, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n7/pdf/scientificamerican08131887-95.pdf.

Anonymous:1887:IBLa

[10071] Anonymous. Improved bolt lock for piston heads. Scientific American, 57(7):98, August 13, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n7/pdf/scientificamerican08131887-
98.pdf.

Anonymous:1887:ISBf

[10072] Anonymous. An improved steam boiler. Scientific American, 57 (7):98, August 13, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n7/pdf/scientificamerican08131887-
98a.pdf.

Anonymous:1887:IIEa

[10073] Anonymous. An improved insulator for electric wires. Scientific American, 57(7):98, August 13, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n7/pdf/scientificamerican08131887-
98b.pdf.

Anonymous:1887:DCA

[10074] Anonymous. A device for catching animals. Scientific American, 57(7):98, August 13, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n7/pdf/scientificamerican08131887-
98c.pdf.

Anonymous:1887:SEW

[10075] Anonymous. A simple and efficient washing machine. Scientific American, 57(7):98, August 13, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n7/pdf/scientificamerican08131887-
98d.pdf.

Anonymous:1887:ICBf

[10076] Anonymous. Improved cutter bar for mowers and reapers. Scientific American, 57(7):99, August 13, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n7/pdf/scientificamerican08131887-
99.pdf.

Anonymous:1887:IRW

[10077] Anonymous. An improved rock washing apparatus. Scientific American, 57(7):99, August 13, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n7/pdf/scientificamerican08131887-
99a.pdf.

Anonymous:1887:IOB

[10078] Anonymous. An improved ox bow. Scientific American, 57(7):99, August 13, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n7/pdf/scientificamerican08131887-99b.pdf.

Anonymous:1887:RLO

[10079] Anonymous. Ripening of liquors by ozone. Scientific American, 57(7):99–100, August 13, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n7/pdf/scientificamerican08131887-
99c.pdf.

Anonymous:1887:BT

[10080] Anonymous. The Bonta telephone. Scientific American, 57(7): 102–103, August 13, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n7/pdf/scientificamerican08131887-
102.pdf.

Benjamin:1887:ESM

[10081] Marcus Benjamin. Edward Sylvester Morse. Scientific American, 57(7):103–104, August 13, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n7/pdf/scientificamerican08131887-
103.pdf.

Seiss:1887:FK

[10082] C. Few Seiss. The fish killer. Scientific American, 57(7):104, August 13, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n7/pdf/scientificamerican08131887-104.pdf.

Anonymous:1887:HAM

[10083] Anonymous. The house of Adelantado Montejo. Scientific American, 57(7):105, August 13, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n7/pdf/scientificamerican08131887-
105.pdf.

Anonymous:1887:EIbx

[10084] Anonymous. Engineering inventions. Scientific American, 57(7): 106, August 13, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n7/pdf/scientificamerican08131887-
106.pdf.

Anonymous:1887:AIbr

[10085] Anonymous. Agricultural invention. Scientific American, 57(7): 106, August 13, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n7/pdf/scientificamerican08131887-
106a.pdf.

Anonymous:1887:MIch

[10086] Anonymous. Miscellaneous inventions. Scientific American, 57 (7):106, August 13, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n7/pdf/scientificamerican08131887-
106b.pdf.

Anonymous:1887:Sr

[10087] Anonymous. Special. Scientific American, 57(7):106, August 13, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n7/pdf/scientificamerican08131887-106c.pdf.

Anonymous:1887:Ico

[10088] Anonymous. To inventors. Scientific American, 57(7):107, August 13, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n7/pdf/scientificamerican08131887-107a.pdf.

Anonymous:1887:BPbz

[10089] Anonymous. Business and personal. Scientific American, 57(8):122, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n8/pdf/scientificamerican08201887-122c.pdf.

Anonymous:1887:BNBbi

[10090] Anonymous. Books: New books and publications. Scientific American, 57(8):122–123, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n8/pdf/scientificamerican08201887-
122d.pdf.

Anonymous:1887:NQbp

[10091] Anonymous. Notes and queries. Scientific American, 57(8):123, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n8/pdf/scientificamerican08201887-123.pdf.

Anonymous:1887:IIcb

[10092] Anonymous. Index of inventions. Scientific American, 57(8): 123–124, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n8/pdf/scientificamerican08201887-
123b.pdf.

Anonymous:1887:PNPbw

[10093] Anonymous. Patent notes: Patents. Scientific American, 57(8):126, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n8/pdf/scientificamerican08201887-126.pdf.

Anonymous:1887:PGC

[10094] Anonymous. Pneumatic gun carriage at sandy hook. Scientific American, 57(8):111, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n8/pdf/scientificamerican08201887-
111.pdf.

Anonymous:1887:AAAa

[10095] Anonymous. The American Association for the Advancement of Science. Scientific American, 57(8):112–113, August 20, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v57/
n8/pdf/scientificamerican08201887-112.pdf.

Anonymous:1887:RPEa

[10096] Anonymous. Revival of patent extensions. Scientific American, 57 (8):113–114, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n8/pdf/scientificamerican08201887-
113.pdf.

Anonymous:1887:IMRc

[10097] Anonymous. Improved mechanism for reciprocating plungers. Scientific American, 57(8):114, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n8/pdf/scientificamerican08201887-
114.pdf.

Anonymous:1887:ICDa

[10098] Anonymous. An improved cistern device. Scientific American, 57(8):114, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n8/pdf/scientificamerican08201887-
114a.pdf.

Anonymous:1887:IRG

[10099] Anonymous. An improved reversing gear for engines. Scientific American, 57(8):114, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n8/pdf/scientificamerican08201887-
114b.pdf.

Anonymous:1887:SAR

[10100] Anonymous. A safety appliance for railroad cars. Scientific American, 57(8):114, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n8/pdf/scientificamerican08201887-
114c.pdf.

Anonymous:1887:IWH

[10101] Anonymous. An improved wardrobe hook and shade fixture. Scientific American, 57(8):115, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n8/pdf/scientificamerican08201887-
115.pdf.

Anonymous:1887:IGc

[10102] Anonymous. An improved governor. Scientific American, 57 (8):115, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n8/pdf/scientificamerican08201887-
115a.pdf.

Anonymous:1887:IGT

[10103] Anonymous. An improved gearing for transmitting power. Scientific American, 57(8):115, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n8/pdf/scientificamerican08201887-
115b.pdf.

Anonymous:1887:SFSb

[10104] Anonymous. A sack-filling spout for flour packers. Scientific American, 57(8):115, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n8/pdf/scientificamerican08201887-
115c.pdf.

Anonymous:1887:ATR

[10105] Anonymous. Attachment for top rail of vehicle seat. Scientific American, 57(8):115, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n8/pdf/scientificamerican08201887-
115d.pdf.

Anonymous:1887:IBBa

[10106] Anonymous. An improved bit brace. Scientific American, 57 (8):116, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n8/pdf/scientificamerican08201887-
116.pdf.

Anonymous:1887:SPB

[10107] Anonymous. Simple plunge battery. Scientific American, 57(8):116, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n8/pdf/scientificamerican08201887-116a.pdf.

Anonymous:1887:IMPb

[10108] Anonymous. Improvement in manufacturing plated ware. Scientific American, 57(8):116, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n8/pdf/scientificamerican08201887-
116b.pdf.

Anonymous:1887:IFPd

[10109] Anonymous. An improved fire-place protector. Scientific American, 57(8):116–117, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n8/pdf/scientificamerican08201887-
116c.pdf.

Anonymous:1887:HMO

[10110] Anonymous. H.M.S. Orlando. Scientific American, 57(8):118, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n8/pdf/scientificamerican08201887-118.pdf.

Benjamin:1887:SPL

[10111] Marcus Benjamin. Samuel Pierpont Langley. Scientific American, 57(8):119–120, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n8/pdf/scientificamerican08201887-
119.pdf.

Anonymous:1887:EST

[10112] Anonymous. Experiments with the “scientific top”. Scientific American, 57(8):121, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n8/pdf/scientificamerican08201887-
121.pdf.

Anonymous:1887:EIby (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n8/pdf/scientificamerican08201887-
122.pdf.

Anonymous:1887:AIbs

[10114] Anonymous. Agricultural inventions. Scientific American, 57(8): 122, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n8/pdf/scientificamerican08201887-
122a.pdf.

Anonymous:1887:MIci

[10115] Anonymous. Miscellaneous inventions. Scientific American, 57 (8):122, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n8/pdf/scientificamerican08201887-
122b.pdf.

Anonymous:1887:Icq

[10116] Anonymous. To inventors. Scientific American, 57(8):123, August 20, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n8/pdf/scientificamerican08201887-123a.pdf.

Anonymous:1887:BPca

[10117] Anonymous. Business and personal. Scientific American, 57(9):139, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n9/pdf/scientificamerican08271887-139d.pdf.

Anonymous:1887:IIcc

[10118] Anonymous. Index of inventions. Scientific American, 57(9): 139–140, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n9/pdf/scientificamerican08271887-
139f.pdf.

Anonymous:1887:PNPbx

[10119] Anonymous. Patent notes: Patents. Scientific American, 57(9):142, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n9/pdf/scientificamerican08271887-142.pdf.

Anonymous:1887:AS

[10120] Anonymous. Advancement of science. Scientific American, 57 (9):128–129, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n9/pdf/scientificamerican08271887-
128.pdf.

Anonymous:1887:PPS

[10121] Anonymous. Positions of the planets in September. Scientific American, 57(9):129, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n9/pdf/scientificamerican08271887-
129.pdf.

Anonymous:1887:BUO

[10122] Anonymous. The blot upon our statute books. Scientific American, 57(9):129, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n9/pdf/scientificamerican08271887-
129a.pdf.

Anonymous:1887:PNr

[10123] Anonymous. Photographic notes. Scientific American, 57(9):129, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n9/pdf/scientificamerican08271887-129b.pdf.

Anonymous:1887:LAR

[10124] Anonymous. Locomotive attachment for removing obstructions from railroad tracks. Scientific American, 57(9):130, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v57/
n9/pdf/scientificamerican08271887-130.pdf.

Anonymous:1887:ICCn

[10125] Anonymous. An improved car coupler. Scientific American, 57 (9):130, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n9/pdf/scientificamerican08271887-
130a.pdf.

Anonymous:1887:IFR

[10126] Anonymous. An improved foot rest. Scientific American, 57(9):130, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n9/pdf/scientificamerican08271887-130b.pdf.

Anonymous:1887:SMC

[10127] Anonymous. A sewing machine cover and stool. Scientific American, 57(9):130, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n9/pdf/scientificamerican08271887-
130c.pdf.

Anonymous:1887:ICu

[10128] Anonymous. An improved churn. Scientific American, 57(9): 130–131, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n9/pdf/scientificamerican08271887-
130d.pdf.

Anonymous:1887:CMJ

[10129] Anonymous. Combined measuring jackets and patterns. Scientific American, 57(9):131, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n9/pdf/scientificamerican08271887-
131.pdf.

Anonymous:1887:IHCb

[10130] Anonymous. An improved horse collar. Scientific American, 57 (9):131, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n9/pdf/scientificamerican08271887-
131a.pdf.

Anonymous:1887:IOd

[10131] Anonymous. An improved odometer. Scientific American, 57 (9):131, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n9/pdf/scientificamerican08271887-
131b.pdf.

Anonymous:1887:MFS

[10132] Anonymous. A machine for forming square tin cans. Scientific American, 57(9):131, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n9/pdf/scientificamerican08271887-
131c.pdf.

Anonymous:1887:IPg

[10133] Anonymous. An improved pump. Scientific American, 57(9): 131–132, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n9/pdf/scientificamerican08271887-
131d.pdf.

Anonymous:1887:IECb

[10134] Anonymous. An improved electrical car brake. Scientific American, 57(9):132, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n9/pdf/scientificamerican08271887-
132.pdf.

Edison:1887:PDM

[10135] Thomas A. Edison. On the pyromagnetic dynamo — a machine for producing electricity directly from fuel. Scientific American, 57(9):133, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n9/pdf/scientificamerican08271887-133.pdf.

Anonymous:1887:MHW

[10136] Anonymous. Millot’s hydraulic wheel. Scientific American, 57 (9):134, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n9/pdf/scientificamerican08271887-
134.pdf.

Baker:1887:FFB

[10137] B. Baker. The Firth of Forth Bridge. Scientific American, 57 (9):134–135, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n9/pdf/scientificamerican08271887-
134a.pdf.

Anonymous:1887:BEP

[10138] Anonymous. Bruijn’s echidna (Proechidna Bruijni), New Guinea ant eater. Scientific American, 57(9):135, August 27, 1887. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v57/
n9/pdf/scientificamerican08271887-135.pdf.

Anonymous:1887:TW

[10139] Anonymous. A tattooed whale. Scientific American, 57(9):135–137, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n9/pdf/scientificamerican08271887-135a.pdf.

Hopkins:1887:MF

[10140] Geo. M. Hopkins. Musical flames. Scientific American, 57(9): 137–138, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n9/pdf/scientificamerican08271887-
137.pdf.

Anonymous:1887:EIbz

[10141] Anonymous. Engineering inventions. Scientific American, 57(9): 139, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n9/pdf/scientificamerican08271887-
139.pdf.

Anonymous:1887:AIbt

[10142] Anonymous. Agricultural inventions. Scientific American, 57(9): 139, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n9/pdf/scientificamerican08271887-
139a.pdf.

Anonymous:1887:MIcj

[10143] Anonymous. Miscellaneous inventions. Scientific American, 57 (9):139, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v57/n9/pdf/scientificamerican08271887-
139b.pdf.

Anonymous:1887:Ss

[10144] Anonymous. Special. Scientific American, 57(9):139, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n9/pdf/scientificamerican08271887-139c.pdf.

Anonymous:1887:Icv

[10145] Anonymous. To inventors. Scientific American, 57(9):139, August 27, 1887. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v57/n9/pdf/scientificamerican08271887-139e.pdf.

Anonymous:1888:Ca

[10146] Anonymous. Correspondence. Scientific American, 58(1):10, January 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n1/pdf/scientificamerican01071888-10.pdf.

Anonymous:1888:BPa

[10147] Anonymous. Business and personal. Scientific American, 58(1):11, January 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n1/pdf/scientificamerican01071888-11c.pdf.

Anonymous:1888:IIa

[10148] Anonymous. Index of inventions. Scientific American, 58(1):11–12, January 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n1/pdf/scientificamerican01071888-11e.pdf.

Anonymous:1888:PNPa

[10149] Anonymous. Patent notes: Patents. Scientific American, 58(1):14, January 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n1/pdf/scientificamerican01071888-14.pdf.

Anonymous:1888:WHM

[10150] Anonymous. Works of the Holly Manufacturing Co., at Lockport, N.Y. Scientific American, 58(1):1, January 7, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v58/
n1/pdf/scientificamerican01071888-1.pdf.

Anonymous:1888:PJI

[10151] Anonymous. Patents by joint inventors. Scientific American, 58(1):2, January 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n1/pdf/scientificamerican01071888-2.pdf.

Anonymous:1888:TBG

[10152] Anonymous. Torpedo boats in a gale. Scientific American, 58(1):2, January 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n1/pdf/scientificamerican01071888-2a.pdf.

Anonymous:1888:CWa

[10153] Anonymous. The celestial world. Scientific American, 58(1):2–3, January 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n1/pdf/scientificamerican01071888-2b.pdf.

Anonymous:1888:PMS

[10154] Anonymous. Patent maple sugar. Scientific American, 58(1):3, January 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n1/pdf/scientificamerican01071888-3.pdf.

Anonymous:1888:PNa

[10155] Anonymous. Photographic notes. Scientific American, 58(1):3, January 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n1/pdf/scientificamerican01071888-3a.pdf.

Anonymous:1888:STa

[10156] Anonymous. A scientific toy. Scientific American, 58(1):5, January 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n1/pdf/scientificamerican01071888-5.pdf.

Anonymous:1888:IRB

[10157] Anonymous. An improved removable barrel head. Scientific American, 58(1):6, January 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n1/pdf/scientificamerican01071888-
6.pdf.

Anonymous:1888:IAS

[10158] Anonymous. An improved air ship. Scientific American, 58(1):6, January 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n1/pdf/scientificamerican01071888-6a.pdf.

Anonymous:1888:ISHa

[10159] Anonymous. An improved sleeve holder. Scientific American, 58(1):6, January 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n1/pdf/scientificamerican01071888-6b.pdf.

Anonymous:1888:IHEa

[10160] Anonymous. An improved hydraulic elevator. Scientific American, 58(1):6, January 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n1/pdf/scientificamerican01071888-
6c.pdf.

Anonymous:1888:MAH

[10161] Anonymous. A method and apparatus for heating cars. Scientific American, 58(1):6, January 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n1/pdf/scientificamerican01071888-
6d.pdf.

Anonymous:1888:WFF

[10162] Anonymous. Wedge ferrule for fishing-rod reels. Scientific American, 58(1):6–7, January 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n1/pdf/scientificamerican01071888-
6e.pdf.

Anonymous:1888:IFEa

[10163] Anonymous. Improved fire extinguisher for car heaters. Scientific American, 58(1):7, January 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n1/pdf/scientificamerican01071888-
7.pdf.

Anonymous:1888:FDG

[10164] Anonymous. Friction driving gear for dynamos. Scientific American, 58(1):8, January 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n1/pdf/scientificamerican01071888-
8.pdf.

Anonymous:1888:IAMa

[10165] Anonymous. Improved attachment for mortising machines. Scientific American, 58(1):8, January 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n1/pdf/scientificamerican01071888-
8a.pdf.

B:1888:FVH

[10166] M. B. Ferdinand Vanderveer Hayden. Scientific American, 58(1):9, January 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n1/pdf/scientificamerican01071888-9.pdf.

Anonymous:1888:EIa

[10167] Anonymous. Engineering inventions. Scientific American, 58(1):11, January 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n1/pdf/scientificamerican01071888-11.pdf.

Anonymous:1888:AIa

[10168] Anonymous. Agricultural inventions. Scientific American, 58(1):11, January 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n1/pdf/scientificamerican01071888-11a.pdf.

Anonymous:1888:MIa

[10169] Anonymous. Miscellaneous inventions. Scientific American, 58(1):11, January 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n1/pdf/scientificamerican01071888-11b.pdf.

Anonymous:1888:Ia

[10170] Anonymous. To inventors. Scientific American, 58(1):11, January 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n1/pdf/scientificamerican01071888-11d.pdf.

Anonymous:1888:BNBa

[10171] Anonymous. Books: New books and publications. Scientific American, 58(10):154, March 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n10/pdf/scientificamerican03101888-
154c.pdf.

Anonymous:1888:BPb

[10172] Anonymous. Business and personal. Scientific American, 58 (10):154–155, March 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n10/pdf/scientificamerican03101888-
154e.pdf.

Anonymous:1888:NQa

[10173] Anonymous. Notes and queries. Scientific American, 58(10):155, March 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n10/pdf/scientificamerican03101888-155.pdf.

Anonymous:1888:IIb

[10174] Anonymous. Index of inventions. Scientific American, 58(10):155–156, March 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n10/pdf/scientificamerican03101888-155b.pdf.

Anonymous:1888:PNPb

[10175] Anonymous. Patent notes: Patents. Scientific American, 58(10):158, March 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n10/pdf/scientificamerican03101888-158.pdf.

Anonymous:1888:MPI

[10176] Anonymous. Minor phenomena of the incandescent electric lamp. Scientific American, 58(10):144, March 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v58/n10/pdf/scientificamerican03101888-
144.pdf.

Anonymous:1888:TMW

[10177] Anonymous. Tests for modern war ships. Scientific American, 58(10):144, March 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n10/pdf/scientificamerican03101888-
144a.pdf.

Anonymous:1888:KES

[10178] Anonymous. The kerosene emulsion for scale insects. Scientific American, 58(10):144–146, March 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n10/pdf/scientificamerican03101888-
144b.pdf.

Anonymous:1888:ESC

[10179] Anonymous. Engines of the steamship Courier. Scientific American, 58(10):146–147, March 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n10/pdf/scientificamerican03101888-
146.pdf.

Anonymous:1888:DCC

[10180] Anonymous. A defective ceiling in the Capitol Building, Albany, N.Y. Scientific American, 58(10):148–149, March 10, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v58/
n10/pdf/scientificamerican03101888-148.pdf.

Anonymous:1888:ICCa

[10181] Anonymous. An improved car coupling. Scientific American, 58(10):149, March 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n10/pdf/scientificamerican03101888-
149.pdf.

Anonymous:1888:IWSa

[10182] Anonymous. An improved weighing scale. Scientific American, 58 (10):149–150, March 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n10/pdf/scientificamerican03101888-
149a.pdf.

Anonymous:1888:FE

[10183] Anonymous. The French Exhibition of 1889. Scientific American, 58(10):150–153, March 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n10/pdf/scientificamerican03101888-
150.pdf.

Anonymous:1888:PNS

[10184] Anonymous. Pallas’ northern sea eagle. Scientific American, 58(10):153, March 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n10/pdf/scientificamerican03101888-153.pdf.

Anonymous:1888:EIb

[10185] Anonymous. Engineering inventions. Scientific American, 58(10):154, March 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n10/pdf/scientificamerican03101888-154.pdf.

Anonymous:1888:AIb

[10186] Anonymous. Agricultural inventions. Scientific American, 58(10):154, March 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n10/pdf/scientificamerican03101888-154a.pdf.

Anonymous:1888:MIb

[10187] Anonymous. Miscellaneous inventions. Scientific American, 58(10):154, March 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n10/pdf/scientificamerican03101888-154b.pdf.

Anonymous:1888:Sa

[10188] Anonymous. Special. Scientific American, 58(10):154, March 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n10/pdf/scientificamerican03101888-154d.pdf.

Anonymous:1888:Ib

[10189] Anonymous. To inventors. Scientific American, 58(10):155, March 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n10/pdf/scientificamerican03101888-155a.pdf.

Anonymous:1888:Cb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n11/pdf/scientificamerican03171888-168.pdf.

Anonymous:1888:BPc

[10191] Anonymous. Business and personal. Scientific American, 58(11):170, March 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n11/pdf/scientificamerican03171888-170c.pdf.

Anonymous:1888:BNBb

[10192] Anonymous. Books: New books and publications. Scientific American, 58(11):170, March 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n11/pdf/scientificamerican03171888-
170d.pdf.

Anonymous:1888:NQb

[10193] Anonymous. Notes and queries. Scientific American, 58(11):171, March 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n11/pdf/scientificamerican03171888-171.pdf.

Anonymous:1888:IIc

[10194] Anonymous. Index of inventions. Scientific American, 58(11):171–172, March 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n11/pdf/scientificamerican03171888-171b.pdf.

Anonymous:1888:PNPc

[10195] Anonymous. Patent notes: Patents. Scientific American, 58(11):174, March 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n11/pdf/scientificamerican03171888-174.pdf.

Anonymous:1888:LOU

[10196] Anonymous. The Lick Observatory of the University of California. Scientific American, 58(11):159, March 17, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v58/
n11/pdf/scientificamerican03171888-159.pdf.

Anonymous:1888:PNb

[10197] Anonymous. Photographic notes. Scientific American, 58(11):161, March 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n11/pdf/scientificamerican03171888-161.pdf.

Sloane:1888:CL

[10198] T. O’conor Sloane. Cohesion of lead. Scientific American, 58 (11):163–164, March 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n11/pdf/scientificamerican03171888-
163.pdf.

Anonymous:1888:DUSa

[10199] Anonymous. A device for unloading and stacking hay. Scientific American, 58(11):164, March 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n11/pdf/scientificamerican03171888-
164.pdf.

Anonymous:1888:ICCb

[10200] Anonymous. An improved collar or cuff button. Scientific American, 58(11):164, March 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n11/pdf/scientificamerican03171888-
164a.pdf.

Anonymous:1888:IGSa

[10201] Anonymous. An improved gate spring. Scientific American, 58(11):164, March 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n11/pdf/scientificamerican03171888-164b.pdf.

Anonymous:1888:IBBa

[10202] Anonymous. An improved brake block for vehicles. Scientific American, 58(11):164, March 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n11/pdf/scientificamerican03171888-
164c.pdf.

Anonymous:1888:ISDa

[10203] Anonymous. An improved stand for displaying goods. Scientific American, 58(11):164, March 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n11/pdf/scientificamerican03171888-
164d.pdf.

Hopkins:1888:SEM

[10204] M. Hopkins. Simple electric motor. Scientific American, 58 (11):165–166, March 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n11/pdf/scientificamerican03171888-
165.pdf.

Anonymous:1888:LPC

[10205] Anonymous. The locks of the Panama Canal. Scientific American, 58(11):166–167, March 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n11/pdf/scientificamerican03171888-
166.pdf.

Anonymous:1888:DFB

[10206] Anonymous. A device to facilitate blasting stumps, logs, etc. Scientific American, 58(11):168, March 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n11/pdf/scientificamerican03171888-
168a.pdf.

Anonymous:1888:MTV

[10207] Anonymous. A machine for truing valve plugs, etc. Scientific American, 58(11):168, March 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n11/pdf/scientificamerican03171888-
168b.pdf.

Anonymous:1888:HL

[10208] Anonymous. The horned lizard. Scientific American, 58(11):169, March 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n11/pdf/scientificamerican03171888-169.pdf.

Anonymous:1888:EIc

[10209] Anonymous. Engineering invention. Scientific American, 58(11):170, March 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n11/pdf/scientificamerican03171888-170.pdf.

Anonymous:1888:AIc

[10210] Anonymous. Agricultural inventions. Scientific American, 58(11):170, March 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n11/pdf/scientificamerican03171888-170a.pdf.

Anonymous:1888:MIc

[10211] Anonymous. Miscellaneous inventions. Scientific American, 58(11):170, March 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n11/pdf/scientificamerican03171888-170b.pdf.

Anonymous:1888:Ic

[10212] Anonymous. To inventors. Scientific American, 58(11):171, March 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n11/pdf/scientificamerican03171888-171a.pdf.

Anonymous:1888:Cc

[10213] Anonymous. Correspondence. Scientific American, 58(12):184, March 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n12/pdf/scientificamerican03241888-184.pdf.

Anonymous:1888:BPd

[10214] Anonymous. Business and personal. Scientific American, 58(12):186, March 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n12/pdf/scientificamerican03241888-186c.pdf.

Anonymous:1888:BNBc

[10215] Anonymous. Books: New books and publications. Scientific American, 58(12):186–187, March 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n12/pdf/scientificamerican03241888-
186d.pdf.

Anonymous:1888:NQc

[10216] Anonymous. Notes and queries. Scientific American, 58(12):187, March 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n12/pdf/scientificamerican03241888-187.pdf.

Anonymous:1888:IId

[10217] Anonymous. Index of inventions. Scientific American, 58(12):187–188, March 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n12/pdf/scientificamerican03241888-187b.pdf.

Anonymous:1888:PNPd

[10218] Anonymous. Patent notes: Patents. Scientific American, 58(12):190, March 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n12/pdf/scientificamerican03241888-190.pdf.

Anonymous:1888:HB

[10219] Anonymous. Henry Bergh. Scientific American, 58(12):176, March 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n12/pdf/scientificamerican03241888-176.pdf.

Anonymous:1888:RSSa

[10220] Anonymous. Rotary steam shovels. Scientific American, 58(12):176, March 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n12/pdf/scientificamerican03241888-176a.pdf.

Anonymous:1888:CSa

[10221] Anonymous. Car stoves. Scientific American, 58(12):176, March 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n12/pdf/scientificamerican03241888-176b.pdf.

Anonymous:1888:LE

[10222] Anonymous. The lesson of an emergency. Scientific American, 58 (12):176–177, March 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n12/pdf/scientificamerican03241888-
176c.pdf.

Anonymous:1888:GSN

[10223] Anonymous. The great storm in New York and vicinity. Scientific American, 58(12):178–180, March 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n12/pdf/scientificamerican03241888-
178.pdf.

Anonymous:1888:ENS

[10224] Anonymous. Edison’s new system of telegraphy. Scientific American, 58(12):180–181, March 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n12/pdf/scientificamerican03241888-
180.pdf.

Anonymous:1888:IBBb

[10225] Anonymous. An improved bridle bit. Scientific American, 58(12):181, March 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n12/pdf/scientificamerican03241888-181.pdf.

Anonymous:1888:IQ

[10226] Anonymous. An improved quilter. Scientific American, 58(12):181, March 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n12/pdf/scientificamerican03241888-181a.pdf.

Anonymous:1888:RCC

[10227] Anonymous. A railroad for common carriages or wagons. Scientific American, 58(12):182, March 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n12/pdf/scientificamerican03241888-
182.pdf.

Anonymous:1888:IPMa

[10228] Anonymous. Improved paint mill. Scientific American, 58(12):182, March 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n12/pdf/scientificamerican03241888-182a.pdf.

Anonymous:1888:HBH

[10229] Anonymous. Heating bodies to a high temperature in a compressed gas. Scientific American, 58(12):183, March 24, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v58/
n12/pdf/scientificamerican03241888-183.pdf.

Anonymous:1888:CK

[10230] Anonymous. Chinese kites. Scientific American, 58(12):185, March 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n12/pdf/scientificamerican03241888-185.pdf.

Anonymous:1888:AId

[10231] Anonymous. Agricultural inventions. Scientific American, 58(12):186, March 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n12/pdf/scientificamerican03241888-186.pdf.

Anonymous:1888:MId

[10232] Anonymous. Miscellaneous inventions. Scientific American, 58(12):186, March 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n12/pdf/scientificamerican03241888-186a.pdf.

Anonymous:1888:Sb

[10233] Anonymous. Special. Scientific American, 58(12):186, March 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n12/pdf/scientificamerican03241888-186b.pdf.

Anonymous:1888:Id

[10234] Anonymous. To inventors. Scientific American, 58(12):187, March 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n12/pdf/scientificamerican03241888-187a.pdf.

Anonymous:1888:BPe

[10235] Anonymous. Business and personal. Scientific American, 58(13):202, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n13/pdf/scientificamerican03311888-202c.pdf.

Anonymous:1888:BNBd

[10236] Anonymous. Books: New books and publications. Scientific American, 58(13):203, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n13/pdf/scientificamerican03311888-
203.pdf.

Anonymous:1888:NQd

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n13/pdf/scientificamerican03311888-203a.pdf.

Anonymous:1888:IIe

[10238] Anonymous. Index of inventions. Scientific American, 58(13):203–204, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n13/pdf/scientificamerican03311888-203c.pdf.

Anonymous:1888:PNPe

[10239] Anonymous. Patent notes: Patents. Scientific American, 58(13):206, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n13/pdf/scientificamerican03311888-206.pdf.

Anonymous:1888:RF

[10240] Anonymous. Repairing a foundation. Scientific American, 58(13):191, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n13/pdf/scientificamerican03311888-191.pdf.

Anonymous:1888:BTC

[10241] Anonymous. The Bell Telephone Cases decided — the patent sustained. Scientific American, 58(13):192–193, March 31, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v58/
n13/pdf/scientificamerican03311888-192.pdf.

Anonymous:1888:CJW

[10242] Anonymous. Chief justice waits. Scientific American, 58(13):193, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n13/pdf/scientificamerican03311888-193.pdf.

Anonymous:1888:IK

[10243] Anonymous. An improved knapsack. Scientific American, 58(13):194, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n13/pdf/scientificamerican03311888-194.pdf.

Anonymous:1888:IRSa

[10244] Anonymous. An improved railroad switch. Scientific American, 58(13):194, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n13/pdf/scientificamerican03311888-
194a.pdf.

Anonymous:1888:IEGa

[10245] Anonymous. An improved extension gauge. Scientific American, 58(13):195, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n13/pdf/scientificamerican03311888-
195.pdf.

Anonymous:1888:NSI

[10246] Anonymous. New sixteen inch swing engine lathe. Scientific American, 58(13):195, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n13/pdf/scientificamerican03311888-
195a.pdf.

Anonymous:1888:IMR

[10247] Anonymous. An improved metallic railway tie. Scientific American, 58(13):195, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n13/pdf/scientificamerican03311888-
195b.pdf.

Anonymous:1888:BSC

[10248] Anonymous. A boiler for steaming and cooking feed. Scientific American, 58(13):195, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n13/pdf/scientificamerican03311888-
195c.pdf.

Anonymous:1888:FAL

[10249] Anonymous. A fire alarm for large buildings. Scientific American, 58(13):196, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n13/pdf/scientificamerican03311888-
196.pdf.

Anonymous:1888:ICCc

[10250] Anonymous. An improved car coupling. Scientific American, 58(13):196, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n13/pdf/scientificamerican03311888-
196a.pdf.

Anonymous:1888:IHDa

[10251] Anonymous. An improved horse detacher. Scientific American, 58(13):196, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n13/pdf/scientificamerican03311888-
196b.pdf.

Anonymous:1888:IVR

[10252] Anonymous. An improved ventilator for railway cars. Scientific American, 58(13):196, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n13/pdf/scientificamerican03311888-
196c.pdf.

Anonymous:1888:FSV

[10253] Anonymous. A fender spar for vessels. Scientific American, 58 (13):196–197, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n13/pdf/scientificamerican03311888-
196d.pdf.

Anonymous:1888:ISHb

[10254] Anonymous. Improved stub holder for check books, etc. Scientific American, 58(13):197, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n13/pdf/scientificamerican03311888-
197.pdf.

Anonymous:1888:BST

[10255] Anonymous. Belt stretching and testing machine. Scientific American, 58(13):197, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n13/pdf/scientificamerican03311888-
197a.pdf.

Anonymous:1888:YTZ

[10256] Anonymous. The young tapir in the Zoological Garden at Cologne. Scientific American, 58(13):199–200, March 31, 1888. CODEN SCAMAC.

ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v58/n13/pdf/scientificamerican03311888-
198.pdf.

Anonymous:1888:MFL

[10257] Anonymous. Magnesium flash lamps. Scientific American, 58 (13):199–200, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n13/pdf/scientificamerican03311888-
199.pdf.

Anonymous:1888:IFTa

[10258] Anonymous. An improved folding tent of lodge. Scientific American, 58(13):200, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n13/pdf/scientificamerican03311888-
200.pdf.

Anonymous:1888:FST

[10259] Anonymous. Farwell’s saw table adjuster. Scientific American, 58(13):200, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n13/pdf/scientificamerican03311888-
200a.pdf.

Anonymous:1888:BFD

[10260] Anonymous. Building fronts on diagonal streets. Scientific American, 58(13):200, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n13/pdf/scientificamerican03311888-
200b.pdf.

Anonymous:1888:NTBa

[10261] Anonymous. A novelty in tooth brushes. Scientific American, 58(13):200, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n13/pdf/scientificamerican03311888-
200c.pdf.

Anonymous:1888:IGa

[10262] Anonymous. An improved grubber. Scientific American, 58(13): 200–201, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n13/pdf/scientificamerican03311888-
200d.pdf.

Anonymous:1888:SAa

[10263] Anonymous. Scientific amusements. Scientific American, 58(13):201, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n13/pdf/scientificamerican03311888-201.pdf.

Anonymous:1888:EId

[10264] Anonymous. Engineering inventions. Scientific American, 58(13):202, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n13/pdf/scientificamerican03311888-202.pdf.

Anonymous:1888:AIe

[10265] Anonymous. Agricultural invention. Scientific American, 58(13):202, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n13/pdf/scientificamerican03311888-202a.pdf.

Anonymous:1888:MIe

[10266] Anonymous. Miscellaneous inventions. Scientific American, 58(13):202, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n13/pdf/scientificamerican03311888-202b.pdf.

Anonymous:1888:Ie

[10267] Anonymous. To inventors. Scientific American, 58(13):203, March 31, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n13/pdf/scientificamerican03311888-203b.pdf.

Anonymous:1888:Cd

[10268] Anonymous. Correspondence. Scientific American, 58(14):217, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n14/pdf/scientificamerican04071888-217.pdf.

Anonymous:1888:BPf

[10269] Anonymous. Business and personal. Scientific American, 58 (14):218–219, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n14/pdf/scientificamerican04071888-
218d.pdf.

Anonymous:1888:NQe

[10270] Anonymous. Notes and queries. Scientific American, 58(14):219, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n14/pdf/scientificamerican04071888-219.pdf.

Anonymous:1888:IIf

[10271] Anonymous. Index of inventions. Scientific American, 58(14):219–220, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n14/pdf/scientificamerican04071888-219b.pdf.

Anonymous:1888:PNPf

[10272] Anonymous. Patent notes: Patents. Scientific American, 58(14):222, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n14/pdf/scientificamerican04071888-222.pdf.

Anonymous:1888:TFC

[10273] Anonymous. Trees and flowers in the Central Park. Scientific American, 58(14):208, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n14/pdf/scientificamerican04071888-
208.pdf.

Anonymous:1888:PPAa

[10274] Anonymous. Position of the planets for April. Scientific American, 58(14):209–210, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n14/pdf/scientificamerican04071888-
209.pdf.

Anonymous:1888:HIR

[10275] Anonymous. Hawley’s improved railway tie. Scientific American, 58(14):210, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n14/pdf/scientificamerican04071888-
210.pdf.

Anonymous:1888:SAAa

[10276] Anonymous. A safety attachment for anchors. Scientific American, 58(14):211, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n14/pdf/scientificamerican04071888-
211.pdf.

Anonymous:1888:ISe

[10277] Anonymous. An improved stacker. Scientific American, 58(14):211, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n14/pdf/scientificamerican04071888-211a.pdf.

Anonymous:1888:ILM

[10278] Anonymous. An improved lawn mower. Scientific American, 58(14):211, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n14/pdf/scientificamerican04071888-211b.pdf.

Anonymous:1888:ISLa

[10279] Anonymous. An improved step ladder. Scientific American, 58(14):211, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n14/pdf/scientificamerican04071888-211c.pdf.

Anonymous:1888:IDEa

[10280] Anonymous. An improved draught equalizer. Scientific American, 58(14):211, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n14/pdf/scientificamerican04071888-
211d.pdf.

Anonymous:1888:ITa

[10281] Anonymous. An improved trap. Scientific American, 58(14):212, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n14/pdf/scientificamerican04071888-212.pdf.

Anonymous:1888:IPWa

[10282] Anonymous. An improved pocket weighing scale. Scientific American, 58(14):212, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n14/pdf/scientificamerican04071888-
212a.pdf.

Anonymous:1888:INLa

[10283] Anonymous. An improved nut lock. Scientific American, 58(14):212, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n14/pdf/scientificamerican04071888-212b.pdf.

Anonymous:1888:ISSa

[10284] Anonymous. An improved sleeve stay. Scientific American, 58(14):212, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n14/pdf/scientificamerican04071888-212c.pdf.

Anonymous:1888:IDC

[10285] Anonymous. An improved door closer. Scientific American, 58(14):212, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n14/pdf/scientificamerican04071888-212d.pdf.

Anonymous:1888:RCF

[10286] Anonymous. The Reed & Carnrick Food Laboratory. Scientific American, 58(14):213, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n14/pdf/scientificamerican04071888-
213.pdf.

Hopkins:1888:SEE

[10287] Geo. M. Hopkins. Some experiments with an electro-magnet. Scientific American, 58(14):214, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n14/pdf/scientificamerican04071888-
214.pdf.

Anonymous:1888:SCa

[10288] Anonymous. Spontaneous combustion. Scientific American, 58 (14):214–215, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n14/pdf/scientificamerican04071888-
214a.pdf.

Anonymous:1888:AC

[10289] Anonymous. An Australian Chlamydosaurus. Scientific American, 58(14):215–216, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n14/pdf/scientificamerican04071888-
215.pdf.

Anonymous:1888:IPI

[10290] Anonymous. An improved pruning implement. Scientific American, 58(14):217, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n14/pdf/scientificamerican04071888-
217a.pdf.

Anonymous:1888:DOT

[10291] Anonymous. A device for operating throttle valves. Scientific American, 58(14):217, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n14/pdf/scientificamerican04071888-
217b.pdf.

Anonymous:1888:AGW

[10292] Anonymous. An automatic grain weigher. Scientific American, 58(14):217, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n14/pdf/scientificamerican04071888-
217c.pdf.

Anonymous:1888:EIe

[10293] Anonymous. Engineering inventions. Scientific American, 58(14):218, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n14/pdf/scientificamerican04071888-218.pdf.

Anonymous:1888:AIf

[10294] Anonymous. Agricultural inventions. Scientific American, 58(14):218, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n14/pdf/scientificamerican04071888-218a.pdf.

Anonymous:1888:MIf

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n14/pdf/scientificamerican04071888-218b.pdf.

Anonymous:1888:Sd

[10296] Anonymous. Special. Scientific American, 58(14):218, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n14/pdf/scientificamerican04071888-218c.pdf.

Anonymous:1888:If

[10297] Anonymous. To inventors. Scientific American, 58(14):219, April 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n14/pdf/scientificamerican04071888-219a.pdf.

Anonymous:1888:Ce

[10298] Anonymous. Correspondence. Scientific American, 58(15):229, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n15/pdf/scientificamerican04141888-229.pdf.

Anonymous:1888:BPg

[10299] Anonymous. Business and personal. Scientific American, 58(15):234, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n15/pdf/scientificamerican04141888-234b.pdf.

Anonymous:1888:BNBe

[10300] Anonymous. Books: New books and publications. Scientific American, 58(15):234, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n15/pdf/scientificamerican04141888-
234c.pdf.

Anonymous:1888:NQf

[10301] Anonymous. Notes and queries. Scientific American, 58(15):234–235, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n15/pdf/scientificamerican04141888-234d.pdf.

Anonymous:1888:IIg

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n15/pdf/scientificamerican04141888-235a.pdf.

Anonymous:1888:PNPg

[10303] Anonymous. Patent notes: Patents. Scientific American, 58(15):238, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n15/pdf/scientificamerican04141888-238.pdf.

Anonymous:1888:VSR

[10304] Anonymous. A Venetian ship railway. Scientific American, 58(15):224, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n15/pdf/scientificamerican04141888-224.pdf.

Anonymous:1888:CWb

[10305] Anonymous. The celestial world. Scientific American, 58(15):224, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n15/pdf/scientificamerican04141888-224a.pdf.

Anonymous:1888:PNc

[10306] Anonymous. Photographic notes. Scientific American, 58(15):225, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n15/pdf/scientificamerican04141888-225.pdf.

Anonymous:1888:BEA

[10307] Anonymous. A bale ejecting attachment for presses. Scientific American, 58(15):226, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n15/pdf/scientificamerican04141888-
226.pdf.

Anonymous:1888:CNY

[10308] Anonymous. The City of New York. Scientific American, 58(15):226, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n15/pdf/scientificamerican04141888-226a.pdf.

Anonymous:1888:DQW

[10309] Anonymous. A device for quieting waves in storms. Scientific American, 58(15):226, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n15/pdf/scientificamerican04141888-
226b.pdf.

Anonymous:1888:IBTa

[10310] Anonymous. An improved belt tightener. Scientific American, 58(15):227, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n15/pdf/scientificamerican04141888-
227.pdf.

Anonymous:1888:IBPa

[10311] Anonymous. An improved balance piston ring packing. Scientific American, 58(15):227, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n15/pdf/scientificamerican04141888-
227a.pdf.

Anonymous:1888:CBP

[10312] Anonymous. A clamp or bridle for paint brushes. Scientific American, 58(15):227, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n15/pdf/scientificamerican04141888-
227b.pdf.

Anonymous:1888:IDA

[10313] Anonymous. An improved draft attachment for plows. Scientific American, 58(15):227, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n15/pdf/scientificamerican04141888-
227c.pdf.

Anonymous:1888:CLA

[10314] Anonymous. A combination lock and alarm. Scientific American, 58(15):227, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n15/pdf/scientificamerican04141888-
227d.pdf.

Anonymous:1888:ISSb

[10315] Anonymous. Improved spring for side bar vehicles. Scientific American, 58(15):228, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n15/pdf/scientificamerican04141888-
228.pdf.

Anonymous:1888:ITD

[10316] Anonymous. An improved tension device for looms. Scientific American, 58(15):228, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n15/pdf/scientificamerican04141888-
228a.pdf.

Anonymous:1888:ISK

[10317] Anonymous. An improved sleigh knee. Scientific American, 58(15):228, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n15/pdf/scientificamerican04141888-228b.pdf.

Anonymous:1888:IDR

[10318] Anonymous. An improved device for renovating garments. Scientific American, 58(15):228, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n15/pdf/scientificamerican04141888-
228c.pdf.

Anonymous:1888:ICCd

[10319] Anonymous. An improved car coupling. Scientific American, 58 (15):228, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n15/pdf/scientificamerican04141888-
228d.pdf.

Anonymous:1888:MBB

[10320] Anonymous. Moving the Brighton Beach Hotel. Scientific American, 58(15):230, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n15/pdf/scientificamerican04141888-
230.pdf.

Anonymous:1888:RAN

[10321] Anonymous. The Roman aqueduct at Nîmes. Scientific American, 58(15):231–232, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n15/pdf/scientificamerican04141888-
231.pdf.

Anonymous:1888:WNM

[10322] Anonymous. The Waltham non-magnetic watch. Scientific American, 58(15):232, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n15/pdf/scientificamerican04141888-
232.pdf.

Hopkins:1888:EIPa

[10323] Geo. M. Hopkins. Experiments illustrating the principle of the dynamo. Scientific American, 58(15):233, April 14, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v58/
n15/pdf/scientificamerican04141888-233.pdf.

Anonymous:1888:IAP

[10324] Anonymous. An improved air pump governor. Scientific American, 58(15):233, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n15/pdf/scientificamerican04141888-
233a.pdf.

Anonymous:1888:EIf

[10325] Anonymous. Engineering invention. Scientific American, 58(15):234, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n15/pdf/scientificamerican04141888-234.pdf.

Anonymous:1888:MIg

[10326] Anonymous. Miscellaneous inventions. Scientific American, 58(15):234, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n15/pdf/scientificamerican04141888-234a.pdf.

Anonymous:1888:Ih

[10327] Anonymous. To inventors. Scientific American, 58(15):235, April 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n15/pdf/scientificamerican04141888-235.pdf.

Anonymous:1888:Cf

[10328] Anonymous. Correspondence. Scientific American, 58(16):245–246, April 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n16/pdf/scientificamerican04211888-245.pdf.

Anonymous:1888:BPh

[10329] Anonymous. Business and personal. Scientific American, 58(16):250, April 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n16/pdf/scientificamerican04211888-250c.pdf.

Anonymous:1888:BNBf

[10330] Anonymous. Books: New books and publications. Scientific American, 58(16):250, April 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n16/pdf/scientificamerican04211888-
250d.pdf.

Anonymous:1888:NQg

[10331] Anonymous. Notes and queries. Scientific American, 58(16):250–251, April 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n16/pdf/scientificamerican04211888-250e.pdf.

Anonymous:1888:IIh

[10332] Anonymous. Index of inventions. Scientific American, 58(16):251–252, April 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n16/pdf/scientificamerican04211888-251a.pdf.

Anonymous:1888:PNPh

[10333] Anonymous. Patent notes: Patents. Scientific American, 58(16):254, April 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n16/pdf/scientificamerican04211888-254.pdf.

Anonymous:1888:CPV

[10334] Anonymous. Central Park vegetation. Scientific American, 58 (16):240–242, April 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n16/pdf/scientificamerican04211888-
240.pdf.

Anonymous:1888:TCH

[10335] Anonymous. A ticket and check holder for railroad cars. Scientific American, 58(16):242, April 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n16/pdf/scientificamerican04211888-
242.pdf.

Anonymous:1888:INYa

[10336] Anonymous. An improved neck yoke fastening. Scientific American, 58(16):242, April 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n16/pdf/scientificamerican04211888-
242a.pdf.

Anonymous:1888:IHDb

[10337] Anonymous. An improved hand drill for miners. Scientific American, 58(16):242, April 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n16/pdf/scientificamerican04211888-
242b.pdf.

Anonymous:1888:ICPa

[10338] Anonymous. Improved cap protector for cartridges. Scientific American, 58(16):243, April 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n16/pdf/scientificamerican04211888-
243.pdf.

Anonymous:1888:IFL

[10339] Anonymous. An improved fastening for lacing cords. Scientific American, 58(16):243, April 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n16/pdf/scientificamerican04211888-
243a.pdf.

Anonymous:1888:IBTb

[10340] Anonymous. An improved butter tub. Scientific American, 58(16):243, April 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n16/pdf/scientificamerican04211888-243b.pdf.

Anonymous:1888:IDF

[10341] Anonymous. An improved door fastener. Scientific American, 58(16):243, April 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n16/pdf/scientificamerican04211888-
243c.pdf.

Anonymous:1888:CTL

[10342] Anonymous. A combination tool for loading cartridges. Scientific American, 58(16):243, April 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n16/pdf/scientificamerican04211888-
243d.pdf.

Anonymous:1888:VWM

[10343] Anonymous. The Volney W. Mason hoisting machinery at the Great Hay Depot of the New York Central Railway, New York City. Scientific American, 58(16):244, April 21, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v58/
n16/pdf/scientificamerican04211888-244.pdf.

Hopkins:1888:EIPb

[10344] M. Hopkins. Experiments illustrating the principle of the dynamo. Scientific American, 58(16):244, April 21, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v58/
n16/pdf/scientificamerican04211888-244a.pdf.

Anonymous:1888:MMA

[10345] Anonymous. Modern military appliances. Scientific American, 58 (16):246–247, April 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n16/pdf/scientificamerican04211888-
246.pdf.

Anonymous:1888:SPS

[10346] Anonymous. A simple and powerful stump puller. Scientific American, 58(16):247–248, April 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n16/pdf/scientificamerican04211888-
247.pdf.

Anonymous:1888:EOE

[10347] Anonymous. Explosion of an ostrich egg. Scientific American, 58(16):249, April 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n16/pdf/scientificamerican04211888-
249.pdf.

Anonymous:1888:EIg

[10348] Anonymous. Engineering inventions. Scientific American, 58(16):250, April 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n16/pdf/scientificamerican04211888-250.pdf.

Anonymous:1888:MIh

[10349] Anonymous. Miscellaneous inventions. Scientific American, 58(16):250, April 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n16/pdf/scientificamerican04211888-250a.pdf.

Anonymous:1888:Se

[10350] Anonymous. Special. Scientific American, 58(16):250, April 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n16/pdf/scientificamerican04211888-250b.pdf.

Anonymous:1888:Ij

[10351] Anonymous. To inventors. Scientific American, 58(16):251, April 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n16/pdf/scientificamerican04211888-251.pdf.

Anonymous:1888:BNBg

[10352] Anonymous. Books: New books and publications. Scientific American, 58(17):266, April 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n17/pdf/scientificamerican04281888-
266c.pdf.

Anonymous:1888:BPi

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n17/pdf/scientificamerican04281888-266d.pdf.

Anonymous:1888:NQh

[10354] Anonymous. Notes and queries. Scientific American, 58(17):267, April 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n17/pdf/scientificamerican04281888-267.pdf.

Anonymous:1888:IIi

[10355] Anonymous. Index of inventions. Scientific American, 58(17):267–268, April 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n17/pdf/scientificamerican04281888-267b.pdf.

Anonymous:1888:PNPi

[10356] Anonymous. Patent notes: Patents. Scientific American, 58(17):270, April 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n17/pdf/scientificamerican04281888-270.pdf.

Anonymous:1888:TSa

[10357] Anonymous. Thomas Silver. Scientific American, 58(17):256, April 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n17/pdf/scientificamerican04281888-256.pdf.

Anonymous:1888:PPMa

[10358] Anonymous. Position of the planets for May. Scientific American, 58(17):256–257, April 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n17/pdf/scientificamerican04281888-
256a.pdf.

Anonymous:1888:MSA

[10359] Anonymous. The manufacture of scientific apparatus. Scientific American, 58(17):258–259, April 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n17/pdf/scientificamerican04281888-
258.pdf.

Anonymous:1888:IHa

[10360] Anonymous. An improved horseshoe. Scientific American, 58(17):259, April 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n17/pdf/scientificamerican04281888-259.pdf.

Anonymous:1888:IESa

[10361] Anonymous. An improved electrical speed indicator. Scientific American, 58(17):259, April 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n17/pdf/scientificamerican04281888-
259a.pdf.

Anonymous:1888:ISA

[10362] Anonymous. An improved spark arrester. Scientific American, 58(17):259, April 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n17/pdf/scientificamerican04281888-
259b.pdf.

Anonymous:1888:IDM

[10363] Anonymous. Improved drilling machine. Scientific American, 58 (17):259–260, April 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n17/pdf/scientificamerican04281888-
259c.pdf.

Anonymous:1888:DPG

[10364] Anonymous. A device for protecting garments. Scientific American, 58(17):260–261, April 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n17/pdf/scientificamerican04281888-
260.pdf.

Anonymous:1888:ERC

[10365] Anonymous. English railway couplings. Scientific American, 58(17):262, April 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n17/pdf/scientificamerican04281888-262.pdf.

Anonymous:1888:NDW

[10366] Anonymous. A newly discovered wild goat (Capra Dorcas). Scientific American, 58(17):263–265, April 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v58/n17/pdf/scientificamerican04281888-
263.pdf.

Anonymous:1888:SR

[10367] Anonymous. A suburban residence. Scientific American, 58(17):265, April 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n17/pdf/scientificamerican04281888-265.pdf.

Anonymous:1888:EIh

[10368] Anonymous. Engineering inventions. Scientific American, 58(17):266, April 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n17/pdf/scientificamerican04281888-266.pdf.

Anonymous:1888:AIg

[10369] Anonymous. Agricultural inventions. Scientific American, 58(17):266, April 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n17/pdf/scientificamerican04281888-266a.pdf.

Anonymous:1888:MIi

[10370] Anonymous. Miscellaneous inventions. Scientific American, 58(17):266, April 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n17/pdf/scientificamerican04281888-266b.pdf.

Anonymous:1888:Il

[10371] Anonymous. To inventors. Scientific American, 58(17):267, April 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n17/pdf/scientificamerican04281888-267a.pdf.

Anonymous:1888:Cg

[10372] Anonymous. Correspondence. Scientific American, 58(18):277, May 5, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n18/pdf/scientificamerican05051888-277a.pdf.

Anonymous:1888:BPj

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n18/pdf/scientificamerican05051888-282d.pdf.

Anonymous:1888:NQi

[10374] Anonymous. Notes and queries. Scientific American, 58(18):282–283, May 5, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n18/pdf/scientificamerican05051888-282e.pdf.

Anonymous:1888:IIj

[10375] Anonymous. Index of inventions. Scientific American, 58(18):283–284, May 5, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n18/pdf/scientificamerican05051888-283a.pdf.

Anonymous:1888:PNPj

[10376] Anonymous. Patent notes: Patents. Scientific American, 58(18):286, May 5, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n18/pdf/scientificamerican05051888-286.pdf.

Anonymous:1888:PTO

[10377] Anonymous. Patent tricks — old and new. Scientific American, 58(18):272, May 5, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n18/pdf/scientificamerican05051888-
272.pdf.

Anonymous:1888:BHP

[10378] Anonymous. The board of health and private stables. Scientific American, 58(18):272–273, May 5, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n18/pdf/scientificamerican05051888-
272a.pdf.

Anonymous:1888:NST

[10379] Anonymous. A new system of teaching geography. Scientific American, 58(18):274, May 5, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n18/pdf/scientificamerican05051888-
274a.pdf.

Anonymous:1888:CRW

[10380] Anonymous. A combined rain water cut-off and filter. Scientific American, 58(18):274, May 5, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n18/pdf/scientificamerican05051888-
274b.pdf.

Anonymous:1888:IBHa

[10381] Anonymous. An improved bag holder. Scientific American, 58(18):274, May 5, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n18/pdf/scientificamerican05051888-274c.pdf.

Anonymous:1888:TTH

[10382] Anonymous. Twelve ton hydraulic wharf crane. Scientific American, 58(18):275–276, May 5, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n18/pdf/scientificamerican05051888-
275.pdf.

Anonymous:1888:ICa

[10383] Anonymous. An improved clasp. Scientific American, 58(18):276, May 5, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n18/pdf/scientificamerican05051888-276.pdf.

Anonymous:1888:IHTa

[10384] Anonymous. Improved hand turning tools. Scientific American, 58(18):276, May 5, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n18/pdf/scientificamerican05051888-
276a.pdf.

Anonymous:1888:RSSb

[10385] Anonymous. The rotary steam snow shovel. Scientific American, 58(18):277, May 5, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n18/pdf/scientificamerican05051888-
277.pdf.

Creighton:1888:DMG

[10386] J. H. Creighton. Daniel McCartney — the great prodigy of memory. Scientific American, 58(18):278, May 5, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v58/
n18/pdf/scientificamerican05051888-278.pdf.

Anonymous:1888:SBP

[10387] Anonymous. The superb bird of paradise. Scientific American, 58(18):278, May 5, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n18/pdf/scientificamerican05051888-
278a.pdf.

Anonymous:1888:GCCa

[10388] Anonymous. The German corvette cruises Irene. Scientific American, 58(18):279–280, May 5, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n18/pdf/scientificamerican05051888-
279.pdf.

Anonymous:1888:IPR

[10389] Anonymous. An improved passenger railway car. Scientific American, 58(18):281, May 5, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n18/pdf/scientificamerican05051888-
281.pdf.

Sloane:1888:EIP

[10390] T. O’Conor Sloane. Experiments illustrating physical and chemical change. Scientific American, 58(18):281, May 5, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v58/
n18/pdf/scientificamerican05051888-281a.pdf.

Anonymous:1888:ICHa

[10391] Anonymous. An improved car-heater. Scientific American, 58(18):281, May 5, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n18/pdf/scientificamerican05051888-281b.pdf.

Anonymous:1888:EIi

[10392] Anonymous. Engineering invention. Scientific American, 58(18):282, May 5, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n18/pdf/scientificamerican05051888-282.pdf.

Anonymous:1888:AIh

[10393] Anonymous. Agricultural inventions. Scientific American, 58(18):282, May 5, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n18/pdf/scientificamerican05051888-282a.pdf.

Anonymous:1888:MIj

[10394] Anonymous. Miscellaneous inventions. Scientific American, 58(18):282, May 5, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n18/pdf/scientificamerican05051888-282b.pdf.

Anonymous:1888:Sf

[10395] Anonymous. Special. Scientific American, 58(18):282, May 5, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n18/pdf/scientificamerican05051888-282c.pdf.

Anonymous:1888:Im

[10396] Anonymous. To inventors. Scientific American, 58(18):283, May 5, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n18/pdf/scientificamerican05051888-283.pdf.

Anonymous:1888:CDC

[10397] Anonymous. The caligraph drop cabinet desk. Scientific American, 58(18):374–274, May 5, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n18/pdf/scientificamerican05051888-
274.pdf.

Anonymous:1888:Ch

[10398] Anonymous. Correspondence. Scientific American, 58(19):293–294, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n19/pdf/scientificamerican05121888-293.pdf.

Anonymous:1888:BPk

[10399] Anonymous. Business and personal. Scientific American, 58(19):298, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n19/pdf/scientificamerican05121888-298c.pdf.

Anonymous:1888:BNBh

[10400] Anonymous. Books: New books and publications. Scientific American, 58(19):298, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n19/pdf/scientificamerican05121888-
298d.pdf.

Anonymous:1888:NQj

[10401] Anonymous. Notes and queries. Scientific American, 58(19):299, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n19/pdf/scientificamerican05121888-299.pdf.

Anonymous:1888:IIk

[10402] Anonymous. Index of inventions. Scientific American, 58(19):299–300, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n19/pdf/scientificamerican05121888-299b.pdf.

Anonymous:1888:PNPk

[10403] Anonymous. Patent notes: Patents. Scientific American, 58(19):302, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n19/pdf/scientificamerican05121888-302.pdf.

Anonymous:1888:LYT

[10404] Anonymous. Launch of the York-Town and Vesuvius. Scientific American, 58(19):287, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n19/pdf/scientificamerican05121888-
287.pdf.

Anonymous:1888:PLS

[10405] Anonymous. A Patent Law for Switzerland. Scientific American, 58(19):288, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n19/pdf/scientificamerican05121888-
288.pdf.

Anonymous:1888:SMG

[10406] Anonymous. Successful moving of a great hotel. Scientific American, 58(19):288, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n19/pdf/scientificamerican05121888-
288a.pdf.

Anonymous:1888:PRP

[10407] Anonymous. Proposed reorganization of public works. Scientific American, 58(19):288, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n19/pdf/scientificamerican05121888-
288b.pdf.

Anonymous:1888:CWc

[10408] Anonymous. The celestial world. Scientific American, 58(19):289, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n19/pdf/scientificamerican05121888-289.pdf.

Anonymous:1888:IRE

[10409] Anonymous. An improved rotary engine. Scientific American, 58 (19):290, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n19/pdf/scientificamerican05121888-
290.pdf.

Anonymous:1888:INLb

[10410] Anonymous. An improved nut lock. Scientific American, 58(19):290, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n19/pdf/scientificamerican05121888-290a.pdf.

Anonymous:1888:ISTa

[10411] Anonymous. An improved stock tether. Scientific American, 58(19):290, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n19/pdf/scientificamerican05121888-290b.pdf.

Anonymous:1888:IBHb

[10412] Anonymous. An improved bag holder. Scientific American, 58(19):290, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n19/pdf/scientificamerican05121888-290c.pdf.

Anonymous:1888:ISBa

[10413] Anonymous. An improved suspender buckle. Scientific American, 58(19):290, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n19/pdf/scientificamerican05121888-
290d.pdf.

Anonymous:1888:IPBa

[10414] Anonymous. An improved paddle-belt. Scientific American, 58(19):291, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n19/pdf/scientificamerican05121888-291.pdf.

Anonymous:1888:IGb

[10415] Anonymous. An improved gate. Scientific American, 58(19):291, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n19/pdf/scientificamerican05121888-291a.pdf.

Anonymous:1888:IAE

[10416] Anonymous. An improved apparatus for examining ores. Scientific American, 58(19):291, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n19/pdf/scientificamerican05121888-
291b.pdf.

Anonymous:1888:CTC

[10417] Anonymous. A combined tie, chair, and fish plate. Scientific American, 58(19):291, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n19/pdf/scientificamerican05121888-
291c.pdf.

Anonymous:1888:IFJ

[10418] Anonymous. An improved fruit jar. Scientific American, 58 (19):291–292, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n19/pdf/scientificamerican05121888-
291d.pdf.

Anonymous:1888:ISIa

[10419] Anonymous. An improved self-inking printing press. Scientific American, 58(19):292, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n19/pdf/scientificamerican05121888-
292.pdf.

Anonymous:1888:IGc

[10420] Anonymous. An improved gate. Scientific American, 58(19):292, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n19/pdf/scientificamerican05121888-292a.pdf.

Anonymous:1888:IPSa

[10421] Anonymous. Improved plate straightening machine. Scientific American, 58(19):294, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n19/pdf/scientificamerican05121888-
294.pdf.

Sloane:1888:DWA

[10422] T. O’Conor Sloane. The decomposition of water and absorption of carbon dioxide by caustic soda. Scientific American, 58(19):295, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n19/pdf/scientificamerican05121888-295.pdf.

Anonymous:1888:VSC

[10423] Anonymous. Virot’s steam carriage. Scientific American, 58 (19):295–296, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n19/pdf/scientificamerican05121888-
295a.pdf.

Anonymous:1888:SMS

[10424] Anonymous. Sectional model of steam engine. Scientific American, 58(19):297, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n19/pdf/scientificamerican05121888-
297.pdf.

Anonymous:1888:PBF

[10425] Anonymous. Preparation of butter-flies for the cabinet. Scientific American, 58(19):297, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n19/pdf/scientificamerican05121888-
297a.pdf.

Anonymous:1888:EIj

[10426] Anonymous. Engineering inventions. Scientific American, 58(19):298, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n19/pdf/scientificamerican05121888-298.pdf.

Anonymous:1888:AIi

[10427] Anonymous. Agricultural invention. Scientific American, 58(19):298, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n19/pdf/scientificamerican05121888-298a.pdf.

Anonymous:1888:MIk

[10428] Anonymous. Miscellaneous inventions. Scientific American, 58(19):298, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n19/pdf/scientificamerican05121888-298b.pdf.

Anonymous:1888:In

[10429] Anonymous. To inventors. Scientific American, 58(19):299, May 12, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n19/pdf/scientificamerican05121888-299a.pdf.

Anonymous:1888:BPl

[10430] Anonymous. Business and personal. Scientific American, 58 (2):26–27, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n2/pdf/scientificamerican01141888-
26d.pdf.

Anonymous:1888:NQk

[10431] Anonymous. Notes and queries. Scientific American, 58(2):27, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n2/pdf/scientificamerican01141888-27.pdf.

Anonymous:1888:BNBi 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n2/pdf/scientificamerican01141888-
27a.pdf.

Anonymous:1888:IIl

[10433] Anonymous. Index of inventions. Scientific American, 58(2):27–28, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n2/pdf/scientificamerican01141888-27c.pdf.

Anonymous:1888:PNPl

[10434] Anonymous. Patent notes: Patents. Scientific American, 58(2):30, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n2/pdf/scientificamerican01141888-30.pdf.

Anonymous:1888:MKG

[10435] Anonymous. Major King’s great magnet. Scientific American, 58 (2):15–16, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n2/pdf/scientificamerican01141888-
15.pdf.

Anonymous:1888:BWP

[10436] Anonymous. The Barbed Wire Patent declared invalid. Scientific American, 58(2):16, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n2/pdf/scientificamerican01141888-
16.pdf.

Anonymous:1888:DGL

[10437] Anonymous. The destruction of the great lumber raft. Scientific American, 58(2):16–17, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n2/pdf/scientificamerican01141888-
16a.pdf.

Anonymous:1888:DPS

[10438] Anonymous. Discovery of platinum in the Sun’s atmosphere. Scientific American, 58(2):17, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n2/pdf/scientificamerican01141888-
17.pdf.

Anonymous:1888:IBCa

[10439] Anonymous. An improved bell cord coupling. Scientific American, 58(2):18, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n2/pdf/scientificamerican01141888-
18.pdf.

Anonymous:1888:IFGa

[10440] Anonymous. An improved furnace grate. Scientific American, 58(2):18, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n2/pdf/scientificamerican01141888-
18a.pdf.

Anonymous:1888:SWM

[10441] Anonymous. The siphon water meter. Scientific American, 58 (2):18, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n2/pdf/scientificamerican01141888-
18b.pdf.

Anonymous:1888:IFa

[10442] Anonymous. An improved ferrule. Scientific American, 58(2):18, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n2/pdf/scientificamerican01141888-18c.pdf.

Anonymous:1888:IMBa

[10443] Anonymous. An improved method of bookbinding. Scientific American, 58(2):18, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n2/pdf/scientificamerican01141888-
18d.pdf.

Anonymous:1888:IAMb

[10444] Anonymous. Improved arrangement of movable shelves. Scientific American, 58(2):19, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n2/pdf/scientificamerican01141888-
19.pdf.

Anonymous:1888:IGP

[10445] Anonymous. An improved gas pipe wrench. Scientific American, 58(2):19, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n2/pdf/scientificamerican01141888-
19a.pdf.

Anonymous:1888:IFRa

[10446] Anonymous. An improved farm and road wagon. Scientific American, 58(2):19, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n2/pdf/scientificamerican01141888-
19b.pdf.

Anonymous:1888:TFT

[10447] Anonymous. Twenty-five ton crane. Scientific American, 58(2): 19–21, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n2/pdf/scientificamerican01141888-
19c.pdf.

Anonymous:1888:SI

[10448] Anonymous. Snakes in India. Scientific American, 58(2):22, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n2/pdf/scientificamerican01141888-22.pdf.

Anonymous:1888:IRLa

[10449] Anonymous. Improved rail loading machine. Scientific American, 58(2):23, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n2/pdf/scientificamerican01141888-
23.pdf.

Anonymous:1888:NP

[10450] Anonymous. Novel puzzle. Scientific American, 58(2):23, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n2/pdf/scientificamerican01141888-23a.pdf.

Sloane:1888:AEW

[10451] T. O’Conor Sloane. Apparatus for the electrolysis of water. Scientific American, 58(2):23–24, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v58/n2/pdf/scientificamerican01141888-
23b.pdf.

Anonymous:1888:GGW

[10452] Anonymous. Group of gas wells at Fort Scott, Kansas. Scientific American, 58(2):25, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n2/pdf/scientificamerican01141888-
25.pdf.

Anonymous:1888:EIk

[10453] Anonymous. Engineering inventions. Scientific American, 58(2): 26, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n2/pdf/scientificamerican01141888-
26.pdf.

Anonymous:1888:AIj

[10454] Anonymous. Agricultural invention. Scientific American, 58(2):26, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n2/pdf/scientificamerican01141888-26a.pdf.

Anonymous:1888:MIl

[10455] Anonymous. Miscellaneous inventions. Scientific American, 58 (2):26, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n2/pdf/scientificamerican01141888-
26b.pdf.

Anonymous:1888:Sg

[10456] Anonymous. Special. Scientific American, 58(2):26, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n2/pdf/scientificamerican01141888-26c.pdf.

Anonymous:1888:Io

[10457] Anonymous. To inventors. Scientific American, 58(2):27, January 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n2/pdf/scientificamerican01141888-27b.pdf.

Anonymous:1888:BPm

[10458] Anonymous. Business and personal. Scientific American, 58(20):314, May 19, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n20/pdf/scientificamerican05191888-314c.pdf.

Anonymous:1888:BNBj

[10459] Anonymous. Books: New books and publications. Scientific American, 58(20):314, May 19, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n20/pdf/scientificamerican05191888-
314d.pdf.

Anonymous:1888:NQl

[10460] Anonymous. Notes and queries. Scientific American, 58(20):314–315, May 19, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n20/pdf/scientificamerican05191888-314e.pdf.

Anonymous:1888:IIm

[10461] Anonymous. Index of inventions. Scientific American, 58(20):315–316, May 19, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n20/pdf/scientificamerican05191888-315a.pdf.

Anonymous:1888:PNPm

[10462] Anonymous. Patent notes: Patents. Scientific American, 58(20):318, May 19, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n20/pdf/scientificamerican05191888-318.pdf.

Anonymous:1888:NTH

[10463] Anonymous. Novel type of high span bridge. Scientific American, 58(20):303, May 19, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n20/pdf/scientificamerican05191888-
303.pdf.

Anonymous:1888:NCB

[10464] Anonymous. The new Copyright Bill now before Congress. Scientific American, 58(20):304, May 19, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n20/pdf/scientificamerican05191888-
304.pdf.

Anonymous:1888:PCB

[10465] Anonymous. Passage of the Copyright Bill in the Senate. Scientific American, 58(20):304–305, May 19, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v58/n20/pdf/scientificamerican05191888-
304a.pdf.

Anonymous:1888:CFa

[10466] Anonymous. City flowers. Scientific American, 58(20):305, May 19, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n20/pdf/scientificamerican05191888-305.pdf.

Anonymous:1888:BAE

[10467] Anonymous. The British African Expedition. Scientific American, 58(20):306, May 19, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n20/pdf/scientificamerican05191888-
306.pdf.

Anonymous:1888:CIB

[10468] Anonymous. The Crosby invalid bed. Scientific American, 58(20):307, May 19, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n20/pdf/scientificamerican05191888-307.pdf.

Anonymous:1888:IHAa

[10469] Anonymous. An improved hot air register. Scientific American, 58(20):307, May 19, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n20/pdf/scientificamerican05191888-
307a.pdf.

Anonymous:1888:ICCe

[10470] Anonymous. An improved cloth cutter. Scientific American, 58(20):307, May 19, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n20/pdf/scientificamerican05191888-307b.pdf.

Anonymous:1888:ICSa

[10471] Anonymous. An improved counter scale. Scientific American, 58 (20):307, May 19, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n20/pdf/scientificamerican05191888-
307c.pdf.

Anonymous:1888:ICSb

[10472] Anonymous. An improved circular saw holder. Scientific American, 58(20):307–308, May 19, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n20/pdf/scientificamerican05191888-
307d.pdf.

Emmons:1888:ESE

[10473] Elmer E. E. Emmons. Experiments in static electricity with the incandescent lamp. Scientific American, 58(20):308–309, May 19, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n20/pdf/scientificamerican05191888-308.pdf.

Anonymous:1888:WES

[10474] Anonymous. Waddington’s electric submarine torpedo boat. Scientific American, 58(20):310–311, May 19, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v58/n20/pdf/scientificamerican05191888-
310.pdf.

Anonymous:1888:TSb

[10475] Anonymous. Thomas Silver. Scientific American, 58(20):311–312, May 19, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n20/pdf/scientificamerican05191888-311.pdf.

Anonymous:1888:EVY

[10476] Anonymous. The engines of the Vesuvius and Yorktown. Scientific American, 58(20):313, May 19, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n20/pdf/scientificamerican05191888-
313.pdf.

Anonymous:1888:EIl

[10477] Anonymous. Engineering invention. Scientific American, 58(20):314, May 19, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n20/pdf/scientificamerican05191888-314.pdf.

Anonymous:1888:AIk

[10478] Anonymous. Agricultural inventions. Scientific American, 58(20):314, May 19, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n20/pdf/scientificamerican05191888-314a.pdf.

Anonymous:1888:MIm

[10479] Anonymous. Miscellaneous inventions. Scientific American, 58(20):314, May 19, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n20/pdf/scientificamerican05191888-314b.pdf.

Anonymous:1888:Ip

[10480] Anonymous. To inventors. Scientific American, 58(20):315, May 19, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n20/pdf/scientificamerican05191888-315.pdf.

Anonymous:1888:Ci

[10481] Anonymous. Correspondence. Scientific American, 58(21):325–326, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n21/pdf/scientificamerican05261888-325.pdf.

Anonymous:1888:BPn

[10482] Anonymous. Business and personal. Scientific American, 58(21):331, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n21/pdf/scientificamerican05261888-331d.pdf.

Anonymous:1888:IIn

[10483] Anonymous. Index of inventions. Scientific American, 58(21):331–332, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n21/pdf/scientificamerican05261888-331e.pdf.

Anonymous:1888:PNPn

[10484] Anonymous. Patent notes: Patents. Scientific American, 58(21):334, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n21/pdf/scientificamerican05261888-334.pdf.

Anonymous:1888:Ir

[10485] Anonymous. The Italia. Scientific American, 58(21):319, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n21/pdf/scientificamerican05261888-319.pdf.

Anonymous:1888:IDEb

[10486] Anonymous. Important decision in the Edison Electric Light Suits. Scientific American, 58(21):320, May 26, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v58/
n21/pdf/scientificamerican05261888-320.pdf.

Anonymous:1888:UPL

[10487] Anonymous. The underground pipes of large cities. Scientific American, 58(21):320, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n21/pdf/scientificamerican05261888-
320a.pdf.

Anonymous:1888:CSB

[10488] Anonymous. The Carnegie Savings Bank. Scientific American, 58(21):320, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n21/pdf/scientificamerican05261888-
320b.pdf.

Anonymous:1888:ELF

[10489] Anonymous. Electric light fires. Scientific American, 58(21):320, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n21/pdf/scientificamerican05261888-320c.pdf.

Anonymous:1888:PP

[10490] Anonymous. Progress of the phonograph. Scientific American, 58(21):320–321, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n21/pdf/scientificamerican05261888-
320d.pdf.

Anonymous:1888:MNa

[10491] Anonymous. Military notes. Scientific American, 58(21):321, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n21/pdf/scientificamerican05261888-321.pdf.

Anonymous:1888:PNd

[10492] Anonymous. Photographic notes. Scientific American, 58(21):321, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n21/pdf/scientificamerican05261888-321a.pdf.

Anonymous:1888:ICG

[10493] Anonymous. An improved cable grip. Scientific American, 58(21):322, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n21/pdf/scientificamerican05261888-322.pdf.

Anonymous:1888:CLB

[10494] Anonymous. Collision of locomotives — Beatrice, Nebraska. Scientific American, 58(21):322, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n21/pdf/scientificamerican05261888-
322a.pdf.

Anonymous:1888:ICCf

[10495] Anonymous. An improved car coupling. Scientific American, 58 (21):322, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n21/pdf/scientificamerican05261888-
322b.pdf.

Anonymous:1888:IRSb

[10496] Anonymous. An improved railroad switch. Scientific American, 58(21):322, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n21/pdf/scientificamerican05261888-
322c.pdf.

Anonymous:1888:IFMa

[10497] Anonymous. Improved feed mechanism for sawing machines. Scientific American, 58(21):323, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n21/pdf/scientificamerican05261888-
323.pdf.

Anonymous:1888:IGd

[10498] Anonymous. An improved gate. Scientific American, 58(21):323, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n21/pdf/scientificamerican05261888-323a.pdf.

Anonymous:1888:IMCa

[10499] Anonymous. An improved music chart. Scientific American, 58(21):323, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n21/pdf/scientificamerican05261888-323b.pdf.

Anonymous:1888:ICRa

[10500] Anonymous. Improved check-row attachment for planters. Scientific American, 58(21):323, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n21/pdf/scientificamerican05261888-
323c.pdf.

Anonymous:1888:IWM

[10501] Anonymous. An improved windmill motor. Scientific American, 58(21):323–324, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n21/pdf/scientificamerican05261888-
323d.pdf.

Anonymous:1888:ILT

[10502] Anonymous. An improved lasting tool. Scientific American, 58(21):324, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n21/pdf/scientificamerican05261888-324.pdf.

Anonymous:1888:IUE

[10503] Anonymous. Improved utensil for extracting grape seeds. Scientific American, 58(21):324, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n21/pdf/scientificamerican05261888-
324a.pdf.

Anonymous:1888:ERH

[10504] Anonymous. Endless rope hauling machine. Scientific American, 58(21):326, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n21/pdf/scientificamerican05261888-
326.pdf.

Anonymous:1888:IMCb

[10505] Anonymous. Indian magic — the celebrated mango tree trick. Scientific American, 58(21):327, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n21/pdf/scientificamerican05261888-
327.pdf.

Anonymous:1888:PFD

[10506] Anonymous. Prize field dogs. Scientific American, 58(21):327–328, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n21/pdf/scientificamerican05261888-327a.pdf.

Anonymous:1888:ISGa

[10507] Anonymous. An improved steering gear. Scientific American, 58 (21):328, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n21/pdf/scientificamerican05261888-
328.pdf.

Anonymous:1888:ITSa

[10508] Anonymous. An improved toilet stand. Scientific American, 58(21):329, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n21/pdf/scientificamerican05261888-329.pdf.

Anonymous:1888:SMB

[10509] Anonymous. Spontaneous motions of bodies on the surface of liquids. Scientific American, 58(21):329–330, May 26, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v58/
n21/pdf/scientificamerican05261888-329a.pdf.

Anonymous:1888:EIm

[10510] Anonymous. Engineering inventions. Scientific American, 58(21):331, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n21/pdf/scientificamerican05261888-331.pdf.

Anonymous:1888:AIl

[10511] Anonymous. Agricultural invention. Scientific American, 58(21):331, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n21/pdf/scientificamerican05261888-331a.pdf.

Anonymous:1888:MIn

[10512] Anonymous. Miscellaneous inventions. Scientific American, 58(21):331, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n21/pdf/scientificamerican05261888-331b.pdf.

Anonymous:1888:Iu

[10513] Anonymous. To inventors. Scientific American, 58(21):331, May 26, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n21/pdf/scientificamerican05261888-331c.pdf.

Anonymous:1888:BPo

[10514] Anonymous. Business and personal. Scientific American, 58(22):346, June 2, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n22/pdf/scientificamerican06021888-346b.pdf.

Anonymous:1888:BNBk

[10515] Anonymous. Books: New books and publications. Scientific American, 58(22):346, June 2, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n22/pdf/scientificamerican06021888-
346c.pdf.

Anonymous:1888:NQm

[10516] Anonymous. Notes and queries. Scientific American, 58(22):346–347, June 2, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n22/pdf/scientificamerican06021888-346d.pdf.

Anonymous:1888:IIo

[10517] Anonymous. Index of inventions. Scientific American, 58(22):347–348, June 2, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n22/pdf/scientificamerican06021888-347a.pdf.

Anonymous:1888:PNPo

[10518] Anonymous. Patent notes: Patents. Scientific American, 58(22):350, June 2, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n22/pdf/scientificamerican06021888-350.pdf.

Anonymous:1888:PPJa

[10519] Anonymous. Position of the planets in June. Scientific American, 58(22):336–338, June 2, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n22/pdf/scientificamerican06021888-
336.pdf.

Anonymous:1888:RMP

[10520] Anonymous. Russian Mountains at Paris. Scientific American, 58(22):338–339, June 2, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n22/pdf/scientificamerican06021888-
338.pdf.

Anonymous:1888:IBSa

[10521] Anonymous. An improved block signal. Scientific American, 58(22):339, June 2, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n22/pdf/scientificamerican06021888-339.pdf.

Anonymous:1888:ICb

[10522] Anonymous. An improved clamp. Scientific American, 58(22):339, June 2, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n22/pdf/scientificamerican06021888-339a.pdf.

Anonymous:1888:ICCg

[10523] Anonymous. An improved car coupling. Scientific American, 58 (22):339, June 2, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n22/pdf/scientificamerican06021888-
339b.pdf.

Anonymous:1888:IPSb

[10524] Anonymous. An improved powder and shot cabinet. Scientific American, 58(22):339–340, June 2, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n22/pdf/scientificamerican06021888-
339c.pdf.

Anonymous:1888:IRR

[10525] Anonymous. An improved railway rail pad. Scientific American, 58(22):340, June 2, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n22/pdf/scientificamerican06021888-
340.pdf.

Anonymous:1888:ILP

[10526] Anonymous. An improved life-preserving chair. Scientific American, 58(22):340, June 2, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n22/pdf/scientificamerican06021888-
340a.pdf.

Anonymous:1888:IHb

[10527] Anonymous. An improved hinge. Scientific American, 58(22):340, June 2, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n22/pdf/scientificamerican06021888-340b.pdf.

Anonymous:1888:IRTa

[10528] Anonymous. Improved receiving-table for printing presses. Scientific American, 58(22):340, June 2, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n22/pdf/scientificamerican06021888-
340c.pdf.

Anonymous:1888:LMC

[10529] Anonymous. The Lidgerwood Manufacturing Co.’s hoisting engines and boilers — the Gorton heater. Scientific American, 58(22):341–342, June 2, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n22/pdf/scientificamerican06021888-341.pdf.

Anonymous:1888:HPC

[10530] Anonymous. 2,500 h.p. Corliss engine. Scientific American, 58(22):342, June 2, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n22/pdf/scientificamerican06021888-342.pdf.

Anonymous:1888:GHC

[10531] Anonymous. George H. Corliss. Scientific American, 58(22):343–344, June 2, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n22/pdf/scientificamerican06021888-343.pdf.

Sloane:1888:ECA

[10532] T. O’conor Sloane. Experiment in capillary attraction. Scientific American, 58(22):345, June 2, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n22/pdf/scientificamerican06021888-
345.pdf.

Anonymous:1888:IBa

[10533] Anonymous. An improved bolt. Scientific American, 58(22):345, June 2, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n22/pdf/scientificamerican06021888-345a.pdf.

Anonymous:1888:INLc

[10534] Anonymous. An improved night lamp. Scientific American, 58(22):345, June 2, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n22/pdf/scientificamerican06021888-345b.pdf.

Anonymous:1888:EIn

[10535] Anonymous. Engineering inventions. Scientific American, 58(22):346, June 2, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n22/pdf/scientificamerican06021888-346.pdf.

Anonymous:1888:MIo

[10536] Anonymous. Miscellaneous intentions. Scientific American, 58(22):346, June 2, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n22/pdf/scientificamerican06021888-346a.pdf.

Anonymous:1888:Iv

[10537] Anonymous. To inventors. Scientific American, 58(22):347, June 2, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n22/pdf/scientificamerican06021888-347.pdf.

Anonymous:1888:Cj

[10538] Anonymous. Correspondence. Scientific American, 58(23):357, June 9, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n23/pdf/scientificamerican06091888-357a.pdf.

Anonymous:1888:BPp

[10539] Anonymous. Business and personal. Scientific American, 58 (23):362–363, June 9, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n23/pdf/scientificamerican06091888-
362b.pdf.

Anonymous:1888:BNBl

[10540] Anonymous. Books: New books and publications. Scientific American, 58(23):363, June 9, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n23/pdf/scientificamerican06091888-
363.pdf.

Anonymous:1888:NQn

[10541] Anonymous. Notes and queries. Scientific American, 58(23):363, June 9, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n23/pdf/scientificamerican06091888-363a.pdf.

Anonymous:1888:IIp

[10542] Anonymous. Index of inventions. Scientific American, 58(23):363–364, June 9, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n23/pdf/scientificamerican06091888-363c.pdf.

Anonymous:1888:PNPp

[10543] Anonymous. Patent notes: Patents. Scientific American, 58(23):365, June 9, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n23/pdf/scientificamerican06091888-365.pdf.

Anonymous:1888:CSc

[10544] Anonymous. The cod have “struck in”. Scientific American, 58 (23):352–354, June 9, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n23/pdf/scientificamerican06091888-
352.pdf.

Anonymous:1888:AWP

[10545] Anonymous. The artesian well in the Place Hébert, Paris. Scientific American, 58(23):354–355, June 9, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v58/n23/pdf/scientificamerican06091888-
354.pdf.

Anonymous:1888:ISGb

[10546] Anonymous. An improved saw gummer. Scientific American, 58 (23):355, June 9, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n23/pdf/scientificamerican06091888-
355.pdf.

Anonymous:1888:ICL

[10547] Anonymous. An improved cartridge loader. Scientific American, 58(23):355–356, June 9, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n23/pdf/scientificamerican06091888-
355a.pdf.

Anonymous:1888:ICDa

[10548] Anonymous. An improved canopy device for chairs. Scientific American, 58(23):356, June 9, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n23/pdf/scientificamerican06091888-
356.pdf.

Anonymous:1888:IOO

[10549] Anonymous. An improved oar and oar lock. Scientific American, 58(23):356, June 9, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n23/pdf/scientificamerican06091888-
356a.pdf.

Anonymous:1888:INLd

[10550] Anonymous. An improved nut lock. Scientific American, 58(23):356, June 9, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n23/pdf/scientificamerican06091888-356b.pdf.

Anonymous:1888:IACa

[10551] Anonymous. An improved adjustable chair and swing. Scientific American, 58(23):356, June 9, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n23/pdf/scientificamerican06091888-
356c.pdf.

Anonymous:1888:UAB

[10552] Anonymous. The U.S. armored battle ship Texas. Scientific American, 58(23):357, June 9, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n23/pdf/scientificamerican06091888-
357.pdf.

Anonymous:1888:REL

[10553] Anonymous. A remarkable escape from lightning. Scientific American, 58(23):358, June 9, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n23/pdf/scientificamerican06091888-
358.pdf.

Anonymous:1888:JCB

[10554] Anonymous. James curtis booth. Scientific American, 58(23):359, June 9, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n23/pdf/scientificamerican06091888-359.pdf.

Anonymous:1888:PMB

[10555] Anonymous. The Pearl mucilage bottle. Scientific American, 58 (23):360–361, June 9, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n23/pdf/scientificamerican06091888-
360.pdf.

Anonymous:1888:RRR

[10556] Anonymous. Restoration of Roman ruin at Reims, France. Scientific American, 58(23):361, June 9, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n23/pdf/scientificamerican06091888-
361.pdf.

Anonymous:1888:EIo

[10557] Anonymous. Engineering invention. Scientific American, 58(23):362, June 9, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n23/pdf/scientificamerican06091888-362.pdf.

Anonymous:1888:MIp

[10558] Anonymous. Miscellaneous inventions. Scientific American, 58(23):362, June 9, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n23/pdf/scientificamerican06091888-362a.pdf.

Anonymous:1888:Iw

[10559] Anonymous. To inventors. Scientific American, 58(23):363, June 9, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n23/pdf/scientificamerican06091888-363b.pdf.

Anonymous:1888:BPq

[10560] Anonymous. Business and personal. Scientific American, 58(24):378, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n24/pdf/scientificamerican06161888-378c.pdf.

Anonymous:1888:BNBm

[10561] Anonymous. Books: New books and publications. Scientific American, 58(24):378, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n24/pdf/scientificamerican06161888-
378d.pdf.

Anonymous:1888:NQo

[10562] Anonymous. Notes and queries. Scientific American, 58(24):378–379, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n24/pdf/scientificamerican06161888-378e.pdf.

Anonymous:1888:IIq

[10563] Anonymous. Index of inventions. Scientific American, 58(24):379–380, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n24/pdf/scientificamerican06161888-379a.pdf.

Anonymous:1888:PNPq

[10564] Anonymous. Patent notes: Patents. Scientific American, 58(24):382, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n24/pdf/scientificamerican06161888-382.pdf.

Anonymous:1888:WGM

[10565] Anonymous. Water gas in Massachusetts, devastation by crickets, and more. Scientific American, 58(24):368–369, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n24/pdf/scientificamerican06161888-368.pdf.

Anonymous:1888:ICCh

[10566] Anonymous. An improved chimney cowl and ventilator. Scientific American, 58(24):370, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n24/pdf/scientificamerican06161888-
370.pdf.

Anonymous:1888:ISGc

[10567] Anonymous. An improved shingling gauge. Scientific American, 58(24):370, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n24/pdf/scientificamerican06161888-
370a.pdf.

Anonymous:1888:DLS

[10568] Anonymous. A device for locking steering gear. Scientific American, 58(24):370, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n24/pdf/scientificamerican06161888-
370b.pdf.

Anonymous:1888:BPBa

[10569] Anonymous. Binns’ patent banding spindles. Scientific American, 58(24):370, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n24/pdf/scientificamerican06161888-
370c.pdf.

Anonymous:1888:IFF

[10570] Anonymous. An improved fence and fence post. Scientific American, 58(24):370, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n24/pdf/scientificamerican06161888-
370d.pdf.

Anonymous:1888:IRSc

[10571] Anonymous. Improved railway switch stand and signal. Scientific American, 58(24):370–371, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n24/pdf/scientificamerican06161888-
370e.pdf.

Anonymous:1888:ITHa

[10572] Anonymous. An improved twine holder. Scientific American, 58 (24):372, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n24/pdf/scientificamerican06161888-
372.pdf.

Anonymous:1888:IFEb

[10573] Anonymous. An improved fire escape. Scientific American, 58(24):372, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n24/pdf/scientificamerican06161888-372a.pdf.

Anonymous:1888:WMH

[10574] Anonymous. The White Mountain hammock chair. Scientific American, 58(24):372, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n24/pdf/scientificamerican06161888-
372b.pdf.

Anonymous:1888:IHEb

[10575] Anonymous. An improved hoof expander. Scientific American, 58(24):372, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n24/pdf/scientificamerican06161888-
372c.pdf.

Anonymous:1888:IWV

[10576] Anonymous. An improved window ventilator. Scientific American, 58(24):372, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n24/pdf/scientificamerican06161888-
372d.pdf.

Keeler:1888:TSI

[10577] James E. Keeler. The thirty-six inch equatorial telescope of the Lick Observatory. Scientific American, 58(24):373–374, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n24/pdf/scientificamerican06161888-373.pdf.

Anonymous:1888:SLO

[10578] Anonymous. Spectroscope for the Lick Observatory. Scientific American, 58(24):374, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n24/pdf/scientificamerican06161888-
374.pdf.

Anonymous:1888:DA

[10579] Anonymous. A dwarf armadillo. Scientific American, 58(24):375–376, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n24/pdf/scientificamerican06161888-375.pdf.

Anonymous:1888:SWT

[10580] Anonymous. Spirally welded tubing. Scientific American, 58(24):377, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n24/pdf/scientificamerican06161888-377.pdf.

Anonymous:1888:EIp

[10581] Anonymous. Engineering invention. Scientific American, 58(24):378, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n24/pdf/scientificamerican06161888-378.pdf.

Anonymous:1888:AIm

[10582] Anonymous. Agricultural invention. Scientific American, 58(24):378, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n24/pdf/scientificamerican06161888-378a.pdf.

Anonymous:1888:MIq

[10583] Anonymous. Miscellaneous inventions. Scientific American, 58(24):378, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n24/pdf/scientificamerican06161888-378b.pdf.

Anonymous:1888:Ix

[10584] Anonymous. To inventors. Scientific American, 58(24):379, June 16, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n24/pdf/scientificamerican06161888-379.pdf.

Anonymous:1888:BPr

[10585] Anonymous. Business and personal. Scientific American, 58(25):394, June 23, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n25/pdf/scientificamerican06231888-394c.pdf.

Anonymous:1888:BNBn

[10586] Anonymous. Books: New books and publications. Scientific American, 58(25):395, June 23, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n25/pdf/scientificamerican06231888-
395.pdf.

Anonymous:1888:NQp

[10587] Anonymous. Notes and queries. Scientific American, 58(25):395, June 23, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n25/pdf/scientificamerican06231888-395a.pdf.

Anonymous:1888:IIr

[10588] Anonymous. Index of inventions. Scientific American, 58(25):395–396, June 23, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n25/pdf/scientificamerican06231888-395c.pdf.

Anonymous:1888:PNPr

[10589] Anonymous. Patent notes: Patents. Scientific American, 58(25):398, June 23, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n25/pdf/scientificamerican06231888-398.pdf.

Anonymous:1888:IOT

[10590] Anonymous. Inside and outside of a torpedo vessel — H.M.S. Rattlesnake. Scientific American, 58(25):383, June 23, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v58/
n25/pdf/scientificamerican06231888-383.pdf.

Anonymous:1888:EG

[10591] Anonymous. The Emperor of Germany. Scientific American, 58 (25):384, June 23, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n25/pdf/scientificamerican06231888-
384.pdf.

Anonymous:1888:ICAa

[10592] Anonymous. The International Congress of Anthropology. Scientific American, 58(25):384, June 23, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n25/pdf/scientificamerican06231888-
384a.pdf.

Anonymous:1888:ACA

[10593] Anonymous. Apparatus for cooling air in theaters. Scientific American, 58(25):384, June 23, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n25/pdf/scientificamerican06231888-
384b.pdf.

Anonymous:1888:FSS

[10594] Anonymous. “Are fast or slow steamers the safest?”. Scientific American, 58(25):384–385, June 23, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n25/pdf/scientificamerican06231888-
384c.pdf.

Anonymous:1888:MNb

[10595] Anonymous. Miscellaneous notes. Scientific American, 58(25):385, June 23, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n25/pdf/scientificamerican06231888-385.pdf.

Anonymous:1888:PNe

[10596] Anonymous. Photographic notes. Scientific American, 58(25):385, June 23, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n25/pdf/scientificamerican06231888-385a.pdf.

Anonymous:1888:BNG

[10597] Anonymous. Burning natural gas. Scientific American, 58(25):386, June 23, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n25/pdf/scientificamerican06231888-386.pdf.

Anonymous:1888:CBC

[10598] Anonymous. Catching bait in Calais Harbor. Scientific American, 58(25):386, June 23, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n25/pdf/scientificamerican06231888-
386a.pdf.

Anonymous:1888:ISCa

[10599] Anonymous. An improved steam coupling for cars. Scientific American, 58(25):387, June 23, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n25/pdf/scientificamerican06231888-
387.pdf.

Anonymous:1888:ICHb

[10600] Anonymous. An improved copy holder. Scientific American, 58(25):387, June 23, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n25/pdf/scientificamerican06231888-387a.pdf.

Anonymous:1888:IFCa

[10601] Anonymous. An improved fan for cars. Scientific American, 58(25):387, June 23, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n25/pdf/scientificamerican06231888-387b.pdf.

Anonymous:1888:ITPa

[10602] Anonymous. An implement for transplanting plants. Scientific American, 58(25):387, June 23, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n25/pdf/scientificamerican06231888-
387c.pdf.

Anonymous:1888:IBCb

[10603] Anonymous. An improved bell collar for cattle. Scientific American, 58(25):387–388, June 23, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n25/pdf/scientificamerican06231888-
387d.pdf.

Anonymous:1888:IAT

[10604] Anonymous. An improved animal trap. Scientific American, 58 (25):388–389, June 23, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n25/pdf/scientificamerican06231888-
388.pdf.

Sloane:1888:PML

[10605] T. O’conor Sloane. Portable magic lantern. Scientific American, 58(25):391–393, June 23, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n25/pdf/scientificamerican06231888-
391.pdf.

Anonymous:1888:EIq

[10606] Anonymous. Engineering inventions. Scientific American, 58(25):394, June 23, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n25/pdf/scientificamerican06231888-394.pdf.

Anonymous:1888:AIn

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n25/pdf/scientificamerican06231888-394a.pdf.

Anonymous:1888:MIr

[10608] Anonymous. Miscellaneous inventions. Scientific American, 58(25):394, June 23, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n25/pdf/scientificamerican06231888-394b.pdf.

Anonymous:1888:Iy

[10609] Anonymous. To inventors. Scientific American, 58(25):395, June 23, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n25/pdf/scientificamerican06231888-395b.pdf.

Anonymous:1888:Cl

[10610] Anonymous. Correspondence. Scientific American, 58(26):405, June 30, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n26/pdf/scientificamerican06301888-405.pdf.

Anonymous:1888:BPs

[10611] Anonymous. Business and personal. Scientific American, 58(26):408, June 30, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n26/pdf/scientificamerican06301888-408c.pdf.

Anonymous:1888:BNBo

[10612] Anonymous. Books: New books and publications. Scientific American, 58(26):408–409, June 30, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n26/pdf/scientificamerican06301888-
408d.pdf.

Anonymous:1888:NQq

[10613] Anonymous. Notes and queries. Scientific American, 58(26):409, June 30, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n26/pdf/scientificamerican06301888-409.pdf.

Anonymous:1888:IIs

[10614] Anonymous. Index of inventions. Scientific American, 58(26):409–412, June 30, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n26/pdf/scientificamerican06301888-409b.pdf.

Anonymous:1888:PNPs

[10615] Anonymous. Patent notes: Patents. Scientific American, 58(26):414, June 30, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n26/pdf/scientificamerican06301888-414.pdf.

Anonymous:1888:AKB

[10616] Anonymous. The Arthur Kill Bridge. Scientific American, 58 (26):399–400, June 30, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n26/pdf/scientificamerican06301888-
399.pdf.

Anonymous:1888:NN

[10617] Anonymous. Naval notes. Scientific American, 58(26):400, June 30, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n26/pdf/scientificamerican06301888-400.pdf.

Anonymous:1888:MCB

[10618] Anonymous. Master car builders in convention. Scientific American, 58(26):401, June 30, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n26/pdf/scientificamerican06301888-
401.pdf.

Anonymous:1888:MHP

[10619] Anonymous. The Mason & Hamlin pianos and organs. Scientific American, 58(26):402–403, June 30, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n26/pdf/scientificamerican06301888-
402.pdf.

Anonymous:1888:IWFa

[10620] Anonymous. An improved wheel fender for carriages. Scientific American, 58(26):403, June 30, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n26/pdf/scientificamerican06301888-
403.pdf.

Anonymous:1888:IRCa

[10621] Anonymous. An improved railway crossing alarm signal. Scientific American, 58(26):403, June 30, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n26/pdf/scientificamerican06301888-
403a.pdf.

Anonymous:1888:IFCb

[10622] Anonymous. An improved folding cot. Scientific American, 58 (26):403–404, June 30, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n26/pdf/scientificamerican06301888-
403b.pdf.

Hovey:1888:BP

[10623] H. C. Hovey. Beetle photography. Scientific American, 58(26):406, June 30, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n26/pdf/scientificamerican06301888-406.pdf.

Anonymous:1888:CSO

[10624] Anonymous. Calming the sea with oil. Scientific American, 58 (26):406–407, June 30, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n26/pdf/scientificamerican06301888-
406a.pdf.

Anonymous:1888:EE

[10625] Anonymous. Execution by electricity. Scientific American, 58(26):407, June 30, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n26/pdf/scientificamerican06301888-407.pdf.

Anonymous:1888:EIr

[10626] Anonymous. Engineering invention. Scientific American, 58(26):408, June 30, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n26/pdf/scientificamerican06301888-408.pdf.

Anonymous:1888:AIo

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n26/pdf/scientificamerican06301888-408a.pdf.

Anonymous:1888:MIs

[10628] Anonymous. Miscellaneous inventions. Scientific American, 58(26):408, June 30, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n26/pdf/scientificamerican06301888-408b.pdf.

Anonymous:1888:Iz

[10629] Anonymous. To inventors. Scientific American, 58(26):409, June 30, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n26/pdf/scientificamerican06301888-409a.pdf.

Anonymous:1888:Cm

[10630] Anonymous. Correspondence. Scientific American, 58(3):37, January 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n3/pdf/scientificamerican01211888-37.pdf.

Anonymous:1888:BNBp

[10631] Anonymous. Books: New books and publications. Scientific American, 58(3):43, January 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n3/pdf/scientificamerican01211888-
43b.pdf.

Anonymous:1888:BPt

[10632] Anonymous. Business and personal. Scientific American, 58(3):43, January 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n3/pdf/scientificamerican01211888-43c.pdf.

Anonymous:1888:NQr

[10633] Anonymous. Notes and queries. Scientific American, 58(3):43, January 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n3/pdf/scientificamerican01211888-43e.pdf.

Anonymous:1888:IIt

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n3/pdf/scientificamerican01211888-44.pdf.

Anonymous:1888:PNPt

[10635] Anonymous. Patent notes: Patents. Scientific American, 58(3):46, January 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n3/pdf/scientificamerican01211888-46.pdf.

Anonymous:1888:BBP

[10636] Anonymous. Bridge building on the Pennsylvania Railway without interrupting traffic. Scientific American, 58(3):31, January 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n3/pdf/scientificamerican01211888-31.pdf.

Anonymous:1888:IPBb

[10637] Anonymous. The Indiana Patent Bill. Scientific American, 58 (3):32, January 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n3/pdf/scientificamerican01211888-
32.pdf.

Anonymous:1888:EEC

[10638] Anonymous. Electric energy from carbon without heat. Scientific American, 58(3):32, January 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n3/pdf/scientificamerican01211888-
32a.pdf.

Anonymous:1888:SMD

[10639] Anonymous. Success of the million-dollar telescope. Scientific American, 58(3):32–33, January 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n3/pdf/scientificamerican01211888-
32b.pdf.

Anonymous:1888:ITb

[10640] Anonymous. An improved tricycle. Scientific American, 58(3):34, January 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n3/pdf/scientificamerican01211888-34.pdf.

Anonymous:1888:IMCc

[10641] Anonymous. An improved music chart. Scientific American, 58 (3):34, January 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n3/pdf/scientificamerican01211888-
34a.pdf.

Hopkins:1888:SLB

[10642] M. Hopkins. Self-luminous buoy. Scientific American, 58(3):34, January 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n3/pdf/scientificamerican01211888-34b.pdf.

Anonymous:1888:AAC

[10643] Anonymous. An alarm attachment for clocks. Scientific American, 58(3):35, January 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n3/pdf/scientificamerican01211888-
35.pdf.

Anonymous:1888:IWSb

[10644] Anonymous. An improved weather strip. Scientific American, 58(3):35, January 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n3/pdf/scientificamerican01211888-
35a.pdf.

Anonymous:1888:IAH

[10645] Anonymous. Improved apparatus for handling brick. Scientific American, 58(3):35, January 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n3/pdf/scientificamerican01211888-
35b.pdf.

Anonymous:1888:ICAb

[10646] Anonymous. An improved car-axle box. Scientific American, 58 (3):35, January 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n3/pdf/scientificamerican01211888-
35c.pdf.

Anonymous:1888:ISMa

[10647] Anonymous. An improved sheet metal mantel. Scientific American, 58(3):35, January 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n3/pdf/scientificamerican01211888-
35d.pdf.

Anonymous:1888:AFM

[10648] Anonymous. The Abbe forging machines. Scientific American, 58(3):38, January 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n3/pdf/scientificamerican01211888-
38.pdf.

Anonymous:1888:RLP

[10649] Anonymous. Road locomotive for postal service. Scientific American, 58(3):38–39, January 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n3/pdf/scientificamerican01211888-
38a.pdf.

Holder:1888:MGC

[10650] J. B. Holder. The Mastodon Giganteus (Cuvier). Scientific American, 58(3):39–40, January 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n3/pdf/scientificamerican01211888-
39.pdf.

Anonymous:1888:ELS

[10651] Anonymous. Electric lighting of the steamships Victoria and Britannia. Scientific American, 58(3):41–42, January 21, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v58/
n3/pdf/scientificamerican01211888-41.pdf.

Anonymous:1888:EIs

[10652] Anonymous. Engineering inventions. Scientific American, 58(3): 43, January 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n3/pdf/scientificamerican01211888-
43.pdf.

Anonymous:1888:MIt

[10653] Anonymous. Miscellaneous inventions. Scientific American, 58 (3):43, January 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n3/pdf/scientificamerican01211888-
43a.pdf.

Anonymous:1888:Ibd

[10654] Anonymous. To inventors. Scientific American, 58(3):43, January 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n3/pdf/scientificamerican01211888-43d.pdf.

Anonymous:1888:Cn

[10655] Anonymous. Correspondence. Scientific American, 58(4):53, January 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n4/pdf/scientificamerican01281888-53.pdf.

Anonymous:1888:BNBq

[10656] Anonymous. Books: New books and publications. Scientific American, 58(4):58, January 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n4/pdf/scientificamerican01281888-
58c.pdf.

Anonymous:1888:BPu

[10657] Anonymous. Business and personal. Scientific American, 58 (4):58–59, January 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n4/pdf/scientificamerican01281888-
58e.pdf.

Anonymous:1888:NQs

[10658] Anonymous. Notes and queries. Scientific American, 58(4):59, January 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n4/pdf/scientificamerican01281888-59.pdf.

Anonymous:1888:IIu

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n4/pdf/scientificamerican01281888-59b.pdf.

Anonymous:1888:PNPu

[10660] Anonymous. Patent notes: Patents. Scientific American, 58(4):62, January 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n4/pdf/scientificamerican01281888-62.pdf.

Anonymous:1888:IWSc

[10661] Anonymous. Iron works of the Syracuse Malleable Iron Company. Scientific American, 58(4):47, January 28, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v58/
n4/pdf/scientificamerican01281888-47.pdf.

Anonymous:1888:CQ

[10662] Anonymous. The copyright question. Scientific American, 58 (4):48, January 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n4/pdf/scientificamerican01281888-
48.pdf.

Anonymous:1888:PPF

[10663] Anonymous. Position of the planets in February. Scientific American, 58(4):49, January 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n4/pdf/scientificamerican01281888-
49.pdf.

Anonymous:1888:PNf

[10664] Anonymous. Photographic notes. Scientific American, 58(4):49, January 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n4/pdf/scientificamerican01281888-49a.pdf.

Anonymous:1888:GRV

[10665] Anonymous. A guard rail for vehicle wheels. Scientific American, 58(4):51, January 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n4/pdf/scientificamerican01281888-
51.pdf.

Anonymous:1888:ITC

[10666] Anonymous. An improved thill coupling. Scientific American, 58(4):51, January 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n4/pdf/scientificamerican01281888-
51a.pdf.

Anonymous:1888:CRR

[10667] Anonymous. A combined railway rail chair and tie. Scientific American, 58(4):51, January 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n4/pdf/scientificamerican01281888-
51b.pdf.

Anonymous:1888:IMFa

[10668] Anonymous. An improved mason’s float handle. Scientific American, 58(4):51, January 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n4/pdf/scientificamerican01281888-
51c.pdf.

Anonymous:1888:IFb

[10669] Anonymous. An improved faucet. Scientific American, 58(4):51, January 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n4/pdf/scientificamerican01281888-51d.pdf.

Anonymous:1888:ATS

[10670] Anonymous. An apparatus for testing seeds. Scientific American, 58(4):51–52, January 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n4/pdf/scientificamerican01281888-
51e.pdf.

Anonymous:1888:IRLb

[10671] Anonymous. An improved ratchet and lever mechanism. Scientific American, 58(4):52, January 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n4/pdf/scientificamerican01281888-
52.pdf.

Anonymous:1888:MTW

[10672] Anonymous. The Morris type writer. Scientific American, 58 (4):54, January 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n4/pdf/scientificamerican01281888-
54.pdf.

Anonymous:1888:ASB

[10673] Anonymous. African sheep in the Berlin Zoological Gardens. Scientific American, 58(4):55–57, January 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v58/n4/pdf/scientificamerican01281888-
55.pdf.

Anonymous:1888:EIt

[10674] Anonymous. Engineering inventions. Scientific American, 58(4): 58, January 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n4/pdf/scientificamerican01281888-
58.pdf.

Anonymous:1888:AIp

[10675] Anonymous. Agricultural inventions. Scientific American, 58 (4):58, January 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n4/pdf/scientificamerican01281888-
58a.pdf.

Anonymous:1888:MIu

[10676] Anonymous. Miscellaneous inventions. Scientific American, 58 (4):58, January 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n4/pdf/scientificamerican01281888-
58b.pdf.

Anonymous:1888:Sh

[10677] Anonymous. Special. Scientific American, 58(4):58, January 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n4/pdf/scientificamerican01281888-58d.pdf.

Anonymous:1888:Ibe

[10678] Anonymous. To inventors. Scientific American, 58(4):59, January 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n4/pdf/scientificamerican01281888-59a.pdf.

Anonymous:1888:BNBr

[10679] Anonymous. Books: New books and publications. Scientific American, 58(5):74, February 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n5/pdf/scientificamerican02041888-
74b.pdf.

Anonymous:1888:BPv

[10680] Anonymous. Business and personal. Scientific American, 58 (5):74, February 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n5/pdf/scientificamerican02041888-
74c.pdf.

Anonymous:1888:NQt

[10681] Anonymous. Notes and queries. Scientific American, 58(5):75, February 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n5/pdf/scientificamerican02041888-75.pdf.

Anonymous:1888:IIv

[10682] Anonymous. Index of inventions. Scientific American, 58(5): 75–76, February 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n5/pdf/scientificamerican02041888-
75b.pdf.

Anonymous:1888:PNPv

[10683] Anonymous. Patent notes: Patents. Scientific American, 58 (5):78, February 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n5/pdf/scientificamerican02041888-
78.pdf.

Anonymous:1888:BPTa

[10684] Anonymous. The book publishing trades and international copyright. Scientific American, 58(5):64–65, February 4, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v58/
n5/pdf/scientificamerican02041888-64.pdf.

Anonymous:1888:PNg

[10685] Anonymous. Photographic notes. Scientific American, 58(5):65, February 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n5/pdf/scientificamerican02041888-65.pdf.

Anonymous:1888:WGW

[10686] Anonymous. Water gauge and water column combined. Scientific American, 58(5):66, February 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n5/pdf/scientificamerican02041888-
66.pdf.

Anonymous:1888:ILBa

[10687] Anonymous. An improved life boat. Scientific American, 58 (5):66, February 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n5/pdf/scientificamerican02041888-
66a.pdf.

Anonymous:1888:SEP

[10688] Anonymous. A simple and efficient pipe cutter. Scientific American, 58(5):66, February 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n5/pdf/scientificamerican02041888-
66b.pdf.

Anonymous:1888:IWC

[10689] Anonymous. An improved water closet box. Scientific American, 58(5):66, February 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n5/pdf/scientificamerican02041888-
66c.pdf.

Anonymous:1888:IRG

[10690] Anonymous. Improved running gear for vehicles. Scientific American, 58(5):66, February 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n5/pdf/scientificamerican02041888-
66d.pdf.

Anonymous:1888:ISGd

[10691] Anonymous. An improved saw gauge. Scientific American, 58 (5):67, February 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n5/pdf/scientificamerican02041888-
67.pdf.

Anonymous:1888:IHP

[10692] Anonymous. An improved horseshoe pad. Scientific American, 58(5):67, February 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n5/pdf/scientificamerican02041888-
67a.pdf.

Anonymous:1888:ICe

[10693] Anonymous. An improved cleat. Scientific American, 58(5):67, February 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n5/pdf/scientificamerican02041888-67b.pdf.

Anonymous:1888:CTI

[10694] Anonymous. A combination toilet implement. Scientific American, 58(5):67, February 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n5/pdf/scientificamerican02041888-
67c.pdf.

Anonymous:1888:CWCa

[10695] Anonymous. A combined whip and cane. Scientific American, 58(5):67, February 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n5/pdf/scientificamerican02041888-
67d.pdf.

Anonymous:1888:ITHb

[10696] Anonymous. Improved tongs for handling nails. Scientific American, 58(5):67–69, February 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n5/pdf/scientificamerican02041888-
67e.pdf.

Anonymous:1888:PBa

[10697] Anonymous. The Poughkeepsie Bridge. Scientific American, 58 (5):70–71, February 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n5/pdf/scientificamerican02041888-
70.pdf.

Anonymous:1888:MCM

[10698] Anonymous. Mathew, the Cuban monkey. Scientific American, 58(5):71–72, February 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n5/pdf/scientificamerican02041888-
71.pdf.

Sloane:1888:PV

[10699] T. O’Conor Sloane. Pascal’s vase. Scientific American, 58(5):73, February 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n5/pdf/scientificamerican02041888-73.pdf.

Higgins:1888:ETW

[10700] F. Higgins. The electrical type writer. Scientific American, 58 (5):73, February 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n5/pdf/scientificamerican02041888-
73a.pdf.

Anonymous:1888:EIu

[10701] Anonymous. Engineering invention. Scientific American, 58(5): 74, February 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n5/pdf/scientificamerican02041888-
74.pdf.

Anonymous:1888:MIv

[10702] Anonymous. Miscellaneous inventions. Scientific American, 58 (5):74, February 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n5/pdf/scientificamerican02041888-
74a.pdf.

Anonymous:1888:Ibf

[10703] Anonymous. To inventors. Scientific American, 58(5):75, February 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n5/pdf/scientificamerican02041888-75a.pdf.

Anonymous:1888:BPw

[10704] Anonymous. Business and personal. Scientific American, 58(6): 91, February 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n6/pdf/scientificamerican02111888-
91c.pdf.

Anonymous:1888:IIw

[10705] Anonymous. Index of inventions. Scientific American, 58(6):91– 92, February 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n6/pdf/scientificamerican02111888-
91e.pdf.

Anonymous:1888:PNPw

[10706] Anonymous. Patent notes: Patents. Scientific American, 58(6): 94, February 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n6/pdf/scientificamerican02111888-
94.pdf.

Anonymous:1888:INYb

[10707] Anonymous. The improvement of New York Harbor. Scientific American, 58(6):79, February 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n6/pdf/scientificamerican02111888-
79.pdf.

Anonymous:1888:SRA

[10708] Anonymous. A ship railway to assist the Panama Canal. Scientific American, 58(6):80, February 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n6/pdf/scientificamerican02111888-
80.pdf.

Anonymous:1888:CWd

[10709] Anonymous. The celestial world. Scientific American, 58(6):81– 82, February 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n6/pdf/scientificamerican02111888-
81.pdf.

Anonymous:1888:ECa

[10710] Anonymous. An electric carriage. Scientific American, 58(6):82, February 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n6/pdf/scientificamerican02111888-82.pdf.

Anonymous:1888:IFEc

[10711] Anonymous. An improved fire extinguisher. Scientific American, 58(6):82, February 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n6/pdf/scientificamerican02111888-
82a.pdf.

Anonymous:1888:ILBb

[10712] Anonymous. An improved life boat. Scientific American, 58 (6):83, February 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n6/pdf/scientificamerican02111888-
83.pdf.

Anonymous:1888:HIC

[10713] Anonymous. Harding’s improved counter. Scientific American, 58(6):83, February 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n6/pdf/scientificamerican02111888-
83a.pdf.

Anonymous:1888:IFCc

[10714] Anonymous. An improved folding clothes frame. Scientific American, 58(6):83, February 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n6/pdf/scientificamerican02111888-
83b.pdf.

Anonymous:1888:IWSd

[10715] Anonymous. An improved window sill. Scientific American, 58 (6):83, February 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n6/pdf/scientificamerican02111888-
83c.pdf.

Anonymous:1888:IPa

[10716] Anonymous. An improved pump. Scientific American, 58(6):83, February 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n6/pdf/scientificamerican02111888-83d.pdf.

Anonymous:1888:ITWa

[10717] Anonymous. An improved two wheeled vehicle. Scientific American, 58(6):83, February 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n6/pdf/scientificamerican02111888-
83e.pdf.

Anonymous:1888:HE

[10718] Anonymous. Hydraulic elevators. Scientific American, 58(6):86– 87, February 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n6/pdf/scientificamerican02111888-
86.pdf.

Anonymous:1888:PAG

[10719] Anonymous. Prof. Asa Gray. Scientific American, 58(6):87–88, February 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n6/pdf/scientificamerican02111888-87.pdf.

Anonymous:1888:RRW

[10720] Anonymous. A remarkable railway wreck. Scientific American, 58 (6):88–89, February 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n6/pdf/scientificamerican02111888-
88.pdf.

Anonymous:1888:SCP

[10721] Anonymous. A simple copying press. Scientific American, 58 (6):89, February 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n6/pdf/scientificamerican02111888-
89.pdf.

Anonymous:1888:ISBb

[10722] Anonymous. An improved steam boiler. Scientific American, 58 (6):89–90, February 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n6/pdf/scientificamerican02111888-
89a.pdf.

Anonymous:1888:EIv

[10723] Anonymous. Engineering inventions. Scientific American, 58(6): 91, February 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n6/pdf/scientificamerican02111888-
91.pdf.

Anonymous:1888:MIw

[10724] Anonymous. Miscellaneous inventions. Scientific American, 58 (6):91, February 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n6/pdf/scientificamerican02111888-
91a.pdf.

Anonymous:1888:Sj

[10725] Anonymous. Special. Scientific American, 58(6):91, February 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n6/pdf/scientificamerican02111888-91b.pdf.

Anonymous:1888:Ibg

[10726] Anonymous. To inventors. Scientific American, 58(6):91, February 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n6/pdf/scientificamerican02111888-91d.pdf.

Anonymous:1888:Co

[10727] Anonymous. Correspondence. Scientific American, 58(7):104, February 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n7/pdf/scientificamerican02181888-104.pdf.

Anonymous:1888:BPx

[10728] Anonymous. Business and personal. Scientific American, 58(7): 107, February 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n7/pdf/scientificamerican02181888-
107c.pdf.

Anonymous:1888:BNBs

[10729] Anonymous. Books: New books and publications. Scientific American, 58(7):107, February 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n7/pdf/scientificamerican02181888-
107d.pdf.

Anonymous:1888:IIx

[10730] Anonymous. Index of inventions. Scientific American, 58(7): 107–108, February 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n7/pdf/scientificamerican02181888-
107f.pdf.

Anonymous:1888:PNPx

[10731] Anonymous. Patent notes: Patents. Scientific American, 58 (7):110, February 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n7/pdf/scientificamerican02181888-
110.pdf.

Anonymous:1888:ARC

[10732] Anonymous. The Annual Report of the Commissioner of Patents. Scientific American, 58(7):96, February 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v58/n7/pdf/scientificamerican02181888-
96.pdf.

Anonymous:1888:DRP

[10733] Anonymous. Decisions relating to patents. Scientific American, 58(7):96–97, February 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n7/pdf/scientificamerican02181888-
96a.pdf.

Anonymous:1888:MNc

[10734] Anonymous. Military notes. Scientific American, 58(7):97, February 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n7/pdf/scientificamerican02181888-97.pdf.

Anonymous:1888:CIC

[10735] Anonymous. A convertible ice creeper and skate. Scientific American, 58(7):98, February 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n7/pdf/scientificamerican02181888-
98.pdf.

Anonymous:1888:NCPa

[10736] Anonymous. Non-conducting pipe coverings. Scientific American, 58(7):98, February 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n7/pdf/scientificamerican02181888-
98a.pdf.

Anonymous:1888:IACb

[10737] Anonymous. An improved ammunition case. Scientific American, 58(7):98, February 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n7/pdf/scientificamerican02181888-
98b.pdf.

Anonymous:1888:IRDa

[10738] Anonymous. An improved rocking and dumping grate. Scientific American, 58(7):99, February 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n7/pdf/scientificamerican02181888-
99.pdf.

Anonymous:1888:HPP

[10739] Anonymous. Hand and power presses for industrial uses. Scientific American, 58(7):99–100, February 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v58/n7/pdf/scientificamerican02181888-
99a.pdf.

Anonymous:1888:ISBc

[10740] Anonymous. An improved steam boiler. Scientific American, 58 (7):100, February 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n7/pdf/scientificamerican02181888-
100.pdf.

Anonymous:1888:ENH

[10741] Anonymous. Erection of the new Harlem River Bridge at 181st Street. Scientific American, 58(7):101, February 18, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v58/
n7/pdf/scientificamerican02181888-101.pdf.

Anonymous:1888:FTP

[10742] Anonymous. Fitting together parts of stoves. Scientific American, 58(7):102, February 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n7/pdf/scientificamerican02181888-
102.pdf.

Anonymous:1888:STb

[10743] Anonymous. A steam tricycle. Scientific American, 58(7):102, February 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n7/pdf/scientificamerican02181888-102a.pdf.

Seiss:1888:C

[10744] C. Few Seiss. The copperhead. Scientific American, 58(7):103, February 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n7/pdf/scientificamerican02181888-103.pdf.

Hopkins:1888:RES

[10745] M. Hopkins. Re-enforcement of sound. Scientific American, 58 (7):105–106, February 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n7/pdf/scientificamerican02181888-
105.pdf.

Anonymous:1888:EIw (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n7/pdf/scientificamerican02181888-
107.pdf.

Anonymous:1888:AIq

[10747] Anonymous. Agricultural inventions. Scientific American, 58(7): 107, February 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n7/pdf/scientificamerican02181888-
107a.pdf.

Anonymous:1888:MIx

[10748] Anonymous. Miscellaneous inventions. Scientific American, 58 (7):107, February 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n7/pdf/scientificamerican02181888-
107b.pdf.

Anonymous:1888:Ibi

[10749] Anonymous. To inventors. Scientific American, 58(7):107, February 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n7/pdf/scientificamerican02181888-107e.pdf.

Anonymous:1888:BPy

[10750] Anonymous. Business and personal. Scientific American, 58(8): 122, February 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n8/pdf/scientificamerican02251888-
122c.pdf.

Anonymous:1888:BNBt

[10751] Anonymous. Books: New books and publications. Scientific American, 58(8):122, February 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n8/pdf/scientificamerican02251888-
122d.pdf.

Anonymous:1888:NQu (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n8/pdf/scientificamerican02251888-
122e.pdf.

Anonymous:1888:IIy

[10753] Anonymous. Index of inventions. Scientific American, 58(8): 123–124, February 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n8/pdf/scientificamerican02251888-
123a.pdf.

Anonymous:1888:PNPy

[10754] Anonymous. Patent notes: Patents. Scientific American, 58 (8):126, February 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n8/pdf/scientificamerican02251888-
126.pdf.

Anonymous:1888:APL

[10755] Anonymous. Amendment of the Patent Laws. Scientific American, 58(8):112, February 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n8/pdf/scientificamerican02251888-
112.pdf.

Anonymous:1888:ABP

[10756] Anonymous. American and British “pirates”. Scientific American, 58(8):112–113, February 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n8/pdf/scientificamerican02251888-
112a.pdf.

Anonymous:1888:IEB

[10757] Anonymous. An improved electric bell. Scientific American, 58 (8):114, February 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n8/pdf/scientificamerican02251888-
114.pdf.

Anonymous:1888:ELG

[10758] Anonymous. Electric lighting of the Great Eastern. Scientific American, 58(8):114, February 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n8/pdf/scientificamerican02251888-
114a.pdf.

Anonymous:1888:IEC

[10759] Anonymous. An improved electric circuit closer. Scientific American, 58(8):114, February 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n8/pdf/scientificamerican02251888-
114b.pdf.

Anonymous:1888:MIWa

[10760] Anonymous. Making insulated wires for electrical uses. Scientific American, 58(8):115–116, February 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n8/pdf/scientificamerican02251888-
115.pdf.

Anonymous:1888:CDU

[10761] Anonymous. A chair for out-door use. Scientific American, 58 (8):116, February 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n8/pdf/scientificamerican02251888-
116.pdf.

Anonymous:1888:ICRb

[10762] Anonymous. An improved car replacer and portable switch. Scientific American, 58(8):116, February 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v58/n8/pdf/scientificamerican02251888-
116a.pdf.

Anonymous:1888:PCK

[10763] Anonymous. A pipe coupling for kitchen boilers. Scientific American, 58(8):116–117, February 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n8/pdf/scientificamerican02251888-
116b.pdf.

Anonymous:1888:BJW

[10764] Anonymous. The birthplace of James Watt. Scientific American, 58(8):117, February 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n8/pdf/scientificamerican02251888-
117.pdf.

Anonymous:1888:GB

[10765] Anonymous. The guns of the Benbow. Scientific American, 58 (8):118–119, February 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n8/pdf/scientificamerican02251888-
118.pdf.

H:1888:LMH

[10766] C. H. H. Lewis Muhlenberg Haupt. Scientific American, 58 (8):119, February 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n8/pdf/scientificamerican02251888-
119.pdf.

Anonymous:1888:U

[10767] Anonymous. Umbromania. Scientific American, 58(8):120–121, February 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n8/pdf/scientificamerican02251888-120.pdf.

Anonymous:1888:EIx

[10768] Anonymous. Engineering inventions. Scientific American, 58(8): 122, February 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n8/pdf/scientificamerican02251888-
122.pdf.

Anonymous:1888:MIy

[10769] Anonymous. Miscellaneous inventions. Scientific American, 58 (8):122, February 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n8/pdf/scientificamerican02251888-
122a.pdf.

Anonymous:1888:Sk

[10770] Anonymous. Special. Scientific American, 58(8):122, February 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n8/pdf/scientificamerican02251888-122b.pdf.

Anonymous:1888:Ibj

[10771] Anonymous. To inventors. Scientific American, 58(8):123, February 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n8/pdf/scientificamerican02251888-123.pdf.

Anonymous:1888:Cp

[10772] Anonymous. Correspondence. Scientific American, 58(9):133, March 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n9/pdf/scientificamerican03031888-133.pdf.

Anonymous:1888:BPz

[10773] Anonymous. Business and personal. Scientific American, 58(9):138, March 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n9/pdf/scientificamerican03031888-138b.pdf.

Anonymous:1888:BNBu

[10774] Anonymous. Books: New books and publications. Scientific American, 58(9):138, March 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n9/pdf/scientificamerican03031888-
138c.pdf.

Anonymous:1888:NQv

[10775] Anonymous. Notes and queries. Scientific American, 58(9):139, March 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n9/pdf/scientificamerican03031888-139.pdf.

Anonymous:1888:IIz

[10776] Anonymous. Index of inventions. Scientific American, 58(9):139–140, March 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n9/pdf/scientificamerican03031888-139a.pdf.

Anonymous:1888:PNPz

[10777] Anonymous. Patent notes: Patents. Scientific American, 58(9):142, March 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n9/pdf/scientificamerican03031888-142.pdf.

Anonymous:1888:ECb

[10778] Anonymous. The electric club. Scientific American, 58(9):127, March 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n9/pdf/scientificamerican03031888-127.pdf.

Anonymous:1888:PPMb

[10779] Anonymous. Position of the planets in March. Scientific American, 58(9):129, March 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n9/pdf/scientificamerican03031888-
129.pdf.

Anonymous:1888:PNh

[10780] Anonymous. Photographic notes. Scientific American, 58(9):129, March 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n9/pdf/scientificamerican03031888-129a.pdf.

Anonymous:1888:PCC

[10781] Anonymous. The Père La Chaise Crematory. Scientific American, 58(9):130, March 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n9/pdf/scientificamerican03031888-
130.pdf.

Anonymous:1888:ILI

[10782] Anonymous. An improved level and inclinometer. Scientific American, 58(9):131, March 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n9/pdf/scientificamerican03031888-
131.pdf.

Anonymous:1888:MGC

[10783] Anonymous. A machine for grinding circular knives. Scientific American, 58(9):131, March 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n9/pdf/scientificamerican03031888-
131a.pdf.

Anonymous:1888:ICCi

[10784] Anonymous. An improved car coupling. Scientific American, 58(9):131, March 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n9/pdf/scientificamerican03031888-131b.pdf.

Anonymous:1888:PVH

[10785] Anonymous. A portable vapor and hot air bath apparatus. Scientific American, 58(9):131, March 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n9/pdf/scientificamerican03031888-
131c.pdf.

Anonymous:1888:HTM

[10786] Anonymous. The Hargreaves thermo-motor. Scientific American, 58(9):134, March 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n9/pdf/scientificamerican03031888-
134.pdf.

Anonymous:1888:LGL

[10787] Anonymous. A live gorilla in London. Scientific American, 58 (9):135–136, March 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n9/pdf/scientificamerican03031888-
135.pdf.

Sloane:1888:AIN

[10788] T. O’conor Sloane. Apparatus for illustrating Newton’s Laws. Scientific American, 58(9):137, March 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v58/n9/pdf/scientificamerican03031888-
137.pdf.

Anonymous:1888:MWR

[10789] Anonymous. A machine for weighing, registering, and bagging grain. Scientific American, 58(9):137, March 3, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v58/
n9/pdf/scientificamerican03031888-137a.pdf.

Anonymous:1888:EIy

[10790] Anonymous. Engineering inventions. Scientific American, 58(9):138, March 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n9/pdf/scientificamerican03031888-138.pdf.

Anonymous:1888:MIz

[10791] Anonymous. Miscellaneous inventions. Scientific American, 58(9):138, March 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v58/n9/pdf/scientificamerican03031888-138a.pdf.

Anonymous:1888:Cr

[10792] Anonymous. Correspondence. Scientific American, 59(1):7, July 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n1/pdf/scientificamerican07071888-7.pdf.

Anonymous:1888:BPbb

[10793] Anonymous. Business and personal. Scientific American, 59(1):11, July 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n1/pdf/scientificamerican07071888-11c.pdf.

Anonymous:1888:BNBv

[10794] Anonymous. Books: New books and publications. Scientific American, 59(1):11, July 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n1/pdf/scientificamerican07071888-
11d.pdf.

Anonymous:1888:NQw

[10795] Anonymous. Notes and queries. Scientific American, 59(1):11, July 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n1/pdf/scientificamerican07071888-11e.pdf.

Anonymous:1888:IIba

[10796] Anonymous. Index of inventions. Scientific American, 59(1):11–12, July 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n1/pdf/scientificamerican07071888-11f.pdf.

Anonymous:1888:PNPba

[10797] Anonymous. Patent notes: Patents. Scientific American, 59(1):13, July 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n1/pdf/scientificamerican07071888-13.pdf.

Anonymous:1888:DI

[10798] Anonymous. A dissolving island. Scientific American, 59(1):2, July 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n1/pdf/scientificamerican07071888-2.pdf.

Anonymous:1888:PPJb

[10799] Anonymous. Position of the planets in July. Scientific American, 59(1): 2, July 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n1/pdf/scientificamerican07071888-2a.pdf.

Anonymous:1888:NYH

[10800] Anonymous. New York Harbor. Scientific American, 59(1):2–3, July 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n1/pdf/scientificamerican07071888-2b.pdf.

Anonymous:1888:IIIa

[10801] Anonymous. An instantaneous interest indicator. Scientific American, 59(1):4, July 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n1/pdf/scientificamerican07071888-
4.pdf.

Anonymous:1888:EEG

[10802] Anonymous. Express engine, Great Northern Railway. Scientific American, 59(1):4, July 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n1/pdf/scientificamerican07071888-
4a.pdf.

Anonymous:1888:IBGa

[10803] Anonymous. An improved block for guard rails, etc. Scientific American, 59(1):5, July 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n1/pdf/scientificamerican07071888-
5.pdf.

Anonymous:1888:IACc

[10804] Anonymous. An improved automatic car brake. Scientific American, 59(1):5, July 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n1/pdf/scientificamerican07071888-
5a.pdf.

Anonymous:1888:ICWa

[10805] Anonymous. An improved clutch for wagon brakes. Scientific American, 59(1):5, July 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n1/pdf/scientificamerican07071888-
5b.pdf.

Anonymous:1888:IWa

[10806] Anonymous. An improved wrench. Scientific American, 59(1):5, July 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n1/pdf/scientificamerican07071888-5c.pdf.

Anonymous:1888:NA

[10807] Anonymous. The new aqueduct. Scientific American, 59(1):6, July 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n1/pdf/scientificamerican07071888-6.pdf.

Anonymous:1888:LGS

[10808] Anonymous. A locomotive with great steaming capacity. Scientific American, 59(1):8, July 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n1/pdf/scientificamerican07071888-
8.pdf.

Anonymous:1888:WBD

[10809] Anonymous. Water buck and dwarf antelope in the Zoological Garden at Cologne. Scientific American, 59(1):9–10, July 7, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v59/
n1/pdf/scientificamerican07071888-9.pdf.

Anonymous:1888:SGH

[10810] Anonymous. “Spiral” garden hose. Scientific American, 59(1):10, July 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n1/pdf/scientificamerican07071888-10.pdf.

Anonymous:1888:EIz

[10811] Anonymous. Engineering inventions. Scientific American, 59(1):11, July 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n1/pdf/scientificamerican07071888-11.pdf.

Anonymous:1888:AIr

[10812] Anonymous. Agricultural inventions. Scientific American, 59(1):11, July 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n1/pdf/scientificamerican07071888-11a.pdf.

Anonymous:1888:MIba

[10813] Anonymous. Miscellaneous inventions. Scientific American, 59(1):11, July 7, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n1/pdf/scientificamerican07071888-11b.pdf.

Anonymous:1888:Ct

[10814] Anonymous. Correspondence. Scientific American, 59(10):149, September 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n10/pdf/scientificamerican09081888-149a.pdf.

Anonymous:1888:BPbc

[10815] Anonymous. Business and personal. Scientific American, 59 (10):154, September 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n10/pdf/scientificamerican09081888-
154c.pdf.

Anonymous:1888:NQx

[10816] Anonymous. Notes and queries. Scientific American, 59(10): 154–155, September 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n10/pdf/scientificamerican09081888-
154d.pdf.

Anonymous:1888:IIbb

[10817] Anonymous. Index of inventions. Scientific American, 59(10): 155–156, September 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n10/pdf/scientificamerican09081888-
155.pdf.

Anonymous:1888:PNPbb

[10818] Anonymous. Patent notes: Patents. Scientific American, 59 (10):158, September 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n10/pdf/scientificamerican09081888-
158.pdf.

Anonymous:1888:DES

[10819] Anonymous. Does excitement shorten life? Scientific American, 59(10):144, September 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n10/pdf/scientificamerican09081888-
144.pdf.

Anonymous:1888:ELCa

[10820] Anonymous. Electrical Lighting Convention notes. Scientific American, 59(10):144–145, September 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n10/pdf/scientificamerican09081888-
144a.pdf.

Anonymous:1888:AIGa

[10821] Anonymous. An all-iron garden hose reel. Scientific American, 59(10):146, September 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n10/pdf/scientificamerican09081888-
146.pdf.

Anonymous:1888:IFCd

[10822] Anonymous. An improved file case. Scientific American, 59 (10):146, September 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n10/pdf/scientificamerican09081888-
146a.pdf.

Anonymous:1888:CFH

[10823] Anonymous. A clamp to facilitate hanging wall paper. Scientific American, 59(10):146, September 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n10/pdf/scientificamerican09081888-
146b.pdf.

Anonymous:1888:IWBa

[10824] Anonymous. An improved wire box strap. Scientific American, 59(10):146, September 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n10/pdf/scientificamerican09081888-
146c.pdf.

Anonymous:1888:WSF

[10825] Anonymous. A window stand for flowers, etc. Scientific American, 59(10):146, September 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n10/pdf/scientificamerican09081888-
146d.pdf.

Anonymous:1888:IWSe

[10826] Anonymous. An improved weather strip. Scientific American, 59 (10):146, September 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n10/pdf/scientificamerican09081888-
146e.pdf.

Anonymous:1888:IPC

[10827] Anonymous. An improved porous cup battery. Scientific American, 59(10):147, September 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n10/pdf/scientificamerican09081888-
147.pdf.

Anonymous:1888:IWd

[10828] Anonymous. An improved wrench. Scientific American, 59(10): 147, September 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n10/pdf/scientificamerican09081888-
147a.pdf.

Anonymous:1888:L

[10829] Anonymous. The Lucigen. Scientific American, 59(10):147– 148, September 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n10/pdf/scientificamerican09081888-
147b.pdf.

Anonymous:1888:TGS

[10830] Anonymous. Thompson’s gravity system for rapid transit in towns and cities. Scientific American, 59(10):149, September 8, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v59/
n10/pdf/scientificamerican09081888-149.pdf.

Anonymous:1888:BM

[10831] Anonymous. Boulogne-Sur-Mer. Scientific American, 59(10): 150, September 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n10/pdf/scientificamerican09081888-
150.pdf.

Coryell:1888:TPO

[10832] John R. Coryell. The termite pest of the Old World. Scientific American, 59(10):151–153, September 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v59/n10/pdf/scientificamerican09081888-
151.pdf.

Anonymous:1888:OFI

[10833] Anonymous. Our first imported locomotive. Scientific American, 59(10):153, September 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n10/pdf/scientificamerican09081888-
153.pdf.

Anonymous:1888:EIba

[10834] Anonymous. Engineering inventions. Scientific American, 59(10): 154, September 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n10/pdf/scientificamerican09081888-
154.pdf.

Anonymous:1888:MIbb

[10835] Anonymous. Miscellaneous inventions. Scientific American, 59 (10):154, September 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n10/pdf/scientificamerican09081888-
154a.pdf.

Anonymous:1888:Ibk

[10836] Anonymous. To inventors. Scientific American, 59(10):154, September 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n10/pdf/scientificamerican09081888-154b.pdf.

Anonymous:1888:Cu

[10837] Anonymous. Correspondence. Scientific American, 59(11):165– 166, September 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n11/pdf/scientificamerican09151888-
165.pdf.

Anonymous:1888:BPbd

[10838] Anonymous. Business and personal. Scientific American, 59 (11):170, September 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n11/pdf/scientificamerican09151888-
170c.pdf.

Anonymous:1888:NQy

[10839] Anonymous. Notes and queries. Scientific American, 59(11): 170–171, September 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n11/pdf/scientificamerican09151888-
170d.pdf.

Anonymous:1888:IIbc

[10840] Anonymous. Index of inventions. Scientific American, 59(11): 171–172, September 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n11/pdf/scientificamerican09151888-
171a.pdf.

Anonymous:1888:PNPbc

[10841] Anonymous. Patent notes: Patents. Scientific American, 59 (11):174, September 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n11/pdf/scientificamerican09151888-
174.pdf.

Anonymous:1888:IPd

[10842] Anonymous. Instantaneous photography. Scientific American, 59 (11):159, September 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n11/pdf/scientificamerican09151888-
159.pdf.

Anonymous:1888:TCN

[10843] Anonymous. Telephone Convention notes. Scientific American, 59 (11):160, September 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n11/pdf/scientificamerican09151888-
160.pdf.

Anonymous:1888:LWS

[10844] Anonymous. The locomotive water scoop. Scientific American, 59 (11):160–161, September 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n11/pdf/scientificamerican09151888-
160a.pdf.

Anonymous:1888:FL

[10845] Anonymous. Ferdinand De Lesseps. Scientific American, 59(11): 162, September 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n11/pdf/scientificamerican09151888-
162.pdf.

Anonymous:1888:RSR

[10846] Anonymous. A regulator for steam radiators. Scientific American, 59 (11):162–163, September 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n11/pdf/scientificamerican09151888-
162a.pdf.

Anonymous:1888:IVC

[10847] Anonymous. An improved vehicle curtain. Scientific American, 59 (11):163, September 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n11/pdf/scientificamerican09151888-
163.pdf.

Anonymous:1888:IVS

[10848] Anonymous. An improved vehicle spindle. Scientific American, 59 (11):163, September 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n11/pdf/scientificamerican09151888-
163a.pdf.

Anonymous:1888:IFTb

[10849] Anonymous. An improved feed trough for stock. Scientific American, 59(11):163, September 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n11/pdf/scientificamerican09151888-
163b.pdf.

Anonymous:1888:IIMa

[10850] Anonymous. An improved ingot manipulator. Scientific American, 59(11):166, September 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n11/pdf/scientificamerican09151888-
166.pdf.

Anonymous:1888:NTP

[10851] Anonymous. New type of photographic portraits. Scientific American, 59(11):166, September 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n11/pdf/scientificamerican09151888-
166a.pdf.

Anonymous:1888:MILa

[10852] Anonymous. The machinery of the Inman Liner City of New York. Scientific American, 59(11):167–169, September 15, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v59/
n11/pdf/scientificamerican09151888-167.pdf.

Anonymous:1888:IEa

[10853] Anonymous. Inoculating an elephant. Scientific American, 59 (11):169, September 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n11/pdf/scientificamerican09151888-
169.pdf.

Anonymous:1888:EIbb

[10854] Anonymous. Engineering inventions. Scientific American, 59(11): 170, September 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n11/pdf/scientificamerican09151888-
170.pdf.

Anonymous:1888:AIs

[10855] Anonymous. Agricultural inventions. Scientific American, 59 (11):170, September 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n11/pdf/scientificamerican09151888-
170a.pdf.

Anonymous:1888:MIbc

[10856] Anonymous. Miscellaneous inventions. Scientific American, 59 (11):170, September 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n11/pdf/scientificamerican09151888-
170b.pdf.

Anonymous:1888:Ibl

[10857] Anonymous. To inventors. Scientific American, 59(11):171, September 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n11/pdf/scientificamerican09151888-171.pdf.

Anonymous:1888:Cv

[10858] Anonymous. Correspondence. Scientific American, 59(12):181, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n12/pdf/scientificamerican09221888-181.pdf.

Anonymous:1888:BPbe

[10859] Anonymous. Business and personal. Scientific American, 59 (12):186, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
186b.pdf.

Anonymous:1888:NQz

[10860] Anonymous. Notes and queries. Scientific American, 59(12): 186–187, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
186c.pdf.

Anonymous:1888:IIbd

[10861] Anonymous. Index of inventions. Scientific American, 59(12): 187–188, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
187a.pdf.

Anonymous:1888:PNPbd

[10862] Anonymous. Patent notes: Patents. Scientific American, 59 (12):190, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
190.pdf.

Anonymous:1888:CSCa

[10863] Anonymous. Compressed steel car wheels. Scientific American, 59 (12):175, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
175.pdf.

Anonymous:1888:WBI

[10864] Anonymous. What is the best insulation? Scientific American, 59(12):176, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
176.pdf.

Anonymous:1888:APF

[10865] Anonymous. Additional postal facilities needed. Scientific American, 59(12):176–177, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
176a.pdf.

Anonymous:1888:IPBc

[10866] Anonymous. An improved punch for bank checks, etc. Scientific American, 59(12):179, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
179.pdf.

Anonymous:1888:IESb

[10867] Anonymous. Improved electrical shoal water indicator. Scientific American, 59(12):179, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
179a.pdf.

Anonymous:1888:ICSc

[10868] Anonymous. An improved cartridge shell cutter. Scientific American, 59(12):179, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
179b.pdf.

Anonymous:1888:ISW

[10869] Anonymous. An improved stop and waste faucet. Scientific American, 59(12):179, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
179c.pdf.

Anonymous:1888:ICCj

[10870] Anonymous. An improved coal chute. Scientific American, 59 (12):179, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
179d.pdf.

Anonymous:1888:ITPb

[10871] Anonymous. An improved tobacco pipe. Scientific American, 59 (12):180, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
180.pdf.

Anonymous:1888:ISSc

[10872] Anonymous. An improved store service apparatus. Scientific American, 59(12):180, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
180a.pdf.

Anonymous:1888:FSM

[10873] Anonymous. A finger shield for musicians. Scientific American, 59(12):180, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
180b.pdf.

Anonymous:1888:ICVa

[10874] Anonymous. An improved cut-off valve gear. Scientific American, 59(12):180, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
180c.pdf.

Anonymous:1888:ILBc

[10875] Anonymous. An improved letter box door plate. Scientific American, 59(12):180–181, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
180d.pdf.

Anonymous:1888:GGF

[10876] Anonymous. A gunboat for the Great French Exposition. Scientific American, 59(12):182–183, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
182.pdf.

Anonymous:1888:IPMb

[10877] Anonymous. An improved pug mill. Scientific American, 59 (12):183, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
183.pdf.

Anonymous:1888:IDG

[10878] Anonymous. An improved drill grinder. Scientific American, 59 (12):183, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
183a.pdf.

Anonymous:1888:IRP

[10879] Anonymous. An improved rail punch. Scientific American, 59(12): 183–184, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
183b.pdf.

Anonymous:1888:PNi

[10880] Anonymous. Photographic notes. Scientific American, 59(12): 184, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
184.pdf.

Hopkins:1888:SEPa

[10881] M. Hopkins. Simple experiments in physics. Scientific American, 59(12):185, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
185.pdf.

Anonymous:1888:EIbc

[10882] Anonymous. Engineering inventions. Scientific American, 59(12): 186, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
186.pdf.

Anonymous:1888:MIbd

[10883] Anonymous. Miscellaneous inventions. Scientific American, 59 (12):186, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n12/pdf/scientificamerican09221888-
186a.pdf.

Anonymous:1888:Ibm

[10884] Anonymous. To inventors. Scientific American, 59(12):187, September 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n12/pdf/scientificamerican09221888-187.pdf.

Anonymous:1888:Cx

[10885] Anonymous. Correspondence. Scientific American, 59(13):197, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n13/pdf/scientificamerican09291888-197.pdf.

Anonymous:1888:BPbf

[10886] Anonymous. Business and personal. Scientific American, 59 (13):202, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
202c.pdf.

Anonymous:1888:NQba

[10887] Anonymous. Notes and queries. Scientific American, 59(13): 202–203, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
202d.pdf.

Anonymous:1888:IIbe

[10888] Anonymous. Index of inventions. Scientific American, 59(13): 203–204, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
203a.pdf.

Anonymous:1888:PNPbe

[10889] Anonymous. Patent notes: Patents. Scientific American, 59 (13):206, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
206.pdf.

Anonymous:1888:PBM

[10890] Anonymous. Proposed bridge with movable floor over the River Bilboa, Spain. Scientific American, 59(13):191, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n13/pdf/scientificamerican09291888-191.pdf.

Anonymous:1888:ELI

[10891] Anonymous. Electric lighting information. Scientific American, 59 (13):192, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
192.pdf.

Anonymous:1888:PPO

[10892] Anonymous. Position of the planets in October. Scientific American, 59(13):192–193, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
192a.pdf.

Anonymous:1888:ICRc

[10893] Anonymous. An improved carpenter’s rule. Scientific American, 59(13):194, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
194.pdf.

Anonymous:1888:IDEc

[10894] Anonymous. An improved draft equalizer. Scientific American, 59 (13):194, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
194a.pdf.

Anonymous:1888:ITHc

[10895] Anonymous. An improved tool holder for grindstones. Scientific American, 59(13):194, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
194b.pdf.

Anonymous:1888:SSCa

[10896] Anonymous. A safety shoe for car trucks. Scientific American, 59(13):194, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
194c.pdf.

Anonymous:1888:IRDb

[10897] Anonymous. An improved reading desk. Scientific American, 59 (13):194, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
194d.pdf.

Anonymous:1888:ISLb

[10898] Anonymous. An improved seal lock. Scientific American, 59 (13):194, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
194e.pdf.

Anonymous:1888:IMCd

[10899] Anonymous. An improvement in mantel construction. Scientific American, 59(13):195, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
195.pdf.

Anonymous:1888:IFWa

[10900] Anonymous. An improved filter for waste oil. Scientific American, 59(13):195, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
195a.pdf.

Anonymous:1888:NMB

[10901] Anonymous. A new method of boat propulsion. Scientific American, 59(13):195, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
195b.pdf.

Anonymous:1888:IGT

[10902] Anonymous. An improved gearing for transmitting motion. Scientific American, 59(13):195, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
195c.pdf.

Anonymous:1888:ISf

[10903] Anonymous. An improved saw. Scientific American, 59(13): 195–196, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
195d.pdf.

Anonymous:1888:EGL

[10904] Anonymous. An electrical gas lighter and extinguisher. Scientific American, 59(13):198, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
198.pdf.

Anonymous:1888:WIR

[10905] Anonymous. The Waterhouse instantaneous regulator. Scientific American, 59(13):198, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
198a.pdf.

Anonymous:1888:ICBa

[10906] Anonymous. An improved charcoal burner. Scientific American, 59(13):198, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
198b.pdf.

Anonymous:1888:GE

[10907] Anonymous. The Glasgow Exhibition. Scientific American, 59(13): 199–201, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
199.pdf.

Anonymous:1888:TNZ

[10908] Anonymous. Two novelties in the Zoological Gardens. Scientific American, 59(13):201, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
201.pdf.

Anonymous:1888:EIbd

[10909] Anonymous. Engineering inventions. Scientific American, 59(13): 202, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
202.pdf.

Anonymous:1888:AIt

[10910] Anonymous. Agricultural inventions. Scientific American, 59 (13):202, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
202a.pdf.

Anonymous:1888:MIbe

[10911] Anonymous. Miscellaneous inventions. Scientific American, 59 (13):202, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n13/pdf/scientificamerican09291888-
202b.pdf.

Anonymous:1888:Ibn

[10912] Anonymous. To inventors. Scientific American, 59(13):203, September 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n13/pdf/scientificamerican09291888-203.pdf.

Anonymous:1888:BPbg

[10913] Anonymous. Business and personal. Scientific American, 59(14):218, October 6, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n14/pdf/scientificamerican10061888-218c.pdf.

Anonymous:1888:BNBw

[10914] Anonymous. Books: New books and publications. Scientific American, 59(14):218, October 6, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n14/pdf/scientificamerican10061888-
218d.pdf.

Anonymous:1888:NQbb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n14/pdf/scientificamerican10061888-218e.pdf.

Anonymous:1888:IIbf

[10916] Anonymous. Index of inventions. Scientific American, 59(14):219–220, October 6, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n14/pdf/scientificamerican10061888-219a.pdf.

Anonymous:1888:PNPbf

[10917] Anonymous. Patent notes: Patents. Scientific American, 59(14):222, October 6, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n14/pdf/scientificamerican10061888-222.pdf.

Anonymous:1888:HBA

[10918] Anonymous. How best to avoid collision at sea. Scientific American, 59(14):208, October 6, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n14/pdf/scientificamerican10061888-
208.pdf.

Anonymous:1888:CLT

[10919] Anonymous. Colored light trials with the insane. Scientific American, 59(14):208–209, October 6, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n14/pdf/scientificamerican10061888-
208a.pdf.

Anonymous:1888:PIB

[10920] Anonymous. Pratt Institute, Brooklyn, N.Y. Scientific American, 59(14):210–211, October 6, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n14/pdf/scientificamerican10061888-
210.pdf.

Anonymous:1888:SHW

[10921] Anonymous. A system to heat walls of buildings. Scientific American, 59(14):212, October 6, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n14/pdf/scientificamerican10061888-
212.pdf.

Anonymous:1888:ITe

[10922] Anonymous. An improved tricycle. Scientific American, 59(14): 212–213, October 6, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n14/pdf/scientificamerican10061888-
212a.pdf.

Anonymous:1888:ECc

[10923] Anonymous. An electric carriage. Scientific American, 59(14):215–216, October 6, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n14/pdf/scientificamerican10061888-215.pdf.

Anonymous:1888:SAS

[10924] Anonymous. Simple attachment for stoves. Scientific American, 59 (14):216–217, October 6, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n14/pdf/scientificamerican10061888-
216.pdf.

Anonymous:1888:IDW

[10925] Anonymous. An improved door or window stop. Scientific American, 59(14):217, October 6, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n14/pdf/scientificamerican10061888-
217.pdf.

Anonymous:1888:IWe

[10926] Anonymous. An improved windmill. Scientific American, 59(14):217, October 6, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n14/pdf/scientificamerican10061888-217a.pdf.

Anonymous:1888:IFd

[10927] Anonymous. An improved fireplace. Scientific American, 59(14):217, October 6, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n14/pdf/scientificamerican10061888-217b.pdf.

Anonymous:1888:EIbe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n14/pdf/scientificamerican10061888-218.pdf.

Anonymous:1888:AIu

[10929] Anonymous. Agricultural inventions. Scientific American, 59(14):218, October 6, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n14/pdf/scientificamerican10061888-218a.pdf.

Anonymous:1888:MIbf

[10930] Anonymous. Miscellaneous inventions. Scientific American, 59(14):218, October 6, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n14/pdf/scientificamerican10061888-218b.pdf.

Anonymous:1888:Ibo

[10931] Anonymous. To inventors. Scientific American, 59(14):219, October 6, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n14/pdf/scientificamerican10061888-219.pdf.

Anonymous:1888:Cy

[10932] Anonymous. Correspondence. Scientific American, 59(15):229–230, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n15/pdf/scientificamerican10131888-229a.pdf.

Anonymous:1888:BPbh

[10933] Anonymous. Business and personal. Scientific American, 59 (15):234, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
234c.pdf.

Anonymous:1888:BNBx

[10934] Anonymous. Books: New books and publications. Scientific American, 59(15):234–235, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
234d.pdf.

Anonymous:1888:NQbc

[10935] Anonymous. Notes and queries. Scientific American, 59(15):235, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n15/pdf/scientificamerican10131888-235.pdf.

Anonymous:1888:IIbg

[10936] Anonymous. Index of inventions. Scientific American, 59(15): 235–236, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
235b.pdf.

Anonymous:1888:PNPbg

[10937] Anonymous. Patent notes: Patents. Scientific American, 59 (15):238, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
238.pdf.

Anonymous:1888:LEG

[10938] Anonymous. A letter envelope gum greatly needed. Scientific American, 59(15):224, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
224.pdf.

Anonymous:1888:PEM

[10939] Anonymous. A proposed electro-motor trial. Scientific American, 59(15):224, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
224a.pdf.

Anonymous:1888:NPS

[10940] Anonymous. The new Philadelphia subways. Scientific American, 59(15):224, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
224b.pdf.

Anonymous:1888:DPL

[10941] Anonymous. The direct production of light. Scientific American, 59(15):224–225, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
224c.pdf.

Anonymous:1888:IBR

[10942] Anonymous. An improved belt rest. Scientific American, 59 (15):226, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
226.pdf.

Anonymous:1888:GEM

[10943] Anonymous. The Great Eastern moving to her last berth. Scientific American, 59(15):226, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
226a.pdf.

Anonymous:1888:AGE

[10944] Anonymous. An automatic gate for elevators. Scientific American, 59(15):226, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
226b.pdf.

Anonymous:1888:IABa

[10945] Anonymous. An improved axle box. Scientific American, 59 (15):227, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
227.pdf.

Anonymous:1888:ICRd

[10946] Anonymous. An improved check row planter. Scientific American, 59(15):227, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
227a.pdf.

Anonymous:1888:SUB

[10947] Anonymous. A seat for use in bath tubs. Scientific American, 59(15):227, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
227b.pdf.

Anonymous:1888:ITSb

[10948] Anonymous. An improved tailor’s square. Scientific American, 59(15):227, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
227c.pdf.

Anonymous:1888:IPH

[10949] Anonymous. An improved paper hanging machine. Scientific American, 59(15):227–228, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
227d.pdf.

Anonymous:1888:ICHc

[10950] Anonymous. An improved cuff holder. Scientific American, 59 (15):228, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
228.pdf.

Anonymous:1888:DUD

[10951] Anonymous. A device for use in darning stockings. Scientific American, 59(15):228, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
228a.pdf.

Anonymous:1888:IBBc

[10952] Anonymous. An improved bosom board. Scientific American, 59(15):228, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
228b.pdf.

Anonymous:1888:UCB

[10953] Anonymous. The U.S. cruiser Baltimore. Scientific American, 59(15):229, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
229.pdf.

Anonymous:1888:CBa

[10954] Anonymous. The centrifugal barrel. Scientific American, 59(15): 230–231, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
230.pdf.

Hopkins:1888:PC

[10955] Geo. M. Hopkins. A pocket camera. Scientific American, 59 (15):231–232, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
231.pdf.

Coryell:1888:G

[10956] John R. Coryell. The glyptodon. Scientific American, 59(15):233, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n15/pdf/scientificamerican10131888-233.pdf.

Anonymous:1888:EIbf

[10957] Anonymous. Engineering intentions. Scientific American, 59(15): 234, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
234.pdf.

Anonymous:1888:AIv

[10958] Anonymous. Agricultural inventions. Scientific American, 59 (15):234, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
234a.pdf.

Anonymous:1888:MIbg

[10959] Anonymous. Miscellaneous intentions. Scientific American, 59 (15):234, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n15/pdf/scientificamerican10131888-
234b.pdf.

Anonymous:1888:Ibq

[10960] Anonymous. To inventors. Scientific American, 59(15):235, October 13, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n15/pdf/scientificamerican10131888-235a.pdf.

Anonymous:1888:Cz

[10961] Anonymous. Correspondence. Scientific American, 59(16):245, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n16/pdf/scientificamerican10201888-245a.pdf.

Anonymous:1888:BPbi

[10962] Anonymous. Business and personal. Scientific American, 59 (16):250, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n16/pdf/scientificamerican10201888-
250b.pdf.

Anonymous:1888:BNBy

[10963] Anonymous. Books: New books and publications. Scientific American, 59(16):250–251, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n16/pdf/scientificamerican10201888-
250c.pdf.

Anonymous:1888:NQbd

[10964] Anonymous. Notes and queries. Scientific American, 59(16):251, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n16/pdf/scientificamerican10201888-251.pdf.

Anonymous:1888:IIbh 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n16/pdf/scientificamerican10201888-
251b.pdf.

Anonymous:1888:PNPbh

[10966] Anonymous. Patent notes: Patents. Scientific American, 59 (16):254, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n16/pdf/scientificamerican10201888-
254.pdf.

Anonymous:1888:AIFa

[10967] Anonymous. American institute fair. Scientific American, 59(16): 240–241, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n16/pdf/scientificamerican10201888-
240.pdf.

Anonymous:1888:HH

[10968] Anonymous. The home of the hop. Scientific American, 59(16):241, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n16/pdf/scientificamerican10201888-241.pdf.

Anonymous:1888:ICCk

[10969] Anonymous. An improved car coupling. Scientific American, 59 (16):242, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n16/pdf/scientificamerican10201888-
242.pdf.

Anonymous:1888:BCF

[10970] Anonymous. A band cutter and feeder for thrashing machines. Scientific American, 59(16):242, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v59/n16/pdf/scientificamerican10201888-
242a.pdf.

Anonymous:1888:ISj

[10971] Anonymous. An improved sled. Scientific American, 59(16):242, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n16/pdf/scientificamerican10201888-242b.pdf.

Anonymous:1888:IWg

[10972] Anonymous. An improved wrench. Scientific American, 59(16):242, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n16/pdf/scientificamerican10201888-242c.pdf.

Anonymous:1888:ISLc

[10973] Anonymous. An improved sash lock. Scientific American, 59 (16):242, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n16/pdf/scientificamerican10201888-
242d.pdf.

Anonymous:1888:ISPa

[10974] Anonymous. An improved screw propeller. Scientific American, 59(16):243, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n16/pdf/scientificamerican10201888-
243.pdf.

Anonymous:1888:IFMb

[10975] Anonymous. An improved fanning mill. Scientific American, 59 (16):243, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n16/pdf/scientificamerican10201888-
243a.pdf.

Anonymous:1888:IWBb

[10976] Anonymous. An improved wagon body. Scientific American, 59 (16):243, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n16/pdf/scientificamerican10201888-
243b.pdf.

Anonymous:1888:HT

[10977] Anonymous. The Howell torpedo. Scientific American, 59(16): 243–244, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n16/pdf/scientificamerican10201888-
243c.pdf.

Anonymous:1888:IFN

[10978] Anonymous. An improved fly net. Scientific American, 59(16):244, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n16/pdf/scientificamerican10201888-244.pdf.

Anonymous:1888:ICDb

[10979] Anonymous. An improved clothes drier. Scientific American, 59 (16):244, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n16/pdf/scientificamerican10201888-
244a.pdf.

Anonymous:1888:IBBd

[10980] Anonymous. An improved bridle for brooms or brushes. Scientific American, 59(16):244, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v59/n16/pdf/scientificamerican10201888-
244b.pdf.

Anonymous:1888:GCCb

[10981] Anonymous. The great chimney at the Clark Thread Works, Kearney, N.J., height 335 feet. Scientific American, 59(16):245, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n16/pdf/scientificamerican10201888-245.pdf.

Anonymous:1888:IAa

[10982] Anonymous. An improved auger. Scientific American, 59(16):245, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n16/pdf/scientificamerican10201888-245b.pdf.

Anonymous:1888:MFF

[10983] Anonymous. Mougin’s “Fort of the Future”. Scientific American, 59(16):246–247, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n16/pdf/scientificamerican10201888-
246.pdf.

Anonymous:1888:PBb

[10984] Anonymous. A prehistoric bear. Scientific American, 59(16): 247–248, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n16/pdf/scientificamerican10201888-
247.pdf.

Anonymous:1888:PRC

[10985] Anonymous. Pavilion of the Republic of Chili at the Paris Exposition of 1889. Scientific American, 59(16):249, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n16/pdf/scientificamerican10201888-249.pdf.

Anonymous:1888:EIbg

[10986] Anonymous. Engineering inventions. Scientific American, 59(16): 250, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n16/pdf/scientificamerican10201888-
250.pdf.

Anonymous:1888:MIbh

[10987] Anonymous. Miscellaneous inventions. Scientific American, 59 (16):250, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n16/pdf/scientificamerican10201888-
250a.pdf.

Anonymous:1888:Ibu

[10988] Anonymous. To inventors. Scientific American, 59(16):251, October 20, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n16/pdf/scientificamerican10201888-251a.pdf.

Anonymous:1888:Cbb

[10989] Anonymous. Correspondence. Scientific American, 59(17):261, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n17/pdf/scientificamerican10271888-261a.pdf.

Anonymous:1888:BPbj

[10990] Anonymous. Business and personal. Scientific American, 59 (17):266, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n17/pdf/scientificamerican10271888-
266c.pdf.

Anonymous:1888:BNBz

[10991] Anonymous. Books: New books and publications. Scientific American, 59(17):267, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n17/pdf/scientificamerican10271888-
267.pdf.

Anonymous:1888:NQbe

[10992] Anonymous. Notes and queries. Scientific American, 59(17):267, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n17/pdf/scientificamerican10271888-267a.pdf.

Anonymous:1888:IIbi

[10993] Anonymous. Index of inventions. Scientific American, 59(17): 267–268, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n17/pdf/scientificamerican10271888-
267c.pdf.

Anonymous:1888:PNPbi

[10994] Anonymous. Patent notes: Patents. Scientific American, 59 (17):270, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n17/pdf/scientificamerican10271888-
270.pdf.

Anonymous:1888:BC

[10995] Anonymous. Bungling chemists. Scientific American, 59(17):256, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n17/pdf/scientificamerican10271888-256.pdf.

Anonymous:1888:ESE

[10996] Anonymous. Electric subway explosions. Scientific American, 59 (17):256, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n17/pdf/scientificamerican10271888-
256a.pdf.

Anonymous:1888:PPN

[10997] Anonymous. Position of the planets in November. Scientific American, 59(17):256–257, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n17/pdf/scientificamerican10271888-
256b.pdf.

Anonymous:1888:SHF

[10998] Anonymous. A simple household freezer. Scientific American, 59(17):258, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n17/pdf/scientificamerican10271888-
258.pdf.

Anonymous:1888:IMFb

[10999] Anonymous. An improved mason’s float. Scientific American, 59(17):258, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n17/pdf/scientificamerican10271888-
258a.pdf.

Anonymous:1888:IWSf

[11000] Anonymous. An improved weather strip. Scientific American, 59(17):258, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n17/pdf/scientificamerican10271888-
258b.pdf.

Anonymous:1888:IRTb

[11001] Anonymous. An improved railroad tie. Scientific American, 59 (17):258, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n17/pdf/scientificamerican10271888-
258c.pdf.

Anonymous:1888:IBv

[11002] Anonymous. An improved binder. Scientific American, 59(17):258, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n17/pdf/scientificamerican10271888-258d.pdf.

Anonymous:1888:IBPb

[11003] Anonymous. An improved bake pan. Scientific American, 59 (17):258, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n17/pdf/scientificamerican10271888-
258e.pdf.

Anonymous:1888:LG

[11004] Anonymous. The Lebel gun. Scientific American, 59(17):259–260, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n17/pdf/scientificamerican10271888-259.pdf.

Anonymous:1888:IMMa

[11005] Anonymous. An improved method of moulding. Scientific American, 59(17):260, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n17/pdf/scientificamerican10271888-
260.pdf.

Anonymous:1888:IAL

[11006] Anonymous. An improved artificial leg. Scientific American, 59 (17):260, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n17/pdf/scientificamerican10271888-
260a.pdf.

Anonymous:1888:NGH

[11007] Anonymous. The new gas holder of the Consolidated Gas Company, of New York. Scientific American, 59(17):261, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n17/pdf/scientificamerican10271888-261.pdf.

Anonymous:1888:EHC

[11008] Anonymous. An elevator for hoisting carriages. Scientific American, 59(17):262, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n17/pdf/scientificamerican10271888-
262.pdf.

Anonymous:1888:BV

[11009] Anonymous. The butting viper. Scientific American, 59(17):262, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n17/pdf/scientificamerican10271888-262a.pdf.

Anonymous:1888:IGW

[11010] Anonymous. An improved gauge for woodworkers. Scientific American, 59(17):262–263, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n17/pdf/scientificamerican10271888-
262b.pdf.

Anonymous:1888:SHP

[11011] Anonymous. Sixteen horse power winding engine. Scientific American, 59(17):263, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n17/pdf/scientificamerican10271888-
263.pdf.

Anonymous:1888:IEE

[11012] Anonymous. An interesting electrical experiment. Scientific American, 59(17):264, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n17/pdf/scientificamerican10271888-
264.pdf.

Fetzes:1888:ICP

[11013] Chas Fetzes. Inexpensive carbon plates for Grenet batteries, etc. Scientific American, 59(17):264, October 27, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v59/
n17/pdf/scientificamerican10271888-264a.pdf.

Anonymous:1888:EM

[11014] Anonymous. Experiment with mirrors. Scientific American, 59 (17):264, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n17/pdf/scientificamerican10271888-
264b.pdf.

Anonymous:1888:EEF

[11015] Anonymous. Experimental electric fuse. Scientific American, 59 (17):264, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n17/pdf/scientificamerican10271888-
264c.pdf.

Anonymous:1888:ICMa

[11016] Anonymous. An improved calculating machine. Scientific American, 59(17):265, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n17/pdf/scientificamerican10271888-
265.pdf.

Anonymous:1888:EIbh

[11017] Anonymous. Engineering inventions. Scientific American, 59(17): 266, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n17/pdf/scientificamerican10271888-
266.pdf.

Anonymous:1888:AIw

[11018] Anonymous. Agricultural invention. Scientific American, 59(17): 266, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n17/pdf/scientificamerican10271888-
266a.pdf.

Anonymous:1888:MIbi

[11019] Anonymous. Miscellaneous inventions. Scientific American, 59 (17):266, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n17/pdf/scientificamerican10271888-
266b.pdf.

Anonymous:1888:Ibw

[11020] Anonymous. To inventors. Scientific American, 59(17):267, October 27, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n17/pdf/scientificamerican10271888-267b.pdf.

Anonymous:1888:Cbd

[11021] Anonymous. Correspondence. Scientific American, 59(18):280– 281, November 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n18/pdf/scientificamerican11031888-
280a.pdf.

Anonymous:1888:BPbk

[11022] Anonymous. Business and personal. Scientific American, 59 (18):282, November 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n18/pdf/scientificamerican11031888-
282c.pdf.

Anonymous:1888:NQbf

[11023] Anonymous. Notes and queries. Scientific American, 59(18): 282–283, November 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n18/pdf/scientificamerican11031888-
282d.pdf.

Anonymous:1888:BNBba

[11024] Anonymous. Books: New books and publications. Scientific American, 59(18):283, November 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n18/pdf/scientificamerican11031888-
283.pdf.

Anonymous:1888:IIbj

[11025] Anonymous. Index of inventions. Scientific American, 59(18): 283–284, November 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n18/pdf/scientificamerican11031888-
283b.pdf.

Anonymous:1888:PNPbj

[11026] Anonymous. Patent notes: Patents. Scientific American, 59 (18):286, November 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n18/pdf/scientificamerican11031888-
286.pdf.

Anonymous:1888:POF

[11027] Anonymous. The Patent Office the friend of the inventor. Scientific American, 59(18):272, November 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v59/n18/pdf/scientificamerican11031888-
272.pdf.

Anonymous:1888:LNT

[11028] Anonymous. Launch of a new type of ferry boat. Scientific American, 59(18):272–273, November 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n18/pdf/scientificamerican11031888-
272a.pdf.

Anonymous:1888:MAF

[11029] Anonymous. The Mackey automatic fire extinguisher. Scientific American, 59(18):274, November 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n18/pdf/scientificamerican11031888-
274.pdf.

Anonymous:1888:IMMb

[11030] Anonymous. An improved mechanical movement. Scientific American, 59(18):274, November 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n18/pdf/scientificamerican11031888-
274a.pdf.

Anonymous:1888:NMM

[11031] Anonymous. A new machine for making excelsior. Scientific American, 59(18):275, November 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n18/pdf/scientificamerican11031888-
275.pdf.

Anonymous:1888:ISHc

[11032] Anonymous. An improved scaffold holder. Scientific American, 59(18):275, November 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n18/pdf/scientificamerican11031888-
275a.pdf.

Anonymous:1888:IBHc

[11033] Anonymous. An improved book holder. Scientific American, 59 (18):275, November 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n18/pdf/scientificamerican11031888-
275b.pdf.

Anonymous:1888:ICCl

[11034] Anonymous. An improved car coupling. Scientific American, 59 (18):275–276, November 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n18/pdf/scientificamerican11031888-
275c.pdf.

Anonymous:1888:DCV

[11035] Anonymous. The dynamite cruiser Vesuvius. Scientific American, 59(18):277, November 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n18/pdf/scientificamerican11031888-
277.pdf.

Coryell:1888:FM

[11036] John R. Coryell. The Florida manatee. Scientific American, 59 (18):278–280, November 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n18/pdf/scientificamerican11031888-
278.pdf.

Anonymous:1888:BCR

[11037] Anonymous. Birmingham cable railway. Scientific American, 59 (18):280, November 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n18/pdf/scientificamerican11031888-
280.pdf.

Anonymous:1888:MNd

[11038] Anonymous. Microscopical notes. Scientific American, 59(18): 281, November 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n18/pdf/scientificamerican11031888-
281.pdf.

Anonymous:1888:EIbi

[11039] Anonymous. Engineering inventions. Scientific American, 59(18): 282, November 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n18/pdf/scientificamerican11031888-
282.pdf.

Anonymous:1888:AIx

[11040] Anonymous. Agricultural invention. Scientific American, 59(18): 282, November 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n18/pdf/scientificamerican11031888-
282a.pdf.

Anonymous:1888:MIbj

[11041] Anonymous. Miscellaneous inventions. Scientific American, 59 (18):282, November 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n18/pdf/scientificamerican11031888-
282b.pdf.

Anonymous:1888:Ibx

[11042] Anonymous. To inventors. Scientific American, 59(18):283, November 3, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n18/pdf/scientificamerican11031888-283a.pdf.

Anonymous:1888:Cbe

[11043] Anonymous. Correspondence. Scientific American, 59(19):293, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n19/pdf/scientificamerican11101888-293.pdf.

Anonymous:1888:BPbl

[11044] Anonymous. Business and personal. Scientific American, 59 (19):299, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n19/pdf/scientificamerican11101888-
299c.pdf.

Anonymous:1888:BNBbb

[11045] Anonymous. Books: New books and publications. Scientific American, 59(19):299, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n19/pdf/scientificamerican11101888-
299d.pdf.

Anonymous:1888:IIbk

[11046] Anonymous. Index of inventions. Scientific American, 59(19): 299–300, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n19/pdf/scientificamerican11101888-
299e.pdf.

Anonymous:1888:PNPbk

[11047] Anonymous. Patent notes: Patents. Scientific American, 59 (19):302, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n19/pdf/scientificamerican11101888-
302.pdf.

Anonymous:1888:NCPb

[11048] Anonymous. The Nebraska City Pontoon Bridge. Scientific American, 59(19):287–288, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n19/pdf/scientificamerican11101888-
287.pdf.

Anonymous:1888:FLL

[11049] Anonymous. Facts as to long life. Scientific American, 59(19): 288, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n19/pdf/scientificamerican11101888-
288.pdf.

Anonymous:1888:QLD

[11050] Anonymous. Quick letter delivery. Scientific American, 59(19): 288–289, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n19/pdf/scientificamerican11101888-
288a.pdf.

Anonymous:1888:CCa

[11051] Anonymous. Car coupling. Scientific American, 59(19):290, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n19/pdf/scientificamerican11101888-290.pdf.

Anonymous:1888:UDM

[11052] Anonymous. Unwin’s dabbing motion for noble combs. Scientific American, 59(19):290, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n19/pdf/scientificamerican11101888-
290a.pdf.

Anonymous:1888:ICF

[11053] Anonymous. An improved chair, fish plate, and rail coupler. Scientific American, 59(19):290, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v59/n19/pdf/scientificamerican11101888-
290b.pdf.

Anonymous:1888:ICWb

[11054] Anonymous. An improved car wheel and axle. Scientific American, 59(19):290, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n19/pdf/scientificamerican11101888-
290c.pdf.

Anonymous:1888:IVM

[11055] Anonymous. Improved ventilating man holes for sewers. Scientific American, 59(19):290, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v59/n19/pdf/scientificamerican11101888-
290d.pdf.

Anonymous:1888:DSS

[11056] Anonymous. A desk signal for schools. Scientific American, 59 (19):291, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n19/pdf/scientificamerican11101888-
291.pdf.

Anonymous:1888:MNe

[11057] Anonymous. Microscopical notes. Scientific American, 59(19): 291, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n19/pdf/scientificamerican11101888-
291a.pdf.

Anonymous:1888:IBSb

[11058] Anonymous. An improved bed stay. Scientific American, 59 (19):291, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n19/pdf/scientificamerican11101888-
291b.pdf.

Anonymous:1888:IGSb

[11059] Anonymous. An improved gig saddle. Scientific American, 59 (19):291, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n19/pdf/scientificamerican11101888-
291c.pdf.

Anonymous:1888:IJa

[11060] Anonymous. An improved jack. Scientific American, 59(19): 291–292, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n19/pdf/scientificamerican11101888-
291d.pdf.

Keeler:1888:RAW

[11061] James E. Keeler. Recent astronomical work at the Lick Observatory. Scientific American, 59(19):294–295, November 10, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v59/
n19/pdf/scientificamerican11101888-294.pdf.

Anonymous:1888:Cbf

[11062] Anonymous. Centenarians. Scientific American, 59(19):295, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n19/pdf/scientificamerican11101888-295.pdf.

Anonymous:1888:CCb

[11063] Anonymous. A Connecticut centenarian. Scientific American, 59 (19):295–297, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n19/pdf/scientificamerican11101888-
295a.pdf.

Anonymous:1888:IABb

[11064] Anonymous. An improved attachment for bicycles. Scientific American, 59(19):297, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n19/pdf/scientificamerican11101888-
297.pdf.

Anonymous:1888:CSM

[11065] Anonymous. Cut-off sawing machine. Scientific American, 59(19): 297–298, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n19/pdf/scientificamerican11101888-
297a.pdf.

Anonymous:1888:EIbj

[11066] Anonymous. Engineering inventions. Scientific American, 59(19): 299, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n19/pdf/scientificamerican11101888-
299.pdf.

Anonymous:1888:AIy

[11067] Anonymous. Agricultural invention. Scientific American, 59(19): 299, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n19/pdf/scientificamerican11101888-
299a.pdf.

Anonymous:1888:MIbk

[11068] Anonymous. Miscellaneous inventions. Scientific American, 59 (19):299, November 10, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n19/pdf/scientificamerican11101888-
299b.pdf.

Anonymous:1888:BPbm

[11069] Anonymous. Business and personal. Scientific American, 59(2):27, July 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n2/pdf/scientificamerican07141888-27b.pdf.

Anonymous:1888:BNBbc

[11070] Anonymous. Books: New books and publications. Scientific American, 59(2):27, July 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n2/pdf/scientificamerican07141888-
27c.pdf.

Anonymous:1888:IIbl

[11071] Anonymous. Index of inventions. Scientific American, 59(2):27–28, July 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n2/pdf/scientificamerican07141888-27e.pdf.

Anonymous:1888:PNPbl

[11072] Anonymous. Patent notes: Patents. Scientific American, 59(2):30, July 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n2/pdf/scientificamerican07141888-30.pdf.

Anonymous:1888:RAS

[11073] Anonymous. Reproduction of articulate speech and other sounds. Scientific American, 59(2):15, July 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n2/pdf/scientificamerican07141888-
15.pdf.

Anonymous:1888:CWe

[11074] Anonymous. The celestial world. Scientific American, 59(2):16, July 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n2/pdf/scientificamerican07141888-16.pdf.

Anonymous:1888:MNf

[11075] Anonymous. Military notes. Scientific American, 59(2):16–17, July 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n2/pdf/scientificamerican07141888-16a.pdf.

Anonymous:1888:ISLd

[11076] Anonymous. An improved ship’s log. Scientific American, 59(2):18, July 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n2/pdf/scientificamerican07141888-18.pdf.

Anonymous:1888:DSL

[11077] Anonymous. A device for stopping leaks in vessels. Scientific American, 59(2):18, July 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n2/pdf/scientificamerican07141888-
18a.pdf.

Anonymous:1888:IEAa

[11078] Anonymous. An improved electrical annunciator. Scientific American, 59(2):18, July 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n2/pdf/scientificamerican07141888-
18b.pdf.

Anonymous:1888:ICO

[11079] Anonymous. An improved can opener. Scientific American, 59(2):18, July 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n2/pdf/scientificamerican07141888-18c.pdf.

Anonymous:1888:CHC

[11080] Anonymous. A clamp for holding covers on fruit cans. Scientific American, 59(2):18, July 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n2/pdf/scientificamerican07141888-
18d.pdf.

Anonymous:1888:ICCm

[11081] Anonymous. An improved car coupling. Scientific American, 59(2):19, July 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n2/pdf/scientificamerican07141888-19.pdf.

Anonymous:1888:DHB

[11082] Anonymous. A device for holding bed clothes in position. Scientific American, 59(2):19, July 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n2/pdf/scientificamerican07141888-
19a.pdf.

Anonymous:1888:RHC

[11083] Anonymous. A removable handle for cross-cut saws. Scientific American, 59(2):19, July 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n2/pdf/scientificamerican07141888-
19b.pdf.

Anonymous:1888:ITPc

[11084] Anonymous. An improved tobacco pipe. Scientific American, 59(2):19, July 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n2/pdf/scientificamerican07141888-19c.pdf.

Anonymous:1888:IDT

[11085] Anonymous. An improved drilling tool. Scientific American, 59(2):19, July 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n2/pdf/scientificamerican07141888-19d.pdf.

Anonymous:1888:CNA

[11086] Anonymous. Corea by native artists. Scientific American, 59(2):20–22, July 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n2/pdf/scientificamerican07141888-20.pdf.

Anonymous:1888:MFN

[11087] Anonymous. The manufacture of fishing nets. Scientific American, 59(2):22, July 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n2/pdf/scientificamerican07141888-
22.pdf.

Anonymous:1888:NSP

[11088] Anonymous. The National system of purifying water by filtration, precipitation, and aeration. Scientific American, 59(2):23–24, July 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n2/pdf/scientificamerican07141888-23.pdf.

Hopkins:1888:SEPb

[11089] M. Hopkins. Simple experiments in physics. Scientific American, 59(2):25–26, July 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n2/pdf/scientificamerican07141888-
25.pdf.

Anonymous:1888:EIbk

[11090] Anonymous. Engineering inventions. Scientific American, 59(2):27, July 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n2/pdf/scientificamerican07141888-27.pdf.

Anonymous:1888:MIbl

[11091] Anonymous. Miscellaneous inventions. Scientific American, 59(2):27, July 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n2/pdf/scientificamerican07141888-27a.pdf.

Anonymous:1888:Iby

[11092] Anonymous. To inventors. Scientific American, 59(2):27, July 14, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n2/pdf/scientificamerican07141888-27d.pdf.

Anonymous:1888:Cbg

[11093] Anonymous. Correspondence. Scientific American, 59(20):309, November 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n20/pdf/scientificamerican11171888-309.pdf.

Anonymous:1888:BPbn

[11094] Anonymous. Business and personal. Scientific American, 59 (20):315, November 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n20/pdf/scientificamerican11171888-
315b.pdf.

Anonymous:1888:IIbm

[11095] Anonymous. Index of inventions. Scientific American, 59(20): 315–316, November 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n20/pdf/scientificamerican11171888-
315d.pdf.

Anonymous:1888:PNPbm

[11096] Anonymous. Patent notes: Patents. Scientific American, 59 (20):318, November 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n20/pdf/scientificamerican11171888-
318.pdf.

Anonymous:1888:PE

[11097] Anonymous. The Paris Exposition of 1889. Scientific American, 59(20):304, November 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n20/pdf/scientificamerican11171888-
304.pdf.

Anonymous:1888:FM

[11098] Anonymous. The football malady. Scientific American, 59(20): 304, November 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n20/pdf/scientificamerican11171888-
304a.pdf.

Anonymous:1888:MD

[11099] Anonymous. “Mosquito defense”. Scientific American, 59(20): 304–305, November 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n20/pdf/scientificamerican11171888-
304b.pdf.

Anonymous:1888:PNj

[11100] Anonymous. Photographic notes. Scientific American, 59(20): 305, November 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n20/pdf/scientificamerican11171888-
305.pdf.

Anonymous:1888:IGe

[11101] Anonymous. An improved gate. Scientific American, 59(20): 306, November 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n20/pdf/scientificamerican11171888-
306.pdf.

Anonymous:1888:ICBb

[11102] Anonymous. Improved curtain bracket and sash lock. Scientific American, 59(20):306, November 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n20/pdf/scientificamerican11171888-
306a.pdf.

Anonymous:1888:ICVb

[11103] Anonymous. An improved cleat for vessels. Scientific American, 59(20):306, November 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n20/pdf/scientificamerican11171888-
306b.pdf.

Anonymous:1888:IHTb

[11104] Anonymous. An improved hand truck. Scientific American, 59 (20):306, November 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n20/pdf/scientificamerican11171888-
306c.pdf.

Anonymous:1888:ISBd

[11105] Anonymous. An improved sled brake. Scientific American, 59 (20):306, November 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n20/pdf/scientificamerican11171888-
306d.pdf.

Anonymous:1888:EOS

[11106] Anonymous. The East Orange Sewage System and Water Works. Scientific American, 59(20):307–308, November 17, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v59/
n20/pdf/scientificamerican11171888-307.pdf.

Anonymous:1888:IFh

[11107] Anonymous. An improved fence. Scientific American, 59(20): 308, November 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n20/pdf/scientificamerican11171888-
308.pdf.

Anonymous:1888:INT

[11108] Anonymous. An improved needle and twine cutter. Scientific American, 59(20):308, November 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n20/pdf/scientificamerican11171888-
308a.pdf.

Anonymous:1888:IHc

[11109] Anonymous. An improved harrow. Scientific American, 59(20): 308, November 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n20/pdf/scientificamerican11171888-
308b.pdf.

Anonymous:1888:BEM

[11110] Anonymous. The Baxter electric motor. Scientific American, 59 (20):310, November 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n20/pdf/scientificamerican11171888-
310.pdf.

Anonymous:1888:ILBd

[11111] Anonymous. An improvement in locomotive boilers. Scientific American, 59(20):310, November 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n20/pdf/scientificamerican11171888-
310a.pdf.

Anonymous:1888:IFEd

[11112] Anonymous. Improved fortable electric light apparatus. Scientific American, 59(20):310–312, November 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v59/n20/pdf/scientificamerican11171888-
310b.pdf.

Anonymous:1888:ICAc

[11113] Anonymous. An improved centering apparatus. Scientific American, 59(20):312, November 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n20/pdf/scientificamerican11171888-
312.pdf.

Anonymous:1888:SSA

[11114] Anonymous. The Sing Sing antelope and the Gorgon antelope. Scientific American, 59(20):313–314, November 17, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v59/
n20/pdf/scientificamerican11171888-313.pdf.

Anonymous:1888:EIbl (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n20/pdf/scientificamerican11171888-
315.pdf.

Anonymous:1888:MIbm

[11116] Anonymous. Miscellaneous inventions. Scientific American, 59 (20):315, November 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n20/pdf/scientificamerican11171888-
315a.pdf.

Anonymous:1888:Ibz

[11117] Anonymous. To inventors. Scientific American, 59(20):315, November 17, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n20/pdf/scientificamerican11171888-315c.pdf.

Anonymous:1888:Cbh

[11118] Anonymous. Correspondence. Scientific American, 59(21):325, November 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n21/pdf/scientificamerican11241988-325.pdf.

Anonymous:1888:BPbo

[11119] Anonymous. Business and personal. Scientific American, 59 (21):330, November 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n21/pdf/scientificamerican11241988-
330c.pdf.

Anonymous:1888:BNBbd

[11120] Anonymous. Books: New books and publications. Scientific American, 59(21):330, November 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n21/pdf/scientificamerican11241988-
330d.pdf.

Anonymous:1888:NQbg

[11121] Anonymous. Notes and queries. Scientific American, 59(21): 330–331, November 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n21/pdf/scientificamerican11241988-
330e.pdf.

Anonymous:1888:IIbn

[11122] Anonymous. Index of inventions. Scientific American, 59(21): 331–332, November 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n21/pdf/scientificamerican11241988-
331.pdf.

Anonymous:1888:PNPbn

[11123] Anonymous. Patent notes: Patents. Scientific American, 59 (21):334, November 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n21/pdf/scientificamerican11241988-
334.pdf.

Anonymous:1888:RMD

[11124] Anonymous. A remarkable marine disaster. Scientific American, 59(21):319, November 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n21/pdf/scientificamerican11241988-
319.pdf.

Anonymous:1888:ACB

[11125] Anonymous. Another collision between ocean steamers. Scientific American, 59(21):320, November 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v59/n21/pdf/scientificamerican11241988-
320.pdf.

Anonymous:1888:TOF

[11126] Anonymous. Torpedo and other fast ships. Scientific American, 59 (21):320–321, November 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n21/pdf/scientificamerican11241988-
320a.pdf.

Anonymous:1888:IWH

[11127] Anonymous. An improved water heater. Scientific American, 59 (21):322, November 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n21/pdf/scientificamerican11241988-
322.pdf.

Anonymous:1888:IRDc

[11128] Anonymous. An improved rock drill. Scientific American, 59 (21):322, November 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n21/pdf/scientificamerican11241988-
322a.pdf.

Anonymous:1888:ITBa

[11129] Anonymous. An improved trunk brace. Scientific American, 59 (21):322, November 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n21/pdf/scientificamerican11241988-
322b.pdf.

Anonymous:1888:IAD

[11130] Anonymous. An improved advertising device. Scientific American, 59(21):322, November 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n21/pdf/scientificamerican11241988-
322c.pdf.

Anonymous:1888:IPWb

[11131] Anonymous. An improved paddle wheel. Scientific American, 59 (21):322, November 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n21/pdf/scientificamerican11241988-
322d.pdf.

Hopkins:1888:SEPc

[11132] M. Hopkins. Simple experiments in physics. Scientific American, 59(21):323, November 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n21/pdf/scientificamerican11241988-
323.pdf.

Anonymous:1888:EPS

[11133] Anonymous. Explosion of a petroleum steamer at Calais, France. Scientific American, 59(21):323–324, November 24, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v59/
n21/pdf/scientificamerican11241988-323a.pdf.

Anonymous:1888:PNk

[11134] Anonymous. Photographic notes. Scientific American, 59(21): 325–326, November 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n21/pdf/scientificamerican11241988-
325a.pdf.

Anonymous:1888:IFS

[11135] Anonymous. Improved floating scouring dam. Scientific American, 59 (21):326–328, November 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n21/pdf/scientificamerican11241988-
326.pdf.

Anonymous:1888:USB

[11136] Anonymous. A urus skeleton (bos primigenius). Scientific American, 59(21):329, November 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n21/pdf/scientificamerican11241988-
329.pdf.

Anonymous:1888:EIbm

[11137] Anonymous. Engineering inventions. Scientific American, 59(21): 330, November 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n21/pdf/scientificamerican11241988-
330.pdf.

Anonymous:1888:AIz

[11138] Anonymous. Agricultural inventions. Scientific American, 59 (21):330, November 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n21/pdf/scientificamerican11241988-
330a.pdf.

Anonymous:1888:MIbn

[11139] Anonymous. Miscellaneous inventions. Scientific American, 59 (21):330, November 24, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n21/pdf/scientificamerican11241988-
330b.pdf.

Anonymous:1888:Cbi

[11140] Anonymous. Correspondence. Scientific American, 59(22):341, December 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n22/pdf/scientificamerican12011888-341a.pdf.

Anonymous:1888:BPbp

[11141] Anonymous. Business and personal. Scientific American, 59(22): 345–346, December 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n22/pdf/scientificamerican12011888-
345b.pdf.

Anonymous:1888:NQbh

[11142] Anonymous. Notes and queries. Scientific American, 59(22): 346–347, December 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n22/pdf/scientificamerican12011888-
346.pdf.

Anonymous:1888:IIbo

[11143] Anonymous. Index of inventions. Scientific American, 59(22): 347–348, December 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n22/pdf/scientificamerican12011888-
347a.pdf.

Anonymous:1888:PNPbo

[11144] Anonymous. Patent notes: Patents. Scientific American, 59 (22):350, December 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n22/pdf/scientificamerican12011888-
350.pdf.

Anonymous:1888:TCE

[11145] Anonymous. Tall chimney engineering. Scientific American, 59 (22):336, December 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n22/pdf/scientificamerican12011888-
336.pdf.

Anonymous:1888:TE

[11146] Anonymous. Theatrical emotion. Scientific American, 59(22):336, December 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n22/pdf/scientificamerican12011888-336a.pdf.

Anonymous:1888:NAB

[11147] Anonymous. New animals and birds in the Central Park. Scientific American, 59(22):336–337, December 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v59/n22/pdf/scientificamerican12011888-
336b.pdf.

Anonymous:1888:PPD

[11148] Anonymous. Position of the plants in December. Scientific American, 59(22):337, December 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n22/pdf/scientificamerican12011888-
337.pdf.

Anonymous:1888:PNl

[11149] Anonymous. Photographic notes. Scientific American, 59(22):337, December 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n22/pdf/scientificamerican12011888-337a.pdf.

Anonymous:1888:ACV

[11150] Anonymous. An air compressor for variable pressures. Scientific American, 59(22):338, December 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n22/pdf/scientificamerican12011888-
338.pdf.

Anonymous:1888:ICSd

[11151] Anonymous. As improved car stop. Scientific American, 59(22):339, December 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n22/pdf/scientificamerican12011888-339.pdf.

Anonymous:1888:ISCb

[11152] Anonymous. Improved straw carrier for thrashing machines. Scientific American, 59(22):339, December 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v59/n22/pdf/scientificamerican12011888-
339a.pdf.

Anonymous:1888:IFEe

[11153] Anonymous. An improved fire escape. Scientific American, 59 (22):339, December 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n22/pdf/scientificamerican12011888-
339b.pdf.

Anonymous:1888:SCD

[11154] Anonymous. A simple churn dash. Scientific American, 59(22): 339–340, December 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n22/pdf/scientificamerican12011888-
339c.pdf.

Anonymous:1888:ISPb

[11155] Anonymous. An improved smoking pipe. Scientific American, 59 (22):340, December 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n22/pdf/scientificamerican12011888-
340.pdf.

Anonymous:1888:MT

[11156] Anonymous. The Monitor terror. Scientific American, 59(22):341, December 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n22/pdf/scientificamerican12011888-341.pdf.

Anonymous:1888:TEP

[11157] Anonymous. The Thorenberg electric plant. Scientific American, 59 (22):342–343, December 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n22/pdf/scientificamerican12011888-
342.pdf.

Anonymous:1888:GS

[11158] Anonymous. A great stone. Scientific American, 59(22):343, December 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n22/pdf/scientificamerican12011888-343.pdf.

Anonymous:1888:COI

[11159] Anonymous. Curious optical illusion. Scientific American, 59(22): 343–344, December 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n22/pdf/scientificamerican12011888-
343a.pdf.

Anonymous:1888:EIbn

[11160] Anonymous. Engineering inventions. Scientific American, 59(22): 345, December 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n22/pdf/scientificamerican12011888-
345.pdf.

Anonymous:1888:MIbo

[11161] Anonymous. Miscellaneous inventions. Scientific American, 59 (22):345, December 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n22/pdf/scientificamerican12011888-
345a.pdf.

Anonymous:1888:Ici

[11162] Anonymous. To inventors. Scientific American, 59(22):347, December 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n22/pdf/scientificamerican12011888-347.pdf.

Anonymous:1888:RPIa

[11163] Anonymous. Recently patented inventions. Scientific American, 59(23):362, December 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n23/pdf/scientificamerican12081888-
362.pdf.

Anonymous:1888:BPbq

[11164] Anonymous. Business and personal. Scientific American, 59 (23):362, December 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n23/pdf/scientificamerican12081888-
362a.pdf.

Anonymous:1888:NQbi

[11165] Anonymous. Notes and queries. Scientific American, 59(23): 362–363, December 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n23/pdf/scientificamerican12081888-
362b.pdf.

Anonymous:1888:IIbp

[11166] Anonymous. Index of inventions. Scientific American, 59(23): 363–364, December 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n23/pdf/scientificamerican12081888-
363a.pdf.

Anonymous:1888:PNPbp

[11167] Anonymous. Patent notes: Patents. Scientific American, 59 (23):366, December 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n23/pdf/scientificamerican12081888-
366.pdf.

Anonymous:1888:SFB

[11168] Anonymous. The screw ferry boat Bergen. Scientific American, 59 (23):351–352, December 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n23/pdf/scientificamerican12081888-
351.pdf.

Anonymous:1888:EPL

[11169] Anonymous. Electricity and plant life. Scientific American, 59 (23):352, December 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n23/pdf/scientificamerican12081888-
352.pdf.

Anonymous:1888:SSW

[11170] Anonymous. “Scenes from a silent world”. Scientific American, 59 (23):352–353, December 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n23/pdf/scientificamerican12081888-
352a.pdf.

Anonymous:1888:CWf

[11171] Anonymous. The celestial world. Scientific American, 59(23):353, December 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n23/pdf/scientificamerican12081888-353.pdf.

Anonymous:1888:CPS

[11172] Anonymous. A combined punching and shearing machine. Scientific American, 59(23):354, December 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v59/n23/pdf/scientificamerican12081888-
354.pdf.

Anonymous:1888:ISMb

[11173] Anonymous. Improved slotting machine. Scientific American, 59 (23):354, December 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n23/pdf/scientificamerican12081888-
354a.pdf.

Anonymous:1888:ICPb

[11174] Anonymous. Insulated coverings for pipes, boilers, etc. Scientific American, 59(23):355, December 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n23/pdf/scientificamerican12081888-
355.pdf.

Anonymous:1888:ISIb

[11175] Anonymous. An improved speed indicator. Scientific American, 59(23):355, December 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n23/pdf/scientificamerican12081888-
355a.pdf.

Anonymous:1888:IFCe

[11176] Anonymous. An improved flywheel clutching device. Scientific American, 59(23):355, December 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n23/pdf/scientificamerican12081888-
355b.pdf.

Anonymous:1888:IGSc

[11177] Anonymous. An improved gas stove. Scientific American, 59(23): 355–357, December 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n23/pdf/scientificamerican12081888-
355c.pdf.

Anonymous:1888:NFS

[11178] Anonymous. The new French system of movable batteries. Scientific American, 59(23):358, December 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v59/n23/pdf/scientificamerican12081888-
358.pdf.

Hopkins:1888:SEPd

[11179] M. Hopkins. Simple experiments in physics. Scientific American, 59 (23):359–360, December 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n23/pdf/scientificamerican12081888-
359.pdf.

Anonymous:1888:RHP

[11180] Anonymous. The rock hopper penguin. Scientific American, 59 (23):361, December 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n23/pdf/scientificamerican12081888-
361.pdf.

Anonymous:1888:RER

[11181] Anonymous. Running elevated railroad trains by electricity. Scientific American, 59(23):361, December 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v59/n23/pdf/scientificamerican12081888-
361a.pdf.

Anonymous:1888:Icj

[11182] Anonymous. To inventors. Scientific American, 59(23):363, December 8, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n23/pdf/scientificamerican12081888-363.pdf.

Anonymous:1888:Cbj

[11183] Anonymous. Correspondence. Scientific American, 59(24):376, December 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n24/pdf/scientificamerican12151888-376.pdf.

Anonymous:1888:RPIb

[11184] Anonymous. Recently patented inventions: Recently patented intentions. Scientific American, 59(24):378, December 15, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v59/
n24/pdf/scientificamerican12151888-378.pdf.

Anonymous:1888:BPbr

[11185] Anonymous. Business and personal. Scientific American, 59 (24):378, December 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n24/pdf/scientificamerican12151888-
378a.pdf.

Anonymous:1888:BNBbe

[11186] Anonymous. Books: New books and publications. Scientific American, 59(24):378, December 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n24/pdf/scientificamerican12151888-
378b.pdf.

Anonymous:1888:NQbj

[11187] Anonymous. Notes and queries. Scientific American, 59(24): 378–379, December 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n24/pdf/scientificamerican12151888-
378c.pdf.

Anonymous:1888:IIbq

[11188] Anonymous. Index of inventions. Scientific American, 59(24): 379–380, December 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n24/pdf/scientificamerican12151888-
379a.pdf.

Anonymous:1888:PNPbq

[11189] Anonymous. Patent notes: Patents. Scientific American, 59 (24):382, December 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n24/pdf/scientificamerican12151888-
382.pdf.

Anonymous:1888:EDEa

[11190] Anonymous. Experiments on death by electricity. Scientific American, 59(24):368, December 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n24/pdf/scientificamerican12151888-
368.pdf.

Anonymous:1888:WHP

[11191] Anonymous. The Webster–Higgins Patent suit. Scientific American, 59(24):368–369, December 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n24/pdf/scientificamerican12151888-
368a.pdf.

Anonymous:1888:IFCf

[11192] Anonymous. An improved freight car. Scientific American, 59 (24):370, December 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n24/pdf/scientificamerican12151888-
370.pdf.

Anonymous:1888:SHB

[11193] Anonymous. A superficial heater for boilers. Scientific American, 59(24):370, December 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n24/pdf/scientificamerican12151888-
370a.pdf.

Anonymous:1888:IPT

[11194] Anonymous. An improved power treadle attachment. Scientific American, 59(24):370, December 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n24/pdf/scientificamerican12151888-
370b.pdf.

Anonymous:1888:ITWb

[11195] Anonymous. An improved toothed wheel. Scientific American, 59 (24):370, December 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n24/pdf/scientificamerican12151888-
370c.pdf.

Anonymous:1888:IBBe

[11196] Anonymous. An improved brace for bedsteads. Scientific American, 59(24):371, December 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n24/pdf/scientificamerican12151888-
371.pdf.

Anonymous:1888:NSS

[11197] Anonymous. The new Spanish steamer Alphonso XII. Scientific American, 59(24):371, December 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n24/pdf/scientificamerican12151888-
371a.pdf.

Anonymous:1888:ISBe

[11198] Anonymous. An improved scaffold bracket. Scientific American, 59 (24):371–373, December 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n24/pdf/scientificamerican12151888-
371b.pdf.

Anonymous:1888:FFP

[11199] Anonymous. Forster’s finger pen and pencil holder. Scientific American, 59(24):373, December 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n24/pdf/scientificamerican12151888-
373.pdf.

Anonymous:1888:IFRb

[11200] Anonymous. An improved fishing reel. Scientific American, 59 (24):373, December 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n24/pdf/scientificamerican12151888-
373a.pdf.

Anonymous:1888:RT

[11201] Anonymous. The Remington typewriter. Scientific American, 59 (24):374–375, December 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n24/pdf/scientificamerican12151888-
374.pdf.

Anonymous:1888:NTBb

[11202] Anonymous. A night trip in a balloon without a car. Scientific American, 59(24):375–376, December 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v59/n24/pdf/scientificamerican12151888-
375.pdf.

Hopkins:1888:SEPe

[11203] M. Hopkins. Simple experiments in physics. Scientific American, 59(24):377, December 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n24/pdf/scientificamerican12151888-
377.pdf.

Anonymous:1888:Ick

[11204] Anonymous. To inventors. Scientific American, 59(24):379, December 15, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n24/pdf/scientificamerican12151888-379.pdf.

Anonymous:1888:Cbk

[11205] Anonymous. Correspondence. Scientific American, 59(25):389, December 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n25/pdf/scientificamerican12221888-389.pdf.

Anonymous:1888:RPIc

[11206] Anonymous. Recently patented inventions. Scientific American, 59(25):394, December 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n25/pdf/scientificamerican12221888-
394.pdf.

Anonymous:1888:BPbs 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n25/pdf/scientificamerican12221888-
394a.pdf.

Anonymous:1888:BNBbf

[11208] Anonymous. Books: New books and publications. Scientific American, 59(25):394, December 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n25/pdf/scientificamerican12221888-
394b.pdf.

Anonymous:1888:NQbk

[11209] Anonymous. Notes and queries. Scientific American, 59(25): 395, December 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n25/pdf/scientificamerican12221888-
395.pdf.

Anonymous:1888:IIbr

[11210] Anonymous. Index of inventions. Scientific American, 59(25): 395–396, December 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n25/pdf/scientificamerican12221888-
395b.pdf.

Anonymous:1888:PNPbr

[11211] Anonymous. Patent notes: Patents. Scientific American, 59 (25):398, December 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n25/pdf/scientificamerican12221888-
398.pdf.

Anonymous:1888:AGS

[11212] Anonymous. The American Geological Society. Scientific American, 59(25):384, December 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n25/pdf/scientificamerican12221888-
384.pdf.

Anonymous:1888:MM

[11213] Anonymous. Madness and murder. Scientific American, 59(25): 384, December 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n25/pdf/scientificamerican12221888-
384a.pdf.

Anonymous:1888:MTF

[11214] Anonymous. Milk and typhoid fever. Scientific American, 59(25): 384–385, December 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n25/pdf/scientificamerican12221888-
384b.pdf.

Anonymous:1888:TEJ

[11215] Anonymous. The total eclipse of January 1, 1889. Scientific American, 59(25):385, December 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n25/pdf/scientificamerican12221888-
385.pdf.

Hopkins:1888:SEPf

[11216] M. Hopkins. Simple experiments in physics. Scientific American, 59(25):386, December 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n25/pdf/scientificamerican12221888-
386.pdf.

Anonymous:1888:ICCn

[11217] Anonymous. An improved car coupling. Scientific American, 59 (25):386, December 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n25/pdf/scientificamerican12221888-
386a.pdf.

Anonymous:1888:ISV

[11218] Anonymous. An improved safe or vault door lock. Scientific American, 59(25):386–387, December 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n25/pdf/scientificamerican12221888-
386b.pdf.

Anonymous:1888:HM

[11219] Anonymous. Homes for the many. Scientific American, 59(25): 387, December 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n25/pdf/scientificamerican12221888-
387.pdf.

Anonymous:1888:DUSb

[11220] Anonymous. Decisions of the United States courts. Scientific American, 59(25):388, December 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n25/pdf/scientificamerican12221888-
388.pdf.

Anonymous:1888:IBTc

[11221] Anonymous. An improved bottom-trip sling. Scientific American, 59 (25):388–389, December 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n25/pdf/scientificamerican12221888-
388a.pdf.

Anonymous:1888:SCS

[11222] Anonymous. Self-charging siphon. Scientific American, 59(25): 389, December 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n25/pdf/scientificamerican12221888-
389a.pdf.

Anonymous:1888:EAD

[11223] Anonymous. Electrical aids to the drama. Scientific American, 59 (25):390–391, December 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n25/pdf/scientificamerican12221888-
390.pdf.

Hovey:1888:RHE

[11224] H. C. Hovey. Royal E. House, the electrician. Scientific American, 59(25):391–392, December 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n25/pdf/scientificamerican12221888-
391.pdf.

Anonymous:1888:DTG

[11225] Anonymous. Diverging torpedo guns. Scientific American, 59 (25):392, December 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n25/pdf/scientificamerican12221888-
392.pdf.

Anonymous:1888:EDEb

[11226] Anonymous. Experiments on death by electricity. Scientific American, 59(25):393, December 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n25/pdf/scientificamerican12221888-
393.pdf.

Anonymous:1888:Icn

[11227] Anonymous. To inventors. Scientific American, 59(25):395, December 22, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n25/pdf/scientificamerican12221888-395a.pdf.

Anonymous:1888:RPId

[11228] Anonymous. Recently patented inventions. Scientific American, 59(26):408, December 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n26/pdf/scientificamerican12291888-
408.pdf.

Anonymous:1888:BPbt

[11229] Anonymous. Business and personal. Scientific American, 59 (26):408, December 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n26/pdf/scientificamerican12291888-
408a.pdf.

Anonymous:1888:BNBbg

[11230] Anonymous. Books: New books and publications. Scientific American, 59(26):408, December 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n26/pdf/scientificamerican12291888-
408b.pdf.

Anonymous:1888:NQbl

[11231] Anonymous. Notes and queries. Scientific American, 59(26): 408–409, December 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n26/pdf/scientificamerican12291888-
408c.pdf.

Anonymous:1888:IIbs

[11232] Anonymous. Index of inventions. Scientific American, 59(26): 409–412, December 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n26/pdf/scientificamerican12291888-
409a.pdf.

Anonymous:1888:PNPbs

[11233] Anonymous. Patent notes: Patents. Scientific American, 59 (26):414, December 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n26/pdf/scientificamerican12291888-
414.pdf.

Anonymous:1888:NDG

[11234] Anonymous. New dynamite gun. Scientific American, 59(26): 399–400, December 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n26/pdf/scientificamerican12291888-
399.pdf.

Anonymous:1888:DLT

[11235] Anonymous. Dogs that learn trades. Scientific American, 59 (26):400, December 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n26/pdf/scientificamerican12291888-
400.pdf.

Anonymous:1888:PRM

[11236] Anonymous. “A possible revolution in medicine”. Scientific American, 59(26):400–401, December 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n26/pdf/scientificamerican12291888-
400a.pdf.

Hopkins:1888:SEPg

[11237] M. Hopkins. Simple experiments in physics. Scientific American, 59(26):402, December 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n26/pdf/scientificamerican12291888-
402.pdf.

Anonymous:1888:ICq

[11238] Anonymous. An improved churn. Scientific American, 59(26): 402, December 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n26/pdf/scientificamerican12291888-
402a.pdf.

Anonymous:1888:IEL

[11239] Anonymous. An improved extension ladder. Scientific American, 59 (26):402–403, December 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n26/pdf/scientificamerican12291888-
402b.pdf.

Anonymous:1888:OMS

[11240] Anonymous. Offset mechanism for sawmill carriages. Scientific American, 59(26):403, December 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n26/pdf/scientificamerican12291888-
403.pdf.

Anonymous:1888:ISSd

[11241] Anonymous. An improved steam steering-gear. Scientific American, 59(26):403, December 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n26/pdf/scientificamerican12291888-
403a.pdf.

Anonymous:1888:GV

[11242] Anonymous. The Garabit Viaduct. Scientific American, 59(26): 403–405, December 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n26/pdf/scientificamerican12291888-
403b.pdf.

Anonymous:1888:SDT

[11243] Anonymous. A simple device for threading needles. Scientific American, 59(26):405, December 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n26/pdf/scientificamerican12291888-
405.pdf.

Anonymous:1888:TCS

[11244] Anonymous. Test of the cast steel breech loading rifle. Scientific American, 59(26):406–407, December 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v59/n26/pdf/scientificamerican12291888-
406.pdf.

Anonymous:1888:GFM

[11245] Anonymous. Gigantic fossil mammals. Scientific American, 59 (26):407, December 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n26/pdf/scientificamerican12291888-
407.pdf.

Anonymous:1888:Ict

[11246] Anonymous. To inventors. Scientific American, 59(26):409, December 29, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n26/pdf/scientificamerican12291888-409.pdf.

Anonymous:1888:Cbl

[11247] Anonymous. Correspondence. Scientific American, 59(3):37, July 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n3/pdf/scientificamerican07211888-37.pdf.

Anonymous:1888:BPbu

[11248] Anonymous. Business and personal. Scientific American, 59(3):42, July 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n3/pdf/scientificamerican07211888-42c.pdf.

Anonymous:1888:BNBbh

[11249] Anonymous. Books: New books and publications. Scientific American, 59(3):42, July 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n3/pdf/scientificamerican07211888-
42d.pdf.

Anonymous:1888:NQbm

[11250] Anonymous. Notes and queries. Scientific American, 59(3):42–43, July 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n3/pdf/scientificamerican07211888-42e.pdf.

Anonymous:1888:IIbt

[11251] Anonymous. Index of inventions. Scientific American, 59(3):43–44, July 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n3/pdf/scientificamerican07211888-43a.pdf.

Anonymous:1888:PNPbt

[11252] Anonymous. Patent notes: Patents. Scientific American, 59(3):46, July 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n3/pdf/scientificamerican07211888-46.pdf.

Anonymous:1888:ICRe

[11253] Anonymous. Improvements in the cable railway of the New York and Brooklyn Bridge. Scientific American, 59(3):31, July 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n3/pdf/scientificamerican07211888-31.pdf.

Anonymous:1888:SSM

[11254] Anonymous. Salida springs and mines, zirconia minerals, and more. Scientific American, 59(3):32, July 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v59/n3/pdf/scientificamerican07211888-
32.pdf.

Anonymous:1888:HS

[11255] Anonymous. Hiram Sibley. Scientific American, 59(3):33, July 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n3/pdf/scientificamerican07211888-33.pdf.

Anonymous:1888:PTB

[11256] Anonymous. The projected tunnel between France and England. Scientific American, 59(3):34, July 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n3/pdf/scientificamerican07211888-
34.pdf.

Anonymous:1888:SSCb

[11257] Anonymous. A safety seal for checks, bonds, etc. Scientific American, 59(3):35, July 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n3/pdf/scientificamerican07211888-
35.pdf.

Anonymous:1888:INLe

[11258] Anonymous. An improved nut lock for valves. Scientific American, 59(3):35, July 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n3/pdf/scientificamerican07211888-
35a.pdf.

Anonymous:1888:IDI

[11259] Anonymous. An improved drawing instrument. Scientific American, 59(3):35, July 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n3/pdf/scientificamerican07211888-
35b.pdf.

Anonymous:1888:WLC

[11260] Anonymous. A waterproof lining for cellars. Scientific American, 59(3):35–36, July 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n3/pdf/scientificamerican07211888-
35c.pdf.

Anonymous:1888:IWi

[11261] Anonymous. An improved wrench. Scientific American, 59(3):36, July 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n3/pdf/scientificamerican07211888-36.pdf.

Anonymous:1888:NC

[11262] Anonymous. New chromatrope. Scientific American, 59(3):36, July 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n3/pdf/scientificamerican07211888-36a.pdf.

Anonymous:1888:TC

[11263] Anonymous. Temminck’s Chelydra. Scientific American, 59(3):38, July 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n3/pdf/scientificamerican07211888-38.pdf.

Hopkins:1888:SEPh

[11264] M. Hopkins. Simple experiments in physics. Scientific American, 59(3):41, July 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n3/pdf/scientificamerican07211888-
41.pdf.

Anonymous:1888:TF

[11265] Anonymous. Tire furnace. Scientific American, 59(3):41, July 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n3/pdf/scientificamerican07211888-41a.pdf.

Anonymous:1888:SAI

[11266] Anonymous. Simple apparatus for illustrating Newton’s Second Law. Scientific American, 59(3):41, July 21, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v59/
n3/pdf/scientificamerican07211888-41b.pdf.

Anonymous:1888:EIbo

[11267] Anonymous. Engineering invention. Scientific American, 59(3):42, July 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n3/pdf/scientificamerican07211888-42.pdf.

Anonymous:1888:AIba

[11268] Anonymous. Agricultural inventions. Scientific American, 59(3):42, July 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n3/pdf/scientificamerican07211888-42a.pdf.

Anonymous:1888:MIbp

[11269] Anonymous. Miscellaneous inventions. Scientific American, 59(3):42, July 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n3/pdf/scientificamerican07211888-42b.pdf.

Anonymous:1888:Icu

[11270] Anonymous. To inventors. Scientific American, 59(3):43, July 21, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n3/pdf/scientificamerican07211888-43.pdf.

Anonymous:1888:BPbv

[11271] Anonymous. Business and personal. Scientific American, 59(4):59, July 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n4/pdf/scientificamerican07281888-59c.pdf.

Anonymous:1888:BNBbi

[11272] Anonymous. Books: New books and publications. Scientific American, 59(4):59, July 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n4/pdf/scientificamerican07281888-
59d.pdf.

Anonymous:1888:IIbu

[11273] Anonymous. Index of inventions. Scientific American, 59(4):59–60, July 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n4/pdf/scientificamerican07281888-59e.pdf.

Anonymous:1888:PNPbu

[11274] Anonymous. Patent notes: Patents. Scientific American, 59(4):62, July 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n4/pdf/scientificamerican07281888-62.pdf.

Anonymous:1888:HST

[11275] Anonymous. High speed in thick weather. Scientific American, 59(4):48, July 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n4/pdf/scientificamerican07281888-48.pdf.

Anonymous:1888:PPAb

[11276] Anonymous. Position of the planets in August. Scientific American, 59(4):48, July 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n4/pdf/scientificamerican07281888-
48a.pdf.

Anonymous:1888:PNm

[11277] Anonymous. Photographic notes. Scientific American, 59(4):49, July 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n4/pdf/scientificamerican07281888-49.pdf.

Anonymous:1888:ICSe

[11278] Anonymous. An improved car stove. Scientific American, 59(4):50, July 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n4/pdf/scientificamerican07281888-50.pdf.

Anonymous:1888:IRSd

[11279] Anonymous. Improved railway switch stand and signal. Scientific American, 59(4):50, July 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n4/pdf/scientificamerican07281888-
50a.pdf.

Anonymous:1888:ISLe

[11280] Anonymous. An improved spirit level. Scientific American, 59(4):50, July 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n4/pdf/scientificamerican07281888-50b.pdf.

Anonymous:1888:ICCo

[11281] Anonymous. An improved car coupling. Scientific American, 59(4):50, July 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n4/pdf/scientificamerican07281888-50c.pdf.

Anonymous:1888:ISCc

[11282] Anonymous. An improved show case. Scientific American, 59(4):50–51, July 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n4/pdf/scientificamerican07281888-50d.pdf.

Anonymous:1888:IHH

[11283] Anonymous. An improved harmonica holder. Scientific American, 59(4):51, July 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n4/pdf/scientificamerican07281888-
51.pdf.

Anonymous:1888:EAS

[11284] Anonymous. An easel with adjustable shelf. Scientific American, 59(4):51, July 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n4/pdf/scientificamerican07281888-
51a.pdf.

Anonymous:1888:IEGb

[11285] Anonymous. An improved end gate for wagons. Scientific American, 59(4):51–52, July 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n4/pdf/scientificamerican07281888-
51b.pdf.

Hovey:1888:MTS

[11286] H. C. Hovey. A marine toboggan slide. Scientific American, 59(4):53, July 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n4/pdf/scientificamerican07281888-53.pdf.

Anonymous:1888:IRDd

[11287] Anonymous. Improved rock drilling machine. Scientific American, 59(4):54, July 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n4/pdf/scientificamerican07281888-
54.pdf.

Hopkins:1888:SEPi

[11288] M. Hopkins. Simple experiments in physics. Scientific American, 59(4):55, July 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n4/pdf/scientificamerican07281888-
55.pdf.

Anonymous:1888:SV

[11289] Anonymous. The spectacled viper. Scientific American, 59(4):55–56, July 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n4/pdf/scientificamerican07281888-55a.pdf.

Anonymous:1888:SCb

[11290] Anonymous. The swinging of clubs. Scientific American, 59(4):57–58, July 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n4/pdf/scientificamerican07281888-57.pdf.

Anonymous:1888:EIbp

[11291] Anonymous. Engineering invention. Scientific American, 59(4):59, July 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n4/pdf/scientificamerican07281888-59.pdf.

Anonymous:1888:AIbb

[11292] Anonymous. Agricultural inventions. Scientific American, 59(4):59, July 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n4/pdf/scientificamerican07281888-59a.pdf.

Anonymous:1888:MIbq

[11293] Anonymous. Miscellaneous inventions. Scientific American, 59(4):59, July 28, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n4/pdf/scientificamerican07281888-59b.pdf.

Anonymous:1888:Cbm

[11294] Anonymous. Correspondence. Scientific American, 59(5):69, August 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n5/pdf/scientificamerican08041888-69.pdf.

Anonymous:1888:BPbw

[11295] Anonymous. Business and personal. Scientific American, 59(5):74, August 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n5/pdf/scientificamerican08041888-74c.pdf.

Anonymous:1888:BNBbj

[11296] Anonymous. Books: New books and publications. Scientific American, 59(5):74, August 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n5/pdf/scientificamerican08041888-
74d.pdf.

Anonymous:1888:NQbn

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n5/pdf/scientificamerican08041888-75.pdf.

Anonymous:1888:IIbv

[11298] Anonymous. Index of inventions. Scientific American, 59(5):75–76, August 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n5/pdf/scientificamerican08041888-75b.pdf.

Anonymous:1888:PNPbv

[11299] Anonymous. Patent notes: Patents. Scientific American, 59(5):76, August 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n5/pdf/scientificamerican08041888-76.pdf.

Anonymous:1888:IWSg

[11300] Anonymous. The Italian war ship Italia. Scientific American, 59(5):63, August 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n5/pdf/scientificamerican08041888-63.pdf.

Anonymous:1888:AWU

[11301] Anonymous. Artesian wells in Utah. Scientific American, 59(5):64, August 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n5/pdf/scientificamerican08041888-64.pdf.

Anonymous:1888:SSL

[11302] Anonymous. Salt at salt lake, utah. Scientific American, 59(5):64, August 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n5/pdf/scientificamerican08041888-64a.pdf.

Anonymous:1888:MNg

[11303] Anonymous. Military notes. Scientific American, 59(5):64–65, August 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n5/pdf/scientificamerican08041888-64b.pdf.

Anonymous:1888:IBDa

[11304] Anonymous. An improved boat detaching apparatus. Scientific American, 59(5):66, August 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n5/pdf/scientificamerican08041888-
66.pdf.

Anonymous:1888:ICx

[11305] Anonymous. An improved clamp. Scientific American, 59(5):66, August 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n5/pdf/scientificamerican08041888-66a.pdf.

Anonymous:1888:IDL

[11306] Anonymous. An improved door latch. Scientific American, 59(5):66, August 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n5/pdf/scientificamerican08041888-66b.pdf.

Anonymous:1888:ISHd

[11307] Anonymous. An improved slaughter house hoist. Scientific American, 59(5):66–67, August 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n5/pdf/scientificamerican08041888-
66c.pdf.

Anonymous:1888:IFWb

[11308] Anonymous. An improved foot warmer for beds. Scientific American, 59(5):67, August 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n5/pdf/scientificamerican08041888-
67.pdf.

Anonymous:1888:BSH

[11309] Anonymous. A barrel stave hammock. Scientific American, 59(5):67, August 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n5/pdf/scientificamerican08041888-67a.pdf.

Anonymous:1888:ICCp

[11310] Anonymous. An improved car coupling. Scientific American, 59(5):67–68, August 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n5/pdf/scientificamerican08041888-
67b.pdf.

Anonymous:1888:IHM

[11311] Anonymous. An improved hoisting machine. Scientific American, 59(5):68, August 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n5/pdf/scientificamerican08041888-
68.pdf.

Anonymous:1888:IBFa

[11312] Anonymous. An improved boiler feeder. Scientific American, 59(5):68, August 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n5/pdf/scientificamerican08041888-68a.pdf.

Anonymous:1888:EES

[11313] Anonymous. An efficient and economical steam launch. Scientific American, 59(5):70, August 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n5/pdf/scientificamerican08041888-
70.pdf.

Anonymous:1888:VMR

[11314] Anonymous. Vibratory motions of rails during the running of trains. Scientific American, 59(5):71–72, August 4, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v59/
n5/pdf/scientificamerican08041888-71.pdf.

Anonymous:1888:EPE

[11315] Anonymous. The Edison phonograph in England. Scientific American, 59(5):72–73, August 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n5/pdf/scientificamerican08041888-
72.pdf.

Anonymous:1888:EIbq

[11316] Anonymous. Engineering inventions. Scientific American, 59(5):74, August 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n5/pdf/scientificamerican08041888-74.pdf.

Anonymous:1888:AIbc

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n5/pdf/scientificamerican08041888-74a.pdf.

Anonymous:1888:MIbr

[11318] Anonymous. Miscellaneous inventions. Scientific American, 59(5):74, August 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n5/pdf/scientificamerican08041888-74b.pdf.

Anonymous:1888:Icy

[11319] Anonymous. To inventors. Scientific American, 59(5):75, August 4, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n5/pdf/scientificamerican08041888-75a.pdf.

Anonymous:1888:BPbx

[11320] Anonymous. Business and personal. Scientific American, 59(6):90, August 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n6/pdf/scientificamerican08111888-90c.pdf.

Anonymous:1888:NQbo

[11321] Anonymous. Notes and queries. Scientific American, 59(6):90–91, August 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n6/pdf/scientificamerican08111888-90d.pdf.

Anonymous:1888:IIbw

[11322] Anonymous. Index of inventions. Scientific American, 59(6):91–92, August 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n6/pdf/scientificamerican08111888-91a.pdf.

Anonymous:1888:PNPbw

[11323] Anonymous. Patent notes: Patents. Scientific American, 59(6):94, August 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n6/pdf/scientificamerican08111888-94.pdf.

Anonymous:1888:GHN

[11324] Anonymous. The Gate House and Northern Terminus of the new aqueduct. Scientific American, 59(6):79–87, August 11, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v59/
n6/pdf/scientificamerican08111888-87.pdf.

Anonymous:1888:E

[11325] Anonymous. Ericsson at 85. Scientific American, 59(6):80, August 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n6/pdf/scientificamerican08111888-80.pdf.

Anonymous:1888:IICa

[11326] Anonymous. “Injurious influences of city life”. Scientific American, 59(6):80–81, August 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n6/pdf/scientificamerican08111888-
80a.pdf.

Anonymous:1888:MPF

[11327] Anonymous. Multiple punching and forging press. Scientific American, 59(6):82, August 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n6/pdf/scientificamerican08111888-
82.pdf.

Anonymous:1888:ICMb

[11328] Anonymous. An improved current motor. Scientific American, 59(6):82, August 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n6/pdf/scientificamerican08111888-
82a.pdf.

Anonymous:1888:IWSh

[11329] Anonymous. An improved weather strip. Scientific American, 59(6):83, August 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n6/pdf/scientificamerican08111888-
83.pdf.

Anonymous:1888:IFGb

[11330] Anonymous. An improved farm gate. Scientific American, 59(6):83, August 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n6/pdf/scientificamerican08111888-83a.pdf.

Anonymous:1888:IMMc

[11331] Anonymous. An improved mechanical movement. Scientific American, 59(6):83, August 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n6/pdf/scientificamerican08111888-
83b.pdf.

Anonymous:1888:IFCg

[11332] Anonymous. An improved flue cleaner. Scientific American, 59 (6):83, August 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n6/pdf/scientificamerican08111888-
83c.pdf.

Anonymous:1888:ITL

[11333] Anonymous. Improved truck for locomotives and street motors. Scientific American, 59(6):83, August 11, 1888. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v59/
n6/pdf/scientificamerican08111888-83d.pdf.

Anonymous:1888:IOC

[11334] Anonymous. An improved oil can. Scientific American, 59(6):84, August 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n6/pdf/scientificamerican08111888-84.pdf.

Anonymous:1888:CSH

[11335] Anonymous. A combined shingling hammer and plane. Scientific American, 59(6):84, August 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n6/pdf/scientificamerican08111888-
84a.pdf.

Anonymous:1888:ICCq

[11336] Anonymous. An improved curry comb. Scientific American, 59 (6):84, August 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n6/pdf/scientificamerican08111888-
84b.pdf.

Anonymous:1888:IPP

[11337] Anonymous. An improved propeller for pleasure boats. Scientific American, 59(6):84–86, August 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n6/pdf/scientificamerican08111888-
84c.pdf.

Anonymous:1888:FB

[11338] Anonymous. The Forth Bridge. Scientific American, 59(6):86, August 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n6/pdf/scientificamerican08111888-86.pdf.

Hopkins:1888:SEPj

[11339] Geo. M. Hopkins. Simple experiments in physics. Scientific American, 59(6):87–88, August 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n6/pdf/scientificamerican08111888-
87a.pdf.

Anonymous:1888:SGS

[11340] Anonymous. Sand grouse (Syrrhaptes Paradoxus). Scientific American, 59(6):89, August 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n6/pdf/scientificamerican08111888-
89.pdf.

Anonymous:1888:EIbr

[11341] Anonymous. Engineering inventions. Scientific American, 59(6):90, August 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n6/pdf/scientificamerican08111888-90.pdf.

Anonymous:1888:AIbd

[11342] Anonymous. Agricultural inventions. Scientific American, 59(6):90, August 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n6/pdf/scientificamerican08111888-90a.pdf.

Anonymous:1888:MIbs

[11343] Anonymous. Miscellaneous inventions. Scientific American, 59 (6):90, August 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n6/pdf/scientificamerican08111888-
90b.pdf.

Anonymous:1888:Icz

[11344] Anonymous. To inventors. Scientific American, 59(6):91, August 11, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n6/pdf/scientificamerican08111888-91.pdf.

Anonymous:1888:BPby

[11345] Anonymous. Business and personal. Scientific American, 59(7):106, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n7/pdf/scientificamerican08181888-106c.pdf.

Anonymous:1888:BNBbk

[11346] Anonymous. Books: New books and publications. Scientific American, 59(7):106–107, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n7/pdf/scientificamerican08181888-
106d.pdf.

Anonymous:1888:NQbp

[11347] Anonymous. Notes and queries. Scientific American, 59(7):107, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n7/pdf/scientificamerican08181888-107.pdf.

Anonymous:1888:IIbx

[11348] Anonymous. Index of inventions. Scientific American, 59(7): 107–108, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n7/pdf/scientificamerican08181888-
107b.pdf.

Anonymous:1888:PNPbx

[11349] Anonymous. Patent notes: Patents. Scientific American, 59(7):110, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n7/pdf/scientificamerican08181888-110.pdf.

Anonymous:1888:IFGc

[11350] Anonymous. Improved floating grain elevator. Scientific American, 59(7):95–96, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n7/pdf/scientificamerican08181888-
95.pdf.

Anonymous:1888:GPH

[11351] Anonymous. Gen. Philip Henry Sheridan. Scientific American, 59(7):96, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n7/pdf/scientificamerican08181888-
96.pdf.

Anonymous:1888:MCA

[11352] Anonymous. The mechanism of the counter attack. Scientific American, 59(7):96, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n7/pdf/scientificamerican08181888-
96a.pdf.

Anonymous:1888:TWP

[11353] Anonymous. Torpedo warfare in practice and theory. Scientific American, 59(7):96–97, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n7/pdf/scientificamerican08181888-
96b.pdf.

Anonymous:1888:IBEa

[11354] Anonymous. An improved balance escapement for watches. Scientific American, 59(7):98, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n7/pdf/scientificamerican08181888-
98.pdf.

Anonymous:1888:ARM

[11355] Anonymous. The automatic rubber mixer. Scientific American, 59(7):98, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n7/pdf/scientificamerican08181888-
98a.pdf.

Anonymous:1888:IWFb

[11356] Anonymous. An improved wire fence. Scientific American, 59 (7):98, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n7/pdf/scientificamerican08181888-
98b.pdf.

Anonymous:1888:RCH

[11357] Anonymous. A removable calk for horseshoes. Scientific American, 59(7):98, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n7/pdf/scientificamerican08181888-
98c.pdf.

Anonymous:1888:IHDc

[11358] Anonymous. An improved heating drum. Scientific American, 59(7):98–99, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n7/pdf/scientificamerican08181888-
98d.pdf.

Anonymous:1888:IACd

[11359] Anonymous. An improved automatic car brake. Scientific American, 59(7):99, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n7/pdf/scientificamerican08181888-
99.pdf.

Anonymous:1888:ISDb

[11360] Anonymous. An improved signaling device for mines. Scientific American, 59(7):99, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n7/pdf/scientificamerican08181888-
99a.pdf.

Anonymous:1888:IDB

[11361] Anonymous. An improved die and dice box. Scientific American, 59(7):99–100, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n7/pdf/scientificamerican08181888-
99b.pdf.

Anonymous:1888:ITSc

[11362] Anonymous. An improved trunk strap coupling. Scientific American, 59(7):101, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n7/pdf/scientificamerican08181888-
101.pdf.

Anonymous:1888:IDH

[11363] Anonymous. An improved device for hiving bees. Scientific American, 59(7):101, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n7/pdf/scientificamerican08181888-
101a.pdf.

Anonymous:1888:CTE

[11364] Anonymous. Combined traction engine and crane. Scientific American, 59(7):102–103, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n7/pdf/scientificamerican08181888-
102.pdf.

Anonymous:1888:MK

[11365] Anonymous. Macronus Kettlewelli. Scientific American, 59(7):103, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n7/pdf/scientificamerican08181888-103.pdf.

Anonymous:1888:JWP

[11366] Anonymous. John Wesley Powell. Scientific American, 59(7): 103–104, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n7/pdf/scientificamerican08181888-
103a.pdf.

Anonymous:1888:DWT

[11367] Anonymous. Design for a windmill tower and water tank. Scientific American, 59(7):105, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n7/pdf/scientificamerican08181888-
105.pdf.

Anonymous:1888:EIbs

[11368] Anonymous. Engineering inventions. Scientific American, 59(7): 106, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n7/pdf/scientificamerican08181888-
106.pdf.

Anonymous:1888:AIbe

[11369] Anonymous. Agricultural inventions. Scientific American, 59(7): 106, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n7/pdf/scientificamerican08181888-
106a.pdf.

Anonymous:1888:MIbt

[11370] Anonymous. Miscellaneous inventions. Scientific American, 59 (7):106, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n7/pdf/scientificamerican08181888-
106b.pdf.

Anonymous:1888:Idd

[11371] Anonymous. To inventors. Scientific American, 59(7):107, August 18, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n7/pdf/scientificamerican08181888-107a.pdf.

Anonymous:1888:Cbn

[11372] Anonymous. Correspondence. Scientific American, 59(8):121, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n8/pdf/scientificamerican08251888-121.pdf.

Anonymous:1888:BPbz

[11373] Anonymous. Business and personal. Scientific American, 59(8):123, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n8/pdf/scientificamerican08251888-123c.pdf.

Anonymous:1888:IIby 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
123e.pdf.

Anonymous:1888:PNPby

[11375] Anonymous. Patent notes: Patents. Scientific American, 59(8):126, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n8/pdf/scientificamerican08251888-126.pdf.

Anonymous:1888:NTBc

[11376] Anonymous. The new “Times” building. Scientific American, 59 (8):111–117, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
117.pdf.

Anonymous:1888:SNF

[11377] Anonymous. Some new French torpedo boats. Scientific American, 59(8):112, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
112.pdf.

Anonymous:1888:CBO

[11378] Anonymous. Collision between ocean steamers. Scientific American, 59(8):112, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
112a.pdf.

Anonymous:1888:SDB

[11379] Anonymous. Sure death to buffalo moths. Scientific American, 59 (8):112–113, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
112b.pdf.

Anonymous:1888:PNn

[11380] Anonymous. Photographic notes. Scientific American, 59(8):113, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n8/pdf/scientificamerican08251888-113.pdf.

Anonymous:1888:IRM

[11381] Anonymous. Improved refrigerating machine. Scientific American, 59(8):114, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
114.pdf.

Anonymous:1888:ILS

[11382] Anonymous. An improved lamp shade. Scientific American, 59 (8):114, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
114a.pdf.

Anonymous:1888:FAE

[11383] Anonymous. A fumigator and animal exterminator. Scientific American, 59(8):114, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
114b.pdf.

Anonymous:1888:DWA

[11384] Anonymous. A dumping wagon with adjustable scraper. Scientific American, 59(8):115, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
115.pdf.

Anonymous:1888:IPDa

[11385] Anonymous. An improved pallet for drying brick. Scientific American, 59(8):115, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
115a.pdf.

Anonymous:1888:IRSe

[11386] Anonymous. An improved railway signal lantern. Scientific American, 59(8):115, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
115b.pdf.

Anonymous:1888:HTW

[11387] Anonymous. A hose thimble for walls of buildings. Scientific American, 59(8):115, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
115c.pdf.

Anonymous:1888:IWj

[11388] Anonymous. An improved windmill. Scientific American, 59 (8):115, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
115d.pdf.

Anonymous:1888:WHS

[11389] Anonymous. Wing’s high speed engines for fans and other uses. Scientific American, 59(8):116, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
116.pdf.

Anonymous:1888:IDS

[11390] Anonymous. An improved drain for stalls. Scientific American, 59(8):116, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
116a.pdf.

Anonymous:1888:EDR

[11391] Anonymous. Eureka diamond reamer. Scientific American, 59 (8):116, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
116b.pdf.

Anonymous:1888:ISHe

[11392] Anonymous. An improved steam heating boiler. Scientific American, 59(8):116, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
116c.pdf.

Anonymous:1888:IWSi

[11393] Anonymous. An improved window shade. Scientific American, 59(8):116, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
116d.pdf.

Anonymous:1888:NRB

[11394] Anonymous. The new Rouen bridge. Scientific American, 59 (8):118, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
118.pdf.

Hopkins:1888:SEPk

[11395] Geo. M. Hopkins. Simple experiments in physics. Scientific American, 59(8):119, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
119.pdf.

Anonymous:1888:MR

[11396] Anonymous. The magic rose. Scientific American, 59(8):119, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n8/pdf/scientificamerican08251888-119a.pdf.

Anonymous:1888:MRB

[11397] Anonymous. The magic rose bush. Scientific American, 59(8): 119–120, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
119b.pdf.

Anonymous:1888:IFCh

[11398] Anonymous. An improved felly-clip. Scientific American, 59 (8):121–122, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
121a.pdf.

Anonymous:1888:EIbt

[11399] Anonymous. Engineering inventions. Scientific American, 59(8): 123, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
123.pdf.

Anonymous:1888:AIbf

[11400] Anonymous. Agricultural invention. Scientific American, 59(8): 123, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
123a.pdf.

Anonymous:1888:MIbu

[11401] Anonymous. Miscellaneous inventions. Scientific American, 59 (8):123, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n8/pdf/scientificamerican08251888-
123b.pdf.

Anonymous:1888:Idj

[11402] Anonymous. To inventors. Scientific American, 59(8):123, August 25, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n8/pdf/scientificamerican08251888-123d.pdf.

Anonymous:1888:BPca

[11403] Anonymous. Business and personal. Scientific American, 59(9): 138, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
138c.pdf.

Anonymous:1888:NQbq

[11404] Anonymous. Notes and queries. Scientific American, 59(9):138– 139, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
138d.pdf.

Anonymous:1888:IIbz

[11405] Anonymous. Index of inventions. Scientific American, 59(9): 139–140, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
139a.pdf.

Anonymous:1888:PNPbz

[11406] Anonymous. Patent notes: Patents. Scientific American, 59 (9):140, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
140.pdf.

Anonymous:1888:IMBb

[11407] Anonymous. Improved means for buoying up vessels. Scientific American, 59(9):127, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
127.pdf.

Anonymous:1888:RR

[11408] Anonymous. A remarkable raft. Scientific American, 59(9): 127, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
127a.pdf.

Anonymous:1888:YFF

[11409] Anonymous. Yellow fever in Florida. Scientific American, 59 (9):128, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
128.pdf.

Anonymous:1888:GAS

[11410] Anonymous. The Great August Storm. Scientific American, 59 (9):128, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
128a.pdf.

Anonymous:1888:ELCb

[11411] Anonymous. The Electric Lighting Convention. Scientific American, 59(9):128, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
128b.pdf.

Anonymous:1888:MAA

[11412] Anonymous. Meeting of the American Association for the Advancement of Science. Scientific American, 59(9):128, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n9/pdf/scientificamerican09011888-128c.pdf.

Anonymous:1888:PPS

[11413] Anonymous. Position of the planets for September. Scientific American, 59(9):129, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
129.pdf.

Anonymous:1888:IKC

[11414] Anonymous. An improved knife cleaner. Scientific American, 59 (9):130, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
130.pdf.

Anonymous:1888:ISCd

[11415] Anonymous. Improved setting of cultivator plow points. Scientific American, 59(9):130, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
130a.pdf.

Anonymous:1888:ICHd

[11416] Anonymous. An improved corn husker. Scientific American, 59 (9):130, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
130b.pdf.

Anonymous:1888:ICCr

[11417] Anonymous. An improved car coupling. Scientific American, 59 (9):130, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
130c.pdf.

Anonymous:1888:IAf

[11418] Anonymous. An improved auger. Scientific American, 59(9): 130, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
130d.pdf.

Anonymous:1888:AWT

[11419] Anonymous. An automatic water trap for cisterns. Scientific American, 59(9):131, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
131.pdf.

Anonymous:1888:IFEf

[11420] Anonymous. An improved fire escape. Scientific American, 59 (9):131, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
131a.pdf.

Anonymous:1888:ILC

[11421] Anonymous. An improved lock-clamp for pipes. Scientific American, 59(9):131, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
131b.pdf.

Anonymous:1888:ISTb

[11422] Anonymous. An improved screw tap. Scientific American, 59 (9):131, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
131c.pdf.

Anonymous:1888:IRCb

[11423] Anonymous. An improved reservoir for cooling ales, etc. Scientific American, 59(9):131, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
131d.pdf.

Anonymous:1888:SSD

[11424] Anonymous. The spiral screw driver. Scientific American, 59 (9):133, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
133.pdf.

Anonymous:1888:ICCs

[11425] Anonymous. An improved cable car clutch. Scientific American, 59(9):133, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
133a.pdf.

Anonymous:1888:ETC

[11426] Anonymous. Economical triple compound semi-portable engine. Scientific American, 59(9):134, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
134.pdf.

Anonymous:1888:PCA

[11427] Anonymous. The plague of crickets in Algeria. Scientific American, 59(9):135–136, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
135.pdf.

Anonymous:1888:PA

[11428] Anonymous. The Polyparium Ambulans. Scientific American, 59 (9):137, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
137.pdf.

Anonymous:1888:EIbu

[11429] Anonymous. Engineering inventions. Scientific American, 59(9): 138, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
138.pdf.

Anonymous:1888:AIbg

[11430] Anonymous. Agricultural inventions. Scientific American, 59(9): 138, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
138a.pdf.

Anonymous:1888:MIbv

[11431] Anonymous. Miscellaneous inventions. Scientific American, 59 (9):138, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v59/n9/pdf/scientificamerican09011888-
138b.pdf.

Anonymous:1888:Idk

[11432] Anonymous. To inventors. Scientific American, 59(9):139, September 1, 1888. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v59/n9/pdf/scientificamerican09011888-139.pdf.

Anonymous:1889:Ca

[11433] Anonymous. Correspondence. Scientific American, 60(1):7, January 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n1/pdf/scientificamerican01051889-7c.pdf.

Anonymous:1889:RPIa

[11434] Anonymous. Recently patented inventions. Scientific American, 60(1):10, January 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n1/pdf/scientificamerican01051889-
10.pdf.

Anonymous:1889:BPa

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n1/pdf/scientificamerican01051889-10a.pdf.

Anonymous:1889:BNBa

[11436] Anonymous. Books: New books and publications. Scientific American, 60(1):10, January 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n1/pdf/scientificamerican01051889-
10b.pdf.

Anonymous:1889:NQa

[11437] Anonymous. Notes and queries. Scientific American, 60(1):10–11, January 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n1/pdf/scientificamerican01051889-10c.pdf.

Anonymous:1889:IIa

[11438] Anonymous. Index of inventions. Scientific American, 60(1):11–12, January 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n1/pdf/scientificamerican01051889-11a.pdf.

Anonymous:1889:PNPa

[11439] Anonymous. Patent notes: Patents. Scientific American, 60(1):14, January 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n1/pdf/scientificamerican01051889-14.pdf.

Anonymous:1889:DEN

[11440] Anonymous. Death by electricity — the new law of New York. Scientific American, 60(1):2, January 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n1/pdf/scientificamerican01051889-
2.pdf.

Anonymous:1889:PPJa

[11441] Anonymous. Position of the planets in January. Scientific American, 60(1):2, January 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n1/pdf/scientificamerican01051889-
2a.pdf.

Anonymous:1889:PNa

[11442] Anonymous. Photographic notes. Scientific American, 60(1):3, January 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n1/pdf/scientificamerican01051889-3.pdf.

Anonymous:1889:ICB

[11443] Anonymous. An improved car brake and coupler. Scientific American, 60(1):4, January 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n1/pdf/scientificamerican01051889-
4.pdf.

Anonymous:1889:ICAa

[11444] Anonymous. An improved crutch attachment. Scientific American, 60(1):4, January 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n1/pdf/scientificamerican01051889-
4a.pdf.

Anonymous:1889:IT

[11445] Anonymous. An improved tent. Scientific American, 60(1):4, January 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n1/pdf/scientificamerican01051889-4b.pdf.

Anonymous:1889:ITCa

[11446] Anonymous. An improved thill coupling. Scientific American, 60(1):4, January 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n1/pdf/scientificamerican01051889-4c.pdf.

Anonymous:1889:IHR

[11447] Anonymous. An improved head-rest for car seats. Scientific American, 60(1):4, January 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n1/pdf/scientificamerican01051889-
4d.pdf.

Anonymous:1889:MG

[11448] Anonymous. The Mannlicher gun. Scientific American, 60(1):5–6, January 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n1/pdf/scientificamerican01051889-5.pdf.

Anonymous:1889:NRR

[11449] Anonymous. New railroad over the Rocky Mountains of Colorado. Scientific American, 60(1):7, January 5, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v60/
n1/pdf/scientificamerican01051889-7.pdf.

Anonymous:1889:SPP

[11450] Anonymous. A suspender pocket for pencils, etc. Scientific American, 60(1):7, January 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n1/pdf/scientificamerican01051889-
7a.pdf.

Anonymous:1889:IVK

[11451] Anonymous. An improved vegetable knife. Scientific American, 60(1):7, January 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n1/pdf/scientificamerican01051889-7b.pdf.

Hopkins:1889:SEPa

[11452] Geo. M. Hopkins. Simple experiments in physics. Scientific American, 60(1):8, January 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n1/pdf/scientificamerican01051889-
8.pdf.

Anonymous:1889:LT

[11453] Anonymous. The Lowth telephone. Scientific American, 60(1):8–9, January 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n1/pdf/scientificamerican01051889-8a.pdf.

Anonymous:1889:GPC

[11454] Anonymous. A gas-propelled carriage. Scientific American, 60(1):9, January 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n1/pdf/scientificamerican01051889-9.pdf.

Anonymous:1889:Ia

[11455] Anonymous. To inventors. Scientific American, 60(1):11, January 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n1/pdf/scientificamerican01051889-11.pdf.

Anonymous:1889:Cb

[11456] Anonymous. Correspondence. Scientific American, 60(10):149, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n10/pdf/scientificamerican03091889-149.pdf.

Anonymous:1889:RPIb

[11457] Anonymous. Recently patented inventions. Scientific American, 60(10):153, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n10/pdf/scientificamerican03091889-
153.pdf.

Anonymous:1889:BPb

[11458] Anonymous. Business and personal. Scientific American, 60(10):153, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n10/pdf/scientificamerican03091889-153a.pdf.

Anonymous:1889:NQb

[11459] Anonymous. Notes and queries. Scientific American, 60(10):153–155, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n10/pdf/scientificamerican03091889-153b.pdf.

Anonymous:1889:IIb

[11460] Anonymous. Index of inventions. Scientific American, 60(10):155–156, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n10/pdf/scientificamerican03091889-155a.pdf.

Anonymous:1889:PNPb

[11461] Anonymous. Patent notes: Patents. Scientific American, 60(10):158, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n10/pdf/scientificamerican03091889-158.pdf.

Anonymous:1889:MST

[11462] Anonymous. A machine to supersede typesetting. Scientific American, 60(10):143, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n10/pdf/scientificamerican03091889-
143.pdf.

Anonymous:1889:MNE

[11463] Anonymous. Meeting of the National Electric Lighting Association. Scientific American, 60(10):144–145, March 9, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v60/
n10/pdf/scientificamerican03091889-144.pdf.

Anonymous:1889:PPMa

[11464] Anonymous. Position of the planets in March. Scientific American, 60(10):145, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n10/pdf/scientificamerican03091889-
145.pdf.

Anonymous:1889:PVA

[11465] Anonymous. Proposed visit of the American Engineering Societies to Europe. Scientific American, 60(10):145, March 9, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v60/
n10/pdf/scientificamerican03091889-145a.pdf.

Anonymous:1889:MRV

[11466] Anonymous. The Mason reducing valve. Scientific American, 60 (10):146, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n10/pdf/scientificamerican03091889-
146.pdf.

Anonymous:1889:IHM

[11467] Anonymous. An improved hoisting machine. Scientific American, 60(10):146, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n10/pdf/scientificamerican03091889-
146a.pdf.

Anonymous:1889:IGBa

[11468] Anonymous. An improved grate bar. Scientific American, 60(10):146, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n10/pdf/scientificamerican03091889-146b.pdf.

Anonymous:1889:IOCa

[11469] Anonymous. An improved oil can nozzle. Scientific American, 60(10):146, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n10/pdf/scientificamerican03091889-
146c.pdf.

Anonymous:1889:SRC

[11470] Anonymous. The Sartell railway car heater. Scientific American, 60(10):146–147, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n10/pdf/scientificamerican03091889-
146d.pdf.

Anonymous:1889:MDP

[11471] Anonymous. A machine to distribute poison on plants. Scientific American, 60(10):147, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n10/pdf/scientificamerican03091889-
147.pdf.

Anonymous:1889:IDCa

[11472] Anonymous. Improved device to close a tent entrance. Scientific American, 60(10):147, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n10/pdf/scientificamerican03091889-
147a.pdf.

Anonymous:1889:FEE

[11473] Anonymous. The French Exposition of 1889 at the Esplanade Des Invalides. Scientific American, 60(10):147–148, March 9, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v60/
n10/pdf/scientificamerican03091889-147b.pdf.

Anonymous:1889:ISMa

[11474] Anonymous. An improved saw mill dog. Scientific American, 60(10):148, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n10/pdf/scientificamerican03091889-
148.pdf.

Anonymous:1889:IWSa

[11475] Anonymous. An improved wood-sawing machine. Scientific American, 60(10):148, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n10/pdf/scientificamerican03091889-
148a.pdf.

Anonymous:1889:ISEa

[11476] Anonymous. An improved steam engine. Scientific American, 60(10):148, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n10/pdf/scientificamerican03091889-
148b.pdf.

Anonymous:1889:ICWa

[11477] Anonymous. An improved car wheel and bearing. Scientific American, 60(10):148, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n10/pdf/scientificamerican03091889-
148c.pdf.

Anonymous:1889:IPRa

[11478] Anonymous. Improved plunger rod for pumps. Scientific American, 60(10):148–149, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n10/pdf/scientificamerican03091889-
148d.pdf.

Anonymous:1889:AIM

[11479] Anonymous. Automatic indicator for magazine guns. Scientific American, 60(10):151–152, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n10/pdf/scientificamerican03091889-
151.pdf.

Hopkins:1889:LP

[11480] M. Hopkins. Lantern pantographs. Scientific American, 60(10):152, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n10/pdf/scientificamerican03091889-152.pdf.

Anonymous:1889:TET

[11481] Anonymous. The top of the Eiffel Tower. Scientific American, 60(10):152, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n10/pdf/scientificamerican03091889-
152a.pdf.

Anonymous:1889:IASa

[11482] Anonymous. An improved air ship. Scientific American, 60(10):152, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n10/pdf/scientificamerican03091889-152b.pdf.

Anonymous:1889:Ib

[11483] Anonymous. To inventors. Scientific American, 60(10):155, March 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n10/pdf/scientificamerican03091889-155.pdf.

Anonymous:1889:RPIc

[11484] Anonymous. Recently patented inventions. Scientific American, 60(11):169, March 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n11/pdf/scientificamerican03161889-
169.pdf.

Anonymous:1889:BPc

[11485] Anonymous. Business and personal. Scientific American, 60(11):169, March 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n11/pdf/scientificamerican03161889-169a.pdf.

Anonymous:1889:NQc

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n11/pdf/scientificamerican03161889-169b.pdf.

Anonymous:1889:IIc

[11487] Anonymous. Index of inventions. Scientific American, 60(11):171–172, March 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n11/pdf/scientificamerican03161889-171a.pdf.

Anonymous:1889:PNPc

[11488] Anonymous. Patent notes: Patents. Scientific American, 60(11):174, March 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n11/pdf/scientificamerican03161889-174.pdf.

Anonymous:1889:OD

[11489] Anonymous. The obstructions in the delaware. Scientific American, 60(11):159–160, March 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n11/pdf/scientificamerican03161889-
159.pdf.

Anonymous:1889:NMP

[11490] Anonymous. Naval and maritime prospects under the new Administration. Scientific American, 60(11):160, March 16, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v60/
n11/pdf/scientificamerican03161889-160.pdf.

Anonymous:1889:NWFa

[11491] Anonymous. Naval war of the future. Scientific American, 60 (11):160–161, March 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n11/pdf/scientificamerican03161889-
160a.pdf.

Anonymous:1889:PIBa

[11492] Anonymous. Proposed increase of the British Navy. Scientific American, 60(11):161, March 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n11/pdf/scientificamerican03161889-
161.pdf.

Anonymous:1889:RJ

[11493] Anonymous. A rival for jute. Scientific American, 60(11):161, March 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n11/pdf/scientificamerican03161889-161a.pdf.

Anonymous:1889:CWa

[11494] Anonymous. The celestial world. Scientific American, 60(11):161, March 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n11/pdf/scientificamerican03161889-161b.pdf.

Anonymous:1889:IPCa

[11495] Anonymous. An improved pipe coupling. Scientific American, 60(11):162, March 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n11/pdf/scientificamerican03161889-
162.pdf.

Anonymous:1889:IBa

[11496] Anonymous. An improved buckle. Scientific American, 60(11):162, March 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n11/pdf/scientificamerican03161889-162a.pdf.

Anonymous:1889:ILBa

[11497] Anonymous. An improved lamp burner. Scientific American, 60(11):162, March 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n11/pdf/scientificamerican03161889-
162b.pdf.

Anonymous:1889:IPPa

[11498] Anonymous. An improved potato planter. Scientific American, 60(11):162, March 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n11/pdf/scientificamerican03161889-
162c.pdf.

Anonymous:1889:IBb

[11499] Anonymous. An improved barrel. Scientific American, 60(11):162, March 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n11/pdf/scientificamerican03161889-162d.pdf.

Anonymous:1889:IMBa

[11500] Anonymous. An improved music or book holder. Scientific American, 60(11):162, March 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n11/pdf/scientificamerican03161889-
162e.pdf.

Anonymous:1889:CJE

[11501] Anonymous. Captain John Ericsson. Scientific American, 60 (11):163–164, March 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n11/pdf/scientificamerican03161889-
163.pdf.

Anonymous:1889:MSCa

[11502] Anonymous. The Manchester Ship Canal. Scientific American, 60 (11):164–165, March 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n11/pdf/scientificamerican03161889-
164.pdf.

Anonymous:1889:IVSa

[11503] Anonymous. An improved vehicle shaft support. Scientific American, 60(11):165, March 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n11/pdf/scientificamerican03161889-
165.pdf.

Anonymous:1889:IRCa

[11504] Anonymous. An improved road cart. Scientific American, 60(11):165, March 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n11/pdf/scientificamerican03161889-165a.pdf.

Anonymous:1889:IPa

[11505] Anonymous. An improved pavement. Scientific American, 60(11):165, March 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n11/pdf/scientificamerican03161889-165b.pdf.

Anonymous:1889:CIP

[11506] Anonymous. The charity institutions of Paris. Scientific American, 60(11):166, March 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n11/pdf/scientificamerican03161889-
166.pdf.

Anonymous:1889:SCB

[11507] Anonymous. A suggestion in canal boat propulsion. Scientific American, 60(11):167, March 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n11/pdf/scientificamerican03161889-
167.pdf.

Hopkins:1889:VD

[11508] M. Hopkins. Vibrations of diaphragms. Scientific American, 60 (11):167–168, March 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n11/pdf/scientificamerican03161889-
167a.pdf.

Anonymous:1889:Ic

[11509] Anonymous. To inventors. Scientific American, 60(11):171, March 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n11/pdf/scientificamerican03161889-171.pdf.

Anonymous:1889:Cc

[11510] Anonymous. Correspondence. Scientific American, 60(12):181, March 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n12/pdf/scientificamerican03231889-181.pdf.

Anonymous:1889:RPId

[11511] Anonymous. Recently patented inventions. Scientific American, 60(12):186, March 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n12/pdf/scientificamerican03231889-
186.pdf.

Anonymous:1889:BPd

[11512] Anonymous. Business and personal. Scientific American, 60(12):186, March 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n12/pdf/scientificamerican03231889-186a.pdf.

Anonymous:1889:NQd

[11513] Anonymous. Notes and queries. Scientific American, 60(12):186–187, March 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n12/pdf/scientificamerican03231889-186b.pdf.

Anonymous:1889:IId

[11514] Anonymous. Index of inventions. Scientific American, 60(12):187–189, March 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n12/pdf/scientificamerican03231889-187a.pdf.

Anonymous:1889:PNPd

[11515] Anonymous. Patent notes: Patents. Scientific American, 60(12):190, March 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n12/pdf/scientificamerican03231889-190.pdf.

Anonymous:1889:LCE

[11516] Anonymous. The late Capt. Ericsson and his work. Scientific American, 60(12):175, March 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n12/pdf/scientificamerican03231889-
175.pdf.

Anonymous:1889:SPW

[11517] Anonymous. Satisfactory prospects for the war vessels. Scientific American, 60(12):176, March 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n12/pdf/scientificamerican03231889-
176.pdf.

Anonymous:1889:PEI

[11518] Anonymous. Progress of electric illumination. Scientific American, 60(12):176–177, March 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n12/pdf/scientificamerican03231889-
176a.pdf.

Anonymous:1889:PAI

[11519] Anonymous. Present aspect of invention. Scientific American, 60(12):177, March 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n12/pdf/scientificamerican03231889-
177.pdf.

Anonymous:1889:PIBb

[11520] Anonymous. The Pratt Institute of Brooklyn. Scientific American, 60(12):177, March 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n12/pdf/scientificamerican03231889-
177a.pdf.

Anonymous:1889:IRM

[11521] Anonymous. An improved routing machine. Scientific American, 60(12):178, March 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n12/pdf/scientificamerican03231889-
178.pdf.

Anonymous:1889:IVCa

[11522] Anonymous. Improved vertical cylinder boring machine. Scientific American, 60(12):178, March 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n12/pdf/scientificamerican03231889-
178a.pdf.

Hopkins:1889:SF

[11523] M. Hopkins. Sensitive flames. Scientific American, 60(12):179, March 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n12/pdf/scientificamerican03231889-179.pdf.

Anonymous:1889:CSH

[11524] Anonymous. A cheap summer house. Scientific American, 60 (12):179–180, March 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n12/pdf/scientificamerican03231889-
179a.pdf.

Anonymous:1889:DRPa

[11525] Anonymous. Decisions relating to patents. Scientific American, 60(12):181, March 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n12/pdf/scientificamerican03231889-
181a.pdf.

Anonymous:1889:ICCa

[11526] Anonymous. An improved car coupling. Scientific American, 60(12):185, March 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n12/pdf/scientificamerican03231889-
185.pdf.

Anonymous:1889:Id

[11527] Anonymous. To inventors. Scientific American, 60(12):187, March 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n12/pdf/scientificamerican03231889-187.pdf.

Anonymous:1889:CCa

[11528] Anonymous. Correspondence: correspondence. Scientific American, 60(13):197, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n13/pdf/scientificamerican03301889-
197.pdf.

Anonymous:1889:BPe

[11529] Anonymous. Business and personal. Scientific American, 60(13):202, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n13/pdf/scientificamerican03301889-202a.pdf.

Anonymous:1889:NQe

[11530] Anonymous. Notes and queries. Scientific American, 60(13):202–203, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n13/pdf/scientificamerican03301889-202b.pdf.

Anonymous:1889:IIe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n13/pdf/scientificamerican03301889-203a.pdf.

Anonymous:1889:PNPe

[11532] Anonymous. Patent notes: Patents. Scientific American, 60(13):206, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n13/pdf/scientificamerican03301889-206.pdf.

Anonymous:1889:MTC

[11533] Anonymous. The Metropolitan Telephone Co.’s New Central Station and Great Switchboard. Scientific American, 60(13):191, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n13/pdf/scientificamerican03301889-191.pdf.

Anonymous:1889:NCP

[11534] Anonymous. The new Commissioner of Patents. Scientific American, 60(13):192, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n13/pdf/scientificamerican03301889-
192.pdf.

Anonymous:1889:WMA

[11535] Anonymous. War material of American designing. Scientific American, 60(13):192, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n13/pdf/scientificamerican03301889-
192a.pdf.

Anonymous:1889:WFC

[11536] Anonymous. Wanted — a first class gun. Scientific American, 60(13):192, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n13/pdf/scientificamerican03301889-
192b.pdf.

Anonymous:1889:PPAa

[11537] Anonymous. Position of the planets in April. Scientific American, 60(13):192–193, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n13/pdf/scientificamerican03301889-
192c.pdf.

Anonymous:1889:IHSa

[11538] Anonymous. Improved handle for sliding doors. Scientific American, 60(13):194, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n13/pdf/scientificamerican03301889-
194.pdf.

Anonymous:1889:IBBa

[11539] Anonymous. An improved brake block. Scientific American, 60(13):194, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n13/pdf/scientificamerican03301889-194a.pdf.

Anonymous:1889:IBc

[11540] Anonymous. An improved buckboard. Scientific American, 60(13):194, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n13/pdf/scientificamerican03301889-194b.pdf.

Anonymous:1889:IKC

[11541] Anonymous. An improved knife cleaner. Scientific American, 60(13):194, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n13/pdf/scientificamerican03301889-
194c.pdf.

Anonymous:1889:IDDa

[11542] Anonymous. Improved device for detaching animals. Scientific American, 60(13):195, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n13/pdf/scientificamerican03301889-
195.pdf.

Anonymous:1889:IHSb

[11543] Anonymous. Improved hood for stoves or ranges. Scientific American, 60(13):195, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n13/pdf/scientificamerican03301889-
195a.pdf.

Anonymous:1889:IECa

[11544] Anonymous. An insulator for electrical conductors. Scientific American, 60(13):195, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n13/pdf/scientificamerican03301889-
195b.pdf.

Anonymous:1889:NNL

[11545] Anonymous. A novel nut lock. Scientific American, 60(13):195, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n13/pdf/scientificamerican03301889-195c.pdf.

Anonymous:1889:IRH

[11546] Anonymous. An improved rotary harrow. Scientific American, 60(13):195, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n13/pdf/scientificamerican03301889-
195d.pdf.

Anonymous:1889:IPN

[11547] Anonymous. An improved plant nest. Scientific American, 60 (13):195–196, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n13/pdf/scientificamerican03301889-
195e.pdf.

Anonymous:1889:AEC

[11548] Anonymous. Armstrong’s envelope cutter. Scientific American, 60(13):196, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n13/pdf/scientificamerican03301889-
196.pdf.

Anonymous:1889:IPPb

[11549] Anonymous. An improved photograph printing frame. Scientific American, 60(13):196, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n13/pdf/scientificamerican03301889-
196a.pdf.

Anonymous:1889:ICDa

[11550] Anonymous. Improved calipers and dividers. Scientific American, 60(13):198, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n13/pdf/scientificamerican03301889-
198.pdf.

Anonymous:1889:NSS

[11551] Anonymous. The new Spanish ship of war Pelayo. Scientific American, 60(13):198, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n13/pdf/scientificamerican03301889-
198a.pdf.

Anonymous:1889:IWTa

[11552] Anonymous. An improved watering trough. Scientific American, 60(13):198, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n13/pdf/scientificamerican03301889-
198b.pdf.

Anonymous:1889:PNb

[11553] Anonymous. Photographic notes. Scientific American, 60(13):201, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n13/pdf/scientificamerican03301889-201.pdf.

Anonymous:1889:RPIe

[11554] Anonymous. Recently patented inventions. Scientific American, 60(13):202, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n13/pdf/scientificamerican03301889-
202.pdf.

Anonymous:1889:Ie

[11555] Anonymous. To inventors. Scientific American, 60(13):203, March 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n13/pdf/scientificamerican03301889-203.pdf.

Anonymous:1889:Cd

[11556] Anonymous. Correspondence. Scientific American, 60(14):213, April 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n14/pdf/scientificamerican04061889-213.pdf.

Anonymous:1889:RPIf

[11557] Anonymous. Recently patented inventions. Scientific American, 60(14):218, April 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n14/pdf/scientificamerican04061889-
218.pdf.

Anonymous:1889:BPf

[11558] Anonymous. Business and personal. Scientific American, 60(14):218, April 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n14/pdf/scientificamerican04061889-218a.pdf.

Anonymous:1889:NQf

[11559] Anonymous. Notes and queries. Scientific American, 60(14):218–219, April 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n14/pdf/scientificamerican04061889-218b.pdf.

Anonymous:1889:IIf

[11560] Anonymous. Index of inventions. Scientific American, 60(14):219–220, April 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n14/pdf/scientificamerican04061889-219a.pdf.

Anonymous:1889:PNPf

[11561] Anonymous. Patent notes: Patents. Scientific American, 60(14):222, April 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n14/pdf/scientificamerican04061889-222.pdf.

Korns:1889:AWJ

[11562] L. F. Korns. The artesian wells of the James River Valley, Dakota. Scientific American, 60(14):207, April 6, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v60/
n14/pdf/scientificamerican04061889-207.pdf.

Anonymous:1889:ESE

[11563] Anonymous. Electric subway explosions. Scientific American, 60 (14):208–209, April 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n14/pdf/scientificamerican04061889-
208.pdf.

Anonymous:1889:IVB

[11564] Anonymous. An improved vehicle brake. Scientific American, 60 (14):210, April 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n14/pdf/scientificamerican04061889-
210.pdf.

Anonymous:1889:IFDa

[11565] Anonymous. An improved fertilizer distributer. Scientific American, 60(14):210, April 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n14/pdf/scientificamerican04061889-
210a.pdf.

Anonymous:1889:IDEa

[11566] Anonymous. An improved draught equalizer. Scientific American, 60(14):210, April 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n14/pdf/scientificamerican04061889-
210b.pdf.

Anonymous:1889:ISSa

[11567] Anonymous. An improved snow shovel and scraper. Scientific American, 60(14):210, April 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n14/pdf/scientificamerican04061889-
210c.pdf.

Anonymous:1889:IFBa

[11568] Anonymous. An improved foot board for wood choppers. Scientific American, 60(14):210, April 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n14/pdf/scientificamerican04061889-
210d.pdf.

Anonymous:1889:CTC

[11569] Anonymous. Chaplin try and center square. Scientific American, 60(14):211, April 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n14/pdf/scientificamerican04061889-
211.pdf.

Anonymous:1889:IBCa

[11570] Anonymous. An improved book carriage. Scientific American, 60(14):211, April 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n14/pdf/scientificamerican04061889-
211a.pdf.

Anonymous:1889:IDPa

[11571] Anonymous. An improved device for poaching eggs. Scientific American, 60(14):211, April 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n14/pdf/scientificamerican04061889-
211b.pdf.

Anonymous:1889:OFC

[11572] Anonymous. Ostrich farming in California. Scientific American, 60(14):211–212, April 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n14/pdf/scientificamerican04061889-
211c.pdf.

Anonymous:1889:PNc

[11573] Anonymous. Photographic notes. Scientific American, 60(14):212, April 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n14/pdf/scientificamerican04061889-212.pdf.

Anonymous:1889:JHB

[11574] Anonymous. John H. B. Latrobe. Scientific American, 60(14):214–215, April 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n14/pdf/scientificamerican04061889-214.pdf.

Anonymous:1889:IDT

[11575] Anonymous. An improved drainage-trap. Scientific American, 60 (14):215, April 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n14/pdf/scientificamerican04061889-
215.pdf.

Anonymous:1889:IIg

[11576] Anonymous. An improved inkstand. Scientific American, 60(14):215, April 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n14/pdf/scientificamerican04061889-215a.pdf.

Anonymous:1889:ISPa

[11577] Anonymous. An improved seal-press. Scientific American, 60 (14):215–217, April 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n14/pdf/scientificamerican04061889-
215b.pdf.

Hopkins:1889:SES

[11578] Geo. M. Hopkins. Some experiments in sound. Scientific American, 60(14):217, April 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n14/pdf/scientificamerican04061889-
217.pdf.

Anonymous:1889:If

[11579] Anonymous. To inventors. Scientific American, 60(14):219, April 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n14/pdf/scientificamerican04061889-219.pdf.

Anonymous:1889:Ce

[11580] Anonymous. Correspondence. Scientific American, 60(15):229, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n15/pdf/scientificamerican04131889-229a.pdf.

Anonymous:1889:RPIg

[11581] Anonymous. Recently patented inventions. Scientific American, 60(15):233, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n15/pdf/scientificamerican04131889-
233.pdf.

Anonymous:1889:BPg

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n15/pdf/scientificamerican04131889-233a.pdf.

Anonymous:1889:BNBb

[11583] Anonymous. Books: New books and publications. Scientific American, 60(15):233, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n15/pdf/scientificamerican04131889-
233b.pdf.

Anonymous:1889:NQg

[11584] Anonymous. Notes and queries. Scientific American, 60(15):234–235, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n15/pdf/scientificamerican04131889-234.pdf.

Anonymous:1889:IIh

[11585] Anonymous. Index of inventions. Scientific American, 60(15):235–236, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n15/pdf/scientificamerican04131889-235a.pdf.

Anonymous:1889:PNPg

[11586] Anonymous. Patent notes: Patents. Scientific American, 60(15):238, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n15/pdf/scientificamerican04131889-238.pdf.

Anonymous:1889:NAC

[11587] Anonymous. A new assistant Commissioner of Patents. Scientific American, 60(15):224, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n15/pdf/scientificamerican04131889-
224.pdf.

Anonymous:1889:LB

[11588] Anonymous. Lager beer. Scientific American, 60(15):224, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n15/pdf/scientificamerican04131889-224a.pdf.

Anonymous:1889:PNd

[11589] Anonymous. Photographic notes. Scientific American, 60(15):225, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n15/pdf/scientificamerican04131889-225.pdf.

Anonymous:1889:ITLa

[11590] Anonymous. Improved tongs for lamp chimneys, etc. Scientific American, 60(15):226, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n15/pdf/scientificamerican04131889-
226.pdf.

Anonymous:1889:IBCb

[11591] Anonymous. An improved basket or carrier. Scientific American, 60(15):226, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n15/pdf/scientificamerican04131889-
226a.pdf.

Anonymous:1889:ISPb

[11592] Anonymous. An improved snow plow. Scientific American, 60(15):226, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n15/pdf/scientificamerican04131889-226b.pdf.

Anonymous:1889:ICCb

[11593] Anonymous. An improved car coupling. Scientific American, 60 (15):226, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n15/pdf/scientificamerican04131889-
226c.pdf.

Anonymous:1889:ISCa

[11594] Anonymous. An improved spout connection for cans, etc. Scientific American, 60(15):226, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n15/pdf/scientificamerican04131889-
226d.pdf.

Anonymous:1889:ISDa

[11595] Anonymous. An improved school desk and seat. Scientific American, 60(15):227, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n15/pdf/scientificamerican04131889-
227.pdf.

Anonymous:1889:ICPa

[11596] Anonymous. An improved coin package. Scientific American, 60 (15):227, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n15/pdf/scientificamerican04131889-
227a.pdf.

Anonymous:1889:IABa

[11597] Anonymous. An improved axle bearing. Scientific American, 60 (15):227, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n15/pdf/scientificamerican04131889-
227b.pdf.

Anonymous:1889:ALC

[11598] Anonymous. An adjustable leg for chairs. Scientific American, 60(15):227, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n15/pdf/scientificamerican04131889-
227c.pdf.

Anonymous:1889:IFBb

[11599] Anonymous. An improved folding bath tub. Scientific American, 60(15):227, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n15/pdf/scientificamerican04131889-
227d.pdf.

Pike:1889:MLS

[11600] Nicolas Pike. The mud laff — (Synanceia verrucosa Cuv.). Scientific American, 60(15):227–228, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v60/n15/pdf/scientificamerican04131889-
227e.pdf.

Anonymous:1889:IPDa

[11601] Anonymous. An improved potato digger. Scientific American, 60(15):228, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n15/pdf/scientificamerican04131889-
228.pdf.

Anonymous:1889:JEC

[11602] Anonymous. The Jessup exercising club. Scientific American, 60(15):228, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n15/pdf/scientificamerican04131889-
228a.pdf.

Anonymous:1889:IDSa

[11603] Anonymous. An improved device for suspending hammocks. Scientific American, 60(15):228, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n15/pdf/scientificamerican04131889-
228b.pdf.

Anonymous:1889:VA

[11604] Anonymous. Views in Alaska. Scientific American, 60(15):229, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n15/pdf/scientificamerican04131889-229.pdf.

Anonymous:1889:ICG

[11605] Anonymous. An improved centrifugal governor. Scientific American, 60(15):230, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n15/pdf/scientificamerican04131889-
230.pdf.

Anonymous:1889:LTG

[11606] Anonymous. Large twin gas engine. Scientific American, 60(15):230, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n15/pdf/scientificamerican04131889-230a.pdf.

Anonymous:1889:ICSa

[11607] Anonymous. An improved cotton seed cleaner. Scientific American, 60(15):230, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n15/pdf/scientificamerican04131889-
230b.pdf.

Anonymous:1889:PEa

[11608] Anonymous. The Paris Exhibition. Scientific American, 60(15): 231–232, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n15/pdf/scientificamerican04131889-
231.pdf.

Anonymous:1889:IPCb

[11609] Anonymous. An improved pump and condenser. Scientific American, 60(15):232, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n15/pdf/scientificamerican04131889-
232.pdf.

Anonymous:1889:Ig

[11610] Anonymous. To inventors. Scientific American, 60(15):235, April 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n15/pdf/scientificamerican04131889-235.pdf.

Anonymous:1889:Cf

[11611] Anonymous. Correspondence. Scientific American, 60(16):245, April 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n16/pdf/scientificamerican04201889-245a.pdf.

Anonymous:1889:RPIh

[11612] Anonymous. Recently patented inventions. Scientific American, 60(16):249, April 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n16/pdf/scientificamerican04201889-
249.pdf.

Anonymous:1889:BPh

[11613] Anonymous. Business and personal. Scientific American, 60(16):249, April 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n16/pdf/scientificamerican04201889-249a.pdf.

Anonymous:1889:BNBc

[11614] Anonymous. Books: New books and publications. Scientific American, 60(16):249–250, April 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n16/pdf/scientificamerican04201889-
249b.pdf.

Anonymous:1889:NQh

[11615] Anonymous. Notes and queries. Scientific American, 60(16):250–251, April 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n16/pdf/scientificamerican04201889-250.pdf.

Anonymous:1889:IIi

[11616] Anonymous. Index of inventions. Scientific American, 60(16):251–252, April 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n16/pdf/scientificamerican04201889-251a.pdf.

Anonymous:1889:PNPh

[11617] Anonymous. Patent notes: Patents. Scientific American, 60(16):254, April 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n16/pdf/scientificamerican04201889-254.pdf.

Anonymous:1889:ESCa

[11618] Anonymous. The electric subways of the City of New York. Scientific American, 60(16):239, April 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n16/pdf/scientificamerican04201889-
239.pdf.

Anonymous:1889:PCH

[11619] Anonymous. Permeability of cements, the Hastings telescope, and more. Scientific American, 60(16):240–241, April 20, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v60/
n16/pdf/scientificamerican04201889-240.pdf.

Hopkins:1889:EMa

[11620] Geo. M. Hopkins. Experiments in magnetism. Scientific American, 60(16):242, April 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n16/pdf/scientificamerican04201889-
242.pdf.

Anonymous:1889:ISU

[11621] Anonymous. Improved structure for use as a silo. Scientific American, 60(16):242, April 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n16/pdf/scientificamerican04201889-
242a.pdf.

Anonymous:1889:IWSb

[11622] Anonymous. An improved wire stretcher. Scientific American, 60(16):242, April 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n16/pdf/scientificamerican04201889-
242b.pdf.

Anonymous:1889:IBSa

[11623] Anonymous. An improved book support. Scientific American, 60(16):243, April 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n16/pdf/scientificamerican04201889-
243.pdf.

Anonymous:1889:IRCb

[11624] Anonymous. An improved rocking chair. Scientific American, 60(16):243, April 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n16/pdf/scientificamerican04201889-
243a.pdf.

Anonymous:1889:IPKa

[11625] Anonymous. An improved pocket knife. Scientific American, 60 (16):243, April 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n16/pdf/scientificamerican04201889-
243b.pdf.

Anonymous:1889:IBd

[11626] Anonymous. An improved binder. Scientific American, 60(16):243, April 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n16/pdf/scientificamerican04201889-243c.pdf.

Anonymous:1889:IPDb

[11627] Anonymous. Improved protector for dikes or levees. Scientific American, 60(16):243, April 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n16/pdf/scientificamerican04201889-
243d.pdf.

Anonymous:1889:IRCc

[11628] Anonymous. An improved rope clamp. Scientific American, 60 (16):243–244, April 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n16/pdf/scientificamerican04201889-
243e.pdf.

Anonymous:1889:PNe

[11629] Anonymous. Photographic notes. Scientific American, 60(16):245, April 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n16/pdf/scientificamerican04201889-245.pdf.

Anonymous:1889:MEC

[11630] Anonymous. Michel Eugene Chevreul. Scientific American, 60 (16):247–248, April 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n16/pdf/scientificamerican04201889-
247.pdf.

Anonymous:1889:RJG

[11631] Anonymous. The Rev. J. G. Wood. Scientific American, 60(16):248, April 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n16/pdf/scientificamerican04201889-248.pdf.

Anonymous:1889:Ih

[11632] Anonymous. To inventors. Scientific American, 60(16):251, April 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n16/pdf/scientificamerican04201889-251.pdf.

Anonymous:1889:RPIi

[11633] Anonymous. Recently patented inventions. Scientific American, 60(17):265, April 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n17/pdf/scientificamerican04271889-
265.pdf.

Anonymous:1889:BPi

[11634] Anonymous. Business and personal. Scientific American, 60(17):265, April 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n17/pdf/scientificamerican04271889-265a.pdf.

Anonymous:1889:BNBd

[11635] Anonymous. Books: New books and publications. Scientific American, 60(17):265, April 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n17/pdf/scientificamerican04271889-
265b.pdf.

Anonymous:1889:NQi

[11636] Anonymous. Notes and queries. Scientific American, 60(17):265–267, April 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n17/pdf/scientificamerican04271889-265c.pdf.

Anonymous:1889:IIj

[11637] Anonymous. Index of inventions. Scientific American, 60(17):267–268, April 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n17/pdf/scientificamerican04271889-267a.pdf.

Anonymous:1889:PNPi

[11638] Anonymous. Patent notes: Patents. Scientific American, 60(17):270, April 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n17/pdf/scientificamerican04271889-270.pdf.

Anonymous:1889:RAE

[11639] Anonymous. Removal of aerial electric wires in New York. Scientific American, 60(17):256, April 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n17/pdf/scientificamerican04271889-
256.pdf.

Anonymous:1889:NSU

[11640] Anonymous. A new suit under electric dynamo construction patents. Scientific American, 60(17):256, April 27, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v60/
n17/pdf/scientificamerican04271889-256a.pdf.

Anonymous:1889:DGA

[11641] Anonymous. Delay in granting applications for patents. Scientific American, 60(17):256, April 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n17/pdf/scientificamerican04271889-
256b.pdf.

Anonymous:1889:PPMb

[11642] Anonymous. Position of the planets in May. Scientific American, 60(17):256–257, April 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n17/pdf/scientificamerican04271889-
256c.pdf.

Anonymous:1889:BPBa

[11643] Anonymous. The Berrie paving block. Scientific American, 60(17):258, April 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n17/pdf/scientificamerican04271889-258.pdf.

Anonymous:1889:SPSa

[11644] Anonymous. Straub’s plan for a sub-river tunnel. Scientific American, 60(17):258, April 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n17/pdf/scientificamerican04271889-
258a.pdf.

Anonymous:1889:IRSa

[11645] Anonymous. An improved railroad snow plow. Scientific American, 60(17):258, April 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n17/pdf/scientificamerican04271889-
258b.pdf.

Anonymous:1889:IOT

[11646] Anonymous. An improved oven thermometer. Scientific American, 60(17):259, April 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n17/pdf/scientificamerican04271889-
259.pdf.

Anonymous:1889:IBPa

[11647] Anonymous. An improved bake pan. Scientific American, 60(17):259, April 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n17/pdf/scientificamerican04271889-259a.pdf.

Anonymous:1889:IGF

[11648] Anonymous. An improved gate fastening. Scientific American, 60(17):259, April 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n17/pdf/scientificamerican04271889-
259b.pdf.

Anonymous:1889:ILCa

[11649] Anonymous. An improved lightning conductor for fences. Scientific American, 60(17):259, April 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n17/pdf/scientificamerican04271889-
259c.pdf.

Anonymous:1889:IBBb

[11650] Anonymous. An improved building block. Scientific American, 60(17):259, April 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n17/pdf/scientificamerican04271889-
259d.pdf.

Anonymous:1889:IRSb

[11651] Anonymous. An improved roller stove foot. Scientific American, 60(17):259, April 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n17/pdf/scientificamerican04271889-
259e.pdf.

Anonymous:1889:IHB

[11652] Anonymous. An improved hydrocarbon burner. Scientific American, 60(17):260–261, April 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n17/pdf/scientificamerican04271889-
260.pdf.

Anonymous:1889:IRFa

[11653] Anonymous. Improvement of the river front of New York City. Scientific American, 60(17):261, April 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n17/pdf/scientificamerican04271889-
261.pdf.

Hopkins:1889:EMb

[11654] Geo. M. Hopkins. Experiments in magnetism. Scientific American, 60(17):262, April 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n17/pdf/scientificamerican04271889-
262.pdf.

Anonymous:1889:RLE

[11655] Anonymous. Remarkable locomotive explosion in Norway. Scientific American, 60(17):262–263, April 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v60/n17/pdf/scientificamerican04271889-
262a.pdf.

Anonymous:1889:TPA

[11656] Anonymous. The twin palms and ancient well of Los Angeles. Scientific American, 60(17):263–264, April 27, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v60/
n17/pdf/scientificamerican04271889-263.pdf.

Sloane:1889:MJS

[11657] T. O’Conor Sloane. Mercurial jet siphon. Scientific American, 60(17):264, April 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n17/pdf/scientificamerican04271889-
264.pdf.

Anonymous:1889:Ij

[11658] Anonymous. To inventors. Scientific American, 60(17):267, April 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n17/pdf/scientificamerican04271889-267.pdf.

Anonymous:1889:Cg

[11659] Anonymous. Correspondence. Scientific American, 60(18):280–281, May 4, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n18/pdf/scientificamerican05041889-280.pdf.

Anonymous:1889:RPIj

[11660] Anonymous. Recently patented inventions. Scientific American, 60(18):282, May 4, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n18/pdf/scientificamerican05041889-
282.pdf.

Anonymous:1889:BPj

[11661] Anonymous. Business and personal. Scientific American, 60(18):282, May 4, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n18/pdf/scientificamerican05041889-282a.pdf.

Anonymous:1889:NQj

[11662] Anonymous. Notes and queries. Scientific American, 60(18):282–283, May 4, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n18/pdf/scientificamerican05041889-282b.pdf.

Anonymous:1889:BNBe

[11663] Anonymous. Books: New books and publications. Scientific American, 60(18):283, May 4, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n18/pdf/scientificamerican05041889-
283.pdf.

Anonymous:1889:IIk

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n18/pdf/scientificamerican05041889-283b.pdf.

Anonymous:1889:PNPj

[11665] Anonymous. Patent notes: Patents. Scientific American, 60(18):286, May 4, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n18/pdf/scientificamerican05041889-286.pdf.

Anonymous:1889:OND

[11666] Anonymous. Opening of a new dry dock at Norfolk, Va. Scientific American, 60(18):272, May 4, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n18/pdf/scientificamerican05041889-
272.pdf.

Anonymous:1889:HQF

[11667] Anonymous. Heavy quick-firing batteries. Scientific American, 60(18):272–273, May 4, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n18/pdf/scientificamerican05041889-
272a.pdf.

Anonymous:1889:IHa

[11668] Anonymous. An improved hinge. Scientific American, 60(18):274, May 4, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n18/pdf/scientificamerican05041889-274.pdf.

Anonymous:1889:IOCb

[11669] Anonymous. An improved oil can. Scientific American, 60(18):274, May 4, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n18/pdf/scientificamerican05041889-274a.pdf.

Anonymous:1889:ICCc

[11670] Anonymous. An improved car coupling. Scientific American, 60 (18):274, May 4, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n18/pdf/scientificamerican05041889-
274b.pdf.

Anonymous:1889:LHV

[11671] Anonymous. A lamp heater for vessels. Scientific American, 60(18):274, May 4, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n18/pdf/scientificamerican05041889-274c.pdf.

Anonymous:1889:ISCb

[11672] Anonymous. An improved seed cotton cleaner. Scientific American, 60(18):274, May 4, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n18/pdf/scientificamerican05041889-
274d.pdf.

Anonymous:1889:IFN

[11673] Anonymous. Improved feed or nose bag for horses. Scientific American, 60(18):274, May 4, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n18/pdf/scientificamerican05041889-
274e.pdf.

Anonymous:1889:EAM

[11674] Anonymous. Envelope addressing machine. Scientific American, 60(18):275, May 4, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n18/pdf/scientificamerican05041889-
275.pdf.

Anonymous:1889:GBH

[11675] Anonymous. A gas burner for heating purposes. Scientific American, 60(18):275, May 4, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n18/pdf/scientificamerican05041889-
275a.pdf.

Anonymous:1889:VLL

[11676] Anonymous. The Vyrnwy Lake of the Liverpool Water Works. Scientific American, 60(18):275–276, May 4, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v60/n18/pdf/scientificamerican05041889-
275b.pdf.

Holman:1889:TBS

[11677] George F. W. Holman. The torpedo boat Stiletto. Scientific American, 60(18):276–277, May 4, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n18/pdf/scientificamerican05041889-
276.pdf.

Anonymous:1889:LOT

[11678] Anonymous. Lake’s ocean tricycle at Atlantic City, N.J. Scientific American, 60(18):278, May 4, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n18/pdf/scientificamerican05041889-
278.pdf.

Anonymous:1889:CAA

[11679] Anonymous. Chinese alligators in the Aquarium at Berlin. Scientific American, 60(18):278, May 4, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n18/pdf/scientificamerican05041889-
278a.pdf.

Hopkins:1889:ASM

[11680] M. Hopkins. Apparatus for soldering and melting. Scientific American, 60(18):279–280, May 4, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n18/pdf/scientificamerican05041889-
279.pdf.

Anonymous:1889:IACa

[11681] Anonymous. An improved automatic clutch and tension machine. Scientific American, 60(18):281, May 4, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v60/
n18/pdf/scientificamerican05041889-281.pdf.

Anonymous:1889:Ik

[11682] Anonymous. To inventors. Scientific American, 60(18):283, May 4, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n18/pdf/scientificamerican05041889-283a.pdf.

Anonymous:1889:Ch

[11683] Anonymous. Correspondence. Scientific American, 60(19):292–293, May 11, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n19/pdf/scientificamerican05111889-292a.pdf.

Anonymous:1889:RPIk

[11684] Anonymous. Recently patented inventions: Recently patented inventions. Scientific American, 60(19):298, May 11, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v60/
n19/pdf/scientificamerican05111889-298.pdf.

Anonymous:1889:BPk

[11685] Anonymous. Business and personal. Scientific American, 60(19):298, May 11, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n19/pdf/scientificamerican05111889-298a.pdf.

Anonymous:1889:BNBf

[11686] Anonymous. Books: New books and publications. Scientific American, 60(19):298, May 11, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n19/pdf/scientificamerican05111889-
298b.pdf.

Anonymous:1889:NQk

[11687] Anonymous. Notes and queries. Scientific American, 60(19):298–299, May 11, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n19/pdf/scientificamerican05111889-298c.pdf.

Anonymous:1889:IIl

[11688] Anonymous. Index of inventions. Scientific American, 60(19):299–300, May 11, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n19/pdf/scientificamerican05111889-299a.pdf.

Anonymous:1889:PNPk

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n19/pdf/scientificamerican05111889-302.pdf.

Anonymous:1889:GSO

[11690] Anonymous. The grand scenery of Oregon. Scientific American, 60(19):287, May 11, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n19/pdf/scientificamerican05111889-
287.pdf.

Anonymous:1889:CIG

[11691] Anonymous. The centennial of the Inauguration of George Washingtion as President of the United States. Scientific American, 60(19):288–289, May 11, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n19/pdf/scientificamerican05111889-288.pdf.

Anonymous:1889:FW

[11692] Anonymous. Floating wrecks. Scientific American, 60(19):289, May 11, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n19/pdf/scientificamerican05111889-289.pdf.

Anonymous:1889:GPS

[11693] Anonymous. Gem pencil sharpener. Scientific American, 60(19):290, May 11, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n19/pdf/scientificamerican05111889-290.pdf.

Anonymous:1889:ITSa

[11694] Anonymous. An improved trousers strap. Scientific American, 60(19):290, May 11, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n19/pdf/scientificamerican05111889-
290a.pdf.

Anonymous:1889:CTS

[11695] Anonymous. A coupon ticket stamp and punch. Scientific American, 60(19):290, May 11, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n19/pdf/scientificamerican05111889-
290b.pdf.

Anonymous:1889:IIDa

[11696] Anonymous. Improved inking device for type writers. Scientific American, 60(19):290, May 11, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n19/pdf/scientificamerican05111889-
290c.pdf.

Anonymous:1889:IFSa

[11697] Anonymous. An improved folding step for vehicles. Scientific American, 60(19):290, May 11, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n19/pdf/scientificamerican05111889-
290d.pdf.

Anonymous:1889:MAO

[11698] Anonymous. A mirror attachment for opera glasses, etc. Scientific American, 60(19):291, May 11, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n19/pdf/scientificamerican05111889-
291.pdf.

Hopkins:1889:ME

[11699] M. Hopkins. A magnet for experimentation. Scientific American, 60(19):291, May 11, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n19/pdf/scientificamerican05111889-
291a.pdf.

Anonymous:1889:ICe

[11700] Anonymous. An improved cultivator. Scientific American, 60(19):291, May 11, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n19/pdf/scientificamerican05111889-291b.pdf.

Anonymous:1889:FBS

[11701] Anonymous. A fountain brush and sponge holder. Scientific American, 60(19):291, May 11, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n19/pdf/scientificamerican05111889-
291c.pdf.

Anonymous:1889:IPb

[11702] Anonymous. An improved padlock. Scientific American, 60(19):291, May 11, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n19/pdf/scientificamerican05111889-291d.pdf.

Anonymous:1889:SPSb

[11703] Anonymous. A steam-propelled sled. Scientific American, 60(19):292, May 11, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n19/pdf/scientificamerican05111889-292.pdf.

Anonymous:1889:TEE

[11704] Anonymous. Triple expansion engine. Scientific American, 60 (19):294–295, May 11, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n19/pdf/scientificamerican05111889-
294.pdf.

Anonymous:1889:VST

[11705] Anonymous. The venomous serpents or Toxicophidia of the United States. Scientific American, 60(19):295–296, May 11, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v60/
n19/pdf/scientificamerican05111889-295.pdf.

Anonymous:1889:CAM

[11706] Anonymous. Caudal appendage in man. Scientific American, 60 (19):296, May 11, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n19/pdf/scientificamerican05111889-
296.pdf.

Anonymous:1889:CPM

[11707] Anonymous. Crude petroleum for manufacturing purposes. Scientific American, 60(19):297, May 11, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n19/pdf/scientificamerican05111889-
297.pdf.

Anonymous:1889:Il

[11708] Anonymous. To inventors. Scientific American, 60(19):299, May 11, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n19/pdf/scientificamerican05111889-299.pdf.

Anonymous:1889:Ci

[11709] Anonymous. Correspondence. Scientific American, 60(2):21, January 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n2/pdf/scientificamerican01121889-21.pdf.

Anonymous:1889:RPIl

[11710] Anonymous. Recently patented inventions. Scientific American, 60(2):25, January 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n2/pdf/scientificamerican01121889-
25.pdf.

Anonymous:1889:BNBg

[11711] Anonymous. Books: New books and publications. Scientific American, 60(2):25–26, January 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n2/pdf/scientificamerican01121889-
25a.pdf.

Anonymous:1889:BPl

[11712] Anonymous. Business and personal. Scientific American, 60(2):26, January 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n2/pdf/scientificamerican01121889-26.pdf.

Anonymous:1889:NQl

[11713] Anonymous. Notes and queries. Scientific American, 60(2):26–27, January 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n2/pdf/scientificamerican01121889-26a.pdf.

Anonymous:1889:IIm

[11714] Anonymous. Index of inventions. Scientific American, 60(2):27–28, January 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n2/pdf/scientificamerican01121889-27a.pdf.

Anonymous:1889:PNPl

[11715] Anonymous. Patent notes: Patents. Scientific American, 60(2):30, January 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n2/pdf/scientificamerican01121889-30.pdf.

Anonymous:1889:SLa

[11716] Anonymous. The strong locomotive. Scientific American, 60(2):15, January 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n2/pdf/scientificamerican01121889-15.pdf.

Anonymous:1889:DED

[11717] Anonymous. The danger of electric distribution. Scientific American, 60(2):16–17, January 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n2/pdf/scientificamerican01121889-
16.pdf.

Anonymous:1889:SIO

[11718] Anonymous. Some interesting observations of Jupiter. Scientific American, 60(2):17, January 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n2/pdf/scientificamerican01121889-
17.pdf.

Anonymous:1889:IKT

[11719] Anonymous. Improved kitchen table and cabinet. Scientific American, 60(2):18, January 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n2/pdf/scientificamerican01121889-
18.pdf.

Anonymous:1889:IDMa

[11720] Anonymous. An improved ditching machine. Scientific American, 60(2):18, January 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n2/pdf/scientificamerican01121889-
18a.pdf.

Anonymous:1889:NBA

[11721] Anonymous. The new British armored war ship Australia. Scientific American, 60(2):19–20, January 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v60/n2/pdf/scientificamerican01121889-
19.pdf.

Anonymous:1889:ILL

[11722] Anonymous. An improved latch and lock. Scientific American, 60(2):20, January 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n2/pdf/scientificamerican01121889-
20.pdf.

Anonymous:1889:IPHa

[11723] Anonymous. An improved pen holder. Scientific American, 60 (2):20, January 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n2/pdf/scientificamerican01121889-
20a.pdf.

Anonymous:1889:IBW

[11724] Anonymous. An improved butter worker. Scientific American, 60(2):20, January 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n2/pdf/scientificamerican01121889-
20b.pdf.

Anonymous:1889:DRPb

[11725] Anonymous. Decisions relating to patents. Scientific American, 60(2):21, January 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n2/pdf/scientificamerican01121889-
21a.pdf.

Anonymous:1889:BWP

[11726] Anonymous. The barbed wire patent. Scientific American, 60 (2):21–22, January 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n2/pdf/scientificamerican01121889-
21b.pdf.

Anonymous:1889:HCS

[11727] Anonymous. A hand cart steam fire engine. Scientific American, 60(2):22, January 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n2/pdf/scientificamerican01121889-
22.pdf.

Hopkins:1889:SEPb

[11728] Geo. M. Hopkins. Simple experiments in physics. Scientific American, 60(2):23, January 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n2/pdf/scientificamerican01121889-
23.pdf.

Anonymous:1889:NDH

[11729] Anonymous. The new Dieulafoy Halls at the Louvre. Scientific American, 60(2):23–24, January 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n2/pdf/scientificamerican01121889-
23a.pdf.

Anonymous:1889:Im

[11730] Anonymous. To inventors. Scientific American, 60(2):27, January 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n2/pdf/scientificamerican01121889-27.pdf.

Anonymous:1889:RPIm

[11731] Anonymous. Recently patented inventions. Scientific American, 60(20):314, May 18, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n20/pdf/scientificamerican05181889-
314.pdf.

Anonymous:1889:BPm

[11732] Anonymous. Business and personal. Scientific American, 60(20):314, May 18, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n20/pdf/scientificamerican05181889-314a.pdf.

Anonymous:1889:NQm

[11733] Anonymous. Notes and queries. Scientific American, 60(20):314–315, May 18, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n20/pdf/scientificamerican05181889-314b.pdf.

Anonymous:1889:IIn

[11734] Anonymous. Index of inventions. Scientific American, 60(20):315–316, May 18, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n20/pdf/scientificamerican05181889-315a.pdf.

Anonymous:1889:PNPm

[11735] Anonymous. Patent notes: Patents. Scientific American, 60(20):318, May 18, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n20/pdf/scientificamerican05181889-318.pdf.

Anonymous:1889:WBN

[11736] Anonymous. The Washington Bridge, New York City. Scientific American, 60(20):303–304, May 18, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n20/pdf/scientificamerican05181889-
303.pdf.

Anonymous:1889:ROV

[11737] Anonymous. A remarkable ocean voyage. Scientific American, 60(20):304, May 18, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n20/pdf/scientificamerican05181889-
304.pdf.

Anonymous:1889:MCa

[11738] Anonymous. The Marine Conference. Scientific American, 60 (20):304–305, May 18, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n20/pdf/scientificamerican05181889-
304a.pdf.

Anonymous:1889:ITCb

[11739] Anonymous. An improved thill coupling. Scientific American, 60 (20):306, May 18, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n20/pdf/scientificamerican05181889-
306.pdf.

Anonymous:1889:ICV

[11740] Anonymous. An improved cut-off valve gear. Scientific American, 60(20):306, May 18, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n20/pdf/scientificamerican05181889-
306a.pdf.

Anonymous:1889:IRTa

[11741] Anonymous. An improved radial tube boiler. Scientific American, 60(20):306, May 18, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n20/pdf/scientificamerican05181889-
306b.pdf.

Anonymous:1889:ITB

[11742] Anonymous. Improved truss for buildings or bridges. Scientific American, 60(20):306, May 18, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n20/pdf/scientificamerican05181889-
306c.pdf.

Anonymous:1889:IKP

[11743] Anonymous. Improved knife for printers’ use. Scientific American, 60(20):306, May 18, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n20/pdf/scientificamerican05181889-
306d.pdf.

Anonymous:1889:IFBc

[11744] Anonymous. An improved folding book case. Scientific American, 60(20):307, May 18, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n20/pdf/scientificamerican05181889-
307.pdf.

Hopkins:1889:MCR

[11745] M. Hopkins. Making carbon rods and plates. Scientific American, 60(20):307, May 18, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n20/pdf/scientificamerican05181889-
307a.pdf.

Anonymous:1889:BCP

[11746] Anonymous. Band cutter platform for thrashing machines. Scientific American, 60(20):307–308, May 18, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v60/n20/pdf/scientificamerican05181889-
307b.pdf.

Anonymous:1889:IFT

[11747] Anonymous. An improved fish trap. Scientific American, 60(20):308, May 18, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n20/pdf/scientificamerican05181889-308.pdf.

Anonymous:1889:TES

[11748] Anonymous. The total eclipse of the Sun at Lick Observatory. Scientific American, 60(20):308, May 18, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v60/
n20/pdf/scientificamerican05181889-308a.pdf.

Anonymous:1889:IAI

[11749] Anonymous. An improved automatic injector. Scientific American, 60(20):308–309, May 18, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n20/pdf/scientificamerican05181889-
308b.pdf.

Anonymous:1889:OEP

[11750] Anonymous. The Otis elevators at the Paris Exposition. Scientific American, 60(20):310–311, May 18, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v60/n20/pdf/scientificamerican05181889-
310.pdf.

Anonymous:1889:ETa

[11751] Anonymous. The Eiffel Tower. Scientific American, 60(20):311, May 18, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n20/pdf/scientificamerican05181889-311.pdf.

Anonymous:1889:UE

[11752] Anonymous. The Universal Exposition of 1889. Scientific American, 60(20):311–313, May 18, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n20/pdf/scientificamerican05181889-
311a.pdf.

Anonymous:1889:PNf

[11753] Anonymous. Photographic notes. Scientific American, 60(20):313, May 18, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n20/pdf/scientificamerican05181889-313.pdf.

Anonymous:1889:In

[11754] Anonymous. To inventors. Scientific American, 60(20):315, May 18, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n20/pdf/scientificamerican05181889-315.pdf.

Anonymous:1889:Cj

[11755] Anonymous. Correspondence. Scientific American, 60(21):325, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n21/pdf/scientificamerican05251889-325.pdf.

Anonymous:1889:RPIn

[11756] Anonymous. Recently patented inventions. Scientific American, 60(21):330, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n21/pdf/scientificamerican05251889-
330.pdf.

Anonymous:1889:BPn

[11757] Anonymous. Business and personal. Scientific American, 60(21):330, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n21/pdf/scientificamerican05251889-330a.pdf.

Anonymous:1889:NQn

[11758] Anonymous. Notes and queries. Scientific American, 60(21):330–331, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n21/pdf/scientificamerican05251889-330b.pdf.

Anonymous:1889:BNBh

[11759] Anonymous. Books: New books and publications. Scientific American, 60(21):331, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n21/pdf/scientificamerican05251889-
331.pdf.

Anonymous:1889:PNPn

[11760] Anonymous. Patent notes: Patents. Scientific American, 60(21):334, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n21/pdf/scientificamerican05251889-334.pdf.

Haines:1889:DRL

[11761] Thornton F. Haines. Dredging and reclaiming land in Glorietta Bay, California. Scientific American, 60(21):319, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n21/pdf/scientificamerican05251889-319.pdf.

Anonymous:1889:V

[11762] Anonymous. The Valkyrie. Scientific American, 60(21):320, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n21/pdf/scientificamerican05251889-320.pdf.

Anonymous:1889:WIB

[11763] Anonymous. W. Irving Bishop, “mind reader”. Scientific American, 60(21):320–321, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n21/pdf/scientificamerican05251889-
320a.pdf.

Anonymous:1889:ALD

[11764] Anonymous. Automatic leak detector for pipes. Scientific American, 60(21):322, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n21/pdf/scientificamerican05251889-
322.pdf.

Anonymous:1889:IHCa

[11765] Anonymous. Improved hatchway closing device. Scientific American, 60(21):322, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n21/pdf/scientificamerican05251889-
322a.pdf.

Anonymous:1889:ESR

[11766] Anonymous. An electric striking and repeating clock. Scientific American, 60(21):322, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n21/pdf/scientificamerican05251889-
322b.pdf.

Anonymous:1889:RHT

[11767] Anonymous. A rivet holding tool for boilers. Scientific American, 60(21):322, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n21/pdf/scientificamerican05251889-
322c.pdf.

Anonymous:1889:DAB

[11768] Anonymous. Drying attachment for brick kilns. Scientific American, 60(21):323, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n21/pdf/scientificamerican05251889-
323.pdf.

Anonymous:1889:IRa

[11769] Anonymous. An improved reinholder. Scientific American, 60(21):323, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n21/pdf/scientificamerican05251889-323a.pdf.

Anonymous:1889:IBCc

[11770] Anonymous. An improved bundle carrier. Scientific American, 60(21):323, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n21/pdf/scientificamerican05251889-
323b.pdf.

Anonymous:1889:IHSc

[11771] Anonymous. An improved hammock support. Scientific American, 60(21):323, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n21/pdf/scientificamerican05251889-
323c.pdf.

Anonymous:1889:IBCd

[11772] Anonymous. An improved barrel cover. Scientific American, 60(21):323, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n21/pdf/scientificamerican05251889-323d.pdf.

Anonymous:1889:ITP

[11773] Anonymous. An improved transplanting pot. Scientific American, 60(21):323, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n21/pdf/scientificamerican05251889-
323e.pdf.

Anonymous:1889:IGSa

[11774] Anonymous. An improved gravel screen. Scientific American, 60 (21):324, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n21/pdf/scientificamerican05251889-
324.pdf.

Anonymous:1889:IHCb

[11775] Anonymous. Improved hose coupling for railway cars. Scientific American, 60(21):324, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n21/pdf/scientificamerican05251889-
324a.pdf.

Anonymous:1889:ITA

[11776] Anonymous. Improved truss for abdominal ruptures. Scientific American, 60(21):324, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n21/pdf/scientificamerican05251889-
324b.pdf.

Anonymous:1889:ISCc

[11777] Anonymous. An improved side curtain for buggies. Scientific American, 60(21):324, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n21/pdf/scientificamerican05251889-
324c.pdf.

Anonymous:1889:ABH

[11778] Anonymous. Air brake hose. Scientific American, 60(21):325, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n21/pdf/scientificamerican05251889-325a.pdf.

Anonymous:1889:FWS

[11779] Anonymous. The French war ship Amiral Baudin. Scientific American, 60(21):326, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n21/pdf/scientificamerican05251889-
326.pdf.

Anonymous:1889:FPB

[11780] Anonymous. Frederick A. P. Barnard. Scientific American, 60 (21):327–328, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n21/pdf/scientificamerican05251889-
327.pdf.

Halcomb:1889:CPL

[11781] N. W. Halcomb. A curiosity in plant life. Scientific American, 60(21):329, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n21/pdf/scientificamerican05251889-
329.pdf.

Hopkins:1889:VM

[11782] Geo. M. Hopkins. Vortex motion. Scientific American, 60(21):329, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n21/pdf/scientificamerican05251889-329a.pdf.

Anonymous:1889:Io

[11783] Anonymous. To inventors. Scientific American, 60(21):331, May 25, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n21/pdf/scientificamerican05251889-331a.pdf.

Anonymous:1889:RPIo

[11784] Anonymous. Recently patented inventions. Scientific American, 60(22):346, June 1, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n22/pdf/scientificamerican06011889-
346.pdf.

Anonymous:1889:BNBi 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n22/pdf/scientificamerican06011889-
346a.pdf.

Anonymous:1889:BPo

[11786] Anonymous. Business and personal. Scientific American, 60(22):347, June 1, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n22/pdf/scientificamerican06011889-347.pdf.

Anonymous:1889:NQo

[11787] Anonymous. Notes and queries. Scientific American, 60(22):347, June 1, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n22/pdf/scientificamerican06011889-347a.pdf.

Anonymous:1889:IIo

[11788] Anonymous. Index of inventions. Scientific American, 60(22):347–348, June 1, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n22/pdf/scientificamerican06011889-347c.pdf.

Anonymous:1889:PNPo

[11789] Anonymous. Patent notes: Patents. Scientific American, 60(22):350, June 1, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n22/pdf/scientificamerican06011889-350.pdf.

Anonymous:1889:AR

[11790] Anonymous. Alaska and its resources. Scientific American, 60(22):335, June 1, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n22/pdf/scientificamerican06011889-335.pdf.

Anonymous:1889:NSA

[11791] Anonymous. The new steamship Augusta Victoria. Scientific American, 60(22):336, June 1, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n22/pdf/scientificamerican06011889-
336.pdf.

Anonymous:1889:TDF

[11792] Anonymous. Talk of dismantling the forts. Scientific American, 60(22):336, June 1, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n22/pdf/scientificamerican06011889-
336a.pdf.

Anonymous:1889:PPJb

[11793] Anonymous. Position of the planets in June. Scientific American, 60(22):336–337, June 1, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n22/pdf/scientificamerican06011889-
336b.pdf.

Anonymous:1889:IGTa

[11794] Anonymous. An improved grip for tramways, etc. Scientific American, 60(22):338, June 1, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n22/pdf/scientificamerican06011889-
338.pdf.

Anonymous:1889:IMLa

[11795] Anonymous. An improved motor for light work. Scientific American, 60(22):338, June 1, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n22/pdf/scientificamerican06011889-
338a.pdf.

Anonymous:1889:ILCb

[11796] Anonymous. An improved lathe chuck. Scientific American, 60(22):338, June 1, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n22/pdf/scientificamerican06011889-338b.pdf.

Anonymous:1889:NSF

[11797] Anonymous. The new Syracuse fan. Scientific American, 60(22):338, June 1, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n22/pdf/scientificamerican06011889-338c.pdf.

Anonymous:1889:CSC

[11798] Anonymous. A collapsible shipping cask. Scientific American, 60(22):339, June 1, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n22/pdf/scientificamerican06011889-
339.pdf.

Anonymous:1889:IDPb

[11799] Anonymous. An improved dinner pail. Scientific American, 60(22):339, June 1, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n22/pdf/scientificamerican06011889-339a.pdf.

Anonymous:1889:IDH

[11800] Anonymous. An improved door hanger. Scientific American, 60 (22):339, June 1, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n22/pdf/scientificamerican06011889-
339b.pdf.

Hovey:1889:JC

[11801] H. C. Hovey. The Jewell Cavern. Scientific American, 60(22):339–341, June 1, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n22/pdf/scientificamerican06011889-339c.pdf.

Sloane:1889:SAB

[11802] T. O’conor Sloane. A simple air bath for laboratory use. Scientific American, 60(22):345, June 1, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n22/pdf/scientificamerican06011889-
345.pdf.

Anonymous:1889:Iq

[11803] Anonymous. To inventors. Scientific American, 60(22):347, June 1, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n22/pdf/scientificamerican06011889-347b.pdf.

Anonymous:1889:RPIp

[11804] Anonymous. Recently patented inventions: Recently patented intentions. Scientific American, 60(23):362, June 8, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v60/
n23/pdf/scientificamerican06081889-362.pdf.

Anonymous:1889:BPp

[11805] Anonymous. Business and personal. Scientific American, 60(23):362, June 8, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n23/pdf/scientificamerican06081889-362a.pdf.

Anonymous:1889:NQp

[11806] Anonymous. Notes and queries. Scientific American, 60(23):362–363, June 8, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n23/pdf/scientificamerican06081889-362b.pdf.

Anonymous:1889:IIp

[11807] Anonymous. Index of inventions. Scientific American, 60(23):363–364, June 8, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n23/pdf/scientificamerican06081889-363a.pdf.

Anonymous:1889:PNPp

[11808] Anonymous. Patent notes: Patents. Scientific American, 60(23):366, June 8, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n23/pdf/scientificamerican06081889-366.pdf.

Anonymous:1889:NLD

[11809] Anonymous. The new London Double Track Railroad Bridge and Draw Span. Scientific American, 60(23):351, June 8, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v60/
n23/pdf/scientificamerican06081889-351.pdf.

Anonymous:1889:ERS

[11810] Anonymous. An expected revolution in steam navigation. Scientific American, 60(23):352–353, June 8, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v60/n23/pdf/scientificamerican06081889-
352.pdf.

Anonymous:1889:IVCb

[11811] Anonymous. An improved vegetable cutter. Scientific American, 60(23):354, June 8, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n23/pdf/scientificamerican06081889-
354.pdf.

Anonymous:1889:IGBb

[11812] Anonymous. An improved guide for band-saws. Scientific American, 60(23):354, June 8, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n23/pdf/scientificamerican06081889-
354a.pdf.

Anonymous:1889:ISFa

[11813] Anonymous. An improved sash fastener. Scientific American, 60 (23):354, June 8, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n23/pdf/scientificamerican06081889-
354b.pdf.

Anonymous:1889:IPBa

[11814] Anonymous. An improved pinch bar. Scientific American, 60(23):354, June 8, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n23/pdf/scientificamerican06081889-354c.pdf.

Anonymous:1889:ISAa

[11815] Anonymous. An improved steam-actuated valve. Scientific American, 60(23):354, June 8, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n23/pdf/scientificamerican06081889-
354d.pdf.

Anonymous:1889:JML

[11816] Anonymous. Jerboas, marmots, and lemmings in the Zoological Garden of Berlin. Scientific American, 60(23):355–356, June 8, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n23/pdf/scientificamerican06081889-355.pdf.

Anonymous:1889:IFSb

[11817] Anonymous. An improved folding step-ladder. Scientific American, 60(23):357, June 8, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n23/pdf/scientificamerican06081889-
357.pdf.

Anonymous:1889:IASb

[11818] Anonymous. An improved air ship. Scientific American, 60(23):358, June 8, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n23/pdf/scientificamerican06081889-358.pdf.

Hopkins:1889:SMV

[11819] Geo. M. Hopkins. A siren for measuring velocities. Scientific American, 60(23):359–361, June 8, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n23/pdf/scientificamerican06081889-
359.pdf.

Sloane:1889:SWW

[11820] T. O’conor Sloane. A scientific will-o’-the-wisp. Scientific American, 60(23):361, June 8, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n23/pdf/scientificamerican06081889-
361.pdf.

Anonymous:1889:Is

[11821] Anonymous. To inventors. Scientific American, 60(23):363, June 8, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n23/pdf/scientificamerican06081889-363.pdf.

Anonymous:1889:Ck

[11822] Anonymous. Correspondence. Scientific American, 60(24):373, June 15, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n24/pdf/scientificamerican06151889-373.pdf.

Anonymous:1889:RPIq

[11823] Anonymous. Recently patented inventions. Scientific American, 60(24):378, June 15, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n24/pdf/scientificamerican06151889-
378.pdf.

Anonymous:1889:BPq

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n24/pdf/scientificamerican06151889-378a.pdf.

Anonymous:1889:NQq

[11825] Anonymous. Notes and queries. Scientific American, 60(24):378–379, June 15, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n24/pdf/scientificamerican06151889-378b.pdf.

Anonymous:1889:IIq

[11826] Anonymous. Index of inventions. Scientific American, 60(24):379–380, June 15, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n24/pdf/scientificamerican06151889-379a.pdf.

Anonymous:1889:PNPq

[11827] Anonymous. Patent notes: Patents. Scientific American, 60(24):382, June 15, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n24/pdf/scientificamerican06151889-382.pdf.

Anonymous:1889:TFM

[11828] Anonymous. The Tilly Foster mine. Scientific American, 60 (24):367–368, June 15, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n24/pdf/scientificamerican06151889-
367.pdf.

Anonymous:1889:AD

[11829] Anonymous. An appalling disaster. Scientific American, 60(24):369, June 15, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n24/pdf/scientificamerican06151889-369.pdf.

Anonymous:1889:IMT

[11830] Anonymous. An improved machine for turning tenons. Scientific American, 60(24):370, June 15, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n24/pdf/scientificamerican06151889-
370.pdf.

Anonymous:1889:IECb

[11831] Anonymous. An improved electric current regulator. Scientific American, 60(24):370, June 15, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n24/pdf/scientificamerican06151889-
370a.pdf.

Anonymous:1889:IDEb

[11832] Anonymous. An improved draught equalizer. Scientific American, 60(24):370, June 15, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n24/pdf/scientificamerican06151889-
370b.pdf.

Anonymous:1889:ACP

[11833] Anonymous. An automatic cable propeller for vessels. Scientific American, 60(24):370, June 15, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n24/pdf/scientificamerican06151889-
370c.pdf.

Anonymous:1889:DEF

[11834] Anonymous. A device for extinguishing fires on ships. Scientific American, 60(24):370, June 15, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n24/pdf/scientificamerican06151889-
370d.pdf.

Anonymous:1889:ICSb

[11835] Anonymous. An improved car seat. Scientific American, 60(24):371, June 15, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n24/pdf/scientificamerican06151889-371.pdf.

Anonymous:1889:ICRa

[11836] Anonymous. An improved compound rail. Scientific American, 60(24):371, June 15, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n24/pdf/scientificamerican06151889-
371a.pdf.

Anonymous:1889:ICCd

[11837] Anonymous. An improved cable car transfer system. Scientific American, 60(24):371, June 15, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n24/pdf/scientificamerican06151889-
371b.pdf.

Anonymous:1889:IECc

[11838] Anonymous. An improved extensible car step. Scientific American, 60(24):371–372, June 15, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n24/pdf/scientificamerican06151889-
371c.pdf.

Anonymous:1889:LBZ

[11839] Anonymous. Lynx in the Berlin Zoological Garden. Scientific American, 60(24):375, June 15, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n24/pdf/scientificamerican06151889-
375.pdf.

Anonymous:1889:ETb

[11840] Anonymous. The Eiffel Tower. Scientific American, 60(24):375–376, June 15, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n24/pdf/scientificamerican06151889-375a.pdf.

Anonymous:1889:AL

[11841] Anonymous. Artificial limbs. Scientific American, 60(24):377, June 15, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n24/pdf/scientificamerican06151889-377.pdf.

Anonymous:1889:Iu

[11842] Anonymous. To inventors. Scientific American, 60(24):379, June 15, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n24/pdf/scientificamerican06151889-379.pdf.

Anonymous:1889:Cl

[11843] Anonymous. Correspondence. Scientific American, 60(25):389, June 22, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n25/pdf/scientificamerican06221889-389.pdf.

Anonymous:1889:RPIr

[11844] Anonymous. Recently patented inventions. Scientific American, 60(25):394, June 22, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n25/pdf/scientificamerican06221889-
394.pdf.

Anonymous:1889:BPr

[11845] Anonymous. Business and personal. Scientific American, 60(25):394, June 22, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n25/pdf/scientificamerican06221889-394a.pdf.

Anonymous:1889:NQr

[11846] Anonymous. Notes and queries. Scientific American, 60(25):394–395, June 22, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n25/pdf/scientificamerican06221889-394b.pdf.

Anonymous:1889:IIr

[11847] Anonymous. Index of inventions. Scientific American, 60(25):395–396, June 22, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n25/pdf/scientificamerican06221889-395a.pdf.

Anonymous:1889:PNPr

[11848] Anonymous. Patent notes: Patents. Scientific American, 60(25):398, June 22, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n25/pdf/scientificamerican06221889-398.pdf.

Anonymous:1889:CSF

[11849] Anonymous. Collapse of the South Fork Dam and destruction of Johnstown. Scientific American, 60(25):383–391, June 22, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v60/
n25/pdf/scientificamerican06221889-383.pdf.

Anonymous:1889:AIW

[11850] Anonymous. American and Italian war ships. Scientific American, 60(25):384, June 22, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n25/pdf/scientificamerican06221889-
384.pdf.

Anonymous:1889:NM

[11851] Anonymous. The new mausoleum. Scientific American, 60(25):384, June 22, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n25/pdf/scientificamerican06221889-384a.pdf.

Anonymous:1889:TSM

[11852] Anonymous. Time signals by the magnesium light. Scientific American, 60(25):385, June 22, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n25/pdf/scientificamerican06221889-
385.pdf.

Anonymous:1889:DRPc

[11853] Anonymous. Decisions relating to patents. Scientific American, 60(25):385, June 22, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n25/pdf/scientificamerican06221889-
385a.pdf.

Anonymous:1889:IEG

[11854] Anonymous. An improved end gate for vehicles. Scientific American, 60(25):386, June 22, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n25/pdf/scientificamerican06221889-
386.pdf.

Anonymous:1889:IBCe

[11855] Anonymous. An improved bolting chest. Scientific American, 60 (25):386, June 22, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n25/pdf/scientificamerican06221889-
386a.pdf.

Anonymous:1889:BSS

[11856] Anonymous. A boot or shoe sole protector. Scientific American, 60(25):386, June 22, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n25/pdf/scientificamerican06221889-
386b.pdf.

Anonymous:1889:ILa

[11857] Anonymous. An improved lock. Scientific American, 60(25):386, June 22, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n25/pdf/scientificamerican06221889-386c.pdf.

Anonymous:1889:ELHa

[11858] Anonymous. Electric lamp hanger. Scientific American, 60(25):386, June 22, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n25/pdf/scientificamerican06221889-386d.pdf.

Anonymous:1889:FOC

[11859] Anonymous. A “fault” in an ocean cable. Scientific American, 60(25):387, June 22, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n25/pdf/scientificamerican06221889-
387.pdf.

Hopkins:1889:ALU

[11860] Geo. M. Hopkins. Aspirators for laboratory use. Scientific American, 60(25):387–388, June 22, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n25/pdf/scientificamerican06221889-
387a.pdf.

Anonymous:1889:PNg

[11861] Anonymous. Photographic notes. Scientific American, 60(25):392–393, June 22, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n25/pdf/scientificamerican06221889-392.pdf.

Anonymous:1889:UHL

[11862] Anonymous. Universal hand lathe. Scientific American, 60(25):393, June 22, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n25/pdf/scientificamerican06221889-393.pdf.

Anonymous:1889:NMS

[11863] Anonymous. New micro-surface gauge. Scientific American, 60(25):393, June 22, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n25/pdf/scientificamerican06221889-393a.pdf.

Anonymous:1889:Iv

[11864] Anonymous. To inventors. Scientific American, 60(25):395, June 22, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n25/pdf/scientificamerican06221889-395.pdf.

Anonymous:1889:RPIs

[11865] Anonymous. Recently patented inventions. Scientific American, 60(26):409, June 29, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n26/pdf/scientificamerican06291889-
409.pdf.

Anonymous:1889:BPs

[11866] Anonymous. Business and personal. Scientific American, 60(26):409, June 29, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n26/pdf/scientificamerican06291889-409a.pdf.

Anonymous:1889:IIs

[11867] Anonymous. Index of inventions. Scientific American, 60(26):409–410, June 29, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n26/pdf/scientificamerican06291889-409b.pdf.

Anonymous:1889:Iw

[11868] Anonymous. Index. Scientific American, 60(26):411–412, June 29, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n26/pdf/scientificamerican06291889-411.pdf.

Anonymous:1889:PNPs

[11869] Anonymous. Patent notes: Patents. Scientific American, 60(26):414, June 29, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n26/pdf/scientificamerican06291889-414.pdf.

Anonymous:1889:WSNa

[11870] Anonymous. The water supply of New York. Scientific American, 60(26):400, June 29, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n26/pdf/scientificamerican06291889-
400.pdf.

Anonymous:1889:DC

[11871] Anonymous. Dynamo and compass. Scientific American, 60(26):400, June 29, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n26/pdf/scientificamerican06291889-400a.pdf.

Anonymous:1889:PPJc

[11872] Anonymous. Position of the planets in July. Scientific American, 60(26):401, June 29, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n26/pdf/scientificamerican06291889-
401.pdf.

Anonymous:1889:IFJ

[11873] Anonymous. An improved flooring jack. Scientific American, 60 (26):402, June 29, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n26/pdf/scientificamerican06291889-
402.pdf.

Anonymous:1889:TB

[11874] Anonymous. Torpedo boats. Scientific American, 60(26):402, June 29, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n26/pdf/scientificamerican06291889-402a.pdf.

Anonymous:1889:IDSb

[11875] Anonymous. An improved drawbridge signal. Scientific American, 60(26):402, June 29, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n26/pdf/scientificamerican06291889-
402b.pdf.

Anonymous:1889:LHS

[11876] Anonymous. A lock-house for switch-stands. Scientific American, 60(26):402, June 29, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n26/pdf/scientificamerican06291889-
402c.pdf.

Anonymous:1889:ISVa

[11877] Anonymous. An improved sunshade for vehicles. Scientific American, 60(26):403, June 29, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n26/pdf/scientificamerican06291889-
403.pdf.

Anonymous:1889:SAFa

[11878] Anonymous. A stovepipe anchor and flue cleaner. Scientific American, 60(26):403, June 29, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n26/pdf/scientificamerican06291889-
403a.pdf.

Anonymous:1889:IACb

[11879] Anonymous. An improved animal catcher. Scientific American, 60(26):403, June 29, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n26/pdf/scientificamerican06291889-
403b.pdf.

Anonymous:1889:ISBa

[11880] Anonymous. An improved spring back for stools, etc. Scientific American, 60(26):403, June 29, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n26/pdf/scientificamerican06291889-
403c.pdf.

Anonymous:1889:CFM

[11881] Anonymous. Crystallizing frames for monograms, designs, etc. Scientific American, 60(26):403, June 29, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v60/n26/pdf/scientificamerican06291889-
403d.pdf.

Anonymous:1889:ICCe

[11882] Anonymous. An improved car coupling. Scientific American, 60 (26):403–404, June 29, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n26/pdf/scientificamerican06291889-
403e.pdf.

Anonymous:1889:SSSa

[11883] Anonymous. A spindle support for spinning machines. Scientific American, 60(26):404, June 29, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n26/pdf/scientificamerican06291889-
404.pdf.

Anonymous:1889:PIC

[11884] Anonymous. Palmer’s improved carpenter’s chisel. Scientific American, 60(26):404, June 29, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n26/pdf/scientificamerican06291889-
404a.pdf.

Anonymous:1889:ICCf

[11885] Anonymous. An improved car coupling. Scientific American, 60 (26):404, June 29, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n26/pdf/scientificamerican06291889-
404b.pdf.

Anonymous:1889:MRI

[11886] Anonymous. Mississippi River improvements. Scientific American, 60(26):405, June 29, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n26/pdf/scientificamerican06291889-
405.pdf.

Anonymous:1889:JD

[11887] Anonymous. The Johnstown Disaster. Scientific American, 60 (26):406–408, June 29, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n26/pdf/scientificamerican06291889-
406.pdf.

Anonymous:1889:Cm

[11888] Anonymous. Correspondence. Scientific American, 60(3):37, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n3/pdf/scientificamerican01191889-37.pdf.

Anonymous:1889:RPIt

[11889] Anonymous. Recently patented inventions: Recently patented intentions. Scientific American, 60(3):42, January 19, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v60/
n3/pdf/scientificamerican01191889-42.pdf.

Anonymous:1889:BPt

[11890] Anonymous. Business and personal. Scientific American, 60(3):42, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n3/pdf/scientificamerican01191889-42a.pdf.

Anonymous:1889:BNBj

[11891] Anonymous. Books: New books and publications. Scientific American, 60(3):42, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n3/pdf/scientificamerican01191889-
42b.pdf.

Anonymous:1889:NQs

[11892] Anonymous. Notes and queries. Scientific American, 60(3):42–43, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n3/pdf/scientificamerican01191889-42c.pdf.

Anonymous:1889:IIt

[11893] Anonymous. Index of inventions. Scientific American, 60(3):43–44, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n3/pdf/scientificamerican01191889-43a.pdf.

Anonymous:1889:PNPt

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n3/pdf/scientificamerican01191889-46.pdf.

Hovey:1889:TNB

[11895] H. C. Hovey. Test of a new bridge by means of elephants, at Bridgeport, Conn. Scientific American, 60(3):31, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n3/pdf/scientificamerican01191889-31.pdf.

Anonymous:1889:CDC

[11896] Anonymous. Coupling a dynamo to a car-axle. Scientific American, 60(3):32, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n3/pdf/scientificamerican01191889-
32.pdf.

Anonymous:1889:TCH

[11897] Anonymous. Are townsmen or countrymen the healthiest? Scientific American, 60(3):32, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n3/pdf/scientificamerican01191889-
32a.pdf.

Anonymous:1889:PSC

[11898] Anonymous. Postal street cars. Scientific American, 60(3):32–33, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n3/pdf/scientificamerican01191889-32b.pdf.

Anonymous:1889:DWR

[11899] Anonymous. The deep wells of Reading, Pa. Scientific American, 60(3):33, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n3/pdf/scientificamerican01191889-
33.pdf.

Anonymous:1889:PNh

[11900] Anonymous. Photographic notes. Scientific American, 60(3): 33–34, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n3/pdf/scientificamerican01191889-
33a.pdf.

Anonymous:1889:ILCc

[11901] Anonymous. Improved lifting cage for mines. Scientific American, 60(3):34, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n3/pdf/scientificamerican01191889-
34.pdf.

Anonymous:1889:IDR

[11902] Anonymous. Improved double-reciprocating engine. Scientific American, 60(3):34, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n3/pdf/scientificamerican01191889-
34a.pdf.

Anonymous:1889:FSP

[11903] Anonymous. A foot-shield for pianos, etc. Scientific American, 60(3):34, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n3/pdf/scientificamerican01191889-
34b.pdf.

Anonymous:1889:ISDb

[11904] Anonymous. An improved screw driver. Scientific American, 60 (3):35, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n3/pdf/scientificamerican01191889-
35.pdf.

Anonymous:1889:IPW

[11905] Anonymous. Improved pulsatory wool washing machine. Scientific American, 60(3):35, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n3/pdf/scientificamerican01191889-
35a.pdf.

Anonymous:1889:NSCa

[11906] Anonymous. A new screw cutting die. Scientific American, 60 (3):35, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n3/pdf/scientificamerican01191889-
35b.pdf.

Anonymous:1889:IRGa

[11907] Anonymous. An improved road grader and ditcher. Scientific American, 60(3):35, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n3/pdf/scientificamerican01191889-
35c.pdf.

Anonymous:1889:BPW

[11908] Anonymous. The Bartlett project for a water supply for lower New York. Scientific American, 60(3):36, January 19, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v60/
n3/pdf/scientificamerican01191889-36.pdf.

Anonymous:1889:ICO

[11909] Anonymous. An improved can opener. Scientific American, 60 (3):36, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n3/pdf/scientificamerican01191889-
36a.pdf.

Anonymous:1889:BI

[11910] Anonymous. A boy’s invention. Scientific American, 60(3):37, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n3/pdf/scientificamerican01191889-37a.pdf.

Anonymous:1889:GRB

[11911] Anonymous. The great railway bridge over the Firth of Forth. Scientific American, 60(3):38–39, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v60/n3/pdf/scientificamerican01191889-
38.pdf.

Anonymous:1889:SLS

[11912] Anonymous. St. louis to be a seaport. Scientific American, 60 (3):39, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n3/pdf/scientificamerican01191889-
39.pdf.

Anonymous:1889:ECF

[11913] Anonymous. Experiments in capillary force. Scientific American, 60(3):39–40, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n3/pdf/scientificamerican01191889-
39a.pdf.

Hopkins:1889:ACR

[11914] M. Hopkins. Apparatus for compounding rectangular vibrations. Scientific American, 60(3):41, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v60/n3/pdf/scientificamerican01191889-
41.pdf.

Anonymous:1889:Ix

[11915] Anonymous. To inventors. Scientific American, 60(3):43, January 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n3/pdf/scientificamerican01191889-43.pdf.

Anonymous:1889:Cn

[11916] Anonymous. Correspondence. Scientific American, 60(4):53, January 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n4/pdf/scientificamerican01261989-53.pdf.

Anonymous:1889:RPIu

[11917] Anonymous. Recently patented inventions. Scientific American, 60(4):57, January 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n4/pdf/scientificamerican01261989-
57.pdf.

Anonymous:1889:BPu

[11918] Anonymous. Business and personal. Scientific American, 60(4):57, January 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n4/pdf/scientificamerican01261989-57a.pdf.

Anonymous:1889:NQt

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n4/pdf/scientificamerican01261989-57b.pdf.

Anonymous:1889:IIu

[11920] Anonymous. Index of inventions. Scientific American, 60(4):59–60, January 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n4/pdf/scientificamerican01261989-59a.pdf.

Anonymous:1889:PNPu

[11921] Anonymous. Patent notes: Patents. Scientific American, 60(4):62, January 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n4/pdf/scientificamerican01261989-62.pdf.

Anonymous:1889:ESCb

[11922] Anonymous. The engines of the steamer Connecticut. Scientific American, 60(4):47, January 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n4/pdf/scientificamerican01261989-
47.pdf.

Anonymous:1889:ECC

[11923] Anonymous. An electrical course at Columbia. Scientific American, 60(4):48, January 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n4/pdf/scientificamerican01261989-
48.pdf.

Anonymous:1889:NEC

[11924] Anonymous. A novel experiment with criminals. Scientific American, 60(4):48, January 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n4/pdf/scientificamerican01261989-
48a.pdf.

Anonymous:1889:DFH

[11925] Anonymous. Dangerous flat houses. Scientific American, 60(4):48, January 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n4/pdf/scientificamerican01261989-48b.pdf.

Anonymous:1889:URH

[11926] Anonymous. The use of radiated heat. Scientific American, 60 (4):48, January 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n4/pdf/scientificamerican01261989-
48c.pdf.

Anonymous:1889:PPF

[11927] Anonymous. Position of the planets in February. Scientific American, 60(4):49, January 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n4/pdf/scientificamerican01261989-
49.pdf.

Anonymous:1889:RTN

[11928] Anonymous. The recent tornado in New York. Scientific American, 60(4):49, January 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n4/pdf/scientificamerican01261989-
49a.pdf.

Anonymous:1889:RHC

[11929] Anonymous. Revolving Hercules crane. Scientific American, 60 (4):50, January 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n4/pdf/scientificamerican01261989-
50.pdf.

Anonymous:1889:IVSb

[11930] Anonymous. An improved vehicle spring. Scientific American, 60(4):51, January 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n4/pdf/scientificamerican01261989-
51.pdf.

Anonymous:1889:IGM

[11931] Anonymous. An improved gas manufacturing apparatus. Scientific American, 60(4):51, January 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n4/pdf/scientificamerican01261989-
51a.pdf.

Anonymous:1889:ISCd

[11932] Anonymous. An improved safety caster. Scientific American, 60(4):51, January 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n4/pdf/scientificamerican01261989-
51b.pdf.

Anonymous:1889:IWI

[11933] Anonymous. Improved wear iron for vehicles. Scientific American, 60(4):51–52, January 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n4/pdf/scientificamerican01261989-
51c.pdf.

Anonymous:1889:MSD

[11934] Anonymous. Mlle Schultze, Doctor of Medicine of the Faculty of Paris. Scientific American, 60(4):54, January 26, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v60/
n4/pdf/scientificamerican01261989-54.pdf.

Hopkins:1889:SEPc

[11935] M. Hopkins. Simple experiments in physics. Scientific American, 60(4):54, January 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n4/pdf/scientificamerican01261989-
54a.pdf.

Hovey:1889:LMO

[11936] H. C. Hovey. The Leander McCormick Observatory of the University of Virginia. Scientific American, 60(4):55–56, January 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n4/pdf/scientificamerican01261989-55.pdf.

Anonymous:1889:TTV

[11937] Anonymous. Trial trip of the Vesuvius. Scientific American, 60(4):56, January 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n4/pdf/scientificamerican01261989-
56.pdf.

Anonymous:1889:Iy

[11938] Anonymous. To inventors. Scientific American, 60(4):59, January 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n4/pdf/scientificamerican01261989-59.pdf.

Anonymous:1889:Co

[11939] Anonymous. Correspondence. Scientific American, 60(5):69, February 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n5/pdf/scientificamerican02021889-69a.pdf.

Anonymous:1889:RPIv

[11940] Anonymous. Recently patented inventions. Scientific American, 60(5):73, February 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n5/pdf/scientificamerican02021889-
73.pdf.

Anonymous:1889:BPv

[11941] Anonymous. Business and personal. Scientific American, 60 (5):73, February 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n5/pdf/scientificamerican02021889-
73a.pdf.

Anonymous:1889:NQu

[11942] Anonymous. Notes and queries. Scientific American, 60(5):73– 75, February 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n5/pdf/scientificamerican02021889-
73b.pdf.

Anonymous:1889:BNBk

[11943] Anonymous. Books: New books and publications. Scientific American, 60(5):75, February 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n5/pdf/scientificamerican02021889-
75.pdf.

Anonymous:1889:IIv (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n5/pdf/scientificamerican02021889-
75b.pdf.

Anonymous:1889:PNPv

[11945] Anonymous. Patent notes: Patents. Scientific American, 60 (5):78, February 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n5/pdf/scientificamerican02021889-
78.pdf.

Anonymous:1889:CAP

[11946] Anonymous. Complaints against the Patent Office. Scientific American, 60(5):64, February 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n5/pdf/scientificamerican02021889-
64.pdf.

Anonymous:1889:ISCe

[11947] Anonymous. An important Supreme Court decision. Scientific American, 60(5):64–65, February 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n5/pdf/scientificamerican02021889-
64a.pdf.

Anonymous:1889:IDU

[11948] Anonymous. Improved device for use in paying help. Scientific American, 60(5):66, February 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n5/pdf/scientificamerican02021889-
66.pdf.

Anonymous:1889:IFDb

[11949] Anonymous. Improved furnace for destroying refuse. Scientific American, 60(5):66, February 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n5/pdf/scientificamerican02021889-
66a.pdf.

Anonymous:1889:IPBb

[11950] Anonymous. An improved pillow bolster. Scientific American, 60(5):66, February 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n5/pdf/scientificamerican02021889-
66b.pdf.

Anonymous:1889:IFCa

[11951] Anonymous. An improved folding car step. Scientific American, 60(5):66–68, February 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n5/pdf/scientificamerican02021889-
66c.pdf.

Sheldon:1889:EB

[11952] Samuel Sheldon. The electric blowpipe. Scientific American, 60 (5):68, February 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n5/pdf/scientificamerican02021889-
68.pdf.

Anonymous:1889:PEb

[11953] Anonymous. The Paris Exhibition. Scientific American, 60(5): 68–69, February 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n5/pdf/scientificamerican02021889-
68a.pdf.

Anonymous:1889:UGY

[11954] Anonymous. U.S. gunboat Yorktown. Scientific American, 60 (5):69, February 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n5/pdf/scientificamerican02021889-
69.pdf.

Anonymous:1889:REL

[11955] Anonymous. Remarkable explosion of a locomotive. Scientific American, 60(5):70, February 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n5/pdf/scientificamerican02021889-
70.pdf.

Anonymous:1889:CV

[11956] Anonymous. A continental villa. Scientific American, 60(5):70, February 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n5/pdf/scientificamerican02021889-70a.pdf.

Hopkins:1889:SEPd

[11957] Geo. M. Hopkins. Simple experiments in physics. Scientific American, 60(5):71–72, February 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n5/pdf/scientificamerican02021889-
71.pdf.

Anonymous:1889:Iz

[11958] Anonymous. To inventors. Scientific American, 60(5):75, February 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n5/pdf/scientificamerican02021889-75a.pdf.

Anonymous:1889:RPIw

[11959] Anonymous. Recently patented inventions. Scientific American, 60(6):90, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
90.pdf.

Anonymous:1889:BPx

[11960] Anonymous. Business and personal. Scientific American, 60 (6):90, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
90a.pdf.

Anonymous:1889:NQv

[11961] Anonymous. Notes and queries. Scientific American, 60(6):90– 91, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
90b.pdf.

Anonymous:1889:BNBl

[11962] Anonymous. Books: New books and publications. Scientific American, 60(6):91, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
91.pdf.

Anonymous:1889:IIw

[11963] Anonymous. Index of inventions. Scientific American, 60(6): 91–92, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
91b.pdf.

Anonymous:1889:PNPw

[11964] Anonymous. Patent notes: Patents. Scientific American, 60 (6):94, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
94.pdf.

Anonymous:1889:RAC

[11965] Anonymous. A rumor about the compass. Scientific American, 60(6):80, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
80.pdf.

Anonymous:1889:RMG

[11966] Anonymous. “Recipes for making gout”. Scientific American, 60(6):80, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
80a.pdf.

Anonymous:1889:HVI

[11967] Anonymous. Hemp vs. iron in ocean cables. Scientific American, 60(6):80, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
80b.pdf.

Anonymous:1889:MGS

[11968] Anonymous. Machine guns in shore defense. Scientific American, 60(6):80, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
80c.pdf.

Anonymous:1889:NWFb

[11969] Anonymous. Naval wars of the future. Scientific American, 60 (6):80–81, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
80d.pdf.

Anonymous:1889:ITLb

[11970] Anonymous. An improved track laying apparatus. Scientific American, 60(6):82, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
82.pdf.

Anonymous:1889:MGP

[11971] Anonymous. The Machine Gallery, Paris Exhibition. Scientific American, 60(6):82, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
82a.pdf.

Anonymous:1889:IMR

[11972] Anonymous. An improved metallic railroad tie. Scientific American, 60(6):82, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
82b.pdf.

Anonymous:1889:NWP

[11973] Anonymous. A new water power developing device. Scientific American, 60(6):83, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
83.pdf.

Anonymous:1889:IWB

[11974] Anonymous. An improved window burglar guard. Scientific American, 60(6):83, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
83a.pdf.

Anonymous:1889:ICf

[11975] Anonymous. An improved cultivator. Scientific American, 60 (6):83, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
83b.pdf.

Anonymous:1889:IMCa

[11976] Anonymous. Improved mill for crushing sugar cane. Scientific American, 60(6):83, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
83c.pdf.

Anonymous:1889:ILSa

[11977] Anonymous. An improved life-saving net. Scientific American, 60(6):83–84, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
83d.pdf.

Hopkins:1889:MPD

[11978] Geo. M. Hopkins. Method of producing designs on wire cloth. Scientific American, 60(6):84, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
84.pdf.

Anonymous:1889:ITSb

[11979] Anonymous. An improved tobacco smoking pipe. Scientific American, 60(6):84, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
84a.pdf.

Anonymous:1889:BEL

[11980] Anonymous. The Ball electric light system. Scientific American, 60(6):85, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
85.pdf.

Anonymous:1889:ISVb

[11981] Anonymous. An improved sewer ventilator. Scientific American, 60(6):86, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
86.pdf.

Anonymous:1889:APP

[11982] Anonymous. The American patent pop safety valve. Scientific American, 60(6):86, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
86a.pdf.

Anonymous:1889:IAP

[11983] Anonymous. An indicator attachment for passenger elevators. Scientific American, 60(6):86, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
86b.pdf.

Anonymous:1889:GRM

[11984] Anonymous. The great rolling mill of the St. Jacques Works. Scientific American, 60(6):87–88, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
87.pdf.

Anonymous:1889:IBe

[11985] Anonymous. An improved bicycle. Scientific American, 60(6):89, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n6/pdf/scientificamerican02091889-89.pdf.

Anonymous:1889:CHC

[11986] Anonymous. Chesley heal, the centenarian. Scientific American, 60(6):89, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n6/pdf/scientificamerican02091889-
89a.pdf.

Anonymous:1889:Ibf

[11987] Anonymous. To inventors. Scientific American, 60(6):91, February 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n6/pdf/scientificamerican02091889-91a.pdf.

Anonymous:1889:RPIx

[11988] Anonymous. Recently patented inventions. Scientific American, 60(7):106, February 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n7/pdf/scientificamerican02161889-
106.pdf.

Anonymous:1889:BPy

[11989] Anonymous. Business and personal. Scientific American, 60(7): 106, February 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n7/pdf/scientificamerican02161889-
106a.pdf.

Anonymous:1889:NQw

[11990] Anonymous. Notes and queries. Scientific American, 60(7):106– 107, February 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n7/pdf/scientificamerican02161889-
106b.pdf.

Anonymous:1889:IIx

[11991] Anonymous. Index of inventions. Scientific American, 60(7): 107–108, February 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n7/pdf/scientificamerican02161889-
107a.pdf.

Anonymous:1889:PNPx

[11992] Anonymous. Patent notes: Patents. Scientific American, 60 (7):110, February 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n7/pdf/scientificamerican02161889-
110.pdf.

Anonymous:1889:NSCb

[11993] Anonymous. The Nicaragua Ship Canal. Scientific American, 60(7):95, February 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n7/pdf/scientificamerican02161889-
95.pdf.

Anonymous:1889:CWb

[11994] Anonymous. The celestial world. Scientific American, 60(7):96, February 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n7/pdf/scientificamerican02161889-96.pdf.

Anonymous:1889:PNAa

[11995] Anonymous. Progress of new armaments. Scientific American, 60(7):96, February 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n7/pdf/scientificamerican02161889-
96a.pdf.

Anonymous:1889:MEL

[11996] Anonymous. Meeting of electric light men. Scientific American, 60(7):96, February 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n7/pdf/scientificamerican02161889-
96b.pdf.

Anonymous:1889:EWG

[11997] Anonymous. Electric wires in gas mains. Scientific American, 60 (7):96–97, February 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n7/pdf/scientificamerican02161889-
96c.pdf.

Anonymous:1889:ILBb

[11998] Anonymous. An improved locomotive boiler. Scientific American, 60(7):98, February 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n7/pdf/scientificamerican02161889-
98.pdf.

Anonymous:1889:EEF

[11999] Anonymous. Elastic engine foundations and suspension of vehicles. Scientific American, 60(7):98, February 16, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v60/
n7/pdf/scientificamerican02161889-98a.pdf.

Anonymous:1889:WSB

[12000] Anonymous. Weeke’s system of building dikes. Scientific American, 60(7):98–99, February 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n7/pdf/scientificamerican02161889-
98b.pdf.

Anonymous:1889:BHB

[12001] Anonymous. A box to hold bricks of ice cream. Scientific American, 60(7):99, February 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n7/pdf/scientificamerican02161889-
99.pdf.

Anonymous:1889:IES

[12002] Anonymous. An improved exhibitor for shades, maps, etc. Scientific American, 60(7):99, February 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n7/pdf/scientificamerican02161889-
99a.pdf.

Anonymous:1889:ICCg

[12003] Anonymous. An improved car coupling. Scientific American, 60(7):99, February 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n7/pdf/scientificamerican02161889-
99b.pdf.

Anonymous:1889:KCB

[12004] Anonymous. Kane’s car brake. Scientific American, 60(7):99, February 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n7/pdf/scientificamerican02161889-99c.pdf.

Anonymous:1889:ASP

[12005] Anonymous. Ancel’s stove pipe and tent support. Scientific American, 60(7):99, February 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n7/pdf/scientificamerican02161889-
99d.pdf.

Anonymous:1889:IHCc

[12006] Anonymous. An improved hose coupling. Scientific American, 60 (7):99–100, February 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n7/pdf/scientificamerican02161889-
99e.pdf.

Anonymous:1889:IFEa

[12007] Anonymous. An improved fire escape. Scientific American, 60 (7):100, February 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n7/pdf/scientificamerican02161889-
100.pdf.

Anonymous:1889:ISCf

[12008] Anonymous. An improved sod cutter and harrow. Scientific American, 60(7):100, February 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n7/pdf/scientificamerican02161889-
100a.pdf.

Anonymous:1889:DRPd

[12009] Anonymous. Decisions relating to patents and trade marks. Scientific American, 60(7):101, February 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v60/n7/pdf/scientificamerican02161889-
101.pdf.

Maxim:1889:MAM

[12010] Hudson Maxim. The Maxim automatic machine gun. Scientific American, 60(7):102–104, February 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n7/pdf/scientificamerican02161889-
102.pdf.

Anonymous:1889:Ibg

[12011] Anonymous. To inventors. Scientific American, 60(7):107, February 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n7/pdf/scientificamerican02161889-107.pdf.

Anonymous:1889:Cp

[12012] Anonymous. Correspondence. Scientific American, 60(8):117, February 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n8/pdf/scientificamerican02231889-117.pdf.

Anonymous:1889:RPIy

[12013] Anonymous. Recently patented inventions. Scientific American, 60(8):122, February 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n8/pdf/scientificamerican02231889-
122.pdf.

Anonymous:1889:BPz

[12014] Anonymous. Business and personal. Scientific American, 60(8): 122, February 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n8/pdf/scientificamerican02231889-
122a.pdf.

Anonymous:1889:NQx

[12015] Anonymous. Notes and queries. Scientific American, 60(8):122– 123, February 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n8/pdf/scientificamerican02231889-
122b.pdf.

Anonymous:1889:BNBm

[12016] Anonymous. Books: New books and publications. Scientific American, 60(8):123, February 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n8/pdf/scientificamerican02231889-
123.pdf.

Anonymous:1889:IIy 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n8/pdf/scientificamerican02231889-
123b.pdf.

Anonymous:1889:PNPy

[12018] Anonymous. Patent notes: Patents. Scientific American, 60 (8):126, February 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n8/pdf/scientificamerican02231889-
126.pdf.

Anonymous:1889:RCP

[12019] Anonymous. Report of the Commissioner of Patents for 1888. Scientific American, 60(8):112, February 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v60/n8/pdf/scientificamerican02231889-
112.pdf.

Anonymous:1889:RNY

[12020] Anonymous. A reply to the New York “World”. Scientific American, 60(8):112, February 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n8/pdf/scientificamerican02231889-
112a.pdf.

Anonymous:1889:OTG

[12021] Anonymous. Official trial of the gunboat Yorktown. Scientific American, 60(8):112–113, February 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n8/pdf/scientificamerican02231889-
112b.pdf.

Anonymous:1889:IWTb

[12022] Anonymous. An improved wire tightener. Scientific American, 60(8):114, February 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n8/pdf/scientificamerican02231889-
114.pdf.

Anonymous:1889:DPE

[12023] Anonymous. A device for piercing ears and inserting earrings. Scientific American, 60(8):114, February 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v60/n8/pdf/scientificamerican02231889-
114a.pdf.

Anonymous:1889:IBSb

[12024] Anonymous. An improved brake shoe for railroad cars. Scientific American, 60(8):114, February 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n8/pdf/scientificamerican02231889-
114b.pdf.

Anonymous:1889:ISHa

[12025] Anonymous. An improved steam heater. Scientific American, 60(8):114, February 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n8/pdf/scientificamerican02231889-
114c.pdf.

Anonymous:1889:DSS

[12026] Anonymous. A device for sharpening stone-cutting tools. Scientific American, 60(8):114, February 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v60/n8/pdf/scientificamerican02231889-
114d.pdf.

Anonymous:1889:IGSb

[12027] Anonymous. An improved guard for step ladders. Scientific American, 60(8):115, February 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n8/pdf/scientificamerican02231889-
115.pdf.

Anonymous:1889:IFEb

[12028] Anonymous. An improved fire escape. Scientific American, 60 (8):115, February 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n8/pdf/scientificamerican02231889-
115a.pdf.

Anonymous:1889:IAa

[12029] Anonymous. An improved axle. Scientific American, 60(8):115, February 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n8/pdf/scientificamerican02231889-115b.pdf.

Anonymous:1889:IRb

[12030] Anonymous. An improved railway. Scientific American, 60(8): 115, February 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n8/pdf/scientificamerican02231889-
115c.pdf.

Anonymous:1889:IPCc

[12031] Anonymous. An improved plow clevis. Scientific American, 60 (8):115, February 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n8/pdf/scientificamerican02231889-
115d.pdf.

Rohrer:1889:GCH

[12032] Karl Rohrer. Gun cotton — its history, manufacture, use. Scientific American, 60(8):116–117, February 23, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v60/
n8/pdf/scientificamerican02231889-116.pdf.

Anonymous:1889:TCS

[12033] Anonymous. Test of the cast steel breech loading rifle of open hearth steel. Scientific American, 60(8):118–119, February 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n8/pdf/scientificamerican02231889-118.pdf.

Sloane:1889:TIP

[12034] T. O’Conor Sloane. Terraced irrigation process of sewage disposal. Scientific American, 60(8):119, February 23, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v60/
n8/pdf/scientificamerican02231889-119.pdf.

Anonymous:1889:J

[12035] Anonymous. The jerboa. Scientific American, 60(8):119–120, February 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n8/pdf/scientificamerican02231889-119a.pdf.

Hopkins:1889:SEPe

[12036] Geo. M. Hopkins. Simple experiments in physics. Scientific American, 60(8):121, February 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n8/pdf/scientificamerican02231889-
121.pdf.

Anonymous:1889:Ibh

[12037] Anonymous. To inventors. Scientific American, 60(8):123, February 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n8/pdf/scientificamerican02231889-123a.pdf.

Anonymous:1889:RPIz

[12038] Anonymous. Recently patented inventions. Scientific American, 60(9):138, March 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n9/pdf/scientificamerican03021889-
138.pdf.

Anonymous:1889:BPbb

[12039] Anonymous. Business and personal. Scientific American, 60(9):138, March 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n9/pdf/scientificamerican03021889-138a.pdf.

Anonymous:1889:BNBn

[12040] Anonymous. Books: New books and publications. Scientific American, 60(9):138, March 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n9/pdf/scientificamerican03021889-
138b.pdf.

Anonymous:1889:NQy

[12041] Anonymous. Notes and queries. Scientific American, 60(9):138, March 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n9/pdf/scientificamerican03021889-138c.pdf.

Anonymous:1889:IIz

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n9/pdf/scientificamerican03021889-138d.pdf.

Anonymous:1889:PNPz

[12043] Anonymous. Patent notes: Patents. Scientific American, 60(9):142, March 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n9/pdf/scientificamerican03021889-142.pdf.

Anonymous:1889:IRAa

[12044] Anonymous. The improvement of the railroad approaches of New York. Scientific American, 60(9):127, March 2, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v60/
n9/pdf/scientificamerican03021889-127.pdf.

Anonymous:1889:DEL

[12045] Anonymous. The Daft electric locomotive. Scientific American, 60(9):128, March 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n9/pdf/scientificamerican03021889-
128.pdf.

Anonymous:1889:GD

[12046] Anonymous. The germ of diphtheria. Scientific American, 60(9):128, March 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n9/pdf/scientificamerican03021889-128a.pdf.

Anonymous:1889:SI

[12047] Anonymous. Strange incendiarism. Scientific American, 60(9):128–129, March 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n9/pdf/scientificamerican03021889-128b.pdf.

Anonymous:1889:IUM

[12048] Anonymous. An improved universal milling machine. Scientific American, 60(9):130, March 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n9/pdf/scientificamerican03021889-
130.pdf.

Anonymous:1889:ITE

[12049] Anonymous. An instrument to test engine crank shafts. Scientific American, 60(9):130, March 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n9/pdf/scientificamerican03021889-
130a.pdf.

Anonymous:1889:LCB

[12050] Anonymous. A luggage carrier for bicycles. Scientific American, 60(9):131, March 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n9/pdf/scientificamerican03021889-
131.pdf.

Anonymous:1889:ICSc

[12051] Anonymous. An improved car-starter. Scientific American, 60(9):131, March 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n9/pdf/scientificamerican03021889-131a.pdf.

Anonymous:1889:IDMb

[12052] Anonymous. An improved dyeing machine. Scientific American, 60(9):131, March 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n9/pdf/scientificamerican03021889-
131b.pdf.

Anonymous:1889:IVa

[12053] Anonymous. An improved velocipede. Scientific American, 60(9):131, March 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n9/pdf/scientificamerican03021889-131c.pdf.

Anonymous:1889:ISBb

[12054] Anonymous. An improved signal bell. Scientific American, 60 (9):131–133, March 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n9/pdf/scientificamerican03021889-
131d.pdf.

Sloane:1889:MET

[12055] T. O’Conor Sloane. Melde’s experiment with the Trevelyan Rocker. Scientific American, 60(9):135, March 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v60/n9/pdf/scientificamerican03021889-
135.pdf.

Anonymous:1889:SAI

[12056] Anonymous. The Scientific American iodide of mercury barometer. Scientific American, 60(9):135–136, March 2, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v60/
n9/pdf/scientificamerican03021889-135a.pdf.

Anonymous:1889:TTL

[12057] Anonymous. Ten-ton locomotive crane. Scientific American, 60(9):136, March 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v60/n9/pdf/scientificamerican03021889-136.pdf.

Anonymous:1889:IAH

[12058] Anonymous. Improved apparatus for heating cars. Scientific American, 60(9):136–137, March 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n9/pdf/scientificamerican03021889-
136a.pdf.

Hopkins:1889:SEPf

[12059] Geo. M. Hopkins. Simple experiments in physics. Scientific American, 60(9):137, March 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v60/n9/pdf/scientificamerican03021889-
137.pdf.

Anonymous:1889:Cq

[12060] Anonymous. Correspondence. Scientific American, 61(1):7, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n1/pdf/scientificamerican07061889-7.pdf.

Anonymous:1889:RPIba

[12061] Anonymous. Recently patented inventions. Scientific American, 61(1): 10, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n1/pdf/scientificamerican07061889-10.pdf.

Anonymous:1889:BPbc

[12062] Anonymous. Business and personal. Scientific American, 61(1):10, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n1/pdf/scientificamerican07061889-10a.pdf.

Anonymous:1889:BNBo

[12063] Anonymous. Books: New books and publications. Scientific American, 61(1):10, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n1/pdf/scientificamerican07061889-
10b.pdf.

Anonymous:1889:NQz

[12064] Anonymous. Notes and queries. Scientific American, 61(1):10–11, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n1/pdf/scientificamerican07061889-10c.pdf.

Anonymous:1889:IIba

[12065] Anonymous. Index of inventions. Scientific American, 61(1):11–12, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n1/pdf/scientificamerican07061889-11a.pdf.

Anonymous:1889:PNPba

[12066] Anonymous. Patent notes: Patents. Scientific American, 61(1):14, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n1/pdf/scientificamerican07061889-14.pdf.

Anonymous:1889:BVF

[12067] Anonymous. Bridal veil falls, snake river. Scientific American, 61(1):1– 2, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n1/pdf/scientificamerican07061889-1.pdf.

Anonymous:1889:WSNb

[12068] Anonymous. To water supply for New York City. Scientific American, 61(1):2–3, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n1/pdf/scientificamerican07061889-
2.pdf.

Sloane:1889:EIW

[12069] T. O’conor Sloane. Experiments on the inexpansibility of water and contraction of ice. Scientific American, 61(1):4, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n1/pdf/scientificamerican07061889-4.pdf.

Anonymous:1889:ICT

[12070] Anonymous. An improved clinical thermometer. Scientific American, 61(1):4, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n1/pdf/scientificamerican07061889-
4a.pdf.

Anonymous:1889:IBAa

[12071] Anonymous. An improved burglar alarm. Scientific American, 61(1):4, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n1/pdf/scientificamerican07061889-4b.pdf.

Anonymous:1889:IPL

[12072] Anonymous. An improved piano lamp bracket. Scientific American, 61(1):4, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n1/pdf/scientificamerican07061889-
4c.pdf.

Anonymous:1889:ITCc

[12073] Anonymous. An improved thill coupling. Scientific American, 61(1):5, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n1/pdf/scientificamerican07061889-5.pdf.

Anonymous:1889:IFCb

[12074] Anonymous. Improved fender for car windows. Scientific American, 61(1):5, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n1/pdf/scientificamerican07061889-
5a.pdf.

Anonymous:1889:SCO

[12075] Anonymous. A spring cover oiler. Scientific American, 61(1):5, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n1/pdf/scientificamerican07061889-5b.pdf.

Anonymous:1889:ICSd

[12076] Anonymous. An improved child’s swing. Scientific American, 61(1):5, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n1/pdf/scientificamerican07061889-5c.pdf.

Anonymous:1889:SAT

[12077] Anonymous. Sieve attachment for thrashing machines. Scientific American, 61(1):5, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n1/pdf/scientificamerican07061889-
5d.pdf.

Anonymous:1889:ILBc

[12078] Anonymous. Improved leather belting. Scientific American, 61(1):6, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n1/pdf/scientificamerican07061889-6.pdf.

Anonymous:1889:NRT

[12079] Anonymous. The new rubber tip Pearl mucilage bottle. Scientific American, 61(1):6, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n1/pdf/scientificamerican07061889-
6a.pdf.

Anonymous:1889:AHS

[12080] Anonymous. An apparatus for heating and steaming grain. Scientific American, 61(1):6, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n1/pdf/scientificamerican07061889-
6b.pdf.

Anonymous:1889:PNi

[12081] Anonymous. Photographic notes. Scientific American, 61(1):7, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n1/pdf/scientificamerican07061889-7a.pdf.

Anonymous:1889:WAN

[12082] Anonymous. The Washington Arch of New York. Scientific American, 61(1):8, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n1/pdf/scientificamerican07061889-
8.pdf.

Anonymous:1889:VBZ

[12083] Anonymous. The vari in the Berlin Zoological Garden. Scientific American, 61(1):9, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n1/pdf/scientificamerican07061889-
9.pdf.

Anonymous:1889:Ibi

[12084] Anonymous. To inventors. Scientific American, 61(1):11, July 6, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n1/pdf/scientificamerican07061889-11.pdf.

Anonymous:1889:RPIbb

[12085] Anonymous. Recently patented inventions. Scientific American, 61(10):154, September 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n10/pdf/scientificamerican09071889-
154.pdf.

Anonymous:1889:BPbd

[12086] Anonymous. Business and personal. Scientific American, 61 (10):154, September 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n10/pdf/scientificamerican09071889-
154a.pdf.

Anonymous:1889:NQba

[12087] Anonymous. Notes and queries. Scientific American, 61(10): 154–155, September 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n10/pdf/scientificamerican09071889-
154b.pdf.

Anonymous:1889:IIbb

[12088] Anonymous. Index of inventions. Scientific American, 61(10): 155–156, September 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n10/pdf/scientificamerican09071889-
155a.pdf.

Anonymous:1889:PNPbb

[12089] Anonymous. Patent notes: Patents. Scientific American, 61 (10):158, September 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n10/pdf/scientificamerican09071889-
158.pdf.

Anonymous:1889:HRT

[12090] Anonymous. The Hudson River Tunnel. Scientific American, 61 (10):143, September 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n10/pdf/scientificamerican09071889-
143.pdf.

Anonymous:1889:GCU

[12091] Anonymous. Government contracts undesirable. Scientific American, 61(10):144–145, September 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n10/pdf/scientificamerican09071889-
144.pdf.

Anonymous:1889:EL

[12092] Anonymous. Elias loomis. Scientific American, 61(10):146, September 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n10/pdf/scientificamerican09071889-146.pdf.

Anonymous:1889:IGA

[12093] Anonymous. An improved game apparatus. Scientific American, 61(10):147, September 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n10/pdf/scientificamerican09071889-
147.pdf.

Anonymous:1889:IGc

[12094] Anonymous. An improved governor. Scientific American, 61 (10):147, September 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n10/pdf/scientificamerican09071889-
147a.pdf.

Anonymous:1889:IWV

[12095] Anonymous. An improved window ventilator. Scientific American, 61(10):147, September 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n10/pdf/scientificamerican09071889-
147b.pdf.

Anonymous:1889:IAM

[12096] Anonymous. An improved astronomical mirror. Scientific American, 61(10):147, September 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n10/pdf/scientificamerican09071889-
147c.pdf.

Anonymous:1889:AAS

[12097] Anonymous. An automatic ash sifter. Scientific American, 61 (10):147–149, September 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n10/pdf/scientificamerican09071889-
147d.pdf.

Anonymous:1889:BBE

[12098] Anonymous. The Boynton bicycle engine. Scientific American, 61(10):150, September 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n10/pdf/scientificamerican09071889-
150.pdf.

Anonymous:1889:IEA

[12099] Anonymous. An improved elevator alarm. Scientific American, 61(10):150, September 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n10/pdf/scientificamerican09071889-
150a.pdf.

Anonymous:1889:PEc

[12100] Anonymous. The Paris Exhibition. Scientific American, 61(10): 153, September 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n10/pdf/scientificamerican09071889-
153.pdf.

Anonymous:1889:Ibj

[12101] Anonymous. To inventors. Scientific American, 61(10):155, September 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n10/pdf/scientificamerican09071889-155.pdf.

Anonymous:1889:Cr

[12102] Anonymous. Correspondence. Scientific American, 61(11):165, September 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n11/pdf/scientificamerican09141889-165a.pdf.

Anonymous:1889:RPIbc

[12103] Anonymous. Recently patented inventions. Scientific American, 61 (11):170, September 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n11/pdf/scientificamerican09141889-
170.pdf.

Anonymous:1889:BPbe

[12104] Anonymous. Business and personal. Scientific American, 61 (11):170, September 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n11/pdf/scientificamerican09141889-
170a.pdf.

Anonymous:1889:NQbb

[12105] Anonymous. Notes and queries. Scientific American, 61(11): 170–171, September 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n11/pdf/scientificamerican09141889-
170b.pdf.

Anonymous:1889:IIbc 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n11/pdf/scientificamerican09141889-
171a.pdf.

Anonymous:1889:PNPbc

[12107] Anonymous. Patent notes: Patents. Scientific American, 61 (11):174, September 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n11/pdf/scientificamerican09141889-
174.pdf.

Anonymous:1889:NJB

[12108] Anonymous. The new jet boat evolution. Scientific American, 61 (11):160, September 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n11/pdf/scientificamerican09141889-
160.pdf.

Anonymous:1889:SMT

[12109] Anonymous. The scientific meeting at Toronto. Scientific American, 61(11):160–161, September 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n11/pdf/scientificamerican09141889-
160a.pdf.

Anonymous:1889:SDP

[12110] Anonymous. The sewage disposal problem. Scientific American, 61(11):161, September 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n11/pdf/scientificamerican09141889-
161.pdf.

Anonymous:1889:IRCd

[12111] Anonymous. An improved road cart. Scientific American, 61 (11):162, September 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n11/pdf/scientificamerican09141889-
162.pdf.

Anonymous:1889:IDCb

[12112] Anonymous. An improved dumping car. Scientific American, 61 (11):162, September 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n11/pdf/scientificamerican09141889-
162a.pdf.

Anonymous:1889:IDG

[12113] Anonymous. An improved drawbridge gate. Scientific American, 61(11):162, September 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n11/pdf/scientificamerican09141889-
162b.pdf.

Anonymous:1889:IBCf

[12114] Anonymous. An improved blow-off cock and fire extinguisher. Scientific American, 61(11):162, September 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v61/n11/pdf/scientificamerican09141889-
162c.pdf.

Anonymous:1889:IRTb

[12115] Anonymous. An improved roofing tank apparatus. Scientific American, 61(11):162–163, September 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n11/pdf/scientificamerican09141889-
162d.pdf.

Anonymous:1889:RRC

[12116] Anonymous. Rheostat and resistance column. Scientific American, 61(11):163, September 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n11/pdf/scientificamerican09141889-
163.pdf.

Anonymous:1889:IOCc

[12117] Anonymous. An improved oil can. Scientific American, 61(11): 163, September 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n11/pdf/scientificamerican09141889-
163a.pdf.

Anonymous:1889:ILSb

[12118] Anonymous. Improved lock and shackle for railway switches. Scientific American, 61(11):163–164, September 14, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v61/
n11/pdf/scientificamerican09141889-163b.pdf.

Anonymous:1889:NAL

[12119] Anonymous. New arc lamp. Scientific American, 61(11):164, September 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n11/pdf/scientificamerican09141889-164.pdf.

Anonymous:1889:LP

[12120] Anonymous. Launch of the Philadelphia. Scientific American, 61 (11):164, September 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n11/pdf/scientificamerican09141889-
164a.pdf.

Anonymous:1889:HRA

[12121] Anonymous. The Harlem River Aqueduct Tunnel and its approaches. Scientific American, 61(11):165, September 14, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v61/
n11/pdf/scientificamerican09141889-165.pdf.

Anonymous:1889:PNj

[12122] Anonymous. Photographic notes. Scientific American, 61(11): 165, September 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n11/pdf/scientificamerican09141889-
165b.pdf.

Anonymous:1889:MEE

[12123] Anonymous. With Mr. Edison on the Eiffel Tower. Scientific American, 61(11):166, September 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n11/pdf/scientificamerican09141889-
166.pdf.

Anonymous:1889:ACD

[12124] Anonymous. An attractive cheap dwelling house. Scientific American, 61(11):167–168, September 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n11/pdf/scientificamerican09141889-
167.pdf.

Anonymous:1889:PEd

[12125] Anonymous. The Paris Exhibition. Scientific American, 61(11): 169, September 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n11/pdf/scientificamerican09141889-
169.pdf.

Anonymous:1889:Ibk

[12126] Anonymous. To inventors. Scientific American, 61(11):171, September 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n11/pdf/scientificamerican09141889-171.pdf.

Anonymous:1889:Cs

[12127] Anonymous. Correspondence. Scientific American, 61(12):181, September 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n12/pdf/scientificamerican09211889-181.pdf.

Anonymous:1889:BPbf

[12128] Anonymous. Business and personal. Scientific American, 61 (12):187, September 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n12/pdf/scientificamerican09211889-
187.pdf.

Anonymous:1889:NQbc

[12129] Anonymous. Notes and queries. Scientific American, 61(12): 187, September 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n12/pdf/scientificamerican09211889-
187a.pdf.

Anonymous:1889:IIbd

[12130] Anonymous. Index of inventions. Scientific American, 61(12): 187–188, September 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n12/pdf/scientificamerican09211889-
187c.pdf.

Anonymous:1889:PNPbd 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n12/pdf/scientificamerican09211889-
190.pdf.

Anonymous:1889:RMA

[12132] Anonymous. The recent meeting of the American Association for the Advancement of Science. Scientific American, 61(12):176–177, September 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n12/pdf/scientificamerican09211889-176.pdf.

Anonymous:1889:ENU

[12133] Anonymous. Eikonogen — a new universal developer for photographic dry plates and bromide paper. Scientific American, 61(12):177, September 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n12/pdf/scientificamerican09211889-177.pdf.

Anonymous:1889:IFPa

[12134] Anonymous. An improved fruit picker. Scientific American, 61 (12):178, September 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n12/pdf/scientificamerican09211889-
178.pdf.

Anonymous:1889:IVH

[12135] Anonymous. An improved valve head for blowing engines. Scientific American, 61(12):178, September 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v61/n12/pdf/scientificamerican09211889-
178a.pdf.

Anonymous:1889:ISBc

[12136] Anonymous. An improved steam boiler. Scientific American, 61 (12):178, September 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n12/pdf/scientificamerican09211889-
178b.pdf.

Anonymous:1889:CCR

[12137] Anonymous. A combined carpenters rule, calipers, etc. Scientific American, 61(12):178, September 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n12/pdf/scientificamerican09211889-
178c.pdf.

Anonymous:1889:ILJa

[12138] Anonymous. An improved lock joint for beams. Scientific American, 61(12):178, September 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n12/pdf/scientificamerican09211889-
178d.pdf.

Anonymous:1889:IPHb

[12139] Anonymous. An improved pen holder. Scientific American, 61 (12):179, September 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n12/pdf/scientificamerican09211889-
179.pdf.

Anonymous:1889:IBBc

[12140] Anonymous. An improved blacking box. Scientific American, 61 (12):179, September 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n12/pdf/scientificamerican09211889-
179a.pdf.

Anonymous:1889:ITW

[12141] Anonymous. An improved two-wheeled vehicle. Scientific American, 61(12):179, September 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n12/pdf/scientificamerican09211889-
179b.pdf.

Anonymous:1889:IMLb

[12142] Anonymous. An improved music leaf turner. Scientific American, 61(12):179, September 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n12/pdf/scientificamerican09211889-
179c.pdf.

Anonymous:1889:IAL

[12143] Anonymous. An improved artificial leg. Scientific American, 61 (12):179, September 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n12/pdf/scientificamerican09211889-
179d.pdf.

Anonymous:1889:BNY

[12144] Anonymous. The Brooklyn Navy Yard. Scientific American, 61 (12):180, September 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n12/pdf/scientificamerican09211889-
180.pdf.

Anonymous:1889:TG

[12145] Anonymous. Tangent galvanometer. Scientific American, 61(12): 181–182, September 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n12/pdf/scientificamerican09211889-
181a.pdf.

Anonymous:1889:PEe

[12146] Anonymous. The Paris Exhibition. Scientific American, 61(12): 182, September 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n12/pdf/scientificamerican09211889-
182.pdf.

Anonymous:1889:FGE

[12147] Anonymous. The First Gallery of the Eiffel Tower. Scientific American, 61(12):183–185, September 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n12/pdf/scientificamerican09211889-
183.pdf.

Anonymous:1889:RPIbd

[12148] Anonymous. Recently patented inventions. Scientific American, 61 (12):186, September 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n12/pdf/scientificamerican09211889-
186.pdf.

Anonymous:1889:Ibl

[12149] Anonymous. To inventors. Scientific American, 61(12):187, September 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n12/pdf/scientificamerican09211889-187b.pdf.

Anonymous:1889:Ct

[12150] Anonymous. Correspondence. Scientific American, 61(13):197, September 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n13/pdf/scientificamerican09281889-197a.pdf.

Anonymous:1889:BPbg

[12151] Anonymous. Business and personal. Scientific American, 61 (13):202, September 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n13/pdf/scientificamerican09281889-
202a.pdf.

Anonymous:1889:NQbd

[12152] Anonymous. Notes and queries. Scientific American, 61(13): 202–203, September 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n13/pdf/scientificamerican09281889-
202b.pdf.

Anonymous:1889:IIbe

[12153] Anonymous. Index of inventions. Scientific American, 61(13): 203–204, September 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n13/pdf/scientificamerican09281889-
203a.pdf.

Anonymous:1889:PNPbe

[12154] Anonymous. Patent notes: Patents. Scientific American, 61 (13):206, September 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n13/pdf/scientificamerican09281889-
206.pdf.

Anonymous:1889:UWP

[12155] Anonymous. Utilizing wave power at the sea beach. Scientific American, 61(13):191, September 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n13/pdf/scientificamerican09281889-
191.pdf.

Anonymous:1889:STN

[12156] Anonymous. Successful trials of the new cruiser Baltimore. Scientific American, 61(13):192, September 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v61/n13/pdf/scientificamerican09281889-
192.pdf.

Anonymous:1889:PPO

[12157] Anonymous. Position of the planets for October. Scientific American, 61(13):192–193, September 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n13/pdf/scientificamerican09281889-
192a.pdf.

Anonymous:1889:IGSc

[12158] Anonymous. An improved governor for steam engines. Scientific American, 61(13):194, September 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n13/pdf/scientificamerican09281889-
194.pdf.

Anonymous:1889:ISAb

[12159] Anonymous. An improved storm apron. Scientific American, 61 (13):194, September 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n13/pdf/scientificamerican09281889-
194a.pdf.

Anonymous:1889:ICPb

[12160] Anonymous. An improved cocoanut parer. Scientific American, 61 (13):194, September 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n13/pdf/scientificamerican09281889-
194b.pdf.

Anonymous:1889:ISPc

[12161] Anonymous. An improved sill plate for railway cars. Scientific American, 61(13):194, September 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n13/pdf/scientificamerican09281889-
194c.pdf.

Anonymous:1889:IBHa

[12162] Anonymous. An improved backband hook. Scientific American, 61(13):195, September 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n13/pdf/scientificamerican09281889-
195.pdf.

Anonymous:1889:RCC

[12163] Anonymous. A railway car for carrying wide stones. Scientific American, 61(13):195, September 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n13/pdf/scientificamerican09281889-
195a.pdf.

Anonymous:1889:JGF

[12164] Anonymous. The John Good force draught fuel gas generator and smoke consumer for steam boilers. Scientific American, 61(13):195, September 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n13/pdf/scientificamerican09281889-195b.pdf.

Anonymous:1889:IKH

[12165] Anonymous. An improved key-hole guard. Scientific American, 61 (13):195, September 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n13/pdf/scientificamerican09281889-
195c.pdf.

Anonymous:1889:IUF

[12166] Anonymous. An improved undulating furnace hearth. Scientific American, 61(13):195–197, September 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v61/n13/pdf/scientificamerican09281889-
195d.pdf.

Anonymous:1889:VT

[12167] Anonymous. The Victor typewriter. Scientific American, 61 (13):197, September 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n13/pdf/scientificamerican09281889-
197.pdf.

Anonymous:1889:W

[12168] Anonymous. The woodcock. Scientific American, 61(13):198– 199, September 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n13/pdf/scientificamerican09281889-
198.pdf.

Anonymous:1889:TFN

[12169] Anonymous. Types of the French Navy — the Condor. Scientific American, 61(13):199–201, September 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v61/n13/pdf/scientificamerican09281889-
199.pdf.

Anonymous:1889:RPIbe

[12170] Anonymous. Recently patented inventions. Scientific American, 61 (13):202, September 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n13/pdf/scientificamerican09281889-
202.pdf.

Anonymous:1889:Ibm

[12171] Anonymous. To inventors. Scientific American, 61(13):203, September 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n13/pdf/scientificamerican09281889-203.pdf.

Anonymous:1889:RPIbf

[12172] Anonymous. Recently patented inventions. Scientific American, 61(14):218, October 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n14/pdf/scientificamerican10051889-
218.pdf.

Anonymous:1889:BPbh

[12173] Anonymous. Business and personal. Scientific American, 61(14):218, October 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n14/pdf/scientificamerican10051889-218a.pdf.

Anonymous:1889:NQbe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n14/pdf/scientificamerican10051889-218b.pdf.

Anonymous:1889:IIbf

[12175] Anonymous. Index of inventions. Scientific American, 61(14):219–220, October 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n14/pdf/scientificamerican10051889-219a.pdf.

Anonymous:1889:PNPbf

[12176] Anonymous. Patent notes: Patents. Scientific American, 61(14):222, October 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n14/pdf/scientificamerican10051889-222.pdf.

Anonymous:1889:APH

[12177] Anonymous. The American Public Health Association. Scientific American, 61(14):208, October 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n14/pdf/scientificamerican10051889-
208.pdf.

Anonymous:1889:PPE

[12178] Anonymous. The professional and practical engineer. Scientific American, 61(14):208, October 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n14/pdf/scientificamerican10051889-
208a.pdf.

Anonymous:1889:NCM

[12179] Anonymous. The nickel and copper mines at Sudbury, Ontario. Scientific American, 61(14):208–209, October 5, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v61/
n14/pdf/scientificamerican10051889-208b.pdf.

Anonymous:1889:DWP

[12180] Anonymous. Damage by water in Peru. Scientific American, 61(14):209, October 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n14/pdf/scientificamerican10051889-209.pdf.

Anonymous:1889:HEP

[12181] Anonymous. A hand exerciser for pianists. Scientific American, 61(14):210, October 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n14/pdf/scientificamerican10051889-
210.pdf.

Anonymous:1889:IASc

[12182] Anonymous. An improved ash sifter. Scientific American, 61(14):210, October 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n14/pdf/scientificamerican10051889-210a.pdf.

Anonymous:1889:ESF

[12183] Anonymous. The Evans system of frictional gearing. Scientific American, 61(14):210, October 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n14/pdf/scientificamerican10051889-
210b.pdf.

Anonymous:1889:ILd

[12184] Anonymous. An improved lamp. Scientific American, 61(14):210–211, October 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n14/pdf/scientificamerican10051889-210c.pdf.

Anonymous:1889:IHL

[12185] Anonymous. An improved high and low water alarm. Scientific American, 61(14):211, October 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n14/pdf/scientificamerican10051889-
211.pdf.

Anonymous:1889:IPI

[12186] Anonymous. An improved price indicator for scales. Scientific American, 61(14):211, October 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n14/pdf/scientificamerican10051889-
211a.pdf.

Anonymous:1889:RLM

[12187] Anonymous. The Remington–Lee military magazine rifle. Scientific American, 61(14):211, October 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n14/pdf/scientificamerican10051889-
211b.pdf.

Anonymous:1889:IBCg

[12188] Anonymous. An improved band cutter and feeder. Scientific American, 61(14):211–212, October 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n14/pdf/scientificamerican10051889-
211c.pdf.

Anonymous:1889:BAC

[12189] Anonymous. Building the armored cruiser Maine. Scientific American, 61(14):212, October 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n14/pdf/scientificamerican10051889-
212.pdf.

Anonymous:1889:IWW

[12190] Anonymous. An improved water wheel. Scientific American, 61(14):213, October 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n14/pdf/scientificamerican10051889-213.pdf.

Anonymous:1889:ISWa

[12191] Anonymous. An improved shutter worker. Scientific American, 61(14):213, October 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n14/pdf/scientificamerican10051889-
213a.pdf.

Anonymous:1889:EPD

[12192] Anonymous. Electro-plating dynamo. Scientific American, 61(14):213, October 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n14/pdf/scientificamerican10051889-213b.pdf.

Anonymous:1889:IWPa

[12193] Anonymous. Improved web printing machine. Scientific American, 61(14):214, October 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n14/pdf/scientificamerican10051889-
214.pdf.

Anonymous:1889:MS

[12194] Anonymous. Mount Stephen. Scientific American, 61(14):215–216, October 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n14/pdf/scientificamerican10051889-215.pdf.

Anonymous:1889:PNk

[12195] Anonymous. Photographic notes. Scientific American, 61(14):216, October 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n14/pdf/scientificamerican10051889-216.pdf.

Pike:1889:SI

[12196] Nicolas Pike. The Seychelles Islands. Scientific American, 61(14):217, October 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n14/pdf/scientificamerican10051889-217.pdf.

Anonymous:1889:Ibn

[12197] Anonymous. To inventors. Scientific American, 61(14):219, October 5, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n14/pdf/scientificamerican10051889-219.pdf.

Anonymous:1889:Cu

[12198] Anonymous. Correspondence. Scientific American, 61(15):229–230, October 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n15/pdf/scientificamerican10121889-229.pdf.

Anonymous:1889:RPIbg

[12199] Anonymous. Recently patented inventions. Scientific American, 61(15):234, October 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n15/pdf/scientificamerican10121889-
234.pdf.

Anonymous:1889:BNBp

[12200] Anonymous. Books: New books and publications. Scientific American, 61(15):234, October 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n15/pdf/scientificamerican10121889-
234a.pdf.

Anonymous:1889:BPbi

[12201] Anonymous. Business and personal. Scientific American, 61 (15):234, October 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n15/pdf/scientificamerican10121889-
234b.pdf.

Anonymous:1889:NQbf

[12202] Anonymous. Notes and queries. Scientific American, 61(15): 234–235, October 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n15/pdf/scientificamerican10121889-
234c.pdf.

Anonymous:1889:IIbg

[12203] Anonymous. Index of inventions. Scientific American, 61(15): 235–236, October 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n15/pdf/scientificamerican10121889-
235a.pdf.

Anonymous:1889:PNPbg

[12204] Anonymous. Patent notes: Patents. Scientific American, 61 (15):238, October 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n15/pdf/scientificamerican10121889-
238.pdf.

Anonymous:1889:HST

[12205] Anonymous. The Halpin–Savage torpedo. Scientific American, 61 (15):223–224, October 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n15/pdf/scientificamerican10121889-
223.pdf.

Anonymous:1889:FHA

[12206] Anonymous. 1,492 feet for the height of the American Tower. Scientific American, 61(15):224, October 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v61/n15/pdf/scientificamerican10121889-
224.pdf.

Anonymous:1889:MNT

[12207] Anonymous. Meeting of the National Telephone Association. Scientific American, 61(15):224–225, October 12, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v61/
n15/pdf/scientificamerican10121889-224a.pdf.

Anonymous:1889:ICAb

[12208] Anonymous. The International Congress of American States. Scientific American, 61(15):225, October 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v61/n15/pdf/scientificamerican10121889-
225.pdf.

Anonymous:1889:IRFb

[12209] Anonymous. An improved replacing frog. Scientific American, 61(15):226, October 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n15/pdf/scientificamerican10121889-
226.pdf.

Anonymous:1889:IBHb

[12210] Anonymous. An improved blast heater. Scientific American, 61 (15):226, October 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n15/pdf/scientificamerican10121889-
226a.pdf.

Anonymous:1889:ISTa

[12211] Anonymous. An improved sketching tripod, or delineator. Scientific American, 61(15):226, October 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v61/n15/pdf/scientificamerican10121889-
226b.pdf.

Anonymous:1889:AS

[12212] Anonymous. An attachment for scissors. Scientific American, 61(15):226, October 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n15/pdf/scientificamerican10121889-
226c.pdf.

Anonymous:1889:ISWb

[12213] Anonymous. An improved swinging window sash. Scientific American, 61(15):226, October 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n15/pdf/scientificamerican10121889-
226d.pdf.

Anonymous:1889:JL

[12214] Anonymous. Jacob’s ladder. Scientific American, 61(15):227, October 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n15/pdf/scientificamerican10121889-227.pdf.

Anonymous:1889:OGE

[12215] Anonymous. An “Otto” gas engine of one-third horse power. Scientific American, 61(15):227–228, October 12, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v61/
n15/pdf/scientificamerican10121889-227a.pdf.

Anonymous:1889:TAE

[12216] Anonymous. Thomas Alva Edison. Scientific American, 61(15): 230–232, October 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n15/pdf/scientificamerican10121889-
230.pdf.

Anonymous:1889:ABC

[12217] Anonymous. Arrangement of battery cells. Scientific American, 61(15):233, October 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n15/pdf/scientificamerican10121889-
233.pdf.

Anonymous:1889:Ibo

[12218] Anonymous. To inventors. Scientific American, 61(15):235, October 12, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n15/pdf/scientificamerican10121889-235.pdf.

Anonymous:1889:Cx

[12219] Anonymous. Correspondence. Scientific American, 61(16):245–246, October 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n16/pdf/scientificamerican10191889-245.pdf.

Anonymous:1889:RPIbh

[12220] Anonymous. Recently patented inventions. Scientific American, 61(16):250, October 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n16/pdf/scientificamerican10191889-
250.pdf.

Anonymous:1889:BPbj

[12221] Anonymous. Business and personal. Scientific American, 61 (16):250, October 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n16/pdf/scientificamerican10191889-
250a.pdf.

Anonymous:1889:BNBq

[12222] Anonymous. Books: New books and publications. Scientific American, 61(16):250, October 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n16/pdf/scientificamerican10191889-
250b.pdf.

Anonymous:1889:NQbg

[12223] Anonymous. Notes and queries. Scientific American, 61(16): 250–251, October 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n16/pdf/scientificamerican10191889-
250c.pdf.

Anonymous:1889:IIbh

[12224] Anonymous. Index of inventions. Scientific American, 61(16): 251–252, October 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n16/pdf/scientificamerican10191889-
251a.pdf.

Anonymous:1889:PNPbh

[12225] Anonymous. Patent notes: Patents. Scientific American, 61 (16):254, October 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n16/pdf/scientificamerican10191889-
254.pdf.

Anonymous:1889:JPT

[12226] Anonymous. The Johnson pneumatic tube. Scientific American, 61(16):239, October 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n16/pdf/scientificamerican10191889-
239.pdf.

Anonymous:1889:ONL

[12227] Anonymous. Opening of the new London Bridge. Scientific American, 61(16):240, October 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n16/pdf/scientificamerican10191889-
240.pdf.

Anonymous:1889:EEC

[12228] Anonymous. The electric execution of criminals. Scientific American, 61(16):240, October 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n16/pdf/scientificamerican10191889-
240a.pdf.

Anonymous:1889:IEL

[12229] Anonymous. The incandescent electric lamp patent. Scientific American, 61(16):240, October 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n16/pdf/scientificamerican10191889-
240b.pdf.

Anonymous:1889:RIC

[12230] Anonymous. The raisin industry of California. Scientific American, 61(16):241, October 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n16/pdf/scientificamerican10191889-
241.pdf.

Anonymous:1889:IPe

[12231] Anonymous. An improved planter. Scientific American, 61(16):242, October 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n16/pdf/scientificamerican10191889-242.pdf.

Anonymous:1889:IFCc

[12232] Anonymous. An improved freight car door. Scientific American, 61(16):242, October 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n16/pdf/scientificamerican10191889-
242a.pdf.

Anonymous:1889:ISWc

[12233] Anonymous. An improved spindle wharve. Scientific American, 61(16):242, October 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n16/pdf/scientificamerican10191889-
242b.pdf.

Anonymous:1889:IFPb

[12234] Anonymous. An improved feathering paddle wheel. Scientific American, 61(16):242, October 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n16/pdf/scientificamerican10191889-
242c.pdf.

Anonymous:1889:IDB

[12235] Anonymous. An improved door bolt. Scientific American, 61 (16):242–243, October 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n16/pdf/scientificamerican10191889-
242d.pdf.

Anonymous:1889:WBD

[12236] Anonymous. William B. Dunning. Scientific American, 61(16): 243–245, October 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n16/pdf/scientificamerican10191889-
243.pdf.

Anonymous:1889:LFD

[12237] Anonymous. Luminous fountains for dining tables. Scientific American, 61(16):246, October 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n16/pdf/scientificamerican10191889-
246.pdf.

Hovey:1889:PDM

[12238] H. C. Hovey. The pits and domes of Mammoth Cave. Scientific American, 61(16):247–248, October 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n16/pdf/scientificamerican10191889-
247.pdf.

Anonymous:1889:MR

[12239] Anonymous. Measurement of resistance. Scientific American, 61 (16):249, October 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n16/pdf/scientificamerican10191889-
249.pdf.

Anonymous:1889:Ibq

[12240] Anonymous. To inventors. Scientific American, 61(16):251, October 19, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n16/pdf/scientificamerican10191889-251.pdf.

Anonymous:1889:Cy

[12241] Anonymous. Correspondence. Scientific American, 61(17):261–262, October 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n17/pdf/scientificamerican10261889-261.pdf.

Anonymous:1889:RPIbi

[12242] Anonymous. Recently patented inventions. Scientific American, 61(17):266, October 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n17/pdf/scientificamerican10261889-
266.pdf.

Anonymous:1889:BPbk

[12243] Anonymous. Business and personal. Scientific American, 61 (17):266, October 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n17/pdf/scientificamerican10261889-
266a.pdf.

Anonymous:1889:BNBr 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n17/pdf/scientificamerican10261889-
266b.pdf.

Anonymous:1889:NQbh

[12245] Anonymous. Notes and queries. Scientific American, 61(17):267, October 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n17/pdf/scientificamerican10261889-267.pdf.

Anonymous:1889:IIbi

[12246] Anonymous. Index of inventions. Scientific American, 61(17): 267–268, October 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n17/pdf/scientificamerican10261889-
267b.pdf.

Anonymous:1889:PNPbi

[12247] Anonymous. Patent notes: Patents. Scientific American, 61 (17):270, October 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n17/pdf/scientificamerican10261889-
270.pdf.

Anonymous:1889:GES

[12248] Anonymous. Glass enameled steel casks for brewers, sugar refiners, glucose makers, etc. Scientific American, 61(17):255, October 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n17/pdf/scientificamerican10261889-255.pdf.

Anonymous:1889:DAA

[12249] Anonymous. Departure of the American Astronomical Expedition. Scientific American, 61(17):256, October 26, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v61/
n17/pdf/scientificamerican10261889-256.pdf.

Anonymous:1889:PYR

[12250] Anonymous. The phonograph not yet ready. Scientific American, 61(17):256, October 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n17/pdf/scientificamerican10261889-
256a.pdf.

Anonymous:1889:NYN

[12251] Anonymous. New York Navy Yard. Scientific American, 61(17):256, October 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n17/pdf/scientificamerican10261889-256b.pdf.

Anonymous:1889:IMCb

[12252] Anonymous. The International Maritime Conference. Scientific American, 61(17):256, October 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n17/pdf/scientificamerican10261889-
256c.pdf.

Anonymous:1889:BEC

[12253] Anonymous. A bridge for the English Channel. Scientific American, 61(17):256–257, October 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n17/pdf/scientificamerican10261889-
256d.pdf.

Anonymous:1889:NTW

[12254] Anonymous. The new type of war ship. Scientific American, 61(17):257, October 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n17/pdf/scientificamerican10261889-
257.pdf.

Anonymous:1889:PPN

[12255] Anonymous. Position of the planets for November. Scientific American, 61(17):257, October 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n17/pdf/scientificamerican10261889-
257a.pdf.

Anonymous:1889:PCU

[12256] Anonymous. Pneumatic conveyers for use in mills. Scientific American, 61(17):258, October 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n17/pdf/scientificamerican10261889-
258.pdf.

Anonymous:1889:WAR

[12257] Anonymous. Wolfe’s automatic railroad switch. Scientific American, 61(17):258, October 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n17/pdf/scientificamerican10261889-
258a.pdf.

Anonymous:1889:PAC

[12258] Anonymous. Producing artificial currents. Scientific American, 61(17):258, October 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n17/pdf/scientificamerican10261889-
258b.pdf.

Anonymous:1889:DFJ

[12259] Anonymous. Dix’s floor jack. Scientific American, 61(17):258, October 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n17/pdf/scientificamerican10261889-258c.pdf.

Anonymous:1889:BWM

[12260] Anonymous. Besemer’s water motor. Scientific American, 61 (17):258, October 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n17/pdf/scientificamerican10261889-
258d.pdf.

Anonymous:1889:MDD

[12261] Anonymous. Multipolar drum dynamo. Scientific American, 61 (17):259–260, October 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n17/pdf/scientificamerican10261889-
259.pdf.

Anonymous:1889:PEf

[12262] Anonymous. The Paris Exhibition. Scientific American, 61(17): 262–264, October 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n17/pdf/scientificamerican10261889-
262.pdf.

Anonymous:1889:MVA

[12263] Anonymous. Mouth vacuum apparatus. Scientific American, 61 (17):264–265, October 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n17/pdf/scientificamerican10261889-
264.pdf.

Holder:1889:HSB

[12264] C. F. Holder. The hard-scaled bass of the Pacific. Scientific American, 61(17):265, October 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n17/pdf/scientificamerican10261889-
265.pdf.

Anonymous:1889:GHC

[12265] Anonymous. George Hammell Cook. Scientific American, 61 (17):265, October 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n17/pdf/scientificamerican10261889-
265a.pdf.

Anonymous:1889:Ibr

[12266] Anonymous. To inventors. Scientific American, 61(17):267, October 26, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n17/pdf/scientificamerican10261889-267a.pdf.

Anonymous:1889:Cz

[12267] Anonymous. Correspondence. Scientific American, 61(18):277, November 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n18/pdf/scientificamerican11021889-277.pdf.

Anonymous:1889:RPIbj

[12268] Anonymous. Recently patented inventions. Scientific American, 61(18):282, November 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n18/pdf/scientificamerican11021889-
282.pdf.

Anonymous:1889:BPbl

[12269] Anonymous. Business and personal. Scientific American, 61 (18):282, November 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n18/pdf/scientificamerican11021889-
282a.pdf.

Anonymous:1889:BNBs

[12270] Anonymous. Books: New books and publications. Scientific American, 61(18):282, November 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n18/pdf/scientificamerican11021889-
282b.pdf.

Anonymous:1889:NQbi

[12271] Anonymous. Notes and queries. Scientific American, 61(18): 282–283, November 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n18/pdf/scientificamerican11021889-
282c.pdf.

Anonymous:1889:IIbj

[12272] Anonymous. Index of inventions. Scientific American, 61(18): 283–284, November 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n18/pdf/scientificamerican11021889-
283a.pdf.

Anonymous:1889:PNPbj

[12273] Anonymous. Patent notes: Patents. Scientific American, 61 (18):286, November 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n18/pdf/scientificamerican11021889-
286.pdf.

Anonymous:1889:NYS

[12274] Anonymous. The New York Signal Service Station. Scientific American, 61(18):271–272, November 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n18/pdf/scientificamerican11021889-
271.pdf.

Anonymous:1889:IMCc

[12275] Anonymous. The International Marine Conference. Scientific American, 61(18):272–273, November 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n18/pdf/scientificamerican11021889-
272.pdf.

Anonymous:1889:MHN

[12276] Anonymous. Mr. Hopkins’ new book, “Experimental Science”. Scientific American, 61(18):273, November 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v61/n18/pdf/scientificamerican11021889-
273.pdf.

Anonymous:1889:ILe

[12277] Anonymous. Incandescent lighting. Scientific American, 61(18): 274, November 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n18/pdf/scientificamerican11021889-
274.pdf.

Anonymous:1889:ISWd

[12278] Anonymous. An improved screwdriver and wrench. Scientific American, 61(18):275, November 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n18/pdf/scientificamerican11021889-
275.pdf.

Anonymous:1889:ILDa

[12279] Anonymous. An improved lathe dog. Scientific American, 61 (18):275, November 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n18/pdf/scientificamerican11021889-
275a.pdf.

Anonymous:1889:IMCd

[12280] Anonymous. An improved meat-carrying apparatus. Scientific American, 61(18):275, November 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n18/pdf/scientificamerican11021889-
275b.pdf.

Anonymous:1889:IVR

[12281] Anonymous. Improved vehicle running gear. Scientific American, 61(18):275, November 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n18/pdf/scientificamerican11021889-
275c.pdf.

Anonymous:1889:IED

[12282] Anonymous. An improved electrical door alarm. Scientific American, 61(18):275–276, November 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n18/pdf/scientificamerican11021889-
275d.pdf.

Anonymous:1889:DCA

[12283] Anonymous. Dernell’s clevis arm for ice flows. Scientific American, 61(18):276, November 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n18/pdf/scientificamerican11021889-
276.pdf.

Anonymous:1889:ICCh

[12284] Anonymous. An improved cotton cultivator. Scientific American, 61(18):276, November 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n18/pdf/scientificamerican11021889-
276a.pdf.

Anonymous:1889:GFE

[12285] Anonymous. Germany’s Floating Exposition. Scientific American, 61(18):278, November 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n18/pdf/scientificamerican11021889-
278.pdf.

Anonymous:1889:PEg

[12286] Anonymous. The Paris Exhibition. Scientific American, 61(18): 279–280, November 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n18/pdf/scientificamerican11021889-
279.pdf.

Anonymous:1889:WCE

[12287] Anonymous. Willard condensing engine. Scientific American, 61 (18):280–281, November 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n18/pdf/scientificamerican11021889-
280.pdf.

Anonymous:1889:Ibt

[12288] Anonymous. To inventors. Scientific American, 61(18):283, November 2, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n18/pdf/scientificamerican11021889-283.pdf.

Anonymous:1889:RPIbk

[12289] Anonymous. Recently patented inventions. Scientific American, 61(19):297, November 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n19/pdf/scientificamerican11091889-
297.pdf.

Anonymous:1889:BPbm

[12290] Anonymous. Business and personal. Scientific American, 61 (19):297, November 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n19/pdf/scientificamerican11091889-
297a.pdf.

Anonymous:1889:BNBt

[12291] Anonymous. Books: New books and publications. Scientific American, 61(19):297–298, November 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n19/pdf/scientificamerican11091889-
297b.pdf.

Anonymous:1889:NQbj

[12292] Anonymous. Notes and queries. Scientific American, 61(19): 298–299, November 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n19/pdf/scientificamerican11091889-
298.pdf.

Anonymous:1889:IIbk

[12293] Anonymous. Index of inventions. Scientific American, 61(19): 299–300, November 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n19/pdf/scientificamerican11091889-
299a.pdf.

Anonymous:1889:PNPbk

[12294] Anonymous. Patent notes: Patents. Scientific American, 61 (19):302, November 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n19/pdf/scientificamerican11091889-
302.pdf.

Anonymous:1889:MSCb

[12295] Anonymous. The Manchester Ship Canal. Scientific American, 61 (19):287–288, November 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n19/pdf/scientificamerican11091889-
287.pdf.

Anonymous:1889:AIF

[12296] Anonymous. The American Institute Fair. Scientific American, 61 (19):288–289, November 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n19/pdf/scientificamerican11091889-
288.pdf.

Anonymous:1889:EBF

[12297] Anonymous. An extinct bird (Fregilupus Varius Boddaert). Scientific American, 61(19):290, November 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v61/n19/pdf/scientificamerican11091889-
290.pdf.

Anonymous:1889:IDSc

[12298] Anonymous. An improved drag saw. Scientific American, 61 (19):290, November 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n19/pdf/scientificamerican11091889-
290a.pdf.

Anonymous:1889:ISMb

[12299] Anonymous. An improved sawing machine. Scientific American, 61(19):290, November 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n19/pdf/scientificamerican11091889-
290b.pdf.

Anonymous:1889:ALE

[12300] Anonymous. An automatic lamp extinguisher. Scientific American, 61(19):291, November 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n19/pdf/scientificamerican11091889-
291.pdf.

Anonymous:1889:ITCd

[12301] Anonymous. An improved thill coupling. Scientific American, 61 (19):291, November 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n19/pdf/scientificamerican11091889-
291a.pdf.

Anonymous:1889:IWPb

[12302] Anonymous. An improved washboard protector. Scientific American, 61(19):291, November 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n19/pdf/scientificamerican11091889-
291b.pdf.

Anonymous:1889:ISSb

[12303] Anonymous. An improved scaffold-supporting standard. Scientific American, 61(19):291, November 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v61/n19/pdf/scientificamerican11091889-
291c.pdf.

Anonymous:1889:IEP

[12304] Anonymous. An improved exhaust pipe for locomotives. Scientific American, 61(19):291–292, November 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v61/n19/pdf/scientificamerican11091889-
291d.pdf.

Anonymous:1889:NI

[12305] Anonymous. The Nampa image. Scientific American, 61(19): 292–293, November 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n19/pdf/scientificamerican11091889-
292.pdf.

Anonymous:1889:MWM

[12306] Anonymous. Modified Wimshurst machine. Scientific American, 61 (19):294–295, November 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n19/pdf/scientificamerican11091889-
294.pdf.

Anonymous:1889:CRD

[12307] Anonymous. A California rabbit drive. Scientific American, 61 (19):295–296, November 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n19/pdf/scientificamerican11091889-
295.pdf.

Anonymous:1889:Ibu

[12308] Anonymous. To inventors. Scientific American, 61(19):299, November 9, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n19/pdf/scientificamerican11091889-299.pdf.

Anonymous:1889:RPIbl

[12309] Anonymous. Recently patented inventions. Scientific American, 61(2):26, July 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n2/pdf/scientificamerican07131889-
26.pdf.

Anonymous:1889:BPbn

[12310] Anonymous. Business and personal. Scientific American, 61(2):26, July 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n2/pdf/scientificamerican07131889-26a.pdf.

Anonymous:1889:BNBu

[12311] Anonymous. Books: New books and publications. Scientific American, 61(2):26–27, July 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n2/pdf/scientificamerican07131889-
26b.pdf.

Anonymous:1889:NQbk

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n2/pdf/scientificamerican07131889-27.pdf.

Anonymous:1889:IIbl

[12313] Anonymous. Index of inventions. Scientific American, 61(2):27–28, July 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n2/pdf/scientificamerican07131889-27b.pdf.

Anonymous:1889:PNPbl

[12314] Anonymous. Patent notes: Patents. Scientific American, 61(2):30, July 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n2/pdf/scientificamerican07131889-30.pdf.

Anonymous:1889:CNN

[12315] Anonymous. Condition and needs of the Navy. Scientific American, 61(2):16–17, July 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n2/pdf/scientificamerican07131889-
16.pdf.

Anonymous:1889:IGTb

[12316] Anonymous. An improved grindstone tool holder. Scientific American, 61(2):18, July 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n2/pdf/scientificamerican07131889-
18.pdf.

Anonymous:1889:CUS

[12317] Anonymous. A chair for use on the sea beach. Scientific American, 61(2):18, July 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n2/pdf/scientificamerican07131889-
18a.pdf.

Anonymous:1889:ISHb

[12318] Anonymous. An improved sash holder. Scientific American, 61(2):18, July 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n2/pdf/scientificamerican07131889-18b.pdf.

Anonymous:1889:IDEc

[12319] Anonymous. An improved draught equalizer. Scientific American, 61(2):18, July 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n2/pdf/scientificamerican07131889-
18c.pdf.

Anonymous:1889:IFBd

[12320] Anonymous. Improved foot best for radiators. Scientific American, 61(2):18, July 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n2/pdf/scientificamerican07131889-
18d.pdf.

Anonymous:1889:NAN

[12321] Anonymous. The new additions to the Navy. Scientific American, 61(2):19–20, July 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n2/pdf/scientificamerican07131889-
19.pdf.

Anonymous:1889:TIW

[12322] Anonymous. Tare’s improved wash boiler. Scientific American, 61(2):20, July 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n2/pdf/scientificamerican07131889-
20.pdf.

Anonymous:1889:IOPa

[12323] Anonymous. An improved oyster pail. Scientific American, 61(2):20, July 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n2/pdf/scientificamerican07131889-20a.pdf.

Anonymous:1889:SET

[12324] Anonymous. Sims–Edison torpedo. Scientific American, 61(2):21, July 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n2/pdf/scientificamerican07131889-21.pdf.

Hopkins:1889:SB

[12325] Geo. M. Hopkins. Secondary battery. Scientific American, 61(2):22, July 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n2/pdf/scientificamerican07131889-22.pdf.

Anonymous:1889:CWG

[12326] Anonymous. Curiously worn gear wheel. Scientific American, 61(2):22, July 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n2/pdf/scientificamerican07131889-22a.pdf.

Pike:1889:MIS

[12327] M. L. Pike. Maternal instinct in spiders. Scientific American, 61(2):23, July 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n2/pdf/scientificamerican07131889-23.pdf.

Anonymous:1889:ISEb

[12328] Anonymous. An improved steam engine. Scientific American, 61(2):23–24, July 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n2/pdf/scientificamerican07131889-
23a.pdf.

Anonymous:1889:AHR

[12329] Anonymous. Accident on the Housatonic Railroad. Scientific American, 61(2):25, July 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n2/pdf/scientificamerican07131889-
25.pdf.

Anonymous:1889:Ibv

[12330] Anonymous. To inventors. Scientific American, 61(2):27, July 13, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n2/pdf/scientificamerican07131889-27a.pdf.

Anonymous:1889:Cba

[12331] Anonymous. Correspondence. Scientific American, 61(20):309, November 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n20/pdf/scientificamerican11161889-309.pdf.

Anonymous:1889:RPIbm

[12332] Anonymous. Recently patented inventions. Scientific American, 61(20):314, November 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n20/pdf/scientificamerican11161889-
314.pdf.

Anonymous:1889:BPbo

[12333] Anonymous. Business and personal. Scientific American, 61 (20):314, November 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n20/pdf/scientificamerican11161889-
314a.pdf.

Anonymous:1889:NQbl

[12334] Anonymous. Notes and queries. Scientific American, 61(20): 314–315, November 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n20/pdf/scientificamerican11161889-
314b.pdf.

Anonymous:1889:IIbm

[12335] Anonymous. Index of inventions. Scientific American, 61(20): 315–316, November 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n20/pdf/scientificamerican11161889-
315a.pdf.

Anonymous:1889:PNPbm

[12336] Anonymous. Patent notes: Patents. Scientific American, 61 (20):318, November 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n20/pdf/scientificamerican11161889-
318.pdf.

Anonymous:1889:ETc

[12337] Anonymous. The electric tachyscope. Scientific American, 61 (20):303, November 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n20/pdf/scientificamerican11161889-
303.pdf.

Anonymous:1889:IIWa

[12338] Anonymous. An important invention wanted. Scientific American, 61(20):304, November 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n20/pdf/scientificamerican11161889-
304.pdf.

Anonymous:1889:NME

[12339] Anonymous. The National Maritime Exhibition. Scientific American, 61(20):304–305, November 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n20/pdf/scientificamerican11161889-
304a.pdf.

Anonymous:1889:PNW

[12340] Anonymous. The purity of the new water supply for New York. Scientific American, 61(20):305, November 16, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v61/
n20/pdf/scientificamerican11161889-305.pdf.

Anonymous:1889:ILEa

[12341] Anonymous. An improved lamp extinguisher. Scientific American, 61(20):306, November 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n20/pdf/scientificamerican11161889-
306.pdf.

Anonymous:1889:IDMc

[12342] Anonymous. An improved dough mixer and kneader. Scientific American, 61(20):306, November 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n20/pdf/scientificamerican11161889-
306a.pdf.

Anonymous:1889:DB

[12343] Anonymous. Dry battery. Scientific American, 61(20):306, November 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n20/pdf/scientificamerican11161889-306b.pdf.

Anonymous:1889:IFH

[12344] Anonymous. An improved fireplace and heater. Scientific American, 61(20):306, November 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n20/pdf/scientificamerican11161889-
306c.pdf.

Anonymous:1889:ISFb

[12345] Anonymous. Improved shield for firemen or stokers. Scientific American, 61(20):307, November 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n20/pdf/scientificamerican11161889-
307.pdf.

Anonymous:1889:IRd

[12346] Anonymous. An improved refrigerator. Scientific American, 61 (20):307, November 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n20/pdf/scientificamerican11161889-
307a.pdf.

Anonymous:1889:IWC

[12347] Anonymous. An improved watch case pendant. Scientific American, 61(20):307, November 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n20/pdf/scientificamerican11161889-
307b.pdf.

Anonymous:1889:IBSc

[12348] Anonymous. An improved boot or shoe heel. Scientific American, 61 (20):307–308, November 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n20/pdf/scientificamerican11161889-
307c.pdf.

Anonymous:1889:IFEc

[12349] Anonymous. An improved fire escape. Scientific American, 61 (20):308, November 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n20/pdf/scientificamerican11161889-
308.pdf.

Anonymous:1889:TTT

[12350] Anonymous. Ten ton traveling crane. Scientific American, 61(20): 311–313, November 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n20/pdf/scientificamerican11161889-
311.pdf.

Anonymous:1889:IRGb

[12351] Anonymous. An improved road grader. Scientific American, 61 (20):313, November 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n20/pdf/scientificamerican11161889-
313.pdf.

Anonymous:1889:Ibx

[12352] Anonymous. To inventors. Scientific American, 61(20):315, November 16, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n20/pdf/scientificamerican11161889-315.pdf.

Anonymous:1889:Cbb

[12353] Anonymous. Correspondence. Scientific American, 61(21):329, November 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n21/pdf/scientificamerican11231889-329.pdf.

Anonymous:1889:RPIbn

[12354] Anonymous. Recently patented inventions. Scientific American, 61(21):330, November 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n21/pdf/scientificamerican11231889-
330.pdf.

Anonymous:1889:BPbp

[12355] Anonymous. Business and personal. Scientific American, 61 (21):330, November 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n21/pdf/scientificamerican11231889-
330a.pdf.

Anonymous:1889:NQbm (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n21/pdf/scientificamerican11231889-
330b.pdf.

Anonymous:1889:IIbn

[12357] Anonymous. Index of inventions. Scientific American, 61(21): 331–332, November 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n21/pdf/scientificamerican11231889-
331a.pdf.

Anonymous:1889:PNPbn

[12358] Anonymous. Patent notes: Patents. Scientific American, 61 (21):332, November 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n21/pdf/scientificamerican11231889-
332.pdf.

Anonymous:1889:MPR

[12359] Anonymous. The Mount Pilatus Railway. Scientific American, 61 (21):319–320, November 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n21/pdf/scientificamerican11231889-
319.pdf.

Anonymous:1889:FCH

[12360] Anonymous. The food and care of horses. Scientific American, 61(21):320, November 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n21/pdf/scientificamerican11231889-
320.pdf.

Anonymous:1889:SAFb

[12361] Anonymous. Safeguards against fires. Scientific American, 61(21): 320–321, November 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n21/pdf/scientificamerican11231889-
320a.pdf.

Anonymous:1889:ISMc

[12362] Anonymous. An improved sandpapering machine. Scientific American, 61(21):322, November 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n21/pdf/scientificamerican11231889-
322.pdf.

Anonymous:1889:APS

[12363] Anonymous. An air purifier for sleeping apartments, sick rooms, etc. Scientific American, 61(21):322, November 23, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v61/
n21/pdf/scientificamerican11231889-322a.pdf.

Anonymous:1889:ICWb

[12364] Anonymous. Improved coal washer and separator. Scientific American, 61(21):322, November 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n21/pdf/scientificamerican11231889-
322b.pdf.

Anonymous:1889:ILAa

[12365] Anonymous. An improved locomotive ash pan cleaner. Scientific American, 61(21):322, November 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n21/pdf/scientificamerican11231889-
322c.pdf.

Anonymous:1889:IBPb

[12366] Anonymous. An improved bottle protector. Scientific American, 61(21):323, November 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n21/pdf/scientificamerican11231889-
323.pdf.

Anonymous:1889:IBCh

[12367] Anonymous. An improved biscuit cutter. Scientific American, 61 (21):323, November 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n21/pdf/scientificamerican11231889-
323a.pdf.

Anonymous:1889:ISMd

[12368] Anonymous. An improved stamp mill. Scientific American, 61 (21):323, November 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n21/pdf/scientificamerican11231889-
323b.pdf.

Anonymous:1889:IDCc

[12369] Anonymous. An improved drive chain. Scientific American, 61 (21):323, November 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n21/pdf/scientificamerican11231889-
323c.pdf.

Anonymous:1889:IRSc

[12370] Anonymous. An improved razor strop. Scientific American, 61 (21):323, November 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n21/pdf/scientificamerican11231889-
323d.pdf.

Anonymous:1889:ISTb

[12371] Anonymous. An improved screw-threading die. Scientific American, 61(21):323, November 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n21/pdf/scientificamerican11231889-
323e.pdf.

Anonymous:1889:CTA

[12372] Anonymous. The Congress of the Three Americas. Scientific American, 61(21):324–326, November 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n21/pdf/scientificamerican11231889-
324.pdf.

Anonymous:1889:GSR

[12373] Anonymous. Girard’s sliding railway. Scientific American, 61(21): 327–329, November 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n21/pdf/scientificamerican11231889-
327.pdf.

Anonymous:1889:Iby

[12374] Anonymous. To inventors. Scientific American, 61(21):331, November 23, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n21/pdf/scientificamerican11231889-331.pdf.

Anonymous:1889:RPIbo

[12375] Anonymous. Recently patented inventions. Scientific American, 61(22):345, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n22/pdf/scientificamerican11301889-
345.pdf.

Anonymous:1889:BPbq

[12376] Anonymous. Business and personal. Scientific American, 61 (22):346, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n22/pdf/scientificamerican11301889-
346.pdf.

Anonymous:1889:NQbn

[12377] Anonymous. Notes and queries. Scientific American, 61(22): 346, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n22/pdf/scientificamerican11301889-
346a.pdf.

Anonymous:1889:IIbo

[12378] Anonymous. Index of inventions. Scientific American, 61(22): 347–348, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n22/pdf/scientificamerican11301889-
347a.pdf.

Anonymous:1889:PNPbo

[12379] Anonymous. Patent notes: Patents. Scientific American, 61 (22):348, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n22/pdf/scientificamerican11301889-
348.pdf.

Anonymous:1889:USB

[12380] Anonymous. The United States of Brazil. Scientific American, 61(22):336, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n22/pdf/scientificamerican11301889-
336.pdf.

Anonymous:1889:SLD

[12381] Anonymous. Stanley’s latest discoveries. Scientific American, 61 (22):336, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n22/pdf/scientificamerican11301889-
336a.pdf.

Anonymous:1889:ONB

[12382] Anonymous. Our new book. Scientific American, 61(22):336, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n22/pdf/scientificamerican11301889-336b.pdf.

Anonymous:1889:AE

[12383] Anonymous. The American Exhibition of 1892. Scientific American, 61(22):336–337, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n22/pdf/scientificamerican11301889-
336c.pdf.

Anonymous:1889:BC

[12384] Anonymous. Brewing chemistry. Scientific American, 61(22): 337, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n22/pdf/scientificamerican11301889-
337.pdf.

Anonymous:1889:PPD

[12385] Anonymous. Position of the planets for December. Scientific American, 61(22):337, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n22/pdf/scientificamerican11301889-
337a.pdf.

Anonymous:1889:STA

[12386] Anonymous. Some thoughts about China. Scientific American, 61 (22):337, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n22/pdf/scientificamerican11301889-
337b.pdf.

Anonymous:1889:JPS

[12387] Anonymous. Johnston’s patent side drum. Scientific American, 61(22):338, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n22/pdf/scientificamerican11301889-
338.pdf.

Anonymous:1889:OI

[12388] Anonymous. An optical illusion. Scientific American, 61(22): 338, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n22/pdf/scientificamerican11301889-
338a.pdf.

Anonymous:1889:IDDb

[12389] Anonymous. An improved dish drier. Scientific American, 61 (22):338, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n22/pdf/scientificamerican11301889-
338b.pdf.

Anonymous:1889:IVT

[12390] Anonymous. An improved violin tail piece. Scientific American, 61(22):338, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n22/pdf/scientificamerican11301889-
338c.pdf.

Anonymous:1889:IMM

[12391] Anonymous. An improved mechanical mixer. Scientific American, 61(22):339, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n22/pdf/scientificamerican11301889-
339.pdf.

Anonymous:1889:ICL

[12392] Anonymous. An improved combination lock. Scientific American, 61(22):339, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n22/pdf/scientificamerican11301889-
339a.pdf.

Anonymous:1889:RAT

[12393] Anonymous. Reversing attachment for type writers. Scientific American, 61(22):339, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n22/pdf/scientificamerican11301889-
339b.pdf.

Anonymous:1889:IOPb

[12394] Anonymous. An improved organ pallet. Scientific American, 61 (22):339, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n22/pdf/scientificamerican11301889-
339c.pdf.

Anonymous:1889:IFI

[12395] Anonymous. An improved family ice machine. Scientific American, 61(22):339–340, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n22/pdf/scientificamerican11301889-
339d.pdf.

Anonymous:1889:H

[12396] Anonymous. Heat. Scientific American, 61(22):340, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n22/pdf/scientificamerican11301889-340.pdf.

Anonymous:1889:NDD

[12397] Anonymous. The new dry dock, Brooklyn Navy Yard. Scientific American, 61(22):341, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n22/pdf/scientificamerican11301889-
341.pdf.

Anonymous:1889:PBE

[12398] Anonymous. The proposed bridge over the English Channel. Scientific American, 61(22):342, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v61/n22/pdf/scientificamerican11301889-
342.pdf.

Anonymous:1889:MB

[12399] Anonymous. The military bicycle. Scientific American, 61(22): 343, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n22/pdf/scientificamerican11301889-
343.pdf.

Pike:1889:HS

[12400] Nicolas Pike. The hoof snake. Scientific American, 61(22):344, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n22/pdf/scientificamerican11301889-344.pdf.

Anonymous:1889:Ibz

[12401] Anonymous. To inventors. Scientific American, 61(22):347, November 30, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n22/pdf/scientificamerican11301889-347.pdf.

Anonymous:1889:Cbc

[12402] Anonymous. Correspondence. Scientific American, 61(23):357, December 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n23/pdf/scientificamerican12071889-357.pdf.

Anonymous:1889:RPIbp

[12403] Anonymous. Recently patented inventions. Scientific American, 61(23):361, December 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n23/pdf/scientificamerican12071889-
361.pdf.

Anonymous:1889:BNBv

[12404] Anonymous. Books: New books and publications. Scientific American, 61(23):361, December 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n23/pdf/scientificamerican12071889-
361a.pdf.

Anonymous:1889:BPbr 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n23/pdf/scientificamerican12071889-
362.pdf.

Anonymous:1889:NQbo

[12406] Anonymous. Notes and queries. Scientific American, 61(23): 362–363, December 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n23/pdf/scientificamerican12071889-
362a.pdf.

Anonymous:1889:IIbp

[12407] Anonymous. Index of inventions. Scientific American, 61(23): 363–364, December 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n23/pdf/scientificamerican12071889-
363a.pdf.

Anonymous:1889:PNPbp

[12408] Anonymous. Patent notes: Patents. Scientific American, 61 (23):366, December 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n23/pdf/scientificamerican12071889-
366.pdf.

Anonymous:1889:SOT

[12409] Anonymous. Shortening the ocean trip half a day. Scientific American, 61(23):352, December 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n23/pdf/scientificamerican12071889-
352.pdf.

Anonymous:1889:PMF

[12410] Anonymous. A plea for the merchant fleet. Scientific American, 61(23):352, December 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n23/pdf/scientificamerican12071889-
352a.pdf.

Anonymous:1889:SSSb

[12411] Anonymous. Saccharin, the sweetest of sweets. Scientific American, 61(23):352–353, December 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n23/pdf/scientificamerican12071889-
352b.pdf.

Anonymous:1889:PNl

[12412] Anonymous. Photographic notes. Scientific American, 61(23):353, December 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n23/pdf/scientificamerican12071889-353.pdf.

Anonymous:1889:IBMa

[12413] Anonymous. An improved brick mould. Scientific American, 61 (23):354, December 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n23/pdf/scientificamerican12071889-
354.pdf.

Anonymous:1889:IGHa

[12414] Anonymous. An improved gate hinge. Scientific American, 61 (23):354, December 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n23/pdf/scientificamerican12071889-
354a.pdf.

Anonymous:1889:ILAb

[12415] Anonymous. An improved lathe attachment. Scientific American, 61(23):354, December 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n23/pdf/scientificamerican12071889-
354b.pdf.

Anonymous:1889:IPDc

[12416] Anonymous. An improved portable dam. Scientific American, 61 (23):354–355, December 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n23/pdf/scientificamerican12071889-
354c.pdf.

Anonymous:1889:NSG

[12417] Anonymous. New Spanish guns. Scientific American, 61(23): 355–356, December 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n23/pdf/scientificamerican12071889-
355.pdf.

Anonymous:1889:BME

[12418] Anonymous. The Boston Maritime Exhibition. Scientific American, 61(23):358–359, December 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n23/pdf/scientificamerican12071889-
358.pdf.

Anonymous:1889:MCb

[12419] Anonymous. A mammoth cactus. Scientific American, 61(23): 359–360, December 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n23/pdf/scientificamerican12071889-
359.pdf.

Anonymous:1889:MSB

[12420] Anonymous. The moving stone of Buenos Ayres. Scientific American, 61(23):360, December 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n23/pdf/scientificamerican12071889-
360.pdf.

Anonymous:1889:MA

[12421] Anonymous. Molecular actions. Scientific American, 61(23):360, December 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n23/pdf/scientificamerican12071889-360a.pdf.

Anonymous:1889:Ich

[12422] Anonymous. To inventors. Scientific American, 61(23):363, December 7, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n23/pdf/scientificamerican12071889-363.pdf.

Anonymous:1889:Cbd

[12423] Anonymous. Correspondence. Scientific American, 61(24):373, December 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n24/pdf/scientificamerican12141889-373.pdf.

Anonymous:1889:RPIbq

[12424] Anonymous. Recently patented inventions. Scientific American, 61(24):378, December 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n24/pdf/scientificamerican12141889-
378.pdf.

Anonymous:1889:BPbs

[12425] Anonymous. Business and personal. Scientific American, 61 (24):378, December 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n24/pdf/scientificamerican12141889-
378a.pdf.

Anonymous:1889:NQbp

[12426] Anonymous. Notes and queries. Scientific American, 61(24): 378, December 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n24/pdf/scientificamerican12141889-
378b.pdf.

Anonymous:1889:IIbq

[12427] Anonymous. Index of inventions. Scientific American, 61(24): 378–379, December 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n24/pdf/scientificamerican12141889-
378d.pdf.

Anonymous:1889:PNPbq

[12428] Anonymous. Patent notes: Patents. Scientific American, 61 (24):380, December 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n24/pdf/scientificamerican12141889-
380.pdf.

Anonymous:1889:NCW

[12429] Anonymous. The new Chicago Water Works — the Lake Tunnel intakes, and caissons. Scientific American, 61(24):367–368, December 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n24/pdf/scientificamerican12141889-367.pdf.

Anonymous:1889:SAE

[12430] Anonymous. The Scientific American — Edición Española. Scientific American, 61(24):368, December 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v61/n24/pdf/scientificamerican12141889-
368.pdf.

Anonymous:1889:EPS

[12431] Anonymous. Ericsson’s physical strength. Scientific American, 61 (24):368–369, December 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n24/pdf/scientificamerican12141889-
368a.pdf.

Anonymous:1889:PTC

[12432] Anonymous. A pocket ticket case. Scientific American, 61(24): 370, December 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n24/pdf/scientificamerican12141889-
370.pdf.

Anonymous:1889:ISDc

[12433] Anonymous. An improved seed drill and fertilizer distributer. Scientific American, 61(24):370, December 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v61/n24/pdf/scientificamerican12141889-
370a.pdf.

Anonymous:1889:ICDb

[12434] Anonymous. An improved car dumping device. Scientific American, 61(24):370, December 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n24/pdf/scientificamerican12141889-
370b.pdf.

Anonymous:1889:IDSd

[12435] Anonymous. Improved device for sanding railroad tracks. Scientific American, 61(24):370, December 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v61/n24/pdf/scientificamerican12141889-
370c.pdf.

Anonymous:1889:UAC

[12436] Anonymous. An unloading attachment for cars. Scientific American, 61(24):370, December 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n24/pdf/scientificamerican12141889-
370d.pdf.

Anonymous:1889:BCF

[12437] Anonymous. The beaver (Castor Fiber). Scientific American, 61 (24):371, December 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n24/pdf/scientificamerican12141889-
371.pdf.

Anonymous:1889:FP

[12438] Anonymous. A feat in photography. Scientific American, 61 (24):372, December 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n24/pdf/scientificamerican12141889-
372.pdf.

Anonymous:1889:HMSa

[12439] Anonymous. How to make a simple electric telephone requiring no battery. Scientific American, 61(24):374–375, December 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n24/pdf/scientificamerican12141889-374.pdf.

Anonymous:1889:LMR

[12440] Anonymous. Lewis Morris Rutherfurd. Scientific American, 61 (24):375–376, December 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n24/pdf/scientificamerican12141889-
375.pdf.

Anonymous:1889:LFF

[12441] Anonymous. The luminous fountains at the French Exposition. Scientific American, 61(24):376–377, December 14, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v61/
n24/pdf/scientificamerican12141889-376.pdf.

Anonymous:1889:Ici

[12442] Anonymous. To inventors. Scientific American, 61(24):378, December 14, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n24/pdf/scientificamerican12141889-378c.pdf.

Anonymous:1889:RPIbr

[12443] Anonymous. Recently patented inventions. Scientific American, 61(25):393, December 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n25/pdf/scientificamerican12211889-
393.pdf.

Anonymous:1889:BPbt

[12444] Anonymous. Business and personal. Scientific American, 61 (25):394, December 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n25/pdf/scientificamerican12211889-
394.pdf.

Anonymous:1889:BNBw

[12445] Anonymous. Books: New books and publication. Scientific American, 61(25):394, December 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n25/pdf/scientificamerican12211889-
394a.pdf.

Anonymous:1889:NQbq

[12446] Anonymous. Notes and queries. Scientific American, 61(25): 394–395, December 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n25/pdf/scientificamerican12211889-
394b.pdf.

Anonymous:1889:IIbr

[12447] Anonymous. Index of inventions. Scientific American, 61(25): 395–396, December 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n25/pdf/scientificamerican12211889-
395a.pdf.

Anonymous:1889:PNPbr

[12448] Anonymous. Patent notes: Patents. Scientific American, 61 (25):398, December 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n25/pdf/scientificamerican12211889-
398.pdf.

Anonymous:1889:NN

[12449] Anonymous. Naval notes. Scientific American, 61(25):384, December 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n25/pdf/scientificamerican12211889-384.pdf.

Anonymous:1889:HTC

[12450] Anonymous. High tension currents in crowded centers. Scientific American, 61(25):384–385, December 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v61/n25/pdf/scientificamerican12211889-
384a.pdf.

Anonymous:1889:ASC

[12451] Anonymous. Amito’s stair climber. Scientific American, 61(25): 386, December 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n25/pdf/scientificamerican12211889-
386.pdf.

Anonymous:1889:IDCd

[12452] Anonymous. An improved dumping car. Scientific American, 61 (25):386, December 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n25/pdf/scientificamerican12211889-
386a.pdf.

Anonymous:1889:IWSc

[12453] Anonymous. An improved window screen. Scientific American, 61(25):387, December 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n25/pdf/scientificamerican12211889-
387.pdf.

Anonymous:1889:ISg

[12454] Anonymous. An improved screwdriver. Scientific American, 61 (25):387, December 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n25/pdf/scientificamerican12211889-
387a.pdf.

Anonymous:1889:ICCi

[12455] Anonymous. An improved chimney cover. Scientific American, 61 (25):387, December 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n25/pdf/scientificamerican12211889-
387b.pdf.

Anonymous:1889:ICRb

[12456] Anonymous. An improved curtain ring attachment. Scientific American, 61(25):387, December 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n25/pdf/scientificamerican12211889-
387c.pdf.

Anonymous:1889:IDF

[12457] Anonymous. An improved device for felling trees. Scientific American, 61(25):387–388, December 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n25/pdf/scientificamerican12211889-
387d.pdf.

Anonymous:1889:PST

[12458] Anonymous. Passenger ships of the transatlantic service. Scientific American, 61(25):389, December 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v61/n25/pdf/scientificamerican12211889-
389.pdf.

Anonymous:1889:STT

[12459] Anonymous. The saber-toothed tiger. Scientific American, 61 (25):390, December 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n25/pdf/scientificamerican12211889-
390.pdf.

Anonymous:1889:UPF

[12460] Anonymous. Utilization of poor fuel. Scientific American, 61 (25):390, December 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n25/pdf/scientificamerican12211889-
390a.pdf.

Anonymous:1889:HMSb

[12461] Anonymous. Henry M. Stanley. Scientific American, 61(25): 391–392, December 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n25/pdf/scientificamerican12211889-
391.pdf.

Anonymous:1889:MFC

[12462] Anonymous. Multnomah Falls on Columbia River. Scientific American, 61(25):392, December 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n25/pdf/scientificamerican12211889-
392.pdf.

Anonymous:1889:Icj

[12463] Anonymous. To inventors. Scientific American, 61(25):395, December 21, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n25/pdf/scientificamerican12211889-395.pdf.

Anonymous:1889:RPIbs

[12464] Anonymous. Recently patented inventions. Scientific American, 61(26):408, December 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n26/pdf/scientificamerican12281889-
408.pdf.

Anonymous:1889:BPbu

[12465] Anonymous. Business and personal. Scientific American, 61 (26):408, December 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n26/pdf/scientificamerican12281889-
408a.pdf.

Anonymous:1889:NQbr

[12466] Anonymous. Notes and queries. Scientific American, 61(26): 408–409, December 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n26/pdf/scientificamerican12281889-
408b.pdf.

Anonymous:1889:BNBx

[12467] Anonymous. Books: New books and publications. Scientific American, 61(26):409, December 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n26/pdf/scientificamerican12281889-
409.pdf.

Anonymous:1889:IIbs

[12468] Anonymous. Index of inventions. Scientific American, 61(26): 409–410, December 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n26/pdf/scientificamerican12281889-
409b.pdf.

Anonymous:1889:Ick

[12469] Anonymous. Index. Scientific American, 61(26):411–412, December 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n26/pdf/scientificamerican12281889-411.pdf.

Anonymous:1889:PNPbs

[12470] Anonymous. Patent notes: Patents. Scientific American, 61 (26):414, December 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n26/pdf/scientificamerican12281889-
414.pdf.

Anonymous:1889:ELT

[12471] Anonymous. The electric-light troubles in New York. Scientific American, 61(26):400, December 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n26/pdf/scientificamerican12281889-
400.pdf.

Anonymous:1889:PPJd

[12472] Anonymous. Position of the planets for January. Scientific American, 61(26):400–401, December 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n26/pdf/scientificamerican12281889-
400a.pdf.

Anonymous:1889:EPN

[12473] Anonymous. Elias Parkman Needham. Scientific American, 61 (26):401, December 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n26/pdf/scientificamerican12281889-
401.pdf.

Anonymous:1889:IRW

[12474] Anonymous. An improved raw water cut-off. Scientific American, 61(26):402, December 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n26/pdf/scientificamerican12281889-
402.pdf.

Anonymous:1889:IACc

[12475] Anonymous. An improved air cooling apparatus. Scientific American, 61(26):402, December 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n26/pdf/scientificamerican12281889-
402a.pdf.

Anonymous:1889:APC

[12476] Anonymous. An automatic paper carrier for drying rooms. Scientific American, 61(26):402, December 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v61/n26/pdf/scientificamerican12281889-
402b.pdf.

Anonymous:1889:IER

[12477] Anonymous. An improved electric railway signal. Scientific American, 61(26):402, December 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n26/pdf/scientificamerican12281889-
402c.pdf.

Anonymous:1889:IRR

[12478] Anonymous. An improved railway rail tie. Scientific American, 61 (26):402–403, December 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n26/pdf/scientificamerican12281889-
402d.pdf.

Anonymous:1889:GTC

[12479] Anonymous. The great traveling crane at Trubia. Scientific American, 61(26):403, December 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n26/pdf/scientificamerican12281889-
403.pdf.

Anonymous:1889:IFCd

[12480] Anonymous. An improved friction clutch. Scientific American, 61 (26):403, December 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n26/pdf/scientificamerican12281889-
403a.pdf.

Anonymous:1889:APW

[12481] Anonymous. Armor plate for war ships. Scientific American, 61 (26):404, December 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n26/pdf/scientificamerican12281889-
404.pdf.

Anonymous:1889:NSR

[12482] Anonymous. A new system of rapid telegraphy. Scientific American, 61(26):405, December 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n26/pdf/scientificamerican12281889-
405.pdf.

Pike:1889:PFI

[12483] Nicolas Pike. Poisonous fish of the Indian Ocean. Scientific American, 61(26):406, December 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n26/pdf/scientificamerican12281889-
406.pdf.

Anonymous:1889:NSW

[12484] Anonymous. The nitrate of soda works and fields of Chile. Scientific American, 61(26):407, December 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v61/n26/pdf/scientificamerican12281889-
407.pdf.

Anonymous:1889:Icm

[12485] Anonymous. To inventors. Scientific American, 61(26):409, December 28, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n26/pdf/scientificamerican12281889-409a.pdf.

Anonymous:1889:RPIbt

[12486] Anonymous. Recently patented inventions. Scientific American, 61(3):42, July 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n3/pdf/scientificamerican07201889-
42.pdf.

Anonymous:1889:BNBy

[12487] Anonymous. Books: New books and publications. Scientific American, 61(3):42, July 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n3/pdf/scientificamerican07201889-
42a.pdf.

Anonymous:1889:BPbv

[12488] Anonymous. Business and personal. Scientific American, 61(3):43, July 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n3/pdf/scientificamerican07201889-43.pdf.

Anonymous:1889:NQbs

[12489] Anonymous. Notes and queries. Scientific American, 61(3):43, July 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n3/pdf/scientificamerican07201889-43a.pdf.

Anonymous:1889:IIbt

[12490] Anonymous. Index of inventions. Scientific American, 61(3):43–45, July 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n3/pdf/scientificamerican07201889-43c.pdf.

Anonymous:1889:PNPbt

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n3/pdf/scientificamerican07201889-46.pdf.

Anonymous:1889:NTB

[12492] Anonymous. Novel torpedo boat designing and maneuvering. Scientific American, 61(3):32, July 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n3/pdf/scientificamerican07201889-
32.pdf.

Anonymous:1889:PBH

[12493] Anonymous. Pennsylvania Board of Health notes. Scientific American, 61(3):32–33, July 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n3/pdf/scientificamerican07201889-
32a.pdf.

Anonymous:1889:MMa

[12494] Anonymous. A mechanical movement. Scientific American, 61(3):34, July 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n3/pdf/scientificamerican07201889-34.pdf.

Anonymous:1889:IWM

[12495] Anonymous. An improved washing machine. Scientific American, 61(3):34, July 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n3/pdf/scientificamerican07201889-
34a.pdf.

Anonymous:1889:IABb

[12496] Anonymous. An improved annealing box. Scientific American, 61(3):34, July 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n3/pdf/scientificamerican07201889-34b.pdf.

Anonymous:1889:BIF

[12497] Anonymous. Barton’s improved fence. Scientific American, 61(3):34, July 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n3/pdf/scientificamerican07201889-34c.pdf.

Anonymous:1889:GHH

[12498] Anonymous. Gerstle’s hinged handle for pans, etc. Scientific American, 61(3):34, July 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n3/pdf/scientificamerican07201889-
34d.pdf.

Anonymous:1889:EUA

[12499] Anonymous. Eberhardt’s universal automatic gear cutting machine. Scientific American, 61(3):35, July 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v61/n3/pdf/scientificamerican07201889-
35.pdf.

Anonymous:1889:IMO

[12500] Anonymous. An improved magnetic ore separator. Scientific American, 61(3):35–36, July 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n3/pdf/scientificamerican07201889-
35a.pdf.

Anonymous:1889:IGHb

[12501] Anonymous. An improves globe holder. Scientific American, 61(3):36, July 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n3/pdf/scientificamerican07201889-36.pdf.

Drake:1889:NUC

[12502] F. T. Drake. The new U.S. cruiser Baltimore. Scientific American, 61(3):36–37, July 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n3/pdf/scientificamerican07201889-
36a.pdf.

Anonymous:1889:MMb

[12503] Anonymous. Maria Mitchell. Scientific American, 61(3):38, July 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n3/pdf/scientificamerican07201889-38.pdf.

Pike:1889:NBF

[12504] Nicolas Pike. A nest-building frog. Scientific American, 61(3):39, July 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n3/pdf/scientificamerican07201889-39.pdf.

Anonymous:1889:IBCi

[12505] Anonymous. Improved bar channeling stone drill. Scientific American, 61(3):39–41, July 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n3/pdf/scientificamerican07201889-
39a.pdf.

Anonymous:1889:Icn

[12506] Anonymous. To inventors. Scientific American, 61(3):43, July 20, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n3/pdf/scientificamerican07201889-43b.pdf.

Anonymous:1889:Cbe

[12507] Anonymous. Correspondence. Scientific American, 61(4):53, July 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n4/pdf/scientificamerican07271889-53.pdf.

Anonymous:1889:RPIbu

[12508] Anonymous. Recently patented inventions. Scientific American, 61(4):58, July 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n4/pdf/scientificamerican07271889-
58.pdf.

Anonymous:1889:BPbw

[12509] Anonymous. Business and personal. Scientific American, 61(4):58, July 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n4/pdf/scientificamerican07271889-58a.pdf.

Anonymous:1889:NQbt

[12510] Anonymous. Notes and queries. Scientific American, 61(4):58–59, July 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n4/pdf/scientificamerican07271889-58b.pdf.

Anonymous:1889:BNBz

[12511] Anonymous. Books: New books and publications. Scientific American, 61(4):59, July 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n4/pdf/scientificamerican07271889-
59.pdf.

Anonymous:1889:IIbu

[12512] Anonymous. Index of inventions. Scientific American, 61(4):59–62, July 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n4/pdf/scientificamerican07271889-59b.pdf.

Anonymous:1889:IIbv

[12513] Anonymous. Index of inventions. Scientific American, 61(4):62, July 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n4/pdf/scientificamerican07271889-62.pdf.

Anonymous:1889:LCA

[12514] Anonymous. Loss of the Campbell air ship. Scientific American, 61(4):47, July 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n4/pdf/scientificamerican07271889-
47.pdf.

Anonymous:1889:FAS

[12515] Anonymous. A fast American steam line to Europe. Scientific American, 61(4):48, July 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n4/pdf/scientificamerican07271889-
48.pdf.

Anonymous:1889:FJS

[12516] Anonymous. Frederick J. Seymour. Scientific American, 61(4):48, July 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n4/pdf/scientificamerican07271889-48a.pdf.

Anonymous:1889:PPAb

[12517] Anonymous. Position of the planets in August. Scientific American, 61(4):49, July 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n4/pdf/scientificamerican07271889-
49.pdf.

Anonymous:1889:HGB

[12518] Anonymous. Hints on glass blowing. Scientific American, 61(4):50, July 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n4/pdf/scientificamerican07271889-50.pdf.

Anonymous:1889:IVSc

[12519] Anonymous. An improved valve for steam engines. Scientific American, 61(4):50, July 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n4/pdf/scientificamerican07271889-
50a.pdf.

Anonymous:1889:INa

[12520] Anonymous. An improved necktie. Scientific American, 61(4):51, July 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n4/pdf/scientificamerican07271889-51.pdf.

Anonymous:1889:EPC

[12521] Anonymous. Electric propulsion for common roads. Scientific American, 61(4):51, July 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n4/pdf/scientificamerican07271889-
51a.pdf.

Anonymous:1889:IHd

[12522] Anonymous. An improved harness. Scientific American, 61(4):51, July 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n4/pdf/scientificamerican07271889-51b.pdf.

Anonymous:1889:ELHb

[12523] Anonymous. Electric lamp hanger. Scientific American, 61(4):51–52, July 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n4/pdf/scientificamerican07271889-51c.pdf.

Holden:1889:RDN

[12524] Edward S. Holden. Recent discoveries in the nebulæ by means of photography. Scientific American, 61(4):54–56, July 27, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v61/
n4/pdf/scientificamerican07271889-54.pdf.

Anonymous:1889:PEh

[12525] Anonymous. The Paris Exhibition. Scientific American, 61(4):57, July 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n4/pdf/scientificamerican07271889-57.pdf.

Anonymous:1889:Icq

[12526] Anonymous. To inventors. Scientific American, 61(4):59, July 27, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n4/pdf/scientificamerican07271889-59a.pdf.

Anonymous:1889:RPIbv

[12527] Anonymous. Recently patented inventions. Scientific American, 61(5):73, August 3, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n5/pdf/scientificamerican08031889-
73.pdf.

Anonymous:1889:BNBba

[12528] Anonymous. Books: New books and publications. Scientific American, 61(5):73, August 3, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n5/pdf/scientificamerican08031889-
73a.pdf.

Anonymous:1889:BPbx

[12529] Anonymous. Business and personal. Scientific American, 61(5):74, August 3, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n5/pdf/scientificamerican08031889-74.pdf.

Anonymous:1889:NQbu

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n5/pdf/scientificamerican08031889-74a.pdf.

Anonymous:1889:IIbw

[12531] Anonymous. Index of inventions. Scientific American, 61(5):75–78, August 3, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n5/pdf/scientificamerican08031889-75a.pdf.

Anonymous:1889:PNPbu

[12532] Anonymous. Patent notes: Patents. Scientific American, 61(5):78, August 3, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n5/pdf/scientificamerican08031889-78.pdf.

Anonymous:1889:TS

[12533] Anonymous. Telemeter system. Scientific American, 61(5):63–64, August 3, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n5/pdf/scientificamerican08031889-63.pdf.

Anonymous:1889:WK

[12534] Anonymous. Working for knowledge. Scientific American, 61(5):64, August 3, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n5/pdf/scientificamerican08031889-64.pdf.

Anonymous:1889:NYW

[12535] Anonymous. The New York Water Supply — how the people suffer from lack of pressure. Scientific American, 61(5):64–67, August 3, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n5/pdf/scientificamerican08031889-64a.pdf.

Anonymous:1889:IBPc

[12536] Anonymous. An improved baling press. Scientific American, 61(5):67, August 3, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n5/pdf/scientificamerican08031889-67.pdf.

Anonymous:1889:ICu

[12537] Anonymous. An improved cultivator. Scientific American, 61(5):67, August 3, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n5/pdf/scientificamerican08031889-67a.pdf.

Anonymous:1889:ISPd

[12538] Anonymous. An improved seed planter and fertilizer distributer. Scientific American, 61(5):67, August 3, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v61/n5/pdf/scientificamerican08031889-
67b.pdf.

Anonymous:1889:ICx

[12539] Anonymous. An improved cultivator. Scientific American, 61(5):67–68, August 3, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n5/pdf/scientificamerican08031889-67c.pdf.

Anonymous:1889:PEj

[12540] Anonymous. To Paris Exhibition. Scientific American, 61(5):68–70, August 3, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n5/pdf/scientificamerican08031889-68.pdf.

Anonymous:1889:LF

[12541] Anonymous. The lightning flash. Scientific American, 61(5):70, August 3, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n5/pdf/scientificamerican08031889-70.pdf.

Anonymous:1889:HPG

[12542] Anonymous. How to perforate glass. Scientific American, 61(5):70–71, August 3, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n5/pdf/scientificamerican08031889-70a.pdf.

Sheldon:1889:RSM

[12543] S. L. Sheldon. The Retsof salt mines at Greigsville, N.Y. Scientific American, 61(5):71–72, August 3, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n5/pdf/scientificamerican08031889-
71.pdf.

Anonymous:1889:SSE

[12544] Anonymous. Simple scientific experiments. Scientific American, 61(5):72, August 3, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n5/pdf/scientificamerican08031889-
72.pdf.

Anonymous:1889:ILJb

[12545] Anonymous. An improved lifting jack. Scientific American, 61(5):72, August 3, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n5/pdf/scientificamerican08031889-72a.pdf.

Anonymous:1889:IWPc

[12546] Anonymous. An improved wall paper protector. Scientific American, 61(5):72, August 3, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n5/pdf/scientificamerican08031889-
72b.pdf.

Anonymous:1889:IFa

[12547] Anonymous. An improved fence. Scientific American, 61(5):72–73, August 3, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n5/pdf/scientificamerican08031889-72c.pdf.

Anonymous:1889:Icy

[12548] Anonymous. To inventors. Scientific American, 61(5):75, August 3, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n5/pdf/scientificamerican08031889-75.pdf.

Anonymous:1889:CCb

[12549] Anonymous. Correspondence: correspondence. Scientific American, 61(6):85–86, August 10, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n6/pdf/scientificamerican08101889-
85a.pdf.

Anonymous:1889:RPIbw

[12550] Anonymous. Recently patented inventions: Recently patented invention. Scientific American, 61(6):90, August 10, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v61/
n6/pdf/scientificamerican08101889-90.pdf.

Anonymous:1889:BPby

[12551] Anonymous. Business and personal. Scientific American, 61(6):90, August 10, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n6/pdf/scientificamerican08101889-90a.pdf.

Anonymous:1889:NQbv

[12552] Anonymous. Notes and queries. Scientific American, 61(6):90–91, August 10, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n6/pdf/scientificamerican08101889-90b.pdf.

Anonymous:1889:IIbx

[12553] Anonymous. Index of inventions. Scientific American, 61(6):91–93, August 10, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n6/pdf/scientificamerican08101889-91a.pdf.

Anonymous:1889:PNPbv

[12554] Anonymous. Patent notes: Patents. Scientific American, 61(6):94, August 10, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n6/pdf/scientificamerican08101889-94.pdf.

Anonymous:1889:DBS

[12555] Anonymous. Dr. Brown-Sequard’s recent experiments. Scientific American, 61(6):80, August 10, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n6/pdf/scientificamerican08101889-
80.pdf.

Anonymous:1889:NGA

[12556] Anonymous. The new German artillery tactics. Scientific American, 61(6):80–82, August 10, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n6/pdf/scientificamerican08101889-
80a.pdf.

Anonymous:1889:SCL

[12557] Anonymous. A spark conductor for locomotives. Scientific American, 61(6):82, August 10, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n6/pdf/scientificamerican08101889-
82.pdf.

Anonymous:1889:IPS

[12558] Anonymous. An improved propeller-shaft bearing. Scientific American, 61(6):82, August 10, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n6/pdf/scientificamerican08101889-
82a.pdf.

Anonymous:1889:ILf

[12559] Anonymous. An improved lock. Scientific American, 61(6):82, August 10, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n6/pdf/scientificamerican08101889-82b.pdf.

Anonymous:1889:IASd

[12560] Anonymous. An improved ash sifter. Scientific American, 61 (6):82–83, August 10, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n6/pdf/scientificamerican08101889-
82c.pdf.

Anonymous:1889:PEk

[12561] Anonymous. The Paris Exhibition. Scientific American, 61(6): 83–84, August 10, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n6/pdf/scientificamerican08101889-
83.pdf.

Anonymous:1889:IPKb

[12562] Anonymous. An improved pocket knife. Scientific American, 61 (6):84, August 10, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n6/pdf/scientificamerican08101889-
84.pdf.

Anonymous:1889:IIEa

[12563] Anonymous. An improved ink eraser. Scientific American, 61(6):84, August 10, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n6/pdf/scientificamerican08101889-84a.pdf.

Anonymous:1889:IDPc

[12564] Anonymous. An improved draughtsman’s protractor. Scientific American, 61(6):84–85, August 10, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n6/pdf/scientificamerican08101889-
84b.pdf.

Anonymous:1889:NUC

[12565] Anonymous. The new U.S. cruiser Philadelphia. Scientific American, 61(6):85, August 10, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n6/pdf/scientificamerican08101889-
85.pdf.

Anonymous:1889:VMH

[12566] Anonymous. Views on mt. hood. Scientific American, 61(6):86–87, August 10, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n6/pdf/scientificamerican08101889-86.pdf.

Anonymous:1889:GB

[12567] Anonymous. Gold beating. Scientific American, 61(6):87–89, August 10, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n6/pdf/scientificamerican08101889-87.pdf.

Anonymous:1889:IFBe

[12568] Anonymous. An improved floor beam connection. Scientific American, 61(6):89, August 10, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n6/pdf/scientificamerican08101889-
89.pdf.

Anonymous:1889:IFW

[12569] Anonymous. An improved feed water heater and condenser. Scientific American, 61(6):89–90, August 10, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v61/n6/pdf/scientificamerican08101889-
89a.pdf.

Anonymous:1889:Icz

[12570] Anonymous. To inventors. Scientific American, 61(6):91, August 10, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n6/pdf/scientificamerican08101889-91.pdf.

Anonymous:1889:Cbf

[12571] Anonymous. Correspondence. Scientific American, 61(7):101, August 17, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n7/pdf/scientificamerican08171889-101.pdf.

Anonymous:1889:RPIbx

[12572] Anonymous. Recently patented inventions. Scientific American, 61(7):106, August 17, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n7/pdf/scientificamerican08171889-
106.pdf.

Anonymous:1889:BNBbb

[12573] Anonymous. Books: New books and publications. Scientific American, 61(7):106, August 17, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n7/pdf/scientificamerican08171889-
106a.pdf.

Anonymous:1889:BPbz

[12574] Anonymous. Business and personal. Scientific American, 61(7):106, August 17, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n7/pdf/scientificamerican08171889-106b.pdf.

Anonymous:1889:NQbw

[12575] Anonymous. Notes and queries. Scientific American, 61(7):106– 107, August 17, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n7/pdf/scientificamerican08171889-
106c.pdf.

Anonymous:1889:IIby

[12576] Anonymous. Index of inventions. Scientific American, 61(7): 107–108, August 17, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n7/pdf/scientificamerican08171889-
107a.pdf.

Anonymous:1889:PNPbw

[12577] Anonymous. Patent notes: Patents. Scientific American, 61(7):110, August 17, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n7/pdf/scientificamerican08171889-110.pdf.

Anonymous:1889:RSL

[12578] Anonymous. A recent search light experiment. Scientific American, 61(7):96, August 17, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n7/pdf/scientificamerican08171889-
96.pdf.

Anonymous:1889:ISWe

[12579] Anonymous. The inadequate supply of water for New York. Scientific American, 61(7):96, August 17, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n7/pdf/scientificamerican08171889-
96a.pdf.

Anonymous:1889:STB

[12580] Anonymous. Speed trials of a British squadron. Scientific American, 61(7):96, August 17, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n7/pdf/scientificamerican08171889-
96b.pdf.

Anonymous:1889:CNR

[12581] Anonymous. Construction of new reservoirs for New York Water Supply. Scientific American, 61(7):97, August 17, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v61/
n7/pdf/scientificamerican08171889-97.pdf.

Anonymous:1889:IUS

[12582] Anonymous. An improved umbrella stand. Scientific American, 61(7):98, August 17, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n7/pdf/scientificamerican08171889-
98.pdf.

Anonymous:1889:ILBd

[12583] Anonymous. An improved locking buckle. Scientific American, 61(7):98, August 17, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n7/pdf/scientificamerican08171889-
98a.pdf.

Anonymous:1889:ICCj

[12584] Anonymous. An improved car coupling. Scientific American, 61 (7):98, August 17, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n7/pdf/scientificamerican08171889-
98b.pdf.

Anonymous:1889:ODF

[12585] Anonymous. Operative dentistry-filling teeth. Scientific American, 61(7):98, August 17, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n7/pdf/scientificamerican08171889-
98c.pdf.

Anonymous:1889:MRB

[12586] Anonymous. Manufacture of rolled bars or rails direct from the molten metal. Scientific American, 61(7):99, August 17, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v61/
n7/pdf/scientificamerican08171889-99.pdf.

Anonymous:1889:RNP

[12587] Anonymous. Robert’s new process of converting crude iron into malleable iron or steel. Scientific American, 61(7):99–100, August 17, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n7/pdf/scientificamerican08171889-99a.pdf.

Anonymous:1889:ISL

[12588] Anonymous. Inauguration of the Statue of Leverrier. Scientific American, 61(7):100, August 17, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n7/pdf/scientificamerican08171889-
100.pdf.

Anonymous:1889:PFP

[12589] Anonymous. To Palace of Food Products at the French Exhibition. Scientific American, 61(7):102, August 17, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v61/
n7/pdf/scientificamerican08171889-102.pdf.

Brown:1889:SDR

[12590] Harold Brown. Sodom Dam and Reservoir. Scientific American, 61(7):103–105, August 17, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n7/pdf/scientificamerican08171889-
103.pdf.

Baker:1889:CML

[12591] Charles D. Baker. The Coco De Mer — Lodoicea Seychellarum. Scientific American, 61(7):105, August 17, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v61/
n7/pdf/scientificamerican08171889-105.pdf.

Anonymous:1889:Ida

[12592] Anonymous. To inventors. Scientific American, 61(7):107, August 17, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n7/pdf/scientificamerican08171889-107.pdf.

Anonymous:1889:RPIby

[12593] Anonymous. Recently patented inventions. Scientific American, 61(8):122, August 24, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n8/pdf/scientificamerican08241889-
122.pdf.

Anonymous:1889:BPca

[12594] Anonymous. Business and personal. Scientific American, 61(8):122, August 24, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n8/pdf/scientificamerican08241889-122a.pdf.

Anonymous:1889:NQbx

[12595] Anonymous. Notes and queries. Scientific American, 61(8):122– 123, August 24, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n8/pdf/scientificamerican08241889-
122b.pdf.

Anonymous:1889:IIbz

[12596] Anonymous. Index of inventions. Scientific American, 61(8): 123–124, August 24, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n8/pdf/scientificamerican08241889-
123a.pdf.

Anonymous:1889:PNPbx

[12597] Anonymous. Patent notes: Patents. Scientific American, 61(8):126, August 24, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n8/pdf/scientificamerican08241889-126.pdf.

Anonymous:1889:CAG

[12598] Anonymous. The City of Aberdeen, a great lumber center, Washington Territory. Scientific American, 61(8):111, August 24, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v61/
n8/pdf/scientificamerican08241889-111.pdf.

Anonymous:1889:AAA

[12599] Anonymous. The American Association for the Advancement of Science. Scientific American, 61(8):112, August 24, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v61/
n8/pdf/scientificamerican08241889-112.pdf.

Anonymous:1889:REE

[12600] Anonymous. Report on explosions in the electric subways. Scientific American, 61(8):112, August 24, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n8/pdf/scientificamerican08241889-
112a.pdf.

Anonymous:1889:POS

[12601] Anonymous. Progress of ocean steam navigation. Scientific American, 61(8):112, August 24, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n8/pdf/scientificamerican08241889-
112b.pdf.

Anonymous:1889:SBS

[12602] Anonymous. Secondary battery sub-stations. Scientific American, 61(8):112–113, August 24, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n8/pdf/scientificamerican08241889-
112c.pdf.

Anonymous:1889:ILg

[12603] Anonymous. An improved lens. Scientific American, 61(8):114, August 24, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n8/pdf/scientificamerican08241889-114.pdf.

Anonymous:1889:ITCe

[12604] Anonymous. An improved thill coupling jack. Scientific American, 61(8):114, August 24, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n8/pdf/scientificamerican08241889-
114a.pdf.

Anonymous:1889:ISHc

[12605] Anonymous. An improved safety harness. Scientific American, 61(8):114, August 24, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n8/pdf/scientificamerican08241889-
114b.pdf.

Anonymous:1889:IPRb

[12606] Anonymous. An improved piston rod packing. Scientific American, 61(8):114, August 24, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n8/pdf/scientificamerican08241889-
114c.pdf.

Anonymous:1889:ICba

[12607] Anonymous. An improved hen coop. Scientific American, 61 (8):114, August 24, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n8/pdf/scientificamerican08241889-
114d.pdf.

Anonymous:1889:SLb

[12608] Anonymous. The scientific lantern. Scientific American, 61(8): 115–117, August 24, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n8/pdf/scientificamerican08241889-
115.pdf.

Anonymous:1889:ISBd

[12609] Anonymous. An improved steel belt lacing. Scientific American, 61(8):117, August 24, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n8/pdf/scientificamerican08241889-
117.pdf.

Anonymous:1889:WSL

[12610] Anonymous. The White Star liner Teutonic. Scientific American, 61(8):117, August 24, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n8/pdf/scientificamerican08241889-
117a.pdf.

Anonymous:1889:ITD

[12611] Anonymous. An improved tile ditcher. Scientific American, 61 (8):118–120, August 24, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n8/pdf/scientificamerican08241889-
118.pdf.

Anonymous:1889:ERB

[12612] Anonymous. The electric rolling bridges of the Paris Exposition. Scientific American, 61(8):120–121, August 24, 1889. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v61/
n8/pdf/scientificamerican08241889-120.pdf.

Anonymous:1889:Idi

[12613] Anonymous. To inventors. Scientific American, 61(8):123, August 24, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n8/pdf/scientificamerican08241889-123.pdf.

Anonymous:1889:Cbg

[12614] Anonymous. Correspondence. Scientific American, 61(9):133, August 31, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n9/pdf/scientificamerican08311889-133.pdf.

Anonymous:1889:RPIbz

[12615] Anonymous. Recently patented inventions. Scientific American, 61(9):138, August 31, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n9/pdf/scientificamerican08311889-
138.pdf.

Anonymous:1889:BPcb

[12616] Anonymous. Business and personal. Scientific American, 61(9):138, August 31, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n9/pdf/scientificamerican08311889-138a.pdf.

Anonymous:1889:NQby

[12617] Anonymous. Notes and queries. Scientific American, 61(9):138– 139, August 31, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n9/pdf/scientificamerican08311889-
138b.pdf.

Anonymous:1889:IIca

[12618] Anonymous. Index of inventions. Scientific American, 61(9): 139–140, August 31, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n9/pdf/scientificamerican08311889-
139a.pdf.

Anonymous:1889:PNPby

[12619] Anonymous. Patent notes: Patents. Scientific American, 61(9):142, August 31, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n9/pdf/scientificamerican08311889-142.pdf.

Anonymous:1889:GWS

[12620] Anonymous. Great war ships. Scientific American, 61(9):128, August 31, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n9/pdf/scientificamerican08311889-128.pdf.

Anonymous:1889:SIE

[12621] Anonymous. The site for the International Exposition of 1892. Scientific American, 61(9):128, August 31, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n9/pdf/scientificamerican08311889-
128a.pdf.

Anonymous:1889:PPS

[12622] Anonymous. Position of the planets for September. Scientific American, 61(9):129, August 31, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n9/pdf/scientificamerican08311889-
129.pdf.

Pike:1889:SWC

[12623] Nicolas Pike. Spider web to cocoon making. Scientific American, 61(9):130, August 31, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n9/pdf/scientificamerican08311889-
130.pdf.

Anonymous:1889:IMBb

[12624] Anonymous. An improved machine for bending and coiling pipe. Scientific American, 61(9):130, August 31, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v61/n9/pdf/scientificamerican08311889-
130a.pdf.

Anonymous:1889:ISi

[12625] Anonymous. An improved siphon. Scientific American, 61(9):131, August 31, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n9/pdf/scientificamerican08311889-131.pdf.

Anonymous:1889:AEO

[12626] Anonymous. Astigmatic eyepiece for optical instruments. Scientific American, 61(9):131, August 31, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n9/pdf/scientificamerican08311889-
131a.pdf.

Anonymous:1889:SPL

[12627] Anonymous. A system of protection from lightning. Scientific American, 61(9):131, August 31, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n9/pdf/scientificamerican08311889-
131b.pdf.

Anonymous:1889:ARR

[12628] Anonymous. An apparatus for recording respirations. Scientific American, 61(9):131–132, August 31, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n9/pdf/scientificamerican08311889-
131c.pdf.

Anonymous:1889:IRSd

[12629] Anonymous. An improved reversible safety stirrup. Scientific American, 61(9):132, August 31, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n9/pdf/scientificamerican08311889-
132.pdf.

Anonymous:1889:IMP

[12630] Anonymous. An improved method of propelling sleds. Scientific American, 61(9):132, August 31, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n9/pdf/scientificamerican08311889-
132a.pdf.

Anonymous:1889:MPT

[12631] Anonymous. The Maccoy pneumatic tool. Scientific American, 61(9):134, August 31, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n9/pdf/scientificamerican08311889-
134.pdf.

Anonymous:1889:PB

[12632] Anonymous. A powerful battery. Scientific American, 61(9):135, August 31, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n9/pdf/scientificamerican08311889-135.pdf.

Benjamin:1889:TCM

[12633] Marcus Benjamin. Thomas Corwin Mendenhall. Scientific American, 61(9):136, August 31, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v61/n9/pdf/scientificamerican08311889-
136.pdf.

Anonymous:1889:PEl

[12634] Anonymous. The Paris Exhibition. Scientific American, 61(9):137, August 31, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n9/pdf/scientificamerican08311889-137.pdf.

Anonymous:1889:Idj

[12635] Anonymous. To inventors. Scientific American, 61(9):139, August 31, 1889. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v61/n9/pdf/scientificamerican08311889-139.pdf.